A Complete Bibliography of Publications in Marine Ecology Progress Series: 1990–1999

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

21 August 2024 Version 1.07

Title word cross-reference

< 10μ [PPW96]. < 5 [RLT^{+}94]. < 6 [SSM91]. > 5μ [DLFL96]. ^{+} [SGR92]. ^{12} [Fry96, GT98, JR92, RTRF90, RHT91, RAB^{+}92, SDFJv91]. ^{13} [FMFH97, FH97a, FW91b, Fry96, GT98, HHSvS96, HW92, HAR95, JRMN^{+}97, JR92, Joh96, JLGMP97, Kal91a, LBM^{+}99, MvdVD^{+}97, MJSDF91, RTRF90, RHT91, RAB^{+}92, RRWG96, RWJ^{+}95, RR97b, RSS94, SDFJv91, TFL95]. ^{137} [KI97]. ^{14} [AMB^{+}96, CCV^{+}96, CGA95, CRLP94, Gar91, HWLC98, Irw91, LL92, MKCL94, NH90, ORL97, RTRF90, RHT91, RAB^{+}92, Ric91, SDFJv91]. 15 [FMFH97, SR95, BG91, FHCC98, FCP98, HW92, HAR95, JRMN^{+}97, MvdVD^{+}97, MHM90, PPCR97, RTRF90, RHT91, RAB^{+}92, Rie98, SRSH99, SDFJv91, TFL95, TCB93, WCLB99]. ^{16} [FD97, GT98, LMK93, YKK^{+}98]. ^{18} [GT98, JLGMP97, Kal91a, RVGC^{+}99]. ^{241} [PGD^{+}93]. ^{3} [WR91]. ^{31} [TR91, TRPS94]. ^{4+} [Lai92, MSK95]. ^{l5} [GCH93]. _{1} [SJMC93]. _{12} [GR94]. _{2} [AGP98, BK96a, BR91a, BR97, BRZ99, CSM^{+}98, CMB^{+}99, CH97b, FGB^{+}96, Gar91, GPDF93, GPF95, GPD^{+}96, KRJ98, Kru93, KCD^{+}95, LL92, LL97, MHRV^{+}99, PJF93, PDP^{+}90, RRWG96, RNHO98, Sie93, SJMC93]. _{2}^{−} [BG93, SJMC93]. _{3} [JLS90, LJS90, PS93]. _{4}^{+} [GMG^{+}92, JLS90, LJS90, VNB97, VBN95]. _{O2} [AGP98]. β [DGO92, MNH92]. : [BR97, FMFH97, Mar99b, Som94, TPR98, TRMW95]. δ [FMFH97, FHCC98, FCP98, HW92, HAR95, JLGMP97, MvdVD^{+}97, MJSDF91, RVGC^{+}99, RR97b, Rie98, RSS94, SRSH99, TFL95]. K [O’D98]. μ [SSM91]. n [DBHM99]. P/B [BJTB96]. S_{∞} [Kar95].

-ATPase [SGR92]. -butyl-tin [DBHM99]. -carbolxylation [DGO92]. -D-glucuronide [MNH92]. -fixing [CSM^{+}98, CMB^{+}99, MHRV^{+}99]. -flux [JLS90]. -K [LCTT92]. -number [Ito97]. -polluted [RBG94a]. -pools [LJS90]. -scale [SLB^{+}92]. -tolerant [Sie93].

/aponica [OINN94].

0 [PDP^{+}90]. 0-group [BK90, EG95, EG97, vBDW91, vW93]. 0-year [LAK99].

1 [JM99, KRFL97, KRF^{+}98]. 118 [WKH^{+}98]. 128 [BS96a]. 135^{o} [PLH91]. 1378 [CCV^{+}96]. 14-3-3 [WKH^{+}98]. 17’ [CD98]. 191 [Ano00]. 1970s [Jos90]. 1980s [FAVV98, Jos90]. 1986 [PD95]. 1988 [GST^{+}91, KLMR91]. 1990s [FAVV98]. 1991 [HC94, SKM^{+}97]. 1992 [GGB94]. 1993 [AW98, KL95, Ven98]. 1995 [Ven98]. 1998 [DSG^{+}99]. 19^{o} [Por94]. 19th [JWL96]. 1A1 [GLBF92, SGR92].

2 [GRGK96, LGR95, PD95]. 2-Ethyl-1-hexanol [Ack97]. 2-ethylhexyl [HFW99]. 210Po [DCG^{+}99]. 228 [COS93]. 228/Ra [COS93]. 23 [SKM^{+}97]. 234Th [LBLA93]. 26’ [MMG97]. 2D [SR98a].

3.4-tribromopyrrole [FT95]. 36^{o} [UC94]. 37^{o} [CD98].

4-bromophenol [LSP99]. 4-dibromophenol [GK97]. 4-methylumbelliferyl- [MNH92]. 47^{o} [Por94].

5-bisphos-phate [Riv90]. 5-pentachlorobiphenyl [WKH^{+}98]. 50^{o} [MH99]. 50’ [CD98]. 54^{o} [MMG97]. 55^{o} [LGH95b]. 58^{o} [LGH95b]. 5” [MMG97].

6-diamidino-2-phenylindole [HP98].

75* [MP96b]. 75^{o} [LGH^{+}93]. 77* [PLG^{+}96]. 77*-81*N [PLG^{+}96].

8 [BCG97]. 81*N [PLG^{+}96].

= [NP90, Sto96]. a-glucosidase [KFPR94]. a-specific [BAM^{+}99]. A. [GS99b, SG96, TK95a, TK97, TK98, WCP^{+}98]. Åland [BDR^{+}96]. Aarhus [HR96b, JLS90, LJS90]. abalone [AAE^{+}96, LK93, Sch93]. abandoned [HKA96]. Abashiri [BTKK99]. abasrer [FMRC93]. abatement [DG96]. Abatus [PF95]. abdominal [TR91]. aberrations [CB92]. abilities [SB97c, VKF^{+}98]. abiotic [GPB^{+}98, GRSV98, MS96a, San93b]. Abnormal [PLIM95]. above [PBGE91, TGvJW95, VWS92]. above-bottom [PBGE91]. Abra [GD96]. Abraiia [Big92]. absence [TLB93]. absolute [SR97]. absorbing [HS98a, JMD^{+}99, Les96, LM97, RR97a, SDC^{+}92]. Absorption [LSH96, AC92, AFZ96, BAM^{+}99, CGA95, DL91b, EAD92, FG93, HS91a, JNJP94, JSGS94, MG96c, NP90, NPB93, Sos96, SL97b, VMS^{+}96, vdL98]. absorptive [MGC95b]. Abundance [CCG^{+}92, DS96, DRP^{+}97, GM99, HKT^{+}96, MPM97, NTK98, PWR96, SB98, Sha98, SSA96, Whi94, ÅP97a, ADA98, AWH93, And94b, AB96a, AWH^{+}99, BC91, Bar97b, BM99b, Bee97, BH97a, BCAAC96, BH93, BWM97, BM95a, BLR94, CT92, Cer98, CGNH98, COK99, CCG96, CLB91, Con98, DPM96, DE90, ELH92, Egg95, FW92, FDZ93, Fow90a, Gas94, GH94b, GAR93, Góm95, Gon92, GF94, GF91, GP98, HCS93, HD93, Hic92, HSW99, HNvB94, Hul97, Hul99, HOM95, JP94, JR92, JFS90, KW97a, KW99a, Kur99, LZYW93, LBH91, LAJ95, LDKR93, LTM^{+}98, LBK90, LB93, MH96a, MB95a, MR97a, Mar97, MC91c, MF90a, MF90b, ME97, MLES94, MM95b, MHM90, MJC97, MJ92b, Nor94, OM90, OMH95a, OMH95b, PDD^{+}99, PS95, PS97, PMOD90, PWNW99, QCM94, RWJ^{+}95, RGE^{+}95]. abundance [RZWY96, RDA97, SRH^{+}98, SKvB98, SR98b, She96, SD94, SBT97, SB96, SDG95, SHO96, SR96, SC93, SC95b, TDBH94, TGBB97, Tok95, Tol98, VH93, VHHS96, VHH^{+}98, VPM96, VD99, WFL95, WH95, WE95, ZF97, ZF99, ZDJ^{+}95, ZR92]. Abundances [Alo90, HS99b, Ken92, LPF94, MMHW96]. abundant [PF90]. abyssal [MF98]. Acanthaster [BM91a, CED91, HK96, HKR96, HR96a, KH92b, SP93]. acanthurid [CA96b]. Acanthuridae [SML97]. Acanthurus [HK96, HR96a, McC99, Pla93, Ris97]. Acartia [BD94, CA96a, CA97, CLW99, DDW90, Dut98, FD98, FR91, GA97a, ICG96, Jeo94b, JT90, KSV96, KSV99, Nor96, RMMB91, RD92, SA92, SK95a, TDD94, UCV98, WDD92]. acaule [Gar99a]. acclimated [HLF^{+}91]. Acclimation [IPT94, Les96, GMK97, MRMR92, MP95, RMMB91]. acclimatization [OSJ93]. accompanying [AOT97]. Accumulation [BC93, CC92b, FBL95, GSC96, GS99b, IWF^{+}97, MKYI98, MBL^{+}98, ND91, PGD^{+}93, ZCL97, DvRN99, Fra92, LGdC^{+}90, MM96, ROM^{+}91, WF97, WJD93]. accumulations [NB96c]. Accuracy [ACC90]. Accurate [TJST98, MM99]. acetate [KKH^{+}94]. acetic [MH93b]. acetylene [LKvRH96, SCLS94]. Acharax [BL97]. Achoerodus [Gil97]. Acid [WJD93, BLKA93, BFMM99, BBCB99, BUCP^{+}92, CCC94, CG93, CWHA99, DKL94, DQT^{+}97, DMP^{+}99, FJR^{+}94, Fuh90, GHTH94, GSS93a, HSN93, IFL97, MHBW90, MCA90, MVA99, MH93b, Naj96, Naj97, NB90, ND96, PM90a, PRM^{+}96a, PRM^{+}96b, PLBP96, PFMC95, PFMG96, SAH98, SPH94, SH92b]. Acid-base [WJD93]. acids [BHQ99, CM99, DL94a, FF92a, FFJ93, FTBR97, HQ99, JMPH90, JKC^{+}93, KK91, KZK^{+}95, LH98, LNSL91, MBK95, NK90, PMOD90, PDB^{+}97, RTLG97, RFGF92, SUH95, SAH98, SDC^{+}92, Sim91, SCF91, WP95]. Acknowledgements [SDA^{+}92c]. acorn [BGBS91, BGW96, SBBG94]. Acoustic [BLJ99, BW96, MA99c, ZGM^{+}96, BMHD93, CSB94, KKW97, LB93]. acoustically [CWH98]. acquisition [GKRD96, KM99b, KR94]. Acropora [Byt90, JLGMP97, MB95b]. across [CK99, CT96, EDS^{+}99, FGP98, FVDL95, GFPD92, HNC92, JJNR95, Mun93, Sim92, TA93, TUN99, VBF94, WC90]. acrylic [DKL94, SPH94]. act [Han91a]. acting [HB97]. Actiniaria [AS95]. Action [LL96, BHQ99, KPS97, MM91a, VWM90a, VWM90b]. activation [WCP^{+}98]. Active [BNW94, GdGMD95, Ker90, DAB^{+}91, HB97, JT93, KM99a, MC91b, ODS^{+}96, PH91b, PL99b]. activities [CK95, DdA95, FCW98, GSBB97, LVP98, MA93, Nau96, RM97a, San93a, TMC98, vDdWKW99]. Activity [GSG^{+}94, LSP99, MNH92, Van92b, WWL97, AAD91, AC92, Bar97b, BM91b, BLM^{+}95, BDNS99, CCC90a, CC97a, CS99, Dem95, DGO92, DE91, EGKE92, ES90, GBC92, GGF99, GDP^{+}98, GCKW95, GSB98, GSS93a, GH92, GDW90, HWLC98, HL90, HRÆ94, Hil91c, HW99, HNL94, JLD^{+}96, KKL^{+}98, KBP98, LBV^{+}99, LS91, MG96a, MRMR92, MRPC98, MHMH96, OJR96, PNB94, PFH98, PMOD90, PH95, PDP^{+}90, RGLS99, RSH93, Rel96a, Rel96b, RA92, RVBL92, Riv90, RMMB91, RBN99, SMTL99, SCJ99b, SO96b, SSV97, SDG95, SPPL92, SCKS90, TGBB97, TB98, TGB99, UNPH93, UMT^{+}91, VAR94, VD94, VL97, WHG99b, Zaj91a, Zel97, dT94a]. activity-related [Dem95]. acuminata [GGC99, SBT97]. Acuña [Ano00]. acuspes [Nor96]. acutifrons [GSMM96]. acutorostrata [KNI95]. acutus [ASSWM97, HKG93, LHL93, SSHM91]. Adamussium [MONR90, VCMP93]. Adaptation [VMHH90, BMD^{+}90, DF95, EBB95, HBD95, IPT94, KB91, LB91, Ter93, ZAN99, vLTW^{+}98]. Adaptations [WHK97, AOOT94, AC96b, JGT96, SC92a, YKL99]. adapted [RT92]. adaptive [CM98b, SP99a]. added [And96, Eno93, HHB93]. addition [HJA99, SK98a]. additional [PM94b, Raw96]. additions [RGS90]. additive [MR97a]. Aden [BS95c, SL91, Uib95]. adenine [RBJ90]. adenosine [HSRHH90]. adina [BB93]. adjacent [HSK^{+}94, JDSB99, Mor96]. adjust [Röp93]. Adriatic [BH92c, BH92d, KH92a, MMM^{+}95, PH91a, BSVK93, CCS98, FMRC93, MH93b, MNH92, NPB94, Naj96, Naj97, PF91a, SH95a, SK94b]. adsorbed [LGV^{+}99a, LH98, NK96]. Adsorption [SAH98]. adspersus [NH98]. Adult [AR91, ST98, THPC96, ÅP97a, ACLS97, AH93, CLW99, DE91, FFT98, FH97b, IHK^{+}98, JAOS90, Ken94a, LFS90, Mar91, MGC95a, Nor92, PNS90, PHT96, RA96, SGP99, TWWE95, TRS91, Tol98, UMTM98, VEW99, VÅCL99, WW96, WH93, van94a, vdL95]. adult-larvae [SGP99].

Adult-larval [AR91]. Adult/juvenile [THPC96]. adults [She95, OMH95a]. advantages [Pec99]. advected [PM92a]. Advection [DLJ91, Hil91a, CC97b, CHDW92, JC95, Kep91, Wor95]. Advection-diffusion-mortality [Hil91a]. advective [PG91]. aeglefinus [Cam97, FQV98]. aequizonata [HCF93]. Aequorea [Pur91]. aeration [JR92]. aerial [Lee90]. Aerobic [HB91a, BMTF92, VDVC94]. Aethotaxis [KZ92]. affect [ARF92, MS93d, Ols93, San94a]. affected [DÁSP97, HMVHH95, RÁSP99]. affecting [CC91a, CBL96a, MJW96, MC93, PMETM96, WD99, ZD99]. affects [BR97, MPCD90, SO96b, WC96, Yam97a]. affinis [DWG98, AJS95, AN97, COS93, HN91, ICG96, Joh97, KEV99, Leh96, LA98, OE91, SE98]. Africa [BA91, BSF95, CBL96a, CBL96b, DL93b, Dye93, Dye98, Las98, WW90]. African [GB97a, GB97b, RFNBHG94, Whi94, vG91]. africanus [RDA97]. after [Beu92, BC99, DvdM92, DE91, GPP98, GJ96c, GBJ94, Jon97a, MYYT96, MGD91, NH95b, PPCL^{+}98]. Ag [CL90]. against [CP95a, Irw91, PCTF95, Raw96]. Agarum [DC91]. Age [Ang91a, Big92, GCC95, HKMK98, HBP^{+}93, HMB93, HMB95, JD90, KA93, LT97, MCM^{+}97, MB95b, NAS95, RR92, RE96, SCMM95, Ber92, BCL93, BL94b, BBS93, CL90, CN92b, DMA^{+}94, HRFC90, Ito97, JSH99, KJGC96, LAJ95, Lop91, MCF90, McC94d, MM95a, MGC91, RCTH99, RGE^{+}95, SF95, SPCF97, VDM^{+}97, WTO96, HR96a]. Age- [Ang91a]. age-class [MGC91]. Age-related [HMB93, HMB95, KA93, MB95b]. Age-specific [GCC95, HBP^{+}93]. Ageing [SSW^{+}96, RFNBHG94, WT98]. ages [SGM98]. aggregate [HKA96]. aggregated [KE95]. aggregates [AV95, ES99, KY97, MH93a, Naj96, PJ99]. Aggregation [Cha95, Rie93, IMB99, CP95b, HlS90, Kep94, LSM92, OC97, RK95b, RS94, Rie91a, Rie91b, SL94, TT98, UKR^{+}91, WGD97]. aggregation-factor-induced [UKR^{+}91]. aggregations [BH92c, BH92d, GP90, KH92a, KEJ98, Sam97, Zel98]. Agile [PD90]. aging [DD97, KFD99]. Aglantha [PGG96]. aglomerular [CDI96]. Agulhas [DL93b]. aided [AAD91]. Aiguillon [RSN^{+}99]. Aiptasia [MPH96, MPCD90]. air [FGB^{+}96, GPDF93, GPF95, GPD^{+}96, JBO97, MM90]. air-sea [FGB^{+}96, GPDF93, GPF95, GPD^{+}96]. airborne [MGEC97]. aircraft [WHIC97]. AK [KRFL97, KRF^{+}98]. AK-1 [KRFL97, KRF^{+}98]. Akajima [HSK^{+}93]. akashiwo [KYI^{+}98, YKK^{+}98]. al. [BAR96, DHB^{+}98, DD97, JNCS90, Käh98, Mil90, PPFP99]. ALA [Bro93]. Alabama [CPB96]. Alaska [DH96, SH92b, AP99b, BM94a, BCN^{+}95, BBK91, CBI91, FAHW99, HS92, HBP^{+}93, HH90, IFL97, KO92, MN99, Neu92]. Alaskan [CH96, FS90, SH97a, SS96a]. albacore [YD90]. albatross [WWL97]. albatrosses [CW95, WWC^{+}99, WW92a]. albocincta [BA97]. albolineatus [KW97b]. Alboran [BMP^{+}95]. alcid [HS94b]. alcyonaceans [Dai90]. alcyonarian [ASN93]. Alcyonium [Gar99a]. Aleutian [CWH98].

Alexandrium [AKM^{+}99, AGGKA96, BLCA90, DRGV91, Dut98, FFF^{+}94, MCC97]. Alfacs [Vid94, VM95]. alga [Åbe96, ÅP97a, AdN91, ATH95, BB97a, BEH93, BH95b, CC98, CC99, Ekl98, FSA^{+}97, GS94a, HN90a, HN90b, Han95, IG96, JPZ^{+}98, KMB97, LLB^{+}95, Mar91, PCLÅ97, PB96a, PKY96, Poo94, VBN95, YKK^{+}98, vTSD97]. algae [AV98, AR99, BCC94, BCC95, BCNC96, BDA^{+}99, DFS94, DKVD95, EP97, FTP99, Fra95, HBM93, Hul97, HHN90, JMR91, JSOG91, JBHH97, KW97a, Kin92, Kin95a, KW99b, Kuo91, LC93, MSD92, MGD91, NB96c, RLT^{+}94, RKS97, SSM91, SCEG93, Ste94, SM90, Tha92a, VEW99, VNB97, WSMM99, Wer97]. Algal [BWRH99, APCL^{+}93, ABB^{+}97, AD90, BCH95, BEH93, Eno93, FM95a, Gra96a, HCS90, HPF92, HS91b, KW98, KW97b, KD96, KM92, LDV95, Leo99, LH96, Lig90, LH99, MBJ94, MGDT90, MG99, NT96a, NK93, NB96b, Ols93, PB96b, PCBA94, RH90b, San93b, SD93a, SMB98, TT90, TSB93, TL95, Tro92, Tro93, VEW99, VDM90, WFL95, WK92a]. algal-bacterial [TSB93]. Algarve [EMR^{+}97]. Algeciras [CNGG96, NCGG96]. Algerian [MSM^{+}98]. algicidal [KYI^{+}98, YKK^{+}98]. Algoa [WW90]. alinae [BL97]. Alkaline [LVP98, GGKJ^{+}98, MA93, vBV90a, vBV90b]. alkalinity [CH97b]. allelochemicals [DBvA91, VT92]. allelopathy [MSC95]. allocation [AZKA99, HHSvS96, HH95, Rin96, SF95, UTB^{+}95]. Allogeneic [FBR^{+}95]. Allometric [Dua91, TTR99]. Allozyme [AW97, JKT90a, Väi90, GS90, JDLL95, QZA95, Way95]. allozymes [HSBP97]. Allozymic [Fev92]. ALOHA [STK99]. alone [GQ97a]. along [AHS97, BBFD98, BP98, BVL92, BMH93, CBB90, CS97, EW91, FPB^{+}94, FF97, FSG^{+}97, FSG^{+}98, GBC92, GGPGG95, GHL^{+}97, JSJMH90, JJNR95, KBWC90, KDMR97, KR92, LSG^{+}99, LM91a, MDC98, McF92, McG92, MGC91, MGC95a, MKN^{+}96, MNO96, O’D98, OSC97, ODS^{+}96, PLH91, PKL97, Pin94, RSH93, RPN95, Rie98, SSR99, SO91, SH92a, SHV91, Tom91, VBL^{+}92, Wes90a, Wes90b]. Alosa [Lim95]. alpina [MJ92a]. alpinus [RSM^{+}96]. altemiflora [Blu93]. alteration [Cro98]. alterations [LH96]. Altered [RO97, BDH94]. alternate [PS95]. alternation [PS94]. alternative [BC91, GB93, PM94a, Ruz96]. alterniflora [ATNW97, CNP95, KNK90, MW97, MH90, Pak92, SCC98, Tur93, TZ99, dG91a]. alvinellid [CJ93a]. Alvinellidae [JDLL95]. Alvinocaris [PDB^{+}97]. always [DGCT93]. Am [PGD^{+}93]. Ambient [GMT99, CNCS93, CCC94, CC97a, JPA^{+}96, NDG^{+}99, RWJ99a, SC90]. ambiseta [SFB92]. amblyrhynchos [WCL97]. amboinensis [Ker96]. ambush [BGH95]. ambush-predator [BGH95]. amencanus [BBS93, WB92b]. America [WT98]. American [BS93, BSB91, BCv^{+}92, JPC97, KCS94, Lim95, MB91, PWS98, RM95, RGT98, SCC94, Tok95, WS91, WTO96, WT98, vCRS92]. americana [ND98, SB96]. americanus [BAH94, BS93, BSB91, BSH^{+}91a, BSH^{+}91b, HF99, HKM92, IW91, JPC97, JC92, KCS94, Kea90, PWS98, WS91, WTO96, WA95]. amines [FF92a]. Amino [PFMC95, BHQ99, CCC94, CM99, CG93, FF92a, FFJ93, FJR^{+}94, Fuh90, HQ99, JKC^{+}93, KK91, LH98, MBK95, NK90, PM90a, PMOD90, RTLG97, RFGF92, SUH95, SAH98, SDC^{+}92, Sim91, SH92b, SCF91, WP95]. aminopeptidase [KFPR94, MA93]. Ammonia [Mat92, CNCS93, CCC94, CC96, CC97a, HS94a, SUH95]. ammonia-N [CNCS93, CCC94]. ammonification [MB97b, OSGB96]. Ammonium [DW95, GTL^{+}91, MLL98, vKVSR97, BCH95, CG93, Dor90, Fly98, Fly99, FTBR97, GSBB97, HDPW93, KK91, LGV^{+}99a, LH98, MBK95, PDP^{+}90, RWJ99a, RFN98, Tam95, TPR98, TTR99, TUN99, TK91, UK98, VH99b, VM95, WYY^{+}98]. ammonium-nitrate [Fly99]. amoeba [CRLP94]. amoebae [MRMLP97]. Among [BSH^{+}91a, AMK99, BCNC96, BA96, CWWH99, FB91, FWWL99, FKC^{+}90, Fra93, HBD95, Hob93, HM97b, HSCC96, JA91, KSH97, LVP98, MACR98, MDRW99, NB98, NTK93, PC96, PB96b, Pud93, RTRF90, Rei90, RH90a, SDFJv91, SL97b, VHVB^{+}98, WC96, WLB^{+}99]. Among-location [BSH^{+}91a]. amorphous [BH92c, BH92d, KH92a, Naj96]. Amounts [GCF96, SOK98]. Ampelisca [She98]. Ampharetidae [MT94]. Amphiascus [GCC95, WC91]. Amphibolis [VWK96]. amphioxus [Sto96]. Amphipholis [GSP99]. amphipod [CPT90, CBT98, CD99, CH94, DL93a, Fra93, GH93, LT93, Leh96, LA98, PP96, PNB94, RK95a, SDC99, SHC96, SE98, TW98, vCRS92]. Amphipoda [Fra93, CV96, GD99, Moo94, She98, SS91]. amphipods [EE98, GH92, HC92b, Joh97, OE91, Poo94, RATA^{+}99, Wer97]. Amphiprion [AMFB99, MF97, RHH97]. Amphisamytha [MT94]. amphitrite [QGQ97, QQ99, WCT95, ACOK95, GQ97a, PGB91, QGQ97, QQ99, WQQ99, WCT95]. Amphiura [GS99b, LJSK96, RNHH97, SLR94, SG96]. Amphiuridae [RNHH97]. amplification [AFZ96, MRF^{+}99]. amplified [CCGM93, HBBT98]. amplitude [BVW^{+}98]. amurensis [BL95, CTSN90, DL91b, DL94a, LWL98, NTS90]. Anadara [Sca93]. Anaerobic [HS90, JBB90, BMTF92, HB91a, MB97b, SO91, Sie93, STS99, VDVC94, VH96]. anaerobiosis [Oes90, SP99c]. analogue [FF92b]. analyses [BDdWC96, Dye98, FCMB94, HM97c, KSRB99, MHvB97, MP96c, OSC97, SM90, TC93]. Analysis [AGW93, BM94a, BCNC96, BP96b, LAJ95, MAC92, PCC^{+}97, URHR98, WTM^{+}96, ASRC99, Aer98, AD93, AMO^{+}99, BSH^{+}94, BKWML93, BPC98, BBCB99, BKB94, BWHC91, CA96b, CH94, CCR98, CNP95, DPO91, DL91a, DL93a, DEP^{+}96, EF97, EJ93, FJR^{+}94, FCCM94, FDW92, FH99, FL92, GGC99, GTWF98, GCC95, GCNSNd94, GBB^{+}98, Har90, HK96, HH94a, HT96, HS97b, HW92, HAR95, JSGS94, JE95, KSE96, KCV94, LMF93, MC91a, MT91, MDD95, MSP^{+}97, Mil98, MO97, NJLD98, Ohm96, OG95, PM94a, PPO^{+}99, PSR^{+}99, PM98, RCS^{+}96, RS94, RRGB96, RK99, SR98a, SSY98, SM95, SM97b, Sym95, TJ90a, Tho91, TJ90b, VWD^{+}95, VK90, WTO96, WGD97, Way95, WMPT94, WCLB99, WJM^{+}91, YB90, YKK^{+}98, ZCPA94, vdBTvDR96, vdVW99, MDTB94]. analyze [LBV^{+}99]. anceps [PF92a]. anchialine [PIC97]. Anchoa [CC91a, CH90, JAOS90, LB93, MH96a, NH95a, PHT96, PND^{+}94, WH95, ZHM91]. anchored [WK98]. anchovies [JAOS90]. anchovy [CC91a, CCS98, CH90, CRH^{+}99, GCR98, LSB95, LB93, MH96a, NH95a, Pal91, PHT96, PND^{+}94, RCC^{+}99, SKPT97, TLT93, TP95, TP97, WH95, ZHM91]. ancient [KML91, RBSO99]. Andaman [DB96b]. anemone [AS94b, BA97, MPH96, MPCD90]. anemonefish [AMFB99, RHH97]. anemonefishes [SJC99]. anemones [AMB99]. anemoneshrimp [FGH95]. angles [PZ91]. anglica [AWdS92, AWH93, AWHH94, HHSvS96]. Anguilia [OINN94]. Anguilla [AHY^{+}99, AOT97, ALOT99, AALT99, CT96, RBN99, TSW97, WT98]. anguillid [AMO^{+}99]. angulata [MR97c]. angustata [Poo94]. angustipes [AHMd90]. animal [MSK95]. animals [JPR92, LD93, SCC94, Tay98]. animones [MGD91]. anisotropy [DPM96]. annelid [GCGS90, VBCRB97]. Annelida [DAE^{+}99, MGDT90, QC94, TD92, TDW96]. Annelida. [Vis90]. annelids [HH99]. Annual [ADR95, Cam91, CG99a, DSW^{+}91, DSV^{+}91, HRÆ94, HSW99, KKL91, KK96, NKKO97, Roa93, SK95c, ZBW92, Fow90b, GAR93, HK93, Ito97, MS97, PRC^{+}92, PB93, RCG99b, RE96, VHHS96, VHH^{+}98]. annularis [MNB94, Szm91, VB93]. annularis. [Van94b, VK94, VB94]. annulata [Cra94]. annulatus [DEF91]. annulipes [ÓN97]. anomalies [ASN95]. Anomura. [JA91]. anophagefferens [MC97b]. ANOSIM [CU99]. anoxia [BFVG92, DvdTCC97, DS93b, FKR90, HS94a, MLC97, MvdZH^{+}97, SGC^{+}98, VDJA98, WW91, WW93]. anoxic [And96, BR91a, DGT^{+}93, HHB93, SD93a]. Antarctic [KWEG94, Øre90, Øre95, ARF92, AMB99, ALB^{+}96, AMB^{+}96, BMHD93, BMSHC98, BGT91, Bra93, BBCB99, Buc91, BBG90, CWHA99, CRP99, DM91, DEF91, DFR95, FD92, FPPM96, GB91, GT92, GKRD96, Gon92, GGD99, HKG93, HVK96, HMVHH95, HSRHH90, HD93, HTCW96, HHE92, HNL94, ISF^{+}98, KL95, KEJ98, KHFP^{+}96, KH98, KZ92, LNQT93, LS91, MWA93, MS93a, MONR90, MRMLP97, MACR98, MBWT98, McC94b, MB97a, MSS95, MP96c, MO95, MA96, NT92, Nev99, NSC^{+}92, Øre91, PWR96, PPM94, Pak95, PM90b, PPMF97, PWPS95, PLS^{+}95, PDP^{+}90, RR97a, RPT^{+}96, Ruz96, SB94, SW98b, SL95, SMB97, SHM^{+}97, TW98, TAD^{+}94, TDH^{+}94, TPW^{+}93, VSQ90, VCMP93, WSCT96, WBP^{+}92, WMD99, WC98, WR97, WCBL94, Wöh96, WHK97]. Antarctica [GWKW97, SK97, ALBS96, ARS93, ADG^{+}95, ASP^{+}96, AS97, AWH^{+}99, Bar95, BH92e, ER90, GP90, HC94, KR97b, KGW93, LAG96, LHL93, Nor92, Nor94, PBVB99, PMS94, RKS97, RAlBW99, SSR99, SSHM91, SBP92, SBP93, VWD^{+}95, VAV99, WtD90, MOM98, Øre91, Øre95, RR97a, SMB97, SS95, SCDH98, VWK96, WMG^{+}94]. antarctica-dominated [SCDH98]. antarcticum [WHK97]. antennatus [DF92, Dem95, MC97a, SMT97]. antennules [BP93]. Anthothoe [BA97]. Anthozoa [AS95, Gar99a, YAG92]. anthropogenic [MV98, NO90, RCH98, RSAD98, SSGD96]. anthropogenically [SSG96]. anti [DCM90]. anti-herbivore [DCM90]. antibodies [VPGV94]. anticyclonic [YSY^{+}99]. Antifer [SBT97]. antifouling [WS92a]. Antifreeze [Wöh96]. Antimicrobial [KKL^{+}98]. Antioxidant [LACM92, AGP98, BAOT96]. antiqua [IWF^{+}97, NU91, NYH92]. Antofagasta [EM99]. Antonio [MC96a]. Ao [KKH^{+}94]. Aphally [BQR99]. Aphanius [Mal99]. aphotic [LBL91]. apical [TDK97b, TDK97a]. apiculata [COG97, DGHT^{+}98]. Aplodactylus [OC95]. Aplysia [dNSR^{+}96]. Apo [RA96]. apochlorotic [NAI^{+}93a]. Apogonidae [FGG97]. aponica [OINN94]. apoptosis [BBR^{+}93]. apostome [CPS91]. aposymbiontic [CGZR94]. apparatus [JPR92]. Apparent [HdWS91, BP97a]. appendicularian [ADBH99, LUA99]. appendicularians [Flo91]. Appendix [SDA^{+}92c, SDA^{+}92b, SDA^{+}92a, SDA^{+}92d]. Application [HF99, FM94, Kir92, KZP93, LGJ99, MMHW96, NK90, PLP92, RVGC^{+}99, SR98a, SL97a, VBCRB97, WH98, vR93a, CC91b, ICG96]. applications [Ano99a, HdWS91, MDD95]. applied [NG96]. Applying [Gar98]. approach [BHT^{+}92, CLM^{+}93, CRGZ98, GSB98, GM98b, LVL^{+}99, LJM99, MM91a, MU99, MCNB98, Mee99, NA95, OE91, PKL97, RSAD98, SW99, SR98a, SD99, TLT93, UMT^{+}91, vR93a]. approaches [GDMS90, KV90]. April [SKM^{+}97, GGB94, PD95, Ven98]. Aptenodytes [ROANM98, WR97]. Apulian [FPB^{+}94]. Aqaba [GAN^{+}94]. aquarium [JPZ^{+}98]. aquatic [BHG94, Bow97, ERC95, HB91b, JTPP92, LGJ99, ML93, MB95a, Min93, MW97, ZCPA94]. aqueous [KRJ98, SGL^{+}97]. Arabian [BSP93, CSM^{+}98, EAT95, Joc95, Mik97, Röp93, Wit99]. Arbacia [RZLB96, SRH^{+}98]. arborescent [HB97]. Arcachon [IF96]. Arcartia [WR92a]. Archaea [FD97]. Archaeomysis [TK95a, TK97, TK98]. Archipelago [CRCS96, PDBA90, Per91, SAvD^{+}96, PM92a, SR93a, BGG^{+}97, DCBG99, GKB^{+}97]. architecture [GS91b, HM91, SSV97]. Arctic [LDV95, SB98, AKL^{+}99, BHHL96, CLKO98, CG99b, CMBM99, CPH^{+}96, GH91, GHHC98, Gra96a, Gra91a, GWWP99, HC92b, HW92, Hob93, KTGC99, KCI98, MBJ94, MG99, PBv^{+}95, PS97, PMOD90, RSM^{+}96, RTRP^{+}98, RNH99, SC92b, ST90, SCEG93, VD94, Wer97]. Arctica [Oes90]. area [AFC96, COG97, DFJ^{+}93, DBN^{+}94, DFR95, EM99, FD92, FS91, FH96, FS98, GGH99, GTL^{+}91, GST^{+}91, GD96, GK96b, HK99, JMD95, Leh96, LA98, MJV92, NLR^{+}93, OB94c, PF92b, RC91b, RHJ92, SRH^{+}98, SC92b, SFP91, TPR98]. areas [FGG95, KS94, KR97a, KNI95, LMS^{+}93, MBF^{+}97, MLDM95, MF95, TJST98, WSP^{+}95]. arenaria [AR91, BDB97, DMM99, LC97a, SGP99]. Arenicola [Ack97, BAOT96, FB94, GR91, HBBM96, HBS98, HSBP97, Hüt90, KF97, Phi94, PM94b, RBA98, RBT96, SP99a, VHG95, WCP^{+}98]. arenicolus [SW91]. areolata [Joh92]. areolated [LCW99]. areolatus [LCW99]. argenteus [PNS90]. Argentina [WSP^{+}95]. argentinus [ASRC99, CA99b, WTR99]. Arginase [CC97a]. argo [HPM92]. Argonauta [HPM92]. Argopecten [AV95, AI92, AMBH98, PBGE91]. argus [AADSCC97, BHH^{+}95, ELMCC90]. Aristeus [DF92, Dem95, MC97a, SMT97]. aristotelis [BRLM90, WFHM99]. arm [Aro91, LL94, PLFSW97]. Armases [Ang95a, Ang95b, CCDA98]. armed [DEF91]. aromatic [MCSV95, ONR98]. Arosa [FP91b]. Arousa [ASRP^{+}96b, ASRP^{+}96a, FGG95]. arquata [RD97]. arranged [VSPLH98]. arrangement [HN90a, HN90b]. arrest [PlMM94]. arrival [CT96]. arsenate [WW92b]. art [Rel96a, Rel96b]. artefacts [McG97b]. Artemia [BFD91, Bur94, VTNA91, ZGB90]. Arternia [BKB92]. arthropods [Mar99a]. artifact [LF93]. artifacts [Mic96, PB94]. Artificial [AADSCC97, BJTB96, JMPH90, Sca96, CES^{+}94, CSE^{+}96, Eml90, Lav97, NB98, Tom91, Van92b]. artisanal [WO94]. arvae [BN91a]. asbestinum [HFR^{+}93, WHW93]. Ascension [HLP^{+}99a, HGB99]. ascent [WH92]. ascidian [BLJ^{+}96, DBvA91, Hur92, NCGG96, PLF90, PR92, PST95, RCG98, RB90, TB92, VY96, VPF98]. ascidians [DY92, WJF94]. ascogiossan [Tro93]. ascoglossan [Tro92, Tro94]. Ascophyllum [Åbe90, Åbe96, ÅP97a, PCLÅ97, VWM90a, VWM90b, VÅCL99]. Asexual [BA97, PWNW99]. Asia [HTLK99]. Asian [CTSN90, KHB91, NTS90]. aspect [MK93b]. Aspects [BAOT96, Sal95, JKL90, Ker90, Rie91a, Rie91b, SGC^{+}98, SFB92, VB93]. aspera [ABB^{+}97, LB96a, Sak97]. assay [GSS93a, NKI^{+}96, Tha92b, UKR^{+}91]. assemblage [ASCW90, AFC96, BCC94, BA91, BSGM94, BG99, Cer98, CWB93a, ECF^{+}97, GGB^{+}99, HFGT94, Hul99, KL95, LDV95, LSG^{+}99, McC94c, RD95, SS96a, Tok95, WB92a, WM92]. Assemblages [BC99, AKM^{+}99, Bia91, Bia92, BK94, BAY91, BSY^{+}99, CM98a, Cha98, CU99, CCG99, CLK90, CBL96b, Con95, CSM99, DF94, FSdM^{+}92, FB91, FP91b, GDvS98, GB91, Gla99, HPF92, HM97b, JJ99, KW98, LJ95, Las98, LW94, MG96b, McG94, MS96a, MS96b, Mil95a, NB90, ÓM92, OB94c, Pel96, RAS92, RBJ90, RSAD98, SSR99, SW99, SLHA97, SCDH98, SS99b, SGMS99, SHV91, SH95b, TWS^{+}95, TRU91, TK91, VWJ96, WA99a, Wit99, YFI91]. assess [BL95, BSW93, Sch99, SW91]. assessed [BMS^{+}97, GA97a, GF94, JLD^{+}96, PM98, PPCR97, SKM95]. Assessing [ASL99, DFB96, CLD99, CA98, DD97, FC90, Jon97b, JHG99, RM96]. Assessment [CBDC96, HHSvS96, SGB90, Way95, BVL92, DCK99, FSCE92, Gas94, Hal98, KFD99, SD99, SM97b, ZC92]. assimilated [LC95]. Assimilation [CJL91, FNF91, FR91, LC95, WFL95, ABY97, AMB^{+}96, BHHL96, BBB^{+}94, CLC^{+}98, HL90, HB94b, LBV^{+}99, SDC99, TT90, VNB97, VBN95]. associated [AWHH94, BMTF92, BHHL96, BQKC97, BP96b, Cer98, CB96, DL93b, DMA^{+}94, FB91, FHCC98, FAHW99, GT92, GCGR90, HDPW93, HKK^{+}95, JBWH97, JMR91, KH92a, LDV95, LS91, LJS90, LF97, NK93, OSZ95, PF90, Reu98a, SO96a, Sal95, SC90, SDC99, SW97, SdC97, SB96, SH99a, SPS98, SR96, SPZ^{+}93, TA90, VMS^{+}96]. Association [BMBD95, CSR^{+}92a, PMS94, Bro98, CMB^{+}99, HHM97, IWF^{+}97, KOSS90, Sym95]. associations [BH97a, McC94c, dG91a]. assortative [Cro99, SMUS^{+}99]. assumption [LF93]. Astarte [Oes90]. Asterias [TLW^{+}96]. Asterina [Kur99]. asteroid [RHH94, RMH96]. Asteroidea [BM91a, DEF91, Geo94]. astreoides [NB98]. asymmetry [SKPT97]. Asynchronous [Cli95]. at-sea [HW99, TMCE98]. At-sea-activity [GGF99]. Atapozoa [PLF90]. athecate [BBG90]. atidal [AKK94]. Atlantic [BP97b, BAK97, CR95, CAATS94, DSG^{+}99, DLFL96, EY90, EDP^{+}98, FP96a, GR93, Moo94, RWJ99a, TM94, ASSWM97, BK98, BCR92, BI99, BBO^{+}95, BKL99, CA99a, Cam96, CMB^{+}99, CD98, CSSPV98, DHM92, DCBC91, DLHB99, FBHH93, FCW98, FH96, FP96a, FP96b, FQ94, GRC^{+}97, GAK99, GTYR97, GPLB96, GMF98, GA97b, HKMK98, HDPW93, HM94a, HB94a, IBB97, Kep91, KW92a, KW92b, KW99a, KTW^{+}97, LGV^{+}99b, LMM98, Li95, LAK99, LSH96, LK99b, MSD98, MC91a, MH99, MD94b, MF96, MF98, MU97, MHB96, PWG^{+}99, PCM^{+}95, PKY96, PTB^{+}97, PF96, Por94, PB98, RSH92, RM96, RT92, SD93b, SH95b, SBC^{+}98, SF99, SM98b, SFF92, SK95c, TWS^{+}95, TFL95, TRS91, TB95, VTGK99, VK90, WTR99, WT98, WASSM97, WSM91, WT99a, WPTB96]. atlanticus [BK92]. atmosphere [RNHO98]. Atmospheric [SSR^{+}97, AFP99]. Atoll [NAW99, NJLD98, CRCS96, CCR91, Cha96, CDG97, CCRB98, CB98, PPCL^{+}98]. ATP [FCC91]. ATPase [LCTT92, SGR92]. atriceps [WFHM99]. attached [CB96, GSG^{+}94, TM94]. attachment [AI92, BHQ99, HQ99, SHW^{+}91]. attenuation [Ole96, PLP92]. attractant [CR98, Kin95a]. attraction [MVdM^{+}99, Nev99]. attributes [BH97a]. atypical [LNQT93]. Audouin [OJR96, OP99]. audouinii [OJR96]. augeneri [Mar91]. auklet [ASA96]. auklets [HH90, HRCW98]. auratus [MMBT97, OC93b]. Aurelia [BS95a, Han97, HL98, OFR94, Ole95]. Aureococcus [MC97b]. aureolum [BN91b, Han95, VPGV94]. aurita [BS95a, Han97, HL98, OFR94, Ole95]. austral [AMB^{+}96, BGG^{+}97, FPPM96, GB91, GKB^{+}97, HD93, MPS95, PP96, PRW^{+}94, RPM^{+}94, SCKS90]. Australasian [Ste94]. Australia [TA90, AU93, AHS97, BMSF95, BK92, BAM92, BRW95, COK99, Cha95, CGP97, CLD99, CM97, FK95b, FTS94, Gra91b, Gru94, HSH94, HB95, HVL96, Jon93, JBO97, Ken92, KUK91, KKW97, LJ95, LCLC96, LBH91, LPW95, MRP90, O’B94b, OR95, PM91a, RHH97, STBR99, She96, SW93, UDLM99, VS92, VHH^{+}96, VHHS96, VHH^{+}98, WH96, WH93, WC90, YBLL96]. Australian [AW97, CGFC95, HHK^{+}98, JJ99, MW99, MS91, Spe95, WTM^{+}96]. australis [BUCP^{+}92, FK95b, LRV94, PM94a, TB99, Way95]. Austrovenus [HTCP96]. Autecology [BUCP^{+}92]. Author [Ano96a, Ano97a, Ano97b, Ano98, Ano99b]. autofluorescence [JM99]. automata [BH98]. automated [SSY98, SM97b]. Automatic [CES^{+}94, CSE^{+}96, KCV94, RDPB99, Tho91]. autometallographical [SM97b]. autotomy [LL94, Smi95]. Autotrophic [HNC92, BBL^{+}95, HSRHH90, KH92a, KMR98]. autotrophs [CVD99]. auxin [MH93b]. Availability [CCH93, ATH95, AH98, BOC90, BR97, BRZ99, CA99a, Cli95, CRP99, DL94a, DDW90, FCVS92, FF97, GPB^{+}98, Gol99, HJH^{+}99, HRCW98, Iri90, JAOS90, Ker94, KJMRC96, Mat92, MS93c, NBM94, Oro96, PKY96, PB92, PLB95, QGQ97, SCIA99, Sch99, SM99a, SDB92, SDB94, TL91, WYY^{+}98]. Available [BVS^{+}91, DL94b, GAC^{+}97, NT95, ZB92]. avara [RCG99b]. Aveiro [AA92, CSM99, QCM94]. avian [RHrsHG99]. Avicennia [SFV96]. avoidance [CP91, DBMK99, WB91]. axenic [Sv93b]. axes [EL90]. axinellae [Gar99a]. Azores [GMF98, MGF98].

B [PGB96, GR93, BCHL93, BK90, Dun99, GHTH94, GR94, Les96, MM96, MSND^{+}99, NKI^{+}96, PCLÅ97, SDH95, BCG97]. B. [DWG98]. baccularia [SK98a, SK98b]. BACI’ [RSAD98]. Bacillario [DRGV91]. Bacillario-phyceae [DRGV91]. Bacillariophyceae [HFN91, TT98]. back [Rv92]. back-calculation [Rv92]. backed [Oro96]. background [JS97]. backscattering [MWA93]. Bacteria [FD97, GWW^{+}98, HW95, MMG97, ADR94, BNK^{+}97, BSVK93, BF95a, BMTF92, BF99, BL96, BH93, BR91b, CGTB95, CLM^{+}91, CJL91, CCCH94, CA95, CPSA96, CWS^{+}93, CKP98, CB96, DS93a, DG91b, ER95a, ER95b, EAC^{+}97, FB95, GdGMD95, GH94b, GCGS90, GAB90a, GAB90b, GSG^{+}94, GR91, GJ96b, GJ96c, GBH96, HG99, HH94b, HBB^{+}99, HSCC96, HBvD92, HHW99, JP96, JMR91, JSOG91, JKC^{+}93, JG90, JG94, KYI^{+}98, KOSS90, KL90, LAA93, LMK93, LP99, MJV92, Moe96a, Moe96b, Moe97c, MNH92, NKKO97, NH90, PML93, PL99b, PS98, PF92c, PMS94, RT92, SM94, SCB92, SFWD90, SCLS94, SKKO97, Sie93, SJMC93, SJCL93, Sim91, SSA96, SKT97, TT99, Tho91, TA90, TK91, TM94, VFMFBE96, VGJ90, VDM90, VGM94, VDT92, WSM91, YST^{+}97, YKK^{+}98, ZF97, ZNH93]. Bacterial [AA92, BJdJ^{+}98, BB92, BBF94, BG94, FPLH92, HA90, Hei91, LBLA93, MBK95, NK96, PWD^{+}91, RZWY96, SAA90, SPPL92, TGBB97, TB98, Tuo97, VWS93, WF94, ZR92, AKM^{+}99, BHG94, BF99, BDNS99, Cam91, CBD96, CA90, CVD99, CK95, CG93, DAE^{+}99, EPL94, Far98, Flo91, GDP^{+}98, GSS93b, HK93, HDK93, HP98, JMR91, JTB99, Kir90, Kir92, KHW^{+}99, KL90, LAG96, Lig90, LSP99, MBM92, MRL96, MBJ94, MB97b, MVdM^{+}99, MWH91, MRP90, MO95, OH95, PA94, PMG^{+}98, PM91a, PMOD90, RSH93, RBJ90, SC90, SDC99, SD94, Sie93, SJMC93, SCA92, SK94b, TSB93, TS99, VB90, VPFL92, VWS92, VDBF96, WJF94, WP94, YFI91, YKK^{+}98, ZF99, vDK90, vDvRK93, vDdWKW99, vR93a]. bacterio [SCKS90]. bacterio-plankton [SCKS90]. bacteriocinogeny [JP94]. bacteriolysis [PM94b]. bacterioneuston [HM97a]. bacteriophage [Moe96a, Moe96b]. bacteriophages [WP94]. Bacterioplankton [Hei92a, HKK^{+}95, HMVHH95, SWT99, BK91, BFJ^{+}93, CLK90, CLB91, CLB92, CCH93, DFR95, DV92, DLHB99, FD92, JPA^{+}96, JLD^{+}96, KK91, KKW^{+}93, LF91, LKF95, LKK92, MH94b, MPS95, PLP^{+}93, PSSP95, SAH98, SPH94, SH99b, TB93]. bacterium [LD94b, MA93, Moe97a, Moe97b, Moe97c]. bacterium- [MA93]. bacterivorous [HR96b, LHL97, MRMLP97, MVdM^{+}99, PF90]. Bacterivory [Dou93a, HBvD92, CVD99, Dou93b, GSS93a, GQ97a, LDV95, LAG96, MLC92]. baffling [FH97a]. bags [Kru93]. Bahamas [ELMR98, SMW^{+}93, TSM^{+}94]. Bahia [ER90, GGDLE96]. bahianus [Ris97]. Baie [SL97a, SH93]. bait [LSM92]. baited [Wit99]. Baixas [FP91b]. Baja [GGDLE96, GGPGG95, Góm95, Lav94, LGGLLHV98, MZCMOCJP91, RGLD^{+}98]. bakeri [MPM97]. Balaenoptera [KNI95]. balance [ATNW97, BLD94, CM98a, CP90, FW97, GWKW97, HHv91, KSMDB97, LL98, LKC^{+}99, MM90, MH94a, NNR95, PB97, RPDR^{+}98, RMG^{+}98, SR95, SDB94]. balanced [GMK97]. balances [HAH^{+}90, HHKS92, HH91a]. balancing [RS94]. Balanion [JH97]. balanoides [BGBS91, BGW96, HT96, HB94a]. Balanus [ACOK95, GQ97a, LB96b, PGB91, QGQ97, QQ99, WQQ99, WCT95]. Balearic [MSM^{+}98]. balearica [HRD92]. baleen [PZCJ99, PM92b]. Ballast [PCCG97, DZ99, Hal98, RHS99, ZD99]. Ballenas [TBAB91]. balloon [GHED^{+}94]. balloon-like [GHED^{+}94]. Balsfjord [HSN93]. balthica [BC90, CL90, DL91b, DL94a, EB92, Gün91, HV91, HvdM97, HvdMBK98, JT97, JJTS97, LWL98, TBY98, WJM^{+}99, ZB92]. Baltic [GW97, HS97a, JTB99, MHvB97, MHRV^{+}99, Oes90, SKvB98, VS91, AR93, AKK94, AH93, AH94, AH99, BCR92, BS95a, BRL97, BDR^{+}96, BR91a, BH93, CJ95, DGT^{+}93, Ejd98, EE98, FAVV98, GJVG96, GS99a, Hei91, Hei92a, Hei95, HTG96, HSB97, JJTS97, JKT90b, Joh97, JBB90, KLN^{+}90, KKL91, KVK98, KJT90, KV90, KBWC90, KDMR97, Kuo91, Leh96, LA98, Lig90, LL92, LKK92, LHK^{+}93, LL97, MJW96, MEK99, MMR99, MMG97, MHD97, NKV94, NW97c, NK92b, ÓMvdB99, RCS^{+}96, RBSO99, RD97, RHJ92, Sch96, Som96, SE98, Tam95, TAJK90, Tuo97, THKN98, Uit96, UHL^{+}97, Väi90, VNM97, WOM^{+}97]. banana [VHH^{+}96, VHH^{+}98, WH93]. band [KPWL93, KPS92]. banding [LCBD94]. bands [CWB93b, RSN90]. Bank [CBTM93, HF99, NLR^{+}93, RC91b, SHA^{+}95a, CEV97, DRC^{+}97, HK99, KR92, MMGIM90, TRU91, TRS91, TRMW95, TS90c]. Banks [HS91b, PSR^{+}99, HG94, WLB^{+}99, MB91]. Bannock [LC91]. Banyuls [VBCRB97]. Barbados [CK99, HS91c, HH91b, RK96, SC96a, Tom91, TH94b]. Barbara [OC97]. barbatus [Kal91b, LA97b]. barbels [LA97b]. bare [FB91]. Barents [AW98, DS96, MS98, PBv^{+}95, TGvJW95, WVMR90]. barite [FKB92, SFB92]. Barkley [Tan98a, Tan98b]. Barnacle [HJHG98, NC90, BGBS91, BGW96, GQ96, GQ97a, HT96, HB94a, MS93c, MBB99a, NFM98, PGB91, Pin94, QGQ97, QQ99, SBBG94, SBC^{+}98, WQQ99, WB99]. barnacles [Leo99, MS93d, PPO^{+}99, SK99]. barren [Row90]. Barrier [AHS97, BEU99, Bel95, BMSF95, BM91a, CL97, Fab95, FDW92, KM92, LC97b, McC96b, McC94d, MD94c, MKYI98, PD90, PD94, RS91, RSS94, SR90a, SRSH99, Sam97, SB99, SU93, Tho92, TcBW96, UDLM99, VDG95, TZ99, WC90]. barriers [YTF97]. Barrow [HCS90]. basal [GW91]. base [WJD93]. based [ATNW97, BSP^{+}99, BMHD93, Bia92, Big92, BAK97, BZ92, CH97a, CNP95, FW97, FQV98, FKT^{+}99, Har90, HHM96, JBH98, KKW97, LB93, ML93, PM92a, RRWG96, RWCM99, RCC^{+}99, SFWD90, TJST98, Tuo97, Vet95, ZH97]. baseline [BGVF92]. Basin [GW97, PS92, SPZ^{+}93, CSB94, DSW^{+}91, EJ93, MSM^{+}98, NvA91, SD93a, CPT90, DWG98, HBC^{+}91, LC91, LM97, OC97, RZWY96]. Basin-scale [PS92]. basins [SH90a]. basis [Tee99]. basking [SM97a]. bassana [GGF99]. batch [NU91, PRM^{+}96a]. bathyal [CSC92, CM98a, Har97, KW92b]. bathyal-like [Har97]. Bathymetric [FB90, MST98, MD91, RGLD^{+}98, VL97, CSC92, MY98]. Bathymodiolus [DWG98]. Bathynectes [GAV93]. Batillus [KSH97]. bating [WD90]. battagliai [Wil92]. Bay [AKM^{+}99, BL94a, BT93, BLJ92, BUCP^{+}92, BAM92, BHH^{+}95, CCG98, CC92a, CT92, CBTM93, CNGG96, CW93, CBL96a, CBL96b, CM97, CPB96, DSV^{+}91, Dur94, FK95b, FKC96, FMH96, FMFH97, FF97, FII95, GB95, GBB^{+}98, HSK^{+}94, HC94, IT96, KS98, KYI^{+}98, KSO^{+}97, LGV^{+}99a, LGV^{+}99b, LBJ^{+}98, LD96, MC96a, MW97, MLTD96, NAI^{+}93a, Nak92, NTK93, NCGG96, NKKO97, PRC^{+}92, PLB95, PT94a, PT94b, Pur92, RR97b, Rie98, RSN^{+}99, RBC^{+}91, RLT^{+}94, RE94a, RE94b, SOW97, SOSV95, SD94, SW93, Sie93, SJMC93, SSP97, SPS98, SH92b, TK95a, TK97, TK98, TDD94, TB93, VBCRB97, Vid94, VM95, VS91, WA99a, WB92a, WW90, WC92, WS92b, ZCL97, ZB94, dG92a, AV95, AI92, AMBH98, CC91a, CH90, CRH^{+}99, FMRC93, HPPZ93, HGS94, JWKH98, JAOS90, MH96a, NH95a, PHT96]. bay [PS92, PSL96, PBGE91, PND^{+}94, RCC^{+}99, SLHA97, TT99, TCB93, TCRB93, UC94, WH95, WA95, ZHM91, ASCW90, ADA98, AD94, BL95, CL90, CTSN90, COK99, CH93, CRH^{+}99, DCK99, DLT^{+}97, DQT^{+}97, DFPG91, FPAH92, FPG95, GGB^{+}99, GFPD92, GCF^{+}95, GSG^{+}94, HFN91, Har94, HP97, HR96b, Hil91c, HPF92, HHW90, JCP93, JMP90, JE95, KSRB99, KD97, KSG^{+}92, KSMDB97, KGO94, LJM99, LOS90, LL98, LB93, MH96a, MDPP91, MDD95, MHBW90, MDC98, ML92, MBK95, MHM90, MNO96, NH95a, NTS90, PND^{+}94, RRGB96, RT93, RCC^{+}99, RSK^{+}99, RAGH94, RHP93, SK94a, SKvB98, SJCL93, SMTL99, SK95b, TD92, TDW96, WH95, WHIC97, WR92a, WR92b, ZHM91, dG91a]. bays [LCK^{+}96]. be [Ano99d, BMH91, EP92, LCC97, LF93, MLP90, TK95b, THCD95]. Beach [CV96, ACR99, BLJ92, BD99, Con94, DL93a, DdA95, GAR93, GRBA96, Har98, McL96]. beaches [GY97, GD99, HGB99, JMC93, JMD95, MM91b, San94b]. beam [KRM94, KS95a, LWK^{+}95]. bearded [Kal91b]. Beaufort [FB98, HC90, SBV98]. become [KCI98]. bed [CJ93b, DBB^{+}95, FS91, HS99b, LLB92, MN92, NHL99, PSPD96, RPDR^{+}98, RSK^{+}99, Row90, SR90b, SM91, SMH97, VWS93, VAD^{+}95, WB94, WP92]. Bedford [SPZ^{+}93]. beds [BBFD98, BH97a, BS93, EM93, FKC96, FMH96, GNJ90, GF94, GI99, HSK^{+}94, Hil97, HN97b, IVM^{+}98, JDSB99, JCH93, KG99, MSD92, MF95, PM91a, SH93, Tok95, TF98, Wor95]. behavior [Bus98, DM91, DP91, DPO91, DO92, Dav94, DEF91, DDW90, EPDC99, Gar99b, Har99, KCS94, MSD98, MY96, MGYV97, MSK95, PM92b, RO99, SCC94, SHFP94, SK98c, SC96b, Van94b, WX97, WRBF97, WDD92].

Behavioral [Bus97, SML97, WFH99, VB93]. behaviors [OF94]. Behaviour [LK99b, AS99a, Aks96, AJL93, And98, APS91, Bur94, CMRF94, CPC98, CJ93a, DB99, DBvA91, FRK92, GH94c, Gon92, GKB^{+}97, HlS90, HSTY98, Hay95, HWL96, HLP^{+}99a, HVL96, HB98, JT90, JAL91, KS94, KSV96, LJSK96, MFD98, MC91b, Moo94, Mun92, PCM^{+}95, RD92, SA92, SAA99, TCRB99, TK95b, TJ90b, UMTM98, VV95, VBCRB97, WFHM99, WH93, Wol93, van94a, vdL95]. Behavioural [BSB91, Ejd98, MF91, MK93a, AAD91, AH98]. belcheri [HSTY98]. Belgian [BLDF95b, BLDF95a]. Belgian-Dutch [BLDF95a]. Belize [Mum99, WC90]. bell [HdWS91]. below [DAS92]. belowground [DMA^{+}98]. Belt [NW97c]. beneath [AMMG91, KMKS96, KTH98, PDP^{+}90]. benefit [MC96b]. benefits [BA94, FGH95]. Benguela [SDFJv91, BGC91, DG96, Gib94, GGB^{+}99, MG96b, PG91, PLP^{+}93, PML93, PBHMI98]. bennettae [CM97]. Benthic [CC92a, CBTM93, CH94, EJ93, GHHC98, GGH99, Hei92b, JLS90, JC97, KHTP99, KHB91, MDC98, vDvRK93, Alo90, AB95, AJL93, AB96a, AB92, AD94, BYO95, BJdJ^{+}98, BCAAC96, BB92, BDR^{+}96, Bra95, BM95b, BDdWC96, BLR94, CCC90a, Cam91, CTEF91, CSR^{+}92a, Cha98, CGA95, CR94, CCG^{+}92, CEV97, CGZ95, CRGZ98, CLKO98, CNP95, DL91a, DL93a, DW97, DSV^{+}91, DL94b, DD99, DLHB99, EJ96, FTW90, FW97, FDZ93, FGP98, Fra95, FCP98, FL90, FAP92, Gar98, GH94b, GFH96, GRGK96, GF99, GD96, GQ97b, GPU^{+}97, GH92, GBGH97, GB98, GI99, GSD96, HA90, HDPW93, HGS94, Hil91c, HAR95, Hol97, HG92b, HS97c, IW91, IT96, JDSB99, Jos90, JJNR95, KSS90, KSG^{+}92, KDAS92, Kro91, KCI98, LD93, LHL97, LJM99, Leh96, LA98, LC93, LGR95, MC96a, MG96b]. benthic [MM91a, MG91a, MSK95, McL96, MMR99, MWT96, MG94, MBB95, MHMH96, MvdZH^{+}97, MSvdZH98, NL94, NG96, NR94, NNB93, OE91, ÓMvdB99, OG95, Pec99, PGB96, PM90b, PBv^{+}95, PZ91, Pir92, PBGE91, PH95, PLP92, Raw96, RCG99a, RB98, Rie98, RSN^{+}99, RTLG97, ROI93, RHJ91, RCN95, SB98, Sch90, SNC^{+}98, SS99a, SHO96, TWWE95, TRU91, THCD95, TGSN99, TD94, UNPH93, UMT^{+}91, Ved98, VM95, WB94, WHG^{+}99a, WHG99b]. benthic-pelagic [Bra95, HAR95, MWT96]. benthic/pelagic [CR94]. benthopelagic [KZ92, MG91a, Sal96]. benthos [DSV^{+}91, FNJ^{+}94, FC90, FH99, FCH99, GCWH90, GP91, LGH95b, LTM^{+}98, MC95, MS94b, RSD^{+}91, SLB^{+}92, Vet95]. benthos-environmental [FNJ^{+}94]. Bering [CHDW92, HCH^{+}96, Bro98, CPS91, CH94, GH92, HHP95, Liv93, RHrsHG99, SBV98, SSA96]. Bermuda [McG92, SC94]. Beryl [MHC92]. betaine [KG95]. better [DGCT93]. Between [IBB97, ASN93, AC92, AWH93, AI92, AMB99, AS94a, AM91, BLKA93, BCAAC96, BCL93, Bia91, Bro90, Bro98, Cal95a, Cal95b, CGNH98, Cha98, CW99b, CBT98, Dai90, DD90, DBV98, DFB96, DCBC91, Dor90, DGHT^{+}98, DLHB99, Dut98, EM93, FGH95, FL90, FH97b, Gil97, GTYR97, GBP^{+}97, GK96b, Gru95, GP98, GSMM96, HSRHH90, Hol90, Ito97, JBB98, JWL96, JKT90b, JBHH97, JC95, KJT90, KYI^{+}98, Kog95, KRJ98, KLAH^{+}94, LFRB97, LGP95, LHL97, LC95, LC97a, ML91a, MB95a, McG94, Mic96, MG94, MBB95, MGF98, MGAE99, MJ97b, NDRP93, OSC97, OSC98, ONR98, PWS98, PB97, PS94, Phi94, PWPK95, RLDM94, RBF92, RTCH96, Rv92, RZLB96, RT98, SM98a, SRH^{+}98, Sal95, SCB92, SFWD90, SFV96, Sci91, Sha95b, SDFJv91, SRRdM99, SRW95a, Ste95, SEGP91]. between [TDBH94, TTW97, TA95, TAJK90, TC93, TPHC94, Tro93, TFFS90, Uib95, UT90, VFMFBE96, Väi90, WA99a, WHG^{+}99a, YFI91, ZF97, vDdWKW99, vRBVB95]. Between-habitat [IBB97, Mic96]. ‘Beyond [RSAD98]. bias [BVV^{+}94, Fre97, FP91b]. biases [Cer98]. bicinctus [MF97]. bicoloratus [MYYT96]. bifasciatum [HH91b, MSP^{+}97]. bifilosa [ICG96, UCV98]. Bight [EDP^{+}98, JLS90, KZK^{+}95, LJS90, MHvB97, RG93, AO91, BP97b, BDD92, BS95a, BKS92, BVL92, CSR^{+}92a, CA95, EGKE92, FC90, GAD^{+}92, HC92a, HFJS91, HHH92, Kar92, KR92, LPF94, LLP94, MC91c, McF92, MMG97, Mol97, PGAJ99, RBD^{+}97, SSR^{+}97, SPBS93, SKM^{+}97, SGR92, Sün97, Tha92a, Tha92b, VBL^{+}92, Vos91, Wil92, vdVW99]. Bilbao [GD96]. binding [DCG^{+}99, WNJ91]. Bio [Sos96, SPS^{+}96, Tha92b, UMT^{+}91]. bio-activity [UMT^{+}91]. bio-assay [Tha92b]. Bio-optical [Sos96, SPS^{+}96]. bioaccumulation [MCSV95, MCFS96, RQB98b, WFL96, WTF95]. bioassay [CLM^{+}93, Jon97b, MC94a]. Bioassaying [RBG94a]. bioassays [BHQ99, HJA99, TvLdJ^{+}98]. Bioavailability [BK94, LP99, VB96, WL97, WSF99, Bou90a, Bou90b, CM99, Mal96, NFT92, SM97b, TBY98]. Bioavailable [SKM95, NMR97]. Biochemical [AOOT94, Add92, KMR98, AJ95, BKWML93, CD96, DDM^{+}99, FSdM^{+}92, GAC^{+}97, GRSV98, MGYV97, NT95, SGB90]. biodeposition [JBB92, RCG94]. ’biodiversity [WC95, CW99b, SEAI96]. bioeconomic [Hal98]. bioenergetic [DDM^{+}99, DL93a, Leh96]. bioenergetics [AH94, LB93, VR93b, VL97]. Bioerosion [BvKvRB96, KML91, PPCL^{+}98, Bak90, Bel95, SR90a]. biofilms [BB92, MR94, NW97a, PL99b]. biofouling [AJ95, MS95, YOH98]. Biogenic [WLL97, AO91, AW98, LK94, LR96, LM98b, SPPL92, ZHF96]. Biogeochemical [CA90, PSE97, SGC^{+}98, TM94, ASRP^{+}96b, CSS93, FNF91, HM97c, MH97, RDT^{+}94, RC95, SCA92]. Biogeochemistry [CJ95, GGD99]. biogeographical [tor93]. biogeography [GDMS90, PT93]. Biohydro [BSP98]. Biohydro-optical [BSP98]. bioindicator [Bou90a, Bou90b]. Bioindicators [BBP^{+}96, NCGG96]. Biokinetics [WTF96]. Biological [GS91a, GPF95, LGL^{+}95, LGMD98, SRJD97, SJ99, SWEC91, SD92, AWVP98, BK90, BP96a, BBT^{+}93, COD^{+}93, CAATS94, DvdM92, DM91, DB96b, HS96, Kep94, MC91a, OE91, Röp93, Spe95, TR95a, UKR^{+}91, VSQ90, ZCL97, AJ95, BSW^{+}95, BWHW95, Rie91b]. Biological/physical [LGMD98]. biologically [WL97]. Biology [Aro92, JT93, Ang95a, Ang95b, BGT91, CV96, GTYR97, McC94b, MT94, NH95a, NKH^{+}99, NAS95, Per90, Pud93, RVv92, SCAT91, Tan98a, Tan98b, US95, VTGK99, WMG^{+}94, SDA^{+}92b]. Bioluminescence [HLBW93, KTW^{+}97, LJ96]. bioluminescent [PTB^{+}97]. Biomagnification [BMSHC98, RH90b]. biomarker [Bro93, KWHM^{+}96, KBA^{+}97, VBZ^{+}94]. biomarkers [BMS^{+}97, BMP^{+}95, BRL97, JL96, WKH^{+}98]. Biomass [FQ94, GMMD93, KFN99, MG96b, OSJ94a, SSR99, SMH97, WV95, YBLL96, AR93, ALB^{+}96, ADG^{+}95, BH92b, BK91, Bra93, BRW95, CCV^{+}96, Cha96, CB98, CJ93a, CA90, CWH98, DPM96, DSJ90a, DMA^{+}98, DD99, FFF^{+}94, Fra92, GJ97a, GJ97b, GBB^{+}98, HP97, HKR96, Hei91, HHSvS96, HSRHH90, HK93, HRD92, HN97b, HNC92, IP94, JL96, JP93a, Jos90, KKL91, KL95, KM92, KKW97, KA93, LHRL97, LDV95, LZYW93, Lig90, LPW95, MP96c, MPS95, Onu96, PS97, PWM95, PR93b, RGLD^{+}98, RH98, RB99, RZWY96, RPM^{+}94, RSD^{+}91, RA96, She94, SNJV92, SCA92, Spr93, SPPL92, SNR^{+}96, Tay98, Tho91, TL91, TS99, US98, VWS92, WH95, ZH97, ZBK^{+}98, dG92a, dC94, vDK90]. Biomass-density [OSJ94a, HN97b]. biomechanics [FSGG96]. biomineralization [LGP95]. biomonitoring [BQR99, MGM96, Reg92]. biooptical [ARS93]. bioroughness [HG92a]. biosensor [NMR97]. biosentinel [BL95]. Biosynthesis [SGK^{+}97, FBMH94, KHBL95]. biota [BF95a, GC93, ZF97]. Biotechnological [AJ95]. biotic [BH97a, FTW90, GRSV98, KW98, MS96a, San93b]. Biotransformation [VT92, RMGB^{+}94]. bioturbating [RCN95]. Bioturbation [TF98, AWdS92, BSG^{+}97, BD98, DHH98, FGKP95, GSB98, HFW99, KF97, RBA98, WA99b]. biperforata [AC96b]. biphenyls [GS99b, WKH^{+}98]. bird [GWWP99, VPM96, GBP^{+}97, HHE92]. birds [Hil97, Liv93, RAB^{+}92, WFHM99]. Birth [Sew94]. bisphos [Riv90]. bisphosphate [PFP93, PCK^{+}96, PCC^{+}97]. Bivalve [Jør96, AJL93, BTKK99, BL97, BVP99, BC90, BKWML93, BL95, CL90, Cra94, CH99, DBN^{+}94, DRS94, EK95, EB92, Har99, HSTY98, HCF93, HvdMBK98, HLP99b, JAL91, KSRB99, NT95, RLDM94, ROM^{+}91, Reu98b, Rin92, RBG94a, RBG94b, SDB92, Urb94, Vis93, WX97, YBG96, YBG^{+}98]. bivalves [AR91, BDB97, BR91b, Coc90, CD98, DL91b, DL94a, EE98, FSCE92, HTCP96, Kam94a, KOSS90, LWL98, LAJ95, MCL90, MLL98, PTW^{+}91, THPC96, UC94]. Bivalvia [BSJPW93, BD96, BSJ97, BH92e, CJL91, GM99, JT97, JJTS97, MW94, Nak94, Oes90, PRC^{+}92, PCBA94, RR92, RM95, SDB94, WJM^{+}99, YKL98a]. Bjornafjorden [GGB94]. Black [FGG97, KCI98, AAE^{+}96, CW95, Mil97, Mil98, Oro96, RM97a, SM97b, WD99, KFS91, Mik95, WHG99b, YSY^{+}99]. black-backed [Oro96]. black-browed [CW95]. black-legged [RM97a]. blade [SDB96]. blades [OB93]. ’bleaching [HHYD92, KRFL97, Mum99, ABB^{+}97, HGS95, Hoe91, Jon97a, JBO97, JHG99, KRF^{+}98, LB96a, MB93]. blennies [BKH99, Har99]. blennioid [Sym95]. blennoid [MZ99]. Block [CC97b]. blood [BPR90, DvCd94, Kal91b, LFD95, MM96]. bloom [ARS93, ADG^{+}95, ASP^{+}96, BH92a, BN91b, BGvDR96, BvBBV96, BD94, CCG98, CSM^{+}98, CMB^{+}99, CCCH94, CJ95, FBHH93, FS90, GJC97, GCF^{+}95, Gra96a, HvB91, IHK^{+}98, JLS90, JSGS94, KLMR91, Käh98, KLN^{+}90, KNM93, KJN97, KR98b, LHK^{+}93, MBJ94, MG97, Mik95, NKH^{+}99, NKB90, NH95b, PWD^{+}91, Rie91a, Rie91b, SCEG93, SD93b, SDG95, SPZ^{+}93, Tam95, TGSN99, VKV95, WYY^{+}99, YHGB96, HKK^{+}95]. bloom-forming [JSGS94]. blooming [HRKR93]. blooms [BEH93, BLM^{+}95, BHH^{+}95, EST95, GCGR90, Hin92, KLR93, KYI^{+}98, LKC^{+}99, LKM^{+}99, NT96a, NGS97, Ols93, PBHMI98, Rie93, SFS95, TNKMK99, VMS^{+}96, WMG91, WPTB96]. Blue [RCG94, ABY97, BA90, BM95b, CWWH99, DFS94, EK95, ELH92, FW91a, FW92, FFR94, FDR^{+}96, FTBR97, Hil97, HK99, HHW90, HH91b, JM99, JSH99, MCF90, ML91b, MUH97, ML92, MZFHC96, Olm94, PODL99, PLBP96, Reu98a, RvMM97, SW93, Smi95, WFHM99, WRBF97, WFH99, ZFFH^{+}94]. blue-eyed [WFHM99]. blue-head [HH91b]. blue-light [JM99]. bluefin [DJY90, DLJ91, JD90, KJGC96, YD90, YBLL96]. bluefish [HC93, JC94, JC95, MC91c]. bluehead [MSP^{+}97]. boat [LOS90]. Bodden [KDMR97]. body [AS94b, FCP98, GPP98, HS97b, HL98, HvdM97, ISF^{+}98, JJTS97, Ker94, Leh96, MD93, Mau92, MM95a, MAAT98, Pen90a, PS95, PCBA94, Sal95, Spr93, WF97, YKL98b]. body-size [HS97b]. Bolinao [ADF96, BBFD98]. bomb [KJGC96]. bone [SDFJv91]. bones [Bar95, BSGM94]. bony [HSLV94]. booming [BLJ^{+}96]. Boonea [GMC91]. Bopyridae [OR95]. Bor [KKH^{+}94]. Bora [KSO^{+}97]. bordering [SR97]. boreal [PPW96, SP99a, Spr93]. borealis [Ber92, HSN93, Oes90, OTF95, RT95]. Boreogadus [GCKW95, SCJ99a]. Boreomysis [Sal96]. boring [BKWML93, MLL98]. borne [SHM^{+}97]. Bornholm [GW97]. bosc [Har99]. bosquianus [Har99]. Boston [Kel98]. both [Han91a]. Bothnia [Hei92a]. bottle [GKRD96]. bottleneck [WS91, ZC92]. bottles [CDGF^{+}93]. bottom [Amb93, Blo91, BDR^{+}96, BMG99, CSB94, CEV97, FSCE92, FAVV98, Gas94, HPF92, HFJS91, JDSB99, JAL91, KD99, KSG^{+}92, KO92, LOS90, LA97b, MS96b, MRS96, MJ97b, NT96c, Ols93, ONR98, PPFP99, PBGE91, PSR^{+}99, RH91, RPL96, SGC^{+}98, SE91, TNSM98, TN98, WFHM99]. bottom-mounted [CSB94]. bottom-up [FAVV98, Gas94, MS96b]. bottom-water [BMG99, PPFP99]. Bottomless [Roz92]. bottoms [Mic96]. boulder [Tak95, Tak96, Tak99]. boulders [SHA^{+}95a]. bound [CSG93, KF97, LAA93, LP99, MFF97, WSF99]. boundary [Arr93, BYO95, CK99, FHZ96, JAL91, LGR95, MG91a, RTLG97, SS99a, TB98, Ved98, VM95, BL94a]. boundary-layer [JAL91]. Bouvet [PLM92]. box [RM94]. brachiatus [GM98b]. Brachidontes [LTM90]. Brachyura [AAWD97, Per90, QCM94, SMUS^{+}99]. Brachyura. [SCAT91]. brachyuran [BW93, Gar99b, JE95, LE90, RS99b]. brackish [AJS95, AA92, BTKK99, CMD^{+}94, CSM99, FMRC93, HS94a, Mal99, MMG97, NK93, SE93, VDA96]. brackish-water [Mal99]. branch [Tro93]. branches [RT98]. branchial [SGR92]. branching [MB95b]. Branchiostoma [Sto96]. Bransfield [KL95, Nor92]. brant [BL94a]. brasiliensis [CV96, GD99]. Brazil [NAW99, dG92a, CV96, CF93b, HSLV94, NL94, Pir92, San94b, dG91a]. breadth [SH99a]. break [MLDM95, MLDM99, SH95b]. breakdown [CWB93a, KFF92]. Breaking [LSGW98]. breaks [MBL^{+}98]. Breeding [KNI95, AAS^{+}91, Ang95a, Ang95b, Cer98, GSC95, HHM97, JCCCB94, LGC^{+}93, LK99b, Oro96, OJR96, RM97a, SMUS^{+}99, SM99a, TMCE98, VF99, WTA^{+}97, WC98, WSP^{+}95]. Bremerhaven [SD92, SDA^{+}92d, SDA^{+}92e]. bresiliid [GPS93]. Brest [DLT^{+}97, DQT^{+}97, HPF92, LJM99, Hil91c]. breve [KMR98, MKV95]. brevicarpalis [FGH95]. brevicula [SH98a]. brevior [DWG98]. Brevoortia [CMBC98, FCW98, MC91a]. Briareum [HFR^{+}93, WHW93]. brightwelli [WH92]. brightwellii [KRJ98]. brine [BFD91, BKB92, KCI98, SBP92, SBP93, ZGB90]. brine-filled [KCI98]. Bristol [RBN99]. Britain [HBBT98]. British [BL94a, HCC^{+}91, HM97b, Mga92, SB96, HTS92, PG90, Sew96, SR98b, SS92, THS94]. brittle [GS99b, MDG98, SG96]. Broad [PMMH95, KPS92]. broad-band [KPS92]. broadly [BA96]. bromophenol [LSP99, SNR^{+}96]. bromophenolosus [GK97]. bromophenols [SPPL92, WLL97]. brood [PF95, SB97b]. brood-protecting [PF95]. brooding [AHS97, CT98, HHG99, MCNB98, Sak97, Sew94]. broth [Moe97b]. browed [CW95]. Browman [BS96a]. ’brown [BH95b, MC97b, Åbe96, ÅP97a, ATH95, CLHF95, DFS94, EP97, HN90a, HN90b, HHN90, KMG98, LB95, LCK^{+}96, O’B94a, Onu96, PCLÅ97, PKY96, Poo94, Ste94, SP99b, SMP97, TC94, dT94a, vTSD97]. bruciei [AWHH94]. Brunswick [PRC^{+}92]. Bryozoa [RM97b, RG97]. Bryozoan [MC96b, HHG99, HH95]. Bryozoan-macroalgal [MC96b]. bryozoans [DC91, Har97, NT94, NR98]. BSD [SM97b]. bubbler [Kog95]. Bubbling [JAB^{+}92, Mar99b]. Buccinum [IWF^{+}97, LSM92, RMH96]. budding [KWL96]. budget [BHT^{+}92, CM93, Far98, FZ91, KW99a, LL98, MR97b, ONH^{+}97, PB93]. budgets [ASRP^{+}96a, Col92, HdWS91, LCW99, MS97, Pre91, RSB^{+}90, SV97, YKL98a, YKL98b]. building [Byt90, FBR^{+}95, HGJ99, LNSL91, LQ97, MNB94, MPB97, VB93, Van94b, VK94, VB94]. Buildup [ZBK^{+}98]. Buoyancy [MV96, NKV94, RC95]. buoyant [Con94, DeV92, MV96]. burdens [MGM96]. burial [BA90, BHHL96, DTA^{+}97, MD94a, WUA^{+}99]. buried [FTP99, Ter97]. Burkholder [PPFP99]. Burrow [SSV97, GS91b, IBB97, MSK95, NA95]. Burrowing [Wol92a, BI99, GSP99, IBB97, TK95a, TK97, TK98]. burrows [Dit96, MK93a, MDRW99, NA95, SNR^{+}96, War90]. bursa [GHED^{+}94, VAD^{+}94, VD95]. butter [KB91]. butyl [DBHM99]. butyltin [MPH96]. Buzzards [BT93, PT94a, PT94b, TB93]. by-catch [KS95a]. bycatch [PSR^{+}99]. Bythograea [EPDC99].

C [BHT^{+}92, Fry96, Gar91, JR92, RTRF90, RHT91, RAB^{+}92, SDFJv91, TR95a, AMB^{+}96, BCG97, BR97, CCV^{+}96, CGA95, CRLP94, FMFH97, FPFR92, FH97a, FW91b, Fry96, Gar91, GT98, HP90, HWLC98, HHSvS96, HW92, HAR95, Irw91, JRMN^{+}97, JR92, Joh96, JLGMP97, Kal91a, LBM^{+}99, LH91, MvdVD^{+}97, Mar99b, MKCL94, MJSDF91, NH90, ORL97, RHT91, RRWG96, RWJ^{+}95, Ric91, RR97b, RSS94, SJMC93, SV97, TFL95, VBN95, RTRF90, RAB^{+}92, SDFJv91]. C-8 [BCG97]. C-CO [Gar91]. C-glucose [Gar91]. C-labeled [LBM^{+}99]. C-protein-labelled [NH90]. C-pumps [TR95a]. C-rich [FW91b]. C-tracer [HWLC98]. C-transfer [BHT^{+}92]. C. [BM99b, BV99, DS93b, FP96a, LAJ95, MS98, PF96]. C/ [Fry96, JR92, RTRF90, RHT91, RAB^{+}92, SDFJv91, GT98]. C1 [Sie93, SJCL93]. Ca [EL90]. Ca/Mg [EL90]. Cabo [CF93b]. CaCO [PS93, PPP^{+}96]. Cacospongia [BPS98]. Cadiz [DSG^{+}99, AD94]. Cadmium [TP96, WLL98, WNJ91, BKB92, CBRD92, LM95, Mal96, MIA90, MAAT98, MBL^{+}98, RK95a, RBA98, Rin92, SDC99, SFF98, SE91, TLW^{+}96, WL97, WLT^{+}95, WTF95]. Cadmium-binding [WNJ91]. caespitosa [BH97b, PMB99]. cage [HAH^{+}90, HH91a, HK92, LW91]. cages [KTH98, LCW99]. Caicos [SB96]. Calanoid [Nag96, SC96b, BS95c, BP96b, HvB91, HSW99, Jeo94b, KVK98, Kle93, KEV99, LGED98, MSS95, MHvB97, MHRV^{+}99, SW98b, SK97, TJ90b, Tis98b, VV95, WR92a]. Calanoida [ASSWM97, CA96a, CA97, KGS95, KBS98, RD92, SA92, SSHM91, SM98b, WASSM97]. Calanoides [ASSWM97, HKG93, LHL93, PPB90, SSHM91]. calanoids [RD91]. Calanus [AS99a, BM99b, BM95a, CH97a, DBMK99, DWSA98, DRC^{+}97, EM99, FP96a, GM98b, HKG93, HL90, Hay95, HH92, HH96, HB98, IHK^{+}98, LPI^{+}95, LAA93, Lop91, LGMD98, MS98, MCW^{+}91, MT93, NGS97, NKH^{+}99, OF94, OC97, PF96, PR93a, PHHH96, PLIM95, RCTH99, SH90a, SSHM91, ST90, SH98b, SM98b, ST99, UMD92, Uye96, WDB98, WSCT96, YSY^{+}99]. Calappidae [Per90]. calcar [Hun93]. calcareous [HHB97, Lew91]. calcification [BFF96, CCRB98, FBR^{+}95, MG97, MA99b]. calcifying [GPF95]. calcite [VKV95]. calcitrans [KRJ98]. Calcium [CBA^{+}99, MDG98, CMBC98, OINN94, ROM^{+}91, TRMW95, van92a]. calculated [KPS92]. calculating [Kar95, MS97]. calculation [RC91b, Rv92]. Caledonia [CG99a]. calendar [MMBT97]. Calibration [KGHTN99]. Calidris [MJ92a, ZBW92, ZB92]. California [CCG98, CA95, FC90, FMFH97, Gen92, GGPGG95, GGDLE96, Góm95, LPF94, LLP94, MZCMOCJP91, PGAJ99, SPBS93, Sha95b, SSP97, SPS98, ASN95, ASA96, AAE^{+}96, Bro94, BUCP^{+}92, CTSN90, DSV^{+}91, ESDK94, EJ93, HBC^{+}91, KL91b, LK93, LRV94, Lav94, LGGLLHV98, NPB93, NTS90, Pin94, RGLD^{+}98, Reu98b, TD91, VPM96, Ven98, VMHH90, Vet96a, Vet96b, WH91, WBLM95]. californianus [DM96]. californica [HRFC90, Laf93]. californicum [Rin92]. californiensis [FAED97]. Callianassa [SSV97]. Callichirus [SF99]. callidegenita [GC95a]. Callinectes [BA90, BM95b, CWWH99, DS93b, EK95, FW91a, FFR94, FDR^{+}96, FTBR97, JSH99, MZFHC96, Olm94, Sha98, WRBF97, WFH99, WD94]. Callionymus [BBT^{+}93]. Calliopius [DL91a, DL93a]. calm [CPC98, KS95c]. Calonectris [GMF98]. Calvi [HRD92]. Calyptraeidae [CP90]. camera [GF94, Tis98a, Wit99]. camouflage [HM91]. Campana [Ano99a]. Campanularia [CGZ95]. camtschaticus [SK98c]. Can [Ano99d, DBMK99, DA93, GQ97a, LCC97, BMH91, DC91, Han91a, LF93, MLP90, PZCJ99, THCD95]. Canada [BH94b, CLGM96, CHS^{+}95, HCS90, HM97b, HS95, JNCS89, LD93, MCF90, MBJ94, MHBW90, MRP98, MHS94, MS94b, MLTD96, PRC^{+}92, RLT^{+}94, SL97a, Sew96, SR98b, SMTL99, SCEG93, SPZ^{+}93, SK95c, TMKL97, Tan98a, Tan98b]. Canadian [LDV95, CLKO98, SC92b, SFF92]. canal [SLHA97, QCM94, TBAB91]. canaliculus [HJH^{+}99]. Cancer [CW93, CC97b, FIA93, Fer99, HBR^{+}91, PWS98, SLSH98, WC92, WS92b]. candlelight [WPB^{+}93]. Cannibalism [BBK91, Fer99, RvMM97, Val91, Ang95b, LJ96, MPvM98, Smi95]. canopies [COG97, DBdCS99, Leo99, VSPLH98]. canopy [BCC92, Con95, JHN99, Ken94a, PBGE91, Ste95, VÅCL99]. Canuella [BDdWC96, CM90b]. canutus [ZBW92, ZB92]. canyons [VD99]. Cap [And90]. capacity [DFS94, FL92, GBGH97, LCLC96, MPB97, MSND^{+}99, PR92, RM94, STS99]. Cape [AAS^{+}91, ADF96, BBFD98, BYO95]. capelin [DL91a, DCBC91, SM99a]. capensis [AAS^{+}91]. capitata [GVGG98]. Capitella [FFH96, GVGG98, HFW99, MFF97, MGDT90, QC94, SFF98, Tsu90, TFFS90]. capitellid [AGP98, LGFS99]. Capitellidae [Tsu90]. capricorni [FB91]. capsules [Raw96]. captive [STBR99]. capture [BLH99, CGZR94, CLW99, JT90, JC94, RM97b, WBGS97]. carbo [GSC95, BRLM90]. carbohydrate [BDM96, FRP^{+}99, vDdWKW99]. carbohydrates [GAB90a, GAB90b, SR98a, WP95]. carbolxylation [DGO92]. Carbon [CCV^{+}96, Dou93b, Fra95, FGB^{+}96, GPD^{+}96, GKRD96, HAH^{+}90, HSK^{+}94, JKFB95, Ken94b, LL98, Mar99b, NBM94, RSB^{+}90, RÁSP99, SBV98, Tur93, AC96a, And96, AR93, ABY97, AMB^{+}96, BRL97, BD98, BCG94, BP96a, BvBBV96, BLR94, BRZ99, CD94, CCH93, DSG^{+}99, DS93a, DL91b, DGO92, DA93, DWG98, DH92, EJ97, FSdM^{+}92, FBHH93, FPGC^{+}98, FPLH92, FCMB94, FCVS92, FZ91, FRP^{+}99, FW91b, GV96, GYL93, GMK97, Gol99, GWKW97, GKMT99, HHB93, HHSvS96, HM96, HCG^{+}93, HSRHH90, HSW99, Hol96, HG94, IG96, JCP93, JLD^{+}96, JMP90, JKC^{+}93, Käh98, KSRB99, KM99a, KSMDB97, KW92a, KJN97, KW99a, KRJL96, KRJ98, KTGC99, KKH^{+}94, KR94, KR98b, KLAH^{+}94, LD96, LR96, LM98b, Lig90, LHK^{+}93, MM90, MRL96, MDPP91, MR97b, MBK95, MDG98, MO97, MJSDF91, MH90, MO95, NBW94]. carbon [NMR97, NH95b, PS97, PGAJ99, PM91a, PSB^{+}97, Pre91, RDPB99, RN98, RSN^{+}99, RSD^{+}91, RTRP^{+}98, SO91, SWT99, SH97c, SRG^{+}99, SAA90, Sim91, SM95, SPS98, TA93, TA95, TGG97, TL95, TT98, Tuo97, URHR98, VGJ90, WF90, YST^{+}97, ZB99a, ZB94, ZA96, ZNH93, NNB93]. Carbon-13 [Fra95]. carbon-to-chlorophyll [GV96]. carbon-to-chlorophyll-a [TGG97]. carbon-to-nitrogen [Käh98, KJN97]. Carbonate [Bel95, KSO^{+}97, DS93a, EF97, EM93, JAB^{+}92, MDG98, SDC90]. carbonate-cemented [JAB^{+}92]. carboxylase [DGO92, PFP93, PCK^{+}96, PCC^{+}97]. carcasses [GGH95, HFG^{+}95, Wit99]. carcharhinid [STBR99]. Carcharhinus [WCL97]. Carcinus [AWD95, AAWD97, BC93, BKH99, CBRD92, DL95, HMHDA94, HN93, LS92, MPvM98, QCM94, QCM97, RN93, SAA99, WJD93, ZN96a, ZN96b, ZAN99]. card [JM93]. cardinal [FGG97]. Caretta [MBB99b]. caribaeorum [FP91a]. Caribbean [Aer98, Ano99d, AADSCC97, Bee97, BKB94, BvKvRB96, CP95a, Cli95, ELMCC90, ELMR98, EL90, FP91a, GMMD93, GMR^{+}94, GDvS98, GBJ94, GT98, HOM95, LBL97, Maz95, MD94b, MLES94, NBS^{+}97, NB98, OMH95a, OMH95b, PF92a, PCTF95, PR93b, RSH93, RS99b, RH90a, RSB93, RH90c, SC90, SR94, Szm91, Tol98, Tom91, VB93, Van94b, VK94, VB94, VPF98, VDT92, WC90]. caridean [SB96]. carinatus [PPB90]. cariosus [NFM98]. Carlo [BMHH94]. carnivores [Gib94, PWR96]. carnivorous [MSS95, Øre91, ØW93, VFMFBE96]. carnivory [PM94a]. Carolina [BYO95, BCS93, BW93, CC92a, CT92, DSW^{+}91, FB98, MP94, O’C93a, PPFP98, PLFSW97, SCCS98, TGG^{+}95, WPG99, AFP99, CBS91, HC90, LBM^{+}99, MF95, RPMB91, SLB^{+}92, SR97, TF98]. Carotenoid [KL92, ABB^{+}97, BDdWC96, GW93, HH92]. carotenoid-labeling [GW93]. carotenoids [AO91, WJM^{+}91]. Carpentaria [BRW95, LPW95, MRP90, OR95, PM91a, VS92, WH93, BLR94, CLD99]. carrying [AN97, RM94, RD91]. Cartesian [CRLP94]. cascade [GB98]. Cascading [BHH^{+}95]. case [DBB^{+}95, EE98, EGC95, Far98, GGF99, GGD99, KMG98, KHK^{+}99, MBB99a, SSR^{+}97, SFG^{+}99, VBL^{+}92]. casitas [AADSCC97]. Cassin [ASA96]. Cassiopea [FFH99, VM98a]. castration [Laf93]. Catalan [CSC92, CS93, TP97]. Catalina [Sha95b]. catastrophes [KHTP99]. catastrophic [CH96, PBVB99, Tru97]. catch [KS95a, MMBT97]. Catchability [VS92]. Catches [KBN^{+}96, CBL96a, Con94, OC93b]. categorisation [CES^{+}94]. category [Åbe90]. catenation [FPFR92]. catenatum [MHT^{+}97]. Catharacta [PCT^{+}97]. caught [CCV^{+}98, FCW98, HTCW96, RTCH96]. Caulerpa [BB97a, CC98, CC99, CJ97, FGR97, JPZ^{+}98, KMB97, MP92]. Caulerpales [MP92]. Causative [SOS^{+}92]. cause [IMB99, CS91, JHG99, KLR93, KGO94]. caused [FPB^{+}94, Flo91, HHH97, LGR95, RPL96, Ved98]. causes [LGP95, PG90, SL95, VS91]. cave [AC96a, GF95, Har97, PIC97]. caveat [MKCL94]. CaveCam [WR98]. cavernosa [LS96a, LA97a, Szm91]. Cavity [RW99]. Cavity-dwelling [RW99]. cavolini [WV95]. Cay [SB96, SJ95]. CCA [WW92b]. CCMP [CCV^{+}96]. Cd [CAATS94, LWL98, SDC99, VCMP93, WQQ99]. cDNA [KWHM^{+}96]. Cell [BMS^{+}97, LBL^{+}99, BBR^{+}93, BLC91, BZ92, BRN^{+}95, BOC90, CC90, DHM92, DRGV91, GGC99, JM99, KRJL96, LMF93, LCC97, MHvB97, Moo92b, MSND^{+}99, MWB^{+}99, NC95, PM90a, PMP^{+}96, RCGSN96, SM97b, SGP94, UYMH95, UGSBA94, WFTH97]. cell-cycle [BOC90]. cell-specific [NC95]. cell-surface [PM90a]. Cellana [WM95]. celled [MJ92b]. Celleporella [HHG99]. cells [AGGKA96, And96, GWW^{+}98, GGKJ^{+}98, KCV94, KHT93, LKF95, LS94, LFD95, MM96, RH90b, SM94, TTR99].

Cellular [GAB90a, GAB90b, BH98, GHTH94, HMB93, HMB95, LMK93, MIA90, MOM98, Moo92a]. Celtic [MJV92, PH95]. cemented [JAB^{+}92]. Central [BEU99, MBR94, AAE^{+}96, AC96c, APS91, BD98, BH93, BVL98, BCBR97, CO92, CSM^{+}98, DGT^{+}93, DB96b, ELMR98, GGH99, JMC93, KL91b, LKC95, Li95, LGH95b, MDC98, McC96b, MP95, NC93, PHL92, PD90, PD94, Roa93, RGE^{+}95, RO93, RA96, SL91, SSV97, Tho92, TcBW96, VNM97, WH91, YHGB96, YGH^{+}97, BR91a]. centres [BVL92]. centric [MP96a, TP96]. Centropages [CPC98, GM98b]. Centrostephanus [AU93]. century [EFSS95, JWL96, RAGH94]. Cephalaspidea [Ito97]. Cephalopod [BSY^{+}99, ROANM98, HPM92]. Cephalopoda [CF93b, SB97b, WTR99]. Cephalopods [RPT^{+}96, RSH92]. Ceramium [Ekl98]. Cerastoderma [AR91, KSRB99, LAJ95, NW97b, RII93]. Ceratium [HN97a]. Ceratoporella [SC90]. Ceriops [SOSV95]. cerithid [SU93]. Cerithidea [Laf93]. certain [BB92, DE91]. cessation [KHTP99]. cetacean [TBAB91]. Cetorhinus [SM97a]. cf [VMS^{+}96, SBT97]. cf. [Jeo94b, Jeo94a, JLYK97, JSK^{+}99, VPGV94]. CFU [ZF99]. Chaetoceros [GT92, KRJ98, KHT93, LW91]. chaetodontid [Fow90a]. chaetognath [DA93, DG96, Øre90, Øre95, Ter93]. chaetognaths [BP97a, BP97b, TNN93]. Chaetomorpha [KJMRC96]. chaicogramma [Liv93]. chain [Ejd98, KZP93, LJM99, MDTB94, RH90b]. chalcogramma [BM94a, BCN^{+}95, BBK91, CBI91, DO92, HHP95, HBP^{+}93, OD90, YB90]. Chaleurs [SL97a, SH93]. challenge [MR94, PB94]. chamber [NG96]. chambers [FGP98, GFH96, HG92b]. Chamelea [RR92]. Champsocephalus [KWEG94]. chances [Mal98]. change [ALB^{+}96, APS91, HLP99b, KR97b, LBH91, LGH95b, SFS95, ST99, VPM96]. Changes [BKS^{+}99, CCG98, DE91, FAVV98, FFJ93, FFF^{+}94, GGB^{+}99, GD96, HLP^{+}99a, Jon97a, LLB92, LOS90, LMT99, MACR98, MV98, NHL99, RKH96, RC95, SGR92, SBP92, SBP93, AWD95, Ano99c, AOT97, AMB^{+}96, AH96, AH98, BLJ^{+}96, IMB99, CA99a, CCC94, CSM99, DP91, DFR95, DRGV91, DL95, Dur94, Dye98, FH99, FCH99, GRBA96, GAC^{+}97, HN90a, HN90b, HDPW93, HL90, HWL96, HB98, HvdM97, HS99b, Hul97, ISF^{+}98, IPT94, Joh99, KW99a, KDHZ98, LGGLLHV98, LGC^{+}93, LGH95b, LGR95, Mac91, MGYV97, MDD95, MLC92, MBWT98, MCRGC90, MGC95a, Mil92, MOM98, MSND^{+}99, Naj97, O’B94a, OINN94, PODL99, PDC97, Pel96, Rei90, RVGC^{+}99, Rv92, RSAD98, RO99, SM94, SH90a, SS96a, SCJ99b, SDB94, TMKL97, TRPS94, TDD94, TFL95, TPHC94, VWS92, VNM97, Wes90a, Wes90b, WB91]. changes [WB92b, vBV95]. changing [BM99a, BK96a, BC99, CCS98, McL96, RCS^{+}96]. Channel [LK93, MDC98, BBL93, HM97c, MH97, RBN99, TDW96]. channels [CSM99]. chaperones [WKH^{+}98]. Characterisation [SL97a, BLJ99, SWK98]. Characteristic [JMR91]. Characteristics [SD96, ADR94, AR99, ARS93, Bar99, BKP98, BL97, BAK97, BM90a, BM90b, BD94, COD^{+}93, CK99, CS97, Dur94, EML^{+}93, HKvB93, HRFC90, IBB97, Ito97, JSGS94, LBL^{+}99, MV96, PFMG96, RBT96, SJ99, SPS^{+}96, SM97a, SC93, Van94b, VK94, VB94, WHG99b, Zaj91a]. Characterization [BDR^{+}96, EPD^{+}99, RCGSN96, GHD92, JSV92, KHT93, MS95, PSE97, SC95b, ZGM^{+}96]. Characterizing [WHIC97]. Charleston [SCCS98]. charr [RSM^{+}96]. Chasmagnathus [BI99, IBB97]. Chasmodes [Har99]. Chattonella [NU91, NYH92]. check [GS99a]. cheilostome [NT94]. cheliped [LS92]. Chelonia [HGBW99]. Chemical [AP97b, BYO95, CCJC91, CDGF^{+}93, GHD92, HAH^{+}90, HHKS92, HH91a, PLF90, PF92a, RP91a, SC92a, SL97b, VPF98, WJF94, AR99, APT99, BPS98, BBT^{+}93, CHFL95, DvdM92, Ejd98, JFP95, KB91, LC90, MMHW96, MB97a, NB96a, Nie96, PNS90, RC95, SJML99, TJST98, WRBF97, WK93, WB99, PCTF95]. chemically [UMT^{+}91]. chemiluminescent [BBO^{+}95]. Chemistry [Ano99a, CMBC98, SDA^{+}92a, CdJS^{+}92, GHD92, HFR^{+}93, Kal91b, MC94a, OSZ95]. chemoautotrophic [BSGM94, DS93a, HCF93]. chemolithoautotrophic [AC96a]. chemoreceptor [LA97b]. chemostat [FF92b, LBL^{+}99]. chemosynthesis [FKT^{+}99]. chemosynthesis-based [FKT^{+}99]. chemotaxis [BF99]. ’CHEMTAX’ [WTM^{+}96, MMHW96]. Chesapeake [CH93, CRH^{+}99, FPAH92, GCF^{+}95, GSG^{+}94, Har94, HP97, HHW90, JMP90, KSG^{+}92, KSMDB97, LB93, MH96a, MDPP91, MDC98, ML92, MHM90, MNO96, NH95a, Pur92, PND^{+}94, RT93, RCC^{+}99, RAGH94, RHP93, SK94a, SD94, Sie93, SJMC93, SJCL93, SK95b, WA99a, WH95, WHIC97, WR92a, WR92b, ZHM91]. chiajei [GS99b, SG96]. chick [GGF99, LMT99]. chick-rearing [GGF99]. chicks [RM96]. Chile [AC96c, BCBR97, CO92, CW93, EM99, GGH99, GW91, JBB92, JMC93, MADM98, NC93, Roa93, RGE^{+}95, UC94, Urb94, WMG^{+}94, WtD90, WC92, WS92b, ZMD95]. Chilean [CT98, GM98b, MP95, RO93, TVG96]. chilensis [CT98, EM99, GM98b, TTW97, TVG96]. chimeric [Mal98]. China [CCG99, Han92, KN99, TNKMK99, DZ99, LL98, ZCL97]. chinensis [CNCS93]. Chinese [Ang91c]. chinstrap [JBB98, ZGM^{+}96]. Chionoecetes [SMUS^{+}99]. chiton [MD98b]. chl [MRP98]. Chlamys [AH96, AH98, Fev92, HBM93]. chlordane [PTW^{+}91]. chloride [CK95]. chlorin [BDdWC96]. Chlorophyll [HH92, HH96, ABB^{+}97, Atk96, BD98, BT93, BAM^{+}99, BDA^{+}99, CCG96, CWS^{+}93, DP91, DAS92, EML^{+}93, FP91b, FL92, GV96, GMK97, GM98a, GTJ92, GF91, Har90, HSRHH90, IP94, JM99, Jon97a, JCMW91, KBP98, MC95, MD94b, PLH91, PLB95, SH98b, TLB93, TGG97, WHIC97, WPG99]. chlorophyll-a [BD98, BT93, CCG96, CWS^{+}93, DAS92, GMK97, GM98a, IP94, PLB95, SH98b, TLB93, WPG99]. Chlorophylls [GR93, BCG97, WJM^{+}91]. Chlorophyta [CJ97, MP92, VSPLH98]. Chlorophyta-Caulerpales [MP92]. chlorophyte [LS96a]. chloroplast [Put90]. chloroplast-retaining [Put90]. chloroticus [CK98, Lam98, LB99a]. chlorpyrifos [RWJ99b]. choanoflagellates [MS93a]. Choice [MJ92a, ARF92, KS99, MA99a, RS95, SMUS^{+}99]. cholinesterase [GBC92]. Chondrosia [CBA^{+}99]. Chondrus [WC96]. choosing [Åbe90]. Chowder [GJ96a]. Christi [LD96]. chromated [WW92b]. Chromatic [NP90, BMD^{+}90, vLTW^{+}98]. chromatid [PMETM96]. chromatograms [LM97]. Chromatographic [GCH93]. chromatography [VLKC94]. Chromis [NKGK92]. chromium [DL91b]. chromophoric [AB96b]. Chromoprotein [JNJP94]. Chromoprotein- [JNJP94]. chromosomal [CB92]. chromulma [JSV92]. Chronic [RBG94b, Rin92]. Chronographic [Min93, HP96]. chronology [KPWL93, WC92]. chronometer [KJGC96]. chroococcoid [CLM^{+}91]. Chrysaora [OPS96, Pur92, PC95, PWNW99]. Chryso [JSV92]. Chryso-chromulma [JSV92]. Chrysochromulina [GR94, KLMR91, NKB90, SC95b]. chrysolophus [HTCW96]. Chrysomonadida [PCG96]. chrysurus [NKGK92]. Chthamalus [PPO^{+}99]. Chukchi [CPH^{+}96, FNJ^{+}94, SBV98, SSA96]. chum [FS98, SOK98]. cialis [GS90]. ciavata [CZG95]. ciliate [ADR94, Atk96, CPS91, CP91, Han91a, Han95, JH97, KA97, LHL97, Put90]. Ciliated [SNJV92, Eml90]. ciliates [BR94, CBDC96, DLHB99, FC94, GF99, JT90, KS95d, Mon96, NC95, PDF97, SSM91, SGP94]. ciliolata [CLHF95]. Ciliophora [KS95d]. cineraria [Cla90]. Ciona [HlS90, PR92, PST95, SHW^{+}91]. Circadian [AKK94, KKYH97]. circatidal [RN93]. circular [SFG^{+}99]. circularisquama [KA97, MUH97]. circulation [ASRP^{+}96b]. cirripede [LB91]. Cirripedia [PGB91]. Citation [Pud93]. Cladocora [PMB99]. Cladophora [TCB93, TCRB93]. cladorhizid [VFMFBE96]. Clam [GJ96a, lYALM93, lYAML93, APCL^{+}93, BBLY98, BD99, CTSN90, CJL91, DdA95, ELH92, GK96b, GI99, JJTS97, KGO94, KG94, KGC92, KB91, LC97a, McF92, NTS90, Ski94, SGP99]. clam-zooxanthellae [BBLY98]. clams [BSP^{+}99, DMM99, GCM98, IVM^{+}98, KL94, ML91b, SL90]. clarifying [MRP98]. clarki [Sew94]. Class [Mil92, Cam96, DCRG92, LW91, MMHW96, MGC91, SW91, vdVW99]. classes [JSGS94, JMP90, KW99b, SDFJv91, SGK^{+}97]. classification [BSP98, CSE^{+}96, NA95, TJST98]. clausi [Dut98, GA97a, RMMB91, SA92]. clavata [CGZR94, CRZG95, MC94b, RCG99a]. cleaner [AMK99, Gru96]. cleaning [Gru95]. Clearance [Ole95, SSM91, SDB92, ICG96, KS95d, MUH97, NH90]. cleave [VDT92]. Cletocamptus [VDA96, VDJA98]. climate [AP99b, KDHZ98]. climatic [TN98]. climbing [Bar97c]. Clinocottus [Mga92]. ’clod [JM93]. Clonal [DCK99, CS91, Gar98, KFF92, MD98a, TDK97b]. clonality [Way95]. clone [RBSO99]. Cloning [KWHM^{+}96]. close [Bra95, HF99, KYI^{+}98]. closed [GKRD96]. closely [SDB96]. closterium [ES99]. clues [SR96]. Clupea [AH94, AH99, Bus96, FF99, Liv93, MHB94, Mun92, SGM98]. clupeoids [Val91]. Clustering [KEJ98]. Clytia [Mat92]. Cnidaria [AT96, BS95a, Gar99a, Han97, Mat92, RCG99a]. cnidarians [Maz95]. CO [Gar91, LL92, Hin92, SFWD90, CDME97, JLYK97, KM99b, PWG^{+}99, BK96a, BR97, BRZ99, FGB^{+}96, GPDF93, GPF95, GPD^{+}96, KRJ98, LL97, PJF93, RRWG96]. Co-existence [SFWD90]. co-limiting [PWG^{+}99]. Co-occurrence [Hin92]. co-occurring [CDME97, JLYK97, KM99b]. coagulation [HK97, Kep91, MB98]. Coast [KR97b, AAE^{+}96, BLDF95b, BLDF95a, BCC95, BCNC96, BCBC98, BA91, BSF95, COK99, DFJ^{+}93, DHM92, Dye93, EBB95, FCH99, FSG^{+}97, FSG^{+}98, GSC96, GAD^{+}92, GGPGG95, Góm95, GD96, HTS92, Ken92, Las98, LSG^{+}99, MDC^{+}96, MP95, NH95b, OB94c, RFNBHG94, RPN95, RHH97, RO93, She98, Spe95, SDG95, SF99, TI96, TT99, THS94, TN98, VF99, VMHH90, WT98]. Coastal [BE97, GPU^{+}97, PSK98, AFP99, AFC96, Alo90, ASP91, AB96a, ALBS96, ALB^{+}96, AMB^{+}96, BK98, BMSHC98, BK94, BB99, BBL^{+}95, Bra93, BRN^{+}95, BVS^{+}91, BGCM95, CSG93, CGTB95, CLM^{+}91, CCG96, CPSA96, CMA95, CR94, DW95, DÁSP97, DLHB99, EJ93, EG95, Eno93, FBMH94, FM95a, FC94, FCCM94, FHZ96, FLDT92, GCH93, GdGMD95, GSB98, GK96a, Gil97, GGDLE96, GM98b, GTJ92, HA90, Han91b, HGS94, Hei95, HTG96, Hol96, IF96, JMP90, KD97, KG95, KFN99, Kin92, KHW^{+}99, KDMR97, KR95, KK96, Lai92, LHRL97, LRV94, LP95, LJM99, LAG96, LFTH92, LSL98, LMS^{+}93, Mac91, MM91a, MB98, MLC92, Mee99, MMR99, MDG98, MH94a, MP96c, MZCMOCJP91, MADM98, MA96, NTK98, NPB93, NG96, NIU^{+}98, NB96b, Nov90, NDRP93, OIO^{+}95, OM99a, ONR98, ODN^{+}95, PH91a]. coastal [PS94, PHHH96, PF91a, RDT^{+}94, RATA^{+}99, RC95, RZLB96, RAlBW99, RDE^{+}90a, RDE^{+}90b, RÁSP99, RHJ92, RTRP^{+}98, RNH99, SJJMH90, SL97a, Sha98, SR98b, She94, SD99, SS99a, SS99b, SK94b, SJÆ91, SC93, TH94a, TNGB95, TT98, TFF94, TFFA99, Uit96, UHL^{+}97, UMD92, UT90, VMHH90, WF94, WOM^{+}97, YNYK94, ZDJ^{+}95, ZN96b, vR93a]. coastline [BP98]. coasts [BBT^{+}93, FPB^{+}94, MCM^{+}97, SZA94]. coatings [SDC99]. Cobalt [NFT92, GR94]. Cobb [SNJV92]. cobble [SR90b]. coccolithophore [BFF96, FBHH93, FBMH94, MG97, NGS97, NBM94, VKV95, VH99b]. coccolithophores [WYY^{+}98]. coccolithophorid [IG96]. Cockburn [LCLC96]. cockle [RII93]. cockles [Sca93]. cod [Cam96, CD96, EG97, FSG^{+}98, GCKW95, GPLB96, GA97b, GW97, GS99a, HSB97, HM94b, JL96, Kal91b, KS99, LAK99, MJW96, MF96, MLDM95, MHB96, Nie93, NKV94, NW97c, PF98, PB98, SCJ99a, SAG92, SFF92, SK95c, TH90, TRMW95, TB95]. Codium [GHED^{+}94, VAD^{+}94, VD95]. coefficient [AFZ96, LSH96]. coefficients [BAM^{+}99, HdWS91]. Coelacanth [PFH98]. coelacanths [Fri97]. Coelenterata [EL90, WV95]. Coelosehs [BLD94]. coenriched [SJMC93]. coeruleoalba [MCM^{+}97]. Coexistence [HK94, HS94b, LT93]. Coherence [JWL96, AMBH98, ELMR98, PS92]. cohesiveness [BSG^{+}97]. Cohort [Han99]. cohorts [Kne93, Lop91, MC91c, TDER97]. COI [BSP^{+}99]. colbecki [MONR90, VCMP93]. colcarva [CLW99]. cold [Beu92, KTGC99, MGD91, NT96b, ODS^{+}96, RT92, SB98, WF90]. cold-water [WF90]. coliform [MNH92]. collagen [SDFJv91]. collect [JCH93]. collected [BG94, CSE^{+}96, DO92, PPR94, RDPB99, RGBJM95, Sha98, ZN96a]. collecting [NR98, Sty97]. collection [Kne91, WD94]. ’colloidal [TSS93, KNM93, SAH98, SKKO97]. colloids [Kep94]. Colne [RNHO98, ONH^{+}97]. Colombia [AvS97, GDvS98]. colonial [BLJ^{+}96, CRGZ98, Hur92, ORL97, SMB97]. Colonies [HSMS99, LGdC^{+}90, Nau96, TFL95, VB94]. colonisation [FTP99, GJVG96, GH93, HHB97]. Colonization [DG93, FKB92, AA92, DSJ90a, DSJ90b, HS98c, PA94, PDC97, SML97, SF94, VDM90]. colonized [BB97a, RCN95]. colony [BM98, FK95a, MB93, Mil98, OP99, PCT^{+}97, VK94, van92a]. Color [BE97, HM91, PLB95, WHIC97]. colour [BA97, SAA99]. colourless [BF95a, FB95]. Columbia [HM97b, SR98b, BL94a, CB96, HTS92, HCC^{+}91, Mga92, PSE97, PG90, Sew96, THS94]. columbiae [Bus98]. Columbian [SS92]. column [BRC99, BH93, BGC94, CCSN93, DAS92, FKR90, GHD92, HHN^{+}96, KNK90, KKK^{+}90, MB97b, Nak92, RGS95, RLT^{+}94, Sch96, Sie93, SJMC93, SJCL93, TD92, TGB99, TL91, TS90c, VWS92, Vis97, WSMM99, WPY98, ZB99b]. Comacchio [SSG96]. comatum [JH97]. Combined [Ano96a, Ano97a, Ano97b, Ano98, Ano99b, QQ97, QQ98, AFP99, JKC^{+}93, MBK95, VK90, WJD93]. combustion [URHR98]. commensal [VTGK99]. Comment [BS96a, BAR96, Mil90, Cus95, DHB^{+}98, Gib94, JNCS90, Kep98, Käh98]. Commentaries [Bro95a, Bro96a, Bro96b]. Commentary [MP94]. comments [BS96b, JRB97]. commercial [GCF96, San94c, San95, Sch93]. commercial-scale [Sch93]. commercially [PDD^{+}99]. commitment [CWB93b]. common [BCHL93, BLH99, COD^{+}93, HP90, IWF^{+}97, LT97, MTP92, RII93, VB96, WW93, WTF96]. communis [Lee90]. communities [Aro98, AWVP98, CA92, CMD^{+}94, CLB91, CPH91, CLKO98, CSB97, DB96a, DY92, DW97, Dit96, EB96, EJ93, EJ96, EPN^{+}99, FCMB94, FP96b, FPP^{+}97, GPB96, GAR93, GRGK96, GF99, GW93, GI99, HG98, HCS90, HGS95, HC94, Hsi95, HBC^{+}91, JLD^{+}96, JP96, KRM94, KE97, KV90, Ken94b, Ken94c, KDR97, KM92, KDHZ98, KO92, KK95, LOS90, Lew97, LVP98, LGR95, LSP99, LBK90, MDD95, MHS94, MG94, MRP90, NLR^{+}93, NS96, NR94, NIU^{+}98, Ols93, OG95, ODS^{+}96, PLP^{+}93, PPM94, PKL97, PCK^{+}96, PCC^{+}97, RB99, RCGSN96, SC90, SH97c, SW98a, SNJV92, SPH94, SRW95b, SM97c, TD91, TPHC91, THCD95, TN98, TTP^{+}95, UMT^{+}91, VDG95, VWD^{+}95, VAV99, VVT97, vLTW^{+}98]. Community [AP99b, BS92, Cal95a, DHH98, Dye98, GPDF93, GGPGG95, HFJS91, KM92, MS93b, PBA^{+}95, PBVB99, RT92, SR90b, AvS97, AGW93, Amb93, AW95, ALB^{+}96, AD94, BKS^{+}99, BS95a, BSGM94, BH92b, BB99, Cal95b, CGTB95, CLM^{+}91, CA93, CH94, Cro98, CP96, CB97, DFB96, FTGW90, FTW90, FGKP95, FGS98, FVDL95, FKT^{+}99, FII95, GJVG96, GB91, GdGMD95, GPD^{+}96, GPS93, GAD^{+}92, GW94, GB98, HS94b, HSH94, HB95, JHN99, JSJMH90, JDSB99, Joc95, JKFB95, KH92a, KSS90, Kea90, KD96, KHK^{+}99, KHW^{+}99, KKW97, KDR^{+}92, LKC95, LB99b, LBJ^{+}98, LGJ99, LF91, LL92, LPW95, LTM90, MCC96a, Mac94, MDD95, MLC92, MP96b, McG95, MS93c, MR91, MZCMOCJP91, MSM^{+}98, MN99, Nai93b, NK93, NTS90, NKB90, NH95b, NHL99, NB96c, Nov90, NDG^{+}99, OM99a, OE91, OSC97]. community [OSC98, OWM92, OW95, PWR96, Per91, PRW^{+}94, PBv^{+}95, PPF95, PD90, PD94, RBF92, RPM^{+}94, RD99, SPB95, SB98, SK94a, SJ99, Sch90, SNC^{+}98, SL91, SDBW95, SK95b, SGW94, SRW95a, SC95a, SSP97, SNR^{+}96, SPDM91, SBP92, SBP93, SS96b, SE93, TI96, UK98, VWS92, WC93, WW92b, Wes90a, Wes90b, WA99b, WOM^{+}97, Ber90]. Community-level [Dye98, Cro98]. Comore [PFH98]. Comparative [BCR92, BU93, BKB94, BGC94, CA92, HRCW98, MCSV95, MU97, MGC95b, NW97a, NC95, OM99b, PPM94, PLBP96, RSB93, VAV99, YKL98a, JH97, LMF93, MC98, NEIR95, ZCPA94]. compared [Fra95, GPU^{+}97, JBK99, NA99, RGT98]. Comparing [CT97, WC93, BBRBK94]. Comparison [Bec98, BAK97, Con94, DO92, Gil97, Hul99, MBM92, McC94c, PM90a, PLC^{+}91, Ric91, RK99, Ruz96, SC90, SFF92, YS96, YFI91, AR99, AW97, Aro91, BTYO94, BCAAC96, BMHH94, BBO^{+}95, BZ92, BCv^{+}92, CSR^{+}92b, CPH91, EPL94, EM93, GKMT99, GP91, HWLC98, HTCW96, Irw91, LGED98, MB95a, MB98, MHC92, MGAE99, PPB90, Row90, SRH^{+}98, SCB92, SSHM91, SCA92, SPDM91, Tuo97, Van92b, Wah97, YY91, vG91, vCRS92, CD96]. Comparisons [FKC^{+}90, WC90, AHT^{+}90, BA90, BAH94, HS91b, KWD93, Las98, ML93, Mic96, PPP^{+}96, RH90a, WTM^{+}96, Web91a]. compartment [BMFR97, BFMM99]. compensation [SFSA99]. compensatory [BCL93, CRH^{+}99, SH99a]. compete [Rin96]. Competence [BDZB98, Dav94]. competing [Cla96, Pur91]. Competition [FDZ93, WHG^{+}99a, BCC96, Bre96, Cha91, Dai90, EA92, LBDD97, NTK93, SFV96, Som96, SJC99, TSB93, WC96]. competitive [BLJ^{+}96, Nan96, TB95]. competitors [GPB^{+}98, HS98b]. complanata [Lew91, Lew92]. complete [BWHC91]. complex [BMG99, ECF^{+}97, Geo94, GBV96, KHK^{+}99, RVT^{+}97, RPN95, WW96]. complexes [NK92a]. complexing [BFD91, VB96]. complexity [Gri94, Hul97, Kin95a, LB99b, LB96b, MM97a, Pen90a]. component [SM90]. Components [LC98, Bus97, FKC96, FMH96, HH95, RN93, WB97, van92a]. Composition [Hol96, Li95, Ano99a, AALT99, ALB^{+}96, BBFD98, BYO95, Bar95, BS95a, BF95a, BRL97, BKL99, BSH^{+}91b, BR97, DDM^{+}99, DWG98, DH92, FSdM^{+}92, FCVS92, Fra94, GPP98, GGB^{+}99, GAR93, GHTH94, GAC^{+}97, GRSV98, GT98, HS91a, HNC92, HSN93, HNL94, JRMN^{+}97, JSV92, JFP95, Kal91b, KL95, Ker94, KPWL93, KWEG94, LMM98, LF91, Leh96, LH96, LC90, LH91, MCA90, MACR98, MVA99, MP96c, MLH96, Naj96, Naj97, NT95, Nie96, OTF95, PPM96, PMMH95, PD95, PWPS95, PSB^{+}97, PFMC95, RC95, SKvB98, SDBW95, SU93, SV97, SCEG93, TDH^{+}94, VK90, WH96, WJD93, WMH96, WF90, ZDJ^{+}95]. compound [LM97, PFMC95, RB90, VY96]. compounds [BK94, BQKC97, DCM90, HS94a, HS98a, IWF^{+}97, JMD^{+}99, KE95, LQ97, Les96, RR97a]. comprehensive [GBB^{+}98, OG95]. computer [AAD91, CC90, JPR92]. computer-aided [AAD91]. computer-controlled [JPR92]. computer-simulated [CC90]. concentrated [MLDM95, Sut92]. concentrates [WP95]. Concentration [JTPP92, CA97, CCG96, DM96, FDZ94, FF92b, GB95, GT96, HN97a, KI97, KS91, LGB91, MSD98, MRMR92, MP92, PCBA94, PMOD90, PWD^{+}91, SDC99, SCMM95, YSL94, YKL98a, van94a]. concentrations [AB95, BC90, Byt90, CLGM96, CAATS94, CNCS93, CCC94, CSB94, DO92, DCM90, EBM93, Fly99, GPP98, GSMM96, HTLK99, Jon97a, KDKS96, KRJ98, LGB91, MHB94, MGM96, MCFS96, NW97b, RWJ99a, RDM91, SGL^{+}97, SPCF97, SPB^{+}99b, SCF91, TLB93, VV95, dNDS98, vDdWKW99]. concept [BBRBK94, EFSS95]. Concerning [Käh98]. conch [Dav94, SL94]. conchilega [PE99]. Concholepas [MADM98]. conchyliatum [VPF98]. conclusions [Hei92b, Kar92, SDA^{+}92e]. Condition [FCW98, And94a, Bro93, CBI91, CD96, HF99, ISF^{+}98, JL96, Ker96, LMTF96, MM93, MM95a, SKPT97, SFF92, WDB98, YSY^{+}99, vW93]. conditions [AP99a, BCN^{+}95, BG94, BFMM99, Bra93, BGH95, BGCM95, BD94, CPC98, CHS^{+}95, CCS98, Dee90, DHKP96, DHM92, DGHT^{+}98, EBM93, FGKP95, GB91, Har97, HGB99, JPR92, KMR98, KGS95, Lav97, LCW99, Moe97c, MH96b, NKV94, OD90, RR99, SH99b, WYY^{+}98, dRAV92]. conducting [GS94a, PN99]. conduits [EAEP98]. confined [HG92b, LF91]. confinement [JAL91]. Confluence [GTL^{+}91, GST^{+}91]. confluens [VDA96, VDJA98]. confocal [BP93]. confoederata [RR99]. congeneric [WB99]. congeners [JA91]. Congo [Bia92]. conifera [HLBW93]. connectedness [Hun93]. Connecting [BM99a]. consequence [IMB99, LGH95b]. Consequences [JS93, KR94, SDB96, BQR99, BPS98, FF92a, FAP92, Ger90, KS99, MBL^{+}98, PPFP98, Pfi97, RC95, RH91, Sos96, VKF^{+}98]. conservation [GK97, LL98, PFH98]. conserving [SO96b]. consideration [SRSH99]. considerations [Bou90a, Bou90b, CNP95, DDM^{+}99, HLH96, KKL^{+}98, Leh96, SC92a]. considering [vdBTvDR96]. consorts [Sie93, SJMC93]. conspecific [BQKC97, Mar91, MLES94]. conspecifics [SFA98]. Conspicuous [VS91]. constant [CA96a, MC97b]. constraints [CM99, GT92, Har96, KWL98]. constricta [WX97]. Construction [WW92b]. Consumer [Bre96, MCC96a]. Consumption [LS96b, Wit99, ZNH93, AAS^{+}91, AJST98, AH93, AH94, BS95b, Bre96, BG93, BK92, CM98a, CW93, CGCLH99, HdWS91, KJLH96, LK94, LB93, MRMLP97, MC98, PDF97, RO99, TMC98, VTNA91, Web91a, WK93, WSS94, YB92]. consumption-food [CM98a]. contact [OS96, Sun96, UYMH95, VM98b]. container [DZ99, ZD99]. containing [KGC92, LC95, Vis93, WPY98]. Contaminant [LME92, LFD95, Ack97, MFF97, Sün97, WDB^{+}95]. Contaminant-induced [LFD95]. Contaminants [SD92, BP94, BBT^{+}93, KCV94, WT99a]. contaminated [Ber90, DvdM92, RA92, SPPL92, SCCS98]. Contamination [Hei95, BQR99, BL95, HC92a, MPH96, PLBP96, SGR92, TLW^{+}96, VBZ^{+}94, WLT^{+}95]. Contemporaneous [JP93b]. contemporary [KML91]. Content [Vos91, ATH95, Bro93, BZSG96, Cle93, CWHA99, DPO91, DCRG92, Dua90, EL90, FFF^{+}94, GCR98, HKA96, HSN93, KW97b, Mar99b, PWPS95, ROM^{+}91, RBB^{+}91, RMGB^{+}94, SH97c, Tee99, UMD92, VCMP93, WPG99, vBV95]. contents [BH92a, LPF94, LMTF96, SPB^{+}99a]. continental [AML95, Bia91, Bia92, CSB97, EJ93, FH96, FII95, Gra94, Gra96b, GDW90, HBC^{+}91, JT93, KE97, KKW97, LFR^{+}97, LKvRH96, LEHvR96, MJV92, MP96b, MSM^{+}98, OG95, Pir92, PH95, RSS94, RGE^{+}95, SR90a, SRSH99, SLB^{+}92, TDBH94, TB98, TGvJW95]. continua [FCP98]. continued [VH99a]. continues [Kin95b]. Continuous [BM99b, MWT96, SK98a, SEGP91, PRM^{+}96b]. continuous-flow [SEGP91]. contraction [FK95a]. contradictions [LC98]. Contrasting [BCN^{+}95, BG93, Tol98, DHH98, ELH92, GD99, Hun93, LGL^{+}95, SH99b, THPC96, TGLP95]. contrasts [JBB98, JMP90]. contribute [DA93, KFPR94, Reu98b, Tro92, UMTM98]. Contribution [IT96, KK91, MM90, NPB94, SS97, SPS98, ADR95, BFJ^{+}93, BVL98, CB98, DSV^{+}91, Dou93b, GK96a, GM98a, IP94, Kin95a, KHW^{+}99, LS96b, ML91a, MLL98, Wai90]. Contributions [KG94, MH90, MSD92]. Control [Hoe91, OM90, TS90a, WCP^{+}98, AWH^{+}99, BNvB90, CK95, DCBG99, EST95, FF92b, Gas94, GPF95, GU98, LGH^{+}93, LR96, MPS96, MCW^{+}91, MLTD96, NK92b, PB96b, PB97, RPMB91, WJF94, WFTH97]. controlled [AB95, JPR92, MAdG^{+}96, NAW99, OD90, RCTH99, VLKC94]. controlling [BKP98, BF95b, CAATS94, Hil91c, Kro91, LD94a, LKK92, PDBA90, SK94b, Sün97, WYY^{+}99]. controls [GEW97, MDC98, McG92, GJ97a]. convection [BWHK99]. Convergence [FP96b, HSLV94]. conversion [RB98]. convict [Pla93]. cooccurring [Jeo94b]. Cooperation [Pak95]. cope [PlMM94]. cope-pods [PlMM94]. Copepod [BNvB90, BH92a, FC94, GGH95, HFG^{+}95, JFP95, SCIA99, SKvB98, Tee99, AJS95, AN97, AS97, AWH^{+}99, BBC^{+}97, BDdWC96, BP93, BD94, CA99a, CN92a, CM90b, CPL^{+}96, CMT94, DP91, DWSA98, DDW90, Dut98, FD98, FR91, GGB^{+}99, GS94b, GCC95, GSMM96, HvB91, HRKR93, HLBW93, HSW99, HN91, Jeo94b, Jeo94a, JKT^{+}98, JT90, Ker90, KSV96, KSV99, KBS98, KEV99, LY97, LGED98, MH93a, MRMR92, MMGIM90, MTB^{+}99, Ohm96, Øre91, ØW93, Øre95, PL99a, PR93a, RD91, RMMB91, SK95a, SW91, SW98b, SC96b, TSI^{+}93, URHR98, US95, Uye96, US98, VV95, Vos91, WDB98, WB94, Web91a, WMH96, WR92a, Wil92, WDD92, YS96]. Copepoda [AS99a, ASSWM97, CA96a, CA97, GM98b, KGS95, KBS98, Met98, SA92, SSHM91, SM98b, TH90, VDA96, VDJA98, WASSM97, WC91]. copepodite [LHL93]. copepods [Aks96, Atk96, BK98, BH91, BS92, BS95c, BM95a, BP96b, CN92b, CGTB95, Car90, CTEF91, CN95, CBDC96, CS99, CLW99, ENSR90, FTS97, FNF91, FH97b, Gra91a, HKG93, HH94a, Hic92, HS97b, ICG96, KB94, KS95c, KS95b, KV99, Kle93, LF97, MLC97, MRMR92, McC92, MSS95, MHvB97, MHRV^{+}99, Nag96, NS96, Nor96, OF94, PC96, PDF97, SK97, ST98, SF91, TDVF99, TWWE95, TMC98, TJ90b, Tis98b, UT90, VKF^{+}98, Web91a, Web91b, WP92]. Copopoda [RD92]. Copper [AHT^{+}90, BDH94, BFD91, DL95, HDB95, HBD95, HMB93, LP99, MAAT98, RBG94b, WJD93, WW92b]. Coral [BM98, GCGR90, McC94a, Adj97, AvS97, Aer98, ABB^{+}97, Ano99d, Aro98, AHS97, Bak90, BLJ^{+}96, BJdJ^{+}98, Bee97, BMH91, BKWML93, BLST90, BLD94, BDA^{+}99, Byt90, CDHD92, CH95, CK99, CT97, CWB93a, Cla96, CG99a, Cli95, Cro99, DB96b, EB96, ELG97, Fab95, FD99, FCC91, FP91a, FPGC^{+}98, FPGJ99, Fow90b, FDW92, FHCC98, FGB^{+}96, Fre97, GWW^{+}98, GPDF93, GYL93, GPD^{+}96, GMO95, GGH95, Gre96, Gri94, GEW97, Gru95, GP98, GT98, HHYD92, HSH94, HB95, HFR^{+}93, HSK^{+}93, HSK^{+}94, Hoe91, Joh99, JS93, JKFB95, JKL90, JM93, JFS90, JFS92, JHG99, JLGMP97, KKL^{+}98, KMG98, KSO^{+}97, KRFL97, KRF^{+}98, KYN94, LS96a, LCBD94, LB96a, LGJ99, LSGW98, LCE97, LL96, Lew91, Lew97, MSC95, MG96c, MA99b, MM97a, McG94, MJ97a, MMD93, MB93, MNB94, MB95b]. coral [MPB97, MD94c, Mil94b, Mil95b, Mil97, Mil98, MLL98, Mum99, MJC97, OBL97, ORL97, PF92a, PR93b, PD90, PD94, RK96, RGLS99, RS99b, RW99, RHB96, RS91, RSS94, Rog90, RMT91, SKM93, Sak97, SRSH99, Sam97, SOW97, SK98a, SK98b, Sch99, SNC^{+}98, SC94, SCLS94, SDC^{+}92, SMB97, SJ95, SC96a, SP99b, SB97c, SEAI96, TH94b, UK98, VB93, Van94b, VK94, VB94, VDG95, WO94, WV95, WHW93, YAG92, YMD94, YFI91, Zel97, Zel98, vTS97]. coral-bacteria [SCLS94]. coral-boring [BKWML93]. coral-dominated [HB95]. coral-dwelling [MJC97, PD90, PD94]. Coral-eating [McC94a]. coral-feeding [MJ97a]. coral-reef [CH95, CWB93a, LCE97, PR93b, SC94]. Corallimorpharia [Cha91]. Corallina [AdN91]. coralline [BCC94, JMR91, JSOG91]. Corallium [ASN93]. corals [BDZB98, BDSL94, CED91, Dai90, ECF^{+}97, FK95a, FBR^{+}95, GBJ94, HB95, HGJ99, Jon97a, JBO97, Jon97b, KML91, KL91c, KPWL93, KWL96, KWL98, KCD^{+}95, LNSL91, LGP95, LGH95a, Lew92, MSC95, MDSB92, MNK^{+}97, MGD91, NB98, RVGC^{+}99, RH90a, Rin96, SCLS94, SHM^{+}97, SP93, Szm91, TTL^{+}96, Tom91, WUA^{+}99, vW98]. Corbicula [BTKK99]. cordatus [PF95]. core [DL93b, FP96b, NB90, ZR92, ZBK^{+}98]. cores [HG92b]. cormorant [GWWP99]. cormorants [BRLM90, BS95b, GSC95]. cornea [ROM^{+}91]. cornus [McC94a]. cornutus [KSH97]. corolla [DeV92]. coronata [MTP92]. Corophium [BF95b, CD99, GH93, GH94b, GH94c, MJ97b, PNB94, SHC96]. Corpus [LD96]. correct [AFZ96, vR93a]. correlates [DSS^{+}98, LCBD94]. Correlation [AWH93, MIA90, JW96]. correlations [Fev92, HKR96, MAC92, MBB95]. correleations [BCL93]. Correspondence [McG94, RBF92, WA99a, SPBS93]. corresponding [KSH97, RPT95]. corrugatus [SDB92, SDB94]. Corsica [HRD92]. Cory [GMF98]. Coryphaenoides [APS91]. Cosmonaut [Pak95]. cosmopolitan [CS91]. cosms [CR94]. cossyphus [Cow91]. Cost [Ang92, Bus98, CWB93b, LL94, RHS99, RI90, RTJ^{+}93, RBT96, Sto97, TWWE95]. cost-effective [RHS99]. Costa [DE90, Bia91, GCGR90]. costatum [BR97, HFN91, TT98]. Costs [AP99a, FGH95, BA94, GMC91]. Cottidae [Mga92]. Coulometric [Irw91]. counter [HCS93, McG97a]. counts [CT97, EAC^{+}97, MHvB97]. coupled [VM95]. Coupling [Sci91, Bra95, DG91b, HAR95, KFF92, KGW93, LJM99, LTM^{+}98, MFF97, Mil94a, MWT96, NNB93, PM90b, VBCRB97, vDdWKW99]. course [HHN93]. courses [DL91b]. Cove [PMMH95, Ole95]. cover [Aer98, BCC92, DGCT93, GA97b, LC93]. covered [FPG95, GFPD92]. CPR [BM95a]. Cr [LWL98, WQQ99]. Crab [Pal97, AAWD97, AP99a, AS99b, AHMd90, Ang91c, Ang95a, Ang95b, Asa91, BAH94, BI99, BLJ92, BM95b, CW93, CCDA98, CM98b, CC97b, DE90, EAEP98, EDP^{+}98, EPDC99, FIA93, Fer99, FW91a, FFT98, FFR94, FDR^{+}96, FTBR97, FGG95, GAV93, HBR^{+}91, HN93, IBB97, JSH99, KW97b, Kog95, Kyo92, Lee97, LH91, ML92, MPvM98, MBL94, MDRW99, O’C90, OJ97, OJ99, Olm94, PDD^{+}99, PODL99, Per90, QCM97, RKH96, RN93, RS99b, RvMM97, SMUS^{+}99, SW93, Smi95, SL90, SK98c, SAA99, SLSH98, TTW97, TA93, WJD93, WE95, WRBF97, WFH99, WBLM95, WC92, Wol92a, Wol92b, Wol93, WG97, ZN96a, ZN96b]. crabs [BS94, Bar97b, Bar97c, Bar97a, Bar99, BA90, BKH99, CBRD92, CWWH99, DS93b, DL95, EK95, ELH92, EAT95, FW92, FSGG96, Gar99b, Har96, HHW90, JA91, Ken92, LMO99, LS92, LK94, LFS90, MRF^{+}99, ML91b, MZFHC96, Mor96, O’C93a, ÓN97, PWS98, RATA^{+}99, RDA97, SH99a, SM99b, War90, Wat98, ZAN99, ZFFH^{+}94]. cracherodii [AAE^{+}96, LK93]. Crambe [UTB^{+}95]. Crangon [Beu92, Weh91, WA95]. crassa [MDRW99]. Crassostrea [BM94b, BN91a, BN95, BH94a, BD96, BV99, Bro90, BA96, Dou93a, Dou93b, DLHB99, GMC91, HSTY98, LHL97, MR97c, MAAT98, PLBP96, RR97b, RCH98, RM95, Tha92b]. crater [ZT90]. created [LTT96, MW97]. creek [CMD^{+}94, HG94, NL94]. creek-banks [HG94]. creekbank [TZ99]. creeks [BNK^{+}97, HKY^{+}97, MBM92]. Crepidula [CP90]. crested [BS95b]. Cretan [DMP^{+}99]. Crete [KE97]. Crisia [NR98]. crispus [WC96]. cristatus [TSI^{+}93]. criteria [FKC^{+}90, SMUS^{+}99]. Critical [KFD99, EFSS95, GQ96, VDVC94]. criticisms [JRB97]. critique [BC91, DD97]. crocea [LB96b]. crop [MGEC97]. Cross [DSG^{+}99, PD90, PD94, RS91, RSS94, SRSH99, YST^{+}97, McC96b, RB90, SR90a, Sha98, SM95, SC93]. Cross-continental [RSS94, SRSH99]. cross-fertilization [RB90]. cross-infectivity [SC93]. Cross-shelf [PD90, PD94, RS91, McC96b, Sha98]. Cross-slope [DSG^{+}99, YST^{+}97]. cross-system [SM95]. crossed [Bro90]. crowding [LC98]. crown [BMBD95, HK96, HKR96, HR96a, JSOG91, RI90, SB99]. crown-filament [RI90]. crown-of-thorns [BMBD95, HK96, HKR96, HR96a, JSOG91, SB99]. Crozet [BGG^{+}97, GKB^{+}97]. cruce [Ber90]. crude [GCKW95]. cruise [Hei91]. crust [OKK95]. Crustacea [AT96, APT99, PD90, PD94, TAD^{+}94, TDH^{+}94, WD90, AS99a, AAWD97, Bra95, BB93, CCDA98, CA99b, CBT98, Fra93, Gen92, Har96, JGT96, MRF^{+}99, MC97a, MHD97, Moo94, PAR98, SlRR94, SS91, SF99, Vet96b, VR98, WW90]. Crustacean [SDBW95, CM98a, CCD94, KKT96a, RK95a]. crustaceans [EBM93, GF90, Jua92, MC98, MZ92, RATA^{+}99, RHP93, WTO96]. cryopelagic [KGW93]. cryopreserved [McF92, MC94a]. Cryptic [WB97, EGC95, HHB97, MTP92, MSCTBE96, WR98]. Cryptodonta [BL97]. Cryptofauna [PD94]. cryptonisci [OR95]. Cryptophyceae [LC90]. Crystal [Nor94]. crystallorophias [KH98, Nor94]. Cs [KI97]. Ctenochaetus [SML97]. Ctenophora [AT96]. ctenophore [CLW99]. ctenophores [CH93, Fal96, PND^{+}94, PC95, SC92b]. Cu [CL90, CAATS94, VCMP93]. cucculata [HSTY98]. Cucumaria [SMTL99]. cucumber [Sew94, SMTL99]. cue [FFH99]. Cues [OJ99, Ejd98, FFR94, FDR^{+}96, Har96, OJ97, SF99, WRBF97, WD94, WLL97, WB99]. cultivated [LFRB97]. cultivation [Pre95]. Culture [ADR94, CHC94, DvRN99, EGMC94, HGS94, HLF^{+}91, LBL^{+}99, Moe97b, MWB^{+}99, PRM^{+}96a, PRM^{+}96b, PLP92, RVGC^{+}99, SK98a, SE93]. cultured [FBMH94, GF94, HDPW93, KGS95, LCW99, LH91, MC99, Ral98]. cultures [BVP91, CLM^{+}91, CH97b, DGO92, LW91, NK90, NU91, RGS90, SJCL93, Sv93b]. cuprea [PL99b]. curlews [RD97]. Current [VPM96, BSS92, Bro95a, CSB94, Ken92, MD93, RHH94, SFG^{+}99, Wor95, ASN95, BGC91, LGGLLHV98, Mil95a, Ven98]. Currents [LSM92, BMBD95, Hil91b, KSH97, RPN95, VBCRB97]. curvata [Met98]. curves [MFD98]. cutlassfishes [KN99]. cuttings [DvdM92, KDR^{+}92]. cuttlefish [RFNBHG94]. CVAA [RH90b]. Cyana [LBL91]. cyanide [JHG99]. cyanobacteria [BHH^{+}95, CLM^{+}91, CR95, GAN^{+}94, KEV99, MHRV^{+}99, NW97a, SJ97, WHG^{+}99a]. Cyanobacterial [KLR93, CMB^{+}99, FDZ93, HKK^{+}95, SFS95, TNP97, Vil91]. cyanobacterium [CSM^{+}98, CCC^{+}90b, KHK^{+}99]. cyanophages [SC93]. cycle [ASSWM97, BS96b, BOC90, Buc91, CN92a, CW99a, CA95, CRZG95, CSS93, CD99, CF93b, DRGV91, DSV^{+}91, GGC99, GM98b, HVL96, HCG^{+}93, JM99, JCCCB94, KKL91, KGW93, KK96, LNQT93, LCC97, LBL^{+}99, LFBH94, MDG98, PMP^{+}96, PRC^{+}92, PGAJ99, PR93a, RDT^{+}94, RCG99a, RCG99b, RSH92, SH90a, TVG96, VNM97, WFHM99, WASSM97]. cycles [ACC90, CC90, CC91b, CGZR94, DL94b, FMRC93, HM97c, LGL^{+}95, MH97, Pec99, Van92b, VS96, Vet96a, vG91]. cyclical [TN98]. cycling [ASRP^{+}96b, ASRP^{+}96a, BU93, BNW94, CK90, Col92, FCCM94, Fuh90, HFJS91, JTB99, KEvN95, LMS^{+}93, MOM98, NH95b, PBA^{+}95, PIC97, PLS^{+}95, RLC92, RT93, RÁSP99, Sim92, SBH^{+}95, TSB93, TCB93, ZCL97, ZB99a, ZB99b]. Cyclone [VDG95]. cyclones [CCG96]. cyclonic [YSY^{+}99]. cyclopoid [BS92, GS94b, NS96, US95, US98]. Cyclopterus [WB91]. Cyclosalpa [MPM97]. cyclostomatous [NR98]. Cyclotella [BEG97]. cydonium [KWHM^{+}96, KBA^{+}97, UKR^{+}91, WKH^{+}98]. cygnus [Jer90]. cylinder [SFG^{+}99]. Cymatocylis [CES^{+}94, WMPT94]. Cymodocea [DSJ90a, DSJ90b, KM99b, MD94a, PDC97, PBDF98, PH91a, TDK97b, TDK97a, VDM^{+}97]. cynoglossus [MRS96]. Cynoscion [CE93, DE94, GT96, TJST98]. CYP1A [GCKW95]. Cyphoma [RH90c, VT92]. cyprid [PPO^{+}99]. Cyprideis [JGT96]. Cyst [MHT^{+}97, AKM^{+}99, NU91]. Cystoseira [BCC92, HRD92]. cysts [IT96]. cytochemical [KDL92, KCV94]. cytochrome [GLBF92]. Cytometric [JM99, BVP91, FL92, LZYW93]. cytometry [CB98, JLD^{+}96, LMF93, Lip95, VK90].

D [MNH92]. D. [Coc90, CLHF95]. dab [BVL92, CB92, CMLE92, EGKE92, GLBF92, HHH92, KDL92, LACM92, LME92, Pro92, RA92, RVv92, SH92a, SGR92, SFPC97, VBL^{+}92]. dabbling [BL94a]. dabs [KR97a]. Daily [GCR98, HN90a, HN90b, HHN93, MC98, NFM98, Sep94, dT94a, Big92, BM94c, FPPM96, LT97, MLC92, MC97a, MKCL94, PPM96, PPMF97, TP95, ZCPA94]. Damage [DF94, MNB94, WUA^{+}99, Aro91, BDSL94, Kai96, LFD95, MPB97, VBZ^{+}94]. damaged [GCM98, Rin96, VB94]. damicornis [AHS97, LL96, MNK^{+}97]. damselfish [MMD93]. damselfishes [WB97]. Dan [SS95]. Danish [JSJMH90, Ole96, SJÆ91, TCB93, TCRB93]. DAPI [HP98]. dark [WR91, ZA96, GJ96c]. darkness [tor93]. data [BBRBK94, BM99b, BMHH94, BJTB96, BM95a, FIL93, JE95, LBV^{+}99, LS92, MJW96, MAC92, OSC97, RWCM99, Rel96a, Rel96b, SRG^{+}99, SDA^{+}92b, SDA^{+}92a, TMCE98, WASSM97, WC93, WHIC97, WTM^{+}96]. date [ZHM91]. dated [Cam97]. davisae [US95, US98]. day [GGB^{+}99, JBB98, SR98b, Van92b]. day-night [Van92b]. daylength [BRZ99]. daylight [SMH97]. daytime [Gen92]. daytoni [Vet96a]. DCMU [FL92]. DCMU-enhanced [FL92]. DDTs [MCM^{+}97]. dead [CNP95, LGH95a]. death [KCI98]. DeBlois [Cus95]. debris [GSC96, LF92, LF94]. debt [VH96]. Decadal [Aro98, ASA96]. Decadal-scale [Aro98]. decade [JJ99]. decapod [CS93, CM98a, GF90, HFGT94, Jua92, MC98, MZ92, WD90]. Decapoda [Har96, MRF^{+}99, SF99, AHMd90, Ang91c, AT96, BB93, CCDA98, CA99b, DF92, Dem95, GS91b, HMHDA94, Jon93, MC97a, PAR98, QCM94, SlRR94]. decapods [LK99a]. decay [HB91b]. December [GST^{+}91]. Decimeter [BN91b]. decisions [MTB^{+}99]. Decline [RAGH94, JRB96, MDC^{+}96, MHB96, VS91]. declines [AAE^{+}96, KGO94]. decomposing [FFH99]. Decomposition [KE95, KS91, LF92, OB93, PH91a, AWH93, BM91b, BM95a, CA95, CBB90, CR94, Lee90, PM91a, VSSM98]. decompression [BTYO94]. decontamination [PLBP96]. Decoupling [BFF96]. decrease [AH98, BAY91, WC95]. decreasing [AH98]. decussatus [PCBA94]. deduced [FRP^{+}99]. Deep [BA90, LBL91, Ter93, VWD^{+}95, AAW98, BGT91, BL97, BSGM94, BVV^{+}94, BL94c, BL96, BK92, CSC92, CA92, Car93, CS93, CM98a, CJ93a, CS97, CSSPV98, DDM^{+}99, DF92, Dem95, EML^{+}93, FD97, FKT^{+}99, GSM^{+}97, GSC96, Gra94, GPU^{+}97, GP90, HKT^{+}96, HSN93, JDLL95, KW92b, KW99a, LK99a, LC95, MY98, Mau92, MC97a, MC98, MD94b, OC97, PZCJ99, PGM^{+}98, Por94, PFMG96, RR99, RT95, RSD^{+}91, RT92, SH90a, STK99, SRG^{+}99, SH90b, SHV91, SH95b, TDBH94, TW98, TGB99, TMC98, Uib95, VFMFBE96, VL97, VNM97, WBGS97, Wit99, YTF97, ZBK^{+}98, LL98]. deep-living [VL97]. Deep-sea [Ter93, VWD^{+}95, AAW98, BL97, BVV^{+}94, BL94c, BL96, BK92, CA92, CS93, CJ93a, CSSPV98, DDM^{+}99, GSM^{+}97, Gra94, GPU^{+}97, GP90, JDLL95, KW99a, LC95, Mau92, MC97a, Por94, RR99, RSD^{+}91, SRG^{+}99, SH90b, SHV91, SH95b, TMC98, VFMFBE96, WBGS97]. deep-water [BGT91, Car93, CM98a, CS97, DF92, Dem95, HSN93, LK99a, MY98, MC98, RT95, TDBH94, VNM97]. deepening [KD99]. deepest [FKT^{+}99]. deepwater [JSOG91]. deethylase [RBN99]. defaunated [Ber90]. defences [WS92a]. defense [AP97b, KB91, MB97a, PF92a, PKY96, SL97b, VPF98]. Defenses [CP95a, PCTF95, BPS98, BLH99, CHFL95, PLF90, PNS90, RP91a]. Defensive [FP91a, Tro94, AMB99, HFR^{+}93, UTB^{+}95]. deferred [SCD^{+}90]. deficiency [Joh97, RVT^{+}97]. deficient [DO92]. define [MJSDF91]. defodiens [WCP^{+}98]. Degradation [TTL^{+}96, AJST98, BL94c, HFW99, KY97, KG95, LLB^{+}95, LMTF96, NK96, Por94, RR99, SB99, SMB98]. degrading [Cla96, MH90]. degrees [JT97]. DEHP [HFW99]. Delaware [BLJ92, JE95, MBK95, RE94a, RE94b]. Delayed [Har96, McC99, BB95]. delicatula [Mik95]. Delisea [dNSR^{+}96]. Delmarva [MBM92, RGS95]. delpinids [CAATS94]. Delta [FH97a, KDAS92, VM95, BNK^{+}97, HKY^{+}97, FMRC93, OP99]. deltaic [RR93]. demand [FW97, PS97, ZNH93]. demands [DL93a]. Demersal [Bia91, GDvS98, HSLV94, MF98, Bia92, CT92, FII95, HSB97, JBK99, MD91, MN99, SHO96]. demethylate [VDT92]. demissa [KNK90]. Demographic [Gra91b, Pak95, Pfi97, Dur94, KFD99, WS91, Zaj91b]. demographics [JRB96]. demography [BD99, DSJ90a, Duf92, DD97, GMR^{+}94, Gra91c, HJHG98, KS98, SLR94, VL97, VDM^{+}97]. demonstrated [BDdWC96, DS93a, DRS94, SB97b]. demonstration [BWL^{+}99]. demosponge [MGYV97]. demosponges [LBL97, MY96]. Demospongiae [MSCTBE96, RCG99b, UTB^{+}95]. dendritica [Tro92]. Dendrobranchiata [DF92, Dem95]. Dendroceratida [RCG99b]. dendrochirote [SMTL99]. Dendronotacea [MTP92]. Denitrification [BR91a, LKvRH96, NG96, PNB94, YB92, ZB94, BBF94, GSBB97, GSB98, JP93b, KLEH91, LMS^{+}93, MB97b, NNR95, OSGB96, RMTB^{+}95, RCN95, THKN98]. Denmark [AWD95, DFJ^{+}93, HR96b, JLS90, JJNR95, LJS90, NNR95, RH91]. Dense [GP90, ARS93, ADG^{+}95, KJMRC96, MDG98, TA93]. densities [CHFL95, JS97, Jon97a, MGC91, TPHC94]. Density [AU93, AD92, BCBC98, ELH92, FB94, LGFS99, ML91b, Mga92, Rv92, RNHH97, Tay98, AMK99, BSH^{+}94, BD99, CO92, CC98, CK99, CPL^{+}96, CWWH99, CHFL95, DHM92, DE94, FD98, Fer99, FW91a, HT96, Hol90, HN97b, HS99b, IPT94, KWD93, Ken94a, KL91c, KPWL93, LMM98, LCBD94, MCRGC90, MGC95a, MKV95, Mye95, NFM98, OSJ94a, PMB99, PF91a, RS91, RD95, SL90, SR97, SL94, TTL^{+}96, VÅCL99, Web91b, Wor95, HKR96]. Density-dependence [AD92]. Density-dependent [AU93, ELH92, LGFS99, ML91b, Rv92, RNHH97, CPL^{+}96, Hol90, KWD93, MGC95a, Mye95, NFM98]. Density-governing [FB94]. density-independent [Mye95]. dentatus [BB95, MT91]. depend [Lop91, VEW99]. Dependence [FQV98, AD92, AJST98, BD99, DS93a, DV92, GdGMD95, KS91, OH95, PSB^{+}97, SH90b]. dependencies [TGG97]. dependency [HDK93, KE95, ML91c]. dependent [AU93, Ang91a, BCBC98, BEG97, CO92, CPL^{+}96, ELH92, FVDL95, GS91a, GYL93, GJ96b, GB98, HPSJ95, Hob93, Hol90, Hov92, JW96, JNJP94, JC94, KWD93, KS99, LM98b, Lew91, LGFS99, LFBH94, ML91b, MGC95a, MTB^{+}99, Mye95, Nak94, NFM98, RT95, Rv92, RNHH97, TDK97b]. depending [Kog95]. depletion [Arr94, BPR90, BDH94, CPH^{+}96, KSG^{+}92, KHT93, LC90, McG97a, ND96, PGR94, San94a, Tam95, TGvJW95, Ved98]. deployed [WE95]. Deposit [SMB97, CM90a, CGA95, CM99, EE98, GAC^{+}97, HAA94, HP98, Joh97, Kam94a, KM96, PM94b, PS98, PTW^{+}91, SE98, TBY98, WSF99]. deposit- [Kam94a]. deposit-feeder [HP98]. deposit-feeders [CGA95, GAC^{+}97]. deposit-feeding [CM90a, EE98, Joh97, SE98, WSF99]. deposited [FKB92, LJS90, OSZ95]. Deposition [GBV96, AFC96, AV98, BSS92, CJ95, WBGS97]. deposits [SM97b]. depressions [KCI98, SF94]. deprivation [SO96b]. Depth [FII95, GYL93, KPWL93, MGYV97, AHY^{+}99, CS97, DY92, DBMK99, EGMC94, FK95a, GEW97, HMWN97, JPA^{+}96, JJNR95, KBWC90, KW92b, Mac91, MSJ92, McG94, Mil95b, MSM^{+}98, MDRW99, Ole96, PGM^{+}98, SSR99, TAD^{+}94, WH91, WHW93, WPB^{+}93, WD94]. Depth- [GYL93]. Depth-gradient [FII95]. Depth-related [KPWL93, MSM^{+}98]. depths [DGT^{+}93, SE93]. depuration [BFMM99]. derasa [KL94]. derive [TMCE98]. derived [AP97b, BGT91, HKA96, Hol96, JWL96, LS96a, MLP90, MKYI98, TGG^{+}95, TL95]. deriving [LZYW93]. describe [JLF95, Sal95]. Description [BM94b, O’D98, Ken94c, Moe97a, TJ99, ZF99]. desertification [FPB^{+}94]. desiccation [BUWB99]. design [BB97b, HP96, Tis98a, BRNH97]. destabilization [RCH98]. destratification [KKW^{+}93]. destruction [HH92, HH96, RAGH94, SH98b, SEAI96]. Destructive [WH91, Hag95, SKM93]. detect [VWJ96]. detected [RCGSN96]. Detecting [SMF91, LBH91, PKL97]. Detection [GLBF92, GGKJ^{+}98, GCWH90, HCS93, Sew96, VPGV94, CLW99, HHYD92, JT90, SKT97]. detects [MRF^{+}99]. deter [PS93]. detergent [UKR^{+}91]. detergent-polluted [UKR^{+}91]. deterioration [MB95b]. Determinants [BCBR97, Kal91b, SMG97, MS93c]. Determination [AFZ96, BK91, HW92, KZP93, KBP98, MO97, MG99, RSN^{+}99, SM97a, Ber92, BA97, BVP91, BBCB99, COS93, Cle93, Fly90a, Fly90b, GCLS94, HL98, HW99, HPSG94, JSH99, MTB^{+}99, PPR94, RDPB99, RR92, SH90b]. determinations [DFB^{+}99, DMA^{+}94, Ric91, Tho91, WFL96]. determine [CF94, KW97a, KJGC96, TB95]. determined [BvB94, CM99, CD96, CWH98, CRLP94, GTWF98, HH94a, JD90, JBWH97, JP96, KSRB99, MFD98, MWA93, MJ91, MHvB97, RRGB96, VMHH90, YB90, YKK^{+}98, Zel98]. determines [Cli95]. Determining [GU98, Bus96, HlS90, LS94, OWM92, PHHH96, SJC99, dNDS98]. deterministic [BH98]. deterrence [GK97, SL97b]. deterrency [PCTF95]. detoxication [VCMP93]. detoxification [BMFR97, JJTS97, OV92, Vis90, Vis91, VT92, WL97, WJM^{+}99]. detrital [AV95, VSSM98]. Detritus [MR97b, Vet95, AWdS92, Alo90, BVS^{+}91, Hol96, JKL90, PF90, SSP97, SPS98, TA90, VD99]. Detritus-based [Vet95]. devastating [CED91]. developing [DBV98, MS93c, ÓM92, RFGF92]. Development [ATNW97, EPDC99, FKC96, FMH96, HHM96, KGS95, MBL94, Put91, WMH96, WB92b, Aks96, And90, AHMd90, Ang91c, ACOK95, CB92, CN92a, CP90, CRZG95, DSJ90a, Gla99, GQ97a, KW98, KS95b, KBS98, LPI^{+}95, LB99a, Lav97, Lav94, Lop91, LF97, MH93a, MS98, OSZ95, PLP^{+}93, PDC97, PlMM94, PLIM95, QQ97, QGQ97, RT95, SK97, TZ99, UCV98, Weh91, WE95, Wil92, Moe97b]. Developmental [Ang95a, Ang95b, GC95a, WB91, CA97, KKL^{+}98, LH91, SKPT97, SGM98, VV95]. device [HP96]. Devon [MCRGC90]. diadinoxanthin [BBL93]. diagenesis [KGHTN99]. dialysis [MA96]. diamidino [HP98]. diamond [NAS95]. diamond-shaped [NAS95]. diapause [Nor96]. Diarrhetic [SPZ^{+}93, HTLK99]. Diatom [FW91b, STK99, And96, BBC^{+}97, BLKA93, BHT98, BUCP^{+}92, BEG97, BR97, CMB^{+}99, CPL^{+}96, CCG99, CCCH94, CJ95, DLT^{+}97, EST95, FSdM^{+}92, GT92, HM97b, KRJL96, KR98b, KHT93, KKT96b, LRV94, LW91, MH93a, Mik95, MLH96, NB90, NH95b, PRM^{+}96a, PRM^{+}96b, Pel96, RC95, Rie91a, Rie91b, SDH95, SHT91, VDM90, Vil91, WH92, WYY^{+}99, vDdWKW99]. diatom-copepod [BBC^{+}97]. diatom-dominated [DLT^{+}97]. diatom-inhabited [vDdWKW99]. diatom/cyanobacterial [CMB^{+}99]. diatoms [BC90, BOC90, BVL98, BZSG96, CP95b, DQT^{+}97, DZ99, Fry96, GH94b, GKRD96, GHD92, Gol99, GKMT99, HH92, JKT^{+}98, KRJ98, LBM^{+}99, MP96a, Mar99b, MV96, MA96, NGS97, NT92, PGB96, PlMM94, RR97a, SHC96, Som94, TP96, Uye96, WFTH97, WYY^{+}98, WHG^{+}99a, ZD99]. diatoxanthin [BBL93]. diazotrophic [COD^{+}93]. dibromophenol [GK97]. dichotoma [HN90a, HN90b]. Dictyosphaeria [LS96a, LA97a]. Dictyota [CLHF95, HN90a, HN90b]. Didemnum [VPF98]. die-off [Dur94]. Diego [LOS90]. Diel [DJY90, DDW90, ES90, FTS97, GDP^{+}98, GPB^{+}98, Hay95, JPA^{+}96, OK98, PM91b, RWCS98, TK97, TS90b, TP95, WW90, WPB^{+}93, ACC90, BCH95, Bur94, CN95, CC90, CC91b, CCV^{+}98, GRBA96, HWL96, HHH97, Hic92, HB98, JP94, OF94, SCJ99b, STS99, WFHM99, WDD92, vBV95, PCK^{+}96]. Diet [BGC91, BK92, CRP99, DEF91, Gib94, GH91, GMF98, MBR94, PGG96, TK98, WH93, ASA96, ACR99, AP97b, BN95, BH94b, BP98, BBCB99, CWHA99, FD98, FSGG96, FPPM96, GH94c, HWH98, HVL96, Joh99, JC92, KWEG94, LJ96, McC92, Mil92, MGF98, O’B94a, ØW93, PCT^{+}97, RAB^{+}92, RCG99a, RCG99b, SH99a, SPCF97, SPB^{+}99b, SLSH98, TJ99, VF99, WCL97]. diet-derived [AP97b]. Dietary [GCKW95, SDC99, Kre93, KHBL95, LM98a, RP91a, RGT98, WB97]. Diets [BRLM90, CSC92, Car93, TFL95, APCL^{+}93, BMSF95, Kle93, RM96, San93a, TDVF99, WMH96, vdL98]. dievengatensis [AWdS92]. difference [PB98]. Differences [FB91, LFRB97, MWA93, SFPC97, ASRP^{+}96a, ASSWM97, CBI91, CRP99, DD99, GS90, IBB97, JKT90b, JDLL95, KFF92, KJT90, LCG97, LA97b, PC96, RTCH96, Rei90, RT98, RÁSP99, She96, SDFJv91, SW98b, TAJK90, TJRG93, UMD92, VF99, WT99a]. Different [JT97, PWS98, RN93, AP99a, BN95, BCC94, BvB94, CBRD92, Cha98, CCC94, CC96, DBMK99, FG93, GVGG98, GY97, GSMM96, HKG93, HB97, HP96, HK99, Joc95, KW98, KGS95, KM92, KW94, KW99b, Lai92, LMF93, LA97b, MF91, MRL96, MGP^{+}96, MMR99, MDRW99, MH96b, NL96, PWG^{+}99, PCG96, Ric91, RK99, SPB95, Sci91, SDFJv91, SBH^{+}95, SS93, Sog92, TC93, THPC96, TVG96, VV95, VTNA91, VCMP93, WHG^{+}99a, WP94, WMH96, dRAV92, vBV95]. Differential [CP95b, GSS93b, HH99, MvdZH^{+}97, Mun93, NB98, SW99, KW97b, PLG^{+}96, PS98]. differentially [RATA^{+}99]. differentiate [TC93]. Differentiation [SFP91, ASL99, AHS97, GVGG98, HBBT98, JKT90a, KSH97, MR97c, Pez91, Pla93, SRRdM99, Väi90]. differently [TPHC91]. Differing [Reu98a, Hul97, TB99, Whi94]. diffusion [BF99, GF95, Hil91a, HdWS91, LEHvR96, TPR98]. diffusion-sedimentation [GF95]. Diffusional [CRCS96]. diffusive [LGR95, RRWG96]. Digenea [BB97a]. digenetic [CUM98]. Digestion [CTC93, CMT94, GS93, KS99, PS98, SMF91]. Digestive [GSS93a, OC95, CM99, HL90, KCV94, LS94, MRMR92, RMMB91, SM97b]. digging [HBR^{+}91]. digital [MGEC97]. digitale [PGG96]. dilatata [CP90]. dilution [ASP91, LKC95, Sim91, TCB93, vR93a]. dimensional [CCR98, GRGK96, GWSG99, ZGM^{+}96]. dimensions [FD98, GW94, WW96]. dimers [BEG97]. dimethyl [DKL94, SGPAA95, SHML98, WK93]. Dimethylsulfide [LD94b, TK96, CLGM96, KQH^{+}96, OKL^{+}96, SGL^{+}97]. dimethylsulfonio [LD94b]. dimethylsulfonio-propionate [LD94b]. Dimethylsulfoniopropionate [TDVF99, KG95, KQH^{+}96, OKL^{+}96, SGL^{+}97, SWK98, TK96, VDT92]. Dimethylsulphide [KK95, vdBTvDR96]. dimethylsulphoniopropionate [vdBTvDR96]. dimethylsulphoniopropionate- [vdBTvDR96]. dimidiatus [Gru96]. dimorphism [ST99]. DIN [ATNW97, CLC^{+}98, RCS^{+}96, WK92a]. dine [WHW93]. dinner [WPB^{+}93]. dino [GCGR90]. dino-flagellate [GCGR90]. dinoflagellate [AKM^{+}99, AGGKA96, ACC90, BLCA90, BLC91, BBG90, BGH95, Bus97, DH90, Dut98, FFJ93, FFF^{+}94, FJR^{+}94, GGC99, Han91a, Han95, Hin92, JH97, Jeo94b, Jeo94a, JLYK97, JSK^{+}99, KA97, KMR98, KHK^{+}99, LJ96, Les96, MCC97, MUH97, MHT^{+}97, MKV95, NYH92, Nie93, Nie96, PBHMI98, RGLS99, SA94, SLC^{+}97, Str91, UYMH95, UGSBA94, VPGV94]. dinoflagellate-cyanobacterium [KHK^{+}99]. dinoflagellates [ALBS96, BW95, CSE^{+}96, DHM92, DZ99, GAN^{+}94, Hal98, Han91b, HGJ99, IPT94, JL94, JSK^{+}99, JNJP94, LJ96, NSH95, NC95, RP91b, SOS^{+}92, ZD99, PT94a]. Dinophyceae [AKM^{+}99, DRGV91, HN97a, IT96, Sci91, Sko96]. Dinophysis [DHM92, GGC99, Han91a, SBT97, SPZ^{+}93]. Dioaccumulation [CF93a]. Diogenes [Asa91]. Diogenidae [Har96]. dioica [LUA99]. Diomedea [WW92a, GMF98]. Diopatra [PL99b]. dioxide [BCG94, FRP^{+}99, NBM94]. Diplodus [MBF^{+}97, SB97a, VHVB^{+}98]. Diplolaimella [AWdS92]. Diplolaimelloides [AWHH94]. diploproa [SDBW95]. Diplosoma [RB90]. Direct [BWL^{+}99, CH97b, JHN99, ML90, OIO^{+}95, RMTB^{+}95, WM90, JBH98]. Direct-interception [ML90]. Directional [MSC95]. directly [KW91]. disadvantages [Pec99]. disc [GW91]. discard [Oro96]. discards [BMSF95, GH94a, GCF96, SM99a, Str99]. discharge [Gri94, GF91, JCMW91, KDR^{+}92, MH94a, ONR98, SR93a, SLHA97, Sim92, YHGB96, YGH^{+}97]. discharges [DvdM92]. Discobotellina [AC96b]. discontinuities [MCL90]. discontinuity [QZA95]. discontinuous [AB92]. discovered [FKT^{+}99]. Discovery [SOW97]. discriminant [JSGS94]. Discriminating [FL90]. Discrimination [BLKA93, EF97, Joh96, KRJL96, KRJ98, SS93, Tee99]. disease [BSW93, BCBR97, LGP95, NBS^{+}97, Pro92, SH97b, VBL^{+}92]. disequilibria [Kal91a]. disequilibrium [GPD^{+}96]. dislodgement [TJRG93]. dislodgment [BG97]. Dispensation [WH98]. Dispersal [KW91, PF92b, ASL99, AMO^{+}99, BM91a, CTSN90, CC99, CT96, GC95b, Har97, Hil91a, Hun93, Ker90, LB91, LFBH94, Mil98, Pla93, PDS97, YTF97, ZN96a]. Dispersion [HD93, Lav94, Bla90, DLJ91, LFS90, RPN95, SF91, Tok95, VBCRB97]. Displacement [SS99b, CHS^{+}95, SlRR94, NTS90]. disposal [SRW95a]. dissimilatory [MB97b]. dissipation [ML93]. Dissolved [DÁSP97, Fuh90, JKC^{+}93, KW92a, ONM93, SCF91, ZB99a, AMB^{+}96, AB96b, BHQ99, BH94a, BH95a, BPH95, BSD^{+}94, BG91, BG93, CM99, CBD96, CCH93, Col92, CG93, DSG^{+}99, FM95a, FPGC^{+}98, GBA^{+}96, HQ99, HDPW93, Hol96, HG94, Irw91, JTB99, KK91, KD99, KS91, KG95, Kir90, Kro91, LH98, MRL96, MS93a, MBK95, MH90, MWH91, MH94b, NKI^{+}96, NK90, PMOD90, PPCR97, RTLG97, RMG^{+}98, SAH98, SS97, SB94, SGPAA95, SR95, SM95, Sv93b, SKT97, TL95, VB96, WSF99, WP95, ZNH93, KH92a]. distance [Mil98, SC94]. distances [Bro90, KW91]. distichus [Ang91a, Ang91b, Ang92, AD92, AD93, PB96a]. distinct [DBV98, WPY98]. distinctness [CW99b, HG98, SOC97, WC95]. distributed [BA96]. Distribution [ACR99, AMO^{+}99, BMP^{+}95, Bro94, CWH98, CHFL95, GH94a, Góm95, Gon92, GM98b, GF94, GC95b, HK99, HD91, HOM95, Jer90, LGV^{+}99a, LT93, LPW95, Nor92, OMH95a, OMH95b, PCM^{+}95, PA94, PCB90, PM94b, RO93, RPT95, Sab90, SS92, SHO96, SCKS90, TWS^{+}95, TB99, WH95, WCL97, AWD95, ABM96, And94b, AB92, AT96, BA93, Bar97b, BQR99, BCC95, BSJPW93, BD96, BLJ99, BFS^{+}98, BF95b, BFJ^{+}93, BS92, BS95c, BWL^{+}99, BAOT96, BRW95, BGH95, CLGM96, Cha91, CN95, CCG99, CBS91, CPT90, CC97b, CWWH99, CWS^{+}93, DM91, DD90, DJY90, DHKP96, DFR95, DC91, ELH92, EAEP98, FH96, FLDT92, FB90, Fow90a, GM99, GRBA96, GRGK96, GJ96a, GSG^{+}94, GF91, GW97, GEW97, HPPZ93, HMWN97, Har94, HF99, HRÆ94, HC90, HS99b]. distribution [HC94, HCI93, IP94, IFL97, JW96, JWKH98, JBK99, JA91, KMKS96, KW97a, KID91, Kur99, LC91, LRV94, LD94a, Mac91, MH96a, MST98, MY98, Mar97, MJV92, MACR98, McC96b, McG92, MF98, MK93b, MBB99a, MLTD96, MPS95, MB91, MJC97, MLDM99, Nor94, Ole96, OD90, ODS^{+}96, Onu96, PDD^{+}99, PS97, PF95, RLDM94, RK96, RGLD^{+}98, RBA98, Rei90, RPN95, RC91a, RII93, RH97, Röp93, RGT98, SB98, SH90a, SDER96, Sea90, SCL95, SO96b, SPPL92, SH93, SR97, SR96, SGM98, TD92, TB98, TcBW96, TRS91, TMCE98, TS90c, Tro92, TSI^{+}93, VK90, Ven98, Vis97, VDA96, WSCT96, WLT^{+}95, WMD99, WP94, WHG99b, WPY98, ZDJ^{+}95, Zho98, KLMR91, LHL93, Sie93, VHH^{+}96]. Distributional [Fal96, Arr93, BM94d]. Distributions [HHE92, Ken92, PL99b, BM98, BPL95, CSC92, CCR98, FTS97, HTCP96, JWL96, KL95, O’D98, SO96a, SJML99, Sch90, TRMW95, VDM^{+}97]. Disturbance [Hol97, AGW93, Amb93, Aro91, AWVP98, BCC96, BAY91, Bre96, CH96, CWB93a, CLKO98, DMM99, FTGW90, FTW90, KR97a, KD96, Lew97, MJ97a, NL94, OE91, OSC98, PKL97, RS99a, RKH96, THCD95, Tsu90, THR^{+}98, WC93]. disturbance-induced [BAY91]. disturbances [BHH^{+}95, FTP99, HPF92, Mum99, SW99, TPHC91, Vid94]. disturbed [SHA^{+}95a]. Ditylum [KRJ98, WH92]. Diurnal [HLF^{+}91, JWKH98, FFJ93, JMP90]. dive [HGBW99, PZCJ99, WPB^{+}93]. diver [CRLP94]. divergence [Mal99, MTP92]. divergens [Jeo94b, Jeo94a, JLYK97]. Divergent [TN90, CBT98]. diverse [JAB^{+}92, SS99b, VVT97]. diversicolor [AOOT94, NEIR95, Rii91, RVBL92, RPL96, SHC96, VR93b, VAR94, Ved98, Vis90]. diversities [GPU^{+}97]. Diversity [Har97, SHV91, SB96, ASRC99, AB96a, AWVP98, CSSPV98, DAE^{+}99, Gra94, HG98, Hul97, KRJ^{+}95, LHRL97, LD94a, LSG^{+}99, LH99, Mar99a, RHT91, Sch90, SH90b, WA99b, WLB^{+}99]. diving [HGBW99, WFHM99, YOH98]. Divinyl [BAM^{+}99, BCG97, GR93]. division [BZ92]. DMS [KS91, KK96, NKKO97, Sv93b, TK96, vdBTvDR96]. DMS-producing [NKKO97]. DMSP [BLM^{+}95, CBDC96, KS91, KK96, LVL^{+}99, NKKO97, Sv93b, SDG95, SD96, SWK98, TDVF99, TK96, WSS94, vdBTvDR96]. DMSP-lyase [SDG95, SD96]. DNA [ACC90, BSP^{+}99, BD91, BFJ^{+}93, BVP91, CC90, CC91b, CCV^{+}98, Cle93, CD96, DFB^{+}99, DCBC91, FM95b, GCR98, GGC99, GCNSNd94, HBBT98, HW91, JPA^{+}96, JM99, JKC^{+}93, LF91, MHB94, MC91d, RBB^{+}91, SB97b, SRS92, VBZ^{+}94, WDB98, YFI91, vBV95]. DnaJ [KBA^{+}97]. DNAs [CCGM93]. Do [KW97a, Mal98, Ols93, PlMM94, Rin96, Röp93, RA96, She96, TN98, Jua92, MC96b, OS96, RCG94, TPHC91, UMTM98, VHH^{+}96, WW92a]. DOC [KNM93, KE95, MGAE99]. documenting [Bra95]. Does [ARF92, CD96, HvdMBK98, JCH93, KMG98, PS93, San94a, DKVD95, STS99]. Dogger [KR92, NLR^{+}93, RC91b]. dogwhelk [DHB^{+}97]. Dolioletta [Dei90]. dolphins [MCM^{+}97, Rei90, Sha95b]. DOM’ [TSS93, AR99, FM95a, SAH98]. domain [HHP95]. dominance [DQT^{+}97, TDK97a]. dominans [NYH92]. dominant [FML^{+}95, SJMC93, SK97, TTR99, VB93]. dominated [CBG^{+}99, DLT^{+}97, HvB91, HB95, Kea90, KG99, LTM90, NW97a, SCDH98, TPHC91, TR95c, ZB98, ZB99a]. domoic [BLKA93, BUCP^{+}92, MHBW90, PRM^{+}96a, PRM^{+}96b]. domoic-acid-producing [BLKA93]. domuncula [BWL^{+}99, MBL^{+}98, MWB^{+}99]. DON [WK92a]. Donacilla [ROM^{+}91]. Donax [Coc90, DdA95]. Doppler [CSB94]. dorid [BSS92]. doses [RPT95]. dosimeter [Dun99]. Dotilla [ÓN97]. Doto [MTP92]. Dotoidae [MTP92]. double [BD91, BS95b, CF94]. double-crested [BS95b]. double-staining [BD91]. double-tracer [CF94]. down [FAVV98, Gas94, MS96b, PH95]. downstream [KDC99]. downward [VKV95]. downwelling [TFF94]. Dr [Ano94, Bia92, PdG96]. Dr. [Tor94]. drag [Wah96, Wah97]. dragonet [BBT^{+}93]. drainage [HG94]. Drastic [AOT97, OINN94]. drawbacks [GGC99]. dredging [BLST90, CP96, RH91, THCD95]. dredgings [SRW95a]. Drescher [GGD99]. Drift [BH97a, Kin92, Kin95a, KS97, Mol97, RBD^{+}97, SDBW95, SH94]. drifting [Arm92, GM99, MC91b, MD93, NIU^{+}98, NB96b, NB96c, WX97]. drill [DvdM92, KDR^{+}92]. drilling [DvdM92, GAD^{+}92, RS92]. driven [BGW96, FM94, RÁSP99, VBCRB97]. drives [MMD93]. driving [CA99a]. droebachiensis [SH97b, SCL95, SHT91, SLHC98]. drogue [GGB^{+}99]. drum [RH97, SHFP94]. Drupella [McC94a]. dry [Mat92]. DSi [RCS^{+}96]. DSP [SPZ^{+}93]. DTAF [SR98a]. dual [MA99c, RRGB96]. ducks [BL94a, YOH98]. due [GMC91, KV99, LGB91, PND^{+}94]. dugong [NA99, Pre95]. dule [BGC94]. dumerilii [ZHBH90]. dumping [MR91, RDM91]. Dunalielia [HPSG94]. Dungeness [EAEP98, FIA93, Fer99, MBL94, SLSH98]. dunlins [MJ92a]. duorarum [MF95]. duration [Cow91, GCM98, GAV93, MT93]. during [ASRC99, ABB^{+}97, AMB^{+}96, BH92a, BL94b, BC90, BMD^{+}90, BN91b, BM91b, BGG^{+}97, Bra93, BW96, BG93, BK96b, CCG98, CCV^{+}96, CBG^{+}99, CB92, CN92a, Car90, CP90, CCCH94, DG91b, DG96, FFF^{+}94, FPFR92, FPPM96, FPP^{+}97, GB91, GCKW95, GGB^{+}99, GTL^{+}91, GST^{+}91, GGPGG95, GGB94, Gos97, GAB90a, GAB90b, GKB^{+}97, HLP^{+}99a, HKK^{+}95, HSCC96, HHM97, HNL94, IHK^{+}98, Jon97a, Jos90, KNM93, KY97, KKW^{+}93, KR98b, LED92, LH91, LL97, LHL93, LK99b, MG96a, MFD98, MYYT96, MBJ94, MC91b, MM95a, MF96, MHRV^{+}99, Mol97, MPS95, Naj97, NSHF93, NSH95, NGS97, NSC^{+}92, NKH^{+}99, NLR^{+}93, NH95b, NNB93, NK92b, ORL97, PLP^{+}93, PP96, PB92, PDC97, PRW^{+}94, PWD^{+}91, PF91a, RBD^{+}97, RC95, Rie91a, Rie91b, Rie93, RFN98, RN98, RPM^{+}94, RDE^{+}90a, RDE^{+}90b, SMUS^{+}99]. during [SGL^{+}97, SJJMH90, SCEG93, SJÆ91, SMH97, SBP92, SBP93, SGP94, SMB98, ST98, SCKS90, SE93, TLB93, Tak96, Tam95, TR91, TRPS94, TMCE98, TP95, VHG95, WH92, WSCT96, WYY^{+}98, WYY^{+}99, WPTB96, WSP^{+}95, WA95, WSS94, YHC^{+}99, ZBK^{+}98, TBAB91]. Durvillaea [SS95, WMG^{+}94]. Dutch [BLDF95a, GAD^{+}92, HvB91, PGB96, SDG95, vDK90, vBDW91, vW93]. dwarf [CJL91]. dwelling [Asa91, Duf92, GF98, MJC97, PD90, PD94, RGLS99, RW99, Ste95]. dye [ER95b, JM99, SR98a]. Dynamic [GMK97, MFD98, Mil94a, NBW94, LP95, SBT97, VNB97]. dynamical [BBRBK94]. Dynamics [AFP99, Beu92, BKK^{+}93, CLB91, FPP^{+}97, GCF^{+}95, GD90, GBB^{+}98, LB96a, MM91b, NSHF93, PLP^{+}93, PDP^{+}90, RHJ92, VDM^{+}97, dC94, AJS95, ADR95, AWHH94, Ang91b, AD93, ALBS96, BC91, Blu93, BHT98, BH98, Cal95a, Cal95b, CPSA96, CM97, CRH^{+}99, DM91, Dee90, DMP^{+}99, DMA^{+}94, Dur94, EJ93, FW92, FS90, FGB^{+}96, FSG^{+}97, GMMD93, GNJ90, GBA^{+}96, Góm95, GHL^{+}97, GBGH97, HM97a, HCC^{+}91, HKY^{+}97, HGS94, HT96, HSCC96, HN97b, HS95, IT96, JFS92, KLMR91, KM92, KKW^{+}93, KSO^{+}97, Kuo91, KYN94, Laf93, LGB91, LAG96, LD96, LA98, LFTH92, LL98, LGFS99, LUA99, LGMD98, MBM92, MBJ94, MDC^{+}96, MW99, MS94a, MR97b, MCRGC90, MGC91, MGC95a, MO97, Mol97, MSD92, MZCMOCJP91, MJ92b, MJ93, NSH95, ND98, Neu92, O’B94b, OSJ94b, OFR94, PB92, PBDF98, PS92]. dynamics [Pir92, RBD^{+}97, RDT^{+}94, RS99a, RGLD^{+}98, RSH93, RWCM99, RAlBW99, RCC^{+}99, SMUS^{+}99, SBBG94, San94c, San94b, SWT99, SRJD97, SR90b, Sch96, SR93a, SAA90, SLR94, ST90, SCDH98, SLHC98, TGG^{+}95, TH90, TR95c, TS99, TH94b, UHL^{+}97, UC94, VDM90, Vid94, VM95, VNM97, WSCT96, WSMM99, WP95, WS92b, YGH^{+}97, YST^{+}97, ZH97, vBDW91]. Dysidea [RCG99b].

E. [Nor94]. Early [EDP^{+}98, GQ96, HM94a, KW98, OTF95, RSH92, RM95, lYALM93, BL94b, CA97, CH95, CC91a, DY92, DLDL93, DRC^{+}97, GC95b, HHM96, HS97c, IW91, Jeo94a, KGHTN99, KY97, KW99b, LTM^{+}98, MCNB98, MS95, OW95, PWS98, QQ97, Ruz96, SW98b, TLT93, WW96, Weh91, ZAN99]. Early-spring [HM94a]. East [KDHZ98, COK99, CGP97, Dye93, GMF98, Ken92, RHH97, Spe95, ALBS96, CCG99, DJY90, DLJ91, HTLK99, LAG96, MP96b, RWJ99a, RAlBW99]. Eastern [DMP^{+}99, HSH94, LC91, BN91a, BBL93, BA96, CBD96, CVD99, CMA95, FP96a, GCGR90, HW95, HCI93, HOM95, KSS90, Kea90, LJ95, LZYW93, LL98, Liv93, MCF90, MHBW90, MG91b, Mun93, OMH95a, OMH95b, PT95, PF96, Rei90, RT92, Sal96, SPZ^{+}93, TDW96, TRU91, TRS91, YD90, YBLL96, GMC91, HB95, MJW96, ZR92]. eat [CGFC95]. eating [McC94a]. ebb [ZN96a]. Ebro [VM95, OP99]. eburnea [NR98]. echinoderm [PF92b]. Echinodermata [BM91a, CJ93b, DEF91, FB97, FG93, Geo94, GSP99, Lam98, LB99a, LJSK96, LGL^{+}95, MDG98, RO93, RNHH97, SMTL99, SCL95, TLW^{+}96, TGLP95, WTF95]. echinoderms [McC94b, SB94]. echinoid [Bak90, PF95, Ste95, WLT^{+}95, WTF95, WTF96]. Echinoida [SLHC98]. Echinoidea [CK98, FB97, FG93, Lam98, LB99a, LTM^{+}98, LGL^{+}95, RO93, TGLP95]. echinoids [YTF97]. Echinomermella [SCL95, SLHC98]. Echinopora [FD99]. echiuran [HAA94]. echmocephalus [KYN94]. echo [MA99c]. Ecklonia [ACLS97, BWB98, SS93, Ste95]. Ecktoma [Ste94]. eco [Fri97, Kin97]. eco-ethics [Fri97, Kin97]. Ecological [CAATS94, CU99, DFS94, HC93, HR96b, LS92, MSND^{+}99, RH91, SHM^{+}97, AB92, BGC91, BSMS95, BM99a, BLDF95b, BLDF95a, BMHH94, BSB91, BM90a, BM90b, BKWML93, BLST90, CPB96, DHS96, Gos97, HM96, KH98, Kin93, LH99, MU99, PN99, PAM96, QZA95, RC95, RBG94b, She94, SL94, UMT^{+}91, VGM94, WC92, tor93]. ecologically [CBT98]. ecologist [DMA^{+}94]. ecologists [MR94]. Ecology [AC96b, Bar97b, Bar97c, Bar97a, Bar99, BH92c, BH92d, DAB^{+}91, DFJ^{+}93, KH92a, MT94, OB94c, Sal96, And94a, ACOK95, ALB^{+}96, AH95, Asa91, AP97b, BPR90, CW93, CGZ95, EB96, EAT95, FGG95, GRC^{+}97, GC93, GCGS90, GFPD92, GTYR97, GMF98, HTS92, HK96, HFGT94, HCH^{+}96, HRCW98, KR97b, KL94, McL96, MBB95, NKGK92, OJR96, PB96a, PIC97, RD97, SFB92, THS94, TP97, WC98, WC92, YD90, Zaj91a, Zaj91b]. ecomorphological [DSS^{+}98]. Ecophysiological [EBB95, GVGG98, KHT93, TMC98, ZF99]. Ecophysiology [Leh96, BCR92, DM96]. Ecosystem [DKW96, HSLV94, PPFP98, SAG92, ZB98, ARS93, ADG^{+}95, BU93, BNK^{+}97, BBL^{+}95, BW95, BP92, CCG98, CCG96, CM93, DSV^{+}91, FMFH97, GPDF93, GGB^{+}99, JAB^{+}92, KD99, KSMDB97, KZK^{+}95, Kro91, MvdVD^{+}97, MR97b, MCA90, MDG98, MG94, MU97, MWH91, ODN^{+}95, PG91, PIC97, PSR^{+}99, RDT^{+}94, RM94, RNH99, SB97a, SCB92, San97, SD99, STNI97, SSGD96, SSG96, SE91, VPFL92, ZF99, ZT90, SDFJv91]. ecosystems [Bow97, CNGG96, DLT^{+}97, DQT^{+}97, GPF95, Gri94, HS98b, PSK98, She94, TH94a, TS90a, VS96]. ecotoxicological [DB96a]. Ectocarpus [Mül91]. Ectoenzymatic [RSH93]. ectoenzyme [CK95]. ectoparasite [Gru95]. ectoparasites [Gru94]. ectoparasitic [GMC91]. ectoprocts [RM97b, RG97]. ectoproteolytic [LBV^{+}99, TGBB97]. ectosymbiotic [HBB^{+}99]. Edaphic [SM90]. Eddy [CB99, FP96b, PWPK95, TT99, ZR92, ZBK^{+}98]. edge [Bia91, MPS95, NT92, PGR94, RBF92, WJM^{+}99]. edifice [SJ99]. edifices [SJML99]. Editorial [Ano95, Ano97c]. edule [AR91, KSRB99, LAJ95, NW97b, RII93]. edulis [BSJPW93, BSJ97, Bou90a, Bou90b, BLCA90, Bro93, CR96, CH99, GS90, HSBH96, Hil97, HDB95, HBD95, HMB93, HS96, HS98c, JKT90b, KJT90, KHBL95, KCV94, LS94, MCRGC90, MGC91, MGC95a, McG97a, ND91, NHL99, PLBP96, RVT^{+}97, RCG94, Reu98a, RSN90, RH90b, SUH95, Sal95, SZA94, SV97, SP99c, TAJK90, TN90, VB96, WW91, WW93, WFL95, WFL96, WF97, WDB^{+}95]. Edward [MRP98, PDBA90, Per91]. edwardsii [CB99, Mac91]. eel [AOT97, CT96, KDC99, ONM93, OINN94, TSW97, WT98]. eelgrass [AZKA99, BSW93, BDM96, BGC94, FKC^{+}90, FMFH97, GNJ90, HHN^{+}96, HN97b, JDSB99, KA93, MRPC98, MNO96, OSJ93, OSJ94b, Ole96, PB92, PBGE91, RCG94, Reu98a, RBSO99, RPDR^{+}98, TS99, VSSM98, Wor95, ZA96, vKVSR97]. eelpout [FRK92, MGM96, MRS96]. eels [AHY^{+}99, ALOT99, AALT99, OINN94, RBN99]. Effect [ABM96, AAW98, BTYO94, BFD91, BKB92, BL96, BBS93, CC97b, EW91, ERC95, FGR97, FPGJ99, FRP^{+}99, GH93, GH94b, GHTH94, GRSV98, Han97, HKvB93, HS91a, JR92, KM96, KBS98, KCV94, KRF^{+}98, LM98a, LJ96, LL92, LMTF96, MCA90, MM97b, MZCMOCJP91, OBL97, OS96, PMP^{+}96, PGM^{+}98, PTW^{+}91, PB98, RK96, RBB^{+}91, RMG^{+}98, RO98, SK99, SP93, TH94a, TT98, TUN99, VB90, VWJ96, VH99b, WFHM99, WTF95, WO94, Web91a, Web91b, WP95, WE95, WDD92, ZA96, van94a, BC93, BGR^{+}96, BMFR97, BFMM99, BLD94, CC98, CV91, FFT98, GPF95, GB90, HTCP96, Joh96, KA97, KI97, LS96a, MG96a, MFF97, MM91a, MD98b, MF98, NKI^{+}96, NPB93, Nie93, OW95, PM94a, QK93, RR97b, RQB98b, STNI97, Ved98, VM98a, VBN95, Vos91, WLL98]. effective [RHS99]. effectiveness [GC95b]. Effects [AKL^{+}99, lYALM93, AdN91, Amb93, AHMd90, Ang91c, AS94b, AS95, AWVP98, BTKK99, BPR90, BP97a, Bee97, BH94a, BH95a, BCC92, BB95, BI99, BD99, BSD^{+}94, Bro93, BP92, BS96b, BRN^{+}95, BH95b, CA96a, CHS^{+}95, CNCS93, CD94, CPH91, CEV97, Con98, Con95, CD99, CLHF95, CP96, Dee90, DdA95, Dit96, DE94, DH92, FD98, FAED97, FSA^{+}97, FK95b, FGKP95, FDR^{+}96, FAHW99, GBA^{+}96, GR94, GT96, Gri94, GP91, HMWN97, HGS94, HDB95, Hol90, HS99b, Hun93, JHG99, JLR90, Kam94b, KSS90, KNK90, Ken94a, KJLH96, KD96, KRJL96, KDR^{+}92, Kur99, Laf93, LA97a, LK94, LF94, Lev91, Lew97, LCTT92, LF97, MSD98, MSC95, MY96, MD93, MA99b, McC96b, MB93, MNK^{+}97, MG91b, MBB99b, MAAT98, MLH96, MH96b, NT94, NCGG96, ND98, NL94, NKB90, NH98, NDG^{+}99, OE91]. Effects [ÓM92, OD90, Oro96, OC93b, PDD^{+}99, PG91, PRM^{+}96a, PRM^{+}96b, PGB91, PCLÅ97, PPP^{+}96, PCBA94, PMG^{+}98, PD93, PZ91, Pur92, QGQ97, RBA98, RAS92, Rin92, RCH98, Ris97, RMT91, RD99, SGL^{+}97, SOW97, SW98a, SDER96, SA94, SR93a, SLHA97, SU93, SS93, Smi95, SS95, SFB92, SD92, SS91, SRW95b, TNP97, TGB99, THR^{+}98, TB97, VDG95, VTNA91, VDJA98, WH96, WSMM99, WRBF97, Wor95, YNYK94, YKL98b, ZFFH^{+}94, vKSGvA99, vLTW^{+}98, AMS99, ASCW90, BM94b, BS94, BSH^{+}94, Bec98, BL94b, BWB98, BSG^{+}97, BH97b, BB93, BM95b, BFVG92, BBT^{+}93, BGC94, CUM98, CH95, Cha91, CK99, CWWH99, CK98, CSSPV98, CHC94, CRH^{+}99, Cro98, DBHM99, DP91, DHB^{+}97, DH90, DB96b, ELG97, EGMC94, Eml90, Eno93, ER90, Fer99, FFR94, FTBR97, FH97a, GC95a, Gla99, GB93]. effects [GCWH90, Gra91b, GW97, GS99b, HN97a, HH96, Iri90, JHN99, Jeo94b, JS93, JKFB95, JBH98, JCMW91, Jos90, KWD93, Kal91a, KDR97, KR95, LLB92, LWL98, Leo99, LL96, LC90, LFS90, LGFS99, LCK^{+}96, ML91a, MPM97, MT91, MEK99, MR97a, MD98b, MCNB98, Mat92, MUH97, MSK95, McL96, MS96a, MP92, MBK95, Mil95b, MM96, MLES94, MJ92b, NC90, NP90, OM99a, OG95, PKY96, Pen90a, PSR^{+}99, PWNW99, QRSC99, QQ97, QQ98, QQ99, RK95b, RM97a, Reu98a, RW96, RI90, RVBL92, RTJ^{+}93, RWJ99b, RD95, SKM93, SM94, San94c, SDC99, SW99, SL90, SM99a, SLHC98, SE91, TDER97, TM97, TPHC94, TA90, Tol98, Tru97, VÅCL99, WTA^{+}97, WFL95, WA99b, WCT95, YAG92, YKL98a, ZDJ^{+}95, vdBTvDR96, BSH^{+}91b, DZ99]. efficacy [Hal98]. efficiencies [CGA95, SPS^{+}91, vdL98]. efficiency [BK91, BAM^{+}99, GGF99, HK97, HL90, HPSG94, JT90, KBP98, LWK^{+}95, PM94b, SDB92, TT90, WC92]. effluent [LFG98, MM91b]. efflux [BHHL96, HS94a]. effort [Åbe96, CZG95]. Egg [Hav93, Ito97, PHT96, RCTH99, UCV98, WR92a, AN97, CA96a, CA99a, CH97a, CP90, CMT94, Dut98, HvdMBK98, JFP95, Mye95, NG99, NKV94, Oro96, PHHH96, Raw96, RD91, SCIA99, SKvB98, SGM98]. egg-carrying [AN97, RD91]. Egg-size [Ito97]. eggs [BBK91, BSH^{+}91b, CPL^{+}96, CMT94, DL91a, DL93a, FTS94, GRSV98, HOM95, Jeo94a, KVK98, LP95, LGED98, MH96a, MJW96, MLC97, MB97a, PND^{+}94, RD91, RFGF92, Sty97]. Eilat [KL91c, KEvN95]. Ekofisk [GCWH90]. Elbe [KE95, KY97]. Eldfisk [GCWH90]. electrochemical [AMMG91]. Electron [CCC90a, LHH96, Rel96a, Rel96b]. Electronic [Ano99c]. electrophoresis [Bro90]. elegans [BHQ99, BQKC97, GQ97a, HQ99, LQ97, MRP^{+}99, QQ97, QQ98, Ter93, TRPS94, TNN93]. element [MRF^{+}99, WFL96, WF97, WD99]. Elemental [CGM95, HNL94, BR97, LH91]. Elements [GK96a, CAATS94, HM97c, LF92, LF94, MR97b, MH97, Pro92, WFL95, ZCL97]. elephant [BBCB99, HD93]. elevated [BAOT96, CC97a, HGS95, JCH93, MA99b]. elevations [HH99]. elongate [AdN91]. elongation [MD98a]. elongatus [KGS95, KBS98]. elucidate [GK96a]. elucidating [GMMD93, MDC^{+}96]. elutriate [Tha92a, Tha92b]. elutriates [Kar92, Wil92]. elvers [WT98]. embayment [DÁSP97, HA90, KCI98, MH94a, Ole96, RÁSP99, SS99a, TS90b]. Embley [HVL96, STBR99, VS92, VHHS96]. embryo [BH94a, BCv^{+}92, Tha92b]. embryo-genesis [BH94a]. embryo-larval [BCv^{+}92]. embryogenesis [BH95a]. embryonic [CB92, LPI^{+}95, PlMM94, PLIM95]. embryos [BSS92, CP90, DE91, HBD95, MM99, MB97a, Pen90b, SB94, SE98]. Emergence [RII93, MBB99b, WB94]. Emerita [AS99b]. emersion [TUN99]. emigration [lYALM93]. Emiliania [BFF96, BEH93, BLM^{+}95, BvBBV96, FBMH94, MG97, MH96b, NT96a, NGS97, NBM94, SWK98, VKV95, VH99b, WSS94, vBV95]. emission [Maz95, RNHO98]. emperor [KR97b, WR97]. emphasis [Gar99a, KS95c, SSGD96, WD90]. empirical [BJTB96, HS97b, ML93, MRP98, Mee99, Sca96, TD94]. emption [BCMC99]. Ems [dC94]. enclosed [CDGF^{+}93, FSdM^{+}92, LJM99, MMM^{+}95, MCA90]. enclosure [QK93]. enclosures [CH90, DE94, Hin92, KLAH^{+}94, RSB^{+}90, WE95, vBV95]. encoding [KWHM^{+}96]. encounter [JCH93, ML91a]. encrasicolus [CCS98, GCR98, Pal91, SKPT97, TP95, TP97]. encrusting [KW97a, MU98, MU99, NT94, RSAD98, UTB^{+}95, UMTM98]. Encystment [NU91, AAW98]. endemic [PF95]. endemism [SNC^{+}98]. endobenthic [CB97, TTW97]. endoevaporitic [OKK95]. Endogenous [KKYH97, ZN96a, ZAN99, Bus96, ZN96b]. endoliths [LGH95a]. endolymph [Kal91b]. endoparasite [CPS91, SCL95]. endophytic [EP97]. Endoscopic [WMB99, WR98]. endosymbionts [DAE^{+}99, LC95, RGLS99, Sch91]. endosymbioses [SFWD90]. endosymbiotic [GCGS90]. Energetic [Bus98, GMC91, Sto97, CRGZ98, MBR94, TWWE95]. Energetics [TT90, CT98, LFRB97, LK99b, NT96b, OFR94, PCG96, PST95, TR95b]. Energy [AAS^{+}91, GSC95, YMD94, BBB^{+}94, CP90, CM93, DvCd94, HC92b, KFF92, KHBL95, LMM98, ML93, MF95, PWPK95, RI90, RTJ^{+}93, RBT96, Rin96, RFGF92, RMG^{+}98, RDE^{+}90a, RDE^{+}90b, Sch91, SO96b, SDB94, ST98, VDVC94, YKL98a, YKL98b]. energy-conserving [SO96b]. England [AML95, LLB92, FCH99, GH94b, GH94c, HG94, NT96c, She98, TNSM98, YBG96, YBG^{+}98]. English [BBL93, HM97c, MH97, TDW96]. Engraulis [CCS98, GCR98, Pal91, SKPT97, TP95, TP97]. enhance [KMG98, SK98b]. Enhanced [lYAML93, CLB92, RTLG97, EJ96, FL92, HRKR93, JL96, MCFS96, Sch99]. enhancement [PLM92, SK98a, Sch93]. enhances [SAH98]. enriched [Cra94, DO92, RATA^{+}99]. Enrichment [NIU^{+}98, Vet96a, BHG94, BMV^{+}96, BGC94, Fra95, GP91, HS97a, Jos90, KR98a, LL92, Mal96, NK93, SW98a, SG96, TvLdJ^{+}98, Tsu90, TFFS90, VD99, WSMM99, Wes90a, Wes90b, WCLB99]. enrichments [BL96, PF92c, VB90]. Ensis [GCM98]. ENSO [GGPGG95, LMT99]. Entacmaea [FGH95]. Enteromorpha [BSMS95, FDZ94]. enteropneust [CM90a]. Enteropneusta [Mil92]. enteropneusts [GK97]. entomon [EB92, HS90, HS94a, San94a, Vis91, VH96]. entrance [KHK^{+}99]. Enumeration [ER95a, EAC^{+}97, VDT92, HP98]. environment [AGW93, Atk96, BL97, BCA^{+}99, Bra93, BDNS99, EM93, FB97, FP96a, GCR98, HM97b, JP94, Kin95a, KZP93, KQH^{+}96, LP95, MBWT98, MHvB97, MG94, MBL94, MBB99b, NT95, NT96b, OK98, PF96, RWJ^{+}95, RHB96, SH90b, TMKL97, VWD^{+}95, WWC^{+}99]. Environmental [CC91a, DvdTCC97, DHM92, HS94b, HH95, KID91, LD94a, LGH^{+}93, PHHH96, AGP98, BMFR97, BFMM99, BP92, BLST90, BDA^{+}99, BAOT96, BGH95, Byt90, CL90, CMBC98, CA93, CKP98, CSSPV98, CSM99, Dee90, DHKP96, DFB96, DKVD95, EBM93, FCC91, FNJ^{+}94, FDR^{+}96, GB91, GAD^{+}92, GCKW95, GC95a, GU98, GT92, GW97, GT98, JW96, Joh96, JHG99, KV90, KWL98, KCV94, LD93, Lim95, LW94, LDKR93, Mar97, MIA90, MGC91, MGC95a, MDTB94, MTB^{+}99, Mil94b, MLTD96, MN99, NCGG96, ND98, RBF92, RAS97, RZLB96, SP99c, TRU91, Tsu90, TSW97, UCV98, VHHS96, VHH^{+}98, VBZ^{+}94, WA99a, WW91, WCBL94, YBG96, YBG^{+}98]. Environmentally [GHMC96, ABY97]. environments [BYO95, BK96a, BDdWC96, CWS^{+}93, Cra94, HSBH96, HB91b, JTPP92, KFPR94, KRJ^{+}95, KS95c, ML93, MCFS96, MH94b, ODN^{+}95, RGS95, SK95a, SCIA99, Sim92, SKT97, VB96, Wor95, vR93a]. enzymatic [AGP98, BL94c, CR94, CDGF^{+}93, HDK93, PH95, TMC98]. enzyme [CV91, FCW98, GSS93a, HL90, MRMR92, RMMB91, VD94, VL97]. enzymes [BPH95, DE91, LACM92, RMGB^{+}94, VT92]. eperlanus [Sep94]. ephyrae [OPS96]. epialgal [HH95]. Epibenthic [MP96b, AB90, CHDW92, Gra91a, HB97, HHW90, KRM94, LB99b, PS97]. epibenthos [SOSV95]. Epibionts [GAV93, MC96b, Wah96]. epibiosis [WS92a, WJF94, WM99]. epibiotic [Gla99, WW92b]. epidemiological [LK93]. Epidemiology [OR95]. epifauna [AB92, MS94a, PSR^{+}99, SFV96, TC94]. epifaunal [DY92, DPM96, OW95]. epilithic [HS97a, KM92, NW97a]. Epinephelus [Egg95, LCW99, SMW^{+}93]. epipelagic [Sea90, Ter93]. epipelic [SHC96]. epiphyte [FF97, LM91a, TL91]. epiphytes [Tro93, WCLB99]. epiphytic [DC91, MSD92, TB99, WSMM99, WP92]. Epiphytism [DC91, MC93]. epipsammon [DB96a]. equalis [DS93a]. equation [AFZ96]. equatorial [FF92b, LKC95, PLH91, PHL92, RAS97, RAS92]. Eretmochelys [HS91c]. Eriocheir [Ang91c]. EROD [BB97b, EGKE92, RBN99]. erosion [MD94a]. Erratum [Ano00]. error [CC90]. erythraeum [CSM^{+}98]. Erythropodium [FP91a]. escape [VR98, WB92b, ZFFH^{+}94]. escapes [MD98b]. esculentus [CLD99, CM97, LWK^{+}95, O’B94a, O’B94b]. Essex [RNHO98]. established [OSC97]. Establishment [MWB^{+}99, BA91]. ester [PDS97]. esters [HKG93]. estimate [CH97b, GGC99, KB94, SR97]. estimated [AHY^{+}99, CCV^{+}96, MA96, NNR95, TBLD90]. Estimates [Cer98, CT92, CO92, CDGF^{+}93, Col98, EP92, Fre97, JMD95, KW99a, LB99a, LC93, MC98, MS93d, PS97, RSD^{+}91, SR98b]. Estimating [BCAAC96, COK99, COG97, GM98a, LWK^{+}95, MLP90, Spr93, vRV96, BBO^{+}95, GV96, Kir92, KKT96b, ML93, MMHW96, NH90, RM94, Sca96]. Estimation [CCV^{+}98, Fly98, GMO95, HKM92, KPS97, MHC92, CGNH98, CB98, CJ93a, HCS93, LT97, LCC97, MC97a, MT93, MGAE99, SH98b, Str99, TN90]. estimations [DGCT93, LHRL97, SM94]. estuanne [KS91]. estuaries [BU93, DKW96, GH94b, LGED98, LKC^{+}99, LKM^{+}99, MV98, MRP98, Mee99, MCFB97, MU97, NO90, ONH^{+}97, Olm94, RNHO98, RHP93, She95, She96, VHH^{+}98, WW92b, Whi94, YB92]. Estuarine [CM98b, EAEP98, NNR95, AB90, ASRP^{+}96b, ASRP^{+}96a, AN97, AT96, APT99, BSD^{+}94, BG93, BPC98, CK90, CMD^{+}94, CPK97, CT96, CS99, CB96, DSW^{+}91, Dee90, DG97, DON^{+}95, FBL95, FFT98, FCVS92, FFR94, Fuh90, GV96, GK96a, Gil97, GU98, GDW90, GP91, HKvB93, Hic92, HdWS91, Hsi97, HH99, ICG96, JAOS90, JCMW91, Kne93, LMT96, LBK90, MRF^{+}99, MPRB93, MGC91, MF90a, MF90b, MHvB97, Mic96, MR97c, MKN^{+}96, NO90, PB96b, PB97, PNB94, Pel96, PT95, RSB^{+}90, Rie98, RWJ99b, RCN95, Sch90, SDC99, SW98a, SS97, SSM91, SJCL93, SD93a, Ski94, Sog92, SKM95, SMP97, TUN99, TJST98, WD90, WD94, WK93, YOH98, YSL94, ZF99]. estuarme [CWS^{+}93, FW92]. Estuary [BPC98, CBS91, Ken94b, KP95, LD93, LL98, MP94, MGM96, PPFP98, RPMB91, AMS99, ABM96, AALT99, BCS93, BW93, CLC^{+}98, CCSN93, CP91, DD90, ERC95, GJ97a, GJ97b, GH94c, HHK^{+}98, Hei92a, HHW90, HK93, HJA99, HN91, IP94, JJ99, JBB92, JSJMH90, KGHTN99, KE95, KY97, KR98a, KKW^{+}93, LDKR93, LE90, LH99, MC96a, MG96a, MCRGC90, MEC92, MAH94, MBK95, MAAT98, NT96c, ONH^{+}97, OL99, PS94, PSE97, PSPD96, QCM97, RBF92, RLC92, RNHO98, STBR99, SWT99, Sew96, SV93a, SH94, SGW94, SGB90, TA95, TGG^{+}95, TL94, TNSM98, UCV98, WK92b, WT99b, dC94, BMG99, DG91b, DG92b, FPLH92, HV91, LD94a, PPFP99, PR93a, RBN99, RCGSN96, SGL^{+}97, VPFL92, VBF94]. ethanol [HKM92]. ethanol-preserved [HKM92]. Ethics [Kin97, Fri97]. ethoxyresorufin [RBN99]. Ethyl [Ack97]. ethylhexyl [HFW99]. ETS [CCC90a, Rel96a, Rel96b]. Eubalaena [HKMK98]. eucaryotes [HSRHH90]. Euchaeta [Øre91, Øre95]. Eudyptes [HTCW96]. euhaline [dG92a]. eukaryotic [CS91, JM99]. Eukrohnia [Øre90, Øre95, TNN93]. eulamellibranch [BDB97]. Eulerian [FGB^{+}96]. Eunicella [WV95]. Euphausia [AS97, AWH^{+}99, BGC91, BMHD93, Buc91, CWHA99, DM91, DWSA98, Góm95, HTCW96, HNL94, KH98, MACR98, NSC^{+}92, Nor94, Pak95, PPMF97, PWPS95, SL95, Tan98a, WMD99]. Euphausiacea [BGC91, Gen92, Lav94, Nor92]. Euphausiid [BK98, CPS91, CSB94, DWSA98, GGPGG95, GGDLE96, KH98, MVA99]. euphausiids [ASA96, Góm95, HCS93]. euphausnds [TLT93]. euphotic [CA90, MHRV^{+}99, RLC92]. europaea [GH93]. Europe [RDT^{+}94, RATA^{+}99]. European [CPT90, AW97, BCv^{+}92, CMD^{+}94, PPO^{+}99, RBN99, Sep94, SSW^{+}96, SCJ99b, TN90, TP97, TSW97, VF99, WFHM99, Wol92b, Wol93, vCRS92]. euryhaline [BL95]. Eurytemora [AJS95, AN97, HN91, ICG96, KEV99]. Eurythenes [CPT90, TW98]. Eurytium [LK94]. Euterpina [GSMM96]. eutrophic [HJA99, SJJMH90, SE93, US95, US98, YOH98]. eutrophication [Ber90, BC90, CSS93, DKW96, GTJ92, GS99b, KDMR97, McG92, Mee99, Nai93b, SO91, SS97, TNGB95, Tom91, vdBTvDR96]. eutrophying [PPFP98]. euxinus [YSY^{+}99]. evacuation [SCJ99a, TT90, WDD92]. evaluated [JKT90b, KJT90, TAJK90]. Evaluating [TGG^{+}95, Tho92, Cer98, CMBC98, SAG92]. Evaluation [BLST90, FL92, HS99a, KG95, MMBT97, PJF93, RC91b, RD92, WMG91, BS92, BBT^{+}93, CCC^{+}90b, ER95b, FDZ94, JM93, LT97, LCBD94, MC91d, PPB90, Put91, Sch93, SJC99, TJ99, WK92b, CC90, KB94]. evanescens [WC96]. Evechinus [CK98, Lam98, LB99a]. event [FHA99, Jon97a, LCC97, Naj97, SML97, SS99b, TBAB91]. events [BM94a, JP93b, LMT99, RAS97, SH97b, WBLM95]. events9 [CED91]. everta [CGZ95]. Evidence [BA91, BMD^{+}90, BS95b, CP91, DG97, GBC92, HDK93, JDLL95, MB97b, MRP^{+}99, Nov90, OSGB96, VMS^{+}96, ADF96, AWdS92, AS94a, AFZ96, BR91a, BKWML93, CR96, Con98, CRH^{+}99, Cro99, DBvA91, FHCC98, GSP99, GA97a, HRKR93, HTCW96, HHG99, HJHG98, JAL91, JPZ^{+}98, LS91, MPS96, MHT^{+}97, MWH91, MAdG^{+}96, MSCTBE96, NT92, ND96, NAW99, OJ97, OJ99, PBA^{+}95, PR93a, PSB^{+}97, Pre95, Sch90, SH97b, SCLS94, SRRdM99, SM90, TTR99, TDK97a, TvLdJ^{+}98, Väi90, VH99a, VDM90, VD95, WM99, WBP^{+}92, WC90, vRBVB95, NO90, RMMB91, RA96]. evolution [ASRP^{+}96a, DLT^{+}97, Hav93, Mar99a, Zaj91b]. evolutionary [HC93, SC92a]. examination [KDR97, PM94a, SOC97, WASSM97, Wes90a, Wes90b]. examined [CCH93]. example [CN95, GCWH90, KGW93, MBL94, RHP93, SCC94, TTP^{+}95]. excavatum [DRGV91]. Excess [CCRB98, FRP^{+}99]. exchange [DZ99, GFH96, GHHC98, HG92a, KSG^{+}92, PMETM96, Pez91, PD93, TUN99, ZD99, vDvRK93]. exchanges [FGB^{+}96, KBWC90, TA93]. excited [JM99]. excludes [RG97]. Excretion [Lig90, MRL96, PPCR97, ADR94, CNCS93, CCC94, CC97a, CDI96, Coc90, Mat92, SA94, vdL98]. excretions [CC96]. excursion [TBLD90]. excystment [Kam94b]. Exe [Ken94b, MCRGC90]. exercise [TR91, TRPS94]. exhaustion [BCH95]. exigua [Hun93]. eximia [Góm95]. existence [SFWD90]. existing [HJHG98, MJW96, NA99, WASSM97]. Exocoetidae [HOM95, OMH95a, OMH95b]. exoculata [PDB^{+}97]. exogenous [Bus96]. exopolymer [ES99, PA94, SDC99, SH99b]. exopolymeric [MB98, Mar99b, SH95a]. exopolysaccharide [SHW^{+}91]. exotic [BL94a, Cro98]. Expansion [BL94a, FK95a]. expectations [OTF95]. expenditure [AAS^{+}91]. experience [CE93, QQ99]. experienced [LD93]. experiment [BBLY98, FGKP95, KS97, KQH^{+}96, LOS90, MA99a, McC99, Mol97, ODN^{+}95, PD95, RBD^{+}97, TGvJW95]. Experimental [AWdS92, BK90, GSP99, GF90, KDR97, PLP92, PSR^{+}99, RS94, Sch93, SJC99, TLW^{+}96, TDK97a, WLT^{+}95, ADF96, AB95, CN92b, CH97a, CF93a, CP96, DRS94, DTA^{+}97, EMR^{+}97, ERC95, FDZ93, FDZ94, GSB98, GA97a, GPLB96, HJH^{+}99, JPR92, KSS90, KKT96a, LKC95, Lew97, MF91, MU99, MT91, MD94a, MPS96, MCNB98, MCA90, Mic96, MAdG^{+}96, PPCL^{+}98, PSK98, Pre95, QRSC99, San97, SK98a, SW99, Sew96, SE91, TL91, VD95, WM99, WK92b, ZC92, dBN^{+}90, Ber90]. experimentalist [PB94]. Experimentally [HLP99b]. experiments [AS99a, AH95, BS94, BSH^{+}94, BSG^{+}97, BMV^{+}96, CMRF94, Car90, CGFC95, EB92, ER95b, EPN^{+}99, GJVG96, GMG^{+}92, GF98, HN90a, HN90b, HRKR93, HMHDA94, HBM93, HS97a, HHN90, LBM^{+}99, LWK^{+}95, OJ99, QK93, RMMB91, SR95, Som96, SF94, THCD95, Vet96a, Wah97, Web91b, Wol92b, San94a]. explain [DBMK99, DC91, EST95, GF95, IVM^{+}98]. Explaining [SBT97]. explicit [HHM96]. Exploitation [Gra91a, JCCCB94, WWC^{+}99, CK98, MU98]. exploited [Ken92, Las98, PS92]. exploration [OG95, WMB99, WR98]. exploratory [RS92]. exploring [Lav97]. Exponential [SM94]. Export [FLDT92, MWH91, HM94a, LR96, RA96]. exposed [ABY97, BEG97, CMRF93, CV98, CCC94, CC97a, Dut98, HJH^{+}99, HS96, HS98c, JMC93, JMD95, JPA^{+}96, JBHH97, LSP99, LFG98, RSAD98, SUH95, WJD93, WK92b]. Exposure [SP99a, BDH94, BBP^{+}96, CL90, DM96, DvCd94, DE91, GPB96, GPP98, GCKW95, GHL^{+}97, HDB95, LS96a, LM91a, MM96, Nan96, PPCL^{+}98, SF95, SMH97, SE91, VDVC94, VGJ90, WLT^{+}95]. exposures [Rin92, TNP97, WTF96]. Expression [KBA^{+}97, BCBR97, HSBP97, MM96, MBL^{+}98, PFP93, PCK^{+}96, PCC^{+}97, UKR^{+}91, WKH^{+}98]. extending [HLP99b]. extension [GT98]. Extensive [CMB^{+}99, BMS^{+}97, CSM^{+}98, GF94, MSCTBE96]. extent [SFS95]. external [ASP91, Bus96, PB96a]. Extracellular [VD94, AR99, GKMT99, GAB90a, GAB90b, Kam94b, OH95]. extractable [JSH99]. Extraction [Lai92, BKS92, CLHF95]. extracts [BCG97, SP93]. extrapolated [MLP90]. Extreme [BGBS91, RSM^{+}96, SP99a]. exudate [CS99]. exudates [MH93a, Mar99b]. exudation [LVL^{+}99]. exulans [WW92a]. Exuma [SMW^{+}93]. Exxon [DF94, JDSB99, vTSD97]. eyed [WFHM99]. eyestalk [SSW^{+}96].

f [Har90]. f-ratio [Har90]. facilitate [PDS97]. facilitated [GH92]. facilitates [HS94b]. facilitatory [WCT95]. factor [BNvB90, Hil91c, LD94a, MH93a, PDBA90, SP99c, UKR^{+}91, WMG91]. factorisation [WTM^{+}96]. Factors [Adj97, BKP98, BF95b, CBL96a, GBH96, LP95, LKK92, MK93b, SK94a, WD99, ADR95, BAM^{+}99, BDA^{+}99, CA99a, CC91a, CAATS94, CPB96, DvRN99, GJ97a, GJ97b, GJVG96, GS91a, GPB^{+}98, GRSV98, Har97, HS96, JA91, KW97a, KW98, MY96, Mar97, McL96, MS96a, MP96c, MAAT98, NB96a, NFM98, OD90, PMETM96, PT95, PHHH96, QZA95, RB98, SJML99, TRU91, VGM94, Yam97b, ZD99]. facultative [NEIR95, WM99, Ang95a]. facultatively [VAR94]. faecal [GS94b, Lee97, RR99, RCTH99, UMD92, Vos91]. faeces [GGB94]. failure [BB97a, ER90, RM97a, Uye96]. Fal [SGW94]. fall [BSP^{+}99, DG91b, DRC^{+}97, LDKR93]. fall-early [DRC^{+}97]. fallout [Cam97]. False [CBL96a, CBL96b]. families [Tol98]. Family [Gre96, Bel95, CN95, CAL96]. fan [CHFL95]. fans [NBS^{+}97]. fanworm [AW97]. far [VHH^{+}96]. farm [HAH^{+}90, HH91a, KTH98, PLP92]. farming [GSBB97, HK92, KHTP99]. fasciata [BSMS95]. fasciatus [Mal99]. fascicularis [FPGC^{+}98]. fast [CMBM99, DAS92, HSBH96, MF96]. fat [ISF^{+}98]. Fate [And96, LAA93, LHK^{+}93, VPFL92, BMP^{+}95, CDME97, Hol96, KF97, KK96, LC95, LGH^{+}93, MFF97, NNB93, OKL^{+}96, PW94]. fates [LM91b]. fats [SM98a]. Fatty [BBCB99, CWHA99, IFL97, KZK^{+}95, LNSL91, Naj96, PFMG96, GHTH94, HSN93, JMPH90, MCA90, MVA99, Naj97, PDB^{+}97]. fauna [Bra95, BMH93, CS93, Dye98, FKC^{+}90, HPF92, HCI93, McL96, MW97, NL94, RHJ91, SM91, SS95]. fauna/flora [HPF92]. Faunal [BSGM94, TRU91, BH91, BLR94, MAC92, TDBH94, FMH96]. Favella [KA97]. Favia [OBL97]. favoured [Som94]. favus [OBL97]. FDC [Tuo97]. Fe [HB94b, MP96a]. Feasibility [FCC91]. features [BM94d, DvdM92, Egg95, HTS92, HHM97, JW96, MR97a, Pal91, RSH92, SFPC97]. February [HC94, DSG^{+}99, KL95]. fecal [BMTF92, BH92d, BBG90, BD94, Dei90, DA93, FD98, MNH92, URHR98]. fecundity [Aks96, Dut98, Iri90, KS95b, PHHH96, QC94, RHH97, SW93, SLHC98, ZHM91, VK94]. fed [ADR94, DO92, EPL94, FG93, KEV99, MHB94, vdL98]. feeder [FAP92, HAA94, HP98, JJNR95, KM96, NEIR95, TBY98, TGvJW95]. feeders [AJL93, CGA95, CM99, DLHB99, FL90, GAC^{+}97, Hil91c, LBL91, LHL97, PM94b, PS98, RCG99a, RW99]. Feeding [And94a, ABY97, BDdWC96, CPC98, CW93, CGZR94, CCS98, CJ93b, DWSA98, FGG95, GRC^{+}97, GPS93, GH94c, GFPD92, HH94a, HHP95, JSK^{+}99, JT90, KW97b, MT91, Met98, MPCD90, NW97b, NGS97, NEIR95, OM99a, OPS96, Øre90, Øre91, PNS90, PPB90, STBR99, TH90, UT90, VEW99, WC98, WC92, YD90, YKL99, AP99a, AMB99, AR91, BP97a, BCN^{+}95, BJdJ^{+}98, BN91a, BGC91, BB95, BvKvRB96, Bus96, Bus97, CBI91, CM90a, CA92, CP90, CW95, CMA95, CE93, CLW99, CH99, DP91, DPO91, DD90, DEF91, DDW90, Dut98, EK95, EE98, FW91a, FPG95, FMRC93, Gib94, GSP99, Gon92, GSS93b, GMF98, GH92, GT96, GS99b, Han95, HN97a, HFW99, Har99, HK96, HL90, HSBH96, HSTY98, HJH^{+}99, HH96, HVL96, HK99, HD91, HSN93, IHK^{+}98, JWKH98, Joh97, JS93]. feeding [JFS92, Jør96, JC92, KR97a, Kam94a, KSRB99, KM96, KEJ98, KNK90, KS95c, KL94, KWEG94, KGC92, Kyo92, LW95, LJSK96, MSD98, MA99a, MRMR92, MJ97a, MSS95, ML90, NYH92, NT95, NR98, Øre95, PS93, PPP^{+}96, PPMF97, PCG96, PTW^{+}91, PM91b, RLDM94, RKH96, RHT91, RCTH99, RCG98, RCG99a, RI90, Rii91, RTJ^{+}93, RPL96, RM97b, RD97, RMMB91, RHH94, RD92, SK95a, STNI97, SM97a, SMTL99, SMB97, SP93, SH99a, ST98, SM99b, SE98, TK98, TNP97, TK95b, TPHC91, Tis98b, TP95, VR93b, VAR94, Ved98, WSF99, WFHM99, WTF95, WH93, WDD92, Wol92b, YKL98a, YKL98b, dT94a, dRAV92, van94a, vdL95, WC92, SB94]. feeding-preference [MA99a]. female [BSH^{+}91b, CH97a, DP91, IHK^{+}98, Rv92, SMUS^{+}99, SCMM95, ST98, ST99, YSY^{+}99]. females [Kyo92, RCTH99, San93a, SB97b]. femorata [Joh97]. fenestrata [ÓN97]. Ferrara [SSG96]. ferredoxin [DEP^{+}96, EPD^{+}99]. Ferrioxamines [SDH95]. ferritin [DCG^{+}99]. ferrugineus [BP99]. fertilisation [BBLY98]. fertilization [Hsi97, KMG98, MSP^{+}97, PB96a, RZLB96, RB90]. fertilize [RCG94]. fetch [OL99]. fetch-limited [OL99]. fiber [AFZ96, KLJ94, LKF95, MGAE99]. fiber-optic [KLJ94]. fiddler [LMO99, O’C90, O’C93a, OJ97, OJ99, Wol92a, Wol92b, Wol93]. Field [HVL96, HN91, KH92b, MC97a, SF94, VAR94, Wol92b, BSH^{+}94, BH91, BCH95, CMRF94, CCV^{+}98, Cla90, CSE^{+}96, DO92, EB92, FDZ94, FQV98, GM98a, GF98, HN90a, HN90b, HHN93, HMHDA94, HBM93, HHN90, JBK99, KTW^{+}97, LJSK96, Lop91, MY98, McC94c, McC99, MBL^{+}98, NU91, NC95, OJ99, PML93, PLFSW97, PTB^{+}97, QK93, SLHA97, Spr93, TGvJW95, TL91, VTGK99, WE95, ZGM^{+}96, ZN96a, vRBVB95]. field-caught [CCV^{+}98]. field-collected [CSE^{+}96, DO92]. field-deployed [WE95]. fields [BP96b, CD98, Eml90, PHT96, PWPK95, PWP^{+}97, Pud93]. fifteen [RM97b]. Fiji [DWG98]. filament [RI90]. filamentous [HBM93, KJMRC96]. filiforme [GMR^{+}94, KS98, SDC90]. filiformis [AGP98, GS99b, LJSK96, RNHH97, SLR94, SG96]. filled [KCI98]. Filling [TR95c]. fills [KHFP^{+}96]. film [Dun99]. films [MS95, WCT95]. Filter [RVBL92, RM97b, VR93b, CR96, Jør96, KNK90, KL94, LBL91, NEIR95, NR98, RPL96, RG97, VAR94]. filter-feeder [NEIR95]. filter-feeders [LBL91]. Filter-feeding [RM97b, VR93b, KNK90, KL94, NR98, RPL96, VAR94]. Filter-net [RVBL92]. filter-pump [CR96, RG97]. Filtering [LCLC96]. filters [AFZ96, BP94, LKF95, MGAE99, PPR94, RDPB99]. Filtration [JTPP92, NK90, PR92, PSPD96, CR96, RI90, RTJ^{+}93, RG97]. Filtration-induced [NK90]. fimbriatum [DC91]. Fine [MC95, GRGK96, IFL97, SD93a]. Fine-scale [MC95, SD93a]. fingerprinting [CGM95]. Finland [KHK^{+}99, PT95, TI96]. finmarchicus [AS99a, BM99b, CH97a, DRC^{+}97, FP96a, HB98, IHK^{+}98, LGMD98, MS98, MCW^{+}91, NGS97, NKH^{+}99, PF96, PR93a, RCTH99, SH90a, ST99, UMD92, WDB98]. finned [AS94a, GRC^{+}97]. fiord [Lam98]. fiords [Mil97]. first [BHT^{+}92, HTLK99, KL91b, QCM97, SMUS^{+}99, SLSH98]. Fish [BGH95, BMG99, GEW97, JFS92, KS94, MSM^{+}98, PPFP99, SZ96, VBL^{+}92, ASCW90, AB90, And94b, AC99, Ano99a, APS91, BP97b, BAH94, BPS98, Bee97, BB97b, BS96a, BA91, BJB95, BH92b, BS95b, BSD^{+}94, BPL95, BRC99, BS96b, BvKvRB96, BK92, Cal95b, CGM95, CF94, CP95a, CK99, CT97, CDI96, CBL96a, CBL96b, Cla96, Cle93, CD96, Con98, Con94, CGFC95, CH93, Cro99, DBMM99, DL91a, DL93a, DFPG91, DG93, DCG^{+}99, DvRN99, ELG97, FA99, FB91, FLDT92, FPG95, Fow90b, FDW92, FH99, FII95, FWF^{+}99, GBC92, GDvS98, GMO95, GPP98, GAR93, GFPD92, GK96a, GB93, GCNSNd94, GF91, Gru94, GRSV98, HSLV94, HPPZ93, HAH^{+}90, HG98, HHB93, HKR96, Har98, HS92, HS99a, HHM96, HHW90, HB98, HH91a, HK92, JJ99, JBWH97, JBK99, KMKS96, KKH98]. fish [Kal91a, KTH98, KHTP99, Ker94, KMG98, KJLH96, KKYH97, Kin92, Kin95a, Kne93, KZ92, LJ95, LP95, LSG^{+}99, Lev91, LCG97, Lew97, LBH91, LMTF96, LBK90, LB93, LBK96, MLP90, ML91a, MD91, Mac94, MG96b, MG91a, MM93, McC94c, MM95a, McC99, MMD92, Mil94a, MD94c, MZ92, MCFS96, MM95b, MLDM99, Mye95, NvA91, NKGK92, OC95, OM99a, PCTF95, PS95, Pro92, RAB^{+}92, RD91, RMGB^{+}94, RD95, RHP93, RA96, RO99, Sab90, SO96a, SB97a, SLHA97, SDFJv91, SGMS99, SO96b, SKPT97, SC96a, SFA98, TH90, Tho92, TB97, VHH^{+}96, Vet92, VDBF96, WA99a, WB92a, WO94, Wöh96, Yam97b, vRBVB95, vRV96]. fish-habitat [HKR96]. fishcagefarm [HHKS92]. fished [FCH99]. fisheries [FPB^{+}94, GCF96, RM97a, Str99]. fishery [BPR90, GTYR97, MW97, PFH98, SOW97, SM99a, WTR99, WO94]. fishes [AMS99, ACR99, Bee97, BL94b, Bro98, Bro94, Cal95a, CH95, CA96b, CAL96, CCG^{+}92, CA98, DKL94, FGG97, Fow90a, Fre97, GRBA96, Gra96b, Gre96, Gri94, Gru95, GP98, HK96, HKR96, HR96a, HS98b, HS99b, JRMN^{+}97, Ker96, LCE97, LH96, MBF^{+}97, MZ99, MA99a, MM97a, McG94, MP92, MRS96, MN99, PPM96, PR93b, RK96, RBF92, RSB93, Röp93, Ruz96, RO97, SM98a, SMG97, SC94, She95, She96, SMW^{+}93, Sog92, Sto97, SB97c, Sym95, TC93, TM97, TSM^{+}94, TcBW96, Tol98, TH94b, VL97, VHVB^{+}98, Whi94, WM97]. fishing [ASRC99, CEV97, EMR^{+}97, Gri94, JHG99, KBN^{+}96, MMBT97, SKM93, YBLL96]. fissifolium [MS94a]. Fitting [BMHH94]. five [CS97, CES^{+}94, GK96b, HDD98, PPB90, PAR98]. fixation [BEU99, BP96a, FHCC98, JP93b, KZP93, MRPC98, MC99, PJF93, SCLS94, SGP94, Vil91]. fixed [BEU99, PSB^{+}97]. fixed-nitrogen [BEU99]. fixing [CSM^{+}98, CMB^{+}99, MHRV^{+}99]. Fjord [AWD95, NNR95, AAWD97, BMH93, HTS92, KR98b, NNB93, OFR94, PR92, SAG92, THS94]. fjords [Ber90, FB90, ONR98, SS92]. FLA [SSM91]. flagellate [GD90, GCGR90, Kuo91, NAI^{+}93a, NU91, NYH92]. flagellated [AAW98]. flagellates [BC90, DLHB99, GSS93b, HR96b, JG90, ML91c, NK92a, NH90, SSM91, TSS93, UT90]. flamingos [GB97a, GB97b]. flat [Beu92, GKMT99, HH99, Hüt90, LH99, MKYI98, NA99, OSZ95, Phi94, RD97, Tsu90, YMD94, dG91a]. flatfish [Moo92b]. flatfishes [EG95, EG97]. flats [BPC98, Dit96, FB94, GBGH97, Hil97, IF96, PGB96, PE99, PCB90]. flattened [JS93]. flavodoxin [DEP^{+}96, EPD^{+}99]. Flemish [And90]. flesus [LFG98, RH90b, RBN99, vBDW91]. Flexibility [O’C90, OJ97, SP99a]. flexuosa [BSMS95]. flexuosus [VR98]. Flipper [CWB93b]. Flipper-bands [CWB93b]. flocculator [WGD97]. flooded [ZT90]. floor [GSC96]. flora [HPF92]. Florida [FKC96, FMH96, BHH^{+}95, DY92, DCK99, Dur94, FF97, GEW97, Har96, KRJ^{+}95, KS98, LBJ^{+}98, LSGW98, Mil95a, MKV95, OM99b, PLB95, RBC^{+}91, RSK^{+}99, Sto96]. floridae [Sto96]. Floridan [AMBH98]. flounder [BP99, BCL93, BB95, BSH^{+}91a, BSH^{+}91b, CA98, HKM92, Kea90, LFG98, MT91, MYYT96, MJ91, McC92, MRS96, WB92b, WA95, YNYK94, vBDW91]. flounders [RH90b, RBN99, SGB90]. Flow [BVP91, Eml90, JTPP92, SBBG94, AP99a, AS94b, AHS97, BG97, BWRH99, BPL95, BP96b, CB98, DBV98, FKB92, FHZ96, FL92, GNJ90, JCP93, JLD^{+}96, JS93, JAL91, KM96, KG99, KLAH^{+}94, LMF93, LZYW93, LHK^{+}93, Lip95, LJSK96, MPRB93, MWT96, NNB93, PJ99, PM98, QRSC99, SK98a, SSY98, SEGP91, VK90, Wah96, WB99, YMD94]. flow-habitat [AS94b]. flowering [MW99]. Flows [Pre91, GBV96, HG92b, SC96b]. Fluctuating [SKPT97, CA96a, LGB91, MC97b]. Fluctuations [ABB^{+}97, AGP98, APT99, BSVK93, BDA^{+}99, CA97, FSG^{+}98, Hic92, MDTB94, NT95, RW96]. fluid [ODS^{+}96]. fluke [RCA^{+}95]. Flume [GNJ90, Kne91, BSG^{+}97, FKB92, GBV96, HG92b, JAL91, LJSK96, SFG^{+}99, Wah97]. fluoranthene [FFH96, KF97, MFF97]. fluorescence [BAK97, DRGV91, FM95a, FL92, HHYD92, HWLC98, HLF^{+}91, HPSG94, KBP98, LZYW93, Maz95, MOM98, MGC95b, MWH91, MC91d, PPB90, SR98a, SH98b, TBLD90]. fluorescent [HP98, LMK93, Moo92b]. fluorescently [ER95b, HBvD92, SSM91]. fluorimetric [Cle93]. fluorometry [BVW^{+}98]. fluoxetine [HLP99b]. flushing [CP91]. Flux [LM98b, RE94a, ATNW97, ANW96, AW98, AMB^{+}96, BNK^{+}97, BD98, BDNS99, DGO92, DA93, DRP^{+}97, GS94b, HHH97, HSW99, IPT94, JLS90, JKFB95, Kro91, LLP94, MDPP91, MN92, NPB94, NU91, NG96, PF91b, Sim92, SEGP91, TT98, TM94, VKV95, WK92a]. Fluxes [SUH95, AC96a, ASRP^{+}96a, BCG94, BD94, CRCS96, CCR91, CCRB98, CG93, CPB96, DBB^{+}95, DSV^{+}91, FGP98, FGB^{+}96, GPDF93, GPF95, GPD^{+}96, GST^{+}91, HAH^{+}90, HHKS92, HDPW93, Hei95, HH91a, KLEH91, MM91a, MPCD90, NT96c, OL99, PRPM^{+}94, RGS95, STK99, Sün97, TDBH94, TA95, TCRB93, TNSM98, Wai90, ZHF96]. flying [JCCCB94]. flyingfish [COS93, HOM95, OMH95a, OMH95b]. Foliar [Pak92]. following [AP99b, BDH94, BMP^{+}95, BA91, Bre96, CMT94, DB99, DF94, ER90, GH94a, JLS90, KHTP99, PBVB99, RD91, SR90b, SM91, SMH97, VGJ90, VBN95, vTSD97]. Food [AH93, CM98a, HSB97, KSRB99, Kea90, LR96, LCK^{+}96, MHRV^{+}99, RRGB96, SO96b, AAS^{+}91, ASN95, AV95, AH94, Atk96, BN95, BMSHC98, Bar97a, BC93, BGG^{+}97, BP92, BBB^{+}94, BRN^{+}95, BGvDR96, BK92, CA96a, CA97, CA99a, CLC^{+}98, CF94, Cli95, CTC93, CHC94, CH99, CNP95, DHH98, Dav96, DL91b, DG97, DDW90, Ejd98, FL90, FAP92, FW91b, GCF96, GPB^{+}98, GC95b, GSMM96, Han91b, HHB93, Han99, HTG96, HW92, ICG96, JCH93, Kam94a, KI97, Ker94, KZK^{+}95, KKK^{+}90, KS95d, KKT96a, KGS95, KBS98, Kur99, LDV95, LM98b, LHK^{+}93, MW94, MCNB98, Mat92, MC98, MRMR92, MV98, McG97a, MM97b, MVdM^{+}99, MHD97, MJSDF91, Nak92, NJLD98, OK98, ONM93, PC96, PH91a, PS93, PCBA94, PM90b, PMG^{+}98, PMOD90, PG90, PWNW99, QQ97, QGQ97, RM97a, RNP^{+}98]. food [RSB^{+}90, RSN^{+}99, RH90b, SCIA99, SB97a, SBBG94, SJ97, SDFJv91, SW98b, ST90, SMP97, SM90, TT90, TR95a, TS90c, UHL^{+}97, UMD92, VV95, VAD^{+}94, Vos91, Wai90, WFL95, WQQ99, WTF96, WR92a, WCLB99, WC92, YKL98a, ZBW92, vW93, Ang95b, BvB94, CW93]. food-chain [RH90b]. Food-web [LR96]. food-webs [PM90b, SM90]. foodweb [Fra95]. foodwebs [FCP98]. Foot [MC91b, SH94]. Foot-raising [MC91b]. forage [HS92, WW92a]. Foraging [BGG^{+}97, BvB94, BBB^{+}94, Bur94, BKH99, DH96, GKB^{+}97, HHM97, HH90, HCH^{+}96, JBB98, MBWT98, Mun92, RHH94, RHrsHG99, San93a, TJ90b, WR97, WSP^{+}95, AWD95, AU93, And98, APS91, BCBC98, BA94, BKB94, CWWH99, CHDW92, GGF99, HHW90, HRCW98, IFL97, KR97b, LADS98, ML91b, MJ92a, Nev99, NKGK92, PM92b, Pre95, PB98, RO97, TMCE98, VKF^{+}98, WWL97, WC98, WB91, ZGM^{+}96]. foraging-predator [WB91]. Foraminifera [AB95, BB92, MHMH96, MvdZH^{+}97, MSvdZH98]. Foraminiferan [AC96b]. forbesi [SB97b]. force [HM97d]. forces [AC96c, ML90]. forcing [ELMR98, HCI93, MCED96]. forest [Cha98, HSK^{+}94, KL91b, NBW94, SFV96, TD91, TDER97, VHH^{+}96, WH91]. forests [Hag95, McG97b, SDER96]. form [Dua91, GHED^{+}94, HCG^{+}93]. formaldehyde [CGA95]. formalin [DFB^{+}99, HKM92]. formalin-preserved [DFB^{+}99, HKM92]. Formation [HH94b, SH95a, SKKO97, AKM^{+}99, DSJ90b, GT92, HKA96, Hol97, KPWL93, MLDM99, NT96a, NDRP93, PW94, PBHMI98, SL94]. formed [KHT93]. former [KDR^{+}92, NTS90]. forming [AV98, JSGS94, YKK^{+}98]. Formosa [Wol92a, JLGMP97]. forms [BLKA93, HS91b, KHT93, SHA^{+}95a, SAA99, WPY98]. formulation [FM94]. forsteri [WR97]. Forth [MGM96]. fortunata [PSB^{+}97]. Fouled [Wah96, WS92a]. Fouling [Ger90, ACOK95, BSS92, NTK93]. found [CPS91, CHFL95, LBL91]. four [BU93, BDA^{+}99, CDME97, DHH98, FTS97, Hul97, KG94, MY96, ØW93, SRJD97, Tay98, WUA^{+}99, vG91]. fractal [GW94, SL97a]. fractals [DPM96]. fraction [PF92c, Sut92, ZF99]. fractionated [GMG^{+}92, GDW90, HHW99, KFS91, LGH^{+}93, RTRF90, RFN98, RN98]. fractionation [BRZ99, Fry96, KM99a, KR98b, RRWG96, WYY^{+}98, WYY^{+}99]. fractions [BG93, DP91, KKL91, LMT96]. Fragilidium [HN97a, JLYK97, JSK^{+}99, Sko96]. fragilis [MDG98]. fragment [Lew91, YKK^{+}98]. fragmentation [CC99]. framework [EW91, Gas94]. France [DLT^{+}97, DQT^{+}97, GBB^{+}98, LJM99, RRGB96, RR97b, Rie98, SBT97, GB95, GSBB97, HPF92, IF96, LGV^{+}99a, LGV^{+}99b, LIPF95, RSN^{+}99, SGC^{+}98, VBCRB97]. franciscana [BFD91, BKB92, ZGB90]. franciscanus [ESDK94, EDS^{+}99, ME97, SDER96]. Francisco [BL95, CCG98, CL90, CTSN90, JCP93, KGO94, NTS90]. Fraser [HCC^{+}91, Sew96]. Free [MS93c, SA92, VVT97, BHQ99, BPH95, CCC94, CDGF^{+}93, CG93, FF92a, Fuh90, HQ99, HH94b, JNCS89, JMPH90, JKC^{+}93, KK91, KDKS96, MBK95, ML91c, NK90, PMOD90, RN93, RFGF92, SUH95, SAH98, Sty97, SCF91, TM94, VTGK99]. Free-living [VVT97, HH94b, JNCS89, KDKS96, ML91c, TM94, VTGK99]. free-running [RN93]. free-spawning [Sty97]. Free-swimming [SA92]. freely [BP96b]. Freezing [Ohm96, CV91, Ken94c]. frenatus [HK96]. French [Adj97, CRCS96, CCRB98, DG93, GPDF93, HGS95, LGH95a, NJLD98, BLDF95b, BLJ99, BBT^{+}93, CCR91, Cha96, CDG97, CB98, DHM92, DLHB99, GSC96, MG96c, PPCL^{+}98, Pla93]. French-Belgian [BLDF95b]. frequencies [BZ92, CHS^{+}95, PMETM96, Sci91]. Frequency [WP94, BM98, BMBD95, BGVF92, FW91a, IHK^{+}98, LCF99, Mac91, MS93d, MA99c, NKH^{+}99, SW98a, TH94a, TDER97, WTM^{+}96, ZHM91, vRV96]. fresh [FS91, MB95a, SCB92]. freshly [FKB92]. freshwater [DLT^{+}97, DQT^{+}97, FPLH92, Fra94, MMM^{+}95, MEK99, RR97b, SLHA97, VBF94, Whi94]. freshwater-saltwater [VBF94]. friderici [Gib94]. Fridtjof [Bia92]. fringe [SR97]. fringing [Nai93b]. Frio [CF93b]. Frisia [KDHZ98]. Frond [GW91, San94c]. frondosa [SMTL99]. fronds [BCBR97]. Front [PPMF97, RPT^{+}96, GLD^{+}91, HNC92, JC97, Mun93, MLDM95, NAW99, NIU^{+}98, RHrsHG99, SO96a]. Frontal [PWR96, WSCT96, CPH91, FLDT92, HRÆ94, JL96, LFTH92, MLDM99, PBHMI98]. Fronts [PPM94, DH96, EAEP98, Fra92, HKK^{+}95]. frozen [HKM92]. fry [SH97a]. Fuca [GJ96a, SRJD97, SJML99]. Fucales [Ang91b, AD93]. Fucus [AKK94, Ang91a, Ang91b, Ang92, AD92, AD93, Arr93, BCR92, CNJ96, DF94, DCM90, HHN90, JBH98, MM90, MEK99, PB96a, PKY96, VS91, WC96, vTSD97]. Full [KW92b]. Full-depth [KW92b]. fulmars [TFL95]. Fulmarus [TFL95]. fulvic [DL94a]. fulvum [KKL^{+}98, KKL^{+}98]. Function [RHB96, ATH95, BSJPW93, BP96a, DBHM99, EBM93, ENSR90, FDZ94, GCKW95, JFP95, LHRL97, MB95b, PPF95, RVGC^{+}99, Roa93, SKvB98, TAD^{+}94]. Functional [DL91a, GHED^{+}94, NA95, DvdTCC97, DSS^{+}98, EP92, MHD97, PKL97, VR98]. functioning [HS91b, SSG96]. functions [Sko96, UTB^{+}95]. Fundulus [Har98]. Fundy [MHBW90, SMTL99]. fundyense [AGGKA96, BLCA90]. Fungi [LGP95]. fungiid [KWL96, KWL98]. Fungiidae [Hoe91]. fungoid [RR99]. funnel [RHB96]. funnel-shaped [RHB96]. further [SRSH99, SOC97]. fuscus [Oro96]. fusiformis [Lav97, TD92, VBCRB97].

Gabon [Bia92]. gadoid [BRLM90, FSG^{+}97, MLDM99]. Gadus [Cam96, GA97b, HSB97, LAK99, MF96, MLDM95, Nie93, NKV94, SFF92, SK95c, TH90, TRMW95]. gains [AP99a, HHv91]. galapagensis [MT94]. Galapagos [FCMB94]. galatheanum [Nie93, Nie96]. Galaxea [FPGC^{+}98]. galbana [Tha92a]. Galeolaria [BH97b]. Galicia [FGG95, BQR99, RQB98b]. Galician [VF99]. gallina [RR92]. galloprovin [GS90]. galloprovin-cialis [GS90]. galloprovincialis [BH95a, BMFR97, BFMM99, BFVG92, BGVF92, CMRF93, LFRB97, LFD95, MUH97, PMETM96, QZA95, SZA94, SRRdM99, SKM95, VBZ^{+}94]. galtoprovincialis [CMRF94]. Galveston [MW97, ZB94]. gamete [AKK94]. gametes [BL97]. Gametogenesis [Sak97, Van94b]. gametophyles [tor93]. gametophytic [HCG^{+}93]. gammarellus [RK95a]. Gammaridae [She98]. Gammarus [SS91, PGD^{+}93, SSW^{+}96, SCJ99b]. ganglia [SSW^{+}96]. gannets [AAS^{+}91, GGF99]. gap [Dye93, Hol97, KHFP^{+}96, TR95c]. Gaps [CCR98]. gardneri [DF94, vTSD97]. Gari [UC94]. gas [Kru93, OG95, Pez91, PD93]. gas-exchange [Pez91]. gas-tight [Kru93]. gassy [DFJ^{+}93]. Gastrointestinal [SCJ99a]. gastropod [BSS92, BA94, Cla90, DS93b, GMC91, Ger90, Ito97, LSM92, MC91b, RS94, RS95, SRW95b, SR96, Tak95, Tak96]. Gastropoda [CP90, GH91, MS91, RH90c, VT92]. gastropods [Bec98, MD93, Pen90a, PFMG96, QK93, Raw96, RQB98b, SU93, Tro94]. gaucha [San94b]. gaudichaudi [PP96]. Gazi [HSK^{+}94, SOSV95]. GBR [Fow90a]. Gdansk [BW95, JJTS97, WOM^{+}97]. gegenbauri [Dei90]. gelatinous [BTYO94, BYO95, HL98, PG91, PWR96, PGG96, RGBC92]. Gelidium [ASL99, DF95, San94c, San95]. gem [lYALM93, lYAML93]. Gemma [lYALM93, lYAML93]. gene [AHS97, DBV98, FD97, PFP93, PCK^{+}96, PCC^{+}97, PM98, UKR^{+}91, WKH^{+}98]. genera [Fab95, Mon96, VL97]. general [MD91, PWM95, HTS92, Moe97a]. generalist [BPS98]. generalized [MR97a]. generation [Hav93, TFF94]. genes [YKK^{+}98]. genesis [BH94a]. Genetic [ASN93, AS94a, AHS97, CMLE92, CBT98, KRJ^{+}95, KSH97, Mal99, MR97c, Mil97, MTP92, MSCTBE96, Pla93, PAM96, TNN93, ASRC99, BA97, DCK99, DBV98, HBBT98, HHG99, HB94a, HSBP97, Hun93, HH95, PM98, RP91b, SRRdM99, TN90, VB93, Way95, JKT90b]. geniculate [BCC94]. Genotoxic [BFVG92]. genotype [GS91a]. genotype-dependent [GS91a]. genotypes [Mil98]. Genotypic [JKT90b, KJT90, TAJK90]. genus [Han91a, KML91, LT93, MBF^{+}97, Mil94b, SB97a, TP96, VHVB^{+}98, Vil91, WMPT94]. geochemistry [TBY98]. Geodia [KWHM^{+}96, KBA^{+}97, UKR^{+}91, WKH^{+}98]. Geographic [Fra93, LCG97, RCA^{+}95, ASL99, CBI91, Fev92, HFR^{+}93, Mar97, MLES94, WT98]. Geographical [Gar98, WT99a, BQR99, MACR98, PF92b, TPR98]. geological [ODS^{+}96, SRJD97]. George [CHDW92, GWKW97, JWKH98]. Georges [CEV97, DRC^{+}97, HF99, MMGIM90, TRU91, TRS91, TRMW95, TS90c]. Georgia [ASP^{+}96, AS97, AWH^{+}99, HCC^{+}91, HM97b, KWEG94, Moo94, WPTB96, BW96, BWM97, BBCB99, CWHA99, DPM96, GBP^{+}97, GDW90, HFJS91, HHE92, ØW93, PP96, Pak92, PWPS95, RWCM99, ROANM98, WMD99, Wat98, YHGB96, YGH^{+}97]. geostatistical [GBB^{+}98]. Gerlache [LHL93, Nor94, Øre90, Øre95]. German [BDD92, BKS92, BVL92, EGKE92, GAD^{+}92, HC92a, HHH92, Kar92, KR92, McF92, Mol97, RBD^{+}97, RG93, SSR^{+}97, SKM^{+}97, SGR92, Sün97, Tha92a, Tha92b, VBL^{+}92, Wil92]. Germany [KDHZ98, KY97]. germinal [SG96]. germling [VÅCL99]. Gersemia [SMB97]. Geukensia [KNK90]. ghost [EMR^{+}97, KBN^{+}96, SF99]. Giant [BHH^{+}92, AHT^{+}90, BBLY98, BL97, CHC94, GHL^{+}97, GK96b, KG94, KL94, PRC^{+}92, PGM^{+}98, SCMM95, SSP97, SH93]. gibberifrons [Ruz96]. gibbosum [VT92]. Gibbula [Cla90]. Gigartinales [GW91]. gigas [BH94a, BV99, Dav94, Dou93a, Dou93b, DLHB99, Fra93, LHL97, MAAT98, PLBP96, RS95, RR97b, SL94, SR96, Tha92b, WSCT96, WASSM97]. Gill [BSJPW93, BD96, EMR^{+}97, LCTT92]. GIS [Gar98, GM98b]. glacial [HS90]. glacialis [HKMK98, MS98, ST90, TFL95]. gladius [GCNSNd94]. glass [AFZ96, LKF95, MGAE99, OINN94]. glass-fiber [AFZ96]. Glenodinium [NP90]. Global [PT93, RB99, HS97b, RHS99, Rud97, SD99]. globally [BK96a]. Globicephala [AS94a, CAATS94, GRC^{+}97]. globiceps [Mga92]. globosa [HSMS99, Kog95]. Glossodoris [AP97b, BPS98, RP91a]. glucose [Gar91]. glucosidase [KFPR94]. glucuronide [MNH92]. glutamate [FF92a]. glutamine [FF92a]. glutamine/glutamate [FF92a]. Glycera [ND98]. glycine [KG95]. Glycogen [BDH94]. glycopeptides [Wöh96]. Glyptocephalus [MRS96]. Gmelin [TRS91]. gnathiid [GP98]. go [OS96]. goatfish [McC94d]. Goban [FH96]. gobies [Har99, MJC97, dT94a]. gobiid [HS99b]. Gobiidae [WM92]. Gobionotothen [Ruz96]. Gobiosoma [Har99]. goby [KYN94, MF91]. gonad [SK97]. gonadal [CRZG95]. Goniastrea [ABB^{+}97, LB96a, Sak97]. good [CH97b, JKT^{+}98, MM93, MMD92]. Gorgonia [CHFL95]. gorgonian [CGZR94, CRZG95, CZG95, FP91a, HFR^{+}93, LKC98, MC94b, MDSB92, RCG99a, RH90c, VT92, WV95, WHW93, YY91]. gorgonians [EL90]. gorgonion [PF92a]. Göteborg [SR93a]. Gotland [MHRV^{+}99]. gouge [KCI98]. governing [FB94, LKC^{+}99, LKM^{+}99]. Gracilariopsis [VBN95]. gracilis [LW91]. gracillima [GSP99, PLFSW97]. gradient [BBFD98, BMH93, CS97, DON^{+}95, EDS^{+}99, FF97, FII95, GBC92, GHL^{+}97, JSJMH90, JJNR95, KBWC90, KDMR97, LM91a, LKK92, McF92, McG92, MKN^{+}96, MNO96, OSC97, PKL97, RSH93, Rie98, SO91, SO96b, SGW94, Tom91, Wes90a, Wes90b]. Gradients [CK99, AGP98, ABM96, BSB91, CBB90, DSS^{+}98, GF95, GAD^{+}92, GFH96, McG94, MGC91, MGC95a, RBF92, Röp93, SH92a, VWJ96, VDBF96, WA99a]. grain [WHG^{+}99a]. Grand [MB91, PSR^{+}99]. Grande [PFH98]. grani [CA96a, CA97]. granulata [BI99, IBB97]. granulometry [DD99]. grapsid [CCDA98]. Grapsidae [AHMd90, Ang91c, Ang95a, Ang95b]. Grapsus [KW97b]. Grass [GF98, KFD99, PH91a]. grasses [VH99a]. grata [WM95]. grave [SH94]. Gravelines [LIPF95]. graveolens [GK97]. gray [GH92, HC92b, WCL97, HFJS91]. Grazer [HB94b]. Grazer-mediated [HB94b]. grazers [BH95b, EP92, KKT96a, TLB93, WLB^{+}99]. Grazing [CLM^{+}91, FF92b, GS93, HvB91, KW99b, MEK99, MPS96, She98, SR94, TMKL97, VGM94, Wer97, ACLS97, BNvB90, BH92a, BCC95, BPH95, BRN^{+}95, Car90, CS99, ER95b, EP92, FP96b, GAK99, Ger90, GMG^{+}92, GSS93a, Hag95, Han95, HH92, HH94a, HH96, JH97, JHN99, JL94, JG94, KSS90, Kuo91, KL90, LPF94, LLP94, LKC95, LF94, MPM97, MH93a, McG95, MEC92, ML91c, NYH92, NGS97, NC95, Per91, PPMF97, PG94, PR92, PPW96, Pre95, RCG99a, RCG99b, RH90b, RPL96, RNH99, SM94, SSM91, Sko96, Str91, SS96b, SMB98, Tee99, VH99a, VPFL92, Ved98, VÅCL99, WH91, WSM91, WR91, WR92b, WSS94, YHGB96, ZDJ^{+}95, LLP94]. grazing-resistant [JG94]. Great [BEU99, GFPD92, MD94c, MKYI98, SU93, GSC95, HBBT98, PCT^{+}97, AHS97, Bel95, BMSF95, BM91a, CC91a, CL97, Fab95, FDW92, KGHTN99, KM92, LC97b, McC96b, McC94d, NT96c, PD90, PD94, RS91, RSS94, SR90a, SRSH99, Sam97, SB99, Tho92, TcBW96, TNSM98, UDLM99, VDG95]. Greater [PR93b, GB97a, GB97b]. grebnitzkii [TK95a]. green [GS94a, Hag95, HJH^{+}99, HLP^{+}99a, HGB99, HGBW99, HTLK99, JPZ^{+}98, KSS90, LC97b, Tro93, MKYI98]. green-lipped [HJH^{+}99]. Greenland [ADBH99, Bra95, LGH^{+}93, MP96b, NH95b, RTRP^{+}98, RNH99, WBGS97]. grey [BH94b]. griseus [LMT99]. groenlandica [LADS98]. grooved [VHHS96]. Gross [Vet92, CCV^{+}96, WC92, vBV95]. ground [EG95, FCH99, GPB^{+}98, MR91, PG90, RGE^{+}95, RDM91, Row90, SGM98]. groundfish [Liv93]. grounds [CLD99, KEJ98, TRU91, WS91, YBLL96]. Groundwater [BDNS99, MH94a, Sim92]. Group [BA94, BK90, EG95, EG97, MSP^{+}97, PKL97, PCC^{+}97, vBDW91, vW93]. group-specific [PCC^{+}97]. grouper [Egg95, Joh99, LCW99, SMW^{+}93]. groups [OIO^{+}95]. grow [Kin95b]. growing [GSS93b, VD95]. grown [LW91, MCA90, MH96b]. Growth [ADA98, AH99, BBLY98, Ber92, BC90, BH92e, BAM92, CA97, CH97a, CA96b, CAL96, CR96, CH90, Cra94, EDS^{+}99, EGMC94, GMR^{+}94, GGDLE96, Han95, HMHDA94, HW91, Hov92, JPC97, JL94, JJNR95, Kne93, LC97a, LH91, LCL95, MD94a, MDC^{+}96, MCED96, MW99, MRMLP97, McF92, Mil95b, MC97b, MC94b, Mon96, MA96, NYH92, NA99, Neu92, NSC^{+}92, Nie96, PST95, RS95, RG93, SF95, ST90, SCEG93, SCDH98, Str91, Tha92a, TR95b, TB92, TGLP95, WF90, dRAV92, AMS99, AKL^{+}99, lYALM93, ADR95, ARDL97, AI92, Ang91a, ACC90, Arr94, AAW98, BFF96, BNvB90, BH94a, BH95a, BP99, BGBS91, BCL93, Big92, BB95, BK91, BM91b, BM94c, BL96, Buc91, BEG97, BRZ99, CCV^{+}96, CL90, Cam96, CN92a, CL97, CC90, CC91b, CD94, CPK97, Che93, CLK90, Cli95, Con98]. growth [CD99, CHC94, DS96, DO92, DSJ90b, DGHT^{+}98, DE94, EPL94, EST95, Ekl98, EP92, FML^{+}95, FPR94, Fev92, FF99, FK95b, FFF^{+}94, FM95b, FQV98, Fow90b, FBR^{+}95, FG93, GS94a, GHMC96, GMMD93, GV96, GJ97a, GCR98, Gar98, Gar99a, GDP^{+}98, GWW^{+}98, GAK99, GMK97, GB90, Gil97, GGC99, GKRD96, GM98a, Gol99, GAB90a, GAB90b, GR94, GBJ94, Han99, Han97, HR96a, HSBH96, HJH^{+}99, HKK^{+}95, HS97b, HDB95, HRD92, HSN93, HDK93, HNL94, JH97, JD90, JR92, KS98, Ker94, KS95b, Kir90, KL91c, KG94, KRJL96, KBS98, KKT96b, Laf93, LS96a, LA97a, LS96b, LC97b, LCC97, LL97, LSB95, MJW96, MC91a, MT91, MYYT96, MD98a, MCM^{+}97, MHC92, MJ91, MC91c, McG97a, MM97b, MEC92, MF96, MAH94, MNB94, Moe96a, Moe96b, MG91b, MGC95b, MV96, MNO96]. growth [MO95, Mun93, Nak92, NT94, NEIR95, NAS95, NBM94, Nor92, OH95, OSJ93, OFR94, OPS96, ORD98, OTF95, PMMH95, PB92, PB96b, PB97, PMB99, PDF97, PMG^{+}98, PD93, PSSP95, PB98, QC94, QQ98, RGLD^{+}98, RR92, RT95, RS94, RCG94, Reu98a, RSN90, RC95, RHB96, RFN98, RPL96, Rv92, Rin92, Roa93, RE96, RT98, RWJ99b, RM95, RH97, Row90, RO97, SK99, SK98a, SK98b, Sci91, SA94, Sep94, Sew94, SD94, SB94, SDC90, SW98b, SG96, Sko96, Sog92, SS91, Sto96, SS96b, SC93, Tak95, TDK97b, Ter97, TS90a, TRS91, TCRB93, TA90, TN90, TVG96, TFFS90, TB95, TB97, US98, VR93b, VWS92, VWS93, VAD^{+}95, VD95, VDM^{+}97, Wah96, Wah97, WSM91, WE95, WDB^{+}95, WtD90, WC92, YNYK94, YAG92, YB90, YY91, ZHM91]. growth [vBV95, vRBVB95, vW93, KJT90, Moe97c]. Growth-dependent [Hov92]. growth-heterozygosity [Fev92]. growth-irradiance [GV96]. growth/optimal [vW93]. gryllus [CPT90, TW98]. grypus [BH94b]. Guild [HHW90, Duf92, Gib94, Sym95]. Gulf [AMBH98, BM94a, BBK91, CBI91, HF99, VS92, WH96, Dee90, MMM^{+}95, AC92, AP99b, BCN^{+}95, Bia91, BFJ^{+}93, BS95c, BW95, Bro94, BLR94, BRW95, CLGM96, CLD99, DSG^{+}99, DE90, DRP^{+}97, FML^{+}95, FAHW99, GDvS98, Gen92, GAN^{+}94, HS92, Hei92a, HD97, HD91, HBP^{+}93, HB94a, JJTS97, JTB99, KL91c, KHK^{+}99, Lev91, LFR^{+}97, LPW95, LGMD98, MMM^{+}95, MMGIM90, MDSB92, MZFHC96, MRP90, NB90, Neu92, OTF95, OR95, PWS98, PH91a, PT95, PM91a, PSSP95, PF91a, SL91, SS96b, SK95c, TMKL97, TK96, Uib95, VH93, WH93, WPG99, WOM^{+}97, YST^{+}97]. gull [Oro96, OJR96, OP99, SM99a]. Gullmarsfjord [AS94b]. Gullmarsfjorden [Han97]. gulls [PE99]. gunnari [KWEG94]. Gut [ICG96, LUA99, BH92a, CM99, CMT94, DPO91, FGG97, HP98, LPF94, LMTF96, PPB90, PM94b, SH98b, TT90, WDD92]. gutless [DAE^{+}99, GCGS90]. guts [KB94]. Gwen [CD98]. Gymnodinium [DH90, FPFR92, JH97, KMR98, MHT^{+}97, MKV95, NSH95, Nie93, Nie96, Str91, VPGV94]. gypsum [OKK95]. Gyre [SK97, DCBG99, STK99]. Gyrodinium [BN91b, Han95, NYH92, UYMH95, VPGV94].

H [PdG96, MKCL94, AGP98, BR91a, SJMC93, WR91]. H24 [Moe97a, Moe97b, Moe97c]. H3 [Moe96b]. H3/1 [Moe96b]. Habitat [AMS99, And97, BCC95, CC98, Cro98, FIA93, LAK99, MJC97, SGP99, SP99c, WM95, AGP98, ARF92, Ang95b, AS94b, BL94a, Bec98, BM94d, BS93, BGG^{+}97, BKB94, BvKvRB96, CSR^{+}92a, CK99, Con94, DBN^{+}94, Egg95, FSCE92, Fer99, FAHW99, Gre96, Gri94, HM91, HBR^{+}91, HKR96, Har98, HFJS91, HH90, IBB97, Joh99, JDLL95, Kin95a, LM98a, LBDD97, LB96b, ML92, Mic96, OM90, ÓM92, PODL99, Pen90a, RK95b, RS95, Rei90, Ris97, RAGH94, RHP93, SDBW95, SU93, Spr93, SJC99, SEAI96, Sym95, THCD95, TJRG93, TB97, ZDJ^{+}95, ZAN99]. Habitat-mediated [LAK99]. habitats [AB90, AADSCC97, Bec98, BCNC96, BDR^{+}96, BvB94, FAED97, FIA93, GVGG98, GSD96, HP96, Kne91, KW94, LJ95, LCLC96, LA97b, LGL^{+}95, MF97, Min93, Pre95, RK95c, RR93, Sog92, Tay98, THPC96, TGLP95, VVT97, WS91, WR98]. habits [CA92, CE93, FW91a, FMRC93, JC92, KI97, Kyo92]. hadal [Fra93, FKT^{+}99]. haddock [Cam97, FQV98]. Hadramout [OB94c]. haemastoma [DS93b]. Haematopus [CGCLH99, McG97a]. haemocyanin [NH98]. Haemolymph [HS94a, SUH95, WJD93]. Halichoerus [BH94b]. Halichondria [RTJ^{+}93, TR95b]. Haliclona [RTJ^{+}93]. Halicryptus [ABN98, Oes90, OV92]. Halifax [JNCS89]. Haliotis [AAE^{+}96, LK93, MS91, Sch93]. Halo [BGC94]. Halo-dule [BGC94]. Haloa [Ito97]. haloclines [VY96]. Halocynthia [RCG98]. Halodule [DT94b, GMR^{+}94, OB93]. halogenated [FWWL99]. Halophila [NA99, Ral98]. halophyte [GH93]. halos [SR94]. hamata [Øre90, Øre95, TNN93]. Haminaea [GC95a]. hamnerorum [CHFL95]. handling [BDH94, EG97, KS99]. hanleyanus [DdA95]. hanseni [BGC91]. haplotypes [PCM^{+}95]. Haptophyceae [BEH93, LGdC^{+}90]. Haptophyta [IG96]. Harbor [Kel98, SCCS98, IFL97]. Harbour [JNCS89, Sch96, BCA^{+}99, BP98, HWH98, TTP^{+}95, WT99a]. hard [lYALM93, lYAML93, HFJS91, LOS90, SL90]. Hardangerfjord [PGG96]. hare [dNSR^{+}96]. haremic [San93a]. harengus [AH94, AH99, Bus96, FF99, MHB94, Mun92, SGM98]. Hariot [PW94]. harmful [CNJ96, RHS99, THS94, ZD99]. harmless [JKT^{+}98]. harp [LADS98]. Harpacticoid [FH97b, BDdWC96, CM90b, GCC95, McC92, ÓM92, ST98, TWWE95, Web91a, Web91b, WP92, WMH96, FS90]. Harpacticoida [VDA96, VDJA98, WC91]. harrisii [FFT98]. harvest [LMT99]. harvestable [ZBW92]. harvesting [BD99]. harvests [MM95b]. hatch [GS99a, WB92b, ZHM91]. hatch-check [GS99a]. hatch-date [ZHM91]. hatchdates [FQV98]. hatchery [SH97a]. Hatching [AN97, Yam97a, YB90, CPL^{+}96, KVK98]. hatchling [MBB99b]. Hatteras [BYO95, MRS96]. Hatteras/Virginia [MRS96]. Hawaii [CGNH98, Gri94, KRJ^{+}95, Sea90, WCL97]. Hawaiian [BK94, BSY^{+}99, Cer98, CGNH98, Rin92, SBC^{+}98]. hawksbill [CL97, HS91c]. HCB [MCM^{+}97]. head [HH91b]. healing [DES92, vW98]. health [KCV94, LFG98, She94]. heat [ADR94, BLD94, KWHM^{+}96, KBA^{+}97, MBL^{+}98]. heat-inducible [KWHM^{+}96]. heat-killed [ADR94]. heat-shock [MBL^{+}98]. heating [KLR93, RHS99]. heavily [FCH99]. Heavy [MONR90, ROM^{+}91, RMGB^{+}94, DE91, HHH92, Reg92, RGBJM95, SGW94, SKM95, SPB^{+}99b, VCMP93, WTF96, ZCL97]. Hediste [Vis90]. Hedophyllum [MD98b]. Height [AI92, HvdM97, HS99b]. Helby [Mga92]. Helena [GGB^{+}99]. helgolandicus [BM99b, FP96a, PF96, PHHH96, PLIM95]. helgolandicus. [LPI^{+}95]. Helice [MDRW99]. helicina [GH91]. Hellenic [CMA95]. help [MPS96]. helpful [CNJ96]. Hematin [Vis93]. Hemiaulus [Vil91]. Hemichordata [Mil92]. hemichordate [FT95]. Hemigrapsus [EDP^{+}98]. Hemisphaeramminidae [AC96b]. Hemocyanin [EBM93, CCC94]. hemoglobin [Vis93]. hemoglobin-containing [Vis93]. Hemolymph [CC96, CC92b, CNCS93, CCC94]. hemprichii [SAvD^{+}96, SMH97]. hentzi [Har99]. Hepatic [EGKE92, HHH92, RA92, GCKW95, SGB90]. hepatocytes [LME92]. hepatopancreas [PGD^{+}93]. herbivore [DCM90, DC91, VEW99]. herbivores [BCC92, McC96b, Tro94]. herbivorous [Cli95, HK96, HKR96, HR96a, KW97b, MA99a, MP92, OC95, OM99a, QK93, SRW95b, Str91, TT90, vRBVB95]. Herbivory [CDM^{+}96, DKL94, ALBS96, Dye93, McG97b, NC94, PCLÅ97, Poo94, Put91, Uit96, VH99a]. herbstii [LK94]. here [OS96]. Herklotsichthys [MBR94]. hermaphroditic [HHG99, RB90]. hermatypic [FCC91, LL96, MNK^{+}97, TTL^{+}96]. hermit [Asa91, Bar97b, Bar97c, Bar97a, Bar99, Har96, RKH96, WG97]. Heron [Gru94]. Herradura [CW93, WC92, WS92b]. herring [AH93, AH94, AH99, Bus96, FF99, FAVV98, FM95b, GHMC96, Liv93, MR97a, Mar97, MHB94, MBR94, Mun92, PG90, SGM98]. Hervey [COK99]. Heterocapsa [JNJP94, KA97, MUH97, NP90, UYMH95]. Heterocarpus [RE96]. Heterogeneity [SLB^{+}92, AB96a, BCC95, BN91b, BB99, CB97, DG96, Gra91b, GB98, KTW^{+}97, LB99b, LB96b, MS96a, PWM95, PWP^{+}97, RNP^{+}98, SNC^{+}98, TF98]. Heterogeneous [RCG99a]. Heteromastus [AGP98]. heteropod [Sea90]. Heteroptera [CR98]. Heterosigma [KYI^{+}98, YKK^{+}98]. Heterotrophic [BW95, CG93, GAN^{+}94, ALBS96, BBL^{+}95, BH93, Bus97, CLM^{+}91, CB96, DGT^{+}93, EPL94, Gas94, GDP^{+}98, Han91b, HSRHH90, HBvD92, HNvB94, JH97, JL94, Jeo94b, Jeo94a, JLYK97, KK91, KG94, LJ96, MF90a, MF90b, NK92a, NYH92, NSH95, NH90, RT92, SCB92, SK94b, SM99b, TT99, TSS93, TK91, VB90, VPFL92, WSM91, ZR92]. heterotrophs [CBD96, CVD99, VAD^{+}94]. Heterotrophy [Lew92, CPH^{+}96, Mil95b]. Heterozoslera [BAM92]. heterozygosity [Fev92, PGB91, TVG96, ZGB90]. Heterozygote [RVT^{+}97]. hexanol [Ack97]. Hg [CAATS94]. HI [BS96a]. hidden [BM98]. hierarchical [CW99b, LAJ95]. hierredda [RFNBHG94]. High [ARS93, ADG^{+}95, CPS91, CGP97, CLKO98, EP97, FBMH94, KDKS96, LDV95, ONH^{+}97, SC92b, SL94, CM93, CHFL95, DLT^{+}97, DQT^{+}97, DAS92, Fly99, GPB96, GBA^{+}96, Gra94, HMWN97, HB95, HH94b, Hin92, HW92, Hob93, IHK^{+}98, KGHTN99, KZ92, LCF99, LGdC^{+}90, MF95, NKH^{+}99, Nor96, PPCL^{+}98, PB92, RHH97, RHB96, RTRP^{+}98, RNH99, SRJD97, SSR^{+}97, SRS92, SCEG93, SH90b, Som94, SFG^{+}99, TA95, URHR98, VLKC94, Vet95, WP95, WHK97, ZAN99]. high-Antarctic [KZ92, WHK97]. high-Arctic [RTRP^{+}98, RNH99]. High-density [SL94]. high-diversity [SH90b]. high-energy [CM93]. high-frequency [LCF99]. high-latitude [Nor96]. high-molecular-weight [GBA^{+}96]. high-nitrate [DLT^{+}97, DQT^{+}97]. high-precision [SFG^{+}99]. high-resolution [DAS92]. high-sedimentation [RHB96]. high-temperature [SRJD97]. higher [Nag96, WC90]. highly [DMP^{+}99, HGBW99, MLDM95, PH91b]. Hill [GJ96a]. hindcasting [TRMW95]. Hippoglossoides [RGT98, SFPC97]. Hirondellea [Fra93]. Hiroshima [AKM^{+}99, KYI^{+}98, NAI^{+}93a]. Hirundichthys [COS93]. Histological [KDL92]. histopathology [Vet92]. historical [SR96]. history [AD90, BL94b, CC91a, CRGZ98, DLDL93, EY90, EDP^{+}98, FDZ94, Gar99a, GC95a, GD99, HH96, HHM96, Hoe91, Kog95, Lim95, LGFS99, MAH94, QZA95, RCTH99, Ruz96, Sal96, San94c, SC92b, SFPC97, SM99b, Tak96, TSW97, WT98, WC91, Zaj91b]. Hokkaido [BTKK99]. holdfast [SS93]. Holdfasts [ACLS97, SS95]. holding [Bus98]. Holocene [GI99, IVM^{+}98]. Holoplankton [SW97]. Holopneustes [Ste95]. Holothuria [CJ93b]. holothurian [TTW97]. holothurians [GP90, UK98]. holothuroid [CJ93b]. Holothuroidea [SMTL99]. holothuroids [SDC^{+}92]. holsatus [LS92]. Homarus [BAH94, BS93, BSB91, BBS93, HF99, IW91, JPC97, JC92, KCS94, PWS98, PGD^{+}93, SSW^{+}96, SCJ99b, WS91, WTO96]. Home [Zel97]. homeostasis [VCMP93]. homolog [KBA^{+}97]. Homoscleromorpha [MSCTBE96]. Hong [DZ99, Lee90, YHC^{+}99]. hook [OC93b]. Hoplostethus [BK92]. Horizontal [MLTD96, LKM^{+}99, WW90]. horse [NT96b]. horseshoe [BLJ92]. Host [AMFB99, BA96, Duf92, CHFL95, FGH95, GP98, KYN94, LM91a, Moe96a, Moe96b, Moe97a, MLL98, SJC99, Ste95, SLHC98, TTW97, Tro92, Tro93, dNSR^{+}96]. host-epiphyte [LM91a]. Host-parasite [BA96]. hour [RGS90]. house [SSP97]. houses [HKA96]. HPLC [BCG97, BPC98, BDdWC96, DEP^{+}96, GB95, HH94a, JSGS94, LM97, MMHW96, MHvB97, RCGSN96, TGG^{+}95, VK90, WJM^{+}91, WTM^{+}96]. HPLC-analysis [VK90]. HPLC-derived [TGG^{+}95]. HPLC-detected [RCGSN96]. HSP70 [WKH^{+}98]. Hudson [DFPG91, FPLH92, FPG95, GFPD92, HCI93, MLTD96, RLT^{+}94, SWT99, VPFL92]. hudsonica [RD92, WDD92]. Hugo [EW91, RMT91]. Human [FPB^{+}94, BBR^{+}93, DdA95, Fri97]. Humans [KUK91, SJ95]. Humboldt [CLJOC98]. Humic [DL94a, CSG93, CGTB95, FTBR97, LP99, MH94b]. humic-bound [LP99]. Humpback [Mik97, BSF95, BFS^{+}98, Cer98, CGNH98, COK99, PCM^{+}95, RCA^{+}95]. hundred [Kin93]. hurricane [JHW95, Mum99, EW91, RMT91]. huxleyi [BFF96, BEH93, BLM^{+}95, BvBBV96, FBMH94, MG97, MH96b, NT96a, NGS97, NBM94, SWK98, VKV95, VH99b, WSS94, vBV95]. hyadesi [GTYR97]. Hyale [LT93]. hyalina [HHG99]. hybrid [GS90, GS91a]. hydrobiological [CCS98]. hydrocarbon [SHO96]. hydrocarbons [GAB90a, GAB90b, MCSV95, ONR98]. hydrochloride [WR91]. Hydroclimatic [LIPF95]. hydrocoral [Lew92]. Hydrodynamic [KSV99, AC96b, ASRP^{+}96b, CHS^{+}95, KCS94, LR96, ML90, PD95, TJ90b]. Hydrodynamical [NK92b, BBL93, KHK^{+}99]. hydrodynamics [AJL93, BM91a, EAEP98, GB98, HlS90, ME97, QRSC99]. hydrogen [AOOT94, BFD91, BAOT96, JGT96, JT97, MSvdZH98, WPG99]. Hydrographic [KS97, BGCM95, KL91a, Pal91, SPBS93, SGMS99, ZF97]. Hydrography [BDD92, FP91b, RHrsHG99, TPW^{+}93, WT99b, TB93]. hydroid [CGZ95, NB96a]. Hydroides [BHQ99, BQKC97, GQ97a, HQ99, LQ97, QQ97, QQ98]. hydroids [HBM93, SL97b]. hydrological [AT96, RPMB91]. hydrolysis [AJST98, HDK93]. hydrolytic [BL96, PH95]. hydromechanical [KV99]. hydromedusae [Mat92, Pen90b]. hydroperiod [CBB90]. hydrostatic [AAW98, BG94, TGB99]. hydrothermal [ASF98, AAW98, CJ93a, CD98, DWG98, EPDC99, FCMB94, GSM^{+}97, GPS93, GJ96a, JDLL95, KZK^{+}95, MT94, PSB^{+}97, PDB^{+}97, PDS97, PFMC95, PFMG96, SRJD97, SJ99, SJML99, VTGK99, VVT97]. hydrozoan [Lew91, Pen90b]. hydrozoans [Pur91]. hyper [MS94b]. hyper-benthos [MS94b]. hyperbolae [BMHH94]. hyperborea [SF95]. hyperboreus [Hay95, MS98]. hyperiid [PP96]. Hypernutrified [RNHO98]. hyperparasitized [Wat98]. hypersaline [CLB91, DT94b]. hyperthermia [HMB95]. hypertrophic [AJS95, DvRN99, Kro91]. hypnoides [McG92]. Hypotheses [MHB96, HS91b, MCW^{+}91, MBB99a]. Hypothesis [TA95, AH95, FPG95, GPLB96, JMR91, JSOG91, LK93, LB95, Lew91, RG97, vW93]. Hypoxia [GJVG96, PD93, ASF98, BDH94, BM94b, BMG99, DvCd94, DvdTCC97, DS93b, FRK92, HMB95, MRS96, ND98, NH98, PPFP98, PPFP99, VDVC94, VH96, WW91, WW93, WK92b]. hypoxia/anoxia [DvdTCC97]. Hypoxic [NR94, RHJ91, FGKP95, KCI98, STS99]. Hypsoblennius [Har99]. Hyrtios [RP91a].

Ia [LC95]. Iberian [CNGG96, SZA94]. Ice [CLKO98, PPMF97, AW98, ALB^{+}96, ARS93, ADG^{+}95, BBG90, CMBM99, DM91, DFR95, DAS92, FPG95, FPP^{+}97, GB91, GC93, GFPD92, GT92, GKRD96, Gra96a, Gra91a, GGD99, HCS90, HW95, KCI98, LDV95, LS91, LK99b, MBJ94, MMG97, MG99, MLTD96, MPS95, NT92, PGR94, PDP^{+}90, PLC^{+}91, RLT^{+}94, RKS97, SM91, SCEG93, SBP92, SBP93, SCKS90, Wer97]. ice-breeding [LK99b]. ice-covered [FPG95, GFPD92]. ice-edge [MPS95, NT92, PGR94]. iceberg [GSD96, PBVB99]. icefish [KWEG94]. Iceland [AH96, AH98, Fev92, HBM93]. Icelandic [LM97]. ichthyofaunal [MF98]. ichthyoplankfon [CH93]. ichthyoplankton [HC90, KID91, SS99b]. ichthyotoxic [VPGV94]. ICPMS [CGM95]. ID [CGM95]. ID-ICPMS [CGM95]. identification [BH93, DLDL93, HQ99, PLFSW97, PPO^{+}99, RP91b]. Identifying [MR97a, MCFB97, JHG99]. identity [Aer98]. II [ASRP^{+}96a, Ang95b, AWH^{+}99, BF95a, BT93, BWHW95, BBB^{+}94, BSH^{+}91b, CW93, CP95a, DvCd94, DQT^{+}97, FS90, FMH96, FGB^{+}96, FP96a, GJ97b, GGH95, GJ96c, HKR96, HLH96, HM97c, HH91a, HOM95, JKT90b, LPI^{+}95, LLP94, LDKR93, LKM^{+}99, Moe96b, Moe97b, MR91, MJ93, NBS^{+}97, NTS90, PRM^{+}96b, PD94, Rel96b, Rie91b, SJMC93, SBP93, THS94, VK94, Web91b, Zaj91b]. Iiella [TK95a, TK97]. III [ACC90, BH92c, BS92, HR96a, Moe97c, OMH95b, PT94a, SJCL93, TAJK90, VB94]. Illex [ASRC99, WTR99]. illuminate [WFTH97]. illumination [TUN99]. illustrations [UMD92]. Ilyanassa [DMM99]. Ilyoplax [SCAT91]. image [KCV94, SR98a, SM97b, Tho91, WGD97]. imagery [RK99]. imaging [BWL^{+}99, GP91, SSY98]. imaging-in-flow [SSY98]. imbricate [HS91c]. imitation [FL90]. imitations [DPO91]. immature [Hag95]. immigration [CM98b]. immuno [Bro90]. immuno-electrophoresis [Bro90]. immunoassays [SMF91]. Immunochemical [SKT97, PLFSW97, TLT93]. Immunofluorescence [DLDL93, BLKA93]. Immunologic [SJCL93]. immunological [VPGV94]. Immunomagnetic [AGGKA96]. immunotypes [SJMC93]. Impact [Arr94, AB96b, BS95a, BGvDR96, DL93a, FHZ96, GB97a, GB97b, GR91, GSD96, HTG96, HVFHH92, HK92, HG92a, KF97, LGR95, MLC97, ÓN97, SW93, TPHC91, VSSM98, WPG99, ZHF96, Atk96, BH92a, BLST90, CM98a, DL91a, Ekl98, FPR94, FW97, GHD92, HHW90, KW99b, KL90, LLP94, LKC95, LME92, Ole95, Øre90, PGG96, Per91, PM92a, PR92, RPL96, Sew96, SSM91, SSGD96, SGB90, THCD95, TGSN99, Ved98, Zaj91b]. Impacts [DLT^{+}97, DQT^{+}97, PPF95, Pre95, SCC94, TA93, FC90, JDSB99, PBVB99]. Imperfect [LKF95]. imperforata [PCG96]. implant [BJB95]. implication [KPS97, RR99, TT98]. Implications [RWJ^{+}95, Aks96, ATNW97, ACOK95, AMB^{+}96, BRLM90, BSMS95, BCL93, BF99, BSB91, BM90a, BM90b, Bra93, BLD94, BBCB99, BvKvRB96, BD94, BHH^{+}95, CLC^{+}98, CGNH98, CT97, CA95, CDGF^{+}93, CLD99, DPM96, FNF91, FHA99, Fra95, GHED^{+}94, GA97a, HC93, HL98, JRMN^{+}97, JHG99, KR97a, KFD99, KH98, KML91, Lee90, LS92, LL98, LBH91, LH99, MC91a, MA93, MGM96, Mil94b, MSND^{+}99, ND98, NKV94, Olm94, PS92, PGAJ99, PDB^{+}97, PLP92, PAM96, QCM97, RTRF90, RHT91, RAB^{+}92, RM97a, Rie98, RPL96, RBG94b, SCA92, SH98b, VT92, WL97, WFH99, ZN96a, ZN96b, ZAN99, tor93, Gon92, CCG98]. Importance [GKMT99, HFW99, Kro91, Liv93, Nan96, NT96a, RGBC92, Sim92, YGH^{+}97, BN95, BEU99, Bre96, CF94, EB96, GF99, HBR^{+}91, Han91b, HR96b, JS97, JG94, Lee97, MPvM98, ML90, MJ92a, NKGK92, NFM98, PB96b, RSN^{+}99, SFF98, Tay98, TPHC94, WW96, WYY^{+}99, WR92a, van92a]. Important [LMM98, AP97b, Bar97a, HM91, HL98, PDD^{+}99, SM90, Weh91]. imposex [BQR99, DHB^{+}98, IWF^{+}97, RQB98a, RQB98b]. impoverished [MF97]. impress [GMC91]. imprinting [AMFB99]. Improved [BCG97, HP98, VLKC94, WJM^{+}91, EGKE92, Mal98, VAR94]. improvement [PPR94]. Improvements [Cle93]. Improving [KRM94, SR95, Str99]. inactive [GdGMD95]. inaequivalvis [Vis93]. inapplicability [TA95]. inbound [FH97b]. incidence [BQR99, DD90]. incident [WTA^{+}97]. include [SJMC93]. including [PPB90, YAG92]. Incorporation [BHT^{+}92, JG90, Kir92, RBJ90, SMP97, AV95, JMP90, Tuo97]. Increase [Jos90, TBAB91, DHM92, HP97, LMK93, SOW97, SNJV92, Wah96]. Increased [PG94, Wah97, AOOT94, CRH^{+}99, KR97a, MV98, NJSS91, SH99a]. increases [McC94a]. increasing [WC95]. increments [Big92, CA96b, Fow90b, LT97, Sep94]. incubation [CDGF^{+}93, GV96, HRKR93, MFD98, MG99, MA96, WSP^{+}95]. incubations [GKRD96, MKCL94, TLB93]. independent [Aro92, HS97b, Mye95, Rv92, RO97]. independently [NNR95]. Index [Ano96a, Ano97a, Ano97b, Ano98, Ano99b, BAY91, BSW93, CWB93a, CD96, COG97, FCC91, HF99, ISF^{+}98, MIA90, Reg92]. Indian [DJY90, DLJ91, Nai93b, PFH98, SPS^{+}96, BFS^{+}98, BSP98, CBG^{+}99, CW99a, DB96b, HFR^{+}93, MACR98, NSC^{+}92, Pak95, PWM95, WSC^{+}99, YD90]. indicate [ASN95, MCFS96]. indicated [FCW98, SCMM95]. indicates [RTLG97]. Indications [KWD93]. indicator [CL90, Cla90, FJR^{+}94, Kai96, MM93, PMETM96, RM97a, SKPT97, WTO96]. indicators [AO91, CNGG96, CW95, CMBC98, FMFH97, FQ94, HJA99, LGC^{+}93, MIA90, SDFJv91, SGR92, WDB98]. indices [BZ92, HR96a, KDL92, SH90b, SKM95, SGB90, SFF92]. indicus [Jon93]. Indies [AGW93, HS91c, HH91b, SC96a, SB96, Tom91]. indigenous [CKP98, EDP^{+}98, MMBT97, Reu98b]. Indirect [CK98, ER90, SM99a, JHN99, JBH98, MD98b, NAW99, RM97a]. Individual [BP99, JBH98, RD95, BCBR97, CN92a, CH97a, Cle93, FQV98, HHM96, Mat92, RSN90, RD92, RCC^{+}99]. individual-based [CH97a, HHM96, RCC^{+}99]. individuals [BA94]. Indo [ALOT99, BDA^{+}99, LCE97]. indole [MH93b]. indole-3-acetic [MH93b]. Indonesia [EM93, SAvD^{+}96]. Indonesian [SMH97]. induce [FBR^{+}95]. induced [BBR^{+}93, BH97b, BAY91, Cla90, DRGV91, GHMC96, JCP93, LY97, LLB^{+}95, LFD95, Moo92b, MSND^{+}99, NK90, NB96b, RO99, Sun96, UKR^{+}91, VWJ96, VM98b, WNJ91, WJM^{+}99]. induces [GCKW95, LW95, SHT91]. Inducible [FFH96, KWHM^{+}96, PKY96]. Induction [BHQ99, BQKC97, HQ99, Uye96, WKH^{+}98, BK90, BEG97, CCJC91, CKP98, HDD98, JMPH90, JP96, JMR91, Ste94]. Indus [BNK^{+}97, HKY^{+}97, KDAS92]. industrialised [SGB90]. inequality [San95]. inermis [CPS91, DS96]. Inexpensive [Sty97]. infauna [FWWL99, LD95b, RCN95, SFV96, Vet96a]. ’infaunal [Tok95, Dit96, ELH92, GCLS94, GCC95, GS99b, KSRB99, LAJ95, ML91b, MCSV95, RNHH97, SG96, THPC96, VWJ96, Vet96a, WX97, WH96, vCRS92]. infected [BCBR97, Mül91]. infecting [SC93, SC95b]. infection [CPS91, EP97, MJ97b, NBS^{+}97, PF91b, SKKO97]. infectivity [SC93]. infer [CGA95]. inference [Mil98]. inferences [FCMB94, GBB^{+}98, KFD99, PN99]. inferred [BM99b, EJ93, GS99a, KPWL93, RG93]. Inferring [FH99, DPO91]. inflow [RR97b]. Influence [ACOK95, BSS92, BM91a, BF99, BPH95, BLC91, CL90, CE93, CLW99, CB97, DM96, DvCd94, DM91, DL94b, EJ97, ES99, FB98, FSGG96, GCM98, GSBB97, GSB98, GFH96, HG99, HCI93, Hüt90, KHW^{+}99, KO92, LB99b, LD95b, LW91, LTM90, MP96a, Mal96, MM95a, ME97, NB96a, ORD98, OJR96, PS95, SOSV95, SPCF97, SPB^{+}99b, SEGP91, Tak99, VDVC94, VDM90, WTR99, dG91a, BB99, FK95a, FTP99, Gar91, Hol97, HM94b, JBK99, KD99, KW97b, KL94, KHK^{+}99, LB96b, MRMR92, MACR98, MV98, MDTB94, MGC95b, Nak92, PCM^{+}95, RCTH99, TN98, VH93, VNM97, WF97, WLT^{+}95, Wol92a, WB99, FP96a]. influenced [HvdMBK98, TCRB99, WHG^{+}99a]. Influences [CGTB95, CCG96, SFA98, HL98, Hsi95, KID91, McG95, Mil94b, SMB98, YBG96, YBG^{+}98]. influencing [Adj97, GS91a, GBH96, HS96, HSB97, KW97a, LP95, MK93b, SJML99, TF98, Yam97b, YGH^{+}97]. Information [Gar98, BM98, BGT91, CU99, CW99a]. infralittoral [SB97a]. infrared [VAR94]. infratidal [Las98]. ingested [Bus96, LAA93, WFL95]. Ingestion [FG93, LHL97, ADBH99, BH92a, BvKvRB96, CGA95, FF99, GS93, HH96, MLP90, MHvB97, PCBA94, SH98b, TDD94]. inhabited [vDdWKW99]. inhabiting [FK95a, Poo94, TRU91]. Inhibition [MH93a, Sut92, UKR^{+}91, AB96b, BP96a, DBvA91, LPI^{+}95, LKvRH96, MSC95]. inhibitor [KG95, LCL95, WK93]. inhibitors [LQ97]. Inhibitory [WCT95]. Initial [vTS97, BM91a, BM91b, GCWH90, MC91b, RD95]. injection [MK93a]. Injury [GBJ94, BBS93, KDL92, Moo92b]. Inland [NSHF93, NSH95]. Inlet [GGD99, HC90, PLFSW97, HvB91, KDMR97, RDA97, US95, US98, HTS92, THS94]. inner [HFN91, KR92, PLB95]. inner-shelf [PLB95]. Innervation [BP93]. Inorganic [Gol99, RT93, AFP99, BC93, BK94, BHHL96, BFD91, BvBBV96, FBHH93, FGB^{+}96, GST^{+}91, IG96, Kro91, LC95, LLB^{+}95, NJSS91, OH95, RPT95, SR95, SD93b, ZNH93]. input [EPN^{+}99, Pel96, SSR^{+}97]. inputs [AC96a, DLT^{+}97, DQT^{+}97, LFR^{+}97, MMM^{+}95, MV98, RCS^{+}96, SJ95, SM97c, Whi94]. insect [CR98]. insensitivity [KB91]. inshore [ALOT99, AALT99, BP98, GBP^{+}97, GA97b, HS99a, KCS94]. insights [Hob93, HAR95, LBJ^{+}98, RSM^{+}96]. instability [AS95, SKPT97]. instriatum [UYMH95]. instruments [And98, LW94]. intake [BGG^{+}97, MCNB98, WC92, BBB^{+}94]. integer [AJS95, MAH94, RWJ99b]. integrated [KEvN95]. integration [BVP99, TDK97b]. integrative [Cha92]. intense [PBA^{+}95]. intensity [HN97a, Kai96, KWEG94, KLJ94, LC90, MSD98, Mic96, PB98, SW98a, Sko96, TP95]. intensive [ASRC99]. Inter [PCT^{+}97, RMH96, TK95a, HLP^{+}99a, HK93, JHN99, SCB92]. Inter- [RMH96, TK95a]. inter-annual [HK93]. Inter-colony [PCT^{+}97]. inter-ecosystem [SCB92]. inter-nesting [HLP^{+}99a]. inter-tidal [JHN99]. interact [PB94]. Interaction [Ste95, Aro92, Cal95b, Dor90, LMS^{+}93, MJ93, QZA95, RO98, vKSGvA99, LGP95]. Interactions [GTYR97, Phi94, Tro93, AvS97, Aer98, AdN91, AMB99, AR91, AS97, AWH^{+}99, BP97b, BBC^{+}97, BCNC96, BH92c, BP92, Bro96a, Bro96b, BA96, CH96, CCSN93, Dai90, EM93, FNJ^{+}94, FD92, FCMB94, Fly99, FBR^{+}95, GBP^{+}97, Gra96b, HPF92, HSCC96, HHE92, JBHH97, KLAH^{+}94, LCK^{+}96, LME92, MC96b, NW97c, PML93, Pen90a, PMG^{+}98, RPT^{+}96, Sch90, SCLS94, SGP99, TR95a, THPC96, UT90, WLB^{+}99, YHGB96, ZC92, dG92a]. Interactive [Gla99, PKY96, RM97a, CWWH99, RD95, VÅCL99]. Interannual [AB90, BWM97, GJC97, ISF^{+}98, MMGIM90, PF98, Tan98a, Tan98b, WWB98, ASN95, CPB96, HM94b, MP96c, LTM^{+}98]. interception [ML90]. Intercolony [VF99]. Interdependence [FAP92]. interdependent [GMG^{+}92]. interface [BCG94, BR91a, CRCS96, DFR95, FGP98, GSBB97, Kro91, RLT^{+}94, TGvJW95, TUN99]. interfaces [Sim92]. interfacia [HG92a]. interfere [RCG94]. Interference [WC96, CWWH99, Hoe91, ML91b]. Intergeneric [Bro90]. interlinked [MvdVD^{+}97]. intermedia [Kyo92]. Intermediary [DvCd94]. intermediate [CLB92, MO95, PBA^{+}95]. intermedium [Gra91b, Gra91c]. intermittent [VBF94]. intermoult [GAV93]. Internal [BMHD93, TCB93, ASP91, KTW^{+}97, LSGW98, LCF99, PPFP98, RSN90, SR90a]. international [CSR^{+}92b, Cha92]. interplay [TPHC94]. interpolation [Zho98]. interpopulation [AHT^{+}90, PB98]. Interpretation [BAY91, FPFR92, WT98, CU99, HF99, HH96, Rel96a, Rel96b, TMC98]. interpretations [WTM^{+}96]. interpreting [FIL93]. Interrelations [AM91]. Interrelationships [SRW95a]. Interseasonal [FTS94]. Interspecific [Dai90, NTK93, PC96, GP98, HK97, SJC99]. interstitial [GF99]. Intertidal [HN93, JMC93, JA91, MBB99a, AOOT94, AGP98, AC96c, AC99, AB96a, BKP98, Bec98, BF95b, BUWB99, BGR^{+}96, BB99, BM94c, BLST90, BDSL94, Cam91, CH96, CV98, CMD^{+}94, CB97, DF94, DKVD95, Dye93, EBB95, EGC95, FA99, FAED97, FIA93, FW91a, FTP99, GB95, GC95b, GQ96, GR91, GBGH97, Gün91, Hic92, HdWS91, HvdM97, HG94, Hul97, Hun93, IF96, JP93b, KLEH91, KD96, Kne91, LFRB97, LGV^{+}99a, LGV^{+}99b, LB96a, Leo99, LGED98, LTM90, MCC96a, MGC91, MM97b, MKN^{+}96, MS93c, MDRW99, NA99, NT94, NC90, NC93, OM99a, OB94c, PB96a, PH91b, Pel96, PZ91, PSPD96, QK93, RS99a, RGLD^{+}98, RK95b, RK95c, RB99, RRGB96, Roz92, SPB95, SS96a, Spr93, Tak96, TBY98, TJRG93, Tru97, WH96, WC96, ZAN99, vDK90, vDdWKW99, vBDW91]. intertidal-height [HvdM97]. intertidally [Yam97a]. intervention [PB94]. intestinalis [HlS90, PR92, PST95, SHW^{+}91]. Intra [PRC^{+}92, ORL97, RMH96]. Intra-annual [PRC^{+}92]. intra-colonial [ORL97]. intra-population [RMH96]. intracapsular [CP90]. Intracellular [LL97, LJS90, NL96, FF92a, FFJ93, FJR^{+}94, LL92, NBM94, Sim91, VMHH90]. Intracolonial [BPS98]. intraguild [GA97a, Val91]. intraovarian [Sew94]. Intraplant [MP92]. Intraspecific [CWWH99, GP98, VB93, TK95a]. introduced [AW97, BB97a, CC98, JPZ^{+}98, KGO94, KMB97, LLB92]. Introduction [Bro95b, SD92, MHT^{+}97, CTSN90]. Intrusions [Gra96b]. inundation [HKvB93, KW94]. invadens [JNCS89]. invader [LFBH94]. Invasion [SBC^{+}98, CTSN90, LE90, NTS90, Reu98b, YTF97]. invasive [AAD91, BB97a, CUM98, CC99]. Inverse [NJLD98, EJ93, vdVW99]. invertebrate [And90, BLH99, DHS96, DL91a, FAHW99, GBV96, JMR91, JS97, MM99, Pec99, PS92, QRSC99, RW96, TD94, WMB99]. invertebrates [AT96, BCAAC96, BC99, CCG^{+}92, CRGZ98, Con95, Fra94, GQ97b, HB97, Hav93, Hol97, HS97c, JRMN^{+}97, MB97a, MCSV95, MHD97, MJ97b, Oes90, RHT91, Rie98, RSN^{+}99, ROI93, SFWD90, Sty97, Tay98, TRU91, WW96]. investigate [GBGH97, RVGC^{+}99, TGvJW95, TSW97]. investigated [BP93, HHN93, HBBT98]. investigating [BH97a, Hil91a]. investigation [AJ95, GPLB96, Ken94c, Sün97, WA99b]. Investigations [Moe97a, Moe97b, Moe97c, FRK92, KR92, VAR94, Moe96a, Moe96b]. investment [Geo94]. involved [SP99c, Ter97]. ions [BFD91, RPT95, van92a]. iradelei [HSTY98]. Ireland [HK99]. Iridaea [GW91]. iridescent [CN95]. iris [Sch93]. Irish [NG99, MTP92, PF98, TH90, TGSN99]. Iron [TvLdJ^{+}98, BMV^{+}96, Bra93, DH90, DEP^{+}96, EPD^{+}99, FF92b, MLH96, MH96b, SDH95, WMG91, dBN^{+}90, vLTW^{+}98]. iron-stress [DEP^{+}96]. irradiance [AdN91, BCH95, CPH91, DL93b, GV96, IPT94, KLJ94, LS91, MFD98, MKV95, MC97b, MP95, Nie96, PLC^{+}91, ZCPA94]. irradians [AV95, AI92, AMBH98, PBGE91]. irradiation [GJ96c]. irrigation [MSK95, PNB94]. irroratus [CC97b, PWS98]. Ischia [CJ93b]. iserti [Cli95]. isetus [KHT93]. Island [Bar97b, Bar97c, Bar97a, Bar99, BH94b, CC91b, CC97b, CJ93b, CHDW92, DG93, GWKW97, GBP^{+}97, Gru94, HSK^{+}93, HH90, HHE92, JFP95, Ker90, KZK^{+}95, MRP98, Mga92, MD94c, MKYI98, MN99, Nai93b, NKGK92, PBVB99, Sha95b, TDD94, TSM^{+}94, TZ99, AALT99, CC97b, HLP^{+}99a, HGB99, HD93, HCI93, JBO97, KSO^{+}97, LCK^{+}96, MMBT97, RA96]. islandica [AH96, AH98, Fev92, HBM93, Oes90]. Islands [DH96, HSH94, HB95, HCH^{+}96, KID91, MSM^{+}98, PLM92, SB96, PPCL^{+}98, BDR^{+}96, BSY^{+}99, BW96, Cer98, CGNH98, CCG^{+}92, CWH98, EW91, KZK^{+}95, LK93, PF95]. Isle [NG99]. Isochrysis [Tha92a]. Isognomon [Rin92]. isolated [AAW98, Fra93, HBBM96, KRJ^{+}95, Lip95, PGB96, VAD^{+}94]. Isolation [SHT91, HQ99, SRS92, SC95b, ZHBH90]. isopod [CCD94, HS90, HS94a, San94a, Vis91]. Isopoda [OR95]. isopods [GP98]. Isotope [Sim91, BRZ99, CLC^{+}98, CMBC98, CNP95, DS93a, DRS94, DWG98, DH92, EF97, FCMB94, FQ94, GTWF98, Hob93, HSCC96, JRMN^{+}97, KPWL93, KRJL96, KRJ98, KR98b, LKvRH96, MO97, MJSDF91, NG96, PIC97, PSB^{+}97, RRGB96, SBV98, SDFJv91, SM90, TC93, VNM97, WYY^{+}98, WF90, vR93a]. isotopes [HM96, KSRB99, RSN^{+}99]. isotopic [BBO^{+}95, CDGF^{+}93, FCVS92, Fra94, GT98, Kal91a, KM99a, SR95, SCC98, WYY^{+}99, WCLB99]. isozymes [SWK98]. Israel [KEvN95]. Israeli [EBB95]. issue [KDC99, Sun96]. issues [BS96b]. Istiophorus [Spe95]. Italy [BCC95, BCNC96, BCBC98, CJ93b, FPB^{+}94, MCM^{+}97, SSG96]. items [CGCLH99]. IV [Asa91, CC90, HHKS92, KH92a, PW94, PT94b, WS92a]. Iwate [YNYK94].

J [Ano00, JNCS90]. Jamaica [SOW97]. Jamaican [AW95, BNW94, NBW94]. James [KKW^{+}93]. January [HC94]. Japan [NSHF93, AKM^{+}99, AM93, BTKK99, FKT^{+}99, FII95, HSK^{+}93, IT96, KYI^{+}98, NAI^{+}93a, Nak92, NTK93, NKKO97, NKGK92, OIO^{+}95, Sak97, TK95a, TK97, TK98, TT99, Ter93, YNYK94]. Japanese [AM93, AOT97, CT96, KDC99, KSH97, OINN94, PLBP96, UT90, YNYK94]. japonica [AHY^{+}99, AOT97, BTKK99, BL94a, CT96, Ito97, TK97, TK98]. japonicus [CC92b, CC97a, LCTT92]. jars [HdWS91]. Jasus [CB99, Mac91]. Java [Hoe91]. Jellett [JNCS90, Mil90]. jellyfish [FFH99, HPM92, OFR94, Ole95, VM98a]. Jensen [DD97]. Jersey [BLJ92]. Jervis [FK95b]. jet [FLDT92, LFTH92, PJ99]. jobs [RD97]. John [EW91, RMT91]. Johor [HTLK99]. journals [Pud93]. ‘Joy’ [VDG95]. Juan [GJ96a, SRJD97, SJML99]. Juncus [CBB90]. June [KLMR91, PD95]. Jurassic [Aro91]. Jurassic-Recent [Aro91]. Jürgen [Tor94]. Juvenile [GQ97b, MF97, ÅP97a, And94a, ACLS97, ACR99, AADSCC97, Arm92, AH96, BS94, BSH^{+}94, BCL93, Beu92, Bro98, Bur94, BHH^{+}95, CHS^{+}95, CMBC98, Con98, CGFC95, ELMCC90, FH97b, GPB^{+}98, GC95b, Gos97, Gra96b, GT96, GA97b, GF91, HBS98, HS92, Hav93, HHK^{+}98, HVL96, HSB97, JL96, KH92b, Ker94, KS99, LJ95, LS96b, LCTT92, LAK99, MBF^{+}97, MT91, MYYT96, MEK99, MJ91, McC92, MLES94, MG91b, MLDM95, Mye95, Nie93, O’B94a, O’B94b, OM99b, PNS90, PBGE91, QQ98, RK95b, RK95c, RGE^{+}95, RSB93, RO97, RO98, RO99, SMF91, Sew94, SDBW95, Smi95, Sog92, SL90, SR97, SCCS98, SFF92, SR94, TNP97, TcBW96, TJST98, TRS91, THPC96, VEW99, Van92b, VS92, VHH^{+}96, VHHS96, VHH^{+}98, WH93, Web91a, Web91b, Whi94, ZAN99, ZFFH^{+}94]. juveniles [BB93, EPDC99, HDB95, HS99b, MPvM98, PODL99, RBB^{+}91, RMG^{+}98, She95, ZC92, OMH95b].

Kähler [Kep98]. Kareius [MYYT96]. Katharina [MD98b]. Kattegat [DFJ^{+}93, HRÆ94, JAB^{+}92, Jos90, JJNR95, KLMR91, RC91a]. Kauai [Cer98, CGNH98]. kDa [KWHM^{+}96]. Kelp [BS93, AHT^{+}90, ACLS97, CO92, DES92, FS91, GHL^{+}97, Hag95, HRFC90, KL91b, LLB92, MD98b, Row90, SDER96, SS93, Ste95, TD91, TDER97, WH91, tor93]. kelps [ACLS97, HDD98, MC93]. Kennebec [WT99b]. Kenya [HSK^{+}94, SOSV95]. Kenyan [McC94a, SFV96, WO94]. Kepkay [Käh98]. keratose [MY98]. Kerguelen [CW95, KID91, PF95]. Kermadec [CCG^{+}92]. Kerteminde [AWD95]. keta [FS98, SOK98]. Keys [GEW97, OM99b]. kHz [MWA93]. kidney [GLBF92]. kidneys [CDI96]. Kiel [MMG97, AO91, BS95a, Sch96, VS91, Vos91]. killed [ADR94]. killifish [Mal99]. kills [BGH95, BMG99, PPFP99, UYMH95]. kilometer [EB96]. kilometer-wide [EB96]. kin [Pen90b]. Kinetic [WFL96, Fly98, Kal91a, NT92, ND96, vR93a]. Kinetics [BSG^{+}97, NB90, BMFR97, BAM^{+}99, BLCA90, CV91, EPL94, HH96, LWL98, LF93, RPT95, SFS95, TI96, WK93]. king [BGG^{+}97, GKB^{+}97, JCCCB94, MG91b, ROANM98, SK98c, GWKW97, HH90, JWKH98]. Kingdom [ONH^{+}97]. Kinne [Ano99c]. Kiribati [MBR94]. kitt [RFGF92]. kittiwakes [RM97a, WTA^{+}97]. km [SLB^{+}92]. knot [ZBW92, ZB92]. Ko [BLST90, LB96a]. Kodiak [MN99]. kokuboi [TK95a, TK97, TK98]. Kong [DZ99, Lee90, YHC^{+}99]. Korea [OIO^{+}95]. Korean [SH98a]. koreni [TDW96]. kowalevskii [FT95, KM96, Mil92]. Kraternaya [KZK^{+}95]. Krill [AS97, AWH^{+}99, RTCH96, RWCM99, BMHD93, BW96, BWM97, Buc91, CWHA99, CRP99, DS96, HVK96, HD93, HTCW96, HHE92, HNL94, KMKS96, KHFP^{+}96, MACR98, Nev99, NSC^{+}92, OK98, Pak95, PPMF97, PM94a, PWPS95, SL95, STS99, TCRB99, TPW^{+}93, WBP^{+}92, WMD99]. Krill-copepod [AS97, AWH^{+}99]. Kuchinoerabu [NKGK92]. Kurile [KZK^{+}95]. Kuroshio [NIU^{+}98]. Kuwait [ASCW90, SCAT91]. Kyushu [AM93].

L [Ano00, PM90a]. L-amino [PM90a]. L. [MC93]. labeled [ER95b, HBvD92, LBM^{+}99, ORL97, SSM91, WR91]. labeling [GW93]. labelled [NH90, ÓMvdB99]. labelling [CCV^{+}96, CRLP94, SD93b]. labels [Car90]. labial [BSJ97]. Labidiaster [DEF91]. labile [CBD96, SM95]. Lability [LMT96]. labio [Tak95, Tak96]. Laboea [Put90]. laboratories [Ric91]. Laboratory [BCH95, Cla90, FTGW90, SlRR94, AS99a, BHQ99, Buc91, BSH^{+}91a, Bur94, CMRF94, CLM^{+}91, EB92, FCW98, FKB92, GPP98, GF98, HHN93, HMHDA94, HBM93, LCW99, MLP90, OE91, OD90, PGM^{+}98, Ral98, RT95, RBB^{+}91, RMMB91, RFGF92, SLHA97, SFG^{+}99, SEGP91, WE95, WNJ91, WA95]. laboratory-cultured [Ral98]. Laboratory-induced [Cla90, WNJ91]. laboratory-reared [FCW98, RBB^{+}91]. Labridae [Gil97, Gre96]. Labroidei [WB97]. Labroides [Gru96]. laceration [AS95]. Lack [GB90, Ste94, Fev92]. lacks [WS92a]. lactate [DvdTCC97]. lactation [LK99b]. Lacuna [MC91b, MD93]. lady [BA90]. laeviusculus [DL91a, DL93a]. laeviusculus-capelin [DL91a]. Lagoon [BEU99, RMTB^{+}95, AD94, BCG94, CRCS96, CCR91, CLB91, CG99a, DG93, DT94b, FCCM94, GSBB97, IF96, JKL90, JFS90, JFS92, LHRL97, McC94d, MG94, MZCMOCJP91, OB93, PLB95, SNC^{+}98, SSGD96, SGC^{+}98, Tho92, TcBW96, YOH98, ZB98, ZB99a, LFD95]. lagoonal [FB97, SSG96]. lagoons [Bak90, Cha96, CDG97, CB98, GB97a, GB97b, McC94a, MMR99, PPCL^{+}98, SRRdM99, TNGB95]. Lagrangian [GPD^{+}96, LSL98]. Laguna [Onu96]. lake [AJS95, BTKK99, Sca93]. lakes [SCA92]. lamarcki [LAJ95]. lamellosa [FD99]. Laminanales [HRFC90]. Laminaria [DFS94, EY90, EP97, SF95]. Laminariales [BWB98, SF95, tor93]. laminarioides [BCBR97, GW91]. lancelet [Sto96]. Land [PM92a, Mee99]. Land-based [PM92a]. land-use [Mee99]. landscape [VDM^{+}97]. landscapes [BH91, FB98, JAB^{+}92]. Lanice [PE99]. lankesteri [HAA94]. lapillus [BQR99, DHB^{+}97, RQB98b]. Laptev [PS97]. Large [AB92, Gow93, SR90a, She95, TDER97, ADR95, AC96b, Arr94, BH97a, CG99a, CP96, DLFL96, DE94, FBHH93, FB91, GHD92, Gol99, JKFB95, KKT96b, Lew97, MEC92, PG91, RT98, RSK^{+}99, RÁSP99, SLB^{+}92, SH97b, SK97, SRW95b, SR96, Tol98, WE95, Wit99]. Large-scale [AB92, SR90a, TDER97, ADR95, Arr94, BH97a, CP96, FBHH93, PG91, SH97b, Tol98]. large-volume [DE94]. largely [JBWH97]. larger [Cal95b]. larger-scale [Cal95b]. Larsen [PF90]. larseni [Ruz96]. Larus [Oro96, OJR96]. larva [CD96, MC91a]. larvacean [HKA96, SSP97]. larvae [AJL93, BN95, BH94a, BH95a, BB95, BM91a, BM94d, BH97b, BLJ92, BSD^{+}94, BRC99, BS96b, BSH^{+}91a, BSH^{+}91b, BLH99, CMRF94, CBI91, Che93, CCV^{+}98, Cle93, CD96, CCS98, CH90, DD90, DJY90, DLJ91, DBvA91, DO92, Dav94, DHS96, DHKP96, DE90, Dou93a, Dou93b, DFPG91, DG93, DE91, Eml90, EPDC99, FF99, FTS94, FM95b, FLDT92, FPG95, FWF^{+}99, GFPD92, GB93, GBV96, GW97, GS99a, GRSV98, HPPZ93, Han99, HMHDA94, HS99a, HK99, HKM92, HDB95, HOM95, JD90, JMR91, JS97, JAL91, Kne93, LT97, Lam98, LP95, Lav94, LB91, LCE97, LB96b, MH96a, MJW96, MGYV97, MCL90, Mar91, MM99, MPS96, MHB94, MB97a, McF92, Mil98, MBL94, MLDM99, OTF95, PF92b, Pen90b, PS95, PCBA94, PND^{+}94, PB98, RLDM94, RT95, RG93, RBB^{+}91]. larvae [RW96, Rin92, RFGF92, Röp93, RE94b, Sep94, SGP99, SHFP94, SLSH98, SHW^{+}91, TD92, TH90, TS90b, TS90c, VY96, WW91, Weh91, WM90, WE95, WCT95, Wil92, YD90, ZN96a, ZC92]. larvae-are [Han99]. Larval [BPL95, Hur92, KCS94, LTM^{+}98, MMD93, MMD92, MD94c, MLDM95, QCM94, RPN95, SdC97, Sto96, WX97, lYAML93, AB90, AS99b, And94a, AR91, AJL93, AHMd90, Ang91c, Ang95a, Ang95b, ACOK95, AH93, BP97b, BM94a, BCN^{+}95, BHQ99, BCL93, BRC99, Bro94, BQKC97, BLH99, Bus96, BCv^{+}92, CCJC91, CB99, CE93, CTC93, CMT94, CRH^{+}99, Cow91, DE94, EAEP98, FCW98, FQV98, GHMC96, GCR98, GM99, Gar99b, GB90, GQ97a, GPLB96, Gra96b, GF91, GP98, HQ99, HF99, Har99, HC93, HlS90, Hav93, Hil91a, HHP95, HBP^{+}93, Hol90, HM97d, HW91, Hov92, Hun93, Ito97, JMPH90, LB99a, LQ97, LH91, LBL97, LMTF96, MRF^{+}99, MLP90, ML91a, MC91a, MY96, MM95a, MC94a, MF96, ME97, MS93c, MBB95, Mun92, Mun93, Mye95, NvA91, Nor92, OD90, Pal97, PGM^{+}98, Pec99]. larval [PHT96, Pla93, PG90, QRSC99, QGQ97, RDA97, RE94a, Sab90, SO96a, SML97, SDER96, Sha98, SGMS99, SKPT97, SC96a, SM99b, TSM^{+}94, TcBW96, TRMW95, UMTM98, VBCRB97, VDBF96, WMB99, Weh91, Whi94, WB91, WB92b, Asa91]. larval-release [Pal97]. Larvatron [JPR92]. larynx [NB96a]. laser [BP93, HHYD92]. Late [DRC^{+}97, BNvB90, HTG96, KLMR91, LCE97, LKK92, LHL93, PP96, PB92, SJJMH90, Sto97, SB97c, ZN96b, ZBW92]. lateral [WHIC97]. lateralis [CJL91]. latitude [HB95, HFGT94, HS98b, KGHTN99, Nor96, RHH97]. latitudes [Cal95a, Dun99, Hay95]. Latitudinal [CA99a, HHB97, EDS^{+}99, TGG97]. Lawrence [CLGM96, FVDL95, HV91, LD93, LD94a, PR93a, RCGSN96, SGL^{+}97, SK95c, TMKL97, VBF94, VDBF96, HD91, HB94a, OTF95]. lawrencianus [SS91]. layer [BYO95, CWH98, DAS92, FHZ96, FM94, GB91, JAL91, KDR97, KR95, LGR95, MG91a, MHRV^{+}99, PLH91, PWP^{+}97, RTLG97, SS99a, TB98, Ved98, VM95]. layers [BA93, GGD99, Har90, Tak99]. laying [Oro96]. Lazarev [FPP^{+}97, Wöh96]. leaching [Tur93]. lead [Bou90a, Bou90b, CF93a, DHM92, FBL95, GPP98, SSR^{+}97, WH96]. lead-210 [CF93a]. leading [DvRN99]. leads [AI92]. Leaf [TW97, COG97, HS99b, PH91a, TB99]. least [HH90]. leaves [AWH93, EAD92, FFH99, SAvD^{+}96, TS99, VSSM98, WD90]. lecithotrophy [Ang95a]. Lectins [MS95]. Lee [TSM^{+}94]. legged [RM97a]. Leggett [Cus95]. leidyi [CLW99, PC95]. Leiostomus [SMF91, SR97, SCCS98]. lemaneiformis [VBN95]. lemon [RFGF92]. length [GCM98, HV91, HSW99, HKM92, Sal95, YKK^{+}98]. length-specific [HV91]. Lengthening [SB99]. lengths [CW99b]. leopardus [Joh99, Sam97, Zel97, Zel98]. Lepeophtheirus [AS99a]. Lepidonotothen [Ruz96]. Leptasterias [Geo94, RHH94, RMH96, RPT95]. leptocephali [AMO^{+}99, Mil95a, ONM93, OINN94]. Leptocheirus [SDC99]. Leptostraca [Vet96b]. leptostracan [Vet96a]. Leptosynapta [Sew94]. Lesion [vW98, MPB97, OBL97]. lesions [NB98]. lesser [Oro96]. Lessonia [CO92, MC93, WMG^{+}94]. Lethal [KA97, KCI98, WTA^{+}97]. Leu [LBV^{+}99]. leucine [Kir92, RBJ90, vR93a]. leucosticta [FSA^{+}97]. Levantine [RZWY96]. level [ASN95, AdN91, Cro98, DL95, Dye98, Ejd98, FC90, FQ94, GT96, KI97, LF91, MZ99, OP99, ODN^{+}95, SCJ99b, Tak95, UKR^{+}91]. levels [AOOT94, BCN^{+}95, BCA^{+}99, Cha98, CCC94, CC96, CW99b, HHN^{+}96, JCH93, MIA90, MGF98, MAAT98, Nag96, PCG96, RGS90, RK99, SS93, SC95a, TL91, VDVC94, WDB^{+}95]. Libinia [Sha98]. lichen [EGC95]. lichen-mediated [EGC95]. Life [FMRC93, Gar99a, GD99, MAH94, SH90a, WC91, AD90, ASSWM97, BG97, BL94b, CC91a, CRGZ98, CD99, CWB93b, DLDL93, EY90, EDP^{+}98, GC95a, GM98b, Gos97, HHM96, Hoe91, IVM^{+}98, Kog95, KZ92, KGW93, KHT93, LGFS99, LFBH94, MF96, MR94, OW95, Pec99, QZA95, RSH92, Ruz96, Sal96, SFPC97, Tak96, TVG96, VS96, WT98, WASSM97, WHK97, Zaj91b]. life-cycle [CD99]. Life-history [Gar99a, CRGZ98, GC95a, LGFS99, QZA95, Ruz96]. life-long [CWB93b]. lift [Roz92]. Light [DSS^{+}98, EAD92, EJ96, KLJ94, MSJ92, RO99, WH98, AS99a, AC92, AZKA99, AdN91, BFF96, BGR^{+}96, BSP93, DRGV91, DFS94, DL94b, DF95, DH92, FFJ93, GS94a, GYL93, GMK97, GM98a, GWKW97, GT96, HMWN97, HN97a, Har94, HS99a, JM99, JNJP94, KKH98, KL94, KMB97, KJMRC96, KR94, KCD^{+}95, LY97, LC90, LZYW93, MSD98, MG96c, Mil95b, MGC95b, MH96b, MJ93, NT96a, NP90, NPB93, NK92b, OSJ93, Ole96, Onu96, OKK95, PS94, PLB95, PB98, RO98, SPS^{+}91, SFS95, Sko96, Ter97, Tho92, VGJ90, VBF94, WFHM99, WR91, WF90, ZA96]. light- [PS94]. light-dependent [GYL93]. Light-enhanced [EJ96]. Light-induced [RO99, DRGV91, LY97]. light-limited [AZKA99, BFF96, GM98a, NK92b]. light-saturated [BGR^{+}96]. light/dark [ZA96]. ligni [Zaj91a, Zaj91b]. lignin [BRL97]. lignin-phenols [BRL97]. lignir [MWH91]. lignocellulose [MH90]. Ligurian [AFC96, GLD^{+}91, LBL91, MVA99, PF92b, PMB99, PD95]. like [GHED^{+}94, Gow93, Har97, LGP95, MAAT98, NAI^{+}93a, SDC^{+}92, TSI^{+}93]. likely [KGO94]. liliana [HTCP96]. Limacina [GH91]. Limanda [BVL92, CB92, CMLE92, EGKE92, HHH92, KDL92, LACM92, LME92, Pro92, RA92, RVv92, SH92a, SGR92, VBL^{+}92]. limb [Smi95]. Limfjord [RH91]. limit [DY92, EY90, Kin95b]. Limitation [Kir90, ADF96, ARDL97, Ang95b, Bra93, DLT^{+}97, DON^{+}95, EPD^{+}99, FPAH92, FAP92, FF92b, GJ97b, GMK97, GAN^{+}94, HJA99, ICG96, KR94, LBJ^{+}98, MP94, MBK95, ND96, ODN^{+}95, PJF93, PRM^{+}96a, PRM^{+}96b, PB96b, PS94, PS92, PSL96, PG90, RPMB91, SA94, SD93b, TNGB95, TP96, VD95, dBN^{+}90]. limitations [TR95a]. limited [AZKA99, AH99, BFF96, BAM92, CHC94, GM98a, Han99, LS96a, LA97a, MCC96a, Moe96a, Moe96b, Moe97c, MHD97, NK92b, OL99, SK98a, SDC90, TR95a]. limiting [GJ97a, GJ97b, HM97c, MH97, NT92, PWG^{+}99, PT95]. Limits [MY98, PSSP95, BF95b, FDZ93, RPM^{+}94]. limosum [LK94]. Limpet [CGCLH99, JHN99, WM95]. limpets [BCNC96, EGC95, JBHH97, JBH98]. Limulus [BLJ92]. Lindaspollene [ES90]. line [BBR^{+}93, MM91b, O’D98]. linear [BJTB96, SM94, SBT97]. lingua [PPW96]. lining [PL99b]. link [DLHB99, FH97b, LHL97, RW99, SH97b]. linkage [JC95]. linkages [WB94]. linked [BGH95, LFR^{+}97, WLB^{+}99]. Linking [BRC99, MN99, TBY98, CA93, JL96, Nag96]. links [MNB94]. linum [KJMRC96]. Liocarcinus [LS92]. Lions [AC92, YST^{+}97]. lipia [LW91]. Lipid [BS96a, DCRG92, JL96, KH98, PSB^{+}97, ASH93, BMS^{+}97, CTEF91, DO92, FBMH94, HKG93, HVK96, HSN93, LMTF96, MACR98, MVA99, OTF95, PWPS95, SM98a, SCEG93, SGK^{+}97, SFF92, WMH96]. lipid-deficient [DO92]. lipid-enriched [DO92]. lipid-storage [CTEF91]. lipids [GAB90a, GAB90b, NPB94, Ohm96, PMMH95, PDB^{+}97, SM98a]. Lipofuscin [WTO96, JSH99, SSW^{+}96, SCMM95]. lipogenic [KHFP^{+}96]. lipophilic [CLHF95]. Lipophrys [BKH99]. lipped [HJH^{+}99]. Liquid [BKS92, VLKC94]. Liquid/solid [BKS92]. Lissarca [BH92e]. listerianum [RB90]. literature [Rel96a]. litho [FPB^{+}94]. litho-phaga [FPB^{+}94]. Lithophaga [BKWML93, FPB^{+}94]. litorale [EGC95]. litter [Lee90, PH91a]. little [WS92a]. Littonna [WS92a]. littoral [AWD95, BCC92, BCC95, BCC96, Dye98, Mac94, MPS96, MAdG^{+}96, WA99a]. littorea [DB99, MIA90, WS92a, Wah96]. Littorina [Cha95, CCGM93, DB99, DBV98, MIA90, SH98a, Wah96]. littorinid [Cha95]. live [HKM92, LGH95a]. liver [DvCd94, DCG^{+}99, GLBF92, KDL92, LACM92, MGM96, Moo92b]. lividus [FB97, FG93, LTM^{+}98, LGL^{+}95, SZ96, SRH^{+}98, TGLP95, WTF95, WTF96]. Living [Fri97, DB96a, HM91, HH94b, JNCS89, KDKS96, LME92, MG96c, MS95, ML91c, TM94, VTGK99, VVT97, VL97]. lixula [SRH^{+}98]. Lizard [Gru94, MD94c]. load [BEU99, NJSS91, vKSGvA99]. loading [Kel98, PPFP98, SBH^{+}95]. loadings [vBV95]. loads [FF97, Gru95]. lobata [LGH95a, RS91]. Lobophora [ATH95]. Lobster [JHW95, BPR90, BS93, BSB91, BBS93, CB99, ELMCC90, ELG97, ELMR98, HF99, JC92, KCS94, Mac91, MLES94, PGD^{+}93, Roa93, RGE^{+}95, RT98, SSW^{+}96, SCJ99b, WS91, WTO96]. lobsters [AM91, AADSCC97, BAH94, BHH^{+}95, IMB99, CA92, IW91, JPC97, Jer90, PWS98, SCC94]. Local [FW92, HFR^{+}93, ADR95, Cal95a, Cal95b, Har97, Hun93, Kep91, LKC^{+}99, SC94]. localities [Las98]. locally [Bar97a, RSN^{+}99]. locate [RPT^{+}96]. location [AP97b, BSH^{+}91a, HS91c, MACR98, MLES94, SO96a]. locations [GU98]. loch [THR^{+}98]. loci [GS90]. locomotion [AS95]. locus [TVG96]. loggerhead [MBB99b]. loisae [DAE^{+}99]. Loliginidae [CF93b, SB97b]. Loligo [CF93b, SB97b]. Long [AT96, BLJ^{+}96, DBHM99, FCH99, FSG^{+}98, Har94, HP97, HWL96, HGB99, HB98, JCMW91, KDHZ98, LGGLLHV98, MP96c, Oes90, Pel96, AS94a, CWB93b, DvCd94, DdA95, DFB^{+}99, Dye98, FSCE92, GRC^{+}97, LB91, MDD95, Onu96, PRC^{+}92, SMT97, SC94, She96, SGW94, SFPC97, TD91, TFL95, VDVC94, VPM96, CC91b, JFP95, LCK^{+}96]. long-distance [SC94]. long-finned [AS94a, GRC^{+}97]. long-range [LB91]. Long-term [AT96, BLJ^{+}96, DBHM99, FCH99, FSG^{+}98, Har94, HP97, HWL96, HGB99, HB98, JCMW91, KDHZ98, LGGLLHV98, MP96c, Oes90, Pel96, DvCd94, DdA95, Dye98, FSCE92, MDD95, Onu96, PRC^{+}92, SMT97, She96, SGW94, TFL95, VDVC94, VPM96]. long-time [DFB^{+}99]. longer [GRBA96]. longevity [BDZB98, COS93, CA96b, CAL96, HKMK98]. longicaudata [DS96]. longicornis [DP91, HvB91, VV95]. longicruris [EY90]. longipes [KGW93]. longiremis [Nor96]. longitudinal [DD90]. look [FS91]. loop [BBRBK94, SFPCR93, TJ90a, WSM91]. López [Ano00]. López-Urrutia [Ano00]. lophophore [RG97]. Lorenzo [WSP^{+}95]. loricae [WMPT94]. Loss [MAdG^{+}96, BHHL96, BLC91, FML^{+}95, FW92, GBH96, HTG96, Jon97b, LWL98, LW95, LKC^{+}99, NBS^{+}97, PLFSW97, SMH97, Tro93, WSM91]. Loss-controlled [MAdG^{+}96]. losses [HHv91, HG94, SGP94]. Louisiana [RBF92, RR93]. Low [KNM93, Atk96, BCNC96, BMBD95, BHHL96, BSD^{+}94, Fly99, GCKW95, HTLK99, HFGT94, HS98b, KGHTN99, KE95, MF95, PB92, RT92, SMH97, TDER97, UKR^{+}91, VHG95, WA99a, Wor95]. low-frequency [BMBD95, TDER97]. low-level [UKR^{+}91]. low-nutrient [RT92]. low-shore [BCNC96]. lower [BF95b, MBM92, MNO96, Nag96, RGS95, RT93, RCGSN96, RPMB91, SWT99, TNSM98, Whi94, WDD92, HV91, PR93a, SGL^{+}97]. lower-higher [Nag96]. lowered [MA99b]. lucens [OF94]. lucinid [DRS94]. Lucinoma [HCF93]. Lucky [VTGK99, CD98]. lugworm [FB94, HSBP97, Hüt90, RBA98, RBT96, VHG95, WCP^{+}98]. luminance [GMT99]. lumpfish [WB91]. lumpus [WB91]. lunans [Per90]. lunar [AKK94, ELMR98, Joh92]. lusitanicum [Dut98]. lutea [MNK^{+}97]. lutein [BCG97]. lutjanid [She95]. Luzon [TNKMK99]. lyase [SDG95, SD96, SWK98, vdBTvDR96]. Lycenchelys [MRS96]. lymphoblastoid [BBR^{+}93]. lyra [BBT^{+}93]. lysianassoid [Moo94]. Lysianassoidea [Fra93]. lysis [BMS^{+}97, BRN^{+}95, CS91]. lysogenic [CKP98, JP96, Moe97a, Moe97b, Moe97c, WF94]. lysogeny [JP94]. lysogeny/bacteriocinogeny [JP94]. Lysosomal [Reg92, KCV94, LS94, LFD95, RCH98]. lysosomes [LME92, SM97b]. lysozyme [GSS93a]. Lytechinus [RSK^{+}99]. lytic [WF94].

m [AMMG91, BK94, CS97, DLFL96, FKT^{+}99, KRM94, PGM^{+}98, SL91, SSM91, TGB99, TFF94, IHK^{+}98, NKH^{+}99, PPW96]. M. [BB93, GS90, HS96, HS98c, LA97b, SZA94, Szm91]. macaroni [HTCW96, HHE92, TMCE98]. maccoyii [DJY90, DLJ91, JD90, KJGC96]. mackerel [KWEG94, MSD98]. Macoma [BC90, CL90, DL91b, DL94a, EB92, GM99, Gün91, HV91, HvdM97, HvdMBK98, JT97, JJTS97, LWL98, TBY98, WJM^{+}99, ZB92]. Macomona [HTCP96]. Macquarie [Sca93]. macro [FC90]. macro-benthos [FC90]. macroalga [BWB98, FDZ94, GHED^{+}94, KJMRC96, LA97a, McG92, SK98a, SK98b, VD95]. macroalgae [AKL^{+}99, AMB99, BK96a, BM95b, BVS^{+}91, DPM96, Dye93, EBB95, FDZ93, FH97a, FHCC98, GW94, GWKW97, GD96, HHN93, HMWN97, HPSJ95, MSJ92, PB96b, PB97, Sch99, SRS92, SP99b, WF90, WtD90]. macroalgal [And94b, BH97a, Hol97, JHN99, KSS90, MC96b, MS94a, OM99a, OB94c, PKL97]. Macrobenthic [FTP99, AGW93, Bre95, Dit96, Kar95, KSE96, KE97, MDD95, MHS94, OSC97, OSC98, SSR99, SFV96, Spr93, TN98, VWJ96, WC93, Wes90a, Wes90b, WA99b, ZC92, dG91a, MR91]. macrobenthos [LTT96, TPHC91]. macroboring [PCB90]. Macrobrachium [Gra91b, Gra91c]. macrocephalus [SPB^{+}99a]. macrocrustaceans [ACR99]. macrocrustracean [GAR93]. Macrocystis [AHT^{+}90, GHL^{+}97, SDER96]. macroepifauna [OM90, ÓM92]. Macrofauna [Ber90, KR92, dG92a, ABN98, CSR^{+}92a, DHH98, GB97a, JLF95, KDR^{+}92, NAW99, ONR98, SS93, SRW95a, SHO96, SRW95b, SE91]. Macrofaunal [LBM^{+}99, CCC90a, CB97, GP91, KDHZ98, Las98, MSK95, RK99]. Macrofouling [QRSC99]. macroinfauna [JMC93]. Macroinfaunal [HBC^{+}91]. macroinvertebrate [AD94, JJ99, RB99, VH93]. macroinvertebrates [RB98]. Macromolecular [SD93b, MG97, NK92a, SGK^{+}97]. macrophyte [LM91b, VD99]. macrophytes [EDSJ95, FGR97, FG93, Han92, RWJ^{+}95, ZCPA94]. macroplankton [MWA93]. macrotidal [DLT^{+}97, DQT^{+}97, RM94]. macrourid [DBMM99]. macrozoobenthic [GJVG96]. macrozoobenthos [FH96]. macrozooplankton [GF91, HC94, PPM94, PD95, TWS^{+}95]. macrozooplankton/ [PPM94]. macrura [KHFP^{+}96, Nor92, Nor94]. mactrid [MCL90]. Mactridae [CJL91]. mactroides [BD99]. maculata [CGFC95]. maculosus [Mga92]. Madagascar [PCB90]. made [AMS99, Ric91]. Madracis [GEW97]. Madre [Onu96]. maenas [AWD95, AAWD97, BC93, BKH99, HMHDA94, HN93, LS92, MPvM98, QCM94, QCM97, RN93, SAA99, WJD93, ZN96a, ZN96b, ZAN99]. maenus [CBRD92]. Magdalena [GGDLE96]. Magellan [TGBB97]. Magellanic [WSP^{+}95]. magellanicus [BSH^{+}94, BSJPW93, BSJ97, CHS^{+}95, CHC94, CH99, EGMC94, MW94, MGP^{+}96, PRC^{+}92, PGM^{+}98, SH93, TRS91, TS90b, TS90c, WSP^{+}95]. magister [FIA93, Fer99, SLSH98]. Magnetic [JBO97]. Magnitude [CDME97, BGvDR96, CDM^{+}96, CC90, GJC97, MD94c, SAH98]. mahonis [PM90b]. main [CA99a]. Maine [HF99, Lev91, LGMD98, MMGIM90, PWS98, TK96, WT99b]. mainland [AAE^{+}96]. maintain [ST98]. maintenance [BCMC99, VAD^{+}95]. majalis [Har98]. Majidae [SMUS^{+}99]. major [AWH^{+}99, CWH98, GBJ94, KHFP^{+}96, RSH92, FB95, SF99]. Malacostraca [Bra95]. Malangen [Fal96]. male [LMO99, TWWE95]. males [AS94a, Kyo92, San93a]. Malformations [SE98]. Mallotus [DL91a, DL93a, DCBC91]. mammals [Bow97, Liv93]. Man [NG99, AMS99]. man-made [AMS99]. management [Hal98, Lee90, SD99]. mandible [KB94]. Manganese [LGdC^{+}90, BDH94]. mangarevense [SP99b]. Mangrove [BWHC91, LJ95, Alo90, BNK^{+}97, Bec98, Cha98, COG97, DE90, DGHT^{+}98, EAT95, FFH99, FHCC98, HKY^{+}97, HSK^{+}94, Hsi95, KHB91, KDAS92, KKH^{+}94, Lee97, MvdVD^{+}97, McG97b, MWH91, NBW94, ÓN97, RMTB^{+}95, SK99, SOSV95, SFV96, TA90, VHH^{+}96, VM98a, War90, WD90]. mangrove-seagrass [MvdVD^{+}97]. mangroves [FH97a, LJ95, VHH^{+}96, ZCL97]. Manicina [Joh92]. manipulated [KLAH^{+}94]. manipulation [OJ97, Web91b]. manipulations [Con98, NvA91]. manna [WSS94]. mantle [BV99, BVP99, GK96b, WJM^{+}99]. many [PN99]. Manzanillo [DZ99]. Mapping [FCP98]. Marennes [RRGB96, GB95, GBB^{+}98, KSRB99, LGV^{+}99a, LGV^{+}99b, RR97b, Rie98]. margarifera [HSTY98]. margaritifera [YKL98a, YKL98b, YKL99]. margin [CSB97, LBM^{+}99, ODS^{+}96, TB98]. Marginal [PPMF97, AW98, DM91, MG91a, MPS95, SCKS90]. Maria [HBC^{+}91]. Mariana [AHY^{+}99]. mariculture [ERC95, KEvN95]. marina [Ack97, AZKA99, BAOT96, BSW93, BDM96, BGC94, CK90, FB94, FKC^{+}90, FMFH97, GNJ90, GR91, HBBM96, HBS98, HHN^{+}96, HN97b, HS99b, HSBP97, Hüt90, KA93, KF97, MRPC98, MNO96, OSJ93, OSJ94a, OSJ94b, Ole96, PF90, PB92, PB93, Phi94, PM94b, RBA98, RCG94, Reu98a, RBT96, RPDR^{+}98, SFV96, SP99a, TS99, VSSM98, VHG95, WCP^{+}98, ZA96, vKVSR97, vKSGvA99]. marinas [GU98]. Marine [BH91, CK99, HKA96, HSH94, HB95, Moe96a, Moe96b, Mül91, PLM92, SC93, WS92a, WLB^{+}99, ZMD95, AAD91, AGW93, AKL^{+}99, Aks96, AWdS92, ADR94, AR99, ASP91, And96, AW97, AB92, AJST98, BKS^{+}99, BMS^{+}97, BMSHC98, BBR^{+}93, BK96a, BCHL93, BA91, BD91, BWB98, BH92c, BH92d, BAY91, Bow97, BHH^{+}92, BEH93, BA94, BOC90, BZSG96, BEG97, BR97, BRZ99, BDNS99, BC99, CNGG96, CCC^{+}90b, CR95, CF93a, CF94, CBA^{+}99, CD94, CM99, CR98, CW95, CLK90, CK95, Cle93, CKP98, CGP97, CGZ95, CD99, CS91, CRLP94, CH97b, CRP99, DBHM99, DPM96, DHS96, DL93a, DeV92, DL94a, DGO92, DWSA98, DL94b, DAE^{+}99, DCG^{+}99, DDW90, DvRN99, EPL94, EBB95, EBM93, EPD^{+}99, FTS97, Far98, FM95a, FWWL99, FTGW90, FKB92, FNF91, FR91, Fly98, Fra94]. marine [FW91b, Fry96, GCH93, GdGMD95, GPF95, GW94, GFPD92, GTWF98, GHTH94, GCLS94, Gol99, GDMS90, GGKJ^{+}98, GSS93b, GS93, GD96, Gos97, GQ97b, GAB90a, GAB90b, GTJ92, Gow93, GCWH90, GWSG99, GF90, GJ96b, GJ96c, GSMM96, HAH^{+}90, HHKS92, HG98, Han92, HHN93, HS98a, HB91a, HP90, HBBM96, HDPW93, HB97, Hav93, HH94b, HLH96, HVFHH92, HBB^{+}99, HHM96, Hin92, HS97b, HW92, HSCC96, HH91a, HMB93, HMB95, HK92, Hol96, HHB97, Hol97, HBvD92, HLP99b, HCI93, HS97c, HHN90, Ito97, JLF95, JMD^{+}99, JP94, JP96, JM99, JMR91, Joh96, JAL91, JBB90, JKC^{+}93, KH92a, Kam94b, KFF92, KFPR94, KI97, KHFP^{+}96, KRJ^{+}95, Kin93, KS95b, KZP93, KOSS90, KDKS96, KRJL96, KWHM^{+}96, Kro91, KLJ94, KHT93, KK96, KQH^{+}96, Laf93, Lai92, LH98, LB99b, LS96b, LMM98, LGC^{+}93]. marine [LD94b, LM95, LHH96, Lip95, Liv93, LW91, LJS90, LSP99, MFF97, MM91a, MP96a, MDD95, MB95a, Mar99a, MSJ92, MM99, MA93, MRMLP97, MRMR92, MBWT98, MB97a, MR97c, Moe97a, Moe97b, Moe97c, MVdM^{+}99, MHD97, ML90, ML91c, MJSDF91, MV96, MH94b, MC91d, MSND^{+}99, MNH92, MBL^{+}98, Mye95, NvA91, NK90, NK92a, NCGG96, Neu92, Nie93, NL96, NJSS91, NR94, NFT92, Nov90, Oes90, OC95, OF94, OWM92, OW95, ODN^{+}95, PJF93, PL99a, PM90a, PMETM96, PW94, PF90, Pec99, PDBA90, PS92, PFP93, PT93, PGAJ99, PR93b, PLS^{+}95, PF92c, Pud93, RK96, RTRF90, RRWG96, RWJ^{+}95, Raw96, RS94, Rel96a, Rel96b, RB98, RB99, RC95, RHS99, RH90b, RTJ^{+}93, RM97b, Riv90, RSAD98, ROI93, RHJ91, RA96, RTRP^{+}98, RNH99, SB98, SRH^{+}98, SCB92, SMG97]. marine [SAH98, SR93b, SW97, SdC97, SKKO97, SRS92, SJCL93, SSY98, Sim92, SAA90, Sim91, SO96b, SRW95a, SC95a, Som94, SDH95, SCC94, SH99a, Sv93b, SNR^{+}96, SH99b, SGP94, Sty97, SC95b, SKT97, SHW^{+}91, TDVF99, TNP97, THCD95, TD94, TK91, VH99a, VB90, VVT97, VGJ90, VS96, VBZ^{+}94, WH92, WFTH97, WW96, WSF99, WFHM99, WH96, WC93, WYY^{+}98, WWC^{+}99, Web91a, Web91b, WWL97, WP94, Whi94, WKH^{+}98, WJM^{+}91, YKK^{+}98, ZT90, Zho98, ZNH93, dNDS98, vR93a]. marinus [BAM^{+}99, BA96]. maritime [BGC94]. marked [PF92b, SM94]. markensis [PDB^{+}97]. marker [BD98, Cam97, MHvB97]. markers [EPD^{+}99, KZK^{+}95, MMHW96, SB97b]. Markov [MDTB94]. marks [WM97]. Marlborough [HJH^{+}99]. marsh [AWHH94, ATNW97, BAH94, Blu93, BI99, BM90a, BM90b, CMD^{+}94, CCSN93, CNP95, DSW^{+}91, FW91a, HKvB93, HHSvS96, HG94, KW94, LK94, LTT96, MH90, NDRP93, O’C93a, OSZ95, Pak92, RBF92, RSN^{+}99, RD95, RR93, SR97, SM90, TA93, TA95, TZ99, dG92a]. marsh-edge [RBF92]. marsh-estuarine [DSW^{+}91]. marsh-surface [RR93]. marsh-water [CCSN93]. marshallae [SM98b]. marshes [MZ92, MW97, OJ99, Roz92, RR93, Tur93]. Martialia [GTYR97]. Maryland [GJC97, GJ97a, GJ97b]. Masfjorden [BA93]. Mass [LK93, RBC^{+}91, WVMR90, ATNW97, BDNS99, FHA99, HAH^{+}90, HHKS92, HGS95, HH91a, HvdM97, KMG98, LL98, LH91, MHRV^{+}99, NNR95, PJ99, SR90b, SR95]. mass-bleaching [HGS95]. mass-cultured [LH91]. Massachusetts [CBTM93, Kel98, PT94a, PT94b, BT93, KD97, TB93]. masses [CWH98, DBMK99, HK99, MO95, TRMW95]. Massisteria [PF90]. Massive [ECF^{+}97, CED91, RS91, vW98]. mat [OKK95, PPF95]. ’match [GPLB96, FPG95]. ’match/mismatch [GPLB96, FPG95]. Matching [HTCP96]. materal [MS93a]. material [AC96a, DSW^{+}91, DL91b, Dei90, Eno93, GS94b, HB91a, HH94b, LMT96, Lee97, MS97, SB94, Sim92, Wai90]. materials [ORL97, WW92b]. maternal [Ito97]. mathematical [DPO91]. mating [SMUS^{+}99]. matrices [AC96c, ABY97]. Matrix [AD90, Åbe90, LSB95, WTM^{+}96]. Mats [BF95a, FB95, Far98, FDZ93, JP93b, KJMRC96, NB96b, PJF93]. matsi [SCL95, SLHC98]. matter [AW98, AJST98, AB96b, BNK^{+}97, BMTF92, BL94c, CSG93, CRCS96, CCR91, CDG97, CBD96, CR94, Col92, DÁSP97, DKVD95, FM95a, GBA^{+}96, GAC^{+}97, HKvB93, HDK93, KE95, Kir90, LM91b, MKN^{+}96, MKYI98, MWH91, NKI^{+}96, NPB94, NNB93, ONM93, PPFP98, PDC97, PLH91, PGR94, PF91a, PSE97, PSPD96, RR97b, Rie98, RTLG97, SCJ99a, SR98a, VKV95]. maturation [GCKW95, QQ98, WCP^{+}98]. maturity [BAY91, DS96, HR96a, MACR98, Roa93, RT98]. Matuta [Per90]. Maurolicus [KKH98]. Mawson [KR97b]. maxilliped [PL99a]. maxima [CB96, ACLS97, YKL98a, YKL98b, YKL99]. ’maximum [vW93, BKK^{+}93, EML^{+}93, FP91b, Har90, MD94b, SD93a, TDD94, VMS^{+}96]. maximus [CCJC91, CTC93, CMT94, HBBT98, SM97a]. Maxmuelteria [HAA94]. May [GGB94, EP92, FBR^{+}95, GK96a, IVM^{+}98, MPS96, Tro92, AW98, KLMR91]. mayeri [BLD94]. Mazzaella [BCBR97]. MCA [LBV^{+}99]. MCA-Leu [LBV^{+}99]. McMurdo [ARS93, ADG^{+}95, PMS94, RKS97, SBP92, SBP93]. Meadow [VAD^{+}95, DSS^{+}98, DCK99, DTA^{+}97, ZB99b]. meadows [FKC^{+}90, HHv91, LCLC96, RH97, SRW95b, VH93]. Meandrina [JS93]. meandrites [JS93]. means [ASP91, BAY91, Bro90, JSGS94, KBP98, RH90b, Tho91, WJM^{+}99]. measure [CCC90a, CW99b, GSS93a, LMK93]. measured [AMB^{+}96, HWLC98, HdWS91, HBvD92, KW91, KLJ94, LCL95, LKvRH96, MNH92, NG96, TCB93]. Measurement [BRNH97, CV91, Kru93, LGJ99, MKCL94, MGEC97, RWJ99a, Rel96a, Rel96b, WR91, ACC90, Bec98, BD91, BG91, CH97b, DEP^{+}96, FF92a, FGB^{+}96, GCH93, GPD^{+}96, Irw91, JM99, KH92b, KS95d, NMR97, QC94, SSR^{+}97, SFS95, SFG^{+}99]. Measurements [vR93a, Add92, BTYO94, BB97b, BAK97, CR95, DRS94, GB95, GRGK96, HPSG94, LBV^{+}99, LZYW93, MMHW96, MM99, Maz95, MA99c, PML93, PLFSW97, RGLS99, RWJ^{+}95, RC91b, SPS^{+}91, SPDM91, TH94a, WGD97, WWGK95, WM90, WK93, WH98, ZCPA94]. measures [HG98, KCV94, LMTF96, LB93, SOC97, WC95]. Measuring [Åbe90, BVW^{+}98, CK95, FGP98, GW93, LKC95, McC94c, Min93, MC91d, RSN90]. meat [Bar95]. Mechanism [LE90, BLM^{+}95, DRGV91, HSB97, TF98, WBP^{+}92, WD90]. Mechanisms [Car90, MD98b, Ano99a, BDB97, BV99, CCC^{+}90b, CM98b, CH94, FB94, HKA96, HG92b, LY97, MCW^{+}91, NR98, OC95, SO96b, WFTH97, vPB90]. mechanistic [HH96]. mediated [DW97, EGC95, HvdM97, HB94b, LR96, LAK99, MJ97b, PMG^{+}98, UMT^{+}91]. mediating [CH96]. mediation [CG93]. mediators [NC90]. Mediomastus [SFB92]. Mediterranean [BG99, DMP^{+}99, KDR97, MSM^{+}98, SFPCR93, TGB99, Vid94, VM95, ASN93, ADA98, ARDL97, BCC95, BCBC98, BCMC99, BR94, BBF94, BBP^{+}96, BGCM95, CSC92, CA92, Car93, CS93, CM98a, CDM^{+}96, CDME97, CC99, CJ97, CVD99, CMA95, CRZG95, CZG95, CS97, CJ93b, DBB^{+}95, Dem95, EBB95, EDSJ95, EML^{+}93, FGR97, FPR94, FHA99, GSC96, GCR98, GF95, GDP^{+}98, GHED^{+}94, GSB98, GLD^{+}91, Har97, JRMN^{+}97, JPZ^{+}98, KTH98, KMB97, KM99b, LC91, LHRL97, LED92, LBL91, LZYW93, Mac94, MBF^{+}97, MZ99, Mal99, MDC^{+}96, MCED96, MCM^{+}97, MPS96, MCNB98, MC97a, MC98, MDTB94, MB97b, MC94b, MAdG^{+}96, MSCTBE96, OSGB96, Oro96, OP99, Pal91, PF92b, PMB99, PDF97, PRPM^{+}94, PAM96, RR92, RZWY96, RGBC92, RMGB^{+}94, RGBJM95, Sab90, SCIA99, SZ96, SB97a, SRH^{+}98, SMT97, SMG97]. Mediterranean [SGPAA95, SHV91, SH95b, SGC^{+}98, TDBH94, TR95a, TB98, TP95, TP97, TB92, UMT^{+}91, UTB^{+}95, UMTM98, VWS92, VD95, VHVB^{+}98, WTF96, YST^{+}97, YTF97, ZR92, ZBK^{+}98]. mediterraneus [DL95]. Medium [MRMR92, RMMB91, ZB92]. medium-sized [ZB92]. Medium-term [RMMB91]. medusa [WK92a]. medusae [BS95a, BG99]. mega [TRU91]. mega-invertebrates [TRU91]. megabenthos [LPW95]. megafauna [AML95, CM98a, CEV97, PS97, Pir92]. megafaunal [VD99]. megalopa [EPDC99]. megalopae [BW93, BM95b, CC97b, FFR94, JE95, LE90, O’C90, OJ97, OJ99, Olm94, WD94, ZN96b]. megalopal [WG97]. megalops [BGC91]. Meganyctiphanes [MVA99, OK98, STS99, TCRB99]. Megaptera [BSF95, BFS^{+}98, Mik97, RCA^{+}95]. Mehuín [GW91, ZMD95]. Meiobenthic [ZC92, AWVP98, Car90, CM90b, GCC95, Hic92, Ken94b, Ker90, LF97, NDG^{+}99, OM90, OE91, PC96, SMF91, SF91]. meiobenthos [BVV^{+}94, ÓN97, SOSV95]. Meiofauna [SE91, ABN98, Bla90, CGFC95, GAD^{+}92, GB97b, GF98, MvdZH^{+}97, NAW99, ÓMvdB99, SW97, SRW95a, SCCS98, SF94, Tho91, VWD^{+}95]. Meiofaunal [FS90, Ken94c, SS96a, SGW94, SPPL92, ZC92]. Melanogrammus [Cam97, FQV98]. melanomacrophage [BVL92]. melanopus [AMFB99]. melanostictus [AM93]. melas [AS94a, CAATS94, GRC^{+}97]. melt [FPP^{+}97]. membrane [LFD95, MGAE99]. membranes [LS94]. Memoriam [Ano94, PdG96, Tor94]. Memorization [RGBJM95]. Menez [CD98]. menhaden [CMBC98, Dee90, FCW98, MC91a]. Menippe [BB93, LFS90]. Menippidae [MRF^{+}99]. menstrualis [CLHF95]. MEPS [Kin93, Kin95b]. Mercenaha [JCH93]. Mercenaria [lYALM93, lYAML93, BM94c, BLC91, BB93]. mercuric [CK95]. Mercury [MCFS96, BMSHC98, BBR^{+}93, BH94a, BH95a, BC93, Bro93, GPP98, MGM96, MGF98, NL96, RH90b, SPCF97]. merguiensis [VHH^{+}96, VHH^{+}98, WH93]. merits [BH92b]. meroplankton [BLDF95b, BLDF95a, SW97]. meroplanktonic [MPS96]. Mertensia [SC92b]. mesh [JLF95, SOW97]. mesh-size [JLF95]. Mesidotea [HS90, Vis91]. Meso [PD95, TcBW96, CR94, HV91]. Meso- [PD95]. meso-cosms [CR94]. Meso-scale [TcBW96]. Mesocosm [WA99b, AB95, Amb93, BBRBK94, CH90, DHS96, EPN^{+}99, FBMH94, KHW^{+}99, KQH^{+}96, Lav97, NGS97]. mesocosms [CPK97, LGJ99, NO90, OKL^{+}96, PGM^{+}98, RCTH99]. Mesodesma [BD99].

Mesodinium [SPDM91]. mesograzers [RWCS98]. mesohaline [CRH^{+}99, RCC^{+}99, WR92a, WR92b]. Mesoherbivory [Tro92]. mesopelagic [BTYO94, Big92, CA95, HDK93, KKH98, MCFS96, Röp93, SSP97]. Mesopodopsis [WW90]. Mesoscale [DG96, HC94, MF90a, MF90b, WSCT96, CLGM96, DM91, KL95, RPT^{+}96, TPW^{+}93]. Mesozooplankton [LPF94, LLP94, Per91, BS95a, BS95c, CMA95, KW99a, PLP^{+}93, PWPK95, Uit96]. messa [Lee97]. messel [YOH98]. metaanalysis [WC93]. Metabolic [ASF98, GDW90, JLD^{+}96, MRS96, WW93, AC92, BYO95, BBLY98, DF95, FCW98, GWKW97, HS90, HNL94, Kal91a, KG94, RMMB91, SB94, TRPS94]. Metabolism [CMA95, HBB^{+}99, LBDD97, Put90, TAD^{+}94, BTYO94, BBL^{+}95, Bre95, CCG98, CCC90a, CJ97, CG99b, DvCd94, DSV^{+}91, EDSJ95, FSdM^{+}92, FPLH92, FFH96, GPDF93, GPD^{+}96, HK92, HFJS91, KH98, KSMDB97, KA93, KHB91, KDAS92, KKH^{+}94, LGJ99, LA98, LC97a, MC99, NH98, PBA^{+}95, RFGF92, SJJMH90, SL91, SSP97, TTR99, TR91, VDVC94, VBN95, ZB98]. metabolite [LSP99, MP92]. metabolites [AP97b, CLHF95, FP91a, PPP^{+}96, SHM^{+}97, SL97b, TT90, TNP97, VH96, dNSR^{+}96]. Metal [GPP98, SM97b, BL95, CM99, MLH96, MH96b, ND91, RATA^{+}99, ROM^{+}91, Reg92, RDM91, RMGB^{+}94, Sca93, SGW94, SKM95, SDA^{+}92a, SPB^{+}99b, TBY98, VCMP93, WL97, WLT^{+}95, WJM^{+}99, ZCL97]. metal/shell [SKM95]. metal/shell-weight [SKM95]. metallo [SGR92]. metallo-thionein [SGR92]. metallothionein [HHH92, MAAT98, VCMP93]. metallothionein-like [MAAT98]. metallothioneins [DCG^{+}99]. metals [DL94a, DE91, FNF91, HHH92, MONR90, RATA^{+}99, RGBC92, RGBJM95, SKM95, WSF99, WH96, WTF96]. metamorphic [BM94b, Dav94, FFH99, MC91b, MD93, O’C90]. Metamorphosis [ALOT99, FFT98, FTBR97, AOT97, BM94b, BHQ99, BH94a, BP99, BM95b, BQKC97, CT96, CCJC91, FDR^{+}96, GB90, HQ99, Har96, JMPH90, McC99, MBL94, SK98c, WB92b]. Metapenaeus [CM97]. metapopulation [Ano99d, OP99]. Metazoan [FGS98, GW94, DD99, KBA^{+}97, Uit96, WR92b]. metazoans [VPFL92]. metazooplankton [KHW^{+}99]. meteorologic [SH97b]. Meteorological [SKM^{+}97]. Methane [Sch96, DAB^{+}91, DBN^{+}94, HK94, JAB^{+}92, JT93, LC95, SFWD90, SJCL93, SD93a, VFMFBE96]. methane- [SFWD90]. methane-oxidizing [SJCL93, VFMFBE96]. methanogens [Sie93, SJMC93]. methanotrophic [LC95, Sch91]. Method [KKT96b, BG91, CA93, EGKE92, FR91, FIL93, GGC99, GW93, HW99, HP98, Kar95, Ken94c, LS94, LKvRH96, MDD95, MM99, MKCL94, PPB90, RDPB99, RH90b, SH98b, URHR98, WBC91, WK98, WCLB99, WJM^{+}91, Zho98, dNDS98]. Methodological [MSP^{+}97, DDM^{+}99]. Methodology [ER95b, LHH96, VGM94, Rel96a]. methods [BMHH94, BBO^{+}95, CRLP94, GP91, HWLC98, Irw91, LMF93, MHC92, McC94c, McG97b, NNR95, PPB90, RM96, Tuo97]. methyl [BBR^{+}93, WR91]. methyl- [WR91]. methylamine [WR91]. methylamines [YSL94]. methylmercury [RPT95, RGT98]. methylotrophic [Sie93, SJMC93]. methylumbelliferyl [MNH92]. Metridia [OF94]. Metridium [AS94b, AS95]. Mexican [AADSCC97, GMMD93]. mexicanum [JLYK97, JSK^{+}99]. Mexico [AMBH98, Bia91, BFJ^{+}93, DZ99, DRP^{+}97, FML^{+}95, Gen92, GGPGG95, Góm95, HD97, LFR^{+}97, MDSB92, MZCMOCJP91, MZFHC96, PSSP95, RGLD^{+}98, RMTB^{+}95, SS96b, VH93, GGDLE96]. Mg [EL90]. micans [Les96]. michelinii [NB98]. micro [BF99, BS95c, HPSJ95, Hei95, LB96b]. micro- [BS95c, HPSJ95]. micro-habitat [LB96b]. micro-organisms [Hei95]. micro-scale [BF99]. microalga [MC97b, UGSBA94]. microalgae [CNP95, Fly90a, Fly90b, FF92a, JMD^{+}99, LS91, MC96a, MMR99, MLTD96, PLC^{+}91, RCN95]. microalgal [ARS93, ADG^{+}95, BSS92, BBO^{+}95, BLR94, CC92a, CBTM93, Cam91, DB96a, GM98a, GBGH97, PC96, PZ91, SPPL92, WNJ91, YKL98a]. microbes [Wai90]. Microbial [ALBS96, ALB^{+}96, BM91b, CDHD92, CLC^{+}98, HSRHH90, LBJ^{+}98, LGH95a, LAG96, MR94, Nau96, RAlBW99, SNR^{+}96, VAD^{+}94, AMB^{+}96, AJST98, BBRBK94, BL94c, BHT^{+}92, BGvDR96, CPH^{+}96, Dav96, FB95, FTGW90, FTW90, GDW90, JTPP92, JP96, JP93b, Kep91, KDR97, LDV95, NMR97, Nov90, OKK95, PJF93, PH91a, PPF95, PMOD90, PH95, PLS^{+}95, RSB^{+}90, RTLG97, SFPCR93, SFSA99, TJ90a, TR95a, TGB99, UNPH93, WSM91, WCT95, YSL94, ZBK^{+}98, NKI^{+}96]. Microbiological [KDMR97, HLH96]. microbiota [BSG^{+}97, KNK90]. microchemistry [AOT97, MCFB97, RSM^{+}96, TSW97]. microcosm [DBdCS99, FGKP95, Mal96, SGL^{+}97]. microcosms [And96, FSdM^{+}92, FDZ93, FDZ94, HB91a, HHB93]. microelectrodes [HdWS91]. Microenvironment [KCD^{+}95, LGdC^{+}90, PH91b]. microfauna [Alo90]. microflagellates [DG91b]. microflora [GB97b, MSD92]. Microhabitat [WM92, AML95, Egg95, Hsi95, Lev91, MU99, MD93, MJ92a, Tol98]. microhabitat-scale [MU99]. microhabitats [GC95b, Kog95]. microhaoitat [MAC92]. microlayer [Gar91, HM97a, HC92a, Kar92, MC94a, Wil92]. Micromesistius [HK99]. micrometre [MS93a]. Micromonas [CS91]. micronekton [BA93, BLJ99, PPM94, PM92a, YBLL96]. micronektonic [TAD^{+}94, TDH^{+}94]. microniches [Sie93]. micronucleus [BGVF92]. microorganisms [BK94, FKB92, LC91, LF93, LGR95, MC91d]. Microphiopholis [PLFSW97]. microphytobenthic [BPC98, GBB^{+}98, SPB95, WH98]. Microphytobenthos [UK98, BKP98, Bla90, BGR^{+}96, HWLC98, HS97a, KBP98, SEGP91, dC94]. Microplankton [FP91b, SS96b, AC92, FPPM96, GGB94, HTS92, PMMH95, PG94, SSY98, SHML98, Uit96]. micropredators [OWM92]. microprobes [KLJ94]. microprofiles [VDA96]. microprotozoan [KKT96a]. Microrefuge [MU98]. Microsatellite [ASRC99, PM98, SB97b]. microscale [Bla90, NMR97]. microscopic [AHT^{+}90, HS91b, SHA^{+}95a]. microscopy [BP93, LHH96, RCGSN96, WTM^{+}96]. microsensor [LGR95, PJ99]. microsensor-caused [LGR95]. microsensors [KCD^{+}95]. Microslomus [RFGF92]. microsomal [RA92]. microstructure [AOT97, JD90, MJ91, RG93]. microtags [BJB95]. microtidal [GPB^{+}98, GY97]. microzones [SR93b]. Microzooplankton [GAK99, EP92, FPR94, FBL95, FP96b, FWF^{+}99, KL92, KW99b, LKC95, MEC92, NGS97, Put91, SM99b]. Mid [DZ99, ZD99, Fab95, HFGT94, KKW97, KW99b, LB93, MU97, NH95a, RT93, ZHM91, CD98, VTGK99]. mid- [Fab95]. mid-Atlantic [MU97, CD98, VTGK99]. mid-Chesapeake [LB93, ZHM91]. mid-continental [KKW97]. Mid-ocean [DZ99, ZD99]. mid-summer [KW99b]. mid-water [HFGT94]. Middle [GJ96a, MRS96, BSVK93, MRF^{+}99, EDP^{+}98]. midsummer [GH91, TI96]. Midwater [GLD^{+}91, HS98b, BYO95, KKW97, SM98a]. midwinter [HW95]. miersii [Ang95a, Ang95b, CCDA98]. migrating [AHY^{+}99, Hei95, KKH98, PM92a]. migration [AS94a, ALOT99, AALT99, BA93, Bar97c, BSF95, BFS^{+}98, CN95, Gar99b, GMT99, Gib94, HPPZ93, Hay95, HWL96, HHH97, HLP^{+}99a, Hil91b, HB98, HM97d, HN93, JWKH98, JBK99, KDC99, MCC97, MGP^{+}96, Olm94, OK98, OF94, PM91b, QCM97, RNHH97, STS99, Tak96, TCRB99, TD92, TS90b, WW90, WPB^{+}93, ZN96a]. Migrations [RSM^{+}96, KNI95, SW91, SK97, TK97]. Migratory [Arm92, AB95, MO97, Sew96]. mikimotoi [NSH95]. mild [Beu92]. milk [BBCB99]. Millepora [Lew91, Lew92]. Miller [JNCS90]. mimicked [WC91]. mimicry [EGC95, HM91]. mimics [GBV96]. mineral [BNW94]. mineralisation [TGSN99]. Mineralization [HHB93, AWdS92, CRCS96, GCH93, GWW^{+}98, GHHC98, HB91a, JBB90, MKN^{+}96, RTRP^{+}98, SH92b, ZB94]. miniature [QRSC99]. Minimal [RG97]. minimum [KR95, Sal96, ACC90, JNJP94, Sci91, SA94, SLC^{+}97]. ’mining [DRS94]. minipellets [Gon92]. minke [ISF^{+}98, KNI95, KEJ98]. minor [HSW99]. minutum [FFF^{+}94]. minutus [MF91]. Mira [QCM94]. mirabilis [GEW97]. Mirocaris [PSB^{+}97]. mismatch [FPG95, GPLB96]. Mississippi [CLB92, CG93, GBA^{+}96, GF91, LFR^{+}97, MSD92, ND96, PBA^{+}95]. mitchilli [CC91a, CH90, JAOS90, LB93, MH96a, PHT96, PND^{+}94, WH95, ZHM91]. mitchitli [NH95a]. mitigation [FKC^{+}90]. Mitochondrial [GCNSNd94, BSP^{+}99, DCBC91, HSBP97]. mitopteryx [KZ92]. Mitotic [BZ92]. mitten [Ang91c]. Mixed [DPM96, BBFD98, BG94, CWH98, DD90, DTA^{+}97, FM94, GCKW95, HN91, LSL98, PWP^{+}97, RLDM94, SL97a, TK91, VAD^{+}95]. mixed-layer [FM94]. mixed-water [BG94]. mixing [BM94a, MG96a, MFF97, MJ93, PSK98, San97, SFS95, Wor95, ZBK^{+}98]. Mixotrophic [HN97a, BR94, JLYK97, JSK^{+}99, NH90]. mixotrophs [CVD99, HR96b]. Mixotrophy [Sko96, SLC^{+}97, FK95a]. mixta [MHD97]. mixture [Han95]. mixtures [HH94a]. Miyako [KSO^{+}97]. MMCs [BVL92].

Mn [HB94b]. Mnemiopsis [CLW99, PC95]. Mobile [TC94, AT96, AML95, BKH99, Hol97, Tak99, Tay98, CPB96]. mobility [Bre96, HTCP96, SMT97, SH99a]. mobilization [AZKA99]. mode [AHS97, BHQ99, BvKvRB96, DCBC91, Hsi97, KZ92, LW95, SMB97, WLT^{+}95]. Model [CN92a, vdBTvDR96, AE91, ATNW97, AH94, BC91, BBRBK94, CH97a, Che93, DD97, FM94, GF95, GMK97, GA97a, GS91b, HHM96, HS97b, HJHG98, JBK99, KSE96, KM99a, KGHTN99, KS97, LBV^{+}99, LSL98, LB93, MDTB94, MJ92b, MJ93, NK96, NJLD98, RM94, RRWG96, RPN95, RCC^{+}99, SAG92, SSR^{+}97, SH98b, TD94, VBCRB97, VNB97]. Modeling [CRH^{+}99, DHKP96, WF97, ZCPA94, GJC97, JNJP94, KD97, LVL^{+}99, NP90, Sos96, GJ97b]. modelled [Ano99d]. Modelling [APT99, BSP93, BH98, CLD99, Dav96, FF99, GBGH97, HJH^{+}99, HM97c, LGB91, MH97, SKM93, SEAI96, WSC^{+}99, Ber92, CL97, LJM99, LC97b, MT93, RVBL92, RBT96, TGG97, YBG^{+}98]. Models [BC91, AD90, BMH91, BMFR97, BFMM99, DRS94, Hob93, KPS92, KPS97, LSB95, ML93, MR97a, QRSC99, Sca96, SW91, Vis97]. moderately [JBHH97]. moderation [MRP98]. modes [GS91b, HKG93, IBB97, JBB98, NA95, WWL97]. Modification [CSS93, RTLG97]. modified [KRM94, UGSBA94]. modifies [BZSG96]. Modiolus [NT96b]. modular [CRGZ98, LC98]. modulated [BVW^{+}98]. Modulation [BBR^{+}93]. mole [AS99b]. Molecular [HHG99, JPZ^{+}98, Moo92a, Moo92b, RP91b, AMO^{+}99, BMS^{+}97, BH93, DCK99, GBA^{+}96, HH94b, KNM93, KE95, LEHvR96, LME92, MS95, RRWG96, SRS92, WP95]. Molenplaat [LH99]. Mollusc [BMH93, DBN^{+}94, JFS90, JFS92]. Mollusca [BSJPW93, BSJ97, DBV98, FPB^{+}94, MW94]. Mollusca. [MS91]. molluscan [Dye93, Jua92, WM90]. molluscs [Arm92, CH99, DRS94, HSTY98, Sea90]. molt [O’C90, WD94]. Molting [FFR94, Lop91, BB93, OJ97, OJ99, RvMM97]. Monachus [FHA99]. monitor [BSW93, WWGK95]. monitored [CLJOC98]. Monitoring [KD97, WCBL94, AAD91, And98, CDGF^{+}93, MP96c, NKKO97, VBL^{+}92]. monk [FHA99]. mono [BCG97, RA92]. mono- [BCG97]. mono-oxygenase [RA92]. monochromatic [GJ96c]. monoclonal [VPGV94]. monoculture [Han95]. monodon [CCC94, Roa93, RGE^{+}95, RT98, SCMM95]. Monodonta [Tak95, Tak96]. monoene [SM98a]. monomethylamine [SJCL93]. monomethylamine- [SJCL93]. Monoporeia [Joh97, Leh96, LA98, OE91, SE98]. monoxide [VGJ90]. Montastrea [MNB94, Szm91, VB93, Van94b, VK94, VB94]. Monte [BMHH94]. Monterey [BUCP^{+}92, SSP97, SPS98]. Monthly [BP98, GB95]. months [DG91b]. Moodley [BAR96]. Moorea [Adj97, DG93, GPDF93, HGS95, LGH95a]. mooring [LOS90]. mordax [DD90]. Moreton [CM97, SW93, WB92a]. morhua [LAK99, Cam96, GA97b, HSB97, MF96, MLDM95, Nie93, NKV94, SFF92, SK95c, TH90, TRMW95]. Morlaix [MDD95]. Morone [Che93]. morphodynamic [GY97]. morphodynamics [GD99]. Morphological [CB92, Mil94b, Uib95, WHW93, HRFC90]. morphologically [VPGV94]. morphology [BWRH99, Dav94, DEF91, DKVD95, JS93, KJT90, KW97b, MDRW99, NA95, SDB96, SC93, TB99, TJRG93, WW96]. morphometric [CO92, SFF92, VB93, WMPT94]. morphometry [BG97, FSGG96]. morphotypes [WP94]. morphs [DBV98, MTP92, NB98]. mortalities [SR90b]. Mortality [MBF^{+}97, PF92c, Aks96, Ang91a, BCN^{+}95, BEH93, BRC99, BHH^{+}95, CN92a, CN92b, Con98, CRH^{+}99, DY92, DLJ91, EDS^{+}99, FHA99, GS91a, GQ96, GQ97b, GS99a, GCGR90, HR96a, Hil91a, HBP^{+}93, Hol90, Hov92, HS97c, Kne93, LK93, LB99a, LCL95, Lop91, LTM^{+}98, MC91a, MCF90, MG91b, MJ97b, Mye95, PG90, PND^{+}94, QC94, RT95, RBC^{+}91, Row90, VWM90a, VWM90b, Web91a, YAG92, ZMD95]. Morus [AAS^{+}91]. mosaic [Hol97, JBH98, KLN^{+}90, SJ99, VDM^{+}97]. motile [Con95, GSS93b]. motion [FH97a, JM93, McG94, MJ92b]. motor [BR91a]. Moult [Buc91, Dem95, CN92a, HVL96, SPB^{+}99b, TCRB99]. Moulting [CN92b, SW98b, CN92a, ZAN99]. mounted [CSB94]. mouth [CCCH94]. move [VHH^{+}96]. movement [AFC96, CHS^{+}95, Fre97, MU99, RWCS98, SA92, Zel98]. Movements [CGNH98, Kog95, ACR99, BKH99, GPB^{+}98, JHW95, MST98]. Mozambique [Bar97b, Bar97c, Bar97a, Bar99]. mRNA [RA92]. mtDNA [PCM^{+}95, PPO^{+}99]. mtDNA-RFLP [PPO^{+}99]. much [Bar95]. mucilage [BMS^{+}97, MH93b, Naj97]. Mucocyte [BD96, BSJPW93]. mucous [MD93]. mucous-thread [MD93]. Mucus [PH91b, BSJPW93, DB99, MH93a, van92a]. mud [DMM99, FAED97, GH93, GBGH97, HH99, MDRW99, Tsu90]. muddy [BDdWC96, GB95, KSRB99, ONH^{+}97, OM90, ÓM92]. mudflat [Cam91, LGV^{+}99b, NDG^{+}99, SCAT91]. mudflats [vDdWKW99]. muelleri [KKH98]. Mulinia [CJL91]. mullet [LS96b, MST98]. Mullidae [LA97b, MM93, McC94d]. Mullus [LA97b, MST98]. Multi [CBS91, FDW92, Sym95, DEF91, HT96]. multi-armed [DEF91]. Multi-scale [FDW92, Sym95, HT96]. Multi-year [CBS91]. Multilocus [ZGB90, QZA95]. Multiphasic [LF93]. Multiple [SMUS^{+}99, SB97b, TVG96, ÅP97a, BJTB96, CNP95, Har96, LMK93, MA99a, MBB99a, MM95b, OV92, PN99, SM90]. multiple-choice [MA99a]. Multiple-locus [TVG96]. multiseries [PRM^{+}96a, PRM^{+}96b]. Multivariate [KV90, Las98, CU99, CA93, DFB96, MDD95, RK99, San93b]. Mumby [Ano99d]. Mundaka [UCV98]. muricid [MADM98]. Muros [FP91b]. Murre [CHDW92]. murres [DH96]. muscle [DvCd94, SDFJv91, TR91]. Musculista [Cro98, Reu98b, YOH98]. mushroom [Hoe91]. Musseis [VH93]. ’mussel [SM97b, JBB92, ABY97, BH95a, BMFR97, BFMM99, BLCA90, CR96, Cro98, FGG95, GS91a, HGS94, HJH^{+}99, Hil97, HMB93, HS95, HS96, JLR90, KNK90, Kre93, KHBL95, LB95, LGB91, LC95, MCF90, MUH97, MCRGC90, MGC91, MGC95a, MM96, MN92, NC90, NT96b, NHL99, PLBP96, PSPD96, QZA95, RVT^{+}97, SRRdM99, SV97, TR95c, Tok95, VB96, VBZ^{+}94, WW91, WW93, WFL95, WFL96, WF97, WD99, vG91]. mussel-dominated [TR95c]. ’mussel-watch [SM97b]. mussels [AC96c, BCNC96, CMRF93, CUM98, DWG98, GS90, HMB95, HTLK99, HS98c, LFRB97, LFD95, McG97a, MRP98, RCG94, Reu98a, RH90b, SKM95, WDB^{+}95]. Mustang [Ker90]. mutual [ML91b]. mutualistic [MLL98]. MXR [MM96]. Mya [AR91, BDB97, DMM99, LC97a, SGP99]. Mycale [PPW96]. mycosporine [SDC^{+}92]. mycosporine-like [SDC^{+}92]. myctophid [PPM96]. myctophids [SM98a]. mydas [HGBW99]. mysid [AJS95, Bus98, HFGT94, RWJ99b, Sal96]. Mysidacea [FB90, MHD97, RWJ99b, VR98, WW90]. Mysidium [Bus98]. mysids [OIO^{+}95, TK95a, TK97, TK98]. Mysis [MHD97]. mytilid [CD98, LC95, LTM90, PFMC95]. Mytilus [ABY97, BH95a, BG97, BSJPW93, BSJ97, BMFR97, BFMM99, Bou90a, Bou90b, BLCA90, BFVG92, BGVF92, CMRF93, CH96, CR96, CH99, DM96, GS90, HGS94, HSBH96, Hil97, HDB95, HBD95, HMB93, HS96, HS98c, JKT90b, KJT90, Kre93, KHBL95, KCV94, LFRB97, LS94, LFD95, MUH97, MCRGC90, MGC91, MGC95a, McG97a, ND91, NHL99, PMETM96, PLBP96, QZA95, RVT^{+}97, RCG94, Reu98a, RSN90, RH90b, SUH95, Sal95, SZA94, SRRdM99, SV97, SKM95, SP99c, TAJK90, Väi90, VB96, VBZ^{+}94, WW91, WW93, WFL95, WFL96, WF97, WDB^{+}95]. Mytitus [CMRF94].

N [BR97, CNCS93, CCC94, FMFH97, MMG97, OH95, RGS90, RTRF90, RHT91, RAB^{+}92, SDFJv91, TCB93, WYY^{+}99, GMG^{+}92, PDB^{+}97, BHHL96, BG91, CSM^{+}98, CMB^{+}99, CD98, CJ95, FFJ93, FMFH97, FPFR92, FHCC98, FCP98, GJ97b, GCH93, HW92, HAR95, JRMN^{+}97, LGH^{+}93, LH91, LGH95b, MP96a, MH99, MvdVD^{+}97, Mar99b, MP96b, MHRV^{+}99, MHM90, PJF93, Por94, PDP^{+}90, PPCR97, RTRF90, RHT91, RAB^{+}92, Rie98, RNHO98, SRSH99, SR95, SV97, Som94, TFL95, VNM97, WYY^{+}98, WCLB99, SDFJv91]. n-4 [PDB^{+}97]. N-cycle [VNM97]. N-cycling [TCB93]. N-starvation [WYY^{+}99]. N-starved [RGS90]. N-tracer [SR95]. N. [NEIR95]. N.W.T [MBJ94]. N/ [RTRF90, RHT91, RAB^{+}92, SDFJv91]. N0 [BG93]. N2 [KZP93]. nagasakiensis [Ker96]. naked [FWF^{+}99, Har99, MRMLP97]. Nam [KKH^{+}94]. nano [HTS92]. nano- [HTS92]. nanofiagellates [GS93]. nanoflagellate [Gas94, HNvB94, LDV95, LAG96, MF90a, MF90b, SK94b, TT99]. nanoflagellate-bacteria [TT99]. Nanoflagellates [CVD99, EPL94, HBvD92, KFPR94, WSM91]. nanomolar [RGS90]. nanophytoplankton [TFFA99]. nanoplankton [DGT^{+}93, FPR94, MPS96, SCB92, ZR92]. Nansen [Bia92]. Narragansett [TDD94]. Nassau [Egg95, SMW^{+}93]. Native [Reu98b, AW97]. Natural [Aks96, Ang91b, JC92, KI97, MHM90, RCG99b, SC92b, AKM^{+}99, AGW93, AGGKA96, AFZ96, BN95, BK94, Byt90, CHS^{+}95, CC91b, CLK90, CTC93, Eml90, ENSR90, FWWL99, FFH99, FJR^{+}94, FKC^{+}90, GW93, HWLC98, HJH^{+}99, HH94a, HVFHH92, HHSvS96, JLF95, JR92, KWHM^{+}96, LB95, LMK93, LKF95, LOS90, Lip95, Mal99, MA99a, MHvB97, MR97c, MZ92, MW97, MWH91, MAAT98, MC99, MBL^{+}98, NB90, OSZ95, PCK^{+}96, PCC^{+}97, RWJ^{+}95, RCG99a, RBJ90, RCH98, SHO96, SDB92, SDB94, SC95b, YKL99, dNDS98, vLTW^{+}98]. naturally [CCJC91, SPPL92, VAD^{+}94]. Nature [Arr93, LD93, LH99, PF91a, CCJC91, Lee90, NBW94, WM99]. naupliar [Jeo94a, Lop91, PLIM95]. nauplii [VTNA91]. Nauticaris [PM90b]. NBS [CH97b]. Neanthes [ORD98, Vis90]. neap [DL94b]. Near [BE97, ASP^{+}96, BH94b, Bro94, CHDW92, FB98, GBP^{+}97, GDW90, HC90, HCI93, HHE92, HCH^{+}96, JAL91, KD99, Ker90, Kin95a, LH98, MM91a, MBJ94, PGR94, RPL96, SH97a, TH94a, UK98, WH96, Wat98, GGH95]. near-bottom [JAL91, KD99, RPL96]. near-shore [GDW90, Kin95a, LH98, MM91a, TH94a, UK98]. Near-surface [BE97]. Nearshore [SC96a, JJ99, LADS98, ME97, RGS95, RQB98b, TMKL97, VDG95, Vet95]. Nebalia [Vet96a, Vet96b]. Necora [FGG95]. necrophagous [CPT90, TW98]. needed [JLF95]. needs [CR96, RG97, SB94]. Negative [BCL93, BRA^{+}99, ASN95, CU99, Leo99, WLL97]. neighboring [Cli95, NT94]. neighbour [O’D98]. neighbouring [CB97]. neighbours [CNJ96]. Nekton [CMD^{+}94, KW94, PRW^{+}94, RR93, Kne91, PWR96, RPM^{+}94, Roz92]. Nella [SS95]. Nematoda [HBB^{+}99, SCL95, SLHC98]. Nematode [SH95b, AWdS92, AWH93, AWHH94, AWVP98, BAY91, Ken94b, ÓM92, SW98a, SW99, SHV91, VAV99]. nematodes [DDM^{+}99, HBB^{+}99, KDKS96, MVdM^{+}99, RRGB96, TA90, VVT97]. Nematonurus [APS91]. Nematoscelis [BGC91]. Nembrotha [PLF90]. nemolymph [Vis93]. nentic [CWS^{+}93]. Neobythites [Uib95]. Neocalanus [TSI^{+}93]. Neomysis [AJS95, MAH94, RWJ99b]. Neotrypaea [FAED97]. Nephrops [BPR90, BDH94]. Neptunea [IWF^{+}97]. nereid [MK93a, MK93b, ZHBH90]. Nereis [AOOT94, LBDD97, NEIR95, ORD98, Rii91, RVBL92, RPL96, SHC96, VR93b, VAR94, Ved98, Vis90, ZHBH90]. neritic [CSB97, DFR95, Dut98, FD92, RMMB91]. Nest [HS91c]. nesting [HLP^{+}99a, HGB99]. Net [Lee90, ASRP^{+}96a, ATNW97, BK91, CCG98, CBL96a, Con94, EMR^{+}97, FW97, KSMDB97, KP95, PJ99, RVBL92, Roz92]. net-jet [PJ99]. net-pens [FW97]. Netherlands [KP95, HdWS91, LH99, MAH94, PSPD96, SH94]. nets [HTCW96, KBN^{+}96, RTCH96, RWCM99]. Network [FCCM94, BJTB96, CES^{+}94, CSE^{+}96]. networks [Sca96, SWEC91, VB90]. neural [BJTB96, CES^{+}94, CSE^{+}96, Sca96, SWEC91]. neuronal [SCMM95]. neurons [KB91]. neurotoxin [KDKS96, PRM^{+}96a, PRM^{+}96b]. Neuse [BCS93, BMG99, CBS91, MP94, PPFP98, PPFP99, RPMB91]. Neustonic [LB91]. neutral [KSE96]. Newfoundland [GA97b, KSS90, Kea90, LADS98, PMMH95, PWD^{+}91, PSR^{+}99, UMD92]. Newly [Row90, EE98, JPC97, Jer90, RS95, RH97, TB95]. Newport [TGG^{+}95]. NH [BG93, GMG^{+}92, JLS90, Lai92, LA97a, LJS90, MSK95, VNB97, VBN95]. NH4 [TCB93]. niche [LB91, MM97a]. nicholsoni [SC90]. Nicoya [DE90]. niger [RO93]. Night [SR98b, JBB98, Van92b]. nights [KKH98]. nigrescens [MC93, WMG^{+}94]. nigrofuscus [HK96, HR96a]. Nile [CTEF91]. Niña [RAS92]. nine [MC98]. Niño [RAS97, TBAB91, MADM98, RAS92, Urb94, YHC^{+}99]. Niños [TD91]. niphobles [Yam97a, Yam97b]. nitidimanus [Asa91]. Nitocra [WMH96]. Nitrate [Col98, Fly99, HCF93, Tam95, BCH95, BHHL96, BBO^{+}95, BGC94, DLT^{+}97, DQT^{+}97, Dor90, Fly98, GSBB97, KGHTN99, MA99b, MB97b, ONH^{+}97, OSGB96, RGS90, RWJ99a, RFN98, Sci91, UGSBA94, VH99b, WYY^{+}98, WWB98]. Nitrification [LGV^{+}99b, MSK95, BBF94, DW97, GSBB97, LMS^{+}93, PBA^{+}95, RCN95, VGJ90]. nitrifying [GJ96b, GJ96c]. nitrite [CC92b, Col98, DvRN99, SA94]. Nitrogen [ASRP^{+}96b, ASRP^{+}96a, BEU99, CK90, CSG93, CCC94, Coc90, CM93, Col92, DvRN99, Fra94, HHKS92, Har90, HCS90, KFF92, KM99b, KEvN95, LS96a, LCW99, Lip95, LMS^{+}93, MC99, NT96c, NK93, PB92, PT95, RPDR^{+}98, SCLS94, SBH^{+}95, TNSM98, VBN95, Vil91, VNM97, WYY^{+}98, WYY^{+}99, vdL98, AFP99, ASP91, ATNW97, ATH95, BG91, BG93, BVS^{+}91, BAM92, CDHD92, CLC^{+}98, CA95, CPH91, DON^{+}95, DH90, EPN^{+}99, Far98, FPGC^{+}98, FCMB94, Fly90a, Fly90b, FFF^{+}94, FCCM94, FHCC98, GCH93, GBA^{+}96, GAK99, GST^{+}91, GAN^{+}94, GTJ92, HS94a, HHB93, HHH97, HPSJ95, HBB^{+}99, HCG^{+}93, HSCC96, JKC^{+}93, JTB99, JP93b, Käh98, KSRB99, KK91, KJN97, Kro91, LC95, LC90, MV98, MRPC98, MBK95, MO97, MHM90, MLH96, MH96b, NBW94, NNR95, NG96, ONH^{+}97, PWG^{+}99, PB93, PB96b]. nitrogen [PB97, PLC^{+}91, PPR94, PPCR97, RDPB99, RSN^{+}99, SBV98, SA94, SS97, SR93a, Sim91, SR95, TW97, Tur93, VB90]. Nitrogen-fixation [Vil91]. Nitrogen-limited [LS96a]. Nitrogen-specific [Lip95]. nitrogenase [CCC^{+}90b]. Nitrogenous [SCD^{+}90, CM90a, CC96, CC97a, RN98]. Nitrous [BH93, KLEH91, PDP^{+}90]. Nitzschia [ES99, MHBW90, Mik95, DRP^{+}97, PRM^{+}96a, PRM^{+}96b]. nm [VMS^{+}96]. NMR [BWL^{+}99, TR91, TRPS94]. No [SRRdM99, CR96, DHB^{+}98, RQB98a, JLS90, LJS90]. nodosa [DSJ90a, DSJ90b, KM99b, MD94a, PDC97, PH91a, TDK97b, TDK97a, VDM^{+}97]. nodosum [Åbe90, Åbe96, ÅP97a, PCLÅ97, VWM90a, VWM90b, VÅCL99]. noise [TSI^{+}93]. noltii [IF96, KM99b, Phi94]. nomadica [LFBH94]. ’non [KEvN95, BDM96, FF92a, GS99a, JNCS89, AAD91, BS95c, EDP^{+}98, GGB^{+}99, Han95, JCCCB94, KEV99, Las98, LCC97, MG96c, Mee99, MS95, PM94a, Reu98b, RHP93, Spr93, SNR^{+}96, Tok95]. non-boreal [Spr93]. non-calanoid [BS95c]. non-copepod [GGB^{+}99]. non-exploited [Las98]. non-flying [JCCCB94]. non-indigenous [EDP^{+}98, Reu98b]. non-invasive [AAD91]. non-living [MG96c, MS95]. Non-occurrence [JNCS89]. ’non-polluting [KEvN95]. non-preferred [PM94a]. non-producing [SNR^{+}96]. Non-protein [FF92a]. Non-random [GS99a]. non-stratified [Mee99]. Non-structural [BDM96]. non-terminal [LCC97]. non-toxic [Han95, KEV99]. non-trivial [Tok95]. non-vegetated [RHP93]. nonhierarchical [BCC96]. nonlinear [EP92]. nonmotile [GSS93b]. Nonparametric [LC97b]. nonspectral [SPS^{+}91]. nontoxic [BLKA93]. Norderney [KDHZ98]. Nordic [STS99]. Nordrügensche [KDMR97]. Normandy [SBT97]. Norsminde [NNR95]. North [BYO95, CC92a, CT92, CBS91, FB98, HC90, MP94, NLR^{+}93, PPFP98, RG93, RC91b, RPMB91, Arr93, CCV^{+}98, GMF98, WPTB96, AFP99, APS91, BM99b, BLDF95b, BLDF95a, BD98, BCS93, BM95a, BVL98, BvBBV96, BCv^{+}92, CAATS94, CSR^{+}92b, Cha92, CBD96, CMLE92, DvdM92, DAB^{+}91, DHH98, DCBC91, DWG98, FGKP95, FGS98, FH99, FP96a, GBC92, GRC^{+}97, GH94a, GCF96, GR93, GLBF92, GCWH90, GCNSNd94, HWH98, HKMK98, Hil97, HB98, HNvB94, HW91, Hov92, JKT90b, JL96, JP93a, JBB90, KJT90, Kep91, KW92a, KW99b, LIPF95, LK96, Li95, LGH95b, LACM92, LMS^{+}93, LKvRH96, LEHvR96, LGdC^{+}90, LSH96, LK99b, MG97, Mar97, MHC92, MC94a, MF98, MMBT97, Mun93, MLDM99, MF95, MHB96, NS96, O’C93a, PCM^{+}95, PLFSW97, PKY96, PF96, RWJ99a, RA92, RNP^{+}98, Rie93, RFN98]. North [RN98, Rv92, RVv92, RS92, SPB^{+}99a, STK99, STNI97, SH92a, SD93b, SH95b, SSV97, SR97, TAJK90, TGG^{+}95, TF98, TSI^{+}93, UNPH93, Väi90, VKV95, VK90, WT98, WSM91, WT99a, WDB^{+}95, WPG99, vDvRK93, vdBTvDR96, vCRS92, vPB90, vdVW99]. north-east [GMF98, RWJ99a]. Northeast [PLG^{+}96, RTRP^{+}98, RNH99, SGK^{+}97, VD94, BSGM94, FBHH93, GAK99, HM94a, IT96, RH98, SJML99, TFL95, ADBH99, BK98, BAK97, Bra95, HPPZ93, HAR95, WWB98]. northeastern [AM91, CK98, FNJ^{+}94, Jon93, Kin92, MLDM99, TK95a, TK97, TK98, TT99, TC94, TTR99, TA90, VHH^{+}98, VH99b, YNYK94]. Northern [Hag95, PH91a, AR93, ANW96, BGC91, BM99b, BMSF95, BDR^{+}96, BLC91, CO92, CW93, CLD99, CH94, DRP^{+}97, DG96, EM99, FML^{+}95, Fal96, FAVV98, GGF99, GH92, HTG96, HD91, HSN93, KKL91, KVK98, KWL98, Kuo91, LHRL97, LP95, Leh96, LA98, LVP98, LL92, LKK92, LHK^{+}93, LL97, LSB95, LFR^{+}97, Mac91, MH93b, McC94d, MDSB92, MB91, MZFHC96, MTP92, MNH92, NPB94, Naj96, Naj97, NNB93, OTF95, PG91, PF91a, RSM^{+}96, RHJ92, RHrsHG99, STBR99, Sam97, SH95a, SS96b, TCRB99, TFL95, Tuo97, THKN98, Uit96, UHL^{+}97, VH93, VHH^{+}96, WBLM95, WC92, WS92b, dT94a, BEU99, BH92c, BH92d, GD96, KH92a]. Northwest [ERC95, HCS90, SCEG93, BCMC99, BR94, FPR94, FQ94, HB94a, LMM98, TDBH94, TP95, TP97, WSC^{+}99, TRS91]. northwestern [BBF94, BSP98, BBP^{+}96, BGCM95, CCS98, GDP^{+}98, MZ99, MB97b, OSGB96, PBv^{+}95, RFNBHG94, RGBC92, SRH^{+}98, TB92, VHVB^{+}98, WHG99b, YST^{+}97, EML^{+}93, LED92, Pal91, Sab90]. norvegica [MVA99, OK98, STS99, SPZ^{+}93, TCRB99]. norvegicus [BPR90, BDH94]. Norway [GGB94, ANW96, BPR90, BA93, BRLM90, BHHL96, ES90, Fal96, FB90, GHHC98, Hag95, HSN93, NC93, PGG96, SCL95, TSB93]. Norwegian [BNvB90, Ber90, FSG^{+}97, FSG^{+}98, Gra94, IHK^{+}98, KR98b, LSG^{+}99, NKH^{+}99, NNB93, OG95, ONR98, SAG92, WBGS97]. Norwegian-Greenland [WBGS97]. notabilis [Nak92, Nak94]. note [CCJC91]. notes [VWK96]. notorcadensis [BH92e]. Nototheniidae [WHK97]. Nova [JNCS89, CHS^{+}95, CSB94, HS95, MHS94, MS94b, SH90a, SH97b, SM91]. novaeangliae [BSF95, BFS^{+}98, Mik97, RCA^{+}95]. novazelandica [RS99a]. Novel [HSBH96, RHS99, Bar97a, CED91, FD97, JCH93, RVGC^{+}99, SMB97, VMS^{+}96, WKH^{+}98]. November [GST^{+}91]. nubbins [vTS97]. Nucella [BQR99, CH96, DHB^{+}97, RQB98b]. Nucella-Mytilus [CH96]. nuclear [BZ92, Cam97]. nucleic [DMP^{+}99]. nuclides [GF90]. Nucula [Cra94]. nudibranch [AP97b, BPS98, BSS92, CHFL95, MTP92, PLF90]. nudibranchs [RP91a]. null [SW91]. number [GCM98, GB93, GSMM96, Ito97, Str99, Tak99]. numbers [BH92b, GK96b, Hei91, MGC95a, Nag96, RS94, SPPL92, TBAB91, ZR92]. Numenius [RD97]. Numerical [BMH91, KDC99, JBK99]. Nurse [CP90]. nurseries [LJ95]. Nursery [RGE^{+}95, Ang95b, Beu92, CLD99, EG95, FS98, GPB^{+}98, Kne93, MF95, TJST98, VDBF96, WS91]. Nutrient [ADF96, Byt90, CM90a, Eno93, FPAH92, FDZ94, HCC^{+}91, HKY^{+}97, HJA99, McG95, MH94a, PB96b, PB97, PGR94, PLS^{+}95, RBD^{+}97, SK98a, SJ95, Som96, SAvD^{+}96, TNGB95, TR95a, TCRB93, ZB99b, AE91, ARDL97, ADG^{+}95, BF99, CRCS96, CCRB98, CPH^{+}96, CPB96, Cra94, DLT^{+}97, DHM92, DAS92, DSV^{+}91, Dua90, DGHT^{+}98, DCBG99, EA92, EM93, FGP98, FRP^{+}99, FF97, FF92b, GMK97, GM98a, HGS94, HTG96, HHv91, HS97a, HFJS91, HG94, Hüt90, JSJMH90, KLMR91, KMR98, Kel98, KW97b, KR98a, KBWC90, KJMRC96, KHT93, LA97a, LL92, LFR^{+}97, MP94, MR97b, MCA90, MRP98, Moe96a, Moe96b, Moe97b, Moe97c, MAdG^{+}96, NT92, OL99, ODN^{+}95, PJF93, PBA^{+}95, PB92, PDC97, PS94, Pre91, RCS^{+}96, RGS95, RLC92, RDE^{+}90a, RDE^{+}90b, RPMB91, RT92, RTRP^{+}98, SRSH99, SK98b, Sch99]. nutrient [SR93a, SD93b, SCD^{+}90, SMH97, SM97c, SEGP91, TS90a, TA90, TL91, VD95, YGH^{+}97, ZB94, vDvRK93, vKSGvA99, PB96b]. nutrient- [GM98a]. nutrient-enriched [Cra94]. nutrient-limitation [GMK97, PS94]. Nutrient-limited [SK98a, LA97a, Moe96a, Moe96b, Moe97c]. nutrient-poor [DHM92, MAdG^{+}96]. nutrient-rich [DHM92, Moe97b]. Nutrients [CPH^{+}96, CBS91, Har94, JCMW91, KH92a, KLAH^{+}94, MNO96, NJSS91, NO90, PWG^{+}99, PPF95, RN98, RNH99, UDLM99, VSQ90, WLB^{+}99, ZNH93, BT93]. nutrition [BS96a, Bus96, DB99, FAVV98, KG94, MCC97, MH96b, SMB97]. Nutritional [APCL^{+}93, FCMB94, KL94, SJ97, SLSH98, YSY^{+}99, CBI91, CP95a, CLM^{+}93, CR96, CD96, DS93a, GAC^{+}97, PHHH96, PSB^{+}97, PDB^{+}97, RG97, WDB98]. nutritive [NT95]. NW [ADF96, ASRP^{+}96b, ASRP^{+}96a, CMRF93, CSM99, DLFL96, FGG95, PF92b, TFFA99, Vid94, VM95, ARDL97, BQR99, BG99, CJ97, GCR98, GF95, GLD^{+}91, KDR97, LBL91, MPS96, MC98, MAdG^{+}96, PMB99, PDF97, Pre91, RATA^{+}99, RMGB^{+}94, RGBJM95, RQB98b, SZ96, SPS^{+}96, SFPCR93, TB98, TGB99, VF99]. Nyctiphanes [Gen92, GGDLE96, Góm95, Lav94, PM94a].

O [AGP98, Ano99c, FPFR92, GRGK96, GT98, JLGMP97, Kal91a, Kru93, KCD^{+}95, LGR95, RVGC^{+}99, RNHO98, RBN99, Sie93, SJMC93]. O-deethylase [RBN99]. O [?]Guzman:1998:SVS. Obituary [Ano96b]. ’objective [DGCT93, Zho98]. obligate [MJ97a]. obligately [Sie93]. obscurus [KJLH96, KKYH97]. observation [SM98b]. observational [Pre95, WM99]. Observations [MSS95, Moo94, Wol93, BMHD93, Bro98, BKH99, CGFC95, GJC97, GNJ90, HBM93, HSBH96, HVL96, KS97, LCK^{+}96, OIO^{+}95, SML97, SLHA97, SD93a, SlRR94, TNKMK99, TJ90b, Tis98a, TL91, WMB99, Wit99, WA95, dBN^{+}90]. observatory [SRJD97]. Observer [TM97]. observers [BCAAC96]. obsoleta [DMM99]. obtained [GKB^{+}97, LM97]. occupancy [TR95c]. occur [Sie93]. Occurrence [Gra96a, JMD^{+}99, JP96, JTB99, KR95, PDB^{+}97, SGPAA95, ZN96b, ASA96, CS91, FWWL99, Hin92, JNCS89, JP94, KFPR94, KVK98, LP95, MHRV^{+}99, TAD^{+}94]. occurring [CDME97, CCJC91, JLYK97, KM99b]. Ocean [KID91, KPS92, PDBA90, Per91, PLM92, VPM96, WSCT96, AP99b, AMMG91, CA90, CCR98, DZ99, GGH95, HFG^{+}95, HDK93, HRCW98, Kel98, LBLA93, NT96b, PA94, PWP^{+}97, PGAJ99, Ris97, Rud97, SCA92, Tan98a, Tan98b, WHIC97, WPTB96, YFI91, ZD99, AMO^{+}99, Arr94, ASSWM97, BSGM94, BFS^{+}98, BK91, BP96a, BAK97, BKL99, BSP98, CBG^{+}99, CR95, CMB^{+}99, CAATS94, CW99a, DSG^{+}99, DJY90, DLJ91, DB96b, EY90, FBHH93, FP96b, FPP^{+}97, GR93, HM94a, KTW^{+}97, LK96, LNQT93, MH99, MACR98, MD94b, MO95, Nai93b, NSC^{+}92, PWG^{+}99, Pak95, PPM96, PLH91, PM92a, PWM95, PTB^{+}97, PFH98, PMOD90, RH98, RWJ99a, RSH92, SB98, SPS^{+}96, Str91, TvLdJ^{+}98, TM94, WASSM97, WSC^{+}99, WWC^{+}99, WWGK95, WWB98, WTM^{+}96, YD90, dBN^{+}90, vLTW^{+}98]. Oceanic [MG91a, AR99, Atk96, BWHK99, BFJ^{+}93, BVP91, CWS^{+}93, DBMK99, GHD92, HKT^{+}96, HHP95, KID91, MLH96, PWPK95, SJCL93, SC93, TSI^{+}93, VH99b, YS96]. oceanica [ADR95, ARDL97, CDM^{+}96, DSS^{+}98, DBdCS99, EAD92, MDC^{+}96, MR97b, PRPM^{+}94, PAM96, PM98, RGBJM95, VWS93]. oceanographic [HHM97, MR97a, Mil94a, RSH92, SH97b, SKM^{+}97, TDER97, WTR99]. oceanography [GTYR97, RPT^{+}96]. oceans [DAS92, LR96, LM98b]. ocellatus [BA90, RH97, SHFP94]. octocoral [Ger90]. Octocorallia [EL90, RCG99a, SC92a, WV95]. octopus [Iri90, And97, Iri90]. Oculina [KRFL97, KRF^{+}98]. Ocypodidae [SCAT91]. odor [AMMG91, FFT98]. odors [BBS93, Nev99, RHH94]. Oegopsida [NAS95]. off [AFP99, ASA96, AI92, ANW96, AM93, BYO95, Bia92, Bra95, BK92, Cer98, CCV^{+}98, CB99, CJ93b, Dur94, FTS94, FCH99, FII95, GSC96, GAD^{+}92, GGH99, GGDLE96, Góm95, GSMM96, Ken92, KKW97, KWL96, KDHZ98, Lav94, LIPF95, MP96b, McG94, MP95, PHHH96, RGE^{+}95, RO93, Sea90, She98, SDG95, Urb94, VMHH90, WPG99, YNYK94, YBLL96, vRBVB95]. offers [RHS99]. offs [WB91]. Offshore [BH94b, WD94, BM94d, BMH93, FFR94, KCS94, LADS98, LMS^{+}93, MH96a, OG95, PDD^{+}99, Pir92, RSN90, Rud97, ZN96a]. offshore-fjord [BMH93]. offshore-to-inshore [KCS94]. offspring [GSMM96]. ohshimai [TK95a, TK97]. Oikopleura [DCRG92, LUA99, UMD92]. oikopleurid [Flo91]. oil [AGW93, Ber90, DvdM92, DF94, ER90, GCKW95, GBJ94, JDSB99, MHC92, OG95, PDD^{+}99, SS95, WTA^{+}97, vTSD97]. oil-contaminated [DvdM92]. oilfields [GCWH90]. Oithona [CLW99, GS94b, NS96, US95, US98]. oivalves [DS93a]. Okadaic [BFMM99]. Okinawa [HSK^{+}93, Sak97]. Olavius [DAE^{+}99]. Oléron [GB95, GBB^{+}98, KSRB99, LGV^{+}99a, LGV^{+}99b, RRGB96, RR97b, Rie98]. Olfactory [Nev99]. oligochaete [WLL98]. oligochete [DAE^{+}99]. Oligocottus [Mga92]. oligotrich [ADR94, KS95d, PDF97]. Oligotrichina [KS95d]. oligotrophic [CVD99, DMP^{+}99, DCBG99, HS98b, RWJ99a, SCIA99, STK99]. olivaceus [YNYK94]. ommastrephid [CW95, GTYR97]. Ommastrephidae [WTR99]. Omnivorous [BN91a, GD90, NEIR95, WLL98]. Omnivory [AS97]. Onagawa [IT96]. Oncaea [BS92, HLBW93, Met98]. Oncorhynchus [FS98, SOK98]. One [Kin93, SH94, BFMM99, CCR98, DBB^{+}95, EPL94, FFH99, BMFR97, Fow90a, SU93]. one- [BFMM99, BMFR97]. one-dimensional [CCR98]. onset [AOT97, EST95, NNB93, Nor96, SK97]. Onshore [MH96a, SMW^{+}93, BM94d, ZN96b]. Onshore-offshore [MH96a, BM94d]. Onslow [CC92a, CT92]. onto [HB97, HH99, LB95]. Ontogenetic [GW97, Joh99, O’B94a, OF94, PODL99, AH98, IMB99, Egg95, FTS97, Gre96, OW95, SK97]. ontogenic [TD92]. Ontogeny [CCD94, CCDA98, CA99b, SHFP94]. onuphid [PL99b]. onychoteuthid [CW99a]. oocyte [ORD98, WCP^{+}98]. oocytes [MGDT90]. Oosterschelde [HdWS91, PSPD96]. oparva [Sal96]. Open [WSCT96, GR93, KLN^{+}90, LGJ99, Leh96, LA98, LCW99, Mik95, PWP^{+}97, RDE^{+}90a, RDE^{+}90b, Str91, Tho92, WPTB96]. open-ocean [WPTB96]. open-sea [Leh96, LA98, LCW99]. open-top [LGJ99]. Operated [SCC94]. operation [MWT96]. Ophidiidae [Uib95]. Ophiothrix [MDG98]. ophiuroid [GSP99, PLFSW97]. Ophiuroidea [LJSK96, MDG98]. ophiuroids [Amb93, Aro91]. opilio [SMUS^{+}99]. opisthobranch [GC95a, Pen90a]. Opisthobranchia [Ito97, MTP92]. Opportunistic [KR97a, FFH96]. opportunity [LW94]. optic [KLJ94]. Optical [MG96c, BSP98, HCS93, KMKS96, MKV95, SPS^{+}96, SPBS93, Sos96]. optimal [ER95a, vW93]. optimization [BB97b]. options [vTS97]. optrodes [GRGK96]. orange [BK92]. Orchestia [RK95a]. Orchomene [Moo94]. orders [SAH98]. ordination [MF98]. ordinations [VWJ96]. Oregon [FC94, JT93, ME97, NC94]. Orest [Ano94]. Organic [AV95, BHHL96, JCP93, KSMDB97, MKN^{+}96, MO95, PM91a, TA95, VD99, AC96a, And96, AR93, AJST98, AB96b, BNK^{+}97, BRL97, BC93, BK94, BL94c, BL96, BG91, BG93, BvBBV96, BLR94, CRCS96, CCR91, CDG97, CBD96, CR94, CCH93, Col92, DSG^{+}99, DFJ^{+}93, DÁSP97, FBHH93, FM95a, FPGC^{+}98, FPLH92, GBA^{+}96, GKMT99, GAC^{+}97, GP91, HB91a, HHB93, HBBM96, HKvB93, HSRHH90, HK94, Hol96, HDK93, HG94, Jos90, KW92a, KE95, Kir90, LM91b, Lig90, MFF97, MRL96, Mal96, MS93a, MKYI98, MH90, MWH91, NKI^{+}96, NK92a, NPB94, NBW94, NMR97, ONM93, PPFP98, PDC97, PLH91, Pel96, PS97, PIC97, Pre91, PF91b, PPR94, PPCR97, RDPB99, RR97b, Rie98, RTLG97, RSD^{+}91, SWT99, SW98a, SR98a, SS97, SB94, SG96, SR95, SBH^{+}95, SM95, SDA^{+}92a]. organic [TL95, Tsu90, TFFS90, VB90, VKV95, Wes90a, Wes90b, YST^{+}97, ZB99a, ZNH93]. Organic-N [BHHL96]. organic-rich [DFJ^{+}93, HK94]. organics [FWWL99]. Organism [VS96, SOS^{+}92, Sch90, WLT^{+}95]. organismic [MIA90]. organisms [AAD91, BJdJ^{+}98, DeV92, DE91, Gar98, GPF95, Hei95, HHB97, KI97, KCI98, MJ92b, NTK93, NT94, Nan96, NFT92, PCB90, PLP92, RHS99, Rog90, RGBC92, Tak99, TB99, WD90, WPY98, Yam96]. organization [MCC96a, MS93b]. organochlorine [PTW^{+}91, WT99a]. organochlorines [MCM^{+}97]. organophosphate [RWJ99b]. organotin [BBR^{+}93, IWF^{+}97, RQB98b]. organotin-induced [BBR^{+}93]. organs [MACR98, Pro92, SDC^{+}92]. orientalis [Jon93]. orientation [KMR98]. Oriented [ORL97]. Origin [Lam98, SM98a, CSG93, FM95a, IVM^{+}98, JPZ^{+}98, Nov90, RTRF90]. originating [FFH99]. Orkney [BW96]. ornata [SP99b]. orthophosphate [ADR94]. Oscillation [FP96a, RAS97, TBAB91]. oscillations [Rud97, TN98]. Oslofjord [Amb93]. osmeriid [DL93a]. Osmerus [DD90, Sep94]. osmoregulation [CCD94, CCDA98, CA99b, LCTT92]. Osmotic [LLB^{+}95]. Osmotic-stress-induced [LLB^{+}95]. ostracod [Hul99]. Ostracoda [JGT96]. ostracods [Hul97]. ostralegus [CGCLH99, McG97a]. Ostrea [Bro90, Bro93, CT98, TN90, TVG96]. Ostreidae [BD96, RM95]. other [BAH94, CGFC95, DJY90, DFS94, Gow93, HS92, LCLC96, MW97, MJ92b, NR98, Pur91, RGLS99, SOS^{+}92, SMW^{+}93, Vet96a, ZF97]. Otolith [Lim95, AOT97, CA96b, CD96, EF97, GS99a, HW91, JD90, Kal91b, MJ91, MCFB97, OINN94, RSM^{+}96, RG93, SFF92, TRMW95, TSW97, WT98, YB90]. otoliths [Ano99a, BK90, CGM95, Cam97, CMBC98, Fow90b, GPP98, GK96a, Kal91b, Kal91a, MC91a, Rv92, Sep94, SKPT97, TJST98, WM97]. Otsuchi [Nak92, TK95a, TK97, TK98]. otter [PSR^{+}99]. otters [KO92]. Ouse [KGHTN99, NT96c, TNSM98]. outbound [FH97b]. outbreaks [BMBD95, CED91, HK96, HKR96, HR96a, SH97b, SB99]. outcome [Nan96]. outcomes [THPC96]. outcrossing [HHG99]. outdoor [HLF^{+}91]. Outer [RSK^{+}99, Fab95, HSLV94, HGS95, HBC^{+}91, KS98, Sha98]. outer-shelf [Fab95]. outflow [Fuh90, MF90a, MF90b]. output [BGBS91, HvdM97, RSB93, SS91, UMTM98, vG91]. Outwelling [TA95, Alo90, HSK^{+}94]. Ovalipes [BA90]. ovalis [NA99, Ral98]. Overall [SDA^{+}92e]. overexploited [SOW97]. overfishing [RAGH94]. Overgrazing [RSK^{+}99]. overgrowth [Aer98]. Overlapping [Bla90]. overlooked [MM97a]. Overwintering [BLJ92, DBMK99, HVK96, TDH^{+}94]. Ovulidae [VT92]. ovum [SC92b]. Owenia [TD92, VBCRB97]. oxalate [CBA^{+}99]. oxic [And96, BR91a, BR91b, FGKP95, NMR97]. oxidases [PM90a]. Oxidation [Sch91, BR91a, KTGC99, PPR94, PPCR97, RDPB99, SD93a]. oxidative [Les96, SJMC93]. oxide [BH93, DHB^{+}97, KLEH91, PDP^{+}90]. oxidizer [PDP^{+}90]. oxidizing [SJCL93, VFMFBE96]. oxolinic [PLBP96]. Oxygen [BCG94, LEHvR96, PBDF98, AB95, AMB^{+}96, AJST98, BPR90, BSD^{+}94, BRC99, Cam91, CCC^{+}90b, CCH93, CPB96, EJ96, FW97, FHZ96, FRP^{+}99, GCLS94, HDPW93, HdWS91, Irw91, Joh97, KD99, KSG^{+}92, KJLH96, KBWC90, KR95, LGJ99, MFD98, MM91a, NH98, RGS95, RLC92, RMG^{+}98, Sal96, San94a, SEGP91, TDBH94, TH94a, TMC98, UNPH93, VTNA91, VH96, VDA96, YB92, ZHF96, vRV96]. oxygen-protective [CCC^{+}90b]. oxygen-stratified [BRC99]. oxygenase [GCKW95, RA92]. Oxyrrhis [WSS94]. oxytetracycline [PLP92, PLBP96]. oyster [BM94b, BN91a, BN95, BH94a, BV99, BPL95, Bro93, BA96, BCv^{+}92, CT98, DHKP96, Dou93a, Dou93b, DLHB99, ERC95, GMC91, LHL97, McF92, PLBP96, RR97b, RM95, RAGH94, SCD^{+}90, Tha92b, TN90, TVG96, dRAV92]. oystercatchers [CGCLH99, McG97a]. oysters [AC99, MAAT98, RCH98, YKL98a, YKL98b, YKL99]. ozone [Arr94].

P [TR95a, AGP98, BRLM90, BR97, CJ95, FMFH97, FPFR92, GJ97b, LH91, OH95, SV97, TR91, TRPS94]. P-limited [TR95a]. P. [CLD99, Hun93, Kep98, Ker96, LWK^{+}95, PW94, TRPS94, TB99, WFHM99, YKL98a, YKL98b, YKL99]. P04 [WK92a]. P450 [GLBF92, SGR92]. P450IA [RA92]. Pachyptila [GBP^{+}97, ASF98, GSM^{+}97]. Pacific [LK96, WWB98, AMO^{+}99, ALOT99, APS91, AMK99, Bak90, BSGM94, BSW^{+}95, BWHW95, BMV^{+}96, BDA^{+}99, BVL98, CCG96, CBD96, Dou93a, Dou93b, DWG98, DCBG99, ERC95, FF92b, GCNSNd94, GCGR90, HPPZ93, JWL96, JW96, Kir90, Kir92, LKC95, LCE97, Liv93, LSB95, MPM97, MBR94, MO97, MLH96, PLH91, PHL92, PFMG96, PM91b, RH98, Rei90, RAS92, RAS97, RH90a, RZLB96, Sal96, SJML99, STK99, STNI97, Str91, TT99, TvLdJ^{+}98, Tok95, Tro94, TSI^{+}93, VH99b, WWC^{+}99, Web91a, Web91b]. pacifica [DWSA98, Tan98a]. pacificus [DWSA98, HL90, LAA93, Lop91, OF94, OC97, Uye96]. pack [GB91, GC93, Gra96a, LS91]. package [NPB93]. pages [Kin93]. Pagrus [MMBT97, OC93b]. Paguridae [Har96]. pagurus [HBR^{+}91]. PAH [ONR98]. pair [MSP^{+}97]. pairing [LKvRH96, NG96]. Pakistan [BNK^{+}97, HKY^{+}97, KDAS92]. Palaemon [NH98, TR91, TRPS94]. Palaemonetes [CA99b, GF98, NH98]. palaemonid [CA99b, Gra91b, Gra91c]. palaeoenvironmental [KML91]. Palatability [MB97a]. pallasi [Liv93]. pallida [AP97b, BPS98, MPH96, MPCD90]. palmata [Byt90, MB95b]. palps [BSJ97]. PAM [BVW^{+}98, HWLC98].

Panama [GCGR90, GBJ94]. pandalid [Car93, CS97]. Pandalus [Ber92, OTF95, RT95]. Pandulus [HSN93]. panel [Nan96]. panicea [RTJ^{+}93, TR95b]. Panopeus [LK94, RDA97]. Panulirus [AADSCC97, BHH^{+}95, ELMCC90, Jer90]. Papagayo [Bia91]. papillosa [RCG98]. Paracaudina [TTW97]. Paracentrotus [FB97, FG93, LTM^{+}98, LGL^{+}95, SZ96, SRH^{+}98, TGLP95, WLT^{+}95, WTF95, WTF96]. paradigm [Mar99a]. paradigms [VL97]. paradox [BBC^{+}97, DvdTCC97, GWWP99, SRSH99]. Paragobiodon [KYN94]. Paraiso [ER90]. paralarvae [Big92, BSY^{+}99]. Paralichthys [BB95, MT91, YNYK94]. Paralithodes [SK98c]. paralytic [BLCA90, BLC91, KOSS90, KB91, SOS^{+}92]. parameters [AT96, BLST90, Bus96, Fly98, GC95a, JSV92, JJNR95, MC95, MH99, TN90]. Paramoeba [JNCS89]. Paramuricea [CGZR94, CRZG95, CZG95, MC94b, RCG99a]. Paranagua [dG91a, dG92a]. Paraphysomonas [PCG96]. Parasagitta [TNN93]. Parasite [Gru96, MJ97b, BA96, SMG97]. Parasites [Spe95, BB97a, SW93]. parasitic [GP98, Laf93, MTB^{+}99]. Parasitism [CUM98, Flo91, GMC91]. parasitized [Wat98]. Parazoanthus [Gar99a]. parenchymella [MGYV97]. Parerythropodium [KKL^{+}98]. parrotfish [BKB94, BvB94, BBB^{+}94, BvKvRB96, Cli95, TT90, TNP97]. parrotfishes [Bel95, OM99b, PPP^{+}96]. part [HM91, HFN91]. Partial [SWK98]. Participants [SDA^{+}92d]. participation [MC91b]. Particle [BSJ97, BDB97, CB96, DBB^{+}95, Kep94, MFF97, PAR98, Rie91a, Rie91b, WBGS97, ANW96, BSJPW93, BD96, BVP99, HKA96, KKL91, KS95d, KF97, LAA93, MW94, McL96, PF91b, RM97b, Tee99, van94a]. Particle-attached [CB96]. particle-bound [KF97, LAA93, MFF97]. particles [AC92, AA92, BN95, DBdCS99, ES99, GS93, Gow93, KL91a, Lip95, MS93a, MB98, Mar99b, MG96c, MJ92b, NAI^{+}93a, NK96, Nov90, PA94, Put91, RTRF90, SH95a, SKKO97, SDB92, SDB94, SH99b, TM94, Vis97, WR91]. particular [ACR99, TH90]. Particulate [CCR91, CDG97, DMP^{+}99, PLH91, SHML98, AC96a, AR93, AW98, BMS^{+}97, BMTF92, BRL97, CBDC96, EJ97, GAC^{+}97, HKvB93, HSRHH90, HDK93, LLP94, MHM90, NNB93, ONM93, PGR94, PF91a, PSE97, PSPD96, PPR94, RDPB99, RR97b, SCJ99a, SR98a, Wai90]. particulates [YKL99]. Partitioning [PBv^{+}95, SB97a, BGC91, GYL93, HS98b, Ken94b, LL92, LL97, Mal96, PAR98, PS98, SWT99, WM95]. parvula [dNSR^{+}96]. Pass [Che93]. Passage [LDV95, CMT94, PMOD90, SC92b]. Passamaquoddy [PRC^{+}92]. passing [NHL99]. Passive [LJSK96, BM91a, HB97, JBK99, Ker90, TPR98]. Patagonian [Iri90]. patagonica [KRFL97, KRF^{+}98]. patagonicus [ROANM98]. Patch [OSJ94b, RS99a, BH98, CCRB98, CWWH99, DLJ91, DSJ90b, JBHH97, KD96, Lew97, NA99, NC90]. patches [BCC94, BF99, CCR98, DSJ90a, Hol97, HS95, NKGK92, NIU^{+}98, Vet95, VDM^{+}97]. patchiness [DV92, LCF99, MAC92, SNJV92]. Patella [CGCLH99]. patens [Pez91, PD93]. paternity [SB97b]. pathlength [AFZ96]. pathology [Moo92a, Moo92b, SH92a, Vet92]. pathways [Ano99a, CBG^{+}99, JRMN^{+}97, KHFP^{+}96, KTGC99, MJSDF91, TC93]. Patinopecten [GF94]. Patiriella [Hun93]. patronus [CMBC98]. Pattern [PF95, AM93, BCMC99, BGVF92, ECF^{+}97, GGB94, HT96, HSW99, JFS90, LT97, LFBH94, MH96a, MBB99b, O’D98, PMB99, RII93, Sab90, SWEC91, SCJ99b, WP92, WWL97]. Patterns [Åbe96, AMK99, AML95, Bak90, BCC96, BCA^{+}99, BCS93, BM95a, EDSJ95, HBR^{+}91, HSH94, HSK^{+}93, LK99a, MJ91, PS98, RS99b, Sea90, SMTL99, TZ99, UTB^{+}95, WB92a, Adj97, Aer98, ASA96, ASL99, ASRP^{+}96b, AS94b, AB92, AT96, BM94a, BM99b, Bel95, BCC92, BF95a, BGW96, BA97, Bla90, BM94d, BPL95, BG93, BPC98, Bur94, Cal95a, CHS^{+}95, Cha98, CU99, CBS91, CC97b, Cli95, CCG^{+}92, CK98, CS97, CPB96, DJY90, Duf92, DG93, ESDK94, ELH92, Fal96, Fow90a, FSG^{+}97, GMR^{+}94, GY97, Góm95, Gra91b, Gre96, GBB^{+}98, HD97, HTCP96, HGBW99, HN97b, HHB97, HS99b, HC94, Hsi95, Hun93, HS98c, JMC93, Joh99, Ker96, KML91, KL91c, KA93, Kre93, LM98a, LT93, LD94a, Lav94, LSG^{+}99, MU98, MYYT96, MGP^{+}96, MCED96, MS94a, MP96b, McG92]. patterns [MMD93, OSJ94a, OSC98, ODS^{+}96, Onu96, PRC^{+}92, PCB90, PD95, Pin94, PF96, PD90, PD94, PM91b, QCM94, RKH96, RH98, RB99, RPN95, RO93, RDA97, RH97, RK99, SR90a, SO91, SMT97, Sch90, SW91, She96, SD93b, SV93a, SK94b, SC96a, SJC99, SP99b, SL94, SR96, SF91, TWS^{+}95, TSM^{+}94, TcBW96, Tok95, Tom91, TN98, UMTM98, Van92b, VL97, VD99, VHVB^{+}98, Web91a, Zel97, Zel98, vDK90, MH94a, NBS^{+}97]. Patterson [PF90]. payability [Tro94]. PCB [WKH^{+}98]. PCBs [MCM^{+}97]. pCO [CH97b]. PCP [KJLH96]. PCR [MRF^{+}99]. pea [TTW97]. pearl [LGP95, YKL98a, YKL98b, YKL99, dRAV92, LL98]. pearl-like [LGP95]. peat [BAH94]. Pecten [CCJC91, HBBT98]. Pectinaria [TDW96]. Pectinidae [DLDL93, PRC^{+}92]. pectinifera [Kur99]. pedal [AS95]. Pegea [RR99]. pelagial [Han91b]. Pelagic [KKW97, LJM99, NNB93, RDE^{+}90a, RDE^{+}90b, SJJMH90, AC92, And94a, BC90, Bra95, BP92, BRN^{+}95, BGvDR96, CR94, Dei90, DCRG92, HM91, HTG96, Hil91a, HSW99, HAR95, IW91, Joc95, JC97, JCCCB94, Kin95a, KKK^{+}90, KBS98, KLAH^{+}94, KKT96b, LD93, Lav97, LM98b, LCE97, MG91a, MWT96, MM95b, OKL^{+}96, PM90b, Pur91, RPT^{+}96, RO97, RNH99, SDFJv91, ST90, SL97b, Sto97, SB97c, SFF92, TcBW96, VPM96, VS96, WB94, WOM^{+}97, WHK97, YFI91]. Pelagic-benthic [NNB93]. pelagicus [SW93]. pellet [BBG90, RCTH99, URHR98]. pellets [BMTF92, BH92d, BD94, DA93, FD98, RR99, UMD92, Vos91]. pellucidum [Bus97]. penaeid [CM97, Van92b, VS92]. Penaeus [CC92b, CNCS93, CCC94, CC97a, CLD99, CM97, HVL96, LCTT92, LWK^{+}95, MG91b, MF95, O’B94a, O’B94b, RMG^{+}98, Sha98, SCMM95, SlRR94, VHH^{+}96, VHHS96, VHH^{+}98, WH93]. penetrability [MJ92a]. Penetration [DB96b, GPB96, KLJ94, MP95, OKK95]. penguin [JCCCB94, ROANM98]. penguins [BGG^{+}97, CWB93b, CLJOC98, GKB^{+}97, HTCW96, HHM97, HHE92, JBB98, KR97b, LGC^{+}93, TMCE98, WR97, WPB^{+}93, WCBL94, WSP^{+}95, ZGM^{+}96]. penicillus [RI90]. Peninsula [CNGG96, MBM92, RGS95, Gon92, AMB^{+}96, Øre90, Øre95, SZA94]. Penplasmic [MA93]. pens [FW97]. pentachlorobiphenyl [WKH^{+}98]. pentachlorophenol [KJLH96]. peptide [FFH99]. peptides [Wöh96]. Peracarid [Bra95]. Peracarida [RWJ99b]. peracaridan [CSM99]. percent [DGCT93]. perception [KSV99, KV99, Mar99a]. perfoliatus [CK90, NK93]. Performance [KSE96, BBFD98, Ekl98, HS99a, PCT^{+}97, RWJ^{+}95, TR91, Tho92, VLKC94]. performed [LBV^{+}99]. Periclimenes [FGH95]. period [ASRC99, BC90, BG93, EFSS95, GQ96, Hav93, HLP^{+}99a, KFS91, LK99b, PB92, SJJMH90, SE93, TUN99, TP95, WSP^{+}95, YB90]. Periodic [HGS95, ND98]. Periodicity [Sam97, BCH95, ELMR98, Joh92]. periods [BL94b, MFD98, RN93, RDA97, Weh91]. periphyton [CPK97, Som96]. periwinkle [DB99, DBV98, WS92a]. Perkinsus [BA96]. permanent [GGB94, GLD^{+}91]. permanently [KTGC99, SB98]. permeable [HG99, HG92a, ZHF96]. Perna [HSTY98, HJH^{+}99, HTLK99, LB95]. peroxide [AOOT94, BAOT96, WPG99]. perplexa [BDdWC96, CM90b]. persistence [Aro98, BG99, Ski94, Tsu90]. Persistent [HTLK99]. perspective [LBK96, She94]. perspectives [MJW96]. Peru [Urb94]. Perumytilus [AC96c]. Peruvian [ODS^{+}96]. pesticide [PTW^{+}91, RWJ99b]. Peters [PdG96]. petrels [SM99a]. Petrolisthes [JA91]. Pfiesteria [BMG99]. pH [CD94, CH97b, Hin92, KCD^{+}95, LGdC^{+}90, MA99b, NBM94]. Phaeocystis [VMS^{+}96, ENSR90, FSdM^{+}92, HSMS99, HvB91, HRKR93, LGdC^{+}90, MOM98, PW94, PMS94, Rie93, RR97a, SCDH98, Sv93b, SDG95, SD96, WVMR90, vBV90a, vBV90b, van92a]. Phaeodactylum [JR92, LLB^{+}95, RGS90, RH90b, SDH95, SHT91]. phaeodarian [Gow93]. Phaeophyceae [Ang91b, AD93, EP97, Mül91, PB96a, SF95]. Phaeophyta [BWB98, HRFC90, KW91, Ken94a, MC93, WMG^{+}94, tor93]. phaeopigment [TLB93]. phaeopigments [BT93]. phaeoplast [HN90a, HN90b]. phaga [FPB^{+}94]. phage [Moe96a, Moe96b, Moe97a, Moe97c]. phage-host [Moe96a, Moe96b, Moe97a]. Phagotrophic [KFPR94, FC94, JG90, SBP93]. Phagotrophy [BBG90]. Phalacrocorax [BRLM90, GSC95, WFHM99]. phanerozoic [Mar99a]. Pharmacokinetics [RGT98]. pharmacological [BQKC97]. phase [BCG97, BKS92, IW91, MC91b, NK92b]. phases [BM94b, JBWH97]. Phasing [BBL^{+}95]. phate [Riv90]. phenanthrene [LF97]. phenol [Ekl98]. Phenology [HRD92, MCW^{+}91]. phenols [BRL97, MWH91]. phenomenon [FGG97]. Phenotypic [ATH95, FB97, Tru97, JKT90b, KJT90, TAJK90]. phenylindole [HP98]. pheromone [Ack97]. pheromones [HBBM96, ZHBH90]. Philippine [ADF96, DTA^{+}97, VAD^{+}95, WUA^{+}99]. Philippines [ADF96, BBFD98, RA96]. Phillip [BAM92]. Philobryidae [BH92e]. phlorotannin [HDD98]. Phlorotannins [LQ97, DCM90, PCLÅ97, PKY96]. Phoca [LADS98]. phocid [LK99b]. Phocoena [WT99a]. Phoenicopterus [GB97a, GB97b]. pholis [BKH99]. phosphatase [GGKJ^{+}98, LVP98, MA93, vBV90a, vBV90b]. Phosphate [Mol97, Vid94, MPCD90, vBV95]. Phosphorus [DON^{+}95, HH91a, SDC90, TSB93, ASP91, BK94, FFF^{+}94, GGKJ^{+}98, JG90, KEvN95, LC90, LW91, Mol97, NK93, Pak92, PMP^{+}96, PT95, PPR94, RDPB99, Tur93]. Phosphorus-limited [SDC90]. phosphorus-related [Mol97]. phosphorus-starved [GGKJ^{+}98]. photo [DRGV91, FM94]. photo-protection [DRGV91]. photo-response [FM94]. Photoacclimation [VSPLH98, BAM^{+}99, CCV^{+}96, RKS97]. Photoadaptation [LSL98, Riv90, VBF94, BHT98, BBL93, JSGS94, TBLD90]. photoalteration [NKI^{+}96]. Photoautotrophic [CJ97]. photodamage [JPA^{+}96]. Photodegradation [NKI^{+}96, TL95]. photographic [PS97]. photography [BGT91]. Photoinduction [HS98a, RR97a]. Photoinhibition [GJ96b, GJ96c, HGJ99, HHN90, FSA^{+}97, HHN93, HPSG94, Les96]. Photoinhibiton [Han92]. photon [IPT94]. photoperiod [ORD98]. Photoperiodic [EST95]. Photophysiology [RKS97, VM98a]. photopigments [OKK95]. photoprotection [HGJ99]. photosynthate [Put90]. Photosynthesis [BSMS95, BK96a, DL93b, FZ91, IG96, LS91, PGAJ99, SPDM91, SE93, AKL^{+}99, BFF96, BGR^{+}96, DL94b, DT94b, FK95a, FPGJ99, Han92, HHN93, HMWN97, HVFHH92, HLF^{+}91, HHN90, JHG99, KL94, KCD^{+}95, Les96, LL96, MG96a, MFD98, MC97b, MNK^{+}97, PLC^{+}91, SK98b, Sut92, TTR99, TFFA99]. Photosynthesis-irradiance [DL93b, LS91, MFD98]. Photosynthetic [BUWB99, CB98, FSdM^{+}92, GWKW97, HWLC98, MH99, Ral98, VDBF96, WH98, ARS93, BKP98, BVW^{+}98, BAK97, BZSG96, CR95, COG97, DFS94, EDSJ95, ES90, FL92, GPB96, GPF95, GBGH97, HD97, HPSG94, IPT94, IP94, JSV92, KBP98, KPS97, LM91b, LGR95, MC95, MKV95, MSND^{+}99, PMP^{+}96, RGLS99, RRWG96, RWJ^{+}95, SPS^{+}91, SPBS93, SBP92]. photosynthetically [PSB^{+}97]. phototactic [SHFP94]. phototransducer [VAR94]. Phototrophic [DGT^{+}93, Joc95, CP91, EJ96, Hei95, KG94]. phototrophs [OKK95]. Phragmatopoma [ZMD95]. Phragmites [Lee90]. phthalate [HFW99]. phtorotannins [Ste94]. Phuket [BLST90, LB96a, ABB^{+}97, KKH^{+}94]. phyceae [DRGV91]. phycobiliprotein [KFF92]. phycoerythrin [VMHH90, WPY98]. phycoerythrin-543 [VMHH90]. phycoerythrin-containing [WPY98]. Phyllospadix [BWRH99, RGLD^{+}98, TW97]. Phylogenetic [DAE^{+}99, BSP^{+}99]. phylogeny [SPB^{+}99b]. phylogeographic [DCBC91]. phyloplankton [BFJ^{+}93]. Phylum [Mil92]. Physeter [SPB^{+}99a]. Physical [BM90a, BM90b, HS96, JKL90, McL96, SJML99, Aro91, BM99a, BH97a, BM94d, Bre96, DM91, FB98, GS91a, HKK^{+}95, HCI93, JBWH97, KW98, KSG^{+}92, LGMD98, MC91a, MY96, MPB97, MG94, MP96c, NB96a, NFM98, OD90, OK98, Röp93, THR^{+}98, Rie91a]. physically [AvS97, LP95]. Physico [DvdM92, APT99, OSZ95]. Physico-chemical [DvdM92, APT99]. physico-chemistry [OSZ95]. physics [HRCW98]. Physiological [BH97b, CT98, DKVD95, EAT95, GT92, JGT96, McG92, NT96b, RK95a, WW91, AAD91, BMHH94, BKL99, Bro93, BLST90, CN92a, CH97a, CR96, ENSR90, FJR^{+}94, HG99, HMB95, LFRB97, LBL^{+}99, MB95b, ML92, NP90, RG97, SP99c]. Physiology [Vis91, BPR90, HMHDA94, MGC95b, MLH96, MH96b, NW97b, PCG96, SAA99, TVG96, TAJK90]. phytal [Hul99, WB94]. Phyto [BWHK99, BVP91, LKK92]. phyto- [LKK92]. Phyto-convection [BWHK99]. phyto-plankton [BVP91]. phytobenthic [KV90]. phytodetritus [AO91, LBM^{+}99, ÓMvdB99, Por94]. phytohormones [MH93b]. Phytopigment [DFB^{+}99]. phytoplankter [COD^{+}93, CS91, GHTH94]. Phytoplankton [Bra93, Cha96, DCBG99, HTS92, JP93a, KR98b, LFTH92, MMM^{+}95, MJV92, MEC92, MD94b, MP95, NPB93, OH95, RDT^{+}94, RPL96, SH97c, THS94, TBLD90, Ved98, WOM^{+}97, WT99b, ABM96, AE91, AW98, ACC90, AFZ96, Arr94, ASP^{+}96, BE97, BCG97, BNvB90, BH92a, BCHL93, BD91, BMD^{+}90, BKK^{+}93, BP96a, BAK97, BCS93, BKL99, BG91, BRN^{+}95, BGvDR96, BRW95, BRZ99, BGCM95, BD94, CCG98, CSG93, CGTB95, CC90, CC91b, CD94, CCCH94, CPH91, CH97b, DM96, DLFL96, DLT^{+}97, DFR95, DGO92, Dor90, EA92, EJ97, EST95, EPD^{+}99, ES90, EPN^{+}99, EP92, FML^{+}95, FSdM^{+}92, FD92, FS91, FPAH92, Fly98, FCVS92, FDZ93, FM94, FVDL95, FF92b, FL92, GV96, GJ97a, GJ97b, GAK99, GMK97, GW93, GGKJ^{+}98, GKMT99, HS98a, HK97, Har94, HP97, HKY^{+}97, HH94a]. phytoplankton [HH96, HRÆ94, HVFHH92, HM94a, HS91a, HJA99, Hol96, HHW99, JLS90, JM99, JSGS94, Joh96, JMP90, Käh98, KLN^{+}90, KKL91, Kam94b, KL95, KM99a, KNM93, KJN97, KY97, KP95, KGC92, KFS91, LVL^{+}99, LMT96, LM95, LGH^{+}93, LMF93, Les96, LZYW93, Lig90, LL92, LHK^{+}93, LCC97, LL97, Lip95, LDKR93, LM97, LJS90, LKC^{+}99, LKM^{+}99, LSH96, MMHW96, MRL96, MH99, MCA90, MS96a, MS96b, MHvB97, MP96c, MSD92, MGAE99, MAdG^{+}96, NK92b, PLP^{+}93, PML93, PM90a, PB96b, PS94, PDBA90, Per91, PPMF97, PLG^{+}96, PFP93, PCK^{+}96, PCC^{+}97, PM94a, PR93a, PSPD96, PPCR97, RRWG96, RC91a, Rie93, RFN98, RN98, Riv90, RZWY96, RCGSN96, RPMB91, Rud97, SPS^{+}96, Sca96, SKvB98, SCA92, SGK^{+}97, Sv93b, SDG95, Sut92, SE93, TMKL97, TNKMK99, TGG97, TGG^{+}95, TS90a]. phytoplankton [TK96, VLKC94, VDM90, Ven98, WGD97, WHIC97, WPTB96, WJM^{+}91, WTM^{+}96, YGH^{+}97, ZDJ^{+}95, vLTW^{+}98, vdBTvDR96, vdL98]. phytoplankton-bacteria [HHW99]. phytoplankton-bacteria-zooplankton [PML93]. phytosociological [FIL93]. piano [KH92b]. Picking [RD97]. pico [DGT^{+}93, FPR94]. pico- [DGT^{+}93]. pico-and [FPR94]. Picoalgae [SJÆ91]. picocyanobacteria [NSHF93]. Picoplankton [MDPP91, CB98, IP94, Joc95, KLAH^{+}94, LHL97]. Picoplanktonic [Kuo91, VK90]. Pigment [BP94, JSV92, AdN91, BE97, BMP^{+}95, BPC98, BZSG96, CCV^{+}96, GB95, GJ97a, GJ97b, GM98a, HS91a, JNJP94, LPF94, LED92, LM97, LH99, SMB98, TGG^{+}95, VK90, VMS^{+}96, WTO96, WTM^{+}96]. pigment-dependent [JNJP94]. pigmentation [BLD94, FSA^{+}97, LC90, RCA^{+}95]. pigmented [HR96b]. pigments [ABM96, ABB^{+}97, BRL97, BDA^{+}99, BDdWC96, BLR94, DRGV91, FFF^{+}94, GHTH94, HH94a, JWL96, KL92, LM91b, MGDT90, MEC92, MD94b, MHvB97, MKV95, MGC95b, RCGSN96, SH97c, VLKC94, VK90]. pilchard [FTS94, van94a, vdL95]. pilchards [EF97]. pilchardus [CCV^{+}98, GMT99]. pilema [LFBH94]. pilot [AS94a, GRC^{+}97, Sha95b]. Pinctada [HSTY98, YKL98a, YKL98b, YKL99]. Pine [SB96]. Pinguipedidae [San93a]. pink [STNI97]. pinnipeds [And98]. Pinnixa [TTW97]. pioneer [GMR^{+}94]. piperitus [GAV93]. pipes [QRSC99]. Pisces [AMFB99, BK92, FMRC93, Gil97, Sam97, San93a, SMF91, SCCS98, Uib95, WM92, WHK97, YD90]. Piscivory [JC94, JC95]. pistillata [FPGJ99, GYL93, KYN94]. pit [HBR^{+}91]. pit-digging [HBR^{+}91]. PJ [Ano99d]. Placida [Tro92]. Placopecten [BSH^{+}94, BSJPW93, BSJ97, CHS^{+}95, CHC94, CH99, EGMC94, MW94, MGP^{+}96, PRC^{+}92, PGM^{+}98, SH93, TRS91, TS90b, TS90c]. plaice [BK90, Bur94, GB90, GPB^{+}98, HW91, Hov92, MB91, NG99, Rv92, RGT98, vPB90, vW93, vdVW99]. plain [PS94, Vet96a]. Plakina [MSCTBE96]. plana [BM91a, RBG94a, RBG94b]. Planar [GRGK96]. planci [CED91, HK96, HKR96, HR96a, SP93]. planktivores [LBK96]. Planktivorous [KB94, KS95c, CH93, HRCW98, NKGK92, RO99]. Planktivory [HM97d, RHJ92, RHrsHG99]. Plankton [BM99b, BT93, JSJMH90, LED92, LCF99, NvA91, NH95b, PT94a, PT94b, TB93, AGGKA96, ASP91, AR93, BN91b, BBL^{+}95, BVP91, BG93, CGTB95, CLM^{+}91, CV98, CB96, FGS98, FIL93, FPP^{+}97, GdGMD95, GDW90, HP90, HCC^{+}91, HCS93, JL96, JS97, KHW^{+}99, Kru93, KK95, LBJ^{+}98, LSGW98, LW94, LCK^{+}96, LTM^{+}98, MRF^{+}99, MLC92, Mau92, Mol97, NTK98, NLR^{+}93, NS96, NHL99, PSK98, PPW96, PWM95, PLS^{+}95, RBD^{+}97, RAlBW99, RGBC92, SK94a, SDFJv91, SK95b, SCDH98, SCKS90, Sun96, Tis98a, VM98b, ZH97, Zho98]. plankton-benthos [LTM^{+}98]. Plankton-pigment [LED92]. Planktonic [AMB^{+}96, PDF97, SD93a, Aks96, BSB91, BG99, BH95b, Bus98, CCC^{+}90b, CR95, Cow91, CP91, CCR98, DBV98, DG92b, FKR90, Fra95, GF95, GBV96, Han91a, HTG96, HS97b, HM97b, HN91, HB94b, JPR92, JPC97, JT90, JG90, JG94, KL91a, KSG^{+}92, KS95b, KSV96, KS95d, KKT96a, KL90, KHT93, LBL91, LIPF95, LMK93, LF94, LVP98, LHK^{+}93, MLC97, Mar91, MHD97, MBL94, Mon96, MZCMOCJP91, MJ92b, NKB90, NJLD98, OF94, PL99a, Pen90b, PWP^{+}97, PR93a, RNP^{+}98, RW96, Sim91, SGP94, TI96, TNN93, UHL^{+}97, UT90, Uye96, VB90, VKF^{+}98, Wai90, WQQ99, Yam96]. planktotrophic [BN91a, Han99]. plant [AWHH94, BBFD98, BRL97, FZ91, HM91, HHSvS96, NK93, PPP^{+}96, SM97c, dG92a, dNSR^{+}96, FKC96]. plant-biomass [dG92a]. plant-part [HM91]. planted [FKC96, FMH96]. plants [BVS^{+}91, DL94b, JRMN^{+}97, LBK90]. planula [Sak97]. planulae [BDZB98]. planulation [Joh92]. plasma [Kal91b]. Plastic [vRBVB95, Kru93]. Plasticity [Mor96, FD99, FB97, WHW93]. plate [EGKE92]. plate-reader [EGKE92]. platelet [ARS93, ADG^{+}95, DAS92, GGD99]. platessa [BK90, GB90, GPB^{+}98, HW91, Hov92, NG99, Rv92, vPB90, vW93, vdVW99]. platessoides [RGT98, SFPC97]. platform [PDD^{+}99]. platforms [AC96c, MHC92, RS99a, RSN90]. Platichthys [LFG98, RH90b, vBDW91]. Platygyra [Mil94b]. Platynereis [ZHBH90]. platypterus [Spe95]. plebeius [GI99]. plebejus [MG91b]. Plectropomus [Joh99, Sam97, Zel97, Zel98]. Plena [dRAV92]. Pleuragramma [WHK97]. Pleurochrysis [IG96]. Pleuroncodes [Roa93, RGE^{+}95, RT98]. Pleuronecles [Rv92, WB92b]. Pleuronectes [BP99, BK90, GB90, GPB^{+}98, HKM92, HW91, Hov92, NG99, RBN99, WA95, vPB90, vW93, vdVW99, Kea90]. Plexaurella [RH90c]. plexaurid [LKC98]. plum [RD97]. plumbeus [CR98]. plume [BBF94, CLB92, CG93, Fuh90, FS98, GBA^{+}96, GFPD92, GF91, HCC^{+}91, JMP90, MF90a, MF90b, ND96, OSGB96, PLP^{+}93, PBA^{+}95, SB96]. plumes [AMMG91, Gra96b]. plumulosus [SDC99]. Plymouth [PHHH96]. pm [RLT^{+}94]. POC [MGAE99]. Pocillopora [AHS97, LL96, MNK^{+}97]. pockmark [DAB^{+}91]. pods [PlMM94]. poecilogonous [GC95a]. poecilogony [MRP^{+}99]. Poecilosclerida [UTB^{+}95]. poecilostomatoid [BS92, HLBW93]. Poecilostomatoida [Met98]. Pogonophora [Sch91, SS92]. pogonophore [DBN^{+}94]. Poigar [AALT99]. Point [LC93, CH97b, DGCT93, RSAD98]. point-source [RSAD98]. poisoning [KB91, SPZ^{+}93]. Polar [ER90, GCKW95, PWR96, PPM94, PPMF97, RPT^{+}96, WSCT96, DAS92, GSD96, HMWN97, SCJ99a]. polaris [RHH94, RMH96, RPT95]. Pollachius [FQV98, MHC92, RK95b, RK95c]. pollack [FSG^{+}98]. pollination [VWK96]. pollock [BM94a, BCN^{+}95, BBK91, CBI91, DO92, FQV98, HS92, HHP95, HBP^{+}93, Liv93, OD90, RK95b, RK95c, RO98, YB90]. pollutant [BBR^{+}93, BBP^{+}96, Moo92b, SGB90]. pollutant-induced [Moo92b]. pollutants [CKP98]. polluted [GD96, MAAT98, RZLB96, RBG94a, SFP91, UKR^{+}91, WH96]. polluting [KEvN95]. Pollution [Cha92, AGW93, Bou90a, Bou90b, BVL92, FC90, GBC92, GAD^{+}92, GCWH90, HBD95, LTM90, McF92, OSC97, PMETM96, PGB91, Reg92, RSAD98, SH92a, SS93, UKR^{+}91, VWJ96, VBL^{+}92, WTA^{+}97]. pollution-induced [VWJ96]. polonium [CF93a, CF94]. polonium-210 [CF93a, CF94]. polycarbonate [BP94]. polycerus [Mar91]. Polychaeta [AOOT94, GVGG98, MGDT90, MRP^{+}99, QC94, San94b, TD92, TDW96, Tsu90, Ved98, Vis90, Zaj91a, Zaj91b, ZMD95]. Polychaete [CSSPV98, Tok95, BHQ99, BH97b, BQKC97, BAOT96, CJ93a, FFH96, GQ97a, Han99, HFW99, HQ99, Hsi95, Hsi97, LQ97, LBDD97, LGFS99, LSP99, MFF97, Mar91, MPS96, MCNB98, MT94, ND98, ORD98, PL99b, QQ97, QQ98, RI90, Rii91, RVBL92, SHC96, SFB92, VR93b, VAR94, WSF99]. polychaetes [FSCE92, HBBM96, JDLL95, LW95, MK93a, MK93b, NEIR95, SdC97, TPHC91, VTGK99]. polychelid [CA92]. polychlorinated [GS99b, WKH^{+}98]. polychromatic [GJ96c]. polycyclic [MCSV95, ONR98]. Polydora [Zaj91a, Zaj91b]. polylepis [GR94, JSV92, KLMR91, NKB90]. polymerase [KZP93]. polymorphic [CCGM93, HBBT98, Van94b, VK94, VB94]. polymorphism [YKK^{+}98]. Polynesia [NJLD98, Adj97, BLJ99, CRCS96, CCR91, Cha96, CDG97, CCRB98, DG93, GPDF93, HGS95, LGH95a, MG96c, PPCL^{+}98, Pla93]. Polynesian [CB98]. polynoid [VTGK99]. Polynya [ADBH99, PLG^{+}96, SGK^{+}97, VD94, Bra95, HAR95]. polyodon [CW93, WC92, WS92b]. polypeptides [WNJ91]. polyphemus [BLJ92]. polyphenolic [ATH95]. polyphosphates [LLB^{+}95]. polyps [LGP95]. polysaccharide [AJST98]. polysaccharides [BMS^{+}97]. Polysulphone [Dun99]. Pomacentridae [AMFB99, Ker96, MF97, MMD92, NKGK92, RHH97, WB97]. Pomacentrus [Ker96]. Pomatomus [HC93, MC91c]. Pomatoschistus [MF91]. Pomeranian [MHvB97, SKvB98]. pond [SE93]. ponds [DLHB99, SCD^{+}90]. Pontes [KML91, KL91c, KPWL93, RS91]. Pontoporeia [Joh97]. pool [CSG93, MH90, SDB92]. pools [ASP91, Ang95a, Ang95b, BCC95, BCC96, BFJ^{+}93, KCI98, Lai92, LGV^{+}99a, LJS90, SHA^{+}95a]. ’poor [SJ97, DHM92, MAdG^{+}96, STS99]. pop [Con94]. populated [SJ95]. Population [AWHH94, Asa91, CV96, Cla96, CD98, Gen92, GHL^{+}97, HBBT98, Jon93, KYN94, LA98, LFS90, LSB95, MCRGC90, MGC91, MGC95a, MZFHC96, NSH95, NH95a, NB96b, O’B94b, OFR94, Pez91, PM98, RVv92, San94b, SCL95, SCAT91, SLHC98, TA90, Tsu90, UC94, WMG^{+}94, WS92b, YKK^{+}98, Zaj91a, Zaj91b, ZH97, vBDW91, AAWD97, Ang91a, Ang91b, AD92, AD93, APT99, BM98, BCA^{+}99, CED91, CGNH98, CPT90, CS97, CRH^{+}99, DCK99, Dee90, Dur94, ER90, Fab95, FW92, FS90, GMO95, Góm95, GCNSNd94, HD91, HB94a, IT96, JRB96, Jeo94b, KL90, Laf93, LCG97, LGFS99, LGMD98, MDC^{+}96, MW99, MA99a, McC94a, MDG98, MS93c, Øre95, PCM^{+}95, PB93, PS92, PFH98, PTW^{+}91, RSM^{+}96, RWCM99, RSN90, RPL96, RT98, RMH96, RCC^{+}99, RNHH97, RAGH94, SBBG94, She98, SLR94]. population [SP99a, SF99, SK95c, Tan98a, Tan98b, TGLP95, UCV98, VDM90, Vet96a, WSCT96, LPF94]. populations [ASN93, ASL99, AW97, Ano99d, Atk96, BLJ^{+}96, BHG94, BJB95, BA97, BVP91, BD99, Bre95, BCBR97, BA96, CK90, CBRD92, CC90, CC91b, CJ93a, CK98, CNJ96, CWHA99, DdA95, DG96, EP97, ENSR90, FML^{+}95, FBMH94, FS91, FB94, FJR^{+}94, Fra93, GS91a, GK96a, Gil97, GGPGG95, HWLC98, HBD95, Hoe91, HvdM97, Hun93, JT97, JTPP92, JKT90b, JFS90, JFS92, KWD93, KJT90, Ken92, KSH97, Lip95, LJSK96, Mal99, MGP^{+}96, MS91, MMGIM90, MDTB94, MR97c, MD94c, MT93, Mil97, MAAT98, MC99, Mum99, NCGG96, NC95, PG91, PFP93, PBHMI98, Pur92, RCTH99, RCA^{+}95, Sab90, SZ96, Sch93, Sie93, SFP91, Ski94, TA93, TAJK90, TD94, Uib95, Väi90, VD99, VDJA98, WS92a, WO94, WC90, WK92b]. porcelain [AP99a, JA91]. Porcellanidae [JA91]. Porcupine [HK99]. pore [LEHvR96, VM95]. Porewater [HG94, OL99, Hüt90]. Porifera [GS94a, MSCTBE96]. Porites [LGH95a, MNK^{+}97, NB98, SR90a, vW98]. Porphyra [FSA^{+}97, HCG^{+}93]. porpoise [BCA^{+}99]. porpoises [WT99a]. Port [BAM92]. portable [Sty97]. Portugal [BPC98, CSM99, QCM94, AA92, EMR^{+}97, FA99, Wol92a]. Portuguese [CSB97]. Portunid [SL90, FSGG96]. Portunidae [HMHDA94]. Portunus [LH91, SW93]. poseidoni [DBN^{+}94, Sch91]. Posidoma [VWS93]. Posidonia [ADR95, ARDL97, CDM^{+}96, DSS^{+}98, DBdCS99, EAD92, FK95b, MDC^{+}96, MR97b, PRPM^{+}94, PAM96, PM98, RGBJM95, TB99, Way95]. position [Bus98, IVM^{+}98, MJSDF91]. position-holding [Bus98]. positions [GU98]. Positive [Leo99, ASN95, Sch90]. positively [MV96]. Possessing [STS99]. Possibility [KP95]. Possible [KOSS90, AMFB99, BRLM90, BFS^{+}98, GS94a, GS94b, Hal98, JDLL95, KR97a, KS99, LB91, LGH95b, McG97a, ONM93, RQB98b, STNI97, SL95, NCGG96]. Post [Fre97, Ker94, MM97a, ZMD95, lYALM93, BLH99, CMRF94, CH95, CTC93, DY92, Egg95, FAED97, GM99, GQ96, HLP^{+}99a, HS97c, JBWH97, MU98, MC91b, MD93, PWS98, Ris97, Sha98, Sto96, WW96]. post-capture [BLH99]. post-larvae [CMRF94]. post-larval [CTC93, GM99, Sha98]. post-metamorphic [MC91b, MD93]. post-nesting [HLP^{+}99a]. Post-settlement [Fre97, Ker94, MM97a, ZMD95, lYALM93, CH95, DY92, Egg95, FAED97, GQ96, HS97c, JBWH97, MU98, PWS98, Ris97, Sto96, WW96]. postlarvae [BBS93, CM98b, EAEP98, ML92, MPvM98, PODL99, TDW96, WRBF97, WFH99]. Postlarval [SK98c, BAH94, DMM99, ELMR98, JC92, MZFHC96, VHHS96, VHH^{+}98, WX97]. postsettlement [BB93]. Potamocorbula [BL95, CTSN90, DL91b, DL94a, LWL98, NTS90]. Potamogeton [CK90, NK93]. Potential [AO91, Far98, Lee97, OWM92, QK93, SPH94, YTF97, Aro98, BNvB90, BM99a, CH94, DHH98, DL91a, DB96b, FH97b, GHD92, Hsi97, JG94, KCS94, Kel98, KZP93, KBA^{+}97, LGH^{+}93, LJSK96, MG96a, MB93, MBL^{+}98, NW97c, Ole95, PGG96, SAG92, Sch91, SSM91, SO96b, TF98, TMCE98, VB90, Wah96, WH91, Zaj91b]. potentially [HSB97, LRV94, Pur91, Str99, THS94]. potentials [BL96, CH93, MSK95]. Potomac [Che93]. pouchetii [VMS^{+}96, ENSR90, PW94, WVMR90]. poutassou [HK99]. practices [SKM93]. practising [MTB^{+}99]. Prainha [CV96]. Praunus [VR98]. prawn [Gra91b, Gra91c, HSN93, RT95]. prawns [CF94, CLD99, CM97, HHK^{+}98, HVL96, LWK^{+}95, MG91b, NH98, O’B94a, O’B94b, SCMM95, TRPS94, Van92b, VS92, VHH^{+}96, VHHS96, VHH^{+}98, WH93]. Pre [BCMC99, JBWH97, JWKH98, JBK99, KB91, SB94]. pre- [JBWH97]. pre-adaptation [KB91]. Pre-emption [BCMC99]. ’Pre-feeding’ [SB94]. pre-settlement [JWKH98, JBK99]. precede [LMT99]. precipitation [WJM^{+}99]. Precision [EP92, SFG^{+}99, TJ99]. Predation [AC99, AJL93, Aro91, BL94b, CED91, CH95, EB92, EE98, EG97, ENSR90, FWF^{+}99, HHK^{+}98, HPM92, Hil97, Jeo94b, Jeo94a, Ken94c, KGO94, Mic96, MPvM98, NC93, OW95, Pen90b, PE99, PG90, Pur91, PND^{+}94, PC95, SHC96, SCCS98, UGSBA94, VKF^{+}98, WA95, YOH98, AJS95, Amb93, ACR99, Aro98, Atk96, BK98, BS94, BSD^{+}94, BRC99, BS96b, CH93, CA98, DL93a, ELH92, Ejd98, EG95, FPB^{+}94, FIA93, FC94, Ger90, GA97a, GB97a, GB97b, GB93, GF98, GEW97, HRKR93, HP96, HB98, IVM^{+}98, JAOS90, JS97, KB94, Lew92, Liv93, McG97b, Min93, ND98, Ole95, Øre90, PLFSW97, PS95, Pur92, RD91, RM97a, ROANM98, RH90c, SZ96, SH97a, Sew96, She98, Ski94, SK94b, SL90, SM99a, SR94, TLT93, Tro94, TB97, Val91, VS96]. predation [WH91, Web91a, Web91b, WC91]. Predator [Bro96a, Bro96b, ELG97, GK97, KV99, Pen90a, TT99, Aro92, BPS98, BBS93, BA94, BGH95, CS99, CWWH99, Con98, DBMK99, DL91a, Fer99, GWSG99, Han91a, HHE92, JLYK97, JC94, JC95, KB94, KLAH^{+}94, OS96, PM92a, San94a, SL97b, TK95b, TPHC94, VKF^{+}98, WWL97, WB91, Web91b]. predators [ASN95, AH96, BSH^{+}94, BAH94, BPS98, BLH99, Bur94, BKH99, CW99a, Con98, CRP99, DL91a, GPB^{+}98, KUK91, LSM92, LMM98, MF91, OK98, PLFSW97, PLF90, PBGE91, RTCH96, RWCM99, Reu98b, RPT^{+}96, RO98, SFA98, TPHC94, War90]. Predatory [SK95a, Bee97, CM98a, CP95a, PGG96, PCTF95, RMH96, ZFFH^{+}94]. predict [HvdMBK98, Hur92]. predictability [FSCE92]. predicted [MB98, THCD95]. Predicting [GTJ92, JRB96, MPB97, Mee99, BJTB96, HTCP96]. Prediction [FW97, TD94]. predictions [AH94, JBK99]. predictor [ICG96, MMD92]. predictors [HRFC90]. predominantly [WM99]. Prefecture [YNYK94]. prefer [Jua92]. preference [ARF92, CGCLH99, JC94, MA99a, SK98c]. preferences [DC91, EK95, Har99, KW97b, PNS90, Tam95, VEW99]. preferred [PM94a]. Preliminary [KDAS92, LC91, Pro92, SRG^{+}99, BLD94, CR98, DBvA91, DAS92, KH92b, LCK^{+}96, RSD^{+}91, SC95b, Tis98a]. preparation [ER95b]. Presence [MH93b, BCH95, GB93, Jeo94b, JM99, SM97a]. present [GI99]. Preservation [SGP94, BP97a, LMF93]. preservative [CK95]. preserved [DFB^{+}99, GBH96, HKM92]. presettlement [Gra96b]. prespawning [MR97a]. pressure [AAW98, BG94, GFH96, Gri94, HvB91, Min93, TGB99, YTF97]. pretreatment [GCLS94]. prevailing [NKV94]. Prevalence [RB90, EP97, HHG99]. Prevalences [Wat98]. prevent [STS99]. prevented [Les96]. prevention [AJ95]. previous [CE93, HDB95, QQ99]. previously [THR^{+}98]. Prey [EK95, GA97a, JH97, KSV96, MJ97a, ML91c, MM95b, PPM96, ARF92, Aro92, Atk96, BCN^{+}95, BS94, Bro96a, Bro96b, Bur94, Bus96, CW93, CWWH99, CGCLH99, CGZR94, Con98, CLW99, DO92, DL91a, DFPG91, ELG97, ENSR90, Fer99, FLDT92, GD90, GSS93b, GBP^{+}97, GB93, GT96, Han91a, HN97a, HH90, HHE92, HRCW98, HB94b, IFL97, JWKH98, Jeo94b, JLYK97, Joh99, JS97, JT90, Jua92, JC94, JC95, KV99, KS99, KO92, LMM98, LK94, LBK96, ML91a, MG91a, MBR94, Min93, MCFS96, Mun92, OS96, PHT96, Pen90a, PS95, PM94a, RHH94, RO99, SM98a, San94a, SMF91, SL90, SM99b, TT99, TK95b, TPHC94, VKF^{+}98, VR98, WWL97, WE95, WPB^{+}93, ZGM^{+}96, ZB92]. Prey-size [ML91c]. preying [ML91b]. Priapulida [ABN98, Oes90, OV92]. Pribilof [DH96, HCH^{+}96]. Primary [BKL99, KFS91, MC96a, MG97, MMR99, MSD92, MRP90, PRPM^{+}94, AMB^{+}96, AB96b, BWHK99, BKP98, BCS93, BSP93, BVL98, CCRB98, CG99a, CDGF^{+}93, CPH^{+}96, DON^{+}95, EW91, GB91, Gib94, GKMT99, Har90, HAR95, HM97c, HHW99, IP94, Irw91, JKL90, JLGMP97, KLMR91, KD97, KR98a, KP95, KPS92, KPS97, LB95, Lee90, Lig90, LSL98, LFR^{+}97, MPRB93, MP94, MH97, MKCL94, MWB^{+}99, PWG^{+}99, PM91a, Ric91, RC91b, RNH99, SCD^{+}90, Sos96, TH94a, TFFA99, WOM^{+}97, WH98]. primary/secondary [LB95]. primitive [BL97]. primmorphs [MWB^{+}99]. Prince [IFL97, MRP98, PDBA90, Per91]. Principles [BVP99, AJ95]. prions [GBP^{+}97]. priority [TB95]. PRISMA [BRNH97, KS97, Sün97]. probabilities [COK99]. probability [BG97]. probably [SJMC93]. probe [BAK97, CTEF91, MS95]. probes [BMS^{+}97, LMK93, LME92, Moo92b, WK98]. probing [PCC^{+}97]. problem [PN99]. problems [Blo91, Yam96]. Procedural [BS94]. procedure [GV96, MS93d, PPR94, SR95]. procedures [BCAAC96, CLHF95, PPCR97]. Procellariiformes [SPB^{+}99b]. process [ATNW97, GSM^{+}97]. process-based [ATNW97]. Processes [LKC^{+}99, LKM^{+}99, BWHW95, Cal95b, CN92a, FB95, FF99, FAVV98, FMFH97, GMG^{+}92, GJ96a, Har97, HB97, HTCP96, JBWH97, KSG^{+}92, KY97, MJW96, MS94a, MB97b, Mil94a, MLTD96, MJ93, NFM98, PG91, RR99, RS99b, Sün97, TCRB99, vTSD97]. processing [BSJ97, BDB97, BVP99, DSW^{+}91, JKL90, LBM^{+}99, MA93]. Prochlorococcus [BAM^{+}99, CCV^{+}96, LCL95, MGC95b, PMP^{+}96]. prochlorophyte [VK90]. produced [AR99, Mar99b, Mül91, RSN^{+}99]. producer [BUCP^{+}92]. producers [KR98a]. producing [BLKA93, BH93, NKKO97, SNR^{+}96]. Production [BD94, CM90b, DFPG91, EM99, FBHH93, HB91b, KJMRC96, LKC98, LD96, OKL^{+}96, SLR94, Sv93b, SH99b, VKV95, VM98a, AD94, AMB^{+}96, ADG^{+}95, BWHK99, BKP98, BK91, BKL99, BH92e, BSP93, BG93, BK96b, BVL98, BBG90, CCG98, CC92a, CBTM93, CBG^{+}99, CA96a, CA99a, Cam91, CV96, CH97a, CDME97, CBA^{+}99, Cha96, CCRB98, CB98, CLB91, CLB92, CVD99, CBB90, CG99a, COG97, CDGF^{+}93, Col98, CPH^{+}96, DKW96, DKL94, DLFL96, DON^{+}95, DL93b, DMA^{+}98, Dut98, EJ97, FLDT92, GSM^{+}97, Gar91, GB91, GGDLE96, GKMT99, HV91, HHSvS96, HM94a, HAR95, HK93, HM97c, HK94, HvdMBK98, HNC92, IP94, Irw91, JKL90, JP93a, JFP95, JKT^{+}98, KLMR91, KD97, KNM93, KE95, Kir92, KHW^{+}99, KP95, KFS91, KK95, KK96, KPS92, KPS97, LAG96, LD94b, Lee90, LGH^{+}93, LSL98]. production [LFR^{+}97, LW91, LKC^{+}99, LB93, LMT99, MC96a, MFD98, MPRB93, MG97, MF97, MMD93, MAH94, MH97, MMR99, Mga92, MDG98, MKCL94, MC94b, MDSB92, MG99, Moe97c, Mol97, MSD92, MWH91, MH94b, MGAE99, MPS95, MAdG^{+}96, Nak92, NG99, OG95, PWG^{+}99, PRM^{+}96a, PRM^{+}96b, PHT96, PHL92, PS94, PRPM^{+}94, PDBA90, PLG^{+}96, PT95, PR93a, PH95, RCTH99, RSN90, Ric91, RC91a, RC91b, RZWY96, RAlBW99, RNH99, SCIA99, SAG92, San94b, Sca96, SKvB98, SFPCR93, SFSA99, SR93b, SD94, SRG^{+}99, SD93a, SHML98, ST90, SK95b, SCD^{+}90, Sos96, Spr93, SSA96, Sto96, TGBB97, TH94a, TFFA99, TS99, TD94, Tuo97, UCV98, UK98, US98, VH93, VPFL92, VWS92, Vet95, Vet96b, VDBF96, WH95, WSC^{+}99, WSMM99, WV95, WR92a, WF94, WOM^{+}97, WH98, YGH^{+}97, ZCPA94, ZR92, vDK90, vDvRK93, vR93a]. production [vdBTvDR96]. production-loss [LKC^{+}99]. productive [MLDM95]. Productivity [EY90, HC92b, TL91, AB96b, BCS93, CRP99, Dua91, FGR97, HFN91, HA90, Har90, Hei91, HHW99, JW96, JLGMP97, KM92, Kro91, Lee90, LK96, Lig90, MP94, MJV92, MP96c, MRP90, PLM92, PZ91, PM91a, SPPL92, Tan98a, Tan98b, Tay98, VWS93, VSQ90, VAD^{+}95, WC90]. products [CTEF91, HV91, Kam94b, dNDS98]. profile [GP91]. profiler [CSB94]. profiles [ADG^{+}95, DAS92, FRP^{+}99, Hüt90, JPA^{+}96, KTH98, KW92b, LEHvR96, LGR95, RS92, TGG^{+}95]. profiling [RK99]. program [MMHW96, WTM^{+}96]. programme [CdJS^{+}92]. programmes [SM97b]. programs [LAJ95]. Progress [RFNBHG94]. progression [BOC90, BSW93, GJ97a, GJ97b]. project [Sün97]. prokaryote [PMP^{+}96]. prokaryotes [SB98]. proline [FFH99]. proline-rich [FFH99]. prolonged [FK95b]. prominence [SNC^{+}98]. Promotion [AKM^{+}99]. propagule [Ken94a, WLB^{+}99]. propagules [KW91]. Proper [MHRV^{+}99, THKN98, KKL91]. properties [BU93, FP91a, GB97b, HS91a, KMKS96, MG96c, MGC95b, NP90, SHT91]. Prophage [CKP98, JP96]. propinquus [HKG93, SSHM91]. propionate [LD94b]. Proportion [AR93, SM94]. Prorocentrum [ACC90, JNJP94, Les96, Sci91, SA94, SLC^{+}97]. Prosobranchia [DBV98]. prosome [HSW99]. protected [HSMS99, MBF^{+}97]. protecting [PF95]. protection [AGP98, BAH94, CK99, CK98, DRGV91]. protective [CCC^{+}90b, GA97b, Raw96]. protein [BBB^{+}94, BZSG96, CNCS93, CCC94, FF92a, KWHM^{+}96, KBA^{+}97, Kre93, KHBL95, MHB94, MM96, MAAT98, MBL^{+}98, NK96, NH90, RBB^{+}91, SR98a, YFI91]. proteinaceous [LMT96]. proteins [DEP^{+}96, SKT97]. Protist [NC94, PF90]. Protista [MvdZH^{+}97, MSvdZH98, PCG96]. protistan [Gon92, GSS93a]. protists [DLHB99, HSCC96, RR99, SS99a, SBP92, SBP93, SLSH98, TMKL97, VPFL92]. Protobranchia [Nak94]. protocol [ER95a]. protoctists [KFN99]. Protoglossus [GK97]. Protoperidinium [Bus97, JL94, Jeo94b, Jeo94a, JLYK97, LJ96]. Protothaca [UC94]. prototype [KEvN95]. protozoa [AAW98, CGTB95, CLM^{+}91, FKR90, HRKR93, KLAH^{+}94, LHL97, NK90, SSM91, CES^{+}94]. Protozoan [KL90, SMB98, BA96, FVDL95, FWF^{+}99, SNJV92]. protozoans [LDV95]. protozooplankton [TMKL97]. proves [BV99]. provide [ACLS97, SR96]. provisioning [CRP99]. proximal [GD96]. Proximate [TDH^{+}94, Yam97b, LMM98]. proximity [Gla99, HF99]. Prydz [HC94]. Prymnesiophyceae [GR94, MSND^{+}99, SD96]. Prymnesiophyte [SC95b]. Pseudo [DRP^{+}97, Kea90, PRM^{+}96a, PRM^{+}96b]. Pseudo-nitzschia [DRP^{+}97, PRM^{+}96a, PRM^{+}96b]. Pseudo-pleuronectes [Kea90]. Pseudocalanus [KGS95, KBS98, Nor96]. pseudocun [KHT93]. pseudocun/isetus [KHT93]. pseudodelicatissima [MHBW90]. pseudofeces [BV99, BVP99]. pseudolamellibranch [BD96]. pseudolysogeny [Moe97b]. Pseudomonas [SHW^{+}91]. pseudonana [DRGV91]. Pseudonitzschia [BLKA93, BUCP^{+}92, LRV94]. Pseudophycis [Kal91b]. Pseudopleuronectes [BSH^{+}91a, BSH^{+}91b]. Pseudopterogorgia [SB96]. Pseudorchestoidea [CV96, GD99]. PSII [KBP98]. PSP [BMFR97, Tee99]. Pteriidae [YKL98a]. Pterogorgia [PF92a]. Pterygophora [HRFC90]. puber [FGG95]. publishing [Ano99c]. Puerto [McG94, YY91]. puffer [KJLH96, KKYH97, Yam97a, Yam97b]. puffin [RM96]. Puffins [HS92]. Puffinus [LMT99]. Puget [KCV94]. pugilator [O’C90, O’C93a]. pugio [GF98]. pugnax [O’C93a, OJ99]. pulcher [Cow91]. pulchra [dNSR^{+}96]. pullastra [APCL^{+}93]. pulse [BVW^{+}98, HPSG94]. pulsed [HHYD92, RGS90, Sci91]. pulsed-laser [HHYD92]. pulses [FRP^{+}99, SK98b, Sch99, VBN95]. pump [BAK97, CR96, DQT^{+}97, JLR90, LSGW98, PR93a, RI90, RVBL92, RBT96, RG97]. pump-modelling [RVBL92]. pumping [RVBL92, RTJ^{+}93, RM97b, VAR94]. pumps [TR95a]. punctata [JWKH98]. punctatus [OC95]. pungens [BLKA93]. purple [ESDK94, SDER96]. purpuratus [AC96c, ESDK94, KL91b, ME97, SDER96]. purpurea [BKWML93]. purpurescens [Ste95]. pusilla [CS91]. pycnocline [BN91b, JJNR95, KD99, SD93a]. pygmaea [JNJP94, NP90]. Pygospio [MRP^{+}99]. Pyrenomonas [LC90]. pyrifera [AHT^{+}90, GHL^{+}97, SDER96].

quadrat [LC93]. quadricolor [FGH95]. quadrimaculatus [MBR94]. quahog [BLC91]. Quality [MM93, AMS99, AdN91, BRC99, CLC^{+}98, CP95a, CSR^{+}92a, Hil91c, KBS98, MW94, McC96b, SJ97]. Quantification [AvS97, Kel98, GGC99, GGKJ^{+}98, HL98]. quantify [MM91a, MJSDF91, NW97a]. Quantifying [ASP91, BGR^{+}96, HK97, ML91a, Rud97, WG97, TLT93]. Quantitative [Blo91, BVL92, CT92, Cro99, DEP^{+}96, Han91b, LA97b, PM94a, Wes90a, Wes90b, dNSR^{+}96, BVV^{+}94, EAC^{+}97, GSS93a, KRM94, KB94, Kne91]. quantitatively [Roz92]. quantum [KBP98, SPBS93, Sos96]. Québec [SL97a, SH93]. queen [Dav94, SL94]. Queensland [CM97, O’B94b, TA90, WB92a]. question [VM98b]. questions [JRB97]. quinquecirrha [OPS96, Pur92, PC95, PWNW99]. Quirimba [Bar97b, Bar97c, Bar97a, Bar99].

R. [JNCS90]. Ra-228 [COS93]. rabbitfish [PNS90, TNP97]. radial [GFH96]. radians [TT90]. radiata [BWB98, SS93, Ste95]. radiate [Ste94]. radiation [AKL^{+}99, BCHL93, BK90, BSS92, BZSG96, BEG97, DB96b, Dun99, FSA^{+}97, Gar91, GPB96, GHTH94, HVFHH92, HMVHH95, JPA^{+}96, Les96, LL96, MSND^{+}99, NDG^{+}99, PCLÅ97, PGB96, Raw96, SGL^{+}97, SDC^{+}92, SL97b, VSSM98]. radiation-absorbing [SDC^{+}92]. radio [ÓMvdB99]. radio-labelled [ÓMvdB99]. radioactive [Car90]. radiocarbon [BAK97, Cam97, KJGC96]. Radiochemical [COS93]. radiography [LCBD94]. Radioisotope [WK93]. radiolarians [Gow93]. radionucleotide [MHT^{+}97]. radionuclide [KI97, LGB91]. radionuclides [WTF96]. radiotracer [FR91]. radiotracers [ND91]. raft [FGG95, LFRB97]. Rafting [DeV92, SM91]. rainbow [DD90]. Rainfall [PWG^{+}99, MPRB93]. rainwater [KDR97, WPG99]. raising [MC91b]. random [CCGM93, DGCT93, GS99a, Vis97]. random-point [DGCT93]. randomly [HBBT98]. range [CT96, JMC93, LB91, NW97b, Zel97]. ranges [TMCE98]. Ranina [Ken92]. RAPD [ASL99, Way95]. RAPDs [CCGM93]. Raphidophyceae [KYI^{+}98, NU91, YKK^{+}98]. Rapid [DRGV91, HDD98, NB96c, RSAD98, SM91, SRS92, BG93, JP93b, Ken94c, RTLG97]. rapid-freezing [Ken94c]. rapidly [NDRP93]. ras [UKR^{+}91]. Rate [GSM^{+}97, PDC97, AWH93, AN97, ACC90, BP99, BWHW95, BEG97, BRZ99, Cam96, CR95, CJ93b, GMK97, GGC99, GT98, Han97, HL90, HH96, ICG96, JD90, KM96, KJLH96, KRJL96, KBS98, KKT96b, LCC97, LJSK96, MC91a, MCL90, MSP^{+}97, Mat92, MUH97, MLH96, Nak92, NNR95, NBM94, Pin94, RCG99a, RCG99b, RTJ^{+}93, RG97, Sci91, SD94, SS91, Tak95, TW98, TT90, TDD94, TN90, US98, Vos91, WFTH97, WDD92, vRV96, vdL95]. Rates [KTGC99, SH92b, ADBH99, ACC90, AAW98, BYO95, BH92a, BVW^{+}98, BCL93, BC90, BBF94, BD94, CCV^{+}96, CA96a, CA97, CA99a, CPS91, CN92b, CL97, CC90, CC91b, CGA95, CA98, DW95, FML^{+}95, FD98, FBMH94, FPR94, FPGJ99, FG93, GHMC96, GCH93, Gil97, GW93, GCLS94, Gru95, Gru96, HS94a, HKA96, HVFHH92, HS97b, HdWS91, JPC97, JL94, JS93, JR92, KH92b, KLEH91, KS95d, LGV^{+}99b, LC97b, Lip95, LCL95, LKvRH96, LGR95, MLP90, ML91a, ML93, MYYT96, MM99, MHC92, MRMLP97, MEC92, MHvB97, MMBT97, MNK^{+}97, MJ97b, MA96, NC95, NH90, OS96, PCBA94, PDF97, PPMF97, PG94, PPB90, RATA^{+}99, RCG98, RSN90, SKvB98, SA94, SSM91, Sog92, SH98b, SPDM91, Str91, Sun96, SCF91, Tam95, TMC98, TGSN99, Tur93, VGM94, VTNA91, VWS92, VWS93]. rates [VM98b, Wah96, WSM91, WK93, WG97, YFI91, vBV95]. rather [WF94]. ratio [CD96, EF97, Fer99, FR91, FF92a, GMK97, Har90, JR92, Käh98, KJN97, Mac91, Mar99b, OH95, TPR98, TBLD90]. ration [FPPM96, MC97a, MC98, PPMF97, TP95]. rations [PPM96, dT94a]. ratios [BJTB96, CMBC98, CCV^{+}98, DS93a, FM95b, FMFH97, GV96, GM98a, HM96, HW91, OINN94, RCS^{+}96, SBV98, SDFJv91, Som94, SCC98, TGG97, TRMW95, WDB98]. rats [NC93]. ray [KPWL93, TPHC91]. razor [GCM98, GI99, IVM^{+}98]. Re [CCC^{+}90b, HS97b, LE90, WASSM97, BS92]. re-analysis [HS97b]. Re-evaluation [CCC^{+}90b, BS92]. re-examination [WASSM97]. re-invasion [LE90]. reaches [WA99a, Whi94]. reaction [BAOT96, KZP93]. reactions [HMB93, HMB95]. reactive [KGHTN99, Pak92]. reactivity [Kep94]. reader [EGKE92]. realisation [BRNH97]. reality [LBK96]. really [Gra94]. reared [BSH^{+}91a, FCW98, RT95, RBB^{+}91, RFGF92, WMH96, WE95]. rearing [GGF99, JPR92, NvA91]. reassessment [Väi90]. Rebuttal [JNCS90]. recently [BB97a, KS94, LBH91]. reciprocal [JKT90b, KJT90, TAJK90]. recognition [AMFB99, LT97, SWEC91]. recolonisation [HBR^{+}91, MS94a]. Recolonization [Dye93, Ber90, LD95b]. Recommendations [Str99]. Reconstruction [DMA^{+}94]. reconstructive [DD97, JRB96, JRB97]. record [HP96, HTLK99]. Recorder [BM99b]. Recovery [BCC94, MR91, RDM91, VGJ90, VH96, vTSD97, BCC96, BA91, DF94, GJ96c, JDSB99, KHTP99, KD96, PBVB99, SM91, SB99, SR95, WUA^{+}99]. recreational [MMBT97]. recruiting [AALT99, TSM^{+}94]. Recruitment [AMBH98, Egg95, FSCE92, HB95, HB97, IW91, LB95, MC91c, MS91, Mye95, OP99, SC94, SL95, TH94b, VWM90a, VWM90b, WS91, vPB90, AB90, ACLS97, Ang91b, BRLM90, Bee97, BH91, BS96a, BCC92, BM91a, BWRH99, Bro95a, CD98, CLD99, DRC^{+}97, ELMR98, EFSS95, FAED97, FW92, FDW92, GK96a, GF91, Hol90, HvdMBK98, HH99, HS97c, HH91b, HSB97, JBWH97, JSOG91, JC95, KW97a, KCS94, LD95a, Leo99, Lev91, LCG97, LTM^{+}98, LGL^{+}95, MZ99, MSC95, MY96, MCNB98, MMD93, MMD92, MS93d, NC90, NB96a, O’C93a, OWM92, OW95, PGM^{+}98, PS92, PSL96, Pfi97, PF98, QC94, Reu98a, RH90a, RSB93, RM95, Sew94, SU93, SFA98, UMTM98, VBCRB97, WTR99, WB92a, WH91, WP92, WLL97, WG97, YBG96, ZN96b, vTSD97]. recruits [Ken94a, LKC98, Mum99, WC96]. rectangular [BMHH94]. Recurrent [Hag95, JBO97, SH97b]. recycled [HSCC96, UK98]. recycling [Eno93, TS90a]. Red [KL91c, YHC^{+}99, AdN91, CTEF91, DH90, DKVD95, ESDK94, EDS^{+}99, Ekl98, FSA^{+}97, Han95, HBM93, IWF^{+}97, JL94, Jeo94b, JSK^{+}99, LJ96, LC93, MST98, MKV95, NU91, NYH92, NSH95, PBHMI98, Pre91, RH97, SDER96, SHFP94, SK98c, TFF94, UYMH95, UGSBA94, UT90, VNB97, VBN95, WMG91, YKK^{+}98, BS92, BS95c, KKL^{+}98, KPWL93, KWL98, LVP98, OBL97, RH90a, SL91, Uib95]. red-tide [Jeo94b, JSK^{+}99, MKV95, UYMH95, UT90]. redfish [And94a]. redistribution [SMH97]. redox [KE95]. reduce [RHS99, Wah96]. Reduced [SH99a, MNK^{+}97, Pel96, SGL^{+}97, Sie93, VVT97, VH96]. reduces [SAH98, Wah97]. Reducing [SR93b, BL97, BR91b, SJMC93, VWJ96]. reductase [BCH95]. Reduction [CS99, GCLS94, AJST98, BEG97, EPN^{+}99, GSBB97, HHB93, HK92, HK94, IF96, KSS90, KHB91, KDAS92, KKH^{+}94, MDC98, Onu96, PM91a, RH90b, RPL96, SCLS94, UNPH93, WW92b]. reductions [SCL95]. reedi [RE96]. Reef [AHS97, BMSF95, Cal95b, SU93, UDLM99, AvS97, And94b, AW95, Aro98, Bak90, BJdJ^{+}98, Bee97, BBLY98, Bel95, BJB95, BPL95, Byt90, Cal95a, CDHD92, CH95, CP95a, CK99, CCRB98, CT97, CWB93a, Cla96, CG99a, Con98, Cro99, DKL94, DG93, EW91, EB96, ELG97, FK95a, FD99, Fow90b, FDW92, FHCC98, FBR^{+}95, Fre97, GWW^{+}98, GPDF93, GMO95, GK96a, Gil97, Gre96, Gri94, Gru94, Gru95, GP98, GBJ94, GT98, HGS95, HGJ99, Joh99, JKL90, JFS90, JFS92, Ker94, Ker96, KMG98, KSO^{+}97, LS96a, LNSL91, LB96a, LGJ99, Lee90, LSGW98, LCE97, Lev91, LH96, LCG97, LBH91, Mac91, MG96c, McC94a, MM93, McC94c, MM95a, MM97a, McC99, McG94, MMD93, MNB94, MPB97, MMD92, Mil94a, MD94c, MKYI98, Nai93b, NKGK92, PCTF95, PCB90, PR93b, RK96]. reef [RGLS99, RHH97, RW99, RSB93, Rog90, RMT91, Sak97, SRSH99, SML97, SOW97, SK98a, SK98b, Sch99, SNC^{+}98, SC94, SB99, SDC^{+}92, SC96a, SFA98, SP99b, Sto97, SB97c, SEAI96, Szm91, Tay98, TC93, TM97, TSM^{+}94, Tol98, TH94b, TB97, UK98, VB93, Van94b, VK94, VB94, WO94, WCL97, WM97, YMD94, YFI91, vRBVB95, vRV96, vTS97, BEU99, Bel95, BM91a, CL97, Fab95, Fow90a, FDW92, HFJS91, KM92, LC97b, McC96b, McC94d, MD94c, MKYI98, PD90, PD94, RS91, RSS94, SR90a, SRSH99, Sam97, SB99, SU93, Tho92, TcBW96, VDG95]. reef-building [Byt90, FBR^{+}95, HGJ99, LNSL91, MNB94, MPB97]. reef-dwelling [RGLS99]. reef-fertilisation [BBLY98]. Reef-fish [Cal95b, Con98]. reef-inhabiting [FK95a]. reefs [Adj97, Aer98, AU93, BLJ^{+}96, BAH94, BCBC98, BMH91, Bro94, BLST90, Cli95, DB96b, Fab95, FGB^{+}96, GPD^{+}96, GGH95, HSH94, HSK^{+}94, JRMN^{+}97, JAB^{+}92, JKFB95, JM93, KUK91, Lew97, RS99b, RW99, Rog90, SKM93, WC90]. refeeding [FFF^{+}94]. reference [ACR99, BS95c, Dai90, KI97, MvdZH^{+}97, MNH92, Nak92, RAS97, SC92b, Tan98a, Tan98b, TH90]. refined [LKC95]. reflect [CD96, MM95b, She96]. refractory [BVS^{+}91]. refuge [AH96, AH98, BA90, LFS90, PBGE91, RHP93, WW96]. Refuges [Ski94, ACLS97, GC95b, War90]. refugia [RvMM97]. refutation [LB95]. regalis [CE93, DE94, GT96, TJST98]. regenerated [SFPCR93]. regenerating [VB94]. Regeneration [KWL96, VSQ90, CBG^{+}99, DW95, GMG^{+}92, GTL^{+}91, GBJ94, HB94b, MB93, MNB94, MPB97, NB98, OBL97, ORL97, Rin96, TA90, TK91, ZB94]. regime [AP99b, CBG^{+}99, DW95]. regimes [Joc95, KW98, Mor96, PLG^{+}96]. Region [Nor92, Alo90, AAW98, BE97, CF93b, DG96, FLDT92, FVDL95, Gib94, HRÆ94, KL95, Lav94, LGGLLHV98, LFTH92, MG91a, PML93, PP96, Pal91, PD90, PD94, RCC^{+}99, TBY98, TvLdJ^{+}98, TZ99, WSC^{+}99]. Regional [ASSWM97, WASSM97, CWHA99, LSH96, SK95b]. regions [AMK99, YSY^{+}99]. regression [BJTB96, LC97b, Sal95, SBT97]. regrowth [GW91]. regularly [ECF^{+}97]. regulates [ELMCC90]. regulating [EA92, JA91, MP96c, SK94a, CPB96]. Regulation [BL94c, GCGS90, MG96a, MPRB93, Ole96, PCK^{+}96, vBV90a, vBV90b, CR96, LM98b, MGYV97, MCRGC90, MS96b, MZFHC96, RG97, SD94, TTL^{+}96]. regulator [KBA^{+}97]. rehabilitation [MF97, vTS97]. rejection [BV99, BVP99]. related [AH96, AH98, CN92b, CWB93a, CWHA99, Dem95, DEP^{+}96, GLD^{+}91, HMB93, HMB95, HN97b, HSBP97, Ito97, KRJ^{+}95, KPWL93, KA93, LGV^{+}99b, LQ97, LOS90, MC91a, MB95b, Mol97, MSM^{+}98, Pel96, PD95, PFMC95, Pud93, RAS97, RFGF92, SPB95, SDB96, SDG95, TL91, Urb94, VPGV94, WBP^{+}92]. Relating [LW94, MKV95]. Relation [TDBH94, Aer98, AGW93, Alo90, AT96, BH92a, BM94d, BLM^{+}95, BLR94, CA97, Cam91, CN92a, CSM99, DDW90, Dye93, GW94, GPP98, GPB^{+}98, GKB^{+}97, HKMK98, HMWN97, HK99, HvdM97, HRCW98, JWKH98, JKL90, KL91a, Ker94, KKK^{+}90, KJMRC96, KDHZ98, LC95, Lig90, LBK90, Mar97, Nie96, NKKO97, Nor96, OK98, PMMH95, PNS90, PHT96, Pla93, PR93a, Pre91, Pro92, RDT^{+}94, RZWY96, RSH92, SH97c, SFS95, SFPCR93, SW98b, SF95, SGW94, SRW95a, SDB92, SDB94, SAA99, SK95c, TDVF99, TD92, TRU91, TK96, TPW^{+}93, TS90c, UCV98, VWD^{+}95, VR98, VDA96, Vos91, WF90, WPB^{+}93, YST^{+}97, vdL95]. relations [GCR98, MRP98, STNI97]. Relationship [AC92, DD90, DCBC91, DGHT^{+}98, FCVS92, Gru95, KRJ98, ML92, MGF98, NDRP93, ONR98, PWPK95, RLDM94, Sha95b, TTW97, TFFS90, ADA98, BD96, BGCM95, DFB96, Dut98, GHTH94, Hol90, HHW99, Ito97, KYI^{+}98, KTW^{+}97, LM91a, LC97a, MD91, MBR94, MADM98, ODS^{+}96, SM98a, Sal95, vdVW99]. Relationships [Cha98, GK96b, OSC97, OSC98, RZLB96, SCB92, ZF97, BSP^{+}99, BC90, BBL93, CO92, CVD99, CBT98, CLD99, DL93b, FP91b, FCP98, FL90, FAP92, GV96, GP98, HSRHH90, HW92, Hob93, HAR95, HSN93, KMR98, LED92, LIPF95, LS91, MvdVD^{+}97, MG94, Pud93, RTRF90, RAS97, San94a, San93b, SDFJv91, SJCL93, TT99, WFTH97]. Relative [BN95, BH92b, CH93, KB91, LC90, BEU99, Bre96, CF94, CWH98, EAC^{+}97, HC92b, HOM95, JBO97, KHBL95, MGYV97, Mye95, OMH95a, OMH95b, Pal91, SFF98, TRMW95, WH95]. relaxation [HDD98, WBLM95]. releasable [And98]. Release [FPGC^{+}98, MGD91, NK92a, WSS94, AS99b, AKK94, BMS^{+}97, Bel95, BLM^{+}95, BG91, BG93, CBDC96, Col98, HG92b, Hur92, JG90, LF92, LF94, LBL97, NK90, OH95, Pak92, Pal97, YNYK94]. released [AHY^{+}99]. Relevance [SHW^{+}91]. relevant [ABY97]. reliability [CM98b]. reliably [MLP90]. relict [HS90]. remains [KB94, SR96]. Remarkable [CTSN90, NTS90]. remineralization [ASP91]. Remote [HHYD92, TNKMK99, MGEC97, WWGK95]. remote-sensing [WWGK95]. Remotely [And98, SCC94]. REMOTS [RS92]. removal [Con95, GWW^{+}98, Gru96, LWK^{+}95, Vis93]. renewal [VNM97]. reniformis [CBA^{+}99]. reorganization [AP99b]. repeated [TNP97]. repetitive [MRF^{+}99]. Replacement [LM95]. replenishment [Mil94a, MD94c]. replete [KMR98]. replicate [Las98]. replicates [KSE96]. Reply [DHB^{+}98, Kep98, JKT90a, PPFP99]. report [CR98]. representative [MDD95]. Repression [Dut98]. Reproduction [CRZG95, DF92, KEV99, KN99, MS98, PR93a, RH90a, SH98a, Ang91b, Ang92, BA97, BH92e, Cli95, CD99, HKMK98, HM94b, KL91b, Kog95, Laf93, LHL93, LF97, MBR94, Moe96a, Moe96b, NGS97, PWNW99, Rin96, Sal95, SW93, Szm91, WH92, WCL97, vRBVB95, VB94]. Reproductive [BL97, CS97, CF93b, DBMM99, Ekl98, FD99, HvdM97, KWL98, LPI^{+}95, PB96a, Per90, PLIM95, RDA97, VTGK99, Van94b, VK94, VB94, Åbe96, AS94b, AHS97, BSMS95, BGBS91, Bre95, BSH^{+}91a, BSH^{+}91b, CZG95, ER90, Geo94, HC93, HM97d, KKL^{+}98, LFG98, MCNB98, MT94, Mor96, Nak94, NKH^{+}99, NAS95, PRC^{+}92, PCT^{+}97, SS91, TCRB99, UTB^{+}95, UMTM98, US95, Uye96, WBP^{+}92, Zaj91a, ZHBH90, vG91]. Reptilian [HGBW99]. requirements [GWKW97, GSC95, HC92b, KK91, KG94, MP96a, MSJ92, NW97c, PB96b, PB97, WtD90]. research [Bro95a, DHS96, EFSS95, Kin93, VWK96]. Researching [LD95a]. Reserve [CK99, HSH94, ZMD95, BA91, CRH^{+}99, Lee90, LD96, RK96, SRH^{+}98, HB95]. reserves [BKS^{+}99, BDM96, PR93b, PDS97, RA96, ST98]. residence [RSD^{+}91]. resident [MAAT98]. resilience [FA99]. Resistance [TJRG93, HBD95, PGB96, Reu98b, WLL98]. resistant [JG94]. Resolute [LDV95, MBJ94, PMOD90, SC92b, SCEG93]. resolution [ARS93, ADG^{+}95, BCG97, Cha98, DAS92, JLF95, MC95, OSC97, OSC98, SSR^{+}97, VWJ96, HCI93]. resolve [KDC99]. Resource [AZKA99, HAA94, SFV96, AH99, BB92, HS98b, HSB97, LK99a, LBJ^{+}98, OM99b, PKY96, PAR98, PS98, UTB^{+}95, WB92a]. resources [CW95, JCCCB94, JCH93, LGC^{+}93, PC96, SB97a, WR92a]. respect [RM95, TK95b]. Respiration [CRLP94, KJN97, TW98, vdL95, AKL^{+}99, AMB^{+}96, CCG98, CLB92, CR96, CG99a, DKW96, EJ96, FPGJ99, FRK92, FZ91, GGH99, GCLS94, HGS94, HCF93, JSJMH90, Käh98, KSG^{+}92, Kep91, Kru93, MM99, MNK^{+}97, NEIR95, PBv^{+}95, PGAJ99, PSSP95, RAlBW99, SO91, SK94a, SFSA99, SK95b, SCF91, SE93, VM98a]. respiratory [Rel96a, Rel96b]. respond [TPHC91]. Response [BHG94, BAR96, DTA^{+}97, EPN^{+}99, FTW90, HK96, HKR96, HR96a, HS97a, KR98a, Mil92, NJSS91, PSL96, AdN91, AB95, ABY97, ASP^{+}96, APT99, BCHL93, BG94, BRZ99, CS99, CJ95, CCS98, DKW96, Dav94, DL91a, DHKP96, Ekl98, EP92, FTGW90, FFJ93, FM94, GJ96b, GT98, Han95, HBM93, JH97, JC97, JP93b, JHW95, Kep91, KWHM^{+}96, LPI^{+}95, LBJ^{+}98, LA98, MD94a, MJ97a, MPH96, MK93a, MHD97, MvdZH^{+}97, NW97b, NR94, PG94, Pez91, PLIM95, PLC^{+}91, PF92c, Ral98, RS99a, RKH96, RMH96, SCIA99, SG96, SP99c, Ter97, TN90, UK98, Van92b, Vet96a, WB92b, WD94]. Responses [AV98, LH96, Mar91, Rog90, UDLM99, VY96, ASF98, BBLY98, Ber90, BCH95, BSB91, BPL95, CRP99, DF95, DKVD95, Ejd98, FS90, GAD^{+}92, GMK97, GT92, HG98, HLF^{+}91, JSGS94, JAOS90, KMB97, KCV94, MF91, MMM^{+}95, MCL90, MCED96, MBWT98, Mon96, MRS96, NvA91, Nai93b, NK93, NB96b, NB96c, Onu96, PPFP98, PSK98, PWD^{+}91, RGS90, RK95a, Reg92, Sch99, Sko96, Tis98b, Tro94, VR98, WW91, WW93, WFH99, WC91, IPT94]. responsibility [Fri97]. responsible [YHC^{+}99]. resting [KHT93]. restricted [Pla93, Tro92]. Restriction [Mau92, YKK^{+}98]. resulting [GPD^{+}96]. results [Amb93, BBRBK94, BRA^{+}99, CN92b, DAS92, DB96b, KH92b, Ken94c, KS97, SMT97, TGG97, THCD95, JCH93]. Resurrection [SH92b]. Resuspension [DMM99, BSG^{+}97, PSPD96, SS99a, Wai90]. retaining [Put90]. retarding [GS94b]. Retention [DLHB99, Bro94, CM99, CB99, DL91b, HTG96, HN91, LKF95, NNR95, SC94, SDB92]. retinol [BCA^{+}99]. retraction [BLD94]. retrieval [WK98]. Retroflection [DL93b]. return [FW91a]. Reunion [Nai93b]. reveal [IFL97, WC95]. revealed [AMO^{+}99, Rv92, vBV95]. reveals [ASA96]. reverse [BCG97, CN95]. review [NR98, Rel96a]. revisited [DvdTCC97, Jør96, MF98, PN99, RVT^{+}97, SC95a, VL97]. reworking [CJ93b]. Reynolds [Nag96]. RFLP [PPO^{+}99]. Rhincalanus [WSCT96, WASSM97]. Rhithropanopeus [FFT98]. Rhizome [MD98a, GMMD93]. Rhizophora [COG97, DGHT^{+}98, ZCL97]. Rhizosoleniacea [KKT96b]. rhizosphere [MRPC98, PBDF98]. Rho [LFBH94]. Rho-pilema [LFBH94]. Rhodamine [MM96]. Rhode [CC97b, TDD94, DG91b, DG92b, GJC97, GJ97a, GJ97b]. Rhodophyta [ASL99, BCBR97, GW91, SDB96]. rhodophytes [KR94]. Rhombosolea [MJ91]. rhombus [NAS95]. Rhone [BBF94, OSGB96]. rhythm [KKYH97, RN93, TK98]. rhythms [Arm92, Pal97, PCK^{+}96, ZN96a, ZN96b, ZAN99]. Ria [QCM94, AA92, CMRF93, CSM99, Pre91, Wol92a, ASRP^{+}96b, ASRP^{+}96a, FGG95, TFFA99, ZF97, ZF99]. ribbed [KNK90]. ribosomal [YKK^{+}98]. Ribulose [PFP93, PCK^{+}96, PCC^{+}97, Riv90]. Rica [Bia91, DE90, GCGR90]. rich [CLC^{+}98, DFJ^{+}93, DHM92, FFH99, FW91b, HK94, Moe97b, SJ97, WP95]. richness [AB96a, CGP97, JMD95, Kar95, SMG97]. Rico [McG94, YY91]. Ridge [SJML99, AHY^{+}99, CD98, GJ96a, SRJD97, VTGK99]. Rift [FCMB94]. Riftia [ASF98, GSM^{+}97]. rig [RS92]. Riga [JTB99]. right [HKMK98]. rigida [VSPLH98]. Rimicaris [PDB^{+}97]. ring [DL93b]. rings [NB90]. Risk [GB93, BL94b, Ejd98, FIA93, Hal98, PS95, RHS99, TB97]. risks [Ano99c]. Rissa [WTA^{+}97]. Risso [Sha95b]. River [AALT99, BMG99, CBS91, DG91b, DG92b, FMRC93, GJC97, GJ97a, GJ97b, GM99, GFPD92, HVL96, LL98, MP94, PPFP98, PPFP99, RPMB91, STBR99, VS92, VHHS96, BWHC91, JCP93, KJLH96, KKYH97, MPRB93, SR93a, TNSM98, YHGB96, YGH^{+}97, BNK^{+}97, BBF94, BCS93, Che93, CLB92, CG93, CB96, FVDL95, GBA^{+}96, GF91, HCC^{+}91, HKY^{+}97, KKW^{+}93, LFR^{+}97, ND96, ONH^{+}97, OSGB96, PBA^{+}95, PSE97, RNHO98, SWT99, Sew96, TGG^{+}95, VDBF96, WK92b]. riverine [CGTB95, FS98]. rivers [SS97]. RMT [PRW^{+}94, RPM^{+}94]. RNA [BD91, CCV^{+}98, Cle93, CD96, FM95b, GCR98, HW91, LMK93, MHB94, MC91d, RBB^{+}91, WDB98, YKK^{+}98]. RNA/DNA [CCV^{+}98, CD96, FM95b, HW91]. Robert [PdG96]. Robust [CL97]. Rocas [NAW99]. Rock [CB99, Ang95a, Ang95b, BCC95, BCC96, CC97b, Jer90, Kal91b, PWS98, SDB92]. rock-pool [SDB92]. rockfish [SDBW95]. rockpools [BCC92]. rocks [JAB^{+}92]. rockweed [VÅCL99]. rocky [AFC96, Bec98, BCMC99, BH98, BKH99, CMRF93, CH96, CV98, Dye93, Dye98, FPB^{+}94, FA99, HS95, HS96, HS98c, JBHH97, JBH98, KW98, KD96, KUK91, Las98, LBH91, LTM90, MHS94, MS94b, MS93c, NW97a, OM99a, PH91b, Pin94, RK95b, RK95c, SB97a, Tay98, TC93, TR95c, Tok95, WMG^{+}94]. rocky-reef [TC93]. rod [FPFR92]. rodgersii [AU93]. rodriguezi [LTM90]. roemerianus [CBB90]. Role [ABN98, And94b, Bow97, CLGM96, CH96, DY92, DB99, FPPM96, FS98, GI99, HS97c, JKT^{+}98, LK96, MF95, NT92, NS96, PM90b, SK94b, SSP97, WH92, War90, AWdS92, AJL93, BWHK99, BM99a, BSJPW93, Bro96a, Bro96b, BDNS99, CC99, CBDC96, CNP95, DQT^{+}97, DG92b, Far98, GS94b, GF91, Han91b, HHB93, Har97, HPF92, HSW99, JBK99, JSOG91, Kel98, KKT96a, LA98, LKM^{+}99, MDG98, NT96c, OKL^{+}96, OWM92, Raw96, RHH94, SPH94, SCD^{+}90, SLSH98, TPR98, TTR99, TNSM98, Tro93, VCMP93, BH92d, MH97]. Roles [HlS90, PF91b, Har94, KSG^{+}92, MCC96a, MGYV97, Mye95, PWG^{+}99, SHM^{+}97, SJC99, TT90, WH91]. Root [DMA^{+}98, Blu93, IF96, TW97]. roots [SAvD^{+}96]. rose [SMT97]. Ross [NT92, SCDH98]. rostrata [WT98]. rotifers [DG91b, DG92b, LHRL97]. rotundata [HLF^{+}91, PBDF98, VSPLH98]. rough [SFPC97]. roughy [BK92]. round [DQT^{+}97]. routes [BFS^{+}98, SFF98]. ROVs [SCC94]. royal [HHM97]. rRNA [FD97, LMK93]. rRNA-targeted [LMK93]. rubens [TLW^{+}96]. ruber [GB97a, GB97b]. Rubisco [MG96a]. rubra [MS91]. rubrum [ASN93, SPDM91]. Ruditapes [PCBA94]. Ruegen [Bus96]. Ruiz [DHB^{+}98]. rule [BV99]. running [RN93]. runoff [JP93b]. Ruppia [BGC94]. ruthemum [LGB91]. ruthemum-106 [LGB91]. RV [Bia92].

S [AGP98, BR91a, FD97, LMK93, MH99, SJMC93, UC94, YKK^{+}98]. S-oxidation [BR91a]. S-tolerant [SJMC93]. S. [ESDK94, FMRC93, ME97, SDER96]. S9 [SHW^{+}91]. Sabella [AW97, RI90]. Sabellariidae [ZMD95]. sabellid [Hsi97]. Sable [BH94b]. Saccoglossus [FT95, GK97, KM96, Mil92]. Saccostrea [Bro90, HSTY98]. Saduna [San94a]. Saduria [EB92, HS90, HS94a, Vis91, VH96]. safety [RS94]. sagax [EF97, van94a, vdL95, vdL98]. Sagitta [Gib94, Ter93]. Sahara [FHA99]. saida [GCKW95, SCJ99a]. sailfish [Spe95]. saithe [MHC92]. Salamanca [GDvS98]. Salicotrm [GH93]. salina [LC90]. salinities [CBRD92, CLB92, PBA^{+}95]. Salinity [HL98, NW97c, ABM96, AHMd90, Ang91c, ACOK95, BTKK99, BKB92, BB93, CCD94, CCC94, CC96, CBB90, FPGJ99, LWL98, LKK92, MT91, MCL90, MNK^{+}97, MAAT98, Nie96, Pez91, PD93, PWNW99, QQ97, QQ98, QQ99, RW96, RD95, SDC99, She96, SS91, TDVF99, WA99a, WJD93, vKSGvA99]. Salmon [SH97a, FW97, FH97b, FS98, HDPW93, SOK98, STNI97, Web91a, Web91b]. salmonis [AS99a]. Salp [ZDJ^{+}95, MDTB94]. Salpa [Lav97]. salps [Lav97, PM91b]. Salt [BAH94, AWHH94, ATNW97, BI99, BM90a, BM90b, CCSN93, CNP95, DSW^{+}91, FT95, FW91a, HKvB93, HHSvS96, HG94, LK94, LTT96, MZ92, MW97, MH90, O’C93a, OJ99, Pak92, RSN^{+}99, RD95, RR93, SR97, SM90, TA93, Tur93, dG92a]. saltatrix [HC93, MC91c]. saltmarsh [KNK90]. salts [Pre91]. saltwater [VBF94]. Salvelinus [RSM^{+}96]. Sample [SH90b, FC90, JMD95, MF98]. Sample-size [SH90b]. sampled [PRW^{+}94, RPM^{+}94]. Sampler [BVV^{+}94, DAS92, Sty97]. samplers [HS92]. samples [AGGKA96, AFZ96, BKS92, BG94, CV91, DFB^{+}99, Gar91, Gow93, GBH96, MRF^{+}99, Ohm96, Tha92a, Tha92b, Tho91]. Sampling [BB97b, RM96, BP97a, Blo91, Cer98, EB96, HS99a, JBK99, KW91, KKW97, LAJ95, LWK^{+}95, MS93d, Roz92, TH94a, Tho92, VS92]. San [BL95, CCG98, CL90, CTSN90, JCP93, KGO94, LOS90, MC96a, NTS90, WSP^{+}95]. sand [Asa91, DB96a, FB91, HBR^{+}91, Hüt90, Kog95, MF91, McL96, MSD92, OSZ95, SW93, TK95a, TK97, TK98, Ter97, Vet96a]. sand-bubbler [Kog95]. sand-burrowing [TK95a, TK97, TK98]. sand-dwelling [Asa91]. sand-living [DB96a]. sandbank [Hic92]. sandflat [BM94c, Cam91, FTP99, Gün91, KSRB99, SPB95, TPHC91, TPHC94]. sandflats [AOOT94]. sandhopper [CV96, GD99]. sandperch [San93a]. Sandsfjord [TSB93]. Sandwich [PLM92]. sandy [ACR99, BLJ92, BD99, DdA95, ER95a, EAC^{+}97, GAR93, GRBA96, GD99, Har98, HP98, JMC93, JMD95, KLJ94, MM91b, Per90, PSR^{+}99, RBA98, San94b, vDvRK93]. sanguineum [DH90]. sanguineus [EDP^{+}98]. sanoessa [Nor92]. sanpaulensis [CF93b]. Santa [HBC^{+}91, OC97, Sha95b]. sapidissima [Lim95]. sapidus [BA90, BM95b, CWWH99, DS93b, EK95, FW91a, FFR94, FDR^{+}96, FTBR97, JSH99, MZFHC96, Olm94, Sha98, WRBF97, WFH99, WD94]. Sapphirinidae [CN95]. Sarcodina [MvdZH^{+}97, MSvdZH98]. Sarcophyton [Fab95]. Sardina [CCV^{+}98, GMT99]. sardine [AM93, GMT99, LSB95, vdL98]. Sardinops [EF97, van94a, vdL95, vdL98]. Sardmops [AM93]. Sargassaceae [KW91, Ken94a]. Sargasso [BMD^{+}90, BK96b, Mil95a, SCF91]. Sargassum [HM91, KW91, Ken94a, MS94a, McC96b, SK98a, SK98b, SL97b, Ste94, SP99b]. sarsi [DS93a, DBN^{+}94]. satellite [BGG^{+}97, GKB^{+}97, JWL96, JCCCB94, MGEC97, RPT^{+}96, WWGK95]. satellite-derived [JWL96]. satellite-tracked [BGG^{+}97, GKB^{+}97]. saturable [LF93]. saturated [BGR^{+}96]. saturating [HPSG94]. Saudi [EAT95]. Savannah [WK92b]. saxatilis [Che93]. saxitoxin [KB91]. saxitoxins [SOK98]. scalar [KLJ94]. Scale [CO92, DV92, JW96, RAS97, ÅP97a, ADR95, ASRP^{+}96b, ASRP^{+}96a, AB92, Aro92, Aro98, Arr94, BH97a, BN91b, BF99, Cal95b, CP96, EB96, FBHH93, FB91, FDW92, GRGK96, GBP^{+}97, GB98, HV91, HT96, IFL97, MC95, ML91a, MU99, MJ97a, ML92, PG91, PG94, PS92, RÁSP99, SR90a, SLB^{+}92, Sch90, SH97b, Sch93, SD93a, Sym95, TDER97, TcBW96, Tol98, VM98b, WB99]. Scale-dependent [CO92, JW96, GB98]. scale-independent [Aro92]. Scale-related [RAS97]. Scales [AB96a, BB99, Ano99d, AMBH98, CB97, HB97, Hul99, HJHG98, JBHH97, KM92, LB99b, MC96a, MP96c, MM95b, RGBJM95, SPB95, Sun96]. Scaling [KS95b, SM97c, TPR98, CPK97, Dua91, ELMCC90, RG97]. scallop [AI92, BSH^{+}94, CHC94, CP96, Fev92, GF94, HBM93, HBBT98, MW94, MGP^{+}96, MONR90, PMMH95, PGM^{+}98, PS92, SH93, TRU91, THCD95, TS90b, TS90c, VCMP93]. scallops [AV95, AMBH98, AH96, AH98, BS94, BSH^{+}94, CHS^{+}95, DLDL93, EGMC94, PRC^{+}92, PSL96, PBGE91, TRS91]. Scandinavia [KLMR91]. Scanner [BE97]. scanning [BP93]. Scapharca [Vis93]. scarcity [BW96]. Scaridae [Bel95, BKB94, CAL96]. Scarlato [Ano94]. Scartichthys [OM99a]. Scarus [BKB94, BvKvRB96, Cli95, HK96]. scatter [LZYW93]. scattering [BA93, CHDW92]. scavenger [MM97b, Wit99]. scavenging [KS94, LBLA93, Moo94]. SCE [PMETM96]. schedule [SPB^{+}99b]. Schelde [KP95]. schizasterid [PF95]. Schooling [KKH98, Cro99]. schools [AS94a]. Sciaenidae [SHFP94]. Sciaenops [RH97, SHFP94]. science [Ano99c]. sclerac [GYL93]. sclerac-tinian [GYL93]. Scleractinia [Cha91, Hoe91, Joh92, Sak97, YAG92]. scleractinian [BLD94, GEW97, HB95, JLGMP97, KCD^{+}95, MSC95, PMB99, RVGC^{+}99, Tom91]. scleractinians [Dai90]. scleractmian [JS93]. sclero [SC90]. sclero-sponge [SC90]. sclerochronology [LCBD94]. Scolelepis [San94b]. Scomber [MSD98]. Scombridae [YD90]. scombrus [MSD98]. Scope [RWJ99b, WDB^{+}95]. Scophthalmus [CTC93, CMT94]. Scopimera [Kog95]. scorpaenid [VL97]. Scotia [CHS^{+}95, HS95, JNCS89, MHS94, MS94b, MPS95, SCKS90, CSB94, GTL^{+}91, GST^{+}91, PRW^{+}94, RPM^{+}94, SH90a, SH97b, dBN^{+}90]. Scotian [DLFL96, MF96, SM91]. Scotland [MGM96]. Scottish [ACR99, EG95, GAR93, GRBA96, GTJ92, SFPC97, THR^{+}98]. scouleri [RGLD^{+}98]. scour [CLKO98]. scouring [GSD96]. Scrippsiella [IT96]. Scrobiculaha [RBG94b]. Scrobicularia [RBG94a]. Scylla [CC96]. Scyllaridae [Jon93]. scyphomedusa [HL98]. scyphomedusae [Bro98, CH93, OPS96, PND^{+}94]. scyphomedusan [LFBH94, Pur92, PC95]. Scyphozoa [BS95a, Han97]. scyphozoan [PWNW99]. SE [dG92a, Ano99a, FPB^{+}94, CAATS94, WQQ99]. Sea [Bar95, BH92e, DMP^{+}99, HM97a, MHRV^{+}99, NSHF93, NLR^{+}93, PF92b, RG93, RC91b, SSHM91, TD91, TS90c, AMB99, AM91, AU93, AS94b, ALB^{+}96, AB96b, AAW98, BS94, BSH^{+}94, BCC95, BCBC98, BL97, BVV^{+}94, BA97, BPH95, BL94c, BL96, Bro96a, Bro96b, BBG90, BK92, CSC92, CHS^{+}95, CA92, CS93, CL97, CJ93a, CSSPV98, CHFL95, DDM^{+}99, DEF91, DAS92, DE91, ESDK94, EDS^{+}99, EGMC94, FGB^{+}96, FD97, GSM^{+}97, GSC96, GB91, GC93, GGF99, GPDF93, GPF95, GPD^{+}96, GT92, GKRD96, GF94, Gra91a, Gra94, GPU^{+}97, GKB^{+}97, GGD99, GP90, Hag95, HS94b, HW95, HW99, HGS95, JDLL95, KLR93, KSS90, Kea90, KL91b, Kep91, KW99a, KO92, Lam98, LB99a, LC95, Leh96, LA98, LCW99, LC97b, MW94, MGP^{+}96, MBJ94, MB98, MA93, Mau92, MC97a]. sea [MC94a, MPH96, ME97, MMG97, MLTD96, MPCD90, MGD91, NBS^{+}97, PH91a, PS93, Por94, PLP92, PDP^{+}90, PLC^{+}91, RR99, RZLB96, RKS97, RSK^{+}99, RSD^{+}91, Row90, SZ96, SRH^{+}98, STK99, SR90b, SH97b, SDER96, SFSA99, Sew94, SRG^{+}99, SMTL99, SPH94, SH90b, SHV91, SH95b, SB96, SHT91, SBP92, SBP93, Ter93, TRU91, TRS91, TMC98, TMCE98, TS90b, Tro92, Tro93, THR^{+}98, VFMFBE96, VWD^{+}95, WTF96, WH91, WWGK95, Wer97, WBGS97, YTF97, dNSR^{+}96, AFC96, AR93, AKK94, AW98, Arm92, AH93, AH94, AH99, BSVK93, BMP^{+}95, BNvB90, BM99b, BS95a, BLDF95b, BLDF95a, BR94, BC90, Beu92, BF95b, BBF94, BRL97, BMD^{+}90, BH92c, BH92d, BDR^{+}96, BD98, BS92, BS95c, BSP93, Bro98, BM95a, BK96b, BG99, BvBBV96, BBP^{+}96, BGCM95, BCv^{+}92]. Sea [CSM^{+}98, CPS91, CS93, CSR^{+}92b, Cha92, CCG99, CMLE92, CCCH94, CJ95, CCS98, CPH^{+}96, CHDW92, CH94, DvdM92, DS96, DAB^{+}91, DHH98, DGT^{+}93, DB96b, Ejd98, EE98, FNJ^{+}94, FPR94, FGKP95, FGS98, FMRC93, FH99, FPP^{+}97, GBC92, GJVG96, GDvS98, GC93, GH94a, GCF96, GDP^{+}98, GLBF92, GLD^{+}91, GCWH90, GH92, GW97, Gün91, GGD99, HWH98, Han92, Hei91, Hei92a, Hei95, HTG96, HW95, Hil97, HS97a, HHP95, HB98, Hoe91, HNvB94, HW91, Hov92, HCH^{+}96, IHK^{+}98, Joc95, JKT90b, JL96, JP93a, JBB90, JTB99, JPZ^{+}98, KLN^{+}90, KH92a, KVK98, KJT90, KV90, KKL^{+}98, KLEH91, KL91c, KPWL93, KDR97, KMB97, KBWC90, KDMR97, KWL98, KFS91, KW99b, Kuo91, KGW93, KN99, LC91, LBL91, LIPF95, LGH^{+}93, Leh96, LA98, LNQT93, LZYW93, LVP98, Lig90, LGH95b]. Sea [Liv93, LACM92, LMS^{+}93, LKvRH96, LEHvR96, LGdC^{+}90, MBF^{+}97, MMM^{+}95, MEK99, MG97, Mar97, MJV92, MHC92, MVA99, MH93b, MC94a, MD94b, MS98, MHvB97, MMR99, MB97b, Mik95, Mik97, Mil95a, MHD97, MPS95, MO95, MTP92, MNH92, Mun93, MLDM99, NPB94, Naj96, NSH95, NG99, NT92, NKH^{+}99, NS96, NKV94, NW97c, NK92b, Oes90, ÓMvdB99, OSGB96, OBL97, PH91a, PMB99, PGB96, PGR94, PE99, Phi94, PRW^{+}94, PBv^{+}95, PS97, PD95, PF98, PH95, RCS^{+}96, RSH93, RHT91, RAB^{+}92, RA92, RNP^{+}98, RH90a, Rie93, RFN98, RN98, Rv92, RVv92, RD97, RZWY96, RPM^{+}94, Röp93, RHJ92, RS92, RHrsHG99, SPB^{+}99a, Sch96, SKvB98, SL91, SH95a, SFPCR93, SGPAA95, SH92a, SCDH98, SH95b, SK94b, Som96, SSV97, SP99c, SCKS90, SE98, SCF91, TNKMK99, TAJK90]. Sea [Ter93, TGB99, TH90, TGvJW95, TGSN99, TP95, TP97, Uib95, UHL^{+}97, UNPH93, Väi90, VKV95, VAV99, VDT92, VNM97, WVMR90, WDB^{+}95, WHG99b, WBGS97, Wit99, Wöh96, YSY^{+}99, ZBW92, dT94a, vDK90, vDvRK93, vdBTvDR96, vCRS92, vPB90, vBDW91, vW93, vdVW99]. sea-grass [PH91a]. sea-ice [GT92, MLTD96, SBP92, SBP93, Wer97]. Sea [?]NorthJorgensen:1990:AMM. sea-surface [GKB^{+}97, MC94a, WWGK95]. seabeds [ZHF96]. Seabird [GH92, CW99a, JCCCB94, O’D98, OJR96, RAS92, RAS97, TBAB91]. seabird-environmental [RAS97]. Seabirds [CW95, ARF92, BMSF95, GH94a, GCF96, Hob93, MGF98, MM95b, Nev99, SH97a, SJ95, SPCF97, Str99, WWGK95, WK98]. seafloor [FAHW99, Gla99]. Seagrass [DSJ90a, DSJ90b, Dua90, Dur94, FF97, Onu96, VH99a, ARDL97, AI92, BBFD98, BL94a, BH97a, BNW94, BM91b, BAM92, CDME97, CLB91, Con94, Con95, CJ93b, DBB^{+}95, Dua91, DMA^{+}94, DTA^{+}97, DMA^{+}98, DBdCS99, DT94b, DH92, DD97, EAD92, EM93, FK95b, FKC96, FMH96, FB98, FZ91, FHCC98, Gra91b, HHv91, HSK^{+}94, HD97, Hol97, HS99b, IF96, JBWH97, JRB96, JCH93, KG99, LD96, LCLC96, LWK^{+}95, MD94a, MDC^{+}96, MD98a, MvdVD^{+}97, MR97b, McG95, MKYI98, MSD92, MRP90, MGEC97, MF95, NA99, OSJ94a, OB93, PB93, PDC97, PRPM^{+}94, PM91a, Pre95, PAM96, PM98, RS99a, RS95, RBC^{+}91, RGBJM95, RH97, RD99, RSK^{+}99, SM91, SDC90, SAvD^{+}96, SMH97, SRW95b, SR96, TDK97a, TF98, UDLM99, VH93, VWS92, VWS93, VWK96, VAD^{+}95, VDM^{+}97, WB94, Way95, WSMM99]. seagrass [WP92, WCLB99, ZB98, ZB99a, ZB99b]. seagrass-associated [JBWH97, SR96]. seagrass-derived [MKYI98]. seagrass-dominated [ZB98, ZB99a]. seagrass-mangrove-coral [FHCC98]. seagrasses [ADF96, BK96a, BVW^{+}98, BH91, BUWB99, FH97a, GMR^{+}94, Gra91c, HM96, KM99b, MCED96, MW99, MA99a, Ter97, TB99]. seal [FHA99]. seals [BH94b, BP98, BBCB99, HWH98, HHE92, IFL97, LADS98, LK99b, MBWT98, RAB^{+}92]. seamount [LD95b, SNJV92]. seamounts [AHY^{+}99, HFG^{+}95]. search [ER95a, RO97]. searching [WWL97]. seas [CG99b, CMBM99, SBV98, Pak95, SSA96, dBN^{+}90]. seaside [SRRdM99]. season [BSH^{+}94, BW96, CHS^{+}95, FTP99, HLP99b, HHM97, KD96, LL97, RM95, SMUS^{+}99, SAA99, TAD^{+}94, TMCE98, VM98a]. Seasonal [AJS95, And90, BR94, BHT98, BDA^{+}99, CPB96, CH99, DDM^{+}99, DFR95, DL95, DE90, DF95, FPR94, GJ97a, GJ97b, GAR93, GT98, HVK96, HWH98, HDPW93, HL90, HCG^{+}93, HD97, HK93, Hul97, JP94, JMP90, JLGMP97, KTH98, KVK98, KD99, KSG^{+}92, KLEH91, KR97b, Kre93, KHBL95, LH98, LFG98, Mac91, MB98, MLC92, MVA99, MB91, Nai93b, NT95, OSJ93, OC97, PMMH95, Rei90, RCG98, RC91a, RBN99, RPMB91, RCN95, RTRP^{+}98, RNH99, SO91, SMT97, SS96a, SSHM91, SFSA99, SV97, SK95b, SCJ99b, TGG97, TR91, TL94, TDD94, TS99, UNPH93, UMD92, US95, US98, VHHS96, VHH^{+}98, VWS92, VD95, WR92b, YSL94, vDK90, AWD95, AGW93, ASA96, CL90, COK99, CLM^{+}93, Cli95, CMBM99, CSM99, DM91, DP91, ELMR98, Fal96, FTS97, FPP^{+}97, Irw91, KW97a]. seasonal [KW98, Kal91b, Kuo91, Lai92, Leh96, MCED96, MF90a, MF90b, MG94, MP96c, Nak92, OB94c, PHHH96, Rud97, SH90a, SD93a, SMTL99, SK94b, SMH97, TMKL97, TH94a, TPHC94, UK98, VF99, VWS93, ZBW92, DLT^{+}97, Zaj91a, Zaj91b, ZD99]. seasonality [ADA98, ADR95, BM99a, BSP93, DBMM99, PPF95, YMD94]. seasonally [GFPD92]. seastars [HD91]. seawater [ASH93, AFZ96, BG94, CDGF^{+}93, GBH96, JBO97, LGB91, MJ93, NK96, PMOD90, PF92c, SC90, Sim92, SGPAA95, SHML98, Sut92, TSS93, UKR^{+}91, WP95, WM90, WF94, WPG99, vdVW99]. seaweed [Åbe90, LC98, San94c, San95]. seaweeds [CLHF95, PNS90, RWCS98, TC94, TPR98, TTR99, WC96, dNDS98]. Sebastes [And94a, SDBW95, VL97]. Sebastolobus [VL97]. Sechelt [HTS92, THS94]. Secondary [MDSB92, Vet96b, BCC96, BS95b, CV96, FP91a, LB95, Lig90, LSP99, MP92, MC94b, MWH91, Nak92, PPP^{+}96, Spr93, SL97b, Sto96, TT90, TNP97, TD94, VH93, Vet95, WV95, dNSR^{+}96]. Section [BRA^{+}99]. Sector [BKL99, WASSM97, ASSWM97, CBG^{+}99, FP96b, MACR98, NSC^{+}92, Pak95, WTM^{+}96]. sedentary [HAA94, MRS96].

Sediment [AFC96, EM93, HBS98, HN97b, KBWC90, OSZ95, RGS95, Wai90, And96, Bel95, Ber90, BNW94, BSG^{+}97, BCG94, BvKvRB96, Cam91, CSR^{+}92b, CRCS96, CLM^{+}93, CG99a, CJ93b, CPB96, CP96, DBHM99, DFJ^{+}93, DGHT^{+}98, DD99, ER95a, EAC^{+}97, FH96, FGP98, FGKP95, GB95, GAD^{+}92, GH94b, GSBB97, GB97b, GCLS94, GF98, GP91, HB91a, HHB93, HGS94, Hei95, HdWS91, HK94, Hsi95, HG92b, JFS92, KTH98, Kar92, KM96, Kro91, Kur99, LH98, LS96a, LD95b, LJS90, LF97, LSP99, MFF97, MSK95, MCSV95, MKYI98, MDRW99, MJ92a, NL94, NG96, NJSS91, Nov90, PBVB99, PNB94, PLP92, RDM91, RSD^{+}91, RT92, RBG94a, RS92, RK99, RCN95, RTRP^{+}98, SNC^{+}98, SDC99, Sch96, SOSV95, SFF98, Sim92, SGW94, SCD^{+}90, SGR92, SPS98, SF94, SEGP91, TDBH94, Tha92a, Tha92b, TL94]. sediment [TB98, TBY98, TUN99, TFFS90, TTP^{+}95, Vid94, WSF99, WUA^{+}99, WHG99b, Wil92, Wol92a, ZHF96, vDvRK93]. sediment- [GF98]. sediment-associated [LF97]. sediment-bound [WSF99]. sediment-derived [LS96a]. sediment-elutriates [Kar92]. Sediment-nutrient [EM93]. Sediment-to-water [Wai90]. Sediment-water [KBWC90, RGS95, CPB96, TDBH94, vDvRK93]. sediment/water [Sim92]. Sedimentary [BLR94, CM99, EM93, LGV^{+}99a, LGV^{+}99b, MR94, SO91]. Sedimentation [CBG^{+}99, BMP^{+}95, BRN^{+}95, FS90, GF95, GGB94, HGS94, JLS90, LA98, LHK^{+}93, NH95b, OKL^{+}96, PF91a, RHB96, Rog90, Vos91, WVMR90]. sedimenting [Nov90]. sediments [AO91, AV98, AJST98, Ber90, BHHL96, Blo91, BL94c, BL96, BD98, BM90a, BM90b, BR91b, BCv^{+}92, CK90, CDHD92, Cha92, CJ95, Eno93, Far98, FTGW90, FKB92, FHZ96, GPB96, GSB98, GHHC98, GR91, HG99, HPF92, HdWS91, HK92, Hol96, HBvD92, HNvB94, HG92a, Hüt90, HP98, JNCS89, JKL90, JFS90, JBB90, Kam94b, KVK98, KGHTN99, KLEH91, KTGC99, KKH^{+}94, KLJ94, Lai92, LM91b, LGED98, LMS^{+}93, LKvRH96, LEHvR96, MKN^{+}96, NBW94, NT96c, NMR97, Por94, RBA98, Rel96a, Rel96b, RMTB^{+}95, RLC92, RT93, SB98, SDC90, SU93, SBH^{+}95, SPPL92, SNR^{+}96, SCCS98, SH92b, TNSM98, TGSN99, THKN98, VVT97, WH96, YB92, ZB94, vDK90, vDvRK93, vCRS92]. see [TK95b]. seed [HS91b]. seeding [BMH91]. seedling [DGHT^{+}98]. seedlings [SK99, ZA96]. seen [TK95b]. seep [DAB^{+}91, DBN^{+}94, JT93, LC95, ODS^{+}96, SHO96]. seepage [AGW93, BDNS99]. seeps [JAB^{+}92]. segmented [Sal95]. seine [CBL96a, Con94, TDW96]. selected [HKR96, Pro92, WTF96]. selecting [MDD95]. Selection [HB94a, MF96, And97, BBS93, BvB94, CT97, EG97, FAED97, FIA93, FSGG96, GD90, GA97b, HBS98, HTCW96, Iri90, JH97, JDLL95, JC94, KS95d, LM98a, LMO99, LB96b, LBK96, MJ97a, MHRV^{+}99, MBR94, MS93c, Pen90b, RK95b, SGP99, SJC99, SK95c, TT90, Tee99, TN90, Tru97, VR98, WG97, ZGB90]. Selective [Har99, LADS98, MVdM^{+}99, Poo94, EG95, GS99b, HM97d, LCL95, SMF91, WFH99]. selectivity [Atk96, BDdWC96, Bus97, HH94a, HHP95, MW94, OM99a, Per91, PAR98]. selenium [BC93, FR91, NL96]. Self [Hsi97, BMH91, FL90, FAP92].

Self-fertilization [Hsi97]. self-seeding [BMH91]. self-thinning [FL90, FAP92]. selmis [Tha92a]. SEM [BKWML93]. Seme [TD92]. Semele [UC94]. semelparous [ORD98]. semi [LJM99, MMM^{+}95, RDPB99, Tho91, Cow91]. semi-automatic [RDPB99, Tho91]. Semi-cossyphus [Cow91]. semi-enclosed [LJM99, MMM^{+}95]. Semibalanus [BGBS91, BGW96, HT96, HB94a, NFM98]. semilunar [ZAN99]. semisulcatus [CLD99, HVL96, LWK^{+}95, VHHS96]. semiterrestrial [AHMd90]. Sendai [FII95]. senescent [OB93]. senhousia [Cro98, Reu98b, YOH98]. senile [AS94b, AS95]. sensing [MGEC97, TNKMK99, WWGK95]. sensitive [Reg92, UKR^{+}91]. Sensitivity [DS93b, DD99, HM97c]. sensory [VKF^{+}98]. sentinel [SKM95]. Separating [CHC94, BL94b, BS92, CCGM93, WCLB99]. separation [AGGKA96]. separations [VLKC94]. Sepia [RFNBHG94]. septemspinosa [Weh91, WA95]. Septifer [WD99]. sequences [BSP^{+}99, FD97, TFF94]. sequencing [BWHC91, WBC91]. sequestered [WL97]. sequestering [KB91]. sequestration [CHFL95]. seriation [CWB93a]. sericea [TCB93, TCRB93]. series [BM99a, BM95a, IHK^{+}98, JE95, KSE96, KHTP99, LIPF95, MC91a, NKH^{+}99, RCS^{+}96]. serpulid [BHQ99, BH97b, BQKC97, HQ99, Mar91]. serra [Coc90]. serranid [She95, She96]. Serranidae [Joh99, Sam97, Zel97, Zel98]. serrata [CC96]. serratum [CCD94]. serratus [Arr93, HHN90, TR91, TRPS94]. Sesarma [AHMd90, Kyo92, Lee97]. sesarmine [Lee97]. sesquipedale [ASL99, DF95, San94c, San95]. sessile [BC99, Nan96, WW96, MD98b]. Seston [BM94c, AA92, ABY97, HJH^{+}99, JFP95, KE95, KY97, LMT96, LBJ^{+}98, MN92, NT95, NW97b, PMMH95, TGvJW95]. set [KBN^{+}96]. setae [PL99a]. setiferus [RMG^{+}98]. Seto [NSHF93, NSH95]. sets [BM95a]. setting [DY92, FB98]. settled [EE98, JPC97, Jer90, LBH91, RS95, RH97, Row90, TB95]. Settlement [BW93, CMRF93, CMRF94, DBvA91, ESDK94, Gün91, HBM93, JSOG91, JE95, MADM98, O’C93a, ROI93, SF99, Tom91, WW96, lYALM93, lYAML93, Asa91, BCC92, BGW96, BBS93, BPL95, BQKC97, CH95, DY92, Egg95, FAED97, Fre97, GQ96, HT96, Hol90, HS96, HS97c, Hur92, JBWH97, JWKH98, JBK99, Ken94a, Ker94, Ker96, LB95, LKC98, LQ97, MZ99, MSC95, MY96, MU98, MYYT96, MJ91, MM93, McC94d, MM95a, MM97a, ML92, ME97, MBB95, NG99, NB96a, NFM98, PWS98, PGM^{+}98, Pin94, QRSC99, RS99b, Ris97, SDER96, SMW^{+}93, SP99b, Sto96, TB95, TB97, VHVB^{+}98, WMB99, WRBF97, WCT95, WM97, WBLM95, WA95, WB99, YBG^{+}98, ZMD95, vTSD97]. settlement-marks [WM97]. settlement-stage [SMW^{+}93]. settling [ABN98, AJL93, SGP99, SK98c]. seven [Gru94]. sevenspine [WA95]. several [BDZB98, GRSV98, GT98, PM90a, TB92]. severe [ASF98]. severity [WC93]. Severn [RBN99]. sewage [Gra96b, Gri94, LTM90, LFG98, MR91, RDM91]. Sex [CR98, HBBM96, ST99, ZHBH90, Ack97, Mac91, MTB^{+}99, SPB^{+}99b].

Sexual [CZG95, Szm91, GCKW95, HV91, HH95, LMO99, MACR98, RT98, WH92, ZGB90]. seychellensis [GS94a]. SGWD [Sim92]. shad [Lim95, MCFB97]. shade [HLF^{+}91, Tak99]. shading [FK95b, Gla99]. shag [WFHM99]. shags [BRLM90, VF99]. Shallow [RHP93, AJS95, AAWD97, AMS99, ASCW90, AFC96, AU93, AD94, BCBC98, CGP97, Con94, DKW96, DFJ^{+}93, DB96b, GPB96, JSJMH90, KD99, KZK^{+}95, KR98a, KDMR97, KCI98, LDKR93, MC95, McC94b, MB97a, MMR99, MH94a, NJSS91, OIO^{+}95, OL99, OFR94, Ole95, PBVB99, PR92, PLB95, PF91a, RDA97, SRW95a, TL94, TCB93, TCRB93, TS90b, WB92a, YY91, YTF97]. shallow-water [AAWD97, ASCW90, KZK^{+}95, MC95, McC94b, MB97a, NJSS91, PBVB99, SRW95a, WB92a, YY91]. shape [MPB97, MJ92b, OBL97]. shaped [Kin93, NAS95, RHB96]. shark [WCL97]. sharks [CSC92, STBR99, SM97a]. shear [BF99]. shearwater [GMF98, LMT99, MO97]. shearwaters [HCH^{+}96, WC98]. shed [DL93b]. Shelf [DLFL96, ANW96, AAW98, AML95, Bia91, Bia92, CSB94, Fab95, FII95, GGDLE96, Gra94, Gra96b, GDW90, HSLV94, HBC^{+}91, JT93, KE97, KD99, LFR^{+}97, LKvRH96, LEHvR96, McC96b, MLDM95, MLDM99, OG95, PLB95, PS97, Pir92, PD90, PD94, RS91, RSS94, RGE^{+}95, SO96a, SH90a, SR90a, SRSH99, Sha98, SS99b, SGMS99, SH95b, TDBH94, TWWE95, VAV99, WR97, WHG99b, MF96]. shelf-slope [SO96a]. Shelikof [BM94a, BBK91, CBI91]. shell [Bar99, Bou90a, Bou90b, DS93a, DBV98, DMM99, FAED97, HV91, KSH97, LC97a, MIA90, RR92, Sal95, SGP99, WG97, SR96]. shell-weight [MIA90, SKM95]. shellfish [BLCA90, BLC91, GSBB97, HSBH96, HTLK99, KOSS90, KB91, RM94, SOS^{+}92, SPZ^{+}93, TPHC91]. shells [EGC95]. Shelter [ELMCC90, And97, BBS93, ELG97, Iri90, MLES94, MPvM98]. shelter-size [ELG97]. sheltered [JHN99, THR^{+}98]. shelters [AADSCC97]. Shetland [BP98, GGF99, SPCF97]. Shields [Raw96]. shift [AP99b, LF91, WLB^{+}99]. shifting [Hol97]. shifts [Cla96, Dav94, Egg95, JKFB95, MM97a, PKL97, UMD92]. Shikoku [AM93]. ship [GH94a, Hal98]. ship-following [GH94a]. shipboard [BTYO94]. shipping [RQB98b]. ships [LW94]. ships-of-opportunity [LW94]. shoal [CHDW92]. shoalgrass [BGC94]. shoaling [RO98]. shock [KWHM^{+}96, KBA^{+}97, MBL^{+}98, MGD91]. shoot [DSJ90a, Dur94, HS99b, KS98, RCG94, Wor95]. shore [AAWD97, Arr93, BCNC96, BCMC99, BKH99, CMRF93, CBRD92, DL95, Dye93, EW91, GJVG96, GDW90, HS95, HS96, HS98c, HN93, JBH98, KW98, Kin95a, LH98, MCC96a, MM91a, MHS94, MS94b, MPvM98, Per90, Pin94, RN93, SAA99, Tak95, Tak96, Tak99, TH94a, TR95c, Tok95, UK98, WJD93, ZN96a, ZN96b, ZAN99]. Shorebird [IVM^{+}98, LL98]. shorebirds [Sew96]. shores [BH98, CV98, JBHH97, KW97a, NW97a, SBC^{+}98, WMG^{+}94, WG97]. Short [Dav94, GGB94, JAOS90, LDKR93, MS94a, Mil98, RGS90, RW96, SK98b, Sch99, SGMS99, SDB94, Tis98b, VM95, AdN91, ASRP^{+}96a, BGR^{+}96, DF95, FRK92, GHMC96, HCH^{+}96, JCMW91, KS98, LD95b, NH98, QGQ97, RÁSP99, San94a, Ste94, WC98, WUA^{+}99, ASRP^{+}96b]. Short-distance [Mil98]. short-shoot [KS98]. short-tailed [HCH^{+}96, WC98]. Short-term [Dav94, GGB94, JAOS90, LDKR93, MS94a, RGS90, RW96, SK98b, Sch99, SDB94, VM95, AdN91, BGR^{+}96, DF95, FRK92, GHMC96, JCMW91, LD95b, NH98, QGQ97, San94a, Ste94, WUA^{+}99]. short-time-scale [ASRP^{+}96a, RÁSP99, ASRP^{+}96b]. shortage [RM97a]. show [BMH91]. showing [CWB93a]. shown [DHB^{+}98, FD97, RQB98a]. shrimp [Beu92, BFD91, BKB92, Bur94, CF93a, CA99b, DF92, Dem95, FAED97, GF98, GS91b, LCTT92, MC97a, NA95, OTF95, PM90b, PSB^{+}97, RE96, SMT97, SlRR94, SSV97, SF99, WLL98, WA95, ZGB90, dT94a]. shrimps [Car93, CS97, Duf92, GPS93, HM91, HS98b, PDB^{+}97, PDS97, PD90, SB96]. shrinkage [NSC^{+}92, SGP94]. Si [CJ95, Som94]. sibling [CCGM93, GVGG98]. Siboglinum [DBN^{+}94, Sch91]. side [NG99]. Siderastrea [GT98]. siderea [GT98]. sieboldi [CRLP94]. Sieve [JLF95]. Siganus [PNS90]. signal [KSV99, LMK93, SB99]. signalling [LMO99]. signals [KV99]. signatum [RH90c]. signature [BBCB99]. signatures [BGCM95, IFL97, TJST98, WCLB99]. Significance [BNK^{+}97, CPSA96, JJTS97, RLC92, SCA92, TFFA99, WB94, BGC91, BL97, CA90, CM98b, CPB96, DL91b, DFS94, GCF96, LBV^{+}99, MU99, PH91a, SH95a, SL94, SM99b, TM94, UMT^{+}91, VGM94, Rel96b]. significantly [DA93]. Signy [PBVB99]. Silica [BK96b, BVL98, CSS93, FKB92, LNQT93]. Silicate [EA92, EJ97, PRM^{+}96a, PRM^{+}96b, TDBH94]. Siliceous [GS94a]. Silicic [ND96, DQT^{+}97, NB90]. Silicon [BOC90, HHN^{+}96, LW91, NT92, ND96, RDT^{+}94]. siliculosus [Mül91]. siliqua [GCM98]. Sillaginodes [JWKH98]. Sillago [CGFC95]. siltation [BBFD98]. silver [AHY^{+}99, FBL95, SM97b]. silverfish [WHK97]. similarities [VVT97]. Similarity [Kam94a]. similis [DS93b]. simillimus [WSCT96]. Simple [HPSG94, SR95, DPO91, FM94, GF95, GBV96, RPN95]. simplex [Gen92, GGDLE96, Góm95, Lav94]. simulate [Vis97]. Simulated [Por94, BD94, CC90, CJ95, DON^{+}95, GT92, PCLÅ97, SO91, SM91, WYY^{+}99]. Simulation [AD93, PML93, BBRBK94, KDC99, Mol97]. simulations [BMHH94, JBH98, LGMD98]. Simultaneous [AAD91, KS95d, RDPB99, TK91, PPR94, SR95]. sinensis [Ang91c, GSC95]. Singapore [HTLK99]. single [CH97b, GGKJ^{+}98, JLD^{+}96, MBL^{+}98, MJ92b, RPT95, SB97b]. single-celled [MJ92b]. single-strand [MBL^{+}98]. Sink [Fra92, DL94a]. Sinking [WFTH97, BMTF92, BH97b, Dei90, ES99, FD98, HKA96, LF94, MLH96, PJ99, PGAJ99, WH92]. sinks [JCP93, ONH^{+}97]. Sinonovacula [WX97]. Sinularia [Fab95]. sinuosa [TB99]. sister [PMETM96]. site [ALB^{+}96, BSH^{+}94, BSS92, CHS^{+}95, DvdM92, GPS93, GAD^{+}92, GBP^{+}97, HS91c, LRV94, PBVB99, SFWD90, Sha98, TWWE95, Yam97b]. sites [AHY^{+}99, BB97a, CSSPV98, Ker90, LBM^{+}99, RATA^{+}99, RA92, SH95b]. situ [DP91, LGB91, LS91, MM91a, VMHH90, ADBH99, ADF96, AH94, BTYO94, BP96a, Bro98, BBL93, DT94b, ER95b, FRK92, GSB98, GGC99, GGH99, HS97a, HS97b, KKT96b, LPI^{+}95, LB99a, LCE97, LBM^{+}99, MPS96, MG99, MP96c, MA96, MAdG^{+}96, NG96, OJ97, PPMF97, PPW96, QC94, RGLS99, RCTH99, RCG98, SPBS93, SFS95, SAA99, TGvJW95, Tis98a, VWK96, VD95, Wah97, vBV95]. situation [NLR^{+}93, SKM^{+}97]. situations [ICG96]. six [EPL94, GGB^{+}99, HS98a, LT93, Moe96a, TJ90b, VHHS96, VHH^{+}98, ZB92]. sixteen [CDG97]. Size [AH96, AH98, DCM90, EG95, FVDL95, HPSJ95, HHW99, IP94, JC95, KL91a, KL91b, Nak94, PLG^{+}96, RFN98, RN98, San95, TB95, Åbe90, Ang91a, AS94b, BM98, BN95, BSH^{+}94, BH92a, BCL93, BL94b, BG93, BvKvRB96, BSH^{+}91a, BSH^{+}91b, Bus96, CC98, CK99, DP91, DD99, Dye93, ELG97, EG97, Fer99, FC90, FCP98, GdGMD95, GMO95, GMG^{+}92, GB93, GDW90, GK96b, GP98, GSMM96, HTCP96, HS97b, Ito97, JJTS97, JH97, JLF95, Joh99, JMP90, JC94, KKL91, Ker96, KD96, KRJL96, KFS91, KW99b, LS92, LWL98, LGH^{+}93, LM98b, Lew91, LBK96, Mac91, MD91, MSD98, MB98, MD98b, MD93, Mau92, MM93, McC94d, MM95a, McG94, McL96, MF98, MLES94, MPvM98, ML91c, MSND^{+}99, Mun92, MJ92b, NC90, NT95, OBL97, OC93b, PA94, Pen90a, PS95, PCBA94, Per91]. size [PAR98, PWM95, PPO^{+}99, PF92c, QK93, RTRF90, RE96, SSR99, San94c, SOW97, SW91, SDFJv91, SH90b, SMB98, Sut92, SK95c, TMKL97, TJRG93, VM98a, VDM^{+}97, WF97, WT98, WBP^{+}92, WP94, WTM^{+}96, YNYK94, YKL98b, van94a]. size- [FCP98]. size-at-age [BCL93, RE96]. size-classes [SDFJv91]. size-dependence [GdGMD95]. Size-dependent [FVDL95, HPSJ95, Nak94, Ang91a, JC94, LM98b, Lew91]. Size-differential [PLG^{+}96]. Size-fractionated [HHW99, RFN98, RN98, GMG^{+}92, GDW90, KFS91, LGH^{+}93, RTRF90]. size-frequency [BM98]. Size-related [AH96, AH98, WBP^{+}92]. Size-selective [EG95]. size-selectivity [Per91]. Size-specific [DCM90, TJRG93]. size-spectra [SSR99]. Size-structured [JC95]. sized [GS93, Jua92, ZB92]. sizes [GS99a, MRL96, MG91b, WHG^{+}99a]. Skagerrak [BKK^{+}93, DBN^{+}94, FSG^{+}97, FSG^{+}98, HRÆ94, Jos90, KL91a, KKK^{+}90, LSG^{+}99, RDE^{+}90a, RDE^{+}90b, SFWD90]. Skagerrak/Kattegat [HRÆ94]. Skeletal [KL91c, LCBD94, GT98, JLGMP97, RVGC^{+}99, RS91]. skeletogenesis [MGYV97]. Skeleton [PMB99, LGP95]. Skeletonema [BR97, HFN91, TT98]. skeletons [LGH95a]. skin [HSMS99]. skipjack [YD90]. skua [PCT^{+}97]. skuas [ER90, PCT^{+}97]. slabberi [WW90]. slimy [SB96]. slope [AML95, Bia91, Bia92, Car93, DSG^{+}99, EJ93, FH96, FII95, HSLV94, HGS95, HBC^{+}91, KKW97, MJV92, MP96b, MSM^{+}98, PH95, SO96a, SLB^{+}92, TDBH94, TGvJW95, VAV99, YST^{+}97, MRS96]. Slow [Fab95, VD95]. slow-growing [VD95]. sludge [MR91, RDM91]. slug [Tro92, Tro93]. Small [GBP^{+}97, Sch90, BJB95, BS95c, Bur94, Con95, FPGJ99, Fow90b, GT92, HVL96, Iri90, JH97, Jua92, Kin92, LWK^{+}95, ML91a, MMM^{+}95, MJ97a, Mee99, MN99, MLDM95, PG94, PR93b, Tay98, Tho92, Tis98b, UC94, UCV98, US95, US98, WB99, HV91]. Small-scale [GBP^{+}97, Sch90, ML91a, MJ97a, PG94, WB99]. small-sized [Jua92]. smaller [BM95a]. smelt [DD90, Sep94]. smelter [ONR98, WH96]. snail [Cha95, DMM99, Gos97, Laf93, McC94a, TJRG93, Tru97, Wah96]. snails [GC95b, MEK99, Wah96, Wah97]. snapper [MMBT97, OC93b]. snow [BH92c, BH92d, DWSA98, HKA96, HLH96, KH92a, LS96b, LHH96, MNH92, PW94, PGAJ99, SMUS^{+}99, SR93b, SW97, SdC97, SAA90]. social [IMB99, RO97]. socially [BvB94]. sodium [FT95]. Soft [GAD^{+}92, SNC^{+}98, SGW94, Amb93, BDZB98, Blo91, BDR^{+}96, CP96, DMM99, Fab95, FK95a, FSCE92, HPF92, JDSB99, JFS90, JFS92, KKL^{+}98, KO92, LC97a, MSC95, Mic96, MJ97b, Ols93, ONR98, PBVB99, SMB97, SHM^{+}97, SGP99, SE91, TN98, TTP^{+}95]. soft-bottom [Amb93, Blo91, BDR^{+}96, FSCE92, HPF92, JDSB99, KO92, MJ97b, Ols93]. Soft-sediment [SNC^{+}98, JFS92, PBVB99, TTP^{+}95]. soft-shell [DMM99, SGP99]. soil [NDRP93, PD93]. Solar [BDSL94, AKL^{+}99, ABB^{+}97, BSS92, DB96b, Dun99, FSA^{+}97, HMVHH95, JPA^{+}96, LB96a, SL97b, VSSM98]. sole [DvdTCC97, LT97, RFGF92]. Solea [DvCd94, DvdTCC97, LT97, RBB^{+}91, VDVC94]. Solemya [KGC92]. Solemyidae [BL97]. Solent [HSW99]. solid [BKS92]. solida [UC94]. solidissima [BDB97]. solitary [CGCLH99, HSH94, HB95]. soluble [CV91, GAB90a, GAB90b, Pak92]. Solute [HG92b, GFH96, GF99, HG92a]. solution [CBDC96]. solutions [Blo91, Hil91a]. Somatic [SG96, HV91, Rv92, UTB^{+}95]. Some [DH90, LS92, SPS^{+}96, AO91, BMFR97, BS96b, Cha91, CSM99, EBB95, HP90, LCE97, McG97b, PMETM96, PWM95, RSB93, SDA^{+}92e, WS92a]. sooty [LMT99, MO97]. sopod [EB92]. sordidus [Coc90]. sorting [LC98, WBP^{+}92]. Sorus [MC93]. Sorus-specific [MC93]. Sound [ARS93, ADG^{+}95, CC97b, IFL97, JFP95, KCV94, LCLC96, Nor94, PMS94, RKS97, RTRP^{+}98, RNH99, SMW^{+}93, SBP92, SBP93, Tan98a, Tan98b, BA93, CHDW92, CC91b, MSD92]. sounder [MA99c]. Sounds [HJH^{+}99]. Source [McC92, Bar97a, DL94a, Kam94a, MHBW90, MLH96, MH96b, RRGB96, RSAD98, SKT97, VWM90a, VWM90b]. Sources [BRL97, CG99b, CMBM99, HAR95, CM90a, CF94, FW91b, GC95b, JCP93, JKC^{+}93, KSRB99, KZK^{+}95, LM91b, ONM93, PWG^{+}99, Rie98, RSN^{+}99, RNHO98, Sch91, Sim91, SDH95]. South [AU93, CC91a, Cha95, CBL96a, CBL96b, ER90, FK95b, GBP^{+}97, HHE92, KUK91, Moo94, PLFSW97, PLM92, Sca93, SCCS98, Tok95, WH96, WPTB96, CGP97, DL93b, EW91, HWH98, JMC93, O’B94b, She98, YB92, ASP^{+}96, AS97, AWH^{+}99, BP97b, BA91, BSF95, BW93, BW96, BWM97, BBCB99, CWHA99, DSW^{+}91, DPM96, DCBG99, Dye98, EM93, GTYR97, Han92, HTLK99, JW96, KWEG94, KN99, Las98, Moo94, ØW93, PP96, PMMH95, PWPS95, RWCM99, ROANM98, TNKMK99, WTR99, WMD99, Wat98, WPTB96]. south-central [JMC93]. south-east [CGP97, HTLK99]. south-western [HWH98]. Southampton [IP94]. southbound [KNI95]. southeast [BE97, BP98, KHB91, KO92, MS91]. Southeastern [RZLB96, BFJ^{+}93, BK92, CCSN93, CPH^{+}96, DFPG91, FPG95, GH94b, GH94c, GFPD92, Gra91b, KKW97, LBH91, MLTD96, NL94, Pir92, RVv92, RZWY96, RLT^{+}94, SSHM91, VAV99, dG91a]. Southern [KID91, PDBA90, Per91, PLM92, RAS97, TBAB91, AM93, AML95, BA91, BD98, BCBR97, CNGG96, CL97, CW99a, DJY90, DLJ91, Dye93, EY90, EMR^{+}97, FTS94, Fow90a, GDvS98, Gib94, GGB^{+}99, GB97a, GB97b, GW91, GP90, HSLV94, Hil97, HM97b, JBB92, JD90, JP93a, JJNR95, KJGC96, KNI95, KDMR97, LC97b, MMR99, MADM98, NCGG96, OIO^{+}95, OB94c, PLP^{+}93, PML93, Pin94, PBHMI98, RS99a, Reu98b, RC91a, Rie93, RS92, SH92a, SK95c, UNPH93, WW90, WMG^{+}94, Whi94, WtD90, WOM^{+}97, YD90, YBLL96, vG91, vdBTvDR96, Arr94, ASSWM97, BK91, BP96a, BKL99, BBCB99, CBG^{+}99, CA95, FC90, FP96b, FPP^{+}97, LPF94, LLP94, LRV94, MO95, NSC^{+}92, Pak95, PPM96, PM92a, PGAJ99, San94b, SKvB98, SPBS93, SDFJv91, TD91, TvLdJ^{+}98, VMHH90, Vet96a, Vet96b, WASSM97, WWC^{+}99]. Southern [WWGK95, WTM^{+}96, dBN^{+}90, vLTW^{+}98, vdVW99]. southwest [BFS^{+}98, RSH92, TNKMK99, TWS^{+}95]. southwestern [AM93, IBB97]. sp [ADA98, BPS98, BEG97, FSdM^{+}92, FFH96, GT92, HFW99, IG96, JH97, LB96b, LGdC^{+}90, MGDT90, NP90, PLF90, PMS94, QC94, SFF98, Sv93b, SDG95, SD96, Str91, SHW^{+}91, Tsu90, TFFS90, UYMH95, vBV90a, vBV90b]. sp. [GHTH94, DWG98]. space [BGC91, CHC94, FL90, FAP92, GWSG99, MS93c, NC90, SB97a, SL97a, TR95c, WR97]. space-time [SL97a]. spaced [ECF^{+}97]. Spacing [LY97]. Spain [ASRP^{+}96b, ASRP^{+}96a, FP91b, FGG95, NCGG96, VM95, Arr93, BQR99, CMRF93, CCV^{+}98, GD96, Pre91, RDA97, RQB98b, TFFA99, UCV98, VF99, ZF97, ZF99]. spallanzanii [AW97]. Spanish [CDM^{+}96, MDC^{+}96]. spanner [Ken92]. sparid [VHVB^{+}98]. Sparidae [SB97a]. Sparisoma [BKB94, BvB94, BBB^{+}94, BvKvRB96, TT90, vRBVB95, vRV96]. Spartina [AWdS92, AWH93, AWHH94, ATNW97, Blu93, BI99, CNP95, HHSvS96, KNK90, MW97, MH90, Pak92, Pez91, PD93, SCC98, Tur93, TZ99, dG91a]. spat [APCL^{+}93, HBM93, PGM^{+}98, RBG94a, RBG94b]. Spatial [AAWD97, ARDL97, BLDF95b, BLDF95a, BPC98, BRW95, Cha91, CPK97, CCSN93, CWS^{+}93, CSM99, FMFH97, Fow90a, FSG^{+}97, Gre96, GF91, Gru94, HV91, Hsi95, HS98c, JJ99, JRMN^{+}97, JBHH97, JFS90, KKW^{+}93, KHK^{+}99, KTW^{+}97, LBH91, MYYT96, MHS94, MS94b, MS96a, MG94, MPS95, Nak92, Pin94, PWM95, PWP^{+}97, PH95, RH98, RNP^{+}98, RT98, SPB95, SLB^{+}92, SV93a, SP99b, SH93, SR97, SGM98, SF91, TRS91, TPW^{+}93, THKN98, VHVB^{+}98, WHG99b, WM97, WBLM95, ÅP97a, AWD95, AGP98, Adj97, ASA96, AV98, Arr94, AH96, AH98, BCNC96, BCC96, BCMC99, BGBS91, BPL95, BWL^{+}99, Cha98, CBL96a, CB97, DG97, DG96, EAEP98, ELMR98, FB91, Gra91b, GBB^{+}98, HB97, HlS90, HTCP96, HNvB94, HCI93, Hul99, HJHG98, JLF95, JBWH97, KM92, LD94a, LAJ95, MZ99]. spatial [MCF90, MRPC98, MGC91, ML92, Mil98, MM95b, MLTD96, O’D98, PF95, RII93, RÁSP99, SMT97, SNC^{+}98, She96, SNJV92, SM97c, TF98, UMTM98, ASRP^{+}96a, MDC98, PF96]. spatially [BWRH99, HHM96]. Spatio [DBV98, KLN^{+}90, LSG^{+}99, CC98, GBGH97, KSE96, KID91, LP95]. Spatio-temporal [DBV98, KLN^{+}90, LSG^{+}99, CC98, GBGH97, KSE96, KID91, LP95]. Spatiotemporal [ELMR98, GST^{+}91, Zho98]. spatuligera [RZLB96]. spawned [MC91c]. spawner [PSL96]. Spawning [AM93, CW99a, CM97, HH91b, Pal91, ZHM91, Zel98, AHY^{+}99, BTKK99, BSH^{+}91b, Cla90, DL93a, Dem95, DCBC91, HSK^{+}93, HLP99b, KMG98, MCFB97, NW97c, PG90, PM91b, RHH97, RSB93, Sak97, Sam97, SHT91, SGM98, Sty97, TP95, Van94b, WCP^{+}98, Yam97a, Yam97b, HOM95]. spawns [MSP^{+}97]. special [BS95c, Dai90, KI97, MvdZH^{+}97, MNH92, SC92b, Tan98a, Tan98b]. specialisation [MJC97]. Specialist [RH90c, BPS98, CHFL95, Tro94]. specialization [GPS93]. Speciation [BKWML93, WMPT94, MTP92, MSCTBE96, NFT92]. Species [ACC90, AALT99, BBFD98, CC90, CC91b, HFN91, LF91, Mil95a, ÓMvdB99, Aer98, AB96a, BC91, BG97, BDZB98, BF95b, BS92, BL95, BDA^{+}99, BRW95, BBT^{+}93, CUM98, COD^{+}93, CDME97, CT97, Cla96, CGP97, CBT98, CGFC95, CCGM93, CES^{+}94, DS96, DES92, DCM90, EK95, EPL94, ELH92, FSG^{+}97, FPPM96, GVGG98, Geo94, GAR93, Gra91b, Gra94, GK96b, Gru94, GRSV98, HS98a, HN97a, HP90, HKR96, HVFHH92, HNC92, Hul97, JMD95, JMD^{+}99, Kar95, KG94, KRJL96, LM98a, LLB92, LT93, LD94a, LS96b, LMM98, LH96, LCC97, LE90, LA97b, LF97, MCC96a, MFF97, MWA93, MDD95, MCED96, Mau92, MC98, MRMR92, MP92, MMD92, MW97, Nev99, NA95, ØW93, PDD^{+}99, PMMH95, PB96b, PAR98, PPO^{+}99, PFMC95, Pur91, RVT^{+}97, RM97a, RTCH96, RMGB^{+}94, SJML99, SMG97]. species [Sch90, SDB96, SW98b, SJCL93, SGPAA95, Sv93b, SPCF97, SPDM91, TK95a, THS94, Tis98b, TB92, Urb94, VEW99, Van92b, VS92, VDM90, Ven98, VPGV94, WH96, WCP^{+}98, WSMM99, WBC91, Wöh96, WB99, YAG92, YY91, YKL98a, ZHBH90, ZB92, vG91, vCRS92, Gra94]. species-sequencing [WBC91]. Species-specific [ACC90, CC90, CC91b, HFN91, LCC97, MCED96, Nev99, SPDM91]. Specific [BGCM95, AS99a, ACC90, BAM^{+}99, CC90, CC91b, CC97a, DCM90, FML^{+}95, GW93, GCC95, HFN91, HV91, HVFHH92, HBP^{+}93, HS97b, KRJL96, LCC97, Lip95, LSB95, MCED96, MC93, MHvB97, NC95, Nev99, ÓMvdB99, PCC^{+}97, SD94, SS99a, Sos96, SPDM91, TJRG93, VWS92, VPGV94, YTF97, DQT^{+}97]. specificity [Yam97b]. specimens [HKM92]. spectra [FM95a, FQ94, MG96b, MB98, Mau92, MKV95, Mun92, NT95, RPM^{+}94, SSR99, ZH97]. Spectral [Maz95, JSV92, JNJP94, KPS92, MG96c, MP95, SPS^{+}91]. spectrally [WPY98]. spectroscopy [HHYD92]. spectrum [CW93, DD99, KPS97, LL96, MJSDF91]. speed [RO97, SlRR94, TW98, vdL95]. speeds [LCE97]. Spencer [WH96]. sperm [JWL96, JW96, KWD93, SPB^{+}99a, WCP^{+}98]. Spermonde [SAvD^{+}96]. Sphaeroma [CCD94]. Spheniscus [WSP^{+}95]. SPI [RK99]. spicules [GS94a, CP95a]. Spiekeroog [Hil97]. spill [DF94, ER90, GBJ94, JDSB99, SS95, vTSD97]. spinifera [Tan98b]. spinipes [WMH96]. spinuligerum [KW91]. spinulosus [ABN98, Oes90, OV92]. spiny [AM91, AADSCC97, BHH^{+}95, IMB99, ELMCC90, ELG97, ELMR98, Mac91, MLES94]. spionid [LW95]. Spionidae [MRP^{+}99, San94b, Zaj91a, Zaj91b]. spiralis [MM90]. Spirobranchus [Mar91]. Spisula [BDB97]. Spitsbergen [ANW96]. split [RN93]. splitnose [SDBW95]. splitting [Tho91]. Sponge [Aer98, DW97, AvS97, BBR^{+}93, BPS98, BWL^{+}99, BHH^{+}95, CBA^{+}99, DBvA91, Duf92, KWHM^{+}96, KBA^{+}97, MU99, MBL^{+}98, MWB^{+}99, MSCTBE96, PPW96, RCG99b, SC90, TR95b, UKR^{+}91, UTB^{+}95, VFMFBE96, WKH^{+}98, WC90]. sponge-dwelling [Duf92]. Sponge-mediated [DW97]. Sponge/coral [Aer98, AvS97]. sponges [BGT91, Bar95, CNGG96, CP95a, Mal98, MY98, MU98, PCTF95, RTJ^{+}93, RP91a, UMTM98, WBGS97]. spore [EST95]. spores [KHT93]. spot [SNC^{+}98, SMF91, SR97, SCCS98]. spp [And94a, ALOT99, AALT99, BEU99, BCC92, CR95, CGCLH99, DH96, DHM92, DFS94, EP97, Far98, GM99, GBP^{+}97, HGS94, HH92, HH96, IT96, JA91, JL94, Ken94a, KPWL93, KN99, LJ96, LK96, MC91b, MD93, Mat92, MC99, NS96, PMP^{+}96, PLF90, RH90c, SH90a, Sha98, SH98b, SC95b, WH91, vW98]. spp.in [Neu92]. sprat [AH93, Mun93, RG93]. Sprattus [Mun93, RG93]. spread [CS91]. spreading [HM97a]. Spring [CCG99, Ven98, YHGB96, BNvB90, BH92a, BMD^{+}90, BK96b, Bus96, CCG98, CCCH94, CJ95, DG92b, DL94b, FSdM^{+}92, FS90, GJC97, GB91, GCF^{+}95, GTL^{+}91, HvB91, HM94a, IHK^{+}98, JLS90, KLN^{+}90, KFS91, KR95, KW99b, KR98b, LKK92, LHK^{+}93, LHL93, MBJ94, MJV92, MC91c, ND96, NKH^{+}99, PWD^{+}91, RBD^{+}97, Rie93, RFN98, RDE^{+}90a, RDE^{+}90b, SCEG93, SD93b, SDG95, SBP92, SBP93, SCKS90, Tam95, TGSN99, WYY^{+}99, WP92, YHC^{+}99, LDKR93]. spring- [MC91c]. spring-neap [DL94b]. spring-summer [SBP92, SBP93]. Spur [FH96]. Spyridea [McG92]. squat [Roa93, RGE^{+}95, RT98]. squid [ASRC99, Big92, GTYR97, MM95b, NAS95, Tho92]. squids [CW95, CW99a]. St [CLGM96, GGB^{+}99, HV91, HD91, HB94a, OTF95, PR93a, SGL^{+}97, SK95c, TMKL97]. St. [CHDW92, EW91, FVDL95, LD93, LD94a, RMT91, RCGSN96, VBF94, VDBF96]. stability [GH94b, LS94, PDBA90, TD91, van92a]. stabilizer [CV91]. Stable [BRZ99, HM96, HSCC96, TC93, Bra93, CNP95, DRS94, DWG98, DH92, EF97, FCMB94, FQ94, GTWF98, GT98, Hob93, JRMN^{+}97, KSRB99, KPWL93, KRJL96, KRJ98, MO97, PIC97, PSB^{+}97, RRGB96, RSN^{+}99, SCC98, SM90, SE93, TTP^{+}95, VNM97, WF90]. stable-isotope [Hob93]. Stage [MTB^{+}99, MT93, NG99, BWB98, CN92b, HSB97, KLMR91, LSB95, Lop91, MACR98, MM95a, PZ91, SMW^{+}93, SLSH98, VEW99, WD94, YTF97]. Stage-dependent [MTB^{+}99]. stage-specific [LSB95, YTF97]. stages [AHT^{+}90, AD90, Ang95a, BCC94, BM91b, BSB91, CA97, DLDL93, Jeo94a, KW94, LH91, LHL93, Mye95, OW95, Pec99, SW98b, Sto97, SB97c, TLT93, TVG96, VV95]. staging [ZBW92]. stain [HP98]. stained [Flo91]. Staining [SR98a, BD91, KW91]. stand [HHSvS96, PDC97, San93b, San95]. standard [HKM92]. standing [BS95a, BSW^{+}95, CVD99, CNP95, HA90, MGEC97, SK94b]. stands [CO92]. star [DEF91, MDG98]. Starfish [Kai96, BMBD95, HK96, HKR96, HR96a, Hun93, JSOG91, Kur99, LL94, RPT95]. starry [McC92]. stars [BS94, GS99b, SG96]. start [Bus96]. starvation [BDH94, BK90, CCV^{+}98, KFF92, PMP^{+}96, RBB^{+}91, Tis98b, UGSBA94, WYY^{+}99, Weh91]. starved [GGKJ^{+}98, MHB94, RGS90]. starving [GSS93b]. state [BFF96, BKL99, ENSR90, Kep91, ML92, Pro92, Rel96a, Rel96b, SFS95, TJ90a]. state-of-the-art [Rel96a, Rel96b]. station [GGB94, VMHH90, STK99]. stations [DHH98, HNC92, SR98b, TDBH94]. Statistical [MDD95, CL97, PN99, YBG96]. statistics [O’D98]. statoliths [Big92, RFNBHG94]. Status [FHA99, Ber90, BAOT96, Cha92, DGHT^{+}98, Fly90a, Fly90b, FJR^{+}94, GTWF98, HG99, Ken92, MMR99, RHH94, TH94a, WBP^{+}92, WJD93]. steady [BFF96, TJ90a]. steady-state [BFF96, TJ90a]. stefanovi [Uib95]. Stellwagen [CBTM93]. stem [GF98]. stem-dwelling [GF98]. Stenella [MCM^{+}97]. step [CW99b]. Stephanocoenia [NB98]. Stephos [KGW93]. stepwise [RH90b]. stepwise-reduction-CVAA [RH90b]. sterna [dRAV92]. sterol [HF99]. stevensi [SCAT91]. Stichaster [LL94]. Stiffness [EL90]. Stilbonematinae [HBB^{+}99]. Still [Dei90, BC99]. Still-water [Dei90]. stimulate [OJ99]. stimulated [CR94, PNB94]. stimulating [AWdS92]. stimulation [DFS94, GWW^{+}98, PWG^{+}99]. stirred [GFH96, HG92b]. Stock [EF97, BRLM90, BS95a, CLD99, CM97, FF92b, MCF90, NW97c, SE93]. stock-recruitment [CLD99]. Stocking [TSM^{+}94]. stocks [BSW^{+}95, CVD99, GST^{+}91, HA90, PDC97, PRPM^{+}94, SK94b]. Stoichiometric [FPFR92]. stoichiometry [LBJ^{+}98]. Stomach [SPB^{+}99a, FH99, WK98]. stone [LFS90, MRF^{+}99, MYYT96]. stones [SHA^{+}95a]. stony [SCLS94, SP93]. stoplight [BvB94, BBB^{+}94]. storage [CTEF91, CLHF95, HKG93, HVK96, KM99b, NL96, Ohm96, PB96b, RGBC92]. Storm [CR94, MC91a, OSZ95, SM99a, Tru97]. storm-deposited [OSZ95]. storm-petrels [SM99a]. Storm-related [MC91a]. Storm-stimulated [CR94]. storms [ST98, TWWE95]. stout [GI99, IVM^{+}98]. straight [RPN95]. strain [SHW^{+}91]. strains [JMD^{+}99, KFF92, SWK98, SC93, VTNA91, WNJ91]. Strait [BM94a, BBK91, CBI91, HCS90, HTLK99, LHL93, Nor92, Nor94, Øre90, Øre95, HCC^{+}91, HM97b, HHW99, HCI93, KL95, TGBB97, TNKMK99, YHGB96, YGH^{+}97]. Stramonita [DS93b]. strand [MBL^{+}98]. stranded [MCM^{+}97, SPB^{+}99a]. Strategies [BG97, DBMM99, EBB95, Hal98, KWL98, PWS98, Ruz96, TDH^{+}94, VKF^{+}98]. strategy [BRNH97, Bur94, HVK96, HS90, HC93, Hoe91, MCNB98, OV92, PB97, RHH94]. stratification [BMP^{+}95, LED92, MH97, PF91a, SJÆ91, TD92, TS90c]. stratified [ABM96, BG94, BRC99, CPH91, FGS98, KR95, Mee99, NLR^{+}93, RLDM94, SL97a, SGPAA95, TJ90a]. stratified-water [BG94]. stream [QCM97, RE94b, WFH99, NB90, WPG99]. streamer [MG91b]. streamflow [Har94]. strength [Cam96, LMK93, MWA93, SDB96, vdVW99]. stress [CNGG96, DH90, DEP^{+}96, FCC91, FK95b, FFF^{+}94, HMWN97, Hoe91, Jon97b, KBA^{+}97, LLB^{+}95, Les96, MIA90, NCGG96, NB96b, QQ99, Ral98, Reg92, TvLdJ^{+}98, VHG95, WC95, WJD93, vLTW^{+}98]. stressed [AvS97]. stressors [KWHM^{+}96, RCH98]. striata [DBV98]. striatus [Egg95, LL94, SMW^{+}93]. Strictly [BMTF92]. strictum [Ekl98]. strigosus [SML97]. Strike [VTGK99, CD98]. striped [Che93, MCM^{+}97]. strobila [Put90]. Strobilidium [Mon96]. Strombidium [ADR94, Mon96]. Strombus [Dav94, RS95, SL94, SR96]. strongly [SdC97, WC96]. Strongylocentrotus [ESDK94, EDS^{+}99, KL91b, ME97, SH97b, SDER96, SCL95, SHT91, SLHC98, WH91]. strontium [CMBC98, Lim95, OINN94, TRMW95]. strontium/calcium [OINN94]. structural [BDM96, Hul97, VWD^{+}95, LHRL97]. Structure [AD94, FP96b, HPF92, HFGT94, HS95, NLR^{+}93, ODS^{+}96, Pir92, SSG96, AAWD97, AGW93, Amb93, And94b, AW95, AWVP98, BKS^{+}99, BU93, BM98, BMHD93, BH92a, Bec98, BLDF95b, BLDF95a, BSGM94, BB99, BS92, BPL95, BP92, BRN^{+}95, BG99, Cal95b, CGNH98, CPT90, CA93, CD96, CD98, CSB97, CB97, DSS^{+}98, DHH98, DCK99, DBV98, DFB96, DCBC91, FGS98, FVDL95, FII95, GB91, GW94, Gen92, Gib94, GGPGG95, GCNSNd94, GB98, HKMK98, HSH94, HTG96, HD91, HHW90, HB94a, Hol97, Hul99, IFL97, Joc95, JKFB95, JBHH97, JFS90, Jon93, KL91a, KE97, KKK^{+}90, KM92, KHW^{+}99, KKW97, KFS91, LLB92, LBJ^{+}98, LK94, LH96, LCG97, Lew97, LAJ95, LPW95, LBK90, MV98, MS93b, Mil97, MS93c, MU97, MSM^{+}98, MN99, NL94, NH95b, NR98, OM99a, OE91]. structure [Øre95, PPM94, PCM^{+}95, PPF95, PWM95, PD90, PD94, PTW^{+}91, PAM96, PM98, RBF92, RAB^{+}92, RNP^{+}98, Rie98, RVBL92, RGE^{+}95, RE96, RD99, SPB95, SB98, SZ96, San95, SRJD97, SS96a, SNC^{+}98, SW99, SCMM95, SGW94, SRW95a, SS96b, TMKL97, TGLP95, VK94, VS96, WA99b, WBLM95, Wol92a, YKK^{+}98, dG91a, vdBTvDR96, BF95a]. structured [JC95]. structures [AMS99, HB97, LGV^{+}99a, LGV^{+}99b, PGD^{+}93, PD95, RT92, SNR^{+}96]. structuring [ABN98, CH94, GJVG96, GI99, JHN99]. studied [Bro90, FR91, HN90a, HN90b, HHN90, KCD^{+}95, LF91, RH90b, TR91]. studies [AMB^{+}96, Ber90, BJB95, BH92b, BT93, Bro95a, CLC^{+}98, DGO92, GJVG96, GU98, Hei92b, HM96, HN91, JS97, KD97, KDMR97, NT92, ND96, Pak95, PRM^{+}96a, PRM^{+}96b, PMMH95, PT94a, PT94b, PJ99, Pro92, Put91, San97, SC95a, SAA99, TRPS94, Tis98a, TL91, TB93]. Study [Bia92, SSGD96, Åbe90, Amb93, ARS93, ADG^{+}95, BU93, BVV^{+}94, Bur94, CF93a, Cha92, CJ93a, CRGZ98, CP96, DBB^{+}95, EE98, EMR^{+}97, FTGW90, FQV98, FGB^{+}96, Gar98, GGF99, GGB^{+}99, GAC^{+}97, GF90, GKB^{+}97, GGD99, Hei91, Irw91, JH97, JRMN^{+}97, KMG98, KKT96a, KHK^{+}99, KDAS92, LC91, LMF93, LSL98, LFD95, MF91, Mal96, MC98, McG97b, MWT96, MT93, MBB99a, MBL^{+}98, NNR95, NEIR95, PIC97, PLP92, PLBP96, Rie98, SGL^{+}97, San93b, SMT97, SSR^{+}97, SFG^{+}99, SEGP91, TGG97, VHHS96, VHH^{+}98, VAV99, VK90, VBL^{+}92, WR92b, YBG96, YBG^{+}98]. studying [QRSC99]. stutchburyi [HTCP96]. stylifera [PL99a]. Stylophora [FPGJ99, GYL93]. stylosa [ZCL97]. Stytophora [KYN94]. Sub [WTA^{+}97, HHW90, MS93a, PM90b, Sie93, SPZ^{+}93, War90]. sub-Antarctic [PM90b]. sub-estuary [HHW90]. Sub-lethal [WTA^{+}97]. sub-micrometre [MS93a]. sub-populations [Sie93]. sub-surface [SPZ^{+}93]. sub-tidal [War90]. Subantarctic [Atk96, BE97, DPM96]. subarctic [BSW^{+}95, BWHW95, BMV^{+}96, DFPG91, Hei92a, Kir90, Kir92, MPM97, MLH96, NT95, PM91b, STNI97, Str91, TSI^{+}93]. subcellular [BAOT96, LMT96, PGD^{+}93]. Suberites [BWL^{+}99, MBL^{+}98, MWB^{+}99]. subglobosum [HN97a, Sko96]. subitaneous [RD91]. subject [Tsu90]. Sublethal [DHB^{+}97, PLFSW97, Aro91, SGB90]. Sublittoral [NAW99, GWKW97, GD96, Har97, Oes90, SEGP91]. submarine [AC96a, GF95, JAB^{+}92, Sim92, VWK96, VD99]. submerged [ERC95]. submergence [RR93]. submerging [NDRP93]. submersed [LBK90, NK93, SM97c]. submersible [HP96, LBL91]. submicron [NK96, SKKO97, Sut92]. Subnose [AMMG91]. Subnose-1 [AMMG91]. subpolar [SP99a]. subsaturating [MC97b]. subseasonal [MP96c]. subsequent [MRL96]. substance [SHT91]. substances [CGTB95, CCJC91, HBBM96, MG96c, MH94b]. Substantial [WPTB96, KGO94]. substrata [GBV96, HOM95, LB95, Tom91]. Substrate [GA97b, Mac94, BHG94, FAED97, JLD^{+}96, KS91, LBV^{+}99, MP96a, MA93, McG94, MNH92, PWD^{+}91, SD94, SL90, SK98c]. substrates [BAH94, GAB90a, GAB90b, PPCL^{+}98]. Substratum [LB96b, AS95, BCC92, BCMC99, GB90, LB99b, MS93c, WMB99, HK96]. Subsurface [MHMH96, PBHMI98, BKK^{+}93, BSP93, FP91b, GSP99, GJ96a, Kar92, NKB90]. subterranea [SSV97]. subterranean [KA93]. subtidal [AU93, BCBC98, Ber90, BA97, CO92, CCG^{+}92, DFJ^{+}93, Gla99, HBR^{+}91, HPF92, HD91, MU98, MTP92, MSCTBE96, OM90, ÓM92, Ols93, San93b, SR90b, SS99a, TC94, Tay98, WB94, WH96, WA99b]. Subtropical [CGFC95, FP96b, HSLV94, PPM94, CCG96, DCBG99, DT94b, GAK99, GR93, HSH94, Hsi95, LJ95, LK96, McG95, NL94, OB93, PFP93, RHH97, RH97, SLHA97, TR95c, VK90, ZB98, ZB99a, ZB99b]. subyearling [Lim95]. success [BSMS95, BSH^{+}91a, BSH^{+}91b, CPL^{+}96, CLW99, Dem95, FAED97, FPG95, HS91c, HSB97, JC94, MZ99, Oro96, Ris97, TB95, Yam97a]. successes [KHTP99]. successful [NW97c, PB96a]. Succession [LTT96, BCC96, BMD^{+}90, FSdM^{+}92, NvA91, Nan96, NK92b, TDER97]. successional [BCC94, BF95a]. successionally [Hag95]. successive [Kne93]. succinea [ZHBH90]. sucrose [AZKA99]. suction [KW91]. suecica [Tha92a]. sufficient [FC90, SH90b]. sugar [WM90]. suggested [TR95a]. Sula [GGF99]. Sulawesi [AALT99, EM93]. sulcatum [ADR94]. sulfamate [FT95]. Sulfate [BR91b, KKH^{+}94, AJST98, HK92, HK94, KHB91, KDAS92, MDC98, UGSBA94]. Sulfate-reducing [BR91b]. Sulfide [OV92, Vis90, VHG95, DKL94, MLC97, MK93a, MK93b, SRJD97, SJML99, SR93b, Vis91, Vis93, VH96, WK93]. sulfidic [GVGG98]. sulfur [HBB^{+}99, HN97b, RT93, SGPAA95, SCC98, VH96]. Sulphate [IF96, HHB93, PM91a, SJMC93, UNPH93]. Sulphide [DRS94, WJM^{+}99, GJVG96, JGT96, JT97, JJTS97, MSvdZH98, VDA96, VDJA98]. Sulphide-induced [WJM^{+}99]. sulphoxide [SHML98]. sulphur [BF95a, DRS94, FB95, NBW94, SFWD90]. sulphur-based [SFWD90]. sulphuretum [VDA96]. summary [Add92, Hei92b, Kar92, Moo92a, SDA^{+}92e, Vet92]. Summer [BS95c, MCF90, ØW93, UHL^{+}97, AMB^{+}96, BB95, BGG^{+}97, CPH^{+}96, DG92b, ES90, FPPM96, GH94a, HTG96, HD93, KKH98, KW99b, LL92, MT91, MC91c, NSHF93, ND96, NSC^{+}92, PWR96, PP96, PB92, PRW^{+}94, RC91b, RN98, RPM^{+}94, SJJMH90, SJÆ91, SBP92, SBP93, TSM^{+}94, WA99a, WSCT96, WPTB96, YGH^{+}97, ZBW92]. summer-recruiting [TSM^{+}94]. summer-spawned [MC91c]. summers [GKB^{+}97]. Summertime [Uit96]. sun [HLF^{+}91, PZ91]. superba [AS97, AWH^{+}99, BMHD93, Buc91, CWHA99, DM91, HTCW96, HNL94, MACR98, NSC^{+}92, Nor94, Pak95, PPMF97, PWPS95, SL95, WMD99]. supplementation [MM97b]. supplied [ASP91]. supply [CA96a, CA97, CM98a, DHH98, FW97, FFJ93, GAK99, Hur92, MCA90, MMD93, MMD92, MD94c, MBB95, MZFHC96, Nak92, SBBG94, Sci91, SW98b, SC96a, TSM^{+}94, Vos91, WWB98, ZBW92]. Support [MH94b, GS94a, JAB^{+}92]. supporting [TC93]. suprabenthic [CSB97]. supratidal [Ang95a, Ang95b]. surf [BA91, CJL91, CBL96a, CBL96b, Coc90, Har98]. surf-zone [BA91, Har98]. Surface [GB91, HC92a, Kep91, AMMG91, BE97, BM99b, CLGM96, CBD96, CPSA96, DEF91, Gar91, Gen92, Gra96b, GAC^{+}97, GBGH97, GKB^{+}97, HM97a, HAA94, KLR93, Kar92, KM96, KW92a, KDR97, LBLA93, McC94c, MC94a, ML90, PM90a, PLH91, PGAJ99, RKS97, RR93, SM97a, SPZ^{+}93, TA93, TPR98, TTR99, WWGK95, WWB98, WPG99, Wil92, dNDS98]. surface-deposit [KM96]. Surface-layer [GB91]. surface-water [Gar91]. surfaces [JMR91, JSOG91, MS95, WW96]. surfgrass [BWRH99, TW97]. surgeonfish [KMG98, Pla93, Ris97, SML97]. surgeonfishes [PPP^{+}96, SR94]. surmuletus [LA97b, MST98]. surrounding [DH96, SNR^{+}96]. Survey [SDC^{+}92, PS97]. surveys [Bia92, GBB^{+}98, KRM94, TM97]. Survival [BSH^{+}94, KS95a, MLES94, Wil92, lYALM93, AI92, BM94a, BH94a, BH95a, BSS92, BB95, BRC99, BB93, BSH^{+}91a, BCBR97, CA97, COK99, Che93, CH90, DE94, ELMCC90, EGMC94, FAED97, FK95b, FQV98, Fre97, GQ96, GCC95, KR97a, KH92b, Ken94a, KEV99, LKC98, Leo99, LCTT92, LF97, MJW96, Mal98, MU98, MT91, McF92, MB93, Min93, MBL94, MNO96, MJ93, NKV94, OTF95, QQ98, RMG^{+}98, SK99, Smi95, SS91, TB97, TB92, VÅCL99, WW96, Weh91, WtD90, WK92b, WC96, YNYK94, YB90, tor93, vTS97]. survivorship [CH95, DES92, HDB95, Lew91, LAK99, RS94, RS95, SDB96, YY91]. Susceptibility [BLH99]. suspended [BSF95, CHC94, HGS94, JCH93, MC96a, NNB93, PF91a, PSE97, SDB92, SDB94, ST98]. Suspension [HSTY98, KGC92, Rii91, RTJ^{+}93, TWWE95, AP99a, AR91, AJL93, BJdJ^{+}98, CH99, DLHB99, FL90, FAP92, HSBH96, HJH^{+}99, Hil91c, JJNR95, Kam94a, KSRB99, LHL97, LJSK96, NT95, PTW^{+}91, RCG99a, RW99, RI90, SK95a, Ved98, YKL98a, YKL98b]. suspension-feeders [DLHB99, LHL97]. Suspension-feeding [HSTY98, AR91, CH99, HSBH96, HJH^{+}99, Kam94a, KSRB99, NT95, RI90, Ved98]. sustain [GQ97a]. Sustainability [She94]. Sustained [SB97c, Sto97, Str99]. sustaining [GK96a]. Svalbard [BHHL96, GHHC98, SB98]. SW [HdWS91, KP95, BI99, CR95, CG99a, LH99, MAH94, PCB90, PSPD96, RDA97, TI96]. swamp [Hsi95, KHB91, KDAS92, MWH91]. swarm [Lav97]. swarms [BMHD93, Gen92, LY97, WBP^{+}92]. swash [MM91b]. Sweden [AS94b, Han97, SR93a]. Swedish [TN98]. swept [WG97]. swim [Fra92]. ’swimmer [SPS98]. Swimming [AHY^{+}99, JAL91, MCL90, VV95, BH97b, BP96b, FGG95, LCE97, LH91, SA92, Sto97, SB97c, TW98, WX97, YS96, ZN96b, vdL95]. switching [GA97a, KSV96, LW95]. sword [GCNSNd94]. sword-fish [GCNSNd94]. Sydney [SS99b, SGMS99]. symbiont [KGC92, Vil91]. symbiont-containing [KGC92]. symbionts [HCF93, WK92a]. Symbiosis [VFMFBE96, BBLY98, FGH95, LGP95, MLL98]. symbiotic [BDA^{+}99, DS93a, GAN^{+}94, HGJ99, IPT94, MPH96, MPCD90, MGD91, PFMC95, PFMG96, RP91b, VM98a]. sympatric [BKB94, MS91, NB98, WWC^{+}99]. Synchronization [LFBH94]. Synchronous [FM95a, Joh92]. synchrony [HM97d, MD94c, RD92]. Synechococcus [ADA98, CCG96, KFF92, LCL95, LBL^{+}99, MGC95b, Neu92, SC93, VMHH90]. synergistic [HLP99b]. Syngnathidae [FMRC93]. Syngnathus [FMRC93]. syntheses [HS97b]. synthesis [ACC90, CC90, CC91b, FRP^{+}99, GW93, LVL^{+}99, LLB^{+}95, MG97, YFI91, vBV90a, vBV90b]. Syringodium [GMR^{+}94, KS98, SDC90]. System [Gar98, LGGLLHV98, MG96b, WH98, ASRP^{+}96b, ASRP^{+}96a, BBL93, BWHC91, CCC90a, DBHM99, DHS96, DL91a, DMP^{+}99, DE90, DvRN99, FHCC98, GS94a, HTS92, HTG96, KSO^{+}97, KEvN95, KK96, LJM99, MM91a, Moe96b, Moe97a, NC94, NJSS91, PBHMI98, PJ99, RM94, Rel96a, Rel96b, SR98a, SSY98, SGW94, SM95, SFG^{+}99, SLHC98, THS94, Tho91, TFF94, TFFA99, VAR94, VWS92, VWS93, WBLM95, VPM96]. systemic [SB99]. systems [JG94, LA97b, MB95a, MWT96, Moe96a, SFG^{+}99, TJ90a].

T. [DS96, DS93a, Jon93, KL94]. tactile [MJ92a]. taenionofus [FMRC93]. TAG [GPS93]. tagal [SOSV95]. Tagelus [GI99]. tags [BK90, MG91b, Spe95]. Tagus [BPC98]. Tahiti [SP99b]. tailed [HCH^{+}96, WC98]. Taiwan [HHW99]. Taiwanese [Dai90]. Takapoto [NJLD98]. take [ZB92]. taken [HTCW96, HH90]. Takifugu [KJLH96, KKYH97, Yam97a, Yam97b]. tale [TDER97]. talitrid [CBT98]. Talitridae [CV96]. talpoida [AS99b]. tamarense [MCC97]. Tampa [FKC96, FMH96]. tangeri [LMO99, RDA97, Wol92a, Wol92b, Wol93]. tapirina [MJ91]. taraikaensis [KA97]. target [BBT^{+}93, PR93b]. targeted [LMK93]. targeting [YKK^{+}98]. targets [BW96]. Tasman [MO95]. Tasmania [WC98, YBLL96]. Tasmanian [MHT^{+}97]. tasmanica [BAM92, Gib94]. tautog [CA98]. taxa [BPC98, CB98, CWH98, FAHW99, GM98a, Joh99, SJMC93]. taxifolia [BB97a, CC98, CC99, CJ97, FGR97, JPZ^{+}98, KMB97]. Taxon [FML^{+}95, MHvB97, SS99a]. Taxon-specific [FML^{+}95, MHvB97, SS99a]. Taxonomic [FC90, HG98, SC95a, Cha98, CW99b, JLF95, JSGS94, Mil94b, MP96c, OSC97, OSC98, RK99, SKvB98, SS93, SOC97, VWJ96, WC95, ZDJ^{+}95]. TBT [BQR99, RBG94a, RBG94b]. technical [PTW^{+}91]. technigues [AAD91, MC91d]. technique [ASP91, BD91, CF94, JRB96, JRB97, JM93, JCH93, LT97, LKC95, LCL95, LKvRH96, Min93, MG99, NG96, RVGC^{+}99, RHS99]. techniques [BZ92, DD97, KFD99, KH92b, LBV^{+}99, NW97a, RTCH96]. technology [vTS97]. technology-options [vTS97]. Tehuantepec [Bia91]. tehuelchus [Iri90]. telemetry [CLJOC98, GKB^{+}97, Zel98]. teleost [DCG^{+}99]. Teleostei [MF97, FGG97, WM92]. Tellinidae [GM99, JT97, JJTS97, WJM^{+}99]. Temora [DP91, HvB91, PL99a, VV95]. temperate [And94b, BKS^{+}99, BBL^{+}95, BHT98, Cha91, Cow91, DKW96, DÁSP97, FWWL99, GAK99, GK96a, GY97, Hay95, HS99a, HK93, IP94, JJ99, JBWH97, JWKH98, JBK99, KGHTN99, KW99a, LB99b, LH96, LCG97, MW99, MEC92, MDG98, Mil95b, OC95, OM99a, OSJ94a, ODN^{+}95, PMB99, RCG98, RCG99a, RCG99b, RPDR^{+}98, RLC92, SK99, SFSA99, SB94, SFA98, Tak99, TMKL97, TB97, UCV98, US95, US98, War90, WP92, WV95, WM92]. temperate-zone [WP92]. Temperature [AJST98, Bre95, KS99, KMB97, PWNW99, RT95, SD94, WtD90, tor93, AGP98, ADA98, Ang91c, ACOK95, BTKK99, BK90, BGR^{+}96, BH97b, BB93, CA99a, CC96, DF95, Fly99, GMK97, GBGH97, Han97, HvdM97, HvdMBK98, HLP99b, HSBP97, HM94b, JBO97, JLR90, KS91, KGS95, KBA^{+}97, KRF^{+}98, LL92, LFBH94, MT91, Mat92, MM95a, Mil95b, MGC95b, MJ97b, Nie96, ORD98, PF98, PWD^{+}91, QQ97, QQ98, QQ99, RVGC^{+}99, RI90, RVBL92, RTJ^{+}93, SRJD97, SK95c, TT98, URHR98, Urb94, VLKC94, VTNA91, Ved98, WR92a, WF90, WDD92, YTF97, vdL95, vdVW99]. temperature-controlled [VLKC94]. Temperature-dependent [KS99, RT95]. temperature-influenced [HvdMBK98]. temperatures [BAOT96, GCKW95, GKB^{+}97, HGS95, SP99a, WHG^{+}99a, WWGK95]. Temporal [ÅP97a, AGP98, APS91, BSVK93, BM99b, BG99, CBI91, CBL96b, CSB94, Gra91c, GAC^{+}97, HC90, HM97b, HNvB94, JBWH97, Ker96, KW99a, LRV94, MZ99, MGM96, MRPC98, MDTB94, PS94, RR97b, SRH^{+}98, TLB93, TSM^{+}94, AWD95, ARDL97, AHT^{+}90, ALB^{+}96, Arr94, BCNC96, BPC98, BGVF92, BRW95, CC98, CPK97, CCSN93, DBV98, FMFH97, Fow90a, FSG^{+}97, Gre96, Gru94, GBGH97, Hsi95, HS98c, Hur92, JJ99, KLN^{+}90, KSE96, KKW^{+}93, KHK^{+}99, KID91, LP95, LSG^{+}99, MCF90, MYYT96, MMD93, MHS94, MS94b, MM95b, MBB99b, Pin94, PF96, SV93a, SP99b, SM97c, SF91, THKN98, VHVB^{+}98, WM97, dG91a]. temporarily [AOOT94, BR91b]. ten [Urb94]. tenacity [BG97, TJRG93]. tension [VTNA91]. Tentacle [NR98]. Tenualosa [MCFB97]. tenuicornis [She98]. tenuiremis [GCC95, WC91]. TEP [ES99, MB98, Mar99b, PA94]. term [AdN91, AT96, BLJ^{+}96, BGR^{+}96, DBHM99, DvCd94, Dav94, DdA95, DF95, Dye98, FSCE92, FRK92, FCH99, FSG^{+}98, GHMC96, GRBA96, GGB94, Har94, HP97, HWL96, HGB99, HB98, JAOS90, JCMW91, KDHZ98, LGGLLHV98, LD95b, LDKR93, MDD95, MS94a, MRMR92, MP96c, NH98, Oes90, Onu96, PRC^{+}92, Pel96, QGQ97, RGS90, RW96, RMMB91, San94a, SMT97, SK98b, Sch99, She96, SGMS99, SGW94, Ste94, SDB94, TD91, TFL95, Tis98b, VDVC94, VPM96, VM95, WUA^{+}99]. terminal [LCC97]. termination [BEH93, KYI^{+}98]. terminology [TS90a]. Terminos [RMTB^{+}95]. terrestrial [CSG93]. terrigenous [EM93]. territorial [WB97]. Territories [HCS90, SCEG93]. tertiolecta [HPSG94]. terubok [MCFB97]. test [AH94, BH91, BGVF92, DD97, ERC95, MY98, PMETM96, PSL96, PKL97, PS98, Sew96, Spr93, vCRS92, vW93, RG97]. Testing [BBRBK94, Lew91, AH95, BKS92, BSG^{+}97, DB96a, Min93, MBB95]. Tests [McG97b, BCv^{+}92, CSR^{+}92b, CU99, LK93, LKC95, PN99]. testudinum [DCK99, DH92, Dur94, FZ91, GMMD93, HD97, KFD99, KG99, LD96, RBC^{+}91, RD99, TL91]. tethered [BSH^{+}94, ND98]. Tethering [AH95, BS94, CA98, HP96, Min93, Smi95, ZFFH^{+}94]. tethers [Eml90]. Tethya [GS94a]. Tetra [Tha92a]. Tetra-selmis [Tha92a]. Tetrapygus [RO93]. Tetraselmis [GHTH94]. tetricus [And97]. tevoroa [KL94]. Texas [LD96, MW97, ZB94, BH95b, Ker90, MC96a, MZ92, Onu96, YB92]. Th [COS93]. Th-228 [COS93]. Th-228/Ra-228 [COS93]. thaca [UC94]. Thailand [ABB^{+}97, BLST90, KKH^{+}94, LB96a, NA99]. Thalassia [DCK99, DH92, Dur94, FZ91, GMMD93, HD97, KFD99, KG99, LD96, RBC^{+}91, RD99, SAvD^{+}96, SMH97, TL91]. thalassinid [SSV97]. Thalassinidea [GS91b, PAR98, SF99]. thalassinidean [FAED97, GS91b, NA95]. Thalassiosira [DRGV91, DEP^{+}96, TP96]. Thalassoma [HH91b, MSP^{+}97]. thallus [DKVD95]. Thau [GSBB97, SGC^{+}98]. Thecata [Mat92]. thecate [JLYK97, JSK^{+}99]. Thecosomata [GH91]. their [AHY^{+}99, BL97, BGCM95, CSM99, DFPG91, EL90, FP91b, FLDT92, FPFR92, GCR98, GH94a, GCF96, GBP^{+}97, HKMK98, HG99, HS91b, HHM97, IFL97, Kin93, LS96b, ML91a, McC94c, MRP98, MBR94, MVdM^{+}99, MLL98, Pro92, RR99, RTRF90, RWJ^{+}95, RP91a, Röp93, RO98, SM98a, Sha98, Sie93, SJMC93, ST98, SKT97, TTL^{+}96, VH93, Wai90, WT98, WSMM99, ZF99, vKSGvA99]. Thelidium [EGC95]. Theme [BRA^{+}99]. Themisto [PP96]. Thenus [Jon93]. Theoretical [GDMS90, AE91]. Theory [MWT96, Bra93, EST95, MB98, SL97a, ZH97, TS90a]. Theragra [BM94a, BCN^{+}95, BBK91, CBI91, DO92, HHP95, HBP^{+}93, Liv93, OD90, YB90]. therapeutic [PLP92]. there [BCHL93, MD91]. thermal [BMP^{+}95, BSB91, HGB99, MBB99b, Ral98, SO96b, WM95]. thermic [dRAV92]. thermoclines [PGM^{+}98]. thermohaline [MH97]. Thermostability [GS90, JDLL95]. thermydron [EPDC99]. these [BC99]. thickness [LGR95]. thin [HSMS99]. thinning [FL90, FAP92]. thionein [SGR92]. Thioploca [Far98]. Thoracica [PGB91]. thorns [BMBD95, HK96, HKR96, HR96a, JSOG91, SB99]. those [MB98, Pud93, RGT98, SM98a]. Thraustochytrid [RR99, KFN99]. thraustochytrids [NTK98]. thread [MD93]. threatened [OJR96]. Three [GWSG99, AMK99, DES92, DCM90, EK95, Ejd98, FP91b, Fra93, FSGG96, FSG^{+}97, GHD92, GKB^{+}97, GP90, HN97a, HS99b, Hul99, Joc95, KR94, LGED98, MZ99, MWA93, MCL90, MLC97, MP92, MMD92, MK93a, MU97, MV96, MH94b, NT94, NA95, PM94b, RP91a, SB97a, SE93, TK95a, Tis98b, Van92b, VS92, WSMM99, Whi94, YAG92, ZGM^{+}96]. Three-dimensional [GWSG99, ZGM^{+}96]. three-trophic-level [Ejd98]. Threshold [PM92b]. thresholds [WDD92]. throughout [JCCCB94, RCG99b]. throughput [LUA99]. Thunnus [DJY90, DLJ91, JD90, KJGC96]. Thy [Nor92]. Thy-sanoessa [Nor92]. Thyasira [DS93a, DBN^{+}94]. thyasirid [DS93a]. thymidine [Kir92, RBJ90, Tuo97]. thymine [BEG97]. Thysanoessa [CPS91, DS96, KHFP^{+}96, Nor94, Tan98b]. Thysanoteuthidae [NAS95]. Thysanoteuthis [NAS95]. Tidal [AWD95, ASCW90, GRBA96, Hic92, RK95b, RE94b, TDW96, vDdWKW99, AAWD97, Arm92, BU93, BNK^{+}97, Beu92, BSG^{+}97, BPC98, Bur94, CC97b, CHDW92, CSM99, DH96, DE90, Dit96, DL94b, FPLH92, FB94, GKMT99, HvB91, HKY^{+}97, Hil97, Hil91b, HH99, JWKH98, JHN99, KW94, LDKR93, LH99, MBM92, MZCMOCJP91, Mor96, Mun93, MS97, NL94, NDG^{+}99, PGB96, PE99, Phi94, PZ91, QCM97, RPN95, RM95, SS99a, SMTL99, SDB92, SDB94, SR97, Tak95, TK97, Van92b, War90, WFH99, ZN96a, ZN96b, ZAN99, dG91a]. tidal-flat [Beu92, GKMT99]. tidally [LSL98]. tide [BH95b, DH90, FPFR92, Han95, JL94, Jeo94b, JSK^{+}99, KG99, LJ96, LCK^{+}96, ML91a, MKV95, MC97b, NU91, NYH92, NSH95, Onu96, PBHMI98, Pre91, SHA^{+}95a, SMH97, SMP97, UYMH95, UT90, VHG95, WMG91, YKK^{+}98, ZN96a]. tide- [KG99]. tidepool [Hul99, MS94b]. tidepools [HS95, MS93b, MHS94, MS96a, MS96b, Mga92]. tides [TFF94, YHC^{+}99]. tied [Vid94]. tiger [CLD99, HHK^{+}98, HVL96, LWK^{+}95, O’B94a, O’B94b, SCMM95, VHHS96]. tight [Kru93]. Tikehau [CRCS96, CCRB98]. Time [BM95b, DL91b, RCS^{+}96, Aks96, ASRP^{+}96b, ASRP^{+}96a, BM99a, BM95a, BSH^{+}91b, Bur94, CM99, DFB^{+}99, Hav93, IHK^{+}98, JE95, KS99, LIPF95, Lop91, MC96a, MC91a, MRMR92, Min93, MP96c, NKH^{+}99, RÁSP99, RSD^{+}91, SL97a, SK98c]. Time-courses [DL91b]. time-scales [MC96a]. time-series [BM99a]. times [EG97, SB99, SCF91]. Timing [AS99b, CT96, HM94b, LBL97, Nor96, RHH97, BP99, GJC97, Kam94a, KPWL93, MBL94, Mor96, Nan96, O’C90, OJ97, Yam97a]. tin [DBHM99]. tinian [GYL93]. tintinnid [KA97, PCCG97, WMPT94]. Tintinnida [CES^{+}94]. Tintinnids [LHRL97, Kam94b, PT93, PT94b]. Tisbe [Wil92]. Tissue [Bar95, BLD94, LW95, BAOT96, CP95a, DS93a, FDZ94, FBR^{+}95, HV91, Hob93, MB93, NBS^{+}97, PLFSW97, SDB96, SV97, WJD93]. tissue-dependent [Hob93]. tissues [BC93, KA93, SDFJv91, VCMP93]. Title [Ano96a, Ano97a, Ano97b, Ano98, Ano99b]. titration [CH97b]. tividus [WLT^{+}95]. Tokyo [HFN91, NKKO97]. Tolerance [Joh97, MSvdZH98, QQ99, CCD94, GJVG96, Hoe91, JT97, OV92, RR99, RHJ91, Urb94, Vis90, VHG95, WNJ91, WM95, tor93]. tolerances [BUWB99]. tolerant [Sie93, SJMC93]. toli [MCFB97]. Tomales [DSV^{+}91, FMFH97]. Tomioka [NTK93]. Tonga [KL94]. tonsa [BD94, CLW99, DDW90, FD98, FR91, Jeo94b, JT90, KSV96, KSV99, SK95a, TDD94, WR92a]. too [PN99]. tool [CA98, GRGK96, MJSDF91, PPO^{+}99, VPGV94]. tools [DMA^{+}94, Sch99]. tooth [GCM98]. Top [MS96b, FAVV98, Gas94, KLAH^{+}94, LGJ99]. Top-down [MS96b, FAVV98, Gas94]. Topographic [GB98, CB97]. topographically [NAW99, WW96]. topography [McC94c, ZHF96]. torosa [JGT96]. torreyi [BWRH99, RGLD^{+}98, TW97]. Total [HSN93, JMD95, RSD^{+}91, DSV^{+}91, GKMT99, MGM96, PL99b]. tough [HSMS99]. toughness [CP95a]. tow [GCM98]. towed [LW94]. Toxic [MUH97, Nie93, AKM^{+}99, AGGKA96, BGH95, BMG99, DHM92, Dut98, Hal98, Han95, JSK^{+}99, JKT^{+}98, KDL92, KEV99, LRV94, MHT^{+}97, Nie96, Ols93]. Toxicant [BWB98, LGFS99]. Toxicity [Kar92, MC94a, SFF98, AHT^{+}90, BKS92, BLC91, BCv^{+}92, CSR^{+}92b, CMLE92, HC92a, KOSS90, MCC97, RBG94a, RBG94b, TP96, vKVSR97, vCRS92]. toxicokinetics [SFF98]. Toxicological [MPH96, SH92a]. toxin [FFF^{+}94, Tee99]. toxins [BLCA90, BLC91, HTLK99, KB91, SOS^{+}92]. trabeculata [CO92, MC93]. Trace [RATA^{+}99, SOK98, Sca93, BL95, CAATS94, RGBC92, SDA^{+}92a, WFL95, WFL96, WF97, WD99, WH96, RDM91]. traced [MvdVD^{+}97]. tracer [BG91, CF94, HWLC98, LM91b, PPCR97, Put91, Sim91, SR95]. tracers [SGMS99]. traces [Lim95]. Trachichthyidae [BK92]. Trachymedusae [PGG96]. tracked [BGG^{+}97, GKB^{+}97, RPT^{+}96]. tracking [AMMG91, JCCCB94]. Trade [GSMM96, AI92, KWL96, WB91, vRBVB95]. Trade-off [GSMM96, AI92, KWL96, vRBVB95]. trade-offs [WB91]. traditional [GP91]. tragula [MM93]. trail [DB99]. trail-following [DB99]. trails [DB99, NA99, PH91b]. training [TM97]. traits [CRGZ98, Gar99a, GC95a, HSBP97, LGFS99, MCNB98, Nak94, Sal96]. trajectory [JD90]. trammel [EMR^{+}97]. transect [BVL92, CT97, KR92, LC93, MAC92, SHV91, VBL^{+}92]. transects [O’D98]. transfer [BHT^{+}92, GF90, LP99, PJ99, RGBC92, RDE^{+}90a, RDE^{+}90b, WLL98, WQQ99]. transfers [SCD^{+}90]. transformation [HH92, SGPAA95]. transformations [CDHD92, DvRN99, KEvN95, OSC97]. transformed [SSG96]. transient [MJV92, VNB97]. Transition [Bus96, Åbe90, FVDL95, Gos97, JBB90, MM97a, SFS95, SBP92, SBP93, VBF94, VDBF96, ZA96]. Transkei [Las98]. Translocation [MGDT90, ZA96]. Transmission [BB97a, MJ97b, LHH96]. transparency [WPY98]. transparent [ES99, MB98, Mar99b, PA94, SH95a]. transplantation [YAG92, vTS97]. transplantations [JKT90b, KJT90, TAJK90, Väi90]. transplanted [FKC^{+}90, MZ92, vTS97]. transplants [McC96b, PSL96]. Transport [Hal98, TD92, ASH93, Bel95, BSJPW93, BD96, BV99, CCC90a, DZ99, EAEP98, FTS94, Fly98, Fly99, Gar99b, GF99, HC93, HG99, HBP^{+}93, JBK99, KM96, KCS94, Lam98, LD95b, LKM^{+}99, MKYI98, MO95, NH98, Olm94, ORL97, PCCG97, QCM97, Rel96a, Rel96b, RHS99, RE94a, RE94b, SC94, Sha98, SMW^{+}93, TDW96, WD90, WFH99, ZD99]. transuranium [GF90]. trap [CV98, DFB^{+}99, Hei95, SOW97, SPS98]. trapezium [Sca93]. traps [HS99a, KCI98, NO90, Tho92]. Trawl [BMSF95, Bia92, KRM94, KS95a, KR97a, KKW97, LWK^{+}95, PSR^{+}99]. trawled [KS94]. trawler [Oro96]. trawling [FAHW99, Kai96, OJR96, PSR^{+}99, RKH96, RH91, THR^{+}98]. trayula [McC94d]. treated [WW92b]. Treatment [Tho91, RHS99]. treatments [PB94]. Tree [Fow90a, SU93, Bar97c]. trematodes [CUM98]. Trench [FKT^{+}99]. Trends [AW95, BvBBV96, Mar97, YAG92, APT99, Bro95a, COK99, DdA95, Har94, JCMW91, MD91, MDC98, MSM^{+}98, PWM95, RS91, RSS94, SR90a, SRSH99]. tri [DBHM99]. tri- [DBHM99]. triacylglycerol [HF99]. triacylglycerol/sterol [HF99]. triacylglycerols [HKG93]. triangle [AMB99]. tribromopyrrole [FT95]. Tributyltin [RQB98a, DHB^{+}97, LOS90, RPT95, RGT98, RBG94a, RBG94b, DHB^{+}98]. tributyltin- [RBG94a]. Trichiurus [KN99]. Trichodesmium [BEU99, CSM^{+}98, CCC^{+}90b, COD^{+}93, CR95, LK96, MC99, Nau96]. Trichosphaerium [CRLP94]. tricornutum [JR92, LLB^{+}95, RGS90, RH90b, SDH95, SHT91]. Tridacna [KL94]. Tridacnidae [GK96b, KG94]. tridactyla [WTA^{+}97]. Tridimensional [AC96c]. Trieste [MMM^{+}95, PH91a, PF91a]. trigger [AGP98]. triggering [CCJC91, WMG91]. triggers [HLP99b]. triglid [LM98a]. trigonura [Big92]. trilobite [BLJ92]. trilopha [MSCTBE96]. Trinidad [AGW93]. triostegus [McC99, Pla93]. triphosphate [HSRHH90]. triplication [Lop91]. tripton [PLB95]. Tritonia [CHFL95]. trituberculatus [LH91]. trivial [Tok95]. Trochopus [CR98]. Trophic [BP97b, CGZ95, DG91b, DG92b, Hob93, KLAH^{+}94, MvdVD^{+}97, McC94b, STNI97, TP97, ASN95, BU93, BB92, BP92, DLHB99, Ejd98, EJ93, FCP98, FQ94, GTWF98, GS91b, Han91b, HW92, HAR95, HFGT94, IBB97, JRMN^{+}97, KI97, LHL97, LJM99, Lee97, LP99, MMR99, MU97, MJSDF91, Nag96, ND98, NA95, PW94, PDB^{+}97, RSH93, RTRF90, RAB^{+}92, Rie98, RPT^{+}96, RPT95, SDBW95, TC93, WLL98, WQQ99]. Trophodynamics [PP96, RW99]. Trophos [PD95]. Trophos-2 [PD95]. tropicai [McC94d]. tropical [AGW93, Alo90, AHMd90, ALOT99, AALT99, ATH95, Bar97b, Bar97c, Bar97a, Bar99, BUWB99, BMSF95, BCG94, BRW95, Cal95a, COD^{+}93, CMB^{+}99, CSSPV98, DeV92, DW97, DE90, Dit96, Duf92, DB96b, Dun99, EM93, FFH99, Fow90b, Ger90, GWWP99, HA90, HSTY98, HHK^{+}98, JWL96, KW97a, KW98, Ker94, Ker96, KTW^{+}97, LFBH94, MCC96a, MY96, MS94a, McC94c, MM95a, McC99, MD94b, MBR94, MCFB97, MRP90, MGD91, NW97a, PLF90, PNS90, PLB95, PM91a, Rei90, RHH97, RBC^{+}91, RT92, SKM93, STBR99, Sal96, She95, She96, SDC90, SFA98, SRW95b, VHH^{+}96, VHH^{+}98, WM95, WM97, YMD94, YFI91, RH90a]. trossulus [ABY97, HS96, HS98c, Kre93]. Trough [Bra95]. trout [Sam97, Zel97, Zel98]. Tuamotu [CRCS96, CCR91, Cha96, CDG97, DCBG99]. tube [ASF98, GSM^{+}97, LQ97, SS92]. tube-building [LQ97]. tube-worms [SS92]. tubes [Mar91, MSK95, PL99b]. tubeworm [GSM^{+}97]. tubeworms [SRG^{+}99]. Tubularia [LB96b, NB96a]. tubulosa [CJ93b]. Tulear [PCB90]. tumida [TTW97]. tumor [BBR^{+}93]. tuna [BM94d, DJY90, DLJ91, JD90, KJGC96, YBLL96]. tunas [YD90]. tunicata [MD98b]. tunicate [Dei90, DCRG92, HlS90, Lav97, RR99]. tunicates [ADBH99, GLD^{+}91]. turban [KSH97]. turbative [SSV97]. Turbid [MRP98, DD90, HSBH96, KP95, MC96a, MG96a]. Turbidity [Mar99a, ASCW90, CB96, GT96, LCTT92, MNO96]. turbidostat [LW91]. Turbinaria [SP99b]. Turbo [KSH97]. turbot [CTC93, CMT94]. Turbulence [Sun96, VM98b, Yam96, BS96b, Bro96a, Bro96b, ML91a, OS96, PGM^{+}98, PG94, PCG96, PMG^{+}98, SH99b, WFH99, YS96]. Turbulence-induced [Sun96, VM98b]. Turbulent [San97, CPC98, HM97b, KS95c, LEHvR96, ML93, PSK98, SA92, SK95a, SC96b, Vis97, Wor95]. turf [AFC96, AV98]. turf-forming [AV98]. Turks [SB96]. Turnover [KQH^{+}96, CF93a, CBD96, Fab95, KKH^{+}94, NBW94, PGAJ99, SCF91, TL94]. turtle [CL97, HGBW99, HS91c, KFD99, LC97b]. turtles [GTWF98, HLP^{+}99a, HGB99, HW99, MBB99b]. TV [BKH99]. two [ASN93, AO91, Bak90, BG97, Ber90, BF95b, BMFR97, BFMM99, BM95a, BG93, COD^{+}93, CSC92, CBRD92, Cha96, CS99, Cla96, CRGZ98, CM97, CCGM93, CSM99, DBV98, FGR97, GH94b, GB97a, GB97b, GRGK96, Gra91c, GAC^{+}97, GBB^{+}98, HKY^{+}97, HNC92, Hun93, ICG96, JNJP94, Joh96, Ker90, Ker96, KL91c, KRJL96, KWL98, KHT93, LM98a, LGV^{+}99a, LS96b, LBM^{+}99, LAJ95, LBL97, LE90, LA97b, LF97, LGL^{+}95, MEK99, MCSV95, MDTB94, MMR99, MTP92, NB90, NR94, PPCL^{+}98, PR93b, PPO^{+}99, PFMC95, PFMG96, RMGB^{+}94, Ruz96, RO97, RO99, SJ97, SW98a, SDB96, SR98b, She96, SW98b, SOC97, SJC99, SAA99, TDER97, TFL95, Tol98, TVG96, TB99, TGLP95, VHH^{+}98, VDM90, WFHM99, WCP^{+}98, WWC^{+}99, WWL97, WC96, WB99, YB92, ZHBH90]. two-compartment [BMFR97, BFMM99]. two-dimensional [GRGK96]. type [BCC92, DZ99, Fer99, FFT98, Hav93, Kur99, SJCL93, SL90]. types [CSR^{+}92a, GWW^{+}98, GY97, Lai92, MF91, MDRW99, SW99]. typical [WS92a]. typicus [CPC98]. tyrannus [FCW98, MC91a]. Tyrrhenian [ASN93].

U. [O’C93a, VSPLH98]. U.S. [MU97]. Ubatuba [Pir92]. Ubiquitous [RQB98b]. Uca [LMO99, O’C90, O’C93a, OJ99, ÓN97, RDA97, Sha98, Wol92a, Wol92b, Wol93]. UK [BP98, HSW99, KGHTN99, Ken94b, PHHH96, RNHO98]. ultimately [KW97a]. Ultra [RLT^{+}94]. Ultra-algae [RLT^{+}94]. Ultraphytoplankton [LZYW93, Li95]. ultrasonic [Zel98]. Ultrastructure [DWG98, HLH96, LHH96, PL99a, BZSG96]. Ultraviolet [PGB96, AS99a, BCHL93, BEG97, Dun99, GPB96, HVFHH92, HMVHH95, JPA^{+}96, MSND^{+}99, MJ93, Raw96, SGL^{+}97, SDC^{+}92]. Ultraviolet-B [PGB96, BCHL93, Dun99, MSND^{+}99]. Ulva [BSMS95, HLF^{+}91, VSPLH98]. umbilicalis [HCG^{+}93]. Unbalanced [CLK90]. uncertainty [DHB^{+}98, RQB98a]. unconsolidated [CDHD92]. Uncoupling [CCCH94, WSM91, LC97a, Sci91]. undatum [IWF^{+}97, LSM92, RMH96]. under-ice [RLT^{+}94, Wer97]. undergoing [RR93]. underlying [DAS92, HSBH96, HW95, HK92, MD98b, MMG97]. undersized [OC93b]. underwater [BGT91, BKH99, Dun99, GF94, RGLS99, SFS95, Sty97, TM97, WR98]. undescribed [LC95]. Undisturbed [LGC^{+}93]. unequal [Ber90]. unfished [THR^{+}98]. unharvested [SH93]. unicells [BSMS95]. unidirectional [QRSC99]. unifasciata [Cha95]. unique [AFZ96, CN95, LJ95]. unit [IPT94]. United [ONH^{+}97]. unpolluted [BCA^{+}99]. unprotected [MBF^{+}97, SRH^{+}98]. unrelated [HB98]. Unusual [Naj97, ZT90]. upcurrent [SML97]. Upeneus [MM93, McC94d]. upon [BLD94, CLW99, ELH92, ENSR90, KA97, KKW97, ML91b]. Upper [ASN95, Urb94, BF95b, Bia92, CWH98, DY92, FII95, HSLV94, NT96c, Onu96, PWP^{+}97, SL91]. upstream [NR98]. upstream-collecting [NR98]. Uptake [BLCA90, CBRD92, LWL98, MS93a, TSS93, AE91, BFD91, BKB92, BBO^{+}95, BLC91, Byt90, Cam91, CM90a, Car90, CD94, CPH91, Col98, DW95, Dor90, Gar91, GST^{+}91, GCLS94, HPSJ95, HBvD92, KM99a, LA97a, Lip95, MRL96, MC99, NB90, ÓMvdB99, PB97, PLC^{+}91, RK95a, RATA^{+}99, RBA98, RRWG96, RWJ99a, RFN98, RN98, Sca93, SAH98, Sci91, SFF98, SAvD^{+}96, SH92b, SCF91, TI96, TPR98, TW97, TK91, UNPH93, WD99, WYY^{+}98, WM90, YSL94, vRV96]. Upwelling [TFF94, ASRP^{+}96b, ASRP^{+}96a, BP92, DW95, DÁSP97, DG96, EM99, FS91, FRP^{+}99, FF92b, Gib94, GGH99, HNC92, Lav94, ME97, NC94, OB94c, PLP^{+}93, PML93, PF92b, PBHMI98, RÁSP99, SS99b, TFFA99, WBLM95, ZF99]. Upwelling-downwelling [TFF94]. upwellings [MADM98]. urban [AMS99]. urceolus [RTJ^{+}93]. urchin [AU93, BCC95, DE91, Kea90, KL91b, Lam98, LB99a, RZLB96, RSK^{+}99, Row90, SZ96, SHT91, WO94]. urchins [AMB99, AM91, BCBC98, ESDK94, EDS^{+}99, Hag95, KSS90, ME97, PS93, Row90, SRH^{+}98, SR90b, SH97b, SDER96, TD91, WH91]. Urea [CA95, TI96, CC96, CPSA96, CG99b, CMBM99, RFN98, TL94, WYY^{+}98]. Uria [DH96]. Urrutia [Ano00]. Uruguay [BD99, DdA95]. Uruguayan [GY97]. USA [ATNW97, AML95, BLJ92, BH95b, CCG98, CPB96, Dur94, EDP^{+}98, GJ97a, GJ97b, LLB92, LSGW98, Pak92, RBC^{+}91, RR93, ASA96, BYO95, BT93, BCS93, BW93, BHH^{+}95, CC92a, CT92, CBTM93, CTSN90, CCSN93, DSW^{+}91, DH96, ESDK94, FKC96, FMH96, FB98, FMFH97, FF97, GJC97, GM99, GEW97, JFP95, KRJ^{+}95, Kel98, KS98, KKW^{+}93, KCV94, LBJ^{+}98, LD96, MBM92, MP94, MW97, NTS90, O’C93a, PPFP98, PLFSW97, PLB95, PT94a, PT94b, Pur92, RGS95, Reu98b, RE94a, RE94b, Sha95b, SD94, Sie93, SJMC93, SJCL93, SSP97, SPS98, SCCS98, TGG^{+}95, TDD94, TB93, VMHH90, WA99a, WPG99, WT99b, ZB94]. Use [BJB95, BL95, Cam97, CNGG96, CCGM93, DB96a, FJR^{+}94, Har98, Iri90, JSH99, KJGC96, LBV^{+}99, LMK93, LM91b, MW97, MBB95, NH90, RK95c, RSN90, SW91, Tha92b, TRMW95, TMCE98, TSW97, WWGK95, And97, AH96, AH98, BL94a, BRL97, BKB94, BvKvRB96, Duf92, DD97, Gre96, GA97b, HM91, HM96, JRB97, KW94, KM99b, LK99a, Mee99, Moo92b, OM99b, PC96, PODL99, Put91, RR93, Sal95, SDBW95, SB94, SKPT97, SJC99, Tol98, WB92a]. used [BKS92, BBT^{+}93, LCBD94, LCC97, McG97b]. Ushisher [ZT90]. Using [GGC99, Vis97, AAD91, BCG97, BLKA93, BMHH94, BH92b, BSG^{+}97, BPC98, BLST90, BP93, BH98, COS93, CGM95, Cha98, CK95, CSB94, DEP^{+}96, EGKE92, EJ93, FCC91, FDZ94, FGP98, FL90, GPLB96, HHYD92, HHSvS96, HCS93, HW92, HLP99b, HRFC90, HP98, KRM94, KCV94, LBM^{+}99, LAJ95, LW94, LSL98, LSB95, MR97a, MB98, MC94a, MDTB94, MMBT97, MCFB97, MO97, MKCL94, MBB99a, MA99c, MGEC97, O’D98, PPO^{+}99, PM98, RDPB99, RFNBHG94, RCGSN96, SMF91, SBT97, SCC94, SH98b, SGB90, URHR98, VLKC94, VAR94, VBL^{+}92, VPGV94, WGD97, WT98, Way95, WR91, WCBL94, WTM^{+}96, vR93a]. USVI [RMT91]. utilisation [Bar99, FHZ96, Mac94, RTLG97, WM92]. utilising [MA99a]. Utility [RPM^{+}94]. Utilization [CGA95, CBD96, FW91a, KY97, MZ92, MN92, RFGF92, RH97, WD90, AFP99, AML95, CLM^{+}91, CCH93, CH99, GH94a, GYL93, GWSG99, Har90, HCS90, HAA94, HSB97, IG96, JLD^{+}96, JL96, Kre93, KHBL95, MBK95, NC90, SPS^{+}91, SAH98, TSS93, VH99b, WOM^{+}97]. UV [AB96b, BK90, BP96a, Gar91, GHTH94, HS98a, Les96, LL96, LM97, MP95, NKI^{+}96, PCLÅ97, RR97a, SL97b]. UV-absorbing [HS98a, Les96, RR97a]. UV-B [BK90, GHTH94, Les96, NKI^{+}96, PCLÅ97]. UVA [JMD^{+}99]. UVA- [JMD^{+}99]. UVB [BZSG96, DB96b, JMD^{+}99, NDG^{+}99]. UVB-absorbing [JMD^{+}99].

V [BH92d, CC91b]. vacuoles [Gow93]. Vainola [JKT90a]. Valdez [DF94, JDSB99, vTSD97]. Valenciennes [OC95]. validate [WWGK95]. Validation [Fow90b, CH97a, LT97, WM97, vdVW99]. validity [HRFC90]. Valley [GJ96a]. value [APCL^{+}93, CLM^{+}93, CA98, GAC^{+}97, NT95, PR93b, SM99b]. values [CU99, Fri97, MLP90, TFL95]. vanhoeffeni [DCRG92, UMD92]. vannamei [SlRR94]. var. [Mar91]. Variability [AC96a, EML^{+}93, FD92, GY97, McC94d, MLDM99, NG99, SPS^{+}91, Sog92, SCC98, AAWD97, AHT^{+}90, BCC94, BCNC96, BCAAC96, BCC96, BP99, BM94c, BCA^{+}99, Bou90a, Bou90b, BWM97, BP92, BSH^{+}91a, CO92, CDM^{+}96, CC98, CCSN93, CBL96a, CSB94, DG97, Fev92, FH97a, GHMC96, GJC97, GC95a, GST^{+}91, Gra91c, HS94b, HV91, HM96, HK93, HS98c, HM94b, JCP93, JBWH97, KTH98, KHK^{+}99, KDHZ98, LD93, LD94a, LAJ95, LDKR93, LTM^{+}98, MH96a, MC96a, MZ99, MF90a, MF90b, MMGIM90, MHS94, MS94b, MG94, MP96c, NFM98, PF98, PHHH96, RÁSP99, SO96a, SRH^{+}98, SPBS93, SL97a, SL95, SGMS99, Sos96, TLB93, TRS91, TPW^{+}93, THKN98, Uib95, VMHH90, VM95, WTR99, WHIC97, WWB98, ZF97, ZF99, dG91a]. variable [BWRH99, HGBW99, KBP98, MFD98, MCC96a, MCA90, MBL94]. variables [BMFR97, CA93, CSM99, DFB96, GT98, Joh96, LW94, MC91a, MDTB94, MN99, RZLB96, UCV98, WCBL94]. variance [RE96]. varians [NH98]. variaoility [Arr94]. Variation [ASA96, Cow91, FA99, FT95, HP90, HBD95, HSBP97, LM91a, MW94, MDRW99, NKV94, PWPS95, SO96a, Tak95, WBP^{+}92, WMD99, ÅP97a, ASN93, AW97, ATH95, BPS98, BLDF95b, BLDF95a, BGBS91, BCL93, BP98, CLM^{+}93, CCV^{+}98, CS91, CH99, DDM^{+}99, DCK99, DKVD95, EB96, ELMR98, Fev92, FTS94, FMFH97, Fra93, GBC92, GYL93, Geo94, GT98, HWH98, HFR^{+}93, HHB97, HH95, Hur92, HJHG98, JJ99, JLF95, JRMN^{+}97, JKT90b, JCMW91, JLGMP97, Kal91b, KI97, KV90, KLEH91, KHBL95, Lai92, LH98, Lev91, LBH91, MC91a, MCF90, MGM96, MC91c, MRPC98, MS91, MP92, Mil94b, Nak92, NA99, Pfi97, PCT^{+}97, PH95, QGQ97, RCG98, RR97b, RSB93, RCA^{+}95, RBN99, RCN95, RNH99, SLB^{+}92, SV97, SK95c, TPR98, TL94, TNN93, UK98, VB93, VHHS96, VHH^{+}98, WT98]. variation [WASSM97, Way95, Zaj91a, Zaj91b, ZBW92, dNSR^{+}96]. Variations [DGO92, KWEG94, Kyo92, LFR^{+}97, LBK90, MS93d, AV98, ARDL97, AB90, BR94, BDA^{+}99, BBL93, CBL96b, CWHA99, DSG^{+}99, ES90, FTS97, FPGJ99, FPFR92, FVDL95, GDP^{+}98, GAR93, GM98a, Gru94, HJH^{+}99, HM97b, HNvB94, Ito97, Leh96, LSH96, MVA99, MGC91, OC97, RTRF90, Riv90, RMH96, SK95b, SM97c, Tan98a, Tan98b, TR91, US98, VWS93, WE95, YSL94, YST^{+}97, YKL99]. variegata [ATH95]. variegatus [RSK^{+}99]. various [Sch91, SW99, WYY^{+}98]. Varying [DS93a, AC96c, DBMK99, DF95, KMKS96, MCNB98]. vascular [BVS^{+}91, LBK90]. vector [PCCG97]. vectors [KV90]. vegetated [Kne91, KW94, RHP93]. vegetation [ERC95, NDRP93, SDBW95]. vegetative [CC99]. Vehicles [SCC94]. veliger [HDB95, PCBA94]. Veligers [MGP^{+}96, Har99]. Veliidae [CR98]. velocity [BH97b, Dei90, ES99, FHZ96, LA97a]. velum [KGC92]. velvet [FGG95]. Veneridae [RR92]. Venerupis [APCL^{+}93, SDB92, SDB94]. Venice [LFD95, SSGD96]. vent [ASF98, AAW98, CJ93a, CD98, DWG98, EPDC99, FCMB94, GSM^{+}97, JDLL95, MT94, PSB^{+}97, PDB^{+}97, PDS97, PFMG96, SRG^{+}99, VTGK99]. ventalina [CHFL95]. ventilation [vRV96]. ventral [RCA^{+}95]. vents [PFMC95]. venturi [KW91]. Verification [PPO^{+}99]. vernal [NK92b]. verrillii [MRS96]. versus [ADR95, BH97b, BJTB96, DvdTCC97, DGCT93, GJ97b, GdGMD95, GSS93b, KWL96, LC93, LEHvR96, MSP^{+}97, NK93, RK99, RO97, SR98b, TW97, WFTH97]. vertebrate [LMM98]. vertebrates [NL96]. Vertical [AB95, ASH93, ANW96, AW98, BM94d, CN95, CPT90, FH96, Gar99b, HPPZ93, Hil91b, KMKS96, KS98, MCC97, Olm94, PDBA90, QCM97, Tak96, TCRB99, VK90, VDA96, WPB^{+}93, BGC91, Bar97c, BS92, BG99, BD94, DD90, DJY90, DHKP96, DRP^{+}97, FTS97, FKR90, GM99, GMT99, Gib94, GS94b, GW97, Hay95, HWL96, HHH97, HB98, JWKH98, JBK99, Kur99, LC91, MC95, MGP^{+}96, OD90, OK98, OF94, RLDM94, RWCS98, Röp93, SB98, SFS95, Sea90, SO96b, STS99, Ter97, TBLD90, TD92, TS90b, TS90c, VBCRB97, Vis97, ZDJ^{+}95, ZN96a, MJV92, TD92]. vertically [Hei95, KKH98, PM92a]. very [GCKW95]. vesicomyid [BSP^{+}99]. vesiculosus [AKK94, BCR92, CNJ96, MEK99, PKY96]. vessel [DZ99, ZD99]. Vestfjorden [Hag95]. vestitum [Ste94]. vetula [BvKvRB96]. vetuta [BKB94]. VHF [CLJOC98]. via [ACC90, CC90, CC91b, CLJOC98, FH97a, HKG93, Hal98]. Viability [LGED98, RD91, BSH^{+}91b, CMT94, HMVHH95, MLC97, PHHH96]. Vibrio [KRFL97, KRF^{+}98]. vibriophages [KRJ^{+}95]. vicinity [HD93]. victoria [Pur91]. video [LC93, Tis98a]. videography [WGD97]. Videoscopic [CJ93a]. videosystem [WR98]. Vigo [CMRF93, FP91b, TFFA99, ZF97, ZF99, Pre91]. villosus [DL91a, DL93a, DCBC91]. vincristine [MM96]. Viral [BEH93, BLM^{+}95, MBJ94, MB95a, MJ92b, MJ93, PF91b, SKKO97]. virens [FQV98, LBDD97, MHC92, NEIR95, ORD98, RK95b, RK95c, Vis90]. virgatus [WD99]. Virgin [EW91]. virgini [ZMD95]. Virginia [GM99, MBM92, ATNW97, Blu93, KKW^{+}93, MRS96]. virginica [BM94b, BN91a, BN95, BD96, BV99, BA96, GMC91, RCH98, RM95]. viride [BKB94, BvB94, BBB^{+}94, BvKvRB96, vRBVB95, vRV96]. viridis [Gil97, HSTY98, HTLK99, OM99a]. virioplankton [Mül91]. Virtual [LBK96, Lav97]. Virus [NAI^{+}93a, GS93, Gow93, MJ93, PF92c, WP95]. Virus-like [NAI^{+}93a, Gow93]. virus-rich [WP95]. virus-sized [GS93]. Viruses [BFJ^{+}93, SC95b, BHH^{+}92, BHT^{+}92, CWS^{+}93, CS91, GS93, HKT^{+}96, HB91b, JP94, SSA96, WF94]. viscera [SOK98]. viscosity [BH97b]. viscosity-induced [BH97b]. visibility [TK95b]. visible [AS99a, Gar91]. Visual [DGCT93, LMO99, CT97, CS99, TM97]. vital [ER95b, KW91]. vitamin [GR94, HP90]. vitro [BMV^{+}96, LFD95]. viviparus [FRK92, MGM96]. vivo [TBLD90, BWL^{+}99, JSGS94, MKV95, VMS^{+}96]. volatile [HBBM96]. volcano [ZT90]. volume [BZSG96, DE94, LC90, TPR98, WFTH97]. Volumes [Ano96a, Ano97a, Ano97b, Ano98, Kin93, LGGLLHV98]. voluntary [vdL95]. volutator [CD99, GH93, GH94b, GH94c, MJ97b, PNB94, SHC96]. Vortex [JTPP92]. VOS [HBBM96]. vs [Gas94]. vulnerability [AH96, VKF^{+}98].

W [PLH91, Por94]. Wadden [Arm92, BC90, Beu92, BF95b, Gün91, KLEH91, PGB96, PE99, Phi94, ZBW92, dT94a, vDK90, vBDW91, vW93]. waders [PE99]. Wairarapa [CB99]. Wales [AU93, Cha95, FK95b, KUK91, Sca93]. walk [Vis97]. wall [UGSBA94]. walleye [BM94a, BCN^{+}95, BBK91, CBI91, DO92, HHP95, HBP^{+}93, Liv93, OD90, RO98, YB90]. walls [CPK97, MSK95, NC90]. wandering [WW92a, WWL97]. wane [HHN^{+}96]. warm [DL93b, FP96b, KSH97, NB90, Tak99, Tan98a, Tan98b, ZR92, ZBK^{+}98]. warm-core [DL93b, FP96b, NB90, ZBK^{+}98]. warming [VPM96]. Wash [YBG^{+}98, YBG96]. Washington [KCV94, HDD98]. waste [Pel96, SR93a]. wasting [BSW93]. watch [SM97b]. Water [ADBH99, Bra95, FKR90, HAR95, JM93, KG99, MO95, PLG^{+}96, SGK^{+}97, VD94, WPY98, AAWD97, AMS99, ASCW90, BTKK99, BGT91, BBR^{+}93, BKS92, BG94, BH97b, BCG94, Bra93, BRC99, BH93, BWL^{+}99, BGC94, BMG99, Car93, CM98a, CF94, CRCS96, CCSN93, CV91, CS97, CPB96, CWH98, DBMK99, Dei90, DFR95, DF92, Dem95, DZ99, DAS92, FFT98, FGP98, FH97a, Gar91, GPB96, GWW^{+}98, GSBB97, GHD92, HS94a, Hal98, HW95, HHN^{+}96, HK99, Hil91c, HSN93, HFGT94, JNCS89, KMKS96, Kar92, KNK90, KZK^{+}95, KFN99, KKK^{+}90, KBWC90, Kro91, KR95, LK99a, LA97a, LEHvR96, MM90, MC95, MY98, Mal99, MG96c, MD93, MC98, McC94b, MB97a, McC96b, MM95a, McG94, MB97b, MMG97, Nak92, NJSS91, PPFP99, PBVB99, PCCG97, PLP92, RT95, RGS95, RHS99, RPL96]. water [RM97b, RLT^{+}94, RHP93, RBG94b, SSR99, SA92, Sch96, SFF98, SR93a, Sie93, SJMC93, SJCL93, Sim92, SHM^{+}97, SRW95a, SEGP91, SE93, TDBH94, Tha92a, Tha92b, TD92, TB98, TGB99, TUN99, TL91, TRMW95, TS90c, VWS92, VM95, Vis97, VDA96, VNM97, Wai90, WB92a, WTF96, WSMM99, WF90, Wil92, YST^{+}97, YY91, YTF97, ZD99, ZB99b, vDvRK93, IP94]. water-borne [SHM^{+}97]. water-column [BGC94, KNK90, WSMM99]. water-sediment [TB98]. waterborne [BC93]. waters [AFP99, AA92, ALBS96, AMB^{+}96, Arm92, AMBH98, BK98, BK94, BSY^{+}99, BHT98, BP98, CA99a, CW95, CBD96, CPSA96, CK95, CMA95, CGP97, DHM92, DB96b, FM95a, FC94, GCH93, GH91, GM98b, GTJ92, Gra96b, GA97b, GDW90, HKT^{+}96, HMVHH95, HSRHH90, HS99a, KD97, KSG^{+}92, KW92a, KS91, KG95, Kin92, LC91, LADS98, LAG96, LBLA93, LSL98, LFR^{+}97, MB95a, MG96c, MHT^{+}97, MS91, MRP98, Mik95, MA96, MAdG^{+}96, NPB93, NK93, OIO^{+}95, Øre91, ONR98, PWG^{+}99, PB96b, PH91a, PT95, PGAJ99, PHHH96, PF91a, RLDM94, RWJ99a, RAS97, RZLB96, RAlBW99, SJJMH90, SCB92, SL97a, SD99, SH98a, SJÆ91, SKM95, Spe95, STS99, SFPC97, TW98, TT98, Tho92, UMD92, UT90, VSQ90, WWB98, WR97, WHK97, WK93, YNYK94, YSL94, YFI91, ZN96b]. watershed [JCMW91, MPRB93, PPFP98]. wave [AC96c, DM96, GHL^{+}97, HS96, HS98c, KG99, LM91a, MM91a, SF95, VWM90a, VWM90b, WG97]. wave-dominated [KG99]. wave-exposed [HS96, HS98c]. wave-swept [WG97]. Wavelength [AS99a, BEG97, GJ96b]. Wavelength-dependent [BEG97, GJ96b]. Wavelength-specific [AS99a]. waves [KTW^{+}97, LSGW98, LCF99]. Wax [HHN^{+}96, PDS97, HKG93]. Wax-ester [PDS97]. weak [GGPGG95]. weakfish [CE93, DE94, GT96, RE94a, RE94b, TJST98]. weakly [AWD95, AAWD97]. Weathership [IHK^{+}98, NKH^{+}99]. web [ASN95, BMSHC98, BP92, BRN^{+}95, BGvDR96, CM98a, Han91b, HTG96, HW92, KKK^{+}90, KKT96a, LDV95, LR96, LHK^{+}93, LCK^{+}96, MV98, NJLD98, PH91a, PMG^{+}98, PMOD90, RNP^{+}98, RSB^{+}90, SDFJv91, ST90, SMP97, UHL^{+}97, UMD92, VAD^{+}94, Wai90, WCLB99]. webs [AV95, CNP95, Dav96, DG97, FW91b, LM98b, MJSDF91, PM90b, SM90, TR95a]. Weddell [GTL^{+}91, GST^{+}91, GGD99, Bar95, BH92e, GC93, GP90, HW95, KGW93, LNQT93, MPS95, PGR94, RHT91, RAB^{+}92, SSHM91, SK97, SCKS90, VAV99, Wöh96, dBN^{+}90]. wedge [DdA95]. Weight [RB98, GBA^{+}96, HH94b, HS97b, HL98, KNM93, KE95, MIA90, Mat92, MAAT98, Sal95, SRS92, SKM95, Spr93, WP95]. weight-specific [HS97b]. Weight-to-weight [RB98]. weighting [BP96a, CW99b]. weir [Kne91]. weissflogii [DEP^{+}96]. well [ASP91, DD90]. well-mixed [DD90]. West [AGW93, HS91c, HH91b, PFH98, SC96a, SB96, Tom91, BCC95, BCNC96, BCBC98, BSF95, GGPGG95, Góm95, HK99, NG99, SAG92, TN98, AHY^{+}99, BCv^{+}92, CPT90, HFR^{+}93, PFMG96, vCRS92]. Western [LCLC96, Nai93b, VS91, AMO^{+}99, BCN^{+}95, BA93, BMP^{+}95, BS95a, BCC95, BCBC98, CSC92, CM98a, CCG96, CWH98, DWG98, ES90, FA99, FHA99, FB90, GRC^{+}97, HWH98, HBBT98, HD97, HBP^{+}93, Jer90, KL95, KLEH91, MBF^{+}97, MC97a, MD94b, MMGIM90, MSM^{+}98, NH95b, OIO^{+}95, Oro96, OP99, PWG^{+}99, RDT^{+}94, RR92, SCIA99, SGPAA95, TMKL97, TSB93, TH90, TGvJW95, TGSN99, WT99a, vDK90, vBDW91, CA92, Car93, CS93, CDME97, CS97, EY90, FTS94, HS97a, MCED96, MW99, MSCTBE96, Oes90, PAM96, SMT97]. Westerschelde [LH99, MAH94, SV93a, SH94]. Westwind [Bra95]. wet [PPR94, PPCR97, RDPB99]. wet-oxidation [PPR94, RDPB99]. wetlands [MS97]. WH7803 [LBL^{+}99]. Whale [GFPD92, AS94a, BSP^{+}99, BSGM94, Cer98, GH92, HC92b, JWL96, JW96, KWD93, RCA^{+}95]. whale-fall [BSP^{+}99]. whales [BSF95, BFS^{+}98, CGNH98, COK99, GRC^{+}97, HKMK98, ISF^{+}98, KNI95, KEJ98, Mik97, PCM^{+}95, PZCJ99, PM92b, SPB^{+}99a, Sha95b]. whelk [IWF^{+}97, RMH96]. where [Kin95b, LD95a, OS96]. which [CS91, DHM92, Han91a, JLYK97, SHT91, VDT92]. while [ST98]. White [TSI^{+}93, SlRR94, SP99c]. White-noise-like [TSI^{+}93]. whiting [CGFC95, HK99, JWKH98]. whole [FZ91]. whose [WPB^{+}93]. Wide [CS91, EB96, NW97b]. Wide-spread [CS91]. Widespread [FK95a, FWWL99, FD97, NBS^{+}97, PF90]. widgeongrass [BGC94]. width [CT97]. wild [BCBR97, FCW98, GRSV98, MLP90, SJCL93]. wild-caught [FCW98]. William [IFL97]. Wind [BGW96, ML93, BM94a, CC97b, ELMR98, FM94, ML91a, RD97, RÁSP99, VBCRB97, YGH^{+}97]. Wind-based [ML93]. Wind-driven [BGW96, FM94, RÁSP99, VBCRB97]. winds [YHGB96]. windy [TTP^{+}95]. Winter [BA93, BFS^{+}98, BSH^{+}91a, BSH^{+}91b, GC93, McG97a, Mik95, Nor94, Øre95, SK97, WP92, BNvB90, BCL93, BSW^{+}95, BWHW95, CGFC95, CA98, DRC^{+}97, GCF^{+}95, HKM92, HvdM97, HNL94, Kea90, KFS91, MPS95, NNB93, Nor96, ØW93, TNKMK99, TWWE95, WR97, WB92b, WA95, ZBK^{+}98, vdVW99, Tor94]. Winter-spring [WP92, KFS91]. winter-temperature [HvdM97]. winter/spring [GCF^{+}95]. winters [Beu92]. wisdom [MS97]. witch [MRS96]. within [ASN93, ADG^{+}95, ASSWM97, Bel95, BA97, CN92b, CBS91, DBdCS99, FBHH93, FH96, GF95, HS94b, HW92, HAR95, KR97a, Ken94b, Ken94c, Kne93, KJMRC96, LY97, Mau92, NA99, OKK95, PPP^{+}96, Pfi97, PDP^{+}90, SPBS93, SB97b, SD93a, SHO96, SFPC97, SGM98, VAD^{+}94, VPM96, VWS93]. within-patch [NA99]. within-plant [PPP^{+}96]. within-reef [Bel95]. within-year [Pfi97]. without [FBR^{+}95, VH96]. wolf [MRS96]. wood [WW92b]. work [FAVV98, LSGW98]. working [MBB99a]. Workshop [SD92, SDA^{+}92d, SDA^{+}92e]. world [Kin93]. worldwide [RCA^{+}95]. worm [AGP98, ASF98, HAA94, RPL96]. worms [SS92, SNR^{+}96]. Wounding [DES92]. wrasse [AMK99, Cow91, Gru96, HH91b, MSP^{+}97, MJ97a]. wrightii [BGC94, DT94b, GMR^{+}94, OB93].

X [KPWL93]. X-ray [KPWL93]. xamachana [FFH99, VM98a]. xanthid [LK94]. Xanthidae [LFS90]. Xanthophyll [MOM98, DRGV91]. xanthophyll-cycle [DRGV91]. xanthophylls [BBL93]. xanthurus [SMF91, SR97, SCCS98]. Xenobiotic [CDI96]. xenogeneic [FBR^{+}95]. Xiphias [GCNSNd94].

Yankich [KZK^{+}95]. yaquinae [APS91]. Year [Cam96, vdVW99, AH94, AH99, CBS91, DvdM92, DBB^{+}95, Dee90, DQT^{+}97, FIA93, GAC^{+}97, JC94, JC95, LAK99, MC91c, Pfi97, RCG99a, VHHS96, VHH^{+}98, ZHM91]. Year-class [Cam96, vdVW99]. year-round [DQT^{+}97]. years [BCN^{+}95, BC99, PPCL^{+}98, SRJD97, Tan98a, Tan98b]. Yellow [Flo91, BD99, MG96c, CCCH94]. Yellow-stained [Flo91]. yellowtail [BP99]. Yemen [OB94c]. yessoensis [GF94]. yield [GTJ92, NU91, SPBS93, Sos96]. yields [CLHF95, She94]. Yingluo [ZCL97]. Yoldia [Nak92, Nak94]. York [CSR^{+}92a, MC91c, CC91a, GM99]. young [AH93, AH94, AH99, Cam96, Dee90, FIA93, JC94, JC95, MC91c, PODL99, YB90, ZHM91, RTRP^{+}98, RNH99]. young-of-the-year [AH94, AH99, Dee90, FIA93, JC94, JC95, MC91c, ZHM91]. YOYO [JM99].

YOYO-1 [JM99].

Zausodes [SW91]. Zealand [AM91, BE97, IMB99, CB99, CK98, HJH^{+}99, Kin92, Lam98, LB99a, Mac91, MMBT97, Mil97, MDRW99, RS99a, Sch93, SPB^{+}99b, TC94, TTR99]. zeaxanthin [BCG97]. zinc [CBRD92, CD99, LM95, MAAT98, MBL^{+}98, RK95a, TP96, VB96]. Zn [CAATS94, HB94b, LWL98, VCMP93]. Zoarces [FRK92, MGM96]. zoea [QCM97, ZN96a]. zoea-1 [ZN96a]. Zonaria [Poo94]. Zonation [CS93, BWHC91, FKR90, GY97, GWKW97, JMC93, JA91, MSJ92, OKK95, RM95, TK95a]. Zone [BE97, PPMF97, WSCT96, AWD95, AW98, BA91, CA90, CBL96a, CBL96b, Coc90, DM91, FW91a, FF92b, FKT^{+}99, Har98, IF96, LBL91, MM91b, ME97, MPS95, MLDM99, NC93, NT92, OM99a, RK95b, Sal96, SCKS90, VBF94, VDBF96, WP92, WOM^{+}97, PWR96]. zones [HHM97]. zoobenthic [NB96c]. zoobenthos [NB96b, San94a]. zooflagellates [ML90]. Zooplanktivory [VR98]. Zooplankton [ASP^{+}96, SL91, And90, BP97b, BTYO94, BYO95, BA93, BSVK93, BH92a, BMTF92, BPH95, CT92, CLGM96, CG99b, CWH98, DPO91, FAVV98, FPPM96, FQ94, GGB^{+}99, GS94b, GGB94, GBP^{+}97, HWL96, HHH97, HL98, JAOS90, KGO94, KW92b, LD94a, LGGLLHV98, LF92, LF94, Lew92, LGH95b, LP99, ML91a, MSD98, MR97a, MA99c, MHM90, NPB94, Naj97, NKGK92, NIU^{+}98, Ole95, PG91, PGG96, PML93, PM92a, Pur92, RH98, RHrsHG99, SBV98, SR98b, SM97a, SFP91, SV93a, SH94, SSP97, SPS98, UMD92, WR91, WR92b, YBLL96, vdL98, BH92c]. zoospore [BWB98]. Zooxanthellae [Jon97b, BBLY98, BZ92, GK96b, KCD^{+}95, RGLS99, RP91b, TTL^{+}96, VM98a]. zooxanthellar [Jon97a]. zooxanthellate [FPGC^{+}98, FPGJ99, GYL93, JLGMP97, PMB99]. Zostera [AZKA99, BL94a, BSW93, BDM96, BGC94, CK90, FB91, FKC^{+}90, FMFH97, GNJ90, HHN^{+}96, HN97b, HS99b, IF96, KA93, KM99b, MRPC98, MNO96, OSJ93, OSJ94a, OSJ94b, Ole96, PB92, PB93, Phi94, RS99a, RCG94, Reu98a, RPDR^{+}98, TS99, VSSM98, ZA96, vKVSR97, vKSGvA99]. zschaui [Moo94].

References

Almeida:1992:BCS

[AA92] M. A. Almeida and F. Alcantara. Bacterial colonization of seston particles in brackish waters (Ria de Aveiro, Portugal). Marine Ecology Progress Series, 89:165–173, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

               tronic). URL https://www.int-res.com/articles/meps/
               89/m089p165.pdf.

Aagaard:1991:SMP

[AAD91] A. Aagaard, B. B. Andersen, and M. H. Depledge. Simultaneous monitoring of physiological and behavioural activity in marine organisms using non-invasive, computer-aided technigues. Marine Ecology Progress Series, 73:277–282, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
73/m073p277.pdf.

Arce:1997:ASC

[AADSCC97] A. M. Arce, W. Aguilar-Dávila, E. Sosa-Cordero, and J. F. Caddy. Artificial shelters (casitas) as habitats for juvenile spiny lobsters Panulirus argus in the Mexican Caribbean. Marine Ecology Progress Series, 158:217–224, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
158/m158p217.pdf.

Altstatt:1996:RDB

[AAE^{+}96] J. M. Altstatt, R. F. Ambrose, J. M. Engle, P. L. Haaker, K. D. Lafferty, and P. T. Raimondi. Recent declines of black abalone Haliotis cracherodii on the mainland coast of central California. Marine Ecology Progress Series, 142:185– 192, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/142/m142p185.pdf.

Arai:1999:SCI

[AALT99] T. Arai, J. Aoyama, D. Limbong, and K. Tsukamoto. Species composition and inshore migration of the tropical eels Anguilla spp. recruiting to the estuary of the Poigar River, Sulawesi Island. Marine Ecology Progress Series, 188:299– 303, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/188/m188p299.pdf.

Adams:1991:EEF

[AAS^{+}91] N. J. Adams, R. W. Abrams, W. R. Siegfried, K. A. Nagy, and I. R. Kaplan. Energy expenditure and food consumption by breeding Cape gannets Morus capensis. Marine Ecology Progress Series, 70:1–9, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/70/m070p001.pdf.

Atkins:1998:EHP

[AAW98] M. S. Atkins, O. R. Anderson, and C. O. Wirsen. Effect of hydrostatic pressure on the growth rates and encystment of flagellated protozoa isolated from a deep-sea hydrothermal vent and a deep shelf region. Marine Ecology Progress Series, 171:85–95, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/171/m171p085.pdf.

Abello:1997:SVP

[AAWD97] P. Abelló, A. Aagaard, C. G. Warman, and M. H. Depledge. Spatial variability in the population structure of the shore crab Carcinus maenas (Crustacea: Brachyura) in a shallowwater, weakly tidal fjord. Marine Ecology Progress Series, 147:97–103, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/147/m147p097.pdf.

Allen:1990:IVL

[AB90] D. M. Allen and D. L. Barker. Interannual variations in larval fish recruitment to estuarine epibenthic habitats. Marine Ecology Progress Series, 63:113–125, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
63/m063p113.pdf.

Ardisson:1992:LSE

[AB92] P.-L. Ardisson and E. Bourget. Large-scale ecological patterns: discontinuous distribution of marine benthic epifauna. Marine Ecology Progress Series, 83:15–34, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
83/m083p015.pdf.

Alve:1995:VMR

[AB95] E. Alve and J. M. Bernhard. Vertical migratory response of benthic Foraminifera to controlled oxygen concentrations in an experimental mesocosm. Marine Ecology Progress Series, 116:137–151, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/116/m116p137.pdf.

Archambault:1996:SCH

[AB96a] P. Archambault and E. Bourget. Scales of coastal heterogeneity and benthic intertidal species richness, diversity and abundance. Marine Ecology Progress Series, 136:111–121, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
136/m136p111.pdf.

Arrigo:1996:ICD

[AB96b] K. R. Arrigo and C. W. Brown. Impact of chromophoric dissolved organic matter on UV inhibition of primary productivity in the sea. Marine Ecology Progress Series, 140: 207–216, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/140/m140p207.pdf.

Ambarsari:1997:FAC

[ABB^{+}97] I. Ambarsari, B. E. Brown, R. G. Barlow, G. Britton, and D. Cummings. Fluctuations in algal chlorophyll and carotenoid pigments during solar bleaching in the coral Goniastrea aspera at Phuket, Thailand. Marine Ecology Progress Series, 159:303–307, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/159/m159p303.pdf.

Aaberg:1990:MSC

[Åbe90] P. Åberg. Measuring size and choosing category size for a transition matrix study of the seaweed Ascophyllum nodosum. Marine Ecology Progress Series, 63:281–287, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
63/m063p281.pdf.

Aaberg:1996:PRE

[Åbe96] P. Åberg. Patterns of reproductive effort in the brown alga Ascophyllum nodosum. Marine Ecology Progress Series, 138: 199–207, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/138/m138p199.pdf.

Ahel:1996:ESG

[ABM96] M. Ahel, R. G. Barlow, and R. F. C. Mantoura. Effect of salinity gradients on the distribution of phytoplankton pigments in a stratified estuary. Marine Ecology Progress Series, 143:289–295, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/143/m143p289.pdf.

Aarnio:1998:RHS

[ABN98] K. Aarnio, E. Bonsdorff, and A. Norkko. Role of Halicryptus spinulosus (Priapulida) in structuring meiofauna and settling macrofauna. Marine Ecology Progress Series, 163:145– 153, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/163/m163p145.pdf.

Arifin:1997:FRC

[ABY97] Z. Arifin and L. I. Bendell-Young. Feeding response and carbon assimilation by the blue mussel Mytilus trossulus exposed to environmentally relevant seston matrices. Marine Ecology Progress Series, 160:241–253, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/160/m160p241.pdf.

Agusti:1992:RBL

[AC92] S. Agusti and A. Cruzado. Relationship between light absorption by pelagic particles and microplankton metabolic activity in the Gulf of Lions. Marine Ecology Progress Series, 85:283–287, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/85/m085p283.pdf.

Airoldi:1996:VFP

[AC96a] L. Airoldi and F. Cinelli. Variability of fluxes of particulate material in a submarine cave with chemolithoautotrophic inputs of organic carbon. Marine Ecology Progress Series, 139:205–217, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/139/m139p205.pdf.

Almers:1996:EHA

[AC96b] H. Almers and T. Cedhagen. Ecology and hydrodynamic adaptations of the large foraminiferan Discobotellina biperforata (Hemisphaeramminidae). Marine Ecology Progress Series, 140:179–186, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/140/m140p179.pdf.

Alvarado:1996:TMM

[AC96c] J. L. Alvarado and J. C. Castilla. Tridimensional matrices of mussels Perumytilus purpuratus on intertidal platforms with varying wave forces in central Chile. Marine Ecology Progress Series, 133:135–141, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/133/m133p135.pdf.

Anderson:1999:PFI

[AC99] M. J. Anderson and S. D. Connell. Predation by fish on intertidal oysters. Marine Ecology Progress Series, 187:203– 211, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/187/m187p203.pdf.

Antia:1990:SSP

[ACC90] A. N. Antia, E. J. Carpenter, and J. Chang. Species-specific phytoplankton growth rates via diel DNA synthesis cycles. III. Accuracy of growth rate measurement in the dinoflagellate Prorocentrum minimum. Marine Ecology Progress Series, 63:273–279, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/63/m063p273.pdf.

Ackman:1997:EHC

[Ack97] R. G. Ackman. 2-ethyl-1-hexanol: contaminant or sex pheromone in Arenicola marina? Marine Ecology Progress Series, 152:311–312, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/152/m152p311.pdf.

Anderson:1997:HAK

[ACLS97] R. J. Anderson, P. Carrick, G. J. Levitt, and A. Share. Holdfasts of adult kelp Ecklonia maxima provide refuges from grazing for recruitment of juvenile kelps. Marine Ecology Progress Series, 159:265–273, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/159/m159p265.pdf.

Anil:1995:ITS

[ACOK95] A. C. Anil, K. Chiba, K. Okamoto, and H. Kurokura. Influence of temperature and salinity on larval development of Balanus amphitrite: implications in fouling ecology. Marine Ecology Progress Series, 118:159–166, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/118/m118p159.pdf.

Ansell:1999:DMD

[ACR99] A. D. Ansell, C. A. Comely, and L. Robb. Distribution, movements and diet of macrocrustaceans on a Scottish sandy beach with particular reference to predation on juvenile fishes. Marine Ecology Progress Series, 176:115–130, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
176/m176p115.pdf.

Ang:1990:MMA

[AD90] P. O. Ang, Jr. and R. E. De Wreede. Matrix models for algal life history stages. Marine Ecology Progress Series, 59: 171–181, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/59/m059p171.pdf.

Ang:1992:DDP

[AD92] P. O. Ang, Jr. and R. E. De Wreede. Density-dependence in a population of Fucus distichus. Marine Ecology Progress Series, 90:169–181, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/90/m090p169.pdf.

Ang:1993:SAD

[AD93] P. O. Ang, Jr. and R. E. De Wreede. Simulation and analysis of the dynamics of a Fucus distichus (Phaeophyceae, Fucales) population. Marine Ecology Progress Series, 93:253– 265, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/93/m093p253.pdf.

Arias:1994:SPB

[AD94] A. M. Arias and P. Drake. Structure and production of the benthic macroinvertebrate community in a shallow lagoon in the Bay of Cadiz. Marine Ecology Progress Series, 115:151–167, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/115/m115p151.pdf.

Agawin:1998:GAS

[ADA98] N. S. R. Agawin, C. M. Duarte, and S. Agustí. Growth and abundance of Synechococcus sp. in a Mediterranean Bay: seasonality and relationship with temperature. Marine Ecology Progress Series, 170:45–53, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/170/m170p045.pdf.

Acuna:1999:LSI

[ADBH99] J. L. Acuña, D. Deibel, A. B. Bochdansky, and E. Hatfield. In situ ingestion rates of appendicularian tunicates in the Northeast Water Polynya (NE Greenland). Marine Ecology Progress Series, 186:149–160, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/186/m186p149.pdf.

Addison:1992:BMS

[Add92] R. F. Addison. Biochemical measurements: summary. Marine Ecology Progress Series, 91:61–63, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/91/m091p061.pdf.

Agawin:1996:NLP

[ADF96] N. S. R. Agawin, C. M. Duarte, and M. D. Fortes. Nutrient limitation of Philippine seagrasses (Cape Bolinao, NW Philippines): in situ experimental evidence. Marine Ecology Progress Series, 138:233–243, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/138/m138p233.pdf.

Arrigo:1995:HRS

[ADG^{+}95] K. R. Arrigo, G. Dieckmann, M. Gosselin, D. H. Robinson, C. H. Fritsen, and C. W. Sullivan. High resolution study of the platelet ice ecosystem in McMurdo Sound, Antarctica: biomass, nutrient, and production profiles within a dense microalgal bloom. Marine Ecology Progress Series, 127:255– 268, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/127/m127p255.pdf.

Adjeroud:1997:FIS

[Adj97] M. Adjeroud. Factors influencing spatial patterns on coral reefs around Moorea, French Polynesia. Marine Ecology Progress Series, 159:105–119, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/159/m159p105.pdf.

Algarra:1991:ELQ

[AdN91] P. Algarra, G. de la Vina, and J. Niell. Effects of light quality and irradiance level interactions on short-term pigment response of the red alga Corallina elongate. Marine Ecology Progress Series, 74:27–32, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/74/m074p027.pdf.

Allali:1994:CCO

[ADR94] K. Allali, J. Dolan, and F. Rassoulzadegan. Culture characteristics and orthophosphate excretion of a marine oligotrich ciliate, Strombidium sulcatum, fed heat-killed bacteria. Marine Ecology Progress Series, 105:159–165, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
105/m105p159.pdf.

Alcoverro:1995:AGD

[ADR95] T. Alcoverro, C. M. Duarte, and J. Romero. Annual growth dynamics of Posidonia oceanica: contribution of large-scale versus local factors to seasonality. Marine Ecology Progress Series, 120:203–210, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/120/m120p203.pdf.

Aksnes:1991:TMN

[AE91] D. L. Aksnes and J. K. Egge. A theoretical model for nutrient uptake in phytoplankton. Marine Ecology Progress Series, 70:65–72, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/70/m070p065.pdf.

Aerts:1998:SCI

[Aer98] L. A. M. Aerts. Sponge/coral interactions in Caribbean reefs: analysis of overgrowth patterns in relation to species identity and cover. Marine Ecology Progress Series, 175:241– 249, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/175/m175p241.pdf.

Airoldi:1996:SDM

[AFC96] L. Airoldi, M. Fabiano, and F. Cinelli. Sediment deposition and movement over a turf assemblage in a shallow rocky coastal area of the Ligurian Sea. Marine Ecology Progress Series, 133:241–251, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/133/m133p241.pdf.

Aguilar:1999:DAC

[AFP99] C. Aguilar, M. L. Fogel, and H. W. Paerl. Dynamics of atmospheric combined inorganic nitrogen utilization in the coastal waters off North Carolina. Marine Ecology Progress Series, 180:65–79, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/180/m180p065.pdf.

Arbones:1996:DPA

[AFZ96] B. Arbones, F. G. Figueiras, and M. Zapata. Determination of phytoplankton absorption coefficient in natural seawater samples: evidence of a unique equation to correct the pathlength amplification on glass-fiber filters. Marine Ecology Progress Series, 137:293–304, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/137/m137p293.pdf.

Aguilera:1996:ISC

[AGGKA96] A. Aguilera, S. Gonzalez-Gil, B. A. Keafer, and D. M. Anderson. Immunomagnetic separation of cells of the toxic dinoflagellate Alexandrium fundyense from natural plankton samples. Marine Ecology Progress Series, 143:255–269, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
143/m143p255.pdf.

Abele:1998:TFS

[AGP98] D. Abele, H. Großpietsch, and H. O. Pörtner. Temporal fluctuations and spatial gradients of environmental P_{O2}, temperature, H_{2}O_{2} and H_{2}S in its intertidal habitat trigger enzymatic antioxidant protection in the capitellid worm Heteromastus filiformis. Marine Ecology Progress Series, 163: 179–191, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/163/m163p179.pdf.

Agard:1993:AMM

[AGW93] J. B. R. Agard, J. Gobin, and R. M. Warwick. Analysis of marine macrobenthic community structure in relation to pollution, natural oil seepage and seasonal disturbance in a tropical environment (Trinidad, West Indies). Marine Ecology Progress Series, 92:233–243, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/92/m092p233.pdf.

Arrhenius:1993:FCL

[AH93] F. Arrhenius and S. Hansson. Food consumption of larval, young and adult herring and sprat in the Baltic Sea. Marine Ecology Progress Series, 96:125–137, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
96/m096p125.pdf.

Arrhenius:1994:LSF

[AH94] F. Arrhenius and S. Hansson. In situ food consumption by young-of-the-year Baltic Sea herring Clupea harengus: a test of predictions from a bioenergetics model. Marine Ecology Progress Series, 110:145–149, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/110/m110p145.pdf.

Aronson:1995:TEH

[AH95] R. B. Aronson and K. L. Heck, Jr. Tethering experiments and hypothesis testing in ecology. Marine Ecology Progress Series, 121:307–309, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/121/m121p307.pdf.

Arsenault:1996:SRC

[AH96] D. J. Arsenault and J. H. Himmelman. Size-related changes in vulnerability to predators and spatial refuge use by juvenile Iceland scallops Chlamys islandica. Marine Ecology Progress Series, 140:115–122, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/140/m140p115.pdf.

Arsenault:1998:SRD

[AH98] D. J. Arsenault and J. H. Himmelman. Size-related decrease in spatial refuge use by Iceland scallops Chlamys islandica: ontogenetic behavioural changes or decreasing refuge availability? Marine Ecology Progress Series, 162:153–161, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
162/m162p153.pdf.

Arrhenius:1999:GBS

[AH99] F. Arrhenius and S. Hansson. Growth of Baltic Sea young-of-the-year herring Clupea harengus is resource limited. Marine Ecology Progress Series, 191:295–299, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
191/m191p295.pdf.

Anger:1990:ESL

[AHMd90] K. Anger, J. Harms, M. Montú, and C. de Bakker. Effects of salinity on the larval development of a semiterrestrial tropical crab, Sesarma angustipes (Decapoda: Grapsidae). Marine Ecology Progress Series, 62:89–94, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/62/m062p089.pdf.

Ayre:1997:GDR

[AHS97] D. J. Ayre, T. P. Hughes, and R. J. Standish. Genetic differentiation, reproductive mode, and gene flow in the brooding coral Pocillopora damicornis along the Great Barrier Reef, Australia. Marine Ecology Progress Series, 159:175– 187, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/159/m159p175.pdf.

Anderson:1990:CTM

[AHT^{+}90] B. S. Anderson, J. W. Hunt, S. L. Turpen, A. R. Coulon, and M. Martin. Copper toxicity to microscopic stages of giant kelp Macrocystis pyrifera: interpopulation comparisons and temporal variability. Marine Ecology Progress Series, 68: 147–156, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/68/m068p147.pdf.

Aoyama:1999:SDM

[AHY^{+}99] J. Aoyama, K. Hissmann, T. Yoshinaga, S. Sasai, T. Uto, and H. Ueda. Swimming depth of migrating silver eels Anguilla japonica released at seamounts of the West Mariana Ridge, their estimated spawning sites. Marine Ecology Progress Series, 186:265–269, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/186/m186p265.pdf.

Ambrose:1992:HAS

[AI92] W. G. Ambrose, Jr. and E. A. Irlandi. Height of attachment on seagrass leads to trade-off between growth and survival in the bay scallop Argopecten irradians. Marine Ecology Progress Series, 90:45–51, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/90/m090p045.pdf.

Abarzua:1995:BIP

[AJ95] S. Abarzua and S. Jakubowski. Biotechnological investigation for the prevention of biofouling. I. Biological and biochemical principles for the prevention of biofouling. Marine Ecology Progress Series, 123:301–312, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/123/m123p301.pdf.

Andre:1993:PSB

[AJL93] C. Andre, P. R. Jonsson, and M. Lindegarth. Predation on settling bivalve larvae by benthic suspension feeders: the role of hydrodynamics and larval behaviour. Marine Ecology Progress Series, 97:183–192, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/97/m097p183.pdf.

Aaser:1995:SDM

[AJS95] H. F. Aaser, E. Jeppesen, and M. Søndergaard. Seasonal dynamics of the mysid Neomysis integer and its predation on the copepod Eurytemora affinis in a shallow hypertrophic brackish lake. Marine Ecology Progress Series, 127:47–56, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
127/m127p047.pdf.

Arnosti:1998:TDM

[AJST98] C. Arnosti, B. B. Jørgensen, J. Sagemann, and B. Thamdrup. Temperature dependence of microbial degradation of organic matter in marine sediments: polysaccharide hydrolysis, oxygen consumption, and sulfate reduction. Marine Ecology Progress Series, 165:59–70, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/165/m165p059.pdf.

Andersson:1994:CLG

[AKK94] S. Andersson, L. Kautsky, and A. Kalvas. Circadian and lunar gamete release in Fucus vesiculosus in the atidal Baltic Sea. Marine Ecology Progress Series, 110:195–201, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
110/m110p195.pdf.

Aguilera:1999:ESR

[AKL^{+}99] J. Aguilera, U. Karsten, H. Lippert, B. Vögele, E. Philipp, D. Hanelt, and C. Wiencke. Effects of solar radiation on growth, photosynthesis and respiration of marine macroalgae from the Arctic. Marine Ecology Progress Series, 191: 109–119, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/191/m191p109.pdf.

Adachi:1999:PCF

[AKM^{+}99] M. Adachi, T. Kanno, T. Matsubara, T. Nishijima, S. Itakura, and M. Yamaguchi. Promotion of cyst formation in the toxic dinoflagellate Alexandrium (Dinophyceae) by natural bacterial assemblages from Hiroshima Bay, Japan. Marine Ecology Progress Series, 191:175–185, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/191/m191p175.pdf.

Aksnes:1996:NMF

[Aks96] D. L. Aksnes. Natural mortality, fecundity and development time in marine planktonic copepods — implications of behaviour. Marine Ecology Progress Series, 131:315–316, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
131/m131p315.pdf.

Archer:1996:MES

[ALB^{+}96] S. D. Archer, R. J. G. Leakey, P. H. Burkill, M. A. Sleigh, and C. J. Appleby. Microbial ecology of sea ice at a coastal Antarctic site: community composition, biomass and temporal change. Marine Ecology Progress Series, 135:179–195, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/135/m135p179.pdf.

Archer:1996:MDC

[ALBS96] S. D. Archer, R. J. G. Leakey, P. H. Burkill, and M. A. Sleigh. Microbial dynamics in coastal waters of East Antarctica: herbivory by heterotrophic dinoflagellates. Marine Ecology Progress Series, 139:239–255, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/139/m139p239.pdf.

Alongi:1990:ABM

[Alo90] D. M. Alongi. Abundances of benthic microfauna in relation to outwelling of mangrove detritus in a tropical coastal region. Marine Ecology Progress Series, 63:53–63, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
63/m063p053.pdf.

Arai:1999:MIM

[ALOT99] T. Arai, D. Limbong, T. Otake, and K. Tsukamoto. Metamorphosis and inshore migration of tropical eels Anguilla spp. in the Indo–Pacific. Marine Ecology Progress Series, 182:283–293, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/182/m182p283.pdf.

Andrew:1991:IBS

[AM91] N. L. Andrew and A. B. MacDiarmid. Interrelations between sea urchins and spiny lobsters in northeastern New Zealand. Marine Ecology Progress Series, 70:211–222, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
70/m070p211.pdf.

Aoki:1993:SPJ

[AM93] I. Aoki and T. Murayama. Spawning pattern of the Japanese sardine Sardmops melanostictus off southern Kyushu and Shikoku southwestern Japan. Marine Ecology Progress Series, 97:127–134, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/97/m097p127.pdf.

Ambrose:1993:EPD

[Amb93] W. G. Ambrose, Jr. Effects of predation and disturbance by ophiuroids on soft-bottom community structure in Oslofjord: results of a mesocosm study. Marine Ecology Progress Series, 97:225–236, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/97/m097p225.pdf.

Aristegui:1996:PPP

[AMB^{+}96] J. Arístegui, M. F. Montero, S. Ballesteros, G. Basterretxea, and K. van Lenning. Planktonic primary production and microbial respiration measured by ^{14}C assimilation and dissolved oxygen changes in coastal waters of the Antarctic Peninsula during austral summer: implications for carbon flux studies. Marine Ecology Progress Series, 132:191–201, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/132/m132p191.pdf.

Amsler:1999:AFT

[AMB99] C. D. Amsler, J. B. McClintock, and B. J. Baker. An Antarctic feeding triangle: defensive interactions between macroalgae, sea urchins, and sea anemones. Marine Ecology Progress Series, 183:105–114, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/183/m183p105.pdf.

Arnold:1998:RBS

[AMBH98] W. S. Arnold, D. C. Marelli, C. P. Bray, and M. M. Harrison. Recruitment of bay scallops Argopecten irradians in Floridan Gulf of Mexico waters: scales of coherence. Marine Ecology Progress Series, 170:143–157, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/170/m170p143.pdf.

Arvedlund:1999:HRP

[AMFB99] M. Arvedlund, M. I. McCormick, D. G. Fautin, and M. Bildsøe. Host recognition and possible imprinting in the anemonefish Amphiprion melanopus (Pisces: Pomacentridae). Marine Ecology Progress Series, 188:207–218, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
188/m188p207.pdf.

Arnal:1999:PCW

[AMK99] C. Arnal, S. Morand, and M. Kulbicki. Patterns of cleaner wrasse density among three regions of the Pacific. Marine Ecology Progress Series, 177:213–220, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
177/m177p213.pdf.

Auster:1995:PMU

[AML95] P. J. Auster, R. J. Malatesta, and S. C. LaRosa. Patterns of microhabitat utilization by mobile megafauna on the southern New England (USA) continental shelf and slope. Marine Ecology Progress Series, 127:77–85, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/127/m127p077.pdf.

Atema:1991:SET

[AMMG91] J. Atema, P. A. Moore, L. P. Madin, and G. A. Gerhardt. Subnose-1. electrochemical tracking of odor plumes at 900 m beneath the ocean surface. Marine Ecology Progress Series, 74:303–306, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/74/m074p303.pdf.

Aoyama:1999:DDA

[AMO^{+}99] J. Aoyama, N. Mochioka, T. Otake, S. Ishikawa, Y. Kawakami, P. Castle, M. Nishida, and K. Tsukamoto. Distribution and dispersal of anguillid leptocephali in the western Pacific Ocean revealed by molecular analysis. Marine Ecology Progress Series, 188:193–200, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/188/m188p193.pdf.

Able:1999:HQS

[AMS99] K. W. Able, J. P. Manderson, and A. L. Studholme. Habitat quality for shallow water fishes in an urban estuary: the effects of man-made structures on growth. Marine Ecology Progress Series, 187:227–235, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/187/m187p227.pdf.

Andersen:1997:HRE

[AN97] C. M. Andersen and T. G. Nielsen. Hatching rate of the eggcarrying estuarine copepod Eurytemora affinis. Marine Ecology Progress Series, 160:283–289, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/160/m160p283.pdf.

Anderson:1990:SDI

[And90] J. T. Anderson. Seasonal development of invertebrate zooplankton on the Flemish Cap. Marine Ecology Progress Series, 67:127–140, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/67/m067p127.pdf.

Anderson:1994:FEC

[And94a] J. T. Anderson. Feeding ecology and condition of larval and pelagic juvenile redfish Sebastes spp. Marine Ecology Progress Series, 104:211–226, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/104/m104p211.pdf.

Anderson:1994:RMS

[And94b] T. W. Anderson. Role of macroalgal structure in the distribution and abundance of a temperate reef fish. Marine Ecology Progress Series, 113:279–290, 1994. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/113/m113p279.pdf.

Andersen:1996:FOC

[And96] F. Ø. Andersen. Fate of organic carbon added as diatom cells to oxic and anoxic marine sediment microcosms. Marine Ecology Progress Series, 134:225–233, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/134/m134p225.pdf.

Anderson:1997:HSS

[And97] T. J. Anderson. Habitat selection and shelter use by Octopus tetricus. Marine Ecology Progress Series, 150:137–148, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
150/m150p137.pdf.

Andrews:1998:RRI

[And98] R. D. Andrews. Remotely releasable instruments for monitoring the foraging behaviour of pinnipeds. Marine Ecology Progress Series, 175:289–294, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/175/m175p289.pdf.

Ang:1991:ASD

[Ang91a] P. O. Ang, Jr. Age- and size-dependent growth and mortality in a population of Fucus distichus. Marine Ecology Progress Series, 78:173–187, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/78/m078p173.pdf.

Ang:1991:NDF

[Ang91b] P. O. Ang, Jr. Natural dynamics of a Fucus distichus (Phaeophyceae, Fucales) population: reproduction and recruitment. Marine Ecology Progress Series, 78:71–85, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
78/m078p071.pdf.

Anger:1991:ETS

[Ang91c] K. Anger. Effects of temperature and salinity on the larval development of the Chinese mitten crab Eriocheir sinensis (Decapoda: Grapsidae). Marine Ecology Progress Series, 72:103–110, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/72/m072p103.pdf.

Ang:1992:CRF

[Ang92] P. O. Ang, Jr. Cost of reproduction in Fucus distichus. Marine Ecology Progress Series, 89:25–35, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
89/m089p025.pdf.

Anger:1995:DBAa

[Ang95a] K. Anger. Developmental biology of Armases miersii (Grapsidae), a crab breeding in supratidal rock pools. I. Facultative lecithotrophy of larval stages. Marine Ecology Progress Series, 117:75–81, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/117/m117p075.pdf.

Anger:1995:DBAb

[Ang95b] K. Anger. Developmental biology of Armases miersii (Grapsidae), a crab breeding in supratidal rock pools. II. Food limitation in the nursery habitat and larval cannibalism. Marine Ecology Progress Series, 117:83–89, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/117/m117p083.pdf.

Anonymous:1994:MDO

[Ano94] Anonymous. In memoriam: Dr Orest A. Scarlato (1920– 1994). Marine Ecology Progress Series, 114:209, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
114/m114p209.pdf.

Anonymous:1995:E

[Ano95] Anonymous. Editorial. Marine Ecology Progress Series, ??:??, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Anonymous:1996:CAT

[Ano96a] Anonymous. Combined author and title index: Volumes 131– 140 (1996). Marine Ecology Progress Series, 140:299–310, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Anonymous:1996:O

[Ano96b] Anonymous. Obituary. Marine Ecology Progress Series, ??:??, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Anonymous:1997:CATa

[Ano97a] Anonymous. Combined author and title index: Volumes 141– 150 (1996–1997). Marine Ecology Progress Series, 150:297– 308, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Anonymous:1997:CATb

[Ano97b] Anonymous. Combined author and title index: Volumes 151– 160 (1997). Marine Ecology Progress Series, 160:295–307, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Anonymous:1997:E

[Ano97c] Anonymous. Editorial. Marine Ecology Progress Series, ??:??, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Anonymous:1998:CAT

[Ano98] Anonymous. Combined author and title index: Volumes 161–170 (1997–1998). Marine Ecology Progress Series, 170: 299–311, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/170/m170p299.pdf.

Anonymous:1999:CSC

[Ano99a] Anonymous. Campana SE chemistry and composition of fish otoliths: pathways, mechanisms and applications. Marine Ecology Progress Series, 188:263–297, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
188/m188p263.pdf.

Anonymous:1999:CAT

[Ano99b] Anonymous. Combined author and title index: 171–180 (1998–1999). Marine Ecology Progress Series, 180:301–312, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Anonymous:1999:KEP

[Ano99c] Anonymous. Kinne O electronic publishing in science: changes and risks. Marine Ecology Progress Series, 180: 1–5, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/180/m180p001.pdf.

Anonymous:1999:MPC

[Ano99d] Anonymous. Mumby PJ can Caribbean coral populations be modelled at metapopulation scales? Marine Ecology Progress Series, 180:275–288, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/180/m180p275.pdf.

Anonymous:2000:ELU

[Ano00] Anonymous. Erratum: López-Urrutia, Á., Acuña, J. L., 191:195–205 (1999). Marine Ecology Progress Series, ??:??, 2000. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). See [LUA99].

Andreassen:1996:VPF

[ANW96] I. Andreassen, E. M. Nöthig, and P. Wassmann. Vertical particle flux on the shelf off northern Spitsbergen, Norway. Marine Ecology Progress Series, 137:215–228, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
137/m137p215.pdf.

Abele-Oeschger:1991:PSC

[AO91] D. Abele-Oeschger. Potential of some carotenoids in two recent sediments of Kiel Bight as biogenic indicators of phytodetritus. Marine Ecology Progress Series, 70:83–92, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
70/m070p083.pdf.

Abele-Oeschger:1994:BAN

[AOOT94] D. Abele-Oeschger, R. Oeschger, and H. Theede. Biochemical adaptations of Nereis diversicolor (Polychaeta) to temporarily increased hydrogen peroxide levels in intertidal sandflats. Marine Ecology Progress Series, 106:101–110, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
106/m106p101.pdf.

Arai:1997:DCO

[AOT97] T. Arai, T. Otake, and K. Tsukamoto. Drastic changes in otolith microstructure and microchemistry accompanying the onset of metamorphosis in the Japanese eel Anguilla japonica. Marine Ecology Progress Series, 161:17–22, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
161/m161p017.pdf.

Aaberg:1997:TMS

[ÅP97a] P. Åberg and H. Pavia. Temporal and multiple scale spatial variation in juvenile and adult abundance of the brown alga Ascophyllum nodosum. Marine Ecology Progress Series, 158:111–119, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/158/m158p111.pdf.

Avila:1997:CEN

[AP97b] C. Avila and V. J. Paul. Chemical ecology of the nudibranch Glossodoris pallida: is the location of diet-derived metabolites important for defense? Marine Ecology Progress Series, 150:171–180, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/150/m150p171.pdf.

Achituv:1999:CGP

[AP99a] Y. Achituv and M. L. Pedrotti. Costs and gains of porcelain crab suspension feeding in different flow conditions. Marine Ecology Progress Series, 184:161–169, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
184/m184p161.pdf.

Anderson:1999:CRG

[AP99b] P. J. Anderson and J. F. Piatt. Community reorganization in the Gulf of Alaska following ocean climate regime shift.

Marine Ecology Progress Series, 189:117–123, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
189/m189p117.pdf.

Albentosa:1993:NVA

[APCL^{+}93] M. Albentosa, A. Perez-Camacho, U. Labarta, R. Beiras, and M. J. Fernandez-Reiriz. Nutritional value of algal diets to clam spat Venerupis pullastra. Marine Ecology Progress Series, 97:261–269, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/97/m097p261.pdf.

Armstrong:1991:TCF

[APS91] J. D. Armstrong, I. G. Priede, and K. L. Smith, Jr. Temporal change in foraging behaviour of the fish Coryphaenoides (Nematonurus) yaquinae in the central North Pacific. Marine Ecology Progress Series, 76:195–199, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
76/m076p195.pdf.

Attrill:1999:MEC

[APT99] M. J. Attrill, M. Power, and R. M. Thomas. Modelling estuarine crustacea population fluctuations in response to physico-chemical trends. Marine Ecology Progress Series, 178: 89–99, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/178/m178p089.pdf.

Andre:1991:ALI

[AR91] C. André and R. Rosenberg. Adult-larval interactions in the suspension-feeding bivalves Cerastoderma edule and Mya arenaria. Marine Ecology Progress Series, 71:227–234, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
71/m071p227.pdf.

Andersson:1993:PPB

[AR93] A. Andersson and Å. Rudehäll. Proportion of plankton biomass in particulate organic carbon in the northern Baltic Sea. Marine Ecology Progress Series, 95:133–139, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
95/m095p133.pdf.

Aluwihare:1999:CCC

[AR99] L. I. Aluwihare and D. J. Repeta. A comparison of the chemical characteristics of oceanic DOM and extracellular DOM produced by marine algae. Marine Ecology Progress Series, 186:105–117, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/186/m186p105.pdf.

Alcoverro:1997:STV

[ARDL97] T. Alcoverro, J. Romero, C. M. Duarte, and N. I. López. Spatial and temporal variations in nutrient limitation of seagrass Posidonia oceanica growth in the NW Mediterranean. Marine Ecology Progress Series, 146:155–161, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
146/m146p155.pdf.

Ainley:1992:DPP

[ARF92] D. G. Ainley, C. A. Ribic, and W. R. Fraser. Does prey preference affect habitat choice in Antarctic seabirds? Marine Ecology Progress Series, 90:207–221, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/90/m090p207.pdf.

Armonies:1992:MRD

[Arm92] W. Armonies. Migratory rhythms of drifting juvenile molluscs in tidal waters of the Wadden Sea. Marine Ecology Progress Series, 83:197–206, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/83/m083p197.pdf.

Aronson:1991:PPD

[Aro91] R. B. Aronson. Predation, physical disturbance, and sublethal arm damage in ophiuroids: a Jurassic-recent comparison. Marine Ecology Progress Series, 74:91–97, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
74/m074p091.pdf.

Aronson:1992:BSI

[Aro92] R. B. Aronson. Biology of a scale-independent predator–prey interaction. Marine Ecology Progress Series, 89:1–13, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
89/m089p001.pdf.

Aronson:1998:DSP

[Aro98] R. B. Aronson. Decadal-scale persistence of predation potential in coral reef communities. Marine Ecology Progress Series, 172:53–60, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/172/m172p053.pdf.

Arrontes:1993:NDB

[Arr93] J. Arrontes. Nature of the distributional boundary of Fucus serratus on the north shore of Spain. Marine Ecology Progress Series, 93:183–193, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/93/m093p183.pdf.

Arrigo:1994:IOD

[Arr94] K. R. Arrigo. Impact of ozone depletion on phytoplankton growth in the Southern Ocean: large-scale spatial and temporal variaoility. Marine Ecology Progress Series, 114: 1–12, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/114/m114p001.pdf.

Arrigo:1993:HRS

[ARS93] K. R. Arrigo, D. H. Robinson, and C. W. Sullivan. High resolution study of the platelet ice ecosystem in McMurdo Sound, Antarctica: photosynthetic and biooptical characteristics of a dense microalgal bloom. Marine Ecology Progress Series, 98:173–185, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/98/m098p173.pdf.

Andersen:1994:GEM

[AS94a] L. W. Andersen and H. R. Siegismund. Genetic evidence for migration of males between schools of the long-finned pilot whale Globicephala melas. Marine Ecology Progress Series, 105:1–7, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/105/m105p001.pdf.

Anthony:1994:EFH

[AS94b] K. R. N. Anthony and I. Svane. Effects of flow-habitat on body size and reproductive patterns in the sea anemone Metridium senile in the Gullmarsfjord, Sweden. Marine Ecology Progress Series, 113:257–269, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/113/m113p257.pdf.

Anthony:1995:ESI

[AS95] K. R. N. Anthony and I. Svane. Effects of substratum instability on locomotion and pedal laceration in Metridium senile (Anthozoa: Actiniaria). Marine Ecology Progress Series, 124:171–180, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/124/m124p171.pdf.

Atkinson:1997:KCI

[AS97] A. Atkinson and R. Snÿder. Krill-copepod interactions at South Georgia, Antarctica. I. Omnivory by Euphausia superba. Marine Ecology Progress Series, 160:63–76, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
160/m160p063.pdf.

Aarseth:1999:WSB

[AS99a] K. A. Aarseth and T. A. Schram. Wavelength-specific behaviour in Lepeophtheirus salmonis and Calanus finmarchicus to ultraviolet and visible light in laboratory experiments (Crustacea: Copepoda). Marine Ecology Progress Series, 186:211–217, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/186/m186p211.pdf.

Amend:1999:TLR

[AS99b] M. Amend and A. Shanks. Timing of larval release in the mole crab Emerita talpoida. Marine Ecology Progress Series, 183:295–300, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/183/m183p295.pdf.

Asakura:1991:PES

[Asa91] A. Asakura. Population ecology of the sand-dwelling hermit crab Diogenes nitidimanus. IV. Larval settlement. Marine Ecology Progress Series, 78:139–146, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
78/m078p139.pdf.

Ainley:1996:VDC

[ASA96] D. G. Ainley, L. B. Spear, and S. G. Allen. Variation in the diet of Cassin’s auklet reveals spatial, seasonal, and decadal occurrence patterns of euphausiids off California, USA. Marine Ecology Progress Series, 137:1–10, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
137/m137p001.pdf.

Abou-Seedo:1990:TTE

[ASCW90] F. Abou-Seedo, D. A. Clayton, and J. M. Wright. Tidal and turbidity effects on the shallow-water fish assemblage of Kuwait Bay. Marine Ecology Progress Series, 65:213– 223, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/65/m065p213.pdf.

Arndt:1998:MRH

[ASF98] C. Arndt, D. Schiedek, and H. Felbeck. Metabolic responses of the hydrothermal vent tube worm Riftia pachyptila to severe hypoxia. Marine Ecology Progress Series, 174:151–158, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
174/m174p151.pdf.

Andersson:1993:VTL

[ASH93] A. Andersson, E. Selstam, and Å. Hagström. Vertical transport of lipid in seawater. Marine Ecology Progress Series, 98: 149–155, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/98/m098p149.pdf.

Alberto:1999:APG

[ASL99] F. Alberto, R. Santos, and J. M. Leitão. Assessing patterns of geographic dispersal of Gelidium sesquipedale (Rhodophyta) through RAPD differentiation of populations. Marine Ecology Progress Series, 191:101–108, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/191/m191p101.pdf.

Abbiati:1993:GVW

[ASN93] M. Abbiati, G. Santangelo, and S. Novelli. Genetic variation within and between two Tyrrhenian populations of the Mediterranean alcyonarian Corallium rubrum. Marine Ecology Progress Series, 95:245–250, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/95/m095p245.pdf.

Ainley:1995:UTL

[ASN95] D. G. Ainley, W. J. Sydeman, and J. Norton. Upper trophic level predators indicate interannual negative and positive anomalies in the California Current food web. Marine Ecology Progress Series, 118:69–79, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/118/m118p069.pdf.

Andersen:1991:QEI

[ASP91] T. Andersen, A. K. L. Schartau, and E. Paasche. Quantifying external and internal nitrogen and phosphorus pools, as well as nitrogen and phosphorus supplied through remineralization, in coastal marine plankton by means of a dilution technique. Marine Ecology Progress Series, 69:67–80, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
69/m069p067.pdf.

Atkinson:1996:ZRP

[ASP^{+}96] A. Atkinson, R. S. Shreeve, E. A. Pakhomov, J. Priddle, S. P. Blight, and P. Ward. Zooplankton response to a phytoplankton bloom near South Georgia, Antarctica. Marine Ecology Progress Series, 144:195–210, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/144/m144p195.pdf.

Adcock:1999:MAG

[ASRC99] G. J. Adcock, P. W. Shaw, P. G. Rodhouse, and G. R. Carvalho. Microsatellite analysis of genetic diversity in the squid Illex argentinus during a period of intensive fishing. Marine Ecology Progress Series, 187:171–178, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
187/m187p171.pdf.

Alvarez-Salgado:1996:NCEb

[ASRP^{+}96a] A. Alvarez-Salgado, G. Rosón, F. F. Pérez, F. G. Figueiras, and A. F. Ríos. Nitrogen cycling in an estuarine upwelling system, the Ría de Arousa (NW Spain). II. Spatial differences in the short-time-scale evolution of fluxes and net budgets. Marine Ecology Progress Series, 135:275–288, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
135/m135p275.pdf.

Alvarez-Salgado:1996:NCEa

[ASRP^{+}96b] X. A. Alvarez-Salgado, G. Rosón, F. F. Pérez, F. G. Figueiras, and Y. Pazos. Nitrogen cycling in an estuarine upwelling system, the Ría de Arousa (NW Spain). I. Short-time-scale patterns of hydrodynamic and biogeochemical circulation. Marine Ecology Progress Series, 135:259–273, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
135/m135p259.pdf.

Atkinson:1997:RDL

[ASSWM97] A. Atkinson, S. B. Schnack-Schiel, P. Ward, and V. Marin. Regional differences in the life cycle of Calanoides acutus (Copepoda: Calanoida) within the Atlantic sector of the Southern Ocean. Marine Ecology Progress Series, 150:99– 111, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/150/m150p099.pdf.

Attrill:1996:LTD

[AT96] M. J. Attrill and R. M. Thomas. Long-term distribution patterns of mobile estuarine invertebrates (Ctenophora, Cnidaria, Crustacea: Decapoda) in relation to hydrological parameters. Marine Ecology Progress Series, 143:25–36, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
143/m143p025.pdf.

Arnold:1995:PVP

[ATH95] T. M. Arnold, C. E. Tanner, and W. I. Hatch. Phenotypic variation in polyphenolic content of the tropical brown alga Lobophora variegata as a function of nitrogen availability. Marine Ecology Progress Series, 123:177–183, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
123/m123p177.pdf.

Atkinson:1996:SCO

[Atk96] A. Atkinson. Subantarctic copepods in an oceanic, low chlorophyll environment: ciliate predation, food selectivity and impact on prey populations. Marine Ecology Progress Series, 130:85–96, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/130/m130p085.pdf.

Anderson:1997:DPB

[ATNW97] I. C. Anderson, C. R. Tobias, B. B. Neikirk, and R. L. Wetzel. Development of a process-based nitrogen mass balance model for a Virginia (USA) Spartina alterniflora salt marsh: implications for net DIN flux. Marine Ecology Progress Series, 159:13–27, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/159/m159p013.pdf.

Andrew:1993:DDF

[AU93] N. L. Andrew and A. J. Underwood. Density-dependent foraging in the sea urchin Centrostephanus rodgersii on shallow subtidal reefs in New South Wales, Australia. Marine Ecology Progress Series, 99:89–98, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/99/m099p089.pdf.

Alber:1995:OAD

[AV95] M. Alber and I. Valiela. Organic aggregates in detrital food webs: incorporation by bay scallops Argopecten irradians.

Marine Ecology Progress Series, 121:117–124, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
121/m121p117.pdf.

Airoldi:1998:RTF

[AV98] L. Airoldi and M. Virgilio. Responses of turf-forming algae to spatial variations in the deposition of sediments. Marine Ecology Progress Series, 165:271–282, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/165/m165p271.pdf.

Aerts:1997:QSC

[AvS97] L. A. M. Aerts and R. W. M. van Soest. Quantification of sponge/coral interactions in a physically stressed reef community, NE Colombia. Marine Ecology Progress Series, 148: 125–134, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/148/m148p125.pdf.

Andres:1995:TCS

[AW95] N. G. Andres and J. D. Witman. Trends in community structure on a Jamaican reef. Marine Ecology Progress Series, 118:305–310, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/118/m118p305.pdf.

Andrew:1997:AVM

[AW97] J. Andrew and R. D. Ward. Allozyme variation in the marine fanworm Sabella spallanzanii: comparison of native European and introduced Australian populations. Marine Ecology Progress Series, 152:131–143, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/152/m152p131.pdf.

Andreassen:1998:VFP

[AW98] I. J. Andreassen and P. Wassmann. Vertical flux of phytoplankton and particulate biogenic matter in the marginal ice zone of the Barents Sea in May 1993. Marine Ecology Progress Series, 170:1–14, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/170/m170p001.pdf.

Aagaard:1995:TSC

[AWD95] A. Aagaard, C. G. Warman, and M. H. Depledge. Tidal and seasonal changes in the temporal and spatial distribution of foraging Carcinus maenas in the weakly tidal littoral zone of Kerteminde Fjord, Denmark. Marine Ecology Progress Series, 122:165–172, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/122/m122p165.pdf.

Alkemade:1992:EER

[AWdS92] R. Alkemade, A. Wielemaker, S. A. de Jong, and A. J. J. Sandee. Experimental evidence for the role of bioturbation by the marine nematode Diplolaimella dievengatensis in stimulating the mineralization of Spartina anglica detritus. Marine Ecology Progress Series, 90:149–155, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
90/m090p149.pdf.

Alkemade:1993:CBN

[AWH93] R. Alkemade, A. Wielemaker, and M. A. Hemminga. Correlation between nematode abundance and decomposition rate of Spartina anglica leaves. Marine Ecology Progress Series, 99: 293–300, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/99/m099p293.pdf.

Atkinson:1999:KCI

[AWH^{+}99] A. Atkinson, P. Ward, A. Hill, A. S. Brierley, and G. C. Cripps. Krill-copepod interactions at South Georgia, Antarctica. II. Euphausia superba as a major control on copepod abundance. Marine Ecology Progress Series, 176:63–79, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
176/m176p063.pdf.

Alkemade:1994:PDD

[AWHH94] R. Alkemade, A. Wielemaker, P. M. J. Herman, and M. A. Hemminga. Population dynamics of Diplolaimelloides bruciei, a nematode associated with the salt marsh plant Spartina anglica. Marine Ecology Progress Series, 105:277– 284, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/105/m105p277.pdf.

Austen:1998:EBD

[AWVP98] M. C. Austen, S. Widdicombe, and N. Villano-Pitacco. Effects of biological disturbance on diversity and structure of meiobenthic nematode communities. Marine Ecology Progress Series, 174:233–246, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/174/m174p233.pdf.

Alcoverro:1999:RAS

[AZKA99] T. Alcoverro, R. C. Zimmerman, D. G. Kohrs, and R. S. Alberte. Resource allocation and sucrose mobilization in lightlimited eelgrass Zostera marina. Marine Ecology Progress Series, 187:121–131, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/187/m187p121.pdf.

Barshaw:1990:DBR

[BA90] D. E. Barshaw and K. W. Able. Deep burial as a refuge for lady crabs Ovalipes ocellatus: comparisons with blue crabs Callinectes sapidus. Marine Ecology Progress Series, 66: 75–79, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/66/m066p075.pdf.

Bennett:1991:ERS

[BA91] B. A. Bennett and C. G. Attwood. Evidence for recovery of a surf-zone fish assemblage following the establishment of a marine reserve on the southern coast of South Africa. Marine Ecology Progress Series, 75:173–181, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
75/m075p173.pdf.

Balino:1993:WDM

[BA93] B. M. Balino and D. L. Aksnes. Winter distribution and migration of the sound scattering layers, zooplankton and micronekton in masfjorden, western Norway. Marine Ecology Progress Series, 102:35–50, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/102/m102p035.pdf.

Brown:1994:GFM

[BA94] K. M. Brown and J. E. Alexander, Jr. Group foraging in a marine gastropod predator: benefits and costs to individuals. Marine Ecology Progress Series, 112:97–105, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
112/m112p097.pdf.

Bushek:1996:HPI

[BA96] D. Bushek and S. K. Allen, Jr. Host-parasite interactions among broadly distributed populations of the eastern oyster Crassostrea virginica and the protozoan Perkinsus marinus. Marine Ecology Progress Series, 139:127–141, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
139/m139p127.pdf.

Billingham:1997:ARG

[BA97] M. R. Billingham and D. J. Ayre. Asexual reproduction and genetic determination of colour patterns within populations of the subtidal sea anemone Anthothoe albocincta. Marine Ecology Progress Series, 156:121–130, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/156/m156p121.pdf.

Barshaw:1994:SMP

[BAH94] D. E. Barshaw, K. W. Able, and K. L. Heck, Jr. Salt marsh peat reefs as protection for postlarval lobsters Homarus americanus from fish and crab predators: comparisons with other substrates. Marine Ecology Progress Series, 106:203– 206, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/106/m106p203.pdf.

Bak:1990:PEB

[Bak90] R. P. M. Bak. Patterns of echinoid bioerosion in two Pacific coral reef lagoons. Marine Ecology Progress Series, 66: 267–272, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/66/m066p267.pdf.

Boyd:1997:CRF

[BAK97] P. W. Boyd, J. Aiken, and Z. Kolber. Comparison of radiocarbon and fluorescence based (pump and probe) measurements of phytoplankton photosynthetic characteristics in the Northeast Atlantic Ocean. Marine Ecology Progress Series, 149:215–226, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/149/m149p215.pdf.

Bulthuis:1992:GSH

[BAM92] D. A. Bulthuis, D. M. Axelrad, and M. J. Mickelson. Growth of the seagrass Heterozoslera tasmanica limited by nitrogen in Port Phillip Bay, Australia. Marine Ecology Progress Series, 89:269–275, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/89/m089p269.pdf.

Bricaud:1999:DCS

[BAM^{+}99] A. Bricaud, K. Allali, A. Morel, D. Marie, M. J. W. Veldhuis, F. Partensky, and D. Vaulot. Divinyl chlorophyll aspecific absorption coefficients and absorption efficiency factors for Prochlorococcus marinus: kinetics of photoacclimation. Marine Ecology Progress Series, 188:21–32, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
188/m188p021.pdf.

Buchner:1996:AAS

[BAOT96] T. Buchner, D. Abele-Oeschger, and H. Theede. Aspects of antioxidant status in the polychaete Arenicola marina: tissue and subcellular distribution, and reaction to environmental hydrogen peroxide and elevated temperatures. Marine Ecology Progress Series, 143:141–150, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/143/m143p141.pdf.

Barthel:1995:TCS

[Bar95] D. Barthel. Tissue composition of sponges from the Weddell Sea, Antarctica: not much meat on the bones. Marine Ecology Progress Series, 123:149–153, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/123/m123p149.pdf.

Bernhard:1996:RCM

[BAR96] J. M. Bernhard, E. Alve, and A. A. Reilly. Response to the comment by Moodley et al. Marine Ecology Progress Series, 145:305–307, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/145/m145p305.pdf. See [MHMH96].

Barnes:1997:ETHc

[Bar97a] D. K. A. Barnes. Ecology of tropical hermit crabs at Quirimba Island, Mozambique: a novel and locally important food source. Marine Ecology Progress Series, 161:299–302, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
161/m161p299.pdf.

Barnes:1997:ETHa

[Bar97b] D. K. A. Barnes. Ecology of tropical hermit crabs at Quirimba Island, Mozambique: distribution, abundance and activity. Marine Ecology Progress Series, 154:133–142, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
154/m154p133.pdf.

Barnes:1997:ETHb

[Bar97c] D. K. A. Barnes. Ecology of tropical hermit crabs at Quirimba Island, Mozambique: vertical migration (tree climbing). Marine Ecology Progress Series, 158:233–240, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
158/m158p233.pdf.

Barnes:1999:ETH

[Bar99] D. K. A. Barnes. Ecology of tropical hermit crabs at Quirimba Island, Mozambique: shell characteristics and utilisation. Marine Ecology Progress Series, 183:241–251, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
183/m183p241.pdf.

Bongers:1991:IDI

[BAY91] T. Bongers, R. Alkemade, and G. W. Yeates. Interpretation of disturbance-induced maturity decrease in marine nematode assemblages by means of the maturity index. Marine Ecology Progress Series, 76:135–142, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/76/m076p135.pdf.

Bernhard:1992:BBT

[BB92] J. M. Bernhard and S. S. Bowser. Bacterial biofilms as a trophic resource for certain benthic Foraminifera. Marine Ecology Progress Series, 83:263–272, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
83/m083p263.pdf.

Brown:1993:ETS

[BB93] S. D. Brown and T. M. Bert. The effects of temperature and salinity on molting and survival of Menippe adina and M. mercenaria (Crustacea, Decapoda) postsettlement juveniles. Marine Ecology Progress Series, 99:41–49, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
99/m099p041.pdf.

Bisbal:1995:EDF

[BB95] G. A. Bisbal and D. A. Bengtson. Effects of delayed feeding on survival and growth of summer flounder Paralichthys dentatus larvae. Marine Ecology Progress Series, 121:301–306, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
121/m121p301.pdf.

Bartoli:1997:TFP

[BB97a] P. Bartoli and C. F. Boudouresque. Transmission failure of parasites (Digenea) in sites colonized by the recently introduced invasive alga Caulerpa taxifolia. Marine Ecology Progress Series, 154:253–260, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/154/m154p253.pdf.

Beliaeff:1997:SDO

[BB97b] B. Beliaeff and T. Burgeot. Sampling design optimization for EROD measurements in fish. Marine Ecology Progress Series, 153:239–246, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/153/m153p239.pdf.

Blanchard:1999:SCH

[BB99] D. Blanchard and E. Bourget. Scales of coastal heterogeneity: influence on intertidal community structure. Marine Ecology Progress Series, 179:163–173, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/179/m179p163.pdf.

Bruggemann:1994:FSPb

[BBB^{+}94] J. H. Bruggemann, J. Begeman, E. M. Bosma, P. Verburg, and A. M. Breeman. Foraging by the stoplight parrotfish Sparisoma viride II. Intake and assimilation of food, protein and energy. Marine Ecology Progress Series, 106:57–71, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
106/m106p057.pdf.

Ban:1997:PDC

[BBC^{+}97] S. Ban, C. Burns, J. Castel, Y. Chaudron, E. Christou, R. Escribano, S. F. Umani, S. Gasparini, F. G. Ruiz, M. Hoffmeyer, A. Ianora, H. K. Kang, M. Laabir, A. Lacoste, A. Miralto, X. Ning, S. Poulet, V. Rodriguez, J. Runge, J. Shi, M. Starr, S. Uye, and Y. Wang. The paradox of diatom-copepod interactions. Marine Ecology Progress Series, 157:287–293, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
157/m157p287.pdf.

Brown:1999:FAS

[BBCB99] D. J. Brown, I. L. Boyd, G. C. Cripps, and P. J. Butler. Fatty acid signature analysis from the milk of Antarctic fur seals and Southern elephant seals from South Georgia: implications for diet determination. Marine Ecology Progress Series, 187:251–263, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/187/m187p251.pdf.

Bianchi:1994:BND

[BBF94] M. Bianchi, P. Bonin, and Feliatra. Bacterial nitrification and denitrification rates in the Rhone River plume (northwestern Mediterranean Sea). Marine Ecology Progress Series, 103:197–202, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/103/m103p197.pdf.

Bach:1998:SCP

[BBFD98] S. S. Bach, J. Borum, M. D. Fortes, and C. M. Duarte. Species composition and plant performance of mixed seagrass beds along a siltation gradient at Cape Bolinao, The Philippines. Marine Ecology Progress Series, 174:247–256, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
174/m174p247.pdf.

Buck:1990:PFP

[BBG90] K. P. Buck, P. A. Bolt, and D. L. Garrison. Phagotrophy and fecal pellet production by an athecate dinoflagellate in Antarctic sea ice. Marine Ecology Progress Series, 60: 75–84, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/60/m060p075.pdf.

Brodeur:1991:CEW

[BBK91] R. D. Brodeur, K. M. Bailey, and S. Kim. Cannibalism on eggs by walleye pollock Theragra chalcogramma in Shelikof Strait, Gulf of Alaska. Marine Ecology Progress Series, 71:207–218, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/71/m071p207.pdf.

Brunet:1993:LSV

[BBL93] C. Brunet, J. M. Brylinski, and Y. Lemoine. In situ variations of the xanthophylls diatoxanthin and diadinoxanthin: photoadaptation and relationships with a hydrodynamical system in the eastern English Channel. Marine Ecology Progress Series, 102:69–77, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/102/m102p069.pdf.

Blight:1995:PAH

[BBL^{+}95] S. P. Blight, T. L. Bentley, D. Lefevre, C. Robinson, R. Rodrigues, J. Rowlands, and P. J. le B. Williams. Phasing of autotrophic and heterotrophic plankton metabolism in a temperate coastal ecosystem. Marine Ecology Progress Series, 128:61–75, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/128/m128p061.pdf.

Belda-Baillie:1998:GMR

[BBLY98] C. A. Belda-Baillie, W. Leggat, and D. Yellowlees. Growth and metabolic responses of the giant clam-zooxanthellae symbiosis in a reef-fertilisation experiment. Marine Ecology Progress Series, 170:131–141, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/170/m170p131.pdf.

Boyd:1995:CIC

[BBO^{+}95] P. W. Boyd, S. J. Bury, N. J. P. Owens, G. J. Savidge, and T. Preston. A comparison of isotopic and chemiluminescent methods of estimating microalgal nitrate uptake in the NE Atlantic. Marine Ecology Progress Series, 116:199–205, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
116/m116p199.pdf.

Burgeot:1996:BPE

[BBP^{+}96] T. Burgeot, G. Bocquéné, C. Porte, J. Dimeet, R. M. Santella, L. M. Garcia de la Parra, A. Píhol-Leszkowicz, C. Raoux, and F. Galgani. Bioindicators of pollutant exposure in the northwestern Mediterranean Sea. Marine Ecology Progress Series, 131:125–141, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/131/m131p125.pdf.

Batel:1993:MOI

[BBR^{+}93] Renato Batel, Nevenka Bihari, Baruch Rinkevich, Jutta Dapper, Heike Schäcke, Heinz C. Schröder, and Werner E. G. Müller. Modulation of organotin-induced apoptosis by the water pollutant methyl mercury in a human lymphoblastoid tumor cell line and a marine sponge. Marine Ecology Progress Series, 93:245–251, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/93/m093p245.pdf.

Baretta-Bekker:1994:TML

[BBRBK94] J. G. Baretta-Bekker, B. Riemann, J. W. Baretta, and E. Koch Rasmussen. Testing the microbial loop concept by comparing mesocosm data with results from a dynamical simulation model. Marine Ecology Progress Series, 106: 187–198, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/106/m106p187.pdf.

Boudreau:1993:EAI

[BBS93] B. Boudreau, E. Bourget, and Y. Simard. Effect of age, injury, and predator odors or settlement and shelter selection by lobster Homarus amencanus postlarvae. Marine Ecology Progress Series, 93:119–129, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/93/m093p119.pdf.

Burgeot:1993:DCL

[BBT^{+}93] T. Burgeot, G. Bocquené, P. Truquet, L. Le Dean, J. C. Poulard, D. Dorel, A. Souplet, and F. Galgani. The dragonet Callionymus lyra, a target species used for evaluation of biological effects of chemical contaminants on French coasts. Marine Ecology Progress Series, 97:309–316, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
97/m097p309.pdf.

Beukema:1990:GRB

[BC90] J. J. Beukema and G. C. Cadee. Growth rates of the bivalve Macoma balthica in the Wadden Sea during a period of eutrophication: relationships with concentrations of pelagic diatoms and flagellates. Marine Ecology Progress Series, 68: 249–256, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/68/m068p249.pdf.

Barange:1991:MSA

[BC91] M. Barangé and B. Campos. Models of species abundance: a critique of and an alternative to the dynamics model. Marine Ecology Progress Series, 69:293–298, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
69/m069p293.pdf.

Bjerregaard:1993:AOI

[BC93] P. Bjerregaard and L. Christensen. Accumulation of organic and inorganic mercury from food in the tissues of Carcinus maenas: effect of waterborne selenium. Marine Ecology Progress Series, 99:271–281, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/99/m099p271.pdf.

Butler:1999:ASM

[BC99] A. J. Butler and R. M. Connolly. Assemblages of sessile marine invertebrates: still changing after all these years? Marine Ecology Progress Series, 182:109–118, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
182/m182p109.pdf.

Borrell:1999:PVR

[BCA^{+}99] A. Borrell, G. Cantos, A. Aguilar, C. Lockyer, A. Brouwer, M. P. Heide-Jørgensen, J. Jensen, and B. Spenkelink. Patterns of variability of retinol levels in a harbour porpoise population from an unpolluted environment. Marine Ecology Progress Series, 185:85–92, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/185/m185p085.pdf.

Benedetti-Cecchi:1996:EAB

[BCAAC96] L. Benedetti-Cecchi, L. Airoldi, M. Abbiati, and F. Cinelli. Estimating the abundance of benthic invertebrates: a comparison of procedures and variability between observers. Marine Ecology Progress Series, 138:93–101, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
138/m138p093.pdf.

Benedetti-Cecchi:1998:DDF

[BCBC98] L. Benedetti-Cecchi, F. Bulleri, and F. Cinelli. Density dependent foraging of sea urchins in shallow subtidal reefs on the west coast of Italy (western Mediterranean). Marine Ecology Progress Series, 163:203–211, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/163/m163p203.pdf.

Buschmann:1997:DDE

[BCBR97] A. H. Buschmann, J. A. Correa, J. Beltrán, and C. A. Retamales. Determinants of disease expression and survival of infected individual fronds in wild populations of Mazzaella laminarioides (Rhodophyta) in central and southern Chile. Marine Ecology Progress Series, 154:269–280, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
154/m154p269.pdf.

Benedetti-Cecchi:1992:ECC

[BCC92] L. Benedetti-Cecchi and F. Cinelli. Effects of canopy cover, herbivores and substratum type on patterns of Cystoseira spp. settlement and recruitment in littoral rockpools. Marine Ecology Progress Series, 90:183–191, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/90/m090p183.pdf.

Benedetti-Cecchi:1994:RPA

[BCC94] L. Benedetti-Cecchi and F. Cinelli. Recovery of patches in an assemblage of geniculate coralline algae: variability at different successional stages. Marine Ecology Progress Series, 110:9–18, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/110/m110p009.pdf.

Benedetti-Cecchi:1995:HHS

[BCC95] L. Benedetti-Cecchi and F. Cinelli. Habitat heterogeneity, sea urchin grazing and the distribution of algae in littoral rock pools on the west coast of Italy (western Mediterranean). Marine Ecology Progress Series, 126:203–212, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
126/m126p203.pdf.

Benedetti-Cecchi:1996:PDR

[BCC96] L. Benedetti-Cecchi and F. Cinelli. Patterns of disturbance and recovery in littoral rock pools: nonhierarchical competition and spatial variability in secondary succession. Marine Ecology Progress Series, 135:145–161, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
135/m135p145.pdf.

Boucher:1994:OCD

[BCG94] G. Boucher, J. Clavier, and C. Garrigue. Oxygen and carbon dioxide fluxes at the water–sediment interface of a tropical lagoon. Marine Ecology Progress Series, 107:185–193, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
107/m107p185.pdf.

Barlow:1997:IRM

[BCG97] R. G. Barlow, D. G. Cummings, and S. W. Gibb. Improved resolution of mono- and divinyl chlorophylls a and b and zeaxanthin and lutein in phytoplankton extracts using reverse phase C-8 HPLC. Marine Ecology Progress Series, 161:303–307, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/161/m161p303.pdf.

Berges:1995:LFR

[BCH95] J. A. Berges, W. P. Cochlan, and P. J. Harrison. Laboratory and field responses of algal nitrate reductase to diel periodicity in irradiance, nitrate exhaustion, and the presence of ammonium. Marine Ecology Progress Series, 124:259–269, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
124/m124p259.pdf.

Behrenfeld:1993:TCR

[BCHL93] M. J. Behrenfeld, J. W. Chapman, J. T. Hardy, and H. Lee II. Is there a common response to ultraviolet-B radiation by marine phytoplankton? Marine Ecology Progress Series, 102:59–68, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/102/m102p059.pdf.

Bertram:1993:NCB

[BCL93] D. R. Bertram, R. O. Chambers, and W. C. Leggett. Negative correleations between larval and juvenile growth rates in winter flounder: implications of compensatory growth for variation in size-at-age. Marine Ecology Progress Series, 96: 209–215, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/96/m096p209.pdf.

Benedetti-Cecchi:1999:PES

[BCMC99] L. Benedetti-Cecchi, M. Menconi, and F. Cinelli. Pre-emption of the substratum and the maintenance of spatial pattern on a rocky shore in the northwest Mediterranean. Marine Ecology Progress Series, 181:13–23, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/181/m181p013.pdf.

Bailey:1995:CYP

[BCN^{+}95] K. M. Bailey, M. F. Canino, J. M. Napp, S. M. Spring, and A. L. Brown. Contrasting years of prey levels, feeding conditions and mortality of larval walleye pollock Theragra chalcogramma in the western Gulf of Alaska. Marine Ecology Progress Series, 119:11–23, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/119/m119p011.pdf.

Benedetti-Cecchi:1996:AST

[BCNC96] L. Benedetti-Cecchi, S. Nuti, and F. Cinelli. Analysis of spatial and temporal variability in interactions among algae, limpets and mussels in low-shore habitats on the west coast of Italy. Marine Ecology Progress Series, 144:87–96, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
144/m144p087.pdf.

Back:1992:CEB

[BCR92] S. Bäck, J. C. Collins, and G. Russel. Comparative ecophysiology of Baltic and Atlantic Fucus vesiculosus. Marine Ecology Progress Series, 84:71–82, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/84/m084p071.pdf.

Boyer:1993:PPP

[BCS93] J. N. Boyer, R. R. Christian, and D. W. Stanley. Patterns of phytoplankton primary productivity in the Neuse River estuary, North Carolina, USA. Marine Ecology Progress Series, 97:287–297, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/97/m097p287.pdf.

Butler:1992:CNA

[BCv^{+}92] R. Butler, P. M. Chapman, P. van den Hurk, B. Roddie, and J. E. Thain. A comparison of North American and West European oyster embryo-larval toxicity tests on North Sea sediments. Marine Ecology Progress Series, 91:245–251, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
91/m091p245.pdf.

Berdalet:1991:NDS

[BD91] E. Berdalet and Q. Dortch. New double-staining technique for RNA and DNA measurement in marine phytoplankton. Marine Ecology Progress Series, 73:295–305, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
73/m073p295.pdf.

Butler:1994:PRC

[BD94] M. Butler and H. G. Dam. Production rates and characteristics of fecal pellets of the copepod Acartia tonsa under simulated phytoplankton bloom conditions: implications for vertical fluxes. Marine Ecology Progress Series, 114:81–91, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
114/m114p081.pdf.

Beninger:1996:MDR

[BD96] P. G. Beninger and S. C. Dufour. Mucocyte distribution and relationship to particle transport on the pseudolamellibranch gill of Crassostrea virginica (Bivalvia: Ostreidae). Marine Ecology Progress Series, 137:133–138, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
137/m137p133.pdf.

Boon:1998:CMB

[BD98] A. R. Boon and G. C. A. Duineveld. chlorophyll-a as a marker for bioturbation and carbon flux in southern and central North Sea sediments. Marine Ecology Progress Series, 162:33–43, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/162/m162p033.pdf.

Brazeiro:1999:EHD

[BD99] A. Brazeiro and O. Defeo. Effects of harvesting and density dependence on the demography of sandy beach populations: the yellow clam Mesodesma mactroides of Uruguay. Marine Ecology Progress Series, 182:127–135, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
182/m182p127.pdf.

Brown:1999:SFE

[BDA^{+}99] B. E. Brown, R. P. Dunne, I. Ambarsari, M. D. A. Le Tissier, and U. Satapoomin. Seasonal fluctuations in environmental factors and variations in symbiotic algae and chlorophyll pigments in four Indo–Pacific coral species. Marine Ecology Progress Series, 191:53–69, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/191/m191p053.pdf.

Beninger:1997:PPM

[BDB97] P. G. Beninger, S. C. Dufour, and J. Bourque. Particle processing mechanisms of the eulamellibranch bivalves Spisula solidissima and Mya arenaria. Marine Ecology Progress Series, 150:157–169, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/150/m150p157.pdf.

Becker:1992:HGB

[BDD92] G. A. Becker, S. Dick, and J. W. Dippner. Hydrography of the German Bight. Marine Ecology Progress Series, 91: 9–18, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/91/m091p009.pdf.

Buffan-Dubau:1996:FSH

[BDdWC96] E. Buffan-Dubau, R. de Wit, and J. Castel. Feeding selectivity of the harpacticoid copepod Canuella perplexa in benthic muddy environments demonstrated by HPLC analyses of chlorin and carotenoid pigments. Marine Ecology Progress Series, 137:71–82, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/137/m137p071.pdf.

Baden:1994:GDA

[BDH94] S. P. Baden, M. H. Depledge, and L. Hagerman. Glycogen depletion and altered copper and manganese handling in Nephrops norvegicus following starvation and exposure to hypoxia. Marine Ecology Progress Series, 103:65–72, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
103/m103p065.pdf.

Burke:1996:NSC

[BDM96] M. K. Burke, W. C. Dennison, and K. A. Moore. Nonstructural carbohydrate reserves of eelgrass Zostera marina. Marine Ecology Progress Series, 137:195–201, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
137/m137p195.pdf.

Bussmann:1999:GSM

[BDNS99] I. Bussmann, P. R. Dando, S. J. Niven, and E. Suess. Groundwater seepage in the marine environment: role for mass flux and bacterial activity. Marine Ecology Progress Series, 178: 169–177, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/178/m178p169.pdf.

Bonsdorff:1996:CSB

[BDR^{+}96] E. Bonsdorff, R. J. Diaz, R. Rosenberg, A. Norkko, and G. R. Cutter, Jr. Characterization of soft-bottom benthic habitats of the Åland Islands, northern Baltic Sea. Marine Ecology Progress Series, 142:235–245, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/142/m142p235.pdf.

Brown:1994:SDI

[BDSL94] B. E. Brown, R. P. Dunne, T. P. Scoffin, and M. D. A. Le Tissier. Solar damage in intertidal corals. Marine Ecology Progress Series, 105:219–230, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/105/m105p219.pdf.

Ben-David-Zaslow:1998:CLP

[BDZB98] R. Ben-David-Zaslow and Y. Benayahu. Competence and longevity in planulae of several species of soft corals. Marine Ecology Progress Series, 163:235–243, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
163/m163p235.pdf.

Banse:1997:NSP

[BE97] K. Banse and D. C. English. Near-surface phytoplankton pigment from the coastal zone color scanner in the subantarctic region southeast of New Zealand. Marine Ecology Progress Series, 156:51–66, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/156/m156p051.pdf.

Beck:1998:CME

[Bec98] M. W. Beck. Comparison of the measurement and effects of habitat structure on gastropods in rocky intertidal and mangrove habitats. Marine Ecology Progress Series, 169: 165–178, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/169/m169p165.pdf.

Beets:1997:EPF

[Bee97] J. Beets. Effects of a predatory fish on the recruitment and abundance of Caribbean coral reef fishes. Marine Ecology Progress Series, 148:11–21, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/148/m148p011.pdf.

Buma:1997:WDI

[BEG97] A. G. J. Buma, A. H. Engelen, and W. W. C. Gieskes. Wavelength-dependent induction of thymine dimers and growth rate reduction in the marine diatom Cyclotella sp. exposed to ultraviolet radiation. Marine Ecology Progress Series, 153:91–97, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/153/m153p091.pdf.

Bratbak:1993:VMM

[BEH93] G. Bratbak, J. K. Egge, and M. Heldal. Viral mortality of the marine alga Emiliania huxleyi (Haptophyceae) and termination of algal blooms. Marine Ecology Progress Series, 93:39–48, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/93/m093p039.pdf.

Bellwood:1995:CTW

[Bel95] D. R. Bellwood. Carbonate transport and within-reef patterns of bioerosion and sediment release by parrotfishes (family Scaridae) on the Great Barrier Reef. Marine Ecology Progress Series, 117:127–136, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/117/m117p127.pdf.

Berge:1990:MRS

[Ber90] J. A. Berge. Macrofauna recolonization of subtidal sediments. Experimental studies on defaunated sediment contaminated with cruce oil in two Norwegian fjords with unequal eutrophication status. I. Community responses. Marine Ecology Progress Series, 66:103–115, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/66/m066p103.pdf.

Bergstrom:1992:GGM

[Ber92] B. Bergstrom. Growth, growth modelling and age determination of Pandalus borealis. Marine Ecology Progress Series, 83:167–183, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/83/m083p167.pdf.

Beukema:1992:DJS

[Beu92] J. J. Beukema. Dynamics of juvenile shrimp Crangon crangon in a tidal-flat nursery of the Wadden Sea after mild and cold winters. Marine Ecology Progress Series, 83:157–165, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
83/m083p157.pdf.

Bell:1999:NFT

[BEU99] P. R. F. Bell, I. Elmetri, and P. Uwins. Nitrogen fixation by Trichodesmium spp. in the central and Northern Great Barrier Reef lagoon: relative importance of the fixed-nitrogen load. Marine Ecology Progress Series, 186:119–126, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
186/m186p119.pdf.

Bernard:1995:MCS

[BF95a] C. Bernard and T. Fenchel. Mats of colourless sulphur bacteria. II. Structure, composition of biota and successional patterns. Marine Ecology Progress Series, 128:171–179, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
128/m128p171.pdf.

Beukema:1995:FCU

[BF95b] J. J. Beukema and E. C. Flach. Factors controlling the upper and lower limits of the intertidal distribution of two Corophium species in the Wadden Sea. Marine Ecology Progress Series, 125:117–126, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/125/m125p117.pdf.

Blackburn:1999:IBD

[BF99] N. Blackburn and T. Fenchel. Influence of bacteria, diffusion and shear on micro-scale nutrient patches, and implications for bacterial chemotaxis. Marine Ecology Progress Series, 189:1–7, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/189/m189p001.pdf.

Blust:1991:EHI

[BFD91] R. Blust, A. Fontaine, and W. Decleir. Effect of hydrogen ions and inorganic complexing on the uptake of copper by the brine shrimp Artemia franciscana. Marine Ecology Progress Series, 76:273–282, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/76/m076p273.pdf.

Balch:1996:DCP

[BFF96] W. M. Balch, J. Fritz, and E. Fernandez. Decoupling of calcification and photosynthesis in the coccolithophore Emiliania huxleyi under steady-state light-limited growth. Marine Ecology Progress Series, 142:87–97, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/142/m142p087.pdf.

Boehme:1993:VBP

[BFJ^{+}93] J. Boehme, M. E. Frischer, S. C. Jiang, C. A. Kellogg, S. Pichard, J. B. Rose, C. Steinway, and J. H. Paul. Viruses, bacterioplankton, and phyloplankton in the southeastern Gulf of Mexico: distribution and contribution to oceanic DNA pools. Marine Ecology Progress Series, 97:1–10, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
97/m097p001.pdf.

Blanco:1999:OAD

[BFMM99] J. Blanco, M. L. Fernández, A. Míguez, and A. Moroño. Okadaic acid depuration in the mussel Mytilus galloprovincialis: one- and two-compartment models and the effect of environmental conditions. Marine Ecology Progress Series, 176:153–163, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/176/m176p153.pdf.

Best:1998:WDP

[BFS^{+}98] P. B. Best, K. P. Findlay, K. Sekiguchi, V. M. Peddemors, B. Rakotonirina, A. Rossouw, and D. Gove. Winter distribution and possible migration routes of humpback whales Megaptera novaeangliae in the southwest Indian Ocean. Marine Ecology Progress Series, 162:287–299, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
162/m162p287.pdf.

Brunetti:1992:GEA

[BFVG92] R. Brunetti, O. Fumagalli, P. Valerio, and M. Gabriele. Genotoxic effects of anoxia on Mytilus galloprovincialis. Marine Ecology Progress Series, 83:71–74, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/83/m083p071.pdf.

Bronk:1991:TMM

[BG91] D. A. Bronk and P. M. Glibert. A ^{15}N tracer method for the measurement of dissolved organic nitrogen release by phytoplankton. Marine Ecology Progress Series, 77:171–182, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
77/m077p171.pdf.

Bronk:1993:CPD

[BG93] D. A. Bronk and P. M. Glibert. Contrasting patterns of dissolved organic nitrogen release by two size fractions of estuarine plankton during a period of rapid NH consumption and N0_{2}^{−} production. Marine Ecology Progress Series, 96: 291–299, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/96/m096p291.pdf.

Bianchi:1994:BRH

[BG94] A. Bianchi and J. Garcin. Bacterial response to hydrostatic pressure in seawater samples collected in mixed-water and stratified-water conditions. Marine Ecology Progress Series, 111:137–141, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/111/m111p137.pdf.

Bell:1997:SLF

[BG97] E. C. Bell and J. M. Gosline. Strategies for life in flow: tenacity, morphometry, and probability of dislodgment of two Mytilus species. Marine Ecology Progress Series, 159: 197–208, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/159/m159p197.pdf.

Buecher:1999:TPV

[BG99] E. Buecher and M. J. Gibbons. Temporal persistence in the vertical structure of the assemblage of planktonic medusae in the NW Mediterranean Sea. Marine Ecology Progress Series, 189:105–115, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/189/m189p105.pdf.

Bertness:1991:ESV

[BGBS91] M. D. Bertness, S. D. Gaines, D. Bermudez, and E. Sanford. Extreme spatial variation in the growth and reproductive output of the acorn barnacle Semibalanus balanoides. Marine Ecology Progress Series, 75:91–100, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/75/m075p091.pdf.

Barange:1991:DFE

[BGC91] M. Barange, M. J. Gibbons, and M. Carola. Diet and feeding of Euphausia hanseni and Nematoscelis megalops (Euphausiacea) in the northern Benguela Current: ecological significance of vertical space partitioning. Marine Ecology Progress Series, 73:173–181, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/73/m073p173.pdf.

Burkholder:1994:CEW

[BGC94] J. M. Burkholder, H. B. Glasgow, Jr., and J. E. Cooke. Comparative effects of water-column nitrate enrichment on eelgrass Zostera marina, shoalgrass Halo-dule wrightii, and widgeongrass Ruppia maritime. Marine Ecology Progress Series, 105:121–138, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/105/m105p121.pdf.

Bustillos-Guzman:1995:SPS

[BGCM95] J. Bustillos-Guzmán, H. Claustre, and J. C. Marty. Specific phytoplankton signatures and their relationship to hydrographic conditions in the coastal northwestern Mediterranean Sea. Marine Ecology Progress Series, 124:247–258, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
124/m124p247.pdf.

Bost:1997:FHF

[BGG^{+}97] C. A. Bost, J. Y. Georges, C. Guinet, Y. Cherel, K. Pütz, J. B. Charrassin, Y. Handrich, T. Zorn, J. Lage, and Y. Le Maho. Foraging habitat and food intake of satellite-tracked king penguins during the austral summer at Crozet Archipelago. Marine Ecology Progress Series, 150:21–33, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
150/m150p021.pdf.

Burkholder:1995:FKL

[BGH95] J. M. Burkholder, H. B. Glasgow, Jr., and C. W. Hobbs. Fish kills linked to a toxic ambush-predator dinoflagellate: distribution and environmental conditions. Marine Ecology Progress Series, 124:43–61, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/124/m124p043.pdf.

Blanchard:1996:QST

[BGR^{+}96] G. F. Blanchard, J. M. Guarini, P. Richard, P. h. Gros, and F. Mornet. Quantifying the short-term temperature effect on light-saturated photosynthesis of intertidal microphytobenthos. Marine Ecology Progress Series, 134:309–313, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
134/m134p309.pdf.

Barthel:1991:NIB

[BGT91] D. Barthel, J. Gutt, and O. S. Tendal. New information on the biology of Antarctic deep-water sponges derived from underwater photography. Marine Ecology Progress Series, 69: 303–307, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/69/m069p303.pdf.

Brussaard:1996:IPB

[BGvDR96] C. P. D. Brussaard, G. J. Gast, F. C. van Duyl, and R. Riegman. Impact of phytoplankton bloom magnitude on a pelagic microbial food web. Marine Ecology Progress Series, 144: 211–221, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/144/m144p211.pdf.

Brunetti:1992:MTT

[BGVF92] R. Brunetti, M. Gabriele, P. Valerio, and O. Fumagalli. The micronucleus test: temporal pattern of baseline frequency in Mytilus galloprovincialis. Marine Ecology Progress Series, 83:75–78, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/83/m083p075.pdf.

Bertness:1996:WDS

[BGW96] M. D. Bertness, S. D. Gaines, and R. A. Wahle. Wind-driven settlement patterns in the acorn barnacle Semibalanus balanoides. Marine Ecology Progress Series, 137:103–110, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
137/m137p103.pdf.

Bell:1991:MLF

[BH91] S. S. Bell and G. R. F. Hicks. Marine landscapes and faunal recruitment: a field test with seagrasses and copepods. Marine Ecology Progress Series, 73:61–68, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/73/m073p061.pdf.

Bautista:1992:CGC

[BH92a] B. Bautista and R. P. Harris. Copepod gut contents, ingestion rates and grazing impact on phytoplankton in relation to size structure of zooplankton and phytoplankton during a spring bloom. Marine Ecology Progress Series, 82:41–50, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
82/m082p041.pdf.

Bianchi:1992:RMU

[BH92b] G. Bianchi and T. Hoisseter. Relative merits of using numbers and biomass in fish community studies. Marine Ecology Progress Series, 85:25–33, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/85/m085p025.pdf.

Bochdansky:1992:EAAa

[BH92c] A. B. Bochdansky and G. J. Herndl. Ecology of amorphous aggregations (marine snow) in the Northern Adriatic Sea. III. Zooplankton interactions with marine snow. Marine Ecology Progress Series, 87:135–146, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/87/m087p135.pdf.

Bochdansky:1992:EAAb

[BH92d] A. B. Bochdansky and G. J. Herndl. Ecology of amorphous aggregations (marine snow) in the Northern Adriatic Sea. V. Role of fecal pellets in marine snow. Marine Ecology Progress Series, 89:297–303, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/89/m089p297.pdf.

Brey:1992:GRP

[BH92e] Thomas Brey and Stefan Hain. Growth, reproduction and production of Lissarca notorcadensis (Bivalvia: Philobryidae) in the Weddell Sea, Antarctica. Marine Ecology Progress Series, 82:219–226, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/82/m082p219.pdf.

Brettar:1993:NOP

[BH93] I. Brettar and M. G. Hofle. Nitrous oxide producing heterotrophic bacteria from the water column of the central Baltic: abundance and molecular identification. Marine Ecology Progress Series, 94:253–265, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/94/m094p253.pdf.

Beiras:1994:EDM

[BH94a] R. Beiras and E. His. Effects of dissolved mercury on embryogenesis, survival, growth and metamorphosis of Crassostrea gigas oyster larvae. Marine Ecology Progress Series, 113: 95–103, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/113/m113p095.pdf.

Bowen:1994:ODG

[BH94b] W. D. Bowen and G. D. Harrison. Offshore diet of grey seals Halichoerus grypus near Sable Island, Canada. Marine Ecology Progress Series, 112:1–11, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/112/m112p001.pdf.

Beiras:1995:EDM

[BH95a] R. Beiras and E. His. Effects of dissolved mercury on embryogenesis, survival and growth of Mytilus galloprovincialis mussel larvae. Marine Ecology Progress Series, 126:185–189, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
126/m126p185.pdf.

Buskey:1995:ETU

[BH95b] E. J. Buskey and C. J. Hyatt. Effects of the Texas (USA) ’brown tide’ alga on planktonic grazers. Marine Ecology Progress Series, 126:285–292, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/126/m126p285.pdf.

Bell:1997:DMA

[BH97a] S. S. Bell and M. O. Hall. Drift macroalgal abundance in seagrass beds: investigating large-scale associations with physical and biotic attributes. Marine Ecology Progress Series, 147:277–283, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/147/m147p277.pdf.

Bolton:1997:PVV

[BH97b] T. F. Bolton and J. N. Havenhand. Physiological versus viscosity-induced effects of water temperature on the swimming and sinking velocity of larvae of the serpulid polychaete Galeolaria caespitosa. Marine Ecology Progress Series, 159:209–218, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/159/m159p209.pdf.

Burrows:1998:MPD

[BH98] M. T. Burrows and S. J. Hawkins. Modelling patch dynamics on rocky shores using deterministic cellular automata. Marine Ecology Progress Series, 167:1–13, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/167/m167p001.pdf.

Berman:1994:RAB

[BHG94] T. Berman, H.-G. Hoppe, and K. Gocke. Response of aquatic bacterial populations to substrate enrichment. Marine Ecology Progress Series, 104:173–184, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/104/m104p173.pdf.

Bratbak:1992:GMV

[BHH^{+}92] G. Bratbak, O. H. Haslund, M. Heldal, A. Naess, and T. Roeggen. Giant marine viruses? Marine Ecology Progress Series, 85:201–202, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/85/m085p201.pdf.

Butler:1995:CDF

[BHH^{+}95] M. J. Butler IV, J. H. Hunt, W. F. Herrnkind, M. J. Childress, R. Bertelsen, W. Sharp, T. Matthews, J. M. Field, and H. G. Marshall. Cascading disturbances in Florida Bay, USA: cyanobacteria blooms, sponge mortality, and implications for juvenile spiny lobsters Panulirus argus. Marine Ecology Progress Series, 129:119–125, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/129/m129p119.pdf.

Blackburn:1996:OLE

[BHHL96] T. H. Blackburn, P. O. J. Hall, S. Hulth, and A. Landén. Organic-N loss by efflux and burial associated with a low efflux of inorganic N and with nitrate assimilation in Arctic sediments (Svalbard, Norway). Marine Ecology Progress Series, 141:283–293, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/141/m141p283.pdf.

Beckmann:1999:ILA

[BHQ99] M. Beckmann, T. Harder, and P. Y. Qian. Induction of larval attachment and metamorphosis in the serpulid polychaete Hydroides elegans by dissolved free amino acids: mode of action in laboratory bioassays. Marine Ecology Progress Series, 190:167–178, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/190/m190p167.pdf.

Bratbak:1992:IVB

[BHT^{+}92] G. Bratbak, M. Heldal, T. F. Thingstad, B. Riemann, and O. H. Haslund. Incorporation of viruses into the budget of microbial c-transfer a first approach. Marine Ecology Progress Series, 83:273–280, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/83/m083p273.pdf.

Broekhuizen:1998:SPD

[BHT98] N. Broekhuizen, M. Hadfield, and A. H. Taylor. Seasonal photoadaptation and diatom dynamics in temperate waters. Marine Ecology Progress Series, 175:227–239, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
175/m175p227.pdf.

Bortolus:1999:ESA

[BI99] A. Bortolus and O. Iribarne. Effects of the SW Atlantic burrowing crab Chasmagnathus granulata on a Spartina salt marsh. Marine Ecology Progress Series, 178:79–88, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
178/m178p079.pdf.

Bianchi:1991:DAC

[Bia91] G. Bianchi. Demersal assemblages of the continental shelf and slope edge between the Gulf of Tehuantepec (Mexico) and the Gulf of Papagayo (Costa Rica). Marine Ecology Progress Series, 73:121–140, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/73/m073p121.pdf.

Bianchi:1992:SDA

[Bia92] G. Bianchi. Study of the demersal assemblages of the continental shelf and upper slope off Congo and Gabon, based on the trawl surveys of the RV ‘Dr Fridtjof Nansen’. Marine Ecology Progress Series, 85:9–23, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/85/m085p009.pdf.

Bigelow:1992:AGP

[Big92] K. A. Bigelow. Age and growth in paralarvae of the mesopelagic squid Abraiia trigonura based on daily growth increments in statoliths. Marine Ecology Progress Series, 82: 31–40, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/82/m082p031.pdf.

Beukers:1995:UIM

[BJB95] J. S. Beukers, G. P. Jones, and R. M. Buckley. Use of implant microtags for studies on populations of small reef fish. Marine Ecology Progress Series, 125:61–66, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/125/m125p061.pdf.

Bak:1998:BSF

[BJdJ^{+}98] R. P. M. Bak, M. Joenje, I. de Jong, D. Y. M. Lambrechts, and G. Nieuwland. Bacterial suspension feeding by coral reef benthic organisms. Marine Ecology Progress Series, 175: 285–288, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/175/m175p285.pdf.

Brey:1996:ANN

[BJTB96] T. Brey, A. Jarre-Teichmann, and O. Borlich. Artificial neural network versus multiple linear regression: predicting P/B ratios from empirical data. Marine Ecology Progress Series, 140:251–256, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/140/m140p251.pdf.

Berghahn:1990:EIB

[BK90] R. Berghahn and M. Karakiri. Experimental induction of biological tags in otoliths of 0-group plaice Pleuronectes platessa by starvation, temperature, and UV-B radiation. Marine Ecology Progress Series, 67:227–233, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
67/m067p227.pdf.

Bjornsen:1991:DBB

[BK91] P. K. Bjørnsen and J. Kuparinen. Determination of bacterioplankton biomass, net production and growth efficiency in the Southern Ocean. Marine Ecology Progress Series, 71: 185–194, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/71/m071p185.pdf.

Bulman:1992:DFC

[BK92] C. M. Bulman and J. A. Koslow. Diet and food consumption of a deep-sea fish, orange roughy Hoplostethus atlanticus (Pisces: Trachichthyidae), off southeastern Australia. Marine Ecology Progress Series, 82:115–129, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
82/m082p115.pdf.

Bjorkman:1994:BIO

[BK94] K. Bjórkman and D. M. Karl. Bioavailability of inorganic and organic phosphorus compounds to natural assemblages of microorganisms m Hawaiian coastal waters. Marine Ecology Progress Series, 111:265–273, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/111/m111p265.pdf.

Beer:1996:PMM

[BK96a] S. Beer and E. Koch. Photosynthesis of marine macroalgae and seagrasses in globally changing CO_{2} environments. Marine Ecology Progress Series, 141:199–204, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
141/m141p199.pdf.

Brzezinski:1996:SPS

[BK96b] M. A. Brzezinski and C. A. Kosman. Silica production in the Sargasso Sea during spring 1989. Marine Ecology Progress Series, 142:39–45, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/142/m142p039.pdf.

Baamstedt:1998:EPC

[BK98] U. Båmstedt and K. Karlson. Euphausiid predation on copepods in coastal waters of the Northeast Atlantic. Marine Ecology Progress Series, 172:149–168, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/172/m172p149.pdf.

Blust:1992:ESU

[BKB92] R. Blust, E. Kockelbergh, and M. Baillieul. Effect of salinity on the uptake of cadmium by the brine shrimp Arternia franciscana. Marine Ecology Progress Series, 84:245–254, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
84/m084p245.pdf.

Bruggemann:1994:CAF

[BKB94] J. H. Bruggemann, M. W. M. Kuyper, and A. M. Breeman. Comparative analysis of foraging and habitat use by the sympatric Caribbean parrotfish Scarus vetuta and Sparisoma viride (Scaridae). Marine Ecology Progress Series, 112: 51–66, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/112/m112p051.pdf.

Burrows:1999:FMP

[BKH99] M. T. Burrows, K. Kawai, and R. N. Hughes. Foraging by mobile predators on a rocky shore: underwater TV observations of movements of blennies Lipophrys pholis and crabs Carcinus maenas. Marine Ecology Progress Series, 187:237–250, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
187/m187p237.pdf.

Bjornsen:1993:DSP

[BKK^{+}93] P. K. Bjørnsen, H. Kaas, H. Kaas, T. G. Nielsen, M. Olesen, and K. Richardson. Dynamics of a subsurface phytoplankton maximum in the Skagerrak. Marine Ecology Progress Series, 95:279–294, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/95/m095p279.pdf.

Bracher:1999:PPP

[BKL99] A. U. Bracher, B. M. A. Kroon, and M. I. Lucas. Primary production, physiological state and composition of phytoplankton in the Atlantic sector of the Southern Ocean. Marine Ecology Progress Series, 190:1–16, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/190/m190p001.pdf.

Barranguet:1998:FCP

[BKP98] C. Barranguet, J. Kromkamp, and J. Peene. Factors controlling primary production and photosynthetic characteristics of intertidal microphytobenthos. Marine Ecology Progress Series, 173:117–126, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/173/m173p117.pdf.

Bening:1992:LSP

[BKS92] J.-C Bening, L. Karbe, and G. Schupfner. Liquid/solid phase extraction of water samples used for toxicity testing in the German Bight. Marine Ecology Progress Series, 91:233– 236, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/91/m091p233.pdf.

Babcock:1999:CCS

[BKS^{+}99] R. C. Babcock, S. Kelly, N. T. Shears, J. W. Walker, and T. J. Willis. Changes in community structure in temperate marine reserves. Marine Ecology Progress Series, 189:125–134, 1999.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
189/m189p125.pdf.

Brickner:1993:SCB

[BKWML93] I. Brickner, E. Kramarsky-Winter, O. Mokady, and Y. Loya. Speciation in the coral-boring bivalve Lithophaga purpurea: evidence from ecological, biochemical and SEM analysis. Marine Ecology Progress Series, 101:139–145, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
101/m101p139.pdf.

Baldwin:1994:ESH

[BL94a] John R. Baldwin and James R. Lovvorn. Expansion of seagrass habitat by the exotic Zostera japonica and its use by dabbling ducks and brant in Boundary Bay, British Columbia. Marine Ecology Progress Series, 103:119–127, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
103/m103p119.pdf.

Bertram:1994:PRD

[BL94b] D. R. Bertram and W. C. Leggett. Predation risk during the early life history periods of fishes: separating the effects of size and age. Marine Ecology Progress Series, 109:105–114, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
109/m109p105.pdf.

Boetius:1994:RME

[BL94c] A. Boetius and K. Lochte. Regulation of microbial enzymatic degradation of organic matter in deep-sea sediments. Marine Ecology Progress Series, 104:299–307, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
104/m104p299.pdf.

Brown:1995:UEB

[BL95] C. L. Brown and S. N. Luoma. Use of the euryhaline bivalve Potamocorbula amurensis as a biosentinel species to assess trace metal contamination in San Francisco Bay.

Marine Ecology Progress Series, 124:129–142, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
124/m124p129.pdf.

Boetius:1996:EOE

[BL96] A. Boetius and K. Lochte. Effect of organic enrichments on hydrolytic potentials and growth of bacteria in deep-sea sediments. Marine Ecology Progress Series, 140:239–250, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
140/m140p239.pdf.

Beninger:1997:RCP

[BL97] P. G. Beninger and M. Le Pennec. Reproductive characteristics of a primitive bivalve from a deep-sea reducing environment: giant gametes and their significance in Acharax alinae (Cryptodonta: Solemyidae). Marine Ecology Progress Series, 157:195–206, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/157/m157p195.pdf.

Blanchard:1990:OMD

[Bla90] G. F. Blanchard. Overlapping microscale dispersion patterns of meiofauna and microphytobenthos. Marine Ecology Progress Series, 68:101–111, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/68/m068p101.pdf.

Bricelj:1991:IDC

[BLC91] V. M. Bricelj, J. H. Lee, and A. D. Cembella. Influence of dinoflagellate cell toxicity on uptake and loss of paralytic shellfish toxins in the northern quahog Mercenaria mercenaria. Marine Ecology Progress Series, 74:33–46, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
74/m074p033.pdf.

Bricelj:1990:UKP

[BLCA90] V. M. Bricelj, J. H. Lee, A. D. Cembella, and D. M. Anderson. Uptake kinetics of paralytic shellfish toxins from the dinoflagellate Alexandrium fundyense in the mussel Mytilus edulis. Marine Ecology Progress Series, 63:177–188, 1990.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
63/m063p177.pdf.

Brown:1994:TRS

[BLD94] B. E. Brown, M. D. A. Le Tissier, and R. P. Dunne. Tissue retraction in the scleractinian coral Coelosehs mayeri, its effect upon coral pigmentation, and preliminary implications for heat balance. Marine Ecology Progress Series, 105:209–218, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/105/m105p209.pdf.

Belgrano:1995:SSEb

[BLDF95a] A. Belgrano, P. Legendre, J. M. Dewarumez, and S. Frontier. Spatial structure and ecological variation of meroplankton on the Belgian-Dutch coast of the North Sea. Marine Ecology Progress Series, 128:51–59, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/128/m128p051.pdf.

Belgrano:1995:SSEa

[BLDF95b] A. Belgrano, P. Legendre, J. M. Dewarumez, and S. Frontier. Spatial structure and ecological variation of meroplankton on the French-Belgian coast of the North Sea. Marine Ecology Progress Series, 128:43–50, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/128/m128p043.pdf.

Bullard:1999:SIL

[BLH99] S. G. Bullard, N. L. Lindquist, and M. E. Hay. Susceptibility of invertebrate larvae to predators: how common are post-capture larval defenses? Marine Ecology Progress Series, 191:153–161, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/191/m191p153.pdf.

Botton:1992:OTL

[BLJ92] M. L. Botton, R. E. Loveland, and T. R. Jacobsen. Overwintering by trilobite larvae of the horseshoe crab Limulus polyphemus on a sandy beach of Delaware Bay (New Jersey, USA). Marine Ecology Progress Series, 88:289–292, 1992.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
88/m088p289.pdf.

Bak:1996:LTC

[BLJ^{+}96] R. P. M. Bak, D. Y. M. Lambrechts, M. Joenje, G. Nieuwland, and M. L. J. Van Veghel. Long-term changes on coral reefs in booming populations of a competitive colonial ascidian. Marine Ecology Progress Series, 133:303–306, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
133/m133p303.pdf.

Bertrand:1999:ACM

[BLJ99] A. Bertrand, R. Le Borgne, and E. Josse. Acoustic characterisation of micronekton distribution in French Polynesia. Marine Ecology Progress Series, 191:127–140, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
191/m191p127.pdf.

Bates:1993:DBD

[BLKA93] S. S. Bates, C. Léger, B. A. Keafer, and D. M. Anderson. Discrimination between domoic-acid-producing and nontoxic forms of the diatom Pseudonitzschia pungens using immunofluorescence. Marine Ecology Progress Series, 100:185– 195, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/100/m100p185.pdf.

Bratbak:1995:VAR

[BLM^{+}95] G. Bratbak, M. Levasseur, S. Michaud, G. Cantin, E. Fernández, B. R. Heimdal, and M. Heldal. Viral activity in relation to Emiliania huxleyi blooms: a mechanism of DMSP release? Marine Ecology Progress Series, 128:133–142, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
128/m128p133.pdf.

Blomqvist:1991:QSS

[Blo91] S. Blomqvist. Quantitative sampling of soft-bottom sediments: problems and solutions. Marine Ecology Progress Series, 72:295–304, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/72/m072p295.pdf.

Burford:1994:SPO

[BLR94] M. A. Burford, B. G. Long, and P. C. Rothlisberg. Sedimentary pigments and organic carbon in relation to microalgal and benthic faunal abundance in the Gulf of Carpentaria. Marine Ecology Progress Series, 103:111–117, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
103/m103p111.pdf.

Brown:1990:EEI

[BLST90] B. E. Brown, M. D. A. Le Tissier, T. P. Scoffin, and A. W. Tudhope. Evaluation of the environmental impact of dredging on intertidal coral reefs at Ko Phuket, Thailand, using ecological and physiological parameters. Marine Ecology Progress Series, 65:273–281, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/65/m065p273.pdf.

Blum:1993:SAR

[Blu93] L. K. Blum. Spartina altemiflora root dynamics in a Virginia marsh. Marine Ecology Progress Series, 102:169–178, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
102/m102p169.pdf.

Bradley:1990:PCSa

[BM90a] P. M. Bradley and J. T. Morris. Physical characteristics of salt marsh sediments: ecological implications. Marine Ecology Progress Series, 61:245–252, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/61/m061p245.pdf.

Bradley:1990:PCSb

[BM90b] P. M. Bradley and J. T. Morris. Physical characteristics of salt marsh sediments: ecological implications. Marine Ecology Progress Series, 61:245–252, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/61/m061p245.pdf.

Black:1991:IHP

[BM91a] K. P. Black and P. J. Moran. Influence of hydrodynamics on the passive dispersal and initial recruitment of larvae of Acanthaster plana (Echinodermata: Asteroidea) on the Great Barrier Reef. Marine Ecology Progress Series, 69:55–65, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
69/m069p055.pdf.

Blum:1991:MGA

[BM91b] L. K. Blum and A. L. Mills. Microbial growth and activity during the initial stages of seagrass decomposition. Marine Ecology Progress Series, 70:73–82, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/70/m070p073.pdf.

Bailey:1994:APL

[BM94a] K. M. Bailey and S. A. Macklin. Analysis of patterns in larval walleye pollock Theragra chalcogramma survival and wind mixing events in Shelikof Strait, Gulf of Alaska. Marine Ecology Progress Series, 113:1–12, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/113/m113p001.pdf.

Baker:1994:DMP

[BM94b] S. M. Baker and R. Mann. Description of metamorphic phases in the oyster Crassostrea virginica and effects of hypoxia on metamorphosis. Marine Ecology Progress Series, 104:91–99, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/104/m104p091.pdf.

Bock:1994:SVD

[BM94c] M. J. Bock and D. C. Miller. Seston variability and daily growth in Mercenaria mercenaria on an intertidal sandflat. Marine Ecology Progress Series, 114:117–127, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
114/m114p117.pdf.

Boehlert:1994:VOO

[BM94d] G. W. Boehlert and B. C. Mundy. Vertical and onshoreoffshore distributional patterns of tuna larvae in relation to physical habitat features. Marine Ecology Progress Series, 107:1–13, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/107/m107p001.pdf.

Broekhuizen:1995:PAC

[BM95a] N. Broekhuizen and E. McKenzie. Patterns of abundance for Calanus and smaller copepods in the North Sea: time series decomposition of two CPR data sets. Marine Ecology Progress Series, 118:103–120, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/118/m118p103.pdf.

Brumbaugh:1995:TMB

[BM95b] R. D. Brumbaugh and J. R. McConaugha. Time to metamorphosis of blue crab Callinectes sapidus megalopae: effects of benthic macroalgae. Marine Ecology Progress Series, 129:113–118, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/129/m129p113.pdf.

Bak:1998:CPS

[BM98] R. P. M. Bak and E. H. Meesters. Coral population structure: the hidden information of colony size-frequency distributions. Marine Ecology Progress Series, 162:301–306, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
162/m162p301.pdf.

Beare:1999:CEP

[BM99a] D. J. Beare and E. McKenzie. Connecting ecological and physical time-series: the potential role of changing seasonality. Marine Ecology Progress Series, 178:307–309, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
178/m178p307.pdf.

Beare:1999:TPS

[BM99b] D. J. Beare and E. McKenzie. Temporal patterns in the surface abundance of Calanus finmarchicus and C. helgolandicus in the northern North Sea (1958–1996) inferred from continuous plankton recorder data. Marine Ecology Progress Series, 190:241–251, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/190/m190p241.pdf.

Black:1995:ALF

[BMBD95] K. Black, P. Moran, D. Burrage, and G. De’ath. Association of low-frequency currents and crown-of-thorns starfish outbreaks. Marine Ecology Progress Series, 125:185–194, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
125/m125p185.pdf.

Bidigare:1990:EPS

[BMD^{+}90] Robert R. Bidigare, John Marra, Tommy D. Dickey, Rudolfo Iturriaga, Karen S. Baker, Raymond C. Smith, and Hasong Pak. Evidence for phytoplankton succession and chromatic adaptation in the Sargasso Sea during spring 1985. Marine Ecology Progress Series, 60:113–122, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
60/m060p113.pdf.

Blanco:1997:PDK

[BMFR97] J. Blanco, A. Moroño, J. Franco, and M. I. Reyero. PSP detoxification kinetics in the mussel Mytilus galloprovincialis. One- and two-compartment models and the effect of some environmental variables. Marine Ecology Progress Series, 158: 165–175, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/158/m158p165.pdf.

Burkholder:1999:FKB

[BMG99] J. M. Burkholder, M. A. Mallin, and H. B. Glasgow, Jr. Fish kills, bottom-water hypoxia, and the toxic Pfiesteria complex in the Neuse River and Estuary. Marine Ecology Progress Series, 179:301–310, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/179/m179p301.pdf. See reply

[PPFP99].

Black:1991:NMS

[BMH91] K. P. Black, P. J. Moran, and L. S. Hammond. Numerical models show coral reefs can be self-seeding. Marine Ecology Progress Series, 74:1–11, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/74/m074p001.pdf.

Buhl-Mortensen:1993:MFA

[BMH93] L. Buhl-Mortensen and T. Heisseter. Mollusc fauna along an offshore-fjord gradient. Marine Ecology Progress Series, 97: 209–224, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/97/m097p209.pdf.

Barange:1993:ISA

[BMHD93] M. Barange, D. G. M. Miller, I. Hampton, and T. T. Dunne. Internal structure of Antarctic krill Euphausia superba swarms based on acoustic observations. Marine Ecology Progress Series, 99:205–213, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/99/m099p205.pdf.

Berges:1994:FEP

[BMHH94] J. A. Berges, D. J. S. Montagnes, C. L. Hurd, and P. J. Harrison. Fitting ecological and physiological data to rectangular hyperbolae: a comparison of methods using Monte Carlo simulations. Marine Ecology Progress Series, 114:175–183, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
114/m114p175.pdf.

Barlow:1995:DSF

[BMP^{+}95] R. G. Barlow, R. F. C. Mantoura, R. D. Peinert, A. E. J. Miller, and T. W. Fileman. Distribution, sedimentation and fate of pigment biomarkers following thermal stratification in the western Alboran Sea. Marine Ecology Progress Series, 125:279–291, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/125/m125p279.pdf.

Baldi:1997:CLR

[BMS^{+}97] F. Baldi, A. Minacci, A. Saliot, L. Mejanelle, P. Mozetic, V. Turk, and A. Malej. Cell lysis and release of particulate polysaccharides in extensive marine mucilage assessed by lipid biomarkers and molecular probes. Marine Ecology Progress Series, 153:45–57, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/153/m153p045.pdf.

Blaber:1995:TDD

[BMSF95] S. J. M. Blaber, D. A. Milton, G. C. Smith, and M. J. Farmer. Trawl discards in the diets of tropical seabirds of the northern Great Barrier Reef, Australia. Marine Ecology Progress Series, 127:1–13, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/127/m127p001.pdf.

Bargagli:1998:BMA

[BMSHC98] R. Bargagli, F. Monaci, J. C. Sanchez-Hernandez, and D. Cateni. Biomagnification of mercury in an Antarctic marine coastal food web. Marine Ecology Progress Series, 169: 65–76, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/169/m169p065.pdf.

Bianchi:1992:SAA

[BMTF92] M. Bianchi, D. Marty, J.-L. Teyssie, and S. W. Fowler. Strictly aerobic and anaerobic bacteria associated with sinking particulate matter and zooplankton fecal pellets. Marine Ecology Progress Series, 88:55–60, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/88/m088p055.pdf.

Boyd:1996:LVI

[BMV^{+}96] P. W. Boyd, D. L. Muggli, D. E. Varela, R. H. Goldblatt, R. Chretien, K. J. Orians, and P. J. Harrison. In vitro iron enrichment experiments in the NE subarctic Pacific. Marine Ecology Progress Series, 136:179–193, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
136/m136p179.pdf.

Baldwin:1991:OFP

[BN91a] B. S. Baldwin and R. I. E. Newell. Omnivorous feeding by planktotrophic arvae of the eastern oyster Crassostrea virginica. Marine Ecology Progress Series, 78:285–301, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
78/m078p285.pdf.

Bjornsen:1991:DSH

[BN91b] P. K. Bjornsen and T. G. Nielsen. Decimeter scale heterogeneity in the plankton during a pycnocline bloom of Gyrodinium aureolum. Marine Ecology Progress Series, 73:263–267, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
73/m073p263.pdf.

Baldwin:1995:RID

[BN95] B. S. Baldwin and R. I. E. Newell. Relative importance of different size food particles in the natural diet of oyster larvae Crassostrea virginica. Marine Ecology Progress Series, 120:135–145, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/120/m120p135.pdf.

Bano:1997:SBF

[BNK^{+}97] N. Bano, M. U. Nisa, N. Khan, M. Saleem, P. J. Harrison, S. I. Ahmed, and F. Azam. Significance of bacteria in the flux of organic matter in the tidal creeks of the mangrove ecosystem of the Indus River delta, Pakistan. Marine Ecology Progress Series, 157:1–12, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/157/m157p001.pdf.

Bathmann:1990:CGP

[BNvB90] U. V. Bathmann, T. T. Noji, and B. von Bodungen. Copepod grazing potential in late winter in the Norwegian Sea — a factor in the control of spring phytoplankton growth? Marine Ecology Progress Series, 60:225–233, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
60/m060p225.pdf.

Blackburn:1994:AMC

[BNW94] T. H. Blackburn, D. B. Nedwell, and W. J. Wiebe. Active mineral cycling in a Jamaican seagrass sediment. Marine Ecology Progress Series, 110:233–239, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
110/m110p233.pdf.

Brzezinski:1990:SAC

[BOC90] M. A. Brzezinski, R. J. Olson, and S. W. Chisholm. Silicon availability and cell-cycle progression in marine diatoms. Marine Ecology Progress Series, 67:83–96, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
67/m067p083.pdf.

Bourgoin:1990:MESa

[Bou90a] B. P. Bourgoin. Mytilus edulis shell as a bioindicator of lead pollution considerations on bioavailability and variability. Marine Ecology Progress Series, 61:253–262, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
61/m061p253.pdf.

Bourgoin:1990:MESb

[Bou90b] B. P. Bourgoin. Mytilus edulis shell as a bioindicator of lead pollution considerations on bioavailability and variability. Marine Ecology Progress Series, 61:253–262, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
61/m061p253.pdf.

Bowen:1997:RMM

[Bow97] W. D. Bowen. Role of marine mammals in aquatic ecosystems. Marine Ecology Progress Series, 158:267–274, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
158/m158p267.pdf.

Brodeur:1992:EEV

[BP92] R. D. Brodeur and W. G. Pearcy. Effects of environmental variability on trophic interactions and food web structure in a pelagic upwelling ecosystem. Marine Ecology Progress Series, 84:101–119, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/84/m084p101.pdf.

Bundy:1993:ICA

[BP93] M. H. Bundy and G.-A. Paffenhofer. Innervation of copepod antennules investigated using laser scanning confocal microscopy. Marine Ecology Progress Series, 102:1–14, 1993.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
102/m102p001.pdf.

Burford:1994:PCP

[BP94] M. A. Burford and P. C. Pollard. Pigment contaminants in polycarbonate filters. Marine Ecology Progress Series, 103: 203–206, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/103/m103p203.pdf.

Boucher:1996:LSB

[BP96a] N. P. Boucher and B. B. Prézelin. An in situ biological weighting function for UV inhibition of phytoplankton carbon fixation in the Southern Ocean. Marine Ecology Progress Series, 144:223–236, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/144/m144p223.pdf.

Bundy:1996:AFF

[BP96b] M. H. Bundy and G. A. Paffenhöfer. Analysis of flow fields associated with freely swimming calanoid copepods. Marine Ecology Progress Series, 133:99–113, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
133/m133p099.pdf.

Baier:1997:ESP

[BP97a] C. T. Baier and J. E. Purcell. Effects of sampling and preservation on apparent feeding by chaetognaths. Marine Ecology Progress Series, 146:37–42, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/146/m146p037.pdf.

Baier:1997:TIC

[BP97b] C. T. Baier and J. E. Purcell. Trophic interactions of chaetognaths, larval fish, and zooplankton in the South Atlantic Bight. Marine Ecology Progress Series, 146:43–53, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
146/m146p043.pdf.

Brown:1998:MVD

[BP98] E. G. Brown and G. J. Pierce. Monthly variation in the diet of harbour seals in inshore waters along the southeast Shetland (UK) coastline. Marine Ecology Progress Series, 167:275–289, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/167/m167p275.pdf.

Benoit:1999:IVG

[BP99] H. P. Benoît and P. Pepin. Individual variability in growth rate and the timing of metamorphosis in yellowtail flounder Pleuronectes ferrugineus. Marine Ecology Progress Series, 184:231–244, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/184/m184p231.pdf.

Brotas:1998:STP

[BPC98] V. Brotas and M. R. Plante-Cuny. Spatial and temporal patterns of microphytobenthic taxa of estuarine tidal flats in the Tagus Estuary (Portugal) using pigment analysis by HPLC. Marine Ecology Progress Series, 171:43–57, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
171/m171p043.pdf.

Bochdansky:1995:IZG

[BPH95] A. B. Bochdansky, S. Puskaric, and G. J. Herndl. Influence of zooplankton grazing on free dissolved enzymes in the sea. Marine Ecology Progress Series, 121:53–63, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
121/m121p053.pdf.

Breitburg:1995:LDS

[BPL95] D. L. Breitburg, M. A. Palmer, and T. Loher. Larval distributions and the spatial patterns of settlement of an oyster reef fish: responses to flow and structure. Marine Ecology Progress Series, 125:45–60, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/125/m125p045.pdf.

Baden:1990:EOD

[BPR90] S. P. Baden, L. Pihl, and R. Rosenberg. Effects of oxygen depletion on the ecology, blood physiology and fishery of the Norway lobster Nephrops norvegicus. Marine Ecology Progress Series, 67:141–155, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/67/m067p141.pdf.

Becerro:1998:IVC

[BPS98] M. A. Becerro, V. J. Paul, and J. Starmer. Intracolonial variation in chemical defenses of the sponge Cacospongia sp. and its consequences on generalist fish predators and the specialist nudibranch predator Glossodoris pallida. Marine Ecology Progress Series, 168:187–196, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/168/m168p187.pdf.

Bryan:1997:ILS

[BQKC97] P. J. Bryan, P. Y. Qian, J. L. Kreider, and F. S. Chia. Induction of larval settlement and metamorphosis by pharmacological and conspecific associated compounds in the serpulid polychaete Hydroides elegans. Marine Ecology Progress Series, 146:81–90, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/146/m146p081.pdf.

Barreiro:1999:AIN

[BQR99] R. Barreiro, M. Quintela, and J. M. Ruiz. Aphally and imposex in Nucella lapillus from Galicia (NW Spain): incidence, geographical distribution and consequences for the biomonitoring of TBT contamination. Marine Ecology Progress Series, 185:229–238, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/185/m185p229.pdf.

Brettar:1991:DCB

[BR91a] I. Brettar and G. Rheinheimer. Denitrification in the Central Baltic: evidence for H_{2}S-oxidation as motor of denitrification at the oxic–anoxic interface. Marine Ecology Progress Series, 77:157–169, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/77/m077p157.pdf.

Bussmann:1991:SRB

[BR91b] I. Bussmann and W. Reichardt. Sulfate-reducing bacteria in temporarily oxic sediments with bivalves. Marine Ecology Progress Series, 78:97–102, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/78/m078p097.pdf.

Bernard:1994:SVM

[BR94] C. Bernard and F. Rassoulzadegan. Seasonal variations of mixotrophic ciliates in the northwest Mediterranean Sea. Marine Ecology Progress Series, 108:295–301, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
108/m108p295.pdf.

Burkhardt:1997:CAA

[BR97] S. Burkhardt and U. Riebesell. CO_{2} availability affects elemental composition (C:N:P) of the marine diatom Skeletonema costatum. Marine Ecology Progress Series, 155:67–76, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/155/m155p067.pdf.

Brandini:1993:PBA

[Bra93] F. P. Brandini. Phytoplankton biomass in an Antarctic coastal environment during stable water conditions — implications for the iron limitation theory. Marine Ecology Progress Series, 93:267–275, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/93/m093p267.pdf.

Brandt:1995:PFC

[Bra95] A. Brandt. Peracarid fauna (Crustacea, Malacostraca) of the Northeast Water Polynya off Greenland: documenting close benthic-pelagic coupling in the Westwind Trough. Marine Ecology Progress Series, 121:39–51, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/121/m121p039.pdf.

Browman:1999:TSN

[BRA^{+}99] Howard I. Browman, M. Ruse, D. Allchin, D. L. Hull, and A. J. Underwood. Theme section: Negative results.

Marine Ecology Progress Series, 191:301–309, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
191/m191p301.pdf.

Breitburg:1999:LWQ

[BRC99] D. L. Breitburg, K. A. Rose, and J. H. Cowan, Jr. Linking water quality to larval survival: predation mortality of fish larvae in an oxygen-stratified water column. Marine Ecology Progress Series, 178:39–54, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/178/m178p039.pdf.

Brey:1995:TRM

[Bre95] T. Brey. Temperature and reproductive metabolism in macrobenthic populations. Marine Ecology Progress Series, 125: 87–93, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/125/m125p087.pdf.

Breitburg:1996:CMR

[Bre96] D. L. Breitburg. Consumer mobility and the relative importance of consumption and competition following physical disturbance. Marine Ecology Progress Series, 138:83–92, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
138/m138p083.pdf.

Bianchi:1997:SCP

[BRL97] T. S. Bianchi, C. Rolff, and C. D. Lambert. Sources and composition of particulate organic carbon in the Baltic Sea: the use of plant pigments and lignin-phenols as biomarkers. Marine Ecology Progress Series, 156:25–31, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
156/m156p025.pdf.

Barrett:1990:DSP

[BRLM90] R. T. Barrett, N. Rov, J. Loen, and W. A. Montevecchi. Diets of shags Phalacrocorax aristotelis and cormorants P carbo in Norway and possible implications for gadoid stock recruitment. Marine Ecology Progress Series, 66:205–218, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
66/m066p205.pdf.

Brussaard:1995:EGS

[BRN^{+}95] C. P. D. Brussaard, R. Riegman, A. A. M. Noordeloos, G. C. Cadée, H. Witte, A. J. Kop, G. Nieuwland, F. C. van Duyl, and R. P. M. Bak. Effects of grazing, sedimentation and phytoplankton cell lysis on the structure of a coastal pelagic food web. Marine Ecology Progress Series, 123:259–271, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
123/m123p259.pdf.

Brockmann:1997:MSP

[BRNH97] U. H. Brockmann, T. Raabe, K. Nagel, and M. Haarich. Measurement strategy of PRISMA: design and realisation. Marine Ecology Progress Series, 156:245–254, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
156/m156p245.pdf.

Brock:1990:IDB

[Bro90] V. Brock. Intergeneric distances between Ostrea, Crassostrea, and Saccostrea, studied by means of crossed immunoelectrophoresis. Marine Ecology Progress Series, 68:59–63, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/68/m068p059.pdf.

Brock:1993:EMP

[Bro93] V. Brock. Effects of mercury on physiological condition and content of the biomarker ALA in the oyster Ostrea edulis. Marine Ecology Progress Series, 96:169–175, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
96/m096p169.pdf.

Brogan:1994:DRL

[Bro94] M. W. Brogan. Distribution and retention of larval fishes near reefs in the Gulf of California. Marine Ecology Progress Series, 115:1–13, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/115/m115p001.pdf.

Browman:1995:CCR

[Bro95a] Howard I. Browman. Commentaries on current research trends in recruitment studies. Marine Ecology Progress Series, 128:305–310, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/128/m128p305.pdf. See comment

[BS96a].

Browman:1995:I

[Bro95b] Howard I. Browman. Introduction. Marine Ecology Progress Series, 128:305, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Browman:1996:PPIa

[Bro96a] Howard I. Browman. Predator–prey interactions in the sea: commentaries on the role of turbulence. Marine Ecology Progress Series, 139:301–312, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/139/m139p301.pdf.

Browman:1996:PPIb

[Bro96b] Howard I. Browman. Predator–prey interactions in the sea: commentaries on the role of turbulence. Marine Ecology Progress Series, 139:301–302, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/139/m139p301.pdf.

Brodeur:1998:LSO

[Bro98] R. D. Brodeur. In situ observations of the association between juvenile fishes and scyphomedusae in the Bering Sea. Marine Ecology Progress Series, 163:11–20, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/163/m163p011.pdf.

Burford:1995:STD

[BRW95] M. A. Burford, P. C. Rothlisberg, and Y. Wang. Spatial and temporal distribution of tropical phytoplankton species and biomass in the Gulf of Carpentaria, Australia. Marine Ecology Progress Series, 118:255–266, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/118/m118p255.pdf.

Burkhardt:1999:SCI

[BRZ99] S. Burkhardt, U. Riebesell, and I. Zondervan. Stable carbon isotope fractionation by marine phytoplankton in response to daylength, growth rate, and CO_{2} availability. Marine Ecology Progress Series, 184:31–41, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/184/m184p031.pdf.

Bottger-Schnack:1992:CSV

[BS92] R. Böttger-Schnack. Community structure and vertical distribution of cyclopoid and poecilostomatoid copepods in the Red Sea. III. Re-evaluation for separating a new species of Oncaea. Marine Ecology Progress Series, 80:301–304, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
80/m080p301.pdf.

Bologna:1993:KBH

[BS93] P. A. X. Bologna and R. S. Steneck. Kelp beds as habitat for American lobster Homarus americanus. Marine Ecology Progress Series, 100:127–134, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/100/m100p127.pdf.

Barbeau:1994:PEP

[BS94] M. A. Barbeau and R. E. Scheibling. Procedural effects of prey tethering experiments: predation of juvenile scallops by crabs and sea stars. Marine Ecology Progress Series, 111:305–310, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/111/m111p305.pdf.

Behrends:1995:IAA

[BS95a] G. Behrends and G. Schneider. Impact of Aurelia aurita medusae (Cnidaria, Scyphozoa) on the standing stock and community composition of mesozooplankton in the Kiel Bight (western Baltic Sea). Marine Ecology Progress Series, 127: 39–45, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/127/m127p039.pdf.

Blackwell:1995:ESC

[BS95b] B. F. Blackwell and J. A. Sinclair. Evidence of secondary consumption of fish by double-crested cormorants. Marine Ecology Progress Series, 123:1–4, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/123/m123p001.pdf.

Bottger-Schnack:1995:SDM

[BS95c] R. Böttger-Schnack. Summer distribution of micro- and small mesozooplankton in the Red Sea and Gulf of Aden, with special reference to non-calanoid copepods. Marine Ecology Progress Series, 118:81–102, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/118/m118p081.pdf.

Bell:1996:LNF

[BS96a] M. V. Bell and J. R. Sargent. Lipid nutrition and fish recruitment: Comment on Browman HI, 128:305–310 (1995). Marine Ecology Progress Series, 134:315–316, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
128/m128p305.pdf. See [Bro95a].

Browman:1996:ETP

[BS96b] Howard I. Browman and A. B. Skiftesvik. Effects of turbulence on the predation cycle of fish larvae: comments on some of the issues. Marine Ecology Progress Series, 139: 309–312, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/139/m139p301.pdf.

Boudreau:1991:BRP

[BSB91] B. Boudreau, Y. Simard, and E. Bourget. Behavioural responses of the planktonic stages of the American lobster Homarus americanus to thermal gradients, and ecological implications. Marine Ecology Progress Series, 76:13–23, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
76/m076p013.pdf.

Breitburg:1994:ELD

[BSD^{+}94] D. L. Breitburg, N. Steinberg, S. DuBeau, C. Cooksey, and E. D. Houde. Effects of low dissolved oxygen on predation on estuarine fish larvae. Marine Ecology Progress Series, 104:235–246, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/104/m104p235.pdf.

Best:1995:SMH

[BSF95] P. B. Best, K. Sekiguchi, and K. P. Findlay. A suspended migration of humpback whales Megaptera novaeangliae on the west coast of South Africa. Marine Ecology Progress Series, 118:1–12, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/118/m118p001.pdf.

Blanchard:1997:KTR

[BSG^{+}97] G. F. Blanchard, P. G. Sauriau, V. Cariou-Le Gall, D. Gouleau, M. J. Garet, and F. Olivier. Kinetics of tidal resuspension of microbiota: testing the effects of sediment cohesiveness and bioturbation using flume experiments. Marine Ecology Progress Series, 151:17–25, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/151/m151p017.pdf.

Bennett:1994:FCS

[BSGM94] B. A. Bennett, C. R. Smith, B. Glaser, and H. L. Maybaum. Faunal community structure of a chemoautotrophic assemblage on whale bones in the deep northeast Pacific Ocean. Marine Ecology Progress Series, 108:205–223, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
108/m108p205.pdf.

Buckley:1991:WFPa

[BSH^{+}91a] L. J. Buckley, A. S. Smigielski, T. A. Halavik, E. M. Caldarone, B. R. Burns, and G. C. Laurence. Winter flounder Pseudopleuronectes americanus reproductive success. I among-location variability in size and survival of larvae reared in the laboratory. Marine Ecology Progress Series, 74:117– 124, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/74/m074p117.pdf.

Buckley:1991:WFPb

[BSH^{+}91b] L. J. Buckley, A. S. Smigielski, T. A. Halavik, E. M. Caldarone, B. R. Burns, and G. C. Laurence. Winter flounder Pseudopleuronectes americanus reproductive success. II. Effects of spawning time and female size on size, composition and viability of eggs and larvae. Marine Ecology Progress Series, 74:125–135, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/74/m074p125.pdf.

Barbeau:1994:SAT

[BSH^{+}94] M. A. Barbeau, R. E. Scheibling, B. G. Hatcher, L. H. Taylor, and A. W. Hennigar. Survival analysis of tethered juvenile sea scallops Placopecten magellanicus in field experiments: effects of predators, scallop size and density, site and season. Marine Ecology Progress Series, 115:243–256, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
115/m115p243.pdf.

Beninger:1997:PPL

[BSJ97] P. G. Beninger and S. D. St-Jean. Particle processing on the labial palps of Mytilus edulis and Placopecten magellanicus (Mollusca: Bivalvia). Marine Ecology Progress Series, 147:117–127, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/147/m147p117.pdf.

Beninger:1993:GFM

[BSJPW93] Peter G. Beninger, Sylvie St-Jean, Yves Poussart, and J. Evan Ward. Gill function and mucocyte distribution in Placopecten magellanicus and Mytilus edulis (Mollusca: Bivalvia): the role of mucus in particle transport. Marine Ecology Progress Series, 98:275–282, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/98/m098p275.pdf.

Beach:1995:PRU

[BSMS95] K. S. Beach, C. M. Smith, T. Michael, and H. W. Shin. Photosynthesis in reproductive unicells of Ulva fasciata and Enteromorpha flexuosa: implications for ecological success.

Marine Ecology Progress Series, 125:229–237, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
125/m125p229.pdf.

Brock:1993:MSS

[BSP93] J. Brock, S. Sathyendranath, and T. Piatt. Modelling the seasonality of subsurface light and primary production in the Arabian Sea. Marine Ecology Progress Series, 101:209– 221, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/101/m101p209.pdf.

Brock:1998:BOC

[BSP98] J. C. Brock, S. Sathyendranath, and T. Platt. Biohydrooptical classification of the northwestern Indian Ocean. Marine Ecology Progress Series, 165:1–15, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
165/m165p001.pdf.

Baco:1999:PRW

[BSP^{+}99] A. R. Baco, C. R. Smith, A. S. Peek, G. K. Roderick, and R. C. Vrijenhoek. The phylogenetic relationships of whalefall vesicomyid clams based on mitochondrial COI DNA sequences. Marine Ecology Progress Series, 182:137–147, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
182/m182p137.pdf.

Biermann:1992:ISR

[BSS92] C. H. Biermann, G. O. Schinner, and R. R. Strathmann. Influence of solar radiation, microalgal fouling, and current on deposition site and survival of embryos of a dorid nudibranch gastropod. Marine Ecology Progress Series, 86:205–215, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
86/m086p205.pdf.

Baranovic:1993:TFZ

[BSVK93] A. Baranovic, M. Solic, T. Vucetic, and N. Krstulovic. Temporal fluctuations of zooplankton and bacteria in the middle Adriatic Sea. Marine Ecology Progress Series, 92:65–75, 1993.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
92/m092p065.pdf.

Burdick:1993:IAM

[BSW93] D. M. Burdick, F. T. Short, and J. Wolf. An index to assess and monitor the progression of wasting disease in eelgrass Zostera marina. Marine Ecology Progress Series, 94: 83–90, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/94/m094p083.pdf.

Boyd:1995:SPWa

[BSW^{+}95] P. W. Boyd, S. Strom, F. A. Whitney, S. Doherty, M. E. Wen, P. J. Harrison, C. S. Wong, and D. E. Varela. The NE subarctic Pacific in winter: I. Biological standing stocks. Marine Ecology Progress Series, 128:11–24, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/128/m128p011.pdf.

Bower:1999:CPA

[BSY^{+}99] J. R. Bower, M. P. Seki, R. E. Young, K. A. Bigelow, J. Hirota, and P. Flament. Cephalopod paralarvae assemblages in Hawaiian Islands waters. Marine Ecology Progress Series, 185:203–212, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/185/m185p203.pdf.

Borkman:1993:PSB

[BT93] D. G. Borkman and J. T. Turner. Plankton studies in Buzzards Bay, Massachusetts USA. II. Nutrients, chlorophyll-a and phaeopigments, 1987 to 1990. Marine Ecology Progress Series, 100:27–34, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/100/m100p027.pdf.

Baba:1999:ETS

[BTKK99] K. Baba, M. Tada, T. Kawajiri, and Y. Kuwahara. Effects of temperature and salinity on spawning of the brackish water bivalve Corbicula japonica in Lake Abashiri, Hokkaido, Japan. Marine Ecology Progress Series, 180:213–221, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
180/m180p213.pdf.

Bailey:1994:EDM

[BTYO94] T. G. Bailey, J. J. Torres, M. J. Youngbluth, and G. P. Owen. Effect of decompression on mesopelagic gelatinous zooplankton: a comparison of in situ and shipboard measurements of metabolism. Marine Ecology Progress Series, 113: 13–27, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/113/m113p013.pdf.

Baird:1993:CST

[BU93] D. Baird and R. E. Ulanowicz. Comparative study on the trophic structure, cycling and ecosystem properties of four tidal estuaries. Marine Ecology Progress Series, 99:221–237, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/99/m099p221.pdf.

Buchholz:1991:MCG

[Buc91] F. Buchholz. Moult cycle and growth of Antarctic krill Euphausia superba in the laboratory. Marine Ecology Progress Series, 69:217–229, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/69/m069p217.pdf.

Buck:1992:ADP

[BUCP^{+}92] K. R. Buck, L. Uttal-Cooke, C. H. Pilskaln, D. L. Roelke, M. C. Villac, G. A. Fryxell, L. Cifuentes, and F. P. Chavez. Autecology of the diatom Pseudonitzschia australis, a domoic acid producer, from Monterey Bay, California. Marine Ecology Progress Series, 84:293–302, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/84/m084p293.pdf.

Burrows:1994:FTS

[Bur94] M. T. Burrows. Foraging time strategy of small juvenile plaice: a laboratory study of diel and tidal behaviour patterns with Artemia prey and shrimp predators. Marine Ecology Progress Series, 115:31–39, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/115/m115p031.pdf.

Busch:1996:TEE

[Bus96] A. Busch. Transition from endogenous to exogenous nutrition: larval size parameters determining the start of external feeding and size of prey ingested by Ruegen spring herring Clupea harengus. Marine Ecology Progress Series, 130:39–46, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
130/m130p039.pdf.

Buskey:1997:BCF

[Bus97] E. J. Buskey. Behavioral components of feeding selectivity of the heterotrophic dinoflagellate Protoperidinium pellucidum. Marine Ecology Progress Series, 153:77–89, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
153/m153p077.pdf.

Buskey:1998:ECP

[Bus98] E. J. Buskey. Energetic cost of position-holding behavior in the planktonic mysid Mysidium columbiae. Marine Ecology Progress Series, 172:139–147, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/172/m172p139.pdf.

Bjork:1999:PTD

[BUWB99] M. Björk, J. Uku, A. Weil, and S. Beer. Photosynthetic tolerances to desiccation of tropical intertidal seagrasses. Marine Ecology Progress Series, 191:121–126, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
191/m191p121.pdf.

Beninger:1999:OPR

[BV99] P. G. Beninger and A. Veniot. The oyster proves the rule: mechanisms of pseudofeces transport and rejection on the mantle of Crassostrea virginica and C. gigas. Marine Ecology Progress Series, 190:179–188, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/190/m190p179.pdf.

Bruggemann:1994:FSPa

[BvB94] J. H. Bruggemann, M. J. H. van Oppen, and A. M. Breeman. Foraging by the stoplight parrotfish Sparisoma viride. I. Food selection in different, socially determined habitats. Marine Ecology Progress Series, 106:41–55, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/106/m106p041.pdf.

Buitenhuis:1996:TIO

[BvBBV96] E. Buitenhuis, J. van Bleijswijk, D. Bakker, and M. Veldhuis. Trends in inorganic and organic carbon in a bloom of Emiliania huxleyi in the North Sea. Marine Ecology Progress Series, 143:271–282, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/143/m143p271.pdf.

Bruggemann:1996:BSI

[BvKvRB96] J. H. Bruggemann, A. M. van Kessel, J. M. van Rooij, and A. M. Breeman. Bioerosion and sediment ingestion by the Caribbean parrotfish Scarus vetula and Sparisoma viride: implications of fish size, feeding mode and habitat use. Marine Ecology Progress Series, 134:59–71, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/134/m134p059.pdf.

Bucke:1992:QAM

[BVL92] D. Bucke, A. D. Vethaak, and T. Lang. Quantitative assessment of melanomacrophage centres (MMCs) in dab Limanda limanda along a pollution transect in the German Bight. Marine Ecology Progress Series, 91:193–196, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
91/m091p193.pdf.

Brzezinski:1998:SPC

[BVL98] M. A. Brzezinski, T. A. Villareal, and F. Lipschultz. Silica production and the contribution of diatoms to new and primary production in the central North Pacific. Marine Ecology Progress Series, 167:89–104, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/167/m167p089.pdf.

Boucher:1991:FCD

[BVP91] N. Boucher, D. Vaulot, and F. Partensky. Flow cytometric determination of phyto-plankton DNA in cultures and oceanic populations. Marine Ecology Progress Series, 71:75–84, 1991.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
71/m071p075.pdf.

Beninger:1999:PPR

[BVP99] P. G. Beninger, A. Veniot, and Y. Poussart. Principles of pseudofeces rejection on the bivalve mantle: integration in particle processing. Marine Ecology Progress Series, 178: 259–269, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/178/m178p259.pdf.

Buchsbaum:1991:ARN

[BVS^{+}91] R. Buchsbaum, I. Valiela, T. Swain, M. Dzierzeski, and S. Allen. Available and refractory nitrogen in detritus of coastal vascular plants and macroalgae. Marine Ecology Progress Series, 72:131–143, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/72/m072p131.pdf.

Bett:1994:SBQ

[BVV^{+}94] B. J. Bett, A. Vanreusel, M. Vincx, T. Soltwedel, O. Pfannkuche, P. J. D. Lambshead, A. J. Gooday, T. Ferrero, and A. Dinet. Sampler bias in the quantitative study of deep-sea meiobenthos. Marine Ecology Progress Series, 104:197–203, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
104/m104p197.pdf.

Beer:1998:MPR

[BVW^{+}98] S. Beer, B. Vilenkin, A. Weil, M. Veste, L. Susel, and A. Eshel. Measuring photosynthetic rates in seagrasses by pulse amplitude modulated (PAM) fluorometry. Marine Ecology Progress Series, 174:293–300, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/174/m174p293.pdf.

Boylan:1993:SBM

[BW93] J. M. Boylan and E. L. Wenner. Settlement of brachyuran megalopae in a South Carolina, USA, estuary. Marine Ecology Progress Series, 97:237–246, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/97/m097p237.pdf.

Bralewska:1995:HDE

[BW95] J. M. Bralewska and Z. Witek. Heterotrophic dinoflagellates in the ecosystem of the Gulf of Gdansk. Marine Ecology Progress Series, 117:241–248, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/117/m117p241.pdf.

Brierley:1996:ATS

[BW96] A. S. Brierley and J. L. Watkins. Acoustic targets at South Georgia and the South Orkney Islands during a season of krill scarcity. Marine Ecology Progress Series, 138:51–61, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
138/m138p051.pdf.

Bidwell:1998:TEZ

[BWB98] J. R. Bidwell, K. W. Wheeler, and T. R. Burridge. Toxicant effects on the zoospore stage of the marine macroalga Ecklonia radiata (Phaeophyta: Laminariales). Marine Ecology Progress Series, 163:259–265, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/163/m163p259.pdf.

Bunt:1991:MSA

[BWHC91] J. S. Bunt, W. T. Williams, J. F. Hunter, and H. J. Clay. Mangrove sequencing analysis of zonation in a complete river system. Marine Ecology Progress Series, 72:289–294, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
72/m072p289.pdf.

Backhaus:1999:PCR

[BWHK99] J. O. Backhaus, H. Wehde, E. N. Hegseth, and J. Kämpf. Phyto-convection: the role of oceanic convection in primary production. Marine Ecology Progress Series, 189:77–92, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
189/m189p077.pdf.

Boyd:1995:SPWb

[BWHW95] P. W. Boyd, F. A. Whitney, P. J. Harrison, and C. S. Wong. The NE subarctic Pacific in winter: II. Biological rate processes. Marine Ecology Progress Series, 128:25–34, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
128/m128p025.pdf.

Bringmann:1999:DDS

[BWL^{+}99] G. Bringmann, K. Wolf, T. Lanz, A. Haase, J. Hiort, P. Proksch, and W. E. G. Müller. Direct demonstration of spatial water distribution in the sponge Suberites domuncula by in vivo NMR imaging. Marine Ecology Progress Series, 189:307–310, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/189/m189p307.pdf.

Brierley:1997:IVK

[BWM97] A. S. Brierley, J. L. Watkins, and A. W. A. Murray. Interannual variability in krill abundance at South Georgia. Marine Ecology Progress Series, 150:87–98, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/150/m150p087.pdf.

Blanchette:1999:AMF

[BWRH99] C. A. Blanchette, S. E. Worcester, D. Reed, and S. J. Holbrook. Algal morphology, flow, and spatially variable recruitment of surfgrass Phyllospadix torreyi. Marine Ecology Progress Series, 184:119–128, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/184/m184p119.pdf.

Bailey:1995:CCM

[BYO95] T. G. Bailey, M. J. Youngbluth, and G. P. Owen. Chemical composition and metabolic rates of gelatinous zooplankton from midwater and benthic boundary layer environments off Cape Hatteras, North Carolina, USA. Marine Ecology Progress Series, 122:121–134, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/122/m122p121.pdf.

Bythell:1990:NUR

[Byt90] J. C. Bythell. Nutrient uptake in the reef-building coral Acropora palmata at natural environmental concentrations. Marine Ecology Progress Series, 68:65–69, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
68/m068p065.pdf.

Brown:1992:MIZ

[BZ92] B. E. Brown and N. P. Zamani. Mitotic indices of zooxanthellae: a comparison of techniques based on nuclear and cell division frequencies. Marine Ecology Progress Series, 89: 99–102, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/89/m089p099.pdf.

Buma:1996:URM

[BZSG96] A. G. J. Buma, H. J. Zemmelink, K. Sjollema, and W. W. C. Gieskes. UVB radiation modifies protein and photosynthetic pigment content, volume and ultrastructure of marine diatoms. Marine Ecology Progress Series, 142:47–54, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
142/m142p047.pdf.

Cho:1990:BSB

[CA90] Byung C. Cho and Farooq Azam. Biogeochemical significance of bacterial biomass in the ocean’s euphotic zone. Marine Ecology Progress Series, 63:253–259, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
63/m063p253.pdf.

Cartes:1992:CFH

[CA92] J. E. Cartes and P. Abello. Comparative feeding habits of polychelid lobsters in the Western Mediterranean deepsea communities. Marine Ecology Progress Series, 84:139– 150, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/84/m084p139.pdf.

Clarke:1993:MLM

[CA93] K. R. Clarke and M. Ainsworth. A method of linking multivariate community structure to environmental variables. Marine Ecology Progress Series, 92:205–219, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
92/m092p205.pdf.

Cho:1995:UDB

[CA95] B. C. Cho and F. Azam. Urea decomposition by bacteria in the Southern California Bight and its implications for the mesopelagic nitrogen cycle. Marine Ecology Progress Series, 122:21–26, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/122/m122p021.pdf.

Calbet:1996:ECF

[CA96a] A. Calbet and M. Alcaraz. Effects of constant and fluctuating food supply on egg production rates of Acartia grani (Copepoda: Calanoida). Marine Ecology Progress Series, 140: 33–39, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/140/m140p033.pdf.

Choat:1996:GLA

[CA96b] J. H. Choat and L. M. Axe. Growth and longevity in acanthurid fishes; an analysis of otolith increments. Marine Ecology Progress Series, 134:15–26, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/134/m134p015.pdf.

Calbet:1997:GSR

[CA97] A. Calbet and M. Alcaraz. Growth and survival rates of early developmental stages of Acartia grani (Copepoda: Calanoida) in relation to food concentration and fluctuations in food supply. Marine Ecology Progress Series, 147:181–186, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
147/m147p181.pdf.

Curran:1998:VTF

[CA98] M. C. Curran and K. W. Able. The value of tethering fishes (winter flounder and tautog) as a tool for assessing predation rates. Marine Ecology Progress Series, 163:45–51, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
163/m163p045.pdf.

Calbet:1999:LCC

[CA99a] A. Calbet and S. Agustí. Latitudinal changes of copepod egg production rates in Atlantic waters: temperature and food availability as the main driving factors. Marine Ecology Progress Series, 181:155–162, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/181/m181p155.pdf.

Charmantier:1999:OOP

[CA99b] G. Charmantier and K. Anger. Ontogeny of osmoregulation in the palaemonid shrimp Palaemonetes argentinus (Crustacea: Decapoda). Marine Ecology Progress Series, 181:125– 129, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/181/m181p125.pdf.

Caurant:1994:EBF

[CAATS94] F. Caurant, J. C. Amiard, C. Amiard-Triquet, and P. G. Sauriau. Ecological and biological factors controlling the concentrations of trace elements (As, Cd, Cu, Hg, Se, Zn) in delpinids Globicephala melas from the North Atlantic Ocean. Marine Ecology Progress Series, 103:207–219, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
103/m103p207.pdf.

Caley:1995:CDT

[Cal95a] M. J. Caley. Community dynamics of tropical reef fishes: local patterns between latitudes. Marine Ecology Progress Series, 129:7–18, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/129/m129p007.pdf.

Caley:1995:RFC

[Cal95b] M. J. Caley. Reef-fish community structure and dynamics: an interaction between local and larger-scale processes? Marine Ecology Progress Series, 129:19–29, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/129/m129p019.pdf.

Choat:1996:GLF

[CAL96] J. H. Choat, L. M. Axe, and D. C. Lou. Growth and longevity in fishes of the family Scaridae. Marine Ecology Progress Series, 145:33–41, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/145/m145p033.pdf.

Cammen:1991:ABP

[Cam91] L. M. Cammen. Annual bacterial production in relation to benthic microalgal production and sediment oxygen uptake in an intertidal sandflat and an intertidal mudflat. Marine Ecology Progress Series, 71:13–25, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/71/m071p013.pdf.

Campana:1996:YCS

[Cam96] S. E. Campana. Year-class strength and growth rate in young Atlantic cod Gadus morhua. Marine Ecology Progress Series, 135:21–26, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/135/m135p021.pdf.

Campana:1997:URN

[Cam97] S. E. Campana. Use of radiocarbon from nuclear fallout as a dated marker in the otoliths of haddock Melanogrammus aeglefinus. Marine Ecology Progress Series, 150:49–56, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
150/m150p049.pdf.

Carman:1990:MUR

[Car90] K. R. Carman. Mechanisms of uptake of radioactive labels by meiobenthic copepods during grazing experiments. Marine Ecology Progress Series, 68:71–83, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/68/m068p071.pdf.

Cartes:1993:DDW

[Car93] J. E. Cartes. Diets of deep-water pandalid shrimps on the Western Mediterranean slope. Marine Ecology Progress Series, 96:49–61, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/96/m096p049.pdf.

Cameron:1992:MCA

[CB92] P. Cameron and J. Berg. Morphological and chromosomal aberrations during embryonic development in dab Limanda limanda. Marine Ecology Progress Series, 91:163–169, 1992.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
91/m091p163.pdf.

Crump:1996:PAB

[CB96] B. C. Crump and J. A. Baross. Particle-attached bacteria and heterotrophic plankton associated with the Columbia River estuarine turbidity maxima. Marine Ecology Progress Series, 138:265–273, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/138/m138p265.pdf.

Cusson:1997:ITH

[CB97] M. Cusson and E. Bourget. Influence of topographic heterogeneity and spatial scales on the structure of the neighbouring intertidal endobenthic macrofaunal community. Marine Ecology Progress Series, 150:181–193, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/150/m150p181.pdf.

Charpy:1998:PPF

[CB98] L. Charpy and J. Blanchot. Photosynthetic picoplankton in French Polynesian atoll lagoons: estimation of taxa contribution to biomass and production by flow cytometry. Marine Ecology Progress Series, 162:57–70, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/162/m162p057.pdf.

Chiswell:1999:RLJ

[CB99] S. M. Chiswell and J. D. Booth. Rock lobster Jasus edwardsii larval retention by the wairarapa eddy off New Zealand. Marine Ecology Progress Series, 183:227–240, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
183/m183p227.pdf.

Cerrano:1999:COP

[CBA^{+}99] C. Cerrano, G. Bavestrello, A. Arillo, U. Benatti, S. Bonpadre, R. Cattaneo-Vietti, L. Gaggero, M. Giovine, L. Leone, G. Lucchetti, and M. Sarà. Calcium oxalate production in the marine sponge Chondrosia reniformis. Marine Ecology Progress Series, 179:297–300, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/179/m179p297.pdf.

Christian:1990:JRP

[CBB90] R. R. Christian, W. L. Bryant, Jr., and M. M. Brinson. Juncus roemerianus production and decomposition along gradients of salinity and hydroperiod. Marine Ecology Progress Series, 68:137–145, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/68/m068p137.pdf.

Cherrier:1996:UTL

[CBD96] J. Cherrier, J. E. Bauer, and E. R. M. Druffel. Utilization and turnover of labile dissolved organic matter by bacterial heterotrophs in eastern North Pacific surface waters. Marine Ecology Progress Series, 139:267–279, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
139/m139p267.pdf.

Christaki:1996:ARC

[CBDC96] U. Christaki, S. Belviso, J. R. Dolan, and M. Corn. Assessment of the role of copepods and ciliates in the release to solution of particulate DMSP. Marine Ecology Progress Series, 141:119–127, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/141/m141p119.pdf.

Cailliau:1999:SPI

[CBG^{+}99] C. Cailliau, S. Belviso, M. Goutx, A. Bedo, Y. Park, and E. Charriaud. Sedimentation pathways in the Indian sector of the Southern Ocean during a production regime dominated by regeneration. Marine Ecology Progress Series, 190: 53–67, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/190/m190p053.pdf.

Canino:1991:TGD

[CBI91] M. F. Canino, K. M. Bailey, and L. S. Incze. Temporal and geographic differences in feeding and nutritional condition of walleye pollock larvae Theragra chalcogramma in Shelikof Strait, Gulf of Alaska. Marine Ecology Progress Series, 79:27–35, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/79/m079p027.pdf.

Clark:1996:FAS

[CBL96a] B. M. Clark, B. A. Bennett, and S. J. Lamberth. Factors affecting spatial variability in seine net catches of fish in the surf zone of False Bay, South Africa. Marine Ecology Progress Series, 131:17–34, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/131/m131p017.pdf.

Clark:1996:TVS

[CBL96b] B. M. Clark, B. A. Bennett, and S. J. Lamberth. Temporal variations in surf zone fish assemblages from False Bay, South Africa. Marine Ecology Progress Series, 131:35–47, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
131/m131p035.pdf.

Chan:1992:UZC

[CBRD92] H. M. Chan, P. Bjerregaard, P. S. Rainbow, and M. H. Depledge. Uptake of zinc and cadmium by two different populations of shore crabs Carcinus maenus at different salinities. Marine Ecology Progress Series, 86:91–97, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
86/m086p091.pdf.

Christian:1991:MYD

[CBS91] R. R. Christian, J. N. Boyer, and D. W. Stanley. Multi-year distribution patterns of nutrients within the Neuse River Estuary, North Carolina. Marine Ecology Progress Series, 71: 259–274, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/71/m071p259.pdf.

Conceicao:1998:GRB

[CBT98] M. B. Conceição, J. D. D. Bishop, and J. P. Thorpe. Genetic relationships between ecologically divergent species of talitrid amphipod (Crustacea). Marine Ecology Progress Series, 165:225–233, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/165/m165p225.pdf.

Cahoon:1993:BMP

[CBTM93] L. B. Cahoon, G. R. Beretich, Jr., C. J. Thomas, and A. M. McDonald. Benthic microalgal production at Stellwagen Bank, Massachusetts Bay, USA. Marine Ecology Progress Series, 102:179–185, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/102/m102p179.pdf.

Chang:1990:SSP

[CC90] J. Chang and E. J. Carpenter. Species-specific phytoplankton growth rates via diel DNA synthesis cycles. IV. Evaluation of the magnitude of error with computer-simulated cell populations. Marine Ecology Progress Series, 65:293–304, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
65/m065p293.pdf.

Castro:1991:EFA

[CC91a] L. R. Castro and R. K. Cowen. Environmental factors affecting the early life history of bay anchovy Anchoa mitchilli in Great South Bay, New York. Marine Ecology Progress Series, 76:235–247, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/76/m076p235.pdf.

Chang:1991:SSP

[CC91b] J. Chang and E. J. Carpenter. Species-specific phytoplankton growth rates via diel DNA synthesis cycles. V. Application to natural populations in Long Island Sound. Marine Ecology Progress Series, 78:115–122, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/78/m078p115.pdf.

Cahoon:1992:BMP

[CC92a] L. B. Cahoon and J. E. Cooke. Benthic microalgal production in Onslow Bay, North Carolina, USA. Marine Ecology Progress Series, 84:185–196, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/84/m084p185.pdf.

Chen:1992:ANH

[CC92b] Jiann-Chu Chen and Sheue-Feng Chen. Accumulation of nitrite in hemolymph of Penaeus japonicus. Marine Ecology Progress Series, 83:305–308, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/83/m083p305.pdf.

Chen:1996:HAU

[CC96] J. C. Chen and P. G. Chia. Hemolymph ammonia and urea and nitrogenous excretions of Scylla serrata at different temperature and salinity levels. Marine Ecology Progress Series, 139:119–125, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/139/m139p119.pdf.

Chen:1997:ASA

[CC97a] J. C. Chen and J. M. Chen. Arginase specific activity and nitrogenous excretion of Penaeus japonicus exposed to elevated ambient ammonia. Marine Ecology Progress Series, 153:197–202, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/153/m153p197.pdf.

Clancy:1997:EWT

[CC97b] M. Clancy and J. S. Cobb. Effect of wind and tidal advection on distribution patterns of rock crab Cancer irroratus megalopae in Block Island Sound, Rhode Island. Marine Ecology Progress Series, 152:217–225, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/152/m152p217.pdf.

Ceccherelli:1998:HES

[CC98] G. Ceccherelli and F. Cinelli. Habitat effect on spatiotemporal variability in size and density of the introduced alga Caulerpa taxifolia. Marine Ecology Progress Series, 163: 289–294, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/163/m163p289.pdf.

Ceccherelli:1999:RVF

[CC99] G. Ceccherelli and F. Cinelli. The role of vegetative fragmentation in dispersal of the invasive alga Caulerpa taxifolia in the Mediterranean. Marine Ecology Progress Series, 182:299–303, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/182/m182p299.pdf.

Cammen:1990:ETS

[CCC90a] L. M. Cammen, S. Corwin, and J. P. Christensen. Electron transport system (ETS) activity as a measure of benthic macrofaunal metabolism. Marine Ecology Progress Series, 65:171–182, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/65/m065p171.pdf.

Carpenter:1990:REN

[CCC^{+}90b] E. J. Carpenter, J. Chang, M. Cottrell, J. Schubauer, H. W. Paerl, B. M. Bebout, and D. G. Capone. Re-evaluation of nitrogenase oxygen-protective mechanisms in the planktonic marine cyanobacterium Trichodesmium. Marine Ecology Progress Series, 65:151–158, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/65/m065p151.pdf.

Chen:1994:NEC

[CCC94] Jiann-Chu Chen, Chung-Tin Chen, and Sha-Yen Cheng. Nitrogen excretion and changes of hemocyanin, protein and free amino acid levels in the hemolymph of Penaeus monodon exposed to different concentrations of ambient ammonia-n at different salinity levels. Marine Ecology Progress Series, 110: 85–94, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/110/m110p085.pdf.

Cho:1994:UBP

[CCCH94] Byung Cheol Cho, Joong-Ki Choi, Chang-Soo Chung, and Gi Hoon Hong. Uncoupling of bacteria and phytoplankton during a spring diatom bloom in the mouth of the Yellow Sea. Marine Ecology Progress Series, 115:181–190, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
115/m115p181.pdf.

Charmantier:1994:OOS

[CCD94] G. Charmantier and M. Charmantier-Daures. Ontogeny of osmoregulation and salinity tolerance in the isopod crustacean Sphaeroma serratum. Marine Ecology Progress Series, 114: 93–102, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/114/m114p093.pdf.

Charmantier:1998:OOG

[CCDA98] G. Charmantier, M. Charmantier-Daures, and K. Anger. Ontogeny of osmoregulation in the grapsid crab Armases miersii (Crustacea, Decapoda). Marine Ecology Progress Series, 164:285–292, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/164/m164p285.pdf.

Cole:1992:APS

[CCG^{+}92] R. G. Cole, R. G. Creese, R. V. Grace, P. Irving, and B. R. Jackson. Abundance patterns of subtidal benthic invertebrates and fishes at the Kermadec Islands. Marine Ecology Progress Series, 82:207–218, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/82/m082p207.pdf.

Chang:1996:ICC

[CCG96] J. Chang, C. C. Chung, and G. C. Gong. Influences of cyclones on chlorophyll-a concentration and Synechococcus abundance in a subtropical western Pacific coastal ecosystem. Marine Ecology Progress Series, 140:199–205, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
140/m140p199.pdf.

Caffrey:1998:CPR

[CCG98] J. M. Caffrey, J. E. Cloern, and C. Grenz. Changes in production and respiration during a spring phytoplankton bloom in San Francisco Bay, California, USA: implications for net ecosystem metabolism. Marine Ecology Progress Series, 172:1–12, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/172/m172p001.pdf.

Chiang:1999:SDD

[CCG99] K. P. Chiang, Y. T. Chen, and G. C. Gong. Spring distribution of diatom assemblages in the East China Sea. Marine Ecology Progress Series, 186:75–86, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/186/m186p075.pdf.

Crossland:1993:URA

[CCGM93] S. Crossland, D. Coates, J. Grahame, and P. J. Mill. Use of random amplified polymorphic DNAs (RAPDs) in separating two sibling species of Littorina. Marine Ecology Progress Series, 96:301–305, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/96/m096p301.pdf.

Coffin:1993:ADO

[CCH93] R. B. Coffin, J. P. Connolly, and P. S. Harris. Availability of dissolved organic carbon to bacterioplankton examined by oxygen utilization. Marine Ecology Progress Series, 101:9–22, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/101/m101p009.pdf.

Chevolot:1991:CIL

[CCJC91] L. Chevolot, J.-C. Cochard, and Yvin J.-C. Chemical induction of larval metamorphosis of Pecten maximus with a note on the nature of naturally occurring triggering substances. Marine Ecology Progress Series, 74:83–89, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
74/m074p083.pdf.

Charpy:1991:POM

[CCR91] L. Charpy and C. J. Charpy-Roubaud. Particulate organic matter fluxes in a Tuamotu atoll lagoon (French Polynesia). Marine Ecology Progress Series, 71:53–63, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
71/m071p053.pdf.

Currie:1998:GPO

[CCR98] W. J. S. Currie, M. R. Claereboudt, and J. C. Roff. Gaps and patches in the ocean: a one-dimensional analysis of planktonic distributions. Marine Ecology Progress Series, 171: 15–21, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/171/m171p015.pdf.

Charpy:1998:EPP

[CCRB98] L. Charpy, C. Charpy-Roubaud, and P. Buat. Excess primary production, calcification and nutrient fluxes of a patch reef (Tikehau atoll, French Polynesia). Marine Ecology Progress Series, 173:139–147, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/173/m173p139.pdf.

Conway:1998:FAE

[CCS98] D. V. P. Conway, S. H. Coombs, and C. Smith. Feeding of anchovy Engraulis encrasicolus larvae in the northwestern Adriatic Sea in response to changing hydrobiological conditions. Marine Ecology Progress Series, 175:35–49, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
175/m175p035.pdf.

Childers:1993:STV

[CCSN93] D. L. Childers, S. Cofer-Shabica, and L. Nakashima. Spatial and temporal variability in marsh-water column interactions in a southeastern USA salt marsh estuary. Marine Ecology Progress Series, 95:25–38, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/95/m095p025.pdf.

Cailliau:1996:CBG

[CCV^{+}96] C. Cailliau, H. Claustre, F. Vidussi, D. Marie, and D. Vaulot. Carbon biomass, and gross growth rates as estimated from 14C pigment labelling, during photoacclimation in Prochlorococcus CCMP 1378. Marine Ecology Progress Series, 145: 209–221, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/145/m145p209.pdf.

Chicharo:1998:ESD

[CCV^{+}98] M. A. Chícharo, L. Chícharo, L. Valdés, E. López-Jamar, and P. Ré. Estimation of starvation and diel variation of the RNA/DNA ratios in field-caught Sardina pilchardus larvae off the north of Spain. Marine Ecology Progress Series, 164:273–283, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/164/m164p273.pdf.

Chen:1994:EPG

[CD94] C. Y. Chen and E. G. Durbin. Effects of pH on the growth and carbon uptake of marine phytoplankton. Marine Ecology Progress Series, 109:83–94, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/109/m109p083.pdf.

Clemmesen:1996:DOS

[CD96] C. Clemmesen and T. Doan. Does otolith structure reflect the nutritional condition of a fish larva? Comparison of otolith structure and biochemical index (RNA/DNA ratio) determined on cod larvae. Marine Ecology Progress Series, 138: 33–39, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/138/m138p033.pdf.

Comtet:1998:PSR

[CD98] T. Comtet and D. Desbruyères. Population structure and recruitment in mytilid bivalves from the Lucky Strike and Menez Gwen hydrothermal vent fields (37^{o}17’ N and 37^{o}50’ N on the Mid-Atlantic Ridge). Marine Ecology Progress Series, 163:165–177, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/163/m163p165.pdf.

Conradi:1999:EZL

[CD99] M. Conradi and M. H. Depledge. Effects of zinc on the life-cycle, growth and reproduction of the marine amphipod Corophium volutator. Marine Ecology Progress Series, 176: 131–138, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/176/m176p131.pdf.

Charpy:1997:POM

[CDG97] L. Charpy, P. Dufour, and N. Garcia. Particulate organic matter in sixteen Tuamotu atoll lagoons (French Polynesia). Marine Ecology Progress Series, 151:55–65, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
151/m151p055.pdf.

Collos:1993:CIE

[CDGF^{+}93] Y. Collos, C. Descolas-Gros, M. Fontugne, A. Mortain-Bertrand, M. J. Chretiennot-Dinet, and M. G. Frikha. Chemical, isotopic and enzymatic monitoring of free and enclosed seawater implications for primary production estimates in incubation bottles. Marine Ecology Progress Series, 93:49–54, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/93/m093p049.pdf.

Capone:1992:MNT

[CDHD92] D. G. Capone, S. E. Dunham, S. G. Horrigan, and L. E. Duguay. Microbial nitrogen transformations in unconsolidated coral reef sediments. Marine Ecology Progress Series, 80:75–88, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/80/m080p075.pdf.

Christiansen:1996:XEF

[CDI96] J. S. Christiansen, R. A. Dalmo, and K. Ingebrigtsen. Xenobiotic excretion in fish with aglomerular kidneys. Marine Ecology Progress Series, 136:303–304, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/136/m136p303.pdf.

Cofino:1992:CP

[CdJS^{+}92] W. P. Cofino, S. A. de Jong, R. Smedes, A. Abarnou, J. P. Boon, I. Oostingh, I. M. Davies, J. Klungsoyr, S. Wilhelmsen, R. J. Law, J. A. Whinnett, D. Schmidt, and S. Wilson. The chemistry programme. Marine Ecology Progress Series, 91:47–56, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/91/m091p047.pdf.

Cebrian:1996:HPO

[CDM^{+}96] J. Cebrián, C. M. Duarte, N. Marbà, S. Enríquez, M. Gallegos, and B. Olesen. Herbivory on Posidonia oceanica: magnitude and variability in the Spanish Mediterranean. Marine Ecology Progress Series, 130:147–155, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
130/m130p147.pdf.

Cebrian:1997:MFP

[CDME97] J. Cebrián, C. M. Duarte, N. Marbà, and S. Enríquez. Magnitude and fate of the production of four co-occurring Western Mediterranean seagrass species. Marine Ecology Progress Series, 155:29–44, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/155/m155p029.pdf.

Connaughton:1993:IPE

[CE93] V. P. Connaughton and C. E. Epifanio. Influence of previous experience on the feeding habits of larval weakfish Cynoscion regalis. Marine Ecology Progress Series, 101:237–241, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
101/m101p237.pdf.

Cameron:1991:PMC

[CED91] A. M. Cameron, R. Endean, and L. M. DeVantier. Predation on massive corals: are devastating population outbreaks of Acanthaster planci novel events9. Marine Ecology Progress Series, 75:251–258, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/75/m075p251.pdf.

Cerchio:1998:EHW

[Cer98] S. Cerchio. Estimates of humpback whale abundance off Kauai, 1989 to 1993: evaluating biases associated with sampling the Hawaiian Islands breeding assemblage. Marine Ecology Progress Series, 175:23–34, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/175/m175p023.pdf.

Culverhouse:1994:ACF

[CES^{+}94] P. F. Culverhouse, R. E. Ellis, R. G. Simpson, R. Williams, R. W. Pierce, and J. T. Turner. Automatic categorisation of five species of Cymatocylis (Protozoa, Tintinnida) by artificial neural network. Marine Ecology Progress Series, 107: 273–280, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/107/m107p273.pdf.

Collie:1997:EBF

[CEV97] J. S. Collie, G. A. Escanero, and P. C. Valentine. Effects of bottom fishing on the benthic megafauna of Georges Bank. Marine Ecology Progress Series, 155:159–172, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
155/m155p159.pdf.

Carvalho:1993:ESD

[CF93a] F. P. Carvalho and S. W. Fowler. An experimental study on the dioaccumulation and turnover of polonium-210 and lead-210 in marine shrimp. Marine Ecology Progress Series, 102:125–133, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/102/m102p125.pdf.

Costa:1993:RCL

[CF93b] P. A. S. Costa and F. C. Fernandes. Reproductive cycle of Loligo sanpaulensis (Cephalopoda: Loliginidae) in the Cabo Frio region, Brazil. Marine Ecology Progress Series, 101: 91–97, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/101/m101p091.pdf.

Carvalho:1994:DTT

[CF94] F. P. Carvalho and S. W. Fowler. A double-tracer technique to determine the relative importance of water and food as sources of polonium-210 to marine prawns and fish. Marine Ecology Progress Series, 103:251–264, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
103/m103p251.pdf.

Cotner:1993:HBM

[CG93] J. B. Cotner, Jr. and W. S. Gardner. Heterotrophic bacterial mediation of ammonium and dissolved free amino acid fluxes in the Mississippi River plume. Marine Ecology Progress Series, 93:75–87, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/93/m093p075.pdf.

Clavier:1999:ASP

[CG99a] J. Clavier and C. Garrigue. Annual sediment primary production and respiration in a large coral reef lagoon (SW New Caledonia). Marine Ecology Progress Series, 191:79–89, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
191/m191p079.pdf.

Conover:1999:SUAa

[CG99b] R. J. Conover and K. R. Gustavson. Sources of urea in Arctic seas: zooplankton metabolism. Marine Ecology Progress Series, 179:41–54, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/179/m179p041.pdf.

Charles:1995:UFI

[CGA95] F. Charles, A. Grémare, and J. M. Amouroux. Utilization of 14C formaldehyde to infer ingestion rates and absorption efficiencies by benthic deposit-feeders. Marine Ecology Progress Series, 127:121–129, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/127/m127p121.pdf.

Coleman:1999:LPS

[CGCLH99] R. A. Coleman, J. D. Goss-Custard, S. E. A. Le V. dit Durell, and S. J. Hawkins. Limpet Patella spp. consumption by oystercatchers Haematopus ostralegus: a preference for solitary prey items. Marine Ecology Progress Series, 183: 253–261, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/183/m183p253.pdf.

Coull:1995:SAJ

[CGFC95] B. C. Coull, J. G. Greenwood, D. R. Fielder, and B. A. Coull. Subtropical Australian juvenile fish eat meiofauna: experiments with winter whiting Sillago maculata and observations on other species. Marine Ecology Progress Series, 125:13–19, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/125/m125p013.pdf.

Campana:1995:EFF

[CGM95] S. E. Campana, J. A. Gagné, and J. W. McLaren. Elemental fingerprinting of fish otoliths using ID-ICPMS. Marine Ecology Progress Series, 122:115–120, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/122/m122p115.pdf.

Cerchio:1998:MHW

[CGNH98] S. Cerchio, C. M. Gabriele, T. F. Norris, and L. M. Herman. Movements of humpback whales between Kauai and Hawaii: implications for population structure and abundance estimation in the Hawaiian Islands. Marine Ecology Progress Series, 175:13–22, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/175/m175p013.pdf.

Coleman:1997:HSR

[CGP97] N. Coleman, A. S. H. Gason, and G. C. B. Poore. High species richness in the shallow marine waters of south-east Australia. Marine Ecology Progress Series, 154:17–26, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
154/m154p017.pdf.

Carlsson:1995:IRH

[CGTB95] P. Carlsson, E. Granéli, P. Tester, and L. Boni. Influences of riverine humic substances on bacteria, protozoa, phytoplankton, and copepods in a coastal plankton community. Marine Ecology Progress Series, 127:213–221, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
127/m127p213.pdf.

Coma:1995:TEB

[CGZ95] R. Coma, J.-M. Gili, and M. Zabala. Trophic ecology of a benthic marine hydroid, Campanularia everta. Marine Ecology Progress Series, 119:211–220, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/119/m119p211.pdf.

Coma:1994:FPC

[CGZR94] R. Coma, J.-M. Gili, M. Zabala, and T. Riera. Feeding and prey capture cycles in the aposymbiontic gorgonian Paramuricea clavata. Marine Ecology Progress Series, 115:257– 270, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/115/m115p257.pdf.

Cowan:1990:GSB

[CH90] J. H. Cowan, Jr. and E. D. Houde. Growth and survival of bay anchovy Anchoa mitchilli larvae in mesocosm enclosures. Marine Ecology Progress Series, 68:47–57, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
68/m068p047.pdf.

Cowan:1993:RPP

[CH93] J. H. Cowan, Jr. and E. D. Houde. Relative predation potentials of scyphomedusae ctenophores and planktivorous fish on ichthyoplankfon in Chesapeake Bay. Marine Ecology Progress Series, 95:55–65, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/95/m095p055.pdf.

Coyle:1994:BAC

[CH94] K. O. Coyle and R. C. Highsmith. Benthic amphipod community in the northern Bering Sea: analysis of potential structuring mechanisms. Marine Ecology Progress Series, 107: 233–244, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/107/m107p233.pdf.

Carr:1995:PEE

[CH95] M. H. Carr and M. A. Hixon. Predation effects on early postsettlement survivorship of coral-reef fishes. Marine Ecology Progress Series, 124:31–42, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/124/m124p031.pdf.

Carroll:1996:RCD

[CH96] M. L. Carroll and R. C. Highsmith. Role of catastrophic disturbance in mediating Nucella-Mytilus interactions in the Alaskan rocky intertidal. Marine Ecology Progress Series, 138:125–133, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/138/m138p125.pdf.

Carlotti:1997:GEP

[CH97a] F. Carlotti and H. J. Hirche. Growth and egg production of female Calanus finmarchicus: an individual-based physiological model and experimental validation. Marine Ecology Progress Series, 149:91–104, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/149/m149p091.pdf.

Crawford:1997:DMP

[CH97b] D. W. Crawford and P. J. Harrison. Direct measurement of pCO_{2} in cultures of marine phytoplankton: how good is the estimate from pH NBS and single point titration of alkalinity? Marine Ecology Progress Series, 158:61–74, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
158/m158p061.pdf.

Cranford:1999:SVF

[CH99] P. J. Cranford and P. S. Hill. Seasonal variation in food utilization by the suspension-feeding bivalve molluscs Mytilus edulis and Placopecten magellanicus. Marine Ecology Progress Series, 190:223–239, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/190/m190p223.pdf.

Chadwick:1991:SDE

[Cha91] N. E. Chadwick. Spatial distribution and the effects of competition on some temperate Scleractinia and Corallimorpharia. Marine Ecology Progress Series, 70:39–48, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
70/m070p039.pdf.

Chapman:1992:PSN

[Cha92] P. M. Chapman. Pollution status of North Sea sediments — an international integrative study. Marine Ecology Progress Series, 91:313–322, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/91/m091p313.pdf.

Chapman:1995:ALS

[Cha95] M. G. Chapman. Aggregation of the littorinid snail Littorina unifasciata in New South Wales, Australia. Marine Ecology Progress Series, 126:191–202, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/126/m126p191.pdf.

Charpy:1996:PBP

[Cha96] L. Charpy. Phytoplankton biomass and production in two Tuamotu atoll lagoons (French Polynesia). Marine Ecology Progress Series, 145:133–142, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/145/m145p133.pdf.

Chapman:1998:RBS

[Cha98] M. G. Chapman. Relationships between spatial patterns of benthic assemblages in a mangrove forest using different levels of taxonomic resolution. Marine Ecology Progress Series, 162:71–78, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/162/m162p071.pdf.

Cote:1994:SEL

[CHC94] J. Coté, J. H. Himmelman, and M. R. Claereboudt. Separating effects of limited food and space on growth of the giant scallop Placopecten magellanicus in suspended culture. Marine Ecology Progress Series, 106:85–91, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
106/m106p085.pdf.

Coyle:1992:MFE

[CHDW92] K. O. Coyle, G. L. Hunt, Jr., M. B. Decker, and T. J. Weingartner. Murre foraging, epibenthic sound scattering and tidal advection over a shoal near St. George Island, Bering Sea. Marine Ecology Progress Series, 83:1–14, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
83/m083p001.pdf.

Chesney:1993:MSG

[Che93] E. J. Chesney. A model of survival and growth of striped pass larvae Morone saxatilis in the Potomac River, 1987. Marine Ecology Progress Series, 92:15–25, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/92/m092p015.pdf.

Cronin:1995:DDS

[CHFL95] G. Cronin, M. E. Hay, W. Fenical, and N. Lindquist. Distribution, density, and sequestration of host chemical defenses by the specialist nudibranch Tritonia hamnerorum found at high densities on the sea fan Gorgonia ventalina. Marine Ecology Progress Series, 119:177–189, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/119/m119p177.pdf.

Carsen:1995:ESS

[CHS^{+}95] A. E. Carsen, B. G. Hatcher, R. E. Scheibling, A. W. Hennigar, and L. H. Taylor. Effects of site and season on movement frequencies and displacement patterns of juvenile sea scallops Placopecten magellanicus under natural hydrodynamic conditions in Nova Scotia, Canada. Marine Ecology Progress Series, 128:225–238, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/128/m128p225.pdf.

Chevaldonne:1993:VSD

[CJ93a] P. Chevaldonné and D. Jollivet. Videoscopic study of deep-sea hydrothermal vent alvinellid polychaete populations: biomass estimation and behaviour. Marine Ecology Progress Series, 95:251–262, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/95/m095p251.pdf.

Coulon:1993:FRS

[CJ93b] P. Coulon and M. Jangoux. Feeding rate and sediment reworking by the holothuroid Holothuria tubulosa (Echinodermata) in a Mediterranean seagrass bed off Ischia Island, Italy. Marine Ecology Progress Series, 92:201–204, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
92/m092p201.pdf.

Conley:1995:BPS

[CJ95] D. J. Conley and R. W. Johnstone. Biogeochemistry of N, P and Si in Baltic Sea sediments: response to a simulated deposition of a spring diatom bloom. Marine Ecology Progress Series, 122:265–276, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/122/m122p265.pdf.

Chisholm:1997:PMC

[CJ97] J. R. M. Chisholm and J. M. Jaubert. Photoautotrophic metabolism of Caulerpa taxifolia (Chlorophyta) in the NW Mediterranean. Marine Ecology Progress Series, 153:113– 123, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/153/m153p113.pdf.

Chalermwat:1991:ABD

[CJL91] K. Chalermwat, T. R. Jacobsen, and R. A. Lutz. Assimilation of bacteria by the dwarf surf clam Mulinia lateralis (Bivalvia: Mactridae). Marine Ecology Progress Series, 71:27–35, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
71/m071p027.pdf.

Caffrey:1990:NCS

[CK90] J. M. Caffrey and W. M. Kemp. Nitrogen cycling in sediments with estuarine populations of Potamogeton perfoliatus and Zostera marina. Marine Ecology Progress Series, 66: 147–160, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/66/m066p147.pdf.

Christian:1995:MBE

[CK95] J. R. Christian and D. M. Karl. Measuring bacterial ectoenzyme activities in marine waters using mercuric chloride as a preservative and a control. Marine Ecology Progress Series, 123:217–224, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/123/m123p217.pdf.

Cole:1998:IEP

[CK98] R. G. Cole and D. Keuskamp. Indirect effects of protection from exploitation: patterns from populations of Evechinus chloroticus (Echinoidea) in northeastern New Zealand. Marine Ecology Progress Series, 173:215–226, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
173/m173p215.pdf.

Chapman:1999:GCR

[CK99] M. R. Chapman and D. L. Kramer. Gradients in coral reef fish density and size across the Barbados marine reserve boundary: effects of reserve protection and habitat characteristics. Marine Ecology Progress Series, 181:81–96, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
181/m181p081.pdf.

Cochran:1998:PII

[CKP98] P. K. Cochran, C. A. Kellogg, and J. H. Paul. Prophage induction of indigenous marine lysogenic bacteria by environmental pollutants. Marine Ecology Progress Series, 164: 125–133, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/164/m164p125.pdf.

Cain:1990:ISG

[CL90] D. J. Cain and S. N. Luoma. Influence of seasonal growth, age, and environmental exposure on Cu and Ag in a bivalve indicator, Macoma balthica, in San Francisco Bay. Marine Ecology Progress Series, 60:45–55, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/60/m060p045.pdf.

Chaloupka:1997:RSM

[CL97] M. Y. Chaloupka and C. J. Limpus. Robust statistical modelling of hawksbill sea turtle growth rates (southern Great Barrier Reef). Marine Ecology Progress Series, 146:1–8, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
146/m146p001.pdf.

Clare:1990:LIS

[Cla90] A. S. Clare. Laboratory-induced spawning of the gastropod Gibbula cineraria as an indicator of field spawning. Marine Ecology Progress Series, 63:303–304, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
63/m063p303.pdf.

Clarke:1996:PST

[Cla96] R. D. Clarke. Population shifts in two competing fish species on a degrading coral reef. Marine Ecology Progress Series, 137:51–58, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/137/m137p051.pdf.

Chin-Leo:1991:DBA

[CLB91] G. Chin-Leo and R. Benner. Dynamics of bacterioplankton abundance and production in seagrass communities of a hypersaline lagoon. Marine Ecology Progress Series, 73: 219–230, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/73/m073p219.pdf.

Chin-Leo:1992:EBP

[CLB92] G. Chin-Leo and R. Benner. Enhanced bacterioplankton production and respiration at intermediate salinities in the Mississippi River plume. Marine Ecology Progress Series, 87: 87–103, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/87/m087p087.pdf.

Caraco:1998:MAD

[CLC^{+}98] N. F. Caraco, G. Lampman, J. J. Cole, K. E. Limburg, M. L. Pace, and D. Fischer. Microbial assimilation of DIN in a nitrogen rich estuary: implications for food quality and isotope studies. Marine Ecology Progress Series, 167:59–71, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
167/m167p059.pdf.

Condie:1999:MRT

[CLD99] S. A. Condie, N. R. Loneragan, and D. J. Die. Modelling the recruitment of tiger prawns Penaeus esculentus and P. semisulcatus to nursery grounds in the Gulf of Carpentaria, northern Australia: implications for assessing stockrecruitment relationships. Marine Ecology Progress Series, 178:55–68, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/178/m178p055.pdf.

Clemmesen:1993:IFD

[Cle93] C. Clemmesen. Improvements in the fluorimetric determination of the RNA and DNA content of individual marine fish larvae. Marine Ecology Progress Series, 100:177–183, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
100/m100p177.pdf.

Cantin:1996:RZM

[CLGM96] G. Cantin, M. Levasseur, M. Gosselin, and S. Michaud. Role of zooplankton in the mesoscale distribution of surface dimethylsulfide concentrations in the Gulf of St. Lawrence, Canada. Marine Ecology Progress Series, 141:103–117, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
141/m141p103.pdf.

Cronin:1995:ESE

[CLHF95] G. Cronin, N. Lindquist, M. E. Hay, and W. Fenical. Effects of storage and extraction procedures on yields of lipophilic metabolites from the brown seaweeds Dictyota ciliolata and D. menstrualis. Marine Ecology Progress Series, 119:265– 273, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/119/m119p265.pdf.

Clifton:1995:AFA

[Cli95] K. E. Clifton. Asynchronous food availability on neighboring Caribbean coral reefs determines seasonal patterns of growth and reproduction for the herbivorous parrotfish Scarus iserti. Marine Ecology Progress Series, 116:39–46, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
116/m116p039.pdf.

Culik:1998:HPM

[CLJOC98] B. M. Culik, G. Luna-Jorquera, H. Oyarzo, and H. Correa. Humboldt penguins monitored via VHF telemetry. Marine Ecology Progress Series, 162:279–286, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
162/m162p279.pdf.

Chin-Leo:1990:UGN

[CLK90] G. Chin-Leo and D. L. Kirchman. Unbalanced growth in natural assemblages of marine bacterioplankton. Marine Ecology Progress Series, 63:1–8, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/63/m063p001.pdf.

Conlan:1998:ISD

[CLKO98] K. E. Conlan, H. S. Lenihan, R. G. Kvitek, and J. S. Oliver. Ice scour disturbance to benthic communities in the Canadian high Arctic. Marine Ecology Progress Series, 166:1–16, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
166/m166p001.pdf.

Caron:1991:GUC

[CLM^{+}91] D. A. Caron, E. L. Lim, G. Miceli, J. B. Waterbury, and F. W. Valois. Grazing and utilization of chroococcoid cyanobacteria and heterotrophic bacteria by protozoa in laboratory cultures and a coastal plankton community. Marine Ecology Progress Series, 76:205–217, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/76/m076p205.pdf.

Cheng:1993:BAS

[CLM^{+}93] I.-J. Cheng, J. S. Levinton, M. McCartney, D. Martinez, and M. J. Weissburg. A bioassay approach to seasonal variation in the nutritional value of sediment. Marine Ecology Progress Series, 94:275–285, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/94/m094p275.pdf.

Costello:1999:IPD

[CLW99] J. H. Costello, R. Loftus, and R. Waggett. Influence of prey detection on capture success for the ctenophore Mnemiopsis leidyi feeding upon adult Acartia tonsa and Oithona colcarva copepods. Marine Ecology Progress Series, 191:207– 216, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/191/m191p207.pdf.

Carey:1990:NUD

[CM90a] D. A. Carey and L. M. Mayer. Nutrient uptake by a depositfeeding enteropneust: nitrogenous sources. Marine Ecology Progress Series, 63:79–84, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/63/m063p079.pdf.

Ceccherelli:1990:PMH

[CM90b] V. U. Ceccherelli and M. Mistri. Production of the meiobenthic harpacticoid copepod Canuella perplexa. Marine Ecology Progress Series, 68:225–234, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/68/m068p225.pdf.

Cockcroft:1993:NBH

[CM93] A. C. Cockcroft and A. McLachlan. Nitrogen budget for a high-energy ecosystem. Marine Ecology Progress Series, 100:287–299, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/100/m100p287.pdf.

Courtney:1997:SSD

[CM97] A. J. Courtney and J. M. Masel. Spawning stock dynamics of two penaeid prawns, Metapenaeus bennettae and Penaeus esculentus, in Moreton Bay, Queensland, Australia. Marine Ecology Progress Series, 148:37–47, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/148/m148p037.pdf.

Cartes:1998:FCB

[CM98a] J. E. Cartes and F. Maynou. Food consumption by bathyal decapod crustacean assemblages in the western Mediterranean: predatory impact of megafauna and the food consumptionfood supply balance in a deep-water food web. Marine Ecology Progress Series, 171:233–246, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/171/m171p233.pdf.

Christy:1998:EIC

[CM98b] J. H. Christy and S. G. Morgan. Estuarine immigration by crab postlarvae: mechanisms, reliability and adaptive significance. Marine Ecology Progress Series, 174:51–65, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
174/m174p051.pdf.

Chen:1999:SMB

[CM99] Z. Chen and L. M. Mayer. Sedimentary metal bioavailability determined by the digestive constraints of marine deposit feeders: gut retention time and dissolved amino acids. Marine Ecology Progress Series, 176:139–151, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
176/m176p139.pdf.

Christou:1995:MFM

[CMA95] E. D. Christou and M. Moraitou-Apostolopoulou. Metabolism and feeding of mesozooplankton in the eastern Mediterranean (Hellenic coastal waters). Marine Ecology Progress Series, 126:39–48, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/126/m126p039.pdf.

Carpenter:1999:EBF

[CMB^{+}99] E. J. Carpenter, J. P. Montoya, J. Burns, M. R. Mulholland, A. Subramaniam, and D. G. Capone. Extensive bloom of a N_{2}-fixing diatom/cyanobacterial association in the tropical Atlantic Ocean. Marine Ecology Progress Series, 185: 273–283, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/185/m185p273.pdf.

Chesney:1998:COJ

[CMBC98] E. J. Chesney, B. M. McKee, T. Blanchard, and L. H. Chan. Chemistry of otoliths from juvenile menhaden Brevoortia patronus: evaluating strontium, strontium: calcium and strontium isotope ratios as environmental indicators. Marine Ecology Progress Series, 171:261–273, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/171/m171p261.pdf.

Conover:1999:SUAb

[CMBM99] R. J. Conover, N. Mumm, P. Bruecker, and S. MacKenzie. Sources of urea in Arctic seas: seasonal fast ice? Marine Ecology Progress Series, 179:55–69, 1999. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/179/m179p055.pdf.

Cattrijsse:1994:NCI

[CMD^{+}94] A. Cattrijsse, E. S. Makwaia, H. R. Dankwa, O. Hamerlynck, and M. A. Hemminga. Nekton communities of an intertidal creek of a European estuarine brackish marsh. Marine Ecology Progress Series, 109:195–208, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/109/m109p195.pdf.

Chipman:1992:GTD

[CMLE92] J. K. Chipman, J. W. Marsh, D. R. Livingstone, and B. Evans. Genetic toxicity in dab Limanda limanda from the North Sea. Marine Ecology Progress Series, 91:121–126, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
91/m091p121.pdf.

Caceres-Martinez:1993:SMM

[CMRF93] J. Caceres-Martinez, J. A. F. Robledo, and A. Figueras. Settlement of mussels Mytilus galloprovincialis on an exposed rocky shore in Ria de Vigo, NW Spain. Marine Ecology Progress Series, 93:195–198, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/93/m093p195.pdf.

Caceres-Martinez:1994:SPL

[CMRF94] J. Caceres-Martinez, J. A. F. Robledo, and A. Figueras. Settlement and post-larvae behaviour of Mytitus galtoprovincialis: field and laboratory experiments. Marine Ecology Progress Series, 112:107–117, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/112/m112p107.pdf.

Conway:1994:DCE

[CMT94] D. V. P. Conway, I. R. B. McFadzen, and P. R. G. Tranter. Digestion of copepod eggs by larval turbot Scophthalmus maximus and egg viability following gut passage. Marine Ecology Progress Series, 106:303–309, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/106/m106p303.pdf.

Carlotti:1992:MCG

[CN92a] F. Carlotti and P. Nival. Model of copepod growth and development: moulting and mortality in relation to physiological processes during an individual moult cycle. Marine Ecology Progress Series, 84:219–233, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/84/m084p219.pdf.

Carlotti:1992:MMR

[CN92b] F. Carlotti and S. Nival. Moulting and mortality rates of copepods related to age within stage: experimental results. Marine Ecology Progress Series, 84:235–243, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
84/m084p235.pdf.

Chae:1995:VDD

[CN95] J. Chae and S. Nishida. Vertical distribution and diel migration in the iridescent copepods of the family Sapphirinidae: a unique example of reverse migration? Marine Ecology Progress Series, 119:111–124, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/119/m119p111.pdf.

Chen:1993:EAA

[CNCS93] Jiann-Chu Chen, Fan-Hua Nan, Sha-Yen Cheng, and Shyn-Shin Sheen. Effects of ambient ammonia on ammonia-n and protein concentrations in hemolymph and ammonia-n excretion of Penaeus chinensis. Marine Ecology Progress Series, 98:203–208, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/98/m098p203.pdf.

Carballo:1996:UMS

[CNGG96] J. L. Carballo, S. A. Naranjo, and J. C. García-Gómez. Use of marine sponges as stress indicators in marine ecosystems at Algeciras Bay (southern Iberian Peninsula). Marine Ecology Progress Series, 135:109–122, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/135/m135p109.pdf.

Creed:1996:NHH

[CNJ96] Joel C. Creed, T. A. Norton, and Joanna M. Kain (Jones). Are neighbours harmful or helpful in Fucus vesiculosus populations? Marine Ecology Progress Series, 133:191–201, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
133/m133p191.pdf.

Currin:1995:RSD

[CNP95] C. A. Currin, S. Y. Newell, and H. W. Paerl. The role of standing dead Spartina alterniflora and benthic microalgae in salt marsh food webs: considerations based on multiple stable isotope analysis. Marine Ecology Progress Series, 121:99–116, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/121/m121p099.pdf.

Camus:1992:SDV

[CO92] P. A. Camus and F. P. Ojeda. Scale-dependent variability of density estimates and morphometric relationships in subtidal stands of the kelp Lessonia trabeculata in northern and central Chile. Marine Ecology Progress Series, 90:193–200, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
90/m090p193.pdf.

Cockcroft:1990:NES

[Coc90] A. C. Cockcroft. Nitrogen excretion by the surf zone bivalves Donax serra and D. sordidus. Marine Ecology Progress Series, 60:57–65, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/60/m060p057.pdf.

Carpenter:1993:TDP

[COD^{+}93] E. J. Carpenter, J. M. O’Neil, R. Dawson, D. G. Capone, P. J. A. Siddiqui, T. Roenneberg, and B. Bergman. The tropical diazotrophic phytoplankter Trichodesmium: biological characteristics of two common species. Marine Ecology Progress Series, 95:295–304, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/95/m095p295.pdf.

Clough:1997:ELA

[COG97] B. F. Clough, J. E. Ong, and W. K. Gong. Estimating leaf area index and photosynthetic production in canopies of the mangrove Rhizophora apiculata. Marine Ecology Progress Series, 159:285–292, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/159/m159p285.pdf.

Chaloupka:1999:ESA

[COK99] M. Chaloupka, M. Osmond, and G. Kaufman. Estimating seasonal abundance trends and survival probabilities of humpback whales in Hervey Bay (east coast Australia). Marine Ecology Progress Series, 184:291–301, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
184/m184p291.pdf.

Collos:1992:NBD

[Col92] Y. Collos. Nitrogen budgets and dissolved organic matter cycling. Marine Ecology Progress Series, 90:201–206, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
90/m090p201.pdf.

Collos:1998:NUN

[Col98] Y. Collos. Nitrate uptake, nitrite release and uptake, and new production estimates. Marine Ecology Progress Series, 171:293–301, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/171/m171p293.pdf.

Connolly:1994:CFC

[Con94] R. M. Connolly. Comparison of fish catches from a buoyant pop net and a beach seine net in a shallow seagrass habitat. Marine Ecology Progress Series, 109:305–309, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
109/m109p305.pdf.

Connolly:1995:ERS

[Con95] R. M. Connolly. Effects of removal of seagrass canopy on assemblages of small, motile invertebrates. Marine Ecology Progress Series, 118:129–137, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/118/m118p129.pdf.

Connell:1998:EPG

[Con98] S. D. Connell. Effects of predators on growth, mortality and abundance of a juvenile reef-fish: evidence from manipulations of predator and prey abundance. Marine Ecology Progress Series, 169:251–261, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/169/m169p251.pdf.

Campana:1993:RDL

[COS93] S. E. Campana, H. A. Oxenford, and J. N. Smith. Radiochemical determination of longevity in flyingfish Hirundichthys affinis using Th-228/Ra-228. Marine Ecology Progress Series, 100:211–219, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/100/m100p211.pdf.

Cowen:1991:VPL

[Cow91] R. K. Cowen. Variation in the planktonic larval duration of the temperate wrasse Semi-cossyphus pulcher. Marine Ecology Progress Series, 69:9–15, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/69/m069p009.pdf.

Chaparro:1990:NEF

[CP90] O. R. Chaparro and K. A. Paschke. Nurse egg feeding and energy balance in embryos of Crepidula dilatata (Gastropoda: Calyptraeidae) during intracapsular development. Marine Ecology Progress Series, 65:183–191, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
65/m065p183.pdf.

Crawford:1991:EAF

[CP91] D. W. Crawford and D. A. Purdie. Evidence for avoidance of flushing from an estuary by a planktonic, phototrophic ciliate. Marine Ecology Progress Series, 79:259–265, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
79/m079p259.pdf.

Chanas:1995:DCS

[CP95a] B. Chanas and J. R. Pawlik. Defenses of Caribbean sponges against predatory reef fish. II. Spicules, tissue toughness, and nutritional quality. Marine Ecology Progress Series, 127:195– 211, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/127/m127p195.pdf.

Crocker:1995:DAD

[CP95b] K. M. Crocker and U. Passow. Differential aggregation of diatoms. Marine Ecology Progress Series, 117:249–257, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
117/m117p249.pdf.

Currie:1996:ESD

[CP96] D. R. Currie and G. D. Parry. Effects of scallop dredging on a soft sediment community: a large-scale experimental study. Marine Ecology Progress Series, 134:131–150, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
134/m134p131.pdf.

Cowan:1996:SIP

[CPB96] J. L. W. Cowan, J. R. Pennock, and W. R. Boynton. Seasonal and interannual patterns of sediment-water nutrient and oxygen fluxes in Mobile Bay, Alabama (USA): regulating factors and ecological significance. Marine Ecology Progress Series, 141:229–245, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/141/m141p229.pdf.

Caparroy:1998:FBC

[CPC98] P. Caparroy, M. T. Pérez, and F. Carlotti. Feeding behaviour of Centropages typicus in calm and turbulent conditions. Marine Ecology Progress Series, 168:109–118, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
168/m168p109.pdf.

Cochlan:1991:EIN

[CPH91] W. P. Cochlan, N. M. Price, and P. J. Harrison. Effects of irradiance on nitrogen uptake by phytoplankton: comparison of frontal and stratified communities. Marine Ecology Progress Series, 69:103–116, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/69/m069p103.pdf.

Cota:1996:NPP

[CPH^{+}96] G. F. Cota, L. R. Pomeroy, W. G. Harrison, E. P. Jones, F. Peters, W. M. Sheldon, Jr., and T. R. Weingartner. Nutrients, primary production and microbial heterotrophy in the southeastern Chukchi Sea: Arctic summer nutrient depletion and heterotrophy. Marine Ecology Progress Series, 135:247–258, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/135/m135p247.pdf.

Chen:1997:STS

[CPK97] C. C. Chen, J. E. Petersen, and W. M. Kemp. Spatial and temporal scaling of periphyton growth on walls of estuarine mesocosms. Marine Ecology Progress Series, 155:1–15, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
155/m155p001.pdf.

Chaudron:1996:HSC

[CPL^{+}96] Y. Chaudron, S. A. Poulet, M. Laabir, A. Ianora, and A. Miralto. Is hatching success of copepod eggs diatom densitydependent? Marine Ecology Progress Series, 144:185–193, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/144/m144p185.pdf.

Capriulo:1991:HAC

[CPS91] G. M. Capriulo, M. J. Pedone, and E. B. Small. High apostome ciliate endoparasite infection rates found in the Bering Sea euphausiid Thysanoessa inermis. Marine Ecology Progress Series, 72:203–204, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/72/m072p203.pdf.

Cho:1996:SBU

[CPSA96] B. C. Cho, M. G. Park, J. H. Shim, and F. Azam. Significance of bacteria in urea dynamics in coastal surface waters. Marine Ecology Progress Series, 142:19–26, 1996. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/142/m142p019.pdf.

Christiansen:1990:VDP

[CPT90] B. Christiansen, O. Pfannkuche, and H. Thiel. Vertical distribution and population structure of the necrophagous amphipod Eurythenes gryllus in the West European Basin. Marine Ecology Progress Series, 66:35–45, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/66/m066p035.pdf.

Chrost:1994:SSE

[CR94] R. J. Chrost and B. Riemann. Storm-stimulated enzymatic decomposition of organic matter in benthic/pelagic coastal meso-cosms. Marine Ecology Progress Series, 108:185–192, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/108/m108p185.pdf.

Carpenter:1995:MPC

[CR95] E. J. Carpenter and T. Roenneberg. The marine planktonic cyanobacteria Trichodesmium spp.: photosynthetic rate measurements in the SW Atlantic Ocean. Marine Ecology Progress Series, 118:267–273, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/118/m118p267.pdf.

Clausen:1996:GFR

[CR96] I. Clausen and H. U. Riisgård. Growth, filtration and respiration in the mussel Mytilus edulis: no evidence for physiological regulation of the filter-pump to nutritional needs. Marine Ecology Progress Series, 141:37–45, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/141/m141p037.pdf.

Cheng:1998:SAM

[CR98] L. Cheng and V. Roussis. Sex attractant in the marine insect Trochopus plumbeus (Heteroptera: Veliidae): a preliminary report. Marine Ecology Progress Series, 170:283–286, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
170/m170p283.pdf.

Craig:1994:GBN

[Cra94] N. I. Craig. Growth of the bivalve Nucula annulata in nutrient-enriched environments. Marine Ecology Progress Series, 104:77–90, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/104/m104p077.pdf.

Charpy-Roubaud:1996:DNF

[CRCS96] C. Charpy-Roubaud, L. Charpy, and G. Sarazin. Diffusional nutrient fluxes at the sediment–water interface and organic matter mineralization in an atoll lagoon (Tikehau, Tuamotu Archipelago, French Polynesia). Marine Ecology Progress Series, 132:181–190, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/132/m132p181.pdf.

Coma:1998:EAS

[CRGZ98] R. Coma, M. Ribes, J. M. Gili, and M. Zabala. An energetic approach to the study of life-history traits of two modular colonial benthic invertebrates. Marine Ecology Progress Series, 162:89–103, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/162/m162p089.pdf.

Cowan:1999:MEI

[CRH^{+}99] J. H. Cowan, Jr., K. A. Rose, E. D. Houde, S. B. Wang, and J. Young. Modeling effects of increased larval mortality on bay anchovy population dynamics in the mesohaline Chesapeake Bay: evidence for compensatory reserve. Marine Ecology Progress Series, 185:133–146, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/185/m185p133.pdf.

Crawford:1994:RMA

[CRLP94] D. W. Crawford, A. Rogerson, and J. Laybourn-Parry. Respiration of the marine amoeba Trichosphaerium sieboldi determined by ^{14}C labelling and Cartesian diver methods. Marine Ecology Progress Series, 112:135–142, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
112/m112p135.pdf.

Crooks:1998:HAC

[Cro98] J. A. Crooks. Habitat alteration and community-level effects of an exotic mussel, Musculista senhousia. Marine Ecology Progress Series, 162:137–152, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/162/m162p137.pdf.

Crook:1999:QEA

[Cro99] A. C. Crook. Quantitative evidence for assortative schooling in a coral reef fish. Marine Ecology Progress Series, 176: 17–23, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/176/m176p017.pdf.

Croxall:1999:DPP

[CRP99] J. P. Croxall, K. Reid, and P. A. Prince. Diet, provisioning and productivity responses of marine predators to differences in availability of Antarctic krill. Marine Ecology Progress Series, 177:115–131, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/177/m177p115.pdf.

Coma:1995:RCG

[CRZG95] R. Coma, M. Ribes, M. Zabala, and J.-M. Giti. Reproduction and cycle of gonadal development in the Mediterranean gorgonian Paramuricea clavata. Marine Ecology Progress Series, 117:173–183, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/117/m117p173.pdf.

Cottrell:1991:WSO

[CS91] M. T. Cottrell and C. A. Suttle. Wide-spread occurrence and clonal variation in viruses which cause lysis of a cosmopolitan, eukaryotic marine phytoplankter Micromonas pusilla. Marine Ecology Progress Series, 78:1–9, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/78/m078p001.pdf.

Cartes:1993:ZDS

[CS93] J. E. Cartes and F. Sarda. Zonation of deep-sea decapod fauna in the Catalan Sea (Western Mediterranean). Marine Ecology Progress Series, 94:27–34, 1993. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/94/m094p027.pdf.

Company:1997:RPP

[CS97] J. B. Company and F. Sardà. Reproductive patterns and population characteristics in five deep-water pandalid shrimps in the Western Mediterranean along a depth gradient (150– 1100 m). Marine Ecology Progress Series, 148:49–58, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
148/m148p049.pdf.

Cieri:1999:RGA

[CS99] M. D. Cieri and D. E. Stearns. Reduction of grazing activity of two estuarine copepods in response to the exudate of a visual predator. Marine Ecology Progress Series, 177:157–163, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
177/m177p157.pdf.

Cochrane:1994:TVE

[CSB94] N. A. Cochrane, D. D. Sameoto, and D. J. Belliveau. Temporal variability of euphausiid concentrations in a Nova Scotia shelf basin using a bottom-mounted acoustic Doppler current profiler. Marine Ecology Progress Series, 107:55–66, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
107/m107p055.pdf.

Cunha:1997:SNS

[CSB97] M. R. Cunha, J. C. Sorbe, and C. Bernardes. On the structure of the neritic suprabenthic communities from the Portuguese continental margin. Marine Ecology Progress Series, 157:119–137, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/157/m157p119.pdf.

Carrasson:1992:DBD

[CSC92] M. Carrasson, C. Stefanescu, and J. E. Carles. Diets and bathymetric distributions of two bathyal sharks of the Catalan deep sea (western Mediterranean). Marine Ecology Progress Series, 82:21–30, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/82/m082p021.pdf.

Culverhouse:1996:ACF

[CSE^{+}96] P. F. Culverhouse, R. G. Simpson, R. Ellis, J. A. Lindley, R. Williams, T. Parisini, B. Reguera, I. Bravo, R. Zoppoli, G. Earnshaw, H. McCall, and G. Smith. Automatic classification of field-collected dinoflagellates by artificial neural network. Marine Ecology Progress Series, 139:281–287, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
139/m139p281.pdf.

Carlsson:1993:NBH

[CSG93] P. Carlsson, A. Z. Segatto, and E. Graneli. Nitrogen bound to humic matter of terrestrial origin a nitrogen pool for coastal phytoplankton? Marine Ecology Progress Series, 97:105– 116, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/97/m097p105.pdf.

Capone:1998:EBF

[CSM^{+}98] D. G. Capone, A. Subramaniam, J. P. Montoya, M. Voss, C. Humborg, A. M. Johansen, R. L. Siefert, and E. J. Carpenter. An extensive bloom of the N_{2}-fixing cyanobacterium Trichodesmium erythraeum in the central Arabian Sea. Marine Ecology Progress Series, 172:281–292, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
172/m172p281.pdf.

Cunha:1999:SSC

[CSM99] M. R. Cunha, J. C. Sorbe, and M. H. Moreira. Spatial and seasonal changes of brackish peracaridan assemblages and their relation to some environmental variables in two tidal channels of the Ria de Aveiro (NW Portugal). Marine Ecology Progress Series, 190:69–87, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/190/m190p069.pdf.

Chang:1992:ABM

[CSR^{+}92a] S. Chang, F. W. Steimle, R. N. Reid, S. A. Fromm, V. S. Zdanowicz, and R. A. Pikanowski. Association of benthic macrofauna with habitat types and quality in the New York Bight. Marine Ecology Progress Series, 89:237–251, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
89/m089p237.pdf.

Chapman:1992:ICS

[CSR^{+}92b] P. M. Chapman, R. C. Swartz, B. Roddie, H. L. Phelps, P. van den Hurk, and R. Butler. An international comparison of sediment toxicity tests in the North Sea. Marine Ecology Progress Series, 91:253–264, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/91/m091p253.pdf.

Conley:1993:MBC

[CSS93] D. J. Conley, C. L. Schelske, and E. F. Stoermer. Modification of the biogeochemical cycle of silica with eutrophication. Marine Ecology Progress Series, 101:179–192, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
101/m101p179.pdf.

Cosson-Sarradin:1998:PDT

[CSSPV98] N. Cosson-Sarradin, M. Sibuet, G. L. J. Paterson, and A. Vangriesheim. Polychaete diversity at tropical Atlantic deep-sea sites: environmental effects. Marine Ecology Progress Series, 165:173–185, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/165/m165p173.pdf.

Cahoon:1992:QED

[CT92] L. B. Cahoon and C. R. Tronzo. Quantitative estimates of demersal zooplankton abundance in Onslow Bay, North Carolina, USA. Marine Ecology Progress Series, 87:197– 200, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/87/m087p197.pdf.

Cheng:1996:TME

[CT96] P. W. Cheng and W. N. Tzeng. Timing of metamorphosis and estuarine arrival across the dispersal range of the Japanese eel Anguilla japonica. Marine Ecology Progress Series, 131:87–96, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/131/m131p087.pdf.

Cheal:1997:CVC

[CT97] A. J. Cheal and A. A. Thompson. Comparing visual counts of coral reef fish: implications of transect width and species selection. Marine Ecology Progress Series, 158:241–248, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
158/m158p241.pdf.

Chaparro:1998:PEB

[CT98] O. R. Chaparro and R. J. Thompson. Physiological energetics of brooding in Chilean oyster Ostrea chilensis. Marine Ecology Progress Series, 171:151–163, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/171/m171p151.pdf.

Conway:1993:DNF

[CTC93] D. V. P. Conway, P. R. G. Tranter, and S. H. Coombs. Digestion of natural food by larval and post-larval turbot Scophthalmus maximus. Marine Ecology Progress Series, 100:221– 231, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/100/m100p221.pdf.

Carman:1991:NRP

[CTEF91] Kevin R. Carman, David Thistle, Stephen C. Ertman, and Michael Foy. Nile red as a probe for lipid-storage products in benthic copepods. Marine Ecology Progress Series, 74: 307–311, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/74/m074p307.pdf.

Carlton:1990:RIS

[CTSN90] J. T. Carlton, J. K. Tompson, L. E. Schemel, and F. H. Nichols. Remarkable invasion of San Francisco Bay (California, USA), by the Asian clam Potamocorbula amurensis. I. Introduction and dispersal. Marine Ecology Progress Series, 66:81–94, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/66/m066p081.pdf.

Chapman:1999:EPM

[CU99] M. G. Chapman and A. J. Underwood. Ecological patterns in multivariate assemblages: information and interpretation of negative values in ANOSIM tests. Marine Ecology Progress Series, 180:257–265, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/180/m180p257.pdf.

Calvo-Ugarteburu:1998:PIS

[CUM98] G. Calvo-Ugarteburu and C. D. McQuaid. Parasitism and invasive species: effects of digenetic trematodes on mussels. Marine Ecology Progress Series, 169:149–163, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
169/m169p149.pdf.

Cushing:1995:CLD

[Cus95] D. H. Cushing. A comment on Leggett & DeBlois. Marine Ecology Progress Series, 128:305–306, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Chrost:1991:MEK

[CV91] R. J. Chróst and B. Velimirov. Measurement of enzyme kinetics in water samples: effect of freezing and soluble stabilizer. Marine Ecology Progress Series, 70:93–100, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
70/m070p093.pdf.

Cardoso:1996:PBS

[CV96] R. S. Cardoso and V. G. Veloso. Population biology and secondary production of the sandhopper Pseudorchestoidea brasiliensis (Amphipoda: Talitridae) at Prainha Beach, Brazil. Marine Ecology Progress Series, 142:111–119, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
142/m142p111.pdf.

Castilla:1998:PTE

[CV98] J. C. Castilla and M. A. Varas. A plankton trap for exposed rocky intertidal shores. Marine Ecology Progress Series, 175:299–305, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/175/m175p299.pdf.

Christaki:1999:NMH

[CVD99] U. Christaki, F. Van Wambeke, and J. R. Dolan. Nanoflagellates (mixotrophs, heterotrophs and autotrophs) in the oligotrophic eastern Mediterranean: standing stocks, bacterivory and relationships with bacterial production. Marine Ecology Progress Series, 181:297–307, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/181/m181p297.pdf.

Cerda:1993:FEC

[CW93] G. Cerda and M. Wolff. Feeding ecology of the crab Cancer polyodon in La Herradura Bay, northern Chile. II. Food spectrum and prey consumption. Marine Ecology Progress Series, 100:119–125, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/100/m100p119.pdf.

Cherel:1995:SIM

[CW95] Y. Cherel and H. Weimerskirch. Seabirds as indicators of marine resources: black-browed albatrosses feeding on ommastrephid squids in Kerguelen waters. Marine Ecology Progress Series, 129:295–300, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/129/m129p295.pdf.

Cherel:1999:SCO

[CW99a] Y. Cherel and H. Weimerskirch. Spawning cycle of onychoteuthid squids in the southern Indian Ocean: new information from seabird predators. Marine Ecology Progress Series, 188:93–104, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/188/m188p093.pdf.

Clarke:1999:TDM

[CW99b] K. R. Clarke and R. M. Warwick. The taxonomic distinctness measure of biodiversity: weighting of step lengths between hierarchical levels. Marine Ecology Progress Series, 184: 21–29, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/184/m184p021.pdf.

Clarke:1993:ISB

[CWB93a] K. R. Clarke, R. M. Warwick, and B. E. Brown. An index showing breakdown of seriation, related to disturbance, in a coral-reef assemblage. Marine Ecology Progress Series, 102:153–160, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/102/m102p153.pdf.

Culik:1993:FBP

[CWB93b] B. M. Culik, R. P. Wilson, and R. Bannasch. Flipperbands on penguins: what is the cost of a life-long commitment? Marine Ecology Progress Series, 98:209–214, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
98/m098p209.pdf.

Coyle:1998:DAD

[CWH98] K. O. Coyle, T. J. Weingartner, and G. L. Hunt, Jr. Distribution of acoustically determined biomass and major zooplankton taxa in the upper mixed layer relative to water masses in the western Aleutian Islands. Marine Ecology Progress Series, 165:95–108, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/165/m165p095.pdf.

Cripps:1999:FAC

[CWHA99] G. C. Cripps, J. L. Watkins, H. J. Hill, and A. Atkinson. Fatty acid content of Antarctic krill Euphausia superba at South Georgia related to regional populations and variations in diet. Marine Ecology Progress Series, 181:177–188, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
181/m181p177.pdf.

Cochlan:1993:SDV

[CWS^{+}93] W. P. Cochlan, J. Wikner, G. R. Steward, D. C. Smith, and F. Azam. Spatial distribution of viruses, bacteria and chlorophyll-a in nentic, oceanic and estuarme environments. Marine Ecology Progress Series, 92:77–87, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
92/m092p077.pdf.

Clark:1999:IIA

[CWWH99] M. E. Clark, T. G. Wolcott, D. L. Wolcott, and A. H. Hines. Intraspecific interference among foraging blue crabs Callinectes sapidus: interactive effects of predator density and prey patch distribution. Marine Ecology Progress Series, 178:69–78, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/178/m178p069.pdf.

Coma:1995:SRE

[CZG95] R. Coma, M. Zabala, and J.-M. Gili. Sexual reproductive effort in the Mediterranean gorgonian Paramuricea ciavata. Marine Ecology Progress Series, 117:185–192, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
117/m117p185.pdf.

Dilling:1993:CCF

[DA93] L. Dilling and A. L. Alldredge. Can chaetognath fecal pellets contribute significantly to carbon flux? Marine Ecology Progress Series, 92:51–58, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/92/m092p051.pdf.

Dando:1991:ENS

[DAB^{+}91] P. R. Dando, M. C. Austen, R. A. Burke, Jr., M. A. Kendall, M. C. Kennicutt II, A. G. Judd, D. C. Moore, S. C. M. O’Hara, R. Schmaljohann, and A. J. Southward. Ecology of a North Sea pockmark with an active methane seep. Marine Ecology Progress Series, 70:49–63, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/70/m070p049.pdf.

Dubilier:1999:PDB

[DAE^{+}99] N. Dubilier, R. Amann, C. Erséus, G. Muyzer, S. Y. Park, O. Giere, and C. M. Cavanaugh. Phylogenetic diversity of bacterial endosymbionts in the gutless marine oligochete Olavius loisae (Annelida). Marine Ecology Progress Series, 178:271–280, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/178/m178p271.pdf.

Dai:1990:ICT

[Dai90] C.-F. Dai. Interspecific competition in Taiwanese corals with special reference to interactions between alcyonaceans and scleractinians. Marine Ecology Progress Series, 60:291–297, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/60/m060p291.pdf.

Dieckmann:1992:HRS

[DAS92] G. S. Dieckmann, K. Arrigo, and C. W. Sullivan. A highresolution sampler for nutrient and chlorophyll-a profiles of the sea ice platelet layer and underlying water column below fast ice in polar oceans: preliminary results. Marine Ecology Progress Series, 80:291–300, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/80/m080p291.pdf.

Doval:1997:DOM

[DÁSP97] M. D. Doval, X. A. Álvarez-Salgado, and F. F. Pérez. Dissolved organic matter in a temperate embayment affected by coastal upwelling. Marine Ecology Progress Series, 157: 21–37, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/157/m157p021.pdf.

Davis:1994:STC

[Dav94] M. Davis. Short-term competence in larvae of queen conch Strombus gigas: shifts in behavior, morphology and metamorphic response. Marine Ecology Progress Series, 104:101– 108, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/104/m104p101.pdf.

Davidson:1996:MMF

[Dav96] K. Davidson. Modelling microbial food webs. Marine Ecology Progress Series, 145:279–296, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/145/m145p279.pdf.

Dahl:1996:USL

[DB96a] B. Dahl and H. Blanck. Use of sand-living microalgal communities (epipsammon) in ecotoxicological testing. Marine Ecology Progress Series, 144:163–173, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/144/m144p163.pdf.

Dunne:1996:PSU

[DB96b] R. P. Dunne and B. E. Brown. Penetration of solar UVB radiation in shallow tropical waters and its potential biological effects on coral reefs; results from the central Indian Ocean and Andaman Sea. Marine Ecology Progress Series, 144:109–118, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/144/m144p109.pdf.

Davies:1999:RMT

[DB99] M. S. Davies and P. Beckwith. Role of mucus trails and trail-following in the behaviour and nutrition of the periwinkle Littorina littorea. Marine Ecology Progress Series, 179:247–257, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/179/m179p247.pdf.

Dauby:1995:PFM

[DBB^{+}95] P. Dauby, A. J. Bale, N. Bloomer, C. Canon, R. D. Ling, A. Norro, J. E. Robertson, A. Simon, J. M. Théate, A. J. Watson, and M. Frankignoulle. Particle fluxes over a Mediterranean seagrass bed: a one year case study. Marine Ecology Progress Series, 126:233–246, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/126/m126p233.pdf.

Duarte:1999:MPW

[DBdCS99] C. M. Duarte, E. Benavent, and M. d. C. Sánchez. The microcosm of particles within seagrass Posidonia oceanica canopies. Marine Ecology Progress Series, 181:289–295, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
181/m181p289.pdf.

Dahllof:1999:LTE

[DBHM99] I. Dahllöf, H. Blanck, P. O. J. Hall, and S. Molander. Longterm effects of tri-n-butyl-tin on the function of a marine sediment system. Marine Ecology Progress Series, 188:1–11, 1999.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
188/m188p001.pdf.

Dale:1999:CPA

[DBMK99] T. Dale, E. Bagøien, W. Melle, and S. Kaartvedt. Can predator avoidance explain varying overwintering depth of Calanus in different oceanic water masses? Marine Ecology Progress Series, 179:113–121, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/179/m179p113.pdf.

DOnghia:1999:RSM

[DBMM99] G. D’Onghia, M. Basanisi, A. Matarrese, and F. Megli. Reproductive strategies in macrourid fish: seasonality or not? Marine Ecology Progress Series, 184:189–196, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
184/m184p189.pdf.

deBaar:1990:ILS

[dBN^{+}90] H. J. W. de Baar, A. G. J. Buma, R. F. Nolting, G. C. Cadee, G. Jacques, and P. J. Treguer. On iron limitation of the Southern Ocean: experimental observations in the Weddell and Scotia Seas. Marine Ecology Progress Series, 65:105–122, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
65/m065p105.pdf.

Dando:1994:MSA

[DBN^{+}94] P. R. Dando, I. Bussmann, S. J. Niven, S. C. M. O’Hara, R. Schmaljohann, and L. J. Taylor. A methane seep area in the Skagerrak, the habitat of the pogonophore Siboglinum poseidoni and the bivalve mollusc Thyasira sarsi. Marine Ecology Progress Series, 107:157–167, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/107/m107p157.pdf.

DeWolf:1998:STG

[DBV98] H. De Wolf, T. Backeljau, and R. Verhagen. Spatio-temporal genetic structure and gene flow between two distinct shell morphs of the planktonic developing periwinkle Littorina striata (Mollusca: Prosobranchia). Marine Ecology Progress Series, 163:155–163, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/163/m163p155.pdf.

Davis:1991:SBA

[DBvA91] A. R. Davis, A. J. Butler, and I. van Altena. Settlement behaviour of ascidian larvae: preliminary evidence for inhibition by sponge allelochemicals. Marine Ecology Progress Series, 72:117–123, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/72/m072p117.pdf.

Durante:1991:EAF

[DC91] K. M. Durante and F. S. Chia. Epiphytism on Agarum fimbriatum: can herbivore preferences explain distribution of epiphytic bryozoans? Marine Ecology Progress Series, 77: 279–287, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/77/m077p279.pdf.

deJonge:1994:DMB

[dC94] V. N. de Jonge and F. Colijn. Dynamics of microphytobenthos biomass in the Ems estuary. Marine Ecology Progress Series, 104:185–196, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/104/m104p185.pdf.

Dodson:1991:RBS

[DCBC91] Julian J. Dodson, James E. Carscadden, Louis Bernatchez, and Françoise Colombani. Relationship between spawning mode and phylogeographic structure in mitochondrial DNA of North Atlantic capelin Mallotus villosus. Marine Ecology Progress Series, 76:103–113, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/76/m076p103.pdf.

Dufour:1999:PNC

[DCBG99] P. Dufour, L. Charpy, S. Bonnet, and N. Garcia. Phytoplankton nutrient control in the oligotrophic South Pacific subtropical gyre (Tuamotu Archipelago). Marine Ecology Progress Series, 179:285–290, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/179/m179p285.pdf.

Durand:1999:BMF

[DCG^{+}99] J. P. Durand, F. P. Carvalho, F. Goudard, J. Pieri, S. W. Fowler, and O. Cotret. 210Po binding to metallothioneins and ferritin in the liver of teleost marine fish. Marine Ecology Progress Series, 177:189–196, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/177/m177p189.pdf.

Davis:1999:CVF

[DCK99] J. L. Davis, D. L. Childers, and D. N. Kuhn. Clonal variation in a Florida Bay Thalassia testudinum meadow: molecular genetic assessment of population structure. Marine Ecology Progress Series, 186:127–136, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/186/m186p127.pdf.

Denton:1990:SSC

[DCM90] A. Denton, A. R. O. Chapman, and J. Markham. Sizespecific concentrations of phlorotannins (anti-herbivore compounds) in three species of Fucus. Marine Ecology Progress Series, 65:103–104, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/65/m065p103.pdf.

Deibel:1992:LLC

[DCRG92] D. Deibel, J. F. Cavaletto, M. Riehl, and W. S. Gardner. Lipid and lipid class content of the pelagic tunicate Oikopleura vanhoeffeni. Marine Ecology Progress Series, 88:297–302, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
88/m088p297.pdf.

Dauvin:1990:RBF

[DD90] J.-C Dauvin and J. J. Dodson. Relationship between feeding incidence and vertical and longitudinal distribution of rainbow smelt larvae (Osmerus mordax) in a turbid well-mixed estuary. Marine Ecology Progress Series, 60:1–12, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
60/m060p001.pdf.

Durako:1997:URA

[DD97] M. J. Durako and C. M. Duarte. On the use of reconstructive aging techniques for assessing seagrass demography: a critique of the model test of Jensen et al. (1996). Marine Ecology Progress Series, 146:297–303, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/146/m146p297.pdf.

Duplisea:1999:SBM

[DD99] D. E. Duplisea and A. Drgas. Sensitivity of a benthic, metazoan, biomass size spectrum to differences in sediment granulometry. Marine Ecology Progress Series, 177:73–81, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
177/m177p073.pdf.

Defeo:1995:EHA

[DdA95] O. Defeo and A. de Alava. Effects of human activities on long-term trends in sandy beach populations: the wedge clam Donax hanleyanus in Uruguay. Marine Ecology Progress Series, 123:73–82, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/123/m123p073.pdf.

Danovaro:1999:SVB

[DDM^{+}99] R. Danovaro, A. Dell’Anno, D. Martorano, P. Parodi, N. D. Marrale, and M. Fabiano. Seasonal variation in the biochemical composition of deep-sea nematodes: bioenergetic and methodological considerations. Marine Ecology Progress Series, 179:273–283, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/179/m179p273.pdf.

Durbin:1990:DFB

[DDW90] A. G. Durbin, E. G. Durbin, and E. Wlodarczyk. Diel feeding behavior in the marine copepod Acartia tonsa in relation to food availability. Marine Ecology Progress Series, 68: 23–45, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/68/m068p023.pdf.

Dittel:1990:STA

[DE90] A. I. Dittel and C. E. Epifanio. Seasonal and tidal abundance of crab larvae in a tropical mangrove system, Gulf of Nicoya, Costa Rica. Marine Ecology Progress Series, 65:25–34, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
65/m065p025.pdf.

Durkina:1991:CAC

[DE91] V. B. Durkina and Z. S. Evtushenko. Changes in activity of certain enzymes in sea urchin embryos and larvae after exposure of adult organisms to heavy metals. Marine Ecology Progress Series, 72:111–115, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/72/m072p111.pdf.

Duffy:1994:ELD

[DE94] Janet T. Duffy and Charles E. Epifanio. Effects of larval density on the growth and survival of weakfish Cynoscion regalis in large-volume enclosures. Marine Ecology Progress Series, 104:227–233, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/104/m104p227.pdf.

Deegan:1990:EEE

[Dee90] L. A. Deegan. Effects of estuarine environmental conditions on population dynamics of young-of-the-year gulf menhaden. Marine Ecology Progress Series, 68:195–205, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
68/m068p195.pdf.

Dearborn:1991:DFB

[DEF91] John H. Dearborn, Kelly C. Edwards, and David B. Fratt. Diet, feeding behavior, and surface morphology of the multiarmed Antarctic sea star Labidiaster annulatus (Echinodermata Asteroidea). Marine Ecology Progress Series, 77:65–84, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/77/m077p065.pdf.

Deibel:1990:SWS

[Dei90] D. Deibel. Still-water sinking velocity of fecal material from the pelagic tunicate Dolioletta gegenbauri. Marine Ecology Progress Series, 62:55–60, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/62/m062p055.pdf.

Demestre:1995:MAR

[Dem95] M. Demestre. Moult activity-related spawning success in the Mediterranean deep-water shrimp Aristeus antennatus (Decapoda: Dendrobranchiata). Marine Ecology Progress Series, 127:57–64, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/127/m127p057.pdf.

Doucette:1996:QAI

[DEP^{+}96] G. J. Doucette, D. L. Erdner, M. L. Peleato, J. J. Hartman, and D. M. Anderson. Quantitative analysis of ironstress related proteins in Thalassiosira weissflogii: measurement of flavodoxin and ferredoxin using HPLC. Marine Ecology Progress Series, 130:269–276, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/130/m130p269.pdf.

DeWreede:1992:WHS

[DES92] R. E. DeWreede, P. Ewanchuk, and F. Shaughnessy. Wounding, healing and survivorship in three kelp species. Marine Ecology Progress Series, 82:259–266, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
82/m082p259.pdf.

DeVantier:1992:RTM

[DeV92] L. M. DeVantier. Rafting of tropical marine organisms on buoyant corolla. Marine Ecology Progress Series, 86:301– 302, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/86/m086p301.pdf.

Demestre:1992:RDW

[DF92] M. Demestre and J.-M. Fortuno. Reproduction of the deep-water shrimp Aristeus antennatus (Decapoda: Dendrobranchiata). Marine Ecology Progress Series, 84:41–51, 1992.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
84/m084p041.pdf.

DeVogelaere:1994:DRI

[DF94] A. P. De Vogelaere and M. S. Foster. Damage and recovery in intertidal Fucus gardneri assemblages following the ‘Exxon Valdez’ oil spill. Marine Ecology Progress Series, 106:263–271, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/106/m106p263.pdf.

Duarte:1995:SAS

[DF95] P. Duarte and J. G. Ferreira. Seasonal adaptation and shortterm metabolic responses of Gelidium sesquipedale to varying light and temperature. Marine Ecology Progress Series, 121:289–300, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/121/m121p289.pdf.

Diniz-Filho:1996:ARB

[DFB96] J. A. F. Diniz-Filho and L. M. Bini. Assessing the relationship between multivariate community structure and environmental variables. Marine Ecology Progress Series, 143:303–306, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
143/m143p303.pdf.

DellAnno:1999:PDD

[DFB^{+}99] A. Dell’Anno, M. Fabiano, S. Bompadre, M. Armeni, L. Leone, and R. Danovaro. Phytopigment and DNA determinations in long-time formalin-preserved trap samples. Marine Ecology Progress Series, 191:71–77, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/191/m191p071.pdf.

Dando:1993:EGO

[DFJ^{+}93] P. R. Dando, T. Fenchel, P. Jensen, S. C. M. O’Hara, S. J. Niven, and U. Schuster. Ecology of gassy, organic-rich sediment in a shallow subtidal area on the Kattegat coast of Denmark. Marine Ecology Progress Series, 100:265–271, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
100/m100p265.pdf.

Drolet:1991:PFL

[DFPG91] R. Drolet, L. Fortier, D. Ponton, and M. Gilbert. Production of fish larvae and their prey in subarctic southeastern Hudson Bay. Marine Ecology Progress Series, 77:105–118, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
77/m077p105.pdf.

Delille:1995:SCP

[DFR95] D. Delille, M. Fiala, and C. Rosiers. Seasonal changes in phytoplankton and bacterioplankton distribution at the ice– water interface in the Antarctic neritic area. Marine Ecology Progress Series, 123:225–233, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/123/m123p225.pdf.

Dring:1994:ESB

[DFS94] M. J. Dring, R. M. Forster, and R. Schmid. Ecological significance of blue light stimulation of photosynthetic capacity in Laminaria spp. and other brown algae. Marine Ecology Progress Series, 113:271–277, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/113/m113p271.pdf.

daCunhaLana:1991:ISA

[dG91a] P. da Cunha Lana and C. Guiss. Influence of Spartina alterniflora on structure and temporal variability of macrobenthic associations in a tidal flat of Paranagua Bay (southeastern Brazil). Marine Ecology Progress Series, 73:231–244, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
73/m073p231.pdf.

Dolan:1991:TCR

[DG91b] J. R. Dolan and C. L. Gallegos. Trophic coupling of rotifers, microflagellates, and bacteria during fall months in the Rhode River Estuary. Marine Ecology Progress Series, 77: 147–156, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/77/m077p147.pdf.

daCunhaLana:1992:MPB

[dG92a] Paulo da Cunha Lana and Claudia Guiss. Macrofauna — plant-biomass interactions in a euhaline salt marsh in Paranagua Bay (SE Brazil). Marine Ecology Progress Series, 80:57–64, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/80/m080p057.pdf.

Dolan:1992:TRP

[DG92b] J. R. Dolan and C. C. Gallegos. Trophic role of planktonic rotifers in the Rhode River Estuary, spring–summer 1991. Marine Ecology Progress Series, 85:187–199, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
85/m085p187.pdf.

Dufour:1993:CPR

[DG93] V. Dufour and R. Galzin. Colonization patterns of reef fish larvae to the lagoon at Moorea Island, French Polynesia. Marine Ecology Progress Series, 102:143–152, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
102/m102p143.pdf.

Duro:1996:MSH

[DG96] A. Duró and J. M. Gili. Mesoscale spatial heterogeneity in chaetognath populations during upwelling abatement in the northern Benguela region. Marine Ecology Progress Series, 140:41–58, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/140/m140p041.pdf.

Deegan:1997:ESV

[DG97] L. A. Deegan and R. H. Garritt. Evidence for spatial variability in estuarine food webs. Marine Ecology Progress Series, 147:31–47, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/147/m147p031.pdf.

Dethier:1993:VVR

[DGCT93] M. N. Dethier, E. S. Graham, S. Cohen, and L. M. Tear. Visual versus random-point percent cover estimations: ’objective’ is not always better. Marine Ecology Progress Series, 96:93–100, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/96/m096p093.pdf.

Duarte:1998:RBS

[DGHT^{+}98] C. M. Duarte, O. Geertz-Hansen, U. Thampanya, J. Terrados, M. D. Fortes, L. Kamp-Nielsen, J. Borum, and S. Boromthanarath. Relationship between sediment conditions and mangrove Rhizophora apiculata seedling growth and nutrient status. Marine Ecology Progress Series, 175:277–283, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
175/m175p277.pdf.

Descolas-Gros:1992:VCA

[DGO92] C. Descolas-Gros and L. Oriol. Variations in carboxylase activity in marine phytoplankton cultures, β-carbolxylation in carbon flux studies. Marine Ecology Progress Series, 85: 163–169, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/85/m085p163.pdf.

Detmer:1993:PHP

[DGT^{+}93] A. E. Detmer, H. C. Giesenhagen, V. M. Trenkel, H. Auf dem Venne, and F. J. Jochem. Phototrophic and heterotrophic pico- and nanoplankton in anoxic depths of the central Baltic Sea. Marine Ecology Progress Series, 99:197–204, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
99/m099p197.pdf.

Doucette:1990:SEI

[DH90] G. J. Doucette and P. J. Harrison. Some effects of iron and nitrogen stress on the red tide dinoflagellate Gymnodinium sanguineum. Marine Ecology Progress Series, 62:293–306, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
62/m062p293.pdf.

Durako:1992:ELS

[DH92] M. J. Durako and M. O. Hall. Effects of light on the stable carbon isotope composition of the seagrass Thalassia testudinum. Marine Ecology Progress Series, 86:99–101, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
86/m086p099.pdf.

Decker:1996:FMU

[DH96] M. B. Decker and G. L. Hunt, Jr. Foraging by murres (Uria spp.) at tidal fronts surrounding the Pribilof Islands, Alaska, USA. Marine Ecology Progress Series, 139:1–10, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
139/m139p001.pdf.

Davies:1997:SET

[DHB^{+}97] I. M. Davies, M. J. C. Harding, S. K. Bailey, A. M. Shanks, and R. Länge. Sublethal effects of tributyltin oxide on the dogwhelk Nucella lapillus. Marine Ecology Progress Series, 158:191–204, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/158/m158p191.pdf.

Davies:1998:RCT

[DHB^{+}98] I. M. Davies, M. J. C. Harding, S. K. Bailey, A. M. Shanks, and R. Länge. Reply to comment ‘Tributyltin and imposex: no uncertainty shown’ by Ruiz et al. Marine Ecology Progress Series, 170:295–297, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/170/m170p295.pdf. See

[RQB98a].

Dauwe:1998:CSB

[DHH98] B. Dauwe, P. M. J. Herman, and C. H. R. Heip. Community structure and bioturbation potential of macrofauna at four North Sea stations with contrasting food supply. Marine Ecology Progress Series, 173:67–83, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/173/m173p067.pdf.

Dekshenieks:1996:MVD

[DHKP96] M. M. Dekshenieks, E. E. Hofmann, J. M. Klinck, and E. N. Powell. Modeling the vertical distribution of oyster larvae in response to environmental conditions. Marine Ecology Progress Series, 136:97–110, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/136/m136p097.pdf.

Delmas:1992:ECW

[DHM92] D. Delmas, A. Herbland, and S. Y. Maestrini. Environmental conditions which lead to increase in cell density of the toxic dinoflagellates Dinophysis spp. in nutrient-rich and nutrientpoor waters of the French Atlantic coast. Marine Ecology Progress Series, 89:53–61, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/89/m089p053.pdf.

Davis:1996:MSE

[DHS96] M. Davis, G. A. Hodgkins, and A. W. Stoner. A mesocosm system for ecological research with marine invertebrate larvae. Marine Ecology Progress Series, 130:97–104, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
130/m130p097.pdf.

Dittmann:1996:EMB

[Dit96] S. Dittmann. Effects of macrobenthic burrows on infaunal communities in tropical tidal flats. Marine Ecology Progress Series, 134:119–130, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/134/m134p119.pdf.

Davis:1990:DPV

[DJY90] T. L. O. Davis, G. P. Jenkins, and J. W. Young. Diel patterns of vertical distribution in larvae of southern bluefin Thunnus maccoyii, and other tuna in the East Indian Ocean. Marine Ecology Progress Series, 59:63–74, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/59/m059p063.pdf.

Dacey:1994:HRF

[DKL94] J. W. H. Dacey, G. M. King, and P. S. Lobel. Herbivory by reef fishes and the production of dimethyl sulfide and acrylic acid. Marine Ecology Progress Series, 112:67–74, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
112/m112p067.pdf.

Dudgeon:1995:PRE

[DKVD95] Steven R. Dudgeon, Janet E. Kübler, Robert L. Vadas, and Ian R. Davison. Physiological responses to environmental variation in intertidal red algae: does thallus morphology matter? Marine Ecology Progress Series, 117:193–206, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
117/m117p193.pdf.

DAvanzo:1996:EPR

[DKW96] C. D’Avanzo, J. N. Kremer, and S. C. Wainright. Ecosystem production and respiration in response to eutrophication in shallow temperate estuaries. Marine Ecology Progress Series, 141:263–274, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/141/m141p263.pdf.

DeBlois:1991:FRP

[DL91a] E. M. DeBlois and W. C. Leggett. Functional response and potential impact of invertebrate predators on benthic fish eggs: analysis of the Calliopius laeviusculus-capelin (Mallotus villosus) predator–prey system. Marine Ecology Progress Series, 69:205–216, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/69/m069p205.pdf.

Decho:1991:TCR

[DL91b] A. W. Decho and S. N. Luoma. Time-courses in the retention of food material in the bivalves Potamocorbula amurensis and Macoma balthica significance to the absorption of carbon and chromium. Marine Ecology Progress Series, 78:303–314, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
78/m078p303.pdf.

DeBlois:1993:IAP

[DL93a] E. M. DeBlois and W. C. Leggett. Impact of amphipod predation on the benthic eggs of marine fish: an analysis of Calliopius laeviusculus bioenergetic demands and predation on the eggs of a beach spawning osmeriid (Mallotus villosus). Marine Ecology Progress Series, 93:205–216, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
93/m093p205.pdf.

Dower:1993:PIR

[DL93b] K. M. Dower and M. I. Lucas. Photosynthesis-irradiance relationships and production associated with a warm-core ring shed from the Agulhas Retroflection south of Africa. Marine Ecology Progress Series, 95:141–154, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
95/m095p141.pdf.

Decho:1994:HFA

[DL94a] A. W. Decho and S. N. Luoma. Humic and fulvic acids: sink or source in the availability of metals to the marine bivalves Macoma balthica and Potamocorbula amurensis? Marine Ecology Progress Series, 108:133–145, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
108/m108p133.pdf.

Dring:1994:ISN

[DL94b] M. J. Dring and K. Luning. Influence of spring-neap tidal cycles on the light available for photosynthesis by benthic marine plants. Marine Ecology Progress Series, 104:131– 137, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/104/m104p131.pdf.

Devescovi:1995:SCC

[DL95] M. Devescovi and C. Lucu. Seasonal changes of the copper level in shore crabs Carcinus mediterraneus. Marine Ecology Progress Series, 120:169–174, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/120/m120p169.pdf.

Demers:1993:IIE

[DLDL93] A. Demers, Y. Lagadeuc, J. J. Dodson, and R. Lemieux. Immunofluorescence identification of early life history stages of scallops (Pectinidae). Marine Ecology Progress Series, 97: 83–89, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/97/m097p083.pdf.

Dauchez:1996:NPP

[DLFL96] S. Dauchez, L. Legendre, L. Fortier, and M. Levasseur. New production and production of large phytoplankton (> 5μ m) on the Scotian Shelf (NW Atlantic). Marine Ecology Progress Series, 135:215–222, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/135/m135p215.pdf.

Dupuy:1999:RCF

[DLHB99] C. Dupuy, S. Le Gall, H. J. Hartmann, and M. Bréret. Retention of ciliates and flagellates by the oyster Crassostrea gigas in French Atlantic coastal ponds: protists as a trophic link between bacterioplankton and benthic suspension-feeders. Marine Ecology Progress Series, 177:165–175, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
177/m177p165.pdf.

Davis:1991:ADM

[DLJ91] T. L. O. Davis, V. Lyne, and G. P. Jenkins. Advection, dispersion and mortality of a patch of southern bluefin tuna larvae Thunnus maccoyii in the East Indian Ocean. Marine Ecology Progress Series, 73:33–45, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/73/m073p033.pdf.

DelAmo:1997:IHNa

[DLT^{+}97] Y. Del Amo, O. Le Pape, P. Tréguer, B. Quéguiner, A. Ménesguen, and A. Aminot. Impacts of high-nitrate freshwater inputs on macrotidal ecosystems. I. Seasonal evolution of nutrient limitation for the diatom-dominated phytoplankton of the Bay of Brest (France). Marine Ecology Progress Series, 161:213–224, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/161/m161p213.pdf.

Daly:1991:IPB

[DM91] K. L. Daly and M. C. Macaulay. Influence of physical and biological mesoscale dynamics on the seasonal distribution and behavior of Euphausia superba in the Antarctic marginal ice zone. Marine Ecology Progress Series, 79:37–66, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
79/m079p037.pdf.

Dahlhoff:1996:IPC

[DM96] E. P. Dahlhoff and B. A. Menge. Influence of phytoplankton concentration and wave exposure on the ecophysiology of Mytilus californianus. Marine Ecology Progress Series, 144: 97–107, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/144/m144p097.pdf.

Duarte:1994:RSD

[DMA^{+}94] C. M. Duarte, N. Marbá, N. Agawin, J. Cebrián, S. Enriquez, M. D. Fortes, M. E. Gallegos, M. Merino, B. Olesen, K. Sand-Jensen, J. Uri, and J. Vermaat. Reconstruction of seagrass dynamics: age determinations and associated tools for the seagrass ecologist. Marine Ecology Progress Series, 107:195–209, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/107/m107p195.pdf.

Duarte:1998:RPB

[DMA^{+}98] C. M. Duarte, M. Merino, N. S. R. Agawin, J. Uri, M. D. Fortes, M. E. Gallegos, N. Marbá, and M. A. Hemminga. Root production and belowground seagrass biomass. Marine Ecology Progress Series, 171:97–108, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/171/m171p097.pdf.

Dunn:1999:RPS

[DMM99] R. Dunn, L. S. Mullineaux, and S. W. Mills. Resuspension of postlarval soft-shell clams Mya arenaria through disturbance by the mud snail Ilyanassa obsoleta. Marine Ecology Progress Series, 180:223–232, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/180/m180p223.pdf.

DellAnno:1999:PNA

[DMP^{+}99] A. Dell’Anno, D. Marrale, A. Pusceddu, M. Fabiano, and R. Danovaro. Particulate nucleic acid dynamics in a highly oligotrophic system: the Cretan Sea (Eastern Mediterranean). Marine Ecology Progress Series, 186:19–30, 1999.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
186/m186p019.pdf.

deNys:1998:NMD

[dNDS98] R. de Nys, S. A. Dworjanyn, and P. D. Steinberg. A new method for determining surface concentrations of marine natural products on seaweeds. Marine Ecology Progress Series, 162:79–87, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/162/m162p079.pdf.

deNys:1996:QVS

[dNSR^{+}96] R. de Nys, P. D. Steinberg, C. N. Rogers, T. S. Charlton, and M. W. Duncan. Quantitative variation of secondary metabolites in the sea hare Aplysia parvula and its host plant, Delisea pulchra. Marine Ecology Progress Series, 130:135–146, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
130/m130p135.pdf.

Davis:1992:CGB

[DO92] M. W. Davis and B. L. Olla. Comparison of growth, behavior and lipid concentrations of walleye pollock Theragra chalcogramma larvae fed lipid-enriched, lipid-deficient and field-collected prey. Marine Ecology Progress Series, 90: 23–30, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/90/m090p023.pdf.

Doering:1995:PNL

[DON^{+}95] P. H. Doering, C. A. Oviatt, B. L. Nowicki, E. G. Klos, and L. W. Reed. Phosphorus and nitrogen limitation of primary production in a simulated estuarine gradient. Marine Ecology Progress Series, 124:271–287, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/124/m124p271.pdf.

Dortch:1990:IBA

[Dor90] Q. Dortch. The interaction between ammonium and nitrate uptake in phytoplankton. Marine Ecology Progress Series, 61: 183–201, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/61/m061p183.pdf.

Douillet:1993:BPO

[Dou93a] P. Douillet. Bacterivory in Pacific oyster Crassostrea gigas larvae. Marine Ecology Progress Series, 98:123–134, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
98/m098p123.pdf.

Douillet:1993:CCT

[Dou93b] P. A. Douillet. Carbon contribution through bacterivory in larvae of the Pacific oyster Crassostrea gigas. Marine Ecology Progress Series, 102:303–314, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/102/m102p303.pdf.

Dam:1991:SFB

[DP91] H. G. Dam and W. T. Peterson. In situ feeding behavior of the copepod Temora longicornis: effects of seasonal changes in chlorophyll size fractions and female size. Marine Ecology Progress Series, 71:113–123, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/71/m071p113.pdf.

Davenport:1996:MFA

[DPM96] J. Davenport, P. J. A. Pugh, and J. McKechnie. Mixed fractals and anisotropy in subantarctic marine macroalgae from South Georgia: implications for epifaunal biomass and abundance. Marine Ecology Progress Series, 136:245–255, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
136/m136p245.pdf.

Dam:1991:SMA

[DPO91] H. G. Dam, W. T. Peterson, and A. Okubo. A simple mathematical analysis of the imitations to inferring feeding behavior of zooplankton from gut content. Marine Ecology Progress Series, 69:41–45, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/69/m069p041.pdf.

DelAmo:1997:IHNb

[DQT^{+}97] Y. Del Amo, B. Quéguiner, P. Tréguer, H. Breton, and L. Lampert. Impacts of high-nitrate freshwater inputs on macrotidal ecosystems. II. Specific role of the silicic acid pump in the year-round dominance of diatoms in the Bay of Brest (France). Marine Ecology Progress Series, 161:225–237, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
161/m161p225.pdf.

delRio-Portilla:1992:GPO

[dRAV92] M. A. del Rio-Portilla, A. D. Re-Araujo, and D. Voltolina. Growth of the pearl oyster Plena sterna under different thermic and feeding conditions. Marine Ecology Progress Series, 89:221–227, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/89/m089p221.pdf.

Durbin:1997:LFE

[DRC^{+}97] E. G. Durbin, J. A. Runge, R. G. Campbell, P. R. Garrahan, M. C. Casas, and S. Plourde. Late fall-early winter recruitment of Calanus finmarchicus on Georges Bank. Marine Ecology Progress Series, 151:103–114, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/151/m151p103.pdf.

Demers:1991:RLI

[DRGV91] S. Demers, S. Roy, R. Gagnon, and C. Vignault. Rapid light-induced changes in cell fluorescence and in xanthophyllcycle pigments of Alexandrium excavatum (Dinophyceae) and Thalassiosira pseudonana (Bacillario-phyceae): a photoprotection mechanism. Marine Ecology Progress Series, 76: 185–193, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/76/m076p185.pdf.

Dortch:1997:AVF

[DRP^{+}97] Q. Dortch, R. Robichaux, S. Pool, D. Milsted, G. Mire, N. N. Rabalais, T. M. Soniat, G. A. Fryxell, R. E. Turner, and M. L. Parsons. Abundance and vertical flux of Pseudo-nitzschia in the northern Gulf of Mexico. Marine Ecology Progress Series, 146:249–264, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/146/m146p249.pdf.

Dando:1994:SML

[DRS94] P. R. Dando, S. A. Ridgway, and B. Spiro. Sulphide ’mining’ by lucinid bivalve molluscs: demonstrated by stable sulphur isotope measurements and experimental models. Marine Ecology Progress Series, 107:169–175, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/107/m107p169.pdf.

Dando:1993:VND

[DS93a] P. R. Dando and B. Spiro. Varying nutritional dependence of the thyasirid oivalves Thyasira sarsi and T. equalis on chemoautotrophic symbiotic bacteria, demonstrated by isotope ratios of tissue carbon and shell carbonate. Marine Ecology Progress Series, 92:151–158, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/92/m092p151.pdf.

Das:1993:SCC

[DS93b] T. Das and W. B. Stickle. Sensitivity of crabs Callinectes sapidus and C. similis and the gastropod Stramonita haemastoma to hypoxia and anoxia. Marine Ecology Progress Series, 98:263–274, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/98/m098p263.pdf.

Dalpadado:1996:AMG

[DS96] P. Dalpadado and H. R. Skjoldal. Abundance, maturity and growth of the krill species Thysanoessa inermis and T. longicaudata in the Barents Sea. Marine Ecology Progress Series, 144:175–183, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/144/m144p175.pdf.

Dafner:1999:CSV

[DSG^{+}99] E. V. Dafner, R. Sempéré, N. González, F. Gomez, and M. Goutx. Cross-slope variations of dissolved organic carbon in the Gulf of Cadiz, NE Atlantic Ocean (February 1998). Marine Ecology Progress Series, 189:301–306, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
189/m189p301.pdf.

Duarte:1990:SCB

[DSJ90a] C. M. Duarte and K. Sand-Jensen. Seagrass colonization: biomass development and shoot demography in Cymodocea nodosa patches. Marine Ecology Progress Series, 67:97–103, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/67/m067p097.pdf.

Duarte:1990:SCP

[DSJ90b] C. M. Duarte and K. Sand-Jensen. Seagrass colonization: patch formation and patch growth in Cymodocea nodosa. Marine Ecology Progress Series, 65:193–200, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
65/m065p193.pdf.

DallaVia:1998:LGM

[DSS^{+}98] J. Dalla Via, C. Sturmbauer, G. Schönweger, E. Sötz, S. Mathekowitsch, M. Stifter, and R. Rieger. Light gradients and meadow structure in Posidonia oceanica: ecomorphological and functional correlates. Marine Ecology Progress Series, 163:267–278, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/163/m163p267.pdf.

Dollar:1991:ACB

[DSV^{+}91] S. J. Dollar, S. V. Smith, S. M. Vink, S. Obrebski, and J. T. Hollibaugh. Annual cycle of benthic nutrient fluxes in Tomales Bay, California, and contribution of the benthos to total ecosystem metabolism. Marine Ecology Progress Series, 79:115–125, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/79/m079p115.pdf.

Dame:1991:AMP

[DSW^{+}91] R. F. Dame, J. D. Spurrier, T. M. Williams, B. Kjerfve, R. G. Zingmark, T. G. Wolaver, T. H. Chrzanowski, H. N. McKellar, and F. J. Vernberg. Annual material processing by a salt marsh-estuarine basin in South Carolina, USA. Marine Ecology Progress Series, 72:153–166, 1991. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/72/m072p153.pdf.

delNorte-Campos:1994:DAF

[dT94a] A. G. C. del Norte-Campos and A. Temming. Daily activity, feeding and rations in gobies and brown shrimp in the northern Wadden Sea. Marine Ecology Progress Series, 115: 41–53, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/115/m115p041.pdf.

Dunton:1994:LSP

[DT94b] K. H. Dunton and D. A. Tomasko. In situ photosynthesis in the seagrass Halodule wrightii in a hypersaline subtropical lagoon. Marine Ecology Progress Series, 107:281–293, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
107/m107p281.pdf.

Duarte:1997:RMP

[DTA^{+}97] C. M. Duarte, J. Terrados, N. S. R. Agawin, M. D. Fortes, S. Bach, and W. J. Kenworthy. Response of a mixed Philippine seagrass meadow to experimental burial. Marine Ecology Progress Series, 147:285–294, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/147/m147p285.pdf.

Duarte:1990:SNC

[Dua90] C. M. Duarte. Seagrass nutrient content. Marine Ecology Progress Series, 67:201–207, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/67/m067p201.pdf.

Duarte:1991:ASS

[Dua91] C. M. Duarte. Allometric scaling of seagrass form and productivity. Marine Ecology Progress Series, 77:289–300, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
77/m077p289.pdf.

Duffy:1992:HUP

[Duf92] J. E. Duffy. Host use patterns and demography in a guild of tropical sponge-dwelling shrimps. Marine Ecology Progress Series, 90:127–138, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/90/m090p127.pdf.

Dunne:1999:PFU

[Dun99] R. P. Dunne. Polysulphone film as an underwater dosimeter for solar ultraviolet-B radiation in tropical latitudes. Marine Ecology Progress Series, 189:53–63, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/189/m189p053.pdf.

Durako:1994:SFB

[Dur94] M. J. Durako. Seagrass die-off in Florida Bay (USA): changes in shoot demographic characteristics and population dynamics in Thalassia testudinum. Marine Ecology Progress Series, 110:59–66, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/110/m110p059.pdf.

Dutz:1998:RFN

[Dut98] J. Dutz. Repression of fecundity in the neritic copepod Acartia clausi exposed to the toxic dinoflagellate Alexandrium lusitanicum: relationship between feeding and egg production. Marine Ecology Progress Series, 175:97–107, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
175/m175p097.pdf.

Duarte:1992:SDB

[DV92] C. M. Duarte and D. Vaque. Scale dependence of bacterioplankton patchiness. Marine Ecology Progress Series, 84: 95–100, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/84/m084p095.pdf.

DallaVia:1994:ILT

[DvCd94] J. Dalla Via, G. van den Thillart, O. Cattani, and A. de Zwaan. Influence of long-term hypoxia exposure on the energy metabolism of Solea solea. II. Intermediary metabolism in blood, liver and muscle. Marine Ecology Progress Series, 111:17–27, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/111/m111p017.pdf.

Daan:1992:PCB

[DvdM92] R. Daan, H. van het Groenewoud, S. A. de Jong, and M. Mulder. Physico-chemical and biological features of a drilling site in the North Sea, 1 year after discharges of oil-contaminated drill cuttings. Marine Ecology Progress Series, 91:37–45, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
91/m091p037.pdf.

DallaVia:1997:EVF

[DvdTCC97] J. Dalla Via, G. van den Thillart, O. Cattani, and P. Cortesi. Environmental versus functional hypoxia/anoxia in sole Solea solea: the lactate paradox revisited. Marine Ecology Progress Series, 154:79–90, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/154/m154p079.pdf.

Dvir:1999:NTF

[DvRN99] O. Dvir, J. van Rijn, and A. Neori. Nitrogen transformations and factors leading to nitrite accumulation in a hypertrophic marine fish culture system. Marine Ecology Progress Series, 181:97–106, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/181/m181p097.pdf.

Dickson:1995:AUR

[DW95] M. L. Dickson and P. A. Wheeler. Ammonium uptake and regeneration rates in a coastal upwelling regime. Marine Ecology Progress Series, 121:239–248, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/121/m121p239.pdf.

Diaz:1997:SMN

[DW97] M. C. Diaz and B. B. Ward. Sponge-mediated nitrification in tropical benthic communities. Marine Ecology Progress Series, 156:97–107, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/156/m156p097.pdf.

Dubilier:1998:USC

[DWG98] N. Dubilier, R. Windoffer, and O. Giere. Ultrastructure and stable carbon isotope composition of the hydrothermal vent mussels Bathymodiolus brevior and B. sp. affinis brevior from the North Fiji Basin, western Pacific. Marine Ecology Progress Series, 165:187–193, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/165/m165p187.pdf.

Dilling:1998:FEE

[DWSA98] L. Dilling, J. Wilson, D. Steinberg, and A. Alldredge. Feeding by the euphausiid Euphausia pacifica and the copepod Calanus pacificus on marine snow. Marine Ecology Progress Series, 170:189–201, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/170/m170p189.pdf.

Dalby:1992:REP

[DY92] J. E. Dalby, Jr. and C. M. Young. Role of early postsettlement mortality in setting the upper depth limit of ascidians in Florida epifaunal communities. Marine Ecology Progress Series, 80:221–228, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/80/m080p221.pdf.

Dye:1993:RIM

[Dye93] A. H. Dye. Recolonization of intertidal macroalgae in relation to gap size and molluscan herbivory on a rocky shore on the east coast of southern Africa. Marine Ecology Progress Series, 95:263–271, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/95/m095p263.pdf.

Dye:1998:CLA

[Dye98] A. H. Dye. Community-level analyses of long-term changes in rocky littoral fauna from South Africa. Marine Ecology Progress Series, 164:47–57, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/164/m164p047.pdf.

Dickman:1999:MOE

[DZ99] M. Dickman and F. Zhang. Mid-ocean exchange of container vessel ballast water. 2: Effects of vessel type in the transport of diatoms and dinoflagellates from Manzanillo, Mexico, to Hong Kong, China. Marine Ecology Progress Series, 176:253–262, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/176/m176p253.pdf.

Egge:1992:SRN

[EA92] J. K. Egge and D. L. Aksnes. Silicate as regulating nutrient in phytoplankton competition. Marine Ecology Progress Series, 83:281–289, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/83/m083p281.pdf.

Epstein:1997:ESS

[EAC^{+}97] S. S. Epstein, D. Alexander, K. Cosman, A. Dompé, S. Gallagher, J. Jarsobski, E. Laning, R. Martinez, G. Panasik, C. Peluso, R. Runde, and E. Timmer. Enumeration of sandy sediment bacteria: Are the counts quantitative or relative? Marine Ecology Progress Series, 151:11–16, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
151/m151p011.pdf.

Enriquez:1992:LAS

[EAD92] S. Enriquez, S. Agusti, and C. M. Duarte. Light absorption by seagrass Posidonia oceanica leaves. Marine Ecology Progress Series, 86:201–204, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/86/m086p201.pdf.

Eggleston:1998:EFC

[EAEP98] D. B. Eggleston, D. A. Armstrong, W. E. Elis, and W. S. Patton. Estuarine fronts as conduits for larval transport: hydrodynamics and spatial distribution of Dungeness crab postlarvae. Marine Ecology Progress Series, 164:73–82, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
164/m164p073.pdf.

Eshky:1995:PEC

[EAT95] A. A. Eshky, R. J. A. Atkinson, and A. C. Taylor. Physiological ecology of crabs from Saudi Arabian mangrove. Marine Ecology Progress Series, 126:83–95, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/126/m126p083.pdf.

Ejdung:1992:PBM

[EB92] G. Ejdung and E. Bonsdorff. Predation on the bivalve Macoma balthica by the sopod Saduria entomon: laboratory and field experiments. Marine Ecology Progress Series, 88: 207–214, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/88/m088p207.pdf.

Edmunds:1996:ISS

[EB96] P. J. Edmunds and J. F. Bruno. The importance of sampling scale in ecology: kilometer-wide variation in coral reef communities. Marine Ecology Progress Series, 143:165–171, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
143/m143p165.pdf.

Einav:1995:EAS

[EBB95] R. Einav, S. Breckle, and S. Beer. Ecophysiological adaptation strategies of some intertidal marine macroalgae of the Israeli Mediterranean coast. Marine Ecology Progress Series, 125:219–228, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/125/m125p219.pdf.

Engel:1993:HCM

[EBM93] D. W. Engel, M. Brouwer, and S. McKenna. Hemocyanin concentrations in marine crustaceans as a function of environmental conditions. Marine Ecology Progress Series, 93: 235–244, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/93/m093p235.pdf.

Endean:1997:MCR

[ECF^{+}97] R. Endean, A. M. Cameron, H. E. Fox, R. Tilbury, and L. Gunthorpe. Massive corals are regularly spaced: pattern in a complex assemblage of corals. Marine Ecology Progress Series, 152:119–130, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/152/m152p119.pdf.

Epifanio:1998:ELH

[EDP^{+}98] C. E. Epifanio, A. I. Dittel, S. Park, S. Schwalm, and A. Fouts. Early life history of Hemigrapsus sanguineus, a non-indigenous crab in the Middle Atlantic Bight (USA). Marine Ecology Progress Series, 170:231–238, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
170/m170p231.pdf.

Ebert:1999:GMR

[EDS^{+}99] T. A. Ebert, J. D. Dixon, S. C. Schroeter, P. E. Kalvass, N. T. Richmond, W. A. Bradbury, and D. A. Woodby. Growth and mortality of red sea urchins Strongylocentrotus franciscanus across a latitudinal gradient. Marine Ecology Progress Series, 190:189–209, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/190/m190p189.pdf.

Enriquez:1995:PPM

[EDSJ95] S. Enríquez, C. M. Duarte, and K. Sand-Jensen. Patterns in the photosynthetic metabolism of Mediterranean macrophytes. Marine Ecology Progress Series, 119:243–252, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
119/m119p243.pdf.

Ejdung:1998:PNS

[EE98] G. Ejdung and R. Elmgren. Predation on newly settled bivalves by deposit-feeding amphipods: a Baltic Sea case study. Marine Ecology Progress Series, 168:87–94, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
168/m168p087.pdf.

Edmonds:1997:SDP

[EF97] J. S. Edmonds and W. J. Fletcher. Stock discrimination of pilchards Sardinops sagax by stable isotope ratio analysis of otolith carbonate. Marine Ecology Progress Series, 152:241–247, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/152/m152p241.pdf.

Ellertsen:1995:CPC

[EFSS95] B. Ellertsen, P. Fossum, P. Solemdal, and S. Sundby. The ‘critical period’ concept — a century of recruitment research.

Marine Ecology Progress Series, 128:306–308, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Ellis:1995:SSP

[EG95] T. Ellis and R. N. Gibson. Size-selective predation of 0-group flatfishes on a Scottish coastal nursery ground. Marine Ecology Progress Series, 127:27–37, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/127/m127p027.pdf.

Ellis:1997:PGF

[EG97] T. Ellis and R. N. Gibson. Predation of 0-group flatfishes by 0-group cod: handling times and size selection. Marine Ecology Progress Series, 149:83–90, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/149/m149p083.pdf.

Espoz:1995:LTL

[EGC95] C. Espoz, G. Guzmán, and J. C. Castilla. The lichen Thelidium litorale on shells of intertidal limpets: a case of lichenmediated cryptic mimicry. Marine Ecology Progress Series, 119:191–197, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/119/m119p191.pdf.

Eggleston:1995:RNG

[Egg95] D. B. Eggleston. Recruitment in Nassau grouper Epinephelus striatus: post-settlement abundance, microhabitat features, and ontogenetic habitat shifts. Marine Ecology Progress Series, 124:9–22, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/124/m124p009.pdf.

Eggens:1992:HEA

[EGKE92] M. Eggens, F. Galgani, J. Klungsoyr, and J. Everts. Hepatic EROD activity in dab Limanda limanda in the German Bight using an improved plate-reader method. Marine Ecology Progress Series, 91:71–75, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/91/m091p071.pdf.

Emerson:1994:GSS

[EGMC94] C. W. Emerson, J. Grant, A. Mallet, and C. Carver. Growth and survival of sea scallops Placopecten magellanicus: effects of culture depth. Marine Ecology Progress Series, 108: 119–132, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/108/m108p119.pdf.

Eldridge:1993:BTD

[EJ93] P. M. Eldridge and G. A. Jackson. Benthic trophic dynamics in California coastal basin and continental slope communities inferred using inverse analysis. Marine Ecology Progress Series, 99:115–135, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/99/m099p115.pdf.

Epping:1996:LEO

[EJ96] E. H. G. Epping and B. B. Jørgensen. Light-enhanced oxygen respiration in benthic phototrophic communities. Marine Ecology Progress Series, 139:193–203, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
139/m139p193.pdf.

Egge:1997:ISP

[EJ97] J. K. Egge and A. Jacobsen. Influence of silicate on particulate carbon production in phytoplankton. Marine Ecology Progress Series, 147:219–230, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/147/m147p219.pdf.

Ejdung:1998:BRC

[Ejd98] G. Ejdung. Behavioural responses to chemical cues of predation risk in a three-trophic-level Baltic Sea food chain. Marine Ecology Progress Series, 165:137–144, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
165/m165p137.pdf.

Ebersole:1995:PPB

[EK95] E. L. Ebersole and V. S. Kennedy. Prey preferences of blue crabs Callinectes sapidus feeding on three bivalve species.

Marine Ecology Progress Series, 118:167–177, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
118/m118p167.pdf.

Eklund:1998:RPG

[Ekl98] B. Eklund. Reproductive performance and growth response of the red alga Ceramium strictum under the impact of phenol. Marine Ecology Progress Series, 167:119–126, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
167/m167p119.pdf.

Esford:1990:SCG

[EL90] L. E. Esford and J. C. Lewis. Stiffness of Caribbean gorgonians (Coelenterata, Octocorallia) and Ca/Mg content of their axes. Marine Ecology Progress Series, 67:189–200, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
67/m067p189.pdf.

Eggleston:1997:PSS

[ELG97] D. B. Eggleston, R. N. Lipcius, and J. J. Grover. Predator and shelter-size effects on coral reef fish and spiny lobster prey. Marine Ecology Progress Series, 149:43–59, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
149/m149p043.pdf.

Eggleston:1992:DDP

[ELH92] D. B. Eggleston, R. N. Lipcius, and A. H. Hines. Densitydependent predation by blue crabs upon infaunal clam species with contrasting distribution and abundance patterns. Marine Ecology Progress Series, 85:55–68, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
85/m085p055.pdf.

Eggleston:1990:SSR

[ELMCC90] David B. Eggleston, Romuald N. Lipcius, David L. Miller, and Luis Coba-Cetina. Shelter scaling regulates survival of juvenile Caribbean spiny lobster Panulirus argus. Marine Ecology Progress Series, 62:79–88, 1990. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/62/m062p079.pdf.

Eggleston:1998:SVP

[ELMR98] D. B. Eggleston, R. N. Lipcius, L. S. Marshall, Jr., and S. G. Ratchford. Spatiotemporal variation in postlarval recruitment of the Caribbean spiny lobster in the central Bahamas: lunar and seasonal periodicity, spatial coherence, and wind forcing. Marine Ecology Progress Series, 174:33–49, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
174/m174p033.pdf.

Erftemeijer:1993:SNI

[EM93] P. L. A. Erftemeijer and J. J. Middelburg. Sediment-nutrient interactions in tropical seagrass beds: a comparison between a terrigenous and a carbonate sedimentary environment in South Sulawesi (Indonesia). Marine Ecology Progress Series, 102:187–198, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/102/m102p187.pdf.

Escribano:1999:PCC

[EM99] R. Escribano and I. McLaren. Production of Calanus chilensis in the upwelling area of Antofagasta, northern Chile. Marine Ecology Progress Series, 177:147–156, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
177/m177p147.pdf.

Emlet:1990:FFA

[Eml90] R. B. Emlet. Flow fields around ciliated larvae: effects of natural artificial tethers. Marine Ecology Progress Series, 63: 211–225, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/63/m063p211.pdf.

Estrada:1993:VDC

[EML^{+}93] M. Estrada, C. Marrasé, M. Latasa, E. Berdalet, M. Delgado, and T. Riera. Variability of deep chlorophyll maximum characteristics in the Northwestern Mediterranean. Marine Ecology Progress Series, 92:289–300, 1993. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/92/m092p289.pdf.

Erzini:1997:ESG

[EMR^{+}97] K. Erzini, C. C. Monteiro, J. Ribeiro, M. N. Santos, M. Gaspar, P. Monteiro, and T. C. Borges. An experimental study of gill net and trammel net ‘ghost fishing’ in the Algarve (southern Portugal). Marine Ecology Progress Series, 158: 257–265, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/158/m158p257.pdf.

Enoksson:1993:NRC

[Eno93] V. Enoksson. Nutrient recycling by coastal sediments: effects of added algal material. Marine Ecology Progress Series, 92:245–254, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/92/m092p245.pdf.

Estep:1990:PCU

[ENSR90] K. W. Estep, J. C. Nejstgaard, H. R. Skjoldal, and F. Rey. Predation by copepods upon natural populations of Phaeocystis pouchetii as a function of the physiological state of the prey. Marine Ecology Progress Series, 67:235–249, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
67/m067p235.pdf.

Evans:1992:PEP

[EP92] G. T. Evans and M. A. Paranjape. Precision of estimates of phytoplankton growth and microzooplankton grazing when the functional response of grazers may be nonlinear. Marine Ecology Progress Series, 80:285–290, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
80/m080p285.pdf.

Ellertsdottir:1997:HPI

[EP97] E. Ellertsdóttir and A. F. Peters. High prevalence of infection by endophytic brown algae in populations of Laminaria spp. (Phaeophyceae). Marine Ecology Progress Series, 146:135–143, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/146/m146p135.pdf.

Erdner:1999:CFF

[EPD^{+}99] D. L. Erdner, N. M. Price, G. J. Doucette, M. L. Peleato, and D. M. Anderson. Characterization of ferredoxin and flavodoxin as markers of iron limitation in marine phytoplankton. Marine Ecology Progress Series, 184:43–53, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
184/m184p043.pdf.

Epifanio:1999:DBM

[EPDC99] C. E. Epifanio, G. Perovich, A. I. Dittel, and S. C. Cary. Development and behavior of megalopa larvae and juveniles of the hydrothermal vent crab Bythograea thermydron. Marine Ecology Progress Series, 185:147–154, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
185/m185p147.pdf.

Eccleston-Parry:1994:CGK

[EPL94] J. D. Eccleston-Parry and B. S. C. Leadbeater. A comparison of the growth kinetics of six marine heterotrophic nanoflagellates fed with one bacterial species. Marine Ecology Progress Series, 105:167–177, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/105/m105p167.pdf.

Escaravage:1999:RPC

[EPN^{+}99] V. Escaravage, T. C. Prins, C. Nijdam, A. C. Smaal, and J. C. H. Peeters. Response of phytoplankton communities to nitrogen input reduction in mesocosm experiments. Marine Ecology Progress Series, 179:187–199, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
179/m179p187.pdf.

Eppley:1990:IEO

[ER90] Z. A. Eppley and M. A. Rubega. Indirect effects of an oil spill: reproductive failure in a population of south polar skuas following the ‘Bahia Paraiso’ oil spill in Antarctica. Marine Ecology Progress Series, 67:1–6, 1990. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/67/m067p001.pdf.

Epstein:1995:ESS

[ER95a] S. S. Epstein and J. Rossel. Enumeration of sandy sediment bacteria: search for optimal protocol. Marine Ecology Progress Series, 117:289–298, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/117/m117p289.pdf.

Epstein:1995:MLS

[ER95b] S. S. Epstein and J. Rossel. Methodology of in situ grazing experiments: evaluation of a new vital dye for preparation of fluorescently labeled bacteria. Marine Ecology Progress Series, 128:143–150, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/128/m128p143.pdf.

Everett:1995:EOM

[ERC95] R. A. Everett, G. M. Ruiz, and J. T. Carlton. Effect of oyster mariculture on submerged aquatic vegetation: an experimental test in a Pacific northwest estuary. Marine Ecology Progress Series, 125:205–217, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/125/m125p205.pdf.

Erga:1990:DVP

[ES90] S. R. Erga and H. R. Skjoldal. Diel variations in photosynthetic activity of summer phytoplankton in lindaspollene, western Norway. Marine Ecology Progress Series, 65: 73–85, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/65/m065p073.pdf.

Engel:1999:ITE

[ES99] A. Engel and M. Schartau. Influence of transparent exopolymer particles (TEP) on sinking velocity of Nitzschia closterium aggregates. Marine Ecology Progress Series, 182: 69–76, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/182/m182p069.pdf.

Ebert:1994:SPR

[ESDK94] T. A. Ebert, S. C. Schroeter, J. D. Dixon, and P. Kalvass. Settlement patterns of red and purple sea urchins (Strongylocentrotus franciscanus and S. purpuratus) in California, USA. Marine Ecology Progress Series, 111:41–52, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
111/m111p041.pdf.

Eilertsen:1995:PCD

[EST95] H. C. Eilertsen, S. Sandberg, and H. Tøllefsen. Photoperiodic control of diatom spore growth; a theory to explain the onset of phytoplankton blooms. Marine Ecology Progress Series, 116:303–307, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/116/m116p303.pdf.

Edmunds:1991:EHH

[EW91] P. J. Edmunds and J. D. Witman. Effect of Hurricane Hugo on the primary framework of a reef along the south shore of St. John, US Virgin Islands. Marine Ecology Progress Series, 78:201–204, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/78/m078p201.pdf.

Egan:1990:PLH

[EY90] B. Egan and C. Yarish. Productivity and life history of Laminaria longicruris at its southern limit in the Western Atlantic Ocean. Marine Ecology Progress Series, 67:263–273, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
67/m067p263.pdf.

Faria:1999:VRR

[FA99] C. Faria and V. Almada. Variation and resilience of rocky intertidal fish in western Portugal. Marine Ecology Progress Series, 184:197–203, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/184/m184p197.pdf.

Fabricius:1995:SPT

[Fab95] K. E. Fabricius. Slow population turnover in the soft coral genera Sinularia and Sarcophyton on mid- and outer-shelf reefs of the Great Barrier Reef. Marine Ecology Progress Series, 126:145–152, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/126/m126p145.pdf.

Feldman:1997:ESS

[FAED97] K. L. Feldman, D. A. Armstrong, D. B. Eggleston, and B. R. Dumbauld. Effects of substrate selection and post-settlement survival on recruitment success of the thalassinidean shrimp Neotrypaea californiensis to intertidal shell and mud habitats. Marine Ecology Progress Series, 150:121–136, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
150/m150p121.pdf.

Freese:1999:ETS

[FAHW99] L. Freese, P. J. Auster, J. Heifetz, and B. Wing. Effects of trawling on seafloor habitat and associated invertebrate taxa in the Gulf of Alaska. Marine Ecology Progress Series, 182:119–126, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/182/m182p119.pdf.

Falkenhaug:1996:DSP

[Fal96] T. Falkenhaug. Distributional and seasonal patterns of ctenophores in Malangen, northern Norway. Marine Ecology Progress Series, 140:59–70, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/140/m140p059.pdf.

Frechette:1992:IFS

[FAP92] M. Frechétte, A. E. Aitken, and L. Page. Interdependence of food and space limitation of a benthic suspension feeder: consequences for self-thinning relationships. Marine Ecology Progress Series, 83:55–62, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/83/m083p055.pdf.

Farias:1998:PRB

[Far98] L. Farías. Potential role of bacterial mats in the nitrogen budget of marine sediments: the case of Thioploca spp. Marine Ecology Progress Series, 170:291–292, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
170/m170p291.pdf.

Flinkman:1998:CNB

[FAVV98] J. Flinkman, E. Aro, I. Vuorinen, and M. Viitasalo. Changes in northern Baltic zooplankton and herring nutrition from 1980s to 1990s: top-down and bottom-up processes at work. Marine Ecology Progress Series, 165:127–136, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
165/m165p127.pdf.

Fossaa:1990:BDM

[FB90] J. H. Fosså and T. Brattegard. Bathymetric distribution of Mysidacea in fjords of western Norway. Marine Ecology Progress Series, 67:7–18, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/67/m067p007.pdf.

Ferrell:1991:DAA

[FB91] D. J. Ferrell and J. D. Bell. Differences among assemblages of fish associated with Zostera capricorni and bare sand over a large spatial scale. Marine Ecology Progress Series, 72: 15–24, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/72/m072p015.pdf.

Flach:1994:DGM

[FB94] E. C. Flach and J. J. Beukema. Density-governing mechanisms in populations of the lugworm Arenicola marina on tidal flats. Marine Ecology Progress Series, 115:139–149, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/115/m115p139.pdf.

Fenchel:1995:MCS

[FB95] T. Fenchel and C. Bernard. Mats of colourless sulphur bacteria. I. Major microbial processes. Marine Ecology Progress Series, 128:161–170, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/128/m128p161.pdf.

Fernandez:1997:PPP

[FB97] C. Fernandez and C. F. Boudouresque. Phenotypic plasticity of Paracentrotus lividus (Echinodermata: Echinoidea) in a lagoonal environment. Marine Ecology Progress Series, 152: 145–154, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/152/m152p145.pdf.

Fonseca:1998:IPS

[FB98] M. S. Fonseca and S. S. Bell. Influence of physical setting on seagrass landscapes near Beaufort, North Carolina, USA. Marine Ecology Progress Series, 171:109–121, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
171/m171p109.pdf.

Fernandez:1993:POI

[FBHH93] E. Fernández, P. Boyd, P. M. Holligan, and D. S. Harbour. Production of organic and inorganic carbon within a large-scale coccolithophore bloom in the northeast Atlantic Ocean. Marine Ecology Progress Series, 97:271–285, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
97/m097p271.pdf.

Fisher:1995:ASL

[FBL95] N. S. Fisher, V. T. Breslin, and M. Levandowsky. Accumulation of silver and lead in estuarine microzooplankton. Marine Ecology Progress Series, 116:207–215, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
116/m116p207.pdf.

Fernandez:1994:HRL

[FBMH94] E. Fernandez, W. M. Balch, E. Maranon, and P. M. Holligan. High rates of lipid biosynthesis in cultured, mesocosm and coastal populations of the coccolithophore Emiliania huxleyi. Marine Ecology Progress Series, 114:13–22, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
114/m114p013.pdf.

Frank:1995:AXI

[FBR^{+}95] U. Frank, I. Brickner, B. Rinkevich, Y. Loya, R. P. M. Bak, Y. Achituv, and M. Ilan. Allogeneic and xenogeneic interactions in reef-building corals may induce tissue growth without calcification. Marine Ecology Progress Series, 124:181– 188, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/124/m124p181.pdf.

Ferraro:1990:TLS

[FC90] S. P. Ferraro and F. A. Cole. Taxonomic level and sample size sufficient for assessing pollution impacts on the Southern California Bight macro-benthos. Marine Ecology Progress Series, 67:251–262, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/67/m067p251.pdf.

Fessenden:1994:CPP

[FC94] L. Fessenden and T. J. Cowles. Copepod predation on phagotrophic ciliates in Oregon coastal waters. Marine Ecology Progress Series, 107:103–111, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/107/m107p103.pdf.

Fang:1991:FUA

[FCC91] L.-S. Fang, Y.-W. Chen, and C.-S. Chen. Feasibility of using ATP as an index for environmental stress on hermatypic coral. Marine Ecology Progress Series, 70:257–262, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
70/m070p257.pdf.

Fores:1994:NAN

[FCCM94] E. Forès, R. R. Christian, F. A. Comin, and M. Menendez. Network analysis on nitrogen cycling in a coastal lagoon. Marine Ecology Progress Series, 106:283–290, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
106/m106p283.pdf.

Frid:1999:LTC

[FCH99] C. L. J. Frid, R. A. Clark, and J. A. Hall. Long-term changes in the benthos on a heavily fished ground off the NE coast of England. Marine Ecology Progress Series, 188:13–20, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
188/m188p013.pdf.

Fisher:1994:NIG

[FCMB94] C. R. Fisher, J. J. Childress, S. A. Macko, and J. M. Brooks. Nutritional interactions in Galapagos Rift hydrothermal vent communities: inferences from stable carbon and nitrogen isotope analyses. Marine Ecology Progress Series, 103:45–55, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/103/m103p045.pdf.

France:1998:MTC

[FCP98] R. France, M. Chandler, and R. Peters. Mapping trophic continua of benthic foodwebs: body size- δ ^{15}N relationships. Marine Ecology Progress Series, 174:301–306, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
174/m174p301.pdf.

Fogel:1992:RCA

[FCVS92] M. L. Fogel, L. A. Cifuentes, D. J. Velinsky, and J. H. Sharp. Relationship of carbon availability in estuarine phytoplankton to isotopic composition. Marine Ecology Progress Series, 82:291–300, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/82/m082p291.pdf.

Fiedler:1998:CLR

[FCW98] T. J. Fiedler, M. E. Clarke, and P. J. Walsh. Condition of laboratory-reared and wild-caught larval Atlantic menhaden Brevoortia tyrannus as indicated by metabolic enzyme activities. Marine Ecology Progress Series, 175:51–66, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
175/m175p051.pdf.

Fiala:1992:VIP

[FD92] M. Fiala and D. Delille. Variability and interactions of phytoplankton and bacterioplankton in the Antarctic neritic area. Marine Ecology Progress Series, 89:135–146, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
89/m089p135.pdf.

Fuhrman:1997:WAN

[FD97] J. A. Fuhrman and A. A. Davis. Widespread archaea and novel bacteria from the deep sea as shown by ^{16}S rRNA gene sequences. Marine Ecology Progress Series, 150:275– 285, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/150/m150p275.pdf.

Feinberg:1998:EDD

[FD98] L. R. Feinberg and H. G. Dam. Effects of diet on dimensions, density and sinking rates of fecal pellets of the copepod Acartia tonsa. Marine Ecology Progress Series, 175:87–96, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
175/m175p087.pdf.

Fan:1999:RPR

[FD99] T. Y. Fan and C. F. Dai. Reproductive plasticity in the reef coral Echinopora lamellosa. Marine Ecology Progress Series, 190:297–301, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/190/m190p297.pdf.

Forward:1996:EEC

[FDR^{+}96] R. B. Forward, Jr., M. C. DeVries, D. Rittschof, D. A. Z. Frankel, J. P. Bischoff, C. M. Fisher, and J. M. Welch. Effects of environmental cues on metamorphosis of the blue crab Callinectes sapidus. Marine Ecology Progress Series, 131:165–177, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/131/m131p165.pdf.

Fowler:1992:MSA

[FDW92] A. J. Fowler, P. J. Doherty, and D. McB. Williams. Multiscale analysis of recruitment of a coral reef fish on the Great Barrier Reef. Marine Ecology Progress Series, 82:131–141, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/82/m082p131.pdf.

Fong:1993:CMB

[FDZ93] P. Fong, R. M. Donohoe, and J. B. Zedler. Competition with macroalgae and benthic cyanobacterial mats limits phytoplankton abundance in experimental microcosms. Marine Ecology Progress Series, 100:97–102, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
100/m100p097.pdf.

Fong:1994:NCT

[FDZ94] P. Fong, R. M. Donohoe, and J. B. Zedler. Nutrient concentration in tissue of the macroalga Enteromorpha as a function of nutrient history: an experimental evaluation using field microcosms. Marine Ecology Progress Series, 106:273–281, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
106/m106p273.pdf.

Fernandez:1999:CDC

[Fer99] M. Fernández. Cannibalism in Dungeness crab Cancer magister: effects of predator–prey size ratio, density, and habitat type. Marine Ecology Progress Series, 182:221–230, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
182/m182p221.pdf.

Fevolden:1992:AVI

[Fev92] S. E. Fevolden. Allozymic variability in the Iceland scallop Chlamys islandica: geographic variation and lack of growthheterozygosity correlations. Marine Ecology Progress Series, 85:259–268, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/85/m085p259.pdf.

Flynn:1992:NPF

[FF92a] K. J. Flynn and K. Flynn. Non-protein free amines in microalgae: consequences for the measurement of intracellular amino acids and of the glutamine/glutamate ratio. Marine Ecology Progress Series, 89:73–79, 1992. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/89/m089p073.pdf.

Frost:1992:GIL

[FF92b] B. W. Frost and N. C. Franzen. Grazing and iron limitation in the control of phytoplankton stock and nutrient concentration: a chemostat analogue of the Pacific equatorial upwelling zone. Marine Ecology Progress Series, 83:291–303, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
83/m083p291.pdf.

Frankovich:1997:SEL

[FF97] T. A. Frankovich and J. W. Fourqurean. Seagrass epiphyte loads along a nutrient availability gradient, Florida Bay, USA. Marine Ecology Progress Series, 159:37–50, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
159/m159p037.pdf.

Fiksen:1999:MGI

[FF99] Ø. Fiksen and A. Folkvord. Modelling growth and ingestion processes in herring Clupea harengus larvae. Marine Ecology Progress Series, 184:273–289, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/184/m184p273.pdf.

Flynn:1994:CTC

[FFF^{+}94] K. Flynn, J. M. Franco, P. Fernandez, B. Reguera, M. Zapata, G. Wood, and K. J. Flynn. Changes in toxin content, biomass and pigments of the dinoflagellate Alexandrium minutum during nitrogen refeeding and growth into nitrogen or phosphorus stress. Marine Ecology Progress Series, 111:99–109, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
111/m111p099.pdf.

Forbes:1996:IMF

[FFH96] V. E. Forbes, T. L. Forbes, and M. Holmer. Inducible metabolism of fluoranthene by the opportunistic polychaete Capitella sp. I. Marine Ecology Progress Series, 132:63–70, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/132/m132p063.pdf.

Fleck:1999:PRP

[FFH99] J. Fleck, W. K. Fitt, and M. G. Hahn. A proline-rich peptide originating from decomposing mangrove leaves is one natural metamorphic cue of the tropical jellyfish Cassiopea xamachana. Marine Ecology Progress Series, 183:115–124, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
183/m183p115.pdf.

Flynn:1993:CDI

[FFJ93] K. Flynn, K. J. Flynn, and K. J. Jones. Changes in dinoflagellate intracellular amino acids in response to diurnal changes in light and N supply. Marine Ecology Progress Series, 100:245–252, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/100/m100p245.pdf.

Forward:1994:MMB

[FFR94] Richard B. Forward, Jr., David A. Z. Frankel, and Dan Rittschof. Molting of megalopae from the blue crab Callinectes sapidus: effects of offshore and estuarine cues. Marine Ecology Progress Series, 113:55–59, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/113/m113p055.pdf.

Fitzgerald:1998:MEC

[FFT98] T. P. Fitzgerald, R. B. Forward, Jr., and R. A. Tankersley. Metamorphosis of the estuarine crab Rhithropanopeus harrisii: effect of water type and adult odor. Marine Ecology Progress Series, 165:217–223, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/165/m165p217.pdf.

Frantzis:1993:IAG

[FG93] A. Frantzis and A. Gremare. Ingestion, absorption, and growth rates of Paracentrotus lividus (Echinodermata: Echinoidea) fed different macrophytes. Marine Ecology Progress Series, 95:169–183, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/95/m095p169.pdf.

Frankignoulle:1996:CFC

[FGB^{+}96] M. Frankignoulle, J. P. Gattuso, R. Biondo, I. Bourge, G. Copin-Montégut, and M. Pichon. Carbon fluxes in coral reefs. II. Eulerian study of inorganic carbon dynamics and measurement of air-sea CO_{2} exchanges. Marine Ecology Progress Series, 145:123–132, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/145/m145p123.pdf.

Freire:1995:FEV

[FGG95] J. Freire and E. González-Gurriarán. Feeding ecology of the velvet swimming crab Necora puber in mussel raft areas of the Ría de Arousa (Galicia, NW Spain). Marine Ecology Progress Series, 119:139–154, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/119/m119p139.pdf.

Fishelson:1997:BGP

[FGG97] L. Fishelson, M. Goren, and O. Gon. Black gut phenomenon in cardinal fishes (Apogonidae, Teleostei). Marine Ecology Progress Series, 161:295–298, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/161/m161p295.pdf.

Fautin:1995:CBS

[FGH95] D. G. Fautin, C. C. Guo, and J. S. Hwang. Costs and benefits of the symbiosis between the anemoneshrimp Periclimenes brevicarpalis and its host Entacmaea quadricolor. Marine Ecology Progress Series, 129:77–84, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/129/m129p077.pdf.

Forster:1995:EBO

[FGKP95] S. Forster, G. Graf, J. Kitlar, and M. Powilleit. Effects of bioturbation in oxic and hypoxic conditions: a microcosm experiment with a North Sea sediment community. Marine Ecology Progress Series, 116:153–161, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/116/m116p153.pdf.

Forja:1998:MNF

[FGP98] J. M. Forja and A. Gómez-Parra. Measuring nutrient fluxes across the sediment–water interface using benthic chambers.

Marine Ecology Progress Series, 164:95–105, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
164/m164p095.pdf.

Ferrer:1997:ECT

[FGR97] E. Ferrer, A. Gómez Garreta, and M. A. Ribera. Effect of Caulerpa taxifolia on the productivity of two Mediterranean macrophytes. Marine Ecology Progress Series, 149: 279–287, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/149/m149p279.pdf.

Fransz:1998:MPS

[FGS98] H. G. Fransz, S. R. Gonzalez, and S. F. Steeneken. Metazoan plankton and the structure of the plankton community in the stratified North Sea. Marine Ecology Progress Series, 175:191–200, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/175/m175p191.pdf.

Flach:1996:VDM

[FH96] E. Flach and C. Heip. Vertical distribution of macrozoobenthos within the sediment on the continental slope of the Goban Spur area (NE Atlantic). Marine Ecology Progress Series, 141:55–66, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/141/m141p055.pdf.

France:1997:DVM

[FH97a] R. L. France and J. G. Holmquist. Delta ^{13}C variability of macroalgae: effects of water motion via baffling by seagrasses and mangroves. Marine Ecology Progress Series, 149: 305–308, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/149/m149p305.pdf.

Fujiwara:1997:HCP

[FH97b] M. Fujiwara and R. C. Highsmith. Harpacticoid copepods: potential link between inbound adult salmon and outbound juvenile salmon. Marine Ecology Progress Series, 158:205– 216, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/158/m158p205.pdf.

Frid:1999:ICN

[FH99] C. L. J. Frid and S. J. Hall. Inferring changes in North Sea benthos from fish stomach analysis. Marine Ecology Progress Series, 184:183–188, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/184/m184p183.pdf.

Forcada:1999:SMM

[FHA99] J. Forcada, P. S. Hammond, and A. Aguilar. Status of the Mediterranean monk seal Monachus monachus in the western Sahara and the implications of a mass mortality event. Marine Ecology Progress Series, 188:249–261, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
188/m188p249.pdf.

France:1998:ENF

[FHCC98] R. France, J. Holmquist, M. Chandler, and A. Cattaneo. δ 15N evidence for nitrogen fixation associated with macroalgae from a seagrass-mangrove-coral reef system. Marine Ecology Progress Series, 167:297–299, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/167/m167p297.pdf.

Forster:1996:IBL

[FHZ96] S. Forster, M. Huettel, and W. Ziebis. Impact of boundary layer flow velocity on oxygen utilisation in coastal sediments. Marine Ecology Progress Series, 143:173–185, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
143/m143p173.pdf.

Fernandez:1993:HSY

[FIA93] M. Fernandez, O. Iribame, and D. Armstrong. Habitat selection by young-of-the-year Dungeness crab Cancer magister and predation risk in intertidal habitats. Marine Ecology Progress Series, 92:171–177, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/92/m092p171.pdf.

Fujita:1995:DGS

[FII95] T. Fujita, T. Inada, and Y. Ishito. Depth-gradient structure of the demersal fish community on the continental shelf and upper slope off Sendai Bay, Japan. Marine Ecology Progress Series, 118:13–23, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/118/m118p013.pdf.

Fromentin:1993:PMI

[FIL93] J.-M. Fromentin, F. Ibanez, and P. Legendre. A phytosociological method for interpreting plankton data. Marine Ecology Progress Series, 93:285–306, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/93/m093p285.pdf.

Flynn:1994:UIA

[FJR^{+}94] K. J. Flynn, K. J. Jones, R. Raine, J. Richard, and K. Flynn. Use of intracellular amino acid analysis as an indicator of the physiological status of natural dinoflagellate populations. Marine Ecology Progress Series, 103:175–186, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
103/m103p175.pdf.

Fabricius:1995:WMR

[FK95a] K. E. Fabricius and D. W. Klumpp. Widespread mixotrophy in reef-inhabiting soft corals: the influence of depth, and colony expansion and contraction on photosynthesis. Marine Ecology Progress Series, 125:195–204, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
125/m125p195.pdf.

Fitzpatrick:1995:EPS

[FK95b] J. Fitzpatrick and H. Kirkman. Effects of prolonged shading stress on growth and survival of seagrass Posidonia australis in Jervis Bay, New South Wales, Australia. Marine Ecology Progress Series, 127:279–289, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/127/m127p279.pdf.

Findlay:1992:CFD

[FKB92] Robert H. Findlay, Stacy L. Kim, and Cheryl Ann Butman. Colonization of freshly deposited barite and silica sediments by marine microorganisms in a laboratory flume flow. Marine Ecology Progress Series, 90:73–88, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/90/m090p073.pdf.

Fonseca:1990:CFA

[FKC^{+}90] Mark S. Fonseca, W. Judson Kenworthy, David R. Colby, Keith A. Rittmaster, and Gordon W. Thayer. Comparisons of fauna among natural and transplanted eelgrass Zostera marina meadows: criteria for mitigation. Marine Ecology Progress Series, 65:251–264, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/65/m065p251.pdf.

Fonseca:1996:DPSa

[FKC96] M. S. Fonseca, W. J. Kenworthy, and F. X. Courtney. Development of planted seagrass beds in Tampa Bay, Florida, USA. I. Plant components. Marine Ecology Progress Series, 132:127–139, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/132/m132p127.pdf.

Fenchel:1990:WCA

[FKR90] T. Fenchel, L. D. Kristensen, and L. Rasmussen. Water column anoxia: vertical zonation of planktonic protozoa. Marine Ecology Progress Series, 62:1–10, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/62/m062p001.pdf.

Fujikura:1999:DCB

[FKT^{+}99] K. Fujikura, S. Kojima, K. Tamaki, Y. Maki, J. Hunt, and T. Okutani. The deepest chemosynthesis-based community yet discovered from the hadal zone, 7326 m deep, in the Japan Trench. Marine Ecology Progress Series, 190:17–26, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
190/m190p017.pdf.

Frechette:1990:DBF

[FL90] M. Fréchette and D. Lefaivre. Discriminating between food and space imitation in benthic suspension feeders using selfthinning relationships. Marine Ecology Progress Series, 65: 15–23, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/65/m065p015.pdf.

Furuya:1992:EPC

[FL92] K. Furuya and W. K. W. Li. Evaluation of photosynthetic capacity in phytoplankton by flow cytometric analysis of DCMU-enhanced chlorophyll fluorescence. Marine Ecology Progress Series, 88:279–287, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/88/m088p279.pdf.

Fortier:1992:EPD

[FLDT92] L. Fortier, M. E. Levasseur, R. Drolet, and J.-C. Therriault. Export production and the distribution of fish larvae and their prey in a coastal jet frontal region. Marine Ecology Progress Series, 85:203–218, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/85/m085p203.pdf.

Flood:1991:YSO

[Flo91] P. R. Flood. Yellow-stained oikopleurid appendicularians are caused by bacterial parasitism. Marine Ecology Progress Series, 71:291–295, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/71/m071p291.pdf.

Flynn:1990:DNSa

[Fly90a] K. J. Flynn. The determination of nitrogen status in microalgae. Marine Ecology Progress Series, 61:297–307, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
61/m061p297.pdf.

Flynn:1990:DNSb

[Fly90b] K. J. Flynn. The determination of nitrogen status in microalgae. Marine Ecology Progress Series, 61:297–307, 1990.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
61/m061p297.pdf.

Flynn:1998:EKP

[Fly98] K. J. Flynn. Estimation of kinetic parameters for the transport of nitrate and ammonium into marine phytoplankton. Marine Ecology Progress Series, 169:13–28, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
169/m169p013.pdf.

Flynn:1999:NTA

[Fly99] K. J. Flynn. Nitrate transport and ammonium-nitrate interactions at high nitrate concentrations and low temperature. Marine Ecology Progress Series, 187:283–287, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
187/m187p283.pdf.

Franks:1994:SNF

[FM94] P. J. S. Franks and J. Marra. A simple new formulation for phytoplankton photo-response and an application in a wind-driven mixed-layer model. Marine Ecology Progress Series, 111:143–153, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/111/m111p143.pdf.

Ferrari:1995:SFS

[FM95a] G. M. Ferrari and M. Mingazzini. Synchronous fluorescence spectra of dissolved organic matter (DOM) of algal origin in marine coastal waters. Marine Ecology Progress Series, 125:305–315, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/125/m125p305.pdf.

Folkvord:1995:RDR

[FM95b] A. Folkvord and E. Moksness. RNA/DNA ratios and growth of herring larvae. Marine Ecology Progress Series, 121:311– 312, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/121/m121p311.pdf.

Fourqurean:1997:STV

[FMFH97] J. W. Fourqurean, T. O. Moore, B. Fry, and J. T. Hollibaugh. Spatial and temporal variation in C:N:P ratios, δ ^{1}5N, and δ ^{13}C of eelgrass Zostera marina as indicators of ecosystem processes, Tomales Bay, California, USA. Marine Ecology Progress Series, 157:147–157, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/157/m157p147.pdf.

Fonseca:1996:DPSb

[FMH96] M. S. Fonseca, D. L. Meyer, and M. O. Hall. Development of planted seagrass beds in Tampa Bay, Florida, USA. II. Faunal components. Marine Ecology Progress Series, 132: 141–156, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/132/m132p141.pdf.

Fahnenstiel:1995:TSG

[FML^{+}95] G. L. Fahnenstiel, M. J. McCormick, G. A. Lang, D. G. Redalje, S. E. Lohrenz, M. Markowitz, B. Wagoner, and H. J. Carrick. Taxon-specific growth and loss rates for dominant phytoplankton populations from the northern Gulf of Mexico. Marine Ecology Progress Series, 117:229–239, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
117/m117p229.pdf.

Franzoi:1993:LCF

[FMRC93] P. Franzoi, R. Maccagnani, R. Rossi, and V. U. Ceccherelli. Life cycles and feeding habits of Syngnathus taenionofus and S. abasrer (Pisces, Syngnathidae) in a brackish bay of the Po River Delta (Adriatic Sea). Marine Ecology Progress Series, 97:71–81, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/97/m097p071.pdf.

Fisher:1991:AMM

[FNF91] N. S. Fisher, C. V. Nolan, and S. W. Fowler. Assimilation of metals in marine copepods and its biogeochemical implications. Marine Ecology Progress Series, 71:37–43, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
71/m071p037.pdf.

Feder:1994:NCS

[FNJ^{+}94] H. M. Feder, A. S. Naidu, S. C. Jewett, J. M. Hameedi, W. R. Johnson, and T. E. Whitledge. The northeastern Chukchi Sea: benthos-environmental interactions. Marine Ecology Progress Series, 111:171–190, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/111/m111p171.pdf.

Fowler:1990:STP

[Fow90a] A. J. Fowler. Spatial and temporal patterns of distribution and abundance of chaetodontid fishes at One Tree Reef, southern GBR. Marine Ecology Progress Series, 64:39–53, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/64/m064p039.pdf.

Fowler:1990:VAG

[Fow90b] A. J. Fowler. Validation of annual growth increments in the otoliths of a small, tropical coral reef fish. Marine Ecology Progress Series, 64:25–38, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/64/m064p025.pdf.

Fenical:1991:DPS

[FP91a] W. Fenical and J. R. Pawlik. Defensive properties of secondary metabolites from the Caribbean gorgonian coral Erythropodium caribaeorum. Marine Ecology Progress Series, 75:1–8, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/75/m075p001.pdf.

Figueiras:1991:MAT

[FP91b] F. G. Figueiras and Y. Pazos. Microplankton assemblages in three Bias Baixas (Vigo, Arosa and Muros, Spain) with a subsurface chlorophyll maximum: their relationships to hydrography. Marine Ecology Progress Series, 76:219–233, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
76/m076p219.pdf.

Fromentin:1996:CEE

[FP96a] J. M. Fromentin and B. Planque. Calanus and environment in the eastern North Atlantic. II. Influence of the North Atlantic Oscillation on C. finmarchicus and C. helgolandicus. Marine Ecology Progress Series, 134:111–118, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
134/m134p111.pdf.

Froneman:1996:SGM

[FP96b] P. W. Froneman and R. Perissinotto. Structure and grazing of the microzooplankton communities of the subtropical convergence and a warm-core eddy in the Atlantic sector of the Southern Ocean. Marine Ecology Progress Series, 135:237–245, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/135/m135p237.pdf.

Fisher:1992:NLP

[FPAH92] T. R. Fisher, E. R. Peele, J. W. Ammerman, and L. W. Harding, Jr. Nutrient limitation of phytoplankton in Chesapeake Bay. Marine Ecology Progress Series, 82:51–63, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
82/m082p051.pdf.

Fanelli:1994:HPA

[FPB^{+}94] G. Fanelli, S. Piraino, G. Belmonte, S. Geraci, and F. Boero. Human predation along Apulian rocky coasts (SE Italy): desertification caused by Lithophaga litho-phaga (Mollusca) fisheries. Marine Ecology Progress Series, 110:1–8, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
110/m110p001.pdf.

Fraga:1992:SVP

[FPFR92] F. Fraga, F. F. Perez, F. G. Figueiras, and A. F. Rios. Stoichiometric variations of N, P, C and O during a Gymnodinium catenation rod tide and their interpretation. Marine Ecology Progress Series, 87:123–134, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/87/m087p123.pdf.

Fortier:1995:MMH

[FPG95] L. Fortier, D. Ponton, and M. Gilbert. The match/mismatch hypothesis and the feeding success of fish larvae in ice-covered southeastern Hudson Bay. Marine Ecology Progress Series, 120:11–27, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/120/m120p011.pdf.

Ferrier-Pages:1998:RDO

[FPGC^{+}98] C. Ferrier-Pagès, J. P. Gattuso, G. Cauwet, J. Jaubert, and D. Allemand. Release of dissolved organic carbon and nitrogen by the zooxanthellate coral Galaxea fascicularis. Marine Ecology Progress Series, 172:265–274, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
172/m172p265.pdf.

Ferrier-Pages:1999:ESV

[FPGJ99] C. Ferrier-Pagès, J. P. Gattuso, and J. Jaubert. Effect of small variations in salinity on the rates of photosynthesis and respiration of the zooxanthellate coral Stylophora pistillata. Marine Ecology Progress Series, 181:309–314, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
181/m181p309.pdf.

Findlay:1992:BMO

[FPLH92] S. Findlay, M. L. Pace, D. Lints, and K. Howe. Bacterial metabolism of organic carbon in the tidal freshwater Hudson Estuary. Marine Ecology Progress Series, 89:147–153, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
89/m089p147.pdf.

Froneman:1997:DPC

[FPP^{+}97] P. W. Froneman, E. A. Pakhomov, R. Perissinotto, R. K. Laubscher, and C. D. McQuaid. Dynamics of the plankton communities of the Lazarev Sea (Southern Ocean) during seasonal ice melt. Marine Ecology Progress Series, 149: 201–214, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/149/m149p201.pdf.

Froneman:1996:RMD

[FPPM96] P. W. Froneman, E. A. Pakhomov, R. Perissinotto, and C. D. McQuaid. Role of microplankton in the diet and daily ration of Antarctic zooplankton species during austral summer. Marine Ecology Progress Series, 143:15–23, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
143/m143p015.pdf.

Ferrier-Pages:1994:SIM

[FPR94] C. Ferrier-Pages and F. Rassoulzadegan. Seasonal impact of the microzooplankton on pico-and nanoplankton growth rates in the northwest Mediterranean Sea. Marine Ecology Progress Series, 108:283–294, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/108/m108p283.pdf.

Fry:1994:BSS

[FQ94] B. Fry and R. B. Quinones. Biomass spectra and stable isotope indicators of trophic level in zooplankton of the northwest Atlantic. Marine Ecology Progress Series, 112:201–204, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/112/m112p201.pdf.

Fortier:1998:DSG

[FQV98] L. Fortier and C. Quiñonez-Velazquez. Dependence of survival on growth in larval pollock Pollachius virens and haddock Melanogrammus aeglefinus: a field study based on individual hatchdates. Marine Ecology Progress Series, 174:1–12, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
174/m174p001.pdf.

Fisher:1991:ASM

[FR91] N. S. Fisher and J. R. Reinfelder. Assimilation of selenium in the marine copepod Acartia tonsa studied with a radiotracer ratio method. Marine Ecology Progress Series, 70: 157–164, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/70/m070p157.pdf.

Franks:1992:SSA

[Fra92] P. J. S. Franks. Sink or swim, accumulation of biomass at fronts. Marine Ecology Progress Series, 82:1–12, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
82/m082p001.pdf.

France:1993:GVA

[Fra93] S. C. France. Geographic variation among three isolated populations of the hadal amphipod Hirondellea gigas (Crustacea: Amphipoda: Lysianassoidea). Marine Ecology Progress Series, 92:277–287, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/92/m092p277.pdf.

France:1994:NIC

[Fra94] R. L. France. Nitrogen isotopic composition of marine and freshwater invertebrates. Marine Ecology Progress Series, 115:205–207, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/115/m115p205.pdf.

France:1995:CEB

[Fra95] R. L. France. Carbon-13 enrichment in benthic compared to planktonic algae: foodweb implications. Marine Ecology Progress Series, 124:307–312, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/124/m124p307.pdf.

Frederick:1997:PSM

[Fre97] J. L. Frederick. Post-settlement movement of coral reef fishes and bias in survival estimates. Marine Ecology Progress Series, 150:65–74, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/150/m150p065.pdf.

Fricke:1997:LCV

[Fri97] H. Fricke. Living coelacanths: values, eco-ethics and human responsibility. Marine Ecology Progress Series, 161: 1–15, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/161/m161p001.pdf.

Fischer:1992:LSI

[FRK92] P. Fischer, K. Rademacher, and U. Kils. In situ investigations on the respiration and behaviour of the eelpout Zoarces viviparus under short-term hypoxia. Marine Ecology Progress Series, 88:181–184, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/88/m088p181.pdf.

Fraga:1999:EUP

[FRP^{+}99] F. Fraga, A. F. Ríos, F. F. Pérez, M. Estrada, and C. Marrasé. Effect of upwelling pulses on excess carbohydrate synthesis as deduced from nutrient, carbon dioxide and oxygen profiles. Marine Ecology Progress Series, 189:65–75, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
189/m189p065.pdf.

Fry:1996:SFM

[Fry96] B. Fry. ^{13}C/^{12}C fractionation by marine diatoms. Marine Ecology Progress Series, 134:283–294, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
134/m134p283.pdf.

Fleeger:1990:MRS

[FS90] J. W. Fleeger and T. C. Shirley. Meiofaunal responses to sedimentation from an Alaskan spring bloom. II. Harpacticoid population dynamics. Marine Ecology Progress Series, 59: 239–247, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/59/m059p239.pdf.

Fielding:1991:FLK

[FS91] P. J. Fielding and L. J. Seiderer. A fresh look at kelp bed phytoplankton populations in an upwelling area. Marine Ecology Progress Series, 72:167–177, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/72/m072p167.pdf.

Fukuwaka:1998:RRP

[FS98] M. Fukuwaka and T. Suzuki. Role of a riverine plume as a nursery area for chum salmon Oncorhynchus keta. Marine Ecology Progress Series, 173:289–297, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
173/m173p289.pdf.

Figueroa:1997:ESR

[FSA^{+}97] F. L. Figueroa, S. Salles, J. Aguilera, C. Jiménez, J. Mercado, B. Viñegla, A. Flores-Moya, and M. Altamirano. Effects of solar radiation on photoinhibition and pigmentation in the red alga Porphyra leucosticta. Marine Ecology Progress Series, 151:81–90, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/151/m151p081.pdf.

Feller:1992:RPB

[FSCE92] Robert J. Feller, Stephen E. Stancyk, Bruce C. Coull, and Donald G. Edwards. Recruitment of polychaetes and bivalves: long-term assessment of predictability in a soft-bottom habitat. Marine Ecology Progress Series, 87:227–238, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
87/m087p227.pdf.

Fernandez:1992:PCM

[FSdM^{+}92] Emilio Fernández, Pablo Serret, Iosu de Madariaga, Derek S. Harbour, and Anthony G. Davies. Photosynthetic carbon metabolism and biochemical composition of spring phytoplankton assemblages enclosed in microcosms: the diatom– Phaeocystis sp. succession. Marine Ecology Progress Series, 90:89–102, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/90/m090p089.pdf.

Fromentin:1997:SPT

[FSG^{+}97] J. M. Fromentin, N. C. Stenseth, J. Gjøsæter, O. N. Bjørnstad, W. Falck, and T. Johannessen. Spatial patterns of the temporal dynamics of three gadoid species along the Norwegian Skagerrak coast. Marine Ecology Progress Series, 155:209–222, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/155/m155p209.pdf.

Fromentin:1998:LTF

[FSG^{+}98] J. M. Fromentin, N. C. Stenseth, J. Gjøsæter, T. Johannessen, and B. Planque. Long-term fluctuations in cod and pollack along the Norwegian Skagerrak coast. Marine Ecology Progress Series, 162:265–278, 1998. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/162/m162p265.pdf.

Freire:1996:IMB

[FSGG96] J. Freire, M. P. Sampedro, and E. González-Gurriarán. Influence of morphometry and biomechanics on diet selection in three portunid crabs. Marine Ecology Progress Series, 137:111–121, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/137/m137p111.pdf.

Fielman:1995:VTS

[FT95] K. T. Fielman and N. M. Targett. Variation of 2,3.4tribromopyrrole and its sodium sulfamate salt in the hemichordate Saccoglossus kowalevskii. Marine Ecology Progress Series, 116:125–136, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/116/m116p125.pdf.

Forward:1997:MBC

[FTBR97] R. B. Forward, Jr., R. A. Tankersley, D. Blondel, and D. Rittschof. Metamorphosis of the blue crab Callinectes sapidus: effects of humic acids and ammonium. Marine Ecology Progress Series, 157:277–286, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/157/m157p277.pdf.

Findlay:1990:LSD

[FTGW90] R. H. Findlay, M. B. Trexler, J. B. Guckert, and D. C. White. Laboratory study of disturbance in marine sediments: response of a microbial community. Marine Ecology Progress Series, 62:121–133, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/62/m062p121.pdf.

Ford:1999:MCD

[FTP99] R. B. Ford, S. F. Thrush, and P. K. Probert. Macrobenthic colonisation of disturbances on an intertidal sandflat: the influence of season and buried algae. Marine Ecology Progress Series, 191:163–174, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/191/m191p163.pdf.

Fletcher:1994:IVT

[FTS94] W. J. Fletcher, R. J. Tregonning, and G. J. Sant. Interseasonal variation in the transport of pilchard eggs and larvae off southern Western Australia. Marine Ecology Progress Series, 111:209–224, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/111/m111p209.pdf.

Falkenhaug:1997:DSO

[FTS97] T. Falkenhaug, K. S. Tande, and T. Semenova. Diel, seasonal and ontogenetic variations in the vertical distributions of four marine copepods. Marine Ecology Progress Series, 149:105–119, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/149/m149p105.pdf.

Findlay:1990:RBM

[FTW90] R. H. Findlay, M. B. Trexler, and D. C. White. Response of a benthic microbial community to biotic disturbance. Marine Ecology Progress Series, 62:135–148, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
62/m062p135.pdf.

Fuhrman:1990:DFA

[Fuh90] J. Fuhrman. Dissolved free amino acid cycling in an estuarine outflow plume. Marine Ecology Progress Series, 66: 197–203, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/66/m066p197.pdf.

Frenette:1995:SDV

[FVDL95] J.-J. Frenette, W. F. Vincent, J. J. Dodson, and C. Lovejoy. Size-dependent variations in phytoplankton and protozoan community structure across the St. Lawrence River transition region. Marine Ecology Progress Series, 120:99–110, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
120/m120p099.pdf.

Fitz:1991:UIZ

[FW91a] H. Carlton Fitz and Richard G. Wiegert. Utilization of the intertidal zone of a salt marsh by the blue crab Callinectes sapidus density, return frequency, and feeding habits. Marine Ecology Progress Series, 76:249–260, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
76/m076p249.pdf.

Fry:1991:DSR

[FW91b] B. Fry and S. C. Wainright. Diatom sources of ^{13}C-rich carbon in marine food webs. Marine Ecology Progress Series, 76:149–157, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/76/m076p149.pdf.

Fitz:1992:LPD

[FW92] H. Carlton Fitz and Richard G. Wiegert. Local population dynamics of estuarme blue crabs: abundance, recruitment and loss. Marine Ecology Progress Series, 87:23–40, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
87/m087p023.pdf.

Findlay:1997:PBI

[FW97] R. H. Findlay and L. Watling. Prediction of benthic impact for salmon net-pens based on the balance of benthic oxygen supply and demand. Marine Ecology Progress Series, 155:147–157, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/155/m155p147.pdf.

Fukami:1999:PNP

[FWF^{+}99] K. Fukami, A. Watanabe, S. Fujita, K. Yamaoka, and T. Nishijima. Predation on naked protozoan microzooplankton by fish larvae. Marine Ecology Progress Series, 185: 285–291, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/185/m185p285.pdf.

Fielman:1999:WON

[FWWL99] K. T. Fielman, S. A. Woodin, M. D. Walla, and D. E. Lincoln. Widespread occurrence of natural halogenated organics among temperate marine infauna. Marine Ecology Progress Series, 181:1–12, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/181/m181p001.pdf.

Fourqurean:1991:PRW

[FZ91] J. W. Fourqurean and J. C. Zieman. Photosynthesis, respiration and whole plant carbon budget of the seagrass Thalassia testudinum. Marine Ecology Progress Series, 69:161– 170, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/69/m069p161.pdf.

Gismervik:1997:PSA

[GA97a] I. Gismervik and T. Andersen. Prey switching by Acartia clausi: experimental evidence and implications of intraguild predation assessed by a model. Marine Ecology Progress Series, 157:247–259, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/157/m157p247.pdf.

Gregory:1997:SSU

[GA97b] R. S. Gregory and J. T. Anderson. Substrate selection and use of protective cover by juvenile Atlantic cod Gadus morhua in inshore waters of Newfoundland. Marine Ecology Progress Series, 146:9–20, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/146/m146p009.pdf.

Goutx:1990:CECa

[GAB90a] M. Goutx, M. Acquaviva, and J.-C. Bertrand. Cellular and extracellular carbohydrates and lipids from marine bacteria during growth on soluble substrates and hydrocarbons. Marine Ecology Progress Series, 61:291–296, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
61/m061p291.pdf.

Goutx:1990:CECb

[GAB90b] M. Goutx, M. Acquaviva, and J.-C. Bertrand. Cellular and extracellular carbohydrates and lipids from marine bacteria during growth on soluble substrates and hydrocarbons. Marine Ecology Progress Series, 61:291–296, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
61/m061p291.pdf.

Gremare:1997:TCB

[GAC^{+}97] A. Grémare, J. M. Amouroux, F. Charles, A. Dinet, C. Riaux-Gobin, J. Baudart, L. Medernach, J. Y. Bodiou, G. Vétion, J. C. Colomines, and P. Albert. Temporal changes in the biochemical composition and nutritional value of the particulate organic matter available to surface deposit-feeders: a two year study. Marine Ecology Progress Series, 150:195–206, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
150/m150p195.pdf.

Gee:1992:SSM

[GAD^{+}92] J. M. Gee, M. Austen, G. De Smet, T. Ferraro, A. McEvoy, S. Moore, D. Van Gausbeki, M. Vincx, and R. M. Warwick. Soft sediment meiofauna community responses to environmental pollution gradients in the German Bight and at a drilling site off the Dutch coast. Marine Ecology Progress Series, 91:289–302, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/91/m091p289.pdf.

Gaul:1999:MGN

[GAK99] W. Gaul, A. N. Antia, and W. Koeve. Microzooplankton grazing and nitrogen supply of phytoplankton growth in the temperate and subtropical northeast Atlantic. Marine Ecology Progress Series, 189:93–104, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/189/m189p093.pdf.

Gordon:1994:HDS

[GAN^{+}94] N. Gordon, D. L. Angel, A. Neorl, N. Kress, and B. Kimor. Heterotrophic dinoflagellates with symbiotic cyanobacteria and nitrogen limitation in the Gulf of Aqaba. Marine Ecology Progress Series, 107:83–88, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/107/m107p083.pdf.

Garabetian:1991:GUC

[Gar91] F. Garabétian. ^{14}C-glucose uptake and ^{14}C-CO_{2} production in surface microlayer and surface-water samples: influence of UV and visible radiation. Marine Ecology Progress Series, 77:21–26, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/77/m077p021.pdf.

Gibson:1993:SAV

[GAR93] R. N. Gibson, A. D. Ansell, and L. Robb. Seasonal and annual variations in abundance and species composition of fish and macrocrustracean communities on a Scottish sandy beach. Marine Ecology Progress Series, 98:89–105, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
98/m098p089.pdf.

Garrabou:1998:AGI

[Gar98] J. Garrabou. Applying a geographical information system (GIS) to the study of the growth of benthic clonal organisms. Marine Ecology Progress Series, 173:227–235, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
173/m173p227.pdf.

Garrabou:1999:LHT

[Gar99a] J. Garrabou. Life-history traits of Alcyonium acaule and Parazoanthus axinellae (Cnidaria, Anthozoa), with emphasis on growth. Marine Ecology Progress Series, 178:193– 204, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/178/m178p193.pdf.

Garrison:1999:VMB

[Gar99b] L. P. Garrison. Vertical migration behavior and larval transport in brachyuran crabs. Marine Ecology Progress Series, 176:103–113, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/176/m176p103.pdf.

Gasol:1994:FAT

[Gas94] J. M. Gasol. A framework for the assessment of top-down vs. bottom-up control of heterotrophic nanoflagellate abundance. Marine Ecology Progress Series, 113:291–300, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
113/m113p291.pdf.

Gili:1993:EID

[GAV93] J.-M. Gili, P. Abelló, and R. Villanueva. Epibionts and intermoult duration in the crab Bathynectes piperitus. Marine Ecology Progress Series, 98:107–113, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
98/m098p107.pdf.

Gibson:1990:LSE

[GB90] R. N. Gibson and R. S. Batty. Lack of substratum effect on the growth and metamorphosis of larval plaice Pleuronectes platessa. Marine Ecology Progress Series, 66:219–223, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
66/m066p219.pdf.

Garrison:1991:SLS

[GB91] David L. Garrison and Kurt R. Buck. Surface-layer sea ice assemblages in Antarctic pack ice during the austral spring: environmental conditions, primary production and community structure. Marine Ecology Progress Series, 75:161–172, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
75/m075p161.pdf.

Gotceitas:1993:RPF

[GB93] V. Gotceitas and J. A. Brown. Risk of predation to fish larvae in the presence of alternative prey: effects of prey size and number. Marine Ecology Progress Series, 98:215–222, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
98/m098p215.pdf.

Gall:1995:MHM

[GB95] V. Cariou-Le Gall and G. F. Blanchard. Monthly HPLC measurements of pigment concentration from an intertidal muddy sediment of Marennes–Oléron Bay, France. Marine Ecology Progress Series, 121:171–179, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/121/m121p171.pdf.

Glassom:1997:IPGa

[GB97a] D. Glassom and G. M. Branch. Impact of predation by greater flamingos Phoenicopterus ruber on the macrofauna of two southern African lagoons. Marine Ecology Progress Series, 149:1–12, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/149/m149p001.pdf.

Glassom:1997:IPGb

[GB97b] D. Glassom and G. M. Branch. Impact of predation by greater flamingos Phoenicopterus ruber on the meiofauna, microflora, and sediment properties of two southern African lagoons. Marine Ecology Progress Series, 150:1–10, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
150/m150p001.pdf.

Guichard:1998:THH

[GB98] F. Guichard and E. Bourget. Topographic heterogeneity, hydrodynamics, and benthic community structure: a scaledependent cascade. Marine Ecology Progress Series, 171: 59–70, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/171/m171p059.pdf.

Gardner:1996:EHM

[GBA^{+}96] W. S. Gardner, R. Benner, R. M. W. Amon, J. B. Cotner, Jr., J. F. Cavaletto, and J. R. Johnson. Effects of high-molecular-weight dissolved organic matter on nitrogen dynamics in the Mississippi River plume. Marine Ecology Progress Series, 133:287–297, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/133/m133p287.pdf.

Guarini:1998:DSP

[GBB^{+}98] J. M. Guarini, G. F. Blanchard, C. Bacher, P. Gros, P. Riera, P. Richard, D. Gouleau, R. Galois, J. Prou, and P. G. Sauriau. Dynamics of spatial patterns of microphytobenthic biomass: inferences from a geostatistical analysis of two comprehensive surveys in Marennes–Oléron Bay (France). Marine Ecology Progress Series, 166:131–141, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/166/m166p131.pdf.

Galgani:1992:EVC

[GBC92] F. Galgani, G. Bocquene, and Y. Cadiou. Evidence of variation in cholinesterase activity in fish along a pollution gradient in the North Sea. Marine Ecology Progress Series, 91: 77–82, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/91/m091p077.pdf.

Guarini:1997:MMS

[GBGH97] J. M. Guarini, G. F. Blanchard, P. Gros, and S. J. Harrison. Modelling the mud surface temperature on intertidal flats to investigate the spatio-temporal dynamics of the benthic microalgal photosynthetic capacity. Marine Ecology Progress Series, 153:25–36, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/153/m153p025.pdf.

Gundersen:1996:FIL

[GBH96] K. Gundersen, G. Bratbak, and M. Heldal. Factors influencing the loss of bacteria in preserved seawater samples. Marine Ecology Progress Series, 137:305–310, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
137/m137p305.pdf.

Guzman:1994:IRG

[GBJ94] H. M. Guzmán, K. A. Burns, and J. B. C. Jackson. Injury, regeneration and growth of Caribbean reef corals after a major oil spill in Panama. Marine Ecology Progress Series, 105:231–241, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/105/m105p231.pdf.

Goss:1997:SSI

[GBP^{+}97] C. Goss, D. G. Bone, J. M. Peck, I. Everson, G. L. Hunt, Jr., and A. W. A. Murray. Small-scale interactions between prions Pachyptila spp. and their zooplankton prey at an inshore site near Bird Island, South Georgia. Marine Ecology Progress Series, 154:41–51, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/154/m154p041.pdf.

Gregoire:1996:DMP

[GBV96] Y. Grégoire, E. Bourget, and J. L. Verrette. Deposition of mimics of planktonic invertebrate larvae on simple and complex substrata in flume flows. Marine Ecology Progress Series, 135:89–100, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/135/m135p089.pdf.

Garrison:1993:WES

[GC93] D. L. Garrison and A. R. Close. Winter ecology of the sea ice biota in Weddell Sea pack ice. Marine Ecology Progress Series, 96:17–31, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/96/m096p017.pdf.

Gibson:1995:DVP

[GC95a] G. D. Gibson and F. S. Chia. Developmental variability in the poecilogonous opisthobranch Haminaea callidegenita: life-history traits and effects of environmental parameters. Marine Ecology Progress Series, 121:139–155, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
121/m121p139.pdf.

Gosselin:1995:DDE

[GC95b] L. A. Gosselin and F. S. Chia. Distribution and dispersal of early juvenile snails: effectiveness of intertidal microhabitats as refuges and food sources. Marine Ecology Progress Series, 128:213–223, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/128/m128p213.pdf.

Green:1995:ASS

[GCC95] A. S. Green, G. T. Chandler, and B. C. Coull. Age-specific survival analysis of an infaunal meiobenthic harpacticoid copepod, Amphiascus tenuiremis. Marine Ecology Progress Series, 129:107–112, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/129/m129p107.pdf.

Glibert:1995:DWS

[GCF^{+}95] P. M. Glibert, D. J. Conley, T. R. Fisher, L. W. Harding, Jr., and T. C. Malone. Dynamics of the 1990 winter/ spring bloom in Chesapeake Bay. Marine Ecology Progress Series, 122:27–43, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/122/m122p027.pdf.

Garthe:1996:ADC

[GCF96] S. Garthe, K. Camphuysen, and R. W. Furness. Amounts of discards by commercial fisheries and their significance as food for seabirds in the North Sea. Marine Ecology Progress Series, 136:1–11, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/136/m136p001.pdf.

Guzman:1990:CMA

[GCGR90] H. M. Guzmán, J. Cortés, P. W. Glynn, and R. H. Richmond. Coral mortality associated with dino-flagellate blooms in the eastern Pacific (Costa Rica and Panama). Marine Ecology Progress Series, 60:299–303, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/60/m060p299.pdf.

Giere:1990:RGA

[GCGS90] O. Giere, N. M. Conway, G. Gastrock, and C. Schmidt. Regulation of gutless annelid ecology by endosymbiotic bacteria. Marine Ecology Progress Series, 68:287–299, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
68/m068p287.pdf.

Gardner:1993:CMN

[GCH93] W. S. Gardner, J. B. Cotner, Jr., and L. R. Herche. Chromatographic measurement of nitrogen mineralization rates in marine coastal waters with ^{l5}N. Marine Ecology Progress Series, 93:65–73, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/93/m093p065.pdf.

George:1995:DCO

[GCKW95] S. G. George, J. S. Christiansen, B. Killie, and J. Wright. Dietary crude oil exposure during sexual maturation induces hepatic mixed function oxygenase (CYP1A) activity at very low environmental temperatures in polar cod Boreogadus saida. Marine Ecology Progress Series, 122:307–312, 1995.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
122/m122p307.pdf.

Golde:1994:RMS

[GCLS94] H. M. Golde, M. P. Crosby, C. R. Lovell, and S. E. Stancyk. Reduction of marine sediment oxygen uptake: pretreatment for the determination of infaunal respiration rates. Marine Ecology Progress Series, 111:203–207, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
111/m111p203.pdf.

Gaspar:1998:ITD

[GCM98] M. B. Gaspar, M. Castro, and C. C. Monteiro. Influence of tow duration and tooth length on the number of damaged razor clams Ensis siliqua. Marine Ecology Progress Series, 169:303–305, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/169/m169p303.pdf.

Grijalva-Chon:1994:MDA

[GCNSNd94] J. M. Grijalva-Chon, K. Numachi, O. Sosa-Nishizaki, and J. de la Rosa-Velez. Mitochondrial DNA analysis of North Pacific sword-fish Xiphias gladius population structure. Marine Ecology Progress Series, 115:15–19, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/115/m115p015.pdf.

Garcia:1998:DLG

[GCR98] A. García, D. Cortés, and T. Ramírez. Daily larval growth and RNA and DNA content of the NW Mediterranean anchovy Engraulis encrasicolus and their relations to the environment. Marine Ecology Progress Series, 166:237–245, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
166/m166p237.pdf.

Gray:1990:DIE

[GCWH90] J. S. Gray, K. R. Clarke, R. M. Warwick, and G. Hobbs. Detection of initial effects of pollution on marine benthos: an example from the ekofisk and eldfisk oilfields, North Sea.

Marine Ecology Progress Series, 66:285–299, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
66/m066p285.pdf.

Goldman:1990:DPS

[GD90] J. C. Goldman and M. R. Dennett. Dynamics of prey selection by an omnivorous flagellate. Marine Ecology Progress Series, 59:183–194, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/59/m059p183.pdf.

Gorostiaga:1996:CSB

[GD96] J. M. Gorostiaga and I. Díez. Changes in the sublittoral benthic marine macroalgae in the polluted area of Abra de Bilbao and proximal coast (Northern Spain). Marine Ecology Progress Series, 130:157–167, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/130/m130p157.pdf.

Gomez:1999:LHS

[GD99] J. Gómez and O. Defeo. Life history of the sandhopper Pseudorchestoidea brasiliensis (Amphipoda) in sandy beaches with contrasting morphodynamics. Marine Ecology Progress Series, 182:209–220, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/182/m182p209.pdf.

Gasol:1995:AVI

[GdGMD95] J. M. Gasol, P. A. del Giorgio, R. Massana, and C. M. Duarte. Active versus inactive bacteria: size-dependence in a coastal marine plankton community. Marine Ecology Progress Series, 128:91–97, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/128/m128p091.pdf.

Golikov:1990:TAM

[GDMS90] A. N. Golikov, M. A. Dolgolenko, N. V. Maximovich, and O. A. Scarlato. Theoretical approaches to marine biogeography. Marine Ecology Progress Series, 63:289–301, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
63/m063p289.pdf.

Gasol:1998:DVB

[GDP^{+}98] J. M. Gasol, M. D. Doval, J. Pinhassi, J. I. Calderón-Paz, N. Guixa-Boixareu, D. Vaqué, and C. Pedrós-Alió. Diel variations in bacterial heterotrophic activity and growth in the northwestern Mediterranean Sea. Marine Ecology Progress Series, 164:107–124, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/164/m164p107.pdf.

Garcia:1998:DFA

[GDvS98] C. B. García, L. O. Duarte, and D. von Schiller. Demersal fish assemblages of the Gulf of Salamanca, Colombia (southern Caribbean Sea). Marine Ecology Progress Series, 174: 13–25, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/174/m174p013.pdf.

Griffith:1990:MAS

[GDW90] P. C. Griffith, D. J. Douglas, and S. C. Wainright. Metabolic activity of size-fractionated microbial plankton in estuarine, near-shore, and continental shelf waters of Georgia. Marine Ecology Progress Series, 59:263–270, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
59/m059p263.pdf.

Gendron:1992:PSD

[Gen92] D. Gendron. Population structure of daytime surface swarms of Nyctiphanes simplex (Crustacea: Euphausiacea) in the Gulf of California, Mexico. Marine Ecology Progress Series, 87:1–6, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/87/m087p001.pdf.

George:1994:LEA

[Geo94] S. B. George. The Leptasterias (Echinodermata: Asteroidea) species complex: variation in reproductive investment. Marine Ecology Progress Series, 109:95–98, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
109/m109p095.pdf.

Gerhart:1990:FGP

[Ger90] D. J. Gerhart. Fouling and gastropod predation: consequences of grazing for a tropical octocoral. Marine Ecology Progress Series, 62:103–108, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/62/m062p103.pdf.

Grottoli-Everett:1997:FPS

[GEW97] A. G. Grottoli-Everett and G. M. Wellington. Fish predation on the scleractinian coral Madracis mirabilis controls its depth distribution in the Florida Keys, USA. Marine Ecology Progress Series, 160:291–293, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/160/m160p291.pdf.

Guary:1990:EST

[GF90] J. C. Guary and S. W. Fowler. Experimental study of the transfer of transuranium nuclides in marine decapod crustaceans. Marine Ecology Progress Series, 60:253–270, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
60/m060p253.pdf.

Grimes:1991:SDA

[GF91] C. B. Grimes and J. H. Finucane. Spatial distribution and abundance of larval and juvenile fish, chlorophyll and macrozooplankton around the Mississippi River discharge plume, and the role of the plume in fish recruitment. Marine Ecology Progress Series, 75:109–119, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/75/m075p109.pdf.

Goshima:1994:DAC

[GF94] S. Goshima and H. Fujiwara. Distribution and abundance of cultured scallop Patinopecten yessoensis in extensive sea beds as assessed by underwater camera. Marine Ecology Progress Series, 110:151–158, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/110/m110p151.pdf.

Garrabou:1995:SDS

[GF95] J. Garrabou and J. Flos. A simple diffusion-sedimentation model to explain planktonic gradients within a NW Mediterranean submarine cave. Marine Ecology Progress Series, 123: 273–280, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/123/m123p273.pdf.

Gregg:1998:GSP

[GF98] C. S. Gregg and J. W. Fleeger. Grass shrimp Palaemonetes pugio predation on sediment- and stem-dwelling meiofauna: field and laboratory experiments. Marine Ecology Progress Series, 175:77–86, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/175/m175p077.pdf.

Glud:1999:ICI

[GF99] R. N. Glud and T. Fenchel. The importance of ciliates for interstitial solute transport in benthic communities. Marine Ecology Progress Series, 186:87–93, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/186/m186p087.pdf.

Glud:1996:IRP

[GFH96] R. N. Glud, S. Forster, and M. Huettel. Influence of radial pressure gradients on solute exchange in stirred benthic chambers. Marine Ecology Progress Series, 141:303–311, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
141/m141p303.pdf.

Gilbert:1992:FEM

[GFPD92] M. Gilbert, L. Fortier, D. Ponton, and R. Drolet. Feeding ecology of marine fish larvae across the great whale river plume in seasonally ice-covered southeastern Hudson Bay. Marine Ecology Progress Series, 84:19–30, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/84/m084p019.pdf.

Gonzalez:1994:STS

[GGB94] H. E. González, S. R. Gonzalez, and G.-J. A. Brummer. Short-term sedimentation pattern of zooplankton, faeces and microplankton at a permanent station in the Bjornafjorden (Norway) during April–May 1992. Marine Ecology Progress Series, 105:31–45, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/105/m105p031.pdf.

Gibbons:1999:CCN

[GGB^{+}99] M. J. Gibbons, N. Gugushe, A. J. Boyd, L. J. Shannon, and B. A. Mitchell-Innes. Changes in the composition of the non-copepod zooplankton assemblage in st Helena Bay (southern Benguela ecosystem) during a six day drogue study. Marine Ecology Progress Series, 180:111–120, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
180/m180p111.pdf.

Gisselson:1999:UCC

[GGC99] L. Å. Gisselson, E. Granéli, and P. Carlsson. Using cell cycle analysis to estimate in situ growth rate of the dinoflagellate Dinophysis acuminata: drawbacks of the DNA quantification method. Marine Ecology Progress Series, 184:55–62, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
184/m184p055.pdf.

Gunther:1999:BAS

[GGD99] S. Günther, M. Gleitz, and G. S. Dieckmann. Biogeochemistry of Antarctic sea ice: a case study on platelet ice layers at Drescher Inlet, Weddell Sea. Marine Ecology Progress Series, 177:1–13, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/177/m177p001.pdf.

Gomez-Gutierrez:1996:GPE

[GGDLE96] J. Gómez-Gutiérrez, R. De Silva-Dávila, and B. Lavaniegos-Espejo. Growth production of the euphausiid Nyctiphanes simplex on the coastal shelf off Bahía Magdalena, Baja California Sur, México. Marine Ecology Progress Series, 138: 309–314, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/138/m138p309.pdf.

Garthe:1999:SAF

[GGF99] S. Garthe, D. Grémillet, and R. F. W. Furness. At-sea-activity and foraging efficiency in chick-rearing northern gannets Sula bassana: a case study in Shetland. Marine Ecology Progress Series, 185:93–99, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/185/m185p093.pdf.

Genin:1995:CCO

[GGH95] A. Genin, G. Gal, and L. Haury. Copepod carcasses in the ocean. II. Near coral reefs. Marine Ecology Progress Series, 123:65–71, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/123/m123p065.pdf.

Glud:1999:BLS

[GGH99] R. N. Glud, J. K. Gundersen, and O. Holby. Benthic in situ respiration in the upwelling area off central Chile. Marine Ecology Progress Series, 186:9–18, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/186/m186p009.pdf.

Gonzalez-Gil:1998:DQA

[GGKJ^{+}98] S. González-Gil, B. A. Keafer, R. V. M. Jovine, A. Aguilera, S. Lu, and D. M. Anderson. Detection and quantification of alkaline phosphatase in single cells of phosphorusstarved marine phytoplankton. Marine Ecology Progress Series, 164:21–35, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/164/m164p021.pdf.

Gomez-Gutierrez:1995:CSE

[GGPGG95] J. Gómez-Gutiérrez, R. Palomares-García, and D. Gendron. Community structure of the euphausiid populations along the west coast of Baja California, Mexico, during the weak ENSO 1986–1987. Marine Ecology Progress Series, 120:41–51, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
120/m120p041.pdf.

Gilmer:1991:DLH

[GH91] R. W. Gilmer and G. R. Harbison. Diet of Limacina helicina (Gastropoda: Thecosomata) in Arctic waters in midsummer. Marine Ecology Progress Series, 77:125–134, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
77/m077p125.pdf.

Grebmeier:1992:SFB

[GH92] J. M. Grebmeier and N. M. Harrison. Seabird feeding on benthic amphipods facilitated by gray whale activity in the northern Bering Sea. Marine Ecology Progress Series, 80: 125–133, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/80/m080p125.pdf.

Gerdol:1993:EAC

[GH93] V. Gerdol and R. G. Hughes. Effect of the amphipod Corophium volutator on the colonisation of mud by the halophyte Salicotrm europaea. Marine Ecology Progress Series, 97:61–69, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/97/m097p061.pdf.

Garthe:1994:DSF

[GH94a] S. Garthe and O. Hüppop. Distribution of ship-following seabirds and their utilization of discards in the North Sea in summer. Marine Ecology Progress Series, 106:1–9, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
106/m106p001.pdf.

Gerdol:1994:ECV

[GH94b] V. Gerdol and R. G. Hughes. Effect of Corophium volutator on the abundance of benthic diatoms, bacteria and sediment stability in two estuaries in southeastern England. Marine Ecology Progress Series, 114:109–115, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
114/m114p109.pdf.

Gerdol:1994:FBD

[GH94c] V. Gerdol and R. G. Hughes. Feeding behaviour and diet of Corophium volutator in an estuary in southeastern England. Marine Ecology Progress Series, 114:103–108, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
114/m114p103.pdf.

Goldman:1992:CCT

[GHD92] J. C. Goldman, D. A. Hansell, and M. R. Dennett. Chemical characterization of three large oceanic diatoms: potential impact on water column chemistry. Marine Ecology Progress Series, 88:257–270, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/88/m088p257.pdf.

Geertz-Hansen:1994:FIF

[GHED^{+}94] O. Geertz-Hansen, S. Enriquez, C. M. Duarte, S. Agusti, D. Vaque, and B. Vidondo. Functional implications of the form of Codium bursa, a balloon-like Mediterranean macroalga. Marine Ecology Progress Series, 108:153–160, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/108/m108p153.pdf.

Glud:1998:BME

[GHHC98] R. N. Glud, O. Holby, F. Hoffmann, and D. E. Canfield. Benthic mineralization and exchange in Arctic sediments (Svalbard, Norway). Marine Ecology Progress Series, 173:237– 251, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/173/m173p237.pdf.

Graham:1997:PDG

[GHL^{+}97] M. H. Graham, C. Harrold, S. Lisin, K. Light, J. M. Watanabe, and M. S. Foster. Population dynamics of giant kelp Macrocystis pyrifera along a wave exposure gradient. Marine Ecology Progress Series, 148:269–279, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
148/m148p269.pdf.

Gallego:1996:EIS

[GHMC96] A. Gallego, M. R. Heath, E. McKenzie, and L. H. Cargill. Environmentally induced short-term variability in the growth rates of larval herring. Marine Ecology Progress Series, 137: 11–23, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/137/m137p011.pdf.

Goes:1994:EUB

[GHTH94] J. I. Goes, N. Handa, S. Taguchi, and T. Hama. Effect of UV-B radiation on the fatty acid composition of the marine phytoplankter Tetraselmis sp.. relationship to cellular pigments. Marine Ecology Progress Series, 114:259–274, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
114/m114p259.pdf.

Gutierrez:1999:RHB

[GI99] J. Gutiérrez and O. Iribarne. Role of Holocene beds of the stout razor clam Tagelus plebeius in structuring present benthic communities. Marine Ecology Progress Series, 185:213– 228, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/185/m185p213.pdf.

Gibbons:1994:DVM

[Gib94] M. J. Gibbons. Diet vertical migration and feeding of Sagitta friderici and Sagitta tasmanica in the southern Benguela upwelling region, with a comment on the structure of the guild of primary carnivores. Marine Ecology Progress Series, 111:225–240, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/111/m111p225.pdf.

Gillanders:1997:CGR

[Gil97] B. M. Gillanders. Comparison of growth rates between estuarine and coastal reef populations of Achoerodus viridis (Pisces: Labridae). Marine Ecology Progress Series, 146: 283–287, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/146/m146p283.pdf.

Grehan:1996:CDS

[GJ96a] A. J. Grehan and S. K. Juniper. Clam distribution and subsurface hydrothermal processes at Chowder Hill (Middle Valley), Juan de Fuca Ridge. Marine Ecology Progress Series, 130:105–115, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/130/m130p105.pdf.

Guerrero:1996:PMNa

[GJ96b] M. A. Guerrero and R. D. Jones. Photoinhibition of marine nitrifying bacteria. I. Wavelength-dependent response. Marine Ecology Progress Series, 141:183–192, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
141/m141p183.pdf.

Guerrero:1996:PMNb

[GJ96c] M. A. Guerrero and R. D. Jones. Photoinhibition of marine nitrifying bacteria. II. Dark recovery after monochromatic or polychromatic irradiation. Marine Ecology Progress Series, 141:193–198, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/141/m141p193.pdf.

Gallegos:1997:SPFa

[GJ97a] C. L. Gallegos and T. E. Jordan. Seasonal progression of factors limiting phytoplankton pigment biomass in the Rhode River estuary, Maryland (USA). I. Controls on phytoplankton growth. Marine Ecology Progress Series, 161:185–198, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
161/m161p185.pdf.

Gallegos:1997:SPFb

[GJ97b] C. L. Gallegos and T. E. Jordan. Seasonal progression of factors limiting phytoplankton pigment biomass in the Rhode River estuary, Maryland (USA). II. Modeling N versus P limitation. Marine Ecology Progress Series, 161:199–212, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
161/m161p199.pdf.

Gallegos:1997:IVS

[GJC97] C. L. Gallegos, T. E. Jordan, and D. L. Correll. Interannual variability in spring bloom timing and magnitude in the Rhode River, Maryland, USA: observations and modeling. Marine Ecology Progress Series, 154:27–40, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
154/m154p027.pdf.

Gamenick:1996:HSS

[GJVG96] I. Gamenick, A. Jahn, K. Vopel, and O. Giere. Hypoxia and sulphide as structuring factors in a macrozoobenthic community on the Baltic Sea shore: colonisation studies and tolerance experiments. Marine Ecology Progress Series, 144: 73–85, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/144/m144p073.pdf.

Gillanders:1996:EOM

[GK96a] B. M. Gillanders and M. J. Kingsford. Elements in otoliths may elucidate the contribution of estuarine recruitment to sustaining coastal reef populations of a temperate reef fish. Marine Ecology Progress Series, 141:13–20, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
141/m141p013.pdf.

Griffiths:1996:RBS

[GK96b] C. L. Griffiths and D. W. Klumpp. Relationships between size, mantle area and zooxanthellae numbers in five species of giant clam (Tridacnidae). Marine Ecology Progress Series, 137:139–147, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/137/m137p139.pdf.

Giray:1997:PDD

[GK97] C. Giray and G. M. King. Predator deterrence and 2,4dibromophenol conservation by the enteropneusts Saccoglossus bromophenolosus and Protoglossus graveolens. Marine Ecology Progress Series, 159:229–238, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
159/m159p229.pdf.

Guinet:1997:FBS

[GKB^{+}97] C. Guinet, M. Koudil, C. A. Bost, J. P. Durbec, J. Y. Georges, M. C. Mouchot, and P. Jouventin. Foraging behaviour of satellite-tracked king penguins in relation to seasurface temperatures obtained by satellite telemetry at Crozet Archipelago, a study during three austral summers. Marine Ecology Progress Series, 150:11–20, 1997. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/150/m150p011.pdf.

Goto:1999:IEO

[GKMT99] N. Goto, T. Kawamura, O. Mitamura, and H. Terai. Importance of extracellular organic carbon production in the total primary production by tidal-flat diatoms in comparison to phytoplankton. Marine Ecology Progress Series, 190: 289–295, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/190/m190p289.pdf.

Gleitz:1996:CAG

[GKRD96] M. Gleitz, H. Kukert, U. Riebesell, and G. S. Dieckmann. Carbon acquisition and growth of Antarctic sea ice diatoms in closed bottle incubations. Marine Ecology Progress Series, 135:169–177, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/135/m135p169.pdf.

Glasby:1999:IES

[Gla99] T. M. Glasby. Interactive effects of shading and proximity to the seafloor on the development of subtidal epibiotic assemblages. Marine Ecology Progress Series, 190:113–124, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
190/m190p113.pdf.

Goksoyr:1992:DCP

[GLBF92] Anders Goksoyr, Håvard E. Larsen, Sonja Blom, and Lars Förlin. Detection of cytochrome P450 1A1 in North Sea dab liver and kidney. Marine Ecology Progress Series, 91: 83–88, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/91/m091p083.pdf.

Gorsky:1991:MTT

[GLD^{+}91] G. Gorsky, N. Lins da Silva, S. Dallot, Ph. Laval, J. C. Braconnot, and L. Prieur. Midwater tunicates: are they related to the permanent front of the Ligurian Sea (NW Mediterranean)? Marine Ecology Progress Series, 74:195–204, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
74/m074p195.pdf.

Goericke:1998:ECM

[GM98a] R. Goericke and J. P. Montoya. Estimating the contribution of microalgal taxa to chlorophyll-a in the field — variations of pigment ratios under nutrient- and light-limited growth. Marine Ecology Progress Series, 169:97–112, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
169/m169p097.pdf.

Gonzalez:1998:DLC

[GM98b] A. Gonzalez and V. H. Marín. Distribution and life cycle of Calanus chilensis and Centropages brachiatus (Copepoda) in Chilean coastal waters: a GIS approach. Marine Ecology Progress Series, 165:109–117, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/165/m165p109.pdf.

Garrison:1999:AVD

[GM99] L. P. Garrison and J. A. Morgan. Abundance and vertical distribution of drifting, post-larval Macoma spp. (Bivalvia: Tellinidae) in the York River, Virginia, USA. Marine Ecology Progress Series, 182:175–185, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/182/m182p175.pdf.

Gale:1991:ECE

[GMC91] L. D. Gale, J. J. Manzi, and M. P. Crosby. Energetic costs to the Eastern oyster Crassostrea virginica due to recent parasitism by the ectoparasitic gastropod Boonea impress. Marine Ecology Progress Series, 79:89–98, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/79/m079p089.pdf.

Granadeiro:1998:DFE

[GMF98] J. P. Granadeiro, L. R. Monteiro, and R. W. Furness. Diet and feeding ecology of Cory’s shearwater Calonectris diomedea in the Azores, north-east Atlantic. Marine Ecology Progress Series, 166:267–276, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/166/m166p267.pdf.

Glibert:1992:NRG

[GMG^{+}92] P. M. Glibert, C. A. Miller, C. Garside, M. R. Roman, and G. B. McManus. NH_{4}^{+} regeneration and grazing: interdependent processes n size-fractionated 15 NH_{4}^{+} experiments. Marine Ecology Progress Series, 82:65–74, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
82/m082p065.pdf.

Geider:1997:DMP

[GMK97] R. J. Geider, H. L. MacIntyre, and T. M. Kana. Dynamic model of phytoplankton growth and acclimation: responses of the balanced growth rate and the chlorophyll-a: carbon ratio to light, nutrient-limitation and temperature. Marine Ecology Progress Series, 148:187–200, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/148/m148p187.pdf.

Gallegos:1993:BDT

[GMMD93] M. E. Gallegos, M. Merino, N. Marba, and C. M. Durate. Biomass and dynamics of Thalassia testudinum in the Mexican Caribbean: elucidating rhizome growth. Marine Ecology Progress Series, 95:185–192, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/95/m095p185.pdf.

Gaudian:1995:ESC

[GMO95] G. Gaudian, P. A. H. Medley, and R. F. G. Ormond. Estimation of the size of a coral reef fish population. Marine Ecology Progress Series, 122:107–113, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/122/m122p107.pdf.

Gallegos:1994:GPD

[GMR^{+}94] M. E. Gallegos, M. Merino, A. Rodriguez, N. Marba, and C. M. Duarte. Growth patterns and demography of pioneer Caribbean seagrasses Halodule wrightii and Syringodium filiforme. Marine Ecology Progress Series, 109:99–104, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
109/m109p099.pdf.

Giannoulaki:1999:ALV

[GMT99] M. Giannoulaki, A. Machias, and N. Tsimenides. Ambient luminance and vertical migration of the sardine Sardina pilchardus. Marine Ecology Progress Series, 178:29–38, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
178/m178p029.pdf.

Gambi:1990:FOF

[GNJ90] M. C. Gambi, A. R. M. Nowell, and P. A. Jumars. Flume observations on flow dynamics in Zostera marina (eelgrass) beds. Marine Ecology Progress Series, 61:159–169, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
61/m061p159.pdf.

Goldman:1999:ICA

[Gol99] J. C. Goldman. Inorganic carbon availability and the growth of large marine diatoms. Marine Ecology Progress Series, 180: 81–91, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/180/m180p081.pdf.

Gomez:1995:DPA

[Góm95] J. G. Gómez. Distribution patterns, abundance and population dynamics of the euphausiids Nyctiphanes simplex and Euphausia eximia off the west coast of Baja California, Mexico. Marine Ecology Progress Series, 119:63–76, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
119/m119p063.pdf.

Gonzalez:1992:DAM

[Gon92] H. E. González. Distribution and abundance of minipellets around the Antarctic peninsula. Implications for protistan feeding behaviour. Marine Ecology Progress Series, 90: 223–236, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/90/m090p223.pdf.

Gosselin:1997:ETD

[Gos97] L. A. Gosselin. An ecological transition during juvenile life in a marine snail. Marine Ecology Progress Series, 157:185– 194, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/157/m157p185.pdf.

Gowing:1993:LVL

[Gow93] M. M. Gowing. Large virus-like particles from vacuoles of phaeodarian radiolarians and from other marine samples. Marine Ecology Progress Series, 101:33–43, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
101/m101p033.pdf.

Gutt:1990:DAT

[GP90] J. Gutt and D. Piepenburg. Dense aggregations of three deep-sea holothurians in the southern Weddell Sea, Antarctica. Marine Ecology Progress Series, 68:277–285, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
68/m068p277.pdf.

Grizzle:1991:EOE

[GP91] R. E. Grizzle and C. A. Penniman. Effects of organic enrichment on estuarine macrofaunal benthos: a comparison of sediment profile imaging and traditional methods. Marine Ecology Progress Series, 74:249–262, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/74/m074p249.pdf.

Grutter:1998:IIR

[GP98] A. S. Grutter and R. Poulin. Intraspecific and interspecific relationships between host size and the abundance of parasitic larval gnathiid isopods on coral reef fishes. Marine Ecology Progress Series, 164:263–271, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/164/m164p263.pdf.

Garcia-Pichel:1996:PUR

[GPB96] F. Garcia-Pichel and B. M. Bebout. Penetration of ultraviolet radiation into shallow water sediments: high exposure for photosynthetic communities. Marine Ecology Progress Series, 131:257–262, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/131/m131p257.pdf.

Gibson:1998:DMJ

[GPB^{+}98] R. N. Gibson, L. Pihl, M. T. Burrows, J. Modin, H. Wennhage, and L. A. Nickell. Diel movements of juvenile plaice Pleuronectes platessa in relation to predators, competitors, food availability and abiotic factors on a microtidal nursery ground. Marine Ecology Progress Series, 165:145–159, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/165/m165p145.pdf.

Gattuso:1996:CFC

[GPD^{+}96] J. P. Gattuso, M. Pichon, B. Delesalle, C. Canon, and M. Frankignoulle. Carbon fluxes in coral reefs. I. Lagrangian measurement of community metabolism and resulting air-sea CO_{2} disequilibrium. Marine Ecology Progress Series, 145: 109–121, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/145/m145p109.pdf.

Gattuso:1993:CMA

[GPDF93] J.-P. Gattuso, M. Pichon, B. Delesalle, and M. Frankignoulle. Community metabolism and air-sea CO_{2} fluxes in a coral reef ecosystem (Moorea, French Polynesia). Marine Ecology Progress Series, 96:259–267, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/96/m096p259.pdf.

Gattuso:1995:BCA

[GPF95] J. P. Gattuso, M. Pichon, and M. Frankignoulle. Biological control of air-sea CO_{2} fluxes: effect of photosynthetic and calcifying marine organisms and ecosystems. Marine Ecology Progress Series, 129:307–312, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/129/m129p307.pdf.

Gotceitas:1996:EIM

[GPLB96] V. Gotceitas, V. Puvanendran, L. L. Leader, and J. A. Brown. An experimental investigation of the ’match/mismatch’ hypothesis using larval Atlantic cod. Marine Ecology Progress Series, 130:29–37, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/130/m130p029.pdf.

Geffen:1998:MCF

[GPP98] A. J. Geffen, N. J. G. Pearce, and W. T. Perkins. Metal concentrations in fish otoliths in relation to body composition after laboratory exposure to mercury and lead. Marine Ecology Progress Series, 165:235–245, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/165/m165p235.pdf.

Gebruk:1993:FSB

[GPS93] A. V. Gebruk, N. V. Pimenov, and A. S. Sawichev. Feeding specialization of bresiliid shrimps in the TAG site hydrothermal community. Marine Ecology Progress Series, 98: 247–253, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/98/m098p247.pdf.

Gray:1997:CDS

[GPU^{+}97] J. S. Gray, G. C. B. Poore, K. I. Ugland, R. S. Wilson, F. Olsgard, and Ø. Johannessen. Coastal and deep-sea benthic diversities compared. Marine Ecology Progress Series, 159: 97–103, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/159/m159p097.pdf.

Gosselin:1996:EPS

[GQ96] L. A. Gosselin and P. Y. Qian. Early post-settlement mortality of an intertidal barnacle: a critical period for survival. Marine Ecology Progress Series, 135:69–75, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
135/m135p069.pdf.

Gosselin:1997:CBA

[GQ97a] L. A. Gosselin and P. Y. Qian. Can bacterivory alone sustain larval development in the polychaete Hydroides elegans and the barnacle Balanus amphitrite? Marine Ecology Progress Series, 161:93–101, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/161/m161p093.pdf.

Gosselin:1997:JMB

[GQ97b] L. A. Gosselin and P. Y. Qian. Juvenile mortality in benthic marine invertebrates. Marine Ecology Progress Series, 146:265–282, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/146/m146p265.pdf.

Grossmann:1991:IAM

[GR91] S. Grossmann and W. Reichardt. Impact of Arenicola marina on bacteria in intertidal sediments. Marine Ecology Progress Series, 77:85–93, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/77/m077p085.pdf.

Goericke:1993:CBD

[GR93] R. Goericke and D. J. Repeta. Chlorophylls a and b and divinyl chlorophylls a and b in the open subtropical North Atlantic Ocean. Marine Ecology Progress Series, 101:307– 313, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/101/m101p307.pdf.

Graneli:1994:ECV

[GR94] E. Granéli and L. Risinger. Effects of cobalt and vitamin B_{12} on the growth of Chrysochromulina polylepis (Prymnesiophyceae). Marine Ecology Progress Series, 113:177–183, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
113/m113p177.pdf.

Grainger:1991:EAS

[Gra91a] E. H. Grainger. Exploitation of Arctic sea ice by epibenthic copepods. Marine Ecology Progress Series, 77:119–124, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
77/m077p119.pdf.

Gray:1991:DPP

[Gra91b] C. A. Gray. Demographic patterns of the palaemonid prawn Macrobrachium intermedium in southeastern Australia: spatial heterogeneity and the effects of species of seagrass. Marine Ecology Progress Series, 75:239–249, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
75/m075p239.pdf.

Gray:1991:TVD

[Gra91c] C. A. Gray. Temporal variability in the demography of the palaemonid prawn Macrobrachium intermedium in two seagrasses. Marine Ecology Progress Series, 75:227–237, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
75/m075p227.pdf.

Gray:1994:DSS

[Gra94] J. S. Gray. Is deep-sea species diversity really so high? Species diversity of the Norwegian continental shelf. Marine Ecology Progress Series, 112:205–209, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/112/m112p205.pdf.

Gradinger:1996:OAB

[Gra96a] R. Gradinger. Occurrence of an algal bloom under Arctic pack ice. Marine Ecology Progress Series, 131:301–305, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
131/m131p301.pdf.

Gray:1996:ISS

[Gra96b] C. A. Gray. Intrusions of surface sewage plumes into continental shelf waters: interactions with larval and presettlement juvenile fishes. Marine Ecology Progress Series, 139: 31–45, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/139/m139p031.pdf.

Gibson:1996:TDL

[GRBA96] R. N. Gibson, L. Robb, M. T. Burrows, and A. D. Ansell. Tidal, diel and longer term changes in the distribution of fishes on a Scottish sandy beach. Marine Ecology Progress Series, 130:1–17, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/130/m130p001.pdf.

Gannon:1997:FEL

[GRC^{+}97] D. P. Gannon, A. J. Read, J. E. Craddock, K. M. Fristrup, and J. R. Nicolas. Feeding ecology of long-finned pilot whales Globicephala melas in the western North Atlantic. Marine Ecology Progress Series, 148:1–10, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/148/m148p001.pdf.

Green:1996:STO

[Gre96] A. L. Green. Spatial, temporal and ontogenetic patterns of habitat use by coral reef fishes (family Labridae). Marine Ecology Progress Series, 133:1–11, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/133/m133p001.pdf.

Glud:1996:PON

[GRGK96] R. N. Glud, N. B. Ramsing, J. K. Gundersen, and I. Klimant. Planar optrodes: a new tool for fine scale measurements of two-dimensional O 2 distribution in benthic communities. Marine Ecology Progress Series, 140:217–226, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
140/m140p217.pdf.

Grigg:1994:ESD

[Gri94] R. W. Grigg. Effects of sewage discharge, fishing pressure and habitat complexity on coral ecosystems and reef fishes in Hawaii. Marine Ecology Progress Series, 103:25–34, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
103/m103p025.pdf.

Guisande:1998:EBA

[GRSV98] C. Guisande, I. Riveiro, A. Solá, and L. Valdés. Effect of biotic and abiotic factors on the biochemical composition of wild eggs and larvae of several fish species. Marine Ecology Progress Series, 163:53–61, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/163/m163p053.pdf.

Grutter:1994:STV

[Gru94] A. S. Grutter. Spatial and temporal variations of the ectoparasites of seven reef fish species from Lizard Island and Heron Island, Australia. Marine Ecology Progress Series, 115: 21–30, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/115/m115p021.pdf.

Grutter:1995:RBC

[Gru95] A. S. Grutter. Relationship between cleaning rates and ectoparasite loads in coral reef fishes. Marine Ecology Progress Series, 118:51–58, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/118/m118p051.pdf.

Grutter:1996:PRR

[Gru96] A. Grutter. Parasite removal rates by the cleaner wrasse Labroides dimidiatus. Marine Ecology Progress Series, 130: 61–70, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/130/m130p061.pdf.

Gardner:1990:TDA

[GS90] J. P. A. Gardner and D. O. F. Skibinski. Thermostability differences of allozyme loci in Mytilus edulis, M. galloprovincialis and hybrid mussels. Marine Ecology Progress Series, 64: 99–105, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/64/m064p099.pdf.

Gardner:1991:BPF

[GS91a] J. P. A. Gardner and D. O. F. Skibinski. Biological and physical factors influencing genotype-dependent mortality in hybrid mussel populations. Marine Ecology Progress Series, 71: 235–243, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/71/m071p235.pdf.

Griffis:1991:MBA

[GS91b] R. B. Griffis and T. H. Suchanek. A model of burrow architecture and trophic modes in thalassinidean shrimp (Decapoda: Thalassinidea). Marine Ecology Progress Series, 79: 171–183, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/79/m079p171.pdf.

Gonzalez:1993:GMN

[GS93] J. M. González and C. A. Suttle. Grazing by marine nanofiagellates on viruses and virus-sized particles: ingestion and digestion. Marine Ecology Progress Series, 94:1–10, 1993.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
94/m094p001.pdf.

Gaino:1994:SST

[GS94a] E. Gaino and M. Sara. Siliceous spicules of Tethya seychellensis (Porifera) support the growth of a green alga: a possible light conducting system. Marine Ecology Progress Series, 108:147–151, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/108/m108p147.pdf.

Gonzalez:1994:PRC

[GS94b] H. E. González and V. Smetacek. The possible role of the cyclopoid copepod Oithona in retarding vertical flux of zooplankton faecal material. Marine Ecology Progress Series, 113:233–246, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/113/m113p233.pdf.

Gronkjaer:1999:NRM

[GS99a] P. Grønkjær and M. Schytte. Non-random mortality of Baltic cod larvae inferred from otolith hatch-check sizes. Marine Ecology Progress Series, 181:53–59, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/181/m181p053.pdf.

Gunnarsson:1999:APB

[GS99b] J. S. Gunnarsson and M. Sköld. Accumulation of polychlorinated biphenyls by the infaunal brittle stars Amphiura filiformis and A. chiajei: effects of eutrophication and selective feeding. Marine Ecology Progress Series, 186:173–185, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
186/m186p173.pdf.

Gilbert:1998:IBD

[GSB98] F. Gilbert, G. Stora, and P. Bonin. Influence of bioturbation on denitrification activity in Mediterranean coastal sediments: an in situ experimental approach. Marine Ecology Progress Series, 163:99–107, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/163/m163p099.pdf.

Gilbert:1997:ISF

[GSBB97] F. Gilbert, P. Souchu, M. Bianchi, and P. Bonin. Influence of shellfish farming activities on nitrification, nitrate reduction to ammonium and denitrification at the water–sediment interface of the Thau lagoon, France. Marine Ecology Progress Series, 151:143–153, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/151/m151p143.pdf.

Gremillet:1995:ERB

[GSC95] D. Grémillet, D. Schmid, and B. Culik. Energy requirements of breeding great cormorants Phalacrocorax carbo sinensis. Marine Ecology Progress Series, 121:1–9, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
121/m121p001.pdf.

Galgani:1996:ADD

[GSC96] F. Galgani, A. Souplet, and Y. Cadiou. Accumulation of debris on the deep sea floor off the French Mediterranean coast. Marine Ecology Progress Series, 142:225–234, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
142/m142p225.pdf.

Gutt:1996:IIS

[GSD96] J. Gutt, A. Starmans, and G. Dieckmann. Impact of iceberg scouring on polar benthic habitats. Marine Ecology Progress Series, 137:311–316, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/137/m137p311.pdf.

Griffith:1994:ADA

[GSG^{+}94] P. Griffith, F.-K. Shiah, K. Gloersen, H. W. Ducklow, and M. Fletcher. Activity and distribution of attached bacteria in Chesapeake Bay. Marine Ecology Progress Series, 108: 1–10, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/108/m108p001.pdf.

Gaill:1997:RPT

[GSM^{+}97] F. Gaill, B. Shillito, F. Ménard, G. Goffinet, and J. J. Childress. Rate and process of tube production by the deep-sea hydrothermal vent tubeworm Riftia pachyptila. Marine Ecology Progress Series, 148:135–143, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/148/m148p135.pdf.

Guisande:1996:TBO

[GSMM96] C. Guisande, J. Sánchez, I. Maneiro, and A. Miranda. Tradeoff between offspring number and offspring size in the marine copepod Euterpina acutifrons at different food concentrations. Marine Ecology Progress Series, 143:37–44, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
143/m143p037.pdf.

Gielazyn:1999:EES

[GSP99] M. L. Gielazyn, S. E. Stancyk, and W. W. Piegorsch. Experimental evidence of subsurface feeding by the burrowing ophiuroid Amphipholis gracillima (Echinodermata). Marine Ecology Progress Series, 184:129–138, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/184/m184p129.pdf.

Gonzalez:1993:DEA

[GSS93a] J. M. González, B. F. Sherr, and E. B. Sherr. Digestive enzyme activity as a quantitative measure of protistan grazing: the acid lysozyme assay for bacterivory. Marine Ecology Progress Series, 100:197–206, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/100/m100p197.pdf.

Gonzalez:1993:DFM

[GSS93b] J. M. González, E. B. Sherr, and B. F. Sherr. Differential feeding by marine flagellates on growing versus starving, and on motile versus nonmotile, bacterial prey. Marine Ecology Progress Series, 102:257–267, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/102/m102p257.pdf.

Goeyens:1991:SVI

[GST^{+}91] L. Goeyens, F. Sörensson, P. Tréguer, J. Morvan, M. Panouse, and F. Dehairs. Spatiotemporal variability of inorganic nitrogen stocks and uptake fluxes in the Scotia–Weddell Confluence area during November and December 1988. Marine Ecology Progress Series, 77:7–19, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/77/m077p007.pdf.

Gleitz:1992:PRS

[GT92] M. Gleitz and D. N. Thomas. Physiological responses of a small Antarctic diatom (Chaetoceros sp.) to simulated environmental constraints associated with sea-ice formation. Marine Ecology Progress Series, 88:271–278, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
88/m088p271.pdf.

Grecay:1996:ETL

[GT96] P. A. Grecay and T. E. Targett. Effects of turbidity, light level and prey concentration on feeding of juvenile weakfish Cynoscion regalis. Marine Ecology Progress Series, 131: 11–16, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/131/m131p011.pdf.

Guzman:1998:SVS

[GT98] H. M. Guzmán and A. W. Tudhope. Seasonal variation in skeletal extension rate and stable isotopic (^{13}C/^{12}C and 18O/^{16}O) composition in response to several environmental variables in the Caribbean reef coral Siderastrea siderea. Marine Ecology Progress Series, 166:109–118, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
166/m166p109.pdf.

Gowen:1992:PME

[GTJ92] R. J. Gowen, P. Tett, and K. J. Jones. Predicting marine eutrophication: the yield of chlorophyll from nitrogen in Scottish coastal waters. Marine Ecology Progress Series, 85: 153–161, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/85/m085p153.pdf.

Goeyens:1991:ARS

[GTL^{+}91] L. Goeyens, P. Tréguer, C. Lancelot, S. Mathot, S. Becquevort, J. Morvan, F. Dehairs, and W. Baeyens. Ammonium regeneration in the Scotia–Weddell Confluence area during spring 1988. Marine Ecology Progress Series, 78:241– 252, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/78/m078p241.pdf.

Godley:1998:TSM

[GTWF98] B. J. Godley, D. R. Thompson, S. Waldron, and R. W. Furness. The trophic status of marine turtles as determined by stable isotope analysis. Marine Ecology Progress Series, 166:277–284, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/166/m166p277.pdf.

Gonzalez:1997:IBO

[GTYR97] A. F. González, P. N. Trathan, C. Yau, and P. G. Rodhouse. Interactions between oceanography, ecology and fishery biology of the ommastrephid squid Martialia hyadesi in the South Atlantic. Marine Ecology Progress Series, 152:205–215, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
152/m152p205.pdf.

Glasby:1998:DPC

[GU98] T. M. Glasby and A. J. Underwood. Determining positions for control locations in environmental studies of estuarine marinas. Marine Ecology Progress Series, 171:1–14, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
171/m171p001.pdf.

Gunther:1991:SMB

[Gün91] Carmen-Pia Günther. Settlement of Macoma balthica on an intertidal sandflat in the Wadden Sea. Marine Ecology Progress Series, 76:73–79, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/76/m076p073.pdf.

Gallegos:1996:IPE

[GV96] C. L. Gallegos and W. N. Vant. An incubation procedure for estimating carbon-to-chlorophyll ratios and growth-irradiance relationships of estuarine phytoplankton. Marine Ecology Progress Series, 138:275–291, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/138/m138p275.pdf.

Gamenick:1998:EDC

[GVGG98] I. Gamenick, Bent Vismann, M. K. Grieshaber, and O. Giere. Ecophysiological differentiation of Capitella capitata (Polychaeta). Sibling species from different sulfidic habitats. Marine Ecology Progress Series, 175:155–166, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
175/m175p155.pdf.

Gomez:1991:FRB

[GW91] I. M. Gomez and R. C. Westermeier. Frond regrowth from basal disc in Iridaea laminarioides (Rhodophyta, Gigartinales) at Mehuín, southern Chile. Marine Ecology Progress Series, 73:83–91, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/73/m073p083.pdf.

Goericke:1993:CLM

[GW93] R. Goericke and N. A. Welschmeyer. The carotenoid-labeling method: measuring specific rates of carotenoid synthesis in natural phytoplankton communities. Marine Ecology Progress Series, 98:157–171, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/98/m098p157.pdf.

Gee:1994:MCS

[GW94] J. M. Gee and R. M. Warwick. Metazoan community structure in relation to the fractal dimensions of marine macroalgae. Marine Ecology Progress Series, 103:141–150, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
103/m103p141.pdf.

Gronkjaer:1997:OEE

[GW97] P. Grønkjær and K. Wieland. Ontogenetic and environmental effects on vertical distribution of cod larvae in the Bornholm Basin, Baltic Sea. Marine Ecology Progress Series, 154:91–105, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/154/m154p091.pdf.

Gomez:1997:PLR

[GWKW97] I. Gómez, G. Weykam, H. Klöser, and C. Wiencke. Photosynthetic light requirements, metabolic carbon balance and zonation of sublittoral macroalgae from King George Island (Antarctica). Marine Ecology Progress Series, 148:281–293, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/148/m148p281.pdf.

Gremillet:1999:TDS

[GWSG99] D. Grémillet, R. P. Wilson, S. Storch, and Y. Gary. Threedimensional space utilization by a marine predator. Marine Ecology Progress Series, 183:263–273, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
183/m183p263.pdf.

Gast:1998:BCR

[GWW^{+}98] G. J. Gast, S. Wiegman, E. Wieringa, F. C. van Duyl, and R. P. M. Bak. Bacteria in coral reef water types: removal of cells, stimulation of growth and mineralization. Marine Ecology Progress Series, 167:37–45, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/167/m167p037.pdf.

Gremillet:1999:TBA

[GWWP99] D. Grémillet, R. P. Wilson, S. Wanless, and G. Peters. A tropical bird in the Arctic (the cormorant paradox). Marine Ecology Progress Series, 188:305–309, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/188/m188p305.pdf.

Gimenez:1997:VZP

[GY97] L. Giménez and B. Yannicelli. Variability of zonation patterns in temperate microtidal Uruguayan beaches with different morphodynamic types. Marine Ecology Progress Series, 160:197–207, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/160/m160p197.pdf.

Gattuso:1993:DLD

[GYL93] J.-P. Gattuso, D. Yellowlees, and M. Lesser. Depth- and light-dependent variation of carbon partitioning and utilization in the zooxanthellate sclerac-tinian coral Stylophora pistillata. Marine Ecology Progress Series, 92:267–276, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
92/m092p267.pdf.

Hansen:1990:BPB

[HA90] J. A. Hansen and D. M. Alongi. Bacterial productivity and benthic standing stocks in a tropical coastal embayment. Marine Ecology Progress Series, 68:301–310, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
68/m068p301.pdf.

Hughes:1994:RUS

[HAA94] D. J. Hughes, A. D. Ansell, and R. J. A. Atkinson. Resource utilization by a sedentary surface deposit feeder, the echiuran worm Maxmuelteria lankesteri. Marine Ecology Progress Series, 112:267–275, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/112/m112p267.pdf.

Hagen:1995:RDG

[Hag95] N. T. Hagen. Recurrent destructive grazing of successionally immature kelp forests by green sea urchins in Vestfjorden, Northern Norway. Marine Ecology Progress Series, 123: 95–106, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/123/m123p095.pdf.

Hall:1990:CFM

[HAH^{+}90] P. O. J. Hall, L. G. Anderson, O. Holby, S. Kollberg, and M.-O. Samuelsson. Chemical fluxes and mass balances in a marine fish cage farm. I carbon. Marine Ecology Progress Series, 61:61–73, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/61/m061p061.pdf.

Hallegraeff:1998:TTD

[Hal98] G. M. Hallegraeff. Transport of toxic dinoflagellates via ships’ ballast water: bioeconomic risk assessment and efficacy of possible ballast water management strategies. Marine Ecology Progress Series, 168:297–309, 1998. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/168/m168p297.pdf.

Hansen:1991:DPD

[Han91a] P. J. Hansen. Dinophysis — a planktonic dinoflagellate genus which can act both as a prey and a predator of a ciliate. Marine Ecology Progress Series, 69:201–204, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
69/m069p201.pdf.

Hansen:1991:QIT

[Han91b] P. J. Hansen. Quantitative importance and trophic role of heterotrophic dinoflagellates in a coastal pelagial food web. Marine Ecology Progress Series, 73:253–261, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
73/m073p253.pdf.

Hanelt:1992:PPM

[Han92] D. Hanelt. Photoinhibiton of photosynthesis in marine macrophytes of the South China Sea. Marine Ecology Progress Series, 82:199–206, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/82/m082p199.pdf.

Hansen:1995:GGR

[Han95] P. J. Hansen. Growth and grazing response of a ciliate feeding on the red tide dinoflagellate Gyrodinium aureolum in monoculture and in mixture with a non-toxic alga. Marine Ecology Progress Series, 121:65–72, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/121/m121p065.pdf.

Hansson:1997:ETG

[Han97] L. J. Hansson. Effect of temperature on growth rate of Aurelia aurita (Cnidaria, Scyphozoa) from Gullmarsfjorden, Sweden. Marine Ecology Progress Series, 161:145–153, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
161/m161p145.pdf.

Hansen:1999:CGP

[Han99] B. W. Hansen. Cohort growth of planktotrophic polychaete larvae-are they food limited? Marine Ecology Progress Series, 178:109–119, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/178/m178p109.pdf.

Harrison:1990:NUC

[Har90] W. G. Harrison. Nitrogen utilization in chlorophyll and primary productivity maximum layers: an analysis based on the f-ratio. Marine Ecology Progress Series, 60:85–90, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
60/m060p085.pdf.

Harding:1994:LTT

[Har94] L. W. Harding, Jr. Long-term trends in the distribution of phytoplankton in Chesapeake Bay: roles of light, nutrients and streamflow. Marine Ecology Progress Series, 104: 267–291, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/104/m104p267.pdf.

Hobson:1995:SPP

[HAR95] K. A. Hobson, W. G. Ambrose, Jr., and P. E. Renaud. Sources of primary production, benthic-pelagic coupling, and trophic relationships within the Northeast Water Polynya: insights from δ ^{13}C and δ ^{15}N analysis. Marine Ecology Progress Series, 128:1–10, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/128/m128p001.pdf.

Harvey:1996:DMF

[Har96] A. W. Harvey. Delayed metamorphosis in Florida hermit crabs: multiple cues and constraints (Crustacea: Decapoda: Paguridae and Diogenidae). Marine Ecology Progress Series, 141:27–36, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/141/m141p027.pdf.

Harmelin:1997:DBM

[Har97] J. G. Harmelin. Diversity of bryozoans in a Mediterranean sublittoral cave with bathyal-like conditions: role of dispersal processes and local factors. Marine Ecology Progress Series, 153:139–152, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/153/m153p139.pdf.

Harvey:1998:USB

[Har98] C. J. Harvey. Use of sandy beach habitat by Fundulus majalis, a surf-zone fish. Marine Ecology Progress Series, 164: 307–310, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/164/m164p307.pdf.

Harding:1999:SFB

[Har99] J. M. Harding. Selective feeding behavior of larval naked gobies Gobiosoma bosc and blennies Chasmodes bosquianus and Hypsoblennius hentzi: preferences for bivalve veligers. Marine Ecology Progress Series, 179:145–153, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
179/m179p145.pdf.

Havenhand:1993:EJP

[Hav93] J. N. Havenhand. Egg to juvenile period, generation time, and the evolution of larval type in marine invertebrates. Marine Ecology Progress Series, 97:247–260, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
97/m097p247.pdf.

Hays:1995:DVM

[Hay95] G. C. Hays. Diel vertical migration behaviour of Calanus hyperboreus at temperate latitudes. Marine Ecology Progress Series, 127:301–304, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/127/m127p301.pdf.

Hansen:1991:AAM

[HB91a] L. S. Hansen and T. H. Blackburn. Aerobic and anaerobic mineralization of organic material in marine sediment microcosms. Marine Ecology Progress Series, 75:283–291, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
75/m075p283.pdf.

Heldal:1991:PDV

[HB91b] M. Heldal and G. Bratbak. Production and decay of viruses in aquatic environments. Marine Ecology Progress Series, 72: 205–212, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/72/m072p205.pdf.

Holm:1994:SPG

[HB94a] E. R. Holm and E. Bourget. Selection and population genetic structure of the barnacle Semibalanus balanoides in the northwest Atlantic and Gulf of St. Lawrence. Marine Ecology Progress Series, 113:247–256, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/113/m113p247.pdf.

Hutchins:1994:GMR

[HB94b] D. A. Hutchins and K. W. Bruland. Grazer-mediated regeneration and assimilation of Fe, Zn and Mn from planktonic prey. Marine Ecology Progress Series, 110:259–269, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
110/m110p259.pdf.

Harriott:1995:RSC

[HB95] V. J. Harriott and S. A. Banks. Recruitment of scleractinian corals in the Solitary Islands Marine Reserve, a high latitude coral-dominated community in Eastern Australia. Marine Ecology Progress Series, 123:155–161, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
123/m123p155.pdf.

Harvey:1997:RMI

[HB97] M. Harvey and E. Bourget. Recruitment of marine invertebrates onto arborescent epibenthic structures: active and passive processes acting at different spatial scales. Marine Ecology Progress Series, 153:203–215, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/153/m153p203.pdf.

Hirst:1998:LTC

[HB98] A. G. Hirst and S. D. Batten. Long-term changes in the diel vertical migration behaviour of Calanus finmarchicus in the North Sea are unrelated to fish predation. Marine Ecology Progress Series, 171:307–310, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/171/m171p307.pdf.

Hentschel:1999:MNS

[HBB^{+}99] U. Hentschel, E. C. Berger, M. Bright, H. Felbeck, and J. A. Ott. Metabolism of nitrogen and sulfur in ectosymbiotic bacteria of marine nematodes (Nematoda, Stilbonematinae). Marine Ecology Progress Series, 183:149–158, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
183/m183p149.pdf.

Hardege:1996:SPM

[HBBM96] J. D. Hardege, M. G. Bentley, M. Beckmann, and C. Müller. Sex pheromones in marine polychaetes: volatile organic substances (VOS) isolated from Arenicola marina. Marine Ecology Progress Series, 139:157–166, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/139/m139p157.pdf.

Heipel:1998:PGD

[HBBT98] D. A. Heipel, J. D. D. Bishop, A. R. Brand, and J. P. Thorpe. Population genetic differentiation of the great scallop Pecten maximus in western Britain investigated by randomly amplified polymorphic DNA. Marine Ecology Progress Series, 162:163–171, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/162/m162p163.pdf.

Hyland:1991:MCS

[HBC^{+}91] J. Hyland, E. Baptiste, J. Campbell, J. Kennedy, R. Kropp, and S. Williams. Macroinfaunal communities of the Santa Maria Basin on the California outer continental shelf and slope. Marine Ecology Progress Series, 78:147–161, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
78/m078p147.pdf.

Hoare:1995:VAP

[HBD95] K. Hoare, A. R. Beaumont, and J. Davenport. Variation among populations in the resistance of Mytilus edulis embryos to copper: adaptation to pollution? Marine Ecology Progress Series, 120:155–161, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/120/m120p155.pdf.

Harvey:1993:SIS

[HBM93] M. Harvey, E. Bourget, and G. Miron. Settlement of Iceland scallop Chlamys islandica spat in response to hydroids and filamentous red algae: field observations and laboratory experiments. Marine Ecology Progress Series, 99:283–292, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
99/m099p283.pdf.

Hinckley:1993:ASM

[HBP^{+}93] S. Hinckley, K. Bailey, S. Picquelle, M. Yoklavich, and P. Stabeno. Age-specific mortality and transport of larval walleye pollock Theragra chalcogramma in the western Gulf of Alaska. Marine Ecology Progress Series, 98:17–29, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
98/m098p017.pdf.

Hall:1991:PRI

[HBR^{+}91] S. J. Hall, D. J. Basford, M. R. Robertson, D. G. Raffaelli, and I. Tuck. Patterns of recolonisation and the importance of pit-digging by the crab Cancer pagurus in a subtidal sand habitat. Marine Ecology Progress Series, 72:93–102, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
72/m072p093.pdf.

Hardege:1998:SSJ

[HBS98] J. D. Hardege, M. G. Bentley, and L. Snape. Sediment selection by juvenile Arenicola marina. Marine Ecology Progress Series, 166:187–195, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/166/m166p187.pdf.

Hondeveld:1992:BHN

[HBvD92] B. J. M. Hondeveld, R. P. M. Bak, and F. C. van Duyl. Bacterivory by heterotrophic nanoflagellates in marine sediments measured by uptake of fluorescently labeled bacteria. Marine Ecology Progress Series, 89:63–71, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/89/m089p063.pdf.

Hettler:1990:TDI

[HC90] W. F. Hettler, Jr. and A. J. Chester. Temporal distribution of ichthyoplankton near Beaufort Inlet, North Carolina. Marine Ecology Progress Series, 68:157–168, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
68/m068p157.pdf.

Hardy:1992:SMC

[HC92a] John T. Hardy and John Cleary. Surface microlayer contamination and toxicity in the German Bight. Marine Ecology Progress Series, 91:203–210, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/91/m091p203.pdf.

Highsmith:1992:PAA

[HC92b] Raymond C. Highsmith and Kenneth O. Coyle. Productivity of Arctic amphipods relative to gray whale energy requirements. Marine Ecology Progress Series, 83:141–150, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
83/m083p141.pdf.

Hare:1993:EEI

[HC93] J. A. Hare and R. K. Cowen. Ecological and evolutionary implications of the larval transport and reproductive strategy of bluefish Pomatomus saltatrix. Marine Ecology Progress Series, 98:1–16, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/98/m098p001.pdf.

Hosie:1994:MDP

[HC94] G. W. Hosie and T. G. Cochran. Mesoscale distribution patterns of macrozooplankton communities in Prydz Bay, Antarctica, January to February 1991. Marine Ecology Progress Series, 106:21–39, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/106/m106p021.pdf.

Harrison:1991:NPD

[HCC^{+}91] P. J. Harrison, P. J. Clifford, W. P. Cochlan, K. Yin, M. A. St. John, P. A. Thompson, M. J. Sibbald, and L. J. Albright. Nutrient and plankton dynamics in the Fraser River plume, Strait of Georgia, British Columbia. Marine Ecology Progress Series, 70:291–304, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/70/m070p291.pdf.

Hentschel:1993:NRC

[HCF93] Ute Hentschel, S. Craig Cary, and Horst Felbeck. Nitrate respiration in chemoautotrophic symbionts of the bivalve Lucinoma aequizonata. Marine Ecology Progress Series, 94: 35–41, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/94/m094p035.pdf.

Hernandez:1993:SCG

[HCG^{+}93] I. Hernández, A. Corzo, R. J. Gordillo, M. D. Robles, E. Saez, J. A. Fernández, and R. X. Niell. Seasonal cycle of the gametophytic form of Porphyra umbilicalis: nitrogen and carbon. Marine Ecology Progress Series, 99:301–311, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
99/m099p301.pdf.

Hunt:1996:FES

[HCH^{+}96] G. L. Hunt, Jr., K. O. Coyle, S. Hoffman, M. B. Decker, and E. N. Flint. Foraging ecology of short-tailed shearwaters near the Pribilof Islands, Bering Sea. Marine Ecology Progress Series, 141:1–11, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/141/m141p001.pdf.

Hudon:1993:IPF

[HCI93] C. Hudon, R. E. Crawford, and R. G. Ingram. Influence of physical forcing on the spatial distribution of marine fauna near Resolution Island (eastern Hudson Strait). Marine Ecology Progress Series, 92:1–14, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/92/m092p001.pdf.

Harrison:1990:NUI

[HCS90] W. G. Harrison, G. F. Cota, and R. E. H. Smith. Nitrogen utilization in ice algal communities of Barrow Strait, Northwest Territories, Canada. Marine Ecology Progress Series, 67:275–283, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/67/m067p275.pdf.

Herman:1993:DAE

[HCS93] A. W. Herman, N. A. Cochrane, and D. D. Sameoto. Detection and abundance estimation of euphausiids using an optical plankton counter. Marine Ecology Progress Series, 94: 165–173, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/94/m094p165.pdf.

Himmelman:1991:DPS

[HD91] J. H. Himmelman and C. Dutil. Distribution, population structure and feeding of subtidal seastars in the northern Gulf of St. Lawrence. Marine Ecology Progress Series, 76: 61–72, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/76/m076p061.pdf.

Hewitt:1993:DAA

[HD93] R. P. Hewitt and D. A. Demer. Dispersion and abundance of Antarctic krill in the vicinity of Elephant Island in the 1992 austral summer. Marine Ecology Progress Series, 99: 29–39, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/99/m099p029.pdf.

Herzka:1997:SPP

[HD97] S. Z. Herzka and K. H. Dunton. Seasonal photosynthetic patterns of the seagrass Thalassia testudinum in the western Gulf of Mexico. Marine Ecology Progress Series, 152: 103–117, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/152/m152p103.pdf.

Hoare:1995:EEP

[HDB95] K. Hoare, J. Davenport, and A. R. Beaumont. Effects of exposure and previous exposure to copper on growth of veliger larvae and survivorship of Mytilus edulis juveniles. Marine Ecology Progress Series, 120:163–168, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/120/m120p163.pdf.

Hammerstrom:1998:RPI

[HDD98] K. Hammerstrom, M. N. Dethier, and D. O. Duggins. Rapid phlorotannin induction and relaxation in five Washington kelps. Marine Ecology Progress Series, 165:293–305, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
165/m165p293.pdf.

Hoppe:1993:EDB

[HDK93] H.-G. Hoppe, H. Ducklow, and B. Karrasch. Evidence for dependency of bacterial growth on enzymatic hydrolysis of particulate organic matter in the mesopelagic ocean. Marine Ecology Progress Series, 93:277–283, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
93/m093p277.pdf.

Hargrave:1993:SCB

[HDPW93] B. T. Hargrave, D. E. Duplisea, E. Pfeiffer, and D. J. Wildish. Seasonal changes in benthic fluxes of dissolved oxygen and ammonium associated with marine cultured Atlantic salmon. Marine Ecology Progress Series, 96:249–257, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
96/m096p249.pdf.

Hofman:1991:ASD

[HdWS91] P. A. G. Hofman, S. A. de Jong, E. J. Wagenvoort, and A. J. J. Sandee. Apparent sediment diffusion coefficients for oxygen and oxygen consumption rates measured with microelectrodes and bell jars: applications to oxygen budgets in estuarine intertidal sediments (Oosterschelde, SW Netherlands). Marine Ecology Progress Series, 69:261–272, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
69/m069p261.pdf.

Heinanen:1991:BNB

[Hei91] A. P. Heinänen. Bacterial numbers, biomass and productivity in the Baltic Sea: a cruise study. Marine Ecology Progress Series, 70:283–290, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/70/m070p283.pdf.

Heinanen:1992:BSE

[Hei92a] A. Heinänen. Bacterioplankton in a subarctic estuary: the Gulf of Bothnia (the Baltic Sea). Marine Ecology Progress Series, 86:123–131, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/86/m086p123.pdf.

Heip:1992:BSS

[Hei92b] C. Heip. Benthic studies: summary and conclusions. Marine Ecology Progress Series, 91:265–268, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
91/m091p265.pdf.

Heiskanen:1995:CST

[Hei95] A. S. Heiskanen. Contamination of sediment trap fluxes by vertically migrating phototrophic micro-organisms in the coastal Baltic Sea. Marine Ecology Progress Series, 122: 45–58, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/122/m122p045.pdf.

Harding:1999:ATS

[HF99] G. C. Harding and A. J. Fraser. Application of the triacylglycerol/sterol condition index to the interpretation of larval lobster Homarus americanus distribution in close proximity to Georges Bank, Gulf of Maine. Marine Ecology Progress Series, 186:239–254, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/186/m186p239.pdf.

Haury:1995:CCO

[HFG^{+}95] L. Haury, C. Fey, G. Gal, A. Hobday, and A. Genin. Copepod carcasses in the ocean. I. Over seamounts. Marine Ecology Progress Series, 123:57–63, 1995. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/123/m123p057.pdf.

Hopkins:1994:STE

[HFGT94] T. L. Hopkins, M. E. Flock, J. V. Gartner, Jr., and J. J. Torres. Structure and trophic ecology of a low latitude mid-water decapod and mysid assemblage. Marine Ecology Progress Series, 109:143–156, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/109/m109p143.pdf.

Hopkinson:1991:CMN

[HFJS91] C. S. Hopkinson, Jr., R. D. Fallon, B.-O. Jansson, and J. P. Schubauer. Community metabolism and nutrient cycling at Gray’s Reef, a hard bottom habitat in the Georgia Bight. Marine Ecology Progress Series, 73:105–120, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
73/m073p105.pdf.

Han:1991:SSP

[HFN91] M.-S. Han, K. Furuya, and T. Nemoto. Species-specific productivity of Skeletonema costatum (Bacillariophyceae) in the inner part of Tokyo Bay. Marine Ecology Progress Series, 79: 267–273, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/79/m079p267.pdf.

Harvell:1993:LGV

[HFR^{+}93] C. Drew Harvell, William Fenical, Vassilios Roussis, Jennifer L. Ruesink, Catherine C. Griggs, and Charles H. Greene. Local and geographic variation in the defensive chemistry of a West Indian gorgonian coral (Briareum asbestinum). Marine Ecology Progress Series, 93:165–173, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
93/m093p165.pdf.

Hansen:1999:IBF

[HFW99] R. Hansen, T. L. Forbes, and P. Westermann. Importance of bioturbation and feeding by the polychaete Capitella sp. I in the degradation of di(2-ethylhexyl)phthalate (DEHP).

Marine Ecology Progress Series, 182:187–199, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
182/m182p187.pdf.

Huettel:1992:IBI

[HG92a] M. Huettel and G. Gust. Impact of bioroughness on interfacia solute exchange in permeable sediments. Marine Ecology Progress Series, 89:253–267, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/89/m089p253.pdf.

Huettel:1992:SRM

[HG92b] M. Huettel and G. Gust. Solute release mechanisms from confined sediment cores in stirred benthic chambers and flume flows. Marine Ecology Progress Series, 82:187–197, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
82/m082p187.pdf.

Howes:1994:PDD

[HG94] B. L. Howes and D. D. Goehringer. Porewater drainage and dissolved organic carbon and nutrient losses through the intertidal creek-banks of a New England salt marsh. Marine Ecology Progress Series, 114:289–301, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/114/m114p289.pdf.

Hall:1998:TDD

[HG98] S. J. Hall and S. P. Greenstreet. Taxonomic distinctness and diversity measures: responses in marine fish communities. Marine Ecology Progress Series, 166:227–229, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
166/m166p227.pdf.

Heise:1999:IPS

[HG99] S. Heise and G. Gust. Influence of the physiological status of bacteria on their transport into permeable sediments. Marine Ecology Progress Series, 190:141–153, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
190/m190p141.pdf.

Hays:1999:LTT

[HGB99] G. C. Hays, B. J. Godley, and A. C. Broderick. Long-term thermal conditions on the nesting beaches of green turtles on Ascension Island. Marine Ecology Progress Series, 185: 297–299, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/185/m185p297.pdf.

Hochscheid:1999:RDH

[HGBW99] S. Hochscheid, B. J. Godley, A. C. Broderick, and R. P. Wilson. Reptilian diving: highly variable dive patterns in the green turtle Chelonia mydas. Marine Ecology Progress Series, 185:101–112, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/185/m185p101.pdf.

Hoegh-Guldberg:1999:PPS

[HGJ99] O. Hoegh-Guldberg and R. J. Jones. Photoinhibition and photoprotection in symbiotic dinoflagellates from reef-building corals. Marine Ecology Progress Series, 183:73–86, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
183/m183p073.pdf.

Hatcher:1994:ESM

[HGS94] A. Hatcher, J. Grant, and B. Schofield. Effects of suspended mussel culture (Mytilus spp.) on sedimentation, benthic respiration and sediment nutrient dynamics in a coastal bay. Marine Ecology Progress Series, 115:219–235, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
115/m115p219.pdf.

Hoegh-Guldberg:1995:PMB

[HGS95] O. Hoegh-Guldberg and B. Salvat. Periodic mass-bleaching and elevated sea temperatures: bleaching of outer reef slope communities in Moorea, French Polynesia. Marine Ecology Progress Series, 121:181–190, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/121/m121p181.pdf.

Hunt:1990:FHP

[HH90] G. L. Hunt, Jr. and N. M. Harrison. Foraging habitat and prey taken by least auklets at King Island, Alaska. Marine Ecology Progress Series, 65:141–150, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/65/m065p141.pdf.

Holby:1991:CFM

[HH91a] O. Holby and P. O. J. Hall. Chemical fluxes and mass balances in a marine fish cage farm. II phosphorus. Marine Ecology Progress Series, 70:263–272, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/70/m070p263.pdf.

HuntvonHerbing:1991:SRB

[HH91b] I. Hunt von Herbing and W. Hunte. Spawning and recruitment of the blue-head wrasse Thalassoma bifasciatum in Barbados, West Indies. Marine Ecology Progress Series, 72: 49–58, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/72/m072p049.pdf.

Head:1992:CCT

[HH92] E. J. H. Head and L. R. Harris. Chlorophyll and carotenoid transformation and destruction by Calanus spp. grazing on diatoms. Marine Ecology Progress Series, 86:229–238, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
86/m086p229.pdf.

Head:1994:FSC

[HH94a] E. J. H. Head and L. R. Harris. Feeding selectivity by copepods grazing on natural mixtures of phytoplankton determined by HPLC analysis of pigments. Marine Ecology Progress Series, 110:75–83, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/110/m110p075.pdf.

Heissenberger:1994:FHM

[HH94b] A. Heissenberger and G. J. Herndl. Formation of high molecular weight material by free-living marine bacteria.

Marine Ecology Progress Series, 111:129–135, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
111/m111p129.pdf.

Hunter:1995:EGC

[HH95] E. Hunter and R. N. Hughes. Environmental and genetic components of variation in sexual allocation by an epialgal bryozoan. Marine Ecology Progress Series, 120:193–201, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
120/m120p193.pdf.

Head:1996:CDC

[HH96] E. J. H. Head and L. R. Harris. Chlorophyll destruction by Calanus spp. grazing on phytoplankton: kinetics, effects of ingestion rate and feeding history, and a mechanistic interpretation. Marine Ecology Progress Series, 135:223–235, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
135/m135p223.pdf.

Hsieh:1999:DRA

[HH99] H. L. Hsieh and C. F. Hsu. Differential recruitment of annelids onto tidal elevations in an estuarine mud flat. Marine Ecology Progress Series, 177:93–102, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/177/m177p093.pdf.

Hansen:1993:MON

[HHB93] L. S. Hansen, M. Holmer, and T. H. Blackburn. Mineralization of organic nitrogen and carbon (fish food) added to anoxic sediment microcosms: role of sulphate reduction. Marine Ecology Progress Series, 102:199–204, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
102/m102p199.pdf.

Holmes:1997:LVP

[HHB97] N. J. Holmes, V. J. Harriott, and S. A. Banks. Latitudinal variation in patterns of colonisation of cryptic calcareous marine organisms. Marine Ecology Progress Series, 155: 103–113, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/155/m155p103.pdf.

Hunt:1992:DPP

[HHE92] G. L. Hunt, Jr., D. Heinemann, and I. Everson. Distributions and predator–prey interactions of macaroni penguins, Antarctic fur seals, and Antarctic krill near Bird Island, South Georgia. Marine Ecology Progress Series, 86:15–30, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
86/m086p015.pdf.

Hoare:1999:MGE

[HHG99] K. Hoare, R. N. Hughes, and A. J. Goldson. Molecular genetic evidence for the prevalence of outcrossing in the hermaphroditic brooding bryozoan Celleporella hyalina. Marine Ecology Progress Series, 188:73–79, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
188/m188p073.pdf.

Hylland:1992:HMH

[HHH92] K. Hylland, C. Haux, and C. Hogstrand. Hepatic metallothionein and heavy metals in dab Limanda limanda from the German Bight. Marine Ecology Progress Series, 91: 89–96, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/91/m091p089.pdf.

Hays:1997:VNF

[HHH97] G. C. Hays, R. P. Harris, and R. N. Head. The vertical nitrogen flux caused by zooplankton diel vertical migration. Marine Ecology Progress Series, 160:57–62, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
160/m160p057.pdf.

Haywood:1998:PJT

[HHK^{+}98] M. D. E. Haywood, D. S. Heales, R. A. Kenyon, N. R. Loneragan, and D. J. Vance. Predation of juvenile tiger prawns in a tropical Australian estuary. Marine Ecology Progress Series, 162:201–214, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/162/m162p201.pdf.

Hall:1992:CFM

[HHKS92] P. O. J. Hall, O. Holby, S. Kollberg, and M.-O. Samuelson. Chemical fluxes and mass balances in a marine fishcagefarm. IV nitrogen. Marine Ecology Progress Series, 89:81–91, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
89/m089p081.pdf.

Hinckley:1996:DSE

[HHM96] S. Hinckley, A. J. Hermann, and B. A. Megrey. Development of a spatially explicit, individual-based model of marine fish early life history. Marine Ecology Progress Series, 139: 47–68, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/139/m139p047.pdf.

Hull:1997:FZR

[HHM97] C. L. Hull, M. A. Hindell, and K. Michael. Foraging zones of royal penguins during the breeding season, and their association with oceanographic features. Marine Ecology Progress Series, 153:217–228, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/153/m153p217.pdf.

Huppertz:1990:PPM

[HHN90] K. Huppertz, D. Hanelt, and W. Nultsch. Photoinhibition of photosynthesis in the marine brown algae Fucus serratus as studied in field experiments. Marine Ecology Progress Series, 66:175–182, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/66/m066p175.pdf.

Hanelt:1993:DCP

[HHN93] D. Hanelt, K. Huppertz, and W. Nultsch. Daily course of photosynthesis and photoinhibition in marine macroalgae investigated in the laboratory and field. Marine Ecology Progress Series, 97:31–37, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/97/m097p031.pdf.

Herman:1996:WWE

[HHN^{+}96] P. M. J. Herman, M. A. Hemminga, P. H. Nienhuis, J. M. Verschuure, and E. G. J. Wessel. Wax and wane of eelgrass Zostera marina and water column silicon levels. Marine Ecology Progress Series, 144:303–307, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/144/m144p303.pdf.

Hillgruber:1995:FSL

[HHP95] N. Hillgruber, L. J. Haldorson, and A. J. Paul. Feeding selectivity of larval walleye pollock Theragra chalcogramma in the oceanic domain of the Bering Sea. Marine Ecology Progress Series, 120:1–10, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/120/m120p001.pdf.

Hemminga:1996:ACA

[HHSvS96] M. A. Hemminga, A. H. L. Huiskes, M. Steegstra, and J. van Soelen. Assessment of carbon allocation and biomass production in a natural stand of the salt marsh plant Spartina anglica using ^{13}C. Marine Ecology Progress Series, 130:169– 178, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/130/m130p169.pdf.

Hemminga:1991:BNL

[HHv91] M. A. Hemminga, P. G. Harrison, and F. van Lent. The balance of nutrient losses and gains in seagrass meadows. Marine Ecology Progress Series, 71:85–96, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/71/m071p085.pdf.

Hines:1990:GSF

[HHW90] A. H. Hines, A. M. Haddon, and L. A. Wiechert. Guild structure and foraging impact of blue crabs and epibenthic fish in a sub-estuary of Chesapeake Bay. Marine Ecology Progress Series, 67:105–126, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/67/m067p105.pdf.

Huang:1999:SFP

[HHW99] B. Huang, H. Hong, and H. Wang. Size-fractionated primary productivity and the phytoplankton-bacteria relationship in the Taiwan Strait. Marine Ecology Progress Series, 183:29–38, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/183/m183p029.pdf.

Hardy:1992:RDC

[HHYD92] J. T. Hardy, F. E. Hoge, J. K. Yungel, and R. E. Dodge. Remote detection of coral ’bleaching’ using pulsed-laser fluorescence spectroscopy. Marine Ecology Progress Series, 88: 247–255, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/88/m088p247.pdf.

Hicks:1992:TDF

[Hic92] G. R. F. Hicks. Tidal and diel fluctuations in abundance of meiobenthic copepods on an intertidal estuarine sandbank. Marine Ecology Progress Series, 87:15–21, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
87/m087p015.pdf.

Hill:1991:ADM

[Hil91a] A. E. Hill. Advection-diffusion-mortality solutions for investigating pelagic larval dispersal. Marine Ecology Progress Series, 70:117–128, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/70/m070p117.pdf.

Hill:1991:VMT

[Hil91b] A. E. Hill. Vertical migration in tidal currents. Marine Ecology Progress Series, 75:39–54, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/75/m075p039.pdf.

Hily:1991:ABS

[Hil91c] C. Hily. Is the activity of benthic suspension feeders a factor controlling water quality in the Bay of Brest? Marine Ecology Progress Series, 69:179–188, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/69/m069p179.pdf.

Hilgerloh:1997:PBB

[Hil97] G. Hilgerloh. Predation by birds on blue mussel Mytilus edulis beds of the tidal flats of spiekeroog (southern North Sea). Marine Ecology Progress Series, 146:61–72, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
146/m146p061.pdf.

Hinga:1992:COD

[Hin92] K. R. Hinga. Co-occurrence of dinoflagellate blooms and high pH in marine enclosures. Marine Ecology Progress Series, 86:181–187, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/86/m086p181.pdf.

Holmboe:1999:NAB

[HJA99] N. Holmboe, H. S. Jensen, and F. Ø. Andersen. Nutrient addition bioassays as indicators of nutrient limitation of phytoplankton in an eutrophic estuary. Marine Ecology Progress Series, 186:95–104, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/186/m186p095.pdf.

Hawkins:1999:MSF

[HJH^{+}99] A. J. S. Hawkins, M. R. James, R. W. Hickman, S. Hatton, and M. Weatherhead. Modelling of suspension-feeding and growth in the green-lipped mussel Perna canaliculus exposed to natural and experimental variations of seston availability in the Marlborough Sounds, New Zealand. Marine Ecology Progress Series, 191:217–232, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/191/m191p217.pdf.

Hyder:1998:BDE

[HJHG98] K. Hyder, M. P. Johnson, S. J. Hawkins, and W. S. C. Gurney. Barnacle demography: evidence for an existing model and spatial scales of variation. Marine Ecology Progress Series, 174:89–99, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/174/m174p089.pdf.

Holmer:1992:IMF

[HK92] M. Holmer and E. Kristensen. Impact of marine fish cage farming on metabolism and sulfate reduction of underlying sediments. Marine Ecology Progress Series, 80:191–201, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
80/m080p191.pdf.

Hoch:1993:SIA

[HK93] M. P. Hoch and D. L. Kirchman. Seasonal and inter-annual variability in bacterial production and biomass in a temperate estuary. Marine Ecology Progress Series, 98:283–295, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
98/m098p283.pdf.

Holmer:1994:CSR

[HK94] M. Holmer and E. Kristensen. Coexistence of sulfate reduction and methane production in an organic-rich sediment. Marine Ecology Progress Series, 107:177–184, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
107/m107p177.pdf.

Hart:1996:RHFa

[HK96] A. M. Hart and D. W. Klumpp. Response of herbivorous fishes to crown-of-thorns starfish Acanthaster planci outbreaks. I. Substratum analysis and feeding ecology of Acanthurus nigrofuscus and Scarus frenatus. Marine Ecology Progress Series, 132:11–19, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/132/m132p011.pdf.

Hansen:1997:QIC

[HK97] J. L. S. Hansen and T. Kiørboe. Quantifying interspecific coagulation efficiency of phytoplankton. Marine Ecology Progress Series, 159:75–79, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/159/m159p075.pdf.

Hillgruber:1999:DFB

[HK99] N. Hillgruber and M. Kloppmann. Distribution and feeding of blue whiting Micromesistius poutassou larvae in relation to different water masses in the Porcupine Bank area, west of Ireland. Marine Ecology Progress Series, 187:213–225, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
187/m187p213.pdf.

Hansen:1996:MSD

[HKA96] J. L. S. Hansen, T. Kiørboe, and A. L. Alldredge. Marine snow derived from abandoned larvacean houses: sinking rates, particle content and mechanisms of aggregate formation. Marine Ecology Progress Series, 141:205–215, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
141/m141p205.pdf.

Hagen:1993:CAC

[HKG93] W. Hagen, G. Kattner, and M. Graeve. Calanoides acutus and Calanus propinquus, Antarctic copepods with different lipid storage modes via wax esters or triacylglycerols. Marine Ecology Progress Series, 97:135–142, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
97/m097p135.pdf.

Heinanen:1995:BGA

[HKK^{+}95] A. Heinänen, K. Kononen, H. Kuosa, J. Kuparinen, and K. Mäkelä. Bacterioplankton growth associated with physical fronts during a cyanobacterial Bloom. Marine Ecology Progress Series, 116:233–245, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/116/m116p233.pdf.

Hjooleifsson:1992:ELS

[HKM92] E. Hjoöleifsson and G. Klein-MacPhee. Estimation of live standard length for winter flounder Pleuronectes americanus larvae from formalin-preserved, ethanol-preserved and frozen specimens. Marine Ecology Progress Series, 82:13–19, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
82/m082p013.pdf.

Hamilton:1998:ASL

[HKMK98] P. K. Hamilton, A. R. Knowlton, M. K. Marx, and S. D. Kraus. Age structure and longevity in North Atlantic right whales Eubalaena glacialis and their relation to reproduction. Marine Ecology Progress Series, 171:285–292, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
171/m171p285.pdf.

Hart:1996:RHFb

[HKR96] A. M. Hart, D. W. Klumpp, and G. R. Russ. Response of herbivorous fishes to crown-of-thorns starfish Acanthaster planci outbreaks. II. Density and biomass of selected species of herbivorous fish and fish-habitat correlations. Marine Ecology Progress Series, 132:21–30, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/132/m132p021.pdf.

Hara:1996:AVD

[HKT^{+}96] S. Hara, I. Koike, K. Terauchi, H. Kamiya, and E. Tanoue. Abundance of viruses in deep oceanic waters. Marine Ecology Progress Series, 145:269–277, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/145/m145p269.pdf.

Hemminga:1993:ESM

[HKvB93] M. A. Hemminga, V. A. Klap, J. van Soelen, and J. J. Boon. Effect of salt marsh inundation on estuarine particulate organic matter characteristics. Marine Ecology Progress Series, 99:153–161, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/99/m099p153.pdf.

Harrison:1997:NPD

[HKY^{+}97] P. J. Harrison, N. Khan, K. Yin, M. Saleem, N. Bano, M. Nisa, S. I. Ahmed, N. Rizvi, and F. Azam. Nutrient and phytoplankton dynamics in two mangrove tidal creeks of the Indus River delta, Pakistan. Marine Ecology Progress Series, 157:13–19, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/157/m157p013.pdf.

Hassett:1990:SCF

[HL90] R. P. Hassett and M. R. Landry. Seasonal changes in feeding rate, digestive enzyme activity, and assimilation efficiency of Calanus pacificus. Marine Ecology Progress Series, 62: 203–210, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/62/m062p203.pdf.

Hirst:1998:SIB

[HL98] A. G. Hirst and C. H. Lucas. Salinity influences body weight quantification in the scyphomedusa Aurelia aurita: important implications for body weight determination in gelatinous zooplankton. Marine Ecology Progress Series, 165:259–269, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
165/m165p259.pdf.

Herring:1993:BPC

[HLBW93] P. J. Herring, M. I. Latz, N. J. Bannister, and E. A. Widder. Bioluminescence of the poecilostomatoid copepod Oncaea conifera. Marine Ecology Progress Series, 94:297–309, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
94/m094p297.pdf.

Henley:1991:DRP

[HLF^{+}91] W. J. Henley, G. Levavasseur, L. A. Franklin, S. T. Lindley, J. Ramus, and C. B. Osmond. Diurnal responses of photosynthesis and fluorescence in Ulva rotundata acclimated to sun and shade in outdoor culture. Marine Ecology Progress Series, 75:19–28, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/75/m075p019.pdf.

Heissenberger:1996:UMS

[HLH96] A. Heissenberger, G. G. Leppard, and G. J. Herndl. Ultrastructure of marine snow. II. Microbiological considerations. Marine Ecology Progress Series, 135:299–308, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
135/m135p299.pdf.

Hays:1999:CBD

[HLP^{+}99a] G. C. Hays, P. Luschi, F. Papi, C. del Seppia, and R. Marsh. Changes in behaviour during the inter-nesting period and post-nesting migration for Ascension Island green turtles. Marine Ecology Progress Series, 189:263–273, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
189/m189p263.pdf.

Honkoop:1999:EES

[HLP99b] P. J. C. Honkoop, P. C. Luttikhuizen, and T. Piersma. Experimentally extending the spawning season of a marine bivalve using temperature change and fluoxetine as synergistic triggers. Marine Ecology Progress Series, 180:297–300, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
180/m180p297.pdf.

Havenhand:1990:RHL

[HlS90] J. N. Havenhand and l. Svane. Roles of hydrodynamics and larval behaviour in determining spatial aggregation in the tunicate Ciona intestinalis. Marine Ecology Progress Series, 68:271–276, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/68/m068p271.pdf.

Hacker:1991:WHA

[HM91] S. D. Hacker and L. P. Madin. Why habitat architecture and color are important to shrimps living in pelagic Sargassum: use of camouflage and plant-part mimicry. Marine Ecology Progress Series, 70:143–155, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/70/m070p143.pdf.

Ho:1994:ESE

[HM94a] C. Ho and J. Marra. Early-spring export of phytoplankton production in the northeast Atlantic Ocean. Marine Ecology Progress Series, 114:197–202, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/114/m114p197.pdf.

Hutchings:1994:TCR

[HM94b] J. A. Hutchings and R. A. Myers. Timing of cod reproduction: interannual variability and the influence of temperature. Marine Ecology Progress Series, 108:21–31, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
108/m108p021.pdf.

Hemminga:1996:SCI

[HM96] M. A. Hemminga and M. A. Mateo. Stable carbon isotopes in seagrasses: variability in ratios and use in ecological studies. Marine Ecology Progress Series, 140:285–298, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
140/m140p285.pdf.

Hale:1997:SSM

[HM97a] M. S. Hale and J. G. Mitchell. Sea surface microlayer and bacterioneuston spreading dynamics. Marine Ecology Progress Series, 147:269–276, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/147/m147p269.pdf.

Hobson:1997:TVA

[HM97b] L. A. Hobson and M. R. McQuoid. Temporal variations among planktonic diatom assemblages in a turbulent environment of the southern Strait of Georgia, British Columbia, Canada. Marine Ecology Progress Series, 150:263–274, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
150/m150p263.pdf.

Hoch:1997:MBC

[HM97c] T. Hoch and A. Ménesguen. Modelling the biogeochemical cycles of elements limiting primary production in the English Channel. II. Sensitivity analyses. Marine Ecology Progress Series, 146:189–205, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/146/m146p189.pdf.

Hovel:1997:PSF

[HM97d] K. A. Hovel and S. G. Morgan. Planktivory as a selective force for reproductive synchrony and larval migration. Marine Ecology Progress Series, 157:79–95, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/157/m157p079.pdf.

Hole:1993:ARC

[HMB93] L. M. Hole, M. N. Moore, and D. Bellamy. Age-related cellular reactions to copper in the marine mussel Mytilus edulis. Marine Ecology Progress Series, 94:175–179, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
94/m094p175.pdf.

Hole:1995:ARC

[HMB95] L. M. Hole, M. N. Moore, and D. Bellamy. Age-related cellular and physiological reactions to hypoxia and hyperthermia in marine mussels. Marine Ecology Progress Series, 122: 173–178, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/122/m122p173.pdf.

Harms:1994:GPC

[HMHDA94] J. Harms, B. Meyer-Harms, R. R. Dawirs, and K. Anger. Growth and physiology of Carcinus maenas (Decapoda, Portunidae) larvae in the field and in laboratory experiments. Marine Ecology Progress Series, 108:107–118, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
108/m108p107.pdf.

Helbling:1995:BVA

[HMVHH95] E. W. Helbling, E. R. Marguet, V. E. Villafañe, and O. Holm-Hansen. Bacterioplankton viability in Antarctic waters as affected by solar ultraviolet radiation. Marine Ecology Progress Series, 126:293–298, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/126/m126p293.pdf.

Hanelt:1997:EHL

[HMWN97] D. Hanelt, B. Melchersmann, C. Wiencke, and W. Nultsch. Effects of high light stress on photosynthesis of polar macroalgae in relation to depth distribution. Marine Ecology Progress Series, 149:255–266, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/149/m149p255.pdf.

Hanelt:1990:DCPa

[HN90a] D. Hanelt and W. Nultsch. Daily changes of the phaeoplast arrangement in the brown alga Dictyota dichotoma as studied in field experiments. Marine Ecology Progress Series, 61:273–279, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/61/m061p273.pdf.

Hanelt:1990:DCPb

[HN90b] D. Hanelt and W. Nultsch. Daily changes of the phaeoplast arrangement in the brown alga Dictyota dichotoma as studied in field experiments. Marine Ecology Progress Series, 61:273–279, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/61/m061p273.pdf.

Hough:1991:FSR

[HN91] A. R. Hough and E. Naylor. Field studies on retention of the planktonic copepod Eurytemora affinis in a mixed estuary. Marine Ecology Progress Series, 76:115–122, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
76/m076p115.pdf.

Hunter:1993:IMS

[HN93] E. Hunter and E. Naylor. Intertidal migration by the shore crab Carcinus maenas. Marine Ecology Progress Series, 101: 131–138, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/101/m101p131.pdf.

Hansen:1997:MFF

[HN97a] P. J. Hansen and T. G. Nielsen. Mixotrophic feeding of Fragilidium subglobosum (Dinophyceae) on three species of Ceratium: effects of prey concentration, prey species and light intensity. Marine Ecology Progress Series, 147:187–196, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
147/m147p187.pdf.

Holmer:1997:SSD

[HN97b] M. Holmer and S. L. Nielsen. Sediment sulfur dynamics related to biomass-density patterns in Zostera marina (eelgrass) beds. Marine Ecology Progress Series, 146:163–171, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
146/m146p163.pdf.

Hood:1992:APB

[HNC92] R. R. Hood, S. Neuer, and T. J. Cowles. Autotrophic production, biomass and species composition at two stations across an upwelling front. Marine Ecology Progress Series, 83:221– 232, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/83/m083p221.pdf.

Huntley:1994:ECM

[HNL94] M. E. Huntley, W. Nordhausen, and M. D. G. Lopez. Elemental composition, metabolic activity and growth of Antarctic krill Euphausia superba during winter. Marine Ecology Progress Series, 107:23–40, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/107/m107p023.pdf.

Hondeveld:1994:TSV

[HNvB94] B. J. M. Hondeveld, G. Nieuwland, F. C. van Duyl, and R. P. M. Bak. Temporal and spatial variations in heterotrophic nanoflagellate abundance in North Sea sediments. Marine Ecology Progress Series, 109:235–243, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
109/m109p235.pdf.

Hobson:1993:TRA

[Hob93] K. A. Hobson. Trophic relationships among high Arctic seabirds: insights from tissue-dependent stable-isotope models. Marine Ecology Progress Series, 95:7–18, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
95/m095p007.pdf.

Hoeksema:1991:CBM

[Hoe91] B. W. Hoeksema. Control of bleaching in mushroom coral populations (Scleractinia: Fungiidae) in the Java Sea: stress tolerance and interference by life history strategy. Marine Ecology Progress Series, 74:225–237, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
74/m074p225.pdf.

Holm:1990:EDD

[Hol90] E. R. Holm. Effects of density-dependent mortality on the relationship between recruitment and larval settlement. Marine Ecology Progress Series, 60:141–146, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
60/m060p141.pdf.

Holmer:1996:CFD

[Hol96] M. Holmer. Composition and fate of dissolved organic carbon derived from phytoplankton detritus in coastal marine sediments. Marine Ecology Progress Series, 141:217–228, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
141/m141p217.pdf.

Holmquist:1997:DGF

[Hol97] J. G. Holmquist. Disturbance and gap formation in a marine benthic mosaic: influence of shifting macroalgal patches on seagrass structure and mobile invertebrates. Marine Ecology Progress Series, 158:121–130, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/158/m158p121.pdf.

Hunte:1995:DRA

[HOM95] W. Hunte, H. A. Oxenford, and R. Mahon. Distribution and relative abundance of flyingfish (Exocoetidae) in the eastern Caribbean. II. Spawning substrata, eggs and larvae. Marine Ecology Progress Series, 117:25–37, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/117/m117p025.pdf.

Hovenkamp:1992:GDM

[Hov92] F. Hovenkamp. Growth-dependent mortality of larval plaice Pleuronectes platessa in the North Sea. Marine Ecology Progress Series, 82:95–101, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/82/m082p095.pdf.

Hapette:1990:VVC

[HP90] A. M. Hapette and S. A. Poulet. Variation of vitamin C in some common species of marine plankton. Marine Ecology Progress Series, 64:69–79, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/64/m064p069.pdf.

Haywood:1996:NDS

[HP96] M. D. E. Haywood and R. C. Pendrey. A new design for a submersible chronographic tethering device to record predation in different habitats. Marine Ecology Progress Series, 143:307–312, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/143/m143p307.pdf.

Harding:1997:LTI

[HP97] L. W. Harding, Jr. and E. S. Perry. Long-term increase of phytoplankton biomass in Chesapeake Bay, 1950–1994. Marine Ecology Progress Series, 157:39–52, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/157/m157p039.pdf.

Hymel:1998:IMB

[HP98] S. N. Hymel and C. J. Plante. Improved method of bacterial enumeration in sandy and deposit-feeder gut sediments using the fluorescent stain 4,6-diamidino-2-phenylindole (DAPI). Marine Ecology Progress Series, 173:299–304, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
173/m173p299.pdf.

Hily:1992:SSA

[HPF92] C. Hily, P. Potin, and J.-Y. Floch. Structure of subtidal algal assemblages on soft-bottom sediments: fauna/flora interactions and role of disturbances in the Bay of Brest, France. Marine Ecology Progress Series, 85:115–130, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
85/m085p115.pdf.

Heeger:1992:PJC

[HPM92] T. Heeger, U. Piatkowski, and H. Moller. Predation on jellyfish by the cephalopod Argonauta argo. Marine Ecology Progress Series, 88:293–296, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/88/m088p293.pdf.

Haldorson:1993:VDM

[HPPZ93] L. Haldorson, M. Prichett, A. J. Paul, and D. Ziemann. Vertical distribution and migration of fish larvae in a Northeast Pacific bay. Marine Ecology Progress Series, 101:67–80, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
101/m101p067.pdf.

Hofstraat:1994:SDP

[HPSG94] J. W. Hofstraat, J. C. H. Peeters, J. F. H. Snel, and C. Geel. Simple determination of photosynthetic efficiency and photoinhibition of Dunalielia tertiolecta by saturating pulse fluorescence measurements. Marine Ecology Progress Series, 103:187–196, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/103/m103p187.pdf.

Hein:1995:SDN

[HPSJ95] M. Hein, M. F. Pedersen, and K. Sand-Jensen. Size-dependent nitrogen uptake in micro- and macroalgae. Marine Ecology Progress Series, 118:247–253, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/118/m118p247.pdf.

Harder:1999:ILA

[HQ99] T. Harder and P. Y. Qian. Induction of larval attachment and metamorphosis in the serpulid polychaete Hydroides elegans by dissolved free amino acids: isolation and identification. Marine Ecology Progress Series, 179:259–271, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
179/m179p259.pdf.

Hart:1996:RHFc

[HR96a] A. M. Hart and G. R. Russ. Response of herbivorous fishes to crown-of-thorns starfish Acanthaster planci outbreaks. III. Age, growth, mortality and maturity indices of Acanthurus nigrofuscus. Marine Ecology Progress Series, 136:25–35, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
136/m136p025.pdf.

Havskum:1996:EIB

[HR96b] H. Havskum and B. Riemann. Ecological importance of bacterivorous, pigmented flagellates (mixotrophs) in the Bay of Aarhus, Denmark. Marine Ecology Progress Series, 137:251– 263, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/137/m137p251.pdf.

Heilmann:1994:ADA

[HRÆ94] J. P. Heilmann, K. Richardson, and G. Ærtebjerg. Annual distribution and activity of phytoplankton in the Skagerrak/Kattegat frontal region. Marine Ecology Progress Series, 112:213–223, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/112/m112p213.pdf.

Hunt:1998:CFE

[HRCW98] G. L. Hunt, Jr., R. W. Russell, K. O. Coyle, and T. Weingartner. Comparative foraging ecology of planktivorous auklets in relation to ocean physics and prey availability. Marine Ecology Progress Series, 167:241–259, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/167/m167p241.pdf.

Hoffmann:1992:PGB

[HRD92] L. Hoffmann, R. Renard, and V. Demoulin. Phenology, growth and biomass of Cystoseira balearica in Calvi (Corsica). Marine Ecology Progress Series, 80:249–254, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
80/m080p249.pdf.

Hymanson:1990:VUM

[HRFC90] Z. P. Hymanson, D. C. Reed, M. S. Foster, and J. W. Carter. The validity of using morphological characteristics as predictors of age in the kelp Pterygophora californica (Laminanales, Phaeophyta). Marine Ecology Progress Series, 59: 295–304, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/59/m059p295.pdf.

Hansen:1993:PBE

[HRKR93] F. C. Hansen, M. Reckermann, W. C. M. Klein Breteler, and R. Riegman. Phaeocystis blooming enhanced by copepod predation on protozoa: evidence from incubation experiments. Marine Ecology Progress Series, 102:51–57, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
102/m102p051.pdf.

Hagerman:1990:AMS

[HS90] L. Hagerman and A. Szaniawska. Anaerobic metabolic strategy of the glacial relict isopod Saduria (Mesidotea) entomon. Marine Ecology Progress Series, 59:91–96, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
59/m059p091.pdf.

Hoepffner:1991:EPC

[HS91a] N. Hoepffner and S. Sathyendranath. Effect of pigment composition on absorption properties of phytoplankton. Marine Ecology Progress Series, 73:11–23, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/73/m073p011.pdf.

Hoffmann:1991:BAM

[HS91b] A. J. Hoffmann and B. Santelices. Banks of algal microscopic forms: hypotheses on their functioning and comparisons with seed banks. Marine Ecology Progress Series, 79: 185–194, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/79/m079p185.pdf.

Horrocks:1991:NSL

[HS91c] J. A. Horrocks and N. McA. Scott. Nest site location and nest success in the hawksbill turtle Eretmochelys imbricate in Barbados, West Indies. Marine Ecology Progress Series, 69:1–8, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/69/m069p001.pdf.

Hatch:1992:PSJ

[HS92] S. A. Hatch and G. A. Sanger. Puffins as samplers of juvenile pollock and other forage fish in the Gulf of Alaska. Marine Ecology Progress Series, 80:1–14, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/80/m080p001.pdf.

Hagerman:1994:HNC

[HS94a] L. Hagerman and A. Szaniawska. Haemolymph nitrogen compounds and ammonia efflux rates under anoxia in the brackish water isopod Saduria entomon. Marine Ecology Progress Series, 103:285–289, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/103/m103p285.pdf.

Haney:1994:EVF

[HS94b] J. Christopher Haney and Amy E. S. Schauer. Environmental variability facilitates coexistence within an alcid community at sea. Marine Ecology Progress Series, 103:221–237, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
103/m103p221.pdf.

Hunt:1995:SDM

[HS95] H. L. Hunt and R. E. Scheibling. Structure and dynamics of mussel patches in tidepools on a rocky shore in Nova Scotia, Canada. Marine Ecology Progress Series, 124:105–115, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
124/m124p105.pdf.

Hunt:1996:PBF

[HS96] H. L. Hunt and R. E. Scheibling. Physical and biological factors influencing mussel (Mytilus trossulus, M. edulis) settlement on a wave-exposed rocky shore. Marine Ecology Progress Series, 142:135–145, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/142/m142p135.pdf.

Hillebrand:1997:REM

[HS97a] H. Hillebrand and U. Sommer. Response of epilithic microphytobenthos of the Western Baltic Sea to in situ experiments with nutrient enrichment. Marine Ecology Progress Series, 160:35–46, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/160/m160p035.pdf.

Hirst:1997:LSW

[HS97b] A. G. Hirst and M. Sheader. Are in situ weight-specific growth rates body-size independent in marine planktonic copepods? A re-analysis of the global syntheses and a new empirical model. Marine Ecology Progress Series, 154:155–165, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
154/m154p155.pdf.

Hunt:1997:REP

[HS97c] H. L. Hunt and R. E. Scheibling. Role of early post-settlement mortality in recruitment of benthic marine invertebrates. Marine Ecology Progress Series, 155:269–301, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
155/m155p269.pdf.

Hannach:1998:PUA

[HS98a] G. Hannach and A. C. Sigleo. Photoinduction of UVabsorbing compounds in six species of marine phytoplankton. Marine Ecology Progress Series, 174:207–222, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
174/m174p207.pdf.

Hopkins:1998:MFS

[HS98b] T. L. Hopkins and T. T. Sutton. Midwater fishes and shrimps as competitors and resource partitioning in low latitude oligotrophic ecosystems. Marine Ecology Progress Series, 164: 37–45, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/164/m164p037.pdf.

Hunt:1998:STV

[HS98c] H. L. Hunt and R. E. Scheibling. Spatial and temporal variability of patterns of colonization by mussels (Mytilus trossulus, M. edulis) on a wave-exposed rocky shore. Marine Ecology Progress Series, 167:155–169, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/167/m167p155.pdf.

Hickford:1999:EPL

[HS99a] M. J. H. Hickford and D. R. Schiel. Evaluation of the performance of light traps for sampling fish larvae in inshore temperate waters. Marine Ecology Progress Series, 186:293– 302, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/186/m186p293.pdf.

Horinouchi:1999:ECS

[HS99b] M. Horinouchi and M. Sano. Effects of changes in seagrass shoot density and leaf height on abundances and distribution patterns of juveniles of three gobiid fishes in a Zostera marina bed. Marine Ecology Progress Series, 183:87–94, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
183/m183p087.pdf.

Hussy:1997:FRU

[HSB97] K. Hüssy, M. A. St. John, and U. Böttcher. Food resource utilization by juvenile Baltic cod Gadus morhua: a mechanism potentially influencing recruitment success at the demersal juvenile stage? Marine Ecology Progress Series, 155: 199–208, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/155/m155p199.pdf.

Hawkins:1996:NOU

[HSBH96] A. J. S. Hawkins, R. F. M. Smith, B. L. Bayne, and M. Héral. Novel observations underlying the fast growth of suspension-feeding shellfish in turbid environments: Mytilus edulis. Marine Ecology Progress Series, 131:179–190, 1996.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
131/m131p179.pdf.

Hummel:1997:VGT

[HSBP97] H. Hummel, A. Sommer, R. H. Bogaards, and H. O. Pörtner. Variation in genetic traits of the lugworm Arenicola marina: temperature related expression of mitochondrial allozymes? Marine Ecology Progress Series, 159:189–195, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
159/m159p189.pdf.

Hoch:1996:SID

[HSCC96] M. P. Hoch, R. A. Snyder, L. A. Cifuentes, and R. B. Coffin. Stable isotope dynamics of nitrogen recycled during interactions among marine bacteria and protists. Marine Ecology Progress Series, 132:229–239, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/132/m132p229.pdf.

Harriott:1994:PCC

[HSH94] V. J. Harriott, S. D. A. Smith, and P. L. Harrison. Patterns of coral community structure of subtropical reefs in the Solitary Islands Marine Reserve, Eastern Australia. Marine Ecology Progress Series, 109:67–76, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/109/m109p067.pdf.

Hsieh:1995:STP

[Hsi95] H. L. Hsieh. Spatial and temporal patterns of polychaete communities in a subtropical mangrove swamp: influences of sediment and microhabitat. Marine Ecology Progress Series, 127:157–167, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/127/m127p157.pdf.

Hsieh:1997:SFP

[Hsi97] H. L. Hsieh. Self-fertilization: a potential fertilization mode in an estuarine sabellid polychaete. Marine Ecology Progress Series, 147:143–148, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/147/m147p143.pdf.

Hayashibara:1993:PCS

[HSK^{+}93] T. Hayashibara, K. Shimoike, T. Kimura, S. Hosaka, A. Heyward, P. Harrison, K. Kudo, and M. Omori. Patterns of coral spawning at Akajima Island, Okinawa, Japan. Marine Ecology Progress Series, 101:253–262, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/101/m101p253.pdf.

Hemminga:1994:COM

[HSK^{+}94] M. A. Hemminga, F. J. Slim, J. Kazungu, G. M. Ganssen, J. Nieuwenhuize, and N. M. Kruyt. Carbon outwelling from a mangrove forest with adjacent seagrass beds and coral reefs (Gazi Bay, Kenya). Marine Ecology Progress Series, 106: 291–301, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/106/m106p291.pdf.

Haimovici:1994:DBF

[HSLV94] Manuel Haimovici, Agnaldo Silva Martins, José Lima de Figueiredo, and P. C. Vieira. Demersal bony fish of the outer shelf and upper slope of the southern Brazil subtropical convergence ecosystem. Marine Ecology Progress Series, 108:59–77, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/108/m108p059.pdf.

Hamm:1999:CPG

[HSMS99] C. E. Hamm, D. A. Simson, R. Merkel, and V. Smetacek. Colonies of Phaeocystis globosa are protected by a thin but tough skin. Marine Ecology Progress Series, 187:101–111, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/187/m187p101.pdf.

Hopkins:1993:TLC

[HSN93] C. C. E. Hopkins, J. R. Sargent, and E. M. Nilssen. Total lipid content, and lipid and fatty acid composition of the deep-water prawn Pandulus borealis from Balsfjord, northern Norway: growth and feeding relationships. Marine Ecology Progress Series, 96:217–228, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/96/m096p217.pdf.

Hewes:1990:MAH

[HSRHH90] C. D. Hewes, E. Sakshaug, F. M. H. Reid, and O. Holm-Hansen. Microbial autotrophic and heterotrophic eucaryotes in Antarctic waters: relationships between biomass and chlorophyll, adenosine triphosphate and particulate organic carbon. Marine Ecology Progress Series, 63:27–35, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
63/m063p027.pdf.

Hawkins:1998:SFB

[HSTY98] A. J. S. Hawkins, R. F. M. Smith, S. H. Tan, and Z. B. Yasin. Suspension-feeding behaviour in tropical bivalve molluscs: Perna viridis, Crassostrea belcheri, Crassostrea iradelei, Saccostrea cucculata and Pinctada margarifera. Marine Ecology Progress Series, 166:173–185, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/166/m166p173.pdf.

Hirst:1999:APC

[HSW99] A. G. Hirst, M. Sheader, and J. A. Williams. Annual pattern of calanoid copepod abundance, prosome length and minor role in pelagic carbon flux in the Solent, UK. Marine Ecology Progress Series, 177:133–146, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/177/m177p133.pdf.

Hills:1996:MSA

[HT96] J. M. Hills and J. C. Thomason. A multi-scale analysis of settlement density and pattern dynamics of the barnacle Semibalanus balanoides. Marine Ecology Progress Series, 138:103–115, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/138/m138p103.pdf.

Hewitt:1996:MPP

[HTCP96] J. E. Hewitt, S. F. Thrush, V. J. Cummings, and R. D. Pridmore. Matching patterns with processes: predicting the effect of size and mobility on the spatial distributions of the bivalves Macomona liliana and Austrovenus stutchburyi. Marine Ecology Progress Series, 135:57–67, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/135/m135p057.pdf.

Hill:1996:CAK

[HTCW96] H. J. Hill, P. N. Trathan, J. P. Croxall, and J. L. Watkins. A comparison of Antarctic krill Euphausia superba caught by nets and taken by macaroni penguins Eudyptes chrysolophus: evidence for selection? Marine Ecology Progress Series, 140:1–11, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/140/m140p001.pdf.

Heiskanen:1996:IPF

[HTG96] A. S. Heiskanen, T. Tamminen, and K. Gundersen. Impact of planktonic food web structure on nutrient retention and loss from a late summer pelagic system in the coastal northern Baltic Sea. Marine Ecology Progress Series, 145: 195–208, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/145/m145p195.pdf.

Holmes:1999:PLC

[HTLK99] M. J. Holmes, S. L. M. Teo, F. C. Lee, and H. W. Khoo. Persistent low concentrations of diarrhetic shellfish toxins in green mussels Perna viridis from the Johor Strait, Singapore: first record of diarrhetic shellfish toxins from South-East Asia. Marine Ecology Progress Series, 181:257–268, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
181/m181p257.pdf.

Haigh:1992:PES

[HTS92] R. Haigh, F. J. R. Taylor, and T. F. Sutherland. Phytoplankton ecology of Sechelt Inlet, a fjord system on the British Columbia coast. I. General features of the nano- and microplankton. Marine Ecology Progress Series, 89:117–134, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/89/m089p117.pdf.

Hull:1997:SCD

[Hul97] S. L. Hull. Seasonal changes in diversity and abundance of ostracods on four species of intertidal algae with differing structural complexity. Marine Ecology Progress Series, 161: 71–82, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/161/m161p071.pdf.

Hull:1999:CTP

[Hul99] S. L. Hull. Comparison of tidepool phytal ostracod abundance and assemblage structure on three spatial scales. Marine Ecology Progress Series, 182:201–208, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/182/m182p201.pdf.

Hunt:1993:ECP

[Hun93] Alison Hunt. Effects of contrasting patterns of larval dispersal on the genetic connectedness of local populations of two intertidal starfish, Patiriella calcar and P. exigua. Marine Ecology Progress Series, 92:179–186, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/92/m092p179.pdf.

Hurlbut:1992:LRS

[Hur92] C. J. Hurlbut. Larval release and supply predict temporal variation in settlement of a colonial ascidian. Marine Ecology Progress Series, 80:215–219, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/80/m080p215.pdf.

Huttel:1990:ILA

[Hüt90] M. Hüttel. Influence of the lugworm Arenicola marina on porewater nutrient profiles of sand flat sediments. Marine Ecology Progress Series, 62:241–248, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
62/m062p241.pdf.

Harvey:1991:SVL

[HV91] M. Harvey and B. Vincent. Spatial variability of lengthspecific production in shell, somatic tissue and sexual products of Macoma balthica in the Lower St. Lawrence Estuary. I. Small and meso scale variability. Marine Ecology Progress Series, 75:55–66, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/75/m075p055.pdf.

Hansen:1991:GPC

[HvB91] F. C. Hansen and W. H. M. van Boekel. Grazing pressure of the calanoid copepod Temora longicornis on a Phaeocystis dominated spring bloom in a Dutch tidal inlet. Marine Ecology Progress Series, 78:123–129, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/78/m078p123.pdf.

Honkoop:1997:ROM

[HvdM97] P. J. C. Honkoop and J. van der Meer. Reproductive output of Macoma balthica populations in relation to wintertemperature and intertidal-height mediated changes of body mass. Marine Ecology Progress Series, 149:155–162, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
149/m149p155.pdf.

Honkoop:1998:DTI

[HvdMBK98] P. J. C. Honkoop, J. van der Meer, J. J. Beukema, and D. Kwast. Does temperature-influenced egg production predict the recruitment in the bivalve Macoma balthica? Marine Ecology Progress Series, 164:229–235, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
164/m164p229.pdf.

Helbling:1992:INU

[HVFHH92] E. W. Helbling, V. Villafane, M. Ferrario, and O. Holm-Hansen. Impact of natural ultraviolet radiation on rates of photosynthesis and on specific marine phytoplankton species. Marine Ecology Progress Series, 80:89–100, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
80/m080p089.pdf.

Hagen:1996:SLS

[HVK96] W. Hagen, E. S. Van Vleet, and G. Kattner. Seasonal lipid storage as overwintering strategy of Antarctic krill. Marine Ecology Progress Series, 134:85–89, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/134/m134p085.pdf.

Heales:1996:FOM

[HVL96] D. S. Heales, D. J. Vance, and N. R. Loneragan. Field observations of moult cycle, feeding behaviour, and diet of small juvenile tiger prawns Penaeus semisulcatus in the Embley River, Australia. Marine Ecology Progress Series, 145:43–51, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
145/m145p043.pdf.

Hovenkamp:1991:GOG

[HW91] F. Hovenkamp and J. J. Witte. Growth, otolith growth and RNA/DNA ratios of larval plaice Pleuronectes platessa in the North Sea 1987 to 1989. Marine Ecology Progress Series, 70:105–116, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/70/m070p105.pdf.

Hobson:1992:DTR

[HW92] K. A. Hobson and H. E. Welch. Determination of trophic relationships within a high Arctic marine food web using δ 13C and δ ^{15}N analysis. Marine Ecology Progress Series, 84: 9–18, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/84/m084p009.pdf.

Helmke:1995:BSI

[HW95] E. Helmke and H. Weyland. Bacteria in sea ice and underlying water of the eastern Weddell Sea in midwinter. Marine Ecology Progress Series, 117:269–287, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/117/m117p269.pdf.

Hochscheid:1999:NMD

[HW99] S. Hochscheid and R. P. Wilson. A new method for the determination of at-sea activity in sea turtles. Marine Ecology Progress Series, 185:293–296, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/185/m185p293.pdf.

Hall:1998:SVD

[HWH98] A. J. Hall, J. Watkins, and P. S. Hammond. Seasonal variation in the diet of harbour seals in the south-western North Sea. Marine Ecology Progress Series, 170:269–281, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
170/m170p269.pdf.

Hays:1996:LTC

[HWL96] G. C. Hays, A. J. Warner, and D. Lefevre. Long-term changes in the diel vertical migration behaviour of zooplankton. Marine Ecology Progress Series, 141:149–159, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
141/m141p149.pdf.

Hartig:1998:PAN

[HWLC98] P. Hartig, K. Wolfstein, S. Lippemeier, and F. Colijn. Photosynthetic activity of natural microphytobenthos populations measured by fluorescence (PAM) and ^{14}C-tracer methods: a comparison. Marine Ecology Progress Series, 166:53–62, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
166/m166p053.pdf.

Iribarne:1997:BHD

[IBB97] O. Iribarne, A. Bortolus, and F. Botto. Between-habitat differences in burrow characteristics and trophic modes in the southwestern Atlantic burrowing crab Chasmagnathus granulata. Marine Ecology Progress Series, 155:137–145, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
155/m155p137.pdf.

Irigoien:1996:GCR

[ICG96] X. Irigoien, J. Castel, and S. Gasparini. Gut clearance rate as predictor of food limitation situations. Application to two estuarine copepods: Acartia bifilosa and Eurytemora affinis. Marine Ecology Progress Series, 131:159–163, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
131/m131p159.pdf.

Isaksen:1996:SRR

[IF96] M. F. Isaksen and K. Finster. Sulphate reduction in the root zone of the seagrass Zostera noltii on the intertidal flats of a coastal lagoon (Arcachon, France). Marine Ecology Progress Series, 137:187–194, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/137/m137p187.pdf.

Iverson:1997:FAS

[IFL97] S. J. Iverson, K. J. Frost, and L. F. Lowry. Fatty acid signatures reveal fine scale structure of foraging distribution of harbor seals and their prey in Prince William Sound, Alaska. Marine Ecology Progress Series, 151:255–271, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
151/m151p255.pdf.

Israel:1996:PIC

[IG96] A. A. Israel and E. L. González. Photosynthesis and inorganic carbon utilization in Pleurochrysis sp. (Haptophyta), a coccolithophorid alga. Marine Ecology Progress Series, 137: 243–250, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/137/m137p243.pdf.

Irigoien:1998:HFT

[IHK^{+}98] Xabier Irigoien, Robert Head, Ulrike Klenke, Bettina Meyer-Harms, Derek Harbour, Barbara Niehoff, Hans-Jürgen Hirche, and Roger Harris. A high frequency time series at Weathership M , Norwegian Sea, during the 1997 spring bloom: feeding of adult female Calanus finmarchicus. Marine Ecology Progress Series, 172:127–137, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
172/m172p127.pdf.

Butler:1999:CCO

[IMB99] M. J. Butler IV, A. B. MacDiarmid, and J. D. Booth. The cause and consequence of ontogenetic changes in social aggregation in New Zealand spiny lobsters. Marine Ecology Progress Series, 188:179–191, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/188/m188p179.pdf.

Iriarte:1994:SDC

[IP94] A. Iriarte and D. A. Purdie. Size distribution of chlorophylla biomass and primary production in a temperate estuary (Southampton Water): the contribution of photosynthetic picoplankton. Marine Ecology Progress Series, 115:283–297, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/115/m115p283.pdf.

Iglesias-Prieto:1994:AAI

[IPT94] R. Iglesias-Prieto and R. K. Trench. Acclimation and adaptation to irradiance in symbiotic dinoflagellates. I. Responses of the photosynthetic unit to changes in photon flux density. Marine Ecology Progress Series, 113:163–175, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
113/m113p163.pdf.

Iribarne:1990:USS

[Iri90] O. O. Iribarne. Use of shelter by the small Patagonian octopus Octopus tehuelchus: availability, selection and effects on fecundity. Marine Ecology Progress Series, 66:251–258, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
66/m066p251.pdf.

Irwin:1991:CMP

[Irw91] B. Irwin. Coulometric measurement of primary production, with comparison against dissolved oxygen and ^{14}C methods in a seasonal study. Marine Ecology Progress Series, 71: 97–102, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/71/m071p097.pdf.

Ichii:1998:ICB

[ISF^{+}98] T. Ichii, N. Shinohara, Y. Fujise, S. Nishiwaki, and K. Matsuoka. Interannual changes in body fat condition index of minke whales in the Antarctic. Marine Ecology Progress Series, 175:1–12, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/175/m175p001.pdf.

Ishikawa:1996:CBC

[IT96] A. Ishikawa and A. Taniguchi. Contribution of benthic cysts to the population dynamics of Scrippsiella spp. (Dinophyceae) in Onagawa Bay, northeast Japan. Marine Ecology Progress Series, 140:169–178, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/140/m140p169.pdf.

Ito:1997:ESN

[Ito97] Kenjoi Ito. Egg-size and -number variations related to maternal size and age, and the relationship between egg size and larval characteristics in an annual marine gastropod, Haloa japonica (Opisthobranchia; Cephalaspidea). Marine Ecology Progress Series, 152:187–195, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/152/m152p187.pdf.

Iribarne:1998:SPM

[IVM^{+}98] O. Iribarne, J. Valero, M. M. Martinez, L. Lucifora, and S. Bachmann. Shorebird predation may explain the origin of Holocene beds of stout razor clams in life position. Marine Ecology Progress Series, 167:301–306, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
167/m167p301.pdf.

Incze:1991:RPE

[IW91] L. S. Incze and R. A. Wahle. Recruitment from pelagic to early benthic phase in lobsters Homarus americanus. Marine Ecology Progress Series, 79:77–87, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/79/m079p077.pdf.

Ide:1997:AOC

[IWF^{+}97] I. Ide, E. P. Witten, J. Fischer, W. Kalbfus, A. Zellner, E. Stroben, and B. Watermann. Accumulation of organotin compounds in the common whelk Buccinum undatum and the red whelk Neptunea antiqua in association with imposex. Marine Ecology Progress Series, 152:197–203, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
152/m152p197.pdf.

Jensen:1991:IZA

[JA91] Gregory C. Jensen and David A. Armstrong. Intertidal zonation among congeners: factors regulating distribution of porcelain crabs Petrolisthes spp. (Anomura. Porcellanidae). Marine Ecology Progress Series, 73:47–60, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
73/m073p047.pdf.

Jensen:1992:BRK

[JAB^{+}92] P. Jensen, I. Aagaard, R. A. Burke, Jr., P. R. Dando, N. O. Jørgensen, A. Kuijpers, T. Laier, S. C. M. O’Hara, and R. Schmaljohann. Bubbling reefs in the Kattegat: submarine landscapes of carbonate-cemented rocks support a diverse ecosystem at methane seeps. Marine Ecology Progress Series, 83:103–112, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/83/m083p103.pdf.

Jonsson:1991:SBM

[JAL91] P. R. Jonsson, C. Andre, and M. Lindegarth. Swimming behaviour of marine bivalve larvae in a flume boundary-layer flow: evidence for near-bottom confinement. Marine Ecology Progress Series, 79:67–76, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/79/m079p067.pdf.

Johnson:1990:STP

[JAOS90] W. S. Johnson, D. M. Allen, M. V. Ogburn, and S. E. Stancyk. Short-term predation responses of adult bay anchovies Anchoa mitchilli to estuarine zooplankton availability. Marine Ecology Progress Series, 64:55–68, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/64/m064p055.pdf.

Jorgensen:1990:AMM

[JBB90] B. B. Jørgensen, M. Bang, and T. H. Blackburn. Anaerobic mineralization in marine sediments from the Baltic Sea-North Sea transition. Marine Ecology Progress Series, 59: 39–54, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/59/m059p039.pdf.

Jaramillo:1992:MBE

[JBB92] E. Jaramillo, C. Bertran, and A. Bravo. Mussel biodeposition in an estuary in southern Chile. Marine Ecology Progress Series, 82:85–94, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/82/m082p085.pdf.

Jansen:1998:FMC

[JBB98] J. K. Jansen, P. L. Boveng, and J. L. Bengtson. Foraging modes of chinstrap penguins: contrasts between day and night. Marine Ecology Progress Series, 165:161–172, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
165/m165p161.pdf.

Johnson:1998:IBS

[JBH98] M. P. Johnson, M. T. Burrows, and S. J. Hawkins. Individual based simulations of the direct and indirect effects of limpets on a rocky shore Fucus mosaic. Marine Ecology Progress Series, 169:179–188, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/169/m169p179.pdf.

Johnson:1997:SSM

[JBHH97] M. P. Johnson, M. T. Burrows, R. G. Hartnoll, and S. J. Hawkins. Spatial structure on moderately exposed rocky shores: patch scales and the interactions between limpets and algae. Marine Ecology Progress Series, 160:209–215, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
160/m160p209.pdf.

Jenkins:1999:RPT

[JBK99] G. P. Jenkins, K. P. Black, and M. J. Keough. The role of passive transport and the influence of vertical migration on the pre-settlement distribution of a temperate, demersal fish: numerical model predictions compared with field sampling. Marine Ecology Progress Series, 184:259–271, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
184/m184p259.pdf.

Jones:1997:RBC

[JBO97] R. J. Jones, R. Berkelmans, and J. K. Oliver. Recurrent bleaching of corals at Magnetic Island (Australia) relative to air and seawater temperature. Marine Ecology Progress Series, 158:289–292, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/158/m158p289.pdf.

Jenkins:1997:TSV

[JBWH97] G. P. Jenkins, K. P. Black, M. J. Wheatley, and D. N. Hatton. Temporal and spatial variability in recruitment of a temperate, seagrass-associated fish is largely determined by physical processes in the pre- and post-settlement phases. Marine Ecology Progress Series, 148:23–35, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/148/m148p023.pdf.

Juinio:1992:NDF

[JC92] M. A. R. Juinio and J. S. Cobb. Natural diet and feeding habits of the postlarval lobster Homarus americanus. Marine Ecology Progress Series, 85:83–91, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/85/m085p083.pdf.

Juanes:1994:PPS

[JC94] F. Juanes and D. O. Conover. Piscivory and prey size selection in young-of-the-year bluefish: predator preference or size-dependent capture success? Marine Ecology Progress Series, 114:59–69, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/114/m114p059.pdf.

Juanes:1995:SSP

[JC95] F. Juanes and D. O. Conover. Size-structured piscivory: advection and the linkage between predator and prey recruitment in young-of-the-year bluefish. Marine Ecology Progress Series, 128:287–304, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/128/m128p287.pdf.

Josefson:1997:BRP

[JC97] A. B. Josefson and D. J. Conley. Benthic response to a pelagic front. Marine Ecology Progress Series, 147:49–62, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
147/m147p049.pdf.

Jouventin:1994:EPR

[JCCCB94] P. Jouventin, D. Capdeville, R. Cuenot-Chaillet, and C. Boiteau. Exploitation of pelagic resources by a non-flying seabird: satellite tracking of the king penguin throughout the breeding cycle. Marine Ecology Progress Series, 106: 11–19, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/106/m106p011.pdf.

Judge:1993:DMM

[JCH93] M. L. Judge, L. D. Coen, and K. L. Heck, Jr. Does Mercenaha mercenaha encounter elevated food levels in seagrass beds? Results from a novel technique to collect suspended food resources. Marine Ecology Progress Series, 92:141–150, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
92/m092p141.pdf.

Jordan:1991:LTT

[JCMW91] T. E. Jordan, D. L. Correll, J. Miklas, and D. E. Weller. Longterm trends in estuarine nutrients and chlorophyll, and shortterm effects of variation in watershed discharge. Marine Ecology Progress Series, 75:121–132, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/75/m075p121.pdf.

Jassby:1993:OCS

[JCP93] A. D. Jassby, J. E. Cloern, and T. M. Powell. Organic carbon sources and sinks in San Francisco Bay: variability induced by river flow. Marine Ecology Progress Series, 95:39–54, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
95/m095p039.pdf.

Jenkins:1990:AGR

[JD90] G. P. Jenkins and T. L. O. Davis. Age, growth rate, and growth trajectory determined from otolith microstructure of southern bluefin tuna Thunnus maccoyii larvae. Marine Ecology Progress Series, 63:93–104, 1990. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/63/m063p093.pdf.

Jollivet:1995:EDA

[JDLL95] D. Jollivet, D. Desbruyères, C. Ladrat, and L. Laubier. Evidence for differences in the allozyme thermostability of deepsea hydrothermal vent polychaetes (Alvinellidae): a possible selection by habitat. Marine Ecology Progress Series, 123:125–136, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/123/m123p125.pdf.

Jewett:1999:VEV

[JDSB99] S. C. Jewett, T. A. Dean, R. O. Smith, and A. Blanchard. ‘Exxon Valdez’ oil spill: impacts and recovery in the soft-bottom benthic community in and adjacent to eelgrass beds. Marine Ecology Progress Series, 185:59–83, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
185/m185p059.pdf.

Jones:1995:SBM

[JE95] M. B. Jones and C. E. Epifanio. Settlement of brachyuran megalopae in Delaware Bay: an analysis of time series data. Marine Ecology Progress Series, 125:67–76, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
125/m125p067.pdf.

Jeong:1994:PHD

[Jeo94a] H. J. Jeong. Predation by the heterotrophic dinoflagellate Protoperidinium cf. divergens on copepod eggs and early naupliar stages. Marine Ecology Progress Series, 114:203– 208, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/114/m114p203.pdf.

Jeong:1994:PEC

[Jeo94b] H. J. Jeong. Predation effects of the calanoid copepod Acartia tonsa on a population of the heterotrophic dinoflagellate Protoperidinium cf. divergens in the presence of cooccurring red-tide dinoflagellate prey. Marine Ecology Progress Series, 111:87–97, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/111/m111p087.pdf.

Jernakoff:1990:DNS

[Jer90] P. Jernakoff. Distribution of newly settled western rock lobsters Panulirus cygnus. Marine Ecology Progress Series, 66: 63–74, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/66/m066p063.pdf.

Jonasdottir:1995:CEP

[JFP95] S. H. Jónasdóttir, D. Fields, and S. Pantoja. Copepod egg production in Long Island Sound, USA, as a function of the chemical composition of seston. Marine Ecology Progress Series, 119:87–98, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/119/m119p087.pdf.

Jones:1990:SPA

[JFS90] G. P. Jones, D. J. Ferrell, and P. F. Sale. Spatial pattern in the abundance and structure of mollusc populations in the soft sediments of a coral reef lagoon. Marine Ecology Progress Series, 62:109–120, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/62/m062p109.pdf.

Jones:1992:FFD

[JFS92] G. P. Jones, D. J. Ferrell, and P. F. Sale. Fish feeding and dynamics of soft-sediment mollusc populations in a coral reef lagoon. Marine Ecology Progress Series, 80:175–190, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
80/m080p175.pdf.

Jurgens:1990:IRP

[JG90] Klaus Jürgens and H. Güde. Incorporation and release of phosphorus by planktonic bacteria and phagotrophic flagellates. Marine Ecology Progress Series, 59:271–284, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
59/m059p271.pdf.

Jurgens:1994:PIG

[JG94] K. Jürgens and H. Güde. The potential importance of grazingresistant bacteria in planktonic systems. Marine Ecology Progress Series, 112:169–188, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/112/m112p169.pdf.

Jahn:1996:PAC

[JGT96] A. Jahn, I. Gamenick, and H. Theede. Physiological adaptations of Cyprideis torosa (Crustacea, Ostracoda) to hydrogen sulphide. Marine Ecology Progress Series, 142:215–223, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
142/m142p215.pdf.

Jakobsen:1997:PSS

[JH97] H. H. Jakobsen and P. J. Hansen. Prey size selection, grazing and growth response of the small heterotrophic dinoflagellate Gymnodinium sp. and the ciliate Balanion comatum — a comparative study. Marine Ecology Progress Series, 158: 75–86, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/158/m158p075.pdf.

Jones:1999:ECC

[JHG99] R. J. Jones and O. Hoegh-Guldberg. Effects of cyanide on coral photosynthesis: implications for identifying the cause of coral bleaching and for assessing the environmental effects of cyanide fishing. Marine Ecology Progress Series, 177: 83–91, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/177/m177p083.pdf.

Jenkins:1999:DIE

[JHN99] S. R. Jenkins, S. J. Hawkins, and T. A. Norton. Direct and indirect effects of a macroalgal canopy and limpet grazing in structuring a sheltered inter-tidal community. Marine Ecology Progress Series, 188:81–92, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/188/m188p081.pdf.

Jury:1995:LMR

[JHW95] S. H. Jury, W. H. Howell, and W. H. Watson III. Lobster movements in response to a hurricane. Marine Ecology Progress Series, 119:305–310, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/119/m119p305.pdf.

Jackson:1999:STV

[JJ99] G. Jackson and G. K. Jones. Spatial and temporal variation in nearshore fish and macroinvertebrate assemblages from a temperate Australian estuary over a decade. Marine Ecology Progress Series, 182:253–268, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/182/m182p253.pdf.

Josefson:1995:GPB

[JJNR95] A. B. Josefson, J. N. Jensen, T. G. Nielsen, and B. Rasmussen. Growth parameters of a benthic suspension feeder along a depth gradient across the pycnocline in the southern Kattegat, Denmark. Marine Ecology Progress Series, 125:107–115, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/125/m125p107.pdf.

Jahn:1997:SBS

[JJTS97] A. Jahn, U. Janas, H. Theede, and A. Szaniawska. Significance of body size in sulphide detoxification in the Baltic clam Macoma balthica (Bivalvia, Tellinidae) in the Gulf of Gdansk. Marine Ecology Progress Series, 154:175–183, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
154/m154p175.pdf.

Jorgensen:1993:DFA

[JKC^{+}93] N. O. G. Jørgensen, N. Kroer, R. B. Coffin, X.-H. Yang, and C. Lee. Dissolved free amino acids, combined amino acids, and DNA as sources of carbon and nitrogen to marine bacteria. Marine Ecology Progress Series, 98:135–148, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
98/m098p135.pdf.

Johnson:1995:CFC

[JKFB95] C. Johnson, D. Klumpp, J. Field, and R. Bradbury. Carbon flux on coral reefs: effects of large shifts in community structure. Marine Ecology Progress Series, 126:123–143, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
126/m126p123.pdf.

Johnstone:1990:PAC

[JKL90] R. W. Johnstone, K. Koop, and A. W. D. Larkum. Physical aspects of coral reef lagoon sediments in relation to detritus processing and primary production. Marine Ecology Progress Series, 66:273–283, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/66/m066p273.pdf.

Johannesson:1990:ADR

[JKT90a] K. Johannesson, N. Kautsky, and M. Tedengren. Allozyme differentiation — a reply to Vainola. Marine Ecology Progress Series, 67:309–310, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/67/m067p309.pdf.

Johannesson:1990:GPD

[JKT90b] K. Johannesson, N. Kautsky, and M. Tedengren. Genotypic and phenotypic differences between Baltic and North Sea populations of Mytilus edulis evaluated through reciprocal transplantations. II. Genetic variation. Marine Ecology Progress Series, 59:211–219, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/59/m059p211.pdf.

Jonasdottir:1998:RDC

[JKT^{+}98] S. H. Jónasdóttir, T. Kiørboe, K. W. Tang, M. St. John, A. W. Visser, E. Saiz, and H. G. Dam. Role of diatoms in copepod production: good, harmless or toxic? Marine Ecology Progress Series, 172:305–308, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/172/m172p305.pdf.

Jeong:1994:GGR

[JL94] H. J. Jeong and M. I. Latz. Growth and grazing rates of the heterotrophic dinoflagellates Protoperidinium spp. on red tide dinoflagellates. Marine Ecology Progress Series, 106: 173–185, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/106/m106p173.pdf.

John:1996:LBL

[JL96] M. A. St John and T. Lund. Lipid biomarkers: linking the utilization of frontal plankton biomass to enhanced condition of juvenile North Sea cod. Marine Ecology Progress Series, 131:75–85, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/131/m131p075.pdf.

Jellett:1996:MAB

[JLD^{+}96] J. F. Jellett, W. K. W. Li, P. M. Dickie, A. Boraie, and P. E. Kepkay. Metabolic activity of bacterioplankton communities assessed by flow cytometry and single carbon substrate utilization. Marine Ecology Progress Series, 136:213–225, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
136/m136p213.pdf.

James:1995:SMS

[JLF95] R. J. James, M. P. Lincoln Smith, and P. G. Fairweather. Sieve mesh-size and taxonomic resolution needed to describe natural spatial variation of marine macrofauna. Marine Ecology Progress Series, 118:187–198, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/118/m118p187.pdf.

Juillet-Leclerc:1997:SVP

[JLGMP97] A. Juillet-Leclerc, J. P. Gattuso, L. F. Montaggioni, and M. Pichon. Seasonal variation of primary productivity and skeletal δ ^{13}C and δ ^{18}O in the zooxanthellate scleractinian coral Acropora formosa. Marine Ecology Progress Series, 157:109–117, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/157/m157p109.pdf.

Jorgensen:1990:ETM

[JLR90] C. B. Jørgensen, P. S. Larsen, and H. U. Riisgård. Effects of temperature on the mussel pump. Marine Ecology Progress Series, 64:89–97, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/64/m064p089.pdf.

Jensen:1990:BNN

[JLS90] M. H. Jensen, E. Lomstein, and J. Sørensen. Benthic NH_{4}^{+} and NO_{3}-flux following sedimentation of a spring phytoplankton bloom in Aarhus Bight, Denmark. Marine Ecology Progress Series, 61:87–96, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/61/m061p087.pdf.

Jeong:1997:FCM

[JLYK97] H. J. Jeong, C. W. Lee, W. H. Yih, and J. S. Kim. Fragilidium cf. mexicanum, a thecate mixotrophic dinoflagellate which is prey for and a predator on co-occurring thecate heterotrophic dinoflagellate Protoperidinium cf. divergens. Marine Ecology Progress Series, 151:299–305, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/151/m151p299.pdf.

Jokiel:1993:WMC

[JM93] P. L. Jokiel and J. I. Morrissey. Water motion on coral reefs: evaluation of the ’clod card’ technique. Marine Ecology Progress Series, 93:175–181, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/93/m093p175.pdf.

Jochem:1999:CMD

[JM99] F. J. Jochem and D. Meyerdierks. Cytometric measurement of the DNA cell cycle in the presence of chlorophyll autofluorescence in marine eukaryotic phytoplankton by the bluelight excited dye YOYO-1. Marine Ecology Progress Series, 185:301–307, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/185/m185p301.pdf.

Jaramillo:1993:IZP

[JMC93] E. Jaramillo, A. McLachlan, and P. Coetzee. Intertidal zonation patterns of macroinfauna over a range of exposed sandy beaches in south-central Chile. Marine Ecology Progress Series, 101:105–118, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/101/m101p105.pdf.

Jaramillo:1995:TSA

[JMD95] E. Jaramillo, A. McLachlan, and J. Dugan. Total sample area and estimates of species richness in exposed sandy beaches. Marine Ecology Progress Series, 119:311–314, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
119/m119p311.pdf.

Jeffrey:1999:OUU

[JMD^{+}99] S. W. Jeffrey, H. S. MacTavish, W. C. Dunlap, M. Vesk, and K. Groenewoud. Occurrence of UVA- and UVB-absorbing compounds in 152 species (206 strains) of marine microalgae. Marine Ecology Progress Series, 189:35–51, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
189/m189p035.pdf.

Jones:1990:SCD

[JMP90] T. W. Jones, T. C. Malone, and S. Pike. Seasonal contrasts in diurnal carbon incorporation by phytoplankton size classes of the coastal plume of Chesapeake Bay. Marine Ecology Progress Series, 68:129–136, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/68/m068p129.pdf.

Jensen:1990:AIL

[JMPH90] R. A. Jensen, D. E. Morse, R. L. Petty, and N. Hooker. Artificial induction of larval metamorphosis by free fatty acids. Marine Ecology Progress Series, 67:55–71, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
67/m067p055.pdf.

Johnson:1991:CBA

[JMR91] C. R. Johnson, D. G. Muir, and A. L. Reysenbach. Characteristic bacteria associated with surfaces of coralline algae: a hypothesis for bacterial induction of marine invertebrate larvae. Marine Ecology Progress Series, 74:281–294, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
74/m074p281.pdf.

Jellett:1989:NOF

[JNCS89] J. F. Jellett, J. A. Novitsky, J. A. Cantley, and R. E. Scheibling. Non-occurrence of free-living Paramoeba invadens in water and sediments of Halifax Harbour, Nova Scotia, Canada. Marine Ecology Progress Series, 56:205– 209, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/56/m056p205.pdf. See comment [Mil90] and

rebuttal [JNCS90].

Jellett:1990:RRJ

[JNCS90] J. F. Jellett, J. A. Novitsky, J. A. Cantley, and R. E. Scheibling. Rebuttal to R. J. Miller’s comment on Jellett et al. (1989). Marine Ecology Progress Series, 63:307–308, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/

63/m063p307.pdf. See [JNCS89, Mil90].

Johnsen:1994:CPD

[JNJP94] G. Johnsen, N. B. Nelson, R. V. M. Jovine, and B. B. Prezelin. Chromoprotein- and pigment-dependent modeling of spectral light absorption in two dinoflagellates, Prorocentrum minimum and Heterocapsa pygmaea. Marine Ecology Progress Series, 114:245–258, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/114/m114p245.pdf.

Jochem:1995:PPC

[Joc95] F. J. Jochem. Phototrophic picoplankton community structure in three different pelagic regimes in the Arabian Sea. Marine Ecology Progress Series, 117:307–314, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
117/m117p307.pdf.

Johnson:1992:SPM

[Joh92] K. G. Johnson. Synchronous planulation of Manicina areolata (Scleractinia) with lunar periodicity. Marine Ecology Progress Series, 87:265–273, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/87/m087p265.pdf.

Johnston:1996:EEV

[Joh96] A. M. Johnston. The effect of environmental variables on ^{13}C discrimination by two marine phytoplankton. Marine Ecology Progress Series, 132:257–263, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/132/m132p257.pdf.

Johansson:1997:TDF

[Joh97] B. Johansson. Tolerance of the deposit-feeding Baltic amphipods Monoporeia affinis and Pontoporeia femorata to oxygen deficiency. Marine Ecology Progress Series, 151:135– 141, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/151/m151p135.pdf.

John:1999:OCD

[Joh99] J. St John. Ontogenetic changes in the diet of the coral reef grouper Plectropomus leopardus (Serranidae): patterns in taxa, size and habitat of prey. Marine Ecology Progress Series, 180:233–246, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/180/m180p233.pdf.

Jones:1993:PST

[Jon93] C. M. Jones. Population structure of Thenus orientalis and T. indicus (Decapoda: Scyllaridae) in northeastern Australia. Marine Ecology Progress Series, 97:143–155, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
97/m097p143.pdf.

Jones:1997:CZD

[Jon97a] R. J. Jones. Changes in zooxanthellar densities and chlorophyll concentrations in corals during and after a bleaching event. Marine Ecology Progress Series, 158:51–59, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
158/m158p051.pdf.

Jones:1997:ZLB

[Jon97b] R. J. Jones. Zooxanthellae loss as a bioassay for assessing stress in corals. Marine Ecology Progress Series, 149:163– 171, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/149/m149p163.pdf.

Jorgensen:1996:BFF

[Jør96] C. B. Jørgensen. Bivalve filter feeding revisited. Marine Ecology Progress Series, 142:287–302, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/142/m142p287.pdf.

Josefson:1990:IBB

[Jos90] A. B. Josefson. Increase of benthic biomass in the Skagerrak– Kattegat during the 1970s and 1980s — effects of organic enrichment? Marine Ecology Progress Series, 66:117–130, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
66/m066p117.pdf.

Joint:1993:PBP

[JP93a] I. Joint and A. Pomroy. Phytoplankton biomass and production in the southern North Sea. Marine Ecology Progress Series, 99:169–182, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/99/m099p169.pdf.

Joye:1993:CNF

[JP93b] S. B. Joye and H. W. Paerl. Contemporaneous nitrogen fixation and denitrification in intertidal microbial mats: rapid response to runoff events. Marine Ecology Progress Series, 94: 267–274, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/94/m094p267.pdf.

Jiang:1994:SDA

[JP94] Sunny C. Jiang and John H. Paul. Seasonal and diel abundance of viruses and occurrence of lysogeny/bacteriocinogeny in the marine environment. Marine Ecology Progress Series, 104:163–172, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/104/m104p163.pdf.

Jiang:1996:OLB

[JP96] S. C. Jiang and J. H. Paul. Occurrence of lysogenic bacteria in marine microbial communities as determined by prophage induction. Marine Ecology Progress Series, 142:27–38, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
142/m142p027.pdf.

Jeffrey:1996:DDP

[JPA^{+}96] W. H. Jeffrey, R. J. Pledger, P. Aas, S. Hager, R. B. Coffin, R. Von Haven, and D. L. Mitchell. Diel and depth profiles of DNA photodamage in bacterioplankton exposed to ambient solar ultraviolet radiation. Marine Ecology Progress Series, 137:283–291, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/137/m137p283.pdf.

James-Pirri:1997:GRP

[JPC97] M. J. James-Pirri and J. S. Cobb. Growth rates of planktonic and newly settled American lobsters Homarus americanus. Marine Ecology Progress Series, 160:233–240, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
160/m160p233.pdf.

Jackson:1992:LCC

[JPR92] C. J. Jackson, R. C. Pendrey, and P. C. Rothlisberg. The larvatron: a computer-controlled apparatus for rearing planktonic animals under experimental conditions. Marine Ecology Progress Series, 86:189–199, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/86/m086p189.pdf.

Jousson:1998:MEA

[JPZ^{+}98] O. Jousson, J. Pawlowski, L. Zaninetti, A. Meinesz, and C. F. Boudouresque. Molecular evidence for the aquarium origin of the green alga Caulerpa taxifolia introduced to the Mediterranean Sea. Marine Ecology Progress Series, 172: 275–280, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/172/m172p275.pdf.

Johnston:1992:EAR

[JR92] A. M. Johnston and J. A. Raven. Effect of aeration rates on growth rates and natural abundance, ^{13}C/^{12}C ratio of Phaeodactylum tricornutum. Marine Ecology Progress Series, 87:295–300, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/87/m087p295.pdf.

Jensen:1996:PPD

[JRB96] S. L. Jensen, B. D. Robbins, and S. S. Bell. Predicting population decline: seagrass demographics and the reconstructive technique. Marine Ecology Progress Series, 136: 267–276, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/136/m136p267.pdf.

Jensen:1997:URT

[JRB97] S. L. Jensen, B. D. Robbins, and S. S. Bell. On the use of the reconstructive technique: criticisms, comments and questions. Marine Ecology Progress Series, 146:305–309, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
146/m146p305.pdf.

Jennings:1997:SVS

[JRMN^{+}97] S. Jennings, O. Reñones, B. Morales-Nin, N. V. C. Polunin, J. Moranta, and J. Coll. Spatial variation in the ^{15}N and 13C stable isotope composition of plants, invertebrates and fishes on Mediterranean reefs: implications for the study of trophic pathways. Marine Ecology Progress Series, 146:109– 116, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/146/m146p109.pdf.

Johnson:1993:CFM

[JS93] A. S. Johnson and K. P. Sebens. Consequences of a flattened morphology: effects of flow on feeding rates of the scleractmian coral Meandrina meandrites. Marine Ecology Progress Series, 99:99–114, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/99/m099p099.pdf.

Johnson:1997:IPD

[JS97] K. B. Johnson and A. L. Shanks. The importance of prey densities and background plankton in studies of predation on invertebrate larvae. Marine Ecology Progress Series, 158:

293–296, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/158/m158p293.pdf.

Johnsen:1994:LVA

[JSGS94] G. Johnsen, O. Samset, L. Granskog, and E. Sakshaug. In vivo absorption characteristics in 10 classes of bloom-forming phytoplankton: taxonomic characteristics and responses to photoadaptation by means of discriminant and HPLC analysis. Marine Ecology Progress Series, 105:149–157, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
105/m105p149.pdf.

Ju:1999:UEL

[JSH99] S. J. Ju, D. H. Secor, and H. R. Harvey. Use of extractable lipofuscin for age determination of blue crab Callinectes sapidus. Marine Ecology Progress Series, 185:171–179, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
185/m185p171.pdf.

Jensen:1990:PCR

[JSJMH90] L. M. Jensen, K. Sand-Jensen, S. Marcher, and M. Hansen. Plankton community respiration along a nutrient gradient in a shallow Danish estuary. Marine Ecology Progress Series, 61:75–85, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/61/m061p075.pdf.

Jeong:1999:FMT

[JSK^{+}99] H. J. Jeong, J. H. Shim, J. S. Kim, J. Y. Park, C. W. Lee, and Y. Lee. Feeding by the mixotrophic thecate dinoflagellate Fragilidium cf. mexicanum on red-tide and toxic dinoflagellates. Marine Ecology Progress Series, 176:263–277, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
176/m176p263.pdf.

Johnson:1991:SCT

[JSOG91] C. R. Johnson, D. C. Sutton, R. R. Olson, and R. Giddins. Settlement of crown-of-thorns starfish: role of bacteria on surfaces of coralline algae and a hypothesis for deepwater recruitment. Marine Ecology Progress Series, 71:143–162, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
71/m071p143.pdf.

Johnsen:1992:PCS

[JSV92] G. Johnsen, E. Sakshaug, and M. Vernet. Pigment composition, spectral characterization and photosynthetic parameters in Chryso-chromulma polylepis. Marine Ecology Progress Series, 83:241–249, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/83/m083p241.pdf.

Jonsson:1990:FBP

[JT90] P. P. Jonsson and P. Tiselius. Feeding behaviour, prey detection and capture efficiency of the copepod Acartia tonsa feeding on planktonic ciliates. Marine Ecology Progress Series, 60:35–44, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/60/m060p035.pdf.

Juhl:1993:BAM

[JT93] A. R. Juhl and G. L. Taghon. Biology of an active methane seep on the Oregon continental shelf. Marine Ecology Progress Series, 102:287–294, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/102/m102p287.pdf.

Jahn:1997:DDT

[JT97] A. Jahn and H. Theede. Different degrees of tolerance to hydrogen sulphide in populations of Macoma balthica (Bivalvia, Tellinidae). Marine Ecology Progress Series, 154:185– 196, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/154/m154p185.pdf.

Jorgensen:1999:OBC

[JTB99] N. O. G. Jørgensen, L. J. Tranvik, and G. M. Berg. Occurrence and bacterial cycling of dissolved nitrogen in the Gulf of Riga, the Baltic Sea. Marine Ecology Progress Series, 191:1–18, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/191/m191p001.pdf.

Jiang:1992:CMP

[JTPP92] S. C. Jiang, J. M. Thurmond, S. L. Pichard, and J. H. Paul. Concentration of microbial populations from aquatic environments by vortex flow filtration. Marine Ecology Progress Series, 80:101–107, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/80/m080p101.pdf.

Juanes:1992:WDD

[Jua92] F. Juanes. Why do decapod crustaceans prefer small-sized molluscan prey? Marine Ecology Progress Series, 87:239– 249, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/87/m087p239.pdf.

Jaquet:1996:SDC

[JW96] N. Jaquet and H. Whitehead. Scale-dependent correlation of sperm whale distribution with environmental features and productivity in the South Pacific. Marine Ecology Progress Series, 135:1–9, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/135/m135p001.pdf.

Jenkins:1998:DTV

[JWKH98] G. P. Jenkins, D. C. Welsford, M. J. Keough, and P. A. Hamer. Diurnal and tidal vertical migration of pre-settlement King George whiting Sillaginodes punctata in relation to feeding and vertical distribution of prey in a temperate bay. Marine Ecology Progress Series, 170:239–248, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
170/m170p239.pdf.

Jaquet:1996:CBC

[JWL96] N. Jaquet, H. Whitehead, and M. Lewis. Coherence between 19th century sperm whale distributions and satellite-derived pigments in the tropical Pacific. Marine Ecology Progress Series, 145:1–10, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/145/m145p001.pdf.

Kraemer:1993:ARP

[KA93] G. P. Kraemer and R. S. Alberte. Age-related patterns of metabolism and biomass in subterranean tissues of Zostera marina (eelgrass). Marine Ecology Progress Series, 95:193– 203, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/95/m095p193.pdf.

Kamiyama:1997:LED

[KA97] T. Kamiyama and S. Arima. Lethal effect of the dinoflagellate Heterocapsa circularisquama upon the tintinnid ciliate Favella taraikaensis. Marine Ecology Progress Series, 160: 27–33, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/160/m160p027.pdf.

Kahler:1998:CRC

[Käh98] P. Kähler. Concerning respiration and the carbon-to-nitrogen ratio of a phytoplankton bloom. Comment on Kepkay et al. (1997). Marine Ecology Progress Series, 163:299–303, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/

163/m163p299.pdf. See [KJN97] and reply [Kep98].

Kaiser:1996:SDI

[Kai96] M. J. Kaiser. Starfish damage as an indicator of trawling intensity. Marine Ecology Progress Series, 134:303–307, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
134/m134p303.pdf.

Kalish:1991:IDF

[Kal91a] J. M. Kalish. ^{13}C and ^{18}O isotopic disequilibria in fish otoliths: metabolic and kinetic effects. Marine Ecology Progress Series, 75:191–203, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/75/m075p191.pdf.

Kalish:1991:DOC

[Kal91b] J. M. Kalish. Determinants of otolith chemistry seasonal variation in the composition of blood plasma, endolymph and otoliths of bearded rock cod Pseudophycis barbatus.

Marine Ecology Progress Series, 74:137–159, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
74/m074p137.pdf.

Kamermans:1994:SFS

[Kam94a] P. Kamermans. Similarity in food source and timing of feeding in deposit- and suspension-feeding bivalves. Marine Ecology Progress Series, 104:63–75, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/104/m104p063.pdf.

Kamiyama:1994:EEP

[Kam94b] T. Kamiyama. Effects of extracellular products from phytoplankton on the excystment of tintinnids from marine sediments. Marine Ecology Progress Series, 105:199–201, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
105/m105p199.pdf.

Karbe:1992:TSM

[Kar92] L. Karbe. Toxicity of surface microlayer, subsurface water and sediment-elutriates from the German Bight: summary and conclusions. Marine Ecology Progress Series, 91:197– 201, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/91/m091p197.pdf.

Karakassis:1995:NMC

[Kar95] I. Karakassis. S_{∞}: a new method for calculating macrobenthic species richness. Marine Ecology Progress Series, 120: 299–303, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/120/m120p299.pdf.

Kvitek:1991:RIB

[KB91] R. G. Kvitek and M. K. Beitler. Relative insensitivity of butter clam neurons to saxitoxin; a pre-adaptation for sequestering paralytic shellfish poisoning toxins as a chemical defense. Marine Ecology Progress Series, 69:47–54, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
69/m069p047.pdf.

Karlson:1994:PPC

[KB94] K. Karlson and U. Bamstedt. Planktivorous predation on copepods. Evaluation of mandible remains in predator guts as a quantitative estimate of predation. Marine Ecology Progress Series, 108:79–89, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/108/m108p079.pdf.

Koziol:1997:EPB

[KBA^{+}97] C. Koziol, R. Batel, E. Arinc, H. C. Schröder, and W. E. G. Müller. Expression of the potential biomarker heat shock protein 70 and its regulator, the metazoan DnaJ homolog, by temperature stress in the sponge Geodia cydonium. Marine Ecology Progress Series, 154:261–268, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
154/m154p261.pdf.

Kaiser:1996:CGF

[KBN^{+}96] M. J. Kaiser, B. Bullimore, P. Newman, K. Lock, and S. Gilbert. Catches in ‘ghost fishing’ set nets. Marine Ecology Progress Series, 145:11–16, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/145/m145p011.pdf.

Kromkamp:1998:DMP

[KBP98] J. Kromkamp, C. Barranguet, and J. Peene. Determination of microphytobenthos PSII quantum efficiency and photosynthetic activity by means of variable chlorophyll fluorescence. Marine Ecology Progress Series, 162:45–55, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
162/m162p045.pdf.

Koski:1998:EFQ

[KBS98] M. Koski, W. Klein Breteler, and N. Schogt. Effect of food quality on rate of growth and development of the pelagic copepod Pseudocalanus elongatus (Copepoda, Calanoida). Marine Ecology Progress Series, 170:169–187, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
170/m170p169.pdf.

Koop:1990:SWO

[KBWC90] K. Koop, W. R. Boynton, F. Wulff, and R. Carman. Sediment-water oxygen and nutrient exchanges along a depth gradient in the Baltic Sea. Marine Ecology Progress Series, 63:65–77, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/63/m063p065.pdf.

Kuhl:1995:MPZ

[KCD^{+}95] M. Kühl, Y. Cohen, T. Dalsgaard, B. B. Jørgensen, and N. P. Revsbech. Microenvironment and photosynthesis of zooxanthellae in scleractinian corals studied with microsensors for O_{2}, pH and light. Marine Ecology Progress Series, 117:159–172, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/117/m117p159.pdf.

Kvitek:1998:BPD

[KCI98] R. G. Kvitek, K. E. Conlan, and P. J. Iampietro. Black pools of death: hypoxic, brine-filled ice gouge depressions become lethal traps for benthic organisms in a shallow Arctic embayment. Marine Ecology Progress Series, 162:1–10, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
162/m162p001.pdf.

Katz:1994:LBH

[KCS94] C. H. Katz, J. S. Cobb, and M. Spaulding. Larval behavior, hydrodynamic transport, and potential offshore-to-inshore recruitment in the American lobster Homarus americanus. Marine Ecology Progress Series, 103:265–273, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
103/m103p265.pdf.

Krishnakumar:1994:EEC

[KCV94] P. K. Krishnakumar, E. Casillas, and U. Varanasi. Effect of environmental contaminants on the health of Mytilus edulis from Puget Sound, Washington, USA. I. Cytochemical measures of lysosomal responses in the digestive cells using automatic image analysis. Marine Ecology Progress Series, 106: 249–261, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/106/m106p249.pdf.

Kim:1996:ESS

[KD96] J. H. Kim and R. E. DeWreede. Effects of size and season of disturbance on algal patch recovery in a rocky intertidal community. Marine Ecology Progress Series, 133:217–228, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
133/m133p217.pdf.

Kelly:1997:MMP

[KD97] J. R. Kelly and P. H. Doering. Monitoring and modeling primary production in coastal waters: studies in Massachusetts Bay 1992–1994. Marine Ecology Progress Series, 148:155– 168, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/148/m148p155.pdf.

Kelly:1999:SDP

[KD99] J. R. Kelly and P. H. Doering. Seasonal deepening of the pycnocline in a shallow shelf ecosystem and its influence on near-bottom dissolved oxygen. Marine Ecology Progress Series, 178:151–168, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/178/m178p151.pdf.

Kristensen:1992:PSB

[KDAS92] E. Kristensen, A. H. Devol, S. I. Ahmed, and M. Saleem. Preliminary study of benthic metabolism and sulfate reduction in a mangrove swamp of the Indus Delta, Pakistan. Marine Ecology Progress Series, 90:287–297, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/90/m090p287.pdf.

Kimura:1999:NSR

[KDC99] S. Kimura, K. Döös, and A. C. Coward. Numerical simulation to resolve the issue of downstream migration of the Japanese eel. Marine Ecology Progress Series, 186:303–306, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
186/m186p303.pdf.

Kroncke:1998:LTC

[KDHZ98] I. Kröncke, J. W. Dippner, H. Heyen, and B. Zeiss. Long-term changes in macrofaunal communities off Norderney (East Frisia, Germany) in relation to climate variability. Marine Ecology Progress Series, 167:25–36, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/167/m167p025.pdf.

Kogure:1996:HCN

[KDKS96] K. Kogure, H. K. Do, D. S. Kim, and Y. Shirayama. High concentrations of neurotoxin in free-living marine nematodes. Marine Ecology Progress Series, 136:147–151, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
136/m136p147.pdf.

Kohler:1992:HCI

[KDL92] A. Kohler, H. Deisemann, and B. Lauritzen. Histological and cytochemical indices of toxic injury in the liver of dab Limanda limanda. Marine Ecology Progress Series, 91:141– 153, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/91/m091p141.pdf.

Koster:1997:MSA

[KDMR97] M. Köster, S. Dahlke, and L. A. Meyer-Reil. Microbiological studies along a gradient of eutrophication in a shallow coastal inlet in the southern Baltic Sea (Nordrügensche Bodden). Marine Ecology Progress Series, 152:27–39, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
152/m152p027.pdf.

Kroncke:1992:EFD

[KDR^{+}92] I. Kroncke, G. C. A. Duineveld, S. Raak, E. Rachor, and R. Daan. Effects of a former discharge of drill cuttings on the macrofauna community. Marine Ecology Progress Series, 91:277–287, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/91/m091p277.pdf.

Klein:1997:EEE

[KDR97] C. Klein, J. R. Dolan, and F. Rassoulzadegan. Experimental examination of the effects of rainwater on microbial communities in the surface layer of the NW Mediterranean Sea. Marine Ecology Progress Series, 158:41–50, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/158/m158p041.pdf.

Kerner:1995:DOM

[KE95] M. Kerner and F. Edelkraut. Decomposition of organic matter in aggregated seston from the Elbe estuary: redox dependency and production of low molecular weight DOC compounds. Marine Ecology Progress Series, 123:281–293, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
123/m123p281.pdf.

Karakassis:1997:CSC

[KE97] I. Karakassis and A. Eleftheriou. The continental shelf of Crete: structure of macrobenthic communities. Marine Ecology Progress Series, 160:185–196, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/160/m160p185.pdf.

Keats:1990:FWF

[Kea90] D. W. Keats. Food of winter flounder Pseudo-pleuronectes americanus in a sea urchin dominated community in eastern Newfoundland. Marine Ecology Progress Series, 60:13–22, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/60/m060p013.pdf.

Kasamatsu:1998:CAM

[KEJ98] F. Kasamatsu, P. Ensor, and G. G. Joyce. Clustering and aggregations of minke whales in the Antarctic feeding grounds. Marine Ecology Progress Series, 168:1–11, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
168/m168p001.pdf.

Kelly:1998:QPR

[Kel98] J. R. Kelly. Quantification and potential role of ocean nutrient loading to Boston Harbor, Massachusetts, USA. Marine Ecology Progress Series, 173:53–65, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/173/m173p053.pdf.

Kennelly:1992:DAC

[Ken92] S. J. Kennelly. Distributions, abundances and current status of exploited populations of spanner crabs Ranina ranina off the east coast of Australia. Marine Ecology Progress Series, 85:227–235, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/85/m085p227.pdf.

Kendrick:1994:EPS

[Ken94a] G. A. Kendrick. Effects of propagule settlement density and adult canopy on survival of recruits of Sargassum spp. (Sargassaceae: Phaeophyta). Marine Ecology Progress Series, 103:129–140, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/103/m103p129.pdf.

Kennedy:1994:CPW

[Ken94b] A. D. Kennedy. Carbon partitioning within meiobenthic nematode communities in the Exe Estuary, UK. Marine Ecology Progress Series, 105:71–78, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/105/m105p071.pdf.

Kennedy:1994:PWM

[Ken94c] A. D. Kennedy. Predation within meiofaunal communities: description and results of a rapid-freezing method of investigation. Marine Ecology Progress Series, 114:71–79, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
114/m114p071.pdf.

Kepkay:1991:SCM

[Kep91] P. E. Kepkay. Surface coagulation and microbial respiration in response to local advection and sea state in the North Atlantic. Marine Ecology Progress Series, 69:143–147, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
69/m069p143.pdf.

Kepkay:1994:PAB

[Kep94] P. E. Kepkay. Particle aggregation and the biological reactivity of colloids. Marine Ecology Progress Series, 109: 293–304, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/109/m109p293.pdf.

Kepkay:1998:RCP

[Kep98] P. E. Kepkay. Reply to the comment of P. Kähler. Marine Ecology Progress Series, 163:303–305, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/

163/m163p303.pdf. See [KJN97, Käh98].

Kern:1990:APA

[Ker90] J. C. Kern. Active and passive aspects of meiobenthic copepod dispersal at two sites near Mustang Island, Texas. Marine Ecology Progress Series, 60:211–223, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
60/m060p211.pdf.

Kerrigan:1994:PSG

[Ker94] B. A. Kerrigan. Post-settlement growth and body composition in relation to food availability in a juvenile tropical reef fish. Marine Ecology Progress Series, 111:7–15, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
111/m111p007.pdf.

Kerrigan:1996:TPS

[Ker96] B. A. Kerrigan. Temporal patterns in size and condition at settlement in two tropical reef fishes (Pomacentridae: Pomacentrus amboinensis and P. nagasakiensis). Marine Ecology Progress Series, 135:27–41, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/135/m135p027.pdf.

Koski:1999:RSC

[KEV99] M. Koski, J. Engström, and M. Viitasalo. Reproduction and survival of the calanoid copepod Eurytemora affinis fed with toxic and non-toxic cyanobacteria. Marine Ecology Progress Series, 186:187–197, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/186/m186p187.pdf.

Krom:1995:NPC

[KEvN95] M. D. Krom, S. Ellner, J. van Rijn, and A. Neori. Nitrogen and phosphorus cycling and transformations in a prototype ’non-polluting’ integrated mariculture system, Eilat, Israel. Marine Ecology Progress Series, 118:25–36, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
118/m118p025.pdf.

Kure:1997:IBA

[KF97] L. K. Kure and T. L. Forbes. Impact of bioturbation by Arenicola marina on the fate of particle-bound fluoranthene. Marine Ecology Progress Series, 156:157–166, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
156/m156p157.pdf.

Kaldy:1999:CAT

[KFD99] J. E. Kaldy, N. Fowler, and K. H. Dunton. Critical assessment of Thalassia testudinum (turtle grass) aging techniques: implications for demographic inferences. Marine Ecology Progress Series, 181:279–288, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/181/m181p279.pdf.

Kana:1992:NSM

[KFF92] T. M. Kana, N. L. Feiwel, and L. C. Flynn. Nitrogen starvation in marine Synechococcus strains: clonal differences in phycobiliprotein breakdown and energy coupling. Marine Ecology Progress Series, 88:75–82, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/88/m088p075.pdf.

Kimura:1999:BTP

[KFN99] H. Kimura, T. Fukuba, and T. Naganuma. Biomass of thraustochytrid protoctists in coastal water. Marine Ecology Progress Series, 189:27–33, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/189/m189p027.pdf.

Karner:1994:PNC

[KFPR94] M. Karner, C. Ferrier-Pages, and F. Rassoulzadegan. Phagotrophic nanoflagellates contribute to occurrence of aglucosidase and aminopeptidase in marine environments. Marine Ecology Progress Series, 114:237–244, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
114/m114p237.pdf.

Krupatkina:1991:PPS

[KFS91] D. K. Krupatkina, Z. Z. Finenko, and A. A. Shalapyonok. Primary production and size-fractionated structure of the Black Sea phytoplankton in the winter-spring period. Marine Ecology Progress Series, 73:25–31, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/73/m073p025.pdf.

Klumpp:1994:CPH

[KG94] D. W. Klumpp and C. L. Griffiths. Contributions of phototrophic and heterotrophic nutrition to the metabolic and growth requirements of four species of giant clam (Tridacnidae). Marine Ecology Progress Series, 115:103–115, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
115/m115p103.pdf.

Kiene:1995:EGB

[KG95] R. P. Kiene and G. Gerard. Evaluation of glycine betaine as an inhibitor of dissolved dimethylsulfoniopropionate degradation in coastal waters. Marine Ecology Progress Series, 128:121–131, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/128/m128p121.pdf.

Koch:1999:WFT

[KG99] E. W. Koch and G. Gust. Water flow in tide- and wavedominated beds of the seagrass Thalassia testudinum. Marine Ecology Progress Series, 184:63–72, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/184/m184p063.pdf.

Krueger:1992:SFP

[KGC92] D. M. Krueger, S. M. Gallager, and C. M. Cavanaugh. Suspension feeding on phytoplankton by Solemya velum, a symbiont-containing clam. Marine Ecology Progress Series, 86:145–151, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/86/m086p145.pdf.

Kelly-Gerreyn:1999:CED

[KGHTN99] B. A. Kelly-Gerreyn, D. J. Hydes, M. Trimmer, and D. B. Nedwell. Calibration of an early diagenesis model for high nitrate, low reactive sediments in a temperate latitude estuary (Great Ouse, UK). Marine Ecology Progress Series, 177: 37–50, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/177/m177p037.pdf.

Kimmerer:1994:PIC

[KGO94] W. J. Kimmerer, E. Gartside, and J. J. Orsi. Predation by an introduced clam as the likely cause of substantial declines in zooplankton of San Francisco Bay. Marine Ecology Progress Series, 113:81–93, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/113/m113p081.pdf.

KleinBreteler:1995:DPE

[KGS95] W. C. M. Klein Breteler, S. R. Gonzalez, and N. Schogt. Development of Pseudocalanus elongatus (Copepoda, Calanoida) cultured at different temperature and food conditions. Marine Ecology Progress Series, 119:99–110, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
119/m119p099.pdf.

Kurbjeweit:1993:LCS

[KGW93] F. Kurbjeweit, R. Gradinger, and J. Weissenberger. The life cycle of Stephos longipes — an example for cryopelagic coupling in the Weddell Sea (Antarctica). Marine Ecology Progress Series, 98:255–262, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/98/m098p255.pdf.

Kaltenbock:1992:EAA

[KH92a] E. Kaltenbock and G. J. Herndl. Ecology of amorphous aggregations (marine snow) in the Northern Adriatic Sea. IV. Dissolved nutrients and the autotrophic community associated with marine snow. Marine Ecology Progress Series, 87: 147–159, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/87/m087p147.pdf.

Keesing:1992:FMS

[KH92b] J. K. Keesing and A. R. Halford. Field measurement of survival rates of juvenile Acanthaster piano, techniques and preliminary results. Marine Ecology Progress Series, 85:107– 114, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/85/m085p107.pdf.

Kattner:1998:LMA

[KH98] G. Kattner and W. Hagen. Lipid metabolism of the Antarctic euphausiid Euphausia crystallorophias and its ecological implications. Marine Ecology Progress Series, 170:203– 213, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/170/m170p203.pdf.

Kristensen:1991:BMS

[KHB91] E. Kristensen, M. Holmer, and N. Bussarawit. Benthic metabolism and sulfate reduction in a southeast Asian mangrove swamp. Marine Ecology Progress Series, 73:93–103, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/73/m073p093.pdf.

Kreeger:1995:SVR

[KHBL95] D. A. Kreeger, A. J. S. Hawkins, B. L. Bayne, and D. M. Lowe. Seasonal variation in the relative utilization of dietary protein for energy and biosynthesis by the mussel Mytilus edulis. Marine Ecology Progress Series, 126:177–184, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
126/m126p177.pdf.

Kattner:1996:AKT

[KHFP^{+}96] G. Kattner, W. Hagen, S. Falk-Petersen, J. R. Sargent, and R. J. Henderson. Antarctic krill Thysanoessa macrura fills a major gap in marine lipogenic pathways. Marine Ecology Progress Series, 134:295–298, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/134/m134p295.pdf.

Kononen:1999:STV

[KHK^{+}99] K. Kononen, M. Huttunen, I. Kanoshina, J. Laanemets, P. Moisander, and J. Pavelson. Spatial and temporal variability of a dinoflagellate-cyanobacterium community under a complex hydrodynamical influence: a case study at the entrance to the Gulf of Finland. Marine Ecology Progress Series, 186:43–57, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/186/m186p043.pdf.

Kuwata:1993:ECT

[KHT93] A. Kuwata, T. Hama, and M. Takahashi. Ecophysiological characterization of two life forms, resting spores and resting cells, of a marine planktonic diatom, Chaetoceros pseudocun/isetus, formed under nutrient depletion. Marine Ecology Progress Series, 102:245–255, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/102/m102p245.pdf.

Karakassis:1999:BRF

[KHTP99] I. Karakassis, E. Hatziyanni, M. Tsapakis, and W. Plaiti. Benthic recovery following cessation of fish farming: a series of successes and catastrophes. Marine Ecology Progress Series, 184:205–218, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/184/m184p205.pdf.

Koshikawa:1999:IPC

[KHW^{+}99] H. Koshikawa, S. Harada, M. Watanabe, K. Kogure, T. Ioriya, K. Kohata, T. Kimura, K. Sato, and T. Akehata. Influence of plankton community structure on the contribution of bacterial production to metazooplankton in a coastal mesocosm. Marine Ecology Progress Series, 186:31–42, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
186/m186p031.pdf.

Kasamatsu:1997:NVR

[KI97] F. Kasamatsu and Y. Ishikawa. Natural variation of radionuclide ^{137}Cs concentration in marine organisms with special reference to the effect of food habits and trophic level. Marine Ecology Progress Series, 160:109–120, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
160/m160p109.pdf.

Koubbi:1991:EIS

[KID91] P. Koubbi, F. Ibanez, and G. Duhamel. Environmental influences on spatio-temporal oceanic distribution of ichthyoplankton around the Kerguelen Islands (Southern Ocean). Marine Ecology Progress Series, 72:225–238, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
72/m072p225.pdf.

Kingsford:1992:DAS

[Kin92] M. J. Kingsford. Drift algae and small fish in coastal waters of northeastern New Zealand. Marine Ecology Progress Series, 80:41–55, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/80/m080p041.pdf.

Kinne:1993:OHM

[Kin93] O. Kinne. One hundred MEPS volumes: their pages have shaped the world of marine ecological research MEPS 100:1. Marine Ecology Progress Series, ??:??, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
100/m100p001.pdf.

Kingsford:1995:DAC

[Kin95a] M. J. Kingsford. Drift algae: a contribution to near-shore habitat complexity in the pelagic environment and an attractant for fish. Marine Ecology Progress Series, 116:297– 301, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/116/m116p297.pdf.

Kinne:1995:MCG

[Kin95b] O. Kinne. MEPS continues to grow–where is the limit? Marine Ecology Progress Series, ??:??, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/125/m125p001.pdf.

Kinne:1997:EEE

[Kin97] O. Kinne. Ethics and eco-ethics. Marine Ecology Progress Series, 153:1–3, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/153/m153p001.pdf.

Kirchman:1990:LBG

[Kir90] D. L. Kirchman. Limitation of bacterial growth by dissolved organic matter in the subarctic Pacific. Marine Ecology Progress Series, 62:47–54, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/62/m062p047.pdf.

Kirchman:1992:ITL

[Kir92] D. L. Kirchman. Incorporation of thymidine and leucine in the subarctic Pacific: application to estimating bacterial production. Marine Ecology Progress Series, 82:301–309, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
82/m082p301.pdf.

Kalish:1996:UBR

[KJGC96] J. M. Kalish, J. M. Johnston, J. S. Gunn, and N. P. Clear. Use of the bomb radiocarbon chronometer to determine age of southern bluefin tuna Thunnus maccoyii. Marine Ecology Progress Series, 143:1–8, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/143/m143p001.pdf.

Kim:1996:EPP

[KJLH96] W. S. Kim, J. K. Jeon, S. H. Lee, and H. T. Huh. Effects of pentachlorophenol (PCP) on the oxygen consumption rate of the river puffer fish Takifugu obscurus. Marine Ecology Progress Series, 143:9–14, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/143/m143p009.pdf.

Krause-Jensen:1996:PWD

[KJMRC96] D. Krause-Jensen, K. McGlathery, S. Rysgaard, and P. B. Christensen. Production within dense mats of the filamentous macroalga Chaetomorpha linum in relation to light and nutrient availability. Marine Ecology Progress Series, 134: 207–216, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/134/m134p207.pdf.

Kepkay:1997:RCN

[KJN97] P. E. Kepkay, J. F. Jellett, and S. E. H. Niven. Respiration and the carbon-to-nitrogen ratio of a phytoplankton bloom. Marine Ecology Progress Series, 150:249–261, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/

150/m150p249.pdf. See comment [Käh98] and reply [Kep98].

Kautsky:1990:GPD

[KJT90] N. Kautsky, K. Johannesson, and M. Tedengren. Genotypic and phenotypic differences between Baltic and North Sea populations of Mytilus edulis evaluated through reciprocal transplantations. I. Growth and morphology. Marine Ecology Progress Series, 59:203–210, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/59/m059p203.pdf.

Keil:1991:CDF

[KK91] R. G. Keil and D. L. Kirchman. Contribution of dissolved free amino acids and ammonium to the nitrogen requirements of heterotrophic bacterioplankton. Marine Ecology Progress Series, 73:1–10, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/73/m073p001.pdf.

Kwint:1995:DPP

[KK95] R. L. J. Kwint and K. J. M. Kramer. Dimethylsulphide production by plankton communities. Marine Ecology Progress Series, 121:227–237, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/121/m121p227.pdf.

Kwint:1996:ACP

[KK96] R. L. J. Kwint and K. J. M. Kramer. Annual cycle of the production and fate of DMS and DMSP in a marine coastal system. Marine Ecology Progress Series, 134:217–224, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
134/m134p217.pdf.

Kristensen:1994:SRA

[KKH^{+}94] E. Kristensen, G. M. King, M. Holmer, G. T. Banta, M. H. Jensen, K. Hansen, and N. Bussarawit. Sulfate reduction, acetate turnover and carbon metabolism in sediments of the Ao Nam Bor mangrove, Phuket, Thailand. Marine Ecology Progress Series, 109:245–255, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/109/m109p245.pdf.

Kaartvedt:1998:SVM

[KKH98] S. Kaartvedt, T. Knutsen, and J. C. Holst. Schooling of the vertically migrating mesopelagic fish Maurolicus muelleri in light summer nights. Marine Ecology Progress Series, 170:287–290, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/170/m170p287.pdf.

Kiorboe:1990:SPF

[KKK^{+}90] T. Kiørboe, H. Kaas, B. Kruse, F. Møhlenberg, P. Tiselius, and G. /Ertebjerg. The structure of the pelagic food web in relation to water column structure in the Skagerrak. Marine Ecology Progress Series, 59:19–32, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/59/m059p019.pdf.

Kahru:1991:ACP

[KKL91] M. Kahru, E. Kaasik, and A. Leeben. Annual cycle of particle size fractions and phytoplankton biomass in the northern Baltic proper. Marine Ecology Progress Series, 69:117– 124, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/69/m069p117.pdf.

Kelman:1998:AAR

[KKL^{+}98] D. Kelman, A. Kushmaro, Y. Loya, Y. Kashman, and Y. Benayahu. Antimicrobial activity of a Red Sea soft coral, Parerythropodium fulvum fulvum: reproductive and developmental considerations. Marine Ecology Progress Series, 169: 87–95, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/169/m169p087.pdf.

Kivi:1996:ESR

[KKT96a] K. Kivi, H. Kuosa, and S. Tanskanen. An experimental study on the role of crustacean and microprotozoan grazers in the planktonic food web. Marine Ecology Progress Series, 136:59–68, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/136/m136p059.pdf.

Kuwata:1996:MEL

[KKT96b] A. Kuwata, H. Kanazawa, and M. Takahashi. Method for estimating the in situ growth rate for a large pelagic diatom of rhizosoleniacea. Marine Ecology Progress Series, 131:219–224, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/131/m131p219.pdf.

Koepfler:1993:STB

[KKW^{+}93] E. T. Koepfler, H. I. Kator, R. L. Wetzel, L. W. Haas, and K. L. Webb. Spatial and temporal bacterioplankton dynamics during destratification of the James River estuary, Virginia, USA. Marine Ecology Progress Series, 102:229–244, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
102/m102p229.pdf.

Koslow:1997:PBC

[KKW97] J. A. Koslow, R. J. Kloser, and A. Williams. Pelagic biomass and community structure over the mid-continental slope off southeastern Australia based upon acoustic and midwater trawl sampling. Marine Ecology Progress Series, 146:21–35, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/146/m146p021.pdf.

Kim:1997:ECR

[KKYH97] W. S. Kim, J. M. Kim, S. K. Yi, and H. T. Huh. Endogenous circadian rhythm in the river puffer fish Takifugu obscurus. Marine Ecology Progress Series, 153:293–298, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
153/m153p293.pdf.

Kuuppo-Leinikki:1990:PGP

[KL90] P. Kuuppo-Leinikki. Protozoan grazing on planktonic bacteria and its impact on bacterial population. Marine Ecology Progress Series, 63:227–238, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/63/m063p227.pdf.

Kahru:1991:SSP

[KL91a] M. Kahru and A. Leeben. Size structure of planktonic particles in relation to hydrographic structure in the Skagerrak. Marine Ecology Progress Series, 76:159–166, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
76/m076p159.pdf.

Kenner:1991:SFR

[KL91b] M. C. Kenner and M. T. Lares. Size at first reproduction of the sea urchin Strongylocentrotus purpuratus in a central California kelp forest. Marine Ecology Progress Series, 76: 303–306, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/76/m076p303.pdf.

Klein:1991:SGD

[KL91c] R. Klein and Y. Loya. Skeletal growth and density patterns of two Pontes corals from the Gulf of Eilat, Red Sea. Marine Ecology Progress Series, 77:253–259, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
77/m077p253.pdf.

Kleppel:1992:CPM

[KL92] G. S. Kleppel and E. J. Lessard. Carotenoid pigments in microzooplankton. Marine Ecology Progress Series, 84:211– 218, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/84/m084p211.pdf.

Klumpp:1994:NEG

[KL94] D. W. Klumpp and J. S. Lucas. Nutritional ecology of the giant clams Tridacna tevoroa and T. derasa from Tonga: influence of light on filter-feeding and photosynthesis. Marine Ecology Progress Series, 107:147–156, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
107/m107p147.pdf.

Kang:1995:APA

[KL95] S. H. Kang and S. H. Lee. Antarctic phytoplankton assemblage in the western Bransfield Strait region, February 1993: composition, biomass, and mesoscale distributions. Marine Ecology Progress Series, 129:253–267, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
129/m129p253.pdf.

Kuuppo-Leinikki:1994:TIC

[KLAH^{+}94] P. Kuuppo-Leinikki, R. Autio, S. Hällfors, H. Kuosa, J. Kuparinen, and R. Pajunleml. Trophic interactions and carbon flow between picoplankton and protozoa in pelagic enclosures manipulated with nutrients and a top predator. Marine Ecology Progress Series, 107:89–102, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/107/m107p089.pdf.

Kleppel:1993:DCC

[Kle93] G. S. Kleppel. On the diets of calanoid copepods. Marine Ecology Progress Series, 99:183–195, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
99/m099p183.pdf.

Kieskamp:1991:SVD

[KLEH91] W. M. Kieskamp, L. Lohse, E. Epping, and W. Helder. Seasonal variation in denitrification rates and nitrous oxide fluxes in intertidal sediments of the western Wadden Sea. Marine Ecology Progress Series, 72:145–151, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
72/m072p145.pdf.

Kuhl:1994:LPL

[KLJ94] M. Kühl, C. Lassen, and B. B. Jorgensen. Light penetration and light intensity in sandy marine sediments measured with irradiance and scalar irradiance fiber-optic microprobes. Marine Ecology Progress Series, 105:139–148, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
105/m105p139.pdf.

Kaas:1991:CPB

[KLMR91] H. Kaas, J. Larsen, F. Mohlenberg, and K. Richardson. The Chrysochromulina polylepis bloom in the Kattegat (Scandinavia) May–June 1988. Distribution, primary production and nutrient dynamics in the late stage of the bloom. Marine Ecology Progress Series, 79:151–161, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/79/m079p151.pdf.

Kahru:1990:STM

[KLN^{+}90] M. Kahru, J.-M. Leppanen, S. Nommann, U. Passow, L. Postel, and S. Schulz. Spatio-temporal mosaic of the phytoplankton spring bloom in the open Baltic Sea in 1986. Marine Ecology Progress Series, 66:301–309, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/66/m066p301.pdf.

Kahru:1993:CBC

[KLR93] M. Kahru, J.-M. Leppanen, and O. Rud. Cyanobacterial blooms cause heating of the sea surface. Marine Ecology Progress Series, 101:1–7, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/101/m101p001.pdf.

Klumpp:1992:CSB

[KM92] D. W. Klumpp and A. D. McKinnon. Community structure, biomass and productivity of epilithic algal communities on the Great Barrier Reef: dynamics at different spatial scales. Marine Ecology Progress Series, 86:77–89, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
86/m086p077.pdf.

Karrh:1996:EFS

[KM96] R. R. Karrh and D. C. Miller. Effect of flow and sediment transport on feeding rate of the surface-deposit feeder Saccoglossus kowalevskii. Marine Ecology Progress Series, 130: 125–134, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/130/m130p125.pdf.

Keller:1999:MCI

[KM99a] K. Keller and F. M. M. Morel. A model of carbon isotopic fractionation and active carbon uptake in phytoplankton. Marine Ecology Progress Series, 182:295–298, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
182/m182p295.pdf.

Kraemer:1999:NAS

[KM99b] G. P. Kraemer and L. Mazzella. Nitrogen acquisition, storage, and use by the co-occurring Mediterranean seagrasses Cymodocea nodosa and Zostera noltii. Marine Ecology Progress Series, 183:95–103, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/183/m183p095.pdf.

Komatsu:1997:TLR

[KMB97] T. Komatsu, A. Meinesz, and D. Buckles. Temperature and light responses of alga Caulerpa taxifolia introduced into the Mediterranean Sea. Marine Ecology Progress Series, 146: 145–153, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/146/m146p145.pdf.

Kiflawi:1998:DMS

[KMG98] M. Kiflawi, A. I. Mazeroll, and D. Goulet. Does mass spawning enhance fertilization in coral reef fish? A case study of the brown surgeonfish. Marine Ecology Progress Series, 172:107–114, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/172/m172p107.pdf.

Kaartvedt:1996:VDF

[KMKS96] S. Kaartvedt, W. Melle, T. Knutsen, and H. R. Skjoldal. Vertical distribution of fish and krill beneath water of varying optical properties. Marine Ecology Progress Series, 136: 51–58, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/136/m136p051.pdf.

Klein:1991:BAC

[KML91] R. Klein, O. Mokady, and Y. Loya. Bioerosion in ancient and contemporary corals of the genus Pontes: patterns and palaeoenvironmental implications. Marine Ecology Progress Series, 77:245–251, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/77/m077p245.pdf.

Kamykowski:1998:BRO

[KMR98] D. Kamykowski, E. J. Milligan, and R. E. Reed. Biochemical relationships with the orientation of the autotrophic dinoflagellate Gymnodinium breve under nutrient replete conditions. Marine Ecology Progress Series, 167:105–117, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
167/m167p105.pdf.

Kwok:1999:RCT

[KN99] K. Y. Kwok and I. H. Ni. Reproduction of cutlassfishes Trichiurus spp. from the South China Sea. Marine Ecology Progress Series, 176:39–47, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/176/m176p039.pdf.

Kneib:1991:FWQ

[Kne91] R. T. Kneib. Flume weir for quantitative collection of nekton from vegetated intertidal habitats. Marine Ecology Progress Series, 75:29–38, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/75/m075p029.pdf.

Kneib:1993:GMS

[Kne93] R. T. Kneib. Growth and mortality in successive cohorts of fish larvae within an estuarine nursery. Marine Ecology Progress Series, 94:115–127, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/94/m094p115.pdf.

Kasamatsu:1995:BAS

[KNI95] F. Kasamatsu, S. Nishiwaki, and H. Ishikawa. Breeding areas and southbound migrations of southern minke whales Balaenoptera acutorostrata. Marine Ecology Progress Series, 119:1–10, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/119/m119p001.pdf.

Kemp:1990:EFF

[KNK90] P. F. Kemp, S. Y. Newell, and C. Krambeck. Effects of filter-feeding by the ribbed mussel Geukensia demissa on the water-column microbiota of a Spartina alterniflora saltmarsh. Marine Ecology Progress Series, 59:119–131, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
59/m059p119.pdf.

Kepkay:1993:LMW

[KNM93] P. E. Kepkay, S. E. H. Niven, and T. G. Milligan. Low molecular weight and colloidal DOC production during a phytoplankton bloom. Marine Ecology Progress Series, 100:233– 244, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/100/m100p233.pdf.

Kvitek:1992:ISO

[KO92] R. G. Kvitek and J. S. Oliver. Influence of sea otters on softbottom prey communities in southeast Alaska. Marine Ecology Progress Series, 82:103–113, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/82/m082p103.pdf.

Koga:1995:MBM

[Kog95] T. Koga. Movements between microhabitats depending on reproduction and life history in the sand-bubbler crab Scopimera globosa. Marine Ecology Progress Series, 117: 65–74, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/117/m117p065.pdf.

Kodama:1990:PAM

[KOSS90] M. Kodama, T. Ogata, S. Sato, and S. Sakamoto. Possible association of marine bacteria with paralytic shellfish toxicity of bivalves. Marine Ecology Progress Series, 61:203–206, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
61/m061p203.pdf.

Kromkamp:1995:PNP

[KP95] J. Kromkamp and J. Peene. Possibility of net phytoplankton primary production in the turbid Schelde Estuary (SW Netherlands). Marine Ecology Progress Series, 121:249–259, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/121/m121p249.pdf.

Kyewalyanga:1992:OPP

[KPS92] M. Kyewalyanga, T. Piatt, and S. Sathyendrath. Ocean primary production calculated by spectral and broad-band models. Marine Ecology Progress Series, 85:171–185, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
85/m085p171.pdf.

Kyewalyanga:1997:EPA

[KPS97] M. N. Kyewalyanga, T. Platt, and S. Sathyendranath. Estimation of the photosynthetic action spectrum: implication for primary production models. Marine Ecology Progress Series, 146:207–223, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/146/m146p207.pdf.

Klein:1993:DRT

[KPWL93] R. Klein, J. Patzold, G. Wefer, and Y. Loya. Depth-related timing of density band formation in Pontes spp. corals from the Red Sea inferred from X-ray chronology and stable isotope composition. Marine Ecology Progress Series, 97:99– 104, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/97/m097p099.pdf.

Kwint:1996:TDD

[KQH^{+}96] R. L. J. Kwint, P. Quist, T. A. Hansen, L. Dijkhuizen, and K. J. M. Kramer. Turnover of dimethylsulfoniopropionate and dimethylsulfide in the marine environment: a mesocosm experiment. Marine Ecology Progress Series, 145:223–232, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
145/m145p223.pdf.

Kroncke:1992:MIA

[KR92] I. Kroncke and E. Rachor. Macrofauna investigations along a transect from the inner German Bight towards the Dogger Bank. Marine Ecology Progress Series, 91:269–276, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
91/m091p269.pdf.

Kubler:1994:CLL

[KR94] J. E. Kübler and J. A. Raven. Consequences of light limitation for carbon acquisition in three rhodophytes. Marine Ecology Progress Series, 110:203–209, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/110/m110p203.pdf.

Kruse:1995:OES

[KR95] B. Kruse and B. Rasmussen. Occurrence and effects of a spring oxygen minimum layer in a stratified coastal water. Marine Ecology Progress Series, 125:293–303, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
125/m125p293.pdf.

Kaiser:1997:OFD

[KR97a] M. J. Kaiser and K. Ramsay. Opportunistic feeding by dabs within areas of trawl disturbance: possible implications for increased survival. Marine Ecology Progress Series, 152:307– 310, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/152/m152p307.pdf.

Kirkwood:1997:SCF

[KR97b] R. Kirkwood and G. Robertson. Seasonal change in the foraging ecology of emperor penguins on the Mawson Coast, Antarctica. Marine Ecology Progress Series, 156:205–223, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/156/m156p205.pdf.

Kinney:1998:RPP

[KR98a] E. H. Kinney and C. T. Roman. Response of primary producers to nutrient enrichment in a shallow estuary. Marine Ecology Progress Series, 163:89–98, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/163/m163p089.pdf.

Kukert:1998:PCI

[KR98b] H. Kukert and U. Riebesell. Phytoplankton carbon isotope fractionation during a diatom spring bloom in a Norwegian fjord. Marine Ecology Progress Series, 173:127–138, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
173/m173p127.pdf.

Kreeger:1993:SPU

[Kre93] D. A. Kreeger. Seasonal patterns in utilization of dietary protein by the mussel Mytilus trossulus. Marine Ecology Progress Series, 95:215–232, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/95/m095p215.pdf.

Kushmaro:1998:ETB

[KRF^{+}98] A. Kushmaro, E. Rosenberg, M. Fine, Y. Ben Haim, and Y. Loya. Effect of temperature on bleaching of the coral Oculina patagonica by Vibrio AK-1. Marine Ecology Progress Series, 171:131–137, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/171/m171p131.pdf.

Kushmaro:1997:BCO

[KRFL97] A. Kushmaro, E. Rosenberg, M. Fine, and Y. Loya. Bleaching of the coral Oculina patagonica by Vibrio AK-1. Marine Ecology Progress Series, 147:159–165, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/147/m147p159.pdf.

Kellogg:1995:GDR

[KRJ^{+}95] C. A. Kellogg, J. B. Rose, S. C. Jiang, J. M. Thurmond, and J. H. Paul. Genetic diversity of related vibriophages isolated from marine environments around Florida and Hawaii, USA. Marine Ecology Progress Series, 120:89–98, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
120/m120p089.pdf.

Korb:1998:RBA

[KRJ98] R. E. Korb, J. A. Raven, and A. M. Johnston. Relationship between aqueous CO_{2} concentrations and stable carbon isotope discrimination in the diatoms Chaetoceros calcitrans and Ditylum brightwellii. Marine Ecology Progress Series, 171:303–305, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/171/m171p303.pdf.

Korb:1996:ECS

[KRJL96] R. E. Korb, J. A. Raven, A. M. Johnston, and J. W. Leftley. Effects of cell size and specific growth rate on stable carbon isotope discrimination by two species of marine diatom. Marine Ecology Progress Series, 143:283–288, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
143/m143p283.pdf.

Kaiser:1994:IQS

[KRM94] M. J. Kaiser, S. I. Rogers, and D. T. McCandless. Improving quantitative surveys of epibenthic communities using a modified 2 m beam trawl. Marine Ecology Progress Series, 106:131–138, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/106/m106p131.pdf.

Krom:1991:IBP

[Kro91] M. D. Krom. Importance of benthic productivity in controlling the flux of dissolved inorganic nitrogen through the sediment–water interface in a hypertrophic marine ecosystem. Marine Ecology Progress Series, 78:163–172, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
78/m078p163.pdf.

Kruse:1993:MPR

[Kru93] B. Kruse. Measurement of plankton O_{2} respiration in gastight plastic bags. Marine Ecology Progress Series, 94:155– 163, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/94/m094p155.pdf.

Kiene:1991:DDD

[KS91] R. P. Kiene and S. K. Service. Decomposition of dissolved DMSP and DMS in estuanne waters: dependence on temperature and substrate concentration. Marine Ecology Progress Series, 76:1–11, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/76/m076p001.pdf.

Kaiser:1994:FSB

[KS94] M. J. Kaiser and B. E. Spencer. Fish scavenging behaviour in recently trawled areas. Marine Ecology Progress Series, 112: 41–49, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/112/m112p041.pdf.

Kaiser:1995:SCB

[KS95a] M. J. Kaiser and B. E. Spencer. Survival of by-catch from a beam trawl. Marine Ecology Progress Series, 126:31–38, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
126/m126p031.pdf.

Kiorboe:1995:SFG

[KS95b] T. Kiørboe and M. Sabatini. Scaling of fecundity, growth and development in marine planktonic copepods. Marine Ecology Progress Series, 120:285–298, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/120/m120p285.pdf.

Kiorboe:1995:PFC

[KS95c] T. Kiørboe and E. Saiz. Planktivorous feeding in calm and turbulent environments, with emphasis on copepods. Marine Ecology Progress Series, 122:135–145, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
122/m122p135.pdf.

Kivi:1995:SMF

[KS95d] K. Kivi and O. Setälä. Simultaneous measurement of food particle selection and clearance rates of planktonic oligotrich ciliates (Ciliophora: Oligotrichina). Marine Ecology Progress Series, 119:125–137, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/119/m119p125.pdf.

Konig:1997:HOM

[KS97] P. König and C. Schrum. Hydrographic observations and model results from a PRISMA drift experiment. Marine Ecology Progress Series, 156:255–261, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/156/m156p255.pdf.

Kenworthy:1998:VGS

[KS98] W. J. Kenworthy and A. C. Schwarzschild. Vertical growth and short-shoot demography of Syringodium filiforme in outer Florida Bay, USA. Marine Ecology Progress Series, 173: 25–37, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/173/m173p025.pdf.

Knutsen:1999:TDD

[KS99] I. Knutsen and A. G. V. Salvanes. Temperature-dependent digestion handling time in juvenile cod and possible consequences for prey choice. Marine Ecology Progress Series, 181: 61–79, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/181/m181p061.pdf.

Karakassis:1996:PNM

[KSE96] I. Karakassis, C. J. Smith, and A. Eleftheriou. Performance of neutral model analysis in a spatio-temporal series of macrobenthic replicates. Marine Ecology Progress Series, 137: 173–179, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/137/m137p173.pdf.

Kemp:1992:SDO

[KSG^{+}92] W. M. Kemp, P. A. Sampou, J. Garber, J. Turtle, and W. R. Boynton. Seasonal depletion of oxygen from bottom waters of Chesapeake Bay: roles of benthic and planktonic respiration and physical exchange processes. Marine Ecology Progress Series, 85:137–152, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/85/m085p137.pdf.

Kojima:1997:GDA

[KSH97] S. Kojima, R. Segawa, and I. Hayashi. Genetic differentiation among populations of the Japanese turban shell Turbo (Batillus) cornutus corresponding to warm currents. Marine Ecology Progress Series, 150:149–155, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/150/m150p149.pdf.

Kemp:1997:OCB

[KSMDB97] W. M. Kemp, E. M. Smith, M. Marvin-DiPasquale, and W. R. Boynton. Organic carbon balance and net ecosystem metabolism in Chesapeake Bay. Marine Ecology Progress Series, 150:229–248, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/150/m150p229.pdf.

Kraines:1997:CDC

[KSO^{+}97] S. Kraines, Y. Suzuki, T. Omori, K. Shitashima, S. Kanahara, and H. Komiyama. Carbonate dynamics of the coral reef system at Bora Bay, Miyako Island. Marine Ecology Progress Series, 156:1–16, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/156/m156p001.pdf.

Kang:1999:FSI

[KSRB99] C. K. Kang, P. G. Sauriau, P. Richard, and G. F. Blanchard. Food sources of an infaunal suspension-feeding bivalve Cerastoderma edule in a muddy sandflat of Marennes–Oléron Bay, as determined by analyses of carbon and nitrogen stable isotopes. Marine Ecology Progress Series, 187:147–158, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
187/m187p147.pdf.

Keats:1990:EER

[KSS90] D. W. Keats, G. R. South, and D. H. Steele. Effects of an experimental reduction in grazing by green sea urchins on a benthic macroalgal community in eastern Newfoundland. Marine Ecology Progress Series, 68:181–193, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
68/m068p181.pdf.

Kiorboe:1996:PSB

[KSV96] T. Kiørboe, E. Saiz, and M. Viitasalo. Prey switching behaviour in the planktonic copepod Acartia tonsa. Marine Ecology Progress Series, 143:65–75, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/143/m143p065.pdf.

Kiorboe:1999:HSP

[KSV99] T. Kiørboe, E. Saiz, and A. Visser. Hydrodynamic signal perception in the copepod Acartia tonsa. Marine Ecology Progress Series, 179:97–111, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/179/m179p097.pdf.

Kostka:1999:RPC

[KTGC99] J. E. Kostka, B. Thamdrup, R. N. Glud, and D. E. Canfield. Rates and pathways of carbon oxidation in permanently cold Arctic sediments. Marine Ecology Progress Series, 180:7–21, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/180/m180p007.pdf.

Karakassis:1998:SVS

[KTH98] I. Karakassis, M. Tsapakis, and E. Hatziyanni. Seasonal variability in sediment profiles beneath fish farm cages in the Mediterranean. Marine Ecology Progress Series, 162:243– 252, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/162/m162p243.pdf.

Kushnir:1997:SHB

[KTW^{+}97] V. M. Kushnir, Yu. N. Tokarev, R. Williams, S. A. Piontkovski, and P. V. Evstigneev. Spatial heterogeneity of the bioluminescence field of the tropical Atlantic Ocean and its relationship with internal waves. Marine Ecology Progress Series, 160:1–11, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/160/m160p001.pdf.

Kingsford:1991:HPR

[KUK91] M. J. Kingsford, A. J. Underwood, and S. J. Kennelly. Humans as predators on rocky reefs in New South Wales, Australia. Marine Ecology Progress Series, 72:1–14, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
72/m072p001.pdf.

Kuosa:1991:PAN

[Kuo91] H. Kuosa. Picoplanktonic algae in the northern Baltic Sea: seasonal dynamics and flagellate grazing. Marine Ecology Progress Series, 73:269–276, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/73/m073p269.pdf.

Kurihara:1999:EST

[Kur99] T. Kurihara. Effects of sediment type and food abundance on the vertical distribution of the starfish Asterina pectinifera. Marine Ecology Progress Series, 181:269–277, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
181/m181p269.pdf.

Kautsky:1990:MAV

[KV90] H. Kautsky and E. Van der Maarel. Multivariate approaches to the variation in phytobenthic communities and environmental vectors in the Baltic Sea. Marine Ecology Progress Series, 60:169–184, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/60/m060p169.pdf.

Kiorboe:1999:PPP

[KV99] T. Kiørboe and A. W. Visser. Predator and prey perception in copepods due to hydromechanical signals. Marine Ecology Progress Series, 179:81–95, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/179/m179p081.pdf.

Katajisto:1998:SOH

[KVK98] T. Katajisto, M. Viitasalo, and M. Koski. Seasonal occurrence and hatching of calanoid eggs in sediments of the northern Baltic Sea. Marine Ecology Progress Series, 163: 133–143, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/163/m163p133.pdf.

Kendrick:1991:DDP

[KW91] G. A. Kendrick and D. I. Walker. Dispersal distances for propagules of Sargassum spinuligerum (Sargassaceae, Phaeophyta) measured directly by vital staining and venturi suction sampling. Marine Ecology Progress Series, 79:133–138, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
79/m079p133.pdf.

Kepkay:1992:DOC

[KW92a] P. E. Kepkay and M. L. Wells. Dissolved organic carbon in North Atlantic surface waters. Marine Ecology Progress Series, 80:275–283, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/80/m080p275.pdf.

Koppelmann:1992:FDZ

[KW92b] R. Koppelmann and H. Weikert. Full-depth zooplankton profiles over the deep bathyal of the NE Atlantic. Marine Ecology Progress Series, 86:263–272, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/86/m086p263.pdf.

Kneib:1994:NUV

[KW94] R. T. Kneib and S. L. Wagner. Nekton use of vegetated marsh habitats at different stages of tidal inundation. Marine Ecology Progress Series, 106:227–238, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/106/m106p227.pdf.

Kaehler:1997:DFI

[KW97a] S. Kaehler and G. A. Williams. Do factors influencing recruitment ultimately determine the distribution and abundance of encrusting algae on seasonal tropical shores? Marine Ecology Progress Series, 156:87–96, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/156/m156p087.pdf.

Kennish:1997:FPH

[KW97b] R. Kennish and G. A. Williams. Feeding preferences of the herbivorous crab Grapsus albolineatus: the differential influence of algal nutrient content and morphology. Marine Ecology Progress Series, 147:87–95, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/147/m147p087.pdf.

Kaehler:1998:EDA

[KW98] S. Kaehler and G. A. Williams. Early development of algal assemblages under different regimes of physical and biotic factors on a seasonal tropical rocky shore. Marine Ecology Progress Series, 172:61–71, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/172/m172p061.pdf.

Koppelmann:1999:TCD

[KW99a] R. Koppelmann and H. Weikert. Temporal changes of deepsea mesozooplankton abundance in the temperate NE Atlantic and estimates of the carbon budget. Marine Ecology Progress Series, 179:27–40, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/179/m179p027.pdf.

Kuipers:1999:GIM

[KW99b] B. R. Kuipers and H. J. Witte. Grazing impact of microzooplankton on different size classes of algae in the North Sea in early spring and mid-summer. Marine Ecology Progress Series, 180:93–104, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/180/m180p093.pdf.

Kahn:1993:IDD

[KWD93] B. Kahn, H. Whitehead, and M. Dillon. Indications of density-dependent effects from comparisons of sperm whale populations. Marine Ecology Progress Series, 93:1–7, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
93/m093p001.pdf.

Kock:1994:VDC

[KWEG94] K.-H. Kock, S. Wilhelms, I. Everson, and J. Groger. Variations in the diet composition and feeding intensity of mackerel icefish Champsocephalus gunnari at South Georgia (Antarctic). Marine Ecology Progress Series, 108:43–57, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
108/m108p043.pdf.

Koziol:1996:CHI

[KWHM^{+}96] C. Koziol, C. Wagner-Hülsmann, A. Mikoc, V. Gamulin, M. Kruse, Z. Pancer, H. Schäcke, and W. E. G. Müller. Cloning of a heat-inducible biomarker, the cDNA encoding the 70 kDa heat shock protein, from the marine sponge Geodia cydonium: response to natural stressors. Marine Ecology Progress Series, 136:153–161, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/136/m136p153.pdf.

Kramarsky-Winter:1996:RVB

[KWL96] E. Kramarsky-Winter and Y. Loya. Regeneration versus budding in fungiid corals: a trade-off. Marine Ecology Progress Series, 134:179–185, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/134/m134p179.pdf.

Kramarsky-Winter:1998:RST

[KWL98] E. Kramarsky-Winter and Y. Loya. Reproductive strategies of two fungiid corals from the northern Red Sea: environmental constraints? Marine Ecology Progress Series, 174: 175–182, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/174/m174p175.pdf.

Kerner:1997:UPS

[KY97] M. Kerner and S. Yasseri. Utilization of phytoplankton in seston aggregates from the Elbe estuary, Germany, during early degradation processes. Marine Ecology Progress Series, 158:87–102, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/158/m158p087.pdf.

Kim:1998:CRB

[KYI^{+}98] M. C. Kim, I. Yoshinaga, I. Imai, K. Nagasaki, S. Itakura, A. Uchida, and Y. Ishida. A close relationship between algicidal bacteria and termination of Heterosigma akashiwo (Raphidophyceae) blooms in Hiroshima Bay, Japan. Marine Ecology Progress Series, 170:25–32, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/170/m170p025.pdf.

Kuwamura:1994:PDG

[KYN94] T. Kuwamura, Y. Yogo, and Y. Nakashima. Population dynamics of goby Paragobiodon echmocephalus and host coral Stytophora pistillata. Marine Ecology Progress Series, 103: 17–23, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/103/m103p017.pdf.

Kyomo:1992:VFH

[Kyo92] J. Kyomo. Variations in the feeding habits of males and females of the crab Sesarma intermedia. Marine Ecology Progress Series, 83:151–155, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/83/m083p151.pdf.

Kunzmann:1992:AMH

[KZ92] A. Kunzmann and C. Zimmermann. Aethotaxis mitopteryx, a high-Antarctic fish with benthopelagic mode of life. Marine Ecology Progress Series, 88:33–40, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/88/m088p033.pdf.

Kharlamenko:1995:FAM

[KZK^{+}95] V. I. Kharlamenko, N. V. Zhukova, S. V. Khotimchenko, V. I. Svetashev, and G. M. Kamenev. Fatty acids as markers of food sources in a shallow-water hydrothermal ecosystem (Kraternaya Bight, Yankich Island, Kurile Islands). Marine Ecology Progress Series, 120:231–241, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/120/m120p231.pdf.

Kirshtein:1993:DNF

[KZP93] J. D. Kirshtein, J. P. Zehr, and H. W. Paerl. Determination of N2 fixation potential in the marine environment: application of the polymerase chain reaction. Marine Ecology Progress Series, 95:305–309, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/95/m095p305.pdf.

Larned:1997:EWV

[LA97a] S. T. Larned and M. J. Atkinson. Effects of water velocity on NH 4 and PO 4 uptake and nutrient-limited growth in the macroalga Dictyosphaeria cavernosa. Marine Ecology Progress Series, 157:295–302, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/157/m157p295.pdf.

Lombarte:1997:QDC

[LA97b] A. Lombarte and H. Aguirre. Quantitative differences in the chemoreceptor systems in the barbels of two species of Mullidae (Mullus surmuletus and M. barbatus) with different bottom habitats. Marine Ecology Progress Series, 150: 57–64, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/150/m150p057.pdf.

Lehtonen:1998:PDR

[LA98] K. K. Lehtonen and A. B. Andersin. Population dynamics, response to sedimentation and role in benthic metabolism of the amphipod Monoporeia affinis in an open-sea area of the northern Baltic Sea. Marine Ecology Progress Series, 168: 71–85, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/168/m168p071.pdf.

Lawrence:1993:FPB

[LAA93] S. G. Lawrence, A. Ahmad, and F. Azam. Fate of particlebound bacteria ingested by Calanus pacificus. Marine Ecology Progress Series, 97:299–307, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/97/m097p299.pdf.

Livingstone:1992:AEL

[LACM92] D. R. Livingstone, S. Archibald, J. K. Chipman, and J. W. Marsh. Antioxidant enzymes in liver of dab Limanda limanda from the North Sea. Marine Ecology Progress Series, 91:97–104, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/91/m091p097.pdf.

Lawson:1998:SFH

[LADS98] J. W. Lawson, J. T. Anderson, E. L. Dalley, and G. B. Stenson. Selective foraging by harp seals Phoca groenlandica in nearshore and offshore waters of Newfoundland, 1993 and 1994. Marine Ecology Progress Series, 163:1–10, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
163/m163p001.pdf.

Lafferty:1993:EPC

[Laf93] K. D. Lafferty. Effects of parasitic castration on growth reproduction and population dynamics of the marine snail Cerithidea californica. Marine Ecology Progress Series, 96: 229–237, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/96/m096p229.pdf.

Leakey:1996:MDC

[LAG96] R. J. G. Leakey, S. D. Archer, and J. Grey. Microbial dynamics in coastal waters of East Antarctica: bacterial production and nanoflagellate bacterivory. Marine Ecology Progress Series, 142:3–17, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/142/m142p003.pdf.

Laima:1992:ESV

[Lai92] M. J. C. Laima. Extraction and seasonal variation of NH^{4+} pools in different types of coastal marine sediments. Marine Ecology Progress Series, 82:75–84, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/82/m082p075.pdf.

Lindegarth:1995:ASV

[LAJ95] M. Lindegarth, M. C. Andre, and P. R. Jonsson. Analysis of the spatial variability in abundance and age structure of two infaunal bivalves, Cerastoderma edule and C. lamarcki, using hierarchical sampling programs. Marine Ecology Progress Series, 116:85–97, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/116/m116p085.pdf.

Lindholm:1999:HMS

[LAK99] J. B. Lindholm, P. J. Auster, and L. S. Kaufman. Habitatmediated survivorship of juvenile (0-year) Atlantic cod Gadus morhua . Marine Ecology Progress Series, 180:247–255, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
180/m180p247.pdf.

Lamare:1998:OTL

[Lam98] M. D. Lamare. Origin and transport of larvae of the sea urchin Evechinus chloroticus (Echinodermata: Echinoidea) in a New Zealand fiord. Marine Ecology Progress Series, 174:107–121, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/174/m174p107.pdf.

Lasiak:1998:MCR

[Las98] T. Lasiak. Multivariate comparisons of rocky infratidal macrofaunal assemblages from replicate exploited and nonexploited localities on the Transkei coast of South Africa. Marine Ecology Progress Series, 167:15–23, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
167/m167p015.pdf.

Lavaniegos:1994:DDP

[Lav94] B. E. Lavaniegos. Dispersion and development patterns in larvae of Nyctiphanes simplex (Euphausiacea) in the upwelling region off Baja California. Marine Ecology Progress Series, 106:207–225, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/106/m106p207.pdf.

Laval:1997:VMA

[Lav97] P. Laval. A virtual mesocosm with artificial salps for exploring the conditions of swarm development in the pelagic tunicate Salpa fusiformis. Marine Ecology Progress Series, 154:1–16, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/154/m154p001.pdf.

LeFevre:1991:NNC

[LB91] J. Le Fèvre and E. Bourget. Neustonic niche for cirripede larvae as a possible adaptation to long-range dispersal. Marine Ecology Progress Series, 74:185–194, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
74/m074p185.pdf.

Luo:1993:BAA

[LB93] J. Luo and S. B. Brandt. Bay anchovy Anchoa mitchilli production and consumption in mid-Chesapeake Bay based on a bioenergetics model and acoustic measures of fish abundance. Marine Ecology Progress Series, 98:223–236, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
98/m098p223.pdf.

Lasiak:1995:RBM

[LB95] T. A. Lasiak and T. C. E. Barnard. Recruitment of the brown mussel Perna perna onto natural substrata: a refutation of the primary/secondary settlement hypothesis. Marine Ecology Progress Series, 120:147–153, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/120/m120p147.pdf.

LeTissier:1996:DSB

[LB96a] M. D. A. A. Le Tissier and B. E. Brown. Dynamics of solar bleaching in the intertidal reef coral Goniastrea aspera at Ko Phuket, Thailand. Marine Ecology Progress Series, 136:235–244, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/136/m136p235.pdf.

Lemire:1996:SHC

[LB96b] M. Lemire and E. Bourget. Substratum heterogeneity and complexity influence micro-habitat selection of Balanus sp. and Tubularia crocea larvae. Marine Ecology Progress Series, 135:77–87, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/135/m135p077.pdf.

Lamare:1999:LSE

[LB99a] M. D. Lamare and M. F. Barker. In situ estimates of larval development and mortality in the New Zealand sea urchin Evechinus chloroticus (Echinodermata: Echinoidea). Marine Ecology Progress Series, 180:197–211, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
180/m180p197.pdf.

Lapointe:1999:ISH

[LB99b] L. Lapointe and E. Bourget. Influence of substratum heterogeneity scales and complexity on a temperate epibenthic marine community. Marine Ecology Progress Series, 189: 159–170, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/189/m189p159.pdf.

Lemieux:1997:MHC

[LBDD97] H. Lemieux, P. U. Blier, F. Dufresne, and G. Desrosiers. Metabolism and habitat competition in the polychaete Nereis virens. Marine Ecology Progress Series, 156:151–156, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
156/m156p151.pdf.

LincolnSmith:1991:SVA

[LBH91] M. P. Lincoln Smith, J. D. Bell, and C. A. Hair. Spatial variation in abundance of recently settled rocky reef fish in southeastern Australia: implications for detecting change. Marine Ecology Progress Series, 77:95–103, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/77/m077p095.pdf.

Lavrentyev:1998:MPR

[LBJ^{+}98] P. J. Lavrentyev, H. A. Bootsma, T. H. Johengen, J. F. Cavaletto, and W. S. Gardner. Microbial plankton response to resource limitation: insights from the community structure and seston stoichiometry in Florida Bay, USA. Marine Ecology Progress Series, 165:45–57, 1998. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/165/m165p045.pdf.

Lubbers:1990:VSE

[LBK90] L. Lubbers, W. R. Boynton, and W. M. Kemp. Variations in structure of estuarine fish communities in relation to abundance of submersed vascular plants. Marine Ecology Progress Series, 65:1–14, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/65/m065p001.pdf.

Luo:1996:VRP

[LBK96] J. Luo, S. B. Brandt, and M. J. Klebasko. Virtual reality of planktivores: a fish’s perspective of prey size selection. Marine Ecology Progress Series, 140:271–283, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
140/m140p271.pdf.

Laval:1991:DPF

[LBL91] Philippe Laval, Jean Claude Braconnot, and Nadja Lins da Silva. Deep planktonic filter-feeders found in the aphotic zone with the Cyana submersible in the Ligurian Sea (NW Mediterranean). Marine Ecology Progress Series, 79:235– 241, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/79/m079p235.pdf.

Lindquist:1997:TLR

[LBL97] N. Lindquist, R. Bolser, and K. Laing. Timing of larval release by two Caribbean demosponges. Marine Ecology Progress Series, 155:309–313, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/155/m155p309.pdf.

Liu:1999:CCP

[LBL^{+}99] H. Liu, R. R. Bidigare, E. Laws, M. R. Landry, and L. Campbell. Cell cycle and physiological characteristics of Synechococcus (WH7803) in chemostat culture. Marine Ecology Progress Series, 189:17–25, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/189/m189p017.pdf.

Lee:1993:BSS

[LBLA93] T. Lee, E. Barg, D. Lal, and F. Azam. Bacterial scavenging of 234Th in surface ocean waters. Marine Ecology Progress Series, 96:109–116, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/96/m096p109.pdf.

Levin:1999:MPP

[LBM^{+}99] L. A. Levin, N. E. Blair, C. M. Martin, D. J. DeMaster, G. Plaia, and C. J. Thomas. Macrofaunal processing of phytodetritus at two sites on the Carolina margin: in situ experiments using ^{13}C-labeled diatoms. Marine Ecology Progress Series, 182:37–54, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/182/m182p037.pdf.

Lamy:1999:UDA

[LBV^{+}99] F. Lamy, M. Bianchi, F. Van Wambeke, R. Sempéré, and V. Talbot. Use of data assimilation techniques to analyze the significance of ectoproteolytic activity measurements performed with the model substrate MCA-leu. Marine Ecology Progress Series, 177:27–35, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/177/m177p027.pdf.

Lewitus:1990:REN

[LC90] A. J. Lewitus and D. A. Caron. Relative effects of nitrogen or phosphorus depletion and light intensity on the pigmentation, chemical composition, and volume of Pyrenomonas salina (Cryptophyceae). Marine Ecology Progress Series, 61: 171–181, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/61/m061p171.pdf.

LaFerla:1991:PSV

[LC91] R. La Ferla and E. Crisafi. Preliminary study on vertical distribution of microorganisms in the Bannock Basin waters (eastern Mediterranean Sea). Marine Ecology Progress Series, 75:309–311, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/75/m075p309.pdf.

Leonard:1993:PQV

[LC93] G. H. Leonard and R. P. Clark. Point quadrat versus video transect estimates of the cover of benthic red algae. Marine Ecology Progress Series, 101:203–208, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
101/m101p203.pdf.

Lee:1995:AIN

[LC95] R. W. Lee and J. J. Childress. Assimilation of inorganic nitrogen by seep mytilid Ia, an undescribed deep-sea mussel containing methanotrophic endosymbionts: fate of assimilated nitrogen and the relation between methane and nitrogen assimilation. Marine Ecology Progress Series, 123: 137–148, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/123/m123p137.pdf.

Lewis:1997:GUR

[LC97a] D. E. Lewis and R. M. Cerrato. Growth uncoupling and the relationship between shell growth and metabolism in the soft shell clam Mya arenaria. Marine Ecology Progress Series, 158:177–189, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/158/m158p177.pdf.

Limpus:1997:NRM

[LC97b] C. Limpus and M. Chaloupka. Nonparametric regression modelling of green sea turtle growth rates (southern Great Barrier Reef). Marine Ecology Progress Series, 149:23–34, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/149/m149p023.pdf.

Lazo:1998:CCM

[LC98] M. L. Lazo and A. R. O. Chapman. Components of crowding in a modular seaweed: sorting through the contradictions. Marine Ecology Progress Series, 174:257–267, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
174/m174p257.pdf.

LeTissier:1994:SCC

[LCBD94] M. D’A. A. Le Tissier, B. Clayton, B. E. Brown, and P. S. Davis. Skeletal correlates of coral density banding and an evaluation of radiography as used in sclerochronology. Marine Ecology Progress Series, 110:29–44, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/110/m110p029.pdf.

Lin:1997:CNT

[LCC97] S. Lin, J. Chang, and E. J. Carpenter. Can a non-terminal event of the cell cycle be used for phytoplankton speciesspecific growth rate estimation? Marine Ecology Progress Series, 151:283–290, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/151/m151p283.pdf.

Leis:1997:LSS

[LCE97] J. M. Leis and B. M. Carson-Ewart. In situ swimming speeds of the late pelagic larvae of some Indo–Pacific coral-reef fishes. Marine Ecology Progress Series, 159:165–174, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
159/m159p165.pdf.

Lennert-Cody:1999:PPH

[LCF99] C. E. Lennert-Cody and P. J. S. Franks. Plankton patchiness in high-frequency internal waves. Marine Ecology Progress Series, 186:59–66, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/186/m186p059.pdf.

Levin:1997:GDR

[LCG97] P. S. Levin, W. Chiasson, and J. M. Green. Geographic differences in recruitment and population structure of a temperate reef fish. Marine Ecology Progress Series, 161:23–35, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
161/m161p023.pdf.

Lonsdale:1996:FWI

[LCK^{+}96] D. J. Lonsdale, E. M. Cosper, W. S. Kim, M. Doall, A. Divadeenam, and S. H. Jonasdottir. Food web interactions in the plankton of Long Island bays, with preliminary observations on brown tide effects. Marine Ecology Progress Series, 134:247–263, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/134/m134p247.pdf.

Liu:1995:GMR

[LCL95] H. Liu, L. Campbell, and M. R. Landry. Growth and mortality rates of Prochlorococcus and Synechococcus measured with a selective inhibitor technique. Marine Ecology Progress Series, 116:277–287, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/116/m116p277.pdf.

Lemmens:1996:FCS

[LCLC96] J. W. T. J. Lemmens, G. Clapin, P. Lavery, and J. Cary. Filtering capacity of seagrass meadows and other habitats of Cockburn Sound, Western Australia. Marine Ecology Progress Series, 143:187–200, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/143/m143p187.pdf.

Lin:1992:ETS

[LCTT92] H.-P. Lin, G. Charmantier, P. Thuet, and J.-P. Trilles. Effects of turbidity on survival, osmoregulation and gill na+- k+ ATPase in juvenile shrimp Penaeus japonicus. Marine Ecology Progress Series, 90:31–37, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/90/m090p031.pdf.

Leung:1999:NBA

[LCW99] K. M. Y. Leung, J. C. W. Chu, and R. S. S. Wu. Nitrogen budgets for the areolated grouper Epinephelus areolatus cultured under laboratory conditions and in open-sea cages. Marine Ecology Progress Series, 186:271–281, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
186/m186p271.pdf.

Laprise:1993:NEV

[LD93] R. Laprise and J. J. Dodson. Nature of environmental variability experienced by benthic and pelagic animals in the St. Lawrence Estuary, Canada. Marine Ecology Progress Series, 94:129–139, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/94/m094p129.pdf.

Laprise:1994:EVF

[LD94a] Ft Laprise and J. J. Dodson. Environmental variability as a factor controlling spatial patterns in distribution and species diversity of zooplankton in the St. Lawrence Estuary. Marine Ecology Progress Series, 107:67–81, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/107/m107p067.pdf.

Ledyard:1994:DPD

[LD94b] K. M. Ledyard and J. W. H. Dacey. Dimethylsulfide production from dimethylsulfonio-propionate by a marine bacterium. Marine Ecology Progress Series, 110:95–103, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
110/m110p095.pdf.

Leggett:1995:RRW

[LD95a] W. C. Leggett and E. M. DeBlois. Researching recruitment — where to now? Marine Ecology Progress Series, 128:309–310, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Levin:1995:IST

[LD95b] L. A. Levin and C. DiBacco. Influence of sediment transport on short-term recolonization by seamount infauna. Marine Ecology Progress Series, 123:163–175, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
123/m123p163.pdf.

Lee:1996:PCR

[LD96] K. S. Lee and K. H. Dunton. Production and carbon reserve dynamics of the seagrass Thalassia testudinum in Corpus Christi Bay, Texas, USA. Marine Ecology Progress Series, 143:201–210, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/143/m143p201.pdf.

Litaker:1993:STE

[LDKR93] W. Litaker, C. S. Duke, B. E. Kenney, and J. Ramus. Shortterm environmental variability and phytoplankton abundance in a shallow tidal estuary II. Spring and fall. Marine Ecology Progress Series, 94:141–154, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/94/m094p141.pdf.

Laurion:1995:MFW

[LDV95] I. Laurion, S. Demers, and A. F. Vézina. The microbial food web associated with the ice algal assemblage: biomass and bacterivory of nanoflagellate protozoans in Resolute Passage (High Canadian Arctic). Marine Ecology Progress Series, 120:77–87, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/120/m120p077.pdf.

Little:1990:MRI

[LE90] K. T. Little and C. E. Epifanio. Mechanism for the reinvasion of an estuary by two species of brachyuran megalopae. Marine Ecology Progress Series, 68:235–242, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
68/m068p235.pdf.

Latasa:1992:PPR

[LED92] M. Latasa, M. Estrada, and M. Delgado. Plankton-pigment relationships in the Northwestern Mediterranean during stratification. Marine Ecology Progress Series, 88:61–73, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
88/m088p061.pdf.

Lee:1990:NAP

[Lee90] S. Y. Lee. Net aerial primary productivity, litter production and decomposition of the reef Phragmites communis in a nature reserve in Hong Kong: management implications. Marine Ecology Progress Series, 66:161–173, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
66/m066p161.pdf.

Lee:1997:PTI

[Lee97] S. Y. Lee. Potential trophic importance of the faecal material of the mangrove sesarmine crab Sesarma messa. Marine Ecology Progress Series, 159:275–284, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/159/m159p275.pdf.

Lehtonen:1996:EBA

[Leh96] K. K. Lehtonen. Ecophysiology of the benthic amphipod Monoporeia affinis in an open-sea area of the northern Baltic Sea: seasonal variations in body composition, with bioenergetic considerations. Marine Ecology Progress Series, 143: 87–98, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/143/m143p087.pdf.

Lohse:1996:OPW

[LEHvR96] L. Lohse, E. H. G. Epping, W. Helder, and W. van Raaphorst. Oxygen pore water profiles in continental shelf sediments of the North Sea: turbulent versus molecular diffusion. Marine Ecology Progress Series, 145:63–75, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/145/m145p063.pdf.

Leonard:1999:PNE

[Leo99] G. H. Leonard. Positive and negative effects of intertidal algal canopies on recruitment and survival of barnacles. Marine Ecology Progress Series, 178:241–249, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
178/m178p241.pdf.

Lesser:1996:APU

[Les96] M. P. Lesser. Acclimation of phytoplankton to UV-B radiation: oxidative stress and photoinhibition of photosynthesis are not prevented by UV-absorbing compounds in the dinoflagellate Prorocentrum micans. Marine Ecology Progress Series, 132:287–297, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/132/m132p287.pdf.

Levin:1991:EMR

[Lev91] P. S. Levin. Effects of microhabitat on recruitment variation in a Gulf of Maine reef fish. Marine Ecology Progress Series, 75:183–189, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/75/m075p183.pdf.

Lewis:1991:TCF

[Lew91] J. B. Lewis. Testing the coral fragment size-dependent survivorship hypothesis for the calcareous hydrozoan Millepora complanata. Marine Ecology Progress Series, 70:101–104, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/70/m070p101.pdf.

Lewis:1992:HCZ

[Lew92] J. B. Lewis. Heterotrophy in corals: zooplankton predation by the hydrocoral Millepora complanata. Marine Ecology Progress Series, 90:251–256, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/90/m090p251.pdf.

Lewis:1997:EEC

[Lew97] A. R. Lewis. Effects of experimental coral disturbance on the structure of fish communities on large patch reefs. Marine Ecology Progress Series, 161:37–50, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/161/m161p037.pdf.

Lee:1991:SCS

[LF91] S. Lee and J. A. Fuhrman. Species composition shift of confined bacterioplankton studied at the level of community DNA. Marine Ecology Progress Series, 79:195–201, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
79/m079p195.pdf.

Lee:1992:DRE

[LF92] B.-G. Lee and N. S. Fisher. Decomposition and release of elements from zooplankton debris. Marine Ecology Progress Series, 88:117–128, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/88/m088p117.pdf.

Logan:1993:MKC

[LF93] B. E. Logan and R. C. Fleury. Multiphasic kinetics can be an artifact of the assumption of saturable kinetics for microorganisms. Marine Ecology Progress Series, 102:115–124, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
102/m102p115.pdf.

Lee:1994:ESZ

[LF94] B.-G. Lee and N. S. Fisher. Effects of sinking and zooplankton grazing on the release of elements from planktonic debris. Marine Ecology Progress Series, 110:271–281, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
110/m110p271.pdf.

Lotufo:1997:ESA

[LF97] G. R. Lotufo and J. W. Fleeger. Effects of sedimentassociated phenanthrene on survival, development and reproduction of two species of meiobenthic copepods. Marine Ecology Progress Series, 151:91–102, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/151/m151p091.pdf.

Lotan:1994:SLC

[LFBH94] A. Lotan, M. Fine, and R. Ben-Hillel. Synchronization of the life cycle and dispersal pattern of the tropical invader scyphomedusan Rho-pilema nomadica is temperature dependent. Marine Ecology Progress Series, 109:59–65, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
109/m109p059.pdf.

Lowe:1995:CIL

[LFD95] D. M. Lowe, V. U. Fossato, and M. H. Depledge. Contaminantinduced lysosomal membrane damage in blood cells of mussels Mytilus galloprovincialis from the Venice Lagoon: an in vitro study. Marine Ecology Progress Series, 129:189–196, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
129/m129p189.pdf.

Lye:1998:SRH

[LFG98] C. M. Lye, C. L. J. Frid, and M. E. Gill. Seasonal reproductive health of flounder Platichthys flesus exposed to sewage effluent. Marine Ecology Progress Series, 170:249–260, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
170/m170p249.pdf.

Lohrenz:1997:VPP

[LFR^{+}97] S. E. Lohrenz, G. L. Fahnenstiel, D. G. Redalje, G. A. Lang, X. Chen, and M. J. Dagg. Variations in primary production of northern Gulf of Mexico continental shelf waters linked to nutrient inputs from the Mississippi River. Marine Ecology Progress Series, 155:45–54, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/155/m155p045.pdf.

Labarta:1997:DPE

[LFRB97] U. Labarta, M. J. Fernández-Reiríz, and J. M. F. Babarro. Differences in physiological energetics between intertidal and raft cultivated mussels Mytilus galloprovincialis. Marine Ecology Progress Series, 152:167–173, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/152/m152p167.pdf.

Lindberg:1990:PER

[LFS90] W. J. Lindberg, T. K. Frazer, and G. R. Stanton. Population effects of refuge dispersion for adult stone crabs (Xanthidae, Menippe). Marine Ecology Progress Series, 66:239–249, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
66/m066p239.pdf.

Levasseur:1992:PDC

[LFTH92] M. Levasseur, L. Fortier, J.-C. Therriault, and P. J. Harrison. Phytoplankton dynamics in a coastal jet frontal region. Marine Ecology Progress Series, 86:283–295, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
86/m086p283.pdf.

LeFur:1991:MSM

[LGB91] J. Le Fur, P. Germain, and Y. Baron. Modelling in situ mussel ruthemum-106 concentration dynamics due to fluctuating seawater radionuclide concentrations. Marine Ecology Progress Series, 75:67–78, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/75/m075p067.pdf.

LeMaho:1993:UBP

[LGC^{+}93] Yyon Le Maho, J.-R. Gendner, Étienne Challet, Charles-André Bost, Janick Gilles, Christophe Verdon, Chritiane Plumeré, Jean-Patrice Robin, and Yves Handrich. Undisturbed breeding penguins as indicators of changes in marine resources. Marine Ecology Progress Series, 95:1–6, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
95/m095p001.pdf.

Lubbers:1990:MAH

[LGdC^{+}90] G. W. Lubbers, W. W. C. Gieskes, P. del Castilho, W. Salomons, and J. Bril. Manganese accumulation in the high pH microenvironment of Phaeocystis sp. (Haptophyceae) colonies from the North Sea. Marine Ecology Progress Series, 59: 285–293, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/59/m059p285.pdf.

Lindley:1998:VCC

[LGED98] J. A. Lindley, C. L. George, S. V. Evans, and P. Donkin. Viability of calanoid copepod eggs from intertidal sediments: a comparison of three estuaries. Marine Ecology Progress Series, 162:183–190, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/162/m162p183.pdf.

Linke-Gamenick:1999:DDE

[LGFS99] I. Linke-Gamenick, V. E. Forbes, and R. M. Sibly. Densitydependent effects of a toxicant on life-history traits and population dynamics of a capitellid polychaete. Marine Ecology Progress Series, 184:139–148, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/184/m184p139.pdf.

Lavaniegos:1998:LTC

[LGGLLHV98] B. E. Lavaniegos, J. Gómez-Gutiérrez, J. R. Lara-Lara, and S. Hernández-Vázquez. Long-term changes in zooplankton volumes in the California Current system–the Baja California region. Marine Ecology Progress Series, 169:55–64, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
169/m169p055.pdf.

Legendre:1993:ECP

[LGH^{+}93] L. Legendre, M. Gosselin, H.-J. Hirche, G. Kattner, and G. Rosenberg. Environmental control and potential fate of size-fractionated phytoplankton production in the Greenland Sea (75^{o}N). Marine Ecology Progress Series, 98:297– 313, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/98/m098p297.pdf.

LeCampion-Alsumard:1995:MES

[LGH95a] T. Le Campion-Alsumard, S. Golubic, and P. Hutchings. Microbial endoliths in skeletons of live and dead corals: Porites lobata (Moorea, French Polynesia). Marine Ecology Progress Series, 117:149–157, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/117/m117p149.pdf.

Lindley:1995:CZC

[LGH95b] J. A. Lindley, J. C. Gamble, and H. G. Hunt. A change in the zooplankton of the central North Sea (55^{o}to 58^{o}N): a possible consequence of changes in the benthos. Marine Ecology Progress Series, 119:299–303, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/119/m119p299.pdf.

Leclercq:1999:MOM

[LGJ99] N. Leclercq, J. P. Gattuso, and J. Jaubert. Measurement of oxygen metabolism in open-top aquatic mesocosms: application to a coral reef community. Marine Ecology Progress Series, 177:299–304, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/177/m177p299.pdf.

Lozano:1995:BCR

[LGL^{+}95] J. Lozano, J. Galera, S. López, X. Turon, and C. Palacín. Biological cycles and recruitment of Paracentrotus lividus (Echinodermata: Echinoidea) in two contrasting habitats. Marine Ecology Progress Series, 122:179–191, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
122/m122p179.pdf.

Lynch:1998:BPS

[LGMD98] D. R. Lynch, W. C. Gentleman, D. J. McGillicuddy, Jr., and C. S. Davis. Biological/physical simulations of Calanus finmarchicus population dynamics in the Gulf of Maine. Marine Ecology Progress Series, 169:189–210, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
169/m169p189.pdf.

LeCampion-Alsumard:1995:FCS

[LGP95] T. Le Campion-Alsumard, S. Golubic, and K. Priess. Fungi in corals: symbiosis or disease? Interaction between polyps and fungi causes pearl-like skeleton biomineralization. Marine Ecology Progress Series, 117:137–147, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
117/m117p137.pdf.

Lorenzen:1995:IMC

[LGR95] J. Lorenzen, R. N. Glud, and N. P. Revsbech. Impact of microsensor-caused changes in diffusive boundary layer thickness on O 2 profiles and photosynthetic rates in benthic communities of microorganisms. Marine Ecology Progress Series, 119:237–241, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/119/m119p237.pdf.

Laima:1999:DAA

[LGV^{+}99a] M. J. C. Laima, M. F. Girard, F. Vouvé, G. F. Blanchard, D. Gouleau, R. Galois, and P. Richard. Distribution of adsorbed ammonium pools in two intertidal sedimentary structures, Marennes–Oléron Bay, France. Marine Ecology Progress Series, 182:29–35, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/182/m182p029.pdf.

Laima:1999:NRR

[LGV^{+}99b] M. J. C. Laima, M. F. Girard, F. Vouvé, P. Richard, G. Blanchard, and D. Gouleau. Nitrification rates related to sedimentary structures in an Atlantic intertidal mudflat, Marennes– Oléron Bay, France. Marine Ecology Progress Series, 191: 33–41, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/191/m191p033.pdf.

Lim:1991:GEC

[LH91] B. K. Lim and K. Hirayama. Growth and elemental composition (C, N, P) during larval developmental stages of mass-cultured swimming crab Portunus trituberculatus. Marine Ecology Progress Series, 78:131–137, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
78/m078p131.pdf.

Levin:1996:RTR

[LH96] P. S. Levin and M. E. Hay. Responses of temperate reef fishes to alterations in algal structure and species composition. Marine Ecology Progress Series, 134:37–47, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
134/m134p037.pdf.

Landen:1998:SVD

[LH98] A. Landén and P. O. J. Hall. Seasonal variation of dissolved and adsorbed amino acids and ammonium in a nearshore marine sediment. Marine Ecology Progress Series, 170: 67–84, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/170/m170p067.pdf.

Lucas:1999:NEI

[LH99] C. H. Lucas and P. M. Holligan. Nature and ecological implications of algal pigment diversity on the molenplaat tidal flat (Westerschelde estuary, SW Netherlands). Marine Ecology Progress Series, 180:51–64, 1999. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/180/m180p051.pdf.

Leppard:1996:UMS

[LHH96] G. G. Leppard, A. Heissenberger, and G. J. Herndl. Ultrastructure of marine snow. I. Transmission electron microscopy methodology. Marine Ecology Progress Series, 135: 289–298, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/135/m135p289.pdf.

Lignell:1993:FPS

[LHK^{+}93] R. Lignell, A.-S. Heiskanen, H. Kuosa, K. Gundersen, P. Kuuppo-Leinikki, R. Pajuniemi, and A. Uitto. Fate of a phytoplankton spring bloom: sedimentation and carbon flow in the planktonic food web in the northern Baltic. Marine Ecology Progress Series, 94:239–252, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
94/m094p239.pdf.

Lopez:1993:CAG

[LHL93] M. D. G. Lopez, M. E. Huntley, and J. T. Lovette. Calanoides acutus in Gerlache Strait, Antarctica. I. Distribution of late copepodite stages and reproduction during spring. Marine Ecology Progress Series, 100:153–165, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
100/m100p153.pdf.

LeGall:1997:IBC

[LHL97] S. Le Gall, M. B. Hassen, and P. Le Gall. Ingestion of a bacterivorous ciliate by the oyster Crassostrea gigas: protozoa as a trophic link between picoplankton and benthic suspensionfeeders. Marine Ecology Progress Series, 152:301–306, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
152/m152p301.pdf.

Lam-Hoai:1997:TRN

[LHRL97] T. Lam-Hoai, C. Rougier, and G. Lasserre. Tintinnids and rotifers in a northern Mediterranean coastal lagoon. Structural diversity and function through biomass estimations. Marine Ecology Progress Series, 152:13–25, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/152/m152p013.pdf.

Li:1995:CUC

[Li95] W. K. W. Li. Composition of ultraphytoplankton in the central North Atlantic. Marine Ecology Progress Series, 122: 1–8, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/122/m122p001.pdf.

Lignell:1990:EOC

[Lig90] R. Lignell. Excretion of organic carbon by phytoplankton: its relation to algal biomass, primary productivity and bacterial secondary productivity in the Baltic Sea. Marine Ecology Progress Series, 68:85–99, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/68/m068p085.pdf.

Limburg:1995:OST

[Lim95] K. E. Limburg. Otolith strontium traces environmental history of subyearling American shad Alosa sapidissima. Marine Ecology Progress Series, 119:25–35, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/119/m119p025.pdf.

Lipschultz:1995:NSU

[Lip95] F. Lipschultz. Nitrogen-specific uptake rates of marine phytoplankton isolated from natural populations of particles by flow cytometry. Marine Ecology Progress Series, 123:245– 258, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/123/m123p245.pdf.

LeFevre-Lehoerff:1995:HRP

[LIPF95] G. Le Fevre-Lehoerff, F. Ibanez, P. Poniz, and J. M. Fromentin. Hydroclimatic relationships with planktonic time series from 1975 to 1992 in the North Sea off Gravelines, France. Marine Ecology Progress Series, 129:269–281, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
129/m129p269.pdf.

Livingston:1993:IPG

[Liv93] P. A. Livingston. Importance of predation by groundfish, marine mammals and birds on walleye pollock Theragra chaicogramma and Pacific herring Clupea pallasi in the eastern Bering Sea. Marine Ecology Progress Series, 102:205– 215, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/102/m102p205.pdf.

Laegdsgaard:1995:MHN

[LJ95] P. Laegdsgaard and C. R. Johnson. Mangrove habitats as nurseries: unique assemblages of juvenile fish in subtropical mangroves in eastern Australia. Marine Ecology Progress Series, 126:67–81, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/126/m126p067.pdf.

Latz:1996:ERT

[LJ96] M. I. Latz and H. J. Jeong. Effect of red tide dinoflagellate diet and cannibalism on the bioluminescence of the heterotrophic dinoflagellates Protoperidinium spp. Marine Ecology Progress Series, 132:275–285, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/132/m132p275.pdf.

LePape:1999:PBT

[LJM99] O. Le Pape, F. Jean, and A. Ménesguen. Pelagic and benthic trophic chain coupling in a semi-enclosed coastal system, the Bay of Brest (France): a modelling approach. Marine Ecology Progress Series, 189:135–147, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/189/m189p135.pdf.

Lomstein:1990:INN

[LJS90] E. Lomstein, M. H. Jensen, and J. Sørensen. Intracellular NH_{4}^{+} and NO_{3}-pools associated with deposited phytoplankton in a marine sediment (Aarhus Bight, Denmark). Marine Ecology Progress Series, 61:97–105, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/61/m061p097.pdf.

Loo:1996:PSF

[LJSK96] L. O. Loo, P. R. Jonsson, M. Sköld, and Ö. Karlsson. Passive suspension feeding in Amphiura filiformis (Echinodermata: Ophiuroidea): feeding behaviour in flume flow and potential feeding rate of field populations. Marine Ecology Progress Series, 139:143–155, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/139/m139p143.pdf.

Lafferty:1993:MMA

[LK93] K. D. Lafferty and A. M. Kuris. Mass mortality of abalone Haliotis cracherodii or the California Channel Islands: tests of epidemiological hypothesis. Marine Ecology Progress Series, 96:239–248, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/96/m096p239.pdf.

Lee:1994:EBS

[LK94] S. Y. Lee and R. T. Kneib. Effects of biogenic structure on prey consumption by the xanthid crabs Eurytium limosum and Panopeus herbstii in a salt marsh. Marine Ecology Progress Series, 104:39–47, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/104/m104p039.pdf.

Letelier:1996:RTS

[LK96] R. M. Letelier and D. M. Karl. Role of Trichodesmium spp. in the productivity of the subtropical North Pacific Ocean. Marine Ecology Progress Series, 133:263–273, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
133/m133p263.pdf.

Labropoulou:1999:PRU

[LK99a] M. Labropoulou and I. Kostikas. Patterns of resource use in deep-water decapods. Marine Ecology Progress Series, 184:171–182, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/184/m184p171.pdf.

Lydersen:1999:BEI

[LK99b] C. Lydersen and K. M. Kovacs. Behaviour and energetics of ice-breeding, North Atlantic phocid seals during the lactation period. Marine Ecology Progress Series, 187:265–281, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
187/m187p265.pdf.

Landry:1995:RDT

[LKC95] M. R. Landry, J. Kirshtein, and J. Constantinou. A refined dilution technique for measuring the community grazing impact of microzooplankton, with experimental tests in the central equatorial Pacific. Marine Ecology Progress Series, 120:53–63, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/120/m120p053.pdf.

Lasker:1998:PSS

[LKC98] H. R. Lasker, K. Kim, and M. A. Coffroth. Production, settlement, and survival of plexaurid gorgonian recruits. Marine Ecology Progress Series, 162:111–123, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
162/m162p111.pdf.

Lucas:1999:PGPa

[LKC^{+}99] L. V. Lucas, J. R. Koseff, J. E. Cloern, S. G. Monismith, and J. K. Thompson. Processes governing phytoplankton blooms in estuaries. I: The local production-loss balance. Marine Ecology Progress Series, 187:1–15, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/187/m187p001.pdf.

Lee:1995:IRN

[LKF95] S. Lee, Y.-C. Kang, and J. A. Fuhrman. Imperfect retention of natural bacterioplankton cells by glass fiber filters. Marine Ecology Progress Series, 119:285–290, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
119/m119p285.pdf.

Lignell:1992:FCP

[LKK92] R. Lignell, S. Kaitala, and H. Kuosa. Factors controlling phyto- and bacterioplankton in late spring on a salinity gradient in the northern Baltic. Marine Ecology Progress Series, 84:121–131, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/84/m084p121.pdf.

Lucas:1999:PGPb

[LKM^{+}99] L. V. Lucas, J. R. Koseff, S. G. Monismith, J. E. Cloern, and J. K. Thompson. Processes governing phytoplankton blooms in estuaries. II: The role of horizontal transport. Marine Ecology Progress Series, 187:17–30, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/187/m187p017.pdf.

Lohse:1996:DRM

[LKvRH96] L. Lohse, H. T. Kloosterhuis, W. van Raaphorst, and W. Helder. Denitrification rates as measured by the isotope pairing method and by the acetylene inhibition technique in continental shelf sediments of the North Sea. Marine Ecology Progress Series, 132:169–179, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/132/m132p169.pdf.

Lignell:1992:ENE

[LL92] R. Lignell and K. Lindqvist. Effect of nutrient enrichment and temperature or intracellular partitioning of ^{14}CO_{2} in a summer phytoplankton community in the northern Baltic. Marine Ecology Progress Series, 86:273–281, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
86/m086p273.pdf.

Lawrence:1994:CAA

[LL94] J. M. Lawrence and A. Larrain. The cost of arm autotomy in the starfish Stichaster striatus. Marine Ecology Progress Series, 109:311–313, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/109/m109p311.pdf.

Lesser:1996:ASE

[LL96] M. P. Lesser and S. Lewis. Action spectrum for the effects of UV radiation on photosynthesis in the hermatypic coral Pocillopora damicornis. Marine Ecology Progress Series, 134:171–177, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/134/m134p171.pdf.

Lindqvist:1997:IPC

[LL97] K. Lindqvist and R. Lignell. Intracellular partitioning of 14 CO_{2} in phytoplankton during a growth season in the northern Baltic. Marine Ecology Progress Series, 152:41–50, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
152/m152p041.pdf.

Li:1998:CDD

[LL98] M. S. Li and S. Y. Lee. Carbon dynamics of Deep Bay, eastern Pearl River Estuary, China. I: A mass balance budget and implications for shorebird conservation. Marine Ecology Progress Series, 172:73–87, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/172/m172p073.pdf.

Lambert:1992:CSN

[LLB92] W. J. Lambert, P. S. Levin, and J. Berman. Changes in the structure of a New England (USA) kelp bed the effects of an introduced species? Marine Ecology Progress Series, 88:303–307, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/88/m088p303.pdf.

Leitao:1995:OSI

[LLB^{+}95] J. M. Leitão, B. Lorenz, N. Bachinski, C. Wilhelm, W. E. G. Müller, and H. C. Schröder. Osmotic-stress-induced synthesis and degradation of inorganic polyphosphates in the alga Phaeodactylum tricornutum. Marine Ecology Progress Series, 121:279–288, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/121/m121p279.pdf.

Landry:1994:MGSb

[LLP94] M. R. Landry, C. J. Lorenzen, and W. K. Peterson. Mesozooplankton grazing in the Southern California Bight. II. Grazing impact and particulate flux. Marine Ecology Progress Series, 115:73–85, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/115/m115p073.pdf.

Levin:1991:VHE

[LM91a] P. S. Levin and A. C. Mathieson. Variation in a host-epiphyte relationship along a wave exposure gradient. Marine Ecology Progress Series, 77:271–278, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/77/m077p271.pdf.

Levinton:1991:UPP

[LM91b] J. S. Levinton and M. McCartney. Use of photosynthetic pigments in sediments as a tracer for sources and fates of macrophyte organic matter. Marine Ecology Progress Series, 78:87–96, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/78/m078p087.pdf.

Lee:1995:RZC

[LM95] J. G. Lee and F. M. M. Morel. Replacement of zinc by cadmium in marine phytoplankton. Marine Ecology Progress Series, 127:305–309, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/127/m127p305.pdf.

Llewellyn:1997:UAC

[LM97] C. A. Llewellyn and R. F. C. Mantoura. A UV absorbing compound in HPLC pigment chromatograms obtained from Icelandic Basin phytoplankton. Marine Ecology Progress Series, 158:283–287, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/158/m158p283.pdf.

Labropoulou:1998:EHS

[LM98a] M. Labropoulou and A. Machias. Effect of habitat selection on the dietary patterns of two triglid species. Marine Ecology Progress Series, 173:275–288, 1998. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/173/m173p275.pdf.

Legendre:1998:FBC

[LM98b] L. Legendre and J. Michaud. Flux of biogenic carbon in oceans: size-dependent regulation by pelagic food webs. Marine Ecology Progress Series, 164:1–11, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
164/m164p001.pdf.

Lowe:1992:CII

[LME92] D. M. Lowe, M. N. Moore, and B. M. Evans. Contaminant impact on interactions of molecular probes with lysosomes in living hepatocytes from dab Limanda limanda. Marine Ecology Progress Series, 91:135–140, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/91/m091p135.pdf.

Lepesteur:1993:CSD

[LMF93] M. Lepesteur, J. M. Martin, and A. Fleury. A comparative study of different preservation methods for phytoplankton cell analysis by flow cytometry. Marine Ecology Progress Series, 93:55–63, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/93/m093p055.pdf.

Lee:1993:UMR

[LMK93] S. Lee, C. Malone, and P. F. Kemp. Use of multiple ^{16}S rRNA-targeted fluorescent probes to increase signal strength and measure cellular RNA from natural planktonic bacteria. Marine Ecology Progress Series, 101:193–201, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
101/m101p193.pdf.

Lawson:1998:IPS

[LMM98] J. W. Lawson, A. M. Magalhães, and E. H. Miller. Important prey species of marine vertebrate predators in the northwest Atlantic: proximate composition and energy density. Marine Ecology Progress Series, 164:13–20, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/164/m164p013.pdf.

Latruffe:1999:VSS

[LMO99] C. Latruffe, P. K. McGregor, and R. F. Oliveira. Visual signalling and sexual selection in male fiddler crabs Uca tangeri. Marine Ecology Progress Series, 189:233–240, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
189/m189p233.pdf.

Lohse:1993:NCN

[LMS^{+}93] L. Lohse, J. F. P. Malschaert, C. P. Slomp, W. Helder, and W. van Raaphorst. Nitrogen cycling in North Sea sediments: interaction of denitrification and nitrification in offshore and coastal areas. Marine Ecology Progress Series, 101:283–296, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/101/m101p283.pdf.

Laursen:1996:LPM

[LMT96] A. K. Laursen, L. M. Mayer, and D. W. Townsend. Lability of proteinaceous material in estuarine seston and subcellular fractions of phytoplankton. Marine Ecology Progress Series, 136:227–234, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/136/m136p227.pdf.

Lyver:1999:CSS

[LMT99] P. O. B. Lyver, H. Moller, and C. Thompson. Changes in sooty shearwater Puffinus griseus chick production and harvest precede ENSO events. Marine Ecology Progress Series, 188:237–248, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/188/m188p237.pdf.

Lochmann:1996:EGC

[LMTF96] S. E. Lochmann, G. L. Maillet, C. T. Taggart, and K. T. Frank. Effect of gut contents and lipid degradation on condition measures in larval fish. Marine Ecology Progress Series, 134:27–35, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/134/m134p027.pdf.

Leynaert:1993:SCA

[LNQT93] A. Leynaert, D. M. Nelson, B. Quéguiner, and P. Treguer. The silica cycle in the Antarctic Ocean: is the Weddell Sea atypical? Marine Ecology Progress Series, 96:1–15, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
96/m096p001.pdf.

Latyshev:1991:FAR

[LNSL91] N. A. Latyshev, N. V. Naumenko, V. I. Svetashev, and Y. Ya. Latypov. Fatty acids of reef-building corals. Marine Ecology Progress Series, 76:295–301, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/76/m076p295.pdf.

Lopez:1991:MMD

[Lop91] M. D. G. Lopez. Molting and mortality depend on age and stage in naupliar Calanus pacificus: triplication for development time of field cohorts. Marine Ecology Progress Series, 75:79–89, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/75/m075p079.pdf.

Lenihan:1990:CHB

[LOS90] H. S. Lenihan, J. S. Oliver, and M. A. Stephenson. Changes in hard bottom communities related to boat mooring and tributyltin in San Diego Bay a natural experiment. Marine Ecology Progress Series, 60:147–159, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/60/m060p147.pdf.

Laprise:1995:FIS

[LP95] R. Laprise and P. Pepin. Factors influencing the spatiotemporal occurrence of fish eggs and larvae in a northern, physically dynamic coastal environment. Marine Ecology Progress Series, 122:73–92, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/122/m122p073.pdf.

Lores:1999:BTT

[LP99] E. M. Lores and J. R. Pennock. Bioavailability and trophic transfer of humic-bound copper from bacteria to zooplankton. Marine Ecology Progress Series, 187:67–75, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
187/m187p067.pdf.

Landry:1994:MGSa

[LPF94] M. R. Landry, W. K. Peterson, and V. L. Fagerness. Mesozooplankton grazing in the Southern California Bight. I. Population abundances and gut pigment contents. Marine Ecology Progress Series, 115:55–71, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/115/m115p055.pdf.

Laabir:1995:RRC

[LPI^{+}95] M. Laabir, S. A. Poulet, A. Ianora, A. Miralto, and A. Cueff. Reproductive response of Calanus helgolandicus. II. In situ inhibition of embryonic development. Marine Ecology Progress Series, 129:97–105, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/129/m129p097.pdf.

Long:1995:DBC

[LPW95] B. G. Long, I. R. Poiner, and T. J. Wassenberg. Distribution, biomass and community structure of megabenthos of the Gulf of Carpentaria, Australia. Marine Ecology Progress Series, 129:127–139, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/129/m129p127.pdf.

Lau:1997:PRC

[LQ97] S. C. K. Lau and P. Y. Qian. Phlorotannins and related compounds as larval settlement inhibitors of the tube-building polychaete Hydroides elegans. Marine Ecology Progress Series, 159:219–227, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/159/m159p219.pdf.

Legendre:1996:FWM

[LR96] L. Legendre and F. Rassoulzadegan. Food-web mediated export of biogenic carbon in oceans: hydrodynamic control. Marine Ecology Progress Series, 145:179–193, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
145/m145p179.pdf.

Lange:1994:TDP

[LRV94] C. B. Lange, F. M. H. Reid, and M. Vernet. Temporal distribution of the potentially toxic diatom Pseudonitzschia australis at a coastal site in Southern California. Marine Ecology Progress Series, 104:309–312, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/104/m104p309.pdf.

Lizotte:1991:PIR

[LS91] M. P. Lizotte and C. W. Sullivan. Photosynthesis-irradiance relationships in microalgae associated with Antarctic pack ice: evidence for in situ activity. Marine Ecology Progress Series, 71:175–184, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/71/m071p175.pdf.

Lee:1992:EIC

[LS92] S. Y. Lee and R. Seed. Ecological implications of cheliped size in crabs: Some data from Carcinus maenas and Liocarcinus holsatus. Marine Ecology Progress Series, 84:151–160, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
84/m084p151.pdf.

Lin:1994:MDS

[LS94] S. Lin and D. J. Steichen, Jr. A method for determining the stability of lysosomal membranes in the digestive cells of Mytilus edulis. Marine Ecology Progress Series, 115:237– 241, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/115/m115p237.pdf.

Larned:1996:NLG

[LS96a] S. T. Larned and J. Stimson. Nitrogen-limited growth in the coral reef chlorophyte Dictyosphaeria cavernosa, and the effect of exposure to sediment-derived nitrogen on growth. Marine Ecology Progress Series, 145:95–108, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
145/m145p095.pdf.

Larson:1996:CMS

[LS96b] E. T. Larson and A. L. Shanks. Consumption of marine snow by two species of juvenile mullet and its contribution to their growth. Marine Ecology Progress Series, 130:19–28, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
130/m130p019.pdf.

Lo:1995:PGN

[LSB95] N. C. H. Lo, P. E. Smith, and J. L. Butler. Population growth of northern anchovy and Pacific sardine using stagespecific matrix models. Marine Ecology Progress Series, 127: 15–26, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/127/m127p015.pdf.

Lekve:1999:STP

[LSG^{+}99] K. Lekve, N. C. Stenseth, J. Gjøsæter, J. M. Fromentin, and J. S. Gray. Spatio-temporal patterns in diversity of a fish assemblage along the Norwegian Skagerrak coast. Marine Ecology Progress Series, 178:17–27, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/178/m178p017.pdf.

Leichter:1998:BIW

[LSGW98] J. J. Leichter, G. Shellenbarger, S. J. Genovese, and S. R. Wing. Breaking internal waves on a Florida (USA) coral reef: a plankton pump at work? Marine Ecology Progress Series, 166:83–97, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/166/m166p083.pdf.

Lutz:1996:ACP

[LSH96] V. A. Lutz, S. Sathyendranath, and E. J. H. Head. Absorption coefficient of phytoplankton: regional variations in the North Atlantic. Marine Ecology Progress Series, 135:197–213, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
135/m135p197.pdf.

Lizon:1998:PPP

[LSL98] F. Lizon, L. Seuront, and Y. Lagadeuc. Photoadaptation and primary production study in tidally mixed coastal waters using a Lagrangian model. Marine Ecology Progress Series, 169:43–54, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/169/m169p043.pdf.

Lapointe:1992:CPA

[LSM92] V. Lapointe and B. Sainte-Marie. Currents, predators, and the aggregation of the gastropod Buccinum undatum around bait. Marine Ecology Progress Series, 85:245–257, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
85/m085p245.pdf.

Lovell:1999:AMS

[LSP99] C. R. Lovell, C. C. Steward, and T. Phillips. Activity of marine sediment bacterial communities exposed to 4-bromophenol, a polychaete secondary metabolite. Marine Ecology Progress Series, 179:241–246, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
179/m179p241.pdf.

Lancellotti:1993:DPC

[LT93] D. A. Lancellotti and R. G. Trucco. Distribution patterns and coexistence of six species of the amphipod genus Hyale. Marine Ecology Progress Series, 93:131–141, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
93/m093p131.pdf.

Lagardere:1997:AEC

[LT97] F. Lagardère and H. Troadec. Age estimation in common sole Solea solea larvae: validation of daily increments and evaluation of a pattern recognition technique. Marine Ecology Progress Series, 155:223–237, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/155/m155p223.pdf.

LopezGappa:1990:ISP

[LTM90] J. J. López Gappa, A. Tablado, and N. H. Magaldi. Influence of sewage pollution on a rocky intertidal community dominated by the mytilid Brachidontes rodriguezi. Marine Ecology Progress Series, 63:163–175, 1990. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/63/m063p163.pdf.

Lopez:1998:LAR

[LTM^{+}98] S. López, X. Turon, E. Montero, C. Palacín, C. M. Duarte, and I. Tarjuelo. Larval abundance, recruitment and early mortality in Paracentrotus lividus (Echinoidea). Interannual variability and plankton-benthos coupling. Marine Ecology Progress Series, 172:239–251, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/172/m172p239.pdf.

Levin:1996:SMC

[LTT96] L. A. Levin, D. Talley, and G. Thayer. Succession of macrobenthos in a created salt marsh. Marine Ecology Progress Series, 141:67–82, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/141/m141p067.pdf.

Lopez-Urrutia:1999:GTD

[LUA99] Á. López-Urrutia and J. L. Acuña. Gut throughput dynamics in the appendicularian Oikopleura dioica. Marine Ecology Progress Series, 191:195–205, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/191/m191p195.pdf. See

erratum [Ano00].

Laroche:1999:DSE

[LVL^{+}99] D. Laroche, A. F. Vézina, M. Levasseur, M. Gosselin, J. Stefels, M. D. Keller, P. A. Matrai, and R. L. J. Kwint. DMSP synthesis and exudation in phytoplankton: a modeling approach. Marine Ecology Progress Series, 180:37–49, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
180/m180p037.pdf.

Li:1998:APA

[LVP98] H. Li, M. J. W. Veldhuis, and A. F. Post. Alkaline phosphatase activities among planktonic communities in the northern Red Sea. Marine Ecology Progress Series, 173:107– 115, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/173/m173p107.pdf.

Lombardi:1991:IPS

[LW91] A. T. Lombardi and P. J. Wangersky. Influence of phosphorus and silicon on lipia class production by the marine diatom Chaetoceros gracilis grown in turbidostat cage cultures. Marine Ecology Progress Series, 77:39–47, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
77/m077p039.pdf.

Lindley:1994:RPA

[LW94] J. A. Lindley and R. Williams. Relating plankton assemblages to environmental variables using instruments towed by ships-of-opportunity. Marine Ecology Progress Series, 107: 245–262, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/107/m107p245.pdf.

Lindsay:1995:TLI

[LW95] S. M. Lindsay and S. A. Woodin. Tissue loss induces switching of feeding mode in spionid polychaetes. Marine Ecology Progress Series, 125:159–169, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/125/m125p159.pdf.

Loneragan:1995:EES

[LWK^{+}95] N. R. Loneragan, Y.-G. Wang, R. A. Kenyon, D. J. Staples, D. J. Vance, and D. S. Heales. Estimating the efficiency of a small beam trawl for sampling tiger prawns Penaeus esculentus and P. semisulcatus in seagrass by removal experiments. Marine Ecology Progress Series, 118:139–148, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
118/m118p139.pdf.

Lee:1998:ULK

[LWL98] B. G. Lee, W. G. Wallace, and S. N. Luoma. Uptake and loss kinetics of Cd, Cr and Zn in the bivalves Potamocorbula amurensis and Macoma balthica: effects of size and salinity. Marine Ecology Progress Series, 175:177–189, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
175/m175p177.pdf.

Leising:1997:SMW

[LY97] A. W. Leising and J. Yen. Spacing mechanisms within lightinduced copepod swarms. Marine Ecology Progress Series, 155:127–135, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/155/m155p127.pdf.

Ahn:1993:EGC

[lYALM93] l. Y. Ahn, G. Lopez, and R. Malouf. Effects of the gem clam Gemma gemma on early post-settlement emigration, growth and survival of the hard clam Mercenaria mercenaria. Marine Ecology Progress Series, 99:61–70, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/99/m099p061.pdf.

Ahn:1993:ELS

[lYAML93] l. Y. Ahn, R. Malouf, and G. Lopez. Enhanced larval settlement of the hard clam Mercenaria mercenaria by the gem clam Gemma gemma. Marine Ecology Progress Series, 99: 51–59, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/99/m099p051.pdf.

Li:1993:UEM

[LZYW93] W. K. W. Li, T. Zohary, Y. Z. Yacobi, and A. M. Wood. Ultraphytoplankton in the eastern Mediterranean Sea: towards deriving phytoplankton biomass from flow cytometric measurements of abundance, fluorescence and light scatter. Marine Ecology Progress Series, 102:79–87, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/102/m102p079.pdf.

Martinez:1993:PAA

[MA93] J. Martinez and F. Azam. Penplasmic aminopeptidase and alkaline phosphatase activities in a marine bacterium- implications for substrate processing in the sea. Marine Ecology Progress Series, 92:89–97, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/92/m092p089.pdf.

Mura:1996:GRD

[MA96] M. P. Mura and S. Agustí. Growth rates of diatoms from coastal Antarctic waters estimated by in situ dialysis incubation. Marine Ecology Progress Series, 144:237–245, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
144/m144p237.pdf.

Mariani:1999:MCF

[MA99a] S. Mariani and T. Alcoverro. A multiple-choice feedingpreference experiment utilising seagrasses with a natural population of herbivorous fishes. Marine Ecology Progress Series, 189:295–299, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/189/m189p295.pdf.

Marubini:1999:ELP

[MA99b] F. Marubini and M. J. Atkinson. Effects of lowered pH and elevated nitrate on coral calcification. Marine Ecology Progress Series, 188:117–121, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/188/m188p117.pdf.

Miyashita:1999:AMZ

[MA99c] K. Miyashita and I. Aoki. Acoustic measurements of zooplankton using a dual frequency echo sounder. Marine Ecology Progress Series, 180:105–109, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/180/m180p105.pdf.

Mouneyrac:1998:ENF

[MAAT98] C. Mouneyrac, J. C. Amiard, and C. Amiard-Triquet. Effects of natural factors (salinity and body weight) on cadmium, copper, zinc and metallothionein-like protein levels in resident populations of oysters Crassostrea gigas from a polluted estuary. Marine Ecology Progress Series, 162:125–135, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
162/m162p125.pdf.

MacDiarmid:1991:SCD

[Mac91] A. B. MacDiarmid. Seasonal changes in depth distribution, sex ratio and size frequency of spiny lobster Jasus edwardsii on a coastal reef in northern New Zealand. Marine Ecology Progress Series, 70:129–141, 1991. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/70/m070p129.pdf.

Malatesta:1992:ATD

[MAC92] R. J. Malatesta, P. J. Auster, and B. P. Carlin. Analysis of transect data for microhaoitat correlations and faunal patchiness. Marine Ecology Progress Series, 87:189–195, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
87/m087p189.pdf.

Macpherson:1994:SUM

[Mac94] E. Macpherson. Substrate utilisation in a Mediterranean littoral fish community. Marine Ecology Progress Series, 114: 211–218, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/114/m114p211.pdf.

Mayzaud:1998:CLC

[MACR98] P. Mayzaud, E. Albessard, and J. Cuzin-Roudy. Changes in lipid composition of the Antarctic krill Euphausia superba in the Indian sector of the Antarctic Ocean: influence of geographical location, sexual maturity stage and distribution among organs. Marine Ecology Progress Series, 173:149– 162, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/173/m173p149.pdf.

Mura:1996:LCP

[MAdG^{+}96] M. P. Mura, S. Agustí, P. A. del Giorgio, J. M. Gasol, D. Vaqué, and C. M. Duarte. Loss-controlled phytoplankton production in nutrient-poor littoral waters of the NW Mediterranean: in situ experimental evidence. Marine Ecology Progress Series, 130:213–219, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/130/m130p213.pdf.

Moreno:1998:SMC

[MADM98] C. A. Moreno, G. Asencio, W. E. Duarte, and V. Marín. Settlement of the muricid Concholepas concholepas and its relationship with El Niño and coastal upwellings in southern Chile. Marine Ecology Progress Series, 167:171–175, 1998.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
167/m167p171.pdf.

Mees:1994:LHG

[MAH94] J. Mees, Z. Abdulkerim, and O. Hamerlynck. Life history, growth and production of Neomysis integer in the Westerschelde estuary (SW Netherlands). Marine Ecology Progress Series, 109:43–57, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/109/m109p043.pdf.

Maloney:1996:IOE

[Mal96] J. Maloney. Influence of organic enrichment on the partitioning and bioavailability of cadmium in a microcosm study. Marine Ecology Progress Series, 144:147–161, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
144/m144p147.pdf.

Maldonado:1998:DCS

[Mal98] M. Maldonado. Do chimeric sponges have improved chances of survival? Marine Ecology Progress Series, 164:301–306, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
164/m164p301.pdf.

Maltagliati:1999:GDN

[Mal99] F. Maltagliati. Genetic divergence in natural populations of the Mediterranean brackish-water killifish Aphanius fasciatus. Marine Ecology Progress Series, 179:155–162, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
179/m179p155.pdf.

Marsden:1991:RPL

[Mar91] J. R. Marsden. Responses of planktonic larvae of the serpulid polychaete Spirobranchus polycerus var. augeneri to an alga, adult tubes and conspecific larvae. Marine Ecology Progress Series, 71:245–251, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/71/m071p245.pdf.

Maravelias:1997:TAG

[Mar97] C. D. Maravelias. Trends in abundance and geographic distribution of North Sea herring in relation to environmental factors. Marine Ecology Progress Series, 159:151–164, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
159/m159p151.pdf.

Marcotte:1999:TAE

[Mar99a] Brian M. Marcotte. Turbidity, arthropods and the evolution of perception: toward a new paradigm of marine phanerozoic diversity. Marine Ecology Progress Series, 191:267–288, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
191/m191p267.pdf.

Mari:1999:CCC

[Mar99b] X. Mari. Carbon content and C:N ratio of transparent exopolymeric particles (TEP) produced by bubbling exudates of diatoms. Marine Ecology Progress Series, 183:59–71, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
183/m183p059.pdf.

Matsakis:1992:AER

[Mat92] S. Matsakis. Ammonia excretion rate of Clytia spp. hydromedusae (Cnidaria, Thecata): effects of individual dry weight, temperature and food availability. Marine Ecology Progress Series, 87:55–63, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/87/m087p055.pdf.

Mauchline:1992:RBS

[Mau92] J. Mauchline. Restriction of body size spectra within species of deep-sea plankton. Marine Ecology Progress Series, 90: 1–8, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/90/m090p001.pdf.

Mazel:1995:SMF

[Maz95] C. H. Mazel. Spectral measurements of fluorescence emission in Caribbean cnidarians. Marine Ecology Progress Series, 120:185–191, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/120/m120p185.pdf.

Morgan:1991:SDA

[MB91] M. J. Morgan and W. B. Brodie. Seasonal distribution of American plaice on the northern Grand Banks. Marine Ecology Progress Series, 75:101–107, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/75/m075p101.pdf.

Meesters:1993:ECB

[MB93] E. H. Meesters and R. P. M. Bak. Effects of coral bleaching on tissue regeneration potential and colony survival. Marine Ecology Progress Series, 96:189–198, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
96/m096p189.pdf.

Maranger:1995:VAA

[MB95a] R. Maranger and D. F. Bird. Viral abundance in aquatic systems: a comparison between marine and fresh waters. Marine Ecology Progress Series, 121:217–226, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
121/m121p217.pdf.

Meesters:1995:ARD

[MB95b] E. H. Meesters and R. P. M. Bak. Age-related deterioration of a physiological function in the branching coral Acropora palmata. Marine Ecology Progress Series, 121:203–209, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
121/m121p203.pdf.

McClintock:1997:PCD

[MB97a] J. B. McClintock and B. J. Baker. Palatability and chemical defense of eggs, embryos and larvae of shallow-water Antarctic marine invertebrates. Marine Ecology Progress Series, 154:121–131, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/154/m154p121.pdf.

Michotey:1997:EAB

[MB97b] V. Michotey and P. Bonin. Evidence for anaerobic bacterial processes in the water column: denitrification and dissimilatory nitrate ammonification in the northwestern Mediterranean Sea. Marine Ecology Progress Series, 160:47–56, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
160/m160p047.pdf.

Mari:1998:SSS

[MB98] X. Mari and A. Burd. Seasonal size spectra of transparent exopolymeric particles (TEP) in a coastal sea and comparison with those predicted using coagulation theory. Marine Ecology Progress Series, 163:63–76, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/163/m163p063.pdf.

Miron:1995:ULS

[MBB95] G. Miron, B. Boudreau, and E. Bourget. Use of larval supply in benthic ecology: testing correlations between larval supply and larval settlement. Marine Ecology Progress Series, 124:301–305, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/124/m124p301.pdf.

Miron:1999:IBD

[MBB99a] G. Miron, B. Boudrea, and E. Bourget. Intertidal barnacle distribution: a case study using multiple working hypotheses. Marine Ecology Progress Series, 189:205–219, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
189/m189p205.pdf.

Moran:1999:ETE

[MBB99b] K. L. Moran, K. A. Bjorndal, and A. B. Bolten. Effects of the thermal environment on the temporal pattern of emergence of hatchling loggerhead turtles Caretta caretta. Marine Ecology Progress Series, 189:251–261, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/189/m189p251.pdf.

Macpherson:1997:MJF

[MBF^{+}97] E. Macpherson, F. Biagi, P. Francour, A. García-Rubies, J. Harmelin, M. Harmelin-Vivien, J. Y. Jouvenel, S. Planes, L. Vigliola, and L. Tunesi. Mortality of juvenile fishes of the genus Diplodus in protected and unprotected areas in the western Mediterranean Sea. Marine Ecology Progress Series, 160:135–147, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/160/m160p135.pdf.

Maranger:1994:VBD

[MBJ94] R. Maranger, D. R. Bird, and S. K. Juniper. Viral and bacterial dynamics in Arctic sea ice during the spring algal bloom near Resolute, N.W.T., Canada. Marine Ecology Progress Series, 111:121–127, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/111/m111p121.pdf.

Middelboe:1995:BUD

[MBK95] M. Middelboe, N. H. Borch, and D. L. Kirchman. Bacterial utilization of dissolved free amino acids, dissolved combined amino acids and ammonium in the Delaware Bay estuary: effects of carbon and nitrogen limitation. Marine Ecology Progress Series, 128:109–120, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/128/m128p109.pdf.

Moloney:1994:DST

[MBL94] C. L. Moloney, L. W. Botsford, and J. L. Largier. Development, survival and timing of metamorphosis of planktonic larvae in a variable environment: the Dungeness crab as an example. Marine Ecology Progress Series, 113:61–79, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
113/m113p061.pdf.

Muller:1998:ACZ

[MBL^{+}98] W. E. G. Müller, R. Batel, M. Lacorn, H. Steinhart, T. Simat, S. Lauenroth, H. Hassanein, and H. C. Schröder. Accumulation of cadmium and zinc in the marine sponge Suberites domuncula and its potential consequences on single-strand breaks and on expression of heat-shock protein: a natural field study. Marine Ecology Progress Series, 167:127–135, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
167/m167p127.pdf.

MacMillin:1992:CBD

[MBM92] K. M. MacMillin, L. K. Blum, and A. L. Mills. Comparison of bacterial dynamics in tidal creeks of the lower Delmarva Peninsula, Virginia, USA. Marine Ecology Progress Series, 86:111–121, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/86/m086p111.pdf.

Milton:1994:DPS

[MBR94] D. A. Milton, S. J. M. Blaber, and N. J. F. Rawlinson. Diet, prey selection and their energetic relationship to reproduction in the tropical herring Herklotsichthys quadrimaculatus in Kiribati, Central Pacific. Marine Ecology Progress Series, 103:239–250, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/103/m103p239.pdf.

McCafferty:1998:FRA

[MBWT98] D. J. McCafferty, I. L. Boyd, T. R. Walker, and R. I. Taylor. Foraging responses of Antarctic fur seals to changes in the marine environment. Marine Ecology Progress Series, 166:285–299, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/166/m166p285.pdf.

Maillet:1991:SRV

[MC91a] Gary L. Maillet and D. M. Checkley, Jr. Storm-related variation in the growth rate of otoliths of larval Atlantic menhaden Brevoortia tyrannus: a time series analysis of biological and physical variables and implications for larva growth and mortality. Marine Ecology Progress Series, 79:1–16, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
79/m079p001.pdf.

Martel:1991:FRB

[MC91b] A. Martel and F.-S. Chia. Foot-raising behaviour and active participation during the initial phase of post-metamorphic drifting in the gastropod Lacuna spp. Marine Ecology Progress Series, 72:247–254, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/72/m072p247.pdf.

McBride:1991:RYY

[MC91c] R. S. McBride and D. O. Conover. Recruitment of young-of-the-year bluefish Pomatomus saltatrix to the New York Bight: variation in abundance and growth of spring- and summerspawned cohorts. Marine Ecology Progress Series, 78:205– 216, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/78/m078p205.pdf.

Mordy:1991:EFT

[MC91d] C. W. Mordy and D. J. Carlson. An evaluation of fluorescence technigues for measuring DNA and RNA in marine microorganisms. Marine Ecology Progress Series, 73:283–293, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
73/m073p283.pdf.

Martinez:1993:SSE

[MC93] Enrique Martinez and Juan A. Correa. Sorus-specific epiphytism affecting the kelps Lessonia nigrescens and L. trabeculata (Phaeophyta). Marine Ecology Progress Series, 96: 83–92, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/96/m096p083.pdf.

McFadzen:1994:TCS

[MC94a] I. R. B. McFadzen and J. J. Cleary. Toxicity and chemistry of the sea-surface microlayer in the North Sea using a cryopreserved larval bioassay. Marine Ecology Progress Series, 103:103–109, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/103/m103p103.pdf.

Mistri:1994:GSP

[MC94b] M. Mistri and V. U. Ceccherelli. Growth and secondary production of the Mediterranean gorgonian Paramuricea clavata. Marine Ecology Progress Series, 103:291–296, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
103/m103p291.pdf.

MacIntyre:1995:FSV

[MC95] H. L. MacIntyre and J. J. Cullen. Fine-scale vertical resolution of chlorophyll and photosynthetic parameters in shallowwater benthos. Marine Ecology Progress Series, 122:227– 237, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/122/m122p227.pdf.

MacIntyre:1996:PPS

[MC96a] H. L. MacIntyre and J. J. Cullen. Primary production by suspended and benthic microalgae in a turbid estuary: timescales of variability in San Antonio Bay, Texas. Marine Ecology Progress Series, 145:245–268, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/145/m145p245.pdf.

Manriquez:1996:BMI

[MC96b] P. H. Manríquez and J. M. Cancino. Bryozoan-macroalgal interactions: do epibionts benefit? Marine Ecology Progress Series, 138:189–197, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/138/m138p189.pdf.

Maynou:1997:FED

[MC97a] F. Maynou and J. E. Cartes. Field estimation of daily ration in deep-sea shrimp Aristeus antennatus (Crustacea: Decapoda) in the western Mediterranean. Marine Ecology Progress Series, 153:191–196, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/153/m153p191.pdf.

Milligan:1997:GPB

[MC97b] A. J. Milligan and E. M. Cosper. Growth and photosynthesis of the ’brown tide’ microalga Aureococcus anophagefferens in subsaturating constant and fluctuating irradiance. Marine Ecology Progress Series, 153:67–75, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/153/m153p067.pdf.

Maynou:1998:DRE

[MC98] F. Maynou and J. E. Cartes. Daily ration estimates and comparative study of food consumption in nine species of deep-water decapod crustaceans of the NW Mediterranean. Marine Ecology Progress Series, 171:221–231, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
171/m171p221.pdf.

Mulholland:1999:NFU

[MC99] M. R. Mulholland and D. G. Capone. Nitrogen fixation, uptake and metabolism in natural and cultured populations of Trichodesmium spp. Marine Ecology Progress Series, 188: 33–49, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/188/m188p033.pdf.

Mayzaud:1990:EVN

[MCA90] P. Mayzaud, H. Claustre, and P. Augier. Effect of variable nutrient supply on fatty acid composition of phytoplankton grown in an enclosed experimental ecosystem. Marine Ecology Progress Series, 60:123–140, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/60/m060p123.pdf.

McCall:1992:SHC

[McC92] J. N. McCall. Source of harpacticoid copepods in the diet of juvenile starry flounder. Marine Ecology Progress Series, 86:41–50, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/86/m086p041.pdf.

McClanahan:1994:CES

[McC94a] T. R. McClanahan. Coral-eating snail Drupella cornus population increases in Kenyan coral reef lagoons. Marine Ecology Progress Series, 115:131–137, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/115/m115p131.pdf.

McClintock:1994:TBA

[McC94b] J. B. McClintock. Trophic biology of Antarctic shallowwater echinoderms. Marine Ecology Progress Series, 111:

191–202, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/111/m111p191.pdf.

McCormick:1994:CFM

[McC94c] M. I. McCormick. Comparison of field methods for measuring surface topography and their associations with a tropical reef fish assemblage. Marine Ecology Progress Series, 112: 87–96, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/112/m112p087.pdf.

McCormick:1994:VAS

[McC94d] M. I. McCormick. Variability in age and size at settlement of the tropicai goatfish Upeneus trayula (Mullidae) in the northern Great Barrier Reef lagoon. Marine Ecology Progress Series, 103:1–15, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/103/m103p001.pdf.

Machado:1996:CSL

[MCC96a] K. R. S. Sauer Machado, A. R. O. Chapman, and R. Coutinho. Consumer species have limited and variable roles in community organization on a tropical intertidal shore. Marine Ecology Progress Series, 134:73–83, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
134/m134p073.pdf.

McCook:1996:EHW

[McC96b] L. J. McCook. Effects of herbivores and water quality on Sargassum distribution on the central Great Barrier Reef: cross-shelf transplants. Marine Ecology Progress Series, 139: 179–192, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/139/m139p179.pdf.

MacIntyre:1997:VMN

[MCC97] J. G. MacIntyre, J. J. Cullen, and A. D. Cembella. Vertical migration, nutrition and toxicity in the dinoflagellate Alexandrium tamarense. Marine Ecology Progress Series, 148:201–216, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/148/m148p201.pdf.

McCormick:1999:DMT

[McC99] M. I. McCormick. Delayed metamorphosis of a tropical reef fish (Acanthurus triostegus): a field experiment. Marine Ecology Progress Series, 176:25–38, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/176/m176p025.pdf.

Marba:1996:GPW

[MCED96] N. Marbà, J. Cebrián, S. Enríquez, and C. M. Duarte. Growth patterns of Western Mediterranean seagrasses: speciesspecific responses to seasonal forcing. Marine Ecology Progress Series, 133:203–215, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/133/m133p203.pdf.

Mallet:1990:SMB

[MCF90] A. L. Mallet, C. E. A. Carver, and K. R. Freeman. Summer mortality of the blue mussel in eastern Canada: spatial, temporal, stock and age variation. Marine Ecology Progress Series, 67:35–41, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/67/m067p035.pdf.

McFadzen:1992:GSC

[McF92] I. R. B. McFadzen. Growth and survival of cryopreserved oyster and clam larvae along a pollution gradient in the German Bight. Marine Ecology Progress Series, 91:215–220, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
91/m091p215.pdf.

Milton:1997:ISE

[MCFB97] D. A. Milton, S. R. Chenery, M. J. Farmer, and S. J. M. Blaber. Identifying the spawning estuaries of the tropical shad, terubok Tenualosa toli, using otolith microchemistry. Marine Ecology Progress Series, 153:283–291, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
153/m153p283.pdf.

Monteiro:1996:MCP

[MCFS96] L. R. Monteiro, V. Costa, R. W. Furness, and R. S. Santos. Mercury concentrations in prey fish indicate enhanced bioaccumulation in mesopelagic environments. Marine Ecology Progress Series, 141:21–25, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/141/m141p021.pdf.

McGlathery:1992:PCD

[McG92] K. J. McGlathery. Physiological controls on the distribution of the macroalga Spyridea hypnoides: patterns along a eutrophication gradient in Bermuda. Marine Ecology Progress Series, 87:173–182, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/87/m087p173.pdf.

McGehee:1994:CBA

[McG94] M. A. McGehee. Correspondence between assemblages of coral reef fishes and gradients of water motion, depth, and substrate size off Puerto Rico. Marine Ecology Progress Series, 105:243–255, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/105/m105p243.pdf.

McGlathery:1995:NGI

[McG95] K. J. McGlathery. Nutrient and grazing influences on a subtropical seagrass community. Marine Ecology Progress Series, 122:239–252, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/122/m122p239.pdf.

McGrorty:1997:WGM

[McG97a] S. McGrorty. Winter growth of mussels Mytilus edulis as a possible counter to food depletion by oystercatchers Haematopus ostralegus. Marine Ecology Progress Series, 153:153– 165, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/153/m153p153.pdf.

McGuinness:1997:TAS

[McG97b] K. A. McGuinness. Tests for artefacts in some methods used to study herbivory and predation in mangrove forests. Marine Ecology Progress Series, 153:37–44, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/153/m153p037.pdf.

Mann:1990:SRR

[MCL90] R. Mann, B. M. Campos, and M. W. Luckenbach. Swimming rate and responses of larvae of three mactrid bivalves to salinity discontinuities. Marine Ecology Progress Series, 68: 257–269, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/68/m068p257.pdf.

McLachlan:1996:PFB

[McL96] A. McLachlan. Physical factors in benthic ecology: effects of changing sand particle size on beach fauna. Marine Ecology Progress Series, 131:205–217, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/131/m131p205.pdf.

Marsili:1997:AGO

[MCM^{+}97] L. Marsili, C. Casini, L. Marini, A. Regoli, and S. Focardi. Age, growth and organochlorines (HCB, DDTs and PCBs) in Mediterranean striped dolphins Stenella coeruleoalba stranded in 1988–1994 on the coasts of Italy. Marine Ecology Progress Series, 151:273–282, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/151/m151p273.pdf.

Martin:1998:EAE

[MCNB98] D. Martin, J. H. Cha, C. Nozais, and M. Bhaud. An experimental approach to the effects of varying recruitment strategy and food intake on early reproductive traits in a brooding Mediterranean polychaete. Marine Ecology Progress Series, 164:147–156, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/164/m164p147.pdf.

McGrorty:1990:PDM

[MCRGC90] S. McGrorty, R. T. Clarke, C. J. Reading, and J. D. Goss-Custard. Population dynamics of the mussel Mytilus edulis: density changes and regulation of the population in the Exe estuary, Devon. Marine Ecology Progress Series, 67:157– 169, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/67/m067p157.pdf.

Meador:1995:CBP

[MCSV95] J. P. Meador, E. Casillas, C. A. Sloan, and U. Varanasi. Comparative bioaccumulation of polycyclic aromatic hydrocarbons from sediment by two infaunal invertebrates. Marine Ecology Progress Series, 123:107–124, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
123/m123p107.pdf.

Miller:1991:PCF

[MCW^{+}91] C. B. Miller, T. J. Cowles, P. H. Wiebe, N. J. Copley, and H. Grigg. Phenology in Calanus finmarchicus; hypotheses about control mechanisms. Marine Ecology Progress Series, 72:79–91, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/72/m072p079.pdf.

Macpherson:1991:BTD

[MD91] E. Macpherson and C. M. Duarte. Bathymetric trends in demersal fish size: is there a general relationship? Marine Ecology Progress Series, 71:103–112, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/71/m071p103.pdf.

Martel:1993:EBS

[MD93] A. Martel and T. Diefenbach. Effects of body size, water current and microhabitat on mucous-thread drifting in post-metamorphic gastropods Lacuna spp. Marine Ecology Progress Series, 99:215–220, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/99/m099p215.pdf.

Marba:1994:GRS

[MD94a] N. Marbà and C. M. Duarte. Growth response of the seagrass Cymodocea nodosa to experimental burial and erosion. Marine Ecology Progress Series, 107:307–311, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
107/m107p307.pdf.

McManus:1994:PPD

[MD94b] G. B. McManus and R. Dawson. Phytoplankton pigments in the deep chlorophyll maximum of the Caribbean Sea and the western tropical Atlantic Ocean. Marine Ecology Progress Series, 113:199–206, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/113/m113p199.pdf.

Milicich:1994:LSC

[MD94c] M. J. Milicich and P. J. Doherty. Larval supply of coral reef fish populations: magnitude and synchrony of replenishment to Lizard Island, Great Barrier Reef. Marine Ecology Progress Series, 110:121–134, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/110/m110p121.pdf.

Marba:1998:RES

[MD98a] N. Marbà and C. M. Duarte. Rhizome elongation and seagrass clonal growth. Marine Ecology Progress Series, 174: 269–280, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/174/m174p269.pdf.

Markel:1998:MUE

[MD98b] R. W. Markel and R. E. DeWreede. Mechanisms underlying the effect of the chiton Katharina tunicata on the kelp Hedophyllum sessile: size escapes and indirect effects. Marine Ecology Progress Series, 166:151–161, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
166/m166p151.pdf.

Marba:1996:GPD

[MDC^{+}96] N. Marbà, C. M. Duarte, J. Cebrián, M. E. Gallegos, B. Olesen, and K. Sand-Jensen. Growth and population dynamics of Posidonia oceanica on the Spanish Mediterranean coast: elucidating seagrass decline. Marine Ecology Progress Series, 137:203–213, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/137/m137p203.pdf.

Marvin-DiPasquale:1998:BSR

[MDC98] M. C. Marvin-DiPasquale and D. G. Capone. Benthic sulfate reduction along the Chesapeake Bay central channel. I. Spatial trends and controls. Marine Ecology Progress Series, 168:213–228, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/168/m168p213.pdf.

Mante:1995:SMS

[MDD95] C. Manté, J.-C. Dauvin, and J.-P. Durbec. Statistical method for selecting representative species in multivariate analysis of long-term changes of marine communities. Applications to a macrobenthic community from the Bay of Morlaix. Marine Ecology Progress Series, 120:243–250, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
120/m120p243.pdf.

Migne:1998:CCP

[MDG98] A. Migné, D. Davoult, and J. P. Gattuso. Calcium carbonate production of a dense population of the brittle star Ophiothrix fragilis (Echinodermata: Ophiuroidea): role in the carbon cycle of a temperate coastal ecosystem. Marine Ecology Progress Series, 173:305–308, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/173/m173p305.pdf.

Malone:1991:PCF

[MDPP91] Thomas C. Malone, Hugh W. Ducklow, Emily R. Peele, and Sharon E. Pike. Picoplankton carbon flux in Chesapeake Bay. Marine Ecology Progress Series, 78:11–22, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
78/m078p011.pdf.

Morrisey:1999:VDM

[MDRW99] D. J. Morrisey, T. H. DeWitt, D. S. Roper, and R. B. Williamson. Variation in the depth and morphology of burrows of the mud crab Helice crassa among different types of intertidal sediment in New Zealand. Marine Ecology Progress Series, 182:231–242, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/182/m182p231.pdf.

Mitchell:1992:SPG

[MDSB92] N. D. Mitchell, M. R. Dardeau, W. W. Schroeder, and A. C. Benke. Secondary production of gorgonian corals in the northern Gulf of Mexico. Marine Ecology Progress Series, 87: 275–281, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/87/m087p275.pdf.

Menard:1994:TFT

[MDTB94] F. Ménard, S. Dallot, G. Thomas, and J. C. Braconnot. Temporal fluctuations of two Mediterranean salp populations from 1967 to 1990. Analysis of the influence of environmental variables using a Markov chain model. Marine Ecology Progress Series, 104:139–152, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/104/m104p139.pdf.

Miller:1997:INH

[ME97] B. A. Miller and R. B. Emlet. Influence of nearshore hydrodynamics on larval abundance and settlement of sea urchins Strongylocentrotus franciscanus and S. purpuratus in the Oregon upwelling zone. Marine Ecology Progress Series, 148: 83–94, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/148/m148p083.pdf.

McManus:1992:PPG

[MEC92] G. B. McManus and M. C. Ederington-Cantrell. Phytoplankton pigments and growth rates, and microzooplankton grazing in a large temperate estuary. Marine Ecology Progress Series, 87:77–85, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/87/m087p077.pdf.

Meeuwig:1999:PCE

[Mee99] J. J. Meeuwig. Predicting coastal eutrophication from landuse: an empirical approach to small non-stratified estuaries. Marine Ecology Progress Series, 176:231–241, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
176/m176p231.pdf.

Malm:1999:GET

[MEK99] T. Malm, R. Engkvist, and L. Kautsky. Grazing effects of two freshwater snails on juvenile Fucus vesiculosus in the Baltic Sea. Marine Ecology Progress Series, 188:63–71, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
188/m188p063.pdf.

Metz:1998:FOC

[Met98] C. Metz. Feeding of Oncaea curvata (Poecilostomatoida, Copepoda). Marine Ecology Progress Series, 169:229–235, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/169/m169p229.pdf.

McManus:1990:MSVa

[MF90a] G. B. McManus and J. A. Fuhrman. Mesoscale and seasonal variability of heterotrophic nanoflagellate abundance in an estuarine outflow plume. Marine Ecology Progress Series, 61:207–213, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/61/m061p207.pdf.

McManus:1990:MSVb

[MF90b] G. B. McManus and J. A. Fuhrman. Mesoscale and seasonal variability of heterotrophic nanoflagellate abundance in an estuarine outflow plume. Marine Ecology Progress Series, 61:207–213, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/61/m061p207.pdf.

Magnhagen:1991:BRD

[MF91] C. Magnhagen and E. Forsgren. Behavioural responses to different types of predators by sand goby Pomatoschistus minutus: an experimental study. Marine Ecology Progress Series, 70:11–16, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/70/m070p011.pdf.

Murphey:1995:RHL

[MF95] P. L. Murphey and M. S. Fonseca. Role of high and low energy seagrass beds as nursery areas for Penaeus duorarum in North Carolina. Marine Ecology Progress Series, 121:91–98, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
121/m121p091.pdf.

Meekan:1996:SFG

[MF96] M. G. Meekan and L. Fortier. Selection for fast growth during the larval life of Atlantic cod Gadus morhua on the Scotian Shelf. Marine Ecology Progress Series, 137:25–37, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
137/m137p025.pdf.

Maroz:1997:JPA

[MF97] A. Maroz and L. Fishelson. Juvenile production of Amphiprion bicinctus (Pomacentridae, teleostei) and rehabilitation of impoverished habitats. Marine Ecology Progress Series, 151:295–297, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/151/m151p295.pdf.

Merrett:1998:DID

[MF98] N. R. Merrett and M. J. R. Fasham. Demersal ichthyofaunal distribution in the abyssal North Atlantic revisited: the effect of sample size on ordination. Marine Ecology Progress Series, 173:267–274, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/173/m173p267.pdf.

Macedo:1998:DBP

[MFD98] M. F. Macedo, J. G. Ferreira, and P. Duarte. Dynamic behaviour of photosynthesis-irradiance curves determined from oxygen production during variable incubation periods. Marine Ecology Progress Series, 165:31–43, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
165/m165p031.pdf.

Madsen:1997:PMP

[MFF97] S. D. Madsen, T. L. Forbes, and V. E. Forbes. Particle mixing by the polychaete Capitella species 1: coupling fate and effect of a particle-bound organic contaminant (fluoranthene) in a marine sediment. Marine Ecology Progress Series, 147:129–142, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/147/m147p129.pdf.

Mauchline:1991:OPP

[MG91a] J. Mauchline and J. D. M. Gordon. Oceanic pelagic prey of benthopelagic fish in the benthic boundary layer of a marginal oceanic region. Marine Ecology Progress Series, 74: 109–115, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/74/m074p109.pdf.

Montgomery:1991:ESS

[MG91b] S. S. Montgomery and C. A. Gray. Effects of sizes of streamer tags on mortality and growth of juvenile eastern king prawns Penaeus plebejus. Marine Ecology Progress Series, 76:33–40, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/76/m076p033.pdf.

Millet:1994:SSV

[MG94] B. Millet and O. Guelorget. Spatial and seasonal variability in the relationships between benthic communities and physical environment in a lagoon ecosystem. Marine Ecology Progress Series, 108:161–174, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/108/m108p161.pdf.

MacIntyre:1996:RRA

[MG96a] H. L. MacIntyre and R. J. Geider. Regulation of Rubisco activity and its potential effect on photosynthesis during mixing in a turbid estuary. Marine Ecology Progress Series, 144:247–264, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/144/m144p247.pdf.

Macpherson:1996:BSB

[MG96b] E. Macpherson and A. Gordoa. Biomass spectra in benthic fish assemblages in the Benguela system. Marine Ecology Progress Series, 138:27–32, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/138/m138p027.pdf.

Maritorena:1996:OPW

[MG96c] S. Maritorena and N. Guillocheau. Optical properties of water and spectral light absorption by living and non-living particles and by yellow substances in coral reef waters of French Polynesia. Marine Ecology Progress Series, 131:245–255, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
131/m131p245.pdf.

Maranon:1997:PPC

[MG97] E. Marañón and N. González. Primary production, calcification and macromolecular synthesis in a bloom of the coccolithophore Emiliania huxleyi in the North Sea. Marine Ecology Progress Series, 157:61–77, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/157/m157p061.pdf.

Mock:1999:DAI

[MG99] T. Mock and R. Gradinger. Determination of Arctic ice algal production with a new in situ incubation technique. Marine Ecology Progress Series, 177:15–26, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/177/m177p015.pdf.

Mgaya:1992:DPC

[Mga92] Y. D. Mgaya. Density and production of Clinocottus globiceps and Oligocottus maculosus (Cottidae) in tidepools at Helby Island, British Columbia. Marine Ecology Progress Series, 85:219–225, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/85/m085p219.pdf.

Moran:1999:CBG

[MGAE99] X. A. G. Morán, J. M. Gasol, L. Arin, and M. Estrada. A comparison between glass fiber and membrane filters for the estimation of phytoplankton POC and DOC production. Marine Ecology Progress Series, 187:31–41, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/187/m187p031.pdf.

McGrorty:1991:PDM

[MGC91] S. McGrorty and J. D. Goss-Custard. Population dynamics of the mussel Mytilus edulis: spatial variations in ageclass densities of an intertidal estuarine population along environmental gradients. Marine Ecology Progress Series, 73: 191–202, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/73/m073p191.pdf.

McGrorty:1995:PDM

[MGC95a] S. McGrorty and J. D. Goss-Custard. Population dynamics of Mytilus edulis along environmental gradients: densitydependent changes in adult mussel numbers. Marine Ecology Progress Series, 129:197–213, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/129/m129p197.pdf.

Moore:1995:CPS

[MGC95b] L. R. Moore, R. Goericke, and S. W. Chisholm. Comparative physiology of Synechococcus and Prochlorococcus: influence of light and temperature on growth, pigments, fluorescence and absorptive properties. Marine Ecology Progress Series, 116:259–275, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/116/m116p259.pdf.

Muscatine:1991:RSA

[MGD91] L. Muscatine, D. Grossman, and J. Doino. Release of symbiotic algae by tropical sea animones and corals after cold shock. Marine Ecology Progress Series, 77:233–243, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
77/m077p233.pdf.

Marsh:1990:TAP

[MGDT90] A. G. Marsh, A. Gremare, R. Dawson, and K. R. Tenore. Translocation of algal pigments to oocytes in Capitella sp. I (Annelida: Polychaeta). Marine Ecology Progress Series, 67: 301–304, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/67/m067p301.pdf.

Mumby:1997:MSS

[MGEC97] P. J. Mumby, E. P. Green, A. J. Edwards, and C. D. Clark. Measurement of seagrass standing crop using satellite and digital airborne remote sensing. Marine Ecology Progress Series, 159:51–60, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/159/m159p051.pdf.

Monteiro:1998:RBM

[MGF98] L. R. Monteiro, J. P. Granadeiro, and R. W. Furness. Relationship between mercury levels and diet in Azores seabirds. Marine Ecology Progress Series, 166:259–265, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
166/m166p259.pdf.

Mathieson:1996:TVT

[MGM96] S. Mathieson, S. G. George, and D. S. McLusky. Temporal variation of total mercury concentrations and burdens in the liver of eelpout Zoarces viviparus from the Forth Estuary, Scotland: implications for mercury biomonitoring. Marine Ecology Progress Series, 138:41–49, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/138/m138p041.pdf.

Manuel:1996:VDP

[MGP^{+}96] J. L. Manuel, S. M. Gallager, C. M. Pearce, D. A. Manning, and R. K. O’Dor. Veligers from different populations of sea scallop Placopecten magellanicus have different vertical migration patterns. Marine Ecology Progress Series, 142:147–163, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/142/m142p147.pdf.

Maldonado:1997:DRP

[MGYV97] M. Maldonado, S. B. George, C. M. Young, and I. Vaquerizo. Depth regulation in parenchymella larvae of a demosponge: relative roles of skeletogenesis, biochemical changes and behavior. Marine Ecology Progress Series, 148:115–124, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
148/m148p115.pdf.

Moran:1990:CDS

[MH90] M. A. Moran and R. E. Hodson. Contributions of degrading Spartina alterniflora lignocellulose to the dissolved organic carbon pool of a salt marsh. Marine Ecology Progress Series, 62:161–168, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/62/m062p161.pdf.

Malej:1993:ICG

[MH93a] A. Malej and R. P. Harris. Inhibition of copepod grazing by diatom exudates: a factor in the development of mucus aggregates? Marine Ecology Progress Series, 96:33–42, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
96/m096p033.pdf.

Mazur:1993:PAI

[MH93b] H. Mazur and E. Homme. Presence of auxin indole-3-acetic acid in the northern Adriatic Sea: phytohormones and mucilage. Marine Ecology Progress Series, 99:163–168, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
99/m099p163.pdf.

Millham:1994:NBS

[MH94a] N. P. Millham and B. L. Howes. Nutrient balance of a shallow coastal embayment: I. Patterns of groundwater discharge. Marine Ecology Progress Series, 112:155–167, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
112/m112p155.pdf.

Moran:1994:SBP

[MH94b] M. A. Moran and R. E. Hodson. Support of bacterioplankton production by dissolved humic substances from three marine environments. Marine Ecology Progress Series, 110:241– 247, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/110/m110p241.pdf.

MacGregor:1996:OOP

[MH96a] J. M. MacGregor and E. D. Houde. Onshore-offshore pattern and variability in distribution and abundance of bay anchovy Anchoa mitchilli eggs and larvae in Chesapeake Bay. Marine Ecology Progress Series, 138:15–25, 1996. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/138/m138p015.pdf.

Muggli:1996:ENS

[MH96b] D. L. Muggli and P. J. Harrison. Effects of nitrogen source on the physiology and metal nutrition of Emiliania huxleyi grown under different iron and light conditions. Marine Ecology Progress Series, 130:255–267, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/130/m130p255.pdf.

Menesguen:1997:MBC

[MH97] A. Ménesguen and T. Hoch. Modelling the biogeochemical cycles of elements limiting primary production in the English Channel: I. Role of thermohaline stratification. Marine Ecology Progress Series, 146:173–188, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/146/m146p173.pdf.

Maranon:1999:PPP

[MH99] E. Marañón and P. M. Holligan. Photosynthetic parameters of phytoplankton from 50^{o}N to 50^{o}S in the Atlantic Ocean. Marine Ecology Progress Series, 176:191–203, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
176/m176p191.pdf.

Mathers:1994:RDP

[MHB94] E. M. Mathers, D. F. Houlihan, and L. J. Burren. RNA, DNA and protein concentrations in fed and starved herring Clupea harengus larvae. Marine Ecology Progress Series, 107: 223–231, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/107/m107p223.pdf.

Myers:1996:HDC

[MHB96] R. A. Myers, J. A. Hutchings, and N. J. Barrowman. Hypotheses for the decline of cod in the North Atlantic. Marine Ecology Progress Series, 138:293–308, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
138/m138p293.pdf.

Martin:1990:NPS

[MHBW90] J. L. Martin, K. Haya, L. E. Burridge, and D. J. Wildish. Nitzschia pseudodelicatissima — a source of domoic acid in the Bay of Fundy, eastern Canada. Marine Ecology Progress Series, 67:177–182, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/67/m067p177.pdf.

Mathers:1992:ESP

[MHC92] E. M. Mathers, D. F. Houlihan, and M. J. Cunningham. Estimation of saithe Pollachius virens growth rates around the beryl oil platforms in the North Sea: a comparison of methods. Marine Ecology Progress Series, 86:31–40, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
86/m086p031.pdf.

Mohammadian:1997:MPI

[MHD97] M. A. Mohammadian, S. Hansson, and B. T. De Stasio. Are marine planktonic invertebrates food limited? The functional response of Mysis mixta (Crustacea, Mysidacea) in the Baltic Sea. Marine Ecology Progress Series, 150:113–119, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
150/m150p113.pdf.

Montoya:1990:NAP

[MHM90] J. P. Montoya, S. G. Horrigan, and J. J. McCarthy. Natural abundance of ^{15}N in particulate nitrogen and zooplankton in the Chesapeake Bay. Marine Ecology Progress Series, 65:35–61, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/65/m065p035.pdf.

Moodley:1996:SAB

[MHMH96] L. Moodley, P. M. J. Herman, J. J. Middelburg, and C. H. R. Heip. Subsurface activity of benthic Foraminifera. Marine Ecology Progress Series, 145:303–304, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/

145/m145p303.pdf. See response [BAR96].

Meyer-Harms:1999:FSC

[MHRV^{+}99] B. Meyer-Harms, M. Reckermann, M. Voß, H. Siegmund, and B. von Bodungen. Food selection by calanoid copepods in the euphotic layer of the Gotland Sea (Baltic Proper) during mass occurrence of N_{2}-fixing cyanobacteria. Marine Ecology Progress Series, 191:243–250, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/191/m191p243.pdf.

Metaxas:1994:STVa

[MHS94] Anna Metaxas, Heather L. Hunt, and Robert E. Scheibling. Spatial and temporal variability of macrobenthic communities in tidepools on a rocky shore in Nova Scotia, Canada. Marine Ecology Progress Series, 105:89–103, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
105/m105p089.pdf.

McMinn:1997:CRE

[MHT^{+}97] A. McMinn, G. M. Hallegraeff, P. Thomson, A. V. Jenkinson, and H. Heijnis. Cyst and radionucleotide evidence for the recent introduction of the toxic dinoflagellate Gymnodinium catenatum into Tasmanian waters. Marine Ecology Progress Series, 161:165–172, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/161/m161p165.pdf.

Meyer-Harms:1997:TSI

[MHvB97] B. Meyer-Harms and B. von Bodungen. Taxon-specific ingestion rates of natural phytoplankton by calanoid copepods in an estuarine environment (Pomeranian Bight, Baltic Sea) determined by cell counts and HPLC analyses of marker pigments. Marine Ecology Progress Series, 153:181–190, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
153/m153p181.pdf.

Marigomez:1990:CCS

[MIA90] I. Marigómez, M. P. Ireland, and E. Angulo. Correlation of cadmium shell-weight index with environmental stress indicators at the cellular and organismic levels in Littorina littorea. Marine Ecology Progress Series, 67:171–176, 1990.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
67/m067p171.pdf.

Micheli:1996:PIE

[Mic96] F. Micheli. Predation intensity in estuarine soft bottoms: between-habitat comparisons and experimental artifacts. Marine Ecology Progress Series, 141:295–302, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
141/m141p295.pdf.

Mikaelyan:1995:WBD

[Mik95] A. S. Mikaelyan. Winter bloom of the diatom Nitzschia delicatula in the open waters of the Black Sea. Marine Ecology Progress Series, 129:241–251, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/129/m129p241.pdf.

Mikhalev:1997:HWM

[Mik97] Y. A. Mikhalev. Humpback whales Megaptera novaeangliae in the Arabian Sea. Marine Ecology Progress Series, 149: 13–21, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/149/m149p013.pdf.

Miller:1990:CJA

[Mil90] R. J. Miller. Comment on Jellett et al. (1989). Marine Ecology Progress Series, 63:305–306, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:/

/www.int-res.com/articles/meps/63/m063p305.pdf. See

[JNCS89].

Miller:1992:RSK

[Mil92] D. C. Miller. Response of Saccoglossus kowalevskii (Phylum Hemichordata, class Enteropneusta) to changes in diet. Marine Ecology Progress Series, 87:41–54, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
87/m087p041.pdf.

Milicich:1994:DCR

[Mil94a] M. J. Milicich. Dynamic coupling of reef fish replenishment and oceanographic processes. Marine Ecology Progress Series, 110:135–144, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/110/m110p135.pdf.

Miller:1994:MVC

[Mil94b] K. J. Miller. Morphological variation in the coral genus Platygyra: environmental influences and taxonomic implications. Marine Ecology Progress Series, 110:19–28, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
110/m110p019.pdf.

Miller:1995:SAL

[Mil95a] M. J. Miller. Species assemblages of leptocephali in the Sargasso Sea and Florida Current. Marine Ecology Progress Series, 121:11–26, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/121/m121p011.pdf.

Miller:1995:GTC

[Mil95b] M. W. Miller. Growth of a temperate coral: effects of temperature, light, depth, and heterotrophy. Marine Ecology Progress Series, 122:217–225, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/122/m122p217.pdf.

Miller:1997:GSB

[Mil97] K. J. Miller. Genetic structure of black coral populations in New Zealand’s fiords. Marine Ecology Progress Series, 161:123–132, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/161/m161p123.pdf.

Miller:1998:SDD

[Mil98] K. J. Miller. Short-distance dispersal of black coral larvae: inference from spatial analysis of colony genotypes. Marine Ecology Progress Series, 163:225–233, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
163/m163p225.pdf.

Minello:1993:CTT

[Min93] T. J. Minello. Chronographic tethering: a technique for measuring prey survival time and testing predation pressure in aquatic habitats. Marine Ecology Progress Series, 101:99– 104, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/101/m101p099.pdf.

May:1991:PSG

[MJ91] H. M. A. May and G. P. Jenkins. Patterns of settlement and growth of juvenile flounder Rhombosolea tapirina determined from otolith microstructure. Marine Ecology Progress Series, 79:203–214, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/79/m079p203.pdf.

Mouritsen:1992:CMT

[MJ92a] K. N. Mouritsen and K. T. Jensen. Choice of microhabitat in tactile foraging dunlins Calidris alpina: the importance of sediment penetrability. Marine Ecology Progress Series, 85:1–8, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/85/m085p001.pdf.

Murray:1992:VDM

[MJ92b] A. G. Murray and G. A. Jackson. Viral dynamics: a model of the effects of size shape, motion and abundance of singlecelled planktonic organisms and other particles. Marine Ecology Progress Series, 89:103–116, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/89/m089p103.pdf.

Murray:1993:VDI

[MJ93] A. G. Murray and G. A. Jackson. Viral dynamics II: a model of the interaction of ultraviolet light and mixing processes on virus survival in seawater. Marine Ecology Progress Series, 102:105–114, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/102/m102p105.pdf.

McIlwain:1997:PSO

[MJ97a] J. L. McIlwain and G. P. Jones. Prey selection by an obligate coral-feeding wrasse and its response to small-scale disturbance. Marine Ecology Progress Series, 155:189–198, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
155/m155p189.pdf.

Mouritsen:1997:PTB

[MJ97b] K. N. Mouritsen and K. T. Jensen. Parasite transmission between soft-bottom invertebrates: temperature mediated infection rates and mortality in Corophium volutator. Marine Ecology Progress Series, 151:123–134, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
151/m151p123.pdf.

Munday:1997:HSD

[MJC97] P. L. Munday, G. P. Jones, and M. J. Caley. Habitat specialisation and the distribution and abundance of coral-dwelling gobies. Marine Ecology Progress Series, 152:227–239, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
152/m152p227.pdf.

Monteiro:1991:TPI

[MJSDF91] P. M. S. Monteiro, A. G. James, A. D. Sholto-Douglas, and J. G. Field. The δ ^{13}C trophic position isotope spectrum as a tool to define and quantify carbon pathways in marine food webs. Marine Ecology Progress Series, 78:33–40, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
78/m078p033.pdf.

Martin-Jezequel:1992:PBT

[MJV92] V. Martin-Jézéquel and C. Videau. Phytoplankton and bacteria over the transient area of the continental slope of the Celtic Sea in spring. I. Vertical distribution and productivity. Marine Ecology Progress Series, 85:289–301, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
85/m085p289.pdf.

MacKenzie:1996:EBC

[MJW96] B. MacKenzie, M. St John, and K. Wieland. Eastern Baltic cod: perspectives from existing data on processes affecting growth and survival of eggs and larvae. Marine Ecology Progress Series, 134:265–281, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/134/m134p265.pdf.

Miron:1993:BRT

[MK93a] G. Miron and E. Kristensen. Behavioural response of three nereid polychaetes to injection of sulfide inside burrows. Marine Ecology Progress Series, 101:147–155, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
101/m101p147.pdf.

Miron:1993:FID

[MK93b] G. Miron and E. Kristensen. Factors influencing the distribution of nereid polychaetes: the sulfide aspect. Marine Ecology Progress Series, 93:143–153, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/93/m093p143.pdf.

Mingelbier:1994:MDP

[MKCL94] M. Mingelbier, B. Klein, M. R. Claereboudt, and L. Legendre. Measurement of daily primary production using 24 h incubations with the ^{14}C method: a caveat. Marine Ecology Progress Series, 113:301–309, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/113/m113p301.pdf.

Middelburg:1996:OMM

[MKN^{+}96] J. J. Middelburg, G. Klaver, J. Nieuwenhuize, A. Wielemaker, W. de Haas, T. Vlug, and J. F. W. A. van der Nat. Organic matter mineralization in intertidal sediments along an estuarine gradient. Marine Ecology Progress Series, 132:157–168, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/132/m132p157.pdf.

Millie:1995:RPP

[MKV95] D. F. Millie, G. J. Kirkpatrick, and B. T. Vinyard. Relating photosynthetic pigments and in vivo optical density spectra to irradiance for the Florida red-tide dinoflagellate Gymnodinium breve. Marine Ecology Progress Series, 120: 65–75, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/120/m120p065.pdf.

Miyajima:1998:ATS

[MKYI98] T. Miyajima, I. Koike, H. Yamano, and H. Iizumi. Accumulation and transport of seagrass-derived organic matter in reef flat sediment of Green Island, Great Barrier Reef. Marine Ecology Progress Series, 175:251–259, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
175/m175p251.pdf.

Monger:1990:DIF

[ML90] B. C. Monger and M. R. Landry. Direct-interception feeding by marine zooflagellates: the importance of surface and hydrodynamic forces. Marine Ecology Progress Series, 65: 123–140, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/65/m065p123.pdf.

MacKenzie:1991:QCS

[ML91a] B. R. MacKenzie and W. C. Leggett. Quantifying the contribution of small-scale turbulence to the encounter rates between larval fish and their zooplankton prey: effects of wind and tide. Marine Ecology Progress Series, 73:149–160, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
73/m073p149.pdf.

Mansour:1991:DDF

[ML91b] R. A. Mansour and R. N. Lipcius. Density-dependent foraging and mutual interference in blue crabs preying upon infaunal clams. Marine Ecology Progress Series, 72:239–246, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
72/m072p239.pdf.

Monger:1991:PSD

[ML91c] B. C. Monger and M. R. Landry. Prey-size dependency of grazing by free-living marine flagellates. Marine Ecology Progress Series, 74:239–248, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/74/m074p239.pdf.

Metcalf:1992:RHS

[ML92] K. S. Metcalf and R. N. Lipcius. Relationship of habitat and spatial scale with physiological state and settlement of blue crab postlarvae in Chesapeake Bay. Marine Ecology Progress Series, 82:143–150, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/82/m082p143.pdf.

MacKenzie:1993:WBM

[ML93] B. R. MacKenzie and W. C. Leggett. Wind-based models for estimating the dissipation rates of turbulent energy in aquatic environments: empirical comparisons. Marine Ecology Progress Series, 94:207–216, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/94/m094p207.pdf.

Marrase:1992:SDC

[MLC92] C. Marrasé, E. L. Lim, and D. A. Caron. Seasonal and daily changes in bacterivory in a coastal plankton community. Marine Ecology Progress Series, 82:281–289, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
82/m082p281.pdf.

Marcus:1997:IAS

[MLC97] N. H. Marcus, R. V. Lutz, and J. P. Chanton. Impact of anoxia and sulfide on the viability of eggs of three planktonic copepods. Marine Ecology Progress Series, 146:291– 295, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/146/m146p291.pdf.

Munk:1995:LSJ

[MLDM95] P. Munk, P. O. Larsson, D. Danielsen, and E. Moksness. Larval and small juvenile cod Gadus morhua concentrated in the highly productive areas of a shelf break front. Marine Ecology Progress Series, 125:21–30, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/125/m125p021.pdf.

Munk:1999:VFZ

[MLDM99] P. Munk, P. O. Larsson, D. S. Danielssen, and E. Moksness. Variability in frontal zone formation and distribution of gadoid fish larvae at the shelf break in the northeastern North Sea. Marine Ecology Progress Series, 177:221–233, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
177/m177p221.pdf.

Mintz:1994:SJC

[MLES94] J. D. Mintz, R. N. Lipcius, D. B. Eggleston, and M. S. Seebo. Survival of juvenile Caribbean spiny lobster; effects of shelter size, geographic location and conspecific abundance. Marine Ecology Progress Series, 112:255–266, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
112/m112p255.pdf.

Muggli:1996:EIN

[MLH96] D. L. Muggli, M. Lecourt, and P. J. Harrison. Effects of iron and nitrogen source on the sinking rate, physiology and metal composition of an oceanic diatom from the subarctic Pacific. Marine Ecology Progress Series, 132:215–227, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
132/m132p215.pdf.

Mokady:1998:ACB

[MLL98] O. Mokady, Y. Loya, and B. Lazar. Ammonium contribution from boring bivalves to their coral host — a mutualistic symbiosis? Marine Ecology Progress Series, 169:295–301, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
169/m169p295.pdf.

MacKenzie:1990:ELF

[MLP90] B. R. MacKenzie, W. C. Leggett, and R. H. Peters. Estimating larval fish ingestion rates: can laboratory derived values be reliably extrapolated to the wild? Marine Ecology Progress Series, 67:209–225, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/67/m067p209.pdf.

Monti:1996:HDS

[MLTD96] D. Monti, L. Legendre, J. C. Therriault, and S. Demers. Horizontal distribution of sea-ice microalgae: environmental control and spatial processes (southeastern Hudson Bay, Canada). Marine Ecology Progress Series, 133:229–240, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
133/m133p229.pdf.

Maberly:1990:CAW

[MM90] S. C. Maberly and T. V. Madsen. Contribution of air and water to the carbon balance of Fucus spiralis. Marine Ecology Progress Series, 62:175–183, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/62/m062p175.pdf.

Malan:1991:SBO

[MM91a] D. E. Malan and A. McLachlan. In situ benthic oxygen fluxes in a near-shore coastal marine system: a new approach to quantify the effect of wave action. Marine Ecology Progress Series, 73:69–81, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/73/m073p069.pdf.

McArdle:1991:DSZ

[MM91b] S. B. McArdle and A. McLachlan. Dynamics of the swash zone and effluent line on sandy beaches. Marine Ecology Progress Series, 76:91–99, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/76/m076p091.pdf.

McCormick:1993:QRF

[MM93] M. I. McCormick and B. W. Molony. Quality of the reef fish Upeneus tragula (Mullidae) at settlement: is size a good indicator of condition? Marine Ecology Progress Series, 98: 45–54, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/98/m098p045.pdf.

McCormick:1995:IWT

[MM95a] M. I. McCormick and B. W. Molony. Influence of water temperature during the larval stage on size, age and body condition of a tropical reef fish at settlement. Marine Ecology Progress Series, 118:59–68, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/118/m118p059.pdf.

Montevecchi:1995:PHS

[MM95b] W. A. Montevecchi and R. A. Myers. Prey harvests of seabirds reflect pelagic fish and squid abundance on multiple spatial and temporal scales. Marine Ecology Progress Series, 117:1–9, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/117/m117p001.pdf.

Minier:1996:RBA

[MM96] C. Minier and M. N. Moore. Rhodamine B accumulation and MXR protein expression in mussel blood cells: effects of exposure to vincristine. Marine Ecology Progress Series, 142:165–173, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/142/m142p165.pdf.

McCormick:1997:PST

[MM97a] M. I. McCormick and L. J. Makey. Post-settlement transition in coral reef fishes: overlooked complexity in niche shifts. Marine Ecology Progress Series, 153:247–257, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
153/m153p247.pdf.

McKillup:1997:EFS

[MM97b] S. C. McKillup and R. V. McKillup. Effect of food supplementation on the growth of an intertidal scavenger. Marine Ecology Progress Series, 148:109–114, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/148/m148p109.pdf.

Marsh:1999:MAM

[MM99] A. G. Marsh and D. T. Manahan. A method for accurate measurements of the respiration rates of marine invertebrate embryos and larvae. Marine Ecology Progress Series, 184:1–10, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/184/m184p001.pdf.

Millar:1997:EIF

[MMBT97] R. B. Millar, J. E. McKenzie, J. D. Bell, and L. D. Tierney. Evaluation of an indigenous fishing calendar using recreational catch rates of snapper Pagrus auratus in the North Island of New Zealand. Marine Ecology Progress Series, 151: 219–224, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/151/m151p219.pdf.

Milicich:1992:LSG

[MMD92] M. J. Milicich, M. G. Meekan, and P. J. Doherty. Larval supply: a good predictor of recruitment of three species of reef fish (Pomacentridae). Marine Ecology Progress Series, 86:153–166, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/86/m086p153.pdf.

Meekan:1993:LPD

[MMD93] M. G. Meekan, M. J. Milicich, and P. J. Doherty. Larval production drives temporal patterns of larval supply and recruitment of a coral reef damselfish. Marine Ecology Progress Series, 93:217–225, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/93/m093p217.pdf.

Mock:1997:BSI

[MMG97] T. Mock, K. M. Meiners, and H. C. Giesenhagen. Bacteria in sea ice and underlying brackish water at 54^{o}26’ 5”n (Baltic Sea, Kiel Bight). Marine Ecology Progress Series, 158: 23–40, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/158/m158p023.pdf.

Meise-Munns:1990:IVC

[MMGIM90] C. Meise-Munns, J. Green, M. Ingham, and D. Mountain. Interannual variability in the copepod populations of Georges Bank and the western Gulf of Maine. Marine Ecology Progress Series, 65:225–232, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/65/m065p225.pdf.

Mackey:1996:CPE

[MMHW96] M. D. Mackey, D. J. Mackey, H. W. Higgins, and S. W. Wright. CHEMTAX — a program for estimating class abundances from chemical markers: application to HPLC measurements of phytoplankton. Marine Ecology Progress Series, 144:265–283, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/144/m144p265.pdf.

Malej:1995:PRF

[MMM^{+}95] A. Malej, P. Mozetic, V. Malacic, S. Terzic, and M. Ahel. Phytoplankton responses to freshwater inputs in a small semienclosed gulf (Gulf of Trieste, Adriatic Sea). Marine Ecology Progress Series, 120:111–121, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/120/m120p111.pdf.

Meyercordt:1999:PPB

[MMR99] J. Meyercordt and L. A. Meyer-Reil. Primary production of benthic microalgae in two shallow coastal lagoons of different trophic status in the southern Baltic Sea. Marine Ecology Progress Series, 178:179–191, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/178/m178p179.pdf.

Muschenheim:1992:USF

[MN92] D. K. Muschenheim and C. R. Newell. Utilization of seston flux over a mussel bed. Marine Ecology Progress Series, 85:131–136, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/85/m085p131.pdf.

Mueter:1999:LCS

[MN99] F. J. Mueter and B. L. Norcross. Linking community structure of small demersal fishes around Kodiak Island, Alaska, to environmental variables. Marine Ecology Progress Series, 190:37–51, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/190/m190p037.pdf.

Meesters:1994:DRL

[MNB94] E. H. Meesters, M. Noordeloos, and R. P. M. Bak. Damage and regeneration: links to growth in the reef-building coral Montastrea annularis. Marine Ecology Progress Series, 112:119–128, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/112/m112p119.pdf.

Mueller-Niklas:1992:AFC

[MNH92] G. Mueller-Niklas and G. J. Herndl. Activity of fecal coliform bacteria measured by 4-methylumbelliferyl-β-D-glucuronide substrate in the northern Adriatic Sea with special reference to marine snow. Marine Ecology Progress Series, 89: 305–309, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/89/m089p305.pdf.

Moberg:1997:ERS

[MNK^{+}97] F. Moberg, M. Nyström, N. Kautsky, M. Tedengren, and P. Jarayabhand. Effects of reduced salinity on the rates of photosynthesis and respiration in the hermatypic corals Porites lutea and Pocillopora damicornis. Marine Ecology Progress Series, 157:53–59, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/157/m157p053.pdf.

Moore:1996:ZME

[MNO96] K. A. Moore, H. A. Neckles, and R. J. Orth. Zostera marina (eelgrass) growth and survival along a gradient of nutrients and turbidity in the lower Chesapeake Bay. Marine Ecology Progress Series, 142:247–259, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/142/m142p247.pdf.

Moriarty:1995:OCT

[MO95] D. J. W. Moriarty and M. J. O’Donohue. Organic carbon transport from the Southern Ocean and bacterial growth in the Antarctic intermediate water masses of the Tasman Sea. Marine Ecology Progress Series, 119:291–297, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
119/m119p291.pdf.

Minami:1997:DMD

[MO97] H. Minami and H. Ogi. Determination of migratory dynamics of the sooty shearwater in the Pacific using stable carbon and nitrogen isotope analysis. Marine Ecology Progress Series, 158:249–256, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/158/m158p249.pdf.

Moebus:1996:MBRa

[Moe96a] K. Moebus. Marine bacteriophage reproduction under nutrient-limited growth of host bacteria. I. Investigations with six phage-host systems. Marine Ecology Progress Series, 144: 1–12, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/144/m144p001.pdf.

Moebus:1996:MBRb

[Moe96b] K. Moebus. Marine bacteriophage reproduction under nutrient-limited growth of host bacteria. II. Investigations with phage-host system [H3:H3/1]. Marine Ecology Progress Series, 144:13–22, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/144/m144p013.pdf.

Moebus:1997:IMLa

[Moe97a] K. Moebus. Investigations of the marine lysogenic bacterium H24. I. General description of the phage-host system. Marine Ecology Progress Series, 148:217–2282, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
148/m148p217.pdf.

Moebus:1997:IMLb

[Moe97b] K. Moebus. Investigations of the marine lysogenic bacterium H24. II. Development of pseudolysogeny in nutrient-rich broth culture. Marine Ecology Progress Series, 148:229–240, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
148/m148p229.pdf.

Moebus:1997:IMLc

[Moe97c] K. Moebus. Investigations of the marine lysogenic bacterium H24. III. Growth of bacteria and production of phage under nutrient-limited conditions. Marine Ecology Progress Series, 148:241–250, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/148/m148p241.pdf.

Moll:1997:PPD

[Mol97] A. Moll. Phosphate and plankton dynamics during a drift experiment in the German Bight: simulation of phosphorusrelated plankton production. Marine Ecology Progress Series, 156:289–297, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/156/m156p289.pdf.

Moisan:1998:XCP

[MOM98] T. A. Moisan, M. Olaizola, and B. G. Mitchell. Xanthophyll cycling in Phaeocystis antarctica: changes in cellular fluorescence. Marine Ecology Progress Series, 169:113–121, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
169/m169p113.pdf.

Montagnes:1996:GRP

[Mon96] D. J. S. Montagnes. Growth responses of planktonic ciliates in the genera Strobilidium and Strombidium. Marine Ecology Progress Series, 130:241–254, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/130/m130p241.pdf.

Mauri:1990:HMA

[MONR90] M. Mauri, E. Orlando, M. Nigro, and F. Regoli. Heavy metals in the Antarctic scallop Adamussium colbecki. Marine Ecology Progress Series, 67:27–33, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/67/m067p027.pdf.

Moore:1992:MCPa

[Moo92a] M. N. Moore. Molecular and cellular pathology: summary. Marine Ecology Progress Series, 91:117–119, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
91/m091p117.pdf.

Moore:1992:MCPb

[Moo92b] M. N. Moore. Molecular cell pathology of pollutant-induced liver injury in flatfish: use of fluorescent probes. Marine Ecology Progress Series, 91:127–133, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/91/m091p127.pdf.

Moore:1994:OBS

[Moo94] P. G. Moore. Observations on the behaviour of the scavenging lysianassoid Orchomene zschaui (Crustacea: Amphipoda) from South Georgia (South Atlantic). Marine Ecology Progress Series, 113:29–38, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/113/m113p029.pdf.

Morgan:1996:PRT

[Mor96] S. G. Morgan. Plasticity in reproductive timing by crabs in adjacent tidal regimes. Marine Ecology Progress Series, 139:105–118, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/139/m139p105.pdf.

Meyer:1992:IVS

[MP92] K. D. Meyer and V. J. Paul. Intraplant variation in secondary metabolite concentration in three species of Caulerpa (Chlorophyta-Caulerpales) and its effects on herbivorous fishes. Marine Ecology Progress Series, 82:249–257, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
82/m082p249.pdf.

Mallin:1994:CPP

[MP94] M. A. Mallin and H. W. Paerl. Commentary on primary productivity and nutrient limitation in the Neuse River Estuary, North Carolina, USA. Marine Ecology Progress Series, 111:311–312, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/111/m111p311.pdf.

Montecino:1995:PAS

[MP95] V. Montecino and G. Pizarro. Phytoplankton acclimation and spectral penetration of UV irradiance off the central Chilean coast. Marine Ecology Progress Series, 121:261–269, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
121/m121p261.pdf.

Maldonado:1996:ISF

[MP96a] M. T. Maldonado and N. M. Price. Influence of N substrate on Fe requirements of marine centric diatoms. Marine Ecology Progress Series, 141:161–172, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/141/m141p161.pdf.

Mayer:1996:ECP

[MP96b] M. Mayer and D. Piepenburg. Epibenthic community patterns on the continental slope off East Greenland at 75* N. Marine Ecology Progress Series, 143:151–164, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
143/m143p151.pdf.

Moline:1996:LTM

[MP96c] M. A. Moline and B. B. Prézelin. Long-term monitoring and analyses of physical factors regulating variability in coastal Antarctic phytoplankton biomass, in situ productivity and taxonomic composition over subseasonal, seasonal and interannual time scales. Marine Ecology Progress Series, 145: 143–160, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/145/m145p143.pdf.

Meesters:1997:PRP

[MPB97] E. H. Meesters, W. Pauchli, and R. P. M. Bak. Predicting regeneration of physical damage on a reef-building coral by regeneration capacity and lesion shape. Marine Ecology Progress Series, 146:91–99, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/146/m146p091.pdf.

Muller-Parker:1990:FAP

[MPCD90] G. Muller-Parker, C. B. Cook, and C. F. D’Elia. Feeding affects phosphate fluxes in the symbiotic sea anemone Aiptasia pallida. Marine Ecology Progress Series, 60:283–290, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
60/m060p283.pdf.

Mercier:1996:TRS

[MPH96] A. Mercier, É. Pelletier, and J. F. Hamel. Toxicological response of the symbiotic sea anemone Aiptasia pallida to butyltin contamination. Marine Ecology Progress Series, 144: 133–146, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/144/m144p133.pdf.

Madin:1997:AGE

[MPM97] L. P. Madin, J. E. Purcell, and C. B. Miller. Abundance and grazing effects of Cyclosalpa bakeri in the subarctic Pacific. Marine Ecology Progress Series, 157:175–183, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
157/m157p175.pdf.

Mallin:1993:REP

[MPRB93] M. A. Mallin, H. W. Paerl, J. Rudek, and P. W. Bates. Regulation of estuarine primary production by watershed rainfall and river flow. Marine Ecology Progress Series, 93:199– 203, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/93/m093p199.pdf.

Mordy:1995:SDB

[MPS95] C. W. Mordy, D. M. Penny, and C. W. Sullivan. Spatial distribution of bacterioplankton biomass and production in the marginal ice-edge zone of the Weddell–Scotia Sea during austral winter. Marine Ecology Progress Series, 122:9–19, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
122/m122p009.pdf.

Martin:1996:GMP

[MPS96] D. Martin, S. Pinedo, and R. Sardá. Grazing by meroplanktonic polychaete larvae may help to control nanoplankton in the NW Mediterranean littoral: in situ experimental evidence. Marine Ecology Progress Series, 143:239–246, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
143/m143p239.pdf.

Moksnes:1998:PPJ

[MPvM98] P. O. Moksnes, L. Pihl, and J. van Montfrans. Predation on postlarvae and juveniles of the shore crab Carcinus maenas: importance of shelter, size and cannibalism. Marine Ecology Progress Series, 166:211–225, 1998. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/166/m166p211.pdf.

Moore:1991:RSS

[MR91] D. C. Moore and G. K. Rodger. Recovery of a sewage sludge dumping ground. II. Macrobenthic community. Marine Ecology Progress Series, 75:301–308, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/75/m075p301.pdf.

Meyer-Reil:1994:MLS

[MR94] L.-A. Meyer-Reil. Microbial life in sedimentary biofilms the challenge to microbial ecologists. Marine Ecology Progress Series, 112:303–311, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/112/m112p303.pdf.

Maravelias:1997:IEO

[MR97a] C. D. Maravelias and D. G. Reid. Identifying the effects of oceanographic features and zooplankton on prespawning herring abundance using generalized additive models. Marine Ecology Progress Series, 147:1–9, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/147/m147p001.pdf.

Mateo:1997:DDS

[MR97b] M. A. Mateo and J. Romero. Detritus dynamics in the seagrass Posidonia oceanica: elements for an ecosystem carbon and nutrient budget. Marine Ecology Progress Series, 151: 43–53, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/151/m151p043.pdf.

Michinina:1997:GDM

[MR97c] S. R. Michinina and L. Rebordinos. Genetic differentiation in marine and estuarine natural populations of Crassostrea angulata. Marine Ecology Progress Series, 154:167–174, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
154/m154p167.pdf.

MaKinster:1999:PAM

[MRF^{+}99] J. G. MaKinster, J. E. Roberts, D. L. Felder, C. A. Chlan, M. Boudreaux, A. L. Bilodeau, and J. E. Neigel. PCR amplification of a middle repetitive element detects larval stone crabs (Crustacea: Decapoda: Menippidae) in estuarine plankton samples. Marine Ecology Progress Series, 188:161–168, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
188/m188p161.pdf.

Malinsky-Rushansky:1996:EDO

[MRL96] N. Z. Malinsky-Rushansky and C. Legrand. Excretion of dissolved organic carbon by phytoplankton of different sizes and subsequent bacterial uptake. Marine Ecology Progress Series, 132:249–255, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/132/m132p249.pdf.

Mayes:1997:GCR

[MRMLP97] D. F. Mayes, A. Rogerson, H. Marchant, and J. Laybourn-Parry. Growth and consumption rates of bacterivorous Antarctic naked marine amoebae. Marine Ecology Progress Series, 160:101–108, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/160/m160p101.pdf.

Mayzaud:1992:MTT

[MRMR92] P. Mayzaud, O. Roche-Mayzaud, and S. Razouls. Medium term time acclimation of feeding and digestive enzyme activity in marine copepods: influence of food concentration and copepod species. Marine Ecology Progress Series, 89: 197–212, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/89/m089p197.pdf.

Moriarty:1990:PBP

[MRP90] D. J. W. Moriarty, D. G. Roberts, and P. C. Pollard. Primary and bacterial productivity of tropical seagrass communities in the Gulf of Carpentaria, Australia. Marine Ecology Progress Series, 61:145–157, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/61/m061p145.pdf.

Meeuwig:1998:TWC

[MRP98] J. J. Meeuwig, J. B. Rasmussen, and R. H. Peters. Turbid waters and clarifying mussels: their moderation of empirical chl: nutrient relations in estuaries in Prince Edward Island, Canada. Marine Ecology Progress Series, 171:139–150, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
171/m171p139.pdf.

Morgan:1999:EPP

[MRP^{+}99] T. S. Morgan, A. D. Rogers, G. L. J. Paterson, L. E. Hawkins, and M. Sheader. Evidence for poecilogony in Pygospio elegans (Polychaeta: Spionidae). Marine Ecology Progress Series, 178:121–132, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/178/m178p121.pdf.

McGlathery:1998:TSV

[MRPC98] K. J. McGlathery, N. Risgaard-Petersen, and P. B. Christensen. Temporal and spatial variation in nitrogen fixation activity in the eelgrass Zostera marina rhizosphere. Marine Ecology Progress Series, 168:245–258, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
168/m168p245.pdf.

Moser:1996:MRH

[MRS96] M. L. Moser, S. W. Ross, and K. J. Sulak. Metabolic responses to hypoxia of Lycenchelys verrillii (wolf eelpout) and Glyptocephalus cynoglossus (witch flounder): sedentary bottom fishes of the Hatteras/Virginia Middle Slope. Marine Ecology Progress Series, 144:57–61, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/144/m144p057.pdf.

McShane:1991:RVS

[MS91] P. E. McShane and M. G. Smith. Recruitment variation in sympatric populations of Haliotis rubra (Mollusca. Gastropoda) in southeast Australian waters. Marine Ecology Progress Series, 73:203–210, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/73/m073p203.pdf.

Marchant:1993:USM

[MS93a]  H. J. Marchant and F. J. Scott.  Uptake of sub-

micrometre particles and dissolved organic materal by Antarctic choanoflagellates. Marine Ecology Progress Series, 92:59–64, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/92/m092p059.pdf.

Metaxas:1993:CSO

[MS93b] A. Metaxas and R. E. Scheibling. Community structure and organization of tidepools. Marine Ecology Progress Series, 98:187–198, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/98/m098p187.pdf.

Minchinton:1993:FSA

[MS93c] T. E. Minchinton and R. E. Scheibling. Free space availability and larval substratum selection as determinants of barnacle population structure in a developing rocky intertidal community. Marine Ecology Progress Series, 95:233–244, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
95/m095p233.pdf.

Minchinton:1993:VSP

[MS93d] T. E. Minchinton and R. E. Scheibling. Variations in sampling procedure and frequency affect estimates of recruitment of barnacles. Marine Ecology Progress Series, 99:83–88, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
99/m099p083.pdf.

Martin-Smith:1994:STD

[MS94a] K. M. Martin-Smith. Short-term dynamics of tropical macroalgal epifauna: patterns and processes in recolonisation of Sargassum fissifolium. Marine Ecology Progress Series, 110:177–185, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/110/m110p177.pdf.

Metaxas:1994:STVb

[MS94b] A. Metaxas and R. E. Scheibling. Spatial and temporal variability of tidepool hyper-benthos on a rocky shore in Nova Scotia, Canada. Marine Ecology Progress Series, 108:175– 184, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/108/m108p175.pdf.

Michael:1995:LPM

[MS95] T. Michael and C. M. Smith. Lectins probe molecular films in biofouling: characterization of early films on non-living and living surfaces. Marine Ecology Progress Series, 119: 229–236, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/119/m119p229.pdf.

Metaxas:1996:SHP

[MS96a] A. Metaxas and R. E. Scheibling. Spatial heterogeneity of phytoplankton assemblages in tidepools: effects of abiotic and biotic factors. Marine Ecology Progress Series, 130:179– 199, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/130/m130p179.pdf.

Metaxas:1996:TBR

[MS96b] A. Metaxas and R. E. Scheibling. Top-down and bottomup regulation of phytoplankton assemblages in tidepools. Marine Ecology Progress Series, 145:161–177, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
145/m145p161.pdf.

Murray:1997:WCA

[MS97] A. L. Murray and T. Spencer. On the wisdom of calculating annual material budgets in tidal wetlands. Marine Ecology Progress Series, 150:207–216, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/150/m150p207.pdf.

Melle:1998:RDC

[MS98] W. Melle and H. R. Skjoldal. Reproduction and development of Calanus finmarchicus, C. glacialis and C. hyperboreus in the Barents Sea. Marine Ecology Progress Series, 169:211–228, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/169/m169p211.pdf.

Maida:1995:ESC

[MSC95] M. Maida, P. W. Sammarco, and J. C. Coll. Effects of soft corals on scleractinian coral recruitment. I: Directional allelopathy and inhibition of settlement. Marine Ecology Progress Series, 121:191–202, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/121/m121p191.pdf.

Muricy:1996:GEE

[MSCTBE96] G. Muricy, A. M. Solé-Cava, J. P. Thorpe, and N. Boury-Esnault. Genetic evidence for extensive cryptic speciation in the subtidal sponge Plakina trilopha (Porifera: Demospongiae: Homoscleromorpha) from the Western Mediterranean. Marine Ecology Progress Series, 138:181–187, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
138/m138p181.pdf.

Moncreiff:1992:PPD

[MSD92] C. A. Moncreiff, M. J. Sullivan, and A. E. Daehnick. Primary production dynamics in seagrass beds of Mississippi Sound: the contributions of seagrass epiphytic algae, sand microflora, and phytoplankton. Marine Ecology Progress Series, 87:161–171, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/87/m087p161.pdf.

Macy:1998:EZS

[MSD98] W. K. Macy, S. J. Sutherland, and E. G. Durbin. Effects of zooplankton size and concentration and light intensity on the feeding behavior of Atlantic mackerel Scomber scombrus. Marine Ecology Progress Series, 172:89–100, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
172/m172p089.pdf.

Markager:1992:LRD

[MSJ92] S. Markager and K. Sand-Jensen. Light requirements and depth zonation of marine macroalgae. Marine Ecology Progress Series, 88:83–92, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/88/m088p083.pdf.

Mayer:1995:NPB

[MSK95] M. S. Mayer, L. Schaffner, and W. M. Kemp. Nitrification potentials of benthic macrofaunal tubes and burrow walls: effects of sediment NH^{4+} and animal irrigation behavior. Marine Ecology Progress Series, 121:157–169, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
121/m121p157.pdf.

Moranta:1998:FCS

[MSM^{+}98] J. Moranta, C. Stefanescu, E. Massutí, B. Morales-Nin, and D. Lloris. Fish community structure and depth-related trends on the continental slope of the Balearic Islands (Algerian basin, western Mediterranean). Marine Ecology Progress Series, 171:247–259, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/171/m171p247.pdf.

Mostajir:1999:EIC

[MSND^{+}99] B. Mostajir, T. Sime-Ngando, S. Demers, C. Belzile, S. Roy, M. Gosselin, J. P. Chanut, S. de Mora, J. Fauchot, F. Vidussi, and M. Levasseur. Ecological implications of changes in cell size and photosynthetic capacity of marine Prymnesiophyceae induced by ultraviolet-B radiation. Marine Ecology Progress Series, 187:89–100, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/187/m187p089.pdf.

Marconato:1997:MAF

[MSP^{+}97] A. Marconato, D. Y. Shapiro, C. W. Petersen, R. R. Warner, and T. Yoshikawa. Methodological analysis of fertilization rate in the bluehead wrasse Thalassoma bifasciatum: pair versus group spawns. Marine Ecology Progress Series, 161: 61–70, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/161/m161p061.pdf.

Metz:1995:OCF

[MSS95] C. Metz and S. B. Schnack-Schiel. Observations on carnivorous feeding in Antarctic calanoid copepods. Marine Ecology Progress Series, 129:71–75, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/129/m129p071.pdf.

Machias:1998:BDM

[MST98] A. Machias, S. Somarakis, and N. Tsimenides. Bathymetric distribution and movements of red mullet Mullus surmuletus. Marine Ecology Progress Series, 166:247–257, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
166/m166p247.pdf.

Moodley:1998:TBF

[MSvdZH98] L. Moodley, B. E. M. Schaub, G. J. van der Zwaan, and P. M. J. Herman. Tolerance of benthic Foraminifera (Protista: Sarcodina) to hydrogen sulphide. Marine Ecology Progress Series, 169:77–86, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/169/m169p077.pdf.

Malloy:1991:FGS

[MT91] K. D. Malloy and T. E. Targett. Feeding, growth and survival of juvenile summer flounder Paralichthys dentatus: experimental analysis of the effects of temperature and salinity. Marine Ecology Progress Series, 72:213–223, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
72/m072p213.pdf.

Miller:1993:SDE

[MT93] C. B. Miller and K. S. Tande. Stage duration estimation for Calanus populations, a modelling study. Marine Ecology Progress Series, 102:15–34, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/102/m102p015.pdf.

McHugh:1994:ERB

[MT94] D. McHugh and V. Tunnicliffe. Ecology and reproductive biology of the hydrothermal vent polychaete Amphisamytha galapagensis (Ampharetidae). Marine Ecology Progress Series, 106:111–120, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/106/m106p111.pdf.

Michaud:1999:SDD

[MTB^{+}99] M. Michaud, F. Thomas, S. Becheikh, A. Raibaut, J. A. Shykoff, and F. Renaud. Stage-dependent decisions in a parasitic copepod practising environmental sex determination. Marine Ecology Progress Series, 185:189–193, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
185/m185p189.pdf.

Morrow:1992:GDC

[MTP92] C. C. Morrow, J. P. Thorpe, and B. E. Picton. Genetic divergence and cryptic speciation in two morphs of the common subtidal nudibranch Doto coronata (Opisthobranchia: Dendronotacea: Dotoidae) from the northern Irish Sea. Marine Ecology Progress Series, 84:53–61, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/84/m084p053.pdf.

Monaco:1997:CET

[MU97] M. E. Monaco and R. E. Ulanowicz. Comparative ecosystem trophic structure of three U.S. mid-Atlantic estuaries. Marine Ecology Progress Series, 161:239–254, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
161/m161p239.pdf.

Maldonado:1998:MES

[MU98] M. Maldonado and M. J. Uriz. Microrefuge exploitation by subtidal encrusting sponges: patterns of settlement and postsettlement survival. Marine Ecology Progress Series, 174: 141–150, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/174/m174p141.pdf.

Maldonado:1999:EAE

[MU99] M. Maldonado and M. J. Uriz. An experimental approach to the ecological significance of microhabitat-scale movement in an encrusting sponge. Marine Ecology Progress Series, 185:239–255, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/185/m185p239.pdf.

Matsuyama:1997:TED

[MUH97] Yukihiki Matsuyama, Takuji Uchida, and Tsuneo Honjo. Toxic effects of the dinoflagellate Heterocapsa circularisquama on clearance rate of the blue mussel Mytilus galloprovincialis. Marine Ecology Progress Series, 146:73–80, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
146/m146p073.pdf.

Muller:1991:MVP

[Mül91] D. G. Müller. Marine virioplankton produced by infected Ectocarpus siliculosus (Phaeophyceae). Marine Ecology Progress Series, 76:101–102, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/76/m076p101.pdf.

Mumby:1999:BHD

[Mum99] P. J. Mumby. Bleaching and hurricane disturbances to populations of coral recruits in Belize. Marine Ecology Progress Series, 190:27–35, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/190/m190p027.pdf.

Munk:1992:FBP

[Mun92] P. Munk. Foraging behaviour and prey size spectra of larval herring Clupea harengus. Marine Ecology Progress Series, 80:149–158, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/80/m080p149.pdf.

Munk:1993:DGL

[Mun93] P. Munk. Differential growth of larval sprat Sprattus sprattus across a tidal front in the eastern North Sea. Marine Ecology Progress Series, 99:17–27, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/99/m099p017.pdf.

Moore:1996:BGC

[MV96] J. K. Moore and T. A. Villareal. Buoyancy and growth characteristics of three positively buoyant marine diatoms. Marine Ecology Progress Series, 132:203–213, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
132/m132p203.pdf.

McClelland:1998:CFW

[MV98] J. W. McClelland and I. Valiela. Changes in food web structure under the influence of increased anthropogenic nitrogen inputs to estuaries. Marine Ecology Progress Series, 168: 259–271, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/168/m168p259.pdf.

Mayzaud:1999:SVL

[MVA99] P. Mayzaud, P. Virtue, and E. Albessard. Seasonal variations in the lipid and fatty acid composition of the euphausiid Meganyctiphanes norvegica from the Ligurian Sea. Marine Ecology Progress Series, 186:199–210, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
186/m186p199.pdf.

Moens:1999:SAM

[MVdM^{+}99] T. Moens, L. Verbeeck, A. de Maeyer, J. Swings, and M. Vincx. Selective attraction of marine bacterivorous nematodes to their bacterial food. Marine Ecology Progress Series, 176:165–178, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/176/m176p165.pdf.

Marguillier:1997:TRI

[MvdVD^{+}97] S. Marguillier, G. van der Velde, F. Dehairs, M. A. Hemminga, and S. Rajagopal. Trophic relationships in an interlinked mangrove-seagrass ecosystem as traced by δ ^{13}C and δ ^{15}N. Marine Ecology Progress Series, 151:115–121, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
151/m151p115.pdf.

Moodley:1997:DRB

[MvdZH^{+}97] L. Moodley, G. J. van der Zwaan, P. M. J. Herman, L. Kempers, and P. van Breugel. Differential response of benthic meiofauna to anoxia with special reference to Foraminifera (Protista: Sarcodina). Marine Ecology Progress Series, 158: 151–163, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/158/m158p151.pdf.

MacDonald:1994:VFQ

[MW94] B. A. MacDonald and J. E. Ward. Variation in food quality and particle selectivity in the sea scallop Placopecten magellanicus (Mollusca: Bivalvia). Marine Ecology Progress Series, 108:251–264, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/108/m108p251.pdf.

Minello:1997:UNC

[MW97] T. J. Minello and J. W. Webb, Jr. Use of natural and created Spartina alterniflora salt marshes by fishery species and other aquatic fauna in Galveston Bay, Texas, USA. Marine Ecology Progress Series, 151:165–179, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/151/m151p165.pdf.

Marba:1999:GFP

[MW99] N. Marbà and D. I. Walker. Growth, flowering, and population dynamics of temperate Western Australian seagrasses. Marine Ecology Progress Series, 184:105–118, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
184/m184p105.pdf.

Madureira:1993:DBS

[MWA93] L. S. P. Madureira, P. Ward, and A. Atkinson. Differences in backscattering strength determined at 120 and 38 kHz for three species of Antarctic macroplankton. Marine Ecology Progress Series, 93:17–24, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/93/m093p017.pdf.

Muller:1999:EPC

[MWB^{+}99] Werner E. G. Müller, Matthias Wiens, Renato Batel, Renate Steffen, Heinz C. Schröder, Radovan Borojevic, and Marcio Reis Custodio. Establishment of a primary cell culture from a sponge: primmorphs from Suberites domuncula. Marine Ecology Progress Series, 178:205–219, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
178/m178p205.pdf.

Moran:1991:EDO

[MWH91] M. A. Moran, R. J. Wicks, and R. E. Hodson. Export of dissolved organic matter from a mangrove swamp ecosystem: evidence from natural fluorescence, dissolved lignir phenols, and bacterial secondary production. Marine Ecology Progress Series, 76:175–184, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/76/m076p175.pdf.

Miller-Way:1996:TOC

[MWT96] T. Miller-Way and R. R. Twilley. Theory and operation of continuous flow systems for the study of benthic-pelagic coupling. Marine Ecology Progress Series, 140:257–269, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
140/m140p257.pdf.

Maldonado:1996:EPF

[MY96] M. Maldonado and C. M. Young. Effects of physical factors on larval behavior, settlement and recruitment of four tropical demosponges. Marine Ecology Progress Series, 138: 169–180, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/138/m138p169.pdf.

Maldonado:1998:LBD

[MY98] M. Maldonado and C. M. Young. Limits on the bathymetric distribution of keratose sponges: a field test in deepwater. Marine Ecology Progress Series, 174:123–139, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
174/m174p123.pdf.

Myers:1995:RMF

[Mye95] R. A. Myers. Recruitment of marine fish: the relative roles of density-dependent and density-independent mortality in the egg, larval, and juvenile stages. Marine Ecology Progress Series, 128:308–309, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Malloy:1996:STP

[MYYT96] K. D. Malloy, Y. Yamashita, H. Yamada, and T. E. Targett. Spatial and temporal patterns of juvenile stone flounder Kareius bicoloratus growth rates during and after settlement. Marine Ecology Progress Series, 131:49–59, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
131/m131p049.pdf.

Minello:1992:UNT

[MZ92] T. J. Minello and R. J. Zimmerman. Utilization of natural and transplanted Texas salt marshes by fish and decapod crustaceans. Marine Ecology Progress Series, 90:273–285, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
90/m090p273.pdf.

Macpherson:1999:TSV

[MZ99] E. Macpherson and U. Zika. Temporal and spatial variability of settlement success and recruitment level in three blennoid fishes in the northwestern Mediterranean. Marine Ecology Progress Series, 182:269–282, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/182/m182p269.pdf.

Morales-Zamorano:1991:ETD

[MZCMOCJP91] L. A. Morales-Zamorano, R. Cajal-Medrano, E. Orellana-Cepeda, and L. C. Jimenez-Perez. Effect of tidal dynamics on a planktonic community in a coastal lagoon of Baja California, Mexico. Marine Ecology Progress Series, 78:229– 239, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/78/m078p229.pdf.

Morgan:1996:PRB

[MZFHC96] S. G. Morgan, R. K. Zimmer-Faust, K. L. Heck, Jr., and L. D. Coen. Population regulation of blue crabs Callinectes sapidus in the northern Gulf of Mexico: postlarval supply. Marine Ecology Progress Series, 133:73–88, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/133/m133p073.pdf.

Nickell:1995:FMB

[NA95] L. A. Nickell and R. J. A. Atkinson. Functional morphology of burrows and trophic modes of three thalassinidean shrimp species, and a new approach to the classification of thalassinidean burrow morphology. Marine Ecology Progress Series, 128:181–197, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/128/m128p181.pdf.

Nakaoka:1999:GSH

[NA99] M. Nakaoka and K. Aioi. Growth of seagrass Halophila ovalis at dugong trails compared to existing within-patch variation in a Thailand intertidal flat. Marine Ecology Progress Series, 184:97–103, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/184/m184p097.pdf.

Naganuma:1996:CCL

[Nag96] T. Naganuma. Calanoid copepods: linking lower-higher trophic levels by linking lower-higher Reynolds numbers. Marine Ecology Progress Series, 136:311–313, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
136/m136p311.pdf.

Nagasaki:1993:VLP

[NAI^{+}93a] K. Nagasaki, M. Ando, I. Imai, S. Itakura, and Y. Ishida. Virus-like particles in an apochlorotic flagellate in Hiroshima Bay, Japan. Marine Ecology Progress Series, 96:307–310, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/96/m096p307.pdf.

Naim:1993:SRF

[Nai93b] O. Naim. Seasonal responses of a fringing reef community to eutrophication (Reunion Island, Western Indian Ocean). Marine Ecology Progress Series, 99:137–151, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
99/m099p137.pdf.

Najdek:1996:FAC

[Naj96] M. Najdek. Fatty acid composition in amorphous aggregates in the northern Adriatic Sea. Marine Ecology Progress Series, 139:219–226, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/139/m139p219.pdf.

Najdek:1997:UCZ

[Naj97] M. Najdek. Unusual changes of zooplankton fatty acid composition in the northern Adriatic during the 1991 mucilage event. Marine Ecology Progress Series, 159:143–150, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
159/m159p143.pdf.

Nakaoka:1992:SSV

[Nak92] M. Nakaoka. Spatial and seasonal variation in growth rate and secondary production of Yoldia notabilis in Otsuchi Bay, Japan, with reference to the influence of food supply from the water column. Marine Ecology Progress Series, 88:215– 223, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/88/m088p215.pdf.

Nakaoka:1994:SDR

[Nak94] M. Nakaoka. Size-dependent reproductive traits of Yoldia notabilis (Bivalvia: Protobranchia). Marine Ecology Progress Series, 114:129–137, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/114/m114p129.pdf.

Nandakumar:1996:ITP

[Nan96] K. Nandakumar. Importance of timing of panel exposure on the competitive outcome and succession of sessile organisms. Marine Ecology Progress Series, 131:191–203, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
131/m131p191.pdf.

Nigmatullin:1995:AGR

[NAS95] C. M. Nigmatullin, A. I. Arkhipkin, and R. M. Sabirov. Age, growth and reproductive biology of diamond-shaped squid Thysanoteuthis rhombus (Oegopsida: Thysanoteuthidae). Marine Ecology Progress Series, 124:73–87, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
124/m124p073.pdf.

Nausch:1996:MAT

[Nau96] M. Nausch. Microbial activities on Trichodesmium colonies. Marine Ecology Progress Series, 141:173–181, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
141/m141p173.pdf.

Netto:1999:SMM

[NAW99] S. A. Netto, M. J. Attrill, and R. M. Warwick. Sublittoral meiofauna and macrofauna of Rocas Atoll (NE Brazil): indirect evidence of a topographically controlled front. Marine Ecology Progress Series, 179:175–186, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
179/m179p175.pdf.

Nelson:1990:KSA

[NB90] D. M. Nelson and M. A. Brzezinski. Kinetics of silicic acid uptake by natural diatom assemblages in two Gulf Stream warmcore rings. Marine Ecology Progress Series, 62:283–292, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
62/m062p283.pdf.

Nellis:1996:IPC

[NB96a] P. Nellis and E. Bourget. Influence of physical and chemical factors on settlement and recruitment of the hydroid Tubularia larynx. Marine Ecology Progress Series, 140:123– 139, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/140/m140p123.pdf.

Norkko:1996:PRC

[NB96b] A. Norkko and E. Bonsdorff. Population responses of coastal zoobenthos to stress induced by drifting algal mats. Marine Ecology Progress Series, 140:141–151, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
140/m140p141.pdf.

Norkko:1996:RZC

[NB96c] A. Norkko and E. Bonsdorff. Rapid zoobenthic community responses to accumulations of drifting algae. Marine Ecology Progress Series, 131:143–157, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/131/m131p143.pdf.

Nagelkerken:1998:DRA

[NB98] I. Nagelkerken and R. P. M. Bak. Differential regeneration of artificial lesions among sympatric morphs of the Caribbean corals Porites astreoides and Stephanocoenia michelinii. Marine Ecology Progress Series, 163:279–283, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
163/m163p279.pdf.

Nimer:1994:CDA

[NBM94] N. A. Nimer, C. Brownlee, and M. J. Merrett. Carbon dioxide availability, intracellular pH and growth rate of the coccolithophore Emiliania huxleyi. Marine Ecology Progress Series, 109:257–262, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/109/m109p257.pdf.

Nagelkerken:1997:WDC

[NBS^{+}97] I. Nagelkerken, K. Buchan, G. W. Smith, K. Bonair, P. Bush, J. Garzón-Ferreira, L. Botero, P. Gayle, C. D. Harvell, C. Heberer, K. Kim, C. Petrovic, L. Pors, and P. Yoshioka. Widespread disease in Caribbean sea fans: II. Patterns of infection and tissue loss. Marine Ecology Progress Series, 160:255–263, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/160/m160p255.pdf.

Nedwell:1994:DNT

[NBW94] D. B. Nedwell, T. H. Blackburn, and W. J. Wiebe. Dynamic nature of the turnover of organic carbon, nitrogen and sulphur in the sediments of a Jamaican mangrove forest. Marine Ecology Progress Series, 110:223–231, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/110/m110p223.pdf.

Navarrete:1990:BWM

[NC90] S. A. Navarrete and J. C. Castilla. Barnacle walls as mediators of intertidal mussel recruitment: effects of patch size on the utilization of space. Marine Ecology Progress Series, 68:113–119, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/68/m068p113.pdf.

Navarrete:1993:PNR

[NC93] S. A. Navarrete and J. C. Castilla. Predation by Norway rats in the intertidal zone of central Chile. Marine Ecology Progress Series, 92:187–199, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/92/m092p187.pdf.

Neuer:1994:PHO

[NC94] S. Neuer and T. J. Cowles. Protist herbivory in the Oregon upwelling system. Marine Ecology Progress Series, 113: 147–162, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/113/m113p147.pdf.

Neuer:1995:CCS

[NC95] S. Neuer and T. J. Cowles. Comparative cell-specific grazing rates in field populations of ciliates and dinoflagellates. Marine Ecology Progress Series, 125:259–267, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
125/m125p259.pdf.

Naranjo:1996:EES

[NCGG96] S. A. Naranjo, J. L. Carballo, and J. C. García-Gómez. Effects of environmental stress on ascidian populations in Algeciras Bay (southern Spain). Possible marine bioindicators? Marine Ecology Progress Series, 144:119–131, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
144/m144p119.pdf.

Nolan:1991:AMR

[ND91] C. Nolan and H. Dahlgaard. Accumulation of metal radiotracers by Mytilus edulis. Marine Ecology Progress Series, 70:165–174, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/70/m070p165.pdf.

Nelson:1996:SAD

[ND96] D. M. Nelson and Q. Dortch. Silicic acid depletion and silicon limitation in the plume of the Mississippi River: evidence from kinetic studies in spring and summer. Marine Ecology Progress Series, 136:163–178, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/136/m136p163.pdf.

Nestlerode:1998:EPE

[ND98] J. A. Nestlerode and R. J. Diaz. Effects of periodic environmental hypoxia on predation of a tethered polychaete, Glycera americana: implications for trophic dynamics. Marine Ecology Progress Series, 172:185–195, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/172/m172p185.pdf.

Nozais:1999:EAU

[NDG^{+}99] C. Nozais, G. Desrosiers, M. Gosselin, C. Belzile, and S. Demers. Effects of ambient UVB radiation in a meiobenthic community of a tidal mudflat. Marine Ecology Progress Series, 189:149–158, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/189/m189p149.pdf.

Nyman:1993:RBV

[NDRP93] J. A. Nyman, R. D. DeLaune, H. H. Roberts, and W. H. Patrick, Jr. Relationship between vegetation and soil formation in a rapidly submerging coastal marsh. Marine Ecology Progress Series, 96:269–279, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/96/m096p269.pdf.

Nielsen:1995:FGR

[NEIR95] A. M. Nielsen, N. T. Eriksen, J. J. L. Iversen, and H. U. Riisgård. Feeding, growth and respiration in the polychaetes Nereis diversicolor (facultative filter-feeder) and N. virens (omnivorous) — a comparative study. Marine Ecology Progress Series, 125:149–158, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/125/m125p149.pdf.

Neuer:1992:GDM

[Neu92] S. Neuer. Growth dynamics of marine Synechococcus spp.in the Gulf of Alaska. Marine Ecology Progress Series, 83:251– 262, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/83/m083p251.pdf.

Nevitt:1999:OFA

[Nev99] G. Nevitt. Olfactory foraging in Antarctic seabirds: a speciesspecific attraction to krill odors. Marine Ecology Progress Series, 177:235–241, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/177/m177p235.pdf.

Noda:1998:DSV

[NFM98] T. Noda, K. Fukushima, and T. Mori. Daily settlement variability of the barnacle Semibalanus cariosus: importance of physical factors and density-dependent processes. Marine Ecology Progress Series, 169:289–293, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
169/m169p289.pdf.

Nolan:1992:CSB

[NFT92] C. V. Nolan, S. W. Fowler, and J.-L. Teyssie. Cobalt speciation and bioavailability in marine organisms. Marine Ecology Progress Series, 88:105–116, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/88/m088p105.pdf.

Nielsen:1996:DCS

[NG96] L. P. Nielsen and R. N. Glud. Denitrification in a coastal sediment measured in situ by the nitrogen isotope pairing technique applied to a benthic flux chamber. Marine Ecology Progress Series, 137:181–186, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/137/m137p181.pdf.

Nash:1999:VSE

[NG99] R. D. M. Nash and A. J. Geffen. Variability in stage I egg production and settlement of plaice Pleuronectes platessa, on the west side of the Isle of Man, Irish Sea. Marine Ecology Progress Series, 189:241–250, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/189/m189p241.pdf.

Nejstgaard:1997:FRC

[NGS97] J. C. Nejstgaard, I. Gismervik, and P. T. Solberg. Feeding and reproduction by Calanus finmarchicus, and microzooplankton grazing during mesocosm blooms of diatoms and the coccolithophore Emiliania huxleyi. Marine Ecology Progress Series, 147:197–217, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/147/m147p197.pdf.

Nygaard:1990:UPL

[NH90] K. Nygaard and D. O. Hessen. Use of ^{14}C-protein-labelled bacteria for estimating clearance rates by heterotrophic and mixotrophic flagellates. Marine Ecology Progress Series, 68: 7–14, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/68/m068p007.pdf.

Newberger:1995:PBB

[NH95a] T. A. Newberger and E. D. Houde. Population biology of bay anchovy Anchoa mitchitli in the mid Chesapeake Bay. Marine Ecology Progress Series, 116:25–37, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/116/m116p025.pdf.

Nielsen:1995:PCS

[NH95b] T. G. Nielsen and B. Hansen. Plankton community structure and carbon cycling on the western coast of Greenland during and after the sedimentation of a diatom bloom. Marine Ecology Progress Series, 125:239–257, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/125/m125p239.pdf.

Nielsen:1998:EST

[NH98] A. Nielsen and L. Hagerman. Effects of short-term hypoxia on metabolism and haemocyanin oxygen transport in the prawns Palaemon adspersus and Palaemonetes varians. Marine Ecology Progress Series, 167:177–183, 1998. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/167/m167p177.pdf.

Noren:1999:CPC

[NHL99] F. Norén, J. Haamer, and O. Lindahl. Changes in the plankton community passing a Mytilus edulis mussel bed. Marine Ecology Progress Series, 191:187–194, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
191/m191p187.pdf.

Nielsen:1993:TEM

[Nie93] M. V. Nielsen. Toxic effect of the marine dinoflagellate Gymnodinium galatheanum on juvenile cod Gadus morhua. Marine Ecology Progress Series, 95:273–277, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
95/m095p273.pdf.

Nielsen:1996:GCC

[Nie96] M. V. Nielsen. Growth and chemical composition of the toxic dinoflagellate Gymnodinium galatheanum in relation to irradiance, temperature and salinity. Marine Ecology Progress Series, 136:205–211, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/136/m136p205.pdf.

Noda:1998:ECZ

[NIU^{+}98] M. Noda, I. Ikeda, S. Ueno, H. Hashimoto, and K. Gushima. Enrichment of coastal zooplankton communities by drifting zooplankton patches from the Kuroshio front. Marine Ecology Progress Series, 170:55–65, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/170/m170p055.pdf.

Niquil:1998:IMA

[NJLD98] N. Niquil, G. A. Jackson, L. Legendre, and B. Delesalle. Inverse model analysis of the planktonic food web of Takapoto Atoll (French Polynesia). Marine Ecology Progress Series, 165:17–29, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/165/m165p017.pdf.

Nilsson:1991:RMS

[NJSS91] P. Nilsson, B. Jonsson, I. L. Swanberg, and K. Sundback. Response of a marine shallow-water sediment system to an increased load of inorganic nutrients. Marine Ecology Progress Series, 71:275–290, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/71/m071p275.pdf.

Nagata:1990:FIR

[NK90] T. Nagata and D. L. Kirchman. Filtration-induced release of dissolved free amino acids: application to cultures of marine protozoa. Marine Ecology Progress Series, 68:1–5, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
68/m068p001.pdf.

Nagata:1992:RMO

[NK92a] T. Nagata and D. L. Kirchman. Release of macromolecular organic complexes by heterotrophic marine flagellates. Marine Ecology Progress Series, 83:233–240, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
83/m083p233.pdf.

Nommann:1992:HCP

[NK92b] S. Nömmann and E. Kaasik. Hydrodynamical control of phytoplankton succession during the vernal light-limited phase in the Baltic Sea. Marine Ecology Progress Series, 84:279– 292, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/84/m084p279.pdf.

Neundorfer:1993:NVP

[NK93] J. V. Neundorfer and W. M. Kemp. Nitrogen versus phosphorus enrichment of brackish waters: responses of the submersed plant Potamogeton perfoliatus and its associated algal community. Marine Ecology Progress Series, 94:71–82, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
94/m094p071.pdf.

Nagata:1996:BDP

[NK96] T. Nagata and D. L. Kirchman. Bacterial degradation of protein adsorbed to model submicron particles in seawater. Marine Ecology Progress Series, 132:241–248, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
132/m132p241.pdf.

Nielsen:1990:ECP

[NKB90] T. G. Nielsen, T. Kiorboe, and P. K. Bjørnsen. Effects of a Chrysochromulina polylepis subsurface bloom on the planktonic community. Marine Ecology Progress Series, 62: 21–35, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/62/m062p021.pdf.

Noda:1992:IZP

[NKGK92] M. Noda, K. Kawabata, K. Gushima, and S. Kakuda. importance of zooplankton patches in foraging ecology of the planktivorous reef fish Chromis chrysurus (Pomacentridae) at Kuchinoerabu Island, Japan. Marine Ecology Progress Series, 87:251–263, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/87/m087p251.pdf.

Niehoff:1999:HFT

[NKH^{+}99] B. Niehoff, U. Klenke, H. J. Hirche, X. Irigoien, R. Head, and R. Harris. A high frequency time series at Weathership M , Norwegian Sea, during the 1997 spring bloom: the reproductive biology of Calanus finmarchicus. Marine Ecology Progress Series, 176:81–92, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/176/m176p081.pdf.

Naganuma:1996:PPM

[NKI^{+}96] T. Naganuma, S. Konishi, T. Inoue, T. Nakane, and S. Sukizaki. Photodegradation or photoalteration? Microbial assay of the effect of UV-B on dissolved organic matter. Marine Ecology Progress Series, 135:309–310, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
135/m135p309.pdf.

Niki:1997:AMD

[NKKO97] T. Niki, M. Kunugi, K. Kohata, and A. Otsuki. Annual monitoring of DMS-producing bacteria in Tokyo Bay, Japan, in relation to DMSP. Marine Ecology Progress Series, 156: 17–24, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/156/m156p017.pdf.

Nissling:1994:VEB

[NKV94] A. Nissling, H. Kryvi, and L. Vallin. Variation in egg buoyancy of Baltic cod Gadus morhua and its implications for egg survival in prevailing conditions in the Baltic Sea. Marine Ecology Progress Series, 110:67–74, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/110/m110p067.pdf.

Netto:1994:ESD

[NL94] S. A. Netto and P. C. Lana. Effects of sediment disturbance on the structure of benthic fauna in a subtropical tidal creek of southeastern Brazil. Marine Ecology Progress Series, 106:239–247, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/106/m106p239.pdf.

Nigro:1996:ISM

[NL96] M. Nigro and C. Leonzio. Intracellular storage of mercury and selenium in different marine vertebrates. Marine Ecology Progress Series, 135:137–143, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/135/m135p137.pdf.

Nielsen:1993:SPC

[NLR^{+}93] T. G. Nielsen, B. Løkkegaard, K. Richardson, R. Bo Pedersen, and L. Hansen. Structure of plankton communities in the Dogger Bank area (North Sea) during a stratified situation. Marine Ecology Progress Series, 95:115–131, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
95/m095p115.pdf.

Neudorfer:1997:MBM

[NMR97] F. Neudörfer and L. A. Meyer-Reil. A microbial biosensor for the microscale measurement of bioavailable organic carbon in oxic sediments. Marine Ecology Progress Series, 147: 295–300, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/147/m147p295.pdf.

Noji:1993:PBC

[NNB93] Thomas T. Noji, Carola I.-M. Noji, and Klaus-Günther Barthel. Pelagic-benthic coupling during the onset of winter in a northern Norwegian fjord. Carbon flow and fate of suspended particulate matter. Marine Ecology Progress Series, 93:89–99, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/93/m093p089.pdf.

Nielsen:1995:ENR

[NNR95] K. Nielsen, L. P. Nielsen, and P. Rasmussen. Estuarine nitrogen retention independently estimated by the denitrification rate and mass balance methods: a study of Norsminde Fjord, Denmark. Marine Ecology Progress Series, 119:275–283, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
119/m119p275.pdf.

Nowicki:1990:ETA

[NO90] Barbara L. Nowicki and Candace A. Oviatt. Are estuaries traps for anthropogenic nutrients? Evidence from estuarine mesocosms. Marine Ecology Progress Series, 66:131– 146, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/66/m066p131.pdf.

Nordhausen:1992:DGL

[Nor92] W. Nordhausen. Distribution and growth of larval and adult Thy-sanoessa macrura (Euphausiacea) in the Bransfield Strait Region, Antarctica. Marine Ecology Progress Series, 83:185–196, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/83/m083p185.pdf.

Nordhausen:1994:WAD

[Nor94] W. Nordhausen. Winter abundance and distribution of Euphausia superba, E. crystallorophias, and Thysanoessa macrura in Gerlache Strait and Crystal Sound, Antarctica.

Marine Ecology Progress Series, 109:131–142, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
109/m109p131.pdf.

Norrbin:1996:TDR

[Nor96] M. F. Norrbin. Timing of diapause in relation to the onset of winter in the high-latitude copepods Pseudocalanus acuspes and Acartia longiremis. Marine Ecology Progress Series, 142:99–109, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/142/m142p099.pdf.

Novitsky:1990:ESP

[Nov90] J. A. Novitsky. Evidence for sedimenting particles as the origin of the microbial community in a coastal marine sediment. Marine Ecology Progress Series, 60:161–167, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
60/m060p161.pdf.

Nelson:1990:CLE

[NP90] N. B. Nelson and B. B. Prézelin. Chromatic light effects and physiological modeling of absorption properties of Heterocapsa pygmaea (= Glenodinium sp.). Marine Ecology Progress Series, 63:37–46, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/63/m063p037.pdf.

Nelson:1993:PLA

[NPB93] N. B. Nelson, B. B. Prezelin, and R. R. Bidigare. Phytoplankton light absorption and the package effect in California coastal waters. Marine Ecology Progress Series, 94: 217–227, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/94/m094p217.pdf.

Najdek:1994:CZL

[NPB94] M. Najdek, S. Puskaric, and A. B. Bochdansky. Contribution of zooplankton lipids to the flux of organic matter in the northern Adriatic Sea. Marine Ecology Progress Series, 111:241–249, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/111/m111p241.pdf.

Nilsson:1994:HRT

[NR94] Hans C. Nilsson and Rutger Rosenberg. Hypoxic response of two marine benthic communities. Marine Ecology Progress Series, 115:209–217, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/115/m115p209.pdf.

Nielsen:1998:TSF

[NR98] C. Nielsen and H. U. Riisgård. Tentacle structure and filterfeeding in Crisia eburnea and other cyclostomatous bryozoans, with a review of upstream-collecting mechanisms. Marine Ecology Progress Series, 168:163–186, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
168/m168p163.pdf.

Nielsen:1996:RCC

[NS96] T. G. Nielsen and M. Sabatini. Role of cyclopoid copepods Oithona spp. in North Sea plankton communities. Marine Ecology Progress Series, 139:79–93, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/139/m139p079.pdf.

Nicol:1992:GSA

[NSC^{+}92] S. Nicol, M. Stolp, T. Cochran, P. Geijsel, and J. Marshall. Growth and shrinkage of Antarctic krill Euphausia superba from the Indian Ocean sector of the Southern Ocean during summer. Marine Ecology Progress Series, 89:175–181, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
89/m089p175.pdf.

Nakamura:1995:PDH

[NSH95] Y. Nakamura, S. Suzuki, and J. Hiromi. Population dynamics of heterotrophic dinoflagellates during a Gymnodinium mikimotoi red tide in the Seto Inland Sea. Marine Ecology Progress Series, 125:269–277, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/125/m125p269.pdf.

Nakamura:1993:DPS

[NSHF93] Y. Nakamura, S. Sasaki, J. Hiromi, and K. Fukami. Dynamics of picocyanobacteria in the Seto Inland Sea (Japan) during summer. Marine Ecology Progress Series, 96:117–124, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
96/m096p117.pdf.

Nelson:1992:RSL

[NT92] D. M. Nelson and P. Treguer. Role of silicon as a limiting nutrient to Antarctic diatoms: evidence from kinetic studies in the Ross Sea ice-edge zone. Marine Ecology Progress Series, 80:255–264, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/80/m080p255.pdf.

Nandakumar:1994:ENO

[NT94] K. Nandakumar and M. Tanaka. Effects of neighboring organisms on the growth of three intertidal encrusting cheilostome bryozoans. Marine Ecology Progress Series, 114:157–163, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/114/m114p157.pdf.

Navarro:1995:SFS

[NT95] J. M. Navarro and R. J. Thompson. Seasonal fluctuations in the size spectra, biochemical composition and nutritive value of the seston available to a suspension-feeding bivalve in a subarctic environment. Marine Ecology Progress Series, 125:95–106, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/125/m125p095.pdf.

Nanninga:1996:ILF

[NT96a] H. J. Nanninga and T. Tyrrell. Importance of light for the formation of algal blooms by Emiliania huxleyi. Marine Ecology Progress Series, 136:195–203, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/136/m136p195.pdf.

Navarro:1996:PEH

[NT96b] J. M. Navarro and R. J. Thompson. Physiological energetics of the horse mussel Modiolus modiolus in a cold ocean environment. Marine Ecology Progress Series, 138:135–148, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
138/m138p135.pdf.

Nedwell:1996:NFT

[NT96c] D. B. Nedwell and M. Trimmer. Nitrogen fluxes through the upper estuary of the Great Ouse, England: the role of the bottom sediments. Marine Ecology Progress Series, 142: 273–286, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/142/m142p273.pdf.

Nandakumar:1993:ICA

[NTK93] K. Nandakumar, M. Tanaka, and T. Kikuchi. Interspecific competition among fouling organisms in Tomioka Bay, Japan. Marine Ecology Progress Series, 94:43–50, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
94/m094p043.pdf.

Naganuma:1998:ATC

[NTK98] T. Naganuma, H. Takasugi, and H. Kimura. Abundance of thraustochytrids in coastal plankton. Marine Ecology Progress Series, 162:105–110, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/162/m162p105.pdf.

Nichols:1990:RIS

[NTS90] F. H. Nichols, J. K. Thompson, and L. E. Schemel. Remarkable invasion of San Francisco Bay (California, USA), by the Asian clam Potamocorbula amurensis. II. Displacement of a former community. Marine Ecology Progress Series, 66:95–101, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/66/m066p095.pdf.

Nakamura:1991:ERT

[NU91] Y. Nakamura and T. Umemori. Encystment of the red tide flagellate Chattonella antiqua (Raphidophyceae): cyst yield in batch cultures and cyst flux in the field. Marine Ecology Progress Series, 78:273–284, 1991. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/78/m078p273.pdf.

Naas:1991:PSR

[NvA91] K. E. Naas, T. van der Meeren, and D. L. Aksnes. Plankton succession and responses to manipulations in a marine basin for larval fish rearing. Marine Ecology Progress Series, 74:161–173, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/74/m074p161.pdf.

Nagarkar:1997:CTQ

[NW97a] S. Nagarkar and G. A. Williams. Comparative techniques to quantify cyanobacteria dominated epilithic biofilms on tropical rocky shores. Marine Ecology Progress Series, 154:281– 291, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/154/m154p281.pdf.

Navarro:1997:FPC

[NW97b] J. M. Navarro and J. Widdows. Feeding physiology of Cerastoderma edule in response to a wide range of seston concentrations. Marine Ecology Progress Series, 152:175–186, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
152/m152p175.pdf.

Nissling:1997:SRS

[NW97c] A. Nissling and L. Westin. Salinity requirements for successful spawning of Baltic and Belt Sea cod and the potential for cod stock interactions in the Baltic Sea. Marine Ecology Progress Series, 152:261–271, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/152/m152p261.pdf.

Nakamura:1992:GGH

[NYH92] Y. Nakamura, Y. Yamazaki, and J. Hiromi. Growth and grazing of a heterotrophic dinoflagellate, Gyrodinium dominans, feeding on a red tide flagellate, Chattonella antiqua. Marine Ecology Progress Series, 82:275–279, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
82/m082p275.pdf.

Opsahl:1993:DSB

[OB93] S. Opsahl and R. Benner. Decomposition of senescent blades of the seagrass Halodule wrightii in a subtropical lagoon. Marine Ecology Progress Series, 94:191–205, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
94/m094p191.pdf.

OBrien:1994:OCD

[O’B94a] C. J. O’Brien. Ontogenetic changes in the diet of juvenile brown tiger prawns Penaeus esculentus. Marine Ecology Progress Series, 112:195–200, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/112/m112p195.pdf.

OBrien:1994:PDJ

[O’B94b] C. J. O’Brien. Population dynamics of juvenile tiger prawns Penaeus esculentus in south Queensland, Australia. Marine Ecology Progress Series, 104:247–256, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
104/m104p247.pdf.

Ormond:1994:EIM

[OB94c] R. F. G. Ormond and S. A. Banaimoon. Ecology of intertidal macroalgal assemblages on the Hadramout coast of southern Yemen, an area of seasonal upwelling. Marine Ecology Progress Series, 105:105–120, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/105/m105p105.pdf.

Oren:1997:ELS

[OBL97] U. Oren, Y. Benayahu, and Y. Loya. Effect of lesion size and shape on regeneration of the Red Sea coral Favia favus. Marine Ecology Progress Series, 146:101–107, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
146/m146p101.pdf.

OConnor:1990:FTM

[O’C90] N. J. O’Connor. Flexibility in timing of the metamorphic molt by fiddler crab megalopae Uca pugilator. Marine Ecology Progress Series, 68:243–247, 1990. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/68/m068p243.pdf.

OConnor:1993:SRF

[O’C93a] N. J. O’Connor. Settlement and recruitment of the fiddler crabs Uca pugnax and U. pugilator in a North Carolina, USA; salt marsh. Marine Ecology Progress Series, 93:227–234, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
93/m093p227.pdf.

Otway:1993:EHS

[OC93b] N. M. Otway and J. R. Craig. Effects of hook size on the catches of undersized snapper Pagrus auratus. Marine Ecology Progress Series, 93:9–15, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/93/m093p009.pdf.

Ojeda:1995:DMA

[OC95] F. P. Ojeda and C. W. Caceres. Digestive mechanisms in Aplodactylus punctatus (Valenciennes): a temperate marine herbivorous fish. Marine Ecology Progress Series, 118:37–42, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/118/m118p037.pdf.

Osgood:1997:SVD

[OC97] K. E. Osgood and D. M. Checkley, Jr. Seasonal variations in a deep aggregation of Calanus pacificus in the Santa Barbara Basin. Marine Ecology Progress Series, 148:59–69, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
148/m148p059.pdf.

Olla:1990:EPF

[OD90] Bori L. Olla and Michael W. Davis. Effects of physical factors on the vertical distribution of larval walleye pollock Theragra chalcogramma under controlled laboratory conditions. Marine Ecology Progress Series, 63:105–112, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
63/m063p105.pdf.

ODriscoll:1998:DSP

[O’D98] R. L. O’Driscoll. Description of spatial pattern in seabird distributions along line transects using neighbour K statistics. Marine Ecology Progress Series, 165:81–94, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
165/m165p081.pdf.

Oviatt:1995:ELE

[ODN^{+}95] C. Oviatt, P. Doering, B. Nowicki, L. Reed, J. Cole, and J. Frithsen. An ecosystem level experiment on nutrient limitation in temperate coastal marine environments. Marine Ecology Progress Series, 116:171–179, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/116/m116p171.pdf.

Olu:1996:SDC

[ODS^{+}96] K. Olu, A. Duperret, M. Sibuet, J. P. Foucher, and A. Fiala-Médioni. Structure and distribution of cold seep communities along the Peruvian active margin: relationship to geological and fluid patterns. Marine Ecology Progress Series, 132:109–125, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/132/m132p109.pdf.

Olafsson:1991:EBD

[OE91] E. Olafsson and R. Elmgren. Effects of biological disturbance by benthic amphipods Monoporeia affinis on meiobenthic community structure: a laboratory approach. Marine Ecology Progress Series, 74:99–107, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/74/m074p099.pdf.

Oeschger:1990:LTA

[Oes90] R. Oeschger. Long-term anaerobiosis in sublittoral marine invertebrates from the Western Baltic Sea: Halicryptus spinulosus (Priapulida), Astarte borealis and Arctica islandica (Bivalvia). Marine Ecology Progress Series, 59:133–143, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
59/m059p133.pdf.

Osgood:1994:ODV

[OF94] K. E. Osgood and B. W. Frost. Ontogenetic diel vertical migration behaviors of the marine planktonic copepods Calanus pacificus and Metridia lucens. Marine Ecology Progress Series, 104:13–25, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/104/m104p013.pdf.

Olesen:1994:PDG

[OFR94] N. J. Olesen, K. Frandsen, and H. U. Riisgård. Population dynamics, growth and energetics of jellyfish Aurelia aurita in a shallow fjord. Marine Ecology Progress Series, 105:9–18, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
105/m105p009.pdf.

Olsgard:1995:CAE

[OG95] F. Olsgard and J. S. Gray. A comprehensive analysis of the effects of offshore oil and gas exploration and production on the benthic communities of the Norwegian continental shelf. Marine Ecology Progress Series, 122:277–306, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
122/m122p277.pdf.

Obernosterer:1995:PER

[OH95] I. Obernosterer and G. J. Herndl. Phytoplankton extracellular release and bacterial growth: dependence on the inorganic n: P ratio. Marine Ecology Progress Series, 116:247–257, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
116/m116p247.pdf.

Ohman:1996:FSC

[Ohm96] M. D. Ohman. Freezing and storage of copepod samples for the analysis of lipids. Marine Ecology Progress Series, 130:295–298, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/130/m130p295.pdf.

Otake:1994:DCO

[OINN94] T. Otake, T. Ishii, M. Nakahara, and R. Nakamura. Drastic changes in otolith strontium/calcium ratios in leptocephali and glass eels of Japanese eel Anguilia /aponica. Marine Ecology Progress Series, 112:189–193, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/112/m112p189.pdf.

Ohtsuka:1995:DOG

[OIO^{+}95] S. Ohtsuka, H. Inagaki, T. Onbe, K. Gushima, and Y. H. Yoon. Direct observations of groups of mysids in shallow coastal waters of western Japan and southern Korea. Marine Ecology Progress Series, 123:33–44, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/123/m123p033.pdf.

OConnor:1997:FTM

[OJ97] N. J. O’Connor and M. L. Judge. Flexibility in timing of molting of fiddler crab megalopae: evidence from in situ manipulation of cues. Marine Ecology Progress Series, 146:55–60, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
146/m146p055.pdf.

OConnor:1999:CSM

[OJ99] N. J. O’Connor and M. L. Judge. Cues in salt marshes stimulate molting of fiddler crab Uca pugnax megalopae: more evidence from field experiments. Marine Ecology Progress Series, 181:131–139, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/181/m181p131.pdf.

Oro:1996:ITA

[OJR96] D. Oro, L. Jover, and X. Ruiz. Influence of trawling activity on the breeding ecology of a threatened seabird, Audouin’s gull Larus audouinii. Marine Ecology Progress Series, 139:19–29, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/139/m139p019.pdf.

Onsrud:1998:DVM

[OK98] M. S. R. Onsrud and S. Kaartvedt. Diel vertical migration of the krill Meganyctiphanes norvegica in relation to physical environment, food and predators. Marine Ecology Progress Series, 171:209–219, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/171/m171p209.pdf.

Oren:1995:EMM

[OKK95] A. Oren, M. Kühl, and U. Karsten. An endoevaporitic microbial mat within a gypsum crust: zonation of phototrophs, photopigments, and light penetration. Marine Ecology Progress Series, 128:151–159, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/128/m128p151.pdf.

Osinga:1996:PFD

[OKL^{+}96] R. Osinga, R. L. J. Kwint, W. E. Lewis, G. W. Kraay, J. D. Lont, H. J. Lindeboom, and F. C. van Duyl. Production and fate of dimethylsulfide and dimethylsulfoniopropionate in pelagic mesocosms: the role of sedimentation. Marine Ecology Progress Series, 131:275–286, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/131/m131p275.pdf.

Oldham:1999:PNF

[OL99] C. E. Oldham and P. S. Lavery. Porewater nutrient fluxes in a shallow fetch-limited estuary. Marine Ecology Progress Series, 183:39–47, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/183/m183p039.pdf.

Olesen:1995:CPJ

[Ole95] N. J. Olesen. Clearance potential of jellyfish Aurelia aurita, and predation impact on zooplankton in a shallow cove. Marine Ecology Progress Series, 124:63–72, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
124/m124p063.pdf.

Olesen:1996:RLA

[Ole96] B. Olesen. Regulation of light attenuation and eelgrass Zostera marina depth distribution in a Danish embayment. Marine Ecology Progress Series, 134:187–194, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
134/m134p187.pdf.

Olmi:1994:VMB

[Olm94] E. J. Olmi III. Vertical migration of blue crab Callinectes sapidus megalopae: implications for transport in estuaries. Marine Ecology Progress Series, 113:39–54, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
113/m113p039.pdf.

Olsgard:1993:DTA

[Ols93] F. Olsgard. Do toxic algal blooms affect subtidal soft-bottom communities? Marine Ecology Progress Series, 102:269– 286, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/102/m102p269.pdf.

Olafsson:1990:CMA

[OM90] E. Olafsson and C. G. Moore. Control of meiobenthic abundance by macroepifauna in a subtidal muddy habitat. Marine Ecology Progress Series, 65:241–249, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
65/m065p241.pdf.

Olafsson:1992:EMD

[ÓM92] E. Ólafsson and C. G. Moore. Effects of macroepifauna on developing nematode and harpacticoid assemblages in a subtidal muddy habitat. Marine Ecology Progress Series, 84: 161–171, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/84/m084p161.pdf.

Ojeda:1999:FSH

[OM99a] F. P. Ojeda and A. A. Muñoz. Feeding selectivity of the herbivorous fish Scartichthys viridis: effects on macroalgal community structure in a temperate rocky intertidal coastal zone. Marine Ecology Progress Series, 184:219–229, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
184/m184p219.pdf.

Overholtzer:1999:CRU

[OM99b] K. L. Overholtzer and P. J. Motta. Comparative resource use by juvenile parrotfishes in the Florida Keys. Marine Ecology Progress Series, 177:177–187, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/177/m177p177.pdf.

Oxenford:1995:DRAa

[OMH95a] H. A. Oxenford, R. Mahon, and W. Hunte. Distribution and relative abundance of flyingfish (Exocoetidae) in the eastern Caribbean. I. Adults. Marine Ecology Progress Series, 117: 11–23, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/117/m117p011.pdf.

Oxenford:1995:DRAb

[OMH95b] H. A. Oxenford, R. Mahon, and W. Hunte. Distribution and relative abundance of flyingfish (Exocoetidae) in the eastern Caribbean. III. Juveniles. Marine Ecology Progress Series, 117:39–47, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/117/m117p039.pdf.

Olafsson:1999:SSU

[ÓMvdB99] E. Ólafsson, H. Modig, and W. J. van de Bund. Species specific uptake of radio-labelled phytodetritus by benthic meiofauna from the Baltic Sea. Marine Ecology Progress Series, 177:63–72, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/177/m177p063.pdf.

Olafsson:1997:IMC

[ÓN97] E. Ólafsson and S. G. M. Ndaro. Impact of the mangrove crabs Uca annulipes and Dotilla fenestrata on meiobenthos. Marine Ecology Progress Series, 158:225–231, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
158/m158p225.pdf.

Ogilvie:1997:HNM

[ONH^{+}97] B. Ogilvie, D. B. Nedwell, R. M. Harrison, A. Robinson, and A. Sage. High nitrate, muddy estuaries as nitrogen sinks: the nitrogen budget of the River Colne estuary (United Kingdom). Marine Ecology Progress Series, 150:217–228, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
150/m150p217.pdf.

Otake:1993:DPO

[ONM93] T. Otake, K. Nogami, and K. Maruyama. Dissolved and particulate organic matter as possible food sources for eel leptocephali. Marine Ecology Progress Series, 92:27–34, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
92/m092p027.pdf.

Oug:1998:RBS

[ONR98] E. Oug, K. Næs, and B. Rygg. Relationship between soft bottom macrofauna and polycyclic aromatic hydrocarbons (PAH) from smelter discharge in Norwegian fjords and coastal waters. Marine Ecology Progress Series, 173:39–52, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
173/m173p039.pdf.

Onuf:1996:SRL

[Onu96] C. P. Onuf. Seagrass responses to long-term light reduction by brown tide in upper Laguna Madre, Texas: distribution and biomass patterns. Marine Ecology Progress Series, 138:219–231, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/138/m138p219.pdf.

Oro:1999:RAG

[OP99] D. Oro and R. Pradel. Recruitment of Audouin’s gull to the Ebro Delta colony at metapopulation level in the western Mediterranean. Marine Ecology Progress Series, 180: 267–273, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/180/m180p267.pdf.

Olesen:1996:FGE

[OPS96] N. J. Olesen, J. E. Purcell, and D. K. Stoecker. Feeding and growth by ephyrae of scyphomedusae Chrysaora quinquecirrha. Marine Ecology Progress Series, 137:149–159, 1996.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
137/m137p149.pdf.

Owens:1995:ECB

[OR95] L. Owens and P. C. Rothlisberg. Epidemiology of cryptonisci (Bopyridae: Isopoda) in the Gulf of Carpentaria, Australia. Marine Ecology Progress Series, 122:159–164, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
122/m122p159.pdf.

Olive:1998:IPT

[ORD98] P. J. W. Olive, S. W. Rees, and A. Djunaedi. Influence of photoperiod and temperature on oocyte growth in the semelparous polychaete Nereis (Neanthes) virens. Marine Ecology Progress Series, 172:169–183, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/172/m172p169.pdf.

Oresland:1990:FPI

[Øre90] V. Øresland. Feeding and predation impact of the chaetognath Eukrohnia hamata in Gerlache Strait, Antarctic Peninsula. Marine Ecology Progress Series, 63:201–209, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
63/m063p201.pdf.

Oresland:1991:FCC

[Øre91] V. Øresland. Feeding of the carnivorous copepod Euchaeta antarctica in Antarctic waters. Marine Ecology Progress Series, 78:41–47, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/78/m078p041.pdf.

Oresland:1995:WPS

[Øre95] V. Øresland. Winter population structure and feeding of the chaetognath Eukrohnia hamata and the copepod Euchaeta antarctica in Gerlache Strait, Antarctic Peninsula. Marine Ecology Progress Series, 119:77–86, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/119/m119p077.pdf.

Oren:1997:OIC

[ORL97] U. Oren, B. Rinkevich, and Y. Loya. Oriented intra-colonial transport of ^{14}C labeled materials during coral regeneration. Marine Ecology Progress Series, 161:117–122, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
161/m161p117.pdf.

Oro:1996:ETD

[Oro96] D. Oro. Effects of trawler discard availability on egg laying and breeding success in the lesser black-backed gull Larus fuscus in the western Mediterranean. Marine Ecology Progress Series, 132:43–46, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/132/m132p043.pdf.

Osborn:1996:ETP

[OS96] T. Osborn and A. Scotti. Effect of turbulence on predator– prey contact rates: where do we go from here? Marine Ecology Progress Series, 139:302–304, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/139/m139p301.pdf.

Olsgard:1997:RBT

[OSC97] F. Olsgard, P. J. Somerfield, and M. R. Carr. Relationships between taxonomic resolution and data transformations in analyses of a macrobenthic community along an established pollution gradient. Marine Ecology Progress Series, 149:173–181, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/149/m149p173.pdf.

Olsgard:1998:RBT

[OSC98] F. Olsgard, P. J. Somerfield, and M. R. Carr. Relationships between taxonomic resolution, macrobenthic community patterns and disturbance. Marine Ecology Progress Series, 172:25–36, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/172/m172p025.pdf.

Omnes:1996:EDN

[OSGB96] P. Omnes, G. Slawyk, N. Garcia, and P. Bonin. Evidence of denitrification and nitrate ammonification in the River Rhone plume (northwestern Mediterranean Sea). Marine Ecology Progress Series, 141:275–281, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/141/m141p275.pdf.

Olesen:1993:SAE

[OSJ93] B. Olesen and K. Sand-Jensen. Seasonal acclimatization of eelgrass Zostera marina growth to light. Marine Ecology Progress Series, 94:91–99, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/94/m094p091.pdf.

Olesen:1994:BDP

[OSJ94a] B. Olesen and K. Sand-Jensen. Biomass-density patterns in the temperate seagrass Zostera marina. Marine Ecology Progress Series, 109:283–291, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/109/m109p283.pdf.

Olesen:1994:PDE

[OSJ94b] B. Olesen and K. Sand-Jensen. Patch dynamics of eelgrass Zostera marina. Marine Ecology Progress Series, 106:147– 156, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/106/m106p147.pdf.

Osgood:1995:SPC

[OSZ95] D. T. Osgood, M. C. F. V. Santos, and J. C. Zieman. Sediment physico-chemistry associated with natural marsh development on a storm-deposited sand flat. Marine Ecology Progress Series, 120:271–283, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/120/m120p271.pdf.

Ouellet:1995:EGL

[OTF95] P. Ouellet, C. T. Taggart, and K. T. Frank. Early growth, lipid composition, and survival expectations of shrimp Pandalus borealis larvae in the northern Gulf of St. Lawrence. Marine Ecology Progress Series, 126:163–175, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
126/m126p163.pdf.

Oeschger:1992:SDT

[OV92] R. Oeschger and R. D. Vetter. Sulfide detoxification and tolerance in Halicryptus spinulosus (Priapulida): a multiple strategy. Marine Ecology Progress Series, 86:167–179, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
86/m086p167.pdf.

Oresland:1993:SWD

[ØW93] V. Øresland and P. Ward. Summer and winter diet of four carnivorous copepod species around South Georgia. Marine Ecology Progress Series, 98:73–78, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/98/m098p073.pdf.

Osman:1995:PEO

[OW95] R. W. Osman and R. B. Whitlatch. Predation on early ontogenetic life stages and its effect on recruitment into a marine epifaunal community. Marine Ecology Progress Series, 117:111–126, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/117/m117p111.pdf.

Osman:1992:PRM

[OWM92] R. W. Osman, R. B. Whitlatch, and R. J. Malatesta. Potential role of micropredators in determining recruitment into a marine community. Marine Ecology Progress Series, 83: 35–43, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/83/m083p035.pdf.

Passow:1994:DSB

[PA94] U. Passow and A. L. Alldredge. Distribution, size and bacterial colonization of transparent exopolymer particles (TEP) in the ocean. Marine Ecology Progress Series, 113:185–198, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/113/m113p185.pdf.

Pakulski:1992:FRS

[Pak92] J. D. Pakulski. Foliar release of soluble reactive phosphorus from Spartina alterniflora in a Georgia (USA) salt marsh. Marine Ecology Progress Series, 90:53–60, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
90/m090p053.pdf.

Pakhomov:1995:DSA

[Pak95] E. A. Pakhomov. Demographic studies of Antarctic krill Euphausia superba in the Cooperation and Cosmonaut Seas (Indian sector of the Southern Ocean). Marine Ecology Progress Series, 119:45–61, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/119/m119p045.pdf.

Palomera:1991:SAE

[Pal91] I. Palomera. Spawning of anchovy Engraulis encrasicolus in the Northwestern Mediterranean relative to hydrographic features in the region. Marine Ecology Progress Series, 79: 215–223, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/79/m079p215.pdf.

Palmer:1997:CLR

[Pal97] J. D. Palmer. Crab larval-release rhythms. Marine Ecology Progress Series, 158:307–308, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/158/m158p307.pdf.

Procaccini:1996:GSS

[PAM96] G. Procaccini, R. S. Alberte, and L. Mazzella. Genetic structure of the seagrass Posidonia oceanica in the Western Mediterranean: ecological implications. Marine Ecology Progress Series, 140:153–160, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/140/m140p153.pdf.

Pinn:1998:PSS

[PAR98] E. H. Pinn, R. J. A. Atkinson, and A. Rogerson. Particle size selectivity and resource partitioning in five species of Thalassinidea (Crustacea: Decapoda). Marine Ecology Progress Series, 169:243–250, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/169/m169p243.pdf.

Pedersen:1992:NDE

[PB92] M. F. Pedersen and J. Borum. Nitrogen dynamics of eelgrass Zostera marina during a late summer period of high growth and low nutrient availability. Marine Ecology Progress Series, 80:65–73, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/80/m080p065.pdf.

Pedersen:1993:ANB

[PB93] M. F. Pedersen and J. Borum. An annual nitrogen budget for a seagrass Zostera marina population. Marine Ecology Progress Series, 101:169–177, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/101/m101p169.pdf.

Peterson:1994:ECW

[PB94] C. H. Peterson and R. Black. An experimentalist’s challenge: when artifacts of intervention interact with treatments. Marine Ecology Progress Series, 111:289–297, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
111/m111p289.pdf.

Pearson:1996:REF

[PB96a] G. A. Pearson and S. H. Brawley. Reproductive ecology, of Fucus distichus (Phaeophyceae): an intertidal alga with successful external fertilization. Marine Ecology Progress Series, 143:211–223, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/143/m143p211.pdf.

Pedersen:1996:NCA

[PB96b] M. F. Pedersen and J. Borum. Nutrient control of algal growth in estuarine waters. Nutrient limitation and the importance of nitrogen requirements and nitrogen storage among phytoplankton and species of macroalgae. Marine Ecology Progress Series, 142:261–272, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
142/m142p261.pdf.

Pedersen:1997:NCE

[PB97] M. F. Pedersen and J. Borum. Nutrient control of estuarine macroalgae: growth strategy and the balance between nitrogen requirements and uptake. Marine Ecology Progress Series, 161:155–163, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/161/m161p155.pdf.

Puvanendran:1998:ELI

[PB98] V. Puvanendran and J. A. Brown. Effect of light intensity on the foraging and growth of Atlantic cod larvae: interpopulation difference? Marine Ecology Progress Series, 167:207–214, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/167/m167p207.pdf.

Pakulski:1995:CMN

[PBA^{+}95] J. D. Pakulski, P. Benner, R. Amon, B. Eadie, and T. Whitledge. Community metabolism and nutrient cycling in the Mississippi River plume: evidence for intense nitrification at intermediate salinities. Marine Ecology Progress Series, 117:207–218, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/117/m117p207.pdf.

Pedersen:1998:ODR

[PBDF98] O. Pedersen, J. Borum, C. M. Duarte, and M. D. Fortes. Oxygen dynamics in the rhizosphere of Cymodocea rotundata. Marine Ecology Progress Series, 169:283–288, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
169/m169p283.pdf.

Pohle:1991:ECA

[PBGE91] David G. Pohle, V. Monica Bricelj, and Zaúl García-Esquivel. The eelgrass canopy an above-bottom refuge from benthic predators for juvenile bay scallops Argopecten irradians. Marine Ecology Progress Series, 74:47–59, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
74/m074p047.pdf.

Pitcher:1998:SDP

[PBHMI98] G. C. Pitcher, A. J. Boyd, D. A. Horstman, and B. A. Mitchell-Innes. Subsurface dinoflagellate populations, frontal blooms and the formation of red tide in the southern Benguela upwelling system. Marine Ecology Progress Series, 172:253– 264, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/172/m172p253.pdf.

Piepenburg:1995:PBC

[PBv^{+}95] D. Piepenburg, T. H. Blackburn, C. F. von Dorrien, J. Gutt, P. O. J. Hall, S. Hulth, M. A. Kendall, K. W. Opalinski, E. Rachor, and M. K. Schmid. Partitioning of benthic community respiration in the Arctic (northwestern Barents Sea). Marine Ecology Progress Series, 118:199–213, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
118/m118p199.pdf.

Peck:1999:CRF

[PBVB99] L. S. Peck, S. Brockington, S. Vanhove, and M. Beghyn. Community recovery following catastrophic iceberg impacts in a soft-sediment shallow-water site at Signy Island, Antarctica. Marine Ecology Progress Series, 186:1–8, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
186/m186p001.pdf.

Purcell:1995:PSC

[PC95] J. E. Purcell and J. H. Cowan, Jr. Predation by the scyphomedusan Chrysaora quinquecirrha on Mnemiopsis leidyi ctenophores. Marine Ecology Progress Series, 129:63–70, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/129/m129p063.pdf.

Pace:1996:IDA

[PC96] M. C. Pace and K. R. Carman. Interspecific differences among meiobenthic copepods in the use of microalgal food resources. Marine Ecology Progress Series, 143:77–86, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
143/m143p077.pdf.

Peyrot-Clausade:1990:DPM

[PCB90] M. Peyrot-Clausade and J.-F. Brunei. Distribution patterns of macroboring organisms on Tulear reef flats (SW Madagascar). Marine Ecology Progress Series, 61:133–144, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
61/m061p133.pdf.

Perez-Camacho:1994:EAF

[PCBA94] A. Pérez-Camacho, R. Belras, and M. Albentosa. Effects of algal food concentration and body size on the ingestion rates of Ruditapes decussatus (Bivalvia) veliger larvae. Marine Ecology Progress Series, 115:87–92, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/115/m115p087.pdf.

Pichard:1997:ARB

[PCC^{+}97] S. L. Pichard, L. Campbell, K. Carder, J. B. Kang, J. Patch, F. R. Tabita, and J. H. Paul. Analysis of ribulose bisphosphate carboxylase gene expression in natural phytoplankton communities by group-specific gene probing. Marine Ecology Progress Series, 149:239–253, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/149/m149p239.pdf.

Pierce:1997:BWV

[PCCG97] R. W. Pierce, J. T. Carlton, D. A. Carlton, and J. B. Geller. Ballast water as a vector for tintinnid transport. Marine Ecology Progress Series, 149:295–297, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/149/m149p295.pdf.

Peters:1996:PIP

[PCG96] F. Peters, J. W. Choi, and T. Gross. Paraphysomonas imperforata (Protista, Chrysomonadida) under different turbulence levels: feeding, physiology and energetics. Marine Ecology Progress Series, 134:235–245, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/134/m134p235.pdf.

Pichard:1996:RRB

[PCK^{+}96] S. L. Pichard, L. Campbell, J. B. Kang, F. R. Tabita, and J. H. Paul. Regulation of ribulose bisphosphate carboxylase gene expression in natural phytoplankton communities. I. Diel rhythms. Marine Ecology Progress Series, 139:257–265, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
139/m139p257.pdf.

Pavia:1997:EUB

[PCLÅ97] H. Pavia, G. Cervin, A. Lindgren, and P. Åberg. Effects of UV-B radiation and simulated herbivory on phlorotannins in the brown alga Ascophyllum nodosum. Marine Ecology Progress Series, 157:139–146, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/157/m157p139.pdf.

Palsboll:1995:DMH

[PCM^{+}95] P. J. Palsbøll, P. J. Clapham, D. K. Mattila, F. Larsen, R. Sears, H. R. Siegismund, J. Sigurjonsson, O. Vasquez, and P. Arctander. Distribution of mtDNA haplotypes in North Atlantic humpback whales: the influence of behaviour on population structure. Marine Ecology Progress Series, 116: 1–10, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/116/m116p001.pdf.

Phillips:1997:ICV

[PCT^{+}97] R. A. Phillips, P. Catry, D. R. Thompson, K. C. Hamer, and R. W. Furness. Inter-colony variation in diet and reproductive performance of great skuas Catharacta skua. Marine Ecology Progress Series, 152:285–293, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/152/m152p285.pdf.

Pawlik:1995:DCS

[PCTF95] J. R. Pawlik, B. Chanas, R. J. Toonen, and W. Fenical. Defenses of Caribbean sponges against predatory reef fish. I. Chemical deterrency. Marine Ecology Progress Series, 127: 183–194, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/127/m127p183.pdf.

Preston:1990:CSP

[PD90] N. P. Preston and P. J. Doherty. Cross-shelf patterns in the community structure of coral-dwelling crustacea in the central region of the Great Barrier Reef I. Agile shrimps. Marine Ecology Progress Series, 66:47–61, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/66/m066p047.pdf.

Pezeshki:1993:ESH

[PD93] S. R. Pezeshki and R. D. DeLaune. Effects of soil hypoxia and salinity on gas exchange and growth of Spartina patens. Marine Ecology Progress Series, 96:75–81, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
96/m096p075.pdf.

Preston:1994:CSP

[PD94] N. P. Preston and P. J. Doherty. Cross-shelf patterns in the community structure of coral-dwelling crustacea in the central region of the Great Barrier Reef. II. Cryptofauna. Marine Ecology Progress Series, 104:27–38, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/104/m104p027.pdf.

Pinca:1995:MMC

[PD95] S. Pinca and S. Dallot. Meso- and macrozooplankton composition patterns related to hydrodynamic structures in the Ligurian Sea (Trophos-2 experiment, April–June 1986). Marine Ecology Progress Series, 126:49–65, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/126/m126p049.pdf.

Pond:1997:OFA

[PDB^{+}97] D. W. Pond, D. R. Dixon, M. V. Bell, A. E. Fallick, and J. R. Sargent. Occurrence of 16:2(n-4) and 18:2(n-4) fatty acids in the lipids of the hydrothermal vent shrimps Rimicaris exoculata and Alvinocaris markensis: nutritional and trophic implications. Marine Ecology Progress Series, 156:167–174, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
156/m156p167.pdf.

Perissinotto:1990:VSC

[PDBA90] R. Perissinotto, C. M. Duncombe Rae, B. P. Boden, and B. R. Allanson. Vertical stability as a controlling factor of the marine phytoplankton production at the Prince Edward Archipelago (Southern Ocean). Marine Ecology Progress Series, 60:205–209, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/60/m060p205.pdf.

Pedersen:1997:RCO

[PDC97] M. F. Pedersen, C. M. Duarte, and J. Cebrián. Rate of changes in organic matter and nutrient stocks during seagrass Cymodocea nodosa colonization and stand development. Marine Ecology Progress Series, 159:29–36, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
159/m159p029.pdf.

Page:1999:EOO

[PDD^{+}99] H. M. Page, J. E. Dugan, D. S. Dugan, J. B. Richards, and D. M. Hubbard. Effects of an offshore oil platform on the distribution and abundance of commercially important crab species. Marine Ecology Progress Series, 185:47–57, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
185/m185p047.pdf.

Perez:1997:POC

[PDF97] M. T. Pérez, J. R. Dolan, and E. Fukai. Planktonic oligotrich ciliates in the NW Mediterranean: growth rates and consumption by copepods. Marine Ecology Progress Series, 155: 89–101, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/155/m155p089.pdf.

Pace:1996:MDR

[PdG96] Michael L. Pace and Paul del Giorgio. In memoriam: Dr Robert H. Peters (1946–1996). Marine Ecology Progress Series, 142:1, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/142/m142p001.pdf.

Priscu:1990:DAO

[PDP^{+}90] J. C. Priscu, M. T. Downes, L. R. Priscu, A. C. Palmisano, and C. W. Sullivan. Dynamics of ammonium oxidizer activity and nitrous oxide (N_{2}0) within and beneath Antarctic sea ice. Marine Ecology Progress Series, 62:37–46, 1990.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
62/m062p037.pdf.

Pond:1997:WER

[PDS97] D. Pond, D. Dixon, and J. Sargent. Wax-ester reserves facilitate dispersal of hydrothermal vent shrimps. Marine Ecology Progress Series, 146:289–290, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/146/m146p289.pdf.

Petersen:1999:PWG

[PE99] B. Petersen and K. M. Exo. Predation of waders and gulls on Lanice conchilega tidal flats in the Wadden Sea. Marine Ecology Progress Series, 178:229–240, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/178/m178p229.pdf.

Pechenik:1999:ADL

[Pec99] J. A. Pechenik. On the advantages and disadvantages of larval stages in benthic marine invertebrate life cycles. Marine Ecology Progress Series, 177:269–297, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/177/m177p269.pdf.

Peletier:1996:LTC

[Pel96] H. Peletier. Long-term changes in intertidal estuarine diatom assemblages related to reduced input of organic waste. Marine Ecology Progress Series, 137:265–271, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
137/m137p265.pdf.

Pennings:1990:PPI

[Pen90a] S. C. Pennings. Predator–prey interactions in opisthobranch gastropods: effects of prey body size and habitat complexity. Marine Ecology Progress Series, 62:95–101, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
62/m062p095.pdf.

Pennington:1990:PHH

[Pen90b] J. T. Pennington. Predation by hydromedusae on hydrozoan embryos and larvae: planktonic kin selection? Marine Ecology Progress Series, 60:247–252, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/60/m060p247.pdf.

Perez:1990:RBS

[Per90] O. S. Perez. Reproductive biology of the sandy shore crab Matuta lunans (Brachyura: Calappidae). Marine Ecology Progress Series, 59:83–89, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/59/m059p083.pdf.

Perissinotto:1991:MSS

[Per91] R. Perissinotto. Mesozooplankton size-selectivity and grazing impact on the phytoplankton community of the Prince Edward Archipelago (Southern Ocean). Marine Ecology Progress Series, 79:243–258, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/79/m079p243.pdf.

Pezeshki:1991:PDS

[Pez91] S. R. Pezeshki. Population differentiation in Spartina patens: gas-exchange response to salinity. Marine Ecology Progress Series, 72:125–130, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/72/m072p125.pdf.

Patterson:1990:MML

[PF90] D. J. Patterson and T. Fenchel. Massisteria marina Larsen & Patterson 1990, a widespread and abundant bacterivorous protist associated with marine detritus. Marine Ecology Progress Series, 62:11–19, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/62/m062p011.pdf.

Posedel:1991:NSS

[PF91a] N. Posedel and J. Faganeli. Nature and sedimentation of suspended particulate matter during density stratification in shallow coastal waters (Gulf of Trieste, northern Adriatic). Marine Ecology Progress Series, 77:135–145, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
77/m077p135.pdf.

Proctor:1991:RVI

[PF91b] L. M. Proctor and J. A. Fuhrman. Roles of viral infection in organic particle flux. Marine Ecology Progress Series, 69:133–142, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/69/m069p133.pdf.

Pawlik:1992:CDP

[PF92a] J. R. Pawlik and W. Fenical. Chemical defense of Pterogorgia anceps, a Caribbean gorgonion coral. Marine Ecology Progress Series, 87:183–188, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/87/m087p183.pdf.

Pedrotti:1992:DEL

[PF92b] M. L. Pedrotti and L. Fenaux. Dispersal of echinoderm larvae in a geographical area marked by upwelling (Ligurian Sea, NW Mediterranean). Marine Ecology Progress Series, 86: 217–227, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/86/m086p217.pdf.

Proctor:1992:MMB

[PF92c] L. M. Proctor and J. A. Fuhrman. Mortality of marine bacteria in response to enrichments of the virus size fraction from seawater. Marine Ecology Progress Series, 87:283–293, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
87/m087p283.pdf.

Poulin:1995:PSD

[PF95] É Poulin and J.-P. Féral. Pattern of spatial distribution of a brood-protecting schizasterid echinoid, Abatus cordatus, endemic to the Kerguelen Islands. Marine Ecology Progress Series, 118:179–186, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/118/m118p179.pdf.

Planque:1996:CEE

[PF96] B. Planque and J. M. Fromentin. Calanus and environment in the eastern North Atlantic. I. Spatial and temporal patterns of C. finmarchicus and C. helgolandicus. Marine Ecology Progress Series, 134:101–109, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/134/m134p101.pdf.

Planque:1998:IVT

[PF98] B. Planque and C. J. Fox. Interannual variability in temperature and the recruitment of Irish Sea cod. Marine Ecology Progress Series, 172:101–105, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/172/m172p101.pdf.

Plante:1998:CPC

[PFH98] R. Plante, H. Fricke, and K. Hissmann. Coelacanth population, conservation and fishery activity at Grande Comore, West Indian Ocean. Marine Ecology Progress Series, 166: 231–236, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/166/m166p231.pdf.

Pfister:1997:DCW

[Pfi97] C. A. Pfister. Demographic consequences of within-year variation in recruitment. Marine Ecology Progress Series, 153: 229–238, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/153/m153p229.pdf.

Pranal:1995:AAR

[PFMC95] V. Pranal, A. Fiala-Médioni, and J. C. Colomines. Amino acid and related compound composition in two symbiotic mytilid species from hydrothermal vents. Marine Ecology Progress Series, 119:155–166, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/119/m119p155.pdf.

Pranal:1996:FAC

[PFMG96] V. Pranal, A. Fiala-Médioni, and J. Guezennec. Fatty acid characteristics in two symbiotic gastropods from a deep hydrothermal vent of the West Pacific. Marine Ecology Progress Series, 142:175–184, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/142/m142p175.pdf.

Pichard:1993:RBC

[PFP93] S. L. Pichard, M. E. Frischer, and J. H. Paul. Ribulose bisphosphate carboxylase gene expression in subtropical marine phytoplankton populations. Marine Ecology Progress Series, 101:55–65, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/101/m101p055.pdf.

Purcell:1990:PFL

[PG90] J. E. Purcell and J. J. Grover. Predation and food limitation as causes of mortality in larval herring at a spawning ground in British Columbia. Marine Ecology Progress Series, 59:55–61, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/59/m059p055.pdf.

Pages:1991:ELS

[PG91] F. Pagès and J. M. Gili. Effects of large-scale advective processes on gelatinous zooplankton populations in the northern Benguela ecosystem. Marine Ecology Progress Series, 75: 205–215, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/75/m075p205.pdf.

Peters:1994:IGR

[PG94] F. Peters and T. Gross. Increased grazing rates of microplankton in response to small-scale turbulence. Marine Ecology Progress Series, 115:299–307, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/115/m115p299.pdf.

Ploug:1999:PRC

[PGAJ99] H. Ploug, H. P. Grossart, F. Azam, and B. B. Jørgensen. Photosynthesis, respiration, and carbon turnover in sinking marine snow from surface waters of Southern California Bight: implications for the carbon cycle in the ocean. Marine Ecology Progress Series, 179:1–11, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/179/m179p001.pdf.

Patarnello:1991:EPH

[PGB91] T. Patarnello, R. Guinez, and B. Battaglia. Effects of pollution on heterozygosity in the barnacle Balanus amphitrite (Cirripedia: Thoracica). Marine Ecology Progress Series, 70: 237–243, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/70/m070p237.pdf.

Peletier:1996:UBR

[PGB96] H. Peletier, W. W. C. Gieskes, and A. G. J. Buma. Ultraviolet-b radiation resistance of benthic diatoms isolated from tidal flats in the Dutch Wadden Sea. Marine Ecology Progress Series, 135:163–168, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/135/m135p163.pdf.

Paquet:1993:ASS

[PGD^{+}93] R. Paquet, R. Goudard, J.-P. Durand, P. Germain, and J. Pieri. Accumulation of ^{241}Am in subcellular structures of the hepatopancreas of the lobster Homarus gammarus. Marine Ecology Progress Series, 95:155–162, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
95/m095p155.pdf.

Pages:1996:DGZ

[PGG96] F. Pagès, H. E. González, and S. R. González. Diet of the gelatinous zooplankton in Hardangerfjord (Norway) and potential predatory impact by Aglantha digitale (Trachymedusae). Marine Ecology Progress Series, 139:69–77, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/139/m139p069.pdf.

Pearce:1998:ETT

[PGM^{+}98] C. M. Pearce, S. M. Gallager, J. L. Manuel, D. A. Manning, R. K. O’Dor, and E. Bourget. Effect of thermoclines and turbulence on depth of larval settlement and spat recruitment of the giant scallop Placopecten magellanicus in 9.5 m deep laboratory mesocosms. Marine Ecology Progress Series, 165:195–215, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/165/m165p195.pdf.

Perez:1994:NDP

[PGR94] Fiz F. Pérez, Figueiras Francisco G, and Aida F. Ríos. Nutrient depletion and particulate matter near the ice-edge in the Weddell Sea. Marine Ecology Progress Series, 112:143– 153, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/112/m112p143.pdf.

Peduzzi:1991:DSS

[PH91a] P. Peduzzi and G. J. Herndl. Decomposition and significance of sea-grass leaf litter (Cymodocea nodosa) for the microbial food web in coastal waters (Gulf of Trieste, Northern Adriatic Sea). Marine Ecology Progress Series, 71:163–174, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
71/m071p163.pdf.

Peduzzi:1991:MTR

[PH91b] P. Peduzzi and G. J. Herndl. Mucus trails in the rocky intertidal. A highly active microenvironment. Marine Ecology Progress Series, 75:267–274, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/75/m075p267.pdf.

Poremba:1995:SVB

[PH95] K. Poremba and H.-G. Hoppe. Spatial variation of benthic microbial production and hydrolytic enzymatic activity down the continental slope of the Celtic Sea. Marine Ecology Progress Series, 118:237–245, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/118/m118p237.pdf.

Pond:1996:ENF

[PHHH96] D. Pond, R. Harris, R. Head, and D. Harbour. Environmental and nutritional factors determining seasonal variability in the fecundity and egg viability of Calanus helgolandicus in coastal waters off Plymouth, UK. Marine Ecology Progress Series, 143:45–63, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/143/m143p045.pdf.

Philippart:1994:IBA

[Phi94] C. J. M. Philippart. Interactions between Arenicola marina and Zostera noltii on a tidal flat in the Wadden Sea. Marine Ecology Progress Series, 111:251–257, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
111/m111p251.pdf.

Pena:1992:NPC

[PHL92] M. A. Pena, W. G. Harrison, and M. R. Lewis. New production in the central equatorial Pacific. Marine Ecology Progress Series, 80:265–274, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/80/m080p265.pdf.

Peebles:1996:EPB

[PHT96] E. B. Peebles, J. R. Hall, and S. G. Tolley. Egg production by the bay anchovy Anchoa mitchilli in relation to adult and larval prey fields. Marine Ecology Progress Series, 131: 61–73, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/131/m131p061.pdf.

Pohlman:1997:SIS

[PIC97] J. W. Pohlman, T. M. Iliffe, and L. A. Cifuentes. A stable isotope study of organic cycling and the ecology of an anchialine cave ecosystem. Marine Ecology Progress Series, 155: 17–27, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/155/m155p017.pdf.

Pineda:1994:STP

[Pin94] J. Pineda. Spatial and temporal patterns in barnacle settlement rate along a southern California rocky shore. Marine Ecology Progress Series, 107:125–138, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
107/m107p125.pdf.

Pires:1992:SDB

[Pir92] A. M. S. Pires. Structure and dynamics of benthic megafauna on the continental shelf offshore of Ubatuba, southeastern Brazil. Marine Ecology Progress Series, 86:63–76, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
86/m086p063.pdf.

Ploug:1999:NJF

[PJ99] H. Ploug and B. B. Jørgensen. A net-jet flow system for mass transfer and microsensor studies of sinking aggregates. Marine Ecology Progress Series, 176:279–290, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
176/m176p279.pdf.

Paerl:1993:ENL

[PJF93] H. W. Paerl, S. B. Joye, and M. Fitzpatrick. Evaluation of nutrient limitation of CO_{2} and N_{2} fixation in marine microbial mats. Marine Ecology Progress Series, 101:297–306, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
101/m101p297.pdf.

Phillips:1997:TFG

[PKL97] J. C. Phillips, G. A. Kendrick, and P. S. Lavery. A test of a functional group approach to detecting shifts in macroalgal communities along a disturbance gradient. Marine Ecology Progress Series, 153:125–138, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/153/m153p125.pdf.

Peckol:1996:IEI

[PKY96] P. Peckol, J. M. Krane, and J. L. Yates. Interactive effects of inducible defense and resource availability on phlorotannins in the North Atlantic brown alga Fucus vesiculosus. Marine Ecology Progress Series, 138:209–217, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
138/m138p209.pdf.

Paffenhofer:1999:USM

[PL99a] G. A. Paffenhöfer and P. A. Loyd. Ultrastructure of setae of the maxilliped of the marine planktonic copepod Temora stylifera. Marine Ecology Progress Series, 178:101–107, 1999.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
178/m178p101.pdf.

Phillips:1999:DTA

[PL99b] T. M. Phillips and C. R. Lovell. Distributions of total and active bacteria in biofilms lining tubes of the onuphid polychaete Diopatra cuprea. Marine Ecology Progress Series, 183: 169–178, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/183/m183p169.pdf.

Planes:1993:GDR

[Pla93] S. Planes. Genetic differentiation in relation to restricted larval dispersal of the convict surgeonfish Acanthurus triostegus in French Polynesia. Marine Ecology Progress Series, 98: 237–246, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/98/m098p237.pdf.

Phlips:1995:CTC

[PLB95] E. J. Phlips, T. C. Lynch, and S. Badylak. chlorophyll-a, tripton, color, and light availability in a shallow tropical innershelf lagoon, Florida Bay, USA. Marine Ecology Progress Series, 127:223–234, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/127/m127p223.pdf.

Pouliquen:1996:CSC

[PLBP96] H. Pouliquen, H. Le Bris, V. Buchet, and L. Pinault. Comparative study on the contamination and decontamination of Japanese oyster Crassostrea gigas and blue mussel Mytilus edulis by oxytetracycline and oxolinic acid. Marine Ecology Progress Series, 133:143–148, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/133/m133p143.pdf.

Priscu:1991:CIR

[PLC^{+}91] J. C. Priscu, M. P. Lizotte, G. F. Cota, A. C. Palmisano, and C. W. Sullivan. Comparison of the irradiance response of photosynthesis and nitrogen uptake by sea ice microalgae. Marine Ecology Progress Series, 70:201–210, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
70/m070p201.pdf.

Paul:1990:CDT

[PLF90] V. J. Paul, N. Lindquist, and W. Fenical. Chemical defenses of the tropical ascidian Atapozoa sp. and its nudibranch predators Nembrotha spp. Marine Ecology Progress Series, 59:109–118, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/59/m059p109.pdf.

Pape-Lindstrom:1997:SPF

[PLFSW97] P. A. Pape-Lindstrom, R. J. Feller, S. E. Stancyk, and S. A. Woodin. Sublethal predation: field measurements of arm tissue loss from the ophiuroid Microphiopholis gracillima and immunochemical identification of its predators in North Inlet, South Carolina, USA. Marine Ecology Progress Series, 156:131–140, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/156/m156p131.pdf.

Pesant:1996:SDR

[PLG^{+}96] S. Pesant, L. Legendre, M. Gosselin, R. E. H. Smith, G. Kattner, and R. O. Ramseier. Size-differential regimes of phytoplankton production in the northeast water polynya (77*- 81*N). Marine Ecology Progress Series, 142:75–86, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
142/m142p075.pdf.

Pefia:1991:POM

[PLH91] M. A. Pefia, M. R. Lewis, and W. G. Harrison. Particulate organic matter and chlorophyll in the surface layer of the equatorial Pacific Ocean along 135^{o}W. Marine Ecology Progress Series, 72:179–188, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/72/m072p179.pdf.

Poulet:1995:RRC

[PLIM95] S. A. Poulet, M. Laabir, A. Ianora, and A. Miralto. Reproductive response of Calanus helgolandicus. I. Abnormal embryonic and naupliar development. Marine Ecology Progress Series, 129:85–95, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/129/m129p085.pdf.

Perissinotto:1992:MPE

[PLM92] R. Perissinotto, R. K. Laubscher, and C. D. McQuaid. Marine productivity enhancement around Bouvet and the South Sandwich Islands (Southern Ocean). Marine Ecology Progress Series, 88:41–53, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/88/m088p041.pdf.

Poulet:1994:DDA

[PlMM94] S. A. Poulet, A. lanora, A. Miralto, and L. Meijer. Do diatoms arrest embryonic development in cope-pods? Marine Ecology Progress Series, 111:79–86, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/111/m111p079.pdf.

Pouliquen:1992:EST

[PLP92] H. Pouliquen, H. Le Bris, and L. Pinault. Experimental study of the therapeutic application of oxytetracycline, its attenuation in sediment and sea water, and implications for farm culture of benthic organisms. Marine Ecology Progress Series, 89:93–98, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/89/m089p093.pdf.

Painting:1993:DBP

[PLP^{+}93] S. J. Painting, M. I. Lucas, W. T. Peterson, P. C. Brown, L. Hutchings, and B. A. Mitchell-lnnes. Dynamics of bacterioplankton, phytoplankton and mesozooplankton communities during the development of an upwelling plume in the southern Benguela. Marine Ecology Progress Series, 100:35–53, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
100/m100p035.pdf.

Priddle:1995:NCA

[PLS^{+}95] J. Priddle, R. Leakey, C. Symon, M. Whitehouse, D. Robins, G. Cripps, E. Murphy, and N. Owens. Nutrient cycling by Antarctic marine microbial plankton. Marine Ecology Progress Series, 116:181–198, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/116/m116p181.pdf.

Palenik:1990:CCS

[PM90a] B. Palenik and F. M. M. Morel. Comparison of cell-surface l-amino acid oxidases from several marine phytoplankton. Marine Ecology Progress Series, 59:195–201, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
59/m059p195.pdf.

Perissinotto:1990:RSA

[PM90b] R. Perissinotto and C. McQuaid. Role of the sub-Antarctic shrimp Nauticaris mahonis in coupling benthic and pelagic food-webs. Marine Ecology Progress Series, 64:81–87, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
64/m064p081.pdf.

Pollard:1991:OCD

[PM91a] Peter C. Pollard and D. J. W. Moriarty. Organic carbon decomposition, primary and bacterial productivity, and sulphate reduction, in tropical seagrass beds of the Gulf of Carpentaria, Australia. Marine Ecology Progress Series, 69:149– 159, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/69/m069p149.pdf.

Purcell:1991:DPM

[PM91b] J. E. Purcell and L. P. Madin. Diel patterns of migration, feeding, and spawning by salps in the subarctic Pacific. Marine Ecology Progress Series, 73:211–217, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
73/m073p211.pdf.

Perissinotto:1992:LBP

[PM92a] R. Perissinotto and C. D. McQuaid. Land-based predator impact on vertically migrating zooplankton and micronekton advected to a Southern Ocean archipelago. Marine Ecology Progress Series, 80:15–27, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/80/m080p015.pdf.

Piatt:1992:TFB

[PM92b] J. F. Piatt and D. A. Methven. Threshold foraging behavior of baleen whales. Marine Ecology Progress Series, 84: 205–210, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/84/m084p205.pdf.

Pilditch:1994:QAC

[PM94a] C. A. Pilditch and S. McClatchie. Quantitative analysis of carnivory in the krill Nyctiphanes australis, with an examination of the effect of non-preferred phytoplankton alternative prey. Marine Ecology Progress Series, 107:41–53, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
107/m107p041.pdf.

Plante:1994:DEB

[PM94b] C. J. Plante and L. M. Mayer. Distribution and efficiency of bacteriolysis in the gut of Arenicola marina and three additional deposit feeders. Marine Ecology Progress Series, 109:183–194, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/109/m109p183.pdf.

Procaccini:1998:PGS

[PM98] G. Procaccini and L. Mazzella. Population genetic structure and gene flow in the seagrass Posidonia oceanica assessed using microsatellite analysis. Marine Ecology Progress Series, 169:133–141, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/169/m169p133.pdf.

Peirano:1999:SGD

[PMB99] A. Peirano, C. Morri, and C. N. Bianchi. Skeleton growth and density pattern of the temperate, zooxanthellate scleractinian Cladocora caespitosa from the Ligurian Sea (NW Mediterranean). Marine Ecology Progress Series, 185:195–201, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
185/m185p195.pdf.

Pasantes:1996:SCE

[PMETM96] J. J. Pasantes, M. J. Martínez-Expósito, A. Torreiro, and J. Méndez. The sister chromatid exchange test as an indicator of marine pollution: some factors affecting SCE frequencies in Mytilus galloprovincialis. Marine Ecology Progress Series, 143:113–119, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/143/m143p113.pdf.

Peters:1998:ETB

[PMG^{+}98] F. Peters, C. Marrasé, J. M. Gasol, M. M. Sala, and L. Arin. Effects of turbulence on bacterial growth mediated through food web interactions. Marine Ecology Progress Series, 172: 293–303, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/172/m172p293.pdf.

Painting:1993:SFM

[PML93] S. J. Painting, C. L. Moloney, and M. I. Lucas. Simulation and field measurements of phytoplankton-bacteria-zooplankton interactions in the southern Benguela upwelling region. Marine Ecology Progress Series, 100:55–69, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
100/m100p055.pdf.

Parrish:1995:SSS

[PMMH95] C. C. Parrish, C. H. McKenzie, B. A. MacDonald, and E. A. Hatfield. Seasonal studies of seston lipids in relation to microplankton species composition and scallop growth in South Broad Cove, Newfoundland. Marine Ecology Progress Series, 129:151–164, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/129/m129p151.pdf.

Pomeroy:1990:MFW

[PMOD90] L. R. Pomeroy, S. A. Macko, P. H. Ostrom, and J. Dunphy. The microbial food web in Arctic seawater concentration of dissolved free amino acids and bacterial abundance and activity in the Arctic Ocean and in Resolute Passage. Marine Ecology Progress Series, 61:31–40, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/61/m061p031.pdf.

Parpais:1996:EPS

[PMP^{+}96] J. Parpais, D. Marie, F. Partensky, P. Morin, and D. Vaulot. Effect of phosphorus starvation on the cell cycle of the photosynthetic prokaryote Prochlorococcus spp. Marine Ecology Progress Series, 132:265–274, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/132/m132p265.pdf.

Putt:1994:ABP

[PMS94] M. Putt, G. Miceli, and D. K. Stoecker. Association of bacteria with Phaeocystis sp. in McMurdo Sound, Antarctica. Marine Ecology Progress Series, 105:179–189, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
105/m105p179.pdf.

Peres-Neto:1999:HMS

[PN99] P. R. Peres-Neto. How many statistical tests are too many? The problem of conducting multiple ecological inferences revisited. Marine Ecology Progress Series, 176:303–306, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
176/m176p303.pdf.

Pelegri:1994:DES

[PNB94] S. P. Pelegri, L. P. Nielsen, and T. H. Blackburn. Denitrification in estuarine sediment stimulated by the irrigation activity of the amphipod Corophium volutator. Marine Ecology Progress Series, 105:285–290, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/105/m105p285.pdf.

Purcell:1994:PMB

[PND^{+}94] J. E. Purcell, D. A. Nemazie, S. E. Dorsey, E. D. Houde, and J. C. Gamble. Predation mortality of bay anchovy Anchoa mitchilli eggs and larvae due to scyphomedusae and ctenophores in Chesapeake Bay. Marine Ecology Progress Series, 114:47–58, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/114/m114p047.pdf.

Paul:1990:FPA

[PNS90] V. J. Paul, S. G. Nelson, and H. R. Sanger. Feeding preferences of adult and juvenile rabbitfish Siganus argenteus in relation to chemical defenses of tropical seaweeds. Marine Ecology Progress Series, 60:23–34, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/60/m060p023.pdf.

Pardieck:1999:OCH

[PODL99] R. A. Pardieck, R. J. Orth, R. J. Diaz, and R. N. Lipcius. Ontogenetic changes in habitat use by postlarvae and young juveniles of the blue crab. Marine Ecology Progress Series, 186:227–238, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/186/m186p227.pdf.

Poore:1994:SHA

[Poo94] A. G. B. Poore. Selective herbivory by amphipods inhabiting the brown alga Zonaria angustata. Marine Ecology Progress Series, 107:113–123, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/107/m107p113.pdf.

Poremba:1994:SDP

[Por94] K. Poremba. Simulated degradation of phytodetritus in deepsea sediments of the NE Atlantic (47^{o}N, 19^{o}W). Marine Ecology Progress Series, 105:291–299, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/105/m105p291.pdf.

Pakhomov:1996:THA

[PP96] E. A. Pakhomov and R. Perissinotto. Trophodynamics of the hyperiid amphipod Themisto gaudichaudi in the South Georgia region during late austral summer. Marine Ecology Progress Series, 134:91–100, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/134/m134p091.pdf.

Peterson:1990:FRC

[PPB90] W. Peterson, S. Painting, and R. Barlow. Feeding rates of Calanoides carinatus: a comparison of five methods including evaluation of the gut fluorescence method. Marine Ecology Progress Series, 63:85–92, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/63/m063p085.pdf.

Pari:1998:BES

[PPCL^{+}98] N. Pari, M. Peyrot-Clausade, T. Le Campion-Alsumard, P. Hutchings, V. Chazottes, S. Golubic, J. Le Campion, and M. F. Fontaine. Bioerosion of experimental substrates on high islands and on atoll lagoons (French Polynesia) after two years of exposure. Marine Ecology Progress Series, 166: 119–130, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/166/m166p119.pdf.

Pujo-Pay:1997:EDO

[PPCR97] M. Pujo-Pay, P. Conan, and P. Raimbault. Excretion of dissolved organic nitrogen by phytoplankton assessed by wet oxidation and ^{15}N tracer procedures. Marine Ecology Progress Series, 153:99–111, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/153/m153p099.pdf.

Pinckney:1995:ISN

[PPF95] J. Pinckney, H. W. Paerl, and M. Fitzpatrick. Impacts of seasonality and nutrients on microbial mat community structure and function. Marine Ecology Progress Series, 123:207– 216, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/123/m123p207.pdf.

Paerl:1998:ERI

[PPFP98] H. W. Paerl, J. L. Pinckney, J. M. Fear, and B. L. Peierls. Ecosystem responses to internal and watershed organic matter loading: consequences for hypoxia in the eutrophying Neuse River Estuary, North Carolina, USA. Marine Ecology Progress Series, 166:17–25, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/166/m166p017.pdf.

Paerl:1999:FKB

[PPFP99] H. W. Paerl, J. L. Pinckney, J. M. Fear, and B. L. Peierls. Fish kills and bottom-water hypoxia in the Neuse River and Estuary: reply to Burkholder et al. Marine Ecology Progress Series, 186:307–309, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/186/m186p307.pdf. See [BMG99].

Pakhomov:1994:CSM

[PPM94] E. A. Pakhomov, R. Perissinotto, and C. D. McQuaid. Comparative structure of the macrozooplankton/ micronekton communities of the subtropical and Antarctic polar fronts. Marine Ecology Progress Series, 111:155–169, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
111/m111p155.pdf.

Pakhomov:1996:PCD

[PPM96] E. A. Pakhomov, R. Perissinotto, and C. D. McQuaid. Prey composition and daily rations of myctophid fishes in the Southern Ocean. Marine Ecology Progress Series, 134:1–14, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/134/m134p001.pdf.

Perissinotto:1997:LSG

[PPMF97] R. Perissinotto, E. A. Pakhomov, C. D. McQuaid, and P. W. Froneman. In situ grazing rates and daily ration of Antarctic krill Euphausia superba feeding on phytoplankton at the Antarctic polar front and the marginal ice zone. Marine Ecology Progress Series, 160:77–91, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/160/m160p077.pdf.

Power:1999:VCS

[PPO^{+}99] A. M. Power, S. Piyapattanakorn, R. M. O’Riordan, A. Iyengar, A. A. Myers, S. J. Hawkins, J. Delany, D. McGrath, and N. Maclean. Verification of cyprid size as a tool in the identification of two European species of Chthamalus barnacles using mtDNA-RFLP analysis. Marine Ecology Progress Series, 191:251–256, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/191/m191p251.pdf.

Pennings:1996:ESM

[PPP^{+}96] S. C. Pennings, M. P. Puglisi, T. J. Pitlik, A. C. Himaya, and V. J. Paul. Effects of secondary metabolites and CaCO 3 on feeding by surgeonfishes and parrotfishes: within-plant comparisons. Marine Ecology Progress Series, 134:49–58, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
134/m134p049.pdf.

Pujo-Pay:1994:IWO

[PPR94] M. Pujo-Pay and P. Raimbault. improvement of the wetoxidation procedure for simultaneous determination of particulate organic nitrogen and phosphorus collected on filters. Marine Ecology Progress Series, 105:203–207, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
105/m105p203.pdf.

Pile:1996:LSG

[PPW96] A. J. Pile, M. R. Patterson, and J. D. Witman. In situ grazing on plankton < 10μm by the boreal sponge Mycale lingua. Marine Ecology Progress Series, 141:95–102, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
141/m141p095.pdf.

Petersen:1992:FCA

[PR92] J. K. Petersen and H. U. Riisgård. Filtration capacity of the ascidian Ciona intestinalis and its grazing impact in a shallow fjord. Marine Ecology Progress Series, 88:9–17, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
88/m088p009.pdf.

Plourde:1993:RPC

[PR93a] S. Plourde and J. A. Runge. Reproduction of the planktonic copepod Calanus finmarchicus in the Lower St. Lawrence Estuary: relation to the cycle of phytoplankton production and evidence for a Calanus pump. Marine Ecology Progress Series, 102:217–227, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/102/m102p217.pdf.

Polunin:1993:GBV

[PR93b] N. V. C. Polunin and C. M. Roberts. Greater biomass and value of target coral-reef fishes in two small Caribbean marine reserves. Marine Ecology Progress Series, 100:167–176, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
100/m100p167.pdf.

Parsons:1992:IAL

[PRC^{+}92] G. J. Parsons, S. M. C. Robinson, R. A. Chandler, L. A. Davidson, M. Lanteigne, and M. J. Dadswell. Intra-annual and long-term patterns in the reproductive cycle of giant scallops Placopecten magellanicus (Bivalvia: Pectinidae) from Passamaquoddy Bay, New Brunswick, Canada. Marine Ecology Progress Series, 80:203–214, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/80/m080p203.pdf.

Prego:1991:FBN

[Pre91] R. Prego. Flows and budgets of nutrient salts and organic carbon in relation to a red tide in the Ria of Vigo (NW Spain). Marine Ecology Progress Series, 79:289–302, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
79/m079p289.pdf.

Preen:1995:IDF

[Pre95] A. Preen. Impacts of dugong foraging on seagrass habitats: observational and experimental evidence for cultivation grazing. Marine Ecology Progress Series, 124:201–213, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
124/m124p201.pdf.

Pan:1996:ESLa

[PRM^{+}96a] Y. Pan, D. V. Subba Rao, K. H. Mann, R. G. Brown, and R. Pocklington. Effects of silicate limitation on production of domoic acid, a neurotoxin, by the diatom Pseudo-nitzschia multiseries. I. Batch culture studies. Marine Ecology Progress Series, 131:225–233, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/131/m131p225.pdf.

Pan:1996:ESLb

[PRM^{+}96b] Y. Pan, D. V. Subba Rao, K. H. Mann, W. K. W. Li, and W. G. Harrison. Effects of silicate limitation on production of domoic acid, a neurotoxin, by the diatom Pseudo-nitzschia multiseries. II. Continuous culture studies. Marine Ecology Progress Series, 131:235–243, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/131/m131p235.pdf.

Protasowicki:1992:PSS

[Pro92] M. Protasowicki. Preliminary studies on selected elements in organs of dab Limanda limanda and their relation to fish disease state. Marine Ecology Progress Series, 91:57–60, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
91/m091p057.pdf.

Pergent:1994:PPS

[PRPM^{+}94] G. Pergent, J. Romero, C. Pergent-Martini, M.-A. Mateo, and C.-F. Boudouresque. Primary production, stocks and fluxes in the Mediterranean seagrass Posidonia oceanica. Marine Ecology Progress Series, 106:139–146, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/106/m106p139.pdf.

Piatkowski:1994:NCS

[PRW^{+}94] U. Piatkowski, P. G. Rodhouse, M. G. White, D. G. Bone, and C. Symon. Nekton community of the Scotia Sea as sampled by the RMT 25 during austral summer. Marine Ecology Progress Series, 112:13–28, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/112/m112p013.pdf.

Peterson:1992:BSC

[PS92] C. H. Peterson and H. C. Summerson. Basin-scale coherence of population dynamics of an exploited marine invertebrate, the bay scallop: implications of recruitment limitation. Marine Ecology Progress Series, 90:257–272, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
90/m090p257.pdf.

Pennings:1993:DCF

[PS93] Steven C. Pennings and JoanMarie Svedberg. Does CaCO_{3} in food deter feeding by sea urchins? Marine Ecology Progress Series, 101:163–167, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/101/m101p163.pdf.

Pennock:1994:TAB

[PS94] J. R. Pennock and J. H. Sharp. Temporal alternation between light- and nutrient-limitation of phytoplankton production in a coastal plain estuary. Marine Ecology Progress Series, 111:275–288, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/111/m111p275.pdf.

Pepin:1995:IBS

[PS95] P. Pepin and T. H. Shears. Influence of body size and alternate prey abundance on the risk of predation to fish larvae. Marine Ecology Progress Series, 128:279–285, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
128/m128p279.pdf.

Piepenburg:1997:PSE

[PS97] D. Piepenburg and M. K. Schmid. A photographic survey of the epibenthic megafauna of the Arctic Laptev Sea shelf: distribution, abundance, and estimates of biomass and organic carbon demand. Marine Ecology Progress Series, 147: 63–75, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/147/m147p063.pdf.

Plante:1998:PDD

[PS98] C. J. Plante and A. G. Shriver. Patterns of differential digestion of bacteria in deposit feeders: a test of resource partitioning. Marine Ecology Progress Series, 163:253–258, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
163/m163p253.pdf.

Pond:1997:LLC

[PSB^{+}97] D. W. Pond, M. Segonzac, M. V. Bell, D. R. Dixon, A. E. Fallick, and J. R. Sargent. Lipid and lipid carbon stable isotope composition of the hydrothermal vent shrimp Mirocaris fortunata: evidence for nutritional dependence on photosynthetically fixed carbon. Marine Ecology Progress Series, 157:221–231, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/157/m157p221.pdf.

Prahl:1997:BCS

[PSE97] F. G. Prahl, L. F. Small, and B. Eversmeyer. Biogeochemical characterization of suspended particulate matter in the Columbia River estuary. Marine Ecology Progress Series, 160:173–184, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/160/m160p173.pdf.

Petersen:1998:CPR

[PSK98] J. E. Petersen, L. P. Sanford, and W. M. Kemp. Coastal plankton responses to turbulent mixing in experimental ecosystems. Marine Ecology Progress Series, 171:23–41, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
171/m171p023.pdf.

Peterson:1996:RBS

[PSL96] C. H. Peterson, H. C. Summerson, and R. A. Luettich, Jr. Response of bay scallops to spawner transplants: a test of recruitment limitation. Marine Ecology Progress Series, 132: 93–107, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/132/m132p093.pdf.

Prins:1996:FRP

[PSPD96] T. C. Prins, A. C. Smaal, A. J. Pouwer, and N. Dankers. Filtration and resuspension of particulate matter and phytoplankton on an intertidal mussel bed in the Oosterschelde estuary (SW Netherlands). Marine Ecology Progress Series, 142:121–134, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/142/m142p121.pdf.

Prena:1999:EOT

[PSR^{+}99] J. Prena, P. Schwinghamer, T. W. Rowell, D. C. Gordon, Jr., K. D. Gilkinson, W. P. Vass, and D. L. McKeown. Experimental otter trawling on a sandy bottom ecosystem of the Grand Banks of Newfoundland: analysis of trawl bycatch and effects on epifauna. Marine Ecology Progress Series, 181:107–124, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/181/m181p107.pdf.

Pomeroy:1995:LGR

[PSSP95] L. R. Pomeroy, J. E. Sheldon, W. M. Sheldon, Jr., and F. Peters. Limits to growth and respiration of bacterioplankton in the Gulf of Mexico. Marine Ecology Progress Series, 117: 259–268, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/117/m117p259.pdf.

Petersen:1995:GEA

[PST95] J. K. Petersen, O. Schou, and P. Thor. Growth and energetics in the ascidian Ciona intestinalis. Marine Ecology Progress Series, 120:175–184, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/120/m120p175.pdf.

Pierce:1993:GBM

[PT93] R. W. Pierce and J. T. Turner. Global biogeography of marine tintinnids. Marine Ecology Progress Series, 94:11–26, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
94/m094p011.pdf.

Pierce:1994:PSBa

[PT94a] R. W. Pierce and J. T. Turner. Plankton studies in Buzzards Bay, Massachusetts, USA. III. Dinoflagellates, 1987 to 1988. Marine Ecology Progress Series, 112:225–234, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
112/m112p225.pdf.

Pierce:1994:PSBb

[PT94b] R. W. Pierce and J. T. Turner. Plankton studies in Buzzards Bay, Massachusetts, USA. IV. Tintinnids, 1987 to 1988. Marine Ecology Progress Series, 112:235–240, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
112/m112p235.pdf.

Pitkanen:1995:NPP

[PT95] H. Pitkänen and T. Tamminen. Nitrogen and phosphorus as production limiting factors in the estuarine waters of the eastern Gulf of Finland. Marine Ecology Progress Series, 129:283–294, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/129/m129p283.pdf.

Piontkovski:1997:BFA

[PTB^{+}97] S. A. Piontkovski, Y. N. Tokarev, E. P. Bitukov, R. Williams, and D. A. Kiefer. The bioluminescent field of the Atlantic Ocean. Marine Ecology Progress Series, 156:33–41, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
156/m156p033.pdf.

Pridmore:1991:EOP

[PTW^{+}91] R. D. Pridmore, S. F. Thrush, R. J. Wilcock, T. J. Smith, J. E. Hewitt, and V. J. Cummings. Effect of the organochlorine pesticide technical chlordane on the population structure of suspension and deposit feeding bivalves. Marine Ecology Progress Series, 76:261–271, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/76/m076p261.pdf.

Pudovkin:1993:CRA

[Pud93] A. I. Pudovkin. Citation relationships among marine biology journals and those in related fields. Marine Ecology Progress Series, 100:207–209, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/100/m100p207.pdf.

Purcell:1991:PAV

[Pur91] J. E. Purcell. Predation by Aequorea victoria on other species of potentially competing pelagic hydrozoans. Marine Ecology Progress Series, 72:255–260, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/72/m072p255.pdf.

Purcell:1992:EPS

[Pur92] J. E. Purcell. Effects of predation by the scyphomedusan Chrysaora quinquecirrha on zooplankton populations in Chesapeake Bay, USA. Marine Ecology Progress Series, 87: 65–76, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/87/m087p065.pdf.

Putt:1990:MPC

[Put90] M. Putt. Metabolism of photosynthate in the chloroplastretaining ciliate Laboea strobila. Marine Ecology Progress Series, 60:271–282, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/60/m060p271.pdf.

Putt:1991:DET

[Put91] M. Putt. Development and evaluation of tracer particles for use in microzooplankton herbivory studies. Marine Ecology Progress Series, 77:27–37, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/77/m077p027.pdf.

Passow:1994:TFP

[PW94] U. Passow and P. Wassmann. On the trophic fate of Phaeocystis pouchetii (Hariot): IV. The formation of marine snow by P. pouchetii. Marine Ecology Progress Series, 104:153– 161, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/104/m104p153.pdf.

Pomeroy:1991:BRT

[PWD^{+}91] L. R. Pomeroy, W. J. Wiebe, D. Deibel, R. J. Thompson, G. T. Rowe, and J. D. Pakulski. Bacterial responses to temperature and substrate concentration during the Newfoundland spring bloom. Marine Ecology Progress Series, 75:143– 159, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/75/m075p143.pdf.

Paerl:1999:RSP

[PWG^{+}99] H. W. Paerl, J. D. Willey, M. Go, B. L. Peierls, J. L. Pinckney, and M. L. Fogel. Rainfall stimulation of primary production in western Atlantic Ocean waters: roles of different nitrogen sources and co-limiting nutrients. Marine Ecology Progress Series, 176:205–214, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/176/m176p205.pdf.

Piontkovski:1995:SHB

[PWM95] S. A. Piontkovski, R. Williams, and T. A. Melnik. Spatial heterogeneity, biomass and size structure of plankton of the Indian Ocean: some general trends. Marine Ecology Progress Series, 117:219–227, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/117/m117p219.pdf.

Purcell:1999:TSF

[PWNW99] J. E. Purcell, J. R. White, D. A. Nemazie, and D. A. Wright. Temperature, salinity and food effects on asexual reproduction and abundance of the scyphozoan Chrysaora quinquecirrha. Marine Ecology Progress Series, 180:187–196, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
180/m180p187.pdf.

Piontkovski:1997:SHP

[PWP^{+}97] S. A. Piontkovski, R. Williams, W. T. Peterson, O. A. Yunev, N. I. Minkina, V. L. Vladimirov, and A. Blinko. Spatial heterogeneity of the planktonic fields in the upper mixed layer of the open ocean. Marine Ecology Progress Series, 148:145–154, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/148/m148p145.pdf.

Piontkovski:1995:RBO

[PWPK95] S. A. Piontkovski, R. Williams, W. Peterson, and V. K. Kosnirev. Relationship between oceanic mesozooplankton and energy of eddy fields. Marine Ecology Progress Series, 128: 35–41, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/128/m128p035.pdf.

Pond:1995:VLC

[PWPS95] D. Pond, J. Watkins, J. Priddle, and J. Sargent. Variation in the lipid content and composition of Antarctic krill Euphausia superba at South Georgia. Marine Ecology Progress Series, 117:49–57, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/117/m117p049.pdf.

Pages:1996:AGC

[PWR96] F. Pagès, M. G. White, and P. G. Rodhouse. Abundance of gelatinous carnivores in the nekton community of the Antarctic Polar Frontal Zone in summer 1994. Marine Ecology Progress Series, 141:139–147, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/141/m141p139.pdf.

Palma:1998:DEP

[PWS98] A. T. Palma, R. A. Wahle, and R. S. Steneck. Different early post-settlement strategies between American lobsters Homarus americanus and rock crabs Cancer irroratus in the Gulf of Maine. Marine Ecology Progress Series, 162:215– 225, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/162/m162p215.pdf.

Pinckney:1991:ETS

[PZ91] J. Pinckney and R. G. Zingmark. Effects of tidal stage and sun angles on intertidal benthic microalgal productivity. Marine Ecology Progress Series, 76:81–89, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/76/m076p081.pdf.

Panigada:1999:HDC

[PZCJ99] S. Panigada, M. Zanardelli, S. Canese, and M. Jahoda. How deep can baleen whales dive? Marine Ecology Progress Series, 187:309–311, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/187/m187p309.pdf.

Qian:1994:LSM

[QC94] P.-Y. Qian and F.-S. Chia. In situ measurement of recruitment, mortality, growth, and fecundity of Capitella sp. (Annelida: Polychaeta). Marine Ecology Progress Series, 111: 53–62, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/111/m111p053.pdf.

Queiroga:1994:LAP

[QCM94] H. Queiroga, J. D. Costlow, and M. H. Moreira. Larval abundance patterns of Carcinus maenas (Decapoda, Brachyura) in Canal de Mira (Ria de Aveiro, Portugal). Marine Ecology Progress Series, 111:63–72, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/111/m111p063.pdf.

Queiroga:1997:VMC

[QCM97] H. Queiroga, J. D. Costlow, and M. H. Moreira. Vertical migration of the crab Carcinus maenas first zoea in an estuary: implications for tidal stream transport. Marine Ecology Progress Series, 149:121–132, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/149/m149p121.pdf.

Qiu:1997:EST

[QGQ97] J. W. Qiu, L. A. Gosselin, and P. Y. Qian. Effects of shortterm variation in food availability on larval development in the barnacle Balanus amphitrite amphitrite. Marine Ecology Progress Series, 161:83–91, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/161/m161p083.pdf.

Quinn:1993:PEE

[QK93] G. P. Quinn and M. J. Keough. Potential effect of enclosure size on field experiments with herbivorous intertidal gastropods. Marine Ecology Progress Series, 98:199–201, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
98/m098p199.pdf.

Qiu:1997:CES

[QQ97] J. W. Qiu and P. Y. Qian. Combined effects of salinity, temperature and food on early development of the polychaete Hydroides elegans. Marine Ecology Progress Series, 152: 79–88, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/152/m152p079.pdf.

Qiu:1998:CES

[QQ98] J. W. Qiu and P. Y. Qian. Combined effects of salinity and temperature on juvenile survival, growth and maturation in the polychaete Hydroides elegans. Marine Ecology Progress Series, 168:127–134, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/168/m168p127.pdf.

Qiu:1999:TBB

[QQ99] J. W. Qiu and P. Y. Qian. Tolerance of the barnacle Balanus amphitrite amphitrite to salinity and temperature stress: effects of previous experience. Marine Ecology Progress Series, 188:123–132, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/188/m188p123.pdf.

Qian:1999:MUF

[QRSC99] P. Y. Qian, D. Rittschof, B. Sreedhar, and F. S. Chia. Macrofouling in unidirectional flow: miniature pipes as experimental models for studying the effects of hydrodynamics on invertebrate larval settlement. Marine Ecology Progress Series, 191:141–151, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/191/m191p141.pdf.

Quesada:1995:MAD

[QZA95] H. Quesada, C. Zapata, and G. Alvarez. A multilocus allozyme discontinuity in the mussel Mytilus galloprovincialis: the interaction of ecological and life-history factors. Marine Ecology Progress Series, 116:99–115, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
116/m116p099.pdf.

Renton:1992:HMM

[RA92] K. W. Renton and R. F. Addison. Hepatic microsomal monooxygenase activity and P450IA mRNA in North Sea dab Limanda limanda from contaminated sites. Marine Ecology Progress Series, 91:65–69, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/91/m091p065.pdf.

Russ:1996:DMR

[RA96] G. R. Russ and A. C. Alcala. Do marine reserves export adult fish biomass? Evidence from Apo Island, central Philippines. Marine Ecology Progress Series, 132:1–9, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
132/m132p001.pdf.

Rau:1992:SWS

[RAB^{+}92] G. H. Rau, D. G. Ainley, J. L. Bengtson, J. J. Torres, and T. L. Hopkins. ^{15}N/^{14}N and ^{13}C/^{12}C in Weddell Sea birds, seals, and fish: implications for diet and trophic structure. Marine Ecology Progress Series, 84:1–8, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
84/m084p001.pdf.

Rothschild:1994:DCB

[RAGH94] B. J. Rothschild, J. S. Ault, P. Goulletquer, and M. Héral. Decline of the Chesapeake Bay oyster population: a century of habitat destruction and overfishing. Marine Ecology Progress Series, 111:29–39, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/111/m111p029.pdf.

Ralph:1998:PRL

[Ral98] P. J. Ralph. Photosynthetic response of laboratory-cultured Halophila ovalis to thermal stress. Marine Ecology Progress Series, 171:123–130, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/171/m171p123.pdf.

Robinson:1999:MDC

[RAlBW99] Carol Robinson, Stephen D. Archer, and Peter J. le B. Williams. Microbial dynamics in coastal waters of East Antarctica: plankton production and respiration. Marine Ecology Progress Series, 180:23–36, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/180/m180p023.pdf.

Ribic:1992:ENN

[RAS92] C. A. Ribic, D. G. Ainley, and L. B. Spear. Effects of El Niño and La Niña on seabird assemblages in the Equatorial Pacific. Marine Ecology Progress Series, 80:109–124, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
80/m080p109.pdf.

Ribic:1997:SRS

[RAS97] C. A. Ribic, D. G. Ainley, and L. B. Spear. Scale-related seabird-environmental relationships in Pacific equatorial waters, with reference to El Niño–Southern Oscillation events. Marine Ecology Progress Series, 156:183–203, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
156/m156p183.pdf.

Roson:1999:CCL

[RÁSP99] G. Rosón, X. A. Álvarez-Salgado, and F. F. Pérez. Carbon cycling in a large coastal embayment, affected by winddriven upwelling: short-time-scale variability and spatial differences. Marine Ecology Progress Series, 176:215–230, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
176/m176p215.pdf.

Rainbow:1999:TMU

[RATA^{+}99] P. S. Rainbow, C. Amiard-Triquet, J. C. Amiard, B. D. Smith, S. L. Best, Y. Nassiri, and W. J. Langston. Trace metal uptake rates in crustaceans (amphipods and crabs) from coastal sites in NW Europe differentially enriched with trace metals. Marine Ecology Progress Series, 183:189–203, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
183/m183p189.pdf.

Rawlings:1996:SAU

[Raw96] T. A. Rawlings. Shields against ultraviolet radiation: an additional protective role for the egg capsules of benthic marine gastropods. Marine Ecology Progress Series, 136:81–95, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
136/m136p081.pdf.

Ryland:1990:PCF

[RB90] J. S. Ryland and J. D. D. Bishop. Prevalence of crossfertilization in the hermaphroditic compound ascidian Diplosoma listerianum. Marine Ecology Progress Series, 61:125– 132, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/61/m061p125.pdf.

Ricciardi:1998:WWC

[RB98] A. Ricciardi and E. Bourget. Weight-to-weight conversion factors for marine benthic macroinvertebrates. Marine Ecology Progress Series, 163:245–251, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/163/m163p245.pdf.

Ricciardi:1999:GPM

[RB99] A. Ricciardi and E. Bourget. Global patterns of macroinvertebrate biomass in marine intertidal communities. Marine Ecology Progress Series, 185:21–35, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/185/m185p021.pdf.

Rasmussen:1998:EBL

[RBA98] A. D. Rasmussen, G. T. Banta, and O. Andersen. Effects of bioturbation by the lugworm Arenicola marina on cadmium uptake and distribution in sandy sediments. Marine Ecology Progress Series, 164:179–188, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/164/m164p179.pdf.

Richard:1991:ESR

[RBB^{+}91] P. Richard, J.-P. Bergeron, M. Boulhic, R. Galois, and J. Person-Le Ruyet. Effect of starvation on RNA, DNA and protein content of laboratory-reared larvae and juveniles of Solea solea. Marine Ecology Progress Series, 72:69–77, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
72/m072p069.pdf.

Robblee:1991:MMT

[RBC^{+}91] M. B. Robblee, T. R. Barber, P. R. Carlson, Jr., M. J. Durako, J. W. Fourqurean, L. K. Muehlstein, D. Porter, L. A. Yarbro, R. T. Zieman, and J. C. Zieman. Mass mortality of the tropical seagrass Thalassia testudinum in Florida Bay (USA). Marine Ecology Progress Series, 71:297–299, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
71/m071p297.pdf.

Raabe:1997:NPD

[RBD^{+}97] T. U. Raabe, U. H. Brockmann, C. D. Dürselen, M. Krause, and H. J. Rick. Nutrient and plankton dynamics during a spring drift experiment in the German Bight. Marine Ecology Progress Series, 156:275–288, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/156/m156p275.pdf.

Rakocinski:1992:CBE

[RBF92] C. F. Rakocinski, D. M. Baltz, and J. W. Fleeger. Correspondence between environmental gradients and the community structure of marsh-edge fishes in a Louisiana estuary. Marine Ecology Progress Series, 80:135–148, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
80/m080p135.pdf.

Ruiz:1994:BTT

[RBG94a] J. M. Ruiz, G. W. Bryan, and P. E. Gibbs. Bioassaying the toxicity of tributyltin-(TBT)-polluted sediment to spat of the bivalve Scrobicularia plana. Marine Ecology Progress Series, 113:119–130, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/113/m113p119.pdf.

Ruiz:1994:CTW

[RBG94b] J. M. Ruiz, G. W. Bryan, and P. E. Gibbs. Chronic toxicity of water tributyltin (TBT) and copper to spat of the bivalve Scrobiculaha plana: ecological implications. Marine Ecology Progress Series, 113:105–117, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/113/m113p105.pdf.

Riemann:1990:ITA

[RBJ90] B. Riemann, R. T. Bell, and N. O. G. Jorgensen. Incorporation of thymidine, adenine and leucine into natural bacterial assemblages. Marine Ecology Progress Series, 65:87–94, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
65/m065p087.pdf.

Rotchell:1999:SVE

[RBN99] J. M. Rotchell, D. J. Bird, and L. C. Newton. Seasonal variation in ethoxyresorufin O-deethylase (EROD) activity in European eels Anguilla anguilla and flounders Pleuronectes flesus from the Severn Estuary and Bristol Channel. Marine Ecology Progress Series, 190:263–270, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
190/m190p263.pdf.

Reusch:1999:AEC

[RBSO99] T. B. H. Reusch, C. Boström, W. T. Stam, and J. L. Olsen. An ancient eelgrass clone in the Baltic. Marine Ecology Progress Series, 183:301–304, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/183/m183p301.pdf.

Riisgaard:1996:LAM

[RBT96] H. U. Riisgård, I. Berntsen, and B. Tarp. The lugworm (Arenicola marina) pump: characteristics, modelling and energy cost. Marine Ecology Progress Series, 138:149–156, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
138/m138p149.pdf.

Richardson:1991:SDP

[RC91a] K. Richardson and A. Christoffersen. Seasonal distribution and production of phytoplankton in the southern Kattegat. Marine Ecology Progress Series, 78:217–227, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
78/m078p217.pdf.

Riegman:1991:EMC

[RC91b] R. Riegman and F. Colijn. Evaluation of measurements and calculation of primary production in the Dogger Bank area (North Sea) in summer 1988. Marine Ecology Progress Series, 69:125–132, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/69/m069p125.pdf.

Richardson:1995:CBC

[RC95] T. L. Richardson and J. J. Cullen. Changes in buoyancy and chemical composition during growth of a coastal marine diatom: ecological and biogeochemical consequences. Marine Ecology Progress Series, 128:77–90, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/128/m128p077.pdf.

Rosenbaum:1995:GVV

[RCA^{+}95] H. C. Rosenbaum, P. J. Clapham, J. Allen, M. Nicole-Jenner, C. Jenner, L. Florez-González, J. R. Urbán, P. G. Ladrón, K. Mori, M. Yamaguchi, and C. S. Baker. Geographic variation in ventral fluke pigmentation of humpback whale Megaptera novaeangliae populations worldwide. Marine Ecology Progress Series, 124:1–7, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/124/m124p001.pdf.

Rose:1999:IBM

[RCC^{+}99] K. A. Rose, J. H. Cowan, Jr., M. E. Clark, E. D. Houde, and S. B. Wang. An individual-based model of bay anchovy population dynamics in the mesohaline region of Chesapeake Bay. Marine Ecology Progress Series, 185:113–132, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
185/m185p113.pdf.

Reusch:1994:BMM

[RCG94] T. B. H. Reusch, A. R. O. Chapman, and J. P. Groger. Blue mussels Mytilus edulis do not interfere with eelgrass Zostera marina but fertilize shoot growth through biodeposition. Marine Ecology Progress Series, 108:265–282, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
108/m108p265.pdf.

Ribes:1998:SVL

[RCG98] M. Ribes, R. Coma, and J. M. Gili. Seasonal variation of in situ feeding rates by the temperate ascidian Halocynthia papillosa. Marine Ecology Progress Series, 175:201–213, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
175/m175p201.pdf.

Ribes:1999:HFB

[RCG99a] M. Ribes, R. Coma, and J. M. Gili. Heterogeneous feeding in benthic suspension feeders: the natural diet and grazing rate of the temperate gorgonian Paramuricea clavata (Cnidaria: Octocorallia) over a year cycle. Marine Ecology Progress Series, 183:125–137, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/183/m183p125.pdf.

Ribes:1999:NDG

[RCG99b] M. Ribes, R. Coma, and J. M. Gili. Natural diet and grazing rate of the temperate sponge Dysidea avara (Demospongiae, Dendroceratida) throughout an annual cycle. Marine Ecology Progress Series, 176:179–190, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/176/m176p179.pdf.

Roy:1996:CPC

[RCGSN96] S. Roy, J. P. Chanut, M. Gosselin, and T. Sime-Ngando. Characterization of phytoplankton communities in the lower St. Lawrence Estuary using HPLC-detected pigments and cell microscopy. Marine Ecology Progress Series, 142:55–73, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
142/m142p055.pdf.

Ringwood:1998:ENA

[RCH98] A. H. Ringwood, D. E. Conners, and J. Hoguet. Effects of natural and anthropogenic stressors on lysosomal destabilization in oysters Crassostrea virginica. Marine Ecology Progress Series, 166:163–171, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/166/m166p163.pdf.

Rysgaard:1995:SVN

[RCN95] S. Rysgaard, P. B. Christensen, and L. P. Nielsen. Seasonal variation in nitrification and denitrification in estuarine sediment colonized by benthic microalgae and bioturbating infauna. Marine Ecology Progress Series, 126:111–121, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
126/m126p111.pdf.

Rahm:1996:TSA

[RCS^{+}96] L. Rahm, D. Conley, P. Sandén, F. Wulff, and P. Stålnacke. Time series analysis of nutrient inputs to the Baltic Sea and changing DSi: DIN ratios. Marine Ecology Progress Series, 130:221–228, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/130/m130p221.pdf.

Rey:1999:EFP

[RCTH99] C. Rey, F. Carlotti, K. Tande, and B. H. Hygum. Egg and faecal pellet production of Calanus finmarchicus females from controlled mesocosms and in situ populations: influence of age and feeding history. Marine Ecology Progress Series, 188:133–148, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/188/m188p133.pdf.

Redden:1991:VSC

[RD91] A. M. Redden and G. R. Daborn. Viability of subitaneous copepod eggs following fish predation on egg-carrying calanoids. Marine Ecology Progress Series, 77:307–310, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
77/m077p307.pdf.

Rodriguez:1992:ESF

[RD92] V. Rodriguez and E. G. Durbin. Evaluation of synchrony of feeding behaviour in individual Acartia hudsonica (Copopoda, Calanoida). Marine Ecology Progress Series, 87:7–13, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/87/m087p007.pdf.

Rowe:1995:IIE

[RD95] C. L. Rowe and W. A. Dunson. Individual and interactive effects of salinity and initial fish density on a salt marsh assemblage. Marine Ecology Progress Series, 128:271–278, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
128/m128p271.pdf.

Rippe:1997:PPJ

[RD97] H. Rippe and V. Dierschke. Picking out the plum jobs: feeding ecology of curlews Numenius arquata in a Baltic Sea wind flat. Marine Ecology Progress Series, 159:239–247, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
159/m159p239.pdf.

Rose:1999:ECS

[RD99] C. D. Rose and C. J. Dawes. Effects of community structure on the seagrass Thalassia testudinum. Marine Ecology Progress Series, 184:83–95, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/184/m184p083.pdf.

Rodriguez:1997:RPL

[RDA97] A. Rodríguez, P. Drake, and A. M. Arias. Reproductive periods and larval abundance patterns of the crabs Panopeus africanus and Uca tangeri in a shallow inlet (SW Spain). Marine Ecology Progress Series, 149:133–142, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
149/m149p133.pdf.

Rosenberg:1990:PNEa

[RDE^{+}90a] R. Rosenberg, E. Dahl, L. Edler, L. Fyrberg, E. Graneli, W. Graneli, Å. Hägström, O. Lindahl, M. O. Matos, K. Pettersson, E. Sahlsten, P. Tiselius, V. Turk, and J. Wikner. Pelagic nutrient and energy transfer during spring in the open and coastal Skagerrak. Marine Ecology Progress Series, 61:215–231, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/61/m061p215.pdf.

Rosenberg:1990:PNEb

[RDE^{+}90b] R. Rosenberg, E. Dahl, L. Edler, L. Fyrberg, E. Graneli, W. Graneli, Å. Hägström, O. Lindahl, M. O. Matos, K. Pettersson, E. Sahlsten, P. Tiselius, V. Turk, and J. Wikner. Pelagic nutrient and energy transfer during spring in the open and coastal Skagerrak. Marine Ecology Progress Series, 61:215–231, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/61/m061p215.pdf.

Rodger:1991:RSS

[RDM91] G. K. Rodger, I. M. Davies, and D. C. Moore. Recovery of a sewage sludge dumping ground. I. Trace metal concentrations in the sediment. Marine Ecology Progress Series, 75:293–299, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/75/m075p293.pdf.

Raimbault:1999:SDP

[RDPB99] P. Raimbault, F. Diaz, W. Pouvesle, and B. Boudjellal. Simultaneous determination of particulate organic carbon, nitrogen and phosphorus collected on filters, using a semiautomatic wet-oxidation method. Marine Ecology Progress Series, 180:289–295, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/180/m180p289.pdf.

Ragueneau:1994:PDR

[RDT^{+}94] O. Ragueneau, E. De Bias Varela, P. Tréguer, B. Quéguiner, and Y. Del Amo. Phytoplankton dynamics in relation to the biogeochemical cycle of silicon in a coastal ecosystem of western Europe. Marine Ecology Progress Series, 106:157– 172, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/106/m106p157.pdf.

Rowe:1994:FTL

[RE94a] P. M. Rowe and C. E. Epifanio. Flux and transport of larval weakfish in Delaware Bay, USA. Marine Ecology Progress Series, 110:115–120, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/110/m110p115.pdf.

Rowe:1994:TST

[RE94b] P. M. Rowe and C. E. Epifanio. Tidal stream transport of weakfish larvae in Delaware Bay, USA. Marine Ecology Progress Series, 110:105–114, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/110/m110p105.pdf.

Roa:1996:ASA

[RE96] R. Roa and B. Ernst. Age structure, annual growth, and variance of size-at-age of the shrimp Heterocarpus reedi. Marine Ecology Progress Series, 137:59–70, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/137/m137p059.pdf.

Regoli:1992:LRS

[Reg92] F. Regoli. Lysosomal responses as a sensitive stress index in biomonitoring heavy metal pollution. Marine Ecology Progress Series, 84:63–69, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/84/m084p063.pdf.

Reilly:1990:SCD

[Rei90] S. B. Reilly. Seasonal changes in distribution and habitat differences among dolphins in the eastern tropical Pacific. Marine Ecology Progress Series, 66:1–11, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
66/m066p001.pdf.

Relexans:1996:MREa

[Rel96a] J. C. Relexans. Measurement of the respiratory electron transport system (ETS) activity in marine sediments: state-of-the-art and interpretation. I. Methodology and review of literature data. Marine Ecology Progress Series, 136:277–287, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
136/m136p277.pdf.

Relexans:1996:MREb

[Rel96b] J. C. Relexans. Measurement of the respiratory electron transport system (ETS) activity in marine sediments: state-of-the-art and interpretation. II. Significance of ETS activity data. Marine Ecology Progress Series, 136:289–301, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
136/m136p289.pdf.

Reusch:1998:DEE

[Reu98a] T. B. H. Reusch. Differing effects of eelgrass Zostera marina on recruitment and growth of associated blue mussels Mytilus edulis. Marine Ecology Progress Series, 167:149–153, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
167/m167p149.pdf.

Reusch:1998:NPC

[Reu98b] T. B. H. Reusch. Native predators contribute to invasion resistance to the non-indigenous bivalve Musculista senhousia in southern California, USA. Marine Ecology Progress Series, 170:159–168, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/170/m170p159.pdf.

Ronnestad:1992:UFA

[RFGF92] I. Rønnestad, R. N. Finn, E. P. Groot, and H. J. Fyhn. Utilization of free amino acids related to energy metabolism of developing eggs and larvae of lemon sole Microslomus kitt reared in the laboratory. Marine Ecology Progress Series, 88: 195–205, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/88/m088p195.pdf.

Riegman:1998:SFUa

[RFN98] R. Riegman, I. A. Flameling, and A. A. M. Noordeloos. Size-fractionated uptake of ammonium, nitrate and urea and phytoplankton growth in the North Sea during spring 1994. Marine Ecology Progress Series, 173:85–94, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
173/m173p085.pdf.

Raya:1994:PTA

[RFNBHG94] C. P. Raya, M. Fernandez-Nunez, E. Balguerias, and C. L. Hernandez-Gonzalez. Progress towards ageing cuttlefish Sepia hierredda from the northwestern African coast using statoliths. Marine Ecology Progress Series, 114:139–147, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/114/m114p139.pdf.

Re:1993:GSS

[RG93] P. Ré and E. Gonçalves. Growth of sprat Sprattus sprattus larvae in the German Bight (North Sea) as inferred by otolith microstructure. Marine Ecology Progress Series, 96: 139–145, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/96/m096p139.pdf.

Riisgaard:1997:MSL

[RG97] H. U. Riisgård and A. Goldson. Minimal scaling of the lophophore filter-pump in ectoprocts (Bryozoa) excludes physiological regulation of filtration rate to nutritional needs. Test of hypothesis. Marine Ecology Progress Series, 156: 109–120, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/156/m156p109.pdf.

Romeo:1992:IGP

[RGBC92] M. Romeo, M. Gnassia-Barelli, and C. Carre. Importance of gelatinous plankton organisms in storage and transfer of trace metals in the northwestern Mediterranean. Marine Ecology Progress Series, 82:267–274, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/82/m082p267.pdf.

Romeo:1995:MHM

[RGBJM95] M. Roméo, M. Gnassia-Barelli, T. Juhel, and A. Meinesz. Memorization of heavy metals by scales of the seagrass Posidonia oceanica, collected in the NW Mediterranean. Marine Ecology Progress Series, 120:211–218, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
120/m120p211.pdf.

Roa:1995:NGA

[RGE^{+}95] R. Roa, V. A. Gallardo, B. Ernst, M. Baltazar, J. I. Canete, and S. Enriquez-Brionnes. Nursery ground, age structure and abundance of juvenile squat lobster Pleuroncodes monodon on the continental shelf off central Chile. Marine Ecology Progress Series, 116:47–54, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/116/m116p047.pdf.

Ramirez-Garcia:1998:BDB

[RGLD^{+}98] P. Ramírez-García, A. Lot, C. M. Duarte, J. Terrados, and N. S. R. Agawin. Bathymetric distribution, biomass and growth dynamics of intertidal Phyllospadix scouleri and Phyllospadix torreyi in Baja California (Mexico). Marine Ecology Progress Series, 173:13–23, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/173/m173p013.pdf.

Ralph:1999:LSU

[RGLS99] P. J. Ralph, R. Gademann, A. W. D. Larkum, and U. Schreiber. In situ underwater measurements of photosynthetic activity of coral zooxanthellae and other reef-dwelling dinoflagellate endosymbionts. Marine Ecology Progress Series, 180:139–147, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/180/m180p139.pdf.

Raimbault:1990:STR

[RGS90] P. Raimbault, V. Gentilhomme, and G. Slawyk. Short-term responses of 24 hour n-starved cultures of Phaeodactylum tricornutum to pulsed additions of nitrate at nanomolar levels. Marine Ecology Progress Series, 63:47–52, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
63/m063p047.pdf.

Reay:1995:SWC

[RGS95] W. G. Reay, D. L. Gallagher, and G. M. Simmons, Jr. Sediment-water column oxygen and nutrient fluxes in nearshore environments of the lower Delmarva Peninsula, USA. Marine Ecology Progress Series, 118:215–227, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
118/m118p215.pdf.

Rouleau:1998:PDD

[RGT98] C. Rouleau, C. Gobeil, and H. Tjälve. Pharmacokinetics and distribution of dietary tributyltin compared to those of methylmercury in the American plaice Hippoglossoides platessoides. Marine Ecology Progress Series, 171:275–284, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
171/m171p275.pdf.

Richmond:1990:RRC

[RH90a] R. H. Richmond and C. L. Hunter. Reproduction and recruitment of corals: comparisons among the Caribbean, the Tropical Pacific, and the Red Sea. Marine Ecology Progress Series, 60:185–203, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/60/m060p185.pdf.

Riisgaard:1990:BMM

[RH90b] H. U. Riisgård and S. Hansen. Biomagnification of mercury in a marine grazing food-chain: algal cells Phaeodactylum tricornutum, mussels Mytilus edulis and flounders Platichthys flesus studied by means of a stepwise-reduction-CVAA method. Marine Ecology Progress Series, 62:259–270, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
62/m062p259.pdf.

Ruesink:1990:SPC

[RH90c] J. L. Ruesink and C. D. Harvell. Specialist predation on the Caribbean gorgonian Plexaurella spp. by Cyphoma signatum (Gastropoda). Marine Ecology Progress Series, 65: 265–272, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/65/m065p265.pdf.

Riemann:1991:ECD

[RH91] B. Riemann and E. Hoffmann. Ecological consequences of dredging and bottom trawling in the Limfjord, Denmark. Marine Ecology Progress Series, 69:171–178, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
69/m069p171.pdf.

Rooker:1997:USS

[RH97] J. R. Rooker and S. A. Holt. Utilization of subtropical seagrass meadows by newly settled red drum Sciaenops ocellatus: patterns of distribution and growth. Marine Ecology Progress Series, 158:139–149, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/158/m158p139.pdf.

Rand:1998:SPZ

[RH98] P. S. Rand and S. G. Hinch. Spatial patterns of zooplankton biomass in the northeast Pacific Ocean. Marine Ecology Progress Series, 171:181–186, 1998. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/171/m171p181.pdf.

Riegl:1996:FFS

[RHB96] B. Riegl, C. Heine, and G. M. Branch. Function of funnelshaped coral growth in a high-sedimentation environment. Marine Ecology Progress Series, 145:87–93, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
145/m145p087.pdf.

Rochette:1994:FSA

[RHH94] R. Rochette, J.-F. Hamel, and J. H. Himmelman. Foraging strategy of the asteroid Leptasterias polaris: role of prey odors, current and feeding status. Marine Ecology Progress Series, 106:93–100, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/106/m106p093.pdf.

Richardson:1997:TSF

[RHH97] D. L. Richardson, P. L. Harrison, and V. J. Harriott. Timing of spawning and fecundity of a tropical and subtropical anemonefish (Pomacentridae: Amphiprion) on a high latitude reef on the east coast of Australia. Marine Ecology Progress Series, 156:175–181, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/156/m156p175.pdf.

Rosenberg:1991:HTM

[RHJ91] R. Rosenberg, B. Hellman, and B. Johansson. Hypoxic tolerance of marine benthic fauna. Marine Ecology Progress Series, 79:127–131, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/79/m079p127.pdf.

Rudstam:1992:DPC

[RHJ92] L. G. Rudstam, S. Hansson, and U. Johansson, S.. Larsson. Dynamics of planktivory in a coastal area of the northern Baltic Sea. Marine Ecology Progress Series, 80:159–173, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
80/m080p159.pdf.

Ruiz:1993:SWR

[RHP93] G. M. Ruiz, A. H. Hines, and M. H. Posey. Shallow water as a refuge habitat for fish and crustaceans in non-vegetated estuaries: an example from Chesapeake Bay. Marine Ecology Progress Series, 99:1–6, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/99/m099p001.pdf.

Russell:1999:FFH

[RHrsHG99] R. W. Russell, N. M. Harrison, and J. r. s. Hunt GL. Foraging at a front: hydrography, zooplankton, and avian planktivory in the northern Bering Sea. Marine Ecology Progress Series, 182:77–93, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/182/m182p077.pdf.

Rigby:1999:NBW

[RHS99] G. R. Rigby, G. M. Hallegraeff, and C. Sutton. Novel ballast water heating technique offers cost-effective treatment to reduce the risk of global transport of harmful marine organisms. Marine Ecology Progress Series, 191:289–293, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
191/m191p289.pdf.

Rau:1991:SSW

[RHT91] G. H. Rau, T. L. Hopkins, and J. J. Torres. ^{15}N/^{14}N and 13C/^{12}C in Weddell Sea invertebrates — implications for feeding diversity. Marine Ecology Progress Series, 77:1–6, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
77/m077p001.pdf.

Riisgaard:1990:CFP

[RI90] H. U. Riisgård and N. M. Ivarsson. The crown-filament pump of the suspension-feeding polychaete Sabella penicillus: filtration, effects of temperature, and energy cost. Marine Ecology Progress Series, 62:249–257, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/62/m062p249.pdf.

Richardson:1991:CPP

[Ric91] K. Richardson. Comparison of ^{14}C primary production determinations made by different laboratories. Marine Ecology Progress Series, 72:189–201, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/72/m072p189.pdf.

Riebesell:1991:PADa

[Rie91a] U. Riebesell. Particle aggregation during a diatom bloom. I. Physical aspects. Marine Ecology Progress Series, 69:273– 280, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/69/m069p273.pdf.

Riebesell:1991:PADb

[Rie91b] U. Riebesell. Particle aggregation during a diatom bloom. II. Biological aspects. Marine Ecology Progress Series, 69: 281–291, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/69/m069p281.pdf.

Riebesell:1993:APD

[Rie93] U. Riebesell. Aggregation of Phaeocystis during phytoplankton spring blooms in the southern North Sea. Marine Ecology Progress Series, 96:281–289, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/96/m096p281.pdf.

Riera:1998:OMS

[Rie98] P. Riera. δ ^{15}N of organic matter sources and benthic invertebrates along an estuarine gradient in Marennes–Oléron Bay (France): implications for the study of trophic structure. Marine Ecology Progress Series, 166:143–150, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
166/m166p143.pdf.

Riisgaard:1991:SFP

[Rii91] H. U. Riisgård. Suspension feeding in the polychaete Nereis diversicolor. Marine Ecology Progress Series, 70:29–37, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
70/m070p029.pdf.

Richardson:1993:EPS

[RII93] C. A. Richardson, I. Ibarrola, and R. J. Ingham. Emergence pattern and spatial distribution of the common cockle Cerastoderma edule. Marine Ecology Progress Series, 99: 71–81, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/99/m099p071.pdf.

Ringwood:1992:ECC

[Rin92] A. H. Ringwood. Effects of chronic cadmium exposures on growth of larvae of an Hawaiian bivalve, Isognomon californicum. Marine Ecology Progress Series, 83:63–70, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
83/m083p063.pdf.

Rinkevich:1996:DRR

[Rin96] B. Rinkevich. Do reproduction and regeneration in damaged corals compete for energy allocation? Marine Ecology Progress Series, 143:297–302, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/143/m143p297.pdf.

Risk:1997:EHS

[Ris97] A. Risk. Effects of habitat on the settlement and postsettlement success of the ocean surgeonfish Acanthurus bahianus. Marine Ecology Progress Series, 161:51–59, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
161/m161p051.pdf.

Rivkin:1990:PMP

[Riv90] R. B. Rivkin. Photoadaptation in marine phytoplankton: variations in ribulose 1,5-bisphos-phate activity. Marine Ecology Progress Series, 62:61–72, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/62/m062p061.pdf.

Rainbow:1995:PRU

[RK95a] P. S. Rainbow and M. K. H. Kwan. Physiological responses and the uptake of cadmium and zinc by the amphipod crustacean Orchestia gammarellus. Marine Ecology Progress Series, 127:87–102, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/127/m127p087.pdf.

Rangeley:1995:TEH

[RK95b] R. W. Rangeley and D. L. Kramer. Tidal effects on habitat selection and aggregation by juvenile pollock Pollachius virens in the rocky intertidal zone. Marine Ecology Progress Series, 126:19–29, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/126/m126p019.pdf.

Rangeley:1995:URI

[RK95c] R. W. Rangeley and D. L. Kramer. Use of rocky intertidal habitats by juvenile pollock Pollachius virens. Marine Ecology Progress Series, 126:9–17, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/126/m126p009.pdf.

Rakitin:1996:EMR

[RK96] A. Rakitin and D. L. Kramer. Effect of a marine reserve on the distribution of coral reef fishes in Barbados. Marine Ecology Progress Series, 131:97–113, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/131/m131p097.pdf.

Rumohr:1999:CMP

[RK99] H. Rumohr and I. Karakassis. Comparison of multivariate patterns: different taxonomic levels in macrofaunal analysis versus sediment profiling imagery (SPI). Marine Ecology Progress Series, 190:125–132, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/190/m190p125.pdf.

Ramsay:1996:CHC

[RKH96] K. Ramsay, M. J. Kaiser, and R. N. Hughes. Changes in hermit crab feeding patterns in response to trawling disturbance. Marine Ecology Progress Series, 144:63–72, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
144/m144p063.pdf.

Robinson:1997:PPS

[RKS97] D. H. Robinson, Z. Kolber, and C. W. Sullivan. Photophysiology and photoacclimation in surface sea ice algae from McMurdo Sound, Antarctica. Marine Ecology Progress Series, 147:243–256, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/147/m147p243.pdf.

Rizzo:1992:SES

[RLC92] W. M. Rizzo, G. L. Lackey, and R. R. Christian. Significance of euphotic sediments to oxygen and nutrient cycling in a temperate estuary. Marine Ecology Progress Series, 86: 51–61, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/86/m086p051.pdf.

Raby:1994:RBF

[RLDM94] D. Raby, Y. Lagadeuc, J. J. Dodson, and M. Mingelbier. Relationship between feeding and vertical distribution of bivalve larvae in stratified and mixed waters. Marine Ecology Progress Series, 103:275–284, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/103/m103p275.pdf.

Robineau:1994:UAP

[RLT^{+}94] B. Robineau, L. Legendre, J.-C. Therriault, L. Fortier, G. Rosenberg, and S. Demers. Ultra-algae (< 5 pm) in the ice, at the ice–water interface and in the under-ice water column (southeastern Hudson Bay, Canada). Marine Ecology Progress Series, 115:169–180, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/115/m115p169.pdf.

Raillard:1994:EBM

[RM94] O. Raillard and A. Ménesguen. An ecosystem box model for estimating the carrying capacity of a macrotidal shellfish system. Marine Ecology Progress Series, 115:117–130, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
115/m115p117.pdf.

Roegner:1995:ERG

[RM95] G. Curtis Roegner and Roger Mann. Early recruitment and growth of the American oyster Crassostrea virginica (Bivalvia: Ostreidae) with respect to tidal zonation and season. Marine Ecology Progress Series, 117:91–101, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
117/m117p091.pdf.

Rodway:1996:SMA

[RM96] M. S. Rodway and W. A. Montevecchi. Sampling methods for assessing the diets of Atlantic puffin chicks. Marine Ecology Progress Series, 144:41–55, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/144/m144p041.pdf.

Regehr:1997:IEF

[RM97a] H. M. Regehr and W. A. Montevecchi. Interactive effects of food shortage and predation on breeding failure of blacklegged kittiwakes: indirect effects of fisheries activities and implications for indicator species. Marine Ecology Progress Series, 155:249–260, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/155/m155p249.pdf.

Riisgaard:1997:FFF

[RM97b] H. U. Riisgård and P. Manríquez. Filter-feeding in fifteen marine ectoprocts (Bryozoa): particle capture and water pumping. Marine Ecology Progress Series, 154:223–239, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
154/m154p223.pdf.

Rosas:1998:EDO

[RMG^{+}98] C. Rosas, E. Martinez, G. Gaxiola, R. Brito, E. Díaz-Iglesia, and L. A. Soto. Effect of dissolved oxygen on the energy balance and survival of Penaeus setiferus juveniles. Marine Ecology Progress Series, 174:67–75, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/174/m174p067.pdf.

Romeo:1994:HMC

[RMGB^{+}94] M. Roméo, A. Mathieu, M. Gnassia-Barelli, A. Romana, and M. Lafaurie. Heavy metal content and biotransformation enzymes in two fish species from the NW Mediterranean. Marine Ecology Progress Series, 107:15–22, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/107/m107p015.pdf.

Rochette:1996:IIP

[RMH96] R. Rochette, J. N. McNeil, and J. H. Himmelman. Interand intra-population variations in the response of the whelk Buccinum undatum to the predatory asteroid Leptasterias polaris. Marine Ecology Progress Series, 142:193–201, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
142/m142p193.pdf.

Roche-Mayzaud:1991:MTA

[RMMB91] O. Roche-Mayzaud, P. Mayzaud, and D. C. Biggs. Mediumterm acclimation of feeding and of digestive and metabolic enzyme activity in the neritic copepod Acartia clausi. I. Evidence from laboratory experiments. Marine Ecology Progress Series, 69:25–40, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/69/m069p025.pdf.

Rogers:1991:EHH

[RMT91] C. S. Rogers, L. N. McLain, and C. R. Tobias. Effects of Hurricane Hugo (1989) on a coral reef in St. John, USVI. Marine Ecology Progress Series, 78:189–199, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
78/m078p189.pdf.

Rivera-Monroy:1995:DDM

[RMTB^{+}95] V. H. Rivera-Monroy, R. R. Twilley, R. G. Boustany, J. W. Day, F. Vera-Herrera, and M. del Carmen Ramirez. Direct denitrification in mangrove sediments in Terminos Lagoon, Mexico. Marine Ecology Progress Series, 126:97–109, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
126/m126p097.pdf.

Reid:1993:DFR

[RN93] D. G. Reid and E. Naylor. Different free-running periods in split components of the circatidal rhythm in the shore crab Carcinus maenas. Marine Ecology Progress Series, 102: 295–302, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/102/m102p295.pdf.

Riegman:1998:SFUb

[RN98] R. Riegman and A. A. M. Noordeloos. Size-fractionated uptake of nitrogenous nutrients and carbon by phytoplankton in the North Sea during summer 1994. Marine Ecology Progress Series, 173:95–106, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/173/m173p095.pdf.

Rysgaard:1999:SVN

[RNH99] S. Rysgaard, T. G. Nielsen, and B. W. Hansen. Seasonal variation in nutrients, pelagic primary production and grazing in a high-Arctic coastal marine ecosystem, Young Sound, Northeast Greenland. Marine Ecology Progress Series, 179: 13–25, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/179/m179p013.pdf.

Rosenberg:1997:DDM

[RNHH97] R. Rosenberg, H. C. Nilsson, K. Hollertz, and B. Hellman. Density-dependent migration in an Amphiura filiformis (Amphiuridae, Echinodermata) infaunal population. Marine Ecology Progress Series, 159:121–131, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/159/m159p121.pdf.

Robinson:1998:HES

[RNHO98] A. D. Robinson, D. B. Nedwell, R. M. Harrison, and B. G. Ogilvie. Hypernutrified estuaries as sources of N_{2}O emission to the atmosphere: the estuary of the River Colne, Essex, UK. Marine Ecology Progress Series, 164:59–71, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
164/m164p059.pdf.

Richardson:1998:SHS

[RNP^{+}98] K. Richardson, T. G. Nielsen, F. B. Pedersen, J. P. Heilmann, B. Løkkegaard, and H. Kaas. Spatial heterogeneity in the structure of the planktonic food web in the North Sea. Marine Ecology Progress Series, 168:197–211, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
168/m168p197.pdf.

Rodriguez:1993:DPT

[RO93] S. R. Rodriguez and F. P. Ojeda. Distribution patterns of Tetrapygus niger (Echinodermata: Echinoidea) off the central Chilean coast. Marine Ecology Progress Series, 101: 157–162, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/101/m101p157.pdf.

Ryer:1997:ASS

[RO97] C. H. Ryer and B. L. Olla. Altered search speed and growth: social versus independent foraging in two pelagic juvenile fishes. Marine Ecology Progress Series, 153:273–281, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
153/m153p273.pdf.

Ryer:1998:ELJ

[RO98] C. H. Ryer and B. L. Olla. Effect of light on juvenile walleye pollock shoaling and their interaction with predators. Marine Ecology Progress Series, 167:215–226, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
167/m167p215.pdf.

Ryer:1999:LIC

[RO99] C. H. Ryer and B. L. Olla. Light-induced changes in the prey consumption and behavior of two juvenile planktivorous fish. Marine Ecology Progress Series, 181:41–51, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
181/m181p041.pdf.

Roa:1993:AGM

[Roa93] R. Roa. Annual growth and maturity function of the squat lobster Pleuroncodes monodon in central Chile. Marine Ecology Progress Series, 97:157–166, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/97/m097p157.pdf.

Rodhouse:1998:CPK

[ROANM98] P. G. Rodhouse, O. Olsson, P. Anker-Nilssen, and A. W. A. Murray. Cephalopod predation by the king penguin Aptenodytes patagonicus from South Georgia. Marine Ecology Progress Series, 168:13–19, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/168/m168p013.pdf.

Rogers:1990:RCR

[Rog90] C. S. Rogers. Responses of coral reefs and reef organisms to sedimentation. Marine Ecology Progress Series, 62:185– 202, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/62/m062p185.pdf.

Rodriguez:1993:SBM

[ROI93] S. R. Rodriguez, F. P. Ojeda, and N. C. Inestrosa. Settlement of benthic marine invertebrates. Marine Ecology Progress Series, 97:193–207, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/97/m097p193.pdf.

Regoli:1991:HMA

[ROM^{+}91] F. Regoli, E. Orlando, M. Mauri, M. Nigro, and G. A. Cognetti. Heavy metal accumulation and calcium content in the bivalve Donacilla cornea. Marine Ecology Progress Series, 74:219–224, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/74/m074p219.pdf.

Ropke:1993:DLM

[Röp93] A. Röpke. Do larvae of mesopelagic fishes in the Arabian Sea adjust their vertical distribution to physical and biological gradients? Marine Ecology Progress Series, 101:223– 235, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/101/m101p223.pdf.

Rowley:1990:NSS

[Row90] R. J. Rowley. Newly settled sea urchins in a kelp bed and urchin barren ground, a comparison of growth and mortality. Marine Ecology Progress Series, 62:229–240, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
62/m062p229.pdf.

Rozas:1992:BLN

[Roz92] L. P. Rozas. Bottomless lift net for quantitatively sampling nekton on intertidal marshes. Marine Ecology Progress Series, 89:287–292, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/89/m089p287.pdf.

Rogers:1991:CDT

[RP91a] S. D. Rogers and V. J. Paul. Chemical defenses of three Glossodoris nudibranchs and their dietary Hyrtios sponges. Marine Ecology Progress Series, 77:221–232, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
77/m077p221.pdf.

Rowan:1991:MGI

[RP91b] R. Rowan and D. A. Powers. Molecular genetic identification of symbiotic dinoflagellates (zooxanthellae). Marine Ecology Progress Series, 71:65–73, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/71/m071p065.pdf.

Risgaard-Petersen:1998:NBT

[RPDR^{+}98] N. Risgaard-Petersen, T. Dalsgaard, S. Rysgaard, P. B. Christensen, J. Borum, K. McGlathery, and L. P. Nielsen. Nitrogen balance of a temperate eelgrass Zostera marina bed. Marine Ecology Progress Series, 174:281–291, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
174/m174p281.pdf.

Riisgaard:1996:PRN

[RPL96] H. U. Riisgård, L. Poulsen, and P. S. Larsen. Phytoplankton reduction in near-bottom water caused by filter-feeding Nereis diversicolor — implications for worm growth and population grazing impact. Marine Ecology Progress Series, 141: 47–54, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/141/m141p047.pdf.

Rodhouse:1994:ULB

[RPM^{+}94] P. G. Rodhouse, U. Piatkowski, E. J. Murphy, M. G. White, and D. G. Bone. Utility and limits of biomass spectra: the nekton community sampled with the RMT 25 in the Scotia Sea during austral summer. Marine Ecology Progress Series, 112:29–39, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/112/m112p029.pdf.

Rudek:1991:SHC

[RPMB91] J. Rudek, H. W. Paerl, M. A. Mallin, and P. W. Bates. Seasonal and hydrological control of phytoplankton nutrient limitation in the lower Neuse River Estuary, North Carolina. Marine Ecology Progress Series, 75:133–142, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
75/m075p133.pdf.

Richards:1995:LDA

[RPN95] S. A. Richards, H. P. Possingham, and B. J. Noye. Larval dispersion along a straight coast with tidal currents: complex distribution patterns from a simple model. Marine Ecology Progress Series, 122:59–71, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/122/m122p059.pdf.

Rouleau:1995:DKT

[RPT95] C. Rouleau, E. Pelletier, and H. Tjälve. Distribution kinetics of trophic single doses of methylmercury, tributyltin, and corresponding inorganic ions in the starfish Leptasterias polaris. Marine Ecology Progress Series, 124:143–158, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
124/m124p143.pdf.

Rodhouse:1996:CMO

[RPT^{+}96] P. G. Rodhouse, P. A. Prince, P. N. Trathan, E. M. C. Hatfield, J. L. Watkins, D. G. Bone, E. J. Murphy, and M. G. White. Cephalopods and mesoscale oceanography at the Antarctic polar front: satellite tracked predators locate pelagic trophic interactions. Marine Ecology Progress Series, 136:37–50, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/136/m136p037.pdf.

Ruiz:1998:TIN

[RQB98a] J. M. Ruiz, M. Quintela, and R. Barreiro. Tributyltin and imposex: no uncertainty shown. Marine Ecology Progress Series, 170:293–294, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/170/m170p293.pdf. See reply

[DHB^{+}98].

Ruiz:1998:UIO

[RQB98b] J. M. Ruiz, M. Quintela, and R. Barreiro. Ubiquitous imposex and organotin bioaccumulation in gastropods Nucella lapillus from Galicia (NW Spain): a possible effect of nearshore shipping. Marine Ecology Progress Series, 164:237–244, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
164/m164p237.pdf.

Ramon:1992:ADS

[RR92] M. Ramón and C. A. Richardson. Age determination and shell growth of Chamelea gallina (Bivalvia: Veneridae) in the western Mediterranean. Marine Ecology Progress Series, 89:15–23, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/89/m089p015.pdf.

Rozas:1993:NUM

[RR93] L. P. Rozas and D. J. Reed. Nekton use of marsh-surface habitats in Louisiana (USA) deltaic salt marshes undergoing submergence. Marine Ecology Progress Series, 96:147–157, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
96/m096p147.pdf.

Riegger:1997:PUA

[RR97a] L. Riegger and D. Robinson. Photoinduction of UV-absorbing compounds in Antarctic diatoms and Phaeocystis antarctica. Marine Ecology Progress Series, 160:13–25, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
160/m160p013.pdf.

Riera:1997:TVP

[RR97b] P. Riera and P. Richard. Temporal variation of δ ^{13}C in particulate organic matter and oyster Crassostrea gigas in Marennes–Oléron Bay (France): effect of freshwater inflow. Marine Ecology Progress Series, 147:105–115, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
147/m147p105.pdf.

Raghukumar:1999:TFP

[RR99] S. Raghukumar and C. Raghukumar. Thraustochytrid fungoid protists in faecal pellets of the tunicate Pegea confoederata, their tolerance to deep-sea conditions and implication in degradation processes. Marine Ecology Progress Series, 190:133–140, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/190/m190p133.pdf.

Riera:1996:FSI

[RRGB96] P. Riera, P. Richard, A. Grémare, and G. Blanchard. Food source of intertidal nematodes in the Bay of Marennes– Oléron (France), as determined by dual stable isotope analysis. Marine Ecology Progress Series, 142:303–309, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
142/m142p303.pdf.

Rau:1996:MPF

[RRWG96] G. H. Rau, U. Riebesell, and D. Wolf-Gladrow. A model of photosynthetic ^{13}C fractionation by marine phytoplankton based on diffusive molecular CO_{2} uptake. Marine Ecology Progress Series, 133:275–285, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/133/m133p275.pdf.

Risk:1991:CST

[RS91] M. J. Risk and P. W. Sammarco. Cross-shelf trends in skeletal density of the massive coral Pontes lobata from the Great Barrier Reef. Marine Ecology Progress Series, 69:195–200, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
69/m069p195.pdf.

Rumohr:1992:RSP

[RS92] H. Rumohr and H. Schomann. REMOTS sediment profiles around an exploratory drilling rig in the southern North Sea. Marine Ecology Progress Series, 91:303–311, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
91/m091p303.pdf.

Ray:1994:EAG

[RS94] M. Ray and A. W. Stoner. Experimental analysis of growth and survivorship in a marine gastropod aggregation: balancing growth with safety in numbers. Marine Ecology Progress Series, 105:47–59, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/105/m105p047.pdf.

Ray:1995:GSH

[RS95] M. Ray and A. W. Stoner. Growth, survivorship, and habitat choice in a newly settled seagrass gastropod, Strombus gigas. Marine Ecology Progress Series, 123:83–94, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
123/m123p083.pdf.

Ramage:1999:PDR

[RS99a] D. L. Ramage and D. R. Schiel. Patch dynamics and response to disturbance of the seagrass Zostera novazelandica on intertidal platforms in southern New Zealand. Marine Ecology Progress Series, 189:275–288, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/189/m189p275.pdf.

Reyns:1999:PPB

[RS99b] N. Reyns and S. Sponaugle. Patterns and processes of brachyuran crab settlement to Caribbean coral reefs. Marine Ecology Progress Series, 185:155–170, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
185/m185p155.pdf.

Roberts:1998:RCE

[RSAD98] D. E. Roberts, A. Smith, P. Ajani, and A. R. Davis. Rapid changes in encrusting marine assemblages exposed to anthropogenic point-source pollution: a ‘Beyond BACI’ approach. Marine Ecology Progress Series, 163:213–224, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
163/m163p213.pdf.

Riemann:1990:CBM

[RSB^{+}90] B. Riemann, H. M. Sørensen, P. K. Bjørnsen, S. J. Horsted, L. Møller Jensen, T. G. Nielsen, and M. Søndergaard. Carbon budgets of the microbial food web in estuarine enclosures. Marine Ecology Progress Series, 65:159–170, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
65/m065p159.pdf.

Robertson:1993:CVS

[RSB93] D. R. Robertson, U. M. Schober, and J. D. Brawn. Comparative variation in spawning output and juvenile recruitment of some Caribbean reef fishes. Marine Ecology Progress Series, 94:105–113, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/94/m094p105.pdf.

Rowe:1991:TSB

[RSD^{+}91] G. Rowe, M. Sibuet, J. Deming, A. Khripounoff, J. Tietjen, S. Macko, and R. Theroux. Total sediment biomass and preliminary estimates of organic carbon residence time in deepsea benthos. Marine Ecology Progress Series, 79:99–114, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
79/m079p099.pdf.

Rodhouse:1992:ELC

[RSH92] P. G. Rodhouse, C. Symon, and E. M. C. Hatfield. Early life cycle of cephalopods in relation to the major oceanographic features of the southwest Atlantic Ocean. Marine Ecology Progress Series, 89:183–195, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/89/m089p183.pdf.

Rath:1993:EAB

[RSH93] J. Rath, C. Schiller, and G. J. Herndl. Ectoenzymatic activity and bacterial dynamics along a trophic gradient in the Caribbean Sea. Marine Ecology Progress Series, 102: 89–96, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/102/m102p089.pdf.

Rose:1999:OLS

[RSK^{+}99] C. D. Rose, W. C. Sharp, W. J. Kenworthy, J. H. Hunt, W. G. Lyons, E. J. Prager, J. F. Valentine, M. O. Hall, P. E. Whitfield, and J. W. Fourqurean. Overgrazing of a large seagrass bed by the sea urchin Lytechinus variegatus in outer Florida Bay. Marine Ecology Progress Series, 190:211–222, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
190/m190p211.pdf.

Radtke:1996:MEN

[RSM^{+}96] R. Radtke, M. Svenning, D. Malone, A. Klementsen, J. Ruzicka, and D. Fey. Migrations in an extreme northern population of Arctic charr Salvelinus alpinus: insights from otolith microchemistry. Marine Ecology Progress Series, 136:13–23, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/136/m136p013.pdf.

Richardson:1990:UIG

[RSN90] C. A. Richardson, R. Seed, and E. Naylor. Use of internal growth bands for measuring individual and population growth rates in Mytilus edulis from offshore production platforms. Marine Ecology Progress Series, 66:259–265, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
66/m066p259.pdf.

Riera:1999:DFS

[RSN^{+}99] P. Riera, L. J. Stal, J. Nieuwenhuize, P. Richard, G. Blanchard, and F. Gentil. Determination of food sources for benthic invertebrates in a salt marsh (Aiguillon Bay, France) by carbon and nitrogen stable isotopes: importance of locally produced sources. Marine Ecology Progress Series, 187: 301–307, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/187/m187p301.pdf.

Risk:1994:CCS

[RSS94] M. J. Risk, P. W. Sammarco, and H. P. Schwarcz. Crosscontinental shelf trends in δ ^{13}C in coral on the Great Barrier Reef. Marine Ecology Progress Series, 106:121–130, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
106/m106p121.pdf.

Ruger:1992:CSC

[RT92] H.-J. Rüger and T. L. Tan. Community structures of cold and low-nutrient adapted heterotrophic sediment bacteria from the deep eastern tropical Atlantic. Marine Ecology Progress Series, 84:83–93, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/84/m084p083.pdf.

Roden:1993:ISC

[RT93] E. E. Roden and J. H. Turtle. Inorganic sulfur cycling in mid and lower Chesapeake Bay sediments. Marine Ecology Progress Series, 93:101–118, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/93/m093p101.pdf.

Rasmussen:1995:TDD

[RT95] T. Rasmussen and K. Tande. Temperature-dependent development, growth and mortality in larvae of the deepwater prawn Pandalus borealis reared in the laboratory. Marine Ecology Progress Series, 118:149–157, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
118/m118p149.pdf.

Roa:1998:SDG

[RT98] R. Roa and F. Tapia. Spatial differences in growth and sexual maturity between branches of a large population of the squat lobster Pleuroncodes monodon. Marine Ecology Progress Series, 167:185–196, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/167/m167p185.pdf.

Reid:1996:KCP

[RTCH96] K. Reid, P. N. Trathan, J. P. Croxall, and H. J. Hill. Krill caught by predators and nets: differences between species and techniques. Marine Ecology Progress Series, 140:13–20, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
140/m140p013.pdf.

Riisgaard:1993:SFM

[RTJ^{+}93] H. U. Riisgård, S. Thomassen, H. Jakobsen, J. M. Weeks, and P. S. Larsen. Suspension feeding in marine sponges Halichondria panicea and Haliclona urceolus: effects of temperature on filtration rate and energy cost of pumping. Marine Ecology Progress Series, 96:177–188, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/96/m096p177.pdf.

Ritzrau:1997:EMU

[RTLG97] W. Ritzrau, L. Thomsen, R. J. Lara, and G. Graf. Enhanced microbial utilisation of dissolved amino acids indicates rapid modification of organic matter in the benthic boundary layer. Marine Ecology Progress Series, 156:43–50, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
156/m156p043.pdf.

Rau:1990:SSV

[RTRF90] G. H. Rau, J.-L. Teyssie, F. Rassoulzadegan, and S. W. Fowler. ^{13}C/^{12}C and ^{15}N/^{14}N variations among sizefractionated marine particles: implications for their origin and trophic relationships. Marine Ecology Progress Series, 59:33–38, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/59/m059p033.pdf.

Rysgaard:1998:SCN

[RTRP^{+}98] S. Rysgaard, B. Thamdrup, N. Risgaard-Petersen, H. Fossing, P. Berg, P. B. Christensen, and T. Dalsgaard. Seasonal carbon and nutrient mineralization in a high-Arctic coastal marine sediment, Young Sound, Northeast Greenland. Marine Ecology Progress Series, 175:261–276, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/175/m175p261.pdf.

Rudjakov:1997:QSP

[Rud97] J. A. Rudjakov. Quantifying seasonal phytoplankton oscillations in the global offshore ocean. Marine Ecology Progress Series, 146:225–230, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/146/m146p225.pdf.

Ruzicka:1996:CTA

[Ruz96] J. J. Ruzicka. Comparison of the two alternative early life-history strategies of the Antarctic fishes Gobionotothen gibberifrons and Lepidonotothen larseni. Marine Ecology Progress Series, 133:29–41, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/133/m133p029.pdf.

Rijnsdorp:1992:DDI

[Rv92] A. D. Rijnsdorp and P. I. van Leeuwen. Density-dependent and independent changes in somatic growth of female North Sea plaice Pleuronecles platessa between 1930 and 1985 as revealed by back-calculation of otoliths. Marine Ecology Progress Series, 88:19–32, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/88/m088p019.pdf.

Riisgaard:1992:FNS

[RVBL92] H. U. Riisgård, A. Vedel, H. Boye, and P. S. Larsen. Filternet structure and pumping activity in the polychaete Nereis diversicolor: effects of temperature and pump-modelling. Marine Ecology Progress Series, 83:79–89, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
83/m083p079.pdf.

Reynaud-Vaganay:1999:NCT

[RVGC^{+}99] S. Reynaud-Vaganay, J. P. Gattuso, J. P. Cuif, J. Jaubert, and A. Juillet-Leclerc. A novel culture technique for scleractinian corals: application to investigate changes in skeletal δ 18O as a function of temperature. Marine Ecology Progress Series, 180:121–130, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/180/m180p121.pdf.

Ryer:1997:CRM

[RvMM97] C. H. Ryer, J. van Montfrans, and K. E. Moody. Cannibalism, refugia and the molting blue crab. Marine Ecology Progress Series, 147:77–85, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/147/m147p077.pdf.

Raymond:1997:HDM

[RVT^{+}97] M. Raymond, R. L. Vääntö, F. Thomas, F. Rousset, T. de Meeüs, and F. Renaud. Heterozygote deficiency in the mussel Mytilus edulis species complex revisited. Marine Ecology Progress Series, 156:225–237, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/156/m156p225.pdf.

Rijnsdorp:1992:PBD

[RVv92] A. D. Rijnsdorp, A. D. Vethaak, and P. I. van Leeuwen. Population biology of dab Limanda limanda in the southeastern North Sea. Marine Ecology Progress Series, 91:19–35, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
91/m091p019.pdf.

Richmond:1996:STF

[RW96] C. E. Richmond and S. A. Woodin. Short-term fluctuations in salinity: effects on planktonic invertebrate larvae. Marine Ecology Progress Series, 133:167–177, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
133/m133p167.pdf.

Richter:1999:CDS

[RW99] C. Richter and M. Wunsch. Cavity-dwelling suspension feeders in coral reefs — a new link in reef trophodynamics.

Marine Ecology Progress Series, 188:105–116, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
188/m188p105.pdf.

Reid:1999:KPD

[RWCM99] K. Reid, J. L. Watkins, J. P. Croxall, and E. J. Murphy. Krill population dynamics at South Georgia 1991–1997, based on data from predators and nets. Marine Ecology Progress Series, 177:103–114, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/177/m177p103.pdf.

Rogers:1998:DVM

[RWCS98] C. N. Rogers, J. E. Williamson, D. G. Carson, and P. D. Steinberg. Diel vertical movement by mesograzers on seaweeds. Marine Ecology Progress Series, 166:301–306, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
166/m166p301.pdf.

Raven:1995:INA

[RWJ^{+}95] J. A. Raven, D. I. Walker, A. M. Johnston, L. L. Handley, and J. E. Kübler. Implications of^{13}C natural abundance measurements for photosynthetic performance by marine macrophytes in their natural environment. Marine Ecology Progress Series, 123:193–205, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/123/m123p193.pdf.

Rees:1999:MNA

[RWJ99a] A. Rees, M. Woodward, and I. Joint. Measurement of nitrate and ammonium uptake at ambient concentrations in oligotrophic waters of the North-East Atlantic Ocean. Marine Ecology Progress Series, 187:295–300, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
187/m187p295.pdf.

Roast:1999:SGE

[RWJ99b] S. D. Roast, J. Widdows, and M. B. Jones. Scope for growth of the estuarine mysid Neomysis integer (Peracarida: Mysidacea): effects of the organophosphate pesticide chlorpyrifos. Marine Ecology Progress Series, 191:233–241, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
191/m191p233.pdf.

Riveros:1996:RBF

[RZLB96] A. Riveros, M. Zuñiga, A. Larrain, and J. Becerra. Relationships between fertilization of the southeastern Pacific sea urchin Arbacia spatuligera and environmental variables in polluted coastal waters. Marine Ecology Progress Series, 134:159–169, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/134/m134p159.pdf.

Robarts:1996:BAB

[RZWY96] R. D. Robarts, T. Zohary, M. J. Waiser, and Y. Z. Yacobi. Bacterial abundance, biomass, and production in relation to phytoplankton biomass in the Levantine Basin of the southeastern Mediterranean Sea. Marine Ecology Progress Series, 137:273–281, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/137/m137p273.pdf.

Saiz:1992:FSB

[SA92] E. Saiz and M. Alcaraz. Free-swimming behaviour of Acartia clausi (Copepoda: Calanoida) under turbulent water movement. Marine Ecology Progress Series, 80:229–236, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
80/m080p229.pdf.

Sciandra:1994:ENL

[SA94] A. Sciandra and R. Amara. Effects of nitrogen limitation on growth and nitrite excretion rates of the dinoflagellate Prorocentrum minimum. Marine Ecology Progress Series, 105: 301–309, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/105/m105p301.pdf.

Simon:1990:BCD

[SAA90] M. Simon, A. L. Alldredge, and F. Azam. Bacterial carbon dynamics on marine snow. Marine Ecology Progress Series, 65:205–211, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/65/m065p205.pdf.

Styrishave:1999:LSS

[SAA99] B. Styrishave, A. Aagaard, and O. Andersen. In situ studies on physiology and behaviour in two colour forms of the shore crab Carcinus maenas in relation to season. Marine Ecology Progress Series, 189:221–231, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/189/m189p221.pdf.

Sabates:1990:DPL

[Sab90] A. Sabatés. Distribution pattern of larval fish populations in the Northwestern Mediterranean. Marine Ecology Progress Series, 59:75–82, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/59/m059p075.pdf.

Salvanes:1992:EME

[SAG92] Anne Gro Vvea Salvanes, Dag L. Aksnes, and Jarl Giske. Ecosystem model for evaluating potential cod production in a west Norwegian fjord. Marine Ecology Progress Series, 90: 9–22, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/90/m090p009.pdf.

Schuster:1998:ADF

[SAH98] S. Schuster, J. M. Arrieta, and G. J. Herndl. Adsorption of dissolved free amino acids on colloidal DOM enhances colloidal DOM utilization but reduces amino acid uptake by orders of magnitude in marine bacterioplankton. Marine Ecology Progress Series, 166:99–108, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/166/m166p099.pdf.

Sakai:1997:GSP

[Sak97] K. Sakai. Gametogenesis, spawning, and planula brooding by the reef coral Goniastrea aspera (Scleractinia) in Okinawa, Japan. Marine Ecology Progress Series, 151:67–72, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
151/m151p067.pdf.

Salkeld:1995:ARA

[Sal95] P. N. Salkeld. Aspects of reproduction associated with the use of a segmented regression to describe the relationship between body weight and shell length of Mytilus edulis. Marine Ecology Progress Series, 124:117–128, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/124/m124p117.pdf.

Saltzman:1996:ELH

[Sal96] Jennifer Saltzman. Ecology and life history traits of the benthopelagic mysid Boreomysis oparva from the eastern tropical Pacific oxygen minimum zone. Marine Ecology Progress Series, 139:95–103, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/139/m139p095.pdf.

Samoilys:1997:PSA

[Sam97] M. A. Samoilys. Periodicity of spawning aggregations of coral trout Plectropomus leopardus (Pisces: Serranidae) on the northern Great Barrier Reef. Marine Ecology Progress Series, 160:149–159, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/160/m160p149.pdf.

Sano:1993:FAD

[San93a] M. Sano. Foraging activities and diets of males and females in a haremic sandperch (Pisces: Pinguipedidae). Marine Ecology Progress Series, 98:55–59, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/98/m098p055.pdf.

Santos:1993:MSB

[San93b] R. Santos. A multivariate study of biotic and abiotic relationships in a subtidal algal stand. Marine Ecology Progress Series, 94:181–190, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/94/m094p181.pdf.

Sandberg:1994:DST

[San94a] E. Sandberg. Does short-term oxygen depletion affect predator–prey relationships in zoobenthos? Experiments with the isopod Saduna entomon. Marine Ecology Progress Series, 103:73–80, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/103/m103p073.pdf.

Santos:1994:PDP

[San94b] P. J. P. Santos. Population dynamics and production of Scolelepis gaucha (Polychaeta: Spionidae) on the sandy beaches of Southern Brazil. Marine Ecology Progress Series, 110:159–165, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/110/m110p159.pdf.

Santos:1994:FDC

[San94c] R. Santos. Frond dynamics of the commercial seaweed Gelidium sesquipedale: effects of size and of frond history. Marine Ecology Progress Series, 107:295–305, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
107/m107p295.pdf.

Santos:1995:SSI

[San95] R. Santos. Size structure and inequality in a commercial stand of the seaweed Gelidium sesquipedale. Marine Ecology Progress Series, 119:253–263, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/119/m119p253.pdf.

Sanford:1997:TME

[San97] L. P. Sanford. Turbulent mixing in experimental ecosystem studies. Marine Ecology Progress Series, 161:265–293, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
161/m161p265.pdf.

Stapel:1996:NUL

[SAvD^{+}96] J. Stapel, T. L. Aarts, B. H. M. van Duynhoven, J. D. de Groot, P. H. W. van den Hoogen, and M. A. Hemminga. Nutrient uptake by leaves and roots of the seagrass Thalassia hemprichii in the Spermonde Archipelago, Indonesia. Marine Ecology Progress Series, 134:195–206, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
134/m134p195.pdf.

Shilling:1994:PFE

[SB94] F. M. Shilling and I. Bosch. ’Pre-feeding’ embryos of Antarctic and temperate echinoderms use dissolved organic material for growth and metabolic needs. Marine Ecology Progress Series, 109:173–181, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/109/m109p173.pdf.

Spotte:1996:DAC

[SB96] S. Spotte and P. M. Bubucis. Diversity and abundance of caridean shrimps associated with the slimy sea plume Pseudopterogorgia americana at Pine Cay, Turks and Caicos Islands, British West Indies. Marine Ecology Progress Series, 133:299–302, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/133/m133p299.pdf.

Sala:1997:PSF

[SB97a] E. Sala and E. Ballesteros. Partitioning of space and food resources by three fish of the genus Diplodus (Sparidae) in a Mediterranean rocky infralittoral ecosystem. Marine Ecology Progress Series, 152:273–283, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/152/m152p273.pdf.

Shaw:1997:MPW

[SB97b] P. W. Shaw and P. R. Boyle. Multiple paternity within the brood of single females of Loligo forbesi (Cephalopoda: Loliginidae), demonstrated with microsatellite DNA markers. Marine Ecology Progress Series, 160:279–282, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
160/m160p279.pdf.

Stobutzki:1997:SSA

[SB97c] I. C. Stobutzki and D. R. Bellwood. Sustained swimming abilities of the late pelagic stages of coral reef fishes. Marine Ecology Progress Series, 149:35–41, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/149/m149p035.pdf.

Sahm:1998:AVD

[SB98] K. Sahm and U. G. Berninger. Abundance, vertical distribution, and community structure of benthic prokaryotes from permanently cold marine sediments (Svalbard, Arctic Ocean). Marine Ecology Progress Series, 165:71–80, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
165/m165p071.pdf.

Seymour:1999:LRR

[SB99] R. M. Seymour and R. H. Bradbury. Lengthening reef recovery times from crown-of-thorns outbreaks signal systemic degradation of the Great Barrier Reef. Marine Ecology Progress Series, 176:1–10, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/176/m176p001.pdf.

Sanford:1994:FFS

[SBBG94] E. Sanford, D. Bermudez, M. D. Bertness, and S. D. Gaines. Flow, food supply and acorn barnacle population dynamics. Marine Ecology Progress Series, 104:49–62, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
104/m104p049.pdf.

Southward:1998:IHS

[SBC^{+}98] A. J. Southward, R. S. Burton, S. L. Coles, P. R. Dando, R. DeFelice, J. Hoover, P. E. Parnell, T. Yamaguchi, and W. A. Newman. Invasion of Hawaiian shores by an Atlantic barnacle. Marine Ecology Progress Series, 165:119–126, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
165/m165p119.pdf.

Sloth:1995:NCS

[SBH^{+}95] N. P. Sloth, H. Blackburn, L. S. Hansen, N. Risgaard-Petersen, and B. A. Lomstein. Nitrogen cycling in sediments with different organic loading. Marine Ecology Progress Series, 116:163–170, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/116/m116p163.pdf.

Stoecker:1992:CSI

[SBP92] D. K. Stoecker, K. R. Buck, and M. Putt. Changes in the sea-ice brine community during the spring-summer transition, McMurdo Sound, Antarctica. I. Photosynthetic protists. Marine Ecology Progress Series, 84:265–278, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
84/m084p265.pdf.

Stoecker:1993:CSI

[SBP93] D. K. Stoecker, K. R. Buck, and M. Putt. Changes in the sea-ice brine community during the spring-summer transition, McMurdo Sound, Antarctica. II. Phagotrophic protists. Marine Ecology Progress Series, 95:103–113, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
95/m095p103.pdf.

Soudant:1997:EDC

[SBT97] D. Soudant, B. Beliaeff, and G. Thomas. Explaining Dinophysis cf. acuminata abundance in Antifer (Normandy, France) using dynamic linear regression. Marine Ecology Progress Series, 156:67–74, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/156/m156p067.pdf.

Schell:1998:CNI

[SBV98] D. M. Schell, B. A. Barnett, and K. A. Vinette. Carbon and nitrogen isotope ratios in zooplankton of the Bering, Chukchi and Beaufort seas. Marine Ecology Progress Series, 162:11–23, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/162/m162p011.pdf.

Santavy:1990:CBC

[SC90] Deborah L. Santavy and Rita R. Colwell. Comparison of bacterial communities associated with the Caribbean sclero-sponge Ceratoporella nicholsoni and ambient seawater. Marine Ecology Progress Series, 67:73–82, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
67/m067p073.pdf.

Sammarco:1992:CAO

[SC92a] P. W. Sammarco and J. C. Coll. Chemical adaptations in the Octocorallia: evolutionary considerations. Marine Ecology Progress Series, 88:93–104, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/88/m088p093.pdf.

Siferd:1992:NHC

[SC92b] T. D. Siferd and R. J. Conover. Natural history of ctenophores in the Resolute Passage area of the Canadian high Arctic with special reference to Mertensia ovum. Marine Ecology Progress Series, 86:133–144, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/86/m086p133.pdf.

Suttle:1993:MCI

[SC93] C. A. Suttle and A. M. Chan. Marine cyanophages infecting oceanic and coastal strains of Synechococcus: abundance, morphology, cross-infectivity and growth characteristics. Marine Ecology Progress Series, 92:99–109, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
92/m092p099.pdf.

Schultz:1994:RCR

[SC94] E. T. Schultz and R. K. Cowen. Recruitment of coral-reef fishes to Bermuda: local retention or long-distance transport? Marine Ecology Progress Series, 109:15–28, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
109/m109p015.pdf.

Somerfield:1995:TLM

[SC95a] P. J. Somerfield and K. R. Clarke. Taxonomic levels, in marine community studies, revisited. Marine Ecology Progress Series, 127:113–119, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/127/m127p113.pdf.

Suttle:1995:VIM

[SC95b] C. A. Suttle and A. M. Chan. Viruses infecting the marine prymnesiophyte Chrysochromulina spp.: isolation, preliminary characterization and natural abundance. Marine Ecology Progress Series, 118:275–282, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/118/m118p275.pdf.

Sponaugle:1996:NPC

[SC96a] S. Sponaugle and R. K. Cowen. Nearshore patterns of coral reef fish larval supply to Barbados, West Indies. Marine Ecology Progress Series, 133:13–28, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/133/m133p013.pdf.

Strickler:1996:CCB

[SC96b] J. R. Strickler and J. H. Costello. Calanoid copepod behavior in turbulent flows. Marine Ecology Progress Series, 139:307–309, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/139/m139p301.pdf.

Simon:1992:SBB

[SCA92] M. Simon, B. C. Cho, and F. Azam. Significance of bacterial biomass in lakes and the ocean: comparison to phytoplankton biomass and biogeochemical implications. Marine Ecology Progress Series, 86:103–110, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/86/m086p103.pdf.

Scanes:1993:TMU

[Sca93] P. Scanes. Trace metal uptake in cockles Anadara trapezium from Lake Macquarie, New South Wales. Marine Ecology Progress Series, 102:135–142, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/102/m102p135.pdf.

Scardi:1996:ANN

[Sca96] M. Scardi. Artificial neural networks as empirical models for estimating phytoplankton production. Marine Ecology Progress Series, 139:289–299, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/139/m139p289.pdf.

Snowden:1991:PBL

[SCAT91] R. J. Snowden, D. A. Clayton, and E. Y. Al-Taher. Population biology of Ilyoplax stevensi (Brachyura. Ocypodidae) on a Kuwait mudflat. Marine Ecology Progress Series, 71: 219–225, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/71/m071p219.pdf.

Sanders:1992:RBB

[SCB92] R. W. Sanders, D. A. Caron, and U.-G. Berninger. Relationships between bacteria and heterotrophic nanoplankton in marine and fresh waters: an inter-ecosystem comparison. Marine Ecology Progress Series, 86:1–14, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
86/m086p001.pdf.

Spanier:1994:IRO

[SCC94] E. Spanier, J. S. Cobb, and M. Clancy. Impacts of remotely operated vehicles (ROVs) on the behavior of marine animals: an example using American lobsters. Marine Ecology Progress Series, 104:257–266, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/104/m104p257.pdf.

Stribling:1998:VSS

[SCC98] J. M. Stribling, J. C. Cornwell, and C. Currin. Variability of stable sulfur isotopic ratios in Spartina alterniflora. Marine Ecology Progress Series, 166:73–81, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/166/m166p073.pdf.

Street:1998:PMJ

[SCCS98] G. T. Street, B. C. Coull, G. T. Chandler, and D. M. Sanger. Predation on meiofauna by juvenile spot Leiostomus xanthurus (Pisces) in contaminated sediments from Charleston Harbor, South Carolina, USA. Marine Ecology Progress Series, 170:261–268, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/170/m170p261.pdf.

Sornin:1990:NNT

[SCD^{+}90] J. M. Sornin, Y. Collos, D. Delmas, M. Feuillet-Girard, and D. Gouleau. Nitrogenous nutrient transfers in oyster ponds: role of sediment in deferred primary production. Marine Ecology Progress Series, 68:15–24, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/68/m068p015.pdf.

Smith:1998:GDP

[SCDH98] Walker O. Smith, Jr., Crai A. Carlson, Hugh W. Ducklow, and Dennis A. Hansell. Growth dynamics of Phaeocystis antarctica-dominated plankton assemblages from the Ross Sea. Marine Ecology Progress Series, 168:229–244, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
168/m168p229.pdf.

Smith:1993:GLC

[SCEG93] R. E. H. Smith, J. R. Cavaletto, B. J. Eadie, and W. S. Gardner. Growth and lipid composition of high Arctic ice algae during the spring bloom at Resolute, Northwest Territories, Canada. Marine Ecology Progress Series, 97:19–29, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
97/m097p019.pdf.

Suttle:1991:DFA

[SCF91] C. A. Suttle, A. M. Chan, and J. A. Fuhrman. Dissolved free amino acids in the Sargasso Sea: uptake and respiration rates, turnover times, and concentrations. Marine Ecology Progress Series, 70:189–199, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/70/m070p189.pdf.

Schaffner:1990:SSO

[Sch90] L. C. Schaffner. Small-scale organism distributions and patterns of species diversity: evidence for positive interactions in an estuarine benthic community. Marine Ecology Progress Series, 61:107–117, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/61/m061p107.pdf.

Schmaljohann:1991:OVP

[Sch91] R. Schmaljohann. Oxidation of various potential energy sources by the methanotrophic endosymbionts of Siboglinum poseidoni (Pogonophora). Marine Ecology Progress Series, 76:143–148, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/76/m076p143.pdf.

Schiel:1993:EEC

[Sch93] D. R. Schiel. Experimental evaluation of commercial-scale enhancement of abalone Haliotis iris populations in New Zealand. Marine Ecology Progress Series, 97:167–181, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
97/m097p167.pdf.

Schmaljohann:1996:MDS

[Sch96] R. Schmaljohann. Methane dynamics in the sediment and water column of Kiel harbour (Baltic Sea). Marine Ecology Progress Series, 131:263–273, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/131/m131p263.pdf.

Schaffelke:1999:STN

[Sch99] B. Schaffelke. Short-term nutrient pulses as tools to assess responses of coral reef macroalgae to enhanced nutrient availability. Marine Ecology Progress Series, 182:305–310, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
182/m182p305.pdf.

Sciandra:1991:CUB

[Sci91] A. Sciandra. Coupling and uncoupling between nitrate uptake and growth rate in Prorocentrum minimum (Dinophyceae) under different frequencies of pulsed nitrate supply. Marine Ecology Progress Series, 72:261–269, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
72/m072p261.pdf.

Saiz:1999:CEP

[SCIA99] E. Saiz, A. Calbet, X. Irigoien, and M. Alcaraz. Copepod egg production in the western Mediterranean: response to food availability in oligotrophic environments. Marine Ecology Progress Series, 187:179–189, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/187/m187p179.pdf.

Saether:1999:GEP

[SCJ99a] B. S. Sæther, J. S. Christiansen, and M. Jobling. Gastrointestinal evacuation of particulate matter in polar cod Boreogadus saida. Marine Ecology Progress Series, 188:201– 205, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/188/m188p201.pdf.

Smith:1999:SCL

[SCJ99b] I. P. Smith, K. J. Collins, and A. C. Jensen. Seasonal changes in the level and diel pattern of activity in the European lobster Homarus gammarus. Marine Ecology Progress Series, 186:255–264, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/186/m186p255.pdf.

Sullivan:1990:DAB

[SCKS90] C. W. Sullivan, G. F. Cota, D. W. Krempin, and W. O. Smith, Jr. Distribution and activity of bacterio-plankton in the marginal ice zone of the Weddell–Scotia Sea during austral spring. Marine Ecology Progress Series, 63:239–252, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
63/m063p239.pdf.

Skadsheim:1995:PRT

[SCL95] A. Skadsheim, H. Christie, and H. P. Leinaas. Population reductions of Strongylocentrotus droebachiensis (Echinodermata) in Norway and the distribution of its endoparasite Echinomermella matsi (Nematoda). Marine Ecology Progress Series, 119:199–209, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/119/m119p199.pdf.

Shashar:1994:NFA

[SCLS94] N. Shashar, Y. Cohen, Y. Loya, and N. Sar. Nitrogen fixation (acetylene reduction) in stony corals: evidence for coralbacteria interactions. Marine Ecology Progress Series, 111:

259–264, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/111/m111p259.pdf.

Sheeny:1995:ASF

[SCMM95] M. Sheeny, E. Cameron, G. Marsden, and J. McGrath. Age structure of female giant tiger prawns Penaeus monodon as indicated by neuronal lipofuscin concentration. Marine Ecology Progress Series, 117:59–63, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/117/m117p059.pdf.

Stebbing:1992:IBW

[SD92] A. R. D. Stebbing and V. Dethlefsen. Introduction to the bremerhaven workshop on biological effects of contaminants. Marine Ecology Progress Series, 91:1–8, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
91/m091p001.pdf.

Sieburth:1993:PMP

[SD93a] J. McN. Sieburth and P. L. Donaghay. Planktonic methane production and oxidation within the algal maximum of the pycnocline: seasonal fine-scale observations in an anoxic estuarine basin. Marine Ecology Progress Series, 100:3–15, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
100/m100p003.pdf.

Smith:1993:MLP

[SD93b] R. E. H. Smith and F. M. L. D’Souza. Macromolecular labelling patterns and inorganic nutrient limitation of a North Atlantic spring bloom. Marine Ecology Progress Series, 92: 111–118, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/92/m092p111.pdf.

Shiah:1994:TSR

[SD94] F.-K. Shiah and H. W. Ducklow. Temperature and substrate regulation of bacterial abundance, production and specific growth rate in Chesapeake Bay, USA. Marine Ecology Progress Series, 103:297–308, 1994. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/103/m103p297.pdf.

Stefels:1996:CDL

[SD96] J. Stefels and L. Dijkhuizen. Characteristics of DMSPlyase in Phaeocystis sp. (Prymnesiophyceae). Marine Ecology Progress Series, 131:307–313, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/131/m131p307.pdf.

Sherman:1999:EAG

[SD99] K. Sherman and A. M. Duda. An ecosystem approach to global assessment and management of coastal waters. Marine Ecology Progress Series, 190:271–287, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
190/m190p271.pdf.

Stebbing:1992:ACO

[SDA^{+}92a] A. R. D. Stebbing, V. Dethlefsen, R. F. Addison, M. Carr, P. M. Chapman, W. P. Cofino, C. Heip, L. Karbe, M. N. Moore, and A. D. Vethaak. Appendix 1: Chemistry: organic and trace metal data. Marine Ecology Progress Series, 91:331–348, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/91/m091p331.pdf.

Stebbing:1992:ABD

[SDA^{+}92b] A. R. D. Stebbing, V. Dethlefsen, R. F. Addison, M. Carr, P. M. Chapman, W. P. Cofino, C. Heip, L. Karbe, M. N. Moore, and A. D. Vethaak. Appendix 2. Biology data. Marine Ecology Progress Series, 91:349–356, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
91/m091p349.pdf.

Stebbing:1992:AA

[SDA^{+}92c] A. R. D. Stebbing, V. Dethlefsen, R. F. Addison, M. Carr, P. M. Chapman, W. P. Cofino, C. Heip, L. Karbe, M. N. Moore, and A. D. Vethaak. Appendix 3. Acknowledgements. Marine Ecology Progress Series, 91:357, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
91/m091p357.pdf.

Stebbing:1992:APB

[SDA^{+}92d] A. R. D. Stebbing, V. Dethlefsen, R. F. Addison, M. Carr, P. M. Chapman, W. P. Cofino, C. Heip, L. Karbe, M. N. Moore, and A. D. Vethaak. Appendix 4. Participants in the Bremerhaven Workshop. Marine Ecology Progress Series, 91:359–361, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/91/m091p359.pdf.

Stebbing:1992:OSS

[SDA^{+}92e] A. R. D. Stebbing, V. Dethlefsen, R. F. Addison, M. Carr, P. M. Chapman, W. P. Cofino, C. Heip, L. Karbe, M. N. Moore, and A. D. Vethaak. Overall summary and some conclusions from the Bremerhaven Workshop. Marine Ecology Progress Series, 91:323–329, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/91/m091p323.pdf.

Stenton-Dozey:1992:CRE

[SDB92] J. M. E. Stenton-Dozey and A. C. Brown. Clearance and retention efficiency of natural suspended particles by the rockpool bivalve Venerupis corrugatus in relation to tidal availability. Marine Ecology Progress Series, 82:175–186, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
82/m082p175.pdf.

Stenton-Dozey:1994:STC

[SDB94] J. M. E. Stenton-Dozey and A. C. Brown. Short-term changes in the energy balance of Venerupis corrugatus (Bivalvia) in relation to tidal availability of natural suspended particles. Marine Ecology Progress Series, 103:57–64, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
103/m103p057.pdf.

Shaughnessy:1996:CMT

[SDB96] F. J. Shaughnessy, R. E. De Wreede, and E. C. Bell. Consequences of morphology and tissue strength to blade survivorship of two closely related Rhodophyta species. Marine Ecology Progress Series, 136:257–266, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/136/m136p257.pdf.

Shaffer:1995:CCC

[SDBW95] J. A. Shaffer, D. C. Doty, R. M. Buckley, and J. E. West. Crustacean community composition and trophic use of the drift vegetation habitat by juvenile splitnose rockfish Sebastes diploproa. Marine Ecology Progress Series, 123:13–21, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
123/m123p013.pdf.

Short:1990:PLG

[SDC90] F. T. Short, W. C. Dennison, and D. G. Capone. Phosphoruslimited growth of the tropical seagrass Syringodium filiforme in carbonate sediments. Marine Ecology Progress Series, 62: 169–174, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/62/m062p169.pdf.

Shick:1992:SUR

[SDC^{+}92] J. M. Shick, W. C. Dunlap, B. E. Chalker, A. T. Banaszak, and T. K. Rosenzweig. Survey of ultraviolet radiationabsorbing mycosporine-like amino acids in organs of coral reef holothuroids. Marine Ecology Progress Series, 90:139– 148, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/90/m090p139.pdf.

Shanks:1997:LPS

[SdC97] A. L. Shanks and K. A. del Carmen. Larval polychaetes are strongly associated with marine snow. Marine Ecology Progress Series, 154:211–221, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/154/m154p211.pdf.

Schlekat:1999:DAC

[SDC99] C. E. Schlekat, A. W. Decho, and G. T. Chandler. Dietary assimilation of cadmium associated with bacterial exopolymer sediment coatings by the estuarine amphipod Leptocheirus plumulosus: effects of Cd concentration and salinity. Marine Ecology Progress Series, 183:205–216, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
183/m183p205.pdf.

Schroeter:1996:EKF

[SDER96] S. C. Schroeter, J. D. Dixon, T. A. Ebert, and J. V. Rankin. Effects of kelp forests Macrocystis pyrifera on the larval distribution and settlement of red and purple sea urchins Strongylocentrotus franciscanus and S. purpuratus. Marine Ecology Progress Series, 133:125–134, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/133/m133p125.pdf.

Sholto-Douglas:1991:SSI

[SDFJv91] A. D. Sholto-Douglas, J. G. Field, A. G. James, and N. J. van der Merwe. ^{13}C/^{12}C and ^{15}N/^{14}N isotope ratios in the Southern Benguela Ecosystem: indicators of food web relationships among different size-classes of plankton and pelagic fish; differences between fish muscle and bone collagen tissues. Marine Ecology Progress Series, 78:23–31, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
78/m078p023.pdf.

Stefels:1995:DLA

[SDG95] J. Stefels, L. Dijkhuizen, and W. W. C. Gieskes. DMSPlyase activity in a spring phytoplankton bloom off the Dutch coast, related to Phaeocystis sp. abundance. Marine Ecology Progress Series, 123:235–243, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/123/m123p235.pdf.

Soria-Dengg:1995:FBI

[SDH95] S. Soria-Dengg and U. Horstmann. Ferrioxamines B and E as iron sources for the marine diatom Phaeodactylum tricornutum. Marine Ecology Progress Series, 127:269–277, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
127/m127p269.pdf.

Sundelin:1991:MES

[SE91] B. Sundelin and R. Elmgren. Meiofauna of an experimental soft bottom ecosystem — effects of macrofauna and cadmium exposure. Marine Ecology Progress Series, 70:245–255, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
70/m070p245.pdf.

Szyper:1993:PCR

[SE93] J. P. Szyper and J. M. Ebeling. Photosynthesis and community respiration at three depths during a period of stable phytoplankton stock in a eutrophic brackish water culture pond. Marine Ecology Progress Series, 94:229–238, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
94/m094p229.pdf.

Sundelin:1998:MED

[SE98] B. Sundelin and A. K. Eriksson. Malformations in embryos of the deposit-feeding amphipod Monoporeia affinis in the Baltic Sea. Marine Ecology Progress Series, 171:165–180, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/171/m171p165.pdf.

Seapy:1990:PVD

[Sea90] R. R. Seapy. Patterns of vertical distribution in epipelagic heteropod molluscs off Hawaii. Marine Ecology Progress Series, 60:235–246, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/60/m060p235.pdf.

Stone:1996:MCR

[SEAI96] L. Stone, E. Eilam, A. Abelson, and M. Ilan. Modelling coral reef biodiversity and habitat destruction. Marine Ecology Progress Series, 134:299–302, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/134/m134p299.pdf.

Sundback:1991:ISM

[SEGP91] K. Sundbäck, V. Enoksson, W. Granéli, and K. Pettersson. Influence of sublittoral microphytobenthos on the oxygen and nutrient flux between sediment and water. A laboratory continuous-flow study. Marine Ecology Progress Series, 74:263–279, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/74/m074p263.pdf.

Sepulveda:1994:DGI

[Sep94] A. Sepulveda. Daily growth increments in the otoliths of European smelt Osmerus eperlanus larvae. Marine Ecology Progress Series, 108:33–42, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/108/m108p033.pdf.

Sewell:1994:BRJ

[Sew94] M. A. Sewell. Birth, recruitment and juvenile growth in the intraovarian brooding sea cucumber Leptosynapta clarki. Marine Ecology Progress Series, 114:149–156, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
114/m114p149.pdf.

Sewell:1996:DIP

[Sew96] M. A. Sewell. Detection of the impact of predation by migratory shorebirds: an experimental test in the Fraser River estuary, British Columbia (Canada). Marine Ecology Progress Series, 144:23–40, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/144/m144p023.pdf.

Sun:1991:STP

[SF91] B. Sun and J. W. Fleeger. Spatial and temporal patterns of dispersion in meiobenthic copepods. Marine Ecology Progress Series, 71:1–11, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/71/m071p001.pdf.

Sun:1994:FEC

[SF94] B. Sun and J. W. Fleeger. Field experiments on the colonization of meiofauna into sediment depressions. Marine Ecology Progress Series, 110:167–175, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/110/m110p167.pdf.

Sjotun:1995:GAL

[SF95] K. Sjøtun and S. Fredriksen. Growth allocation in Laminaria hyperborea (Laminariales, Phaeophyceae) in relation to age and wave exposure. Marine Ecology Progress Series, 126:213–222, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/126/m126p213.pdf.

Strasser:1999:SCA

[SF99] K. M. Strasser and D. L. Felder. Settlement cues in an Atlantic coast population of the ghost shrimp Callichirus major (Crustacea: Decapoda: Thalassinidea). Marine Ecology Progress Series, 183:217–225, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/183/m183p217.pdf.

Steele:1998:IPC

[SFA98] M. A. Steele, G. E. Forrester, and G. R. Almany. Influences of predators and conspecifics on recruitment of a tropical and a temperate reef fish. Marine Ecology Progress Series, 172:115–125, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/172/m172p115.pdf.

Starczak:1992:EBA

[SFB92] V. R. Starczak, C. M. Fuller, and C. A. Butman. Effects of barite on aspects of the ecology of the polychaete Mediomastus ambiseta. Marine Ecology Progress Series, 85:269–282, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/85/m085p269.pdf.

Suthers:1992:CLO

[SFF92] I. M. Suthers, A. Fraser, and K. T. Frank. Comparison of lipid, otolith and morphometric condition indices of pelagic juvenile cod Gadus morhua from the Canadian Atlantic. Marine Ecology Progress Series, 84:31–40, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
84/m084p031.pdf.

Selck:1998:TTC

[SFF98] H. Selck, V. E. Forbes, and T. L. Forbes. Toxicity and toxicokinetics of cadmium in Capitella sp. I: relative importance of water and sediment as routes of cadmium uptake.

Marine Ecology Progress Series, 164:167–178, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
164/m164p167.pdf.

Springer:1999:HPC

[SFG^{+}99] B. Springer, M. Friedrichs, G. Graf, J. Nittikowski, and W. Queisser. A high-precision current measurement system for laboratory flume systems: a case study around a circular cylinder. Marine Ecology Progress Series, 183:305–310, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
183/m183p305.pdf.

Siokou-Frangou:1991:DZP

[SFP91] I. Siokou-Frangou and E. Papathanassiou. Differentiation of zooplankton populations in a polluted area. Marine Ecology Progress Series, 76:41–51, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/76/m076p041.pdf.

Stratoudakis:1997:DLH

[SFPC97] Y. Stratoudakis, R. J. Fryer, G. J. Pierce, and R. M. Cook. Differences in life history features of long rough dab Hippoglossoides platessoides within Scottish waters. Marine Ecology Progress Series, 158:303–306, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/158/m158p303.pdf.

Selmer:1993:NRP

[SFPCR93] Johanne-Sophie Selmer, Christine Ferrier-Pages, Christiane Cellario, and Fereidoun Rassoulzadegan. New and regenerated production in relation to the microbial loop in the NW Mediterranean Sea. Marine Ecology Progress Series, 100: 71–83, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/100/m100p071.pdf.

Schubert:1995:LSM

[SFS95] H. Schubert, R. M. Forster, and S. Sagert. In situ measurement of state transition in cyanobacterial blooms: kinetics and extent of the state change in relation to underwater light and vertical mixing. Marine Ecology Progress Series, 128:99–108, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/128/m128p099.pdf.

Serret:1999:SCM

[SFSA99] P. Serret, E. Fernández, J. A. Sostres, and R. Anadón. Seasonal compensation of microbial production and respiration in a temperate sea. Marine Ecology Progress Series, 187: 43–57, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/187/m187p043.pdf.

Schrijvers:1996:RCB

[SFV96] J. Schrijvers, H. Fermon, and M. Vincx. Resource competition between macrobenthic epifauna and infauna in a Kenyan Avicennia marina mangrove forest. Marine Ecology Progress Series, 136:123–135, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/136/m136p123.pdf.

Schmaljohann:1990:CEM

[SFWD90] R. Schmaljohann, E. Faber, M. J. Whiticar, and P. R. Dando. Co-existence of methane- and sulphur-based endosymbioses between bacteria and invertebrates at a site in the Skagerrak. Marine Ecology Progress Series, 61:119–124, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
61/m061p119.pdf.

Skold:1996:SGG

[SG96] M. Sköld and J. S. G. Gunnarsson. Somatic and germinal growth of the infaunal brittle stars Amphiura filiformis and A. chiajei in response to organic enrichment. Marine Ecology Progress Series, 142:203–214, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/142/m142p203.pdf.

Sulaiman:1990:ASP

[SGB90] N. Sulaiman, S. George, and M. D. Burke. Assessment of sublethal pollutant impact on flounders in an industrialised estuary using hepatic biochemical indices. Marine Ecology Progress Series, 68:207–212, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/68/m068p207.pdf.

Souchu:1998:BAB

[SGC^{+}98] P. Souchu, A. Gasc, Y. Collos, A. Vaquer, H. Tournier, B. Bibent, and J. M. Deslous-Paoli. Biogeochemical aspects of bottom anoxia in a Mediterranean lagoon (Thau, France). Marine Ecology Progress Series, 164:135–146, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
164/m164p135.pdf.

Smith:1997:BML

[SGK^{+}97] R. E. H. Smith, M. Gosselin, G. Kattner, L. Legendre, and S. Pesant. Biosynthesis of macromolecular and lipid classes by phytoplankton in the northeast water polynya. Marine Ecology Progress Series, 147:231–242, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/147/m147p231.pdf.

Sakka:1997:ERU

[SGL^{+}97] A. Sakka, M. Gosselin, M. Levasseur, S. Michaud, P. Monfort, and S. Demers. Effects of reduced ultraviolet radiation on aqueous concentrations of dimethylsulfoniopropionate and dimethylsulfide during a microcosm study in the Lower St. Lawrence Estuary. Marine Ecology Progress Series, 149: 227–238, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/149/m149p227.pdf.

Stratoudakis:1998:SDD

[SGM98] Y. Stratoudakis, A. Gallego, and J. A. Morrison. Spatial distribution of developmental egg ages within a herring Clupea harengus spawning ground. Marine Ecology Progress Series, 174:27–32, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/174/m174p027.pdf.

Smith:1999:STV

[SGMS99] K. A. Smith, M. T. Gibbs, J. H. Middleton, and I. M. Suthers. Short term variability in larval fish assemblages of the Sydney shelf: tracers of hydrographic variability. Marine Ecology Progress Series, 178:1–15, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/178/m178p001.pdf.

Stoecker:1994:PMP

[SGP94] D. K. Stoecker, D. J. Gifford, and M. Putt. Preservation of marine planktonic ciliates: losses and cell shrinkage during fixation. Marine Ecology Progress Series, 110:293–299, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
110/m110p293.pdf.

Snelgrove:1999:HSA

[SGP99] P. V. R. Snelgrove, J. Grant, and C. A. Pilditch. Habitat selection and adult-larvae interactions in settling larvae of soft-shell clam Mya arenaria. Marine Ecology Progress Series, 182:149–159, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/182/m182p149.pdf.

Simo:1995:OTD

[SGPAA95] R. Simó, J. O. Grimalt, C. Pedrós-Alió, and J. Albaigés. Occurrence and transformation of dissolved dimethyl sulfur species in stratified seawater (western Mediterranean Sea). Marine Ecology Progress Series, 127:291–299, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
127/m127p291.pdf.

Stagg:1992:CBK

[SGR92] R. Stagg, A. Goksoyr, and G. Rodger. Changes in branchial Na^{+}, K^{+}-ATPase, metallo-thionein and P450 1A1 in dab Limanda limanda in the German Bight: indicators of sediment contamination? Marine Ecology Progress Series, 91: 105–115, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/91/m091p105.pdf.

Somerfield:1994:SSM

[SGW94] P. J. Somerfield, J. M. Gee, and R. M. Warwick. Soft sediment meiofaunal community structure in relation to a longterm heavy metal gradient in the fal estuary system. Marine Ecology Progress Series, 105:79–88, 1994. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/105/m105p079.pdf.

Sameoto:1990:LCD

[SH90a] D. D. Sameoto and A. W. Herman. Life cycle and distribution of Calanus finmarchicus in deep basins on the Nova Scotia shelf and seasonal changes in Calanus spp. Marine Ecology Progress Series, 66:225–237, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/66/m066p225.pdf.

Soetaert:1990:SSD

[SH90b] K. Soetaert and C. Heip. Sample-size dependence of diversity indices and the determination of sufficient sample size in a high-diversity deep-sea environment. Marine Ecology Progress Series, 59:305–307, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/59/m059p305.pdf.

Simpson:1992:TPD

[SH92a] M. G. Simpson and T. H. Hutchinson. Toxicological pathology of dab Limanda limanda along pollution gradients in the southern North Sea. Marine Ecology Progress Series, 91: 155–161, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/91/m091p155.pdf.

Sugai:1992:RAA

[SH92b] S. F. Sugai and S. M. Henrichs. Rates of amino acid uptake and mineralization in Resurrection Bay (Alaska) sediments. Marine Ecology Progress Series, 88:129–141, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
88/m088p129.pdf.

Stokesbury:1993:SDG

[SH93] K. D. E. Stokesbury and J. H. Himmelman. Spatial distribution of the giant scallop Placopecten magellanicus in unharvested beds in the Baie des Chaleurs, Quebec. Marine Ecology Progress Series, 96:159–168, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/96/m096p159.pdf.

Soetaert:1994:OFG

[SH94] K. Soetaert and P. M. J. Herman. One foot in the grave: zooplankton drift into the Westerschelde estuary (The Netherlands). Marine Ecology Progress Series, 105:19–29, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
105/m105p019.pdf.

Schuster:1995:FST

[SH95a] S. Schuster and G. J. Herndl. Formation and significance of transparent exopolymeric particles in the northern Adriatic Sea. Marine Ecology Progress Series, 124:227–236, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
124/m124p227.pdf.

Soetaert:1995:NAD

[SH95b] K. Soetaert and C. Heip. Nematode assemblages of deep-sea and shelf break sites in the North Atlantic and Mediterranean Sea. Marine Ecology Progress Series, 125:171–183, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
125/m125p171.pdf.

Scheel:1997:SFP

[SH97a] D. Scheel and K. R. Hough. Salmon fry predation by seabirds near an Alaskan hatchery. Marine Ecology Progress Series, 150:35–48, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/150/m150p035.pdf.

Scheibling:1997:ROD

[SH97b] R. E. Scheibling and A. W. Hennigar. Recurrent outbreaks of disease in sea urchins Strongylocentrotus droebachiensis in Nova Scotia: evidence for a link with large-scale meteorologic and oceanographic events. Marine Ecology Progress Series, 152:155–165, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/152/m152p155.pdf.

Schluter:1997:PPR

[SH97c] L. Schlüter and H. Havskum. Phytoplankton pigments in relation to carbon content in phytoplankton communities. Marine Ecology Progress Series, 155:55–65, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/155/m155p055.pdf.

Son:1998:RLB

[SH98a] M. H. Son and S. Y. Hong. Reproduction of Littorina brevicula in Korean waters. Marine Ecology Progress Series, 172:215–223, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/172/m172p215.pdf.

Stevens:1998:MCD

[SH98b] C. J. Stevens and E. J. H. Head. A model of chlorophyll-a destruction by Calanus spp. and implications for the estimation of ingestion rates using the gut fluorescence method. Marine Ecology Progress Series, 171:187–198, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
171/m171p187.pdf.

Stachowicz:1999:RMA

[SH99a] J. J. Stachowicz and M. Hay. Reduced mobility is associated with compensatory feeding and increased diet breadth of marine crabs. Marine Ecology Progress Series, 188:169– 178, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/188/m188p169.pdf.

Stoderegger:1999:PEP

[SH99b] K. E. Stoderegger and G. J. Herndl. Production of exopolymer particles by marine bacterioplankton under contrasting turbulence conditions. Marine Ecology Progress Series, 189: 9–16, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/189/m189p009.pdf.

Santelices:1995:BMF

[SHA^{+}95a] B. Santelices, A. J. Hoffmann, D. Aedo, M. Bobadilla, and R. Otaíza. A bank of microscopic forms on disturbed boulders and stones in tide pools. Marine Ecology Progress Series, 129:215–228, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/129/m129p215.pdf.

Shane:1995:RBP

[Sha95b] S. H. Shane. Relationship between pilot whales and Risso’s dolphins at Santa Catalina Island, California, USA. Marine Ecology Progress Series, 123:5–11, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/123/m123p005.pdf.

Shanks:1998:APL

[Sha98] A. L. Shanks. Abundance of post-larval Callinectes sapidus, Penaeus spp., Uca spp., and Libinia spp. collected at an outer coastal site and their cross-shelf transport. Marine Ecology Progress Series, 168:57–69, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/168/m168p057.pdf.

Smith:1996:PED

[SHC96] D. Smith, R. G. Hughes, and E. J. Cox. Predation of epipelic diatoms by the amphipod Corophium volutator and the polychaete Nereis diversicolor. Marine Ecology Progress Series, 145:53–61, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/145/m145p053.pdf.

Sherman:1994:SBY

[She94] K. Sherman. Sustainability, biomass yields, and health of coastal ecosystems: an ecological perspective. Marine Ecology Progress Series, 112:277–301, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/112/m112p277.pdf.

Sheaves:1995:LLS

[She95] M. Sheaves. Large lutjanid and serranid fishes in tropical estuaries: Are they adults or juveniles? Marine Ecology Progress Series, 129:31–41, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/129/m129p031.pdf.

Sheaves:1996:DSD

[She96] M. Sheaves. Do spatial differences in the abundance of two serranid fishes in estuaries of tropical Australia reflect longterm salinity patterns? Marine Ecology Progress Series, 137: 39–49, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/137/m137p039.pdf.

Sheader:1998:GPP

[She98] M. Sheader. Grazing predation on a population of Ampelisca tenuicornis (Gammaridae: Amphipoda) off the south coast of England. Marine Ecology Progress Series, 164:253–262, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
164/m164p253.pdf.

Stearns:1994:OPB

[SHFP94] D. E. Stearns, G. J. Holt, R. B. Forward, Jr., and P. L. Pickering. Ontogeny of phototactic behavior in red drum larvae (Sciaenidae: Sciaenops ocellatus). Marine Ecology Progress Series, 104:1–11, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/104/m104p001.pdf.

Slattery:1997:ERW

[SHM^{+}97] M. Slattery, M. T. Hamann, J. B. McClintock, T. L. Perry, M. P. Puglisi, and W. Y. Yoshida. Ecological roles for waterborne metabolites from Antarctic soft corals. Marine Ecology Progress Series, 161:133–144, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/161/m161p133.pdf.

Simo:1998:PDS

[SHML98] R. Simó, A. D. Hatton, G. Malin, and P. S. Liss. Particulate dimethyl sulphoxide in seawater: production by microplankton. Marine Ecology Progress Series, 167:291–296, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
167/m167p291.pdf.

Steichen:1996:DAB

[SHO96] D. J. Steichen, Jr., S. J. Holbrook, and C. W. Osenberg. Distribution and abundance of benthic and demersal macrofauna within a natural hydrocarbon seep. Marine Ecology Progress Series, 138:71–82, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/138/m138p071.pdf.

Starr:1991:IPS

[SHT91] M. Starr, J. H. Himmelman, and J.-C. Therriault. Isolation and properties of a substance from the diatom Phaeodactylum tricornutum which induces spawning in the sea urchin Strongylocentrotus droebachiensis. Marine Ecology Progress Series, 79:275–287, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/79/m079p275.pdf.

Soetaert:1991:DNA

[SHV91] K. Soetaert, C. Heip, and M. Vincx. Diversity of nematode assemblages along a Mediterranean deep-sea transect. Marine Ecology Progress Series, 75:275–282, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
75/m075p275.pdf.

Szewzyk:1991:REM

[SHW^{+}91] U. Szewzyk, C. Holmström, M. Wrangstadh, M.-O. Samuelsson, J. S. Maki, and S. Kjelleberg. Relevance of the exopolysaccharide of marine Pseudomonas sp. strain S9 for the attachment of Ciona intestinalis larvae. Marine Ecology Progress Series, 75:259–265, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/75/m075p259.pdf.

Sieburth:1993:CBWa

[Sie93] J. McN. Sieburth. C1 bacteria in the water column of Chesapeake Bay, USA. I. Distribution of sub-populations of O_{2}- tolerant, obligately anaerobic, methylotrophic methanogens that occur in microniches reduced by their bacterial consorts. Marine Ecology Progress Series, 95:67–80, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
95/m095p067.pdf.

Simon:1991:IDI

[Sim91] M. Simon. Isotope dilution of intracellular amino acids as a tracer of carbon and nitrogen sources of marine planktonic bacteria. Marine Ecology Progress Series, 74:295–301, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
74/m074p295.pdf.

Simmons:1992:ISG

[Sim92] G. M. Simmons, Jr. Importance of submarine groundwater discharge (SGWD) and seawater cycling to material flux across sediment/water interfaces in marine environments. Marine Ecology Progress Series, 84:173–184, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
84/m084p173.pdf.

Smith:1995:NIS

[SJ95] J. Staunton Smith and C. R. Johnson. Nutrient inputs from seabirds and humans on a populated coral cay. Marine Ecology Progress Series, 124:189–200, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/124/m124p189.pdf.

Schmidt:1997:NQT

[SJ97] K. Schmidt and S. H. Jónasdóttir. Nutritional quality of two cyanobacteria: How rich is ’poor’ food? Marine Ecology Progress Series, 151:1–10, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/151/m151p001.pdf.

Sarrazin:1999:BCH

[SJ99] J. Sarrazin and S. K. Juniper. Biological characteristics of a hydrothermal edifice mosaic community. Marine Ecology Progress Series, 185:1–19, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/185/m185p001.pdf.

Sondergaard:1991:PDC

[SJÆ91] M. Sondergaard, L. M. Jensen, and G. Ærtebjerg. Picoalgae in Danish coastal waters during summer stratification. Marine Ecology Progress Series, 79:139–149, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
79/m079p139.pdf.

Srinivasan:1999:EER

[SJC99] M. Srinivasan, G. P. Jones, and M. J. Caley. Experimental evaluation of the roles of habitat selection and interspecific competition in determining patterns of host use by two anemonefishes. Marine Ecology Progress Series, 186:283– 292, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/186/m186p283.pdf.

Sieburth:1993:CBWc

[SJCL93] J. McN. Sieburth, P. W. Johnson, V. M. Church, and D. C. Laux. C1 bacteria in the water column of Chesapeake Bay USA. III. Immunologic relationships of the type species of marine monomethylamine- and methane-oxidizing bacteria to wild estuarine and oceanic cultures. Marine Ecology Progress Series, 95:91–102, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/95/m095p091.pdf.

Sand-Jensen:1990:PME

[SJJMH90] K. Sand-Jensen, L. M. Jensen, S. Marcher, and M. Hansen. Pelagic metabolism in eutrophic coastal waters during a late summer period. Marine Ecology Progress Series, 65:63–72, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/65/m065p063.pdf.

Sieburth:1993:CBWb

[SJMC93] John McN. Sieburth, Paul W. Johnson, Alberto J. L. Macario, and Everly Conway de Macario. C_{1} bacteria in the water column of Chesapeake Bay, USA. II. The dominant O_{2}^{−} and H_{2}S-tolerant methylotrophic methanogens, coenriched with their oxidative and sulphate reducing bacterial consorts, are all new immunotypes and probably include new taxa. Marine Ecology Progress Series, 95:81–89, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/95/m095p081.pdf.

Sarrazin:1999:PCF

[SJML99] J. Sarrazin, S. K. Juniper, G. Massoth, and P. Legendre. Physical and chemical factors influencing species distributions on hydrothermal sulfide edifices of the Juan de Fuca Ridge, northeast Pacific. Marine Ecology Progress Series, 190: 89–112, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/190/m190p089.pdf.

Sampou:1994:FRP

[SK94a] P. Sampou and W. M. Kemp. Factors regulating plankton community respiration in Chesapeake Bay. Marine Ecology Progress Series, 110:249–258, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/110/m110p249.pdf.

Solid:1994:RPC

[SK94b] M. Solid and N. Krstulovic. Role of predation in controlling bacterial and heterotrophic nanoflagellate standing stocks in the coastal Adriatic Sea: seasonal patterns. Marine Ecology Progress Series, 114:219–235, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/114/m114p219.pdf.

Saiz:1995:PSF

[SK95a] E. Saiz and T. Kiørboe. Predatory and suspension feeding of the copepod Acartia tonsa in turbulent environments. Marine Ecology Progress Series, 122:147–158, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
122/m122p147.pdf.

Smith:1995:SRV

[SK95b] E. M. Smith and W. M. Kemp. Seasonal and regional variations in plankton community production and respiration for Chesapeake Bay. Marine Ecology Progress Series, 116:217– 231, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/116/m116p217.pdf.

Swain:1995:AVT

[SK95c] D. P. Swain and D. L. Kramer. Annual variation in temperature selection by Atlantic cod Gadus morhua in the southern Gulf of St. Lawrence, Canada, and its relation to population size. Marine Ecology Progress Series, 116:11–23, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
116/m116p011.pdf.

Spiridonov:1997:WOM

[SK97] V. A. Spiridonov and K. N. Kosobokova. Winter ontogenetic migrations and the onset of gonad development in large dominant calanoid copepods in the Weddell Gyre (Antarctica). Marine Ecology Progress Series, 157:233–246, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
157/m157p233.pdf.

Schaffelke:1998:NLG

[SK98a] B. Schaffelke and D. W. Klumpp. Nutrient-limited growth of the coral reef macroalga Sargassum baccularia and experimental growth enhancement by nutrient addition in continuous flow culture. Marine Ecology Progress Series, 164: 199–211, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/164/m164p199.pdf.

Schaffelke:1998:STN

[SK98b] B. Schaffelke and D. W. Klumpp. Short-term nutrient pulses enhance growth and photosynthesis of the coral reef macroalga Sargassum baccularia. Marine Ecology Progress Series, 170:95–105, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/170/m170p095.pdf.

Stevens:1998:PSB

[SK98c] B. G. Stevens and J. Kittaka. Postlarval settling behavior, substrate preference, and time to metamorphosis for red king crab Paralithodes camtschaticus. Marine Ecology Progress Series, 167:197–206, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/167/m167p197.pdf.

Satumanatpan:1999:EBS

[SK99] S. Satumanatpan and M. J. Keough. Effect of barnacles on the survival and growth of temperate mangrove seedlings. Marine Ecology Progress Series, 181:189–199, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
181/m181p189.pdf.

Skilleter:1994:RPP

[Ski94] G. A. Skilleter. Refuges from predation and the persistence of estuarine clam populations. Marine Ecology Progress Series, 109:29–42, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/109/m109p029.pdf.

Shibata:1997:FSC

[SKKO97] A. Shibata, K. Kogure, I. Koike, and K. Ohwada. Formation of submicron colloidal particles from marine bacteria by viral infection. Marine Ecology Progress Series, 155:303–307, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
155/m155p303.pdf.

Saila:1993:MED

[SKM93] S. B. Saila, V. Lj Kocic, and J. W. McManus. Modelling the effects of destructive fishing practices on tropical coral reefs. Marine Ecology Progress Series, 94:51–60, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
94/m094p051.pdf.

Soto:1995:BHM

[SKM95] M. Soto, M. Kortabitarte, and I. Marigómez. Bioavailable heavy metals in estuarine waters as assessed by metal/shellweight indices in sentinel mussels Mytilus galloprovincialis. Marine Ecology Progress Series, 125:127–136, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
125/m125p127.pdf.

Schrum:1997:MOS

[SKM^{+}97] C. Schrum, P. König, K. Michaelsen, U. Niemeier, and T. Pohlmann. Meteorological and oceanographic situation in the German Bight from 23 to 29 April 1991. Marine Ecology Progress Series, 156:263–273, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/156/m156p263.pdf.

Skovgaard:1996:MFS

[Sko96] A. Skovgaard. Mixotrophy in Fragilidium subglobosum (Dinophyceae): growth and grazing responses as functions of light intensity. Marine Ecology Progress Series, 143:247–253, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
143/m143p247.pdf.

Somarakis:1997:FAO

[SKPT97] S. Somarakis, I. Kostikas, N. Peristeraki, and N. Tsimenides. Fluctuating asymmetry in the otoliths of larval anchovy Engraulis encrasicolus and the use of developmental instability as an indicator of condition in larval fish. Marine Ecology Progress Series, 151:191–203, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/151/m151p191.pdf.

Suzuki:1997:IDD

[SKT97] S. Suzuki, K. Kogure, and E. Tanoue. Immunochemical detection of dissolved proteins and their source bacteria in marine environments. Marine Ecology Progress Series, 158:1–9, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
158/m158p001.pdf.

Schmidt:1998:CEP

[SKvB98] K. Schmidt, P. Kähler, and B. von Bodungen. Copepod egg production rates in the Pomeranian Bay (Southern Baltic Sea) as a function of phytoplankton abundance and taxonomic composition. Marine Ecology Progress Series, 174: 183–195, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/174/m174p183.pdf.

Sponaugle:1990:PCP

[SL90] S. Sponaugle and P. Lawton. Portunid crab predation on juvenile hard clams: effects of substrate type and prey density. Marine Ecology Progress Series, 67:43–53, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
67/m067p043.pdf.

Schneider:1991:ZCM

[SL91] G. Schneider and J. Lenz. Zooplankton community metabolism in the upper 200 m of the central Red Sea and the Gulf of Aden. Marine Ecology Progress Series, 77:301–306, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/77/m077p301.pdf.

Stoner:1994:HDA

[SL94] A. W. Stoner and J. Lally. High-density aggregation in queen conch Strombus gigas: formation, patterns, and ecological significance. Marine Ecology Progress Series, 106:73–84, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
106/m106p073.pdf.

Siegel:1995:RAK

[SL95] V. Siegel and V. Loeb. Recruitment of Antarctic krill Euphausia superba and possible causes for its variability. Marine Ecology Progress Series, 123:45–56, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/123/m123p045.pdf.

Seuront:1997:CST

[SL97a] L. Seuront and Y. Lagadeuc. Characterisation of spacetime variability in stratified and mixed coastal waters (Baie des Chaleurs, Québec, Canada): application of fractal theory. Marine Ecology Progress Series, 159:81–95, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
159/m159p081.pdf.

Stachowicz:1997:CDA

[SL97b] J. J. Stachowicz and N. Lindquist. Chemical defense among hydroids on pelagic Sargassum: predator deterrence and absorption of solar UV radiation by secondary metabolites. Marine Ecology Progress Series, 155:115–126, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
155/m155p115.pdf.

Schaff:1992:SHB

[SLB^{+}92] T. Schaff, L. Levin, N. Blair, D. DeMaster, R. Pope, and S. Boehme. Spatial heterogeneity of benthos on the Carolina continental slope: large (100 km)-scale variation. Marine Ecology Progress Series, 88:143–160, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
88/m088p143.pdf.

Stoecker:1997:MDP

[SLC^{+}97] D. K. Stoecker, A. Li, D. W. Coats, D. E. Gustafson, and M. K. Nannen. Mixotrophy in the dinoflagellate Prorocentrum minimum. Marine Ecology Progress Series, 152:1–12, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
152/m152p001.pdf.

Serafy:1997:EFC

[SLHA97] J. E. Serafy, K. C. Lindeman, T. E. Hopkins, and J. S. Ault. Effects of freshwater canal discharge on fish assemblages in a subtropical bay: field and laboratory observations. Marine Ecology Progress Series, 160:161–172, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
160/m160p161.pdf.

Stien:1998:PDE

[SLHC98] A. Stien, H. P. Leinaas, O. Halvorsen, and H. Christie. Population dynamics of the Echinomermella matsi (Nematoda)– Strongylocentrotus droebachiensis (Echinoida) system: effects on host fecundity. Marine Ecology Progress Series, 163:193– 201, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/163/m163p193.pdf.

Skold:1994:PDD

[SLR94] M. Sköld, L.-O. Loo, and R. Rosenberg. Production, dynamics and demography of an Amphiura filiformis population. Marine Ecology Progress Series, 103:81–90, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
103/m103p081.pdf.

Solis-lbarra:1994:LOD

[SlRR94] R. Solis-lbarra and S. Rendon-Rodriguez. Laboratory observations on displacement speed of the white shrimp Penaeus vannamei (Crustacea: Decapoda). Marine Ecology Progress Series, 103:309–310, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/103/m103p309.pdf.

Sulkin:1998:NRP

[SLSH98] S. Sulkin, J. Lehto, S. Strom, and D. Hutchinson. Nutritional role of protists in the diet of first stage larvae of the Dungeness crab Cancer magister. Marine Ecology Progress Series, 169:237–242, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/169/m169p237.pdf.

Sullivan:1990:EAI

[SM90] M. J. Sullivan and C. A. Moncreiff. Edaphic algae are an important component of salt marsh food-webs: evidence from multiple stable isotope analyses. Marine Ecology Progress Series, 62:149–159, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/62/m062p149.pdf.

Schneider:1991:RRF

[SM91] F. I. Schneider and K. H. Mann. Rapid recovery of fauna following simulated ice rafting in a Nova Scotian seagrass bed. Marine Ecology Progress Series, 78:57–70, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
78/m078p057.pdf.

Salat:1994:ELE

[SM94] J. Salat and C. Marrasé. Exponential and linear estimations of grazing on bacteria: effects of changes in the proportion of marked cells. Marine Ecology Progress Series, 104:205–209, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/104/m104p205.pdf.

Sondergaard:1995:CSA

[SM95] M. Søndergaard and M. Middelboe. A cross-system analysis of labile dissolved organic carbon. Marine Ecology Progress Series, 118:283–294, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/118/m118p283.pdf.

Sims:1997:DZC

[SM97a] D. W. Sims and D. A. Merrett. Determination of zooplankton characteristics in the presence of surface feeding basking sharks Cetorhinus maximus. Marine Ecology Progress Series, 158:297–302, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/158/m158p297.pdf.

Soto:1997:MBA

[SM97b] M. Soto and I. Marigómez. Metal bioavailability assessment in ’mussel-watch’ programmes by automated image analysis of autometallographical black silver deposits (BSD) in digestive cell lysosomes. Marine Ecology Progress Series, 156: 141–150, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/156/m156p141.pdf.

Sturgis:1997:SNI

[SM97c] R. B. Sturgis and L. Murray. Scaling of nutrient inputs to submersed plant communities: temporal and spatial variations. Marine Ecology Progress Series, 152:89–102, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
152/m152p089.pdf.

Saito:1998:OMF

[SM98a] H. Saito and M. Murata. Origin of the monoene fats in the lipid of midwater fishes: relationship between the lipids of myctophids and those of their prey. Marine Ecology Progress Series, 168:21–33, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/168/m168p021.pdf.

Sundt:1998:AOC

[SM98b] R. C. Sundt and W. Melle. Atlantic observation of Calanus marshallae (Copepoda; Calanoida). Marine Ecology Progress Series, 166:207–210, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/166/m166p207.pdf.

Stenhouse:1999:IEA

[SM99a] I. J. Stenhouse and W. A. Montevecchi. Indirect effects of the availability of capelin and fishery discards: gull predation on breeding storm-petrels. Marine Ecology Progress Series, 184:303–307, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/184/m184p303.pdf.

Sulkin:1999:SFH

[SM99b] S. D. Sulkin and G. L. McKeen. The significance of feeding history on the value of heterotrophic microzooplankton as prey for larval crabs. Marine Ecology Progress Series, 186:219–225, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/186/m186p219.pdf.

Slattery:1997:DFN

[SMB97] M. Slattery, J. B. McClintock, and S. S. Bowser. Deposit feeding: a novel mode of nutrition in the Antarctic colonial soft coral Gersemia antarctica. Marine Ecology Progress Series, 149:299–304, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/149/m149p299.pdf.

Strom:1998:PSI

[SMB98] S. L. Strom, T. A. Morello, and K. J. Bright. Protozoan size influences algal pigment degradation during grazing. Marine Ecology Progress Series, 164:189–197, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
164/m164p189.pdf.

Scholz:1991:DSD

[SMF91] D. S. Scholz, L. L. Matthews, and R. J. Feller. Detecting selective digestion of meiobenthic prey by juvenile spot Leiostomus xanthurus (Pisces) using immunoassays. Marine Ecology Progress Series, 72:59–67, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/72/m072p059.pdf.

Sasal:1997:DPS

[SMG97] P. Sasal, S. Morand, and J. F. Guégan. Determinants of parasite species richness in Mediterranean marine fishes. Marine Ecology Progress Series, 149:61–71, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/149/m149p061.pdf.

Stapel:1997:BLN

[SMH97] J. Stapel, R. Manuntun, and M. A. Hemminga. Biomass loss and nutrient redistribution in an Indonesian Thalassia hemprichii seagrass bed following seasonal low tide exposure during daylight. Marine Ecology Progress Series, 148: 251–262, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/148/m148p251.pdf.

Smith:1995:ELA

[Smi95] L. D. Smith. Effects of limb autotomy and tethering on juvenile blue crab survival from cannibalism. Marine Ecology Progress Series, 116:65–74, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/116/m116p065.pdf.

Sancho:1997:BOU

[SML97] G. Sancho, D. Ma, and P. Lobel. Behavioral observations of an upcurrent reef colonization event by larval surgeonfish Ctenochaetus strigosus (Acanthuridae). Marine Ecology Progress Series, 153:311–315, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/153/m153p311.pdf.

Street:1997:IBT

[SMP97] G. T. Street, P. A. Montagna, and P. L. Parker. Incorporation of brown tide into an estuarine food web. Marine Ecology Progress Series, 152:67–78, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/152/m152p067.pdf.

Sarda:1997:SSM

[SMT97] F. Sardà, F. Maynou, and L. Talló. Seasonal and spatial mobility patterns of rose shrimp Aristeus antennatus in the Western Mediterranean: results of a long-term study. Marine Ecology Progress Series, 159:133–141, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
159/m159p133.pdf.

Singh:1999:PST

[SMTL99] R. Singh, B. A. MacDonald, M. L. H. Thomas, and P. Lawton. Patterns of seasonal and tidal feeding activity in the dendrochirote sea cucumber Cucumaria frondosa (Echinodermata: Holothuroidea) in the Bay of Fundy, Canada. Marine Ecology Progress Series, 187:133–145, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
187/m187p133.pdf.

Sainte-Marie:1999:MCC

[SMUS^{+}99] B. Sainte-Marie, N. Urbani, J. M. Sévigny, F. Hazel, and U. Kuhnlein. Multiple choice criteria and the dynamics of assortative mating during the first breeding season of female snow crab Chionoecetes opilio (Brachyura, Majidae). Marine Ecology Progress Series, 181:141–153, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
181/m181p141.pdf.

Shenker:1993:OTS

[SMW^{+}93] J. M. Shenker, E. D. Maddox, E. Wishinski, A. Pearl, S. R. Thorrold, and N. Smith. Onshore transport of settlementstage Nassau grouper Epinephelus striatus and other fishes in Exuma Sound, Bahamas. Marine Ecology Progress Series, 98:31–43, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/98/m098p031.pdf.

Schlacher:1998:SSB

[SNC^{+}98] T. A. Schlacher, P. Newell, J. Clavier, M. A. Schlacher-Hoenlinger, C. Chevillon, and J. Britton. Soft-sediment benthic community structure in a coral reef lagoon — the prominence of spatial heterogeneity and ‘spot endemism’. Marine Ecology Progress Series, 174:159–174, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
174/m174p159.pdf.

Sime-Ngando:1992:CPC

[SNJV92] T. Sime-Ngando, K. Juniper, and A. Vézina. Ciliated protozoan communities over Cobb Seamount: increase in biomass and spatial patchiness. Marine Ecology Progress Series, 89: 37–51, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/89/m089p037.pdf.

Steward:1996:MBC

[SNR^{+}96] C. C. Steward, S. C. Nold, D. B. Ringelberg, D. C. White, and C. R. Lovell. Microbial biomass and community structures in the burrows of bromophenol producing and non-producing marine worms and surrounding sediments. Marine Ecology Progress Series, 133:149–165, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/133/m133p149.pdf.

Sampou:1991:SPS

[SO91] P. Sampou and C. A. Oviatt. Seasonal patterns of sedimentary carbon and anaerobic respiration along a simulated eutrophication gradient. Marine Ecology Progress Series, 72: 271–282, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/72/m072p271.pdf.

Sabates:1996:VLF

[SO96a] A. Sabatés and M. P. Olivar. Variation of larval fish distributions associated with variability in the location of a shelfslope front. Marine Ecology Progress Series, 135:11–20, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
135/m135p011.pdf.

Sogard:1996:FDA

[SO96b] S. M. Sogard and B. L. Olla. Food deprivation affects vertical distribution and activity of a marine fish in a thermal gradient: potential energy-conserving mechanisms. Marine Ecology Progress Series, 133:43–55, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/133/m133p043.pdf.

Somerfield:1997:FET

[SOC97] P. J. Somerfield, F. Olsgard, and M. R. Carr. A further examination of two new taxonomic distinctness measures. Marine Ecology Progress Series, 154:303–306, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
154/m154p303.pdf.

Sogard:1992:VGR

[Sog92] S. M. Sogard. Variability in growth rates of juvenile fishes in different estuarine habitats. Marine Ecology Progress Series, 85:35–53, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/85/m085p035.pdf.

Sato:1998:TAS

[SOK98] S. Sato, T. Ogata, and M. Kodama. Trace amounts of saxitoxins in the viscera of chum salmon Oncorhynchus keta. Marine Ecology Progress Series, 175:295–298, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
175/m175p295.pdf.

Sommer:1994:MDF

[Som94] U. Sommer. Are marine diatoms favoured by high Si:N ratios? Marine Ecology Progress Series, 115:309–315, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
115/m115p309.pdf.

Sommer:1996:NCE

[Som96] U. Sommer. Nutrient competition experiments with periphyton from the Baltic Sea. Marine Ecology Progress Series, 140:161–167, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/140/m140p161.pdf.

Sakamoto:1992:COP

[SOS^{+}92] S. Sakamoto, T. Ogata, S. Sato, M. Kodama, and T. Takeuchi. Causative organism of paralytic shellfish toxins other than dinoflagellates. Marine Ecology Progress Series, 89:229–235, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/89/m089p229.pdf.

Sosik:1996:BOM

[Sos96] H. M. Sosik. Bio-optical modeling of primary production: consequences of variability in quantum yield and specific absorption. Marine Ecology Progress Series, 143:225–238, 1996.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
143/m143p225.pdf.

Schrijvers:1995:IEM

[SOSV95] J. Schrijvers, J. Okondo, M. Steyaert, and M. Vincx. Influence of epibenthos on meiobenthos of the Ceriops tagal mangrove sediment at Gazi Bay, Kenya. Marine Ecology Progress Series, 128:247–259, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/128/m128p247.pdf.

Sary:1997:EIT

[SOW97] Z. Sary, H. A. Oxenford, and J. D. Woodley. Effects of an increase in trap mesh size on an overexploited coral reef fishery at Discovery Bay, Jamaica. Marine Ecology Progress Series, 154:107–120, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/154/m154p107.pdf.

Sonoda:1993:ESC

[SP93] K. Sonoda and V. J. Paul. Effect of stony corals extracts on feeding by Acanthaster planci. Marine Ecology Progress Series, 102:161–168, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/102/m102p161.pdf.

Sommer:1999:EAM

[SP99a] A. Sommer and H. O. Pörtner. Exposure of Arenicola marina to extreme temperatures: adaptive flexibility of a boreal and a subpolar population. Marine Ecology Progress Series, 181:215–226, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/181/m181p215.pdf.

Stiger:1999:STP

[SP99b] V. Stiger and C. E. Payri. Spatial and temporal patterns of settlement of the brown macroalgae Turbinaria ornata and Sargassum mangarevense in a coral reef on Tahiti. Marine Ecology Progress Series, 191:91–100, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
191/m191p091.pdf.

Sukhotin:1999:HFI

[SP99c] A. A. Sukhotin and H. O. Pörtner. Habitat as a factor involved in the physiological response to environmental anaerobiosis of White Sea Mytilus edulis. Marine Ecology Progress Series, 184:149–160, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/184/m184p149.pdf.

Saburova:1995:SSI

[SPB95] Maria A. Saburova, Igor G. Polikarpov, and Igor V. Burkovsky. Spatial structure of an intertidal sandflat microphytobenthic community as related to different spatial scales. Marine Ecology Progress Series, 129:229–239, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
129/m129p229.pdf.

Santos:1999:SCS

[SPB^{+}99a] M. B. Santos, G. J. Pierce, P. R. Boyle, R. J. Reid, H. M. Ross, I. A. P. Patterson, C. C. Kinze, S. Tougaard, R. Lick, U. Piatkowski, and V. Hernández-García. Stomach contents of sperm whales Physeter macrocephalus stranded in the North Sea 1990–1996. Marine Ecology Progress Series, 183:281–294, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/183/m183p281.pdf.

Stewart:1999:IPD

[SPB^{+}99b] F. M. Stewart, R. A. Phillips, J. A. Bartle, J. Craig, and D. Shooter. Influence of phylogeny, diet, moult schedule and sex on heavy metal concentrations in New Zealand procellariiformes. Marine Ecology Progress Series, 178:295–305, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
178/m178p295.pdf.

Schofield:1993:LSP

[SPBS93] O. Schofield, B. B. Prezélin, Ft. R. Bidigare, and R. C. Smith. In situ photosynthetic quantum yield. Correspondence to hydrographic and optical variability within the Southern California Bight. Marine Ecology Progress Series, 93:25–37, 1993.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
93/m093p025.pdf.

Stewart:1997:ISA

[SPCF97] F. M. Stewart, R. A. Phillips, P. Catry, and R. W. Furness. Influence of species, age and diet on mercury concentrations in Shetland seabirds. Marine Ecology Progress Series, 151:237–244, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/151/m151p237.pdf.

Stoecker:1991:PMR

[SPDM91] D. K. Stoecker, M. Putt, L. H. Davis, and A. E. Michaels. Photosynthesis in Mesodinium rubrum: species-specific measurements and comparison to community rates. Marine Ecology Progress Series, 73:245–252, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/73/m073p245.pdf.

Speare:1995:PBT

[Spe95] P. Speare. Parasites as biological tags for sailfish Istiophorus platypterus from east coast Australian waters. Marine Ecology Progress Series, 118:43–50, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/118/m118p043.pdf.

Slezak:1994:PRA

[SPH94] D. M. Slezak, S. Puskaric, and G. J. Herndl. Potential role of acrylic acid in bacterioplankton communities in the sea. Marine Ecology Progress Series, 105:191–197, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
105/m105p191.pdf.

Steward:1992:BNA

[SPPL92] C. C. Steward, J. Pinckney, Y. Piceno, and C. R. Lovell. Bacterial numbers and activity, microalgal biomass and productivity, and meiofaunal distribution in sediments naturally contaminated with biogenic bromophenols. Marine Ecology Progress Series, 90:61–71, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/90/m090p061.pdf.

Sprung:1993:EMS

[Spr93] M. Sprung. Estimating macrobenthic secondary production from body weight and biomass: a field test in a non-boreal intertidal habitat. Marine Ecology Progress Series, 100:103– 109, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/100/m100p103.pdf.

Schofield:1991:VSN

[SPS^{+}91] O. Schofield, B. B. Prezelin, R. C. Smith, P. M. Stegmann, N. B. Nelson, M. R. Lewis, and K. S. Baker. Variability in spectral and nonspectral measurements of photosynthetic light utilization efficiencies. Marine Ecology Progress Series, 78:253–271, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/78/m078p253.pdf.

Sathyendranath:1996:SBO

[SPS^{+}96] S. Sathyendranath, T. Platt, V. Stuart, B. D. Irwin, M. J. W. Veldhuis, G. W. Kraay, and W. G. Harrison. Some biooptical characteristics of phytoplankton in the NW Indian Ocean. Marine Ecology Progress Series, 132:299–311, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
132/m132p299.pdf.

Steinberg:1998:CZA

[SPS98] D. K. Steinberg, C. H. Pilskaln, and M. W. Silver. Contribution of zooplankton associated with detritus to sediment trap ’swimmer’ carbon in Monterey Bay, California, USA. Marine Ecology Progress Series, 164:157–166, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
164/m164p157.pdf.

SubbaRao:1993:DSP

[SPZ^{+}93] D. V. Subba Rao, Y. Pan, V. Zitko, G. Bugden, and K. Mackeigan. Diarrhetic shellfish poisoning (DSP) associated with a sub-surface bloom of Dinophysis norvegica in Bedford Basin, eastern Canada. Marine Ecology Progress Series, 97:117– 126, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/97/m097p117.pdf.

Sammarco:1990:LSP

[SR90a] P. W. Sammarco and M. J. Risk. Large-scale patterns in internal bioerosion of Porites: cross continental shelf trends on the Great Barrier Reef. Marine Ecology Progress Series, 59:145–156, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/59/m059p145.pdf.

Scheibling:1990:CDS

[SR90b] R. E. Scheibling and B. G. Raymond. Community dynamics on a subtidal cobble bed following mass mortalities of sea urchins. Marine Ecology Progress Series, 63:127–145, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
63/m063p127.pdf.

Selmer:1993:END

[SR93a] Johanne-Sophie Selmer and Lars Rydberg. Effects of nutrient discharge by river water and waste water on the nitrogen dynamics in the archipelago of Göteborg, Sweden. Marine Ecology Progress Series, 92:119–133, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/92/m092p119.pdf.

Shanks:1993:RMS

[SR93b] A. L. Shanks and M. L. Reeder. Reducing microzones and sulfide production in marine snow. Marine Ecology Progress Series, 96:43–47, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/96/m096p043.pdf.

Sweatman:1994:GHP

[SR94] H. Sweatman and D. R. Robertson. Grazing halos and predation on juvenile Caribbean surgeonfishes. Marine Ecology Progress Series, 111:1–6, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/111/m111p001.pdf.

Slawyk:1995:SPS

[SR95] G. Slawyk and P. Raimbault. Simple procedure for simultaneous recovery of dissolved inorganic and organic nitrogen in ^{1}5N-tracer experiments and improving the isotopic mass balance. Marine Ecology Progress Series, 124:289–299, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
124/m124p289.pdf.

Stoner:1996:SRP

[SR96] A. W. Stoner and M. Ray. Shell remains provide clues to historical distribution and abundance patterns in a large seagrass-associated gastropod (Strombus gigas). Marine Ecology Progress Series, 135:101–108, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/135/m135p101.pdf.

Stokesbury:1997:SDA

[SR97] K. D. E. Stokesbury and S. W. Ross. Spatial distribution and an absolute density estimate of juvenile spot Leiostomus xanthurus in the tidal fringe bordering a North Carolina salt marsh. Marine Ecology Progress Series, 149:289–294, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
149/m149p289.pdf.

Schumann:1998:SPO

[SR98a] R. Schumann and D. Rentsch. Staining particulate organic matter with DTAF — a fluorescence dye for carbohydrates and protein: a new approach and application of a 2D image analysis system. Marine Ecology Progress Series, 163: 77–88, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/163/m163p077.pdf.

Shaw:1998:NVD

[SR98b] W. Shaw and C. L. K. Robinson. Night versus day abundance estimates of zooplankton at two coastal stations in British Columbia, Canada. Marine Ecology Progress Series, 175:143–153, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/175/m175p143.pdf.

Shillito:1999:PDC

[SRG^{+}99] B. Shillito, J. Ravaux, F. Gaill, J. Delachambre, E. Thiébaut, and J. J. Childress. Preliminary data on carbon production of deep-sea vent tubeworms. Marine Ecology Progress Series, 183:275–279, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/183/m183p275.pdf.

Sala:1998:TVA

[SRH^{+}98] E. Sala, M. Ribes, B. Hereu, M. Zabala, V. Alvà, R. Coma, and J. Garrabou. Temporal variability in abundance of the sea urchins Paracentrotus lividus and Arbacia lixula in the northwestern Mediterranean: comparison between a marine reserve and an unprotected area. Marine Ecology Progress Series, 168:135–145, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/168/m168p135.pdf.

Sarrazin:1997:BGD

[SRJD97] J. Sarrazin, V. Robigou, S. K. Juniper, and J. R. Delaney. Biological and geological dynamics over four years on a hightemperature sulfide structure at the Juan de Fuca Ridge hydrothermal observatory. Marine Ecology Progress Series, 153: 5–24, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/153/m153p005.pdf.

Skalamera:1999:NEG

[SRRdM99] J. P. Skalamera, F. Renaud, M. Raymond, and T. de Meeûs. No evidence for genetic differentiation of the mussel Mytilus galloprovincialis between lagoons and the seaside. Marine Ecology Progress Series, 178:251–258, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
178/m178p251.pdf.

Shivji:1992:RIH

[SRS92] M. S. Shivji, S. O. Rogers, and M. J. Stanhope. Rapid isolation of high molecular weight DNA from marine macroalgae. Marine Ecology Progress Series, 84:197–203, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
84/m084p197.pdf.

Sammarco:1999:CCS

[SRSH99] P. W. Sammarco, M. J. Risk, H. P. Schwarcz, and J. M. Heikoop. Cross-continental shelf trends in coral δ ^{15}N on the Great Barrier Reef: further consideration of the reef nutrient paradox. Marine Ecology Progress Series, 180:131–138, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
180/m180p131.pdf.

Somerfield:1995:ICS

[SRW95a] P. J. Somerfield, H. L. Rees, and R. M. Warwick. Interrelationships in community structure between shallow-water marine meiofauna and macrofauna in relation to dredgings disposal. Marine Ecology Progress Series, 127:103–112, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
127/m127p103.pdf.

Stoner:1995:ELH

[SRW95b] A. W. Stoner, M. Ray, and J. M. Waite. Effects of a large herbivorous gastropod on macrofauna communities in tropical seagrass meadows. Marine Ecology Progress Series, 121: 125–137, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/121/m121p125.pdf.

Steele:1991:ESS

[SS91] D. H. Steele and V. J. Steele. Effects of salinity on the survival, growth rate, and reproductive output of Gammarus lawrencianus (Crustacea, Amphipoda). Marine Ecology Progress Series, 78:49–56, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/78/m078p049.pdf.

Southward:1992:DPT

[SS92] A. J. Southward and E. C. Southward. Distribution of Pogonophora (tube-worms) in British Columbian fjords. Marine Ecology Progress Series, 82:227–233, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
82/m082p227.pdf.

Smith:1993:EPH

[SS93] S. D. A. Smith and R. D. Simpson. Effects of pollution on holdfast macrofauna of the kelp Ecklonia radiata: discrimination at different taxonomic levels. Marine Ecology Progress Series, 96:199–208, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/96/m096p199.pdf.

Smith:1995:EVN

[SS95] S. D. A. Smith and R. D. Simpson. Effects of the ‘Nella Dan’ oil spill on the fauna of Durvillaea antarctica holdfasts. Marine Ecology Progress Series, 121:73–89, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
121/m121p073.pdf.

Schizas:1996:SCS

[SS96a] N. V. Schizas and T. C. Shirley. Seasonal changes in structure of an Alaskan intertidal meiofaunal assemblage. Marine Ecology Progress Series, 133:115–124, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/133/m133p115.pdf.

Strom:1996:MGG

[SS96b] S. L. Strom and M. W. Strom. Microplankton growth, grazing, and community structure in the northern Gulf of Mexico. Marine Ecology Progress Series, 130:229–240, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
130/m130p229.pdf.

Seitzinger:1997:CDO

[SS97] S. P. Seitzinger and R. W. Sanders. Contribution of dissolved organic nitrogen from rivers to estuarine eutrophication. Marine Ecology Progress Series, 159:1–12, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
159/m159p001.pdf.

Shimeta:1999:TST

[SS99a] J. Shimeta and J. D. Sisson. Taxon-specific tidal resuspension of protists into the subtidal benthic boundary layer of a coastal embayment. Marine Ecology Progress Series, 177: 51–62, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/177/m177p051.pdf.

Smith:1999:DDI

[SS99b] K. A. Smith and I. M. Suthers. Displacement of diverse ichthyoplankton assemblages by a coastal upwelling event on the Sydney shelf. Marine Ecology Progress Series, 176: 49–62, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/176/m176p049.pdf.

Steward:1996:APB

[SSA96] G. F. Steward, D. C. Smith, and F. Azam. Abundance and production of bacteria and viruses in the Bering and Chukchi Seas. Marine Ecology Progress Series, 131:287–300, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
131/m131p287.pdf.

Sorokin:1996:SFA

[SSG96] Yu. I. Sorokin, P. Yu. Sorokin, and A. Gnes. Structure and functioning of the anthropogenically transformed Comacchio lagoonal ecosystem (Ferrara, Italy). Marine Ecology Progress Series, 133:57–71, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/133/m133p057.pdf.

Sorokin:1996:SEL

[SSGD96] Yu. I. Sorokin, P. Yu. Sorokin, O. Giovanardi, and L. Dalla Venezia. Study of the ecosystem of the lagoon of Venice, with emphasis on anthropogenic impact. Marine Ecology Progress Series, 141:247–261, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/141/m141p247.pdf.

Schnack-Schiel:1991:SCC

[SSHM91] S. B. Schnack-Schiel, W. Hagen, and E. Mizdalski. Seasonal comparison of Calanoides acutus and Calanus propinquus (Copepoda: Calanoida) in the southeastern Weddell Sea, Antarctica. Marine Ecology Progress Series, 70:17–27, 1991.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
70/m070p017.pdf.

Sherr:1991:CRM

[SSM91] E. B. Sherr, B. F. Sherr, and J. McDaniel. Clearance rates of < 6 μ m fluorescently labeled algae (FLA) by estuarine protozoa. potential grazing impact of flagellates and ciliates. Marine Ecology Progress Series, 69:81–92, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
69/m069p081.pdf.

Steinberg:1997:RMZ

[SSP97] D. K. Steinberg, M. W. Silver, and C. H. Pilskaln. Role of mesopelagic zooplankton in the community metabolism of giant larvacean house detritus in Monterey Bay, California, USA. Marine Ecology Progress Series, 147:167–179, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
147/m147p167.pdf.

Schlunzen:1997:AIL

[SSR^{+}97] K. H. Schlünzen, T. Stahlschmidt, A. Rebers, U. Niemeier, M. Kriews, and W. Dannecker. Atmospheric input of lead into the German Bight — a high resolution measurement and model case study. Marine Ecology Progress Series, 156: 299–309, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/156/m156p299.pdf.

Saiz-Salinas:1999:BSS

[SSR99] J. I. Saiz-Salinas and A. Ramos. Biomass size-spectra of macrobenthic assemblages along water depth in Antarctica. Marine Ecology Progress Series, 178:221–227, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
178/m178p221.pdf.

Stamhuis:1997:BAT

[SSV97] E. J. Stamhuis, C. E. Schreurs, and J. J. Videler. Burrow architecture and turbative activity of the thalassinid shrimp Callianassa subterranea from the central North Sea.

Marine Ecology Progress Series, 151:155–163, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
151/m151p155.pdf.

Sheehy:1996:AEL

[SSW^{+}96] M. R. J. Sheehy, P. M. J. Shelton, J. F. Wickins, M. Belchier, and E. Gaten. Ageing the European lobster Homarus gammarus by the lipofuscin in its eyestalk ganglia. Marine Ecology Progress Series, 143:99–111, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/143/m143p099.pdf.

Sieracki:1998:IFS

[SSY98] C. K. Sieracki, M. E. Sieracki, and C. S. Yentsch. An imagingin-flow system for automated analysis of marine microplankton. Marine Ecology Progress Series, 168:285–296, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
168/m168p285.pdf.

Slagstad:1990:GPD

[ST90] D. Slagstad and K. S. Tande. Growth and production dynamics of Calanus glacialis in an Arctic pelagic food web. Marine Ecology Progress Series, 63:189–199, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
63/m063p189.pdf.

Suderman:1998:AFH

[ST98] K. Suderman and D. Thistle. Adult female harpacticoid copepods maintain their energy reserves by feeding while suspended during storms. Marine Ecology Progress Series, 164: 245–252, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/164/m164p245.pdf.

Svensen:1999:SCF

[ST99] C. Svensen and K. Tande. Sex change and female dimorphism in Calanus finmarchicus. Marine Ecology Progress Series, 176:93–102, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/176/m176p093.pdf.

Salini:1999:FCT

[STBR99] J. P. Salini, M. Tonks, S. J. M. Blaber, and J. Ross. Feeding of captive, tropical carcharhinid sharks from the Embley River estuary, northern Australia. Marine Ecology Progress Series, 184:309–314, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/184/m184p309.pdf.

Steinberg:1994:LST

[Ste94] P. D. Steinberg. Lack of short-term induction of phtorotannins in the Australasian brown algae Ecktoma radiate and Sargassum vestitum. Marine Ecology Progress Series, 112: 129–133, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/112/m112p129.pdf.

Steinberg:1995:IBC

[Ste95] P. D. Steinberg. Interaction between the canopy dwelling echinoid Holopneustes purpurescens and its host kelp Ecklonia radiata. Marine Ecology Progress Series, 127:169–181, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
127/m127p169.pdf.

Scharek:1999:DFD

[STK99] R. Scharek, L. M. Tupas, and D. M. Karl. Diatom fluxes to the deep sea in the oligotrophic North Pacific gyre at Station ALOHA. Marine Ecology Progress Series, 182:55–67, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
182/m182p055.pdf.

Shiomoto:1997:TRS

[STNI97] A. Shiomoto, K. Tadokoro, K. Nagasawa, and Y. Ishida. Trophic relations in the subarctic North Pacific ecosystem: possible feeding effect from pink salmon. Marine Ecology Progress Series, 150:75–85, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/150/m150p075.pdf.

Stokes:1996:LSP

[Sto96] M. D. Stokes. Larval settlement, post-settlement growth and secondary production of the Florida lancelet (= amphioxus)

Branchiostoma floridae. Marine Ecology Progress Series, 130: 71–84, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/130/m130p071.pdf.

Stobutzki:1997:ECS

[Sto97] I. C. Stobutzki. Energetic cost of sustained swimming in the late pelagic stages of reef fishes. Marine Ecology Progress Series, 152:249–259, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/152/m152p249.pdf.

Strom:1991:GGR

[Str91] S. L. Strom. Growth and grazing rates of the herbivorous dinoflagellate Gymnodinium sp. from the open subarctic Pacific Ocean. Marine Ecology Progress Series, 78:103–113, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
78/m078p103.pdf.

Stratoudakis:1999:RIE

[Str99] Y. Stratoudakis. Recommendations for improving the estimation of the number of seabirds potentially sustained by fisheries discards. Marine Ecology Progress Series, 176:307– 309, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/176/m176p307.pdf.

Spicer:1999:PPA

[STS99] J. I. Spicer, M. A. Thomasson, and J. O. Strömberg. Possessing a poor anaerobic capacity does not prevent the diel vertical migration of Nordic krill Meganyctiphanes norvegica into hypoxic waters. Marine Ecology Progress Series, 185: 181–187, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/185/m185p181.pdf.

Styan:1997:IPS

[Sty97] C. A. Styan. Inexpensive and portable sampler for collecting eggs of free-spawning marine invertebrates underwater. Marine Ecology Progress Series, 150:293–296, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
150/m150p293.pdf.

Skilleter:1993:EHC

[SU93] G. A. Skilleter and A. J. Underwood. Effects of habitat composition on recruitment of cerithid gastropods in sediments at One Tree Reef, Great Barrier Reef. Marine Ecology Progress Series, 93:155–163, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/93/m093p155.pdf.

Sadok:1995:FHA

[SUH95] S. Sadok, R. Uglow, and S. Haswell. Fluxes of haemolymph ammonia and free amino acids in Mytilus edulis exposed to ammonia. Marine Ecology Progress Series, 129:177–187, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
129/m129p177.pdf.

Sundby:1996:TIC

[Sun96] S. Sundby. Turbulence-induced contact rates in plankton: the issue of scales. Marine Ecology Progress Series, 139: 305–307, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/139/m139p301.pdf.

Sundermann:1997:PPI

[Sün97] J. Sündermann. The PRISMA project: an investigation of processes controlling contaminant fluxes in the German Bight. Marine Ecology Progress Series, 156:239–243, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
156/m156p239.pdf.

Suttle:1992:IPP

[Sut92] C. A. Suttle. Inhibition of photosynthesis in phytoplankton by the submicron size fraction concentrated from seawater. Marine Ecology Progress Series, 87:105–112, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
87/m087p105.pdf.

Soetaert:1993:STP

[SV93a] K. Soetaert and P. Van Rijswijk. Spatial and temporal patterns of the zooplankton in the Westerschelde estuary. Marine Ecology Progress Series, 97:47–59, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/97/m097p047.pdf.

Stefels:1993:PDD

[Sv93b] J. Stefels and W. H. M. van Boekel. Production of DMS from dissolved DMSP in axenic cultures of the marine phytoplankton species Phaeocystis sp. Marine Ecology Progress Series, 97:11–18, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/97/m097p011.pdf.

Smaal:1997:SVC

[SV97] A. C. Smaal and A. P. M. A. Vonck. Seasonal variation in C, N and P budgets and tissue composition of the mussel Mytilus edulis. Marine Ecology Progress Series, 153:167–179, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
153/m153p167.pdf.

Service:1991:UNM

[SW91] S. K. Service and K. Walters. Use of null models to assess patterns of size class migrations in the copepod Zausodes arenicolus. Marine Ecology Progress Series, 69:17–23, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
69/m069p017.pdf.

Shields:1993:IPR

[SW93] J. D. Shields and R. E. I. Wood. Impact of parasites on the reproduction and fecundity of the blue sand crab Portunus pelagicus from Moreton Bay, Australia. Marine Ecology Progress Series, 92:159–170, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/92/m092p159.pdf.

Shanks:1997:HMM

[SW97] A. L. Shanks and K. Walters. Holoplankton, meroplankton, and meiofauna associated with marine snow. Marine Ecology Progress Series, 156:75–86, 1997. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/156/m156p075.pdf.

Schratzberger:1998:EIF

[SW98a] M. Schratzberger and R. M. Warwick. Effects of the intensity and frequency of organic enrichment on two estuarine nematode communities. Marine Ecology Progress Series, 164:83–94, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/164/m164p083.pdf.

Shreeve:1998:MGE

[SW98b] R. S. Shreeve and P. Ward. Moulting and growth of the early stages of two species of Antarctic calanoid copepod in relation to differences in food supply. Marine Ecology Progress Series, 175:109–119, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/175/m175p109.pdf.

Schratzberger:1999:DEV

[SW99] M. Schratzberger and R. M. Warwick. Differential effects of various types of disturbances on the structure of nematode assemblages: an experimental approach. Marine Ecology Progress Series, 181:227–236, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/181/m181p227.pdf.

Simpson:1991:BPR

[SWEC91] R. Simpson, R. Williams, R. Ellis, and P. F. Culverhouse. Biological pattern recognition by neural networks. Marine Ecology Progress Series, 79:303–308, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/79/m079p303.pdf.

Steinke:1998:PCD

[SWK98] M. Steinke, G. V. Wolfe, and G. O. Kirst. Partial characterisation of dimethylsulfoniopropionate (DMSP) lyase isozymes in 6 strains of Emiliania huxleyi. Marine Ecology Progress Series, 175:215–225, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/175/m175p215.pdf.

Sanudo-Wilhelmy:1999:BDO

[SWT99] S. A. Sañudo-Wilhelmy and G. T. Taylor. Bacterioplankton dynamics and organic carbon partitioning in the lower Hudson River estuary. Marine Ecology Progress Series, 182: 17–27, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/182/m182p017.pdf.

Syms:1995:MSA

[Sym95] C. Syms. Multi-scale analysis of habitat association in a guild of blennioid fishes. Marine Ecology Progress Series, 125: 31–43, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/125/m125p031.pdf.

Sala:1996:FPS

[SZ96] E. Sala and M. Zabala. Fish predation and the structure of the sea urchin Paracentrotus lividus populations in the NW Mediterranean. Marine Ecology Progress Series, 140: 71–81, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/140/m140p071.pdf.

Sanjuan:1994:MGM

[SZA94] A. Sanjuan, C. Zapata, and G. Alvarez. Mytilus galloprovincialis and M. edulis on the coasts of the Iberian Peninsula. Marine Ecology Progress Series, 113:131–146, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
113/m113p131.pdf.

Szmant:1991:SRC

[Szm91] A. M. Szmant. Sexual reproduction by the Caribbean reef corals Montastrea annularis and M. cavernosa. Marine Ecology Progress Series, 74:13–25, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/74/m074p013.pdf.

Tietjen:1990:PGE

[TA90] J. H. Tietjen and D. M. Alongi. Population growth and effects of nematodes on nutrient regeneration and bacteria associated with mangrove detritus from northeastern Queensland (Australia). Marine Ecology Progress Series, 68:169–179, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
68/m068p169.pdf.

Taylor:1993:IDC

[TA93] D. I. Taylor and B. R. Allanson. Impacts of dense crab populations on carbon exchanges across the surface of a salt marsh. Marine Ecology Progress Series, 101:119–129, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
101/m101p119.pdf.

Taylor:1995:OCF

[TA95] D. I. Taylor and B. R. Allanson. Organic carbon fluxes between a high marsh and estuary, and the inapplicability of the outwelling hypothesis. Marine Ecology Progress Series, 120:263–270, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/120/m120p263.pdf.

Torres:1994:MAM

[TAD^{+}94] J. J. Torres, A. V. Aarset, J. Donnelly, T. L. Hopkins, T. M. Lancraft, and D. G. Ainley. Metabolism of Antarctic micronektonic crustacea as a function of depth of occurrence and season. Marine Ecology Progress Series, 113:207–219, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
113/m113p207.pdf.

Tedengren:1990:GPD

[TAJK90] M. Tedengren, C. André, K. Johannesson, and N. Kautsky. Genotypic and phenotypic differences between Baltic and North Sea populations of Mytilus edulis evaluated through reciprocal transplantations. III. Physiology. Marine Ecology Progress Series, 59:221–227, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/59/m059p221.pdf.

Takada:1995:VGR

[Tak95] Y. Takada. Variation of growth rate with tidal level in the gastropod Monodonta labio on a boulder shore. Marine Ecology Progress Series, 117:103–110, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/117/m117p103.pdf.

Takada:1996:VMD

[Tak96] Y. Takada. Vertical migration during the life history of the intertidal gastropod Monodonta labio on a boulder shore. Marine Ecology Progress Series, 130:117–123, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
130/m130p117.pdf.

Takada:1999:ISN

[Tak99] Y. Takada. Influence of shade and number of boulder layers on mobile organisms on a warm temperate boulder shore. Marine Ecology Progress Series, 189:171–179, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
189/m189p171.pdf.

Tamminen:1995:NAD

[Tam95] T. Tamminen. Nitrate and ammonium depletion rates and preferences during a Baltic spring bloom. Marine Ecology Progress Series, 120:123–133, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/120/m120p123.pdf.

Tanasichuk:1998:IVPa

[Tan98a] R. W. Tanasichuk. Interannual variations in the population biology and productivity of Euphausia pacifica in Barkley Sound, Canada, with special reference to the 1992 and 1993 warm ocean years. Marine Ecology Progress Series, 173:163– 180, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/173/m173p163.pdf.

Tanasichuk:1998:IVPb

[Tan98b] R. W. Tanasichuk. Interannual variations in the population biology and productivity of Thysanoessa spinifera in Barkley Sound, Canada, with special reference to the 1992 and 1993 warm ocean years. Marine Ecology Progress Series, 173:181–195, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/173/m173p181.pdf.

Taylor:1998:DBP

[Tay98] R. B. Taylor. Density, biomass and productivity of animals in four subtidal rocky reef habitats: the importance of small mobile invertebrates. Marine Ecology Progress Series, 172:37–51, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/172/m172p037.pdf.

Turon:1992:GSS

[TB92] X. Turon and M. A. Becerro. Growth and survival of several ascidian species from the northwestern Mediterranean. Marine Ecology Progress Series, 82:235–247, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
82/m082p235.pdf.

Turner:1993:PSB

[TB93] J. T. Turner and D. G. Borkman. Plankton studies in Buzzards Bay, Massachusetts USA. I. Hydrography and bacterioplankton, 1987 to 1990. Marine Ecology Progress Series, 100:17–26, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/100/m100p017.pdf.

Tupper:1995:SPS

[TB95] M. Tupper and R. G. Boutilier. Size and priority at settlement determine growth and competitive success of newly settled Atlantic cod. Marine Ecology Progress Series, 118: 295–300, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/118/m118p295.pdf.

Tupper:1997:EHS

[TB97] M. Tupper and R. G. Boutilier. Effects of habitat on settlement, growth, predation risk and survival of a temperate reef fish. Marine Ecology Progress Series, 151:225–236, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
151/m151p225.pdf.

Tholosan:1998:BDA

[TB98] O. Tholosan and A. Bianchi. Bacterial distribution and activity at the water-sediment boundary layer on NW Mediterranean continental margin. Marine Ecology Progress Series, 168:273–283, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/168/m168p273.pdf.

Trautman:1999:DEO

[TB99] D. A. Trautman and M. A. Borowitzka. Distribution of the epiphytic organisms on Posidonia australis and P. sinuosa, two seagrasses with differing leaf morphology. Marine Ecology Progress Series, 179:215–229, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/179/m179p215.pdf.

Tershy:1991:ICS

[TBAB91] B. R. Tershy, D. Breese, and S. Alvarez-Borrego. Increase in cetacean and seabird numbers in the Canal de Ballenas during an El Niño–Southern Oscillation event. Marine Ecology Progress Series, 69:299–302, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/69/m069p299.pdf.

Therriault:1990:PPV

[TBLD90] J.-C Therriault, D. Booth, L. Legendre, and S. Demers. Phytoplankton photoadaptation to vertical excursion as estimated by an in vivo fluorescence ratio. Marine Ecology Progress Series, 60:97–111, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/60/m060p097.pdf.

Thomas:1998:LSG

[TBY98] C. A. Thomas and L. I. Bendell-Young. Linking the sediment geochemistry of an intertidal region to metal bioavailability in the deposit feeder Macoma balthica. Marine Ecology Progress Series, 173:197–213, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/173/m173p197.pdf.

Thomas:1993:SIA

[TC93] C. J. Thomas and L. B. Cahoon. Stable isotope analyses differentiate between different trophic pathways supporting rocky-reef fishes. Marine Ecology Progress Series, 95: 19–24, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/95/m095p019.pdf.

Taylor:1994:MES

[TC94] R. B. Taylor and R. G. Cole. Mobile epifauna on subtidal brown seaweeds in northeastern New Zealand. Marine Ecology Progress Series, 115:271–282, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/115/m115p271.pdf.

Thybo-Christesen:1993:ICM

[TCB93] M. Thybo-Christesen and T. H. Blackburn. Internal ncycling, measured by ^{15}NH4+ dilution, in Cladophora sericea in a shallow Danish bay. Marine Ecology Progress Series, 100:283–286, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/100/m100p283.pdf.

Thorrold:1996:MSD

[TcBW96] S. R. Thorrold and D. M. c. B. Williams. Meso-scale distribution patterns of larval and pelagic juvenile fishes in the central Great Barrier Reef lagoon. Marine Ecology Progress Series, 145:17–31, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/145/m145p017.pdf.

Thybo-Christesen:1993:NFG

[TCRB93] M. Thybo-Christesen, M. B. Rasmussen, and T. H. Blackburn. Nutrient fluxes and growth of Cladophora sericea in a shallow Danish bay. Marine Ecology Progress Series, 100: 273–281, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/100/m100p273.pdf.

Tarling:1999:VMB

[TCRB99] G. A. Tarling, J. Cuzin-Roudy, and F. Buchholz. Vertical migration behaviour in the northern krill Meganyctiphanes norvegica is influenced by moult and reproductive processes. Marine Ecology Progress Series, 190:253–262, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
190/m190p253.pdf.

Tegner:1991:SUN

[TD91] M. J. Tegner and P. K. Dayton. Sea urchins, El Niños, and the long term stability of Southern California kelp forest communities. Marine Ecology Progress Series, 77:49–63, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
77/m077p049.pdf.

Thiebaut:1992:TOF

[TD92] E. Thiébaut and Y. Dauvin, J.-C, Lagadeuc. Transport of Owenia fusiformis larvae (Annelida: Polychaeta) in the Bay of Seme. I. Vertical distribution in relation to water column stratification and ontogenic vertical migration. Marine Ecology Progress Series, 80:29–39, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/80/m080p029.pdf.

Tumbiolo:1994:EMP

[TD94] M. L. Tumbiolo and J. A. Downing. An empirical model for the prediction of secondary production in marine benthic invertebrate populations. Marine Ecology Progress Series, 114:165–174, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/114/m114p165.pdf.

Tahey:1994:RBS

[TDBH94] T. M. Tahey, G. C. A. Duineveld, E. M. Berghuis, and W. Helder. Relation between sediment-water fluxes of oxygen and silicate and faunal abundance at continental shelf, slope and deep-water stations in the northwest Mediterranean. Marine Ecology Progress Series, 104:119–130, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
104/m104p119.pdf.

Thompson:1994:SCM

[TDD94] A. M. Thompson, E. G. Durbin, and A. G. Durbin. Seasonal changes in maximum ingestion rate of Acartia tonsa in Narragansett Bay, Rhode Island, USA. Marine Ecology Progress Series, 108:91–105, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/108/m108p091.pdf.

Tegner:1997:LSL

[TDER97] M. J. Tegner, P. K. Dayton, P. B. Edwards, and K. L. Riser. Large-scale, low-frequency oceanographic effects on kelp forest succession: a tale of two cohorts. Marine Ecology Progress Series, 146:117–134, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/146/m146p117.pdf.

Torres:1994:PCO

[TDH^{+}94] J. J. Torres, J. Donnelly, T. L. Hopkins, T. M. Lancraft, A. V. Aarset, and D. G. Ainley. Proximate composition and overwintering strategies of Antarctic micronektonic crustacea. Marine Ecology Progress Series, 113:221–232, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
113/m113p221.pdf.

Terrados:1997:EEA

[TDK97a] J. Terrados, C. M. Duarte, and W. J. Kenworthy. Experimental evidence for apical dominance in the seagrass Cymodocea nodosa. Marine Ecology Progress Series, 148:263–268, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
148/m148p263.pdf.

Terrados:1997:AGC

[TDK97b] J. Terrados, C. M. Duarte, and W. J. Kenworthy. Is the apical growth of Cymodocea nodosa dependent on clonal integration? Marine Ecology Progress Series, 158:103–110, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
158/m158p103.pdf.

Tang:1999:DDM

[TDVF99] K. W. Tang, H. G. Dam, P. T. Visscher, and T. D. Fenn. Dimethylsulfoniopropionate (DMSP) in marine copepods and its relation with diets and salinity. Marine Ecology Progress Series, 179:71–79, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/179/m179p071.pdf.

Thiebaut:1996:TTP

[TDW96] E. Thiébaut, J. C. Dauvin, and Z. Wang. Tidal transport of Pectinaria koreni postlarvae (Annelida: Polychaeta) in the Bay of Seine (eastern English Channel). Marine Ecology Progress Series, 138:63–70, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/138/m138p063.pdf.

Teegarden:1999:CGS

[Tee99] G. J. Teegarden. Copepod grazing selection and particle discrimination on the basis of PSP toxin content. Marine Ecology Progress Series, 181:163–176, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/181/m181p163.pdf.

Terazaki:1993:DSA

[Ter93] M. Terazaki. Deep-sea adaptation of the epipelagic chaetognath Sagitta elegans in the Japan Sea. Marine Ecology Progress Series, 98:79–88, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/98/m098p079.pdf.

Terrados:1997:LIV

[Ter97] J. Terrados. Is light involved in the vertical growth response of seagrasses when buried by sand? Marine Ecology Progress Series, 152:295–299, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/152/m152p295.pdf.

Townsend:1998:BPM

[TF98] E. C. Townsend and M. S. Fonseca. Bioturbation as a potential mechanism influencing spatial heterogeneity of North Carolina seagrass beds. Marine Ecology Progress Series, 169: 123–132, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/169/m169p123.pdf.

Tilstone:1994:UDS

[TFF94] G. H. Tilstone, F. G. Figueiras, and F. Fraga. Upwellingdownwelling sequences in the generation of red tides m a coastal upwelling system. Marine Ecology Progress Series, 112:241–253, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/112/m112p241.pdf.

Tilstone:1999:SNP

[TFFA99] G. H. Tilstone, F. G. Figueiras, E. G. Fermín, and B. Arbones. Significance of nanophytoplankton photosynthesis and primary production in a coastal upwelling system (Ría de Vigo, NW Spain). Marine Ecology Progress Series, 183:13–27, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
183/m183p013.pdf.

Tsutsumi:1990:RBG

[TFFS90] H. Tsutsumi, S. Fukunaga, N. Fujita, and M. Sumida. Relationship between growth of Capitella sp. and organic enrichment of the sediment. Marine Ecology Progress Series, 63: 157–162, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/63/m063p157.pdf.

Thompson:1995:DLT

[TFL95] D. R. Thompson, R. W. Furness, and S. A. Lewis. Diets and long-term changes in δ ^{15}N and δ ^{13}C values in northern fulmars Fulmarus glacialis from two northeast Atlantic colonies. Marine Ecology Progress Series, 125:13–11, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
125/m125p003.pdf.

Tholosan:1999:EHP

[TGB99] O. Tholosan, J. Garcin, and A. Bianchi. Effects of hydrostatic pressure on microbial activity through a 2000 m deep water column in the NW Mediterranean Sea. Marine Ecology Progress Series, 183:49–57, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/183/m183p049.pdf.

Talbot:1997:BAP

[TGBB97] V. Talbot, L. Giuliano, V. Bruni, and M. Bianchi. Bacterial abundance, production and ectoproteolytic activity in the Strait of Magellan. Marine Ecology Progress Series, 154: 293–302, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/154/m154p293.pdf.

Tester:1995:EPD

[TGG^{+}95] P. A. Tester, M. E. Geesey, C. Guo, H. W. Paerl, and D. F. Millie. Evaluating phytoplankton dynamics in the Newport River estuary (North Carolina, USA) by HPLC-derived pigment profiles. Marine Ecology Progress Series, 124:237–245, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/124/m124p237.pdf.

Taylor:1997:SLD

[TGG97] A. H. Taylor, R. J. Geider, and F. J. H. Gilbert. Seasonal and latitudinal dependencies of phytoplankton carbonto-chlorophyll-a ratios: results of a modelling study. Marine Ecology Progress Series, 152:51–66, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/152/m152p051.pdf.

Turon:1995:GPS

[TGLP95] X. Turon, G. Giribet, S. López, and C. Palacín. Growth and population structure of Paracentrotus lividus (Echinodermata: Echinoidea) in two contrasting habitats. Marine Ecology Progress Series, 122:193–204, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/122/m122p193.pdf.

Trimmer:1999:SBI

[TGSN99] M. Trimmer, R. J. Gowen, B. M. Stewart, and D. B. Nedwell. The spring bloom and its impact on benthic mineralisation rates in western Irish Sea sediments. Marine Ecology Progress Series, 185:37–46, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/185/m185p037.pdf.

Thomsen:1995:LSE

[TGvJW95] L. Thomsen, G. Graf, K. von Juterzenka, and U. Witte. An in situ experiment to investigate the depletion of seston above an interface feeder field on the continental slope of the western Barents Sea. Marine Ecology Progress Series, 123:295–300, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/123/m123p295.pdf.

Thompson:1990:FDF

[TH90] A. B. Thompson and R. T. Harrop. Feeding dynamics of fish larvae on Copepoda in the western Irish Sea, with particular reference to cod Gadus morhua. Marine Ecology Progress Series, 68:213–223, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/68/m068p213.pdf.

Taylor:1994:ESF

[TH94a] C. D. Taylor and B. L. Howes. Effect of sampling frequency on measurements of seasonal primary production and oxygen status in near-shore coastal ecosystems. Marine Ecology Progress Series, 108:193–203, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/108/m108p193.pdf.

Tupper:1994:RDC

[TH94b] M. Tupper and W. Hunte. Recruitment dynamics of coral reef fishes in Barbados. Marine Ecology Progress Series, 108: 225–235, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/108/m108p225.pdf.

Thain:1992:GAI

[Tha92a] J. Thain. Growth of the algae Isochrysis galbana and Tetraselmis suecica in water and sediment elutriate samples from the German Bight. Marine Ecology Progress Series, 91:229– 231, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/91/m091p229.pdf.

Thain:1992:UOC

[Tha92b] J. Thain. Use of the oyster Crassostrea gigas embryo bioassay on water and sediment elutriate samples from the German Bight. Marine Ecology Progress Series, 91:211–213, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
91/m091p211.pdf.

Thrush:1995:IHD

[THCD95] S. F. Thrush, J. E. Hewitt, V. J. Cummings, and P. K. Dayton. The impact of habitat disturbance by scallop dredging on marine benthic communities: what can be predicted from the results of experiments? Marine Ecology Progress Series, 129:141–150, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/129/m129p141.pdf.

Tuominen:1998:STV

[THKN98] L. Tuominen, A. Heinänen, J. Kuparinen, and L. P. Nielsen. Spatial and temporal variability of denitrification in the sediments of the northern Baltic proper. Marine Ecology Progress Series, 172:13–24, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/172/m172p013.pdf.

Thomsen:1991:TSS

[Tho91] L. Thomsen. Treatment and splitting of samples for bacteria and meiofauna biomass determinations by means of a semiautomatic image analysis system. Marine Ecology Progress Series, 71:301–306, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/71/m071p301.pdf.

Thorrold:1992:EPL

[Tho92] S. R. Thorrold. Evaluating the performance of light traps for sampling small fish and squid in open waters of the central Great Barrier Reef lagoon. Marine Ecology Progress Series, 89:277–285, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/89/m089p277.pdf.

Thrush:1996:AJI

[THPC96] S. F. Thrush, J. E. Hewitt, R. D. Pridmore, and V. J. Cummings. Adult/juvenile interactions of infaunal bivalves: contrasting outcomes in different habitats. Marine Ecology Progress Series, 132:83–92, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/132/m132p083.pdf.

Tuck:1998:EPT

[THR^{+}98] I. D. Tuck, S. J. Hall, M. R. Robertson, E. Armstrong, and D. J. Basford. Effects of physical trawling disturbance in a previously unfished sheltered Scottish sea loch. Marine Ecology Progress Series, 162:227–242, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/162/m162p227.pdf.

Taylor:1994:PES

[THS94] F. J. R. Taylor, R. Haigh, and T. F. Sutherland. Phytoplankton ecology of Sechelt Inlet, a fjord system on the British Columbia coast. II. Potentially harmful species. Marine Ecology Progress Series, 103:151–164, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/103/m103p151.pdf.

Tamminen:1996:UUK

[TI96] T. Tamminen and A. Irmisch. Urea uptake kinetics of a midsummer planktonic community on the SW coast of Finland. Marine Ecology Progress Series, 130:201–211, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
130/m130p201.pdf.

Tiselius:1998:LSV

[Tis98a] P. Tiselius. An in situ video camera for plankton studies: design and preliminary observations. Marine Ecology Progress Series, 164:293–299, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/164/m164p293.pdf.

Tiselius:1998:STF

[Tis98b] P. Tiselius. Short term feeding responses to starvation in three species of small calanoid copepods. Marine Ecology Progress Series, 168:119–126, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/168/m168p119.pdf.

Taylor:1990:SSA

[TJ90a] A. H. Taylor and I. Joint. A steady-state analysis of the ‘microbial loop’ in stratified systems. Marine Ecology Progress Series, 59:1–17, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/59/m059p001.pdf.

Tiselius:1990:FBS

[TJ90b] P. Tiselius and P. R. Jonsson. Foraging behaviour of six calanoid copepods: observations and hydrodynamic analysis. Marine Ecology Progress Series, 66:23–33, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
66/m066p023.pdf.

Tirasin:1999:EPD

[TJ99] M. E. Tirasin and T. Jørgensen. An evaluation of the precision of diet description. Marine Ecology Progress Series, 182:243–252, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/182/m182p243.pdf.

Trussell:1993:RDH

[TJRG93] G. C. Trussell, A. S. Johnson, S. G. Rudolph, and E. S. Gilfillan. Resistance to dislodgement: habitat and size-specific differences in morphology and tenacity in an intertidal snail. Marine Ecology Progress Series, 100:135–144, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
100/m100p135.pdf.

Thorrold:1998:ACJ

[TJST98] S. R. Thorrold, C. M. Jones, P. K. Swart, and T. E. Targett. Accurate classification of juvenile weakfish Cynoscion regalis to estuarine nursery areas based on chemical signatures in otoliths. Marine Ecology Progress Series, 173:253–265, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
173/m173p253.pdf.

Tupas:1991:SUR

[TK91] L. Tupas and I. Koike. Simultaneous uptake and regeneration of ammonium by mixed assemblages of heterotrophic marine bacteria. Marine Ecology Progress Series, 70:273–282, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
70/m070p273.pdf.

Takahashi:1995:IIZ

[TK95a] Kazutaka Takahashi and Kouichi Kawaguchi. Inter- and intraspecific zonation in three species of sand-burrowing mysids, Archaeomysis kokuboi, A. grebnitzkii and Iiella ohshimai, in Otsuchi Bay, northeastern Japan. Marine Ecology Progress Series, 116:75–84, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/116/m116p075.pdf.

Thetmeyer:1995:SSV

[TK95b] H. Thetmeyer and U. Kils. To see and not be seen: the visibility of predator and prey with respect to feeding behaviour. Marine Ecology Progress Series, 126:1–8, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
126/m126p001.pdf.

Townsend:1996:DDD

[TK96] D. W. Townsend and M. D. Keller. Dimethylsulfide (DMS) and dimethylsulfoniopropionate (DMSP) in relation to phytoplankton in the Gulf of Maine. Marine Ecology Progress Series, 137:229–241, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/137/m137p229.pdf.

Takahashi:1997:DTM

[TK97] K. Takahashi and K. Kawaguch. Diel and tidal migrations of the sand-burrowing mysids, Archaeomysis kokuboi, A. japonica and Iiella ohshimai, in Otsuchi Bay, northeastern Japan. Marine Ecology Progress Series, 148:95–107, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
148/m148p095.pdf.

Takahashi:1998:DFR

[TK98] K. Takahashi and K. Kawaguchi. Diet and feeding rhythm of the sand-burrowing mysids Archaeomysis kokuboi and A. japonica in Otsuchi Bay, northeastern Japan. Marine Ecology Progress Series, 162:191–199, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/162/m162p191.pdf.

Tomasko:1991:PBT

[TL91] D. A. Tomasko and B. E. Lapointe. Productivity and biomass of Thalassia testudinum as related to water column nutrient availability and epiphyte levels: field observations and experimental studies. Marine Ecology Progress Series, 75: 9–17, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/75/m075p009.pdf.

Therkildsen:1994:SVS

[TL94] M. S. Therkildsen and B. Aa. Lomstein. Seasonal variation in sediment urea turnover in a shallow estuary. Marine Ecology Progress Series, 109:77–82, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/109/m109p077.pdf.

Thomas:1995:PAD

[TL95] D. N. Thomas and R. J. Lara. Photodegradation of algal derived dissolved organic carbon. Marine Ecology Progress Series, 116:309–310, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/116/m116p309.pdf.

Taguchi:1993:TVC

[TLB93] S. Taguchi, E. A. Laws, and R. R. Bidigare. Temporal variability in chlorophyll-a and phaeopigment concentrations during incubations in the absence of grazers. Marine Ecology Progress Series, 101:45–53, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/101/m101p045.pdf.

Theilacker:1993:IAQ

[TLT93] G. H. Theilacker, N. C. H. Lo, and A. W. Townsend. An immunochemical approach to quantifying predation by euphausnds on the early stages of anchovy. Marine Ecology Progress Series, 92:35–50, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/92/m092p035.pdf.

Temara:1996:ECC

[TLW^{+}96] A. Temara, G. Ledent, M. Warnau, H. Paucot, M. Jangoux, and P. Dubois. Experimental cadmium contamination of Asterias rubens (Echinodermata). Marine Ecology Progress Series, 140:83–90, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/140/m140p083.pdf.

Turley:1994:BSA

[TM94] C. M. Turley and P. J. Mackie. Biogeochemical significance of attached and free-living bacteria and the flux of particles in the NE Atlantic Ocean. Marine Ecology Progress Series, 115:191–203, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/115/m115p191.pdf.

Thompson:1997:OET

[TM97] A. A. Thompson and B. D. Mapstone. Observer effects and training in underwater visual surveys of reef fishes. Marine Ecology Progress Series, 154:53–63, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/154/m154p053.pdf.

Thuesen:1998:EIO

[TMC98] E. V. Thuesen, C. B. Miller, and J. J. Childress. Ecophysiological interpretation of oxygen consumption rates and enzymatic activities of deep-sea copepods. Marine Ecology Progress Series, 168:95–107, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/168/m168p095.pdf.

Trathan:1998:USD

[TMCE98] P. N. Trathan, E. J. Murphy, J. P. Croxall, and I. Everson. Use of at-sea distribution data to derive potential foraging ranges of macaroni penguins during the breeding season. Marine Ecology Progress Series, 169:263–275, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
169/m169p263.pdf.

Tamigneaux:1997:GPS

[TMKL97] E. Tamigneaux, M. Mingelbier, B. Klein, and L. Legendre. Grazing by protists and seasonal changes in the size structure of protozooplankton and phytoplankton in a temperate nearshore environment (western Gulf of St. Lawrence, Canada). Marine Ecology Progress Series, 146:231–247, 1997.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
146/m146p231.pdf.

Toro:1990:DSG

[TN90] J. E. Toro and G. F. Newkirk. Divergent selection for growth rate in the European oyster Ostrea edulis response to selection and estimation of genetic parameters. Marine Ecology Progress Series, 62:219–227, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/62/m062p219.pdf.

Tunberg:1998:DCO

[TN98] B. G. Tunberg and W. G. Nelson. Do climatic oscillations influence cyclical patterns of soft bottom macrobenthic communities on the Swedish west coast? Marine Ecology Progress Series, 170:85–94, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/170/m170p085.pdf.

Taylor:1995:NLE

[TNGB95] D. Taylor, S. Nixon, S. Granger, and B. Buckley. Nutrient limitation and the eutrophication of coastal lagoons. Marine Ecology Progress Series, 127:235–244, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
127/m127p235.pdf.

Tang:1999:RSO

[TNKMK99] D. L. Tang, I. H. Ni, D. R. Kester, and F. E. Müller-Karger. Remote sensing observations of winter phytoplankton blooms southwest of the Luzon Strait in the South China Sea. Marine Ecology Progress Series, 191:43–51, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/191/m191p043.pdf.

Thuesen:1993:GVP

[TNN93] E. V. Thuesen, K. Numachi, and T. Nemoto. Genetic variation in the planktonic chaetognaths Parasagitta elegans and Eukrohnia hamata. Marine Ecology Progress Series, 101:243– 251, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/101/m101p243.pdf.

Thacker:1997:ERE

[TNP97] R. W. Thacker, D. G. Nagle, and V. J. Paul. Effects of repeated exposures to marine cyanobacterial secondary metabolites on feeding by juvenile rabbitfish and parrotfish. Marine Ecology Progress Series, 147:21–29, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/147/m147p021.pdf.

Trimmer:1998:NFT

[TNSM98] M. Trimmer, D. B. Nedwell, D. B. Sivyer, and S. J. Malcolm. Nitrogen fluxes through the lower estuary of the river Great Ouse, England: the role of the bottom sediments. Marine Ecology Progress Series, 163:109–124, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
163/m163p109.pdf.

Tokeshi:1995:PAD

[Tok95] M. Tokeshi. Polychaete abundance and dispersion patterns in mussel beds: a non-trivial ’infaunal’ assemblage on a Pacific South American rocky shore. Marine Ecology Progress Series, 125:137–147, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/125/m125p137.pdf.

Tolimieri:1998:CEM

[Tol98] N. Tolimieri. Contrasting effects of microhabitat use on largescale adult abundance in two families of Caribbean reef fishes. Marine Ecology Progress Series, 167:227–239, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
167/m167p227.pdf.

Tomascik:1991:SPC

[Tom91] T. Tomascik. Settlement patterns of Caribbean scleractinian corals on artificial substrata along a eutrophication gradient, Barbados, West Indies. Marine Ecology Progress Series, 77:261–269, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/77/m077p261.pdf.

tornDieck:1993:TTS

[tor93] I. torn Dieck. Temperature tolerance and survival in darkness of kelp gametophyles (Laminariales, Phaeophyta): ecological and biogeographical implications. Marine Ecology Progress Series, 100:253–264, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/100/m100p253.pdf.

Toro:1994:MDJ

[Tor94] Jorge E. Toro. In memoriam; Dr. Jürgen E. Winter (1937– 1994). Marine Ecology Progress Series, 112:211, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
112/m112p211.pdf.

Tudela:1995:DFI

[TP95] S. Tudela and I. Palomera. Diel feeding intensity and daily ration in the anchovy Engraulis encrasicolus in the northwest Mediterranean Sea during the spawning period. Marine Ecology Progress Series, 129:55–61, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/129/m129p055.pdf.

Tortell:1996:CTZ

[TP96] P. D. Tortell and N. M. Price. Cadmium toxicity and zinc limitation in centric diatoms of the genus Thalassiosira. Marine Ecology Progress Series, 138:245–254, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
138/m138p245.pdf.

Tudela:1997:TEE

[TP97] S. Tudela and I. Palomera. A trophic ecology of the European anchovy Engraulis encrasicolus in the Catalan Sea (northwest Mediterranean). Marine Ecology Progress Series, 160:121–134, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/160/m160p121.pdf.

Thrush:1991:IRF

[TPHC91] S. F. Thrush, R. D. Pridmore, J. E. Hewitt, and V. J. Cummings. Impact of ray feeding disturbances on sandflat macrobenthos: do communities dominated by polychaetes or shellfish respond differently? Marine Ecology Progress Series, 69:245–252, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/69/m069p245.pdf.

Thrush:1994:IPS

[TPHC94] S. F. Thrush, R. D. Pridmore, J. E. Hewitt, and V. J. Cummings. The importance of predators on a sandflat: interplay between seasonal changes in prey densities and predator effects. Marine Ecology Progress Series, 107:211–222, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
107/m107p211.pdf.

Taylor:1998:SAU

[TPR98] R. B. Taylor, J. T. A. Peek, and T. A. V. Rees. Scaling of ammonium uptake by seaweeds to surface area: volume ratio: geographical variation and the role of uptake by passive diffusion. Marine Ecology Progress Series, 169:143–148, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
169/m169p143.pdf.

Trathan:1993:SVA

[TPW^{+}93] P. N. Trathan, J. Priddle, J. L. Watkins, D. G. M. Miller, and A. W. A. Murray. Spatial variability of Antarctic krill in relation to mesoscale hydrography. Marine Ecology Progress Series, 98:61–71, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/98/m098p061.pdf.

Thebault:1991:SVP

[TR91] M. T. Thebault and J. P. Raffin. Seasonal variations in Palaemon serratus abdominal muscle metabolism and performance during exercise, as studied by ^{31}P NMR. Marine Ecology Progress Series, 74:175–183, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/74/m074p175.pdf.

Thingstad:1995:NLM

[TR95a] T. F. Thingstad and F. Rassoulzadegan. Nutrient limitations, microbial food webs, and biological c-pumps: suggested interactions in a p-limited Mediterranean. Marine Ecology Progress Series, 117:299–306, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/117/m117p299.pdf.

Thomassen:1995:GES

[TR95b] S. Thomassen and H. U. Riisgård. Growth and energetics of the sponge Halichondria panicea. Marine Ecology Progress Series, 128:239–246, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/128/m128p239.pdf.

Tokeshi:1995:FGD

[TR95c] M. Tokeshi and L. Romero. Filling a gap: dynamics of space occupancy on a mussel-dominated subtropical rocky shore. Marine Ecology Progress Series, 119:167–176, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
119/m119p167.pdf.

Townsend:1995:UOS

[TRMW95] D. W. Townsend, R. L. Radtke, D. P. Malone, and J. P. Wallinga. Use of otolith strontium: calcium ratios for hindcasting larval cod Gadus morhua distributions relative to water masses on Georges Bank. Marine Ecology Progress Series, 119:37–44, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/119/m119p037.pdf.

Trowbridge:1992:MAS

[Tro92] C. D. Trowbridge. Mesoherbivory: the ascoglossan sea slug Placida dendritica may contribute to the restricted distribution of its algal host. Marine Ecology Progress Series, 83: 207–220, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/83/m083p207.pdf.

Trowbridge:1993:IBA

[Tro93] C. D. Trowbridge. Interactions between an ascogiossan sea slug and its green algal host: branch loss and role of epiphytes. Marine Ecology Progress Series, 101:263–272, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
101/m101p263.pdf.

Trowbridge:1994:DRP

[Tro94] C. D. Trowbridge. Defensive responses and payability of specialist herbivores: predation on NE Pacific ascoglossan gastropods. Marine Ecology Progress Series, 105:61–70, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
105/m105p061.pdf.

Thebault:1994:NSM

[TRPS94] M. T. Thebault, J. P. Raffin, R. Pichon, and A. Smine. ^{31}P NMR studies of the metabolic changes in the prawns Palaemon serratus and P. elegans during exercise. Marine Ecology Progress Series, 111:73–78, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/111/m111p073.pdf.

Thouzeau:1991:SVD

[TRS91] G. Thouzeau, G. Robert, and S. J. Smith. Spatial variability in distribution and growth of juvenile and adult sea scallops Placopecten magellanicus (Gmelin) on eastern Georges Bank (Northwest Atlantic). Marine Ecology Progress Series, 74:205–218, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/74/m074p205.pdf.

Thouzeau:1991:FAB

[TRU91] G. Thouzeau, G. Robert, and R. Ugarte. Faunal assemblages of benthic mega-invertebrates inhabiting sea scallop grounds from eastern Georges Bank, in relation to environmental factors. Marine Ecology Progress Series, 74:61–82, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
74/m074p061.pdf.

Trussell:1997:PSI

[Tru97] G. C. Trussell. Phenotypic selection in an intertidal snail: effects of a catastrophic storm. Marine Ecology Progress Series, 151:73–79, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/151/m151p073.pdf.

Thingstad:1990:CPG

[TS90a] T. F. Thingstad and E. Sakshaug. Control of phytoplankton growth in nutrient recycling ecosystems. Theory and terminology. Marine Ecology Progress Series, 63:261–272, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
63/m063p261.pdf.

Tremblay:1990:DVM

[TS90b] M. J. Tremblay and M. Sinclair. Diel vertical migration of sea scallop larvae Placopecten magellanicus in a shallow embayment. Marine Ecology Progress Series, 67:19–25, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
67/m067p019.pdf.

Tremblay:1990:SSL

[TS90c] M. J. Tremblay and M. Sinclair. Sea scallop larvae Placopecten magellanicus on Georges Bank: vertical distribution in relation to water column stratification and food. Marine Ecology Progress Series, 61:1–15, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/61/m061p001.pdf.

Tornblom:1999:SDB

[TS99] E. Törnblom and M. Søndergaard. Seasonal dynamics of bacterial biomass and production on eelgrass Zostera marina leaves. Marine Ecology Progress Series, 179:231–240, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
179/m179p231.pdf.

Thingstad:1993:PCA

[TSB93] Frede Thingstad, Evy F. Skjoldal, and Rolf A. Bohne. Phosphorus cycling and algal-bacterial competition in Sandsfjord, western Norway. Marine Ecology Progress Series, 99:239– 259, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/99/m099p239.pdf.

Tsuda:1993:WNL

[TSI^{+}93] A. Tsuda, H. Sugisaki, T. Ishimaru, T. Saino, and T. Sato. White-noise-like distribution of the oceanic copepod Neocalanus cristatus in the subarctic North Pacific. Marine Ecology Progress Series, 97:39–46, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/97/m097p039.pdf.

Thorrold:1994:TPL

[TSM^{+}94] S. R. Thorrold, J. M. Shenker, R. Mojica, Jr., E. D. Maddox, and E. Wishinski. Temporal patterns in the larval supply of summer-recruiting reef fishes to Lee Stocking Island, Bahamas. Marine Ecology Progress Series, 112:75–86, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
112/m112p075.pdf.

Tranvik:1993:UUC

[TSS93] L. J. Tranvik, E. B. Sherr, and B. F. Sherr. Uptake and utilization of ’colloidal DOM’ by heterotrophic flagellates in seawater. Marine Ecology Progress Series, 92:301–309, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
92/m092p301.pdf.

Tsutsumi:1990:PPC

[Tsu90] H. Tsutsumi. Population persistence of Capitella sp (Polychaeta; Capitellidae) on a mud flat subject to environmental disturbance by organic enrichment. Marine Ecology Progress Series, 63:147–156, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/63/m063p147.pdf.

Tzeng:1997:UOM

[TSW97] W. N. Tzeng, K. P. Severin, and H. Wickström. Use of otolith microchemistry to investigate the environmental history of European eel Anguilla anguilla. Marine Ecology Progress Series, 149:73–81, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/149/m149p073.pdf.

Targett:1990:EFS

[TT90] T. E. Targett and N. M. Targett. Energetics of food selection by the herbivorous parrotfish Sparisoma radians: roles of assimilation efficiency, gut evacuation rate, and algal secondary metabolites. Marine Ecology Progress Series, 66:13–21, 1990.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
66/m066p013.pdf.

Thornton:1998:ETA

[TT98] D. C. O. Thornton and B. Thake. Effect of temperature on the aggregation of Skeletonema costatum (Bacillariophyceae) and the implication for carbon flux in coastal waters. Marine Ecology Progress Series, 174:223–231, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/174/m174p223.pdf.

Tanaka:1999:PPE

[TT99] T. Tanaka and A. Taniguchi. Predator–prey eddy in heterotrophic nanoflagellate-bacteria relationships in a bay on the northeastern Pacific coast of Japan. Marine Ecology Progress Series, 179:123–134, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/179/m179p123.pdf.

Titlyanov:1996:DZR

[TTL^{+}96] E. A. Titlyanov, T. V. Titlyanova, V. A. Leletkin, J. Tsukahara, R. van Woesik, and K. Yamazato. Degradation of zooxanthellae and regulation of their density in hermatypic corals. Marine Ecology Progress Series, 139:167–178, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
139/m139p167.pdf.

Turner:1995:SSC

[TTP^{+}95] S. J. Turner, S. F. Thrush, R. D. Pridmore, J. E. Hewitt, V. J. Cummings, and M. Maskery. Are soft-sediment communities stable? An example from a windy harbour. Marine Ecology Progress Series, 120:219–230, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/120/m120p219.pdf.

Taylor:1999:AED

[TTR99] M. W. Taylor, R. B. Taylor, and T. A. V. Rees. Allometric evidence for the dominant role of surface cells in ammonium metabolism and photosynthesis in northeastern New Zealand seaweeds. Marine Ecology Progress Series, 184:73–81, 1999.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
184/m184p073.pdf.

Takeda:1997:RBP

[TTW97] S. Takeda, S. Tamura, and M. Washio. Relationship between the pea crab Pinnixa tumida and its endobenthic holothurian host Paracaudina chilensis. Marine Ecology Progress Series, 149:143–154, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/149/m149p143.pdf.

Thornton:1999:EIE

[TUN99] D. C. O. Thornton, G. J. C. Underwood, and D. B. Nedwell. Effect of illumination and emersion period on the exchange of ammonium across the estuarine sediment–water interface. Marine Ecology Progress Series, 184:11–20, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
184/m184p011.pdf.

Tuomi:1997:BCP

[Tuo97] P. Tuomi. Bacterial carbon production in the northern Baltic: a comparison of thymidine incorporation and FDC based methods. Marine Ecology Progress Series, 153:59–66, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
153/m153p059.pdf.

Turner:1993:CNP

[Tur93] R. E. Turner. Carbon, nitrogen, and phosphorus leaching rates from Spartina alterniflora salt marshes. Marine Ecology Progress Series, 92:135–140, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/92/m092p135.pdf.

Toro:1996:MLH

[TVG96] J. E. Toro, A. M. Vergara, and R. Gallegillos. Multiplelocus heterozygosity, physiology and growth at two different stages in the life cycle of the Chilean oyster Ostrea chilensis. Marine Ecology Progress Series, 134:151–158, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
134/m134p151.pdf.

Timmermans:1998:ISP

[TvLdJ^{+}98] K. R. Timmermans, M. A. van Leeuwe, J. T. M. de Jong, R. M. L. McKay, R. F. Nolting, H. J. Witte, J. van Ooyen, M. J. W. Swagerman, H. Kloosterhuis, and H. J. W. de Baar. Iron stress in the Pacific region of the Southern Ocean: evidence from enrichment bioassays. Marine Ecology Progress Series, 166:27–41, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/166/m166p027.pdf.

Terrados:1997:LVR

[TW97] J. Terrados and S. L. Williams. Leaf versus root nitrogen uptake by the surfgrass Phyllospadix torreyi. Marine Ecology Progress Series, 149:267–277, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/149/m149p267.pdf.

Takeuchi:1998:RRS

[TW98] I. Takeuchi and K. Watanabe. Respiration rate and swimming speed of the necrophagous amphipod Eurythenes gryllus from Antarctic deep waters. Marine Ecology Progress Series, 163:285–288, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/163/m163p285.pdf.

Tarling:1995:DPM

[TWS^{+}95] G. A. Tarling, P. Ward, M. Sheader, J. A. Williams, and C. Symon. Distribution patterns of macrozooplankton assemblages in the southwest Atlantic. Marine Ecology Progress Series, 120:29–40, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/120/m120p029.pdf.

Thistle:1995:SWS

[TWWE95] D. Thistle, G. L. Weatherly, A. Wonnacott, and S. C. Ertman. Suspension by winter storms has an energetic cost for adult male benthic harpacticoid copepods at a shelf site. Marine Ecology Progress Series, 125:77–86, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/125/m125p077.pdf.

Tyler:1999:PDC

[TZ99] A. C. Tyler and J. C. Zieman. Patterns of development in the creekbank region of a barrier island Spartina alterniflora marsh. Marine Ecology Progress Series, 180:161–177, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
180/m180p161.pdf.

Urban:1994:PDB

[UC94] H.-J. Urban and B. Campos. Population dynamics of the bivalves Gari solida, Semele solida and Protothaca thaca from a small bay in Chile at 36^{o}S. Marine Ecology Progress Series, 115:93–102, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/115/m115p093.pdf.

Uriarte:1998:EPA

[UCV98] I. Uriarte, U. Cotano, and F. Villate. Egg production of Acartia bifilosa in the small temperate estuary of Mundaka, Spain, in relation to environmental variables and population development. Marine Ecology Progress Series, 166:197–205, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
166/m166p197.pdf.

Udy:1999:RSN

[UDLM99] J. W. Udy, W. C. Dennison, W. J. Lee Long, and L. J. McKenzie. Responses of seagrass to nutrients in the Great Barrier Reef, Australia. Marine Ecology Progress Series, 185: 257–271, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/185/m185p257.pdf.

Ucko:1994:PDR

[UGSBA94] M. Ucko, S. Geresh, B. Simon-Berkovitch, and S. Arad Malis. Predation by a dinoflagellate on a red microalga with a cell wall modified by sulfate and nitrate starvation. Marine Ecology Progress Series, 104:293–298, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/104/m104p293.pdf.

Uitto:1997:SDC

[UHL^{+}97] A. Uitto, A. S. Heiskanen, R. Lignell, R. Autio, and R. Pajuniemi. Summer dynamics of the coastal planktonic food web in the northern Baltic Sea. Marine Ecology Progress Series, 151:27–41, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/151/m151p027.pdf.

Uiblein:1995:MVB

[Uib95] F. Uiblein. Morphological variability between populations of Neobythites stefanovi (Pisces: Ophidiidae) from the deep Red Sea and the Gulf of Aden. Marine Ecology Progress Series, 124:23–29, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/124/m124p023.pdf.

Uitto:1996:SHC

[Uit96] A. Uitto. Summertime herbivory of coastal mesozooplankton and metazoan microplankton in the northern Baltic. Marine Ecology Progress Series, 132:47–56, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/132/m132p047.pdf.

Uthicke:1998:MCP

[UK98] S. Uthicke and D. W. Klumpp. Microphytobenthos community production at a near-shore coral reef: seasonal variation and response to ammonium recycled by holothurians. Marine Ecology Progress Series, 169:1–11, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/169/m169p001.pdf.

Ugarkovic:1991:IAF

[UKR^{+}91] D. Ugarkovic, B. Kurelec, A. Robitzki, W. E. G. Muiler, and H. C. Schroder. Inhibition of aggregation-factor-induced ras gene expression in the sponge Geodia cydonium by detergentpolluted seawater a sensitive biological assay for low-level detergent pollution. Marine Ecology Progress Series, 71:253– 258, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/71/m071p253.pdf.

Urban:1992:SDC

[UMD92] J. L. Urban, C. H. McKenzie, and D. Deibel. Seasonal differences in the content of Oikopleura vanhoeffeni and Calanus finmarchicus faecal pellets: illustrations of zooplankton food web shifts in coastal Newfoundland waters. Marine Ecology Progress Series, 84:255–264, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/84/m084p255.pdf.

Uriz:1991:AES

[UMT^{+}91] M. J. Uriz, D. Martin, X. Turon, E. Ballesteros, R. Hughes, and C. Acebal. An approach to the ecological significance of chemically mediated bio-activity in Mediterranean benthic communities. Marine Ecology Progress Series, 70:175– 188, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/70/m070p175.pdf.

Uriz:1998:HDR

[UMTM98] M. J. Uriz, M. Maldonado, X. Turon, and R. Martí. How do reproductive output, larval behaviour, and recruitment contribute to adult spatial patterns in Mediterranean encrusting sponges? Marine Ecology Progress Series, 167:137–148, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
167/m167p137.pdf.

Upton:1993:SBM

[UNPH93] A. C. Upton, D. B. Nedwell, R. J. Parkes, and S. M. Harvey. Seasonal benthic microbial activity in the southern North Sea; oxygen uptake and sulphate reduction. Marine Ecology Progress Series, 101:273–281, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/101/m101p273.pdf.

Urban:1994:UTT

[Urb94] H.-J. Urban. Upper temperature tolerance of ten bivalve species off Peru and Chile related to El Niño. Marine Ecology Progress Series, 107:139–145, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/107/m107p139.pdf.

Urban-Rich:1998:ACF

[URHR98] J. Urban-Rich, D. A. Hansell, and M. R. Roman. Analysis of copepod fecal pellet carbon using a high temperature combustion method. Marine Ecology Progress Series, 171: 199–208, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/171/m171p199.pdf.

Uye:1995:SRB

[US95] S. Uye and K. Sano. Seasonal reproductive biology of the small cyclopoid copepod Oithona davisae in a temperate eutrophic inlet. Marine Ecology Progress Series, 118:121– 128, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/118/m118p121.pdf.

Uye:1998:SVB

[US98] S. Uye and K. Sano. Seasonal variations in biomass, growth rate and production rate of the small cyclopoid copepod Oithona davisae in a temperate eutrophic inlet. Marine Ecology Progress Series, 163:37–44, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/163/m163p037.pdf.

Uye:1990:FIB

[UT90] S. Uye and K. Takamatsu. Feeding interactions between planktonic copepods and red-tide flagellates from Japanese coastal waters. Marine Ecology Progress Series, 59:97–107, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/59/m059p097.pdf.

Uriz:1995:PRA

[UTB^{+}95] M. J. Uriz, X. Turon, M. A. Becerro, J. Galera, and J. Lozano. Patterns of resource allocation to somatic, defensive, and reproductive functions in the Mediterranean encrusting sponge Crambe crambe (Demospongiae, Poecilosclerida). Marine Ecology Progress Series, 124:159–170, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
124/m124p159.pdf.

Uye:1996:IRF

[Uye96] S. Uye. Induction of reproductive failure in the planktonic copepod Calanus pacificus by diatoms. Marine Ecology Progress Series, 133:89–97, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/133/m133p089.pdf.

Uchida:1995:RTD

[UYMH95] T. Uchida, M. Yamaguchi, Y. Matsuyama, and T. Honjo. The red-tide dinoflagellate Heterocapsa sp. kills Gyrodinium instriatum by cell contact. Marine Ecology Progress Series, 118:301–303, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/118/m118p301.pdf.

Viejo:1999:IEA

[VÅCL99] R. M. Viejo, P. Åberg, G. Cervin, and M. Lindegarth. The interactive effects of adult canopy, germling density and grazing on germling survival of the rockweed Ascophyllum nodosum. Marine Ecology Progress Series, 187:113–120, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
187/m187p113.pdf.

Vaque:1994:MHW

[VAD^{+}94] D. Vaqué, S. Agusti, C. M. Duarte, S. Enriquez, and O. Geertz-Hansen. Microbial heterotrophs within Codium bursa: a naturally isolated microbial food web. Marine Ecology Progress Series, 109:275–282, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/109/m109p275.pdf.

Vermaat:1995:MMG

[VAD^{+}95] J. E. Vermaat, N. S. R. Agawin, C. M. Duarte, M. D. Fortes, N. Marbà, and J. S. Uri. Meadow maintenance, growth and productivity of a mixed Philippine seagrass bed. Marine Ecology Progress Series, 124:215–225, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/124/m124p215.pdf.

Vainola:1990:ADB

[Väi90] R. Väinölä. Allozyme differentiation between Baltic and North Sea Mytilus populations: a reassessment of evidence from transplantations. Marine Ecology Progress Series, 67: 305–308, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/67/m067p305.pdf.

ValdesSzeinfeld:1991:CIP

[Val91] E. Valdes Szeinfeld. Cannibalism and intraguild predation in clupeoids. Marine Ecology Progress Series, 79:17–26, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
79/m079p017.pdf.

vanBoekel:1992:PCM

[van92a] W. H. M. van Boekel. Phaeocystis colony mucus components and the importance of calcium ions for colony stability. Marine Ecology Progress Series, 87:301–305, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
87/m087p301.pdf.

Vance:1992:APJ

[Van92b] D. J. Vance. Activity patterns of juvenile penaeid prawns in response to artificial tidal and day-night cycles: a comparison of three species. Marine Ecology Progress Series, 87: 215–226, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/87/m087p215.pdf.

vanderLingen:1994:EPS

[van94a] C. D. van der Lingen. Effect of particle size and concentration on the feeding behaviour of adult pilchard Sardinops sagax. Marine Ecology Progress Series, 109:1–13, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
109/m109p001.pdf.

VanVeghel:1994:RCPa

[Van94b] M. L. J. Van Veghel. Reproductive characteristics of the polymorphic Caribbean reef building coral Montastrea annularis. I. Gametogenesis and spawning behavior. Marine Ecology Progress Series, 109:209–219, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/109/m109p209.pdf.

Vedel:1994:FIP

[VAR94] A. Vedel, B. B. Andersen, and H. U. Riisgård. Field investigations of pumping activity of the facultatively filterfeeding polychaete Nereis diversicolor using an improved infrared phototransducer system. Marine Ecology Progress Series, 103:91–101, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/103/m103p091.pdf.

Vanhove:1999:CSN

[VAV99] S. Vanhove, W. Arntz, and M. Vincx. Comparative study of the nematode communities on the southeastern Weddell Sea shelf and slope (Antarctica). Marine Ecology Progress Series, 181:237–256, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/181/m181p237.pdf.

VanWambeke:1990:EON

[VB90] F. Van Wambeke and M. A. Bianchi. Effect of organic nitrogen enrichments on marine planktonic networks and heterotrophic bacterial potential. Marine Ecology Progress Series, 63:17–25, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/63/m063p017.pdf.

VanVeghel:1993:IVD

[VB93] M. L. J. Van Veghel and R. P. M. Bak. Intraspecific variation of a dominant Caribbean reef building coral, Montastrea annularis: genetic, behavioral and morphometric aspects. Marine Ecology Progress Series, 92:255–265, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
92/m092p255.pdf.

VanVeghel:1994:RCPc

[VB94] M. L. J. Van Veghel and R. P. M. Bak. Reproductive characteristics of the polymorphic Caribbean reef building coral Montastrea annularis. III. Reproduction in damaged and regenerating colonies. Marine Ecology Progress Series, 109: 229–233, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/109/m109p229.pdf.

Vercauteren:1996:BDZ

[VB96] K. Vercauteren and R. Blust. Bioavailability of dissolved zinc to the common mussel Mytilus edulis in complexing environments. Marine Ecology Progress Series, 137:123–132, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
137/m137p123.pdf.

Verdier-Bonnet:1997:MLD

[VBCRB97] C. Verdier-Bonnet, F. Carlotti, C. Rey, and M. Bhaud. A model of larval dispersion coupling wind-driven currents and vertical larval behaviour: application to the recruitment of the annelid Owenia fusiformis in Banyuls Bay, France. Marine Ecology Progress Series, 160:217–231, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
160/m160p217.pdf.

vanderVeer:1991:PDI

[vBDW91] H. W. van der Veer, M. J. N. Bergman, R. Dapper, and J. IJ. Witte. Population dynamics of an intertidal 0-group flounder Platichthys flesus population in the western Dutch Wadden Sea. Marine Ecology Progress Series, 73:141–148, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
73/m073p141.pdf.

Vincent:1994:PIL

[VBF94] W. F. Vincent, N. Bertrand, and J.-J. Frenette. Photoadaptation to intermittent light across the St. Lawrence Estuary freshwater-saltwater transition zone. Marine Ecology Progress Series, 110:283–292, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/110/m110p283.pdf.

Vethaak:1992:FDM

[VBL^{+}92] A. D. Vethaak, D. Bucke, T. Lang, P. W. Wester, J. Jol, and M. Carr. Fish disease monitoring along a pollution transect: a case study using dab Limanda limanda in the German Bight. Marine Ecology Progress Series, 91:173–192, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
91/m091p173.pdf.

Vergara:1995:NAF

[VBN95] J. J. Vergara, K. T. Bird, and F. X. Niell. Nitrogen assimilation following NH_{4}^{+} pulses in the red alga Gracilariopsis lemaneiformis: effect on C metabolism. Marine Ecology Progress Series, 122:253–263, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/122/m122p253.pdf.

vanBoekel:1990:RAPa

[vBV90a] W. H. M. van Boekel and M. J. W. Veldhuis. Regulation of alkaline phosphatase synthesis in Phaeocystis sp. Marine Ecology Progress Series, 61:281–289, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/61/m061p281.pdf.

vanBoekel:1990:RAPb

[vBV90b] W. H. M. van Boekel and M. J. W. Veldhuis. Regulation of alkaline phosphatase synthesis in Phaeocystis sp. Marine Ecology Progress Series, 61:281–289, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/61/m061p281.pdf.

vanBleijswijk:1995:LSG

[vBV95] J. D. L. van Bleijswijk and M. J. W. Veldhuis. In situ gross growth rates of Emiliania huxleyi in enclosures with different phosphate loadings revealed by diel changes in DNA content. Marine Ecology Progress Series, 121:271–277, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
121/m121p271.pdf.

Vukmirovic:1994:DDM

[VBZ^{+}94] M. Vukmirovic, N. Bihari, R. K. Zahn, W. E. G. Müller, and R. Batel. DNA damage in marine mussel Mytilus galloprovincialis as a biomarker of environmental contamination. Marine Ecology Progress Series, 109:165–171, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
109/m109p165.pdf.

Viarengo:1993:CZC

[VCMP93] A. Viarengo, L. Canesi, A. Mazzucotelli, and E. Ponzano. Cu, Zn and Cd content in different tissues of the Antarctic scallop Adamussium colbecki: role of metallothionein in heavy metal homeostasis and detoxication. Marine Ecology Progress Series, 95:163–168, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/95/m095p163.pdf.

vandenHurk:1992:CNA

[vCRS92] P. van den Hurk, P. M. Chapman, B. Roddie, and R. C. Swartz. A comparison of North American and West European infaunal amphipod species in a toxicity test on North Sea sediments. Marine Ecology Progress Series, 91:237–243, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
91/m091p237.pdf.

Vetter:1994:EEA

[VD94] Y.-A. Vetter and J. W. Deming. Extracellular enzyme activity in the Arctic northeast water polynya. Marine Ecology Progress Series, 114:23–34, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/114/m114p023.pdf.

Vidondo:1995:SGC

[VD95] B. Vidondo and C. M. Duarte. Seasonal growth of Codium bursa, a slow-growing Mediterranean macroalga: in situ experimental evidence of nutrient limitation. Marine Ecology Progress Series, 123:185–191, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/123/m123p185.pdf.

Vetter:1999:OEM

[VD99] E. W. Vetter and P. K. Dayton. Organic enrichment by macrophyte detritus, and abundance patterns of megafaunal populations in submarine canyons. Marine Ecology Progress Series, 186:137–148, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/186/m186p137.pdf.

Vopel:1996:VDC

[VDA96] K. Vopel, J. Dehmlow, and G. Arlt. Vertical distribution of Cletocamptus confluens (Copepoda, Harpacticoida) in relation to oxygen and sulphide microprofiles of a brackish water sulphuretum. Marine Ecology Progress Series, 141:129– 137, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/141/m141p129.pdf.

Vincent:1996:PBP

[VDBF96] W. F. Vincent, J. J. Dodson, N. Bertrand, and J. J. Frenette. Photosynthetic and bacterial production gradients in a larval fish nursery: the St. Lawrence River transition zone. Marine Ecology Progress Series, 139:227–238, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
139/m139p227.pdf.

vandenBerg:1996:MSA

[vdBTvDR96] A. J. van den Berg, S. M. Turner, F. C. van Duyl, and P. Ruardij. Model structure and analysis of dimethylsulphide (DMS) production in the southern North Sea, considering phytoplankton dimethylsulphoniopropionate- (DMSP) lyase and eutrophication effects. Marine Ecology Progress Series, 145:233–244, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/145/m145p233.pdf.

vanDuyl:1999:TCB

[vDdWKW99] F. C. van Duyl, B. de Winder, A. J. Kop, and U. Wollenzien. Tidal coupling between carbohydrate concentrations and bacterial activities in diatom-inhabited intertidal mudflats. Marine Ecology Progress Series, 191:19–32, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
191/m191p019.pdf.

VanWoesik:1995:ECJ

[VDG95] R. Van Woesik, L. M. De Vantier, and J. S. Glazebrook. Effects of Cyclone ‘Joy’ on nearshore coral communities of the Great Barrier Reef. Marine Ecology Progress Series, 128: 261–270, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/128/m128p261.pdf.

Vopel:1998:EAS

[VDJA98] K. Vopel, J. Dehmlow, M. Johansson, and G. Arlt. Effects of anoxia and sulphide on populations of Cletocamptus confluens (Copepoda, Harpacticoida). Marine Ecology Progress Series, 175:121–128, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/175/m175p121.pdf.

vanDuyl:1990:SPB

[vDK90] F. C. van Duyl and A. J. Kop. Seasonal patterns of bacterial production and biomass in intertidal sediments of the western Dutch Wadden Sea. Marine Ecology Progress Series, 59:249–261, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/59/m059p249.pdf.

vanderLingen:1995:RRA

[vdL95] C. D. van der Lingen. Respiration rate of adult pilchard Sardinops sagax in relation to temperature, voluntary swimming speed and feeding behaviour. Marine Ecology Progress Series, 129:41–54, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/129/m129p041.pdf.

vanderLingen:1998:NEA

[vdL98] C. D. van der Lingen. Nitrogen excretion and absorption efficiencies of sardine Sardinops sagax fed phytoplankton and zooplankton diets. Marine Ecology Progress Series, 175: 67–76, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/175/m175p067.pdf.

Vaque:1990:IAP

[VDM90] D. Vaqué, C. M. Duarte, and C. Marrasé. Influence of algal population dynamics on phytoplankton colonization by bacteria: evidence from two diatom species. Marine Ecology Progress Series, 65:201–203, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/65/m065p201.pdf.

Vidondo:1997:DLM

[VDM^{+}97] B. Vidondo, C. M. Duarte, A. L. Middelboe, K. Stefansen, T. Lützen, and S. L. Nielsen. Dynamics of a landscape mosaic: size and age distributions, growth and demography of seagrass Cymodocea nodosa patches. Marine Ecology Progress Series, 158:131–138, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/158/m158p131.pdf.

Visscher:1992:EBW

[VDT92] P. T. Visscher, M. R. Diaz, and B. F. Taylor. Enumeration of bacteria which cleave or demethylate dimethylsulfoniopropionate in the Caribbean Sea. Marine Ecology Progress Series, 89:293–296, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/89/m089p293.pdf.

VandenThillart:1994:ILT

[VDVC94] G. Van den Thillart, J. Dalla Via, G. Vitali, and P. Cortesi. Influence of long-term hypoxia exposure on the energy metabolism of Solea solea. I. Critical 02 levels for aerobic and anaerobic metabolism. Marine Ecology Progress Series, 104:109–117, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/104/m104p109.pdf.

vanDuyl:1993:BBP

[vDvRK93] Fleur C. van Duyl, Wim van Raaphorst, and Arjen J. Kop. Benthic bacterial production and nutrient sediment-water exchange in sandy North Sea sediments. Marine Ecology Progress Series, 100:85–95, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/100/m100p085.pdf.

vanderVeer:1999:YCS

[vdVW99] H. W. van der Veer and J. I. J. Witte. Year-class strength of plaice Pleuronectes platessa in the Southern Bight of the North Sea: a validation and analysis of the inverse relationship with winter seawater temperature. Marine Ecology Progress Series, 184:245–257, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/184/m184p245.pdf.

Vedel:1998:PDB

[Ved98] A. Vedel. Phytoplankton depletion in the benthic boundary layer caused by suspension-feeding Nereis diversicolor (Polychaeta): grazing impact and effect of temperature. Marine Ecology Progress Series, 163:125–132, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
163/m163p125.pdf.

Venrick:1998:SCC

[Ven98] E. L. Venrick. Spring in the California Current: the distribution of phytoplankton species, April 1993 and April 1995. Marine Ecology Progress Series, 167:73–88, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
167/m167p073.pdf.

Vethaak:1992:GPH

[Vet92] A. D. Vethaak. Gross pathology and histopathology in fish: summary. Marine Ecology Progress Series, 91:171–172, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
91/m091p171.pdf.

Vetter:1995:DBP

[Vet95] E. W. Vetter. Detritus-based patches of high secondary production in the nearshore benthos. Marine Ecology Progress Series, 120:251–262, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/120/m120p251.pdf.

Vetter:1996:EEI

[Vet96a] E. W. Vetter. Enrichment experiments and infaunal population cycles on a Southern California sand plain: response of the leptostracan Nebalia daytoni and other infauna. Marine Ecology Progress Series, 137:83–93, 1996. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/137/m137p083.pdf.

Vetter:1996:SPS

[Vet96b] E. W. Vetter. Secondary production of a Southern California Nebalia (Crustacea: Leptostraca). Marine Ecology Progress Series, 137:95–101, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/137/m137p095.pdf.

VanAlstyne:1999:FPJ

[VEW99] K. L. Van Alstyne, J. M. Ehlig, and S. L. Whitman. Feeding preferences for juvenile and adult algae depend on algal stage and herbivore species. Marine Ecology Progress Series, 180:179–185, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/180/m180p179.pdf.

Velando:1999:ISD

[VF99] A. Velando and J. Freire. Intercolony and seasonal differences in the breeding diet of European shags on the Galician coast (NW Spain). Marine Ecology Progress Series, 188:225–236, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/188/m188p225.pdf.

Vacelet:1996:SBM

[VFMFBE96] J. Vacelet, A. Fiala-Médioni, C. R. Fisher, and N. Boury-Esnault. Symbiosis between methane-oxidizing bacteria and a deep-sea carnivorous cladorhizid sponge. Marine Ecology Progress Series, 145:77–85, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/145/m145p077.pdf.

vanErkomSchurink:1991:CRC

[vG91] C. van Erkom Schurink and C. L. Griffiths. A comparison of reproductive cycles and reproductive output in four southern African mussel species. Marine Ecology Progress Series, 76:123–134, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/76/m076p123.pdf.

Vanzella:1990:RNM

[VGJ90] Alessandra Vanzella, Maria A. Guerrero, and Ronald D. Jones. Recovery of nitrification in marine bacteria following exposure to carbon monoxide or light. Marine Ecology Progress Series, 60:91–95, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/60/m060p091.pdf.

Vaques:1994:GRB

[VGM94] D. Vaqués, J. M. Gasol, and C. Marrasé. Grazing rates on bacteria: the significance of methodology and ecological factors. Marine Ecology Progress Series, 109:263–274, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
109/m109p263.pdf.

Valentine:1993:MSM

[VH93] John R. Valentine and K. L. Heck, Jr. Musseis in seagrass meadows: their influence on macroinvertebrate abundance and secondary production in the northern Gulf of Mexico. Marine Ecology Progress Series, 96:63–74, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
96/m096p063.pdf.

Vismann:1996:RHS

[VH96] Bent Vismann and L. Hagerman. Recovery from hypoxia with and without sulfide in Saduria entomon: oxygen debt, reduced sulfur and anaerobic metabolites. Marine Ecology Progress Series, 143:131–139, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/143/m143p131.pdf.

Valentine:1999:SHE

[VH99a] J. F. Valentine and K. L. Heck, Jr. Seagrass herbivory: evidence for the continued grazing of marine grasses. Marine Ecology Progress Series, 176:291–302, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
176/m176p291.pdf.

Varela:1999:EAN

[VH99b] D. E. Varela and P. J. Harrison. Effect of ammonium on nitrate utilization by Emiliania huxleyi, a coccolithophore from the oceanic northeastern Pacific. Marine Ecology Progress Series, 186:67–74, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/186/m186p067.pdf.

Volkel:1995:SST

[VHG95] S. Völkel, K. Hauschild, and M. K. Grieshaber. Sulfide stress and tolerance in the lugworm Arenicola marina during low tide. Marine Ecology Progress Series, 122:205–215, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
122/m122p205.pdf.

Vance:1996:HFD

[VHH^{+}96] D. J. Vance, M. D. E. Haywood, D. S. Heales, R. A. Kenyon, N. R. Loneragan, and R. C. Pendrey. How far do prawns and fish move into mangroves? Distribution of juvenile banana prawns Penaeus merguiensis and fish in a tropical mangrove forest in northern Australia. Marine Ecology Progress Series, 131:115–124, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/131/m131p115.pdf.

Vance:1998:SAV

[VHH^{+}98] D. J. Vance, M. D. E. Haywood, D. S. Heales, R. A. Kenyon, and N. R. Loneragan. Seasonal and annual variation in abundance of postlarval and juvenile banana prawns Penaeus merguiensis and environmental variation in two estuaries in tropical northeastern Australia: a six year study. Marine Ecology Progress Series, 163:21–36, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/163/m163p021.pdf.

Vance:1996:SAV

[VHHS96] D. J. Vance, M. D. E. Haywood, D. S. Heales, and D. J. Staples. Seasonal and annual variation in abundance of postlarval and juvenile grooved tiger prawns Penaeus semisulcatus and environmental variation in the Embley River, Australia: a six year study. Marine Ecology Progress Series, 135:43–55, 1996.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
135/m135p043.pdf.

Vigliola:1998:STP

[VHVB^{+}98] L. Vigliola, M. L. Harmelin-Vivien, F. Biagi, R. Galzin, A. Garcia-Rubies, J. G. Harmelin, J. Y. Jouvenel, L. Le Direach-Boursier, E. Macpherson, and L. Tunesi. Spatial and temporal patterns of settlement among sparid fishes of the genus Diplodus in the northwestern Mediterranean. Marine Ecology Progress Series, 168:45–56, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/168/m168p045.pdf.

Vidal:1994:PDT

[Vid94] M. Vidal. Phosphate dynamics tied to sediment disturbances in Alfacs Bay (NW Mediterranean). Marine Ecology Progress Series, 110:211–221, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/110/m110p211.pdf.

Villareal:1991:NFC

[Vil91] T. A. Villareal. Nitrogen-fixation by the cyanobacterial symbiont of the diatom genus Hemiaulus. Marine Ecology Progress Series, 76:201–204, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/76/m076p201.pdf.

Vismann:1990:SDT

[Vis90] Bent Vismann. Sulfide detoxification and tolerance in Nereis (Hediste) diversicolor and Nereis (Neanthes) virens (Annelida. Polychaeta). Marine Ecology Progress Series, 59:229– 238, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/59/m059p229.pdf.

Vismann:1991:PSD

[Vis91] Bent Vismann. Physiology of sulfide detoxification in the isopod Saduria (Mesidotea) entomon. Marine Ecology Progress Series, 76:283–293, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/76/m076p283.pdf.

Vismann:1993:HSR

[Vis93] Bent Vismann. Hematin and sulfide removal in nemolymph of the hemoglobin-containing bivalve Scapharca inaequivalvis. Marine Ecology Progress Series, 98:115–122, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
98/m098p115.pdf.

Visser:1997:URW

[Vis97] A. W. Visser. Using random walk models to simulate the vertical distribution of particles in a turbulent water column. Marine Ecology Progress Series, 158:275–281, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
158/m158p275.pdf.

Veldhuis:1990:VDP

[VK90] M. J. W. Veldhuis and G. W. Kraay. Vertical distribution and pigment composition of a picoplanktonic prochlorophyte in the subtropical North Atlantic: a combined study of HPLCanalysis of pigments and flow cytometry. Marine Ecology Progress Series, 68:121–127, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/68/m068p121.pdf.

VanVeghel:1994:RCPb

[VK94] M. L. J. Van Veghel and M. E. H. Kahmann. Reproductive characteristics of the polymorphic Caribbean reef building coral Montastrea annularis. II. Fecundity and colony structure. Marine Ecology Progress Series, 109:221–227, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
109/m109p221.pdf.

Viitasalo:1998:PVP

[VKF^{+}98] M. Viitasalo, T. Kiørboe, J. Flinkman, L. W. Pedersen, and A. W. Visser. Predation vulnerability of planktonic copepods: consequences of predator foraging strategies and prey sensory abilities. Marine Ecology Progress Series, 175:129–142, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
175/m175p129.pdf.

vanKatwijk:1999:ESN

[vKSGvA99] M. M. van Katwijk, G. H. W. Schmitz, A. P. Gasseling, and P. H. van Avesaath. Effects of salinity and nutrient load and their interaction on Zostera marina. Marine Ecology Progress Series, 190:155–165, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/190/m190p155.pdf.

VanderWal:1995:PDF

[VKV95] P. Van der Wal, R. S. Kempers, and M. J. W. Veldhuis. Production and downward flux of organic matter and calcite in a North Sea bloom of the coccolithophore Emiliania huxleyi. Marine Ecology Progress Series, 126:247–265, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
126/m126p247.pdf.

vanKatwijk:1997:ATE

[vKVSR97] M. M. van Katwijk, L. H. T. Vergeer, G. H. W. Schmitz, and J. G. M. Roelofs. Ammonium toxicity in eelgrass Zostera marina. Marine Ecology Progress Series, 157:159–173, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
157/m157p159.pdf.

Vetter:1997:BDE

[VL97] R. D. Vetter and E. A. Lynn. Bathymetric demography, enzyme activity patterns, and bioenergetics of deep-living scorpaenid fishes (genera Sebastes and Sebastolobus): paradigms revisited. Marine Ecology Progress Series, 155:173–188, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
155/m155p173.pdf.

VanHeukelem:1994:ISP

[VLKC94] L. Van Heukelem, A. J. Lewitus, T. M. Kana, and N. E. Craft. Improved separations of phytoplankton pigments using temperature-controlled high performance liquid chromatography. Marine Ecology Progress Series, 114:303–313, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
114/m114p303.pdf.

vanLeeuwe:1998:EIS

[vLTW^{+}98] M. A. van Leeuwe, K. R. Timmermans, H. J. Witte, G. W. Kraay, M. J. W. Veldhuis, and H. J. W. de Baar. Effects of iron stress on chromatic adaptation by natural phytoplankton communities in the Southern Ocean. Marine Ecology Progress Series, 166:43–52, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/166/m166p043.pdf.

Vidal:1995:STP

[VM95] M. Vidal and J.-A. Morgui. Short-term pore water ammonium variability coupled to benthic boundary layer dynamics in Alfacs Bay, Spain (Ebro Delta, NW Mediterranean). Marine Ecology Progress Series, 118:229–236, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
118/m118p229.pdf.

Verde:1998:PRP

[VM98a] E. A. Verde and L. R. McCloskey. Production, respiration, and photophysiology of the mangrove jellyfish Cassiopea xamachana symbiotic with zooxanthellae: effect of jellyfish size and season. Marine Ecology Progress Series, 168: 147–162, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/168/m168p147.pdf.

Visser:1998:TIC

[VM98b] A. W. Visser and B. R. MacKenzie. Turbulence-induced contact rates of plankton: the question of scale. Marine Ecology Progress Series, 166:307–310, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/166/m166p307.pdf.

Vernet:1990:ASS

[VMHH90] M. Vernet, B. G. Mitchell, and O. Holm-Hansen. Adaptation of Synechococcus in situ determined by variability in intracellular phycoerythrin-543 at a coastal station off the Southern California coast, USA. Marine Ecology Progress Series, 63:9–16, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/63/m063p009.pdf.

Vernet:1996:ENP

[VMS^{+}96] M. Vernet, B. G. Mitchell, E. Sakshaug, G. Johnsen, R. Iturriaga, and P. Wassmann. Evidence for a novel pigment with in vivo absorption maximum at 708 nm associated with phaeocystis cf. pouchetii blooms. Marine Ecology Progress Series, 133:253–262, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/133/m133p253.pdf.

Vergar:1997:DMT

[VNB97] J. J. Vergar, F. X. Niell, and K. T. Bird. A dynamic model of transient NH_{4}^{+} assimilation in red algae. Marine Ecology Progress Series, 148:295–307, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/148/m148p295.pdf.

Voss:1997:NSI

[VNM97] M. Voß, G. Nausch, and J. P. Montoya. Nitrogen stable isotope dynamics in the central Baltic Sea: influence of deep-water renewal on the N-cycle changes. Marine Ecology Progress Series, 158:11–21, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/158/m158p011.pdf.

Voss:1991:CCF

[Vos91] M. Voss. Content of copepod faecal pellets in relation to food supply in Kiel Bight and its effect on sedimentation rate. Marine Ecology Progress Series, 75:217–225, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
75/m075p217.pdf.

vanderVeer:1990:RMN

[vPB90] H. W. van der Veer, L. Pihl, and M. J. N. Bergman. Recruitment mechanisms in North Sea plaice Pleuronectes platessa. Marine Ecology Progress Series, 64:1–12, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
64/m064p001.pdf.

Vervoort:1998:CDC

[VPF98] H. C. Vervoort, J. R. Pawlik, and W. Fenical. Chemical defense of the Caribbean ascidian Didemnum conchyliatum.

Marine Ecology Progress Series, 164:221–228, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
164/m164p221.pdf.

Vaque:1992:FBP

[VPFL92] D. Vaqué, M. L. Pace, S. Findlay, and D. Lints. Fate of bacterial production in a heterotrophic ecosystem: grazing by protists and metazoans in the Hudson Estuary. Marine Ecology Progress Series, 89:155–163, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/89/m089p155.pdf.

Vrieling:1994:DID

[VPGV94] E. G. Vrieling, L. Peperzak, W. W. C. Gieskes, and M. Veenhuis. Detection of the ichthyotoxic dinoflagellate Gyrodinium (cf.) aureolum and morphologically related Gymnodinium species using monoclonal antibodies: a specific immunological tool. Marine Ecology Progress Series, 103:165–174, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
103/m103p165.pdf.

Veit:1996:OWL

[VPM96] R. R. Veit, P. Pyle, and J. A. McGowan. Ocean warming and long-term change in pelagic bird abundance within the California Current System. Marine Ecology Progress Series, 139:11–18, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/139/m139p011.pdf.

vanLooij:1993:MBP

[vR93a] A. van Looij and B. Riemann. Measurements of bacterial production in coastal marine environments using leucine: application of a kinetic approach to correct for isotope dilution. Marine Ecology Progress Series, 102:97–104, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
102/m102p097.pdf.

Vedel:1993:FFP

[VR93b] A. Vedel and H. U. Riisgård. Filter-feeding in the polychaete Nereis diversicolor: growth and bioenergetics. Marine Ecology Progress Series, 100:145–152, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/100/m100p145.pdf.

Viitasalo:1998:ZPF

[VR98] M. Viitasalo and M. Rautio. Zooplanktivory by Praunus flexuosus (Crustacea: Mysidacea): functional responses and prey selection in relation to prey escape responses. Marine Ecology Progress Series, 174:77–87, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/174/m174p077.pdf.

vanRooij:1995:PGH

[vRBVB95] J. M. van Rooij, J. H. Bruggemann, J. J. Videler, and A. M. Breeman. Plastic growth of the herbivorous reef fish Sparisoma viride: field evidence for a trade-off between growth and reproduction. Marine Ecology Progress Series, 122:93– 105, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/122/m122p093.pdf.

vanRooij:1996:EOU

[vRV96] J. M. van Rooij and J. J. Videler. Estimating oxygen uptake rate from ventilation frequency in the reef fish Sparisoma viride. Marine Ecology Progress Series, 132:31–41, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
132/m132p031.pdf.

Vogt:1991:CDF

[VS91] H. Vogt and W. Schramm. Conspicuous decline of Fucus in Kiel Bay (Western Baltic): what are the causes? Marine Ecology Progress Series, 69:189–194, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
69/m069p189.pdf.

Vance:1992:CST

[VS92] D. J. Vance and D. J. Staples. Catchability and sampling of three species of juvenile penaeid prawns in the Embley River, Gulf of Carpentaria, Australia. Marine Ecology Progress Series, 87:201–213, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/87/m087p201.pdf.

Verity:1996:OLC

[VS96] P. G. Verity and V. Smetacek. Organism life cycles, predation, and the structure of marine pelagic ecosystems. Marine Ecology Progress Series, 130:277–293, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/130/m130p277.pdf.

Vergara:1998:PUR

[VSPLH98] J. J. Vergara, M. Sebastián, J. L. Pérez-Lloréns, and I. Hernández. Photoacclimation of Ulva rigida and U. rotundata (Chlorophyta) arranged in canopies. Marine Ecology Progress Series, 165:283–292, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/165/m165p283.pdf.

Verlencar:1990:RNB

[VSQ90] X. N. Verlencar, K. Somasunder, and S. Z. Qasim. Regeneration of nutrients and biological productivity in Antarctic waters. Marine Ecology Progress Series, 61:41–59, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
61/m061p041.pdf.

Vahatalo:1998:ISR

[VSSM98] A. Vähätalo, M. Søndergaard, L. Schlüter, and S. Markager. Impact of solar radiation on the decomposition of detrital leaves of eelgrass Zostera marina. Marine Ecology Progress Series, 170:107–117, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/170/m170p107.pdf.

Vrolijk:1992:BEC

[VT92] N. H. Vrolijk and N. M. Targett. Biotransformation enzymes in Cyphoma gibbosum (Gastropoda: Ovulidae) — implications for detoxification of gorgonian allelochemicals. Marine Ecology Progress Series, 88:237–246, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
88/m088p237.pdf.

VanDover:1999:RBF

[VTGK99] C. L. Van Dover, J. Trask, J. Gross, and A. Knowlton. Reproductive biology of free-living and commensal polynoid polychaetes at the lucky strike hydrothermal vent field (Mid-Atlantic Ridge). Marine Ecology Progress Series, 181:201– 214, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/181/m181p201.pdf.

Varo:1991:ETO

[VTNA91] I. Varó, A. C. Taylor, J. C. Navarro, and F. Amat. Effects of temperature and oxygen tension on oxygen consumption rates of nauplii of different Artemia strains. Marine Ecology Progress Series, 76:25–31, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/76/m076p025.pdf.

vanTreeck:1997:ISC

[vTS97] P. van Treeck and H. Schuhmacher. Initial survival of coral nubbins transplanted by a new coral transplantation technology-options for reef rehabilitation. Marine Ecology Progress Series, 150:287–292, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/150/m150p287.pdf.

vanTamelen:1997:RPB

[vTSD97] P. G. van Tamelen, M. S. Stekoll, and L. Deysher. Recovery processes of the brown alga Fucus gardneri following the ‘Exxon Valdez’ oil spill: settlement and recruitment. Marine Ecology Progress Series, 160:265–277, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
160/m160p265.pdf.

VanDuren:1995:SBD

[VV95] L. A. Van Duren and J. J. Videler. Swimming behaviour of developmental stages of the calanoid copepod Temora longicornis at different food concentrations. Marine Ecology Progress Series, 126:153–161, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/126/m126p153.pdf.

Vanreusel:1997:FLM

[VVT97] A. Vanreusel, I. Van den Bossche, and F. Thiermann. Freeliving marine nematodes from hydrothermal sediments: similarities with communities from diverse reduced habitats. Marine Ecology Progress Series, 157:207–219, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
157/m157p207.pdf.

vanderVeer:1993:MGO

[vW93] H. W. van der Veer and J. IJ. Witte. The ’maximum growth/optimal food condition’ hypothesis: a test for 0group plaice Pleuronectes platessa in the Dutch Wadden Sea. Marine Ecology Progress Series, 101:81–90, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
101/m101p081.pdf.

vanWoesik:1998:LHM

[vW98] R. van Woesik. Lesion healing on massive Porites spp. corals. Marine Ecology Progress Series, 164:213–220, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
164/m164p213.pdf.

Vanhove:1995:DSM

[VWD^{+}95] S. Vanhove, J. Wittoeck, G. Desmet, B. Van den Berghe, R. L. Herman, R. P. M. Bak, G. Nieuwland, J. H. Vosjan, A. Boldrin, S. Rabitti, and M. Vincx. Deep-sea meiofauna communities in Antarctica: structural analysis and relation with the environment. Marine Ecology Progress Series, 127:65–76, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/127/m127p065.pdf.

Vanderklift:1996:ERT

[VWJ96] M. A. Vanderklift, T. J. Ward, and C. A. Jacoby. Effect of reducing taxonomic resolution on ordinations to detect pollution-induced gradients in macrobenthic infaunal assemblages. Marine Ecology Progress Series, 136:137–145, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
136/m136p137.pdf.

Verduin:1996:LSS

[VWK96] J. J. Verduin, D. I. Walker, and J. Kuo. In situ submarine pollination in the seagrass Amphibolis antarctica: research notes. Marine Ecology Progress Series, 133:307–309, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
133/m133p307.pdf.

Vadas:1990:RANa

[VWM90a] R. L. Vadas, W. A. Wright, and S. L. Miller. Recruitment of Ascophyllum nodosum: wave action as a source of mortality. Marine Ecology Progress Series, 61:263–272, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
61/m061p263.pdf.

Vadas:1990:RANb

[VWM90b] R. L. Vadas, W. A. Wright, and S. L. Miller. Recruitment of Ascophyllum nodosum: wave action as a source of mortality. Marine Ecology Progress Series, 61:263–272, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
61/m061p263.pdf.

Velimirov:1992:SCS

[VWS92] B. Velimirov and M. Walenta-Simon. Seasonal changes in specific growth rates, production and biomass of a bacterial community in the water column above a Mediterranean seagrass system. Marine Ecology Progress Series, 80:237–248, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
80/m080p237.pdf.

Velimirov:1993:BGR

[VWS93] B. Velimirov and M. Walenta-Simon. Bacterial growth rates and productivity within a seagrass system: seasonal variations in a Posidoma oceanica bed. Marine Ecology Progress Series, 96:101–107, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/96/m096p101.pdf.

Vazquez:1996:RCA

[VY96] E. Vázquez and C. M. Young. Responses of compound ascidian larvae to haloclines. Marine Ecology Progress Series, 133:179–190, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/133/m133p179.pdf.

Witting:1995:PSB

[WA95] D. A. Witting and K. W. Able. Predation by sevenspine bay shrimp Crangon septemspinosa on winter flounder Pleuronectes americanus during settlement: laboratory observations. Marine Ecology Progress Series, 123:23–31, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
123/m123p023.pdf.

Wagner:1999:CBE

[WA99a] C. M. Wagner and H. M. Austin. Correspondence between environmental gradients and summer littoral fish assemblages in low salinity reaches of the Chesapeake Bay, USA. Marine Ecology Progress Series, 177:197–212, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
177/m177p197.pdf.

Widdicombe:1999:MIE

[WA99b] S. Widdicombe and M. C. Austen. Mesocosm investigation into the effects of bioturbation on the diversity and structure of a subtidal macrobenthic community. Marine Ecology Progress Series, 189:181–193, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/189/m189p181.pdf.

Wahl:1996:FSF

[Wah96] M. Wahl. Fouled snails in flow: potential of epibionts on Littorina littorea to increase drag and reduce snail growth rates. Marine Ecology Progress Series, 138:157–168, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
138/m138p157.pdf.

Wahl:1997:IDR

[Wah97] M. Wahl. Increased drag reduces growth of snails: comparison of flume and in situ experiments. Marine Ecology Progress Series, 151:291–293, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/151/m151p291.pdf.

Wainright:1990:SWF

[Wai90] S. C. Wainright. Sediment-to-water fluxes of particulate material and microbes by resuspension and their contribution to the planktonic food web. Marine Ecology Progress Series, 62:271–281, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/62/m062p271.pdf.

Warren:1990:RBR

[War90] J. H. Warren. Role of burrows as refuges from sub-tidal predators of temperate mangrove crabs. Marine Ecology Progress Series, 67:295–299, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/67/m067p295.pdf.

Ward:1997:RVL

[WASSM97] P. Ward, A. Atkinson, S. B. Schnack-Schiel, and A. W. A. Murray. Regional variation in the life cycle of Rhincalanus gigas (Copepoda: Calanoida) in the Atlantic sector of the Southern Ocean — re-examination of existing data (1928 to 1993). Marine Ecology Progress Series, 157:261–275, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
157/m157p261.pdf.

Watters:1998:PPH

[Wat98] G. Watters. Prevalences of parasitized and hyperparasitized crabs near South Georgia. Marine Ecology Progress Series, 170:215–229, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/170/m170p215.pdf.

Waycott:1995:AGV

[Way95] M. Waycott. Assessment of genetic variation and clonality in the seagrass Posidonia australis using RAPD and allozyme analysis. Marine Ecology Progress Series, 116:289–295, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
116/m116p289.pdf.

Williams:1991:DCF

[WB91] P. J. Williams and J. A. Brown. Developmental changes in foraging-predator avoidance trade-offs in larval lumpfish Cyclopterus lumpus. Marine Ecology Progress Series, 76: 53–60, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/76/m076p053.pdf.

Warburton:1992:PRR

[WB92a] K. Warburton and S. J. M. Blaber. Patterns of recruitment and resource use in a shallow-water fish assemblage in Moreton Bay, Queensland. Marine Ecology Progress Series, 90:113–126, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/90/m090p113.pdf.

Williams:1992:DCE

[WB92b] P. J. Williams and J. A. Brown. Development changes in the escape response of larval winter flounder Pleuronecles amencanus from hatch through metamorphosis. Marine Ecology Progress Series, 88:185–193, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/88/m088p185.pdf.

Walters:1994:SCE

[WB94] K. Walters and S. S. Bell. Significance of copepod emergence to benthic, pelagic, and phytal linkages in a subtidal seagrass bed. Marine Ecology Progress Series, 108:237–249, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
108/m108p237.pdf.

Wilson:1997:CDC

[WB97] S. Wilson and D. R. Bellwood. Cryptic dietary components of territorial damselfishes (Pomacentridae, Labroidei). Marine Ecology Progress Series, 153:299–310, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
153/m153p299.pdf.

Wright:1999:ISS

[WB99] J. R. Wright and A. J. Boxshall. The influence of smallscale flow and chemical cues on the settlement of two congeneric barnacle species. Marine Ecology Progress Series, 183:179–187, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/183/m183p179.pdf.

Williams:1991:YAM

[WBC91] W. T. Williams, J. S. Bunt, and H. J. Clay. Yet another method of species-sequencing. Marine Ecology Progress Series, 72:283–287, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/72/m072p283.pdf.

Witte:1997:PCD

[WBGS97] U. Witte, T. Brattegard, G. Graf, and B. Springer. Particle capture and deposition by deep-sea sponges from the Norwegian-Greenland Sea. Marine Ecology Progress Series, 154:241–252, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/154/m154p241.pdf.

Wing:1995:SSR

[WBLM95] S. R. Wing, L. W. Botsford, J. L. Largier, and L. E. Morgan. Spatial structure of relaxation events and crab settlement in the northern California upwelling system. Marine Ecology Progress Series, 128:199–211, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/128/m128p199.pdf.

Watkins:1992:VRS

[WBP^{+}92] J. L. Watkins, F. Buchholz, J. Priddle, D. J. Morris, and C. Ricketts. Variation in reproductive status of Antarctic krill swarms; evidence for a size-related sorting mechanism? Marine Ecology Progress Series, 82:163–174, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
82/m082p163.pdf.

Wilkinson:1990:CSP

[WC90] C. R. Wilkinson and A. C. Cheshire. Comparisons of sponge populations across the Barrier Reefs of Australia and Belize: evidence for higher productivity in the Caribbean. Marine Ecology Progress Series, 67:285–294, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
67/m067p285.pdf.

Woods:1991:LHR

[WC91] R. E. Woods and B. C. Coull. Life history responses of Amphiascus tenuiremis (Copepoda, Harpacticoida) to mimicked predation. Marine Ecology Progress Series, 79:225–234, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
79/m079p225.pdf.

Wolff:1992:FEC

[WC92] M. Wolff and G. Cerda. Feeding ecology of the crab Cancer polyodon in La Herradura Bay, northern Chile. I. Feeding chronology, food intake, and gross growth and ecological efficiency. Marine Ecology Progress Series, 89:213–219, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
89/m089p213.pdf.

Warwick:1993:CSD

[WC93] R. M. Warwick and K. R. Clarke. Comparing the severity of disturbance: a metaanalysis of marine macrobenthic community data. Marine Ecology Progress Series, 92:221–232, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
92/m092p221.pdf.

Warwick:1995:NBM

[WC95] R. M. Warwick and K. R. Clarke. New ’biodiversity’ measures reveal a decrease in taxonomic distinctness with increasing stress. Marine Ecology Progress Series, 129:301–305, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
129/m129p301.pdf.

Worm:1996:ICA

[WC96] B. Worm and A. R. O. Chapman. Interference competition among two intertidal seaweeds: Chondrus crispus strongly affects survival of Fucus evanescens recruits. Marine Ecology Progress Series, 145:297–301, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/145/m145p297.pdf.

Weimerskirch:1998:FES

[WC98] H. Weimerskirch and Y. Cherel. Feeding ecology of shorttailed shearwaters: breeding in Tasmania and foraging in the Antarctic? Marine Ecology Progress Series, 167:261– 274, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/167/m167p261.pdf.

Wilson:1994:MAE

[WCBL94] R. P. Wilson, B. M. Culik, R. Bannasch, and J. Lage. Monitoring Antarctic environmental variables using penguins. Marine Ecology Progress Series, 106:199–202, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
106/m106p199.pdf.

Wetherbee:1997:DRD

[WCL97] B. M. Wetherbee, G. L. Crow, and C. G. Lowe. Distribution, reproduction and diet of the gray reef shark Carcharhinus amblyrhynchos in Hawaii. Marine Ecology Progress Series, 151:181–189, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/151/m151p181.pdf.

Winning:1999:EMS

[WCLB99] M. A. Winning, R. M. Connolly, N. R. Loneragan, and S. E. Bunn. ^{15}N enrichment as a method of separating the isotopic signatures of seagrass and its epiphytes for food web analysis. Marine Ecology Progress Series, 189:289–294, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
189/m189p289.pdf.

Watson:1998:COM

[WCP^{+}98] G. J. Watson, P. S. Cadman, L. A. Paterson, M. G. Bentley, and M. F. Auckland. Control of oocyte maturation, sperm activation and spawning in two lugworm species: Arenicola marina and A. defodiens. Marine Ecology Progress Series, 175:167–176, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/175/m175p167.pdf.

Wieczorek:1995:IFE

[WCT95] S. K. Wieczorek, A. S. Clare, and C. D. Todd. Inhibitory and facilitatory effects of microbial films on settlement of Balanus amphitrite amphitrite larvae. Marine Ecology Progress Series, 119:221–228, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/119/m119p221.pdf.

Wehrtmann:1990:UBM

[WD90] I. S. Wehrtmann and A. I. Dittel. Utilization of bating mangrove leaves as a transport mechanism of estuarine organisms, with emphasis on decapod crustacea. Marine Ecology Progress Series, 60:67–73, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/60/m060p067.pdf.

Wolcott:1994:OMC

[WD94] D. L. Wolcott and M. C. De Vries. Offshore megalopae of Callinectes sapidus: depth of collection, molt stage and response to estuarine cues. Marine Ecology Progress Series, 109:157–163, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/109/m109p157.pdf.

Wang:1999:FAT

[WD99] W. X. Wang and R. C. H. Dei. Factors affecting trace element uptake in the black mussel Septifer virgatus . Marine Ecology Progress Series, 186:161–172, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/186/m186p161.pdf.

Widdows:1995:SGC

[WDB^{+}95] J. Widdows, P. Donkin, M. D. Brinsley, S. V. Evans, P. N. Salkeld, A. Franklin, R. J. Law, and M. J. Waldock. Scope for growth and contaminant levels in North Sea mussels Mytilus edulis. Marine Ecology Progress Series, 127:131–148, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
127/m127p131.pdf.

Wagner:1998:RDR

[WDB98] M. Wagner, E. Durbin, and L. Buckley. RNA: DNA ratios as indicators of nutritional condition in the copepod Calanus finmarchicus. Marine Ecology Progress Series, 162: 173–181, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/162/m162p173.pdf.

Wlodarczyk:1992:ETL

[WDD92] E. Wlodarczyk, A. G. Durbin, and E. G. Durbin. Effect of temperature on lower feeding thresholds, gut evacuation rate, and diel feeding behavior in the copepod Acartia hudsonica. Marine Ecology Progress Series, 85:93–106, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
85/m085p093.pdf.

Welch:1995:EVP

[WE95] J. M. Welch and C. E. Epifanio. Effect of variations in prey abundance on growth and development of crab larvae reared in the laboratory and in large field-deployed enclosures. Marine Ecology Progress Series, 116:55–64, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
116/m116p055.pdf.

Webb:1991:EPJa

[Web91a] D. G. Webb. Effect of predation by juvenile Pacific salmon on marine harpacticoid copepods. I. Comparisons of patterns of copepod mortality with patterns of salmon consumption. Marine Ecology Progress Series, 72:25–36, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
72/m072p025.pdf.

Webb:1991:EPJb

[Web91b] D. G. Webb. Effect of predation by juvenile Pacific salmon on marine harpacticoid copepods. II. Predator density manipulation experiments. Marine Ecology Progress Series, 72: 37–47, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/72/m072p037.pdf.

Wehrtmann:1991:HIS

[Weh91] I. S. Wehrtmann. How important are starvation periods in early larval development for survival of Crangon septemspinosa larvae? Marine Ecology Progress Series, 73:183– 190, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/73/m073p183.pdf.

Werner:1997:GAU

[Wer97] I. Werner. Grazing of Arctic under-ice amphipods on seaice algae. Marine Ecology Progress Series, 160:93–99, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
160/m160p093.pdf.

Weston:1990:QEMa

[Wes90a] D. P. Weston. Quantitative examination of macrobenthic community changes along an organic enrichment gradient. Marine Ecology Progress Series, 61:233–244, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
61/m061p233.pdf.

Weston:1990:QEMb

[Wes90b] D. P. Weston. Quantitative examination of macrobenthic community changes along an organic enrichment gradient. Marine Ecology Progress Series, 61:233–244, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
61/m061p233.pdf.

Wiencke:1990:GSC

[WF90] C. Wiencke and G. Fischer. Growth and stable carbon isotope composition of cold-water macroalgae in relation to light and temperature. Marine Ecology Progress Series, 65:283– 292, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/65/m065p283.pdf.

Wilcox:1994:BVC

[WF94] R. M. Wilcox and J. A. Fuhrman. Bacterial viruses in coastal seawater: lytic rather than lysogenic production. Marine Ecology Progress Series, 114:35–45, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/114/m114p035.pdf.

Wang:1997:MIB

[WF97] W. X. Wang and N. S. Fisher. Modeling the influence of body size on trace element accumulation in the mussel Mytilus edulis. Marine Ecology Progress Series, 161:103–115, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
161/m161p103.pdf.

Welch:1999:BRB

[WFH99] J. M. Welch, R. B. Forward, Jr., and P. A. Howd. Behavioral responses of blue crab Callinectes sapidus postlarvae to turbulence: implications for selective tidal stream transport. Marine Ecology Progress Series, 179:135–143, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
179/m179p135.pdf.

Wanless:1999:EDL

[WFHM99] S. Wanless, S. K. Finney, M. P. Harris, and D. J. McCafferty. Effect of the diel light cycle on the diving behaviour of two bottom feeding marine birds: the blue-eyed shag Phalacrocorax atriceps and the European shag P. aristotelis. Marine Ecology Progress Series, 188:219–224, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
188/m188p219.pdf.

Wang:1995:ATE

[WFL95] W. X. Wang, N. S. Fisher, and S. N. Luoma. Assimilation of trace elements ingested by the mussel Mytilus edulis: effects of algal food abundance. Marine Ecology Progress Series, 129:165–176, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/129/m129p165.pdf.

Wang:1996:KDT

[WFL96] W. X. Wang, N. S. Fisher, and S. N. Luoma. Kinetic determinations of trace element bioaccumulation in the mussel Mytilus edulis. Marine Ecology Progress Series, 140:91– 113, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/140/m140p091.pdf.

Waite:1997:SRV

[WFTH97] A. Waite, A. Fisher, P. A. Thompson, and P. J. Harrison. Sinking rate versus cell volume relationships illuminate sinking rate control mechanisms in marine diatoms. Marine Ecology Progress Series, 157:97–108, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/157/m157p097.pdf.

Worcester:1997:QHC

[WG97] S. E. Worcester and S. D. Gaines. Quantifying hermit crab recruitment rates and megalopal shell selection on wave-swept shores. Marine Ecology Progress Series, 157:307–310, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
157/m157p307.pdf.

Waite:1997:NMP

[WGD97] A. Waite, S. Gallager, and H. G. Dam. New measurements of phytoplankton aggregation in a flocculator using videography and image analysis. Marine Ecology Progress Series, 155: 77–88, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/155/m155p077.pdf.

Watanabe:1991:DGS

[WH91] J. M. Watanabe and C. Harrold. Destructive grazing by sea urchins Strongylocentrotus spp. in a central California kelp forest: potential roles of recruitment, depth, and predation. Marine Ecology Progress Series, 71:125–141, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
71/m071p125.pdf.

Waite:1992:RSA

[WH92] A. Waite and P. J. Harrison. Role of sinking and ascent during sexual reproduction in the marine diatom Ditylum brightwelli. Marine Ecology Progress Series, 87:113–122, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
87/m087p113.pdf.

Wassenberg:1993:DFB

[WH93] T. J. Wassenberg and B. J. Hill. Diet and feeding behaviour of juvenile and adult banana prawns Penaeus merguiensis in the Gulf of Carpentaria, Australia. Marine Ecology Progress Series, 94:287–295, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/94/m094p287.pdf.

Wang:1995:DRA

[WH95] S. B. Wang and E. D. Houde. Distribution, relative abundance, biomass and production of bay anchovy Anchoa mitchilli in the Chesapeake Bay. Marine Ecology Progress Series, 121:27–38, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/121/m121p027.pdf.

Ward:1996:ETM

[WH96] T. J. Ward and P. A. Hutchings. Effects of trace metals on infaunal species composition in polluted intertidal and subtidal marine sediments near a lead smelter, Spencer Gulf, South Australia. Marine Ecology Progress Series, 135:123– 135, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/135/m135p123.pdf.

Wolfstein:1998:PLD

[WH98] K. Wolfstein and P. Hartig. The photosynthetic light dispensation system: application to microphytobenthic primary production measurements. Marine Ecology Progress Series, 166: 63–71, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/166/m166p063.pdf.

Watermann:1999:CBB

[WHG^{+}99a] F. Watermann, H. Hillebrand, G. Gerdes, W. E. Krumbein, and U. Sommer. Competition between benthic cyanobacteria and diatoms as influenced by different grain sizes and temperatures. Marine Ecology Progress Series, 187:77–87, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
187/m187p077.pdf.

Wijsman:1999:SDS

[WHG99b] J. W. M. Wijsman, P. M. J. Herman, and M. T. Gomoiu. Spatial distribution in sediment characteristics and benthic activity on the northwestern Black Sea shelf. Marine Ecology Progress Series, 181:25–39, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/181/m181p025.pdf.

Whitfield:1994:ALJ

[Whi94] A. K. Whitfield. Abundance of larval and 0+ juvenile marine fishes in the lower reaches of three southern African estuaries with differing freshwater inputs. Marine Ecology Progress Series, 105:257–267, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/105/m105p257.pdf.

Weiss:1997:CLV

[WHIC97] G. M. Weiss, L. W. Harding, Jr., E. C. Itsweire, and J. W. Campbell. Characterizing lateral variability of phytoplankton chlorophyll in Chesapeake Bay with aircraft ocean color data. Marine Ecology Progress Series, 149:183–199, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
149/m149p183.pdf.

Wohrmann:1997:AAS

[WHK97] A. P. A. Wöhrmann, W. Hagen, and A. Kunzmann. Adaptations of the Antarctic silverfish Pleuragramma antarcticum (Pisces: Nototheniidae) to pelagic life in high-Antarctic waters. Marine Ecology Progress Series, 151:205–218, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
151/m151p205.pdf.

West:1993:MPG

[WHW93] J. M. West, C. D. Harvell, and A.-M. Walls. Morphological plasticity in a gorgonian coral (Briareum asbestinum) over a depth dine. Marine Ecology Progress Series, 94:61–69, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
94/m094p061.pdf.

Williams:1992:SDC

[Wil92] T. D. Williams. Survival and development of copepod larvae Tisbe battagliai in surface microlayer, water and sediment elutriates from the German Bight. Marine Ecology Progress Series, 91:221–228, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/91/m091p221.pdf.

Witte:1999:CLC

[Wit99] U. Witte. Consumption of large carcasses by scavenger assemblages in the deep Arabian Sea: observations by baited camera. Marine Ecology Progress Series, 183:139–147, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
183/m183p139.pdf.

Weeks:1993:ABS

[WJD93] J. M. Weeks, F. B. Jensen, and M. H. Depledge. Acid-base status, haemolymph composition and tissue copper accumulation in the shore crab Carcinus maenas exposed to combined copper and salinity stress. Marine Ecology Progress Series, 97:91–98, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/97/m097p091.pdf.

Wahl:1994:CCB

[WJF94] M. Wahl, P. R. Jensen, and W. Fenical. Chemical control of bacterial epibiosis on ascidians. Marine Ecology Progress Series, 110:45–57, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/110/m110p045.pdf.

Wright:1991:IHM

[WJM^{+}91] S. W. Wright, S. W. Jeffrey, R. F. C. Mantoura, C. A. Llewellyn Wright, T. Bjørnland, D. Repeta, and N. Welschmeyer. Improved HPLC method for the analysis of chlorophylls and carotenoids from marine phytoplankton. Marine Ecology Progress Series, 77:183–196, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/77/m077p183.pdf.

Windoffer:1999:SIM

[WJM^{+}99] R. Windoffer, A. Jahn, F. Meyberg, J. Krieger, and O. Giere. Sulphide-induced metal precipitation in the mantle edge of Macoma balthica (Bivalvia, Tellinidae) — a means of detoxification. Marine Ecology Progress Series, 187:159–170, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
187/m187p159.pdf.

Wilkerson:1992:DDP

[WK92a] F. P. Wilkerson and P. Kremer. DIN, DON and P04 flux by a medusa with algal symbionts. Marine Ecology Progress Series, 90:237–250, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/90/m090p237.pdf.

Winn:1992:ESE

[WK92b] R. N. Winn and D. M. Knott. An evaluation of the survival of experimental populations exposed to hypoxia in the Savannah River estuary. Marine Ecology Progress Series, 88: 161–179, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/88/m088p161.pdf.

Wolfe:1993:RCI

[WK93] G. V. Wolfe and R. P. Kiene. Radioisotope and chemical inhibitor measurements of dimethyl sulfide consumption rates and kinetics in estuarine waters. Marine Ecology Progress Series, 99:261–269, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/99/m099p261.pdf.

Wilson:1998:MRA

[WK98] R. P. Wilson and M. A. M. Kierspel. A method for retrieval of anchored stomach probes from seabirds. Marine Ecology Progress Series, 163:295–297, 1998. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/163/m163p295.pdf.

Wiens:1998:IGE

[WKH^{+}98] M. Wiens, C. Koziol, H. M. A. Hassanein, R. Batel, H. C. Schröder, and W. E. G. Müller. Induction of gene expression of the chaperones 14-3-3 and HSP70 by PCB 118 (2,3’,4,4’,5pentachlorobiphenyl) in the marine sponge Geodia cydonium: novel biomarkers for polychlorinated biphenyls. Marine Ecology Progress Series, 165:247–257, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Wallace:1997:BBS

[WL97] W. G. Wallace and G. R. Lopez. Bioavailability of biologically sequestered cadmium and the implications of metal detoxification. Marine Ecology Progress Series, 147:149–157, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
147/m147p149.pdf.

Worm:1999:MDS

[WLB^{+}99] B. Worm, H. K. Lotze, C. Boström, R. Engkvist, V. Labanauskas, and U. Sommer. Marine diversity shift linked to interactions among grazers, nutrients and propagule banks. Marine Ecology Progress Series, 185:309–314, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
185/m185p309.pdf.

Woodin:1997:BBN

[WLL97] S. A. Woodin, S. M. Lindsay, and D. E. Lincoln. Biogenic bromophenols as negative recruitment cues. Marine Ecology Progress Series, 157:303–306, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/157/m157p303.pdf.

Wallace:1998:CRO

[WLL98] W. G. Wallace, G. R. Lopez, and J. S. Levinton. Cadmium resistance in an oligochaete and its effect on cadmium trophic transfer to an omnivorous shrimp. Marine Ecology Progress Series, 172:225–237, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/172/m172p225.pdf.

Warnau:1995:ECC

[WLT^{+}95] M. Warnau, G. Ledent, A. Temara, M. Jangoux, and P. Dubois. Experimental cadmium contamination of the echinoid Paracentrotus tividus: influence of exposure mode and distribution of the metal in the organism. Marine Ecology Progress Series, 116:117–124, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/116/m116p117.pdf.

Welborn:1990:DMS

[WM90] J. R. Welborn and D. T. Manahan. Direct measurements of sugar uptake from seawater into molluscan larvae. Marine Ecology Progress Series, 65:233–239, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
65/m065p233.pdf.

Wilkins:1992:MUA

[WM92] H. K. A. Wilkins and A. A. Myers. Microhabitat utilisation by an assemblage of temperate Gobiidae (Pisces: Teleostei). Marine Ecology Progress Series, 90:103–112, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
90/m090p103.pdf.

Williams:1995:HPT

[WM95] G. A. Williams and D. Morritt. Habitat partitioning and thermal tolerance in a tropical limpet, Cellana grata. Marine Ecology Progress Series, 124:89–103, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
124/m124p089.pdf.

Wilson:1997:STV

[WM97] D. T. Wilson and M. I. McCormick. Spatial and temporal validation of settlement-marks in the otoliths of tropical reef fishes. Marine Ecology Progress Series, 153:259–271, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
153/m153p259.pdf.

Wahl:1999:PFN

[WM99] M. Wahl and O. Mark. The predominantly facultative nature of epibiosis: experimental and observational evidence. Marine Ecology Progress Series, 187:59–66, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/187/m187p059.pdf.

Walters:1999:EOI

[WMB99] L. J. Walters, G. Miron, and E. Bourget. Endoscopic observations of invertebrate larval substratum exploration and settlement. Marine Ecology Progress Series, 182:95–108, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
182/m182p095.pdf.

Watkins:1999:VDA

[WMD99] J. L. Watkins, A. W. A. Murray, and H. I. Daly. Variation in the distribution of Antarctic krill Euphausia superba around South Georgia. Marine Ecology Progress Series, 188: 149–160, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/188/m188p149.pdf.

Wells:1991:EIT

[WMG91] M. L. Wells, L. M. Mayer, and R. R. L. Guillard. Evaluation of iron as a triggering factor for red tide blooms. Marine Ecology Progress Series, 69:93–102, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/69/m069p093.pdf.

Westermeier:1994:PBD

[WMG^{+}94] R. Westermeier, D. G. Muller, I. Gomez, P. Rivera, and H. Wenzel. Population biology of Durvillaea antarctica and Lessonia nigrescens (Phaeophyta) on the rocky shores of southern Chile. Marine Ecology Progress Series, 110:187– 194, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/110/m110p187.pdf.

Weiss:1996:DLC

[WMH96] G. M. Weiss, G. B. McManus, and H. R. Harvey. Development and lipid composition of the harpacticoid copepod Nitocra spinipes reared on different diets. Marine Ecology Progress Series, 132:57–61, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/132/m132p057.pdf.

Williams:1994:STG

[WMPT94] R. Williams, H. McCall, R. W. Pierce, and J. T. Turner. Speciation of the tintinnid genus Cymatocylis by morphometric analysis of the loricae. Marine Ecology Progress Series, 107:263–272, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/107/m107p263.pdf.

Wikfors:1991:CBP

[WNJ91] G. H. Wikfors, A. Neeman, and P. J. Jackson. Cadmiumbinding polypeptides in microalgal strains with laboratoryinduced cadmium tolerance. Marine Ecology Progress Series, 79:163–170, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/79/m079p163.pdf.

Watson:1994:EAF

[WO94] M. Watson and R. F. G. Ormond. Effect of an artisanal fishery on the fish and urchin populations of a Kenyan coral reef. Marine Ecology Progress Series, 109:115–129, 1994. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
109/m109p115.pdf.

Wohrmann:1996:AGP

[Wöh96] A. P. A. Wöhrmann. Antifreeze glycopeptides and peptides in Antarctic fish species from the Weddell Sea and the Lazarev Sea. Marine Ecology Progress Series, 130:47–59, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
130/m130p047.pdf.

Wolfrath:1992:BFC

[Wol92a] B. Wolfrath. Burrowing of the fiddler crab Uca tangeri in the Ria Formosa in Portugal and its influence on sediment structure. Marine Ecology Progress Series, 85:237–243, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
85/m085p237.pdf.

Wolfrath:1992:FEF

[Wol92b] B. Wolfrath. Field experiments on feeding of European fiddler crab Uca tangeri. Marine Ecology Progress Series, 90: 39–43, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/90/m090p039.pdf.

Wolfrath:1993:OBE

[Wol93] B. Wolfrath. Observations on the behaviour of the European fiddler crab Uca tangeri. Marine Ecology Progress Series, 100:111–118, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/100/m100p111.pdf.

Witek:1997:PPP

[WOM^{+}97] Z. Witek, S. Ochocki, M. Maciejowska, M. Pastuszak, J. Nakonieczny, B. Podgórska, J. M. Kownacka, T. Mackiewicz, and M. Wrzesinska-Kwiecien. Phytoplankton primary production and its utilization by the pelagic community in the coastal zone of the Gulf of Gdansk (southern Baltic). Marine Ecology Progress Series, 148:169–186, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
148/m148p169.pdf.

Worcester:1995:EEB

[Wor95] S. E. Worcester. Effects of eelgrass beds on advection and turbulent mixing in low current and low shoot density environments. Marine Ecology Progress Series, 126:223–232, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
126/m126p223.pdf.

Webb:1992:WSR

[WP92] D. G. Webb and T. R. Parsons. Winter-spring recruitment pattern of epiphytic harpacticoid copepods in a temperatezone seagrass bed. Marine Ecology Progress Series, 82:151– 162, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/82/m082p151.pdf.

Weinbauer:1994:FSD

[WP94] M. G. Weinbauer and P. Peduzzi. Frequency, size and distribution of bacteriophages in different marine bacterial morphotypes. Marine Ecology Progress Series, 108:11–20, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
108/m108p011.pdf.

Weinbauer:1995:EVR

[WP95] M. G. Weinbauer and P. Peduzzi. Effect of virus-rich high molecular weight concentrates of seawater on the dynamics of dissolved amino acids and carbohydrates. Marine Ecology Progress Series, 127:245–253, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/127/m127p245.pdf.

Wilson:1993:DDD

[WPB^{+}93] Ft. P. Wilson, K. Puetz, C. A. Bost, B. M. Culik, R. Bannasch, T. Reins, and D. Adelung. Diel dive depth in penguins in relation to diel vertical migration of prey: whose dinner by candlelight? Marine Ecology Progress Series, 94: 101–104, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/94/m094p101.pdf.

Willey:1999:IRH

[WPG99] J. D. Willey, H. W. Paerl, and M. Go. Impact of rainwater hydrogen peroxide on chlorophyll-a content of surface Gulf Stream seawater off North Carolina, USA. Marine Ecology Progress Series, 178:145–150, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/178/m178p145.pdf.

Whitehouse:1996:SOO

[WPTB96] M. J. Whitehouse, J. Priddle, P. N. Trathan, and M. A. Brandon. Substantial open-ocean phytoplankton blooms to the north of South Georgia, South Atlantic, during summer. Marine Ecology Progress Series, 140:187–197, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
140/m140p187.pdf.

Wood:1998:WCT

[WPY98] A. M. Wood, D. A. Phinney, and C. S. Yentsch. Water column transparency and the distribution of spectrally distinct forms of phycoerythrin-containing organisms. Marine Ecology Progress Series, 162:25–31, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/162/m162p025.pdf.

Wang:1999:TTC

[WQQ99] W. X. Wang, J. W. Qiu, and P. Y. Qian. The trophic transfer of Cd, Cr, and Se in the barnacle Balanus amphitrite from planktonic food. Marine Ecology Progress Series, 187: 191–201, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/187/m187p191.pdf.

White:1991:MZG

[WR91] J. R. White and M. R. Roman. Measurement of zooplankton grazing using particles labeled in light and dark with [methyl-3H] methylamine hydrochloride. Marine Ecology Progress Series, 71:45–52, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/71/m071p045.pdf.

White:1992:EPC

[WR92a] J. R. White and M. R. Roman. Egg production by the calanoid copepod Arcartia tonsa in the mesohaline Chesapeake Bay: the importance of food resources and temperature. Marine Ecology Progress Series, 86:239–249, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
86/m086p239.pdf.

White:1992:SSG

[WR92b] J. R. White and M. R. Roman. Seasonal study of grazing by metazoan zooplankton in the mesohaline Chesapeake Bay. Marine Ecology Progress Series, 86:251–261, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
86/m086p251.pdf.

Wienecke:1997:FSE

[WR97] B. C. Wienecke and G. Robertson. Foraging space of emperor penguins Aptenodytes forsteri in Antarctic shelf waters in winter. Marine Ecology Progress Series, 159:249–263, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
159/m159p249.pdf.

Wunsch:1998:CEU

[WR98] M. Wunsch and C. Richter. The CaveCam — an endoscopic underwater videosystem for the exploration of cryptic habitats. Marine Ecology Progress Series, 169:277–282, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
169/m169p277.pdf.

Welch:1997:ECC

[WRBF97] J. M. Welch, D. Rittschof, T. M. Bullock, and R. B. Forward, Jr. Effects of chemical cues on settlement behavior of blue crab Callinectes sapidus postlarvae. Marine Ecology Progress Series, 154:143–153, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/154/m154p143.pdf.

Wahle:1991:RHN

[WS91] R. A. Wahle and R. S. Steneck. Recruitment habitats and nursery grounds of the American lobster Homarus americanus: a demographic bottleneck? Marine Ecology Progress Series, 69:231–243, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/69/m069p231.pdf.

Wahl:1992:MEI

[WS92a] M. Wahl and H. Sönnichsen. Marine epibiosis. IV. The periwinkle Littonna littorea lacks typical antifouling defences — why are some populations so little fouled? Marine Ecology Progress Series, 88:225–235, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/88/m088p225.pdf.

Wolff:1992:PDC

[WS92b] M. Wolff and M. Soto. Population dynamics of Cancer polyodon in La Herradura Bay, northern Chile. Marine Ecology Progress Series, 85:69–81, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/85/m085p069.pdf.

Watts:1999:MNP

[WSC^{+}99] L. J. Watts, S. Sathyendranath, C. Caverhill, H. Maass, T. Platt, and N. J. P. Owens. Modelling new production in the northwest Indian Ocean region. Marine Ecology Progress Series, 183:1–12, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/183/m183p001.pdf.

Ward:1996:MDP

[WSCT96] P. Ward, R. S. Shreeve, G. C. Cripps, and P. N. Trathan. Mesoscale distribution and population dynamics of Rhincalanus gigas and Calanus simillimus in the Antarctic polar open ocean and polar frontal zone during summer. Marine Ecology Progress Series, 140:21–32, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/140/m140p021.pdf.

Wang:1999:BDS

[WSF99] W. X. Wang, I. Stupakoff, and N. S. Fisher. Bioavailability of dissolved and sediment-bound metals to a marine depositfeeding polychaete. Marine Ecology Progress Series, 178: 281–293, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/178/m178p281.pdf.

Weisse:1991:UML

[WSM91] T. Weisse and U. Scheffel-Moser. Uncoupling the microbial loop: growth and grazing loss rates of bacteria and heterotrophic nanoflagellates in the North Atlantic. Marine Ecology Progress Series, 71:195–205, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/71/m071p195.pdf.

Wear:1999:EWC

[WSMM99] D. J. Wear, M. J. Sullivan, A. D. Moore, and D. F. Millie. Effects of water-column enrichment on the production dynamics of three seagrass species and their epiphytic algae. Marine Ecology Progress Series, 179:201–213, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
179/m179p201.pdf.

Wilson:1995:FAM

[WSP^{+}95] R. P. Wilson, J. A. Scolaro, G. Peters, S. Laurenti, M. Kierspel, H. Gallelli, and J. Upton. Foraging areas of Magellanic penguins Spheniscus magellanicus breeding at San Lorenzo, Argentina, during the incubation period. Marine Ecology Progress Series, 129:1–6, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/129/m129p001.pdf.

Wolfe:1994:RCD

[WSS94] G. V. Wolfe, E. B. Sherr, and B. F. Sherr. Release and consumption of DMSP from Emiliania huxleyi during grazing by Oxyrrhis manna. Marine Ecology Progress Series, 111: 111–119, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/111/m111p111.pdf.

Wang:1998:IGV

[WT98] C. H. Wang and W. N. Tzeng. Interpretation of geographic variation in size of American eel Anguilla rostrata elvers on the Atlantic coast of North America using their life history and otolith ageing. Marine Ecology Progress Series, 168: 35–43, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/168/m168p035.pdf.

Westgate:1999:GDO

[WT99a] A. J. Westgate and K. A. Tolley. Geographical differences in organochlorine contaminants in harbour porpoises Phocoena phocoena from the western North Atlantic. Marine Ecology Progress Series, 177:255–268, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/177/m177p255.pdf.

Wong:1999:PHK

[WT99b] M. W. Wong and D. W. Townsend. Phytoplankton and hydrography of the kennebec estuary, Maine, USA. Marine Ecology Progress Series, 178:133–144, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
178/m178p133.pdf.

Walton:1997:SLE

[WTA^{+}97] P. Walton, C. M. R. Turner, G. Austin, M. D. Burns, and P. Monaghan. Sub-lethal effects of an oil pollution incident on breeding kittiwakes Rissa tridactyla. Marine Ecology Progress Series, 155:261–268, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/155/m155p261.pdf.

Wiencke:1990:TRG

[WtD90] C. Wiencke and I. torn Dieck. Temperature requirements for growth and survival of macroalgae from Antarctica and southern Chile. Marine Ecology Progress Series, 59:157–170, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
59/m059p157.pdf.

Warnau:1995:EFC

[WTF95] M. Warnau, J. L. Teyssié, and S. W. Fowler. Effect of feeding on cadmium bioaccumulation in the echinoid Paracentrotus lividus (Echinodermata). Marine Ecology Progress Series, 126:305–309, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/126/m126p305.pdf.

Warnau:1996:BSH

[WTF96] M. Warnau, J. L. Teyssié, and S. W. Fowler. Biokinetics of selected heavy metals and radionuclides in the common Mediterranean echinoid Paracentrotus lividus: sea water and food exposures. Marine Ecology Progress Series, 141: 83–94, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/141/m141p083.pdf.

Wright:1996:APA

[WTM^{+}96] S. W. Wright, D. P. Thomas, H. J. Marchant, H. W. Higgins, M. D. Mackey, and D. J. Mackey. Analysis of phytoplankton of the Australian sector of the Southern Ocean: comparisons of microscopy and size frequency data with interpretations of pigment HPLC data using the ’CHEMTAX’ matrix factorisation program. Marine Ecology Progress Series, 144:285–298, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/144/m144p285.pdf.

Wahle:1996:LIA

[WTO96] R. A. Wahle, O. Tully, and V. O’Donovan. Lipofuscin as an indicator of age in crustaceans: analysis of the pigment in the American lobster Homarus americanus. Marine Ecology Progress Series, 138:117–123, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/138/m138p117.pdf.

Waluda:1999:IOV

[WTR99] C. M. Waluda, P. N. Trathan, and P. G. Rodhouse. Influence of oceanographic variability on recruitment in the Illex argentinus (Cephalopoda: Ommastrephidae) fishery in the South Atlantic. Marine Ecology Progress Series, 183:159– 167, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/183/m183p159.pdf.

Wesseling:1999:DRF

[WUA^{+}99] I. Wesseling, A. J. Uychiaoco, P. M. Aliño, T. Aurin, and J. E. Vermaat. Damage and recovery of four Philippine corals from short-term sediment burial. Marine Ecology Progress Series, 176:11–15, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/176/m176p011.pdf.

Weinbauer:1995:BSP

[WV95] M. G. Weinbauer and B. Velimirov. Biomass and secondary production of the temperate gorgonian coral Eunicella cavolini (Coelenterata: Octocorallia). Marine Ecology Progress Series, 121:211–216, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/121/m121p211.pdf.

Wassmann:1990:MSP

[WVMR90] P. Wassmann, M. Vernet, B. G. Mitchell, and F. Rey. Mass sedimentation of Phaeocystis pouchetii in the Barents Sea. Marine Ecology Progress Series, 66:183–195, 1990. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
66/m066p183.pdf.

Webb:1990:DHM

[WW90] P. Webb and T. H. Wooldridge. Diel horizontal migration of Mesopodopsis slabberi (Crustacea: Mysidacea) in Algoa Bay, southern Africa. Marine Ecology Progress Series, 62: 73–77, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/62/m062p073.pdf.

Wang:1991:PRM

[WW91] W. X. Wang and J. Widdows. Physiological responses of mussel larvae Mytilus edulis to environmental hypoxia and anoxia. Marine Ecology Progress Series, 70:223–236, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
70/m070p223.pdf.

Weimerskirch:1992:WDW

[WW92a] H. Weimerskirch and R. P. Wilson. When do wandering albatrosses Diomedea exulans forage? Marine Ecology Progress Series, 86:297–300, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/86/m086p297.pdf.

Wers:1992:CME

[WW92b] J. S. Wers and P. Weis. Construction materials in estuaries: reduction in the epibiotic community on chromated copper arsenate (CCA) treated wood. Marine Ecology Progress Series, 83:45–53, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/83/m083p045.pdf.

Wang:1993:MRC

[WW93] W. X. Wang and J. Widdows. Metabolic responses of the common mussel Mytilus edulis to hypoxia and anoxia. Marine Ecology Progress Series, 95:205–214, 1993. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
95/m095p205.pdf.

Walters:1996:SEP

[WW96] L. J. Walters and D. S. Wethey. Settlement and early postsettlement survival of sessile marine invertebrates on topographically complex surfaces: the importance of refuge dimensions and adult morphology. Marine Ecology Progress Series, 137:161–171, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/137/m137p161.pdf.

Whitney:1998:IVN

[WWB98] F. A. Whitney, C. S. Wong, and P. W. Boyd. Interannual variability in nitrate supply to surface waters of the Northeast Pacific Ocean. Marine Ecology Progress Series, 170:15–23, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
170/m170p015.pdf.

Waugh:1999:EME

[WWC^{+}99] S. M. Waugh, H. Weimerskirch, Y. Cherel, U. Shankar, P. A. Prince, and P. M. Sagar. Exploitation of the marine environment by two sympatric albatrosses in the Pacific Southern Ocean. Marine Ecology Progress Series, 177:243–254, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
177/m177p243.pdf.

Weimerskirch:1995:USM

[WWGK95] H. Weimerskirch, R. P. Wilson, C. Guinet, and M. Koudil. Use of seabirds to monitor sea-surface temperatures and to validate satellite remote-sensing measurements in the Southern Ocean. Marine Ecology Progress Series, 126:299–303, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps/126/m126p299.pdf.

Weimerskirch:1997:APF

[WWL97] H. Weimerskirch, R. P. Wilson, and P. Lys. Activity pattern of foraging in the wandering albatross: a marine predator with two modes of prey searching. Marine Ecology Progress Series, 151:245–254, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/151/m151p245.pdf.

Wang:1997:LSP

[WX97] W. X. Wang and Z. Z. Xu. Larval swimming and postlarval drifting behavior in the infaunal bivalve Sinonovacula constricta. Marine Ecology Progress Series, 148:71–81, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
148/m148p071.pdf.

Waser:1998:NIF

[WYY^{+}98] N. A. Waser, K. Yin, Z. Yu, K. Tada, P. J. Harrison, D. H. Turpin, and S. E. Calvert. Nitrogen isotope fractionation during nitrate, ammonium and urea uptake by marine diatoms and coccolithophores under various conditions of N availability. Marine Ecology Progress Series, 169:29–41, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
169/m169p029.pdf.

Waser:1999:NIF

[WYY^{+}99] N. A. Waser, Z. Yu, K. Yin, B. Nielsen, P. J. Harrison, D. H. Turpin, and S. E. Calvert. Nitrogen isotopic fractionation during a simulated diatom spring bloom: importance of n-starvation in controlling fractionation. Marine Ecology Progress Series, 179:291–296, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/179/m179p291.pdf.

Yap:1992:TGM

[YAG92] H. T. Yap, P. M. Alino, and E. D. Gomez. Trends in growth and mortality of three coral species (Anthozoa: Scleractinia), including effects of transplantation. Marine Ecology Progress Series, 83:91–101, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/83/m083p091.pdf.

Yamazaki:1996:TPP

[Yam96] H. Yamazaki. Turbulence problems for planktonic organisms. Marine Ecology Progress Series, 139:304–305, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
139/m139p301.pdf.

Yamahira:1997:HSA

[Yam97a] K. Yamahira. Hatching success affects the timing of spawning by the intertidally spawning puffer Takifugu niphobles. Marine Ecology Progress Series, 155:239–248, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
155/m155p239.pdf.

Yamahira:1997:PFI

[Yam97b] K. Yamahira. Proximate factors influencing spawning site specificity of the puffer fish Takifugu niphobles. Marine Ecology Progress Series, 147:11–19, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/147/m147p011.pdf.

Yoklavich:1990:HPG

[YB90] M. M. Yoklavich and K. M. Bailey. Hatching period, growth and survival of young walleye pollock Theragra chalcogramma as determined from otolith analysis. Marine Ecology Progress Series, 64:13–23, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/64/m064p013.pdf.

Yoon:1992:DOC

[YB92] W. B. Yoon and R. Benner. Denitrification and oxygen consumption in sediments of two south Texas estuaries. Marine Ecology Progress Series, 90:157–167, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
90/m090p157.pdf.

Young:1996:SSE

[YBG96] E. F. Young, G. R. Bigg, and A. Grant. A statistical study of environmental influences on bivalve recruitment in the Wash, England. Marine Ecology Progress Series, 143:121–129, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
143/m143p121.pdf.

Young:1998:MSE

[YBG^{+}98] E. F. Young, G. R. Bigg, A. Grant, P. Walker, and J. Brown. A modelling study of environmental influences on bivalve settlement in The Wash, England. Marine Ecology Progress Series, 172:197–214, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/172/m172p197.pdf.

Young:1996:BZM

[YBLL96] J. W. Young, R. W. Bradford, T. D. Lamb, and V. D. Lyne. Biomass of zooplankton and micronekton in the southern bluefin tuna fishing grounds off eastern Tasmania, Australia. Marine Ecology Progress Series, 138:1–14, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
138/m138p001.pdf.

Young:1990:FEL

[YD90] J. W. Young and T. L. O. Davis. Feeding ecology of larvae of southern bluefin, albacore and skipjack tunas (Pisces: Scombridae) in the eastern Indian Ocean. Marine Ecology Progress Series, 61:17–29, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/61/m061p017.pdf.

Yoshinaga:1991:CDP

[YFI91] I. Yoshinaga, K. Fukami, and Y. Ishida. Comparison of DNA and protein synthesis rates of bacterial assemblages between coral reef waters and pelagic waters in tropical ocean. Marine Ecology Progress Series, 76:167–174, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
76/m076p167.pdf.

Yin:1997:IWR

[YGH^{+}97] K. Yin, R. H. Goldblatt, P. J. Harrison, M. A. St. John, P. J. Clifford, and R. J. Beamish. Importance of wind and river discharge in influencing nutrient dynamics and phytoplankton production in summer in the central Strait of Georgia. Marine Ecology Progress Series, 161:173–183, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
161/m161p173.pdf.

Yin:1999:RTD

[YHC^{+}99] K. Yin, P. J. Harrison, J. Chen, W. Huang, and P. Y. Qian. Red tides during spring 1998 in Hong Kong: is El Niño responsible? Marine Ecology Progress Series, 187:289– 294, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/187/m187p289.pdf.

Yin:1996:SBC

[YHGB96] K. Yin, P. J. Harrison, R. H. Goldblatt, and R. J. Beamish. Spring bloom in the central Strait of Georgia: interactions of river discharge, winds and grazing. Marine Ecology Progress Series, 138:255–263, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/138/m138p255.pdf.

Yoshinaga:1998:PSA

[YKK^{+}98] I. Yoshinaga, M. C. Kim, N. Katanozaka, I. Imai, A. Uchida, and Y. Ishida. Population structure of algicidal marine bacteria targeting the red tide forming alga Heterosigma akashiwo (Raphidophyceae), determined by restriction fragment length polymorphism analysis of the bacterial ^{16}S ribosomal RNA genes. Marine Ecology Progress Series, 170:33–44, 1998. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
170/m170p033.pdf.

Yukihira:1998:CEM

[YKL98a] H. Yukihira, D. W. Klumpp, and J. S. Lucas. Comparative effects of microalgal species and food concentration on suspension feeding and energy budgets of the pearl oysters Pinctada margaritifera and P. maxima (Bivalvia: Pteriidae). Marine Ecology Progress Series, 171:71–84, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/171/m171p071.pdf.

Yukihira:1998:EBS

[YKL98b] H. Yukihira, D. W. Klumpp, and J. S. Lucas. Effects of body size on suspension feeding and energy budgets of the pearl oysters Pinctada margaritifera and P. maxima. Marine Ecology Progress Series, 170:119–130, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/170/m170p119.pdf.

Yukihira:1999:FAP

[YKL99] H. Yukihira, D. W. Klumpp, and J. S. Lucas. Feeding adaptations of the pearl oysters Pinctada margaritifera and P. maxima to variations in natural particulates. Marine Ecology Progress Series, 182:161–173, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/182/m182p161.pdf.

Yap:1994:EFS

[YMD94] H. T. Yap, A. R. F. Montebon, and R. M. Dizon. Energy flow and seasonality in a tropical coral reef flat. Marine Ecology Progress Series, 103:35–43, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/103/m103p035.pdf.

Yamashita:1994:ERS

[YNYK94] Y. Yamashita, S. Nagahora, H. Yamada, and D. Kitagawa. Effects of release size on survival and growth of Japanese flounder Paralichthys olivaceus in coastal waters off iwate prefecture, northeastern Japan. Marine Ecology Progress Series, 105:269–276, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/105/m105p269.pdf.

Yamamuro:1998:PDD

[YOH98] M. Yamamuro, N. Oka, and J. Hiratsuka. Predation by diving ducks on the biofouling messel Musculista senhousia in a eutrophic estuarine lagoon. Marine Ecology Progress Series, 174:101–106, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/174/m174p101.pdf.

Yamazaki:1996:COT

[YS96] H. Yamazaki and K. D. Squires. Comparison of oceanic turbulence and copepod swimming. Marine Ecology Progress Series, 144:299–301, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/144/m144p299.pdf.

Yang:1994:SVC

[YSL94] X.-H. Yang, M. I. Scranton, and C. Lee. Seasonal variations in concentration and microbial uptake of methylamines in estuarine waters. Marine Ecology Progress Series, 108: 303–312, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/108/m108p303.pdf.

Yoro:1997:CSV

[YST^{+}97] S. C. Yoro, R. Sempéré, C. Turley, M. A. Unanue, X. Durrieu de Madron, and M. Bianchi. Cross-slope variations of organic carbon and bacteria in the Gulf of Lions in relation to water dynamics (northwestern Mediterranean). Marine Ecology Progress Series, 161:255–264, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/161/m161p255.pdf.

Yuneva:1999:NCF

[YSY^{+}99] T. V. Yuneva, L. S. Svetlichny, O. A. Yunev, Z. A. Romanova, A. E. Kideys, F. Bingel, Z. Uysal, A. Yilmaz, and G. E. Shulman. Nutritional condition of female Calanus euxinus from cyclonic and anticyclonic regions of the Black Sea. Marine Ecology Progress Series, 189:195–204, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
189/m189p195.pdf.

Young:1997:PDS

[YTF97] C. M. Young, P. A. Tyler, and L. Fenaux. Potential for deep sea invasion by Mediterranean shallow water echinoids: pressure and temperature as stage-specific dispersal barriers. Marine Ecology Progress Series, 154:197–209, 1997. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
154/m154p197.pdf.

Yoshioka:1991:CSG

[YY91] P. M. Yoshioka and B. B. Yoshioka. A comparison of the survivorship and growth of shallow-water gorgonian species of Puerto Rico. Marine Ecology Progress Series, 69:253– 260, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/69/m069p253.pdf.

Zimmerman:1996:ELD

[ZA96] R. C. Zimmerman and R. S. Alberte. Effect of light/dark transition on carbon translocation in eelgrass Zostera marina seedlings. Marine Ecology Progress Series, 136:305–309, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
136/m136p305.pdf.

Zajac:1991:PEPa

[Zaj91a] R. N. Zajac. Population ecology of Polydora ligni (Polychaeta: Spionidae). I. Seasonal variation in population characteristics and reproductive activity. Marine Ecology Progress Series, 77:197–206, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/77/m077p197.pdf.

Zajac:1991:PEPb

[Zaj91b] R. N. Zajac. Population ecology of Polydora ligni (Polychaeta: Spionidae). II. Seasonal demographic variation and its potential impact on life history evolution. Marine Ecology Progress Series, 77:207–220, 1991. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/77/m077p207.pdf.

Zeng:1999:ETS

[ZAN99] C. Zeng, P. Abello, and E. Naylor. Endogenous tidal and semilunar moulting rhythms in early juvenile shore crabs Carcinus maenas: implications for adaptation to a high intertidal habitat. Marine Ecology Progress Series, 191:257– 266, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/191/m191p257.pdf.

Zwarts:1992:WKC

[ZB92] L. Zwarts and A.-M. Blomert. Why knot Calidris canutus take medium-sized Macoma balthica when six prey species are available. Marine Ecology Progress Series, 83:113–128, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
83/m083p113.pdf.

Zimmerman:1994:DNR

[ZB94] A. R. Zimmerman and R. Benner. Denitrification, nutrient regeneration and carbon mineralization in sediments of Galveston Bay, Texas, USA. Marine Ecology Progress Series, 114:275–288, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/114/m114p275.pdf.

Ziegler:1998:EMS

[ZB98] S. Ziegler and R. Benner. Ecosystem metabolism in a subtropical, seagrass-dominated lagoon. Marine Ecology Progress Series, 173:1–12, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/173/m173p001.pdf.

Ziegler:1999:DOC

[ZB99a] S. Ziegler and R. Benner. Dissolved organic carbon cycling in a subtropical seagrass-dominated lagoon. Marine Ecology Progress Series, 180:149–160, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/180/m180p149.pdf.

Ziegler:1999:NCW

[ZB99b] S. Ziegler and R. Benner. Nutrient cycling in the water column of a subtropical seagrass meadow. Marine Ecology Progress Series, 188:51–62, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/188/m188p051.pdf.

Zohary:1998:BMB

[ZBK^{+}98] T. Zohary, S. Brenner, M. D. Krom, D. L. Angel, N. Kress, W. K. W. Li, A. Neori, and Y. Z. Yacobi. Buildup of microbial biomass during deep winter mixing in a Mediterranean warmcore eddy. Marine Ecology Progress Series, 167:47–57, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
167/m167p047.pdf.

Zwarts:1992:ASV

[ZBW92] L. Zwarts, A.-M. Blomert, and J. H. Wanink. Annual and seasonal variation in the food supply harvestable by knot Calidris canutus staging in the Wadden Sea in late summer. Marine Ecology Progress Series, 83:129–139, 1992. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
83/m083p129.pdf.

Zobrist:1992:MIM

[ZC92] Erik C. Zobrist and Bruce C. Coull. Meiobenthic interactions with macrobenthic larvae and juveniles: an experimental assessment of the meiofaunal bottleneck. Marine Ecology Progress Series, 88:1–8, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/88/m088p001.pdf.

Zheng:1997:ABC

[ZCL97] W. J. Zheng, X. Y. Chen, and P. Lin. Accumulation and biological cycling of heavy metal elements in Rhizophora stylosa mangroves in Yingluo Bay, China. Marine Ecology Progress Series, 159:293–301, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/159/m159p293.pdf.

Zimmerman:1994:MDP

[ZCPA94] R. C. Zimmerman, A. Cabello-Pasini, and R. S. Alberte. Modeling daily production of aquatic macrophytes from irradiance measurements: a comparative analysis. Marine Ecology Progress Series, 114:185–196, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/114/m114p185.pdf.

Zhang:1999:MOE

[ZD99] F. Zhang and M. Dickman. Mid-ocean exchange of container vessel ballast water. 1: Seasonal factors affecting the transport of harmful diatoms and dinoflagellates. Marine Ecology Progress Series, 176:243–251, 1999. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/176/m176p243.pdf.

Zeldis:1995:SGE

[ZDJ^{+}95] J. R. Zeldis, C. S. Davis, M. R. James, S. L. Ballara, W. E. Booth, and F. H. Chang. Salp grazing: effects on phytoplankton abundance, vertical distribution and taxonomic composition in a coastal habitat. Marine Ecology Progress Series, 126:267–283, 1995. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/126/m126p267.pdf.

Zeller:1997:HRA

[Zel97] D. C. Zeller. Home range and activity patterns of the coral trout Plectropomus leopardus (Serranidae). Marine Ecology Progress Series, 154:65–77, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/154/m154p065.pdf.

Zeller:1998:SAP

[Zel98] D. C. Zeller. Spawning aggregations: patterns of movement of the coral trout Plectropomus leopardus (Serranidae) as determined by ultrasonic telemetry. Marine Ecology Progress Series, 162:253–263, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/162/m162p253.pdf.

Zdanowski:1997:RBA

[ZF97] M. K. Zdanowski and F. G. Figueiras. Relationships between the abundance of bacteria and other biota and the hydrographic variability in the Ría de Vigo, Spain. Marine Ecology Progress Series, 147:257–267, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/147/m147p257.pdf.

Zdanowski:1999:CBF

[ZF99] M. K. Zdanowski and F. G. Figueiras. CFU bacterial fraction in the estuarine upwelling ecosystem of Ría de Vigo, Spain: variability in abundance and their ecophysiological description. Marine Ecology Progress Series, 182:1–15, 1999. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
182/m182p001.pdf.

Zimmer-Faust:1994:ETP

[ZFFH^{+}94] R. K. Zimmer-Faust, D. R. Fielder, K. L. Heck, Jr., L. D. Coen, and S. G. Morgan. Effects of tethering on predatory escape by juvenile blue crabs. Marine Ecology Progress Series, 111:299–303, 1994. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/111/m111p299.pdf.

Zapata:1990:MHS

[ZGB90] C. Zapata, G. Gajardo, and J. A. Beardmore. Multilocus heterozygosity and sexual selection in the brine shrimp Artemia franciscana. Marine Ecology Progress Series, 62: 211–217, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/62/m062p211.pdf.

Zamon:1996:ACT

[ZGM^{+}96] J. E. Zamon, C. H. Greene, E. Meir, D. A. Demer, R. P. Hewitt, and S. Sexton. Acoustic characterization of the threedimensional prey field of foraging chinstrap penguins. Marine Ecology Progress Series, 131:1–10, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/131/m131p001.pdf.

Zhou:1997:PDT

[ZH97] M. Zhou and M. E. Huntley. Population dynamics theory of plankton based on biomass spectra. Marine Ecology Progress Series, 159:61–73, 1997. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/159/m159p061.pdf.

Zeeck:1990:SPR

[ZHBH90] E. Zeeck, J. Hardege, and H. Bartels-Hardege. Sex pheromones and reproductive isolation in two nereid species, Nereis succinea and Platynereis dumerilii. Marine Ecology Progress Series, 67:183–188, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/67/m067p183.pdf.

Ziebis:1996:IBS

[ZHF96] W. Ziebis, M. Huettel, and S. Forster. Impact of biogenic sediment topography on oxygen fluxes in permeable seabeds. Marine Ecology Progress Series, 140:227–237, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
140/m140p227.pdf.

Zastrow:1991:SFH

[ZHM91] C. E. Zastrow, E. D. Houde, and L. G. Morin. Spawning, fecundity, hatch-date frequency and young-of-the-year growth of bay anchovy Anchoa mitchilli in mid-Chesapeake Bay. Marine Ecology Progress Series, 73:161–171, 1991. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
73/m073p161.pdf.

Zhou:1998:OIM

[Zho98] M. Zhou. An objective interpolation method for spatiotemporal distribution of marine plankton. Marine Ecology Progress Series, 174:197–206, 1998. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps/174/m174p197.pdf.

Zamorano:1995:PSM

[ZMD95] J. H. Zamorano, C. A. Moreno, and W. E. Duarte. Postsettlement mortality in Phragmatopoma virgini (Polychaeta: Sabellariidae) at the Mehuín Marine Reserve, Chile. Marine Ecology Progress Series, 127:149–155, 1995. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
127/m127p149.pdf.

Zeng:1996:ETR

[ZN96a] C. Zeng and E. Naylor. Endogenous tidal rhythms of vertical migration in field collected zoea-1 larvae of the shore crab Carcinus maenas: implications for ebb tide offshore dispersal. Marine Ecology Progress Series, 132:71–82, 1996. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
132/m132p071.pdf.

Zeng:1996:OCW

[ZN96b] C. Zeng and E. Naylor. Occurrence in coastal waters and endogenous tidal swimming rhythms of late megalopae of the shore crab Carcinus maenas: implications for onshore recruitment. Marine Ecology Progress Series, 136:69–79, 1996. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/
136/m136p069.pdf.

Zweifel:1993:CDO

[ZNH93] Ulla Li Zweifel, Bo Norrman, and Åke Hägström. Consumption of dissolved organic carbon by marine bacteria and demand for inorganic nutrients. Marine Ecology Progress Series, 101:23–32, 1993. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/101/m101p023.pdf.

Zohary:1992:BNB

[ZR92] T. Zohary and R. D. Robarts. Bacterial numbers, bacterial production, and heterotrophic nanoplankton abundance in a warm core eddy in the Eastern Mediterranean. Marine Ecology Progress Series, 84:133–137, 1992. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/84/m084p133.pdf.

Zhirmunsky:1990:UME

[ZT90] A. V. Zhirmunsky and V. G. Tarasov. Unusual marine ecosystem in the flooded crater of Ushisher volcano. Marine Ecology Progress Series, 65:95–102, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps/65/m065p095.pdf.